A Complete Bibliography of Scientific American: 1870–1879

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 November 2023 Version 1.04

Title word cross-reference

$20 [845]. $80 [10821].

000-Pounder [917].

10 [4695]. 15 [1914]. 1870 [1494, 1195]. 1871 [396, 2573]. 1873 [3268, 3475]. 1876 [3777]. 1878 [11543].

22d [1195].

30-Ton [10960]. 300 [11925, 12362, 12552].

42D [10820].

A.R.M.M.A [4357]. Abandonment [10883]. Abbe [4703, 7191].

Abbreviating [13344]. Ability [1174]. Able [9749]. Abolish [1079]. Abolition [10685]. Aborigines [13134, 13121]. Abridgement [2566]. Abroad [647, 727, 1069, 10958, 11185, 11550, 11951, 13223, 13328]. Absinthe [13417]. Absolute [6573]. Absorb [5765]. Absorbent [8241]. Absorption [2314]. Abstract [2064]. Absurd [3648]. Abuse [13539]. Academy [5069, 5095, 5691, 6451, 5719, 7066, 9275, 9311, 8488, 8566, 8748, 8844, 9430, 9566, 9643, 9727, 9785, 9893, 9993, 10069, 10097, 10126, 11462, 12160, 10983, 11088, 11428, 12225, 13544, 12737, 12892, 12940, 13607, 13664, 5215, 10774, 13556]. Acalypha [8400]. Accident [3218, 4826, 5977, 9208, 10506, 11501, 11981, 12803]. accidents [12225]. Accomplished [173, 5410]. Account [1047]. Accounted [13703, 2715]. Accounts [3570]. Accumulating [5999]. Accumulator [11183]. Accuracy [969]. Accurate [2524, 6740, 11597]. Achenbach [573]. Acheron [13156]. Acid [607, 811, 1600, 1851, 1975, 2649, 3444, 4202, 6524, 7818, 7757, 7611, 7381, 8383, 8651, 8106, 10593, 11163, 11439, 13008, 13046, 273]. Acid-From [13008]. Acidemetry [7904]. Acids [6355]. Acme [9054, 11221]. Acoustic [5544, 12665]. Acoustics [13172]. Across [8213, 13483, 3274, 4491]. Acting [764, 3238, 3338, 5227, 4705, 3826, 4056, 5720, 7139, 8836, 10593, 11221]. Actino [3931]. Actino-Chemistry [3931]. Action [731, 770, 859, 1322, 2129, 2906, 2886, 3177, 6620, 7812, 9614, 13360, 13333]. Active [12250]. Activity [3340]. Adam [10881]. Adams [3270, 11679]. Adaptability [2462, 2488]. Adding [7561, 8871]. Addition [5040, 12150, 7935]. Additional [2251]. Address [939, 959, 1798, 1832, 3368, 6184, 6212, 10250]. Adee [9921]. Adhesion [7746]. Adirondacks [1114]. Adjust [10838, 9564]. Adjustable [1138, 690, 2014, 2107, 1938, 1318, 3329, 4470, 4498, 5045, 6218, 5580, 5578, 5330, 7138, 8138, 8456, 10892, 10042, 11140, 11482]. Adjuster [11452]. Adjusting [545, 1893, 1541, 2122, 2467, 9908]. Adjutant [11738]. Administering [11715]. Administration [1146, 2079]. Admiral [4771]. Adoption [2870]. Adriatic [12594]. Adulteration [1454, 1354, 3237, 6057, 8326, 12354, 12519, 640]. Adulterations [349, 4502, 9398]. Advance [2848, 3483]. Advancement [821, 835, 939, 959, 3243, 3370, 3402, 3504, 4809, 4835, 4862, 6110, 6144, 6169, 6257, 6311, 7441, 7470, 9061, 9984, 13197]. Advances [504]. Advantages [508, 3298, 3477, 9202, 11561]. Advertise [4045]. Advertisers [3534, 3563, 4998, 6563, 13441]. Advertising [319, 5374, 5451, 5501, 5528, 5558, 5586, 5642, 6068, 6146, 11751, 11477]. Advice [2970, 3861, 218]. Advocates [13274]. Æchmea [5830, 9440]. Aerial [57, 6382, 11446]. Aero [2363]. Aeronaut [361]. Aeronautical [674, 674]. Aeronautics [75, 9397]. Aeronauts [4820]. Aerotherapy [8824]. Aesthetics [6448, 7506]. Affairs [947, 506, 537, 563, 586, 600, 625]. Affect [1100, 11419, 11502]. Affecting [748, 9981]. Affixer [4245]. Affixing [1842]. Affluence [12697]. Africa [1057, 3368, 7237, 8183, 10249, 11457, 13691]. African [2796, 5442, 8349, 9942, 11566]. After [6550, 7356, 10318]. Again [4743, 4893, 6515, 5601, 5708, 7719, 7810, 8623, 9333, 9337]. Against [18, 3723, 10140, 3271, 3131, 12485]. Agassiz [2503, 2433, 5218, 4033, 5599, 5405, 5241, 5611, 5387]. Age [1645, 7155, 8955, 12845]. Aged [2100]. Agency [1997]. Agent [816, 725, 1975, 7580]. Agents [3856]. Ago [1428, 12850]. Agricultural [112, 463, 126, 2415, 1374, 2182, 2627, 3646, 4302, 3961, 6308, 7046, 7703, 7735, 7768, 7858, 7916, 8522, 8661, 7944, 7972, 8851, 8881, 8941, 8971, 9003, 9035, 9065, 9101, 9125, 9159, 9191, 9257, 9292, 9324, 9352, 9382, 9415, 9482, 9518, 8002, 8034, 8058, 8089, 8117, 8144, 8171, 8204, 8230, 8290, 8312, 8373, 8406, 8435, 8464, 8496, 8528, 8580, 8610, 8637, 8665, 8753, 8782, 8817, 10598, 9690, 9570, 9597, 10426, 10455, 10487, 10518, 10551, 10585, 10646, 10673, 10703, 10785, 10812, 10840, 9624, 9655, 9683]. Agricultural [9711, 9788, 9814, 9842, 9868, 9895, 9928, 9946, 9998, 10021, 10050, 10101, 10160, 10183, 10214, 10238, 10266, 10296, 10328, 10363, 10395, 12049, 12631, 12710, 13449, 13463, 13520, 13562, 13622, 13672, 13697, 12585, 12657, 12777, 12804, 12854, 12920, 12962, 13026, 13110, 13182, 13204, 13313, 13347, 13414]. Agriculture [703, 297, 1037]. Agua [13319]. Ahead [11257]. Aid [1093, 1643, 5921, 12736, 8774]. Aids [3162]. Ailanthus [907]. Ailment [10902]. Ailment-Malarial [10902]. Air [524, 774, 533, 269, 124, 1225, 1461, 1766, 2541, 2867, 3658, 3034, 2904, 2694, 2518, 2565, 3009, 2529, 4726, 4094, 4092, 4459, 6247, 5878, 5929, 6335, 6504, 5938, 7333, 6971, 8866, 9181, 8905, 8480, 8626, 9474, 8101, 10248, 10912, 10041, 10837, 10228, 11858, 11619, 11082, 12420, 12766, 13693, 12387, 773, 273]. Alabama [2182]. Alarm [938, 384, 1117, 2059, 3383, 3277, 2640, 2708, 2459, 4639, 7817, 9371, 10778, 10728, 9530, 11807, 11863, 13177, 9797]. Alarms [606, 8368, 8287]. Alaska [1444, 386, 3413, 4260, 6215]. Albany [3335, 3297, 3013, 3044, 5493]. Albert [2590, 5246]. Albinism [5195]. Albion [5732]. Albumen [142]. Alchemists [2923]. Alcohol [4525, 6270, 6424, 6620, 9273]. Aleurometer [10450]. Alexander [720]. Alexandra [7184, 9965]. Alexis [2351]. Algerian [11352]. Alhambra [12051, 12537]. Alhydric [10063]. Alignments [6740]. Alkaline [8198]. Alkaloid [582]. Allan [5411, 10722]. Allanthus [12719]. Alleged [9960, 10747, 11615]. Alleges [12668]. Alleghany [570]. Allegheny [2492]. Allen [1494, 8560, 10535]. Alley [7340]. Allies [503, 1636, 785, 875, 1591]. Allison [10594]. Alloy [1228, 3507]. Alloys [348, 1357, 2985, 2519, 2500, 5323, 10342, 10876, 12310]. Almanac [8138]. Almond [8716]. Almost [1034]. Alocasias [7965]. Aloe [9059]. Aloft [11345]. Already [2503]. Altheide [8626]. Alum [12203, 13092]. Alumina [5278, 5304]. Aluminum [3523, 12653]. Am [3192]. Amalgam [6649, 12249]. Amalgamating [8958]. Amalgamation [2695, 11750]. Amalgamator [10926, 12066, 12795, 13466]. Amalgams [6965]. Amanians [3196]. Amateur [7577, 6816, 10718, 12255, 12231, 12389, 12468, 12588, 12695, 12912, 13018, 13077, 13130, 13196, 3786]. Amazon [1394, 5023, 5049, 5281, 5309, 5335]. Amazons [4935, 4964, 5023, 5049, 7029, 7028]. Ambition [10034]. Ambulance [9783]. Amend [454]. Amended [696]. Amending [10900]. Amendment [6545, 12282]. Amendments [6714, 7605, 10904]. America [854, 2351, 2473, 3177, 2553, 3529, 2563, 5085, 5514, 5710, 9137, 8952, 10573, 12068, 13319, 13134, 13379]. American [422, 1494, 6110, 6144, 9061, 13337, 821, 835, 75, 9, 706, 17, 881, 467, 16, 723, 1013, 1030, 1056, 840, 751, 219, 908, 928, 951, 970, 995, 1015, 176, 665, 748, 305, 36, 59, 653, 682, 709, 736, 779, 797, 823, 849, 866, 887, 912, 933, 953, 974, 998, 1018, 1039, 1060, 1084, 1109, 1131, 1153, 1178, 81, 105, 132, 164, 181, 208, 234, 259, 284, 304, 329, 355, 378, 407, 431, 459, 476, 497, 519, 541, 565, 588, 609, 629, 150, 127, 21]. American [583, 1441, 2270, 1996, 1275, 2072, 1192, 2215, 2184, 2286, 1324, 2104, 2133, 2154, 2174, 2201, 2226, 2251, 2271, 2373, 1248, 1204, 1233, 1883, 1908, 1926, 1951, 1980, 2004, 2026, 2050, 2075, 2093, 2116, 2137, 2158, 2180, 2231, 2255, 2281, 2304, 2330, 2355, 2376, 2404, 1257, 1282, 1310, 1337, 1364, 1386, 1420, 1448, 1472, 1501, 1525, 1558, 1581, 1605, 1628, 1652, 1677, 1704, 1730, 1758, 1783, 1811, 1837, 1861, 2113, 3243, 3370, 3402, 2719, 2699, 2520, 2672, 3303, 3448, 3022, 3778, 2502, 3507, 3427]. American [3457, 3480, 3512, 3679, 3369, 3708, 3308, 3660, 2452, 2478, 3073, 3098, 3118, 3143, 3169, 3200, 3228, 3252, 3287, 3317, 3345, 3374, 3405, 3433, 3460, 3487, 3514, 3549, 3575, 3611, 3635, 3666, 3695, 3730, 3762, 3794, 2508, 2534, 2557, 2579, 2607, 2633, 2659, 2678, 2701, 2728, 2750, 2776, 2802, 2828, 2852, 2872, 2893, 2911, 2930, 2955, 2979, 3000, 3027, 3049, 3706, 4809, 4835, 4862, 5069, 5095, 4425, 4492, 4946, 5067, 4269, 4837, 4797, 4810, 4907, 4930, 4958, 3820, 3845, 4554, 4582, 4617, 4647]. American [4681, 4707, 4733, 4759, 4782, 4813, 4838, 4865, 4894, 4916, 4943, 4968, 5000, 5024, 5051, 5074, 5097, 5125, 5149, 5180, 5203, 5232, 3873, 3904, 3932, 3962, 3990, 4019, 4048, 4082, 4107, 4135, 4162, 4190, 4217, 4248, 4274, 4305, 4330, 4359, 4387, 4412, 4441, 4474, 4503, 4532, 4549, 4257, 4717, 4598, 4893, 6214, 5848, 6494, 6273, 6209, 5768, 5540, 6217, 6281, 6312, 6392, 6457, 5641, 6169, 5257, 5284, 5966, 5990, 6016, 6041, 6070, 6096, 6121, 6147, 6174, 6200, 6232, 6259, 6284, 6315, 6341, 6366]. American [6393, 6414, 6437, 6458, 6482, 6506, 6534, 6565, 5313, 5337, 5365, 5395, 5426, 5453, 5476, 5503, 5529, 5559, 5588, 5621, 5643, 5672, 5697, 5724, 5753, 5777, 5805, 5833, 5859, 5886, 5911, 5940, 7441, 7470, 7751, 6841, 7911, 7556, 7591, 7607, 7635, 7662, 7783, 7292, 6739, 7659, 6588, 6612, 7247, 7273, 7293, 7317, 7348, 7372, 7395, 7420, 7443, 7472, 7495, 7518, 7543, 7568, 7593, 7623, 7649, 7676, 7677, 7702, 7734, 7766, 7795, 7828, 7856, 7886, 6633, 7914, 6661, 6685, 6706, 6729, 6758, 6783, 6808]. American [6832, 6857, 6883, 6900, 6926, 6955, 6984, 7009, 7036, 7059, 7086, 7118, 7142, 7167, 7197, 7225, 7454, 6966, 9045, 9112, 8561, 8588, 8537, 9093, 8216, 9154, 8843, 9255, 7941, 7969, 8694, 8721, 8751, 8781, 8814, 8847, 8879, 8910, 8938, 8969, 9001, 9030, 9062, 9098, 9124, 9157, 9187, 9221, 9256, 9288, 9321, 9350, 9381, 9412, 9444, 9481, 7999, 9514, 8030, 8056, 8087, 8113, 8142, 8170, 8201, 8229, 8259, 8289, 8311, 8338, 8370, 8403, 8433, 8463, 8493, 8523, 8547, 8575, 8606, 8634, 8663, 9236, 10580]. American [9907, 10062, 10526, 9882, 10613, 10958, 10310, 10769, 10349, 9991, 10529, 10308, 9981, 10280, 9542, 9567, 10265, 10292, 10327, 10359, 10391, 10421, 10452, 10484, 10514, 10546, 10581, 10617, 10645, 10671, 10700, 10731, 10756, 10782, 10809, 10839, 9596, 9623, 9654, 9680, 9708, 9734, 9758, 9786, 9838, 9865, 9894, 9924, 9943, 9972, 9994, 10016, 10046, 10070, 10099, 10128, 10155, 10179, 10209, 10237, 11912, 12049, 11462, 11905, 12201, 11974, 11063, 12033, 11826, 12213, 11185, 11550, 11907, 11998, 11540, 11922, 12070, 13197, 12697, 13327, 12305, 13456, 12462, 12483]. American [12514, 12574, 12609, 12787, 12818, 12851, 12873, 12908, 13485, 13519, 13529, 13590, 13611, 13643, 13671, 13699, 13225, 13119, 13581, 13307, 12460, 13275, 13299, 12304, 12669, 12816, 13163, 12734, 12640, 12617, 12491, 12614, 12646, 12680, 12746, 12767, 12837, 12887, 12944, 12966, 12996, 13043, 13071, 13101, 13126, 12677, 13448, 13506, 12501, 13083, 864, 141, 806, 291, 1959, 12419, 12333, 12355, 12384, 12432, 680, 1099, 122, 1121, 1150, 1176, 2419, 1201, 1229, 3131, 4588, 4624, 5568, 6376, 7845, 7876, 8327, 7958, 8271, 8382, 8414, 8472, 8536, 10530, 10845, 10895, 10947]. American [11497, 11663, 11776, 12280, 10999, 11022, 11051, 11384, 11411, 11441, 11466, 12263, 12453, 12478, 12507, 13630, 12543, 13690]. Americans [754, 3645, 7717, 7576, 11881]. Ames [1252]. Ammonia [774, 2109, 5608, 5635, 11858]. Ammoniumnitrodiphenylamin [9400]. Among [5195, 5006, 13618, 471, 296, 662, 5738, 11880]. Amsterdam [809, 2859]. Anæsthetic [12880]. Anæsthetics [8863, 13360]. Analysis [842, 4698, 5091, 5348, 6914, 7929]. Analyzed [9665]. Anatifa [6976]. Anatomical [12618]. Anatomy [11598]. Anchor [70, 2037, 8842, 13054]. Anchorage [7952]. Ancient [596, 128, 113, 1395, 1725, 1877, 3645, 4991, 6719, 7758, 8862, 8323, 8764, 10751, 10573, 11955, 12656, 13651, 13084, 12539]. Anderson [743, 2141, 10475]. Andes [5281, 5309, 5335]. Andrews [11531]. Aneroid [969]. Angamar [13661]. Angle [6385, 10316]. Angles [2564]. Anglo [422]. Aniline [4568, 10290, 10656, 12765, 190]. Animal [726, 661, 3455, 3247, 3892, 4960, 5878, 5929, 6822, 8366, 8773, 8450, 8330, 10714, 12288, 11888, 2616]. Animalism [7244]. Animals [471, 662, 5230, 6211, 6029, 6426, 7554, 9236, 9578, 9637, 6748, 251]. Animation [7580]. Annals [8244]. Annam [11948]. Annexation [4065]. Annexe [8920]. Annihilated [6909]. Annotta [9379]. Annual [1382, 2182, 2566, 3178, 2569, 3999, 1532, 9728, 11132]. Anode [11887]. Anolis [12501]. Anomalies [10596]. Anstie [6620]. Answer [2177, 1782, 1810, 2451, 3097, 3199, 3431, 3547, 3574, 2775, 4895, 4917, 5530, 5754]. Answered [2824, 4455]. Answers [35, 58, 681, 708, 735, 755, 778, 796, 822, 848, 865, 886, 911, 932, 952, 973, 997, 1038, 1058, 1083, 1108, 1130, 1152, 1181, 80, 104, 133, 163, 209, 235, 258, 283, 303, 328, 354, 377, 406, 430, 458, 477, 496, 518, 540, 564, 587, 611, 628, 652, 1203, 1231, 1835, 1860, 1881, 1905, 1924, 1949, 1978, 2002, 2024, 2048, 2073, 2091, 2134, 2156, 2206, 2229, 2253, 2279, 2301, 2328, 2353, 2402, 2427, 1256, 1281, 1308, 1335, 1362, 1384, 1415, 1446, 1471, 1499, 1523, 1556, 1580]. Answers [1603, 1627, 1651, 1676, 1703, 1728, 1757, 1751, 2477, 3071, 3117, 3142, 3168, 3227, 3251, 3286, 3316, 3344, 3373, 3404, 3459, 3485, 3513, 3610, 3634, 3664, 3694, 3729, 2507, 3760, 3793, 2533, 2556, 2578, 2606, 2632, 2658, 2677, 2700, 2727, 2749, 2801, 2827, 2851, 2871, 2892, 2910, 2929, 2954, 2978, 2999, 3026, 3818, 3846, 4504, 4534, 4555, 4583, 4619, 4649, 4683, 4709, 4734, 4761, 4783, 4814, 4839, 4866, 4944, 4969, 5001, 5025, 5052, 5075, 5098, 5126, 5150, 5181, 3874, 5204, 5233, 3905, 3935, 3963, 3991, 4020]. Answers [4049, 4083, 4108, 4136, 4163, 4191, 4219, 4250, 4276, 4307, 4332, 4360, 4388, 4414, 4443, 4476, 5258, 5285, 5967, 5991, 6017, 6042, 6071, 6097, 6122, 6148, 6175, 6201, 6233, 6260, 6285, 6316, 6342, 6367, 6394, 6415, 6438, 6459, 6483, 6507, 6535, 6566, 5314, 5338, 5366, 5396, 5427, 5454, 5477, 5504, 5560, 5589, 5622, 5644, 5673, 5698, 5725, 5778, 5806, 5834, 5860, 5912, 5941, 6589]. Ant [7439, 8779, 11537, 11983, 4855]. Ant-Eater [7439]. Antediluvians [11595]. Antelope [10418, 12916, 13259]. Antelopes [7790]. Anthon [5093]. Anthony [607]. Anthracene [8355]. Anti [1243, 2057, 3835, 3805, 3986, 6999, 8357, 13531, 12800]. Anti-Freezing [3986]. Anti-Friction [2057, 12800]. Anti-Nuisance [8357]. Anti-Primer [6999]. Anti-Sea-Sick [3835, 3805]. Anti-Snoring [1243]. Anti-Vaccination [13531]. Antimony [2498]. Antique [9437]. Antiques [6081]. Antiquities [10153]. Antiquity [2549, 3593, 7068]. Antiseptic [11163, 1025]. Antlered [6089]. Ants [5686, 6800, 6715, 6870, 12623]. Anvil [5616, 8806, 12286]. Any [273, 3582, 6873]. Anybody [8984]. Apartments [273]. Apatite [1380]. Apid [6524]. Apothecaries [3944]. Apparatus [1120, 388, 158, 1140, 980, 783, 149, 146, 943, 1742, 1712, 2192, 2338, 2097, 2382, 1780, 2237, 3623, 3621, 3650, 3235, 3331, 2495, 2843, 4858, 4976, 5174, 4123, 5199, 4485, 4567, 4852, 4215, 5113, 6255, 5982, 5420, 5567, 6327, 6973, 7049, 7824, 7822, 7575, 7641, 7533, 7186, 7765, 8739, 9178, 8104, 9309, 8746, 8368, 9318, 8958, 9433, 7990, 9083, 8283, 8192, 8993, 8041, 8072, 8690, 8106, 9503, 8592, 9095, 8992, 9508, 9307, 10346, 10539, 10662, 10165, 10511, 9679, 10254, 10723, 9914, 10288, 11715, 11531, 11190, 11954, 11214]. Apparatus [11149, 11097, 11274, 11733, 11885, 13698, 12591, 12533, 12712, 13105, 13537, 13650, 12396, 9471, 11958, 12366, 12665, 5305, 273, 11764]. Apparent [8508]. Apparitions [7329]. Appeal [201, 705, 2080]. Appeared [2312]. Apple [7440]. Apples [3225, 6033, 6189]. Appliances [48, 1023, 2165, 3364, 4401, 10614, 11142, 12050, 11374]. Applicable [2641]. Applicants [1832, 1900]. Application [832, 1104, 56, 2415, 1890, 1914, 1970, 3781, 2845, 4698, 6324, 5379, 1736, 1767, 1791, 1843, 1869, 1166, 3644, 921]. Applications [61, 82, 1106, 1180, 134, 162, 183, 211, 261, 302, 379, 475, 521, 539, 590, 631, 655, 684, 707, 734, 777, 798, 825, 910, 955, 976, 996, 1059, 1232, 1283, 2136, 2178, 2232, 2257, 2280, 2303, 2332, 2357, 2377, 2401, 2426, 1309, 1387, 1418, 1500, 1526, 1557, 1583, 1606, 1653, 1680, 1706, 1729, 1760, 1784, 1836, 1862, 1885, 1906, 1928, 1953, 1982, 2028, 2051, 2118, 1867, 1920, 2454, 2476, 2506, 3019, 8651, 8162, 8733, 8657, 10604, 11992, 11342, 12701]. Applied [2696, 4234, 4030, 7904, 6653, 9823, 13659, 2493]. Applying [8746]. Appointment [3340, 4399]. Appreciated [730]. Appreciation [3215, 4810].

Apprentice [2200]. Apprenticeship [2129]. Approach [11310]. Approaching [1996]. Appropriation [8416]. Aquaria [7990]. Aquarium [4531, 6027, 5856, 6390, 6657, 7394, 9477, 8967, 9305, 9535, 11232, 12472, 5176, 12341]. Aqueduct [8303]. Arabesque [13497]. Arabia [4157]. Arabian [2923]. Aranjuez [10095, 10231]. Arbitration [12522]. Arbor [1820]. Arbors [690]. Arch [2816, 8743]. Archeological [6035]. Archer [10780]. Arches [11092]. Archibald [1512]. Arching [2623]. Architect [4798]. Architects [6245]. Architectural [2323, 8331]. Architecture [945, 1931, 7393, 8736, 9770]. Arctic [7433, 9510, 9304, 11912, 11664, 12359, 1797, 4660, 4424, 4231, 4570, 4878, 7162, 7070, 7308, 7281, 9406, 9825, 9582, 11159]. Argand [11558]. Argentine [561]. Argonaut [12234]. Arguments [11447]. Aristocracy [8156]. Arithmetical [6632]. Arizona [3281, 9410, 13013]. Ark [11716]. Armor [5790, 6775, 11059]. Armored [3130, 3163]. Armory [2706]. Arms [2287, 3397, 4452]. Army [5686, 5654, 9435, 3975, 6476]. Arrangement [6161, 7107, 10837]. Arrest [2015]. Arrested [10767]. Arrester [9648]. Arrow [5903, 13496]. Arsenic [3977, 3887, 6308, 13254, 13438, 2714, 2740, 2763]. Art [1007, 960, 505, 552, 1493, 2719, 2654, 2898, 2946, 2770, 4955, 3834, 3866, 3890, 3926, 3972, 4776, 4345, 5738, 6491, 8212, 9608, 11583, 12736, 1987, 12791]. Artesian [1145, 2889, 5604, 7311]. Article [2146]. Articles [1454, 1936, 2295, 4397, 7705, 7737, 7767, 7799, 7831, 7857, 7888, 7915, 7943, 7971, 8003, 8060, 8091, 8115, 8233, 8261, 8315, 8374, 8407, 8436, 8495, 8524, 8549, 8579, 8608, 8668, 9781]. Articulated [4781]. Artificial [494, 991, 700, 65, 946, 1531, 1974, 1873, 1574, 1592, 2274, 2203, 2411, 1711, 2501, 2470, 2442, 3309, 3951, 5082, 4956, 4966, 5160, 5450, 7729, 7485, 9368, 8478, 8954, 9438, 10324, 10207, 12159, 11582, 11803, 11487, 11762, 13613, 12832, 12486, 12594, 45, 8359, 8275]. Artificially [10981]. Artillery [919, 990, 6090, 12201]. Artisan [873, 895, 784, 831, 7876, 9496, 12263]. Artisans [8563, 11880, 12669]. Artist [9496]. Artists [12772]. Arts [1046, 2626, 2621, 2770, 2465, 2689, 2789, 2884, 3984, 8383, 9721, 11927, 11526]. Asa [12949]. Asbestos [2617, 8420]. Ascension [7137]. Ascent [5987, 5857, 13319]. Aschenbrenner [5155]. Ash [1351, 11530, 10449, 5245]. Ash-Colored [10449]. Ashantees [5738]. Ashtabula [9852]. Asia [13120]. Asiatic [11690]. Asparagus [9676]. Aspect [5214]. Asphalt [176, 2045]. Asphalte [508, 2471]. Assay [1532]. Assignments [4373]. Assimilated [726]. Associated [3105]. Association [821, 835, 812, 939, 959, 2149, 3243, 3370, 3402, 2502, 3504, 4809, 4835, 4862, 6110, 6144, 6257, 6311, 7441, 7470, 7813, 9093, 9061, 13197, 13581, 13307, 12856, 13359, 2182, 3368, 5013, 5039, 4997, 5380, 6169, 6184, 7566, 10571, 10606, 10615, 11905, 11926, 13327, 13608, 13337]. Associations [471]. Assyria [12974]. Astonishing [11048]. Astor [5898]. Astounding [11211]. Astronomers [7996]. Astronomical [3676, 3688, 3341, 4200, 3956, 4074, 4181, 4328, 4436, 4544, 5050, 5146, 5632, 5253, 5386, 5495, 5585, 5618, 5690, 5801, 5933, 6386, 6475, 5392, 6584, 6705, 6827, 6920, 7034, 7131, 7265, 7698, 7826, 7178, 7960, 8111, 8251, 8367, 8489, 8627, 8749, 9287, 9443, 9539, 9653, 9777, 9891, 9992, 10091, 10235, 10381, 10472, 10644, 10755, 10891, 10436, 10994, 11295, 13338, 12132, 11086, 11124, 11846, 11866, 11891, 11917, 11941, 11962, 11979, 12010, 12029, 12077, 12098, 12212, 12237, 12266, 12290, 11176, 11200, 11223, 11247, 11263, 11288, 11331, 11348, 11382, 11406, 11427, 11456, 11517, 11528]. Astronomical [11568, 11588, 11638, 11669, 11695, 11713, 11739, 11765, 11789, 11818]. Astronomy [2888, 4746, 5084, 4030]. Aswail [13132]. Asylum [2100, 1941, 4242]. Asymmetry [1426]. Atchison [8454]. Athens [2211]. Atkinson [8396]. Atlantic [3007, 3500, 2469, 4575, 4230, 5824, 8133, 13637, 12878]. Atlantis [10307, 11837]. Atmosphere [8804, 8982, 10852]. Atmospheres [7260]. Atmospheric [2926, 3710, 2721, 5225, 7621, 12420, 11691]. Atomic [2870, 2385]. Atomizer [5274]. Atoms [843, 8649]. Attaching [1936, 1542, 4078]. Attachment [2920, 4394, 4804, 4915, 4131, 4519, 4729, 4129, 5044, 4676, 5638, 5998, 5545, 7643, 7076, 6607, 9058, 9235, 8253, 9371, 9920, 11270, 12179, 12943, 13435, 12729, 12705, 13412, 13106, 11783]. Attack [511, 8648]. Attain [5871]. Attainment [178]. Attempt [1079, 201]. Attempts [176, 3494]. Attend [1052]. Attorney [6587]. Attraction [13543, 2491]. Auction [3366, 10375]. Auger [2265, 5419]. Augers [7434]. Augsburg [5240]. August [3178, 3283, 4695]. Auguste [5324]. Auk [340]. Aurora [3566, 3245, 3836, 7535, 1216]. Auroras [3682]. Australia [526, 1957]. Australian [4099, 5421, 8211, 10230, 11759, 12074, 12761, 13012, 12708]. Austria [873, 3448, 3883, 3861, 4717, 12210]. Austrian [4661, 3852]. Autographic [10094]. Automata [6211]. Automatic [1169, 1043, 1294, 3623, 3183, 3266, 2487, 3124, 4858, 4014, 4645, 5017, 4913, 4450, 5221, 4757, 4873, 4466, 5249, 6300, 6373, 6428, 5721, 5552, 6596, 6852, 9054, 9116, 8159, 8135, 8471, 8741, 8510, 8744, 8187, 8365, 10918, 9524, 10444, 10502, 10695, 10712, 10575, 10797, 11199, 11485, 11796, 11710, 11587, 11915, 13589, 12683, 1190, 1028, 11702]. Automatically [10746]. Automaton [5222, 7750]. Autophone [12929]. Autopsy [5405, 5486]. Auxiliary [4792, 7786]. Aveling [12470]. Avenue [453, 3775]. Average [1381]. Avoid [603]. Awake [2012]. Awakening [1489]. Award [207, 4793]. Awarded [17]. Awards [4797, 12091, 9173, 9303, 9240]. Awl [1662, 7339]. Awning [6780, 12422]. Axe [10118]. Axes [3296]. Axle [2339, 2290, 6828, 9617]. Axled [4781]. Axles [6240, 7085, 8739, 11344]. Axletrees [1764]. Azaleas [7392]. Azores [13104].

B [1143, 2823, 1119]. Babbage [2412]. Babcock [7643]. Baby [5165, 6574, 10013]. Back [2648, 4717]. Backs [3103]. Backus [8486]. Bacon [1034]. Bacteria [6796]. Bad [8348, 8795, 8619, 9810, 13429]. Badly [3742]. Baer [1919]. Baffle [5362]. Bag [2794, 10227, 11928]. Bag-Worm [11928].

Baggage [12744]. Bailey [11190]. Baillairge [2962]. Bain [2166]. Baker [365, 6721, 6608, 10636, 11432]. Baking [12203, 13092]. Balance [275, 246, 3011, 3948, 5355, 5639, 10610, 11855, 13247, 12742, 12741]. Balanced [1426, 1619, 5920, 8630]. Balancing [950, 3858, 3917, 10343]. Baldwin [2546, 4269, 7266]. Bale [4294, 5308, 6803, 7411, 9242, 9314, 8832, 11347]. Bales [13277]. Balestrieri [8691]. Ball [2057, 3868, 9212, 10808, 10261, 10574, 12159, 11592]. Ballard [412]. Ballasting [7838, 10204]. Balloon [1093, 2736, 4820, 4823, 4600, 4616, 4634, 4879, 4979, 4844, 6168, 5987, 6504, 6061, 5857, 7137, 10915, 11954, 11865, 11929, 13093, 550]. Balloon-Apparatus [11954]. Ballooning [7072]. Balloons [927, 4646, 6672, 9463]. Baltimore [2810, 4629, 10006, 13429, 12732]. Bamboo [8016]. Banana [7687]. Bananas [30]. Band [874, 1081, 3941, 4115, 5553, 6778, 8762, 8050, 8081, 10806, 11409, 12799, 12302]. Band-Sawing [10806]. Banded [10941]. Bands [6803]. Baneful [13417]. Banian [10194]. Banish [12045]. Bank [23, 8185, 10662, 12146, 12117]. Bankrupt [1994]. Banks [12284]. Baobab [8384, 11405]. Baptismal [13472]. Bar [274, 1049, 4282, 4429, 5668, 6059, 6726, 7043, 7964, 8743, 10939]. Barbarians [9782]. Barber [8192, 4123, 13445]. Barff [12792]. Bark [11939]. Barker [5420]. Barn [8682]. Barnacles [389]. Barnes [1488]. Barnum [5550]. Barometer [3180, 4236]. Barometers [969, 6186]. Barometric [8766, 11177]. Baromotor [10691]. Baron [7468, 2275, 4678, 4292, 9500, 2060]. Barracouta [118]. Barrel [5391, 6692, 7190, 7994, 8491, 8513, 7985, 12046, 12473, 12774, 13510, 13570]. Barrels [3300]. Barrow [6412, 548, 5280]. Barrow-in-Furness [6412]. Bars [2413, 3561, 7241, 8512, 8567, 10838]. Bartholdi [11886]. Bartholomew [2467, 5114]. Bartlett [659, 1612, 8869]. Bas [10943]. Base [4746, 10808]. Basilica [9619]. Basin [11093, 11454]. Basins [395, 11371]. Basis [417]. Basket [8965, 12095]. Bass [1101]. Bastard [11516]. Bastie [7210, 7214]. Bat [10386]. Batchelder [12517]. Bath [422, 4641, 5449, 6583, 8025, 10794, 11265]. Bathometer [1702, 9834]. Baths [1670, 4927, 11044, 13555]. Batteries [5411, 6360, 7107, 6670, 7413, 3390]. Battering [5434]. Battery [968, 88, 1991, 2744, 4579, 5956, 7057, 7765, 8838, 10828, 11030]. Battle [22, 1576]. Bauer [12003]. Baugh [11169]. Bavaria [11826]. Bawlkin [12595]. Baxter [2191, 5786, 11091]. Bay [6912]. be [75, 726, 695, 1852, 1353, 2012, 3494, 2594, 3851, 3827, 8181, 11476, 11108, 11969, 12718]. Beach [2743, 2769]. Beacon [6214]. Beads [5897]. Beam [2037, 10857]. Beams [401, 4643]. Beans [11302]. Bear [6356, 11537]. Beard [11122]. Bearded [8646, 12007]. Bearer [3561]. Bearing [2984, 5448, 7340, 8350, 11434, 12764, 12800]. Beast [483, 11294]. Beat [6644]. Beater [11144]. Beaudet [10643]. Beautiful [7207, 8935, 8520, 10125, 10992, 11188]. Beauty [178]. Bebout [2292]. Because [3192]. Beck [545]. Beckwith [2691, 6970]. Become [1994, 7876, 12263]. becomes [1538]. Becquerel [11570]. Bed [2242, 3301, 2868, 10420, 11511, 13017]. Bedell [8284, 9057]. Bedstead [193, 10286]. Bedsteads [9920]. Bee [598, 1037, 3330, 4527, 8100, 763]. Bee-Eating [598]. Beef [9015, 13305]. Beehive [3807, 4160, 4543, 6003, 11692, 12055]. Been [3282, 7747, 1148]. Beer [7467, 10538]. Bees [8908, 12983]. Beet [538, 1100, 698, 2693, 2850]. Beet-Root [1100, 698]. Beetle [2944, 4572, 11984, 12649, 486, 10609]. Beetles [2964, 875, 1772, 1871, 10867, 10911]. Beetling [12532]. Beets [12933]. Before [939, 2312, 2168, 454, 959, 21, 79, 103, 131, 474, 2774, 2676, 2821, 3852, 3886, 6184, 8132]. Begonia [6881]. Begun [4660]. Behavior [5106, 9215]. Belgian [4462, 5215, 6722, 12221, 13271]. Belgium [784, 12393]. Bell [871, 360, 2261, 2968, 4665, 5932, 5667, 7691, 10728, 10625, 9530]. Bells [871, 360, 2985, 8364]. Below [10004]. Belt [1917, 1483, 1554, 1691, 6060, 7464, 6752, 7133]. Belt-Geared [7133]. Belting [2864, 3510, 4422]. Belts [1917, 2490, 4227, 4262, 4497, 10031, 13632, 12993]. Bench [5330, 12233, 11282, 12317]. Bender [8988]. Bending [2757, 7465, 8802, 10316]. Benefit [7955]. Benevolence [3220]. Bengal [5683, 9461]. Bennett [3024]. Benoît [11696]. Bent [204, 3870]. Benton [8518]. Benzine [2653]. Benzole [2653]. Bergen [6765, 10059]. Berlin [4079, 12472]. Bernard [11479]. Berried [8519]. Berry [12302]. Berryman [3061, 11056]. Bert [12880]. Berth [13536]. Berthelot [8921]. Bessemer [367, 723, 359, 2163, 1400, 3805, 7130, 6651, 6681, 13246, 13356]. Best [4218, 4249, 4275, 4306, 4331, 4478, 4170, 6743, 10956, 11284, 11258, 12520]. Beton [11325, 10616]. Better [125, 5157, 7018, 12790, 12716]. Between [5514, 6007, 7177, 8476, 9911, 11989, 1047, 1092, 2653, 2592, 6503, 7360, 12669, 12817, 2385]. Bevel [7435, 9372, 10776]. Beverage [11369]. Bias [9968]. Bickford [1735]. Bicycle [11473, 11717, 12762, 12365]. Biddle [6453]. Big [2774]. Bigelow [13688]. Bigots [2395]. Bilge [2459, 7412]. Bill [454, 7428, 7164, 9827, 11925, 12385, 12607, 12484, 12603, 12362, 12552, 8215, 12458]. Billed [10914]. Billiard [9118, 10260, 10261, 10504]. Bills [8099]. Binary [3597, 3393, 3476]. Binder [3263, 4394, 8165, 12910]. Binders [12557]. Binding [9502, 11885, 11930]. Bird [5584, 6980, 8350, 12031, 11814, 12008, 11911, 12106]. Birds [3558, 5195, 4730, 7166, 6727, 9347, 8997, 10611, 11620, 1510]. Birmingham [2899]. Birthday [5165]. Birthwort [10968, 11404]. Biscuit [2902]. Bisulphide [2625, 4453, 4792, 5487]. Bit [668, 3129, 2920, 8323, 12390]. Bitricycle [366]. Bits [4429, 7488, 7967]. Bixa [9379]. Black [1101, 1897, 5122, 3957, 7696, 7589, 9210, 11220, 11496, 8425, 156, 3195, 2756]. Blackboard [11031]. Blackboards [1048]. Blacking [2919, 4521, 4073]. Blacksmith [3239, 5121]. Blackstone [6116]. Blade [1790]. Blades [11409]. Blaisdell [6656]. Blake [8457, 13534, 13486]. Blank [3294]. Blast [269, 3722, 3388, 4268, 3839, 4510, 4853, 5381, 9172, 8710, 8867, 11646, 13559, 921]. Blasting [2409, 2272, 4132, 12252, 12766]. Blatchford [79, 103, 131, 474].

Bleaching [725, 1999, 2055, 8106]. Blessings [9216]. Blight [5022]. Blind [2338, 7819, 7438, 7110, 7338, 9581]. Blinder [6477]. Blindness [9426, 10166, 11134, 11367]. Blister [10867, 10911]. Block [1661, 3720, 7816, 9149, 11453]. Blocks [3309, 7391, 9175, 9241]. Blood [1254, 4283, 6136, 9497, 9501, 11687, 12459, 11018]. Bloodless [5902]. Blooming [7261]. Blooms [10324]. Blotter [1617]. Blow [4000, 6411, 921]. Blow-Off [6411]. Blow-Pipe [921]. Blower [289, 3186, 3721, 4510, 6642, 10692, 11430, 13162]. Blowers [1998]. Blowing [1792, 2237, 4123]. Blowpipe [1643, 7263, 8161, 9439, 8191]. Blue [8445, 10166, 9722, 9748, 9772, 9879, 12181]. Bluffs [9369]. Boadmasters [13557]. Board [983, 1438, 4260, 10496, 10345, 203, 3801, 8274]. Boarding [2244, 2916]. Boardman [1482]. Boards [3023]. Boat [4, 3623, 3190, 2574, 3079, 3757, 4858, 4431, 4771, 4454, 3867, 4102, 8993, 10321, 11716, 12003, 11239, 11536, 13698, 13365, 12502, 13665, 13605]. Boat-Lowering [8993, 13698]. Boating [2567]. Boats [1622, 3473, 3127, 3332, 5044, 9268, 8459, 10931, 10930, 11529, 1166, 2807]. Bobac [12681]. Bock [2243]. Bodies [535, 1852, 7551]. Bodily [10717]. Body [1334, 6995, 11499]. Bogardus [5662]. Bogie [6254]. Bogus [6738]. Bohart [7467]. Boiled [10035]. Boiler [1169, 399, 1124, 938, 314, 860, 423, 469, 15, 1992, 2084, 1594, 2043, 1445, 3754, 2810, 3742, 3392, 3056, 2756, 3785, 3717, 3739, 4597, 5110, 3853, 4041, 4272, 4097, 4027, 5065, 4131, 4069, 3806, 4003, 5732, 6327, 5814, 6309, 6058, 5422, 6336, 6228, 5636, 5568, 5410, 7407, 7231, 7103, 7325, 7148, 7366, 7530, 7404, 7383, 9028, 9394, 8710, 8398, 8897, 10467, 10777, 9750, 9918, 10594, 9607, 11560, 11085, 11633, 11553, 10987, 11262, 11401, 12943, 12363, 12416, 13287, 12729, 12373, 12914, 830, 10]. Boiler-Feed [938]. Boilers [603, 87, 1117, 197, 67, 445, 15, 2293, 1573, 1345, 1544, 1621, 2059, 1648, 2042, 1515, 2017, 1196, 1315, 2012, 3399, 2734, 2517, 2858, 3607, 3414, 2598, 3338, 2592, 4597, 4061, 4438, 4801, 6001, 5959, 6086, 7469, 6893, 7047, 7579, 7787, 8249, 8352, 8395, 8196, 9967, 10117, 10670, 11370, 11298, 11989, 11659, 11585, 12705, 313, 1715, 485]. Boilers-How [603]. Boiling [7614, 10596, 11431]. Boland [11073]. Bolt [334, 981, 2032, 1638, 4470, 4703, 4339, 6277, 5681, 6723, 6948, 7002, 9308, 12289, 11070, 12654]. Bolting [12888]. Bolts [2756, 8160, 11267, 12886]. Bombay [1911, 3417]. Bon [11167]. Bone [2936, 7099, 8425]. Bones [6185]. Bonser [5901]. Bonus [845]. Book [1965, 2927, 5437, 6795, 9346, 10383, 10862, 12527]. Book-Making [6795]. Bookbinder [9136]. Bookbinding [10256]. Books [3, 41, 936, 978, 1001, 1022, 1064, 1088, 1112, 1134, 1156, 64, 85, 109, 185, 213, 239, 264, 288, 310, 333, 382, 410, 437, 481, 569, 612, 634, 656, 758, 802, 827, 852, 891, 915, 733, 2325, 1236, 1261, 2359, 2380, 1287, 1314, 1340, 1424, 1475, 1505, 1562, 1586, 1610, 1633, 1683, 1763, 1788, 1815, 1841, 1955, 2008, 2030, 2121, 2162, 2235, 2259, 2308, 2334, 1417, 1469, 2458, 2482, 3255, 3293, 3325, 3382, 3439, 3467, 3521, 3555, 3580, 3638, 3670, 3735, 2510, 2536]. Books [2562, 2584, 2613, 2639, 2683, 2729, 2752, 2781, 2803, 2832, 2855, 2876, 2915, 2935, 2957, 2982, 3078, 3122, 3149, 3175, 3205, 3234, 3822, 3876, 5078, 5184, 3908, 3937, 3994, 4023, 4052, 4087, 4111, 4165, 4221, 4253, 4310, 4335, 4363, 4417, 4446, 4537, 4558, 4653, 4712, 4738, 4786, 4869, 4898, 4972, 5237, 5263, 6205, 6289, 6346, 6419, 6463, 5290, 5343, 5370, 5400, 5431, 5481, 5508, 5533, 5593, 5625, 5677, 5702, 5758, 5782, 5810, 5838, 5864, 5889, 5916, 5945, 6021, 6075, 6101, 6126, 6173, 6568, 6638]. Books [6665, 6733, 6787, 6812, 6861, 6930, 6987, 7039, 7200, 7321, 7447, 7521, 7597, 7627, 7710, 7741, 7774, 7803, 7864, 7920, 7976, 9389, 8007, 8068, 8122, 8149, 8237, 8266, 8377, 8440, 8555, 8582, 8641, 8698, 8727, 8757, 8855, 8887, 8946, 9039, 9071, 9165, 9196, 9229, 9296, 9329, 9520, 9602, 10921, 10948, 9628, 9686, 9713, 9738, 9845, 9871, 9900, 9951, 10023, 10054, 10076, 10105, 10133, 10186, 10243, 10300, 10398, 10491, 10554, 10588, 10677, 10707, 10789, 10815, 11000, 11076, 12139, 12164, 12190, 12242, 11153, 11179, 11251, 11306]. Books [11334, 11386, 11413, 11443, 11468, 11491, 11575, 11602, 11629, 11676, 11744, 11768, 11792, 11820, 11851, 11875, 11898, 11945, 12038, 12084, 12350, 12378, 13374, 13395, 13421, 13476, 13501, 13549, 13575, 13600, 13626, 13655, 12428, 12510, 12546, 12570, 12691, 12783, 12841, 12868, 12899, 12952, 13002, 13031, 13060, 13087, 13114, 13139, 13166, 13192, 13215, 13242, 13267, 13294, 13322, 13350]. Boomer [1612]. Boot [5071, 4925, 5827, 8104, 10258, 10173, 10940, 12443]. Bootjack [2900, 6721]. Booton [1220]. Boots [4611, 6434, 9083]. Borax [10580, 10687]. Borden [5349, 10779]. Bore [3703, 10543, 10471]. Borealis [3566, 1216]. Borer [1269, 5434, 7109, 13383]. Boring [416, 668, 1802, 1568, 1238, 3617, 3969, 4931, 5198, 6410, 5423, 5688, 7488, 7692, 7839, 8167, 7931, 10832, 10802, 10936, 11783, 1772, 1871]. Born [9081]. Boston [12089, 983, 3633, 3979, 5108, 11994, 12032]. Botanical [12009]. Botany [2528]. Botch [9938]. Bots [9540]. Bottle [2765, 4071, 5579, 7455, 6943, 6701, 8397, 9919, 12291, 13512, 12750]. Bottled [12473]. Bottles [7363, 9310]. Bottom [3301, 5791, 2261]. Boulogne [11892]. Bound [9511]. Boundary [6839]. Bouquet [7113]. Bouquets [6118]. Bourne [1073]. Bournemouth [2335]. Bow [2081, 4669, 5723, 7853]. Bower [10248]. Bowls [395]. Box [1136, 271, 392, 1947, 1694, 2290, 2984, 4521, 4073, 5143, 4380, 4755, 4467, 6644, 6702, 7289, 6828, 8807, 9277, 8768, 10326, 9617, 9532, 11709, 11711, 11452, 12772, 12621, 10540]. Boxes [3300, 10524]. Boy [2316, 12980]. Boyden [369]. Boynton [1359, 4530]. Boys [2247, 12771]. Bozerian [11265]. Brace [111, 4609]. Braces [2920, 5471]. Brachiospongia [7185]. Bracing [1102]. Bracket [5120]. Bradley [4130]. Brain [4928, 13438]. Brainerd [2735]. Brains [7484]. Brake [1722, 2622, 3019, 3111, 4959, 4887, 4299, 5249, 5441, 5379, 7727, 8831, 9205, 8771, 8480, 8332, 8744]. Brakes [2706, 4060]. Bramble [9251]. Branch [12579]. Branching [65]. Brandon [3481]. Brass [501, 1379, 5986]. Brasses [9908]. Brassey [1465]. Brayton [8503, 9551]. Brazil [5145, 7166, 12033, 12147]. Bread [761, 8385]. Breadmakers [2113]. Breadstuff [2188]. Break [201]. Breakage [10857]. Breaker [13136]. Breaking [343]. Breakwater [722, 2411, 3470, 3211]. Breast [1097, 3389]. Breast-Strap [1097]. Brechbiel [7369]. Breech [26, 2079, 2009, 3397, 8957, 8998, 9018, 12006, 12774]. Breech-Loading [26, 8957, 8998, 9018, 12006, 12774]. Breed [6310]. Breeding [12311, 13618]. Breton [6806]. Brewer [13011]. Brewers [764, 1395]. Brewing [3749]. Brewster [4528]. Brick [2037, 2347, 3108, 2857, 2684, 4779, 3850, 6227, 10955, 9980, 10766, 11319, 11463, 11168, 12740, 13125, 2314, 2341]. Brick-Making [10955]. Bricks [370, 350, 335, 2623, 4666, 6143, 6385, 7280]. Bridal [4159]. Bridge [1047, 688, 1143, 626, 1410, 1566, 2735, 2756, 3013, 3276, 2983, 4106, 4807, 4375, 5073, 4384, 4403, 3929, 4015, 5246, 6466, 5361, 6007, 6542, 5329, 6961, 7177, 8476, 8714, 8444, 8981, 7987, 8247, 10912, 9911, 10510, 10112, 11815, 11197, 11515, 11569, 13483, 570, 922, 635, 836, 1792, 1563, 2311, 1547, 1809, 1771, 1674, 1408, 3769, 3195, 2687, 3092, 3524, 4326, 5493, 5346, 5745, 6489, 6006, 7807, 6737, 7410, 7953, 7952, 9636, 10371, 10338, 9909, 9767, 9743, 11471, 11310, 12087, 12855, 12670, 13641]. Bridges [159, 1960, 3443, 3270, 7631, 10539, 13582]. Bridging [4643]. Bridgman [2441]. Bridle [3129, 3891, 6477, 12390]. Bridles [7967]. Brief [11210]. Briesen [2338]. Brigham [13688]. Bright [3194]. Brighter [19]. Brighton [484, 5176, 12341]. Brine [13305]. Bristles [3103]. Bristol [2463, 6025]. Britain [2188, 6244]. Britannic [5988]. British [6257, 6311, 872, 1147, 939, 959, 1595, 3184, 3368, 5013, 5039, 4997, 6012, 6184, 7433, 7566, 6603, 6962, 7655, 6946, 9510, 9347, 9304, 8875, 9965, 10306, 10571, 10606, 10615, 11926, 11486, 12283, 13342]. Broadbook [6655]. Broadway [202, 244, 4259, 4287]. Broken [8457]. Bromine [13428]. Bronze [919, 171, 1830, 4100, 5111, 7454, 10962, 12342, 13391, 12321]. Brooklyn [1047, 635, 3769, 2783, 4065, 6609, 7177, 8476, 9911, 11325, 12817]. Brooks [8136]. Broom [2144]. Brother [9050]. Brotherhood [5569]. Brothers [2494, 10663, 12738, 2544]. Brown [5610, 10386, 12893, 1200, 3807, 9524, 10666]. Browne [10383]. Brownhill [11319]. Brownian [11554]. Brugmansia [10698]. Brush [3103, 2919, 4073, 6038, 7138, 8165, 11549, 11613, 5799]. Brusher [7243]. Brushes [10325]. Brussels [2943, 6781]. Brutal [1125]. Brutes [10013, 11232]. Bubble [3707, 13162]. Buckboard [11240]. Bucket [2216, 4705, 7135, 7025, 9182, 10042]. Buckeye [6596, 11485]. Buckle [11838]. Buddha [8049, 8080]. Buffalo [12715, 11552]. Buffer [1986, 2418]. Bug [5715, 7283, 11381]. Bugs [9283]. Build [346, 2574, 6940, 10933]. Builder [2951, 4798, 13595]. Builders [6245]. Building [370, 33, 226, 2346, 2220, 2246, 2268, 1267, 3085, 2907, 2686, 3020, 3528, 3584, 4693, 6298, 6406, 6874, 7706, 7797, 7830, 7889, 7918, 6907, 9110, 9441, 9175, 9241, 8693, 8162, 8717, 8001, 8032, 9066, 9100, 9128, 9160, 9188, 9223, 9260, 9291, 9322, 9353, 9414, 9447, 8114, 8145, 8173, 8202, 8232, 8262, 8313, 8371, 8405, 8434, 8465, 8526, 8550, 8578, 8609, 8635, 8666, 8723, 8816, 8849, 8883, 8912, 8970, 9032, 7987, 8590, 10931, 10570, 9569, 9658, 10548, 10586, 10647, 10674, 10701, 10758, 10786, 9684, 9709]. Building [9759, 9791, 9816, 9867, 9897, 9926, 9948, 9996, 10018, 10049, 10131, 10159, 10184, 10210, 10240, 10268, 10295, 10330, 10362, 10394, 10425, 10456, 10930, 11879, 11786, 11565, 12905, 13041, 8687]. Buildings [1122, 154, 76, 1855, 2166, 2298, 2442, 3724, 2652, 3357, 2575, 2624, 2548, 2642, 3745, 3499, 3851, 5047, 3801, 6435, 5567, 7479, 7541, 7024, 7004, 7778, 6781, 8543, 8591, 8541, 9576, 10887, 9853, 9747, 11015]. Built [75, 6108]. Bull [9957, 9466]. Bulletin [3505]. Bullets [6378, 7272]. Bullion [11139]. Bulls [6271]. Bumper [5823]. Bunching [3103]. Bung [3128]. Bunker [3014]. Bunkers [6161]. Bunsen [4613, 7765, 7640, 7880]. Buoy [8741]. Buoys [417, 4986, 11949]. Burbot [9284]. Burden [1857, 1306]. Burdict [2032]. Bureau [1439, 1695, 1848, 1868, 1891, 1933, 1958, 3950, 8353, 13173]. Burette [3015]. Burglar [7817, 10778, 11807]. Burglarproof [9370]. Burial [5739, 6246, 5684]. Burlington [3748]. Burmese [2309]. Burned [11030]. Burner [7022, 7640, 8711, 8333, 10753, 10087, 12157, 11558, 7880]. Burning [3606, 3971, 4524, 6306, 5473, 8685, 10670, 12576, 12655]. Burnishing [6198]. Burns [6910]. Burr [10859, 12613]. Burring [12238]. Bursting [5776]. Bushing [690, 7994]. Business [2, 40, 568, 591, 613, 633, 657, 686, 712, 739, 759, 782, 801, 828, 853, 869, 890, 916, 937, 957, 979, 1002, 1021, 1042, 1063, 1087, 1113, 1135, 63, 1158, 86, 108, 137, 167, 186, 214, 238, 263, 287, 309, 332, 358, 383, 411, 436, 462, 480, 500, 523, 544, 2184, 1184, 1207, 1866, 1888, 1910, 1930, 1956, 1984, 2007, 2031, 2054, 2078, 2096, 2120, 2140, 2161, 2185, 2210, 2236, 2260, 2284, 2307, 2352, 2360, 2381, 2407, 1237, 1260, 1286, 1313, 1341, 1367, 1390, 1423]. Business [1451, 1476, 1506, 1529, 1561, 1585, 1609, 1632, 1657, 1682, 1709, 1734, 1762, 1787, 1816, 1840, 1997, 2430, 2457, 3030, 3052, 3075, 3100, 3120, 3146, 3172, 3203, 3231, 3256, 3291, 3322, 3349, 3379, 3410, 3437, 3464, 3491, 3518, 3552, 3578, 3614, 3639, 3671, 3700, 2483, 3736, 3766, 2511, 2537, 2559, 2581, 2610, 2636, 2661, 2681, 2703, 2731, 2753, 2778, 2804, 2830, 2853, 2874, 2895, 2912, 2932, 2958, 2981, 3002, 2570, 3797, 3823, 4538, 4559, 4590, 4626, 4654, 4686, 4713, 4739, 4764, 4787, 4817, 4842, 4870, 4899]. Business [4920, 4947, 4973, 5004, 5029, 5055, 5079, 5101, 5129, 5153, 5185, 3848, 5208, 3877, 3909, 3938, 3966, 3995, 4024, 4053, 4088, 4112, 4139, 4166, 4195, 4222, 4254, 4279, 4311, 4336, 4364, 4391, 4418, 4447, 4480, 4507, 4491, 3833, 5238, 5264, 5946, 5971, 5995, 6022, 6046, 6076, 6102, 6127, 6152, 6180, 6206, 6237, 6264, 6290, 6320, 6347, 6371, 6398, 6420, 6442, 6464, 6487, 6511, 6539, 5291, 5319, 5344, 5371, 5401, 5432, 5458, 5482, 5509, 5534, 5564, 5594, 5626, 5648, 5678, 5703, 5729, 5759, 5783, 5811, 5839, 5865]. Business [5890, 5917, 6472, 6569, 6593, 7228, 7251, 7277, 7297, 7322, 7352, 7376, 7399, 7424, 7448, 7476, 7499, 7522, 7546, 7572, 7598, 7628, 7652, 7681, 7711, 7742, 7775, 7804, 7835, 7865, 7893, 6616, 6639, 6666, 6689, 6710, 6734, 6762, 6788, 6813, 6835, 6862, 6887, 6904, 6931, 6958, 6988, 7012, 7040, 7062, 7089, 7120, 7145, 7170, 7201, 7921, 7949, 8671, 8699, 8728, 8758, 8788, 8820, 8856, 8888, 8916, 8947, 8977, 9008, 9040, 9072, 9106, 9132, 9166, 9197, 9230, 9264, 9297, 9330, 9360, 9390, 9421, 9453, 7977, 9489]. Business [8008, 8038, 8064, 8069, 8123, 8150, 8177, 8208, 8238, 8267, 8293, 8318, 8342, 8378, 8410, 8441, 8468, 8500, 8531, 8556, 8583, 8613, 8642, 9521, 9546, 10216, 10244, 10273, 10301, 10334, 10366, 10399, 10428, 10458, 10492, 10520, 10555, 10589, 10621, 10652, 10678, 10708, 10736, 10762, 10790, 10816, 10846, 10871, 10896, 10922, 9573, 10949, 9603, 9629, 9660, 9687, 9714, 9739, 9763, 9793, 9818, 9846, 9872, 9901, 9930, 9952, 9976, 10000, 10024, 10055, 10077, 10106, 10134, 10161, 10187, 10844, 10869, 10894, 10920, 10946, 10972, 10973, 11001, 11545, 11574, 11601]. Business [11628, 11655, 11675, 11699, 11719, 11743, 11767, 11791, 11819, 11850, 11874, 11897, 11918, 11944, 11964, 11985, 12015, 12037, 12061, 12083, 12110, 12138, 11023, 12163, 12189, 12217, 12241, 12272, 11052, 11075, 11100, 11126, 11152, 11178, 11201, 11224, 11250, 11277, 11305, 11333, 11357, 11385, 11412, 11442, 11467, 11490, 11519, 10998, 11021, 11050, 11099, 11125, 11151, 12298, 12325, 13001, 13030, 13059, 13086, 13113, 13138, 13165, 13191, 13214, 13241, 13266, 13293, 13321, 13349, 13373, 13394, 13420, 13451, 13475, 13500, 13525, 13548, 13574, 13599, 13625, 13654, 12349, 13682, 12377, 12401]. Business [12427, 12455, 12479, 12509, 12545, 12569, 12599, 12635, 12661, 12690, 12723, 12754, 12782, 12810, 12840, 12867, 12898, 12925, 12951, 12976]. Bussell [924]. Bust [1830]. Bustard [4099]. Butchers [6915]. Butler [2107]. Butt [765]. Butter [349, 2882, 4984, 4956, 4966, 5450, 5522, 5717, 5851, 5546, 9368, 10346, 10226, 12872]. Butter-Worker [10226]. Butterflies [11324]. Butterfly [638]. Butteris [3239]. Button [1935, 3750, 7113, 7053]. Button-Hole [1935]. Buttons [8715, 10559]. Buzzard [9096].

Cab [11344]. Cabinet [11956, 10993, 11869, 13497, 12834, 13565, 13646]. Cable [803, 2891, 4317, 4575, 6330, 5824, 10371, 10338, 12521, 13637, 12878]. Cables [783, 398, 11471, 11096]. Cabs [4691, 11862]. Cage [10035]. Cailletet [11157]. Cairo [2038]. Caisson [635, 1792, 1563, 2311, 1547, 1674]. Cake [365, 13567]. Cakes [143]. Calabash [46]. Caladium [7619]. Calais [5376, 11722]. Calais-Douvées [11722]. Calamar [12863]. Calamity [2197, 5776]. Calcareous [12806]. Calculating [8596, 10012, 13412]. Calculation [7686]. Calculator [7337]. Calender [11484]. Calf [10172]. Calfskin [4911]. Calico [4382]. Calicoes [11982]. California [2882, 3707, 3384, 7986, 12099, 11424, 12708, 4351]. Californian [10096]. Caliper [9538, 10385]. Calipers [9528]. Calk [5550, 12391]. Calks [1118]. Call [11893]. Called [2228]. Calles [2363]. Callipers [11116]. Calls [12936]. Caloric [12125]. Calorimeter [9022]. Cam [5706, 11685]. Camacho [9315]. Cambridge [3925]. Came [3810]. Camera [5795, 11763, 11932]. Cameras [9581]. Cameron [3556, 7982, 10881]. Camp [6725, 2124]. Campbell [8792]. Camphor [4904]. Cams [8765]. Can [726, 3130, 3851, 6156, 6052, 6357, 6454, 6715, 7540, 8486, 8185, 10319, 11364, 11848, 12452, 13584, 12493, 3893, 8453]. Canada [3045, 3050, 3072, 3096, 3115, 3144, 3170, 3201, 3229, 3253, 3288, 3320, 3347, 3375, 3407, 5261, 5288, 5317, 5341, 9771]. Canadian [18, 2947, 3244, 3156, 4085, 4193, 4736, 5206, 5235, 5260, 5287, 5969, 5993, 6019, 6044, 6073, 6099, 6124, 6150, 6178, 6203, 6235, 6262, 6287, 6318, 6344, 6369, 6396, 6417, 6440, 6461, 6485, 6509, 6537, 5316, 5368, 5398, 5429, 5456, 5479, 5506, 5562, 5591, 5624, 5646, 5675, 5700, 5727, 5756, 5780, 5808, 5836, 5862, 5914, 5943, 6591, 6614, 7474, 7497, 7519, 7545, 7570, 7595, 7625, 7651, 7679, 7709, 7740, 7772, 7801, 7862, 7892, 6636, 6663, 6687, 6708, 6731, 6760, 6785, 6810, 6834, 6859, 6885, 6902, 6928, 7249]. Canadian [7275, 7295, 7319, 7350, 7374, 7397, 7422, 7445, 7947, 7974, 8006, 8036, 8062, 8093, 8120, 8147, 8206, 8235, 11480]. Canal [54, 672, 896, 1596, 1613, 1622, 3190, 3151, 3395, 3079, 2807, 3273, 3425, 3473, 3974, 5378, 6648, 8459, 13271, 12393, 12931, 13144, 1166, 3450, 346, 7, 14, 1718, 1921, 1641, 1666, 1545, 1751, 1854, 1696, 2891, 2469, 2849, 4774, 4472, 4580, 12221, 11825, 12985, 13620]. Canals [280, 701, 1725, 1890, 1914, 3441, 3567, 2993, 2466, 5350, 1736, 1767, 1791, 1843, 1869]. Canary [12578]. Cancelled [5413]. Canceller [10382]. Cancer [1948]. Candicans [11219]. Candle [6242, 10088, 13467, 13276]. Candles [5490, 12821]. Candlestick [11889, 13639]. Candy [5360]. Cane [5389, 8106, 10858, 12523]. Canned [8864]. Cannon [4327, 5266, 7272]. Cannot [10033]. Canopus [5358]. Cans [720, 579, 1710, 12361]. Can’t [11969]. Caoutchouc [4980, 4796]. Cap [720, 7538, 9177]. Capabilities [2444]. Capable [10797]. Cape [9177, 12715]. Capercailzie [10863]. Capillary [5547]. Capital [1719, 13067]. Capitalists [6402]. Capitol [961, 3044, 5293, 8348]. Capri [1004]. Capron [245]. Caps [3716, 2644]. Capstan [9529]. Capsuling [8422]. Capt. [1143, 13459]. Captain [511, 187, 5412, 7585, 11474, 12056, 13200]. Captive [11954, 11865, 11929, 13093]. Car [334, 1055, 1947, 1722, 2263, 3365, 3430, 3297, 2972, 2991, 3111, 4014, 4645, 5017, 4501, 4204, 4959, 4576, 3959, 5048, 5823, 5303, 5953, 5379, 7085, 7008, 7140, 7843, 6945, 7309, 6619, 7786, 6753, 8866, 9142, 8807, 8831, 9205, 8744, 8860, 10613, 10607, 10665, 10373, 10593, 10312, 11672, 11913, 11668, 11362, 13661, 13530, 13536, 12446, 13701, 13124, 13383, 12205]. Car-Washing [7309]. Caravans [1737]. Carbolic [1600, 11163, 273]. Carbon [2625, 4453, 4792, 5487, 11030, 11551, 12202, 11706, 12197, 13606, 12554]. Carbonates [8198]. Carbonating [9914]. Carbonic [1851, 2649, 4202, 7818, 7757, 7611, 8651]. Carbonite [5191]. Carbureted [1189]. Carbureter [12773]. Carbureting [1742]. Carburetter [3789, 6627]. Card [3534, 3563, 5188, 5212, 6514, 6544, 7808, 7840, 7869, 9396, 9425, 9459, 9494]. Carding [2122]. Care [1290]. Careless [98, 5224]. Carelessness [6187]. Cargoes [5322]. Carillon [7557]. Carnes [11093]. Carnivorous [7634, 10263]. Carolina [3527, 6517, 10355, 12501, 216, 10888]. Caron [5799]. Carp [988]. Carpenter [855, 3330, 12303]. Carpet [220, 8166, 10416, 9613, 11984, 12649]. Carpeting [1136, 1519, 9279]. Carpets [2859, 3530]. Carr [11618]. Carre [388, 4522]. Carriage [942, 2123, 1790, 3107, 3987, 5089, 6240, 7008, 7706, 7797, 7830, 7889, 7918, 9254, 8951, 8428, 8001, 8032, 9066, 9100, 9128, 9160, 9188, 9223, 9260, 9291, 9322, 9353, 9414, 9447, 8114, 8145, 8173, 8202, 8232, 8262, 8313, 8371, 8405, 8434, 8465, 8526, 8550, 8578, 8609, 8635, 8666, 8723, 8816, 8849, 8883, 8912, 8970, 9032, 9621, 9569, 9658, 10548, 10586, 10647, 10674, 10701, 10758, 10786, 9684, 9709, 9759, 9791, 9816, 9867, 9897, 9926, 9948, 9996, 10018, 10049, 10131, 10159, 10184, 10210, 10240, 10268, 10295, 10330]. Carriage [10362, 10394, 10425, 10456, 3870]. Carriages [187, 2648, 7667, 8139]. Carrier [6407, 7779, 6756, 8835, 12744]. Carrigeen [10252]. Carry [318]. Carrying [6776, 10249]. Cars [650, 983, 2105, 2339, 1743, 2273, 2290, 1496, 3596, 3674, 5014, 4702, 7610, 8714, 8895, 10434, 9751, 13382, 13662]. Cart [4270, 7218, 8153]. Carthagena [2411]. Carton [1740]. Cartridge [6752, 8806, 13444]. Cartridges [7272]. Carvalho [11171]. Carver [11712]. Carving [1114, 7379, 8656, 13435, 12882, 1429]. Case [156, 627, 1220, 1426, 2349, 1480, 2945, 4286, 4127, 5614, 6657, 7394, 8994, 9346, 7932, 10855, 10862, 10561, 12965, 12489, 1832, 11007]. Casehardening [10318]. Casement [5473, 5460]. Caserta [5883]. Cases [1947, 2000, 2071, 2132, 2670, 2676, 2937, 2977, 6981, 12153]. Cash [207, 11130, 12199]. Cask [10538, 12294]. Caspian [2756]. Casson [12339]. Cassons [12233]. Cast [69, 1293, 2816, 2756, 12717]. Castalia [8248]. Caster [11592, 10989]. Casters [3651, 5248]. Casting [1096, 1431, 3261, 5616, 5955, 7388, 7668]. Castings [5484, 6293, 5485, 6274]. Castor [9378]. Casts [3248]. Casualties [3704]. Cat [9971]. Cataleptic [7462]. Catalpa [11625, 12835]. Catarrh [12597]. Catastrophism [10250, 10253]. Catch [3363, 4851, 6505]. Catcher [2743, 5900]. Cathedral [6924]. Cattle [5014, 6310, 6806, 6682, 6942, 7405, 8356, 9590, 13023, 12706, 12414]. Caucasian [8197]. Caucasus [3528]. Causation [8923]. Cause [4175, 5734, 7453, 7304, 6996, 9638, 11463, 12907, 13673]. Caused [2870]. Causes [1876, 3187, 3396, 8826, 11502, 1216]. Cave [2288, 5957, 599, 599]. Caveat [5517]. Caveats [4342, 5261, 5288, 5317, 5341]. Cavender [7342]. Cavern [12367, 12564]. Cecropia [4704]. Celebrated [11071]. Celebration [1842, 8793]. Celestial [2459]. Cellar [7045]. Cellars [1693]. Cells [7413, 7485]. Celluloid [7529, 12701]. Cement [10977, 2862]. Cements [2462, 2488]. Cenis [1272, 1238, 2126, 1223, 2541, 2687, 6466]. Centennial [7658, 8966, 1774, 3777, 5132, 6435, 5695, 5682, 5328, 5847, 5269, 6157, 7046, 7479, 7541, 7824, 7359, 6840, 9173, 9303, 9268, 8543, 8591, 8215, 9365, 9027, 9092, 9122, 9340, 9405, 8792, 9110, 8347, 8587, 8541, 8015, 8624, 8650, 8680, 8709, 8737, 8767, 8796, 8865, 8898, 8687, 8535, 8565, 9119, 8504, 9123, 9152, 9184, 9253, 9286, 9319, 9349, 9380, 9411, 9442, 9479, 9254, 8989, 8937, 8563, 8418, 9316, 9300, 9044, 8920, 8432, 8212, 9466, 9364, 9247, 9240, 9285, 8564, 8589, 8361, 8509, 8540, 8660, 9214, 8590, 8712, 8839]. Centennial [9169, 8012, 7981, 9558, 9959]. Center [3686, 6850, 9110]. Centering [918, 2147, 10202, 11317]. Central [1011, 1492, 7897, 7206, 11361, 13319, 966, 415, 1798, 1328, 4273, 5710, 9942]. Centrifugal [3422, 5220, 8220, 11376, 13333]. Century [1034, 267]. Cere [265]. Cereals [7781, 11450]. Cerebral [10823]. Ceremony [1798]. Cereus [11219]. Certain [174, 2085, 3589]. Cerulean [2195]. Cesnola [8544]. Cesspools [4976]. Chætodon [12370]. Chaffee [3542]. Chaffin [8778].

Chain [347, 1872, 2756, 4421, 4675, 6429, 7017, 9097, 10063, 10608, 10042, 11623, 13642]. Chair [1486, 1392, 1487, 1765, 2868, 2522, 4199, 4677, 4125, 4520, 5497, 9281, 13445, 13593, 12894]. Chalk [10262]. Challenge [5140, 10467]. Chamber [7489, 5169]. Chameleon [12969, 12501]. Champion [2242]. Chance [4998, 6107, 7301, 11824, 12048, 12833]. Chandelier [8625, 45]. Chandler [203, 4432]. Change [3771, 8925]. Changed [3585, 9274]. Changes [748, 1595, 3187, 3158, 5063, 6084, 6962, 13175, 6871]. Changing [3922, 7562]. Channel [6363, 7130, 6946, 7552, 8248, 12283, 12054]. Chanoine [13660]. Chaos [5737]. Chapell [11072]. Chapman [2718, 12260]. Chapter [549, 1959]. Character [4514, 5575]. Characteristics [6472, 10280]. Charcoal [1521, 3892, 6576]. Chard [12065]. Charge [3131]. Charger [10227]. Charges [6867, 6755]. Chariot [6619]. Charitable [11902]. Charity [12330]. Charlatanism [400]. Charlatans [400]. Charles [8247, 2412, 1720]. Chartreuse [8506]. Charts [1688]. Chas [744]. Chas-Sepot [744]. Chase [1989, 11707]. Chasing [6791, 8454]. Chatsworth [5499]. Cheap [949, 1123, 3560, 3166, 3262, 3136, 5216, 4011, 6552, 5975, 6423, 9202, 9047, 11610]. Check [7724, 7909, 8595, 8873, 10574, 11371, 12467, 12526]. Checker [4932]. Checking [1986]. Checks [10662]. Cheeked [10697]. Cheerful [6563]. Cheese [4699, 8027, 9504]. Chelodine [11119]. Chemical [906, 338, 843, 2152, 1277, 1194, 2627, 3781, 2586, 3399, 2906, 2941, 3981, 4370, 4303, 4325, 4877, 4694, 5551, 5439, 7781, 7707, 7736, 7769, 7796, 7832, 7860, 7890, 8921, 8809, 8033, 8057, 8973, 9031, 9063, 8088, 8118, 8143, 8174, 8203, 8231, 8260, 8314, 8339, 8372, 8497, 8548, 8577, 8607, 8664, 8695, 8724, 8784, 8815, 8848, 8880, 8911, 8942, 9273, 9650, 9586, 9651, 11030, 12769, 7419]. Chemicals [275, 3237, 9027, 9092, 9122, 8770]. Chemistry [1013, 1030, 1056, 472, 901, 643, 1175, 620, 205, 1805, 2627, 2445, 2716, 2800, 4174, 3931, 3841, 6793, 7871, 8605, 9335, 8603, 13380, 12849, 12937, 12960, 13096, 13236, 13149]. Chemists [1054]. Chersydrus [10941]. Chess [10377]. Chest [8254, 10064, 12474]. Chicago [615, 2312, 2197, 2276, 2310, 2907, 4302, 4490, 5999, 10771]. Chichester [12335]. Chicken [10353]. Chickens [5792]. Chickering [12043]. Chicle [12906, 13635]. Chief [7, 1143, 626, 1695, 1809, 5243]. Child [4182]. Children [2527, 6771, 10280]. Chili [12200]. Chilian [6067, 6775]. Chill [666]. Chilling [3261]. Chimney [1293, 1188, 5388, 6950, 8840, 8745, 11318]. Chimneyless [10612]. Chimneys [268]. Chimpanzee [12618]. China [9858, 11211, 12264, 13357, 13381]. Chinese [113, 5231, 6977, 8461, 8962, 9476, 11879, 11234]. Chip [45]. Chironectes [3159]. Chisel [7074]. Chloral [7331, 718]. Chloralum [1025]. Chlorination [11895]. Chlorine [13330]. Chloroform [965, 5123, 6403, 11008]. Chlorophyl [13439]. Cholera [2130, 6000, 10821, 12285, 11690]. Choosing [3628]. Chopper [2666, 4780, 12498]. Christian [3713]. Christie [5668]. Christmas [10956]. Chromatrope [7792, 7935]. Chromis [8226].

Chromium [71]. Chromos [12073]. Chromosphere [2786, 2813]. Chromostroboscope [8334]. Chronograph [5852]. Chronometer [7480]. Chrysalids [2010]. Chubbuck [1219]. Chuck [595, 1824, 4914, 4377, 4910, 5607, 8716, 9115, 10202, 10543, 9729, 11317]. Chucks [8876, 11290, 11343]. Church [5800, 2335, 5279, 9863]. Churchyards [3246]. Churn [2143, 4728, 5498, 8517, 13202]. Churning [4984, 5113]. Churns [1962]. Cider [8027, 11117, 13174]. Cigar [138, 3642, 4755, 7369, 9174, 10326]. Cigar-Cutting [10326]. Cigarette [11538]. Cimex [7021]. Cinchona [11043]. Cincinnati [584, 1008, 3038, 3367, 3106, 11811, 13145]. Circle [917, 6653, 13219, 10978]. Circles [6873]. Circuit [627, 79, 103, 131, 405, 2168, 12879]. Circular [1071, 690, 1847, 3295, 5015, 5553, 6922, 7460, 7370]. Circular-Saw [690]. Circulating [4465, 11370]. Circulation [1889, 4283, 8878, 9501, 313]. Circumvention [10499]. Cistern [1348, 2317]. Cities [2268, 1670, 4062, 4201, 4598, 6030, 12356]. Citric [13008]. City [650, 507, 536, 555, 1381, 2390, 2399, 2424, 2449, 2479, 2505, 2532, 2555, 2820, 3309, 2721, 3424, 4806, 6108, 6161, 6160, 5574, 6401, 7328, 7483, 9495, 8098, 9770, 9922, 11311, 1520, 1803, 2313, 2939, 3240, 4273, 4876, 4491, 3895, 6502, 6523, 6503, 6062, 6408, 6430, 6450, 6500, 6553, 7502, 6755, 7125, 6579, 6677, 6720, 8793, 8298, 10176, 10820, 10854, 12173]. Civil [601, 534, 3643, 5885, 5908, 10982]. Civilization [605, 2085, 7176]. Clads [219, 6775]. Claimed [371]. Clam [11432]. Clamond [5956]. Clamorous [10509]. Clamp [169, 2102, 2546, 3894, 6252, 6218, 7390, 7416, 8902, 10576, 10570, 9559, 12254]. Clare [674]. Clarence [10250]. Clark [274, 8625]. Class [2088]. Classes [5059, 7193, 9480]. Classification [856, 3340]. Claude [11479]. Claw [4610, 5668]. Claws [11196]. Clay [5278, 5304, 12440]. Clean [1353]. Cleaned [12957]. Cleaner [1188, 3179, 2523, 4104, 7101, 9394, 8427, 8402, 9750]. Cleaning [1536, 13593]. Cleanliness [2724]. Cleansing [7778]. Clearance [11527]. Clearing [5638, 12180]. Clearness [2841]. Cleopatra [8309, 10883, 11300, 11735]. Clerk [1671]. Clerks [6548]. Clerkships [1832]. Cleveland [4836, 6183]. Clever [516]. Clifton [5800]. Climate [1444, 3187, 3158, 3585, 7155]. Climates [13431]. Climax [13696]. Climbing [1618]. Clinical [7335]. Clipper [12369]. Clipping [13384]. Clock [3055, 4066, 5869, 5769, 6924, 9432, 10603, 11484, 12179, 11967, 11138, 13522, 13047, 13407, 13354]. Clocks [9499, 9371, 9555, 10201, 12524]. Clockwork [12973]. Close [602, 1172, 1804, 2419, 9365]. Closet [1665, 4729, 7316, 7760, 8357, 12093]. Closets [2241, 2437]. Closing [1343, 2387, 5798, 5545, 8357]. Cloth [2020, 6522, 5827, 8277, 8421, 11733, 12704, 12888, 12947]. Cloth-Folding [8277]. Cloth-Singeing [6522]. Clothes [503, 442, 195, 2040, 2839, 7000, 8136, 9024, 11261]. Clothes-Moth [503]. Clothing [12238, 12021]. Cloud [6242]. Clouds [5160, 7456]. Club [1371, 12009, 2899]. Clutch [4890, 7666, 9775, 12798]. Co [296, 1853, 6244, 12330, 626, 4286]. Co-Operation [296, 6244]. Co-Operative [1853, 12330]. Co. [7692, 9364, 9282]. Coal [533, 343, 281, 1700, 2382, 1351, 2992, 3150, 2542, 2983, 4132, 5192, 4451, 6306, 6495, 6161, 6193, 6374, 6298, 6910, 7900, 9047, 8800, 9141, 9696, 11032, 11782, 10993, 13508, 12678, 273]. Coal-Breaking [343]. Coals [5245]. Coast [818, 8211, 9707, 12436]. Coat [4379, 10892, 9674]. Coates [1849]. Coating [747, 3588]. Cobalt [8632, 241, 9892]. Cobbler [10542]. Cobweb [7440]. Coca [9252]. Cochineal [1636, 1591]. Cochliostema [7491]. Cochrane [13070, 1342]. Cock [10779, 11220]. Cockatoos [7696]. Cockchafer [527, 547]. Cockeye [9731]. Cockney [12045]. Cockroaches [549]. Cocks [1161]. Cocoa [441]. Cocoons [2010]. Cocuyo [5175]. Cod [1352, 3413, 8692]. Cod-Liver [1352]. Code [5217]. Codeine [4289]. Codfish [483]. Codling [11148]. Coefficient [10876]. Coentre [7822]. Coffee [3391, 2795, 3446, 3006, 3005, 4157, 4017, 5196, 5846, 9670, 10575, 11914, 12957]. Coffeepot [574]. Coffer [7015]. Coffin [5383]. Coffins [7387]. Cog [2490]. Coggia [5895]. Coil [1104, 7414, 7793, 1753]. Coin [11581, 11555]. Coinage [596, 10878, 1532]. Coincidence [11572]. Coining [8048, 8079]. Coke [7989]. Col. [626]. Colburn [466, 4097]. Cold [450, 2187, 1463, 1477, 2595, 3449, 3668, 3697, 3732, 3764, 2961, 7116, 11233, 13431, 12581]. Cold-Rolled [450]. Coleman [7367, 11407]. Collapse [603, 4061, 8216]. Collapsing [3853, 4431, 11536]. Collar [2194, 3536, 4609, 10285]. Collars [3021]. Collecting [2734]. Collection [10632, 8544]. Collections [11028]. College [154, 5006, 6523, 577, 883, 179, 678, 1875]. Colleges [902, 4395, 4033, 7657]. Collegiate [7332]. Collender [10260]. Collins [8594, 8595]. Collision [8703]. Collisions [11029, 13580]. Collmann [11611]. Collodion [3533]. Colmer [11244]. Colonies [1853, 6131]. Colonizing [11902]. Colony [10376, 11501]. Color [6982, 7315, 9426, 8923, 9554, 9581, 9906, 11134, 11367, 2715]. Color-Blind [9581]. Color-Blindness [11134]. Colorado [2189, 2944, 2688, 2668, 5715, 7283, 6625, 6744, 11035, 13069, 10609]. Colored [3247, 6198, 6578, 9402, 9581, 10449]. Colorimetric [9892]. Coloring [2187, 2643, 7729]. Colorman [11475]. Colors [992, 1588, 1478, 2460, 3977, 4568, 3887, 4116, 5769, 12765, 13254, 10090]. Colossal [1830, 9535]. Colt [2706]. Columbia [577, 12365]. Column [2059, 7615, 9316, 7993]. Columns [2756, 3880]. Comb [2761]. Comba [2687]. Combination [420, 4528, 5071, 3958, 4271, 5116, 6105, 6323, 5660, 6304, 7824, 7192, 9281, 9346, 9509, 9085, 9090, 10913, 11641, 12989]. Combined [384, 810, 1482, 1617, 1616, 1537, 1935, 2196, 1292, 1714, 1486, 2418, 2816, 3088, 3153, 3295, 3525, 3383, 3718, 4073, 5199, 4677, 5228, 4937, 5172, 4888, 4889, 4938, 5094, 4933, 5449, 5615, 6657, 7464, 7789, 6702, 7394, 7664, 6723, 7540, 8625, 9084, 8798, 10858, 10503, 10043, 10533, 11403, 12124, 11318, 11478, 13512, 12498, 13151, 13209, 13362]. Combining [11471]. Combs [4239]. Combustibility [1995]. Combustibles [4208]. Combustion [2725, 3189, 6769, 9141, 8618, 2493]. Come [11267]. Comes [4796]. Comet [2722, 5656, 5895, 5301, 2439]. Cometary [5950]. Comets [2224, 3141, 4514]. Comfrey [9649]. Coming [4565, 4806, 4290, 4320, 6106, 7070, 9046, 8242, 8181, 9907]. Commemorate [2211]. Commencement [1896]. Commencements [1875]. Commerce [2626, 11391, 141]. Commercial [7212, 12358, 5245]. Commission [2514, 7404, 651, 4063, 5410, 7096, 8602, 12285, 11262]. Commissioner [49, 2145, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4936, 4967, 5042, 4607, 4642, 11132, 816, 1078, 180, 1382, 1856, 1644, 1650, 2150, 1247, 2146, 2566, 2569, 3161, 3692, 3675, 3999, 4566, 5010, 4399, 5255, 5803, 5751, 6601, 7409, 7532, 6792, 9156, 9857, 10716, 9556, 12069, 13302, 4992, 5016]. Commissioners [402, 966, 323, 3979]. Commissionership [6297]. Commodore [9585]. Common [388, 660, 1572, 2723, 3089, 4572, 4980, 6891, 10902, 12211, 11797, 11968, 11541, 12905]. Commonplace [178]. Communication [6460, 5913, 5942, 8818, 10637, 10750, 11544, 11896, 11755, 11781, 11238, 11689, 13546]. Communications [3432, 3548, 3665, 3761, 3792, 3819, 3906, 4684, 4710, 4735, 4762, 4784, 4815, 4840, 4867, 4896, 4918, 4945, 4970, 5002, 5026, 5053, 5076, 5099, 5127, 5151, 5182, 5205, 5234, 3964, 3992, 4021, 4050, 4084, 4109, 4137, 4164, 4192, 4251, 4277, 4308, 4333, 4361, 4389, 4415, 4444, 4477, 4505, 4535, 4556, 4584, 4620, 4650, 5259, 5286, 5968, 5992, 6018, 6043, 6072, 6098, 6123, 6149, 6176, 6202, 6234, 6261, 6286, 6317, 6343, 6368, 6395, 6416, 6439, 6484, 6508, 6536, 6567, 5315, 5339, 5367, 5397, 5428, 5455, 5478, 5505]. Communications [5531, 5561, 5590, 5623, 5645, 5674, 5699, 5726, 5755, 5779, 5807, 5835, 5861, 5887, 6590, 6613, 7248, 7274, 7294, 7318, 7349, 7373, 7396, 7421, 7444, 7473, 7496, 7544, 7569, 7594, 7624, 7650, 7678, 7708, 7739, 7771, 7800, 7833, 7861, 7891, 6635, 7919, 6662, 6686, 6707, 6730, 6759, 6784, 6809, 6833, 6858, 6884, 6901, 6927, 6956, 6985, 7010, 7037, 7060, 7087, 7119, 7143, 7168, 7198, 7226, 7946, 7973, 8696, 8725, 8755, 8787, 8852, 8884, 8913, 8943, 8975, 9005, 9036, 9068, 9103, 9129, 9162, 9192, 9226, 9261, 9293]. Communications [9326, 9356, 9386, 9418, 9450, 9487, 8005, 9519, 8035, 8061, 8092, 8119, 8146, 8175, 8205, 8234, 8264, 8291, 8316, 8340, 8375, 8408, 8438, 8467, 8498, 8529, 8553, 8581, 8611, 8639, 8669, 9545, 9572, 10272, 10299, 10333, 10365, 10397, 10427, 10457, 10490, 10519, 10553, 10587, 10620, 10651, 10676, 10706, 10735, 10761, 10788, 10814, 10868, 10893, 10919, 10945, 10971, 9601, 9627, 9659, 9685, 9712, 9737, 9762, 9792, 9817, 9844, 9870, 9899, 9929, 9950, 9975, 9999, 10022, 10053, 10075, 10104, 10132, 10185, 10215, 10242, 9560, 10148, 10175, 10199, 10234]. Communications [10283, 10351, 10380, 10414, 10441, 10470, 10505, 10536, 10572, 10605, 10693, 10804, 10830, 10860, 10910, 10935, 10963, 9584, 9615, 9645, 9671, 9701, 9726, 9752, 9776, 9806, 9830, 9861, 9886, 9917, 9941, 9963, 9989, 10009, 10039, 10066, 10089, 10119, 10997, 11020, 11942, 11963, 12013, 12035, 12082, 12108, 12136, 12162, 12188, 12215, 12269, 12296, 11049, 11098, 11249, 11276, 11304, 11332, 11356, 11383, 11410, 11440, 11464, 11489, 11518, 11573, 11600, 11627, 11742, 11766, 11790, 11849, 11872, 10996, 11708, 11730, 11840, 11014, 11039, 11067, 11115, 11143, 11170, 11192, 11215, 11292, 11321, 11346, 11373]. Communications [11399, 11426, 11455, 11503, 11533, 11614, 11643, 12323, 12346, 13085, 13111, 13137, 13189, 13212, 13239, 13264, 13291, 13320, 13348, 13371, 13393, 13418, 13450, 13473, 13498, 13523, 13623, 13653, 13681, 12399, 12425, 12451, 12476, 12506, 12542, 12567, 12598, 12633, 12660, 12688, 12721, 12752, 12780, 12808, 12839, 12865, 12897, 12924, 12950, 12975, 13000, 13028, 13057, 2315]. Commutation [909]. Companies [203, 12067, 2267]. Company [2699, 10934, 13284, 1143, 1592, 2223, 5197, 4577, 4244, 6349, 9395, 9561]. Comparative [969, 9220]. Compared [219]. Comparison [388, 428]. Compartment [5653]. Compass [3709, 5931, 7460, 13177]. Compasses [5322]. Compensated [3864]. Compensating [9309]. Compensation [767, 2996]. Compensator [810]. Competition [1301, 4717, 11393]. Competitive [1832]. Competitor [4812, 12306]. Complete [11016, 9435]. Completed [1223, 13009]. Completion [5168, 7527]. Composed [8922]. Composite [4578]. Composition [2519, 2782]. Compound [1720, 1620, 1990, 4199, 3880, 3867, 6248, 5774, 6561, 7669, 7148, 9117, 8868, 10907, 9639, 10032, 9862, 10594, 12003, 11398]. Compounding [7938]. Compounds [44]. Compress [5032]. Compressed [508, 774, 1461, 2904, 2504, 6247, 8866, 12766, 12387, 773]. Compresses [7511]. Compressing [4779, 10525]. Compression [8047, 8078, 11527, 11097, 11191]. Compressor [9474, 10248, 10041, 13693]. Compressors [2541]. Compulsory [2651]. Comstock [13482]. Concavo [4996]. Concentrated [3741, 3589]. Concentric [4629]. Conception [6555]. Concern [13027, 13056]. Concerning [253, 3712, 3624, 5159, 4153, 5927, 6738, 8804]. Conch [13558]. Concluded [1890, 1914]. Concluding [3954]. Conclusions [9990]. Concrete [858, 2346, 3657, 4693, 4774, 3982, 10413, 13425]. Condemnation [2390]. Condensation [8245]. Condensed [1508, 7049, 7174]. Condenser [1010, 168, 5033, 3985, 11693, 11218, 12318, 12968]. Condensers [6324]. Condensing [11666]. Condiments [339]. Condition [1047, 2372, 4231, 4315, 6908, 8589, 10601]. Conditions [546, 174, 3396, 5632, 5974]. Conduct [2942, 4854]. Conducting [1507]. Conduction [578]. Conductor [6573]. Conductors [9468]. Cone [690, 8519]. Cone-Berried [8519]. Coney [13093]. Confectioner [11475]. Confectioners [12410]. Conference [11971]. Confidentially [11060]. Configuration [4268]. Confirmations [6673]. Confiscating [11229]. Confiscation [12304]. Conflagration [3633, 12817]. Conflict [347]. Conformator [11205]. Congo [10628]. Congress [454, 301, 730, 623, 2774, 2670, 2676, 2821, 3219, 3272, 3914, 4821, 3886, 3833, 4794, 4552, 4874, 3812, 5768, 5461, 8302, 8132, 10743, 11476, 11422, 12120, 11665, 12931]. Congressional [2887, 5538, 11231]. Conical [1010]. Conjurors [9635]. Conklin [395]. Conn [5293]. Connected [2686]. Connecticut [10464].

Connecting [334, 4199, 8628, 9908]. Connection [2745, 1753]. Connections [3880, 6755]. Consanguineous [8473]. Consciousness [3690]. Conservatism [7606]. Conservatories [8218]. Conservatory [359, 3586, 5200, 6011, 7032]. Considered [493, 2085]. Conspiracies [13403]. Constancy [7406]. Constant [12012]. Constipation [11724]. Constituents [11450]. Constitute [2172]. Constitutes [2110, 2146]. Constitution [12314, 6531, 6557, 6611]. Constructed [3742]. Constructing [4853, 10927]. Construction [147, 445, 3151, 3385, 3726, 4578, 4384, 4403, 5092, 4496, 4520, 6219, 6561, 5329, 6961, 7187, 7665, 8218, 9087, 10887, 10373, 11726, 13484]. Constructive [11927]. Consumer [72, 2743]. Consumption [989, 13431, 12907]. Contagious [4178]. Containing [5245]. Contamination [13068]. Contemporary [10560]. Contents [8692]. Contest [6313, 7292, 10599]. Continent [3274, 8213]. Continue [1074]. Continued [6056]. Continuous [5019, 12391]. Contraction [7749, 10466]. Contractor [1465]. Contractors [6830]. Contradictions [3600]. Contrast [971, 12669]. Contrasted [9]. Contributions [9045, 12201]. Controller [13006]. Controlling [8346, 9307]. Controversy [2153, 8878, 11275]. Convection [578]. Convenient [12772]. Convention [5885, 5908, 5352, 12437, 13011]. Conventions [5267]. Converted [413, 439]. Convertible [669]. Convex [4996]. Conveyance [1300]. Conveyer [5629]. Conveying [2382]. Convict [11393]. Coofer [12643]. Coogan [2244]. Cook [12056]. Cooking [48, 5199, 4484, 4117, 5331, 5962, 7183, 7489, 11596]. Cool [8794]. Cooler [7467, 12824, 12338]. Coolers [10143]. Coolest [3279]. Cooling [3928, 6357, 5938, 10358]. Coop [10353]. Cooper [1464, 4756, 12516]. Copal [2995]. Cope [1594]. Copies [3458, 3486, 3515, 3546, 3573, 3609, 6376]. Copper [770, 1357, 1379, 3311, 5384, 8518, 8350, 10342, 11895, 11758, 11970, 6528]. Copper-Bearing [8350]. Copson [2144]. Copying [1104, 1480, 2941, 12181, 13564]. Copyright [385]. Copyrighting [13358]. Copyrights [4413, 4442, 4475, 4228, 5927]. Coquette [11938]. Coral [5830, 5737, 8487]. Coralline [3005]. Cord [4751, 4665, 10717]. Cordonnier [10542]. Cores [1096]. Cork [1717, 3012, 4038, 6434, 8368]. Cork-Soled [6434]. Corking [6701]. Corks [429, 3012]. Corkscrew [13512]. Corliss [42, 8587, 8892, 13326]. Corn [985, 2144, 1452, 3087, 2883, 5071, 3830, 4179, 5297, 7342, 8108, 10855, 9754, 10775, 11808, 12467]. Corn-husking [1452]. Cornell [2336, 6549, 8255, 8989]. Corner [5143, 11254]. Cornice [12988]. Cornwell [1842]. Coroner [2131]. Corporation [1827]. Corps [8324]. Corpuscular [8654]. Correct [5436, 9936]. Correcting [11590]. Corrections [1715]. Correctness [6645]. Correlation [78]. Correspondence [680, 1, 39, 632, 685, 711, 738, 757, 781, 800, 826, 851, 868, 889, 914, 935, 956, 977, 1000, 1020, 1041, 1062, 1086, 1111, 1133, 1155, 62, 84, 107, 136, 166, 184, 212, 237, 262, 286, 308, 331, 357, 381, 409, 435, 461, 479, 499, 522, 543, 567, 729, 1835, 1627, 1782, 1182, 1206, 1761, 1786, 1814, 1839, 1865, 1887, 1909, 1929, 1954, 1983, 2006, 2029, 2053, 2077, 2095, 2119, 2139, 2160, 2183, 2209, 2234, 2258, 2283, 1235, 2306, 2333, 2358, 2379, 2406, 1259, 1285, 1312, 1339, 1366, 1389]. Correspondence [1422, 1450, 1474, 1504, 1528, 1560, 1584, 1608, 1631, 1656, 1681, 1708, 1733, 1922, 3547, 2429, 2455, 3099, 3119, 3145, 3171, 3202, 3230, 3254, 3289, 3321, 3348, 3378, 3409, 3436, 3463, 3490, 3517, 3551, 3577, 3613, 3637, 3669, 3698, 2481, 3733, 3765, 2509, 2535, 2558, 2580, 2609, 2635, 2660, 2679, 2702, 2730, 2751, 2777, 2829, 2873, 2894, 2931, 2956, 3001, 3028, 3051, 3074, 4649, 5150, 3796, 3821, 4362, 4390, 4416, 4445, 4479, 4506, 4536, 4557, 4589, 4625, 4652, 4685, 4711, 4737, 4763, 4785, 3847, 4816, 4841, 4868, 4897, 4919]. Correspondence [4971, 5003, 5028, 5054, 5077, 5100, 5128, 5152, 5183, 5207, 3875, 3907, 3936, 3965, 3993, 4022, 4051, 4086, 4110, 4138, 4194, 4220, 4252, 4278, 4309, 4334, 5236, 5262, 5837, 5863, 5888, 5915, 5944, 5970, 5994, 6020, 6045, 6074, 6100, 6125, 6151, 6179, 6204, 6236, 6263, 6288, 6319, 6345, 6370, 6397, 6418, 6441, 5289, 6462, 6486, 6510, 6538, 5318, 5342, 5369, 5399, 5430, 5457, 5480, 5507, 5532, 5563, 5592, 5647, 5676, 5701, 5728, 5757, 5781, 5809, 5568, 6592, 6615, 7199, 7227, 7250, 7276, 7296, 7320, 7351, 7375, 7398, 7423, 7446]. Correspondence [7475, 7498, 7520, 7571, 7596, 7626, 7680, 7773, 7802, 6637, 7834, 7863, 6664, 6688, 6709, 6732, 6761, 6786, 6811, 6860, 6886, 6903, 6929, 6957, 6986, 7011, 7038, 7061, 7088, 7144, 7169, 7948, 7975, 8640, 8670, 8697, 8726, 8756, 8819, 8854, 8886, 8915, 8945, 9007, 9038, 9070, 9105, 9131, 9164, 9195, 9228, 9263, 9295, 8037, 9328, 9359, 9388, 9420, 9452, 9488, 8063, 8067, 8121, 8148, 8176, 8207, 8236, 8265, 8292, 8317, 8341, 8376, 8409, 8439, 8499, 8530, 8554, 8612, 11873, 11943, 12014, 12060, 12109, 12137, 12216, 12240, 12270, 12297, 12324]. Correspondence [13164, 13190, 13213, 13240, 13265, 13292, 13372, 13419, 13474, 13499, 13524, 13547, 13573, 13598, 13624, 12348, 12376, 12400, 12426, 12454, 12508, 12568, 12634, 12689, 12722, 12753, 12781, 12809, 12866, 13029, 13058, 13112]. Correspondents [35, 58, 681, 708, 735, 755, 778, 796, 822, 848, 865, 886, 911, 932, 952, 973, 997, 1038, 1058, 1083, 1108, 1130, 1152, 1181, 80, 104, 133, 163, 209, 235, 258, 283, 303, 328, 354, 377, 406, 430, 458, 477, 496, 518, 540, 564, 587, 611, 628, 652, 1203, 1231, 1860, 1881, 1905, 1924, 1949, 1978, 2002, 2024, 2048, 2073, 2091, 2134, 2156, 2177, 2206, 2229, 2253, 2279, 2301, 2328, 2353, 2402, 2427, 1256, 1281, 1308, 1335, 1362, 1384, 1415, 1446, 1471, 1499, 1523, 1556, 1580]. Correspondents [1603, 1651, 1676, 1703, 1728, 1757, 1810, 2451, 2477, 3071, 3097, 3117, 3142, 3168, 3199, 3227, 3251, 3286, 3316, 3344, 3373, 3404, 3431, 3459, 3485, 3513, 3574, 3610, 3634, 3664, 3694, 3729, 2507, 3760, 3793, 2533, 2556, 2578, 2606, 2632, 2658, 2677, 2700, 2727, 2749, 2775, 2801, 2827, 2851, 2871, 2892, 2910, 2929, 2954, 2978, 2999, 3026, 3818, 3846, 4504, 4534, 4555, 4583, 4619, 4683, 4709, 4734, 4761, 4783, 4814, 4839, 4866, 4895, 4917, 4944, 4969, 5001, 5025, 5052, 5075, 5098, 5126, 5181, 3874, 5204]. Correspondents [5233, 3935, 3963, 3991, 4020, 4049, 4083, 4108, 4136, 4163, 4191, 4219, 4250, 4276, 4307, 4332, 4360, 4388, 4414, 4443, 4476, 5258, 5285, 5967, 6042, 6071, 6097, 6122, 6148, 6175, 6201, 6233, 6260, 6285, 6316, 6342, 6367, 6394, 6415, 6438, 6459, 6483, 6507, 6535, 6566, 5314, 5338, 5366, 5396, 5427, 5454, 5477, 5504, 5530, 5560, 5589, 5622, 5644, 5673, 5698, 5725, 5754, 5778, 5806, 5834, 5860, 5912, 5941, 6589, 3905, 5991, 6017]. corroding [1239]. Corrosion [11560]. Corrugated [10912, 11085]. Corti [6006]. Corundum [3527, 3602, 11952]. Cosmical [3187, 11437, 3648]. Cosmograph [10727]. Cosmographer [12921]. Cost [1520, 2486, 6377, 7733, 9220, 8351, 7968, 9636, 12174, 11990, 13382, 13541, 12666]. Costs [7660, 9798, 11750]. Costume [960]. Cot [4754]. Cotta [2760, 10720]. Cottage [4439, 7393, 7114]. Cottages [4354]. Cotton [697, 870, 2188, 1615, 1218, 1238, 2108, 2062, 3153, 3420, 2435, 4780, 5920, 5447, 5960, 5976, 5308, 7133, 7173, 8023, 8600, 10666, 10525, 11805, 11296, 11118, 13277, 2860, 140, 101, 130, 143, 189]. Cottons [12276]. Cottrell [7643]. Couch [10475]. Cough [13673]. Counter [5617, 7672, 7639, 9377]. Counterbalanced [4807]. Counterfeit [1003, 11581]. Countershaft [7464]. Countersink [8167]. Countries [3114]. Country [3560, 3246, 4439, 4354, 8652, 10629, 13284]. County [2390]. Coupler [4645, 3959, 8934]. Couplers [10593]. Coupling [334, 918, 1916, 2262, 2441, 3365, 3430, 3590, 3134, 3421, 3242, 4014, 5017, 5173, 5048, 8740, 9458, 10608, 10607, 11122, 11072, 11269, 11668, 13064, 13097, 13335]. Couplings [7085, 7189]. Court [627, 79, 103, 131, 405, 474, 2168, 3426, 5064, 7854, 8720, 13147, 432, 11007, 11131, 12684]. Courts [3693, 3728, 3759, 3791, 2692, 3817, 3872, 5179, 5202, 3933, 3989, 4047, 4081, 4134, 4161, 4188, 4216, 4247, 4304, 4329, 4358, 4386, 4411, 4473, 4533, 4581, 4680, 4706, 4732, 4758, 4811, 4864, 5148, 5254, 5282, 6069, 6120, 6199, 6231, 6258, 6283, 6314, 6365, 6533, 6564, 5311, 5336, 5364, 5394, 5425, 5452, 5475, 5527, 5557, 5587, 5619, 5696, 5752, 5804, 5832, 5858, 5910, 6015, 6040, 6586, 6634, 7622, 7701, 7794, 6660, 6684, 6728, 6831, 6856, 6882, 6925, 6954, 6983, 7035, 7058, 7117, 7196, 7224, 7246, 7347, 7371, 7442]. Courts [7471, 7494, 7542, 7567, 7592, 7940, 7998, 8029, 8055, 8086, 8169, 8200, 8228, 8258, 8288, 8310, 8337, 8369, 8462, 8492, 8546, 8574, 8662, 8750, 8813, 8846, 8909, 9155, 9320, 9513, 9541, 9622, 9757, 9864, 9923, 10015, 10045, 10127, 10154, 10178, 10208, 10236, 10264, 10291, 10748, 10800, 10826, 10903, 10970, 12046, 11650, 5989]. Cousins [806]. Coventry [10473]. Cover [900, 10383]. Covered [10559]. Coverer [6241]. Covering [6336, 8249, 13277, 10]. Covers [12527]. Cow [4995, 6300, 7646, 7821, 9703, 11071, 13495, 7929]. Cowl [6950, 8840]. Cowper [12652]. Cows [8322, 10744]. Crab [8049, 8080, 12185, 11890, 12714, 12775]. Crab-Spider [12775]. Cracked [8364]. Crackers [3081]. Cracow [1659]. Cradle [4677, 8135, 9920]. Crafty [12185]. Craik [840]. Crane [4716, 5743, 6854, 7386, 7684, 8221, 10281]. Cranes [5090, 5363, 6505]. Crank [6117, 5877, 10905]. Crawshay [12958]. Creaks [11816]. Cream [5613, 9533, 11144, 12533, 13052]. Cream-Raising [12533]. Credulity [13042]. Creeper [5985, 12184, 12492]. Cremation [5694, 7232, 9500, 9732, 5684]. Creosoted [12485]. Crescent [6678].

Crested [7589]. Cribbing [5742]. Cricket [10357]. Crime [2799, 6296, 10437]. Criminal [6159, 7017, 9577]. Criminals [6268, 7550]. Crinitum [7007]. Crises [12358]. Crisis [4924]. Criticism [3834]. Crocheted [2644]. Crocker [1916]. Crook [7632]. Crooked [11787]. Crookes [1966, 2269, 2266, 7508, 8654, 2315]. Crop [7384, 10252]. Crops [753, 1746]. Cross [4643, 9935, 13618, 11693]. Cross-Breeding [13618]. Cross-Bridging [4643]. Crosshead [9180, 10543]. Crossing [6839, 12054]. Croton [8746]. Croup [10223]. Crow [6059]. Crowned [2202]. Crows [8214]. Crucibles [6409, 10690]. Crude [3483, 4905, 4982, 4875]. Cruelty [251]. Crusher [1321, 8457, 12183, 11069, 13534, 12645]. Crushing [1939]. Crust [7124]. Crutching [6251]. Cryolite [1333]. Crystal [5227]. Crystallization [2834, 6521, 7158, 3043]. Crystallized [8676]. Crystals [3865, 5406]. Cuba [11975, 11995]. Cuban [3387]. Cube [9592]. Cuca [10959]. Cucumbers [7730]. Cuff [13076]. Culinary [4041, 4272, 7824]. Culprits [2015]. Cultivating [1752]. Cultivation [248, 70, 1746, 4723, 6338, 6600, 6623, 7877, 11043, 11355]. Cultivator [924, 669, 4780, 4608, 8713, 10534, 12498]. Culture [1485, 3082, 4569, 5109, 6494, 7619, 7430, 6821, 9465, 8100, 8631, 10627, 9858, 11396, 11206, 12435, 13040]. Cummings [5018, 4603]. Cup [2125, 4528, 8365, 9086, 9088, 10262, 11810, 12917]. Cupressus [7618]. Cups [12065]. Curare [13273]. Curculio [664, 3154]. Cure [1948, 1443, 6013, 5663, 7094, 10851, 10821, 12597]. Curing [9433, 13305]. Curiosities [4984, 9340, 9405, 9774, 10901, 11740]. Curiosity [2787]. Curious [471, 93, 206, 1426, 2858, 4148, 5749, 5872, 5963, 7158, 6998, 7811, 7019, 7674, 9268, 9432, 9477, 9060, 7984, 10714, 10263, 9880, 10969, 10014, 10954, 10835, 10936, 10979, 11697, 11833, 12277, 13134, 12644, 12445, 11018, 10867, 10911]. Curly [10418]. Curly-Haired [10418]. Currant [5523]. Currency [322, 8325]. Current [733, 3712, 6267, 8190, 13206, 12625]. Currents [8904, 11488]. Currier [2795]. Curtain [7721, 10257]. Curtis [11270]. Cushion [8107, 10504]. Cushioned [4130, 13534]. Cut [429, 1262, 4923, 6037, 6881, 8653, 9524, 9808, 11702, 12859, 13208]. Cut-Leaved [6881]. Cut-Off [4923, 6037, 8653, 9524, 9808, 12859, 13208, 1262, 11702]. Cutter [334, 1617, 1935, 1847, 1741, 3750, 3014, 2496, 6337, 6193, 5717, 6277, 5360, 6560, 6723, 6948, 6654, 7053, 8594, 10412, 10440, 11641, 12254, 12289, 11070, 12127, 13233, 13256, 763]. Cutters [5681, 9470]. Cutting [809, 1985, 1378, 1770, 3012, 2809, 3753, 3388, 4038, 4902, 4428, 4406, 5998, 6064, 5470, 6680, 7043, 7883, 9312, 9280, 9020, 9278, 8253, 10326, 10014, 10496, 9730, 9618, 12292, 10985, 11783, 12257, 11732, 13284, 12858, 12967, 10964, 1589]. Cuttings [1974]. Cycads [7343]. Cyclodes [13022]. Cyclone [9461, 7604]. Cycloscope [13253]. Cylinder [1208, 3982, 3896, 5041, 6433, 7239, 8245]. Cylinders [950, 418, 4486, 8803, 10471, 12238]. Cylindrical [6804]. Cymbals [2985]. Cyprus [3834].

D [271, 2127, 2145]. Daily [4401]. Dairy [9590, 12246, 13619]. Dalgairnsae [10668]. Dallas [7342]. Dam [2318, 7015, 13660]. Damages [1148, 12461]. Damp [6109]. Damper [1043, 7077]. Dampers [803]. Danbury [4914]. Danger [603]. Dangerous [1550, 4118, 7069, 8736, 8679, 7930, 9015, 11395, 11137]. Dangers [468, 1438, 2033, 3365, 3430, 6403, 5539, 7331, 8826, 8155, 8896, 12667]. Daniel [11110]. Danish [11906]. Danube [3418]. Darien [54, 672, 346, 1545]. Dark [13279]. Darkness [5544]. Darwin [8473, 8647, 9935]. Darwinian [5387]. Darwinism [2018, 9301]. Dashboard [7270]. Dasher [2143, 5498]. Dastardly [4743, 4689, 8648]. Date [9718]. Dates [9718]. Daubs [11341]. Davenport [2544, 2494]. David [1332, 5736, 12922]. Davidson [6252]. Davis [1055, 5882, 6066]. Dawn [604]. Dawson [7505]. Day [1271, 1317, 4003, 5769, 2172, 9784]. Days [1355, 6106]. Dayton [3789, 5706]. Dead [6030, 5711, 7554, 10434, 10740, 12171]. Deadly [7072]. Deaf [2597]. Dealings [9398]. Dean [12738]. Dear [6423]. Death [510, 819, 2127, 2250, 1465, 1942, 2079, 3711, 3604, 2953, 3024, 3498, 2823, 4292, 4374, 5218, 5409, 5325, 7455, 7045, 7356, 8449, 9029, 9585, 11233, 1054, 7084]. Debate [5465]. Debt [562]. Decapitation [9580]. Decay [438, 2341]. December [1195, 2573]. Deception [7209, 7235, 7257, 7285, 7302, 7357, 7432, 7661, 8623, 9722, 9748, 9772, 10481, 9829, 11496, 12555]. Decision [1043, 11, 404, 1644, 1650, 1280, 2099, 2115, 1755, 1522, 2145, 2146, 2175, 3161, 2836, 2448, 3661, 2647, 4373, 5598, 6770, 12258, 12046, 11007, 12196, 12684, 13604]. Decisions [923, 233, 375, 2150, 2023, 1414, 2168, 2277, 2300, 2173, 3692, 3693, 3728, 3759, 3791, 3727, 3790, 3817, 3872, 5179, 5202, 3933, 3989, 4047, 4081, 4134, 4161, 4188, 4216, 4247, 4304, 4329, 4358, 4386, 4411, 4473, 4533, 4581, 4680, 4706, 4732, 4758, 4811, 4864, 5148, 3816, 3871, 3903, 3988, 4018, 4046, 4080, 4133, 4246, 4385, 4440, 3844, 5255, 5803, 5751, 5254, 5282, 6069, 6120, 6199, 6231, 6258, 6283, 6314, 6365, 6533, 6564, 5311, 5336, 5364, 5394, 5425, 5452, 5475, 5527, 5557, 5587, 5619, 5696, 5752, 5804]. Decisions [5832, 5858, 5910, 5989, 6015, 6040, 5393, 5502, 5526, 5556, 6586, 6634, 7622, 7701, 7794, 6660, 6684, 6728, 6831, 6856, 6882, 6925, 6954, 6983, 7035, 7058, 7117, 7196, 7224, 7246, 7347, 7371, 7442, 7471, 7494, 7542, 7567, 7592, 6757, 6782, 6807, 9156, 7940, 7998, 8029, 8055, 8086, 8169, 8200, 8228, 8258, 8288, 8310, 8337, 8369, 8462, 8492, 8546, 8574, 8662, 8720, 8750, 8813, 8846, 8909, 9155, 9320, 9513, 9541, 9622, 9757, 9864, 9923, 10015, 10045, 10127, 10154, 10178, 10208, 10236, 10264, 10291, 10748, 10800, 10826, 10903, 10970]. Decisions [10340, 10465, 10512, 10630, 10661, 10689, 10770, 10879, 11010, 11084, 11650, 11303, 11328, 11571, 11589, 11754, 11953, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 13491]. Decline [4769, 11421, 11661]. Decorated [5087]. Decoration [904, 2323, 3080, 5216, 7913, 9151, 11271]. Decorations [732]. Decorator [12699]. Decortication [9833]. Decoy [6917]. Decoys [13447]. Decrease [3164]. Deep [894, 417, 5123, 4852, 4433, 4962, 7558, 7361, 11780]. Deep-Sea [894]. Deer [6899, 9476, 12803, 6089]. Defeat [2079, 8302, 13070]. Defeated [12603].

Defect [2111]. Defective [9527]. Defence [10339, 10224]. Defences [8211, 10083]. Defense [818, 6671, 10566]. Defensive [5790]. Deflection [401, 456]. Deformed [387]. Deformities [9698]. Defunct [675]. Degerfors [6008]. Degradation [10372]. Degrees [769, 2741]. Delamater [792]. Delaware [10059, 11825]. Delay [9638]. Delhi [5869]. Delivered [1648, 1715]. Delusion [7810, 13183]. Demagnetization [13364, 13386, 13408]. Demand [838]. Demenge [12416]. Demerits [9341]. Demming [4989]. Demoniacal [6352]. Demonstration [4830, 4861, 4892]. Demonstrations [8983]. Denied [3538]. Denmark [895]. Denn [365]. Denney [9472, 10938]. Dennis [2080]. Density [12356]. Dental [3306, 8423, 8738, 11425, 12496]. Dentistry [1550, 6195, 6229, 6250, 6280, 6307, 9644, 884]. Denudation [10835]. Denunciation [673]. Deodorizing [4849, 7575]. Department [703, 1494, 1539, 3646, 3680, 5021, 7506, 8387, 4876, 4801, 12089]. Departure [4006]. Dependence [2245]. Depositing [12560]. Deposition [1379, 6532]. Deposits [4451, 5516, 362]. Depot [453, 72, 2619, 7897, 7206, 8011]. Depreciation [751]. Depression [161, 5709]. Depth [11967]. Derrick [4829]. Descent [1726]. Desert [5630, 6051]. Design [962, 97, 1744, 3094, 2770, 3868, 5200, 7097, 6946, 7032, 12903]. Designed [187]. Designs [12423]. Desk [8396, 9346, 9913, 12881]. Desmognath [12893]. Despaquls [10370]. Despatch [3386, 6907]. Destination [3834]. Destroyed [10660, 11096]. Destroyer [5447, 12114, 12406]. Destruction [1074, 1942, 11726]. Destructive [859]. Detachable [395, 5668, 6432, 9051, 11481]. Detacher [6606, 10320, 13640]. Detaching [3623, 4858, 13698]. Details [13484]. Detect [1975]. Detection [3299, 8632]. Detective [6159, 6575, 12932]. Deterioration [533]. Determination [2741, 3627, 2696, 3510, 2763]. Determine [12584]. Determined [12024]. Determining [2345, 1648, 6973, 11960, 1715]. Detoxicated [9306]. Detroit [4151, 5572, 12855]. Develop [3452]. Development [812, 6405, 9045, 8347, 8961, 10767]. Device [1071, 1118, 87, 1243, 1343, 1536, 1936, 3010, 2989, 4857, 5230, 6168, 5742, 6499, 7334, 6879, 7754, 6776, 7055, 8905, 8715, 8684, 9988, 10504, 11239, 13023, 13560, 13052, 13316]. Devices [9246, 8168]. Devil [4777, 5549]. Dexter [9987]. Dexterity [11974]. Diagnosis [12589]. Diagometer [11940]. Diagonal [10171]. Diagram [6161]. Diagrams [11205, 5718, 5664]. Dial [6851]. Diamagnetism [762]. Diamond [809, 1057, 3707, 2809, 4780, 4902, 4077, 6155, 5408, 6669, 8484, 9457, 12254, 10986]. Diamonds [1995, 2796, 3281, 4234, 5134, 6050, 5442, 7719, 11566, 11329]. Diapason [11999, 11988]. Died [12607]. Dies [10014, 12886]. Diet [5766]. Difference [2653, 7360]. Different [2908, 6974, 6855]. Differential [317, 11507]. Difficulties [11976, 13557]. Diffusion [9078]. Digester [4117]. Digestive [4485, 4567]. Digger [11936, 13152, 12417]. Diggers [1530]. Digging [11114]. Dilemma [13302]. Dillon [10608]. Dime [12716]. Dimensions [4268]. Dimness [10409]. Dining [1487, 154]. Dinner [6949, 9401, 11403, 11242].

Dioptric [7820]. Dip [4432]. Diphtheria [7870, 10851, 12372]. Diplograph [10633]. Diploma [13143]. Direct [3238, 2845, 3826, 4562, 5720, 8836, 5302, 9836]. Direct-Acting [3238, 8836]. Direction [3922]. Director [11361]. Disappearance [9706]. Disaster [1403, 2799, 6197, 12247]. Disasters [3364]. Disastrous [6054, 7604, 9723, 11062]. Discharges [990]. Discharging [2992, 8800, 8363]. Discord [8775]. Discount [2067, 10438]. Discourage [12282, 12385]. Discouragement [11799]. Discouraging [3537]. Discovered [772, 7751, 9501, 11983, 12848, 8028, 8112]. Discoveries [1660, 3423, 3629, 3624, 2553, 3981, 4031, 3952, 5009, 5348, 8921, 8127, 13134, 13172, 12974]. Discovery [2245, 1407, 4925, 5106, 5734, 6135, 5602, 6091, 7508, 7901, 8605, 7956, 9428, 10687, 10406, 10498, 2786, 2813]. Discussion [1890, 1914, 2421, 1870, 5538, 5711, 7358, 13036]. Disease [1876, 2079, 3455, 3194, 2799, 3596, 2499, 3655, 6162, 7715, 10825, 9610, 12329, 2240]. Diseased [10167]. Diseases [4178, 6170, 7157, 10479, 13041]. Disengaging [5606]. Disgraced [11108]. Disgust [10595]. Dish [1003, 900, 7191]. Dish-Washing [1003]. Disinfectant [273]. Disinfectants [2421, 3453]. Disinfection [12857]. Disintegrating [3422]. Disintegrator [6362, 10350]. Disk [924]. Disks [11018]. Dispersal [9637]. Display [7046]. Disposal [1993, 636]. Disrespect [11312]. Dissected [9174]. Dissection [8131, 9580]. Dissipation [920]. Dissociation [1198, 13330, 2464]. Dissolving [161]. Distance [1077, 4772, 13276]. Distances [1055, 10882, 9826]. Distant [7961]. Distillation [2845, 11782]. Distilled [6855]. Distilleries [3444]. Distilling [8278]. Distinguished [510, 361, 1054]. Distributer [1220]. Distribution [7127, 8487, 9578, 11952]. Distributor [3108]. District [627, 474, 2168, 2699, 79, 103, 131, 474]. Ditcher [7092]. Ditching [9012, 11273]. Dittman [6972]. Divider [5228]. Dividers [986]. Dividing [9089]. Diving [158, 2261, 3888, 5932, 7691]. Divinity [11860]. Divisibility [843, 12979]. Division [6632]. Dixon [4375]. Do [2170, 3449, 2550, 5083, 6550, 8794, 9399, 11524, 1099, 1533, 5765, 7452, 10040, 9749, 13171]. Doane [3151]. Dobbin [11299]. Dock [651, 5197, 5363, 11270, 12560]. Docking [3441, 8282]. Docks [3657, 3309, 4826, 6473, 6652]. Doctoring [1749]. Doctors [91, 3944]. Dodge [719, 4286, 6939]. Dodlet [7056]. Dodo [7056]. Does [2079, 3109]. Dog [646, 5070, 6970, 11297, 11594]. Dogs [3690, 6771]. Doing [902, 1900]. Dollar [1026, 3535, 3675, 4995, 5084, 5378, 9938]. Dollars [1613, 1622, 4513, 6162, 8827, 8042, 8073, 13530]. Dome [12488]. Domestic [148, 670, 1642, 3152, 3113, 3469, 4753, 5850, 9562, 10629, 10667]. Dominica [11431]. Done [1099, 2594, 8793, 11881, 11160]. Donkey [7882, 10379]. Donnadieu [12938]. Don’t [6574]. Door [981, 810, 1919, 1265, 1540, 2418, 2167, 2768, 3331, 8595, 8682, 9000, 8693, 9988, 10728, 11112, 12235, 12654, 12526]. Doremus [1142, 1191]. Dormoy [12339]. Dorn [1847]. Dorr [10892]. Dosimeter [11244]. Double [111, 2059, 2865, 3328, 4705, 4314, 4483, 6254, 5720, 7906, 10721, 11221, 11194, 12319]. Double-Acting [4705]. Doubletree [5771]. Doubtful [12438]. Dough [6337, 8594, 12745]. Doum [11647]. Douro [11815]. Douvées [11722]. Dover [5040, 5376]. Dovetail [2585]. Dovetailing [5893, 11375, 12882]. Down [154, 76, 201, 2311, 6911, 4891]. Downfall [1358, 3536]. Dr [2266, 5037]. Dr. [516, 1142, 509, 1966, 2269, 2347, 1191, 3398, 3931, 5610, 5409, 5348, 5852, 9269, 10567, 12850]. Draft [903, 2903, 6850, 9054, 9200]. Drafting [4999]. Drafts [7231]. Draftsman [6354]. Draftsmen [1138]. Drag [13078, 13409]. Dragon [13224, 13235]. Drain [7778, 11027]. Drainage [4435, 12946]. Drainer [7191]. Draining [8299]. Drains [11150]. Draper [2347, 3931, 4034, 5348, 5463, 10917, 10990, 12850]. Draught [11531]. Draughtsman [11165]. Draughty [12177]. Draw [2527, 3930, 8798]. Drawbridge [8014]. Drawing [341, 879, 2013, 3755, 5415, 7728, 7842, 7957, 10824, 11031, 12292, 13537]. Drawings [1104, 300, 3625, 3160, 2710]. Drawn [9990]. Dray [4462]. Dreadnought [11047]. Dreams [804]. Dredger [11618]. Dredging [829, 1796, 3127, 4852, 8986, 10909, 10723, 9612]. Dress [1246, 7102]. Dresser [2141, 9457]. Dressing [311, 1570, 4234, 6209, 8195, 10028, 10986, 11652, 13286, 13702]. Drew [10636]. Drill [223, 1643, 1846, 2285, 1824, 3496, 4914, 4890, 4429, 4483, 4910, 5362, 6117, 5423, 8716, 9345, 9115, 10752, 10122, 11087, 11839, 12467, 12530]. Drill-Proof [5362]. Drilling [964, 2936, 4213, 9142, 8296, 8225, 10432, 12255, 12129, 11066, 11671, 13091]. Drills [12255]. Drink [1316, 4383, 11137, 7782]. Drinking [4657, 7331, 11502, 13068, 12917]. Drinks [689]. Drive [5976, 7780, 8735]. Driver [4129, 13614]. Driving [4227, 4183, 4172, 7459, 7149, 8483]. Drop [2878, 5828, 8836, 9371, 11267]. Drowned [11762]. Drug [1671]. Drugging [7528, 8622]. Drugs [468, 2079]. Drunk [10167]. Dry [124, 10041, 11270]. Dryer [195, 1611, 3005, 5821, 7000, 8136]. Drying [418, 2906, 2595, 2845, 2694, 2886, 2671, 5199, 11166, 13308]. Dual [9958]. Dualin [371, 247]. Duck [10914]. Duck-Billed [10914]. Duckham [721]. Ducking [3057]. Due [2646, 4143, 4057, 10223]. Dueber [8994]. Dugdale [442, 2040]. Duluth [1669]. Dumb [243, 2597]. Dumper [9531]. Dumping [4576, 6093, 8153, 9809]. Dundee [6602]. Dunn [7219]. Dunning [5173]. Duplex [3918, 4198, 7255, 7761, 10251, 11264, 13693, 12368]. Duplication [9592]. Dupré [6620]. Durability [2362, 7128]. Durable [570]. Durand [3015]. During [7280, 5765]. Dust [256, 1902, 2936, 9462, 11032, 12019]. Dustless [11530]. Dustpan [10411]. Duties [10716]. Dwelling [993, 1381, 9835]. Dwellings [7694]. Dyeing [56, 1712, 2165, 4382]. Dyers [1397, 1435, 1460, 1509]. Dyes [719, 10290, 10656]. Dynamic [398, 2363, 3016, 4965]. Dynamical [1222]. Dynamics [671]. Dynamite [8103, 11992]. Dynamo [8996]. Dynamo-Electric [8996]. Dynamometer [10348, 11855, 11929].

Eads [1143, 11474, 13200]. Eagle [9756, 12535, 220, 12286]. Ear [2443, 7135, 9182, 12394, 13381]. Earle [5985]. Early [4293, 3920, 4124, 3804, 4368, 6340, 6364, 10464, 13274, 12371]. Earth [320, 1665, 2241, 2437, 5172, 4341, 6084, 6670, 7699, 7282, 7124, 7493, 8804, 8394, 9338, 9171, 8705, 13038, 11657, 11809]. Earthenware [3559, 5244]. Earthquake [1017, 6967, 6916]. Earthquakes [13545]. Earwigs [4885]. Easel [11271]. Easily [12366]. East [1792, 1563, 2409, 2948, 6880, 9478, 635, 1792, 2311, 1547, 1987, 1809, 1410, 1674, 1408, 3769, 3092, 3524, 5346, 7953, 7952, 9636, 10371, 10338, 9767, 9743, 11471, 11310, 12087, 12670, 13091]. Eastern [627, 2168, 1215, 1663, 8213]. Easy [2522, 5217, 10943]. Eat [250, 7504]. Eater [7439]. Eaters [11983]. Eating [598, 6351, 7526, 10124, 10942]. Eats [10417]. Ebony [12474]. Eburine [10482]. Eccentric [47, 2434, 4207, 3810]. Eccentrics [1893]. Echinodermata [12997]. Echo [9333]. Eclipse [1441, 1195, 2573, 4027, 5949, 9690, 11451, 11835, 11323, 11798, 11812, 11834, 11978, 11736, 1329]. Economic [8658]. Economics [3739]. Economizer [2903, 5689, 10912, 9699]. Economy [726, 265, 228, 352, 445, 3469, 3739, 6351, 5634, 11193, 11626, 12576]. Edges [10964]. Edible [5873, 12919]. Edison [11683, 7956, 7984, 11827, 11609, 11728, 11674, 11828, 11667, 11706, 12769, 13432, 13583, 12850, 12627, 12628, 13366, 12742, 13533, 8112]. Edition [2928, 4352, 6376, 11497, 11996, 11663, 11776]. Editor [9856]. Editorial [25, 403, 419, 457, 517, 642, 1602, 1675, 1880, 1904, 1977, 2047, 2155, 2190, 2545, 2603]. Edmund [3456]. Educated [9082, 11232]. Education [581, 594, 694, 179, 2392, 1546, 3298, 3477, 2597, 2443, 2587, 5034, 4977, 5437, 5983, 7430, 7481, 6992, 8214, 9218, 8839, 10251, 10982, 12025]. Educator [12280, 13630]. Educators [11028]. Edward [1695]. Edwin [3542]. Eels [11983, 12311]. Effect [990, 812, 1827, 1463, 2652, 2961, 2822, 4148, 4643, 5929, 5766, 12067, 2616, 5750]. Effective [5652, 5685]. Effects [388, 1148, 124, 1995, 4200, 5878, 5925, 8904, 8446, 8507, 9935, 11488, 11567, 13417, 13482]. Efficiency [397, 2345, 6621, 1006]. Effluvia [1625]. Efflux [4747]. Efforts [3451, 6819, 12875]. Egg [271, 1535, 4381, 4404, 4434, 4324, 4350, 4267, 4296, 8960, 9088, 11553, 13678, 12158]. Eggs [1701, 9990]. Eggstraordinary [11882]. Egypt [2038, 2966, 6279, 8718]. Egyptian [5722, 11296, 117]. Eight [3138, 3018, 2994, 3046, 3064, 3091, 5214, 4093]. Eighty [7328, 7601]. Eighty-Ton [7601]. Ejector [13553]. Elastic [1392, 5036, 4122, 4101, 5815]. Elasticity [1673, 12047]. Eldridge [7139]. Electric [455, 898, 958, 1777, 1825, 1722, 2744, 3040, 4604, 4378, 4345, 4614, 4579, 5956, 5900, 5497, 5606, 6670, 7287, 7001, 7245, 7902, 7156, 7925, 8381, 7956, 8305, 8838, 8833, 8025, 9428, 8996, 10088, 10407, 10604, 10601, 10579, 9530, 10139, 10341, 12116, 12157, 12067, 11135, 11921, 12155, 12168, 12182, 12088, 11036, 12002, 12194, 12197, 12232, 12305, 13583, 12850, 12627, 12628, 13469, 12538, 13513, 12876, 13118, 13024, 12332, 12463, 13015, 13206, 12626, 13044, 13315, 13361, 12979, 12788, 12315, 12625, 12504, 12553, 12392, 11958, 13487, 12343, 504]. Electrical [1067, 1998, 3277, 2791, 3039, 4522, 5033, 4452, 4007, 5581, 7515, 9060, 10539, 10725, 9910, 11999, 13177, 13432, 13669, 11523, 11657, 11809, 11988, 8335, 13533, 12089]. Electricity [2286, 2225, 3782, 3712, 7901, 7509, 6573, 7954, 8162, 9274, 10466, 11662, 13109, 95]. Electro [803, 1379, 1405, 2261, 1462, 1747, 1806, 2941, 2646, 2746, 2835, 5355, 5384, 7217, 6629, 8921, 9315, 8423, 10462, 10563, 10562, 10195, 11030, 11074, 12769, 4965, 6532]. Electro-Chemical [8921, 11030, 12769]. Electro-Deposition [6532]. Electro-Dynamic [4965]. Electro-Magnetic [803, 1462, 2646, 2835, 6629, 9315, 8423]. Electro-Magnetism [11074]. Electro-Metallurgy [1405, 7217]. Electro-Motor [1747, 1806]. Electro-Plating [5355, 10462, 10562]. Electro-Pneumatic [2261]. Electro-Silicic [10563]. Electro-Static [10195]. Electrolytic [982]. Electromagnetic [1944, 11807]. Electromagnetism [5131]. Electromotograph...A [6135]. Electroscope [11314]. Electrostatic [7414]. Elegans [5775]. Elegant [12834]. Element [8347]. Elemental [11707]. Elements [2771, 3899, 12382]. Elephant [13050]. Elevated [425, 3838, 11658, 12144, 11080, 12128, 11012, 11254, 13541, 13484, 13586, 13455]. Elevation [5815, 10198]. Elevations [7961]. Elevator [364, 1027, 372, 1072, 2216, 2125, 3108, 3522, 4501, 3987, 5629, 5876, 7761, 7362, 7163, 7868, 8835, 9986, 10176, 12245, 12802]. Elevators [2122, 2878, 2998]. Elizabethan [12397]. Elliott [8746]. Ellipsograph [3086, 7586, 12885]. Ellis [4297]. Elmer [6116]. Eloquence [5161]. Elwell [11910]. Emaciation [1750]. Emanations [11150]. Emanuel [5800]. Emasculation [11339]. Embalming [10740]. Embankment [1356, 2384]. Embossed [6301]. Embroidered [12947]. Embryology [10567]. Emerson [4077]. Emery [115, 3140, 2990, 4498, 4577, 8107, 8329, 10343, 9987, 11753, 11390, 12799, 12829, 1720]. Emigrants [218]. Emigration [3451, 10529, 12577]. Emile [1991]. Emission [7818]. Emperor [8242, 8474, 8825]. Empire [11530]. Employed [7234]. Employees [6326, 11108]. Employer [7234]. Employers [5515]. Employment [1827, 1577, 3976]. Emptying [4976, 7990]. Enameled [6301]. Enameling [1369, 8421]. Encke [2439]. Encourage [11284]. Encouragement [2970, 9666, 11831]. End [1547, 3770, 3138, 4511, 4821, 5870, 7208, 8677, 9493, 10929, 12278]. Ended [3978]. Endless [1072, 1554, 2834]. Ends [1764, 10471, 13425]. Energetic [8474]. Energy [920, 525, 637, 3063, 13067, 13129]. Enforce [3494]. Enforcement [5517]. Eng [5499]. Engage [12071]. Engine [1073, 42, 1026, 463, 127, 2191, 1849, 1889, 1530, 1289, 1892, 1494, 1738, 3238, 2684, 2898, 3468, 3685, 3739, 4450, 4629, 4756, 3842, 4437, 5104, 3827, 3867, 4044, 4058, 4458, 4406, 4169, 4524, 3896, 5041, 4297, 4459, 6248, 5930, 6400, 5850, 6228, 6329, 5634, 5740, 6444, 5829, 6433, 6596, 7261, 7075, 7849, 7639, 7245, 7534, 8503, 9315, 8704, 9249, 8423, 8951, 8180, 8381, 9055, 8868, 9431, 8892, 9094, 9276, 9395, 8047, 8078, 8567, 9239, 8345, 9524, 10858, 9562, 9690, 10483, 10433, 10838, 10543, 10907, 9639, 10474, 10197]. Engine [9808, 10507, 10568, 10228, 9966, 10966, 10029, 10594, 11485, 11376, 12124, 11323, 11666, 11796, 12001, 11041, 12028, 11194, 11037, 11091, 12261, 11291, 12253, 12143, 11281, 11191, 11522, 12178, 11748, 11217, 12125, 13559, 12416, 13127, 12406, 12709, 13326, 12648, 13075, 12768, 12889, 13021, 13645, 9733]. Engineer [466, 626, 1222, 5243, 8324, 9909, 10982, 1143, 1809, 2015, 4801, 6160, 13387]. Engineering [112, 326, 945, 466, 96, 601, 534, 837, 2111, 1324, 2393, 1193, 3394, 2943, 2468, 2459, 3898, 4012, 4679, 5141, 5270, 5357, 6243, 7577, 7559, 7732, 6765, 7704, 7738, 7770, 7798, 7829, 7859, 7887, 7917, 7945, 7970, 8000, 8031, 8059, 8090, 8263, 8404, 8437, 8466, 8494, 8527, 8552, 8638, 8754, 8786, 8940, 8972, 9004, 9099, 9222, 9325, 9383, 9448, 9483, 9516, 9543, 9600, 10422, 10453, 10485, 10516, 10549, 10582, 10618, 10649, 10672, 10704, 10734, 10760, 10783, 10810, 10842, 9626, 9736, 9760, 9790, 9812, 9841]. Engineering [9869, 9945, 9974, 9997, 10020, 10048, 10072, 10100, 10129, 10156, 10180, 10213, 10239, 10270, 10293, 10329, 10360, 10396, 12622, 12770, 12825, 12883, 12915, 12991, 13055, 13135, 13181, 13203, 13237, 13262, 13290, 13309, 13345, 13442, 13468, 13493, 13509, 13572, 13680, 13163, 12529, 12959, 11351, 2011]. Engineers [691, 3643, 2976, 5885, 5908, 6830, 7926, 5569]. Engines [17, 228, 618, 92, 324, 1050, 150, 47, 1322, 1720, 1620, 2218, 1491, 1858, 1315, 3597, 3393, 3476, 3278, 2667, 2738, 4771, 4008, 4155, 4225, 3949, 5652, 6520, 6324, 5347, 5537, 6353, 6080, 6086, 7758, 7487, 7047, 7284, 7645, 8587, 7938, 8458, 8471, 8929, 8134, 8186, 9169, 10085, 11540, 11398, 11425, 11264, 11862, 11585, 12368, 12728, 12859, 13039, 12496]. England [2149, 6412, 11976, 968, 1024, 698, 1092, 484, 2372, 2335, 2361, 1768, 1940, 2043, 2062, 3658, 3537, 3626, 2615, 2463, 2918, 3918, 4891, 4242, 5040, 4145, 5280, 6473, 5800, 6272, 5380, 6025, 6542, 6011, 7181, 7785, 7290, 7071, 9029, 8102, 9991, 10529, 12068, 12042, 13430, 13036, 12306, 12522]. English [9, 219, 750, 146, 1465, 1324, 2170, 3085, 3212, 3834, 3961, 5085, 4987, 4549, 4717, 6363, 6057, 5440, 5385, 7165, 7428, 7358, 7552, 9112, 9504, 9398, 9406, 10308, 9827, 11344, 12276, 12669, 12371, 13513, 12816, 5671]. Engraving [713, 120, 160, 2396, 3388, 7286, 8454, 10370, 10448, 10934, 10387, 13426]. Engravings [276, 1964, 3310]. Enjoyments [4171]. Enlarging [1009]. Enormous [7384]. Enough [2343, 11136]. Enslaves [1719]. Entailment [7017]. Enter [988]. Enterprise [2699, 8706, 11975]. Enterprises [12874]. Enterprising [844]. Entertainment [4661]. Entirely [8922]. Entitled [1890, 1914]. Entomology [3914]. Envelope [1899, 4499]. Envelopes [1513]. Eothen [11664]. Epidemic [3568, 9139, 9879]. Epidemics [2799, 6296]. Equation [1437]. Equilibrio [10475]. Equine [12916]. Equivalent [5741]. Equtlibrium [12012]. Erastus [13688]. Eremacausis [5684]. Ericsson [511, 187, 6527, 13459]. Erie [1718, 1921, 1641, 1666, 1751, 1854, 1696, 2891, 2849, 13620]. Eruption [10282, 11539]. Escape [3502, 4424, 4612, 10535, 10506, 10712, 10011, 11213, 11915, 11688, 13410]. Escapement [119]. Esourial [3606]. Especially [413, 439]. Essex [4891]. Established [748, 13027, 13056]. Establishment [666, 6794, 12434]. Establishments [13284]. Estes [1889]. Estimate [12845]. Estimating [4262, 6331, 11450, 8425, 8222, 8250]. Estimation [6524, 5245]. Etc [395, 972, 943, 1055, 1488, 2338, 1962, 2323, 2985, 2811, 3006, 3388, 5071, 4990, 5826, 8596, 8249, 9310, 10913, 11371, 11450, 12363, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 2493, 13491]. Etc. [939, 959]. Ether [2153, 1976, 3207, 2625, 6597, 6599, 11612]. Etheric [7925]. Etiology [11690]. Eureka [7588]. Europe [305, 3022, 3778, 3682, 4823, 4911, 4600, 4616, 4634, 4879, 4979, 5514, 5439, 7751, 12606, 12791, 10609]. European [766, 817, 4204, 9981]. Eurydice [11957]. Evans [980]. Evaporating [3235]. Evaporation [3564, 4934, 9961, 273]. Evaporative [2345, 2293]. Evaporator [4530, 4009]. Evens [2105]. Ever [11256, 2079]. Every [325, 4812, 7876, 11363, 12263]. Everybody [5847]. Evidence [6839, 11837]. Evil [4226]. Evolution [3659, 7685, 7505, 10567, 10684, 11248, 12413]. Ex [1856]. Ex-Commissioner [1856]. Exact [2345, 4316]. Examination [174, 3397, 9909]. Examinations [1832]. Examiners [2952]. Example [2198, 6693, 12733]. Excavating [1796, 3441, 7575, 11729]. Excavation [1356]. Excavations [5060, 7527]. Excavator [1139, 1587, 10995, 12393]. Excelsior [6713, 6655]. Excess [2715]. Exchanges [2276]. Excitements [9915]. Excreta [2616]. Excursion [983]. Excuses [7872]. Executioner [7954]. Executions [4423]. Exercising [1140, 1371]. Exeter [13127]. Exhaust [6445, 11402, 11141]. Exhauster [13065]. Exhibit [8989, 9364, 7981, 9907, 11063, 11486, 8966]. Exhibited [5575]. Exhibiting [2192]. Exhibition [1125, 396, 561, 509, 583, 493, 2420, 2297, 2967, 3475, 8387, 12090, 12032, 12631, 13143, 12836, 13340, 13401, 12759, 12337, 584, 127, 2286, 1774, 1494, 3469, 3095, 2572, 3829, 4150, 6435, 5695, 6413, 6436, 5294, 8792, 8587, 8541, 8015, 8535, 8565, 9254, 9019, 8432, 9055, 8361, 9169, 9883, 11879, 11480, 11929, 11906, 11931, 11666, 12156, 12097, 11786, 11624, 11644, 12172, 11565, 11948, 12210, 11845, 11772, 11605, 11908, 12076, 11977, 13145, 12371, 12630, 12320]. Exhibitions [12761, 13012, 6012]. Exhibitors [3860, 3861, 4717, 6840, 10062, 12070]. Exhibits [4425, 8937, 12049]. Existence [7240]. Existing [6755]. Expander [9640, 11195]. Expanding [1076]. Expansion [7749, 10876]. Expected [3483, 7409]. Expects [1099]. Expedition [6551, 6608, 9406, 10719, 11231, 13119, 13226, 13342, 1441, 2503, 2709, 4660, 4570, 5542, 7433, 9510, 9304, 9582, 10497, 10577, 11912, 11664, 12441, 12359]. Expeditions [1797, 9634, 10957, 1329, 7070]. Experience [1856, 7383, 12980]. Experiences [1848, 1868, 1891, 1933, 1958]. Experiment [3163, 7158, 8392, 10036, 10200, 4965]. Experimental [401, 2720, 6839, 7535, 7407, 7019, 1961]. Experimenting [11924]. Experiments [906, 697, 414, 227, 452, 2347, 2266, 2389, 2586, 2517, 3155, 3226, 3260, 3497, 4396, 5134, 6516, 8654, 9060, 7984, 10794, 9828, 10979, 11959, 11157, 12176, 12850, 13459, 12487, 718, 13364, 13386, 13408, 8335, 11188, 9990, 2366]. Expert [5270]. Expiring [12251]. Explained [3218]. Explains [3743].

Explode [1969]. Exploration [894, 2322, 5023, 5049, 6051, 5281, 5309, 5335, 13120]. Explorations [2416, 4424, 4231, 6035, 7633, 11159]. Explorers [4878, 7281]. Exploring [1702, 2503, 2709, 9406]. Explosion [603, 1074, 2291, 1199, 2131, 2062, 1972, 3754, 3056, 5110, 4097, 7843, 9046, 9144, 8897, 9016, 9824, 9910, 11062, 11262, 12827, 13689, 2012, 2015]. Explosions [399, 1124, 1144, 15, 1992, 2084, 1445, 3717, 4143, 4059, 4003, 6109, 5636, 5568, 5410, 7509, 7407, 9800, 12019, 11083, 7404]. Explosive [247, 790, 2858, 6378, 6303, 5442, 10032, 11948]. Explosives [5539, 7927]. Export [11497, 11996, 11776, 11663]. Exportation [9991]. Expose [2544, 2494]. Exposed [7108]. Exposition [2880, 3268, 4798, 3838, 10834, 6659, 6683, 706, 530, 43, 3724, 3038, 3367, 3746, 3626, 3482, 3342, 3740, 3499, 3743, 3594, 3652, 4425, 4318, 4793, 4490, 4659, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4859, 4806, 3809, 4799, 4545, 4573, 4601, 4343, 4607, 4642, 4801, 5090, 4399, 4001, 4456, 6067, 7809, 9268, 9365, 9110, 8347, 8624, 8650, 8680, 8709, 8737, 8767, 8796, 8865, 8898, 9300, 9044, 8920, 9907, 10310, 9576, 10743, 11063, 11352, 11015, 10980, 11608, 11694]. Expositions [3308]. Exposure [533, 514]. Extended [1857, 2252]. Extending [301]. Extension [61, 82, 1106, 1180, 134, 162, 183, 211, 261, 302, 379, 475, 521, 539, 590, 631, 655, 684, 707, 734, 777, 798, 825, 910, 955, 976, 996, 1059, 608, 679, 1232, 1283, 2136, 2178, 2232, 2257, 2280, 2303, 2332, 2357, 2377, 2401, 2426, 1309, 1387, 1418, 1500, 1526, 1557, 1583, 1606, 1653, 1680, 1706, 1729, 1760, 1784, 1836, 1862, 1885, 1906, 1928, 1953, 1982, 2028, 2051, 2118, 1612, 2147, 1947, 2123, 2000, 2071, 2132, 1391, 1832, 2454, 2476, 2506, 2670, 2676, 3570, 3372, 3538]. Extension [2713, 5228, 5874, 5379, 9460, 8906, 10627, 13484, 13379]. Extensions [2349, 1900]. Extent [3827, 4451]. Exterminating [10389]. Extermination [6944, 8198]. Exterminator [4699]. Exterminators [11294]. External [8830]. Externally [1318]. Extinct [9236]. Extinguish [2296]. Extinguisher [3266, 2881, 2530, 3947, 7156, 8273, 10695, 12209]. Extinguishing [128, 1851, 1896, 2097, 3621, 5567, 8159, 10165]. Extirpation [664]. Extortion [2082]. Extract [728]. Extracting [1826, 1773, 2737, 4068]. Extraction [13404]. Extractor [958, 1717, 6526, 7466, 9394, 9206, 12616]. Extracts [56, 2628]. Extraordinary [3396, 2517, 5357, 11788]. Eye [2392, 1662, 3511, 3946, 9438, 9855]. Eyes [50, 1368, 1290, 8478, 10409]. Eyesight [1167]. Ezekiel [4748]. Ezra [6549].

F [2823, 4004, 1119]. F.R.S [939, 959]. F.R.S. [2472]. Fabric [6112]. Fabrication [6081]. Fabrics [3582, 10967, 13687]. Facade [11906, 12210, 11845]. Facades [11948]. Faced [6754]. Faces [2249, 2698]. Facetious [440]. Facsimile [5387]. Factor [10437]. Factories [1291, 4092]. Factory [13606]. Facts [667, 55, 1394, 1272, 1666, 2571, 3510, 6577, 8804, 8990, 9272, 11541, 12644, 787]. Faience [9705, 13211, 13416, 12702, 12758]. Fail [11403]. Failings [4848].

Failure [14, 1252, 4879, 4979, 5514, 5874, 10306, 13250]. Fair [1013, 1030, 1056, 908, 928, 951, 970, 995, 1015, 1996, 2104, 2133, 2154, 2174, 2201, 2226, 2251, 2271, 1248, 2182, 3427, 3457, 3480, 3512, 3679, 3608, 4907, 4930, 4958, 6217, 6281, 6312, 6392, 6457, 7556, 7591, 7607, 7635, 7662, 7783, 9154, 12089, 12118, 12246, 13619, 12730, 559, 1008, 2072, 3369, 4492, 9255, 12213, 13456]. Fairbairn [6130]. Fairbanks [12256]. Fairlie [2218]. Fairmount [11760, 11982, 6356]. Fairport [1074]. Fairs [1052]. Fall [154, 2619, 1271, 7602]. Fallacies [2542, 6137]. Falling [4375]. False [11863]. Family [1735, 1355, 7439, 7223, 6754, 7697, 8052, 8083, 9890, 10925, 12891]. Famine [5683, 10597, 11211, 11234]. Fan [289, 3183, 4757, 10502, 11710]. Fancy [3305]. Fangs [11509]. Fans [546]. Far [7551, 13584]. Faraday [102, 374]. Fare [1947]. Farewell [8825]. Farm [3361, 4547, 4203, 5762, 7616, 6580, 10532, 11508, 12470]. Farmer [742, 4661, 5819, 8707, 12520, 3601, 6890, 12392]. Farmers [4569, 5787, 9082, 9981, 12734]. Farming [12099, 11457]. Farrell [2101]. Fashion [12895]. Fast [491, 3395, 10311, 10005, 11364]. Fastener [810, 1919, 2418, 2768, 4294, 4989, 4751, 5308, 5935, 7139, 9024, 8163, 10726, 12228, 12744]. Fastening [8682, 10778, 10504, 13023]. Fastenings [731, 2811, 2665, 7364, 9097, 10371, 11452]. Faster [4119]. Fat [6190]. Fatal [1199, 7137]. Fate [10061]. Father [3400, 8243, 8328]. Fattening [5792]. Faucet [1823, 1427, 5798, 5545, 6142, 7994, 8308, 10313, 10875, 13512]. Faucets [9058]. Favor [2577, 2605, 2630]. Fawcett [11090]. Fay [7692, 9364]. Fear [2130]. Fearful [2131, 3568, 4003]. Feat [7585, 11788]. Feathering [3863, 5903, 10150]. Feature [126]. Features [2553]. Fecundation [4963]. Fee [12461]. Feeble [661]. Feed [938, 3061, 3414, 4438, 4790, 6375, 7103, 7325, 7427, 7513, 8388, 9407, 8187, 9509, 10412, 9798, 11056, 11298, 11495, 11322, 12943, 13287, 12448]. Feeder [1169, 593, 1594, 3986, 4466, 4670, 5547, 7173, 9408, 10987]. Feeders [12363]. Feeding [3496, 5014, 5019, 4722, 10752]. Feeler [7340]. Fees [4585, 4621, 4651, 6177]. Feet [5328, 387]. Feldspar [576]. Fellah [117]. Felt [3236, 2909, 12344]. Felting [8128]. Female [880, 1832]. Females [1577]. Fence [1267, 5272, 5747, 6918, 7879, 11626, 11737, 12587, 13285]. Fences [13460]. Fender [12713]. Fennec [12748]. Ferdinand [7]. Fermentation [229]. Fern [10125]. Fernery [8521]. Ferocity [8366]. Ferries [1092]. Ferrule [6220]. Ferry [3757, 7215, 8133]. Fertilization [3951, 9935]. Fertilizer [2936]. Fertilizers [3781, 3899, 8809]. Fetta [3562]. Fever [4827, 5421, 6294, 8924, 9139, 12760, 12789, 12589, 13221]. Fevers [2889]. Few [2970, 4032, 4795, 12853, 13540, 13336, 1715, 2628]. Fiber [8384, 11095]. Fibers [4076, 4039, 7687, 6659, 6683]. Ficaria [7344]. Ficht [8872]. Field [919, 4257, 6118, 7582, 11975, 11973, 11861, 12642, 2066]. Fields [22, 7026, 9013, 12678]. Fifteen [7603, 8827]. Fifteen-Inch [7603]. Fifth [3775, 3924, 4575]. Fifty [9938]. Figure [6084]. Figures [1272, 2571, 6645, 8804, 8297, 11697]. Filamentous [8431]. File [1985, 2219, 5794, 7164, 7755, 8484, 7937, 11646, 10988, 12528, 11731]. Files [8287, 11841]. Filing [2718, 4215, 4156, 12027]. Filler [8510]. Filling [6965, 12249]. Fillings [6649]. Fills [7716]. Films [10090]. Filter [1348, 2867, 3157, 4438, 3985, 4068, 8220]. Filtering [2317, 5174]. Filters [148, 7907]. Filth [7715]. Filtration [6676, 2217]. Financial [4924]. Finches [7312]. Find [12561]. Finding [10876]. Fine [171, 9429, 12717]. Fineness [13687]. Fingers [692]. Finishing [904, 6581, 6624, 12986]. Fire [350, 116, 335, 2312, 1343, 2097, 2197, 2327, 1346, 2310, 1858, 2287, 1539, 3649, 3775, 2881, 3680, 2708, 2530, 3217, 3745, 3502, 3979, 3851, 4452, 4260, 4550, 4008, 3842, 4986, 4056, 4240, 4612, 4175, 3947, 4639, 6373, 5999, 6109, 5567, 6520, 6225, 5388, 5896, 7487, 7289, 7241, 8368, 8273, 8539, 8929, 8046, 8077, 8951, 8430, 9169, 8743, 10535, 10407, 10433, 10506, 10658, 10695, 10712, 10011, 10165, 10660, 9530, 11540, 11639, 11213, 12028, 11915, 12143, 11688, 12342, 13123, 13481, 13370, 12471, 13589, 13410, 13039, 12817, 12558, 7419, 12124]. Fire-Extinguishing [10165]. Fire-Flies [13123]. Fire-Fly [7419]. Fire-Grate [4240]. Firearm [12395]. Firearms [2664, 2712, 2969, 8957, 8998, 9018]. Firefly [4987]. Fireless [3241, 4954, 5407, 5926, 7181, 10682]. Fireman [6844]. Fireplace [1095, 5412, 12397]. Fireproof [2298, 2220, 2246, 2228, 2442, 2907, 3726, 4461, 4285, 5092, 6953, 9370, 11112, 12096]. Fires [128, 1851, 1896, 2222, 2296, 3689, 4147, 8159, 9498, 9982]. Firework [11591]. Firing [1573, 1006]. Firkin [4349]. Firm [7643]. Firmenich [11633]. Firmin [10926, 12066]. First [153, 2473, 3529, 4293, 4757, 8962, 11812, 11327, 12332, 2880]. Fish [42, 3164, 3082, 5231, 4777, 4569, 5494, 5549, 7646, 7695, 6917, 9477, 8601, 8140, 8430, 8602, 10542, 10037, 9536, 10033, 11209, 11741, 13618, 13489, 12265, 13083]. Fisher [1078, 1856, 1832, 11068]. Fisheries [1215, 12578, 13507, 402]. Fishery [386, 7096]. Fishes [8601, 8717, 10780]. Fishhooks [528, 9246]. Fishing [1685, 3413, 4724, 12594]. Fisk [2526]. Fiske [20]. Fitter [13610]. Fitters [13336]. Five [3675, 3308, 5616, 13530]. Fixation [11691]. Fixed [307]. Fixing [514]. Fixture [7721, 10257]. Flag [2590]. Flagging [2274]. Flake [6845]. Flame [3105, 7640, 11870]. Flames [6242, 9694]. Flamingo [7620]. Flange [10776, 13335]. Flanger [12964]. Flanging [7254]. Flaring [2013]. Flask [3088, 3998]. Flasks [7757]. Flat [1373, 2044, 7637]. Flavoring [4383]. Flax [5166]. Flesh [6773]. Fleury [7588]. Flexible [3590, 4215]. Flies [13224, 13123]. Flight [2422, 2747, 9828, 1510]. Flights [9478]. Flint [4525, 13134, 13121]. Float [6824, 8518]. Floating [3470, 4008, 4829, 5411, 6520, 7487, 7641, 6652, 10420, 10753, 12463, 12560]. Floato [722]. Floato-Static [722]. Flood [156, 5541, 8460]. Floods [6054]. Floodway [7525, 7848]. Floor [438, 4461, 4643, 6252, 7390, 7416]. Flooring [169]. Floors [9279]. Flora [11599]. Florakin [7590]. Floral [732, 5216]. Florence [6034, 9922]. Florida [3425, 13495]. Floss [12469]. Floue [9308]. Flour [1638, 3422, 8254, 13434, 12888, 13157]. Flouring [10829]. Flow [10531, 12044, 5750]. Flower [2916, 6034, 7700, 8965, 8511, 9507, 10324, 10420, 9695, 9675, 10886, 12999, 875].

Flowered [9251]. Flowering [8659]. Flowers [65, 4963, 11894]. Flowing [1012, 4769]. Flues [3853, 11085]. Fluids [790, 1300, 2599, 3327, 5460, 6921]. Fluor [449, 624]. Fluorescence [4410]. Fluosilicic [8383]. Flush [3355]. Fluting [988, 1373, 6802, 13209]. Fly [663, 4316, 4300, 5900, 6895, 7083, 10502, 10323, 13689, 10864, 7419]. Flying [75, 808, 1794, 1808, 1690, 2383, 2410, 2917, 6741, 9317, 9463, 10669, 9588, 13235, 12747, 13518]. Fob [3261]. Fodder [9207, 10440]. Foes [8707, 9563]. Fog [2081, 2459, 5670, 8368, 10444]. Foil [2973]. Folding [2386, 2868, 4125, 8277, 10862, 11112, 11843]. Folding-Door [11112]. Followed [1942]. Folly [1942, 13531]. Folsom [6141]. Font [5547, 13472]. Food [502, 483, 1614, 4383, 4464, 6270, 6057, 5631, 6773, 9398, 10500, 9641, 12519]. Foot [3295, 3941, 6105, 12101, 12230, 12316]. Footed [5922]. Forage [6190]. Force [1966, 2269, 2266, 2070, 2103, 2340, 2698, 3047, 3063, 4510, 6270, 5610, 5543, 5353, 7216, 7902, 7925, 8281, 9665, 11563, 8028, 1961]. Forceps [12714]. Forces [78, 455, 6599]. Foreign [748, 905, 1160, 36, 59, 653, 682, 709, 736, 754, 779, 797, 823, 849, 866, 887, 912, 933, 953, 974, 998, 1018, 1039, 1060, 1084, 1109, 1131, 1153, 1178, 81, 105, 132, 164, 181, 208, 234, 259, 284, 304, 329, 355, 378, 407, 431, 459, 476, 497, 519, 541, 565, 588, 609, 629, 557, 2184, 1279, 1411, 1459, 1204, 1233, 1883, 1908, 1926, 1980, 2004, 2026, 2050, 2075, 2093, 2116, 2137, 2158, 2180, 2231, 2255, 2304, 2330, 2355, 2376, 2404, 1257, 1282]. Foreign [1310, 1337, 1364, 1386, 1420, 1448, 1472, 1501, 1525, 1558, 1581, 1605, 1628, 1652, 1677, 1704, 1730, 1758, 1783, 1811, 1837, 1861, 3319, 3346, 3376, 3408, 2452, 2478, 3073, 3098, 3118, 3143, 3169, 3200, 3228, 3252, 3287, 3317, 3345, 3374, 3405, 3433, 3460, 3487, 3514, 3549, 3575, 3611, 3635, 3666, 3695, 3730, 3762, 3794, 2508, 2534, 2557, 2579, 2607, 2633, 2659, 2678, 2701, 2728, 2750, 2776, 2802, 2828, 2852, 2872, 2893, 2911, 2930, 2955, 2979, 3000, 3027, 3049, 4810, 3820, 3845, 4554, 4582]. Foreign [4617, 4647, 4681, 4707, 4733, 4759, 4782, 4813, 4838, 4865, 4894, 4916, 4943, 4968, 5000, 5024, 5074, 5097, 5125, 5149, 5180, 5203, 5232, 3873, 3904, 3932, 3962, 3990, 4019, 4048, 4082, 4107, 4135, 4162, 4190, 4217, 4248, 4274, 4305, 4330, 4359, 4387, 4412, 4441, 4474, 4503, 4532, 5257, 5284, 5966, 5990, 6041, 6070, 6096, 6121, 6147, 6174, 6200, 6232, 6259, 6284, 6315, 6341, 6366, 6393, 6414, 6437, 6458, 6482, 6506, 6534, 6565, 5313, 5337, 5365, 5395, 5426, 5453, 5476, 5503, 5529, 5559]. Foreign [5588, 5621, 5643, 5672, 5697, 5724, 5753, 5777, 5805, 5833, 5859, 5886, 5911, 5940, 6840, 7660, 7717, 6755, 6588, 6612, 7247, 7273, 7293, 7317, 7348, 7372, 7395, 7420, 7443, 7472, 7495, 7518, 7543, 7568, 7593, 7623, 7649, 7676, 7677, 7702, 7734, 7766, 7795, 7828, 7856, 7886, 6633, 7914, 6661, 6685, 6706, 6729, 6758, 6783, 6808, 6832, 6857, 6883, 6900, 6926, 6955, 6984, 7009, 7036, 7059, 7086, 7118, 7142, 7167, 7197, 7225, 7733, 8937, 7941, 7969, 8694, 8721, 8751, 8781, 8814, 8847, 8879, 8910]. Foreign [8938, 8969, 9001, 9030, 9062, 9098, 9124, 9157, 9187, 9221, 9256, 9288, 9321, 9350, 9381, 9412, 9444, 9481, 7999, 9514, 8030, 8056, 8087, 8113, 8142, 8170, 8201, 8229, 8259, 8289, 8311, 8338, 8370, 8403, 8433, 8463, 8493, 8523, 8547, 8575, 8606, 8634, 8663, 7968, 10526, 9542, 9567, 10265, 10292, 10327, 10359, 10391, 10421, 10452, 10484, 10514, 10546, 10581, 10617, 10645, 10671, 10700, 10731, 10756, 10782, 10809, 10839, 9596, 9623, 9654, 9680, 9708, 9734, 9758, 9786, 9838, 9865, 9894, 9924, 9943, 9972, 9994]. Foreign [10016, 10046, 10070, 10099, 10128, 10155, 10179, 10209, 10237, 11751, 11497, 12100, 1951, 2281, 5051, 6016]. Foreigners [3094]. Foremanizing [1148]. Foremen [157, 1170]. Forest [6821, 12435, 292]. Forests [2270, 12556]. Forge [6335, 11552]. Forgetfulness [11006]. Forging [4703, 10634, 10439, 12590]. Forgings [1023]. Fork [669, 4888, 4355, 8778, 9207, 10533, 10092, 13151, 13514]. Forks [12151]. Form [3150, 4070, 4602, 7211, 8521, 12554, 8393, 4180, 7901, 8191, 9528, 13315, 13361, 13341, 12396]. Formation [4451, 6050, 9505, 11584, 3338]. Formations [93, 206]. Formed [1410, 7747]. Former [2571, 5546, 13379]. Formic [5787]. Forms [8876]. Formulas [4316]. Forney [12144]. Forster [10926, 12066]. Fortifications [944, 893]. Fortunes [10745]. Forty [4995, 6005]. Forty-Four [6005]. Fossil [4961]. Fossils [440]. Foster [669, 2262]. Foul [4094]. Found [1353, 551]. Foundation [13586]. Foundations [3110, 4487]. Founder [5115, 6412, 5403]. Founderies [3454]. Founders [2082, 6183]. Fountain [742, 1911, 5227, 4531, 4657, 8658, 8899, 11325, 11777, 12706, 5799, 5957]. Fountains [10095, 10231]. Four [560, 4781, 6089, 5922, 6005, 6560, 13279]. Four-Antlered [6089]. Four-Cutter [6560]. Fourth [453, 3889, 12383]. Fowl [13447, 11123, 10667]. Fowler [1893, 1845, 12054]. Fowling [4519]. Fox [12748, 7538, 13518]. Fracture [6056, 5575]. Fragments [1932]. Frame [5773, 9142, 10887, 9559, 11532]. Framed [6562]. Frames [3559, 2819]. France [784, 922, 1100, 807, 1092, 786, 1970, 5008, 6693, 6773, 10560, 11907, 11160, 13225, 13275, 13161]. Francis [5500, 3484]. Francisco [3223]. Franco [882]. Franklin [2547, 6413, 6436]. Frazer [3543]. Free [748, 2518, 3978, 8418, 12731]. Freeland [899]. Freezer [5613, 9533]. Freezing [3986, 6323, 5333, 6824, 13481]. Freight [4491, 10311]. Freights [6552]. French [366, 176, 227, 144, 23, 27, 43, 744, 1262, 3504, 4781, 4659, 6556, 5736, 6476, 6131, 8563, 9399, 10310, 9576, 10834, 9884, 11409, 12090, 11752, 12180, 12154, 13416, 13434, 13143, 12613, 876]. Frere [1911]. Frescoes [3037]. Fresh [1055, 5231, 3803, 7562, 7240, 10033]. Friction [702, 618, 671, 2057, 2793, 2490, 5220, 3832, 4631, 4804, 5828, 7666, 7584, 9113, 9970, 10085, 9775, 11564, 12800]. Frictional [2788, 2817, 2840, 2861]. Friedmann [13553]. Friend [4005, 12520]. Friends [5326, 6695, 8327]. Fringed [10968]. Frog [10124, 12747, 10509]. Front [1692, 3240, 7050, 9853]. Fronts [7004]. Frost [5070, 12760, 10202]. Frozen [1852, 3481, 6866, 7363]. Fruit [579, 1710, 4355, 5043, 4990, 5845, 10526, 13358, 12361, 12493, 13122]. Fruits [1055]. Fuel [227, 645, 1442, 2501, 2504, 3971, 3830, 4905, 4982, 5161, 4880, 5082, 5191, 4541, 4173, 5473, 5689, 6474, 5843, 10912, 10082, 11582, 11032, 12072, 12186].

Fuels [11681, 12145]. Fuller [7192]. Fulling [7065, 11042]. Fully [748]. Fumigator [10123]. Fundamental [3352]. Funds [11422]. Funeral [6550]. Fungi [13249]. Fungoid [13673]. Fungus [5787, 13381]. Furnace [2367, 3839, 5047, 5694, 6272, 5361, 5986, 5981, 7514, 9408, 7964, 8710, 8094, 10741, 10670, 12287, 11038, 11619, 12170, 11642, 12339, 12655]. Furnaces [124, 4268, 4853, 6219, 7231, 7289, 8905, 8867, 1006]. Furness [6412, 5280]. Furniture [482, 488, 5248, 8390, 10989]. Furrowing [1570, 485]. Further [2266, 11837]. Fuses [8287]. Fusible [8398]. Fusing [12202]. Future [2245, 2444, 2601, 6471, 6496, 5897, 10436, 9745, 9634, 13163, 13691, 13380, 13532].

G [12608]. G. [1639]. Gage [2083, 1440, 1971, 4889, 4940, 4214, 7435, 9404, 8834]. Gages [1833, 1298, 2651, 3008, 6804, 7583]. Gail [5349]. Gain [6324, 6375]. Gaining [5418]. Gall [191]. Gallahue [8704]. Galleries [1304, 6491]. Gallery [4238]. Galleys [7905]. Gallium [11377]. Galls [6033]. Galmann [368]. Galton [5412, 8043, 8074]. Galvanic [6360, 7107]. Galvano [8025]. Galvano-Electric [8025]. Galvanometer [3011, 9468]. Game [11167]. Gampert [11734]. Gang [200, 5980, 5881, 7269, 9534, 11931, 11653]. Gap [9956]. Gaps [7747]. Garbage [12655]. Garden [4460, 6118, 6479, 8358, 11311, 12229, 11760, 11982, 12830, 13457, 6390]. Gardener [2901]. Gardening [5554]. Gardens [90, 4157, 5722, 7165, 8026, 5883, 7854]. Gardner [2881, 4637, 5693]. Garments [8828]. Gart [12555]. Gartsherrie [6049]. Gary [12668, 12563, 12551, 12671]. Gas [859, 645, 1168, 670, 110, 121, 746, 203, 345, 199, 1742, 1275, 1189, 1342, 1780, 2391, 2388, 1822, 2055, 1239, 2343, 3658, 3789, 2542, 3060, 5093, 4905, 4604, 5120, 4880, 4633, 3913, 4454, 4875, 5068, 5061, 6373, 6428, 6165, 7183, 7554, 7022, 7638, 7186, 7263, 6627, 7642, 7123, 7810, 9057, 8625, 8874, 8099, 8041, 8072, 8164, 8711, 9245, 8333, 8867, 10568, 9967, 10772, 11199, 11679, 11693, 12067, 12072, 11559, 11371, 12261, 12123, 11291, 12253, 11718, 12576, 13606, 13065, 13108, 12487, 45, 9348, 12186]. Gas-Propelled [4454]. Gaseous [11150, 12314]. Gases [1076, 298, 5011, 11082, 11157, 11214, 2440]. Gasket [10965]. Gasoline [1199]. Gaspipe [10286]. Gasteropods [12945]. Gate [1297, 2057, 2409, 3361, 4913, 4547, 8460, 11508, 1623, 1658, 2784, 3897, 4241, 4845, 5060, 7527, 7402, 9046, 9172, 9121, 9144]. Gates [8168, 9052]. Gathering [6876, 11937, 11243]. Gatling [968, 3109, 2664, 5373, 8926, 11345, 12961]. Gauge [10722, 10779, 10205, 10229, 12175]. Gauges [10149]. Gazette [3182]. Gear [941, 1692, 4775, 5998, 5273, 5606, 6680, 9097, 9377, 8572, 8253, 10315, 11611, 11652, 11016, 13286]. Gear-Cutting [8253]. Gear-Dressing [11652, 13286]. Geared [5220, 7133, 10696, 11564]. Gearing [2793, 2490, 4155, 9143, 9595, 2788, 2817, 2840, 2861]. Gears [3922]. Gecmen [1689]. Geissler [12673]. Gelatin [4488]. Gelatine [13329]. Gem [5978]. General [1454, 1889, 2271, 2570, 6161, 6587, 1967]. Generating [9309]. Generation [793, 4744, 6767, 11275]. Generator [1368, 2055, 5949, 5597, 11370, 11045, 11718, 13432, 13583, 13561, 13533]. Genesis [5600, 7044, 11207]. Geneva [11971]. Genius [11922, 12419]. Gentlemen [443]. Geographical [4319, 5223, 8487, 9578, 11230]. Geologic [6084]. Geological [1407, 2766, 4745, 6275, 5816, 6091, 7633, 10852, 12438]. Geology [10253, 1959]. Geometrical [2623, 8605, 7934, 9527, 10447]. George [173, 2102]. Georgia [559, 3527, 3608]. Geowth [13673]. Geranium [8197]. Gerard [12921]. Gerhart [2622]. Germ [5142, 5177, 5201]. German [6090, 6937, 9237, 10141, 10609]. Germany [3451, 3746, 2737, 6908, 11639, 13149]. Germinating [10511]. Germination [6658]. Get [8794]. Geysers [13104]. Ghent [12393]. Giant [7849, 10513, 9804, 13368, 12624, 13260, 11748]. Giants [1355]. Gibbon [10697]. Gibbs [2842]. Giffard [3302, 6945, 7667, 11149]. Gift [1464, 4206, 7871, 10956]. Gigantic [2221, 1355, 2972, 3842, 5511, 7215, 11595]. Gila [13679]. Gilberds [5522]. Gilbert [11080, 11454, 11254]. Gillott [5005]. Gin [3153, 4156, 7173, 140]. Gipsy [1963]. Girard [12]. Girders [456, 2816, 4643]. Girdlers [1871]. Glacial [3177, 5710, 7304]. Glance [674]. Glandular [10825]. Glasgow [3357]. Glass [204, 325, 443, 741, 940, 771, 2396, 2212, 1325, 2323, 3388, 3584, 3953, 6521, 5465, 7210, 7214, 7128, 7259, 8862, 9300, 10166, 9722, 9748, 9772, 9879, 10448, 10564, 9557, 11592, 12224, 10984, 12610, 12658, 13621]. Glass-Makers [443]. Glasses [6578, 9589, 217, 242, 272]. Glassmaking [9774]. Glassware [12267, 13651, 13542]. Glatton [3184]. Glazed [370]. Gleanings [3866, 3890, 3926]. Gleason [874]. Glen [4159]. Glimpses [10307]. Glissant [12]. Globe [4940]. Glossing [2244, 2916]. Glottidis [9652]. Glove [11806]. Gloves [10419, 12021, 68]. Glucose [12409]. Glue [6585, 10545, 10288]. Glue-Heating [10288]. Gluten [10450]. Glycerin [1074, 877, 772, 1801, 7360, 9186]. Goal [7685]. Goat [11698, 13259]. Goblet [10358]. Godefroy [10087]. Godfrey [11038]. Goethe [9553]. Going [807, 4667]. Gold [789, 1826, 1968, 2171, 1404, 3356, 7026, 9437, 8505, 11255, 11555, 11704, 13633, 3043, 4351]. Golden [12031]. Gone [4005]. Gongs [2985]. Good [750, 1123, 2631, 7876, 9746, 11610, 12734]. Goodridge [3266]. Goods [3359, 10116, 12033]. Goose [11065]. Gordon [3024]. Gorilla [9633, 10478]. Gosa [13069]. Gothard [2393]. Gothic [13346]. Gourd [7112, 8399]. Goux [2437]. Govern [546]. Government [878, 1803, 3397, 3787, 2784, 5065, 3883, 3897, 4241, 3828, 5921, 6546, 11033, 11061, 1439, 3703]. Governor [1073, 2717, 4637, 5119, 4043, 6373, 5651, 6196, 6037, 6825, 7030, 7815, 8560, 8841, 8907, 9403, 10507, 11707, 11041, 12178, 13065, 12625]. Governors [5764, 6525, 6558]. Grab [3705]. Grader [9905]. Gradients [5849, 7814]. Grafting [2844, 13674]. Grahamite [2045]. Grain [1867, 2263, 3005, 2878, 2487, 4394, 4104, 6079, 7868, 8427, 9916, 10347, 10176, 9673, 11172, 13378, 12888, 12364]. Graining [1396, 1478]. Grammar [750]. Grampus [13317]. Granary [6025]. Grand [316, 2393, 2880, 2674, 4836, 4812, 7897, 7206, 7186, 12118, 13143, 3308].

Grande [8506]. Grange [7616]. Granges [4768]. Granite [240, 2471, 12134]. Grant [3070, 10092]. Granular [7362]. Grape [513, 1698, 6162]. Grapes [3803, 7723, 8809]. Graphic [4844, 10803]. Graphite [1897, 3048, 3957]. Grapnel [6253, 7691]. Grappling [5494, 5905]. Grasshopper [9964, 7154, 8538, 10578]. Grasshoppers [785]. Grate [1049, 116, 421, 1368, 3561, 4240, 5638, 6726, 7964, 8685, 10939, 11989]. Gratuitous [4032]. Gravel [9021]. Graves’ [10319]. Gravestones [10168]. Gravitation [5931, 8990]. Gravity [4605, 11106]. Gray [12776, 9969]. Grayhound [6951]. Grease [8365]. Great [1147, 1047, 453, 917, 635, 2321, 2188, 1465, 2420, 2197, 2222, 1779, 1900, 3394, 2971, 2486, 2821, 2469, 4769, 4646, 5073, 5084, 4606, 3840, 3886, 3800, 4692, 5999, 6244, 5541, 5602, 5516, 5387, 7585, 7126, 7380, 7617, 6609, 7177, 8587, 9461, 8476, 9044, 8097, 9399, 10282, 10279, 10344, 9911, 10341, 10878, 11929, 12294, 11057, 13660, 12775, 13606, 12414, 12447, 12540, 12759, 13495, 13083, 340, 1659, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4799, 4607, 4642, 4801, 7809]. Greatest [4633]. Greatly [3364]. Grecian [8862]. Greece [895, 8764]. Greek [2211, 4177, 12656, 12917]. Greeley [3711, 3747, 3902]. Green [12501, 2382, 6776, 9013, 10855, 12595, 12410, 190]. Greenbacks [3415, 3445, 5413]. Greenhouse [2711]. Greenhouses [2438, 6358, 7912]. Greenland [13317]. Greeting [1842]. Grid [12822]. Gridiron [365]. Grievance [2082]. Grinder [3686, 3140, 4498, 4170, 4577, 5484, 7510, 8807, 8329, 9987, 11587]. Grinders [9561]. Grinding [741, 1214, 3958, 5552, 7671, 9424, 9345, 8051, 8082, 10858, 10639, 10773, 10805, 10859, 11753, 12183, 11512, 13179]. Grindstone [9316]. Grip [7417]. Grist [2782]. Griswold [3604, 3306]. Grocer [4127]. Groof [6383]. Grooming [8681]. Grooved [1431]. Grooves [10985]. Grooving [5143, 6410, 7714]. Gross [816]. Grosvenor [3362]. Ground [3868, 9611]. Grounds [5846]. Group [1494]. Grouse [7223]. Grove [2472, 2474]. Growing [7006]. Grown [4989]. Growth [52, 3659, 3247, 8487]. Grubbing [6605]. Gruey [12374]. Guano [2921, 8954]. Guard [1265, 2762, 5388, 11068, 12057, 12342, 12528]. Guardiola [3006, 9670]. Guards [198, 12117]. Guatemala [12582, 218]. Guava [337]. Guelder [7966]. Guernsey [6310]. Guess [1776]. Guide [3921, 5612, 7937, 8567, 10838, 10988]. Guillotined [414]. Guinea [11983]. Gulf [6055, 7779, 2715]. Gull [9183, 10230]. Gum [2995, 12906]. Gummer [5496, 7159]. Gumming [293, 8808, 13184]. Gun [697, 968, 942, 917, 744, 1218, 1406, 2009, 2108, 2062, 2899, 3109, 3501, 3703, 4198, 3987, 5373, 6476, 7221, 7389, 7884, 8926, 8903, 10908, 9779, 10962, 11345, 13696, 12961, 12525, 13252]. Gunboat [7458]. Gunboats [187, 4]. Gunmaking [2758]. Gunnery [4981]. Gunpowder [1969, 3008, 7509, 9216, 12312]. Guns [1986, 6603, 7601, 7126, 7909, 10032, 11061, 12774]. Gurnard [8431]. Gutta [892]. Gutta-Percha [892]. Gutter [12586]. Gymnast [5229]. Gynomania [9593]. Gyroscope [4043, 12374, 8393, 11523]. Gyroscopic [12374].

H [939, 959, 2127, 1989, 3498, 3093, 3116, 3165, 3197, 3224, 3249, 3285, 3315, 3343, 3403, 3419, 3539, 3428, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4936, 4967, 5042, 4607, 4642, 10094, 12307]. H. [37, 353, 376, 434]. Habit [1327, 13647]. Habitations [9942]. Habits [595, 2085]. Hadley [3662]. Haeckel [8257, 12413]. Hail [1213, 9505]. Hail-Storm [1213]. Hair [719, 1712, 1818, 3059, 2863, 4123, 9343, 11983, 12313]. Haired [10418]. Hairy [2309, 6754, 12030, 12264]. Hale [7761]. Half [3705, 4511, 13276]. Hall [794, 154, 2390, 4891, 5115, 8432]. Hallidie [2739]. Halls [9579]. Halter [10608]. Halyards [2590]. Hamilton [5745, 364, 5019, 5707, 6422]. Hammer [1570, 4130, 4993, 6272, 5828, 6335, 5855, 6975, 8836, 10960, 12942]. Hammerless [13696]. Hammers [215]. Hampton [7854]. Hancock [12466]. Hand [985, 964, 27, 1518, 1402, 1846, 3295, 4887, 4405, 6193, 6117, 5789, 7266, 7789, 7141, 7218, 9312, 8951, 9509, 8225, 8860, 9475, 10751, 10014, 10746, 11839, 12248, 12496]. Hand-Drilling [964]. Handed [8734]. Handiwork [178]. Handle [532, 4430, 7138, 9051]. Handled [753]. Handles [3300, 4078]. Handling [6921]. Hands [3125]. Hanger [4523, 7464, 11140]. Hanging [1872, 3541]. Hangings [3139, 9052]. Hannahs [1634]. Happy [13]. Haptonstall [8484]. Harbor [5040, 6671, 11892, 12436]. Harbors [8904]. Harcourt [2212]. Hard [980, 3388, 5083, 7657, 7841, 11859, 12449, 8774, 1589]. Hardened [10564]. Hardening [277, 152, 13675]. Harding [1890, 1914, 8107]. Hardware [2951, 5322]. Hardy [7392]. Hare [9563]. Hark [12153]. Harlem [3269, 3895]. Harlow [4891]. Harmony [8775]. Harness [1097, 3389, 10721, 9731]. Harnessed [12847]. Harpy [9756]. Harrington [7221]. Harris [764, 4751, 5522, 8892, 13326]. Harrison [10829]. Harrow [924, 5797, 11299]. Harrows [110]. Hartford [5293]. Hartz [6644]. Harvard [884, 20, 6053]. Harvest [7928, 11259]. Harvester [11118]. Harvesters [4394, 4717]. Harvesting [2618, 2593, 3714]. Harvey [1200, 10196]. Hassler [2322]. Haste [9401]. Hastie [12001]. Hat-Making [8019]. Hatchet [4610, 4183, 6455, 5774]. Hatchway [6852]. Hatfield [12608]. Hats [3236, 12021, 12344]. Hauling [10468]. Haupt [2522]. Havana [4159]. Haven [4873]. Havre [5189]. Hawaii [10282]. Hawks [8140]. Hay [870, 3650, 4937, 4827, 5960, 7242, 8778, 8835, 8223, 8924, 10092, 12963]. Hayden [6752]. Hayford [11212]. Hayfork [1436]. Haymaking [13613]. Haze [256]. Head [414, 1874, 3720, 3620, 4349, 6089, 5391, 7084, 8280, 9310, 10635, 11453, 11934, 11673]. Headed [11847]. Heading [2032, 4339, 6991]. Heads [5903, 12698]. Health [89, 2275, 1645, 1625, 2499, 3524, 2822, 2866, 5124, 4201, 4435, 5819, 7564, 7813, 8794, 9480, 10290, 13581, 13608, 1091, 203, 1438]. Healthy [10856]. Hear [10916, 13584]. Heard [2814]. Hearing [4211]. Heart [9614]. Hearth [6225]. Heat [578, 325, 514, 324, 815, 1852, 2368, 2340, 1315, 3455, 2845, 2741, 2965, 2986, 2707, 2696, 2725, 3624, 2903, 4370, 3913, 4658, 4903, 6405, 5741, 7779, 7762, 6769, 8705, 10798, 10600, 10254, 11596, 11274, 10977, 13482, 6114, 2740, 2417, 2385, 2364, 2464, 2493, 2440, 11233]. Heat-Distillation [2845]. Heated [4459, 8826]. Heater [3061, 2438, 3054, 4790, 4068, 7325, 7427, 7694, 7644, 7513, 8388, 9407, 9245, 10795, 11056, 11495, 11044]. Heaters [6375]. Heating [988, 1122, 124, 2575, 2624, 2548, 2858, 5199, 4644, 5420, 6445, 6447, 7024, 7469, 7022, 6802, 9120, 8905, 10288, 9668, 11989, 13227, 13560, 13662]. Heats [3114]. Heavy [1023, 2163, 1986, 1507, 2748, 7899]. Heberlein [5441]. Hecla [11539]. Hedge [11697]. Heel [4611, 6278, 10940, 12443]. Heer [8483]. Heinrich [11110, 12673]. Heirs [695]. Held [10035]. Helicoidal [4996]. Helioscope [11236]. Hell [1623, 1658, 2409, 2784, 3897, 4241, 4845, 5060, 7527, 7402, 9046, 9172, 9121, 9144]. Hellgrammite [10177]. Helped [11969]. Helping [10169]. Hematite [7785]. Hemi [9725]. Hemi-Plunger [9725]. Hemiptera [9283]. Hemlock [3209]. Hemmer [1612]. Hencoop [2386]. Henderson [3831]. Hennaman [7614]. Henri [11149]. Henry [1185, 3456, 5348, 8605, 1825, 1306, 5301, 9333, 11525]. Her [9563]. Herald [3368, 11258]. Herculean [4542]. Hercules [7436, 13665]. Here [11951]. Hereditary [10437]. Heredity [8043, 8074]. Hermetical [12093]. Hermit [12185]. Heroes [11950]. Herr [2219]. Herreshoff [12674]. Herschel [1817, 5463]. Hertfordshire [7290]. Heslop [12709]. Hesperians [693]. Hewitt [4751]. Hickman [985]. Hide [9116]. High [1117, 268, 1322, 1995, 2059, 2741, 3627, 2525, 2640, 2696, 4059, 4214, 5328, 5871, 8099, 10041, 11579, 11120, 11488]. High-Level [11579]. Highgate [2543]. Highland [6682]. Hight [847, 7104]. Highton [2112]. Highways [8042, 8073]. Higley [12984]. Hildreth [1938]. Hilf [5469]. Hill [10059, 363, 4950, 5573, 8836, 10862]. Hills [6765]. Hillside [6625]. Himmer [2744]. Hinge [1745, 2057, 3355, 5880, 7819, 10643]. Hinges [8168, 11209]. Hint [3319, 3346, 3376, 3408, 7846, 12415]. Hints [2319, 1299, 1864, 1886, 2179, 3358, 3318, 3377, 3406, 3434, 3461, 3488, 3516, 3550, 3576, 3612, 3636, 2952, 4032, 5216, 4795, 4618, 4648, 5415, 6245, 6402, 7195, 7161, 8794, 11355, 13610]. Hippopotamu [5165]. Hippopotamus [9644]. Hire [1745]. Hirn [803]. Historical [2520]. History [944, 965, 675, 893, 926, 505, 1136, 512, 529, 1495, 1493, 2202, 3109, 2515, 4293, 6340, 6364, 7258, 9880, 10390, 10415, 10451, 11210, 11028, 12305, 12749, 12862, 12998, 13133, 13160, 13261, 13415, 13571, 13617, 13648, 2786, 2813, 340, 3154, 9435, 10867, 10911, 467, 2421, 5848]. Hitch [221]. Hitching [1661, 1244]. Hixson [4160]. Hoax [516]. Hoboken [2011, 6039]. Hod [1072, 8800]. Hoe [669, 4460, 9282, 12990, 4822]. Hoffman [10538]. Hog [12394, 12285]. Hogs [10541]. Hoisting [6, 2382, 1567, 5140, 5973, 5829, 6919, 10085, 10570, 10966, 11145, 11338]. Holborn [5058]. HolbrooK [7819]. Hold [2648]. Hold-Back [2648]. Holden [11228, 6065]. Holder [573, 1162, 614, 2102, 1245, 2106, 3181, 2919, 4073, 4672, 4933, 4126, 5275, 5748, 5615, 5524, 6478, 7367, 7883, 7821, 7110, 7338, 8960, 8256, 10261, 10833, 11868, 13076, 13678]. Holders [9470, 12337]. Holding [1592, 9988]. Hole [325, 198, 1935, 1269, 3750, 7113, 7053]. Holes [1917, 12129]. Holidays [7846]. Holland [784]. Hollow [147, 2265].

Holloway [13389]. Holly [2327, 11796, 13227]. Holmes [6692]. Holtz [2791]. Holyoke [13579]. Home [317, 315, 3417, 7669, 6936, 8688, 12696, 13223]. Home-Made [317, 7669, 8688]. Homes [2169, 3560, 4354, 7193, 8652, 8706, 10856, 12778]. Homesteads [2672]. Hon. [1252, 1832, 1332, 2145, 3604, 5736]. Honesty [10499]. Honey [1037, 9096, 12409, 12580, 12616, 12561, 4855]. Honey-Making [4855]. Honor [2646, 2674, 4057, 12516]. Honors [12091]. Hood [7138]. Hook [2292, 1537, 3502, 4379, 6220, 6997, 9724, 12499, 863]. Hoop [7190, 9242, 13510]. Hoosac [1011, 175, 1492, 5168]. Hoosic [3774]. Hop [10534]. Hops [7844]. Horace [3711, 3747, 3902]. Horizon [8508]. Horizontal [1530, 3911, 6248, 8020, 10907, 11666, 11104, 11535, 13434]. Horn [5670, 11381]. Hornbill [11785]. Horned [11071, 11173]. Hornwort [10642]. Horograph [12973]. Horse [675, 318, 2243, 1661, 1436, 1637, 3153, 3568, 3596, 3655, 3059, 2614, 4369, 4609, 3867, 5600, 5844, 6773, 6606, 7242, 7306, 7153, 6799, 7504, 8389, 8681, 8021, 10092, 10320, 10173, 10285, 11843, 11037, 13384, 13640, 13662, 2213, 12024]. Horse-Power [1436]. Horsenail [3619]. Horses [221, 438, 3303, 3129, 6477, 5742, 6167, 7528, 7334, 7055, 8622, 11715]. Horseshoe [1118, 1857, 4187, 5550, 6432, 5772, 8593, 12391]. Horseshoes [556, 7967, 8802]. Horticultural [8687, 8432]. Horticulture [3781, 7114, 7673]. Horton [5607]. Hose [1343, 3134, 3421, 9058, 8951, 8428, 9458, 11269, 13064, 13670]. Hosmer [12169, 12331, 12672]. Hospital [6602, 7665, 3960]. Hospitals [273]. Hot [1766, 2694, 2438, 3722, 3282, 4091, 5894, 6504, 6194, 7024, 7231, 8710, 8828, 10228, 11619, 13663, 921]. Hot-Air [10228]. Hot-Pressing [4091]. Hotbed [3559]. Hotchkiss [5266, 5828]. Hotel [615, 3775, 9982, 11316, 2038]. Hothouses [7912]. Hotte [574]. Hotz [5798]. Hour [3138, 3018, 2994, 3046, 3064, 3091, 5214, 4093, 6005, 11683]. Hours [2172, 13279]. House [1123, 316, 1381, 3153, 2547, 3748, 4079, 6323, 5554, 5334, 6447, 7706, 7797, 7830, 7889, 7918, 6953, 9441, 8001, 8032, 9066, 9100, 9128, 9160, 9188, 9223, 9260, 9291, 9322, 9353, 9414, 9447, 8114, 8145, 8173, 8202, 8232, 8262, 8313, 8371, 8405, 8434, 8465, 8526, 8550, 8578, 8609, 8635, 8666, 8723, 8816, 8849, 8883, 8912, 8970, 9032, 9569, 9658, 10548, 10586, 10647, 10674, 10701, 10758, 10786, 9684, 9709, 9759, 9791, 9816, 9867, 9897, 9926, 9948, 9996, 10018, 10049, 10131, 10159, 10184, 10210, 10240, 10268, 10295]. House [10330, 10362, 10394, 10425, 10456, 11396, 12905, 13595, 12484, 5499, 11694, 12552]. House-Builder [13595]. Household [4407, 7232, 7195, 7161, 7705, 7737, 7767, 7799, 7831, 7857, 7888, 7915, 8168, 7943, 7971, 8003, 8060, 8091, 8115, 8233, 8261, 8315, 8374, 8407, 8436, 8495, 8524, 8549, 8579, 8608, 8668, 8636, 8722, 8785, 8850, 8882, 8939, 9002, 9033, 9064, 9102, 9127, 9158, 9189, 9224, 9259, 9290, 9323, 9351, 9385, 9413, 9445, 9485, 9517, 9571, 9599, 10733, 10759, 10787, 10813, 10843, 9657, 9681, 9710, 9815, 9843, 9896, 9947, 10017, 10052, 10074, 10102, 10157, 10211, 10269, 10297, 10332, 10364, 10393, 10424, 10489, 10515, 10547, 10583]. Household [10648, 12797, 12531]. Houses [993, 3928, 4849, 5582, 7069, 7033, 9336, 10927]. Howgate [10376]. Howson [11038]. Hub [1790, 4546, 6066, 7109]. Hubbell [245]. Hubs [7967]. Hudson [1403, 2618, 2567, 2593, 6503, 7807, 6896, 9784, 13427]. Hughes [7468, 11615, 11775, 12741]. Hull [5197]. Hulling [1867, 3391, 3006, 4802]. Human [1518, 2863, 8257, 10082, 10768, 10479]. Humboldt [2289]. Humorous [1624]. Humpback [12536]. Hundred [560, 1613, 1622, 1428, 3675, 4513, 5616, 6005, 5378, 6376, 7328, 7884, 9779]. Hundred-Ton [9779]. Hungarian [12294]. Hungary [12210]. Hunting [9611]. Hurdle [12587]. Husker [3087]. Husking [7342, 10775, 1452]. Hutchinson [1049]. Huxley [939, 959, 1185, 6269, 6053, 9269, 9137, 9201]. Hyacinthe [3400]. Hybrid [10969]. Hydrant [5882, 7136, 12995]. Hydrants [6866, 12260]. Hydrate [718]. Hydraulic [1144, 745, 1986, 1822, 2462, 2488, 4833, 4261, 6453, 5357, 6256, 6333, 5770, 7149, 6878, 7300, 8920, 8842, 8354, 9704, 10468, 9966, 9664, 10979, 11424, 12156, 11241, 11741, 11217, 12822, 13465]. Hydriodic [13046]. Hydrium [2773]. Hydro [942, 2363, 7156, 11183]. Hydro-Aero-Dynamic [2363]. Hydro-Electric [7156]. Hydro-Pneumatic [942, 11183]. Hydrobromic [11439]. Hydrocarbon [8503, 9551]. Hydrogen [490, 448, 1742, 1189, 3262, 2773, 3105, 8922, 11149, 11097, 12396]. Hydrology [7237]. Hydrometers [3644]. Hydromotor [13399]. Hydrophobia [646, 6013, 5894, 12148, 13273]. Hydrophobia-Practical [6013]. Hydrostatic [721, 197, 2592, 3826, 9318]. Hygiene [5142, 5177, 5201]. Hygrometer [9583]. Hygroscopes [9642]. Hymenocalyx [6530]. Hymenopters [717]. Hymer [10993]. Hyposulphite [10699]. Hypothesis [2938].

I. [5104]. Ibex [13616]. Ibis [12596]. Ice [93, 206, 388, 676, 627, 1123, 192, 1173, 1031, 1410, 1878, 1795, 1346, 1765, 2741, 3207, 2574, 2618, 2515, 2567, 2593, 3062, 2565, 3009, 3714, 4688, 4847, 5608, 6323, 5985, 5613, 5984, 7155, 7380, 6876, 7609, 8780, 9342, 10931, 10927, 9784, 9533, 10403, 10287, 10850, 10544, 11612, 11228, 11144, 12591, 13136, 12492, 12486, 13505, 8359, 8275, 11816]. Ice-Making [2515, 5608, 7609, 11612, 12591]. Icelanders [6215]. Ices [11816]. Idea [2001, 7505]. Ideas [553, 7531, 12814]. Ideation [9609]. Identity [2707]. Idle [1271]. Ignazio [10669]. Igniting [1780]. Ignition [4208, 3884, 6576]. Ignorance [1942, 2673, 7899]. Iguana [7697]. II [7195, 8561, 11326, 12937, 10911]. III [4381, 8588, 11354, 12960]. Ilion [2808]. Illinois [1143, 1771]. Illuminating [859, 746, 790, 2542, 4905, 4875, 8283]. Illumination [3656, 3855, 6491, 9220, 8617, 9428, 10139, 10341, 11222, 12343]. Illuminator [12850]. Illusion [6998]. Illustrated [2173, 3083, 8839, 10796]. Illustrating [2925]. Images [514]. Imagination [11927, 1045, 1066, 1090]. Imhof [11195]. Imitate [176]. Imitation [10720]. Immense [6794].

Immigration [1695]. Impact [535, 4172, 8963]. Impatience [12904]. Imperfection [2599]. Imperial [10766]. Impermeability [6973]. Impey [2250]. Implement [341, 1214, 7342, 13231, 13674]. Implements [3961, 9887, 13134, 13121]. Importance [2571, 5374, 5451, 5501, 5528, 5558, 5586, 5642, 6068, 6146, 7633]. Important [1043, 604, 791, 923, 748, 1644, 1650, 2150, 2396, 1280, 2099, 2115, 2168, 2173, 2146, 3605, 3538, 2836, 3708, 2924, 4373, 5106, 5598, 6962, 6770, 11007, 12490, 9964]. Imported [665, 2188]. Importers [791]. Imposition [4548]. Impossibilities [2772]. Impracticabilities [446]. Impracticable [6325]. Impregnating [1481]. Impressions [11812]. Improve [12640]. Improved [524, 545, 274, 365, 874, 368, 245, 574, 980, 334, 193, 1097, 343, 195, 870, 829, 981, 1003, 1027, 713, 964, 6, 341, 1089, 1159, 119, 240, 597, 1163, 116, 463, 92, 418, 223, 900, 958, 438, 532, 1136, 783, 892, 501, 760, 896, 766, 444, 810, 465, 289, 200, 614, 420, 294, 1023, 47, 270, 1512, 2243, 2144, 1665, 2101, 1893, 1689, 1639, 2339, 1426, 1351, 1188, 1348, 1452, 1238, 1265, 1540, 1318, 1391, 1587, 1297, 1867, 1846, 1244, 1436, 1530, 1542, 2013, 1567, 2264, 1295]. Improved [1790, 1217, 1323, 1823, 2285, 1640, 2186, 1289, 1368, 1457, 1611, 1541, 1246, 1638, 1824, 2148, 1744, 1664, 1820, 2317, 2265, 2122, 2125, 2290, 1766, 3129, 2691, 2718, 3151, 3562, 2622, 2522, 2857, 3128, 3079, 3005, 3420, 3586, 3755, 3086, 3150, 3414, 3355, 3561, 2844, 3722, 3421, 2614, 2640, 3620, 2807, 3259, 3447, 2878, 2684, 2987, 3361, 3214, 3213, 3788, 3392, 3179, 3709, 2898, 3235, 3468, 3685, 3501, 3054, 3185, 3493, 3715, 3784, 3674, 3326, 3333, 3622, 3526, 2540, 2712, 2794, 2666, 3364, 5155, 3807, 5114, 5018]. Improved [5173, 4960, 5143, 3839, 4339, 3850, 4931, 5113, 4959, 4890, 4428, 4675, 5196, 4938, 4484, 4608, 4728, 4198, 4009, 4804, 4510, 5047, 4125, 5092, 4547, 5070, 3863, 4724, 4180, 4300, 4790, 4104, 4460, 4609, 4546, 4091, 3911, 4688, 5020, 5082, 4939, 4700, 4380, 4669, 4915, 4282, 4576, 4990, 4430, 4010, 4340, 4495, 4668, 4213, 4887, 4013, 4437, 4994, 4298, 4056, 3959, 5032, 3927, 4941, 4142, 4500, 4808, 4832, 5119, 4923, 4131, 4299, 4240, 4638, 4803, 4805, 3827, 5473, 6528, 6252, 5882, 6066, 5985, 6141, 5693, 5607]. Improved [6065, 6218, 5721, 5580, 6335, 5552, 5578, 5274, 6227, 5448, 5553, 5823, 5660, 6374, 6434, 6059, 5331, 5447, 6267, 5893, 6354, 6155, 6093, 5771, 6362, 5497, 5361, 5248, 5689, 5494, 5986, 5651, 5484, 5638, 5998, 5418, 6165, 5905, 5960, 6117, 5973, 5855, 6455, 5717, 5332, 6064, 5666, 5773, 5744, 5468, 6240, 6079, 5307, 5250, 6038, 6277, 5360, 5629, 5272, 5747, 5612, 5881, 6452, 5659, 5306, 5814, 5854, 6221, 5762, 6104, 6309, 6410, 6411, 5545, 6251, 5930, 6526, 5496, 6361, 5536, 6010, 5609, 5582, 6388, 5471, 5273]. Improved [5275, 5300, 5583, 6058, 6163, 5422, 5748, 5414, 5821, 5876, 6196, 6224, 5419, 5640, 6389, 5880, 5548, 6478, 5772, 6142, 6241, 6302, 6166, 6721, 6970, 6656, 6972, 7219, 7221, 6780, 7076, 7561, 7333, 6822, 6948, 6961, 7135, 7103, 7164, 7325, 7133, 7190, 6943, 6803, 7489, 6850, 7015, 7173, 6725, 6922, 7148, 7074, 7270, 7575, 6879, 6949, 7728, 6669, 7362, 7001, 7286, 7163, 7586, 7289, 7390, 7416, 7666, 7065, 7254, 7584, 7427, 7525, 7848, 6726, 7022, 6605, 7671, 7269, 7109, 7218, 6606, 7242, 6878, 6919, 7306, 7136]. Improved [7609, 6846, 7549, 7852, 6797, 7054, 7615, 6851, 7031, 7280, 7341, 7243, 7466, 7490, 7760, 7391, 6654, 7189, 6853, 6571, 7451, 6750, 6828, 7309, 7672, 6642, 7005, 7366, 7262, 7288, 7587, 7823, 7043, 7883, 6825, 7159, 7417, 7714, 6802, 6991, 7694, 6798, 6978, 6572, 7025, 7048, 6875, 7081, 7027, 7337, 6675, 7853, 7239, 7389, 7663, 6872, 6979, 7336, 7078, 7613, 7510, 7910, 7644, 7537, 6923, 7727, 7379, 6898, 6745, 7052, 7101, 7724, 7030, 7082, 7438, 7564, 7905, 7820, 7817, 7821, 6999, 7339, 7411, 7077, 7111, 6997]. Improved [6950, 6753, 7000, 8396, 8716, 8136, 8486, 8778, 8255, 8994, 9472, 8872, 8138, 8958, 8135, 9433, 7990, 9083, 8456, 8806, 8871, 9181, 8471, 8741, 8283, 8330, 8510, 8192, 8165, 9175, 9394, 8682, 9182, 7994, 9242, 8993, 8050, 8081, 8630, 9372, 8167, 8807, 8831, 8218, 8840, 9346, 8277, 8652, 9024, 9180, 8346, 8023, 8656, 8166, 8873, 8600, 9087, 9117, 8153, 8905, 9000, 8986, 9012, 9249, 8423, 9370, 9408, 7964, 8460, 8046, 8077, 8899, 8308, 8388, 9407, 8041, 8072, 8427, 8842, 8710, 8835, 8951, 8280, 8223, 8336]. Improved [8139, 8491, 8513, 8390, 9424, 9149, 8421, 9205, 8195, 8189, 9250, 9467, 9235, 8397, 8249, 8019, 8928, 8429, 8270, 8459, 9025, 8512, 8569, 9234, 8747, 7992, 8307, 9206, 8453, 8422, 9020, 8180, 8690, 8776, 9332, 9314, 8094, 8572, 7963, 8772, 8516, 8106, 8714, 7932, 7993, 8190, 8740, 8800, 8906, 7937, 8163, 8164, 8808, 8841, 8907, 9021, 9404, 9310, 8797, 9503, 8279, 8570, 8988, 7936, 8571, 9076, 9277, 9339, 9344, 9506, 9150, 8901, 8927, 8802, 9345, 9470, 8109, 8482, 8160, 8629, 8837, 8329, 8024, 7931, 8801, 8771]. Improved [9279, 8304, 8933, 8987, 8995, 8362, 8452, 8599, 9243, 8626, 9118, 8428, 8108, 9458, 8188, 9282, 8517, 8256, 8745, 8455, 9056, 8573, 10535, 9921, 10666, 10779, 10643, 10260, 10636, 10092, 10319, 10538, 10444, 10502, 10695, 10712, 10662, 10778, 10261, 10496, 10574, 9980, 10832, 10204, 10802, 10353, 10607, 9639, 10257, 10416, 9703, 10012, 10172, 10413, 10887, 9674, 9613, 10171, 10379, 10721, 10411, 10412, 10440, 10504, 10011, 10165, 10085, 10532, 10037, 10313, 10875, 10205, 10511, 9956, 10545, 9916, 10347, 10173, 9731, 10285, 10534, 10468, 10041, 9533, 10403, 9729]. Improved [10833, 9648, 9679, 10384, 10525, 10474, 10635, 10638, 10314, 10116, 10152, 10744, 10576, 10634, 10939, 10501, 10909, 10741, 9534, 10229, 10795, 10908, 10197, 9905, 10315, 10121, 10524, 10508, 10728, 10227, 10610, 10064, 10665, 9808, 10507, 9617, 10724, 10937, 10967, 9700, 9876, 10865, 9724, 10827, 10861, 9672, 10445, 10568, 9889, 10443, 10410, 9986, 10726, 9968, 10777, 10122, 10806, 10965, 10354, 10663, 9559, 10151, 10111, 11073, 11093, 11618, 11693, 11407, 11299, 11090, 11710, 11587, 11558, 12289, 11282, 11347, 11692, 11432, 11317, 11484, 11370, 12183, 11375, 11507, 11165, 12025, 11213]. Improved [11042, 11298, 11495, 11564, 11512, 11653, 12229, 11619, 11269, 11145, 11338, 11430, 11350, 11343, 11461, 11166, 11041, 11532, 11418, 11423, 11372, 11069, 11117, 11709, 11320, 11435, 11563, 11436, 12208, 11557, 11511, 11613, 11711, 11553, 12028, 11371, 11297, 11349, 11453, 11645, 11672, 11425, 11301, 12227, 11702, 11402, 11535, 11452, 11556, 11460, 11670, 11649, 10989, 11240, 11118, 11668, 10987, 10988, 11141, 11113, 11168, 11172, 12795, 13054, 12774, 13287, 12523, 13565, 13202, 12654, 13387, 12829, 12655, 12493, 12528, 12986, 13179, 13064, 13670, 13675, 12995, 12591, 13283, 12675, 12444, 12318, 12828]. Improved [12884, 13131, 13185, 13646, 12679, 12858, 13152, 12448, 13178, 12497, 13232, 12994, 13282, 13233, 13536, 12886, 13614, 12964, 13229, 13568, 12386, 13097, 12465, 12422, 12988, 12802, 12586, 12685, 13100, 13099, 12826]. Improvement [198, 34, 311, 482, 1065, 110, 321, 215, 2241, 1773, 1513, 1507, 2396, 1799, 2369, 2274, 1821, 1743, 1211, 1822, 1575, 3389, 3605, 2950, 3599, 2461, 2924, 2798, 2790, 4666, 4067, 5068, 3945, 5061, 6330, 5961, 5988, 6447, 7402, 6648, 7187, 7355, 6938, 8475, 8045, 8076, 11763, 12073, 11161, 12074, 13642, 12494, 13184, 13010, 13464, 12735, 10616]. Improvements [723, 943, 557, 2218, 2228, 1431, 2020, 1746, 2437, 3275, 3630, 2890, 4204, 5169, 3974, 5118, 4593, 5980, 5633, 5959, 5953, 5952, 6994, 7050, 7259, 7878, 8929, 7989, 8157, 9270, 8892, 8134, 10341, 10030, 11474, 11829, 11283, 12018, 11255, 11621]. Improves [8732, 10752, 11508]. Improving [4201]. Improvised [6327]. Impulses [5404]. Impurities [3299]. Inaugural [939, 959]. Inauguration [2126, 1807, 3308]. Incendiary [8956, 13554]. Inch [3703, 7603, 7563]. Incident [2060]. Incline [12292]. Inclined [3223, 4158]. Income [49]. Incoming [12069]. Inconsistency [8014]. Increase [6867]. Increased [450, 5817]. Increasing [877, 4185, 10531, 12200]. Incrustation [1345, 3607]. Incrustations [314, 860, 11463]. Incubator [8336]. Indefinite [10797]. Independence [2211, 11880]. Independent [1089, 1159, 11290, 11832]. Index [592, 1157, 1789, 2408, 2485, 2513, 3123, 3148, 3174, 3206, 3233, 3258, 3292, 3324, 3351, 3381, 3412, 3440, 3466, 3492, 3520, 3554, 3581, 3616, 3641, 3673, 3702, 3738, 3767, 2539, 2561, 2583, 2612, 2638, 2663, 2682, 2705, 2733, 2755, 2780, 2806, 2833, 2856, 2877, 2897, 2914, 2934, 2960, 3004, 3031, 3077, 3102, 3033, 3768, 3799, 3825, 4561, 4592, 4628, 4656, 4687, 4715, 4741, 4766, 4789, 4819, 4843, 4872, 4901, 4922, 4949, 4975, 5027, 5031, 5057, 5081, 5103, 5130, 5154, 5186, 5209, 3849, 3879, 3940, 3968, 3997]. Index [4026, 4055, 4090, 4114, 4141, 4168, 4197, 4224, 4256, 4281, 4313, 4338, 4366, 4393, 4420, 4449, 4482, 4508, 4540, 4509, 5210, 5239, 5265, 5918, 5947, 5972, 5996, 6023, 6047, 6077, 6103, 6128, 6153, 6181, 6207, 6238, 6265, 6291, 6321, 6348, 6372, 6399, 6421, 6443, 6465, 6488, 6512, 6540, 5292, 5320, 5345, 5372, 5402, 5433, 5459, 5483, 5510, 5535, 5565, 5595, 5627, 5649, 5679, 5704, 5730, 5760, 5784, 5812, 5840, 5866, 5891, 5867, 6541, 6570, 6595, 7230, 7253, 7279, 7299, 7324, 7354, 7378, 7401, 7426, 7450, 7478]. Index [7501, 7524, 7548, 7574, 7600, 7630, 7654, 7683, 7713, 7744, 7777, 7806, 7837, 7867, 7895, 6618, 6641, 6668, 6691, 6712, 6736, 6764, 6790, 6815, 6837, 6864, 6889, 6906, 6933, 6960, 6990, 7014, 7042, 7064, 7091, 7122, 7147, 7172, 7203, 7205, 7896, 7923, 7951, 8644, 8673, 8701, 8730, 8760, 8790, 8822, 8858, 8890, 8918, 8949, 8979, 9010, 9042, 9074, 9108, 9134, 9168, 9199, 9232, 9266, 9299, 9362, 9392, 9455, 7979, 8010, 8040, 8066, 8071, 8125, 8152, 8179, 8210, 8240, 8269, 8295, 8320, 8344, 8380, 8412, 8443, 8470]. Index [8502, 8533, 8558, 8585, 8615, 8675, 9491, 10952, 9523, 9548, 10246, 10275, 10303, 10336, 10368, 10401, 10430, 10460, 10494, 10522, 10557, 10591, 10623, 10654, 10680, 10710, 10738, 10764, 10792, 10818, 10848, 10873, 10898, 10951, 9605, 9631, 9662, 9716, 9741, 9765, 9795, 9820, 9848, 9874, 9903, 9932, 9954, 9978, 10002, 10026, 10057, 10079, 10108, 10136, 10163, 10189, 10218, 9881, 10975, 11003, 11853, 11877, 11900, 11920, 11947, 11966, 11987, 12017, 12040, 12063, 12086, 12112, 12141, 12166, 12192, 12219, 12244, 12274, 11025, 11054, 11078, 11102, 11128, 11155, 11181, 11203]. Index [11226, 11253, 11279, 11308, 11336, 11493, 11547, 11577, 11678, 11701, 11721, 11746, 11770, 11794, 11822, 11359, 11388, 11415, 11631, 12271, 12300, 12327, 13004, 13033, 13062, 13116, 13141, 13168, 13194, 13217, 13244, 13269, 13296, 13324, 13352, 13376, 13397, 13423, 13453, 13478, 13503, 13527, 13551, 13577, 13602, 13628, 13657, 13684, 12352, 12380, 12403, 12430, 12457, 12481, 12512, 12548, 12572, 12601, 12637, 12663, 12693, 12725, 12756, 12785, 12812, 12843, 12870, 12901, 12927, 12954, 12978, 13089, 12336]. India [1947, 1911, 2087, 3417, 4225, 5869, 6167, 5954, 7429, 11120, 11147, 11817]. Indian [114, 2953, 2948, 6880, 9462, 11486]. Indiana [4859]. Indicate [10539]. Indicating [5347, 12207]. Indications [9589]. Indicator [1342, 4605, 6967, 6916, 7054, 6913, 9249, 8196, 10724, 12471, 13094, 11657, 11809]. Indicators [4370]. Indigestible [10527]. Indigo [11803, 12582]. Indirect [2626]. Indoor [6479, 7913]. Induced [3712]. Induction [1104, 4988, 7414, 7793, 8381, 10195, 12742, 12741, 1753]. Industrial [584, 583, 1301, 3367, 2880, 3678, 2974, 3565, 3482, 3095, 3342, 3308, 4490, 3829, 5487, 6131, 6243, 6513, 6908, 9178, 8543, 8651, 8212, 8420, 10604, 9782, 12090, 11894, 11637, 11593, 13145, 12836, 12630, 13301, 12575, 12731, 13402]. Industries [113, 2365, 6502, 5439, 8845, 8877, 9801, 11660, 12462, 12483, 12514, 12574, 12609, 12787, 12818, 12851, 12873, 12908, 13485, 13519, 13529, 13590, 13611, 13643, 13671, 13699, 12640, 12333, 12355, 12384, 12432]. Industry [604, 839, 2373, 2719, 3308, 3660, 2770, 4902, 6049, 5403, 7020, 6643, 6704, 8184, 8100, 9983, 9811, 11661, 11635, 11640, 12736, 12411, 6183]. Inebriety [9610]. Inertia [278]. Inertiæ [278]. Inexact [10278]. Inexhaustible [5714]. Inexplicable [279]. Inextinguishable [2815, 4986]. Infant [1487]. Infantile [1432]. Infants [2644]. Infectious [4178]. Infernal [8044, 8075]. Infinitesimal [4229]. Influence [857, 788, 1322, 1477, 2626, 3247, 5322, 6084, 6298, 8809, 10798, 10601, 10717, 11437, 11177]. Influenced [2652, 2570]. Information [2995, 2531, 2554, 3360, 3416, 3545, 3571, 3663, 3684, 3783, 2576, 2602, 2629, 2656, 2675, 2697, 2726, 2742, 2767, 2792, 2818, 2838, 2885, 2922, 2949, 2975, 3025, 3069, 3132, 3167, 3198, 3284, 3314, 3334, 3843, 3869, 4850, 4881, 4929, 4957, 4985, 5012, 5038, 5066, 5096, 5112, 5137, 5163, 5194, 5226, 3900, 4036, 4075, 4098, 4121, 4152, 4186, 4212, 4291, 4321, 4347, 4376, 4426, 4457, 4493, 4518, 4553, 4636, 4664, 4697, 4731, 4749, 4773, 4828, 5252, 5276, 6139, 6191, 6328, 6380, 6427, 6449, 6481]. Information [6497, 5310, 5327, 5354, 5382, 5416, 5445, 5466, 5491, 5520, 5605, 5637, 5665, 5687, 5713, 5796, 5831, 5875, 5909, 5928, 5958, 5979, 6031, 6087, 6113, 5746, 6650, 6674, 7675, 7689, 7720, 7752, 7784, 7847, 6699, 6747, 6766, 6801, 6849, 6897, 6941, 6969, 7023, 7079, 7115, 7129, 7180, 7238, 7268, 7305, 7345, 7408, 7457, 7486, 7536, 7581, 7608, 7636, 9079, 7933, 7997, 8018, 8053, 8084, 8110, 8199, 8276, 8401, 8451, 8514, 8545, 8597, 10866]. Infringement [1898, 2937, 2977]. Infusorial [1599, 9171]. Ing [2686]. Ingenious [8, 2394, 10985, 11686]. Ingots [7388, 7668]. Ingram [10796, 12097]. Inhabitants [13134, 218]. Inhabited [4341, 3973, 4884, 4908, 12763]. Inhaling [1120]. Initial [5213]. Injecting [13305]. Injection [5123]. Injector [3885, 10117, 11451, 11505, 12369]. Injectors [7076, 6898]. Injure [11524]. Injured [4926]. Injurious [1625, 10290]. Injustice [4144, 4284, 11438]. Ink [114, 1164, 10435]. Inkstand [5714]. Inlaid [9279, 12947]. Inland [7016]. Inorganic [2771]. Inquiries [10844, 10869, 10894, 10920, 10946, 10972, 10998, 11021, 11050, 11099, 11125, 11151]. Insane [11208]. Insanity [9666]. Insect [1636, 5614, 7526, 8903, 11833, 11430]. Insect-Eating [7526]. Insects [416, 665, 4963, 6029, 7326, 9798, 12708, 13368, 13014, 1591, 10641]. Insecurity [12174]. Inserted [9282, 12293]. Inserting [3103]. Insoles [9083]. Inspected [2012]. Inspection [423, 469, 1544, 1621, 2042, 2810, 4069, 6001, 6893, 7383, 11401, 830]. Inspector [12888]. Inspirator [12466]. Instability [7282]. Instantaneous [5661, 7158]. Instinct [3618]. Instinctive [7686]. Institute [706, 17, 1013, 1030, 1056, 908, 928, 951, 970, 995, 1015, 150, 127, 648, 21, 583, 1996, 2072, 1192, 2286, 2104, 2133, 2154, 2174, 2201, 2226, 2251, 2271, 1248, 1494, 3427, 3457, 3480, 3512, 3679, 2621, 3369, 2976, 2465, 2689, 2789, 2884, 4492, 4907, 4930, 4958, 6217, 6281, 6312, 6392, 6457, 6413, 6436, 7556, 7591, 7607, 7635, 7662, 7783, 6739, 9022, 9154, 9255, 12213, 13456, 12731, 4382, 4295, 4323, 4410, 4351, 3712, 5736]. Instituted [1148]. Institution [3643, 2463]. Institutions [8908]. Instruction [1543]. Instructors [8861, 9080]. Instrument [456, 958, 1818, 3306, 3918, 4076, 4962, 4149, 4999, 6331, 5581, 7134, 7582, 11016, 13669, 13283, 13341, 12584, 13592]. Instruments [606, 1702, 1569, 8227, 10964]. Insulation [982]. Insurance [6001, 8539, 7955]. Intellectual [4171]. Intelligence [100, 172, 1035, 1070, 1127, 1149, 232, 327, 391, 451, 473, 515, 558, 571, 616, 649, 677, 704, 716, 752, 775, 795, 820, 846, 885, 931, 994, 1014, 1361, 1413, 2326, 2350, 2374, 2398, 1455, 1498, 1552, 1578, 1601, 1626, 1647, 1672, 1727, 1778, 1800, 1831, 1879, 1923, 1945, 2021, 2046, 2090, 2142, 2204, 2227, 2299, 2447, 2475]. Intelligent [1409, 13272]. Intending [3860]. Intense [1477]. Intensity [11365]. Intentions [12534, 13566]. Inter [4490, 9204, 11499]. Inter-Mercurial [9204, 11499]. Inter-State [4490]. Interest [2251, 3626, 5192, 13356, 12764, 10528]. Interest-Bearing [12764]. Interesting [361, 691, 2000, 2071, 2132, 1401, 2145, 2012, 2503, 2707, 2553, 5009, 5387, 10206, 11959, 12997, 291]. Interests [6894, 8347]. Interfere [7152]. Interference [2016, 10090]. Interior [2711, 3080, 7493, 10716]. Intermittent [9366]. Internal [2079]. International [1301, 2149, 2967, 3219, 3272, 3475, 2572, 3862, 6012, 6313, 7809, 7098, 8535, 8565, 9119, 8387, 10098, 10599, 12246, 12120, 13619, 12836, 396, 3268, 4063, 12931]. Interoceanic [3450]. Interplanetary [94]. Interpretation [9366]. Interstellar [7716]. Intervals [3058]. Intolerance [3631]. Intrenching [1268]. Introduce [2568, 2608, 2634, 2657, 4189, 5424, 5474, 5620, 6014, 6095, 6172]. Introducing [4218, 4249, 4275, 4306, 4331, 4478]. Introduction [2087, 4062]. Invalid [13017]. Invalidity [2022]. Invalids [193]. Invent [12092]. Invented [7985]. Inventing [4955]. Invention [366, 305, 126, 905, 1160, 647, 727, 1069, 96, 1857, 1353, 1622, 1279, 1411, 1459, 1221, 2796, 3372, 2568, 2608, 2634, 2657, 2626, 3478, 3714, 2925, 4189, 4120, 4596, 5424, 5474, 5620, 6014, 6095, 6172, 7234, 7454, 9044, 8244, 9839, 9927, 9949, 10051, 10073, 10182, 10339, 10877, 10390, 10415, 10451, 10797, 12282, 11284, 11799, 12067, 11680, 11206, 11973, 11903, 12697, 12445, 13530, 12415, 12308, 8217]. Inventions [788, 553, 1034, 1964, 1646, 2066, 2064, 3022, 3778, 3452, 2485, 2513, 3123, 3148, 3174, 3206, 3233, 3258, 3292, 3324, 3351, 3381, 3412, 3440, 3466, 3492, 3520, 3554, 3581, 3616, 3641, 3673, 3702, 3738, 3767, 2539, 2561, 2583, 2612, 2638, 2663, 2682, 2705, 2733, 2755, 2780, 2806, 2833, 2856, 2877, 2897, 2914, 2934, 2960, 3004, 3031, 3077, 3102, 2686, 3689, 3595, 2890, 4218, 4249, 4275, 4306, 4331, 4478, 4810, 3799, 3825, 4561, 4592, 4628, 4656, 4687, 4715, 4741, 4766, 4789, 4819, 4843, 4872, 4901, 4922, 4949]. Inventions [4975, 5027, 5031, 5057, 5081, 5103, 5130, 5154, 5186, 5209, 3849, 3879, 3940, 3968, 3997, 4026, 4055, 4090, 4114, 4141, 4168, 4197, 4224, 4256, 4281, 4313, 4338, 4366, 4393, 4420, 4449, 4482, 4508, 4540, 5239, 5265, 5918, 5947, 5972, 5996, 6023, 6047, 6077, 6103, 6128, 6153, 6181, 6207, 6238, 6265, 6291, 6321, 6348, 6372, 6399, 6421, 6443, 6465, 6488, 6512, 6540, 5292, 5320, 5345, 5372, 5402, 5433, 5459, 5483, 5510, 5535, 5565, 5595, 5627, 5649, 5679, 5704, 5730, 5760, 5784, 5812, 5840, 5866, 5891, 5515, 5786]. Inventions [7751, 7154, 6570, 6595, 7230, 7253, 7279, 7299, 7324, 7354, 7378, 7401, 7426, 7450, 7478, 7501, 7524, 7548, 7574, 7600, 7630, 7654, 7683, 7713, 7744, 7777, 7806, 7837, 7867, 7895, 6618, 6641, 6668, 6691, 6712, 6736, 6764, 6790, 6815, 6837, 6864, 6889, 6906, 6933, 6960, 6990, 7014, 7042, 7064, 7091, 7122, 7147, 7172, 7203, 7764, 7827, 7703, 7735, 7768, 7858, 7916, 7707, 7736, 7769, 7796, 7832, 7860, 7890, 7704, 7738, 7770, 7798, 7829, 7859, 7887, 7917, 7706, 7797, 7830, 7889, 7918, 7828, 7674, 7923, 7951]. Inventions [8644, 8673, 8701, 8730, 8760, 8790, 8822, 8858, 8890, 8918, 8949, 8979, 9010, 9042, 9074, 9108, 9134, 9168, 9199, 9232, 9266, 9299, 9362, 9392, 9455, 7979, 8010, 8040, 8066, 8071, 8125, 8152, 8179, 8210, 8240, 8269, 8295, 8320, 8344, 8380, 8412, 8443, 8470, 8502, 8533, 8558, 8585, 8615, 7944, 7972, 8851, 8881, 8941, 8971, 9003, 9035, 9065, 9101, 9125, 9159, 9191, 9257, 9292, 9324, 9352, 9382, 9415, 9482, 9518, 8002, 8034, 8058, 8089, 8117, 8144, 8171, 8204, 8230, 8290, 8312, 8373, 8406, 8435, 8464]. Inventions [8496, 8528, 8580, 8610, 8637, 8665, 8753, 8782, 8817, 8033, 8057, 8973, 9031, 9063, 8088, 8118, 8143, 8174, 8203, 8231, 8260, 8314, 8339, 8372, 8497, 8548, 8577, 8607, 8664, 8695, 8724, 8784, 8815, 8848, 8880, 8911, 8942, 8636, 8722, 8785, 8850, 8882, 8939, 9002, 9033, 9064, 9102, 9127, 9158, 9189, 9224, 9259, 9290, 9323, 9351, 9385, 9413, 9445, 9485, 9517, 7945, 7970, 8000, 8031, 8059, 8090, 8263, 8404, 8437, 8466, 8494, 8527, 8552, 8638, 8754, 8786, 8940, 8972, 9004, 9099, 9222, 9325, 9383, 9448]. Inventions [9483, 9516, 9126, 9161, 9190, 9225, 9258, 9289, 9355, 9384, 9417, 9449, 9484, 9515, 8001, 8032, 9066, 9100, 9128, 9160, 9188, 9223, 9260, 9291, 9322, 9353, 9414, 9447, 8114, 8145, 8173, 8202, 8232, 8262, 8313, 8371, 8405, 8434, 8465, 8526, 8550, 8578, 8609, 8635, 8666, 8723, 8816, 8849, 8883, 8912, 8970, 9032, 8127, 8738, 10958, 9523, 9548, 10246, 10275, 10303, 10336, 10368, 10401, 10430, 10460, 10494, 10522, 10557, 10591, 10623, 10654, 10680, 10710, 10738, 10764, 10792, 10818, 10848, 10873, 10898, 10951, 9605, 9631, 9662]. Inventions [9716, 9741, 9765, 9795, 9820, 9848, 9874, 9903, 9932, 9954, 9978, 10002, 10026, 10057, 10079, 10108, 10136, 10163, 10189, 10218, 10870, 9570, 9597, 10426, 10455, 10487, 10518, 10551, 10585, 10646, 10673, 10703, 10785, 10812, 10840, 9624, 9655, 9683, 9711, 9788, 9814, 9842, 9868, 9895, 9928, 9946, 9998, 10021, 10050, 10101, 10160, 10183, 10214, 10238, 10266, 10296, 10328, 10363, 10395, 9571, 9599, 10733, 10759, 10787, 10813, 10843, 9657, 9681, 9710, 9815, 9843, 9896, 9947, 10017, 10052, 10074, 10102, 10157, 10211, 10269, 10297, 10332, 10364, 10393]. Inventions [10424, 10489, 10515, 10547, 10583, 10648, 10081, 9543, 9600, 10422, 10453, 10485, 10516, 10549, 10582, 10618, 10649, 10672, 10704, 10734, 10760, 10783, 10810, 10842, 9626, 9736, 9760, 9790, 9812, 9841, 9869, 9945, 9974, 9997, 10020, 10048, 10072, 10100, 10129, 10156, 10180, 10213, 10239, 10270, 10293, 10329, 10360, 10396, 9544, 9568, 10271, 10298, 10331, 10361, 10392, 10423, 10454, 10486, 10517, 10550, 10584, 10619, 10650, 10675, 10705, 10732, 10757, 10784, 10811, 10841, 9598, 9625, 9656, 9682, 9735, 9761, 9789, 9813, 9840, 9866, 9898, 9925, 9944]. Inventions [9973, 9995, 10019, 10047, 10071, 10103, 10130, 10158, 10181, 10212, 10241, 9787, 10267, 10294, 10488, 10552, 10702, 9569, 9658, 10548, 10586, 10647, 10674, 10701, 10758, 10786, 9684, 9709, 9759, 9791, 9816, 9867, 9897, 9926, 9948, 9996, 10018, 10049, 10131, 10159, 10184, 10210, 10240, 10268, 10295, 10330, 10362, 10394, 10425, 10456, 10686, 10775, 10436, 11826, 11524, 11729, 10975, 11003, 11853, 11877, 11900, 11920, 11947, 11966, 11987, 12017, 12040, 12063, 12086, 12112, 12141, 12166, 12192, 12219, 12244, 12274, 11025, 11054, 11078, 11102, 11128, 11155, 11181, 11203]. Inventions [11226, 11253, 11279, 11308, 11336, 11493, 11547, 11577, 11678, 11701, 11721, 11746, 11770, 11794, 11822, 11417, 11882, 11229, 12710, 13449, 13463, 13520, 13562, 13622, 13672, 13697, 12385, 12622, 12770, 12825, 12883, 12915, 12991, 13055, 13135, 13181, 13203, 13237, 13262, 13290, 13309, 13345, 13442, 13468, 13493, 13509, 13572, 13680, 13540, 12797, 13357, 12300, 12327, 13004, 13033, 13062, 13116, 13141, 13168, 13194, 13217, 13244, 13269, 13296, 13324, 13352, 13376, 13397, 13423, 13453, 13478, 13503, 13527, 13551, 13577, 13602, 13628, 13657, 13684, 12352, 12380, 12403, 12430, 12457, 12481, 12512]. Inventions [12548, 12572, 12601, 12637, 12663, 12693, 12725, 12756, 12785, 12812, 12843, 12870, 12901, 12927, 12954, 12978, 12347, 13636, 12500, 13388, 13406, 13411, 13433, 13436, 13461, 13471, 13488, 13490, 13517, 13521, 13535, 13538, 13563, 13587, 13597, 13612, 13615, 13638, 13644, 13667, 13677, 13695, 13700, 12682, 12707, 12801, 12820, 12890, 12911, 13019, 13079, 13107, 13187, 13210, 13228, 13255, 13258, 13280, 13289, 13312, 13318, 13339, 13363, 13367, 13385, 12585, 12657, 12777, 12804, 12854, 12920, 12962, 13026, 13110, 13182, 13328, 12617, 13204, 13313, 13347, 13414, 12711, 12593, 12703, 12739, 12794]. Inventions [12861, 12941, 12970, 12992, 13020, 13048, 13074, 13102, 13153, 13188, 13207, 13234, 12544, 12565, 12620, 13591]. Inventive [880, 5351, 9882]. Inventor [695, 714, 2261, 2061, 1355, 2899, 3508, 4692, 7722, 7876, 8691, 9880, 11312, 11472, 12263, 13053, 12643, 11976]. Inventors [296, 730, 153, 639, 2319, 1864, 1886, 2179, 1918, 3448, 2970, 3318, 3377, 3406, 3434, 3461, 3488, 3516, 3550, 3576, 3612, 3636, 2890, 2996, 4207, 4032, 4623, 4618, 4648, 6107, 6402, 6325, 5380, 7044, 8678, 9114, 9981, 10169, 10742, 11824, 12295, 12034, 12059, 12081, 12107, 12135, 12161, 12187, 12214, 12239, 12268, 12048, 11438, 11257, 11800, 12833, 12733, 12540, 12322, 12345, 12375, 12398, 12424, 12450, 12475, 12505, 12541, 12566, 12632, 12659, 12687, 12720, 12751, 12779, 12807, 12838, 12864, 12896, 12923, 140]. Inversion [8830]. Invertebrate [4095]. Investigate [6519]. Investigation [5921, 13360, 13220, 1961]. Investigations [2942, 4745, 4720, 6192, 8705, 10342, 9614, 11488, 11377]. Investigators [2970, 11856]. Investors [3435, 3462, 3489, 3572, 4587, 4622]. Invisible [347, 4034, 5135, 3924, 7983]. Iodine [13343]. Iowa [6489, 12678]. Irate [679]. Ireland [2239]. Iridescent [3310, 12224]. Iridium [6843]. Iris [6032, 7617, 8659]. Irish [7292]. Iron [919, 219, 688, 1115, 69, 943, 614, 792, 1293, 2196, 1749, 1463, 1673, 1477, 1572, 1373, 2238, 2082, 1960, 2549, 2848, 2816, 2948, 3385, 3020, 3185, 3532, 3592, 3454, 2756, 3660, 2924, 2961, 2589, 4946, 3831, 4384, 4403, 4496, 5164, 5035, 5246, 6183, 6495, 5709, 6008, 6049, 5469, 6562, 5438, 6452, 5323, 5329, 5435, 5321, 6274, 7785, 7153, 7043, 6802, 6798, 6751, 7879, 7051, 6966, 8876, 9270, 8900, 10316, 10912, 9853, 10373, 9983, 9607, 12149, 11737, 11640, 11681, 11773, 11859, 12792, 13209, 12717, 12449, 12579, 13286, 13037]. Iron [13285, 12877, 13483, 141, 29, 5302, 5305]. Iron-Clads [219]. Iron-Framed [6562]. Ironclad [4802, 7825, 9965, 10146, 11421, 11174]. Ironclads [6039, 11726]. Ironing [1262, 12647]. Ironwork [9145]. Irregularity [9338]. Irrepressible [622]. Irrespirable [11142]. Irrigation [2033, 6744, 7429, 9219, 10113]. Irrigator [4938, 9373]. Irving [1830]. Isa [12728]. Isabella [4640]. Island [1004, 1215, 13093]. Islands [8487]. Issue [1890, 1914]. Issued [2248]. Issuing [9638]. Isthmus [54, 12985]. Iswage [5045]. Italian [530, 2261, 3678, 5613, 11845]. Italy [4774, 6034, 5883, 9922]. Item [11395]. Items [3250, 3304, 3401, 3429, 787]. Itself [1719, 3109]. IV [4404, 4434]. Ivan [516]. Ivory [1565, 1429]. Ivy [10404, 12184]. Iwakuni [10510].

J [3539, 7632]. J. [271, 7692, 7632, 9364]. Jabiru [11759]. Jacana [12211]. Jack [1453, 5071, 4542, 6453, 6226, 5330, 8255, 8868, 8455, 11461, 11556]. Jacket [12104]. Jacobianum [7491]. Jacquard [10721]. Jaded [12698].

Jaguar [8843]. Jaguars [7995]. Jamaica [12235]. James [1143, 3662, 10801, 3024, 2526, 5662, 5403]. Jamin [5187, 7346]. Japan [2066, 2214, 2620, 3751, 2514, 5219, 5816, 8845, 8877, 10510, 10387]. Japanese [2898, 2655, 4730, 5021, 5247, 8965, 11786, 13391, 13259, 12805]. Jar [4120]. Jarring [6056]. Jaw [5607, 11290]. Jean [1724]. Jenkins [5774]. Jersey [2223, 1571, 2766, 6503, 8591]. Jet [7412, 8797]. Jetties [9246, 9238, 11057, 13425]. Jetty [8097, 10309]. Jew [13381]. Jewelry [1531, 13469]. Jig [3786, 4348]. Job [2774, 2887, 2821, 3886]. John [1817, 3604, 3506, 3543, 11474, 361, 6456, 9466, 10917]. Johnson [1541, 2761, 10777]. Joining [7413]. Joint [3336, 6853, 7963, 10121, 10990]. Jointer [5842]. Jointing [2785]. Joints [7695]. Joist [368]. Jonah [5215]. Jonathan [2080, 9050]. Jones [2712, 6278, 7438, 13081]. Joseph [1799, 11525, 1825, 5005, 3557]. Joslyn [2079]. Journal [8461, 10524, 12800]. Journalism [844, 74, 1324, 11787]. Journalists [10853]. Journals [3832, 5877, 13560]. Journey [222]. Jr [2080]. Jr. [2526]. Jubea [5907]. Jubilee [2971, 12521]. Judd [794]. Judge [2168, 3776, 13632, 79, 103, 131, 474]. Judges [1494]. Judson [4377]. Jugglery [11189]. Juice [8106]. July [1914, 11798]. Jump [8139]. Junction [2756]. June [703, 1213, 11663]. Jupiter [3773, 2764, 5793]. Juryman [12122]. Justice [2472, 6002]. Juvenile [7722].

K. [5104]. Kagu [6952]. Kanchil [12860]. Kansas [688, 4106, 4806, 13035]. Kaolin [3036, 6968]. Kay [421, 12686]. Keely [7209, 7235, 7257, 7285, 7302, 7357, 7432, 7661, 8623, 10481, 10438, 9829, 11363, 11636]. Keep [12433]. Keeping [2239, 9881]. Keeps [5654]. Keg [8491, 8513]. Kehoe [2794]. Keith [2461]. Kelly [668, 1832]. Kemlo [528]. Kensington [2642]. Kentucky [8154, 599]. Kenwood [2543]. Kenyon [10042]. Keokuk [5745, 6489]. Keramics [11340]. Kernel [5297]. Kerosene [1376, 2058, 3208, 3041, 8684, 8896, 9824, 12718]. Kerosene-Testing [8684]. Kessinger [1638]. Ketchum [2745]. Kettle [900, 7663]. Key [198, 810, 2387, 3753, 4070, 7367, 7139, 10503, 11847]. Key-Hole [198]. Keys [13084]. Khedive [6279]. Kid [10419]. Kieserite [3857]. Killed [12408]. Kills [6379]. Kiln [1689, 2857]. Kind [13284]. Kindling [3942]. King [154, 4099, 6472, 1892, 7905, 10250, 12080]. Kingfisher [10044]. Kinney [11183]. Kinzel [6210]. Kircher [4153]. Kiss [6574]. Kisses [12372]. Kit [9618]. Kitchen [603, 5136, 8874, 8390]. Kite [6083, 7589]. Klinkerfues [1780]. Knapp [1373]. Kneader [6337, 12745]. Kneading [11073]. Knebworth [7290]. Knife [2523, 7369, 6702, 7664, 9084, 8798, 9424, 9561, 11587]. Knife-Grinding [9424]. Knight [3456, 7564]. Knitted [2859, 3359]. Knitter [10925]. Knitting [1735, 5606, 6628]. Knives [10388]. Knob [9000, 9988]. Knobs [1542]. Knot [9210]. Knots [1007, 1479]. Know [1273, 3881]. Knowledge [750, 2336, 3215, 2696, 4096, 4040, 4258, 5437, 10944]. Knowles [8020, 7959]. Known [8875]. Kolapore [11120]. Konz [10989]. Korting [7412]. Krupp [917, 7884, 13015]. Kumis [862].

L [3758, 5489]. L. [274]. Label [10833]. Labeling [1710]. Labels [5927]. Labor [1024, 1032, 226, 2336, 1719, 2149, 2088, 2320, 2724, 2587, 6425, 6772, 9271, 10372, 11513, 11804, 12173, 12666, 13223]. Laboratory [5611, 5735, 9508, 5117]. Laborer [11524]. Laborers [2372]. Laboring [5059]. Lac [11984]. Lace [11939]. Lace-Bark [11939]. Lackawanna [10059]. Ladder [1072, 2713, 8906]. Ladders [4550]. Ladie [1818]. Ladle [11090]. Lady [6797, 10476]. Lafitte [1724]. Lagetto [11939]. Laird [6456]. Laisne [922]. Lake [3626, 11431, 3311, 5512]. Lame [5230]. Lammergeyer [8646]. Lamp [34, 1136, 2215, 2101, 3266, 3262, 3041, 5221, 7001, 7540, 7156, 8745, 11558, 12209, 12207, 12002, 12194, 13513, 13015, 12626, 13044, 13315, 13361, 13487]. Lampblack [12000]. Lamps [1278, 2599, 3137, 9824, 12197, 12332]. Lampwick [8959]. Lancashire [5280]. Lancets [10864]. Land [1346, 2148, 3441, 7647, 6821]. Landing [11735]. Landore [8964, 2361]. Lands [9810, 10004]. Langhorne [444]. Langley [11725]. Language [6029, 715]. Lantern [663, 32, 3508, 4072, 6224, 10323, 11403, 13147, 4830, 4861, 4892]. Lanterns [10884]. Laon [922]. Lap [11844]. Lapidary [170]. Lard [4990, 5717, 8027, 12418]. Large [2252, 4829, 4716, 4455, 6272, 6647, 10281, 10032, 11289, 11264, 11149, 5245]. Largest [1406, 1355, 6543, 9675, 12199, 13284]. Larynx [11198]. Lascell [1849]. Last [1890, 1805, 2061, 3419, 2445, 7508, 10061]. Lasting [668, 1662]. Latch [6141]. Late [2312, 1967, 2069, 2212, 2526, 5037, 6456, 12790]. Latent [4658, 2417, 2464, 2493, 2440]. Latest [3423, 3629, 3479, 4424, 6551, 9211, 10306, 13366, 2315]. Lathe [3686, 4546, 4377, 3912, 6105, 7789, 7255, 9280, 8304, 9278, 10913, 9956, 9729, 11980, 11290, 11343, 11732, 12316]. Lathes [5998, 5748, 9377, 9235, 8253, 10861, 9730, 11783, 13435, 13106]. Lathing [8189]. Latter [2571]. Lattice [12342]. Laughing [345, 3060, 12487]. Launch [4987, 5574, 6847, 9934, 12674]. Launches [10005, 2963]. Launching [635, 1674, 6197]. Laundries [12412]. Laundry [8445]. Laurent [11651]. Lava [5541]. Law [1148, 767, 249, 622, 18, 696, 1671, 1438, 1383, 1419, 1468, 1940, 2248, 2723, 2615, 2474, 3045, 3050, 3072, 3096, 3115, 3144, 3170, 3201, 3229, 3253, 3288, 3320, 3347, 3375, 3407, 3771, 2918, 5214, 4145, 6545, 5927, 5485, 7358, 6937, 7875, 6714, 8300, 8829, 10141, 10877, 11084, 11650, 11830, 11933, 11339, 11107, 2810, 3263, 11194]. Law-Decisions [11650]. Lawn [11380, 12675]. Lawrence [6429]. Laws [454, 401, 748, 201, 2022, 1595, 1564, 1913, 4205, 4172, 6738, 6962, 7073, 7605, 8990, 9237, 10900, 10904, 9693, 13403, 12640, 13405, 12848]. Laxey [3840]. Laying [69, 4575, 6879, 7759, 13310]. Lazy [3192]. Lead [770, 659, 1228, 1897, 2369, 4721, 3957, 10277, 12667]. Leader [3713]. Leading [466]. Leads [2757, 3042]. Leaf [2061, 763]. Leaf-Cutter [763].

Leaks [9918]. Leaping [4371]. Learn [3930]. Learning [9878]. Leasing [11590]. Leather [980, 1974, 2244, 3510, 2916, 4227, 4422, 9116, 8421, 8195, 10116, 13632, 68]. Leather-Dressing [8195]. Leaved [6881]. Leavenworth [4106]. Leaves [5900, 9234]. Lechner [11813]. Leclanche [88]. Lectularius [7021]. Lecture [1067, 2707, 2642, 3529, 3889, 3924, 3837, 3954, 3865, 4324, 4350, 8392, 10200]. Lectures [20, 21, 1192, 5392, 9137, 4382, 4295, 4323, 4410, 4351]. Leech [1873]. Left [5789, 9312, 10014]. Leg [1217]. Legal [11229, 12718]. Legged [6085, 12340]. Leggett [2145]. Legislation [1409, 2043]. Legislature [5951, 10142]. Legs [488]. Leguminous [11654]. Leipsic [4174]. Leisure [2247]. Leland [8934]. Lemman [2057]. Lemon [13008]. Lemuria [9769]. Length [8628]. Lens [4147, 12826, 11397]. Lenses [13298]. Leona [11698]. Leopard [4606]. Lescale [11915]. Lesseps [7]. Lesson [2197, 2907, 9247, 9825, 11234]. Lessons [10824]. Letter [1136, 2261, 3093, 3116, 3165, 3197, 3224, 3249, 3285, 3315, 3343, 3403, 3419, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4936, 4967, 5042, 4607, 4642, 4801, 5387, 11775, 13230, 4992, 5016]. Letter-Box [1136]. Letters [972, 1016, 1036, 1080, 1105, 1128, 3090, 3428, 2628, 4566, 8297, 11624]. Level [1541, 7406, 10004, 11579]. Leveler [9118]. Leveling [9250]. Lever [119, 424, 1937, 7852, 11685]. Leverrier [9335, 10659]. Lewis [2496, 6846]. Lewiston [6737]. Leyden [4120]. Liberty [11886]. Libraries [2905, 12434, 551]. Library [2939, 2463, 5898]. Libyan [6051]. Lice [11991]. License [12956, 12461]. Lick [5084]. Lid [8357]. Lidgerwood [9055]. Lieber [3484]. Liebig [728, 229, 2275, 4678, 4292, 4374, 2060]. Life [1246, 1697, 2065, 2061, 2113, 3595, 4431, 4986, 4543, 4615, 4102, 5044, 5632, 5878, 5929, 6275, 5661, 5733, 7068, 6805, 7533, 7102, 7437, 7898, 7356, 7150, 8763, 9177, 8485, 10573, 10613, 9958, 10565, 10321, 9985, 10081, 10768, 11142, 11233, 11245, 11949, 11420, 12583, 2060, 10867, 10911, 11836]. Life-History [10867, 10911]. Life-Preserving [6805, 7102, 9177, 8485]. Life-Saving [3595, 7533, 10081]. Lifeboats [6222]. Lift [7564, 8842, 13306, 12388]. Lifter [900, 6210, 8683, 13570]. Lifters [12337]. Lifting [1443, 4807, 4542, 8255, 8872, 8455, 11461]. Light [946, 257, 1225, 1777, 1846, 2009, 1593, 1970, 2343, 3658, 2965, 2986, 2707, 2815, 3247, 2748, 3503, 2591, 4148, 3913, 4615, 4614, 3889, 3802, 3837, 3865, 4632, 5544, 5734, 6106, 7287, 7900, 7754, 7216, 7820, 7762, 8478, 8654, 8963, 8763, 9200, 9274, 8281, 9464, 9049, 10563, 10604, 10601, 9826, 10805, 9906, 9854, 12116, 12157, 11475, 12067, 11135, 12155, 12088, 11036, 11394, 12232, 12305, 12627, 12628, 12876, 13118, 12463, 12583, 12979, 12315, 12553, 12392, 45, 2205, 11958, 413, 439]. Lighter [4604, 6303, 7001, 13589]. Lightest [254]. Lighthouse [7885]. Lighthouses [5067, 5641, 6630]. Lighting [110, 4072, 7123, 8333, 12159, 11199, 11921, 12832, 12538, 12788]. Lightning [1359, 1828, 1943, 2826, 3126, 3216, 4719, 4795, 5072, 5640, 6134, 7899, 7507, 7555, 8458, 9468, 9017, 8363, 9800, 10170, 9747, 9773, 9531, 13222, 13480, 13094]. Lights [1780, 6214, 6358]. Like [841]. Likely [1100]. Lily [10513]. Lime [4068, 6524, 11991, 13008]. Limes [1342]. Limit [1221, 5815, 12223]. Limits [546, 9465, 10944, 10192]. Linda [8297]. Line [2839, 3500, 4746, 5514, 9084, 9024, 9564, 12668, 12671, 13009]. Lined [1874, 10143, 10781]. Lines [6546, 8417, 12742]. Lining [9083, 11429]. Link [9149]. Linnæus [9341]. Linseed [1687]. Lions [10096, 8779]. Lip [395]. Lipped [10476]. Liquefaction [11082, 11157, 11214, 11094]. Liquefied [2649]. Liquid [227, 7757, 7907, 7549, 7611, 10151, 12933]. Liquids [6355, 11274, 2417]. Liquors [1226, 12473]. List [38, 60, 654, 683, 710, 737, 756, 780, 799, 824, 850, 867, 888, 913, 934, 954, 975, 999, 1019, 1040, 1061, 1085, 1110, 1132, 1154, 1179, 83, 106, 135, 165, 182, 210, 236, 260, 285, 306, 330, 356, 380, 408, 433, 460, 478, 498, 520, 542, 566, 589, 610, 630, 1903, 1205, 1234, 1884, 1927, 1952, 2005, 2027, 2052, 2076, 2094, 2117, 2138, 2159, 2181, 2208, 2233, 2256, 2282, 2305, 2331, 2356, 2378, 2405, 2428, 1258, 1284, 1311, 1338, 1365, 1388, 1421]. List [1449, 1473, 1503, 1527, 1559, 1582, 1630, 1655, 1679, 1707, 1732, 1759, 1785, 1813, 1838, 1863, 3354, 2453, 2480]. Lists [1981, 1607, 11465]. Literature [881]. Lithographic [1208, 10028]. Lithographs [9581]. Litigation [1303, 4951, 5598, 7211, 7071]. Little [7056, 7849, 8413, 9804, 11748, 13248, 9754]. Live [491, 5711, 10463, 10033, 11800, 13539]. Lively [1870]. Liver [1352]. Liverpool [2878, 12515]. Living [6186, 8827, 11990, 12666]. Livingstone [3368, 5409]. Lizard [10355, 9890]. Lizards [9317]. LL.D [939, 959]. Lloyd [5048, 7641]. Loader [4270]. Loading [26, 2079, 2009, 3397, 8957, 8998, 9018, 12006, 12774, 12712]. Loads [1005]. Local [3680, 4069, 11831]. Localization [13438]. Localizing [12936]. Locating [13222]. Lock [899, 528, 444, 1716, 2167, 3085, 3501, 4198, 4499, 4933, 5307, 6389, 10121, 11372, 12004, 13185, 13232, 13230]. Lock-Out [3085]. Lockage [280]. Locking [1049, 363, 321, 1217, 6141, 6361, 6972, 7002, 13306]. Locking-Time [13306]. Locks [896, 9594, 13306]. Locomotive [805, 714, 837, 2218, 3241, 3622, 3056, 4781, 4954, 5569, 6254, 5407, 5762, 5926, 6581, 6624, 7289, 7658, 7513, 9205, 9466, 8381, 8221, 8134, 9376, 9648, 10281, 10631, 9664, 12144, 12128, 12470, 12444, 12823, 10259, 485]. Locomotives [941, 2339, 1442, 2904, 2473, 4269, 3949, 6306, 5877, 7725, 7655, 6580, 7814, 9180, 9149, 8956, 9374, 10682, 10532, 10696, 12363]. Locust [6004, 6094, 12447, 9478, 9990, 9964, 10145]. Locusts [806]. Log [5070, 10384, 11955]. Loiseau [2504]. Lomaria [10668]. Lomellina [4774]. London [674, 1987, 2939, 12302, 396, 958, 154, 1946, 1931, 1670, 3619, 2543, 3475, 2641, 3748, 5225, 4614, 5655, 6583, 8477, 8706, 10408, 12631, 12515]. Long [1074, 560, 1055, 1215, 2909, 6056, 6339, 12579, 11119]. Long-Continued [6056]. Long-Necked [11119]. Longest [12846]. Lontin [10579]. Look [3191, 5242]. Looking [217, 242, 272]. Looking-Glasses [217, 242, 272]. Looks [10749]. Loom [1089, 1159, 3305, 3582, 2674, 9044, 10416, 10721]. Looms [8141]. Looper [8166]. Loose [3362, 6065, 10776, 11267]. Lordon [6224]. Loss [2098, 4372, 5189, 7899, 6647]. Losses [3827]. Lost [3984, 9721, 9769, 10034]. Loughbridge [8480]. Louis [1143, 836, 1563, 2311, 1771, 2433, 7410, 13641, 5073, 5241, 11905]. Louisiana [5, 1699, 2880]. Louisville [9565]. Lounge [2868, 5449, 6725]. Lovegrove [11037]. Low [938, 384, 1117, 2059, 2525, 4061, 4143, 4214, 7054, 8287, 9463, 10477]. Low-Water [938, 384]. Lowe [2216]. Lowell [155, 224, 282, 2690]. Lower [6426, 7986]. Lowering [6722, 8993, 11175, 13698]. Lowland [6682]. Lozenge [4932]. Lubricant [9679, 9888, 8774]. Lubricant-Testing [9679, 9888]. Lubricants [12852]. Lubricating [3362, 2984, 7754, 2366]. Lubricator [1548, 5020, 9467, 11350]. Lubricators [2370, 8876]. Lubricene [12065]. Lucetti [5613]. Lucida [11932]. Lucifer [98]. Luck [2061]. Ludlow [121]. Luedtge [13045]. Lumber [6970, 9492, 11166]. Luminous [9694]. Luna [1534]. Lunatic [1941, 4242]. Lunch [10795]. Lung [9176]. Luray [12367, 12564]. Lyall [9044]. Lyceum [512, 529, 1495, 2421]. Lycopodium [11959]. Lyman [1332]. Lymans [12075]. Lymph [1801]. Lynn [3986]. Lyons [3838]. Lyre [11814].

M [2145, 5104]. M. [6383, 9505, 10828, 11214, 11570, 11479, 11149, 11246, 11488, 11621, 11514]. Mabey [9614]. Macaroni [1171]. Macfie [622]. Mache [3080, 1740]. Machine [524, 988, 138, 721, 608, 764, 343, 829, 1003, 713, 964, 240, 668, 254, 65, 1114, 679, 1031, 740, 215, 293, 2032, 1735, 1802, 2244, 1796, 2147, 1998, 1794, 1808, 1402, 1452, 1867, 1567, 1432, 1790, 1323, 1795, 2186, 1457, 1208, 2394, 1710, 1186, 2413, 1764, 2349, 1939, 1319, 2691, 2769, 2706, 3295, 3391, 2718, 2887, 3446, 3207, 2821, 2791, 3259, 2684, 3493, 3753, 2757, 2916, 3103, 3068, 2756, 3040, 3353, 2785, 2585, 3294, 3242, 3133, 2997, 3651, 3617, 2497, 4302, 4703, 4797, 4991, 4779, 3969, 4522]. Machine [3958, 4421, 5140, 4603, 4925, 4314, 4156, 3886, 5143, 4339, 3850, 4931, 4091, 3911, 4688, 4213, 4142, 4808, 4638, 4799, 4345, 4301, 4722, 4847, 4951, 3859, 4064, 5198, 4184, 4427, 6209, 6112, 5770, 5721, 5552, 6227, 5893, 6155, 5418, 5973, 6064, 5666, 5773, 6240, 6079, 6277, 6104, 6410, 6251, 5536, 6010, 6388, 6163, 6278, 6293, 5877, 6522, 5470, 5333, 5708, 6349, 6560, 5688, 6092, 7254, 6605, 6919, 7609, 7490, 6571, 7451, 7309, 7714, 6978, 7336, 7613, 7910, 7379, 7052, 7692, 6799, 6918, 7906, 6963, 6716, 6628]. Machine [8454, 8869, 8762, 9136, 9142, 9472, 8302, 8872, 8050, 8081, 8277, 8656, 9012, 9424, 8421, 8195, 8189, 8928, 8270, 9025, 8512, 9234, 9020, 8808, 7936, 8571, 9339, 9345, 8732, 7931, 8413, 9247, 8296, 8096, 9509, 9402, 8832, 8930, 8354, 9395, 9089, 8051, 8082, 8996, 10316, 10955, 10938, 10082, 9704, 9980, 10766, 10832, 10802, 10012, 10171, 10285, 10403, 10638, 10314, 10116, 10152, 10634, 10501, 10508, 10796, 10064, 10937, 10827, 9672, 10579, 10439, 10499, 9755, 9612, 10806, 10028, 9888, 10639, 10694, 10850, 10094, 10432, 10060, 9910, 9692, 9588]. Machine [11319, 11073, 11130, 11612, 11228, 11375, 11145, 11418, 11423, 11349, 11813, 12182, 11246, 11095, 11805, 11114, 10986, 12131, 11884, 12078, 12102, 11652, 11237, 11066, 11671, 11368, 11168, 12092, 12799, 12858, 12448, 13178, 13286, 12685, 13081, 12882, 12987, 12562, 12704, 12532, 13125, 13078, 13409, 13384, 12967, 12615, 13158, 13458, 13676, 12489]. Machine-Made [13458]. Machine-Patent [679]. Machinery [904, 741, 745, 676, 892, 838, 225, 1964, 2066, 1378, 3387, 2758, 2515, 2924, 3443, 3858, 3917, 4187, 4371, 4401, 6247, 6161, 5792, 5322, 5976, 7621, 7911, 7838, 6692, 6991, 7418, 7881, 7182, 7839, 6582, 8522, 8661, 8368, 8986, 8387, 8491, 8513, 8019, 8780, 9076, 8109, 8933, 8987, 9364, 7942, 8004, 8116, 8172, 8525, 8551, 8576, 8667, 8752, 8783, 8974, 9034, 9067, 9354, 9416, 9446, 9486, 8158, 8660, 7991, 8102, 10958, 10258, 10462, 9670, 10468, 9550, 9678, 9805, 10632, 11501, 11549, 11538, 12046, 11753, 11338, 11634]. Machinery [11513, 11804, 11353, 12740, 13080, 12642, 13462, 12819]. Machines [855, 311, 418, 1592, 1238, 1962, 2383, 2410, 1241, 1331, 1326, 2936, 2822, 3265, 4676, 4893, 4863, 6562, 5606, 5385, 7557, 6741, 7286, 7391, 7754, 8044, 8075, 8305, 9085, 7924, 10308, 10496, 10576, 10916, 10725, 9884, 11977, 13508, 13308, 2122]. Machinists [6830]. Madder [56, 4694]. Made [963, 23, 350, 317, 749, 1171, 1003, 1646, 1368, 2288, 1331, 3296, 3236, 3208, 3691, 3021, 3530, 3332, 4401, 5602, 5824, 7669, 8478, 8688, 11476, 11982, 13331, 12778, 13458, 12366]. Madras [3801]. Magazine [881, 2009, 5794, 11061]. Magic [32, 9650, 10884, 13147, 4830, 4861, 4892]. Magnesia [6524]. Magnesian [11991]. Magnesium [2714]. Magnet [5187, 7346]. Magnetic [803, 1462, 2646, 2746, 2835, 3682, 5741, 6629, 9315, 8423, 8391, 12555, 12563, 12934, 13038]. Magnetic-Motor [12555]. Magnetism [762, 3267, 3337, 7517, 7382, 11074, 12644, 13364, 13386, 13408]. Magnetized [13072]. Magnetizing [13037]. Magneto [3040, 4345, 5581, 7245, 8381, 8305, 9960, 10579, 12182]. Magneto-Electric [3040, 4345, 7245, 8381, 8305, 10579, 12182]. Magneto-Electrical [5581]. Magneto-Photography [9960]. Magnetometer [13205]. Magnifying [5614]. Magnum [5577]. Magrath [1872]. Maguey [11937]. Mahan [2127]. Mai [5356]. Mail [3978, 3970, 6160, 6647, 6868, 9850, 9877, 11109]. Main [4427, 6092, 9110]. Maine [2414]. Mains [1822]. Maintained [705]. Maintenance [10769]. Make [1123, 3262, 3725, 3716, 6423, 6873, 10856, 10446, 10544, 813]. Maker [2899, 3811, 12699]. Makers [443]. Makes [6189]. Making [1075, 5, 474, 2214, 3642, 3356, 2515, 3619, 2771, 5155, 4421, 4103, 4353, 5061, 5008, 5608, 5633, 5984, 7609, 7872, 7638, 6918, 6795, 8457, 8027, 9300, 8019, 9214, 8385, 10955, 10338, 10508, 10827, 9532, 10564, 9678, 11549, 11538, 11612, 11591, 11829, 12018, 11116, 11773, 12740, 13134, 12591, 13080, 12735, 4855, 5302, 3870, 8719, 1120]. Malarial [2889, 10902]. Male [12618]. Malleable [3532]. Mallet [5332, 3780]. Malo [6007]. Malt [1689]. Mammalian [440]. Mammoth [2288, 1208, 4314, 4115, 10207, 13516, 599]. Man [502, 483, 661, 492, 1726, 1467, 2012, 3593, 3628, 4961, 6108, 6269, 6028, 7750, 8952, 8181, 10061, 9823, 10669, 11363, 11313, 13249, 3424, 4791, 8182, 12194, 12553]. Management [16, 619, 6219, 7047]. Manatee [7646]. Manchester [3531]. Mandrel [12590]. Manganese [44]. Manganiferous [3592]. Mangle [12828]. Manhattan [3188]. Manias [9593, 10114]. Manipulation [6823, 10014]. Manipulative [9692]. Mankind [553]. Manner [13134]. Manora [3211]. Mansion [5094]. Mantis [987, 10888]. Manual [11513]. Manufactories [2685, 2783, 2217]. Manufactory [290]. Manufacture [1100, 240, 1115, 698, 335, 513, 746, 114, 862, 940, 659, 992, 607, 1508, 2369, 2411, 1210, 1701, 1768, 1521, 2034, 2164, 1538, 2146, 2549, 2902, 3387, 3081, 3630, 3359, 3384, 3264, 3313, 3139, 2973, 5082, 4956, 4966, 4239, 4875, 4742, 4673, 4367, 5068, 5450, 5244, 5978, 6143, 5635, 7280, 6838, 6865, 7259, 9504, 7989, 8355, 8103, 9217, 10559, 10196, 10469, 10690, 10277, 10008, 10305, 10657, 9781, 12043, 11341, 11027, 11390, 11817, 12440, 13010, 13326, 12418, 12344, 12469, 13170, 12702, 12758, 5211, 10616, 8275, 143, 189, 95]. Manufacturer [7876, 12263]. Manufacturers [791, 748, 2198, 6830]. Manufactures [1947, 2859, 2626, 4837, 7690, 10464, 11185, 11831, 12071, 11861]. Manufacturing [614, 1855, 1994, 1592, 2114, 8347]. Manures [2152, 2627]. Many [2172, 11801, 11483]. Maps [144]. Mara [13082]. Marble [8482, 12793]. Marbleizing [2020]. Marcus [3542]. Mare [2391]. Marey [5852]. Margarin [9811]. Marginata [8400]. Marine [482, 691, 2111, 3704, 3364, 4833, 4265, 8967, 9053, 11398, 11041]. Mariner [1342]. Marino [11948]. Maritime [530]. Mark [1650, 2148, 2145, 2151, 2836, 6869, 12258, 12196, 12147, 13604]. Marker [4179, 6779, 11808]. Market [3188, 5844, 13539, 12414, 10480]. Markets [6034, 10526, 12067]. Marking [11645]. Markrud [8332]. Marks [2615, 7359, 7097, 6935, 12042, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 13631, 13692, 13491]. Marmoset [10206, 10991]. Marquard [2203]. Marriage [3400]. Marriages [8473]. Mars [4884, 4908, 10498, 10729]. Marsh [2714, 2740]. Martin [9967]. Marvels [5570]. Mashing [764]. masking [5799]. Mass [155, 1271, 1291, 12089]. Massachusetts [2621, 323, 1829, 3603, 2465, 2689, 2789, 2884, 5776, 9022, 10807, 11580]. Massey [2059]. Master [2200, 2502, 8621]. Masters [6404]. Mastersun [3156]. Mastodon [2614]. Match [597, 5120, 4939, 9119]. Matcher [8988, 12559]. Matches [98, 702]. Matching [6571]. Maté [216]. Material [1481, 2268, 3385]. Materialism [9170]. Materials [2245, 4396, 4103, 5999, 6874, 9091, 11954, 5750]. Mathematics [4770]. Mathews [12729]. Matson [5615]. Matter [726, 843, 1377, 6597, 9691, 11437, 12314, 12383, 2715]. Matterhorn [2081]. Matters [253, 7362, 8132, 11665, 11392]. Matthew [4004, 12020]. Mattieu [1400]. Mattress [193, 28, 6805]. Mattresses [1488, 1754]. Maury [4004]. Maxims [1091]. Maxwell [1594]. May [695, 703, 1852, 3827, 11543, 8181, 5851, 9118]. Mayer [10876]. Mc’Kennan [2168]. McMillan [8711]. McNeil [465]. Me [7505]. Meal [8254, 143].

Means [3399, 11206, 11450, 13404, 8222, 8250, 4830, 4861, 4892]. Measure [1174, 3039, 9968, 10151, 12461]. Measured [4772, 8394]. Measurement [1012, 373, 3058, 5213, 4564, 6295]. Measures [8925]. Measuring [456, 149, 1702, 1823, 1427, 2841, 2725, 2812, 5121, 5228, 4076, 6874, 6746, 6653, 11274, 11733, 12104, 12058, 13652]. Meat [666, 728, 526, 194, 2666, 7031, 7233, 9991]. Meats [53, 1055, 8864]. Mechanic [1046, 2316, 1273, 2502, 6471, 7876, 10308, 12263, 1008]. Mechanical [879, 692, 470, 299, 447, 585, 724, 930, 96, 601, 534, 2415, 2245, 1324, 2395, 1437, 1962, 1249, 1850, 1769, 1915, 2011, 1222, 2182, 2365, 2886, 3511, 2772, 2521, 3531, 3632, 3681, 4571, 3901, 6448, 6561, 5904, 6243, 7842, 7850, 7704, 7738, 7770, 7798, 7829, 7859, 7887, 7917, 6993, 8482, 7957, 8681, 8863, 7945, 7970, 8000, 8031, 8059, 8090, 8263, 8404, 8437, 8466, 8494, 8527, 8552, 8638, 8754, 8786, 8940, 8972, 9004, 9099, 9222, 9325, 9383, 9448, 9483, 9516, 8768, 8799, 9148, 8281, 9048, 10014, 10824]. Mechanical [9828, 10226, 9543, 9600, 10422, 10453, 10485, 10516, 10549, 10582, 10618, 10649, 10672, 10704, 10734, 10760, 10783, 10810, 10842, 9626, 9736, 9760, 9790, 9812, 9841, 9869, 9945, 9974, 9997, 10020, 10048, 10072, 10100, 10129, 10156, 10180, 10213, 10239, 10270, 10293, 10329, 10360, 10396, 11729, 12032, 11198, 12149, 11779, 11006, 11417, 11922, 12697, 13388, 13406, 13433, 13461, 13488, 13521, 13538, 13563, 13597, 13615, 13644, 13677, 13700, 13255, 13289, 13312, 13363, 12593, 12703, 12739, 12794, 12861, 12941, 12970, 12992, 13020, 13048, 13074, 13102, 13153, 13188, 13207]. Mechanical [13234, 1006, 12419]. Mechanically [1648, 3338, 1715]. Mechanics [2415, 2325, 2372, 1199, 1299, 2169, 1464, 1543, 4928, 5570, 8621, 10529, 12255, 12231, 12288, 11473, 12389, 12468, 12588, 12695, 12912, 13018, 13077, 13130, 13196, 12089]. Mechanism [2122, 6302, 8773, 8653, 11811, 11703, 5825, 6063, 5853, 5879, 5934, 6009, 6115, 6164, 6223, 6276, 6334, 6387, 6431, 6501, 6559, 6947, 6626, 6679, 6724, 6777, 6826, 6877, 7003, 7080, 7132, 7188, 7267, 7310, 7365, 7415, 7463, 7512, 7560, 7612, 7670, 7726, 7788, 7851, 7908, 7962, 8022, 8105, 8194, 8252, 8306, 8360, 8426, 8490, 8542, 8598, 8655, 8742, 8805, 8870, 8932, 9023, 9147, 9179, 9248, 9313, 9375, 9436, 9537, 9591, 9646, 9702, 9753, 9807, 9860, 9939, 10010, 10067, 10120, 10174, 10233, 10284, 10442, 10537, 10664, 10831, 10885, 11040, 11216]. Medal [2674, 4836]. Medals [151]. Meddlers [11526]. Media [11142]. Mediæval [9145, 8900]. Medical [1409, 4368, 5064, 6082, 5884, 5939, 5964, 6036, 6138, 6230, 6282, 7314, 7855, 8227, 9667, 9608, 11105, 3390]. Medicament [8193]. Medicaments [8422]. Medicinal [11894]. Medicine [2225, 11715, 13014]. Medicines [4264, 10527]. Medium [10805, 94]. Meeting [674, 3368, 2621, 4997, 5691, 7813, 9275, 9311, 10774]. Megaphone [11827]. Meidinger [7057]. Meigs [7820]. Mein [7815]. Mekarski [12387]. Melbourne [1957, 526]. Melendy [810, 2768]. Melon [7730]. Melter [5190]. Melting [2741, 5986]. Member [5162, 5601]. Members [5817, 9890]. Memorial [5599]. Memorials [12103]. Memory [13278]. Men [1827, 1973, 3804, 4095, 6448, 6404, 9079, 8706, 7981, 10688, 13247, 2149, 7874, 8983, 8012]. Menace [12606]. Mending [7603, 8364]. Menhaden [1215]. Menlo [13128].

Mental [812, 1750, 1895, 5109, 5544, 5974, 11924, 12874]. Mentone [4961]. Mercantile [9480]. Mercator [12921]. Merchandise [6756]. Merchant [8449]. Merchants [12989]. Mercurial [969, 749, 1161, 7539, 9204, 11499]. Mercury [8864, 5472, 11543, 6971, 11499]. Mercy [8096]. Mere [11572]. Meridian [13219, 10978]. Meritens [12182]. Merits [9341]. Merriman [1246, 7102]. Mershon [10122]. Merthyr [5164]. Mery [11017]. Mesembryanthemums [7492]. Mesquite [11871]. Message [3740]. Messrs. [6039]. Metal [871, 490, 1096, 2013, 3014, 2773, 3881, 4322, 6064, 6293, 8739, 9025, 8512, 8274, 10314, 10639, 9885, 9778, 9831, 11761]. Metal-Punching [9025]. Metal-Testing [8274]. Metaline [5277, 6739, 8657]. Metallic [731, 448, 765, 579, 1392, 1874, 3880, 4380, 4602, 7615, 11092, 13460, 13311]. Metallurgical [6133]. Metallurgy [1405, 2199, 3920, 5897, 7217, 6843]. Metals [452, 1970, 1770, 2695, 2460, 2834, 2756, 3588, 4367, 5106, 10798, 10601, 10876, 12584, 3043, 6532, 5575, 5305, 6871]. Meteoric [4720, 10600]. Meteorite [10683]. Meteorites [3017, 8936]. Meteorological [1093, 6082]. Meteorology [2444]. Meteors [464, 3178, 3191, 3283, 5086, 10718, 12696]. Meter [2019, 2193, 5955, 5525, 7584, 13492, 13098]. Meters [673, 1012, 4062, 7907]. Method [511, 1047, 268, 2345, 1893, 1542, 2013, 1648, 2590, 2807, 3780, 3749, 3441, 3473, 3533, 4078, 4201, 5473, 5436, 7280, 7695, 7465, 7668, 6653, 8249, 8429, 8459, 8803, 9343, 8364, 10943, 10525, 10744, 10220, 10256, 10203, 10876, 10985, 11166, 11960, 11175, 13222, 12373, 8222, 8250, 8632, 8719, 11590, 1715]. Methods [5938, 10936]. Metric [51, 1686, 4063, 6719]. Metrological [5571]. Metropolis [2113, 11109]. Metropolitan [552, 1634, 3972, 12128, 12791]. Metz [893]. Mew [10220]. Mexican [7980, 11599, 12906, 12630, 13545]. Mexico [6055, 12236, 4855]. Mezzana [6006]. Mezzana-Corti [6006]. M’Gee [2194]. M’Grath [4463]. Miasma [10223]. Mica [6517, 8386]. Michael [271, 102, 374]. Michigan [6894]. Micro [11609, 11916]. Micro-Tasimeter [11609]. Micro-Telephone [11916]. Micrometer [9538, 10385]. Microphone [11615, 11680, 12011, 12133]. Microphones [12884]. Microscope [1167, 1517, 4039, 6159, 7669, 9138, 11397, 9348]. Microscopes [9883]. Microscopic [5856, 5406]. Microscopical [174]. Microscopy [9984, 11617]. Middle [2668]. Middlings [8569, 11910, 11007]. Might [9399]. Migration [12577]. Migrations [9637]. Migratory [806, 11324]. Mild [99, 12149]. Mildew [8351]. Mile [231]. Mileage [909]. Miles [560, 6005, 13276]. Military [832, 1507, 8747, 10614]. Milk [842, 72, 174, 1508, 1616, 2421, 3299, 4700, 4502, 6357, 5766, 6454, 7929, 8326, 11687, 12824, 13052, 13333]. Milker [6300, 9703]. Milking [1162, 7341, 8322, 12092]. Milking-Stool [1162]. Mill [2106, 3422, 2988, 2782, 5018, 7261, 7616, 7065, 7992, 10858, 10829, 10773, 10859, 9673, 11169, 11476, 12183, 11042, 11117, 11297, 11453, 11289, 11757, 11046, 12523, 13179, 13131, 12302, 13511, 13157, 12364]. Millener [2713]. Millennium [841]. Miller [840, 7909, 8428]. Millers [12611]. Milling [6349, 9235, 8732, 10064, 12987, 13106]. Millington [11241]. Million [3535, 3705, 5084, 11245]. Millionaire [8449]. Mills [1226, 3720, 4775, 7509, 7671, 10805, 11512, 13434, 12613]. Millstone [8307, 9457, 10986]. Millstones [311, 246, 269, 11682]. Millwright [840]. Mind [482, 1518, 8272, 9635, 11634, 11326, 11354, 11293, 11400, 11459, 11622]. Minds [11498]. Mine [1659, 4132, 10692]. Miner [4429]. Mineral [726, 1497, 2489, 4118, 11429, 11013, 12490, 362]. Mineralogy [1175]. Minerals [4663, 5954, 9285]. Miners [6722]. Mines [110, 577, 5954, 6517, 7785, 7900, 10687, 10198, 12050, 5359, 1439, 13482]. Miniature [5139, 8860]. Mining [1057, 1598, 2688, 2641, 2976, 4395, 6256, 6625, 6851, 8020, 10474, 12099, 11424, 11813, 11883, 11993, 11951, 11580, 11904, 12309, 4351]. Minn [2735]. Minnesota [6004]. Minute [231]. Miracles [4824]. Mirrors [77, 12805]. Misapplied [8708]. Miscellaneous [3250, 3304, 7764, 7827, 7707, 7736, 7769, 7796, 7832, 7860, 7890, 9246, 8033, 8057, 8973, 9031, 9063, 8088, 8118, 8143, 8174, 8203, 8231, 8260, 8314, 8339, 8372, 8497, 8548, 8577, 8607, 8664, 8695, 8724, 8784, 8815, 8848, 8880, 8911, 8942, 9126, 9161, 9190, 9225, 9258, 9289, 9355, 9384, 9417, 9449, 9484, 9515, 9544, 9568, 10271, 10298, 10331, 10361, 10392, 10423, 10454, 10486, 10517, 10550, 10584, 10619, 10650, 10675, 10705, 10732, 10757, 10784, 10811, 10841, 9598, 9625, 9656, 9682, 9735, 9761, 9789, 9813, 9840]. Miscellaneous [9866, 9898, 9925, 9944, 9973, 9995, 10019, 10047, 10071, 10103, 10130, 10158, 10181, 10212, 10241, 12500, 12534, 13411, 13436, 13471, 13490, 13517, 13535, 13566, 13587, 13612, 13638, 13667, 13695, 12682, 12707, 12801, 12820, 12890, 12911, 13019, 13079, 13107, 13187, 13210, 13228, 13258, 13280, 13318, 13339, 13367, 13385]. Misconduct [816]. Misfortune [2202]. Misjudged [8678]. Miss [1842, 12331, 12672]. Missing [4878]. Mission [6895]. Mississippi [2735, 6994, 8097, 9238, 10309, 11474, 13425, 13146, 12265, 9478]. Missouri [1799, 1497, 4106, 5438]. Mistake [9201]. Mistakes [813, 12905]. Mitchell [1638, 5597, 11835]. Miter [5773, 10638, 13256]. Mitering [5666]. Mitigate [124]. Mitrailleur [1289]. Mitrailleuse [776, 4209, 4408]. Mitts [4122]. Mixer [7549, 10413]. Mixing [1592, 3982, 5744]. Mixtures [1103, 161, 5925]. Mo [1970, 1799, 8247]. Mo-Tay [1970]. Mo. [4806]. Moabite [1710]. Moat [6389]. Mode [3035, 4249, 4275, 4306, 4331, 4478, 4833, 4520, 4265, 6722, 6942, 8282, 10565, 9691, 9918, 11477, 12277, 11106, 11273, 4218]. Model [2238, 5788, 7271, 7565, 9213, 12604]. Models [3366, 2594, 6868, 10685, 10660, 11187, 11260]. Modern [596, 113, 901, 1046, 1725, 2085, 1877, 3387, 3773, 4824, 4401, 4258, 5641, 5722, 6090, 5334, 5439, 7655, 7631, 6982, 7178, 8893, 9579, 9608, 12201, 11111, 12874, 12959, 2493, 10699]. Modes [128, 6599]. Modification [201, 13487]. Modifications [12133]. Modulus [12047]. Moist [10828]. Moistening [5230]. Moisture [7749, 2314]. Molasses [720]. Mold [3261, 5088]. Molder [1684]. Molding [1114, 215, 3454, 6293, 6560, 7490, 8803]. Moldings [4234, 8270]. Mole [10357]. Molecular [575, 3352, 3564, 5063, 12849, 12937, 12960, 13096, 13236, 6871]. Molecules [12930]. Moles [4851, 10389]. Moll [8626]. Molten [13037]. Monaco [11948]. Money [789, 3691, 3725, 3811, 6992, 7841, 8768, 8505]. Monitor [3184, 8413, 10262]. Monopolies [52, 2388, 1274, 5461]. Monopolists [5708]. Monopoly [3859, 4064, 6546, 8302]. Monson [5419]. Monster [10356, 13679]. Monsters [11487]. Monstrosity [1033, 11123]. Mont [1272, 1238, 2126, 1223, 2541, 6466, 2687]. Month [351, 2631]. Months [1805, 1193, 2445, 2468, 13007]. Monument [2211, 2289]. Moon [3479, 8508, 9013, 8515, 8982, 12250, 11562, 11256, 12763, 2491]. Moonlight [4323]. Moore [7816]. Mop [1163, 3620]. Moral [1053, 1548]. Morbid [5974, 11498]. Mordants [4568]. Morgan [6602]. Morin [452]. Morse [43, 1798, 1807, 1842, 2823, 2837, 1119]. Mortar [1593, 3108, 6143, 11991]. Mortising [1802, 7692, 8928, 12131]. Morton [3712, 12088]. Moscow [3482, 3809]. Mosquito [3083, 11532]. Moss [6339]. Most [2394, 1355]. Motay [3855]. Moth [503, 4606, 4640, 7440, 7084, 11148, 13595, 1534]. Mother [4596, 11928]. Moths [8351, 8875, 617, 11261]. Motion [1089, 1159, 575, 1098, 1116, 1137, 1165, 1639, 1187, 1212, 1242, 1266, 1296, 1320, 1347, 1375, 1399, 1430, 3352, 2674, 3472, 3583, 3647, 3687, 3719, 2564, 4857, 3922, 5922, 5521, 5822, 6060, 7075, 6995, 7304, 9044, 9338, 9856, 10565, 9691, 11477, 11106, 12024, 12119, 11542, 13316, 13666, 12496]. Motive [1076, 2109, 5763, 7216, 7105, 13430]. Motor [1462, 1747, 1806, 4753, 4882, 3916, 5657, 6010, 6140, 5965, 6878, 6745, 7209, 7235, 7257, 7285, 7302, 7357, 7432, 7661, 7611, 6629, 8503, 8866, 8995, 8623, 9464, 9551, 10481, 10438, 9829, 11363, 12169, 12555, 12563, 12551, 12331, 12672, 12909, 13492, 13609, 12387, 773]. Motors [1051, 1944, 1998, 3152, 11559, 12156, 12934]. Mould [10345]. Moulding [9672, 11353, 12882]. Mount [1114, 5518, 11539]. Mountain [805, 5776, 7747, 9723, 12610, 361, 9964]. Mountains [6746, 7688, 8955]. Moustache [2762]. Mouth [427, 6994, 8097]. Mouths [601]. Movable [5468, 7004]. Movement [1024, 1962, 1915, 5733, 5904, 8799, 9148, 11554]. Movements [299, 447, 585, 724, 930, 4346, 9048, 10642]. Movers [5764, 5268, 5298]. Moving [70, 1819, 6995]. Mower [2309, 4105, 12675]. Mowers [6850, 11380, 11608]. Mowing [4603, 7418]. Mr. [359, 255, 1890, 1914, 1464, 2112, 3498, 2474, 2472, 7508, 6776, 7956, 7984, 12054, 11667, 12608, 12671, 8112]. Mt [11981]. Much [2343, 3275]. Mucilage [8397]. Mud [4927, 9394, 11035, 13555]. Muff [8280]. Mule [774, 10635]. Muller [6303]. Multiple [8296, 11009]. Multiplication [6632, 13344]. Multiplier [393, 5688]. Multiplying [8811]. Munchausen [2722]. Munificent [4206]. Munson [5640]. Muntz [8592]. Murchison [2250]. Murder [1376]. Murraysville [13606]. Muscle [11634]. Muscular [5733, 10466]. Museum [467, 552, 1987, 1493, 3972, 5094, 5848, 6119, 9347, 12791, 2939, 11361]. Museums [2770, 12791]. Mushroom [8184]. Mushrooms [9859]. Music [3474, 9554]. Musical [2971, 3180, 3058, 4059, 9241, 8968, 8604, 11811, 11869, 11074, 12318, 13592, 2267]. Musk [12313, 12860]. Mussel [12919, 1618]. Mutual [9580]. Muzzle [10172]. Muzzling [646]. My [2098, 2568, 2608, 2634, 2657, 4189, 5620, 6095, 6172, 5424, 5474, 6014]. Myers [6400]. Mygales [8693]. Mysteries [2422]. Mysterious [10603, 10201, 11083, 13522]. Mystery [580, 3885, 4345, 7693].

N [11171]. N.J. [2011, 10059]. N.Y [626, 6737]. N.Y. [244, 1306, 3044, 2808, 4159, 7807]. Nail [4187, 4183, 6554, 7466, 9206, 10634]. Nails [341]. Names [11483]. Naphthaline [1302]. Napper [7243]. Napping [3259]. Narica [13186]. Narrow [2083, 1440, 1833, 1971, 1298, 3136]. Narwhal [390]. Nascent [3899]. Nash [8485]. Nation [3125, 12071]. National [561, 2197, 7066, 8130, 9851, 11448, 13173, 5691, 6451, 5719, 7100, 9275, 9311, 8488, 8402, 12160, 12281, 11235, 13544, 13556, 13608]. Native [665, 2655, 7576]. Natural [467, 512, 529, 794, 1495, 2421, 1493, 2542, 2553, 5848, 7258, 8867, 10944, 9642, 11028, 11184, 12749, 12862, 12998, 13133, 13160, 13261, 13415, 13571, 13617, 13648, 340, 3154, 9435]. Naturalist [5387]. Nature [502, 605, 97, 1521, 3423, 3352, 2946, 3141, 4776, 4345, 7403, 8487, 10221, 8028, 94, 8112, 372, 6819]. Nautilus [12234]. Naval [11164, 945, 6603, 6621, 8324, 10306, 11500]. Navigable [2736, 5932]. Navigation [57, 1596, 1718, 1921, 1394, 4293, 3974, 6648, 9825]. Navvy [10191]. Navy [2601, 3975, 3808, 5243]. Near [2543, 4106, 6542, 10138, 11136, 4891]. Nebraska [4205]. Nebula [10683]. Nebulæ [5147]. Nebular [2938, 7996]. Necessary [750, 124, 10436]. Necessities [2085, 1697, 11500]. Necessity [1053, 1878, 4596, 4512]. Neck [1744, 2811, 12494]. Neck-Yoke [1744]. Necked [5122, 12596, 11119]. Necklace [10477]. Needed [3275, 3595, 3364, 6545, 6795, 7605, 8045, 8076, 11829, 11059, 13272]. Needham [11869]. Needle [855, 744, 1714, 2521, 4466, 5937, 5548, 10576, 11300, 11641, 11735]. Needles [8309, 10469, 95]. Negatives [13329]. Neglected [5631, 8184, 8100, 11635, 12117]. Neighbor [8679]. Nelaton [5037]. Nelson [1894, 3776, 5251]. Neomorpha [11961]. Neplus [5794]. Nerve [5610, 5543, 9915]. Nervous [5404]. Nessle [7264, 7313]. Nest [2391, 8717, 11997]. Nest-Building [8717]. Nests [6294]. Net [11532]. Neuralgic [10031]. Neuropters [1934]. Neutral [12668, 12671]. Nevada [9369, 10687]. Never [12790]. Newark [3095, 3342, 6912]. Newfoundland [5549]. Newly [8028]. Newport [6132]. News [338, 3479, 3505, 6936, 9211, 10906]. Newspaper [3220, 4401, 5794]. Newspapers [24, 10796]. Next [13171]. Niagara [3195, 5720, 6737, 9909]. Nickel [691, 2461, 5170, 8991, 11315, 11887, 8632, 9892]. Nickel-Plating [691]. Nicolai [1960]. Nicolaieff [12560]. Night [5765, 7554, 12207]. Niin [196, 139]. Nile [3929, 11197]. Niles [9431]. Nine [3309, 3867]. Ninety [769]. Ninth [583]. Nippers [4428, 12257]. Nitric [1975]. Nitro [1074, 772, 7360].

Nitro-Glycerin [1074, 772, 7360]. Nitrogen [11691]. Nitroglycerin [8568]. Nitrous [3060]. No [1890, 1914, 1903, 4979, 8129, 11070, 11925, 12462, 12483, 12514, 12574, 12609, 12787, 12818, 12851, 12873, 12908, 13485, 13519, 13529, 13590, 13611, 13643, 13671, 13699, 12849, 12937, 12960, 13096, 13236, 12552, 1736, 1767, 1791, 1869, 13364, 13386, 13408, 12333, 12355, 12384, 12432]. No.4 [1843]. Nobleman [2953]. Noel [8108]. Noise [12048]. Noisome [1625]. Non [1239, 4044, 6378, 6303, 6824, 6573, 10908, 7880]. Non-Conductor [6573]. Non-corroding [1239]. Non-Explosive [6378, 6303]. Non-Freezing [6824]. Non-Radiating [4044]. Non-Recoil [10908]. Non-Retreating [7880]. Nonsense [138, 2083]. Nordenskjold [7982, 13226]. Normal [6523]. North [3527, 4491, 6517, 10915, 13066]. North-Northwest [13066]. Northamptonshire [12103]. Northeast [12955]. Northern [2735]. Northward [13379]. Northwest [2222, 13066, 13331]. Norway [4750, 8356]. Norwegian [9863]. Nose [607, 3653, 6170]. Note [12146, 12671, 4965, 5245, 11835]. Notes [23, 1003, 658, 687, 929, 2383, 2410, 1901, 1183, 1288, 1425, 3676, 3688, 3212, 2431, 2456, 3053, 3076, 3101, 3121, 3147, 3173, 3204, 3232, 3257, 3290, 3323, 3350, 3380, 3411, 3438, 3465, 3519, 3553, 3579, 3615, 3640, 3672, 3701, 3737, 2484, 2512, 2538, 2560, 2582, 2611, 2637, 2662, 2680, 2704, 2732, 2754, 2779, 2805, 2831, 2854, 2875, 2896, 2913, 2933, 2959, 2980, 3003, 3029, 2432, 3032, 3699, 3734, 3193, 3956, 4074, 4181, 4328, 4436, 4544, 5050, 5146, 4303, 4325, 3898, 4012, 4679, 3798, 3824, 4591, 4627]. Notes [4655, 4714, 4740, 4765, 4788, 4818, 4871, 4900, 4921, 4948, 4974, 5030, 5056, 5080, 5102, 3878, 3910, 3939, 3967, 3996, 4025, 4054, 4089, 4113, 4140, 4167, 4196, 4223, 4255, 4280, 4312, 4337, 4365, 4392, 4419, 4448, 4481, 4539, 4560, 4545, 4573, 4601, 5223, 4435, 4383, 4464, 4502, 4526, 5253, 5386, 5495, 5585, 5618, 5690, 5801, 5933, 6386, 6475, 5551, 5884, 5939, 5964, 6036, 6138, 6230, 6282, 5566, 5596, 6322, 5628, 5650, 5680, 5705, 5731, 5761, 5785, 5813, 5841, 5868, 5892, 5919, 5948, 5997, 6024, 6048, 6078]. Notes [6129, 6154, 6182, 6208, 6239, 6266, 6292, 6584, 6705, 6827, 6920, 7034, 7131, 7162, 7265, 7698, 7826, 7566, 7314, 6594, 6617, 7252, 7278, 7298, 7323, 7353, 7377, 7400, 7425, 7449, 7477, 7500, 7523, 7547, 7573, 7599, 7629, 7653, 7682, 7712, 7743, 7776, 7805, 7836, 7866, 7894, 6640, 6667, 6690, 6711, 6735, 6763, 6789, 6814, 6836, 6863, 6888, 6905, 6932, 6959, 6989, 7013, 7041, 7063, 7090, 7121, 7146, 7171, 7202, 7229, 7204, 7960, 8111, 8251, 8367, 8489, 8627, 8749, 9287, 9443, 9028, 9123, 9152, 9184, 9253, 9286]. Notes [9319, 9349, 9380, 9411, 9442, 9479, 9019, 9255, 8603, 7922, 7950, 8672, 8700, 8729, 8759, 8789, 8821, 8857, 8889, 8917, 8948, 8978, 9009, 9041, 9073, 9107, 9133, 9167, 9198, 9231, 9265, 9298, 9331, 9361, 9391, 9422, 9454, 7978, 9490, 8009, 8039, 8065, 8070, 8124, 8151, 8178, 8209, 8239, 8268, 8294, 8319, 8343, 8379, 8411, 8442, 8469, 8501, 8532, 8557, 8584, 8614, 8643, 9091, 8534, 8559, 9456, 8586, 8616, 8645, 8674, 8702, 8731, 8761, 8791, 8823, 8859, 8891, 8919, 8950, 8980, 9011, 9043, 9075, 9109, 9135]. Notes [9233, 9267, 9363, 9393, 9423, 9153, 9185, 9539, 9653, 9777, 9891, 9992, 10091, 10235, 10381, 10472, 10644, 10755, 10891, 10800, 10826, 10903, 10970, 10340, 10465, 10512, 10630, 10661, 10689, 10770, 10879, 9953, 10217, 10817, 9522, 9547, 10274, 10302, 10335, 10367, 10400, 10429, 10459, 10493, 10521, 10556, 10590, 10622, 10653, 10679, 10709, 10737, 10763, 10791, 10847, 10872, 10897, 10923, 10950, 9574, 9604, 9630, 9661, 9688, 9715, 9740, 9764, 9794, 9819, 9847, 9873, 9902, 9931, 9977, 10001, 10025, 10056, 10078, 10107, 10135, 10162, 10188, 10245, 9549, 9575, 10276]. Notes [10304, 10337, 10369, 10402, 10431, 10461, 10495, 10523, 10558, 10592, 10624, 10655, 10681, 10711, 10739, 10765, 10793, 10819, 10849, 10874, 10899, 10924, 10953, 9606, 9632, 9663, 9689, 9717, 9742, 9766, 9796, 9821, 9875, 9904, 9933, 9955, 9979, 10003, 10027, 10058, 10080, 10109, 10137, 10164, 10190, 10219, 10247, 10994, 11010, 11644, 11084, 11650, 11933, 11303, 11328, 11571, 11589, 11754, 11953, 11053, 11077, 11180, 11252, 11278, 11656, 12165, 10974, 11002, 11769, 11793, 11821, 11852, 11876, 11899, 11919, 11946, 11965, 11986, 12016, 12039, 12062, 12085, 12111]. Notes [12140, 12191, 12218, 12243, 12273, 11024, 11101, 11127, 11154, 11202, 11225, 11307, 11335, 11358, 11387, 11414, 11444, 11469, 11492, 11520, 11546, 11576, 11603, 11630, 11677, 11700, 11720, 11745, 10976, 11004, 11771, 11795, 11823, 11854, 11878, 12041, 12064, 12113, 12142, 12167, 12193, 12220, 12275, 11026, 11055, 11079, 11103, 11129, 11156, 11182, 11204, 11227, 11280, 11309, 11337, 11360, 11389, 11416, 11445, 11470, 11494, 11521, 11548, 11578, 11604, 11632, 11747, 13338, 12764, 12309, 12702, 12758, 12749, 12862, 12998, 13133, 13160, 13261, 13415, 13571, 13617, 13648, 12571, 12869, 12900]. Notes [12926, 12953, 12977, 13003, 12299, 12326, 13193, 13216, 13243, 13268, 13295, 13323, 13351, 13375, 13396, 13422, 13452, 13477, 13502, 13526, 13550, 13576, 13601, 13627, 13656, 13683, 12351, 12379, 12402, 12429, 12456, 12480, 12511, 12547, 12600, 12636, 12662, 12692, 12724, 12755, 12784, 12811, 12842, 13032, 13061, 13088, 13115, 13140, 13167, 12786, 12301, 12328, 13195, 13218, 13245, 13270, 13297, 13325, 13353, 13377, 13398, 13424, 13454, 13479, 13504, 13528, 13552, 13578, 13603, 13629, 13658, 13685, 12353, 12381, 12404, 12431, 12482, 12513, 12549, 12573, 12602, 12638, 12664, 12694]. Notes [12726, 12757, 12813, 12844, 12928, 13005, 13034, 13063, 13090, 13117, 13142, 13169, 12796, 12913, 13103, 12727, 11534, 2039, 2089, 2342, 1091, 5117, 12132, 11086, 11124, 11846, 11866, 11891, 11917, 11941, 11962, 11979, 12010, 12029, 12077, 12098, 12212, 12237, 12266, 12290, 11176, 11200, 11223, 11247, 11263, 11288, 11331, 11348, 11382, 11406, 11427, 11456, 11517, 11528, 11568, 11588, 11638, 11669, 11695, 11713, 11739, 11765, 11818, 11789]. Notice [4563, 4594, 4630, 9850, 9877, 9719, 9744, 9768, 9799, 9837]. Notices [1551]. Novel [27, 3712, 3035, 3505, 3757, 4833, 4610, 4011, 4070, 4265, 4266, 4058, 5987, 5962, 6228, 5959, 6454, 5904, 5937, 6406, 6946, 6917, 7055, 7639, 7186, 9474, 8713, 8715, 9664, 10321, 9985, 11992, 11888, 11623, 12158, 11913, 11045, 12390, 12744, 13510, 12443, 12338, 13467, 12526, 13316, 12648, 13340, 12909, 12442, 12768, 13049, 12464, 13124, 13561, 13645, 13666, 13231]. Novels [12716]. Novelties [12853]. Novelty [2110, 5681]. November [3710, 5086, 10718]. Nozzle [3893, 8346, 8453, 11141, 13670]. Nuisance [8357, 11658, 12815]. Nullify [13403]. Number [2252, 2349, 1900, 8504, 10911, 5825, 6063, 6947]. Numbers [9502]. Nursery [5539]. Nursing [7455]. Nut [444, 5307, 6277, 6972, 10121, 11372, 11070, 12004, 13185]. Nut-Lock [444]. Nutkaensis [7618]. Nuts [899, 321, 8160, 11267].

Oak [1396, 1478, 6033, 5356, 13346]. Oakes [1252]. Oakley [5523]. Obelisk [10713]. Obelisks [10745, 10836]. Obituary [369, 466, 1306, 1332, 1551, 5299]. Object [1250, 4768]. Objection [4455, 12956]. Objections [2018, 13039]. Objects [2251, 7961]. Obligations [2923, 9553]. Obnoxious [7875, 8300, 8829]. Observation [7067, 11903]. Observations [4153, 7178, 8766, 10718, 10436, 11812, 9964, 11543]. Observatory [4238, 12488, 426, 2492, 8126, 10978, 12281, 11164]. Obstructions [1623, 1658]. Obtain [171, 3667, 3696, 3731, 3763, 3795, 4586, 4618, 4648, 4708, 4760, 5261, 5288, 5317, 5341, 5256, 5283, 5312, 5340, 9357, 9451, 12036, 12452, 12477]. Obtained [894, 123, 1878, 2503]. Obtaining [8429]. Occasional [2039, 2089, 2342]. Occupation [3628, 9029]. Occupations [9480]. Occurrence [11952]. Ocean [1669, 1723, 4200, 4317, 3916, 7406, 7551, 6755, 9333, 8481, 10004, 12579, 6610, 6631, 6332]. Oceanic [1889, 2322, 8997, 8878]. Oceans [3135]. Octopus [12472]. Ocular [6998, 8923]. Odd [4103]. Odiorne [1822]. Odorless [4729, 7760]. Œdema [9652]. Oera [8297]. Off [4923, 6411, 6037, 8653, 9524, 9808, 12859, 13208, 863, 1262, 7759, 10807, 11702]. Offal [4849]. Offense [4854]. Offer [11048, 12074]. Offered [6162]. Office [630, 1957, 3182, 4400, 6480, 7502, 6756, 11131, 1079, 253, 37, 353, 376, 434, 923, 506, 537, 563, 586, 600, 625, 233, 375, 404, 300, 1856, 1832, 2313, 1755, 2023, 1414, 1793, 1900, 1344, 3366, 3625, 2787, 2647, 3727, 3790, 3401, 3429, 3280, 2952, 2594, 3816, 3871, 3903, 3988, 4018, 4046, 4080, 4133, 4246, 4385, 4440, 6468, 6490, 6515, 6548, 5393, 5502, 5526, 5556, 6002, 5295, 6350, 6601, 6757, 6782, 6807, 6770, 6817, 8416, 9728, 10658, 10660, 10340, 10465, 10512, 10630, 10661, 10689, 10770, 10879, 11422]. Office [11303, 11328, 11571, 11589, 11754, 11953, 12150, 11108, 11187, 11260, 11366, 12152, 12334]. Officer [1848, 1868, 1891, 1933, 1958, 3340]. Offices [5197, 3801]. Official [38, 60, 654, 683, 710, 737, 756, 780, 799, 824, 850, 867, 888, 913, 934, 954, 975, 999, 1019, 1040, 1061, 1085, 1110, 1132, 1154, 1179, 83, 106, 135, 165, 182, 210, 236, 260, 285, 306, 330, 356, 380, 408, 433, 460, 478, 498, 520, 542, 566, 589, 610, 203, 43, 1205, 1234, 1884, 1927, 1952, 1981, 2005, 2027, 2052, 2076, 2094, 2117, 2138, 2159, 2181, 2208, 2233, 2256, 2282, 2305, 2331, 2356, 2378, 2405, 2428, 1258, 1284, 1311, 1338, 1365, 1388]. Official [1421, 1449, 1473, 1503, 1527, 1559, 1582, 1607, 1630, 1655, 1679, 1707, 1732, 1759, 1785, 1813, 1838, 1863, 2453, 2480, 3708, 11648, 12172]. Ohio [1008, 1544, 3754, 7175]. Oil [743, 720, 149, 146, 614, 1568, 1352, 1215, 2020, 1478, 1687, 3208, 3384, 2588, 4769, 4742, 4466, 3893, 5975, 5403, 5516, 7540, 8486, 8625, 9378, 8746, 8453, 9086, 8183, 10319, 10753, 11434, 12145, 11506, 13646, 13009, 1434, 101, 130, 189]. Oil-Stone [614]. Oiler [2264, 5659, 10918, 12442]. Oiling [690, 1186, 9617, 11672]. Oils [1438, 2058, 4118, 8283, 13404, 2366]. Old [1645, 1684, 1549, 3278, 4925, 4288, 4005, 4028, 6381, 7328, 8633, 9880, 9940, 12361, 13027, 13056, 12778, 11501]. Older [11716]. Oldest [7855]. Oleo [9811]. Oleo-Margarin [9811]. Oleomargarin [5712]. Olive [3384]. Oliver [11847]. Olmstead [1722]. Omnibus [2036, 5303, 6848]. Once [4874]. One [221, 318, 2243, 1613, 1622, 1428, 2645, 4513, 4003, 5999, 6108, 5328, 5378, 6376, 7328, 7388, 7984, 9077, 13284]. One-Horse [318]. Onion [13225]. Onion-Smut [13225]. Only [4341]. Ontario [10577]. Open [116, 2294]. Opener [12158]. Opening [1643, 3369, 2905, 4343, 8564, 8477, 13146, 12955]. Opera [316, 7186]. Operation [296, 6244]. Operations [2415, 2409, 4232, 6468, 6243, 6632, 11993]. Operative [1853, 12330]. Opinion [6587]. Opium [582, 8155, 8286]. Opponent [679]. Opponents [2266]. Opportunities [7717, 13172]. Opportunity [7648, 9114, 10742, 11497]. Oppose [12121]. Opposite [6765]. Opposition [11184]. Optical [8830, 13183, 8775, 11188]. Optically [6145]. Optics [8812]. Oracle [5897]. Oracles [8764]. Orange [794]. Oration [2088]. Orchid [6032, 5334, 10992]. Orchids [7006, 10730]. Orchis [12891]. Ord [6037]. Order [1895, 5010, 10098]. Ordinary [670, 10985]. Ordnance [2163, 1507, 5193, 12712]. Ore [1939, 3447, 10926, 11069, 12645, 5302, 6391]. Oregon [5541]. Orellana [9379]. Oremation [5658]. Oreodoxa [5167]. Ores [2189, 1826, 2695, 2739, 5438, 5826, 11255, 11970, 11704, 12584]. Organ [1998, 2237, 5279, 8157, 8162, 11811]. Organic [1377, 2771, 2800, 7124, 13380]. Organization [3340]. Organizations [4226]. Organography [11598]. Organs [2109, 9940]. Oriental [7673, 12535, 13555]. Origin [553, 320, 2065, 3017, 5408, 7116, 7898, 7688, 7124, 9049, 11922, 13458]. Original [9095]. Originated [12878]. Origination [5603]. Orleans [405, 1438]. Ornament [9146]. Ornamental [204, 2462, 2488, 4322, 6880, 7490, 7112, 8399]. Ornamenting [12877]. Ornithological [9146, 9698]. Orthopters [785]. Orton [5023, 5049, 5281, 5309, 5335, 10801, 12302]. Oscillating [4042, 6651, 10939, 11867, 12078]. Osmium [8676]. Ostrich [11457]. Other [508, 621, 1481, 1780, 1710, 10539, 9593, 10749, 12414, 5305, 428, 1702, 3019, 2625, 2649, 7754, 12149, 11834, 3043, 273]. Otocyon [13413]. Otter [13569]. Otto [9427, 11291, 12253, 11717]. Ought [2172, 2012, 1273]. Our [280, 729, 322, 120, 243, 160, 1890, 1914, 1922, 2276, 1290, 3657, 3675, 2923, 3928, 4451, 3808, 4040, 4258, 5156, 5326, 6294, 5271, 6085, 7633, 6695, 6621, 7100, 7656, 8347, 8185, 8301, 8017, 8894, 8925, 10749, 9983, 9851, 11210, 11060, 11640, 11500, 11996, 11504, 11011, 11034, 11064, 12100, 11561, 11256, 12414, 12640, 10867, 207, 753, 902, 3164, 3585, 2596, 4033, 6496, 9512, 11660, 11800, 10911]. Ours [621]. Ourselves [5417, 6697]. Out-Door [3331]. Outlook [11883]. Outrage [4743, 4689]. Ovals [6873]. Oven [3722]. Over-Culture [5109]. Over-Density [12356]. Overhead [5363, 5743, 10501]. Overheated [4208].

Overshoe [9214]. Overstrain [10539]. Owl [8285, 12535]. Own [11363]. Owner [2651]. Owners [4597, 5380]. Owns [1646]. Ox [4669, 191]. Oxalic [10593]. Oxide [659, 3060]. Oxidizing [3856]. Oxygen [1120, 3780, 5767, 5765, 10406, 11097, 11764, 11094]. Oxyhydric [3656, 3855]. Oxyhydrogen [1970, 3137]. Oyster [6494, 5522]. Oysters [250, 6600, 6623]. Ozone [5436, 8869]. Ozonized [1976].

PA [6678, 10112, 13606, 570, 8981]. Pacific [415, 2735, 4517, 983, 615, 3007, 3970, 3980, 6160, 6332, 7330, 9707, 12436]. Package [4932]. Packer [666, 12949]. Packing [92, 418, 1219, 1318, 1694, 2617, 2847, 4854, 5721, 7192, 10346, 13284, 12621, 13201]. Pad [11838]. Paddle [3863, 3970, 10150, 12265]. Padlock [4271, 6361]. Page [3494, 4748, 12922]. Paget [1963]. Pah [13069]. Pah-Gosa [13069]. Paid [9938]. Pail [1162, 1616, 4010, 5522, 6949]. Pail-Holder [1162]. Pain [6426, 10221]. Paine [1747, 1806]. Paint [1245, 3048, 6038, 10884, 177]. Paint-Pot [1245]. Painting [12205]. Paints [640]. Palace [7854, 983, 4798, 10834, 11352]. Palæotherium [5577]. Palette [2940]. Palier [12]. Palissy [13470]. Palm [5167, 5145, 5907, 7207, 9718, 10640, 11647, 13288, 441, 9500]. Palmer [2785, 4993]. Palmieri [11940]. Palms [8935, 7028]. Palmstedt [819]. Panama [13300, 13155]. Pane [325]. Paneling [1114, 8656, 12882]. Pangenesis [8647]. Panic [5059, 4952, 6350]. Panics [9336]. Paniculata [9440]. Pantagraph [3983, 4128]. Pantograph [13150]. Pap [3883]. Paper [674, 418, 322, 33, 789, 699, 1617, 1890, 1914, 2219, 2020, 2214, 1226, 3536, 3716, 3021, 3541, 3332, 3139, 4103, 4722, 4353, 5008, 5465, 8016, 8806, 8362, 9402, 9088, 10197, 10657, 9532, 10038, 10932, 10593, 10086, 12234, 12254, 11320, 12146, 11089, 12858, 12488]. Paper-Pulp [10197]. Papers [466, 5141, 6198, 11725]. Papier [1740, 3080]. Papin [4117, 9733]. Papua [12447]. Papyrus [10038]. Parachute [6383]. Paradise [5584, 10611, 9769, 11620, 12031, 12008]. Paradox [11637, 13248]. Paradoxes [11158]. Paraffin [5403, 5924]. Paraffine [839]. Paraguay [216]. Parallel [5822, 12262, 12319, 12497]. Parasites [8450]. Parasol [9051, 12623]. Parchment [10593]. Pardee [5115]. Paris [1970, 947, 316, 426, 948, 1358, 2034, 3469, 7186, 8968, 8126, 9907, 10062, 10743, 10138, 12049, 11063, 12091, 11352, 11998, 11015, 11879, 11480, 11929, 11906, 11931, 11666, 12156, 12097, 11786, 11624, 10978, 11644, 12172, 11565, 11948, 12210, 11845, 11772, 10980, 11605, 11908, 12076, 12120, 11608, 12070, 11694, 11977, 12371, 12876, 12320, 5211]. Parish [2148]. Park [1798, 2668, 4273, 6011, 6609, 7290, 11325, 11361, 11760, 11982, 966, 1328, 6356, 11772, 13128]. Parkes [2737]. Parks [732, 4876, 11311]. Parliament [4079]. Parlor [4531]. Parrot [8285, 13083]. Partially [8829]. Parting [1818]. Partition [11802]. Partnerships [6870]. Parts [1322]. Party [4424]. Pashiuba [5145]. Pass [12484, 12552]. Passage [10141, 13331, 12955]. Passed [8215, 12458]. Passenger [4501, 13661]. Passes [8418]. Past [1305, 1330, 1360, 1372, 1412, 1466, 9667]. Patch [2883]. Patent [1043, 454, 816, 608, 37, 353, 376, 434, 764, 223, 271, 421, 622, 748, 18, 11, 201, 765, 658, 687, 385, 623, 733, 696, 679, 1095, 474, 1857, 1219, 2195, 2147, 1947, 2170, 2123, 1634, 1745, 1280, 2099, 1832, 1383, 1419, 1468, 1997, 1595, 1564, 1940, 1183, 1288, 1425, 1694, 1946, 2168, 1522, 2277, 2300, 1913, 2173, 2175, 2774, 2670, 3538, 2474, 3045, 3050, 3072, 3096, 3115, 3144, 3170, 3201, 3229, 3253, 3288, 3320, 3347, 3375, 3407, 2644, 3771, 3668, 3697, 3732, 3764, 2448, 3661, 2937]. Patent [2977, 2918, 2432, 3032, 3699, 3734, 3494, 2710, 2991, 3752, 3353, 2945, 4703, 4528, 4432, 4925, 5161, 4821, 4751, 4377, 4463, 4205, 3833, 4327, 4794, 4552, 4847, 4264, 4145, 4670, 4951, 4755, 4644, 3844, 4585, 4621, 4651, 4467, 4349, 4874, 6545, 5901, 5538, 5598, 5874, 5798, 5360, 5881, 5814, 5774, 5851, 5380, 5267, 5391, 5579, 5461, 5566, 5596, 6322, 5628, 5650, 5680, 5705, 5731, 5761, 5785, 5813, 5841, 5868, 5892, 5919, 5948, 5997, 6024, 6048, 6078, 6129, 6154, 6182, 6208, 6239, 6266, 6292, 5352, 5906]. Patent [5842, 6177, 5513, 7210, 6738, 6743, 7428, 7358, 6937, 7819, 6768, 6962, 7815, 7211, 7755, 6619, 7780, 6714, 7071, 7204, 7175, 7051, 7563, 8518, 8735, 8253, 9245, 8132, 9237, 8534, 8559, 9456, 8586, 8616, 8645, 8674, 8702, 8731, 8761, 8791, 8823, 8859, 8891, 8919, 8950, 8980, 9011, 9043, 9075, 9109, 9135, 9233, 9267, 9363, 9393, 9423, 8246, 8694, 8259, 10900, 9728, 10248, 10658, 9827, 10903, 10970, 10904, 10141, 9549, 9575, 10276, 10304, 10337, 10369, 10402, 10431, 10461, 10495, 10523, 10558, 10592, 10624, 10655, 10681, 10711]. Patent [10739, 10765, 10793, 10819, 10849, 10874, 10899, 10924, 10953, 9606, 9632, 9663, 9689, 9717, 9742, 9766, 9796, 9821, 9875, 9904, 9933, 9955, 9979, 10003, 10027, 10058, 10080, 10109, 10137, 10164, 10190, 10219, 10247, 9803, 9693, 11210, 11056, 11516, 11169, 11476, 11010, 11084, 11650, 11830, 11933, 11339, 11652, 12120, 11665, 10976, 11004, 11771, 11795, 11823, 11854, 11878, 12041, 12064, 12113, 12142, 12167, 12193, 12220, 12275, 11026, 11055, 11079, 11103, 11129, 11156, 11182, 11204, 11227, 11280, 11309, 11337, 11360, 11389, 11416, 11445, 11470, 11494, 11521]. Patent [11548, 11578, 11604, 11632, 11747, 12121, 12279, 11561, 13403, 12452, 12640, 12607, 12484, 12603, 12786, 12301, 12328, 13195, 13218, 13245, 13270, 13297, 13325, 13353, 13377, 13398, 13424, 13454, 13479, 13504, 13528, 13552, 13578, 13603, 13629, 13658, 13685, 12353, 12381, 12404, 12431, 12482, 12513, 12549, 12573, 12602, 12638, 12664, 12694, 12726, 12757, 12813, 12844, 12928, 13005, 13034, 13063, 13090, 13117, 13142, 13169, 13405, 13036, 12815, 13328, 12552, 291, 1079, 253, 923, 506, 537, 563, 586, 600, 625, 233, 375, 404, 300, 1856, 1832, 1755, 2023]. Patent [1414, 2168, 1793, 1900, 1344, 3366, 3625, 2787, 2647, 3727, 3790, 3182, 3401, 3429, 3280, 3340, 2952, 2594, 3816, 3871, 3903, 3988, 4018, 4046, 4080, 4133, 4246, 4385, 4440, 6468, 6490, 6515, 6548, 5393, 5502, 5526, 5556, 6002, 5295, 6480, 6350, 6601, 6757, 6782, 6807, 6770, 6817, 6598, 7503, 8416, 8353, 10660, 10340, 10465, 10512, 10630, 10661, 10689, 10770, 10879, 11422, 11303, 11328, 11571, 11589, 11754, 11953, 12150, 11007, 11131, 11108, 11187, 11260, 11366, 12152, 12458, 12334]. Patentability [2146]. Patentable [2110]. Patented [2295, 2890, 4397, 5953, 12544]. Patentee [6977, 12856]. Patentees [1903, 3319, 3346, 3376, 3408, 9552]. Patents [61, 82, 1106, 1180, 134, 162, 183, 211, 261, 302, 379, 475, 521, 539, 590, 631, 655, 684, 707, 734, 777, 798, 825, 910, 955, 976, 996, 1059, 301, 962, 97, 748, 38, 60, 654, 683, 710, 737, 756, 780, 799, 824, 850, 867, 888, 913, 934, 954, 975, 999, 1040, 1061, 1085, 1110, 1132, 1154, 1179, 83, 106, 135, 165, 182, 210, 236, 260, 285, 306, 330, 356, 380, 408, 433, 460, 478, 498, 520, 542, 566, 589, 610, 630, 658, 687]. Patents [36, 59, 653, 682, 736, 754, 779, 797, 823, 849, 866, 887, 912, 933, 953, 974, 998, 1018, 1039, 1060, 1084, 1109, 1131, 1153, 1178, 81, 105, 132, 164, 181, 208, 234, 259, 284, 304, 329, 355, 378, 407, 431, 459, 476, 497, 519, 541, 565, 588, 609, 629, 1232, 1283, 2136, 2178, 2232, 2257, 2280, 2303, 2332, 2357, 2377, 2401, 2426, 1309, 1387, 1418, 1500, 1526, 1557, 1583, 1606, 1653, 1680, 1706, 1729, 1760, 1784, 1836, 1862, 1885, 1906, 1928, 1953, 1982]. Patents [2028, 2051, 2118, 1374, 2184, 1650, 2022, 2252, 1205, 1234, 1884, 1927, 1981, 2005, 2027, 2076, 2094, 2117, 2138, 2159, 2181, 2208, 2233, 2256, 2282, 2305, 2331, 2356, 2378, 2405, 2428, 1258, 1284, 1311, 1338, 1365, 1388, 1421, 1449, 1473, 1503, 1527, 1559, 1582, 1607, 1630, 1655, 1679, 1707, 1732, 1759, 1785, 1813, 1838, 1863, 1183, 1288, 1425, 1393, 1204, 1233, 1883, 1908, 1926, 1980, 2004, 2026, 2075, 2093, 2116, 2137, 2158, 2180, 2231, 2255, 2304, 2330, 2355, 2376, 2404, 1257, 1282]. Patents [1310, 1337, 1364, 1386, 1420, 1448, 1472, 1501, 1525, 1558, 1581, 1605, 1628, 1652, 1704, 1730, 1758, 1811, 1837, 1861, 2016, 1899, 2248, 2173, 2454, 2476, 2506, 2947, 3244, 3458, 3486, 3515, 3546, 3573, 3609, 3094, 3160, 3570, 3372, 3319, 3346, 3376, 3408, 2570, 3725, 2686, 2453, 2480, 3708, 2927, 3751, 2432, 3032, 3699, 3734, 2669, 2452, 2478, 3073, 3098, 3118, 3143, 3169, 3200, 3228, 3252, 3287, 3317, 3345, 3374, 3405, 3433, 3460, 3487, 3514, 3549, 3575, 3611, 3635, 3666, 3695, 3730, 3762, 3794]. Patents [2508, 2534, 2557, 2579, 2607, 2633, 2659, 2678, 2701, 2728, 2750, 2776, 2802, 2828, 2852, 2872, 2893, 2911, 2930, 2955, 2979, 3000, 3027, 3049, 3667, 3696, 3731, 3763, 3795, 3975, 4085, 4193, 4736, 5206, 5235, 5007, 5158, 5178, 5107, 3820, 3845, 4554, 4582, 4617, 4647, 4681, 4707, 4733, 4759, 4782, 4813, 4838, 4865, 4894, 4916, 4943, 4968, 5000, 5024, 5074, 5097, 5125, 5149, 5180, 5203, 5232, 3873, 3904, 3932, 3962, 3990, 4019, 4048, 4082, 4107, 4135, 4162, 4190, 4217, 4248, 4274, 4305, 4330]. Patents [4359, 4387, 4412, 4441, 4474, 4503, 4532, 4586, 4618, 4648, 4708, 4760, 3934, 4682, 5105, 5260, 5287, 5969, 5993, 6019, 6044, 6073, 6099, 6124, 6150, 6178, 6203, 6235, 6262, 6287, 6318, 6344, 6369, 6396, 6417, 6440, 6461, 6485, 6509, 6537, 5316, 5368, 5398, 5429, 5456, 5479, 5506, 5562, 5591, 5624, 5646, 5675, 5700, 5727, 5756, 5780, 5808, 5836, 5862, 5914, 5943, 5261, 5288, 5317, 5341, 5566, 5596, 6322, 5628, 5650, 5680, 5705, 5731, 5761, 5785, 5813, 5841, 5868, 5892, 5919, 5948, 5997, 6024, 6048, 6078]. Patents [6129, 6154, 6182, 6208, 6239, 6266, 6292, 5257, 5284, 5966, 5990, 6041, 6070, 6096, 6121, 6147, 6174, 6200, 6232, 6259, 6284, 6315, 6341, 6366, 6393, 6414, 6437, 6458, 6482, 6506, 6534, 6565, 5313, 5337, 5365, 5395, 5426, 5453, 5476, 5503, 5529, 5559, 5588, 5621, 5672, 5697, 5724, 5753, 5777, 5805, 5833, 5859, 5886, 5911, 5940, 5256, 5283, 5312, 5340, 6591, 6614, 7474, 7497, 7519, 7545, 7570, 7595, 7625, 7651, 7679, 7709, 7740, 7772, 7801, 7862, 7892, 6636, 6663, 6687, 6708, 6731, 6760, 6785]. Patents [6810, 6834, 6859, 6885, 6902, 6928, 7249, 7275, 7295, 7319, 7350, 7374, 7397, 7422, 7445, 7097, 7660, 7717, 7204, 6622, 6588, 6612, 7247, 7273, 7293, 7317, 7348, 7372, 7395, 7420, 7443, 7472, 7495, 7518, 7543, 7568, 7593, 7623, 7649, 7676, 7677, 7702, 7734, 7766, 7795, 7856, 7886, 6633, 7914, 6661, 6685, 6706, 6729, 6758, 6783, 6808, 6832, 6857, 6883, 6900, 6926, 6955, 6984, 7009, 7036, 7059, 7086, 7118, 7142, 7167, 7197, 7225, 7733, 6587, 9202, 7947, 7974, 8006, 8036, 8062, 8093, 8120, 8147, 8206]. Patents [8235, 8534, 8559, 9456, 8586, 8616, 8645, 8674, 8702, 8731, 8761, 8823, 8859, 8891, 8919, 8950, 8980, 9011, 9043, 9075, 9109, 9135, 9233, 9267, 9363, 9393, 9423, 8054, 8085, 7941, 7969, 8721, 8751, 8781, 8814, 8847, 8879, 8910, 8938, 8969, 9001, 9030, 9062, 9098, 9124, 9157, 9187, 9221, 9256, 9288, 9321, 9350, 9381, 9412, 9444, 9481, 7999, 9514, 8030, 8056, 8087, 8113, 8142, 8170, 8201, 8229, 8289, 8311, 8338, 8370, 8403, 8433, 8463, 8493, 8523, 8547, 8575, 8606, 8634, 8663, 7968, 9357, 9451]. Patents [8447, 8853, 8885, 8914, 8944, 8976, 9006, 9037, 9069, 9104, 9130, 9163, 9193, 9194, 9227, 9262, 9294, 9327, 9358, 9387, 9419, 9638, 9549, 9575, 10276, 10304, 10337, 10369, 10402, 10431, 10461, 10495, 10523, 10558, 10592, 10624, 10655, 10681, 10711, 10739, 10765, 10793, 10819, 10849, 10874, 10899, 10924, 10953, 9606, 9632, 9663, 9689, 9717, 9742, 9766, 9796, 9821, 9875, 9904, 9933, 9955, 9979, 10003, 10027, 10058, 10080, 10109, 10137, 10164, 10190, 10219, 10247, 9542, 9567, 10265, 10292, 10327, 10359, 10391, 10421, 10452, 10484, 10514, 10546]. Patents [10581, 10617, 10645, 10671, 10700, 10731, 10756, 10809, 10839, 9596, 9623, 9654, 9680, 9708, 9734, 9758, 9786, 9838, 9865, 9894, 9924, 9943, 9972, 9994, 10016, 10046, 10070, 10099, 10128, 10155, 10179, 10209, 10237, 10375, 9771, 11610, 11419, 11465, 11184, 10976, 11004, 11771, 11795, 11823, 11854, 11878, 12041, 12064, 12113, 12142, 12167, 12193, 12220, 12275, 11026, 11055, 11079, 11103, 11129, 11156, 11182, 11204, 11227, 11280, 11309, 11337, 11360, 11389, 11416, 11445, 11470, 11494, 11521, 11548, 11578, 11604, 11632, 11747, 12171, 11162, 12042, 12036, 12304]. Patents [13070, 12903, 12301, 12328, 13195, 13218, 13245, 13270, 13297, 13325, 13353, 13377, 13398, 13424, 13454, 13479, 13504, 13528, 13552, 13578, 13603, 13629, 13658, 13685, 12353, 12381, 12404, 12431, 12482, 12513, 12549, 12573, 12602, 12638, 12664, 12694, 12726, 12757, 12813, 12844, 12928, 13005, 13034, 13063, 13090, 13117, 13142, 13169, 12357, 12531, 12871, 12902, 12491, 12614, 12646, 12680, 12746, 12767, 12837, 12887, 12944, 12966, 12996, 13043, 13071, 13101, 13126, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 12529, 12642, 12408, 12677, 13448, 12477, 13491, 1019, 709, 1952]. patents [2052, 1951, 2050, 2281, 1677, 1783, 5051, 6016, 5643, 10782, 180, 1382, 1644, 1650, 2150, 1247, 2145, 2146, 2566, 2569, 3161, 3692, 3999, 5010, 6297, 5751, 6601, 7409, 7532, 6792, 9156, 9857, 10716, 9556, 11132, 12069, 13302]. Patents-Reduction [7660]. Patents-Splendid [7717]. Paterfamilias [8226]. Path [486]. Patient [414]. Patients [2797]. Patina [171]. Patron [7876, 12263]. Patrons [8327]. Pattern [11193]. Patterns [2013, 3530].

Patteson [11668]. Patton [2666]. Pauper [12022]. Pauperism [10437]. Pavement [412, 1094, 2045, 2540, 3752, 2782, 4565, 13515]. Pavements [2274, 1224, 2310, 3471, 4983, 5462, 10771, 11752]. Pavilion [3868, 4834, 4860, 8591, 10065]. Pavilions [8543]. Paving [858]. Pay [346, 392, 2994, 6326, 8353]. Payen [2069]. Payne [1462, 2167]. Pays [12435]. Peace [10249]. Peacemakers [11391]. Peach [1089, 7384]. Peael [13558]. Peanut [1434]. Peanuts [1434]. Pearly [9059]. Peart [2991]. Peas [6381]. Peat [524, 1442, 4541, 6474]. Pebbling [2244, 2916]. Peculiar [126, 6299, 13095]. Peculiarities [12618]. Pecuniary [1408]. Pedal [12407]. Pedden [7821]. Pedespeed [289]. Pedigree [8257]. Pedometer [12743]. Peerless [11281]. Peg [8]. Peking [6160, 5574]. Pen [814, 2102, 3356, 5547, 8899, 11981, 13024, 12650, 12504, 12973]. Pencil [3042, 7561, 8871, 12177, 12405]. Penciled [12394]. Pencils [129, 6820, 10435]. Pending [454]. Pendula [5144]. Pendulum [6525, 6558, 11104, 12374]. Pendulums [12179]. Penetrate [11967]. Penguin [4099, 8052, 8083]. Penguins [10232]. Penholders [6158]. Penikese [7016]. Pennell [2585]. Pennsylvania [627, 2168, 3602, 4769, 9492, 11434]. Penny [200]. Pens [1538, 4848, 10435, 11777, 13458]. People [491, 1052, 1932, 6299, 9938, 11245, 11969]. People’s [4725]. Pepper [6702, 3539]. Perambulating [4754]. Percentage [1648, 1715]. Perception [6026, 8830]. Percha [892]. Perfect [4515]. Perfecting [7643]. Perfume [9919]. Perfumes [315]. Peril [8828, 8620]. Perils [2108]. Periodicals [551]. Peripolar [4988]. Perishable [570]. Permanent [5469, 10006, 11082]. Permeator [5019]. Pernot [6700]. Pernyi [4671]. Perpetual [1098, 1116, 1137, 1165, 1187, 1212, 1242, 1266, 1296, 1320, 1347, 1375, 1399, 1430, 9856, 11542]. Persia [11948]. Persian [4834, 9835, 10065]. Person [9077]. Personal [2, 40, 568, 591, 613, 633, 657, 686, 712, 739, 759, 782, 801, 828, 853, 869, 890, 916, 937, 957, 979, 1002, 1021, 1042, 1063, 1087, 1113, 1135, 63, 1158, 86, 108, 137, 167, 186, 214, 238, 263, 287, 309, 332, 358, 383, 411, 436, 462, 480, 500, 523, 544, 2184, 1184, 1207, 1866, 1888, 1910, 1930, 1956, 1984, 2007, 2031, 2054, 2078, 2096, 2120, 2140, 2161, 2185, 2210, 2236, 2260, 2284, 2307, 2352, 2360, 2381, 2407, 1237, 1260, 1286, 1313, 1341, 1367, 1390, 1423]. Personal [1451, 1476, 1506, 1529, 1561, 1585, 1609, 1632, 1657, 1682, 1709, 1734, 1762, 1787, 1816, 1840, 2430, 2457, 3030, 3052, 3075, 3100, 3120, 3146, 3172, 3203, 3231, 3256, 3291, 3322, 3349, 3379, 3410, 3437, 3464, 3491, 3518, 3552, 3578, 3614, 3639, 3671, 3700, 2483, 3736, 3766, 2511, 2537, 2559, 2581, 2610, 2636, 2661, 2681, 2703, 2731, 2753, 2778, 2804, 2830, 2853, 2874, 2895, 2912, 2932, 2958, 2981, 3002, 2724, 3797, 3823, 4538, 4559, 4590, 4626, 4654, 4686, 4713, 4739, 4764, 4787, 4817, 4842, 4870, 4899, 4920]. Personal [4947, 4973, 5004, 5029, 5055, 5079, 5101, 5129, 5153, 5185, 3848, 5208, 3877, 3909, 3938, 3966, 3995, 4024, 4053, 4088, 4112, 4139, 4166, 4195, 4222, 4254, 4279, 4311, 4336, 4364, 4391, 4418, 4447, 4480, 4507, 5238, 5264, 5946, 5971, 5995, 6022, 6046, 6076, 6102, 6127, 6152, 6180, 6206, 6237, 6264, 6290, 6320, 6347, 6371, 6398, 6420, 6442, 6464, 6487, 6511, 6539, 5291, 5319, 5344, 5371, 5401, 5432, 5458, 5482, 5509, 5534, 5564, 5594, 5626, 5648, 5678, 5703, 5729, 5759, 5783, 5811, 5839, 5865, 5890, 5917, 6569]. Personal [6593, 7228, 7251, 7277, 7297, 7322, 7352, 7376, 7399, 7424, 7448, 7476, 7499, 7522, 7546, 7572, 7598, 7628, 7652, 7681, 7711, 7742, 7775, 7804, 7835, 7865, 7893, 6616, 6639, 6666, 6689, 6710, 6734, 6762, 6788, 6813, 6835, 6862, 6887, 6904, 6931, 6958, 6988, 7012, 7040, 7062, 7089, 7120, 7145, 7170, 7201, 7921, 7949, 8671, 8699, 8728, 8758, 8788, 8820, 8856, 8888, 8916, 8947, 8977, 9008, 9040, 9072, 9106, 9132, 9166, 9197, 9230, 9264, 9297, 9330, 9360, 9390, 9421, 9453, 7977, 9489, 8008, 8038, 8064, 8069]. Personal [8123, 8150, 8177, 8208, 8238, 8267, 8293, 8318, 8342, 8378, 8410, 8441, 8468, 8500, 8531, 8556, 8583, 8613, 8642, 9521, 9546, 10216, 10244, 10273, 10301, 10334, 10366, 10399, 10428, 10458, 10492, 10520, 10555, 10589, 10621, 10652, 10678, 10708, 10736, 10762, 10790, 10816, 10846, 10871, 10896, 10922, 9573, 10949, 9603, 9629, 9660, 9687, 9714, 9739, 9763, 9793, 9818, 9846, 9872, 9901, 9930, 9952, 9976, 10000, 10024, 10055, 10077, 10106, 10134, 10161, 10187, 10973, 11001, 11545, 11574, 11601, 11628, 11655, 11675, 11699, 11719, 11743, 11767, 11791, 11819, 11850]. Personal [11874, 11897, 11918, 11944, 11964, 11985, 12015, 12037, 12061, 12083, 12110, 12138, 11023, 12163, 12189, 12217, 12241, 12272, 11052, 11075, 11100, 11126, 11152, 11178, 11201, 11224, 11250, 11277, 11305, 11333, 11357, 11385, 11412, 11442, 11467, 11490, 11519, 12298, 12325, 13001, 13030, 13059, 13086, 13113, 13138, 13165, 13191, 13214, 13241, 13266, 13293, 13321, 13349, 13373, 13394, 13420, 13451, 13475, 13500, 13525, 13548, 13574, 13599, 13625, 13654, 12349, 13682, 12377, 12401, 12427, 12455, 12479, 12509, 12545, 12569, 12599, 12635, 12661, 12690, 12723, 12754, 12782, 12810, 12840, 12867, 12898]. Personal [12925, 12951, 12976]. Personalities [9077]. Persons [318]. Perspective [13537]. Pertaining [12531]. Peru [12539, 5359, 10528]. Peruvian [10153]. Pest [8931, 9964]. Pestilence [11950]. Pestle [3006]. Petek [12643]. Peter [2335, 9619, 12516]. Peters [614]. Petroleum [1354, 1438, 3483, 3217, 3603, 4905, 4982, 4875, 5011, 5965, 7514, 7638, 9274, 8867, 10225, 12222, 12170, 12115, 12195, 8774]. Pettit [5500]. Pharmaceutical [9471]. Pharmacy [2086]. Phase [4318, 7902]. Phelps [1262, 4286]. Phenic [273]. Phenomena [1067, 2858, 3564, 3105, 5900, 6658, 7530, 10768, 2493, 8112]. Phenomenon [9078]. Philadelphia [10112, 1774, 4398, 8444, 11380, 11760, 11982, 13532]. Philander [12984]. Philanthropic [1467]. Philippi [1824]. Phillips [11240]. Philosophers [5899]. Philosophical [8227, 2463]. Philosophy [230, 5381]. Phineas [3758]. Phoenix [7850]. Phonetics [10803]. Phonograph [10928, 11607, 11286, 11714, 11842, 12939]. Phonomoter [11728]. Phosphides [8222, 8250]. Phosphor [4100, 5111]. Phosphor-Bronze [4100, 5111]. Phosphoric [2591, 6524]. Phosphorus [8222, 8250]. Photo [967, 6556, 6993, 9581, 10370, 10934, 8770]. Photo-Chemicals [8770]. Photo-Engraving [10370, 10934]. Photo-Lithographs [9581].

Photo-Mechanical [6993]. Photo-Relief [967]. Photograph [1304, 3945]. Photographer [9913]. Photographers [2009]. Photographic [509, 1901, 3358, 2869, 2790, 4238, 8690, 9503, 11834]. Photographing [7983, 13304]. Photographs [1009, 255, 6301, 9855, 9697, 13176]. Photography [832, 926, 5135, 5791, 8812, 9960, 10406, 10030, 10822, 11135, 11295, 11342, 12832, 13175, 13118, 13426]. Photometer [3773, 4030, 3955]. Phrenology [1870, 9111]. Phylloxera [6162, 6944, 7303, 8198, 13161, 9833]. Physical [78, 73, 4315, 5109, 6424, 6599, 7356, 6658, 9465, 8446, 8952, 9078, 8507, 10280, 12366]. Physician [5389]. Physicians [9552]. Physics [1543]. Physiological [5925, 11567, 11177]. P’i [6646]. P’i-Ryi [6646]. Piano [12043, 12371]. Pianoforte [7558]. Pianos [4926]. Piccadilly [3748]. Pichiciago [10889]. Pick [2106, 5018]. Picker [3087]. Picket [5747]. Pickets [6918]. Pictet [10850, 11214]. Picture [11562]. Pictures [7018]. Pictus [3159]. Pie [6779]. Piece [4257, 6910, 10014, 12496]. Pieces [4519]. Pier [1563, 2311, 12579, 3769]. Pierced [1096]. Pierre [1740]. Piers [1748]. Pig [614, 4670]. Pigeon [4571]. Pigeons [6407]. Piking [6161]. Pile [4172]. Piles [394, 13281]. Pilgrim [2100]. Pilot [13665]. Pilotage [13634]. Pin [2194, 10503, 10905]. Pinchers [4941]. Pine [1773, 5011, 7029, 11433, 292]. Pineapple [7687]. Pinewood [45]. Pinking [6452]. Pinks [9677]. Pins [3264, 5877, 95]. Pioneers [13067]. Pipe [465, 1664, 1239, 3525, 3590, 4432, 5173, 4237, 6277, 5470, 6853, 8456, 9020, 8934, 10043, 9558, 10277, 10147, 9536, 11407, 11072, 11435, 11195, 11027, 12127, 12967, 13009, 921]. Pipes [69, 1012, 1300, 1536, 1642, 3084, 3851, 4631, 3923, 7049, 6866, 7465, 7778, 7988, 12044, 12440, 12886]. Piracy [11615]. Piscicultural [6818]. Pisciculture [1571]. Pistol [524, 8482]. Piston [1322, 1219, 2617, 2518, 3302, 2847, 9056, 11709, 13201]. Piston-Speed [1322]. Pistons [92, 1318]. Pit [2991]. Pitch [273]. Pitcher [627, 5936]. Pitman [2745]. Pits [6356, 7699]. Pitting [485]. Pittsburg [13660, 12982]. Pittsburgh [570, 4097, 6502, 6678, 8981]. Place [493, 1289, 3279, 5663, 8743]. Places [5630]. Plague [6094, 12518]. Plain [9970]. Plan [300, 1613, 3680, 4093, 6406, 6953, 12045, 13662, 12330]. Plane [2564, 4939, 5612, 11282, 12994, 12317, 12303]. Planer [4282, 4340, 5045, 5707, 6422, 7048, 7748, 8776, 8988, 8479, 9561, 10985, 11460, 12559]. Planer-Toothed [5045]. Planet [9204, 3773, 2764, 4884, 4908, 11857]. Planetary [7260, 11584]. Planets [4315, 3973, 6749, 10882]. Planing [4951, 5536, 6163, 6560, 7391, 6571, 9509, 10171, 10694, 12489]. Planking [1456]. Plans [651, 4124]. Plant [1565, 6657, 7620, 7913, 6981, 8384, 9378, 9252, 10754, 11326, 11354, 11293, 11400, 11459, 11622, 11355, 267]. Plantains [30]. Plante [9505, 11246, 11488]. Planter [985, 3715, 3927, 6654, 8023, 8108, 12467]. Plants [984, 2983, 4980, 4934, 5160, 5749, 6190, 5406, 5765, 7729, 7244, 7634, 7526, 6870, 6829, 8272, 8811, 9215, 8810, 10263, 9935, 11396, 11654]. Plaster [2034, 3248, 5211]. Plate [741, 2396, 3677, 4611, 6209, 7881, 10777, 9561, 13586, 13470]. Platen [12175]. Plates [2756, 7695, 9885, 9778, 9831, 8719]. Platform [1554, 12012]. Plating [691, 1635, 2461, 3523, 3588, 5355, 5384, 10462, 10562, 11315, 11059]. Platinized [217, 242, 272]. Platinum [1228, 11329, 13332]. Platinum-Tipped [11329]. Platt [13650]. Platypus [10914]. Play [10715]. Playing [8482]. Playthings [7930]. Plea [243, 251, 3089]. Pleasant [5519]. Pleasonton [2195, 8654]. Plenum [6750]. Plethysmograph [8686]. Pleuro [12641]. Pleuro-Pneumonia [12641]. Pliers [1295]. Plow [221, 1023, 1947, 3329, 4179, 5172, 4915, 4430, 4495, 4327, 4243, 5250, 7269, 10345, 9912, 12180, 11931, 11653, 11808, 13516]. Plowing [1746, 1752, 4562, 13109]. Plows [2865, 6305, 6607, 11907]. Pluck [13067]. Plucky [2061]. Plug [8046, 8077, 8398]. Plugging [3306]. Plumb [1533, 1541]. Plumbago [1897, 3957, 10690]. Plumber [6187, 7928]. Plumbing [5224]. Plunger [4705, 9725]. Pluviometer [8224]. Pneumatic [942, 560, 202, 244, 2261, 2926, 3386, 3522, 3443, 3019, 4487, 3987, 4702, 6527, 6246, 6907, 10909, 10198, 11183, 12050, 11967, 11862, 13047, 12524, 9822]. Pneumonia [12641]. Pocket [2264, 2900, 8798, 9176, 11949]. Pocketbook [10954, 13443]. Podometer [7385]. Poetry [1393]. Pogpisil [7721]. Point [8363, 10596, 12433, 12438]. Pointed [1662, 9773]. Points [6274, 9800, 12589, 95]. Poison [7811, 10747, 10404, 12184, 8931]. Poisoned [8864]. Poisoning [579, 2498, 4721, 10902, 12667]. Poisonous [1148, 3752, 4092, 6355, 8828, 12021, 11654]. Poisons [4599, 11541, 13496]. Polar [2294, 3629, 6551, 5769, 10376, 9634, 13119, 13342, 13150]. Polaris [4424, 4372, 4906, 5542]. Polarized [3889]. Pole [7236, 10915]. Poles [4602, 4210, 13038]. Police [2042]. Policy [9746, 13405]. Polisher [2523, 9987]. Polishing [741, 240, 2196, 3391, 2551, 8270, 10171, 12799]. Politics [6003, 7175]. Polychrome [6493]. Polyscope [11935]. Pomerania [12247]. Pompeii [10068]. Pond [5906]. Ponsard [943]. Pony [8776]. Pooley [2487]. Poor [1826, 4791]. Poppy [8286]. Popular [6137, 11285]. Popularizing [1721]. Population [9802, 12356]. Porcelain [5247, 6841, 6838, 6865, 8689]. Porcupine [12030, 11983, 12421]. Port [8188]. Portable [463, 993, 345, 2166, 2191, 1289, 2414, 2055, 1480, 2367, 2684, 2936, 2530, 3037, 3474, 2782, 4756, 4576, 4468, 4524, 5629, 5272, 5659, 6329, 6111, 7394, 7459, 7645, 6604, 8690, 9332, 8221, 8354, 10829, 10887, 9534, 10795, 10966, 10029, 11117, 11091, 11757, 13232, 13075, 13157, 3390]. Porter [4771, 12470]. Portland [2862]. Portraits [3945]. Portsmouth [6473]. Portugal [895, 4011, 11815]. Position [6974, 11960]. Positive [230, 2674, 4465, 9044]. Possessions [6352]. Possibilities [3632, 3681, 3901]. Possible [664, 6447, 7996, 12735]. Post [765, 1661, 1244, 2313, 1269, 3558, 4400, 5446, 7502, 6756, 10861, 13285, 1957]. Post-Hole [1269]. Postage [3415, 3445, 3181, 4245, 10382, 10686]. Postal [3065, 5062, 5089, 6867, 7098, 7875, 6575, 7605, 8300, 8829, 10098, 11283]. Posterity [9720]. Posts [1136, 2589]. Pot [1245, 2795, 5196, 7614, 10545, 10575, 11914]. Potash [949, 576]. Potato [1530, 2944, 5022, 5715, 6241, 7283, 6654, 11936, 13152, 12417, 8931, 10609].

Potato-Beetle [10609]. Potatoes [8809]. Potin [8025]. Pots [7700, 10886]. Pottery [8689, 11271, 12267, 12130, 12776, 12948, 12699]. Pottery-Maker [12699]. Pouch [6572]. Pounder [917, 7884]. Powder [8903, 11430, 12252]. Powdered [2983]. Powders [12203, 13092]. Powell [7817, 11118]. Power [774, 279, 1076, 373, 397, 1046, 1051, 2109, 2293, 1846, 1436, 1567, 1432, 1249, 1850, 1461, 2706, 3153, 2486, 2614, 2908, 3265, 3772, 3047, 3063, 3221, 5159, 4227, 4262, 3941, 4369, 3827, 4564, 3867, 4993, 4060, 4348, 5652, 5685, 5763, 5976, 5951, 5746, 7852, 7189, 7451, 7105, 7284, 7300, 8389, 9509, 8241, 8352, 8141, 8225, 9699, 10142, 12292, 11037, 12101, 12230, 11756, 11353, 11662, 13430, 12448, 12847, 13154, 1072]. Powerful [7386, 7684, 7927]. Powers [10716, 10192]. Pox [2645, 2797, 2552, 2650, 4028]. Practical [840, 1864, 1886, 2179, 1543, 3318, 3377, 3406, 3434, 3461, 3488, 3516, 3550, 3576, 3612, 3636, 2531, 2554, 3360, 3416, 3545, 3571, 3663, 3684, 3783, 2576, 2602, 2629, 2656, 2675, 2697, 2726, 2742, 2767, 2792, 2818, 2838, 2885, 2922, 2949, 2975, 3025, 3069, 3132, 3167, 3198, 3284, 3314, 3334, 4795, 3861, 5214, 4093, 5178, 3843, 3869, 4850, 4881, 4929, 4957, 4985, 5012, 5038, 5066, 5096, 5112, 5137, 5163, 5194, 5226, 3900, 4036, 4075, 4098, 4121, 4152, 4186, 4212, 4291, 4321, 4347, 4376, 4426]. Practical [4457, 4493, 4518, 4553, 4636, 4664, 4697, 4731, 4749, 4773, 4828, 4618, 4648, 5415, 6013, 6448, 5983, 6028, 5252, 5276, 6139, 6191, 6328, 6380, 6427, 6449, 6481, 6497, 5310, 5327, 5354, 5382, 5416, 5445, 5466, 5491, 5520, 5605, 5637, 5665, 5687, 5713, 5796, 5831, 5875, 5909, 5928, 5958, 5979, 6031, 6087, 6113, 7094, 6650, 6674, 7675, 7689, 7720, 7752, 7784, 7847, 6699, 6747, 6766, 6801, 6849, 6897, 6941, 6969, 7023, 7079, 7115, 7129, 7180, 7238, 7268, 7305, 7345, 7408, 7457, 7486, 7536, 7581, 7608]. Practical [7636, 8733, 8446, 9079, 9406, 8419, 7933, 7997, 8018, 8053, 8084, 8110, 8199, 8276, 8401, 8451, 8514, 8545, 8597, 10684, 12762, 12979, 13364, 13386, 13408, 787, 10866, 5825, 6063, 5853, 5879, 5934, 6009, 6115, 6164, 6223, 6276, 6334, 6387, 6431, 6501, 6559, 6947, 6626, 6679, 6724, 6777, 6826, 6877, 7003, 7080, 7132, 7188, 7267, 7310, 7365, 7415, 7463, 7512, 7560, 7612, 7670, 7726, 7788, 7851, 7908, 7962, 8022, 8105, 8194, 8252, 8306, 8360, 8426, 8490, 8542, 8598, 8655, 8742, 8805, 8870, 8932, 9023, 9147, 9179, 9248, 9313, 9375, 9436, 9537, 9591, 9646, 9702, 9753, 9807, 9860, 9939, 10010, 10067, 10120, 10174, 10233, 10284, 10442]. Practical [10537, 10664, 10831, 10885, 11040, 11216]. Practice [9, 4368, 6132, 7431, 8013, 12334]. Prairie [11594, 12980]. Prall [3124]. Praxinoscope [12619]. Preached [9138]. Precaution [11448]. Precautions [1634]. Precious [2695]. Precipitation [2740]. Precipitator [2368]. Prehensile [12030]. Prehensile-Tailed [12030]. Prehistoric [9410, 13134]. Prejudice [11648]. Prejudices [1877]. Premier [11458]. Premiums [4893, 4863, 12771]. Preparation [2519, 5166, 4488, 4541, 4464, 8991, 10593, 11681]. Prepare [6157]. Prepared [4796, 10480]. Preparing [1687, 10346, 11095]. Prescott [3544]. Prescriptions [1409]. Presence [482, 12584, 8632]. Present [1047, 817, 1305, 1330, 1360, 1372, 1412, 1466, 2800, 4040, 4231, 6496, 8588, 8589].

Preservation [55, 192, 53, 194, 1553, 7844, 9878, 12556, 29]. Preserve [13174]. Preserver [1246, 3062, 9985]. Preserving [1148, 526, 2298, 1481, 5661, 6805, 7102, 7723, 7580, 9177, 8485, 8424, 11212, 11741, 12792, 12877]. President [939, 959, 2015, 3740]. Press [845, 870, 501, 593, 1238, 3034, 3153, 2941, 3420, 2878, 3826, 4822, 4990, 5240, 5920, 5960, 5952, 7643, 7133, 7728, 8792, 12292, 12097, 11756, 13346, 12986, 12679, 13081, 13230, 12963, 13334, 13314]. Presses [5085, 11761]. Pressing [4091, 6227, 13178]. Pressure [367, 87, 1144, 2059, 1196, 3008, 2640, 3110, 3186, 2592, 3047, 3063, 4889, 5460, 5878, 5929, 5763, 7149, 6676, 7391, 6642, 6599, 7534, 10722, 10229, 11796, 11507, 11264, 11741, 11177]. Pressures [2908]. Prevalent [10114]. Prevent [438, 2296, 3451, 3338, 10578]. Preventable [7715]. Preventing [3399, 5742, 6499, 7055, 11029, 13560]. Prevention [251, 3689, 2797, 6013, 5663, 8099, 9498, 9577, 11150]. Preventive [4497, 5568, 12993]. Prevost [1991]. Price [733, 2848, 4769, 6298, 2782, 9458]. Prices [11419]. Prickly [9649]. Priestley [3557]. Primary [581]. Prime [1051, 1648, 3916, 5764, 5268, 5298, 1715, 3806]. Primer [6999]. Priming [6499]. Primitive [12699]. Prince [2351]. Principal [651]. Principle [1102, 424]. Principles [153, 2415]. Printer [1989, 2757, 7905]. Printers [1164]. Printing [593, 3034, 2547, 2710, 5085, 4822, 4401, 5240, 6112, 5295, 6493, 5952, 7643, 6993, 6622, 8792, 9402, 8660, 10796, 11162, 4382]. Printing-Firm [7643]. Prison [3219, 6446]. Prisons [7187]. Privileges [3978]. Prize [160, 6953, 13143]. Prizes [17, 207, 4793, 5118, 4846, 4978, 11998, 12256]. Probabilities [8688]. Probable [13172]. Problem [1147, 4, 536, 1322, 1718, 1921, 1353, 2088, 3352, 2747, 4290, 4320, 5872, 6698, 7934, 8633, 10279, 10224, 11606]. Problems [12575]. Proboscis [10864]. Procedure [1047]. Proceedings [1946, 6169, 5719, 5352, 6110, 6144, 6257, 6311, 7066, 9061, 13337]. Process [991, 367, 1148, 514, 554, 614, 967, 1592, 1999, 1490, 1687, 1400, 2737, 3388, 5155, 3855, 3831, 5608, 5984, 6979, 8422, 8424, 10370, 10562, 10740, 11679, 12181, 11212, 11758, 11229, 12792, 13564, 13081, 12877, 11691, 5302, 1589]. Processes [9178, 9217]. Proctor [1639, 5392, 5444]. Produce [11185, 12414]. Produced [161, 1995, 2908, 4059, 5063, 6109, 10981]. Producing [10943, 11597, 10964]. Product [11139, 13633]. Production [508, 949, 1023, 2598, 7762, 8281, 10580, 11149, 11487, 13428]. Productions [11550]. Products [1444, 1700, 1277, 6308, 6495, 5635, 218, 292]. Prof. [12880, 12742]. Profanity [5036]. Profession [1222]. Professor [510, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4936, 4967, 5042, 4607, 4642, 819, 203, 20, 43, 1067, 1798, 2127, 2069, 1842, 1902, 1844, 3712, 2707, 3093, 3116, 3165, 3197, 3224, 3249, 3285, 3315, 3343, 3403, 3419, 3090, 3539, 3662, 3506, 3543, 2433, 2837, 3428, 3529, 3743, 5218, 4006, 5023, 5049, 4033, 3889, 3924, 3802, 3837, 3811, 3954, 3865, 5405, 5281, 5309, 5335, 5387, 5383, 6053, 5392, 5444, 6184, 6212, 7765, 9269, 8605, 9137, 9201, 9969, 10801, 9614, 10876, 10094, 10768, 11835]. Professor [11775, 11171, 11706, 11525, 11725, 12088, 11940, 12741, 2315]. Profitableness [9219]. Profits [319]. Profuse [7453]. Prognosis [12589]. Progress [1047, 926, 305, 643, 905, 1160, 647, 727, 1069, 992, 96, 73, 620, 601, 2245, 1194, 1545, 1279, 1411, 1459, 2128, 2344, 2199, 1253, 1193, 3678, 2974, 3565, 3308, 3660, 2888, 3046, 3064, 3091, 3082, 3774, 3309, 2716, 2468, 3841, 4691, 5060, 5007, 4723, 5034, 4551, 4016, 5131, 5439, 5269, 5926, 5268, 5298, 5655, 5488, 5351, 7606, 7659, 7177, 6793, 6818, 7483, 6964, 7125, 7020, 8561, 8588, 8537, 8387, 8361, 8509, 8540, 8097, 9238, 9882, 9586, 9651, 9667, 10139, 10564, 10309, 10530, 11230, 11105, 11295, 12068]. Progress [11773, 11804, 12100, 11660, 12087, 11366, 12222, 12236, 11993, 11057, 12788, 13455, 13128, 13148, 12359, 12605, 13301, 13007, 13686]. Project [2221, 11285]. Projected [701]. Projectile [2712, 2969, 13668, 12971]. Projectiles [5213, 7272]. Projection [10090]. Projects [326]. Promising [11973]. Promote [12640]. Promoter [7, 11774]. Promotion [3882]. Promotions [3280, 3340]. Proof [1246, 5362]. Propagation [9577, 9915]. Propelled [4454]. Propeller [1845, 1990, 4293, 4996, 4833, 4675, 4668, 5723, 6295, 5988, 7436, 7753, 7878, 9200, 13251]. Propellers [3221, 11175]. Propelling [1622, 2807, 3473, 8459, 9669, 11270, 11239, 13105, 13039]. Proper [1433, 8628]. Properties [2192, 1516, 4696, 3857, 9906, 12149, 241, 3589]. Property [2064, 2173]. Prophesies [9745]. Prophets [8154]. Proportion [324, 4316]. Proportions [546, 3923, 4549, 7579, 6774]. Proposed [201, 300, 295, 1595, 1774, 2297, 1248, 2211, 2469, 2927, 3771, 2820, 5040, 4062, 4145, 5485, 5735, 6962, 6737, 7756, 7552, 6912, 10904, 10915, 9732, 11474, 12150, 11515, 11339, 11438, 11569, 13120, 13342, 13300, 11691, 3705]. Proprietors [1170]. Propulsion [1613, 2273, 2926, 2993, 4833, 4265, 5603, 10220]. Prospect [9237, 6609, 11325]. Prospective [12478]. Prospects [1408, 3537, 3854, 6425, 6793, 7330, 6717, 9397, 12087, 13040, 10145]. Prospectus [12453, 12507]. Prosper [5301]. Protect [8185]. Protected [10908, 10688]. Protecting [7334, 12171]. Protection [2327, 2373, 3216, 2797, 5160, 3980, 5896, 6134, 9747, 12284, 12485, 13480]. Protective [12856]. Protector [368, 1097, 2261, 3784, 4527, 4909, 4356, 6872, 10383, 10940, 11934, 11909, 12527]. Protest [18, 2651]. Protoplasm [861]. Protractor [12052]. Prove [1533]. Provided [5084]. Provision [12918]. Pruning [1537, 1349, 5306, 6655, 6823, 13674]. Prussia [831, 807, 2086]. Prussian [882, 744, 3756, 876]. Pseudo [4146]. Psychic [1966, 2269, 2266, 2070, 2103, 2698, 5353]. Psycho [7108, 7693]. Ptarmigan [11220]. Public [732, 562, 89, 2321, 2390, 2298, 1197, 2905, 2797, 2866, 4548, 4657, 7813, 6781, 11134, 12023, 13581, 12434, 13608]. Publication [2337, 3467, 3708, 2669]. Publications [3, 41, 936, 978, 1001, 1022, 1064, 1088, 1112, 1134, 1156, 64, 85, 109, 185, 213, 239, 264, 288, 310, 333, 382, 410, 437, 481, 569, 612, 634, 656, 758, 802, 827, 852, 891, 915, 1236, 1261, 2359, 2380, 1287, 1314, 1340, 1424, 1475, 1505, 1562, 1586, 1610, 1633, 1683, 1763, 1788, 1815, 1841, 1955, 2008, 2030, 2121, 2162, 2235, 2259, 2308, 2334, 1417, 1469, 2458, 2482, 3255, 3293, 3325, 3382, 3439, 3521, 3555, 3580, 3638, 3670, 3735, 2510, 2536, 2562, 2584, 2613]. Publications [2639, 2683, 2729, 2752, 2781, 2803, 2832, 2855, 2876, 2915, 2935, 2957, 2982, 3078, 3122, 3149, 3175, 3205, 3234, 3822, 3876, 5078, 5184, 3908, 3937, 3994, 4023, 4052, 4087, 4111, 4165, 4221, 4253, 4310, 4335, 4363, 4417, 4446, 4537, 4558, 4653, 4712, 4738, 4786, 4869, 4898, 4972, 5237, 5263, 6205, 6289, 6346, 6419, 6463, 5290, 5343, 5370, 5400, 5431, 5481, 5508, 5533, 5593, 5625, 5677, 5702, 5758, 5782, 5810, 5838, 5864, 5889, 5916, 5945, 6021, 6075, 6101, 6126, 6173, 6568, 6638, 6665, 6733, 6787]. Publications [6812, 6861, 6930, 6987, 7039, 7200, 7321, 7447, 7521, 7597, 7627, 7710, 7741, 7774, 7803, 7864, 7920, 7976, 9389, 8007, 8068, 8122, 8149, 8237, 8266, 8377, 8440, 8555, 8582, 8641, 8698, 8727, 8757, 8855, 8887, 8946, 9039, 9071, 9165, 9196, 9229, 9296, 9329, 9520, 9602, 10921, 10948, 9628, 9686, 9713, 9738, 9845, 9871, 9900, 9951, 10023, 10054, 10076, 10105, 10133, 10186, 10243, 10300, 10398, 10491, 10554, 10588, 10677, 10707, 10789, 10815, 11000, 11076, 12139, 12164, 12190, 12242, 11153, 11179, 11251, 11306, 11334, 11386, 11413]. Publications [11443, 11468, 11491, 11575, 11602, 11629, 11676, 11744, 11768, 11792, 11820, 11851, 11875, 11898, 11945, 12038, 12084, 12350, 12378, 13374, 13395, 13421, 13476, 13501, 13549, 13575, 13600, 13626, 13655, 12428, 12510, 12546, 12570, 12691, 12783, 12841, 12868, 12899, 12952, 13002, 13031, 13060, 13087, 13114, 13139, 13166, 13192, 13215, 13242, 13267, 13294, 13322, 13350]. Publisher [3563, 5188, 5212, 4563, 4594, 4630, 6514, 6544, 7808, 7840, 7869, 9396, 9425, 9459, 9494, 9877, 9719, 9744, 9768, 9799, 9837]. Publishers [9850, 3534]. Publishing [300]. Puddle [4953]. Puddler [6700]. Puddlers [12982]. Puddling [943, 2987, 2924, 10741, 9550, 11038, 11779]. Pugnacious [7553]. Pull [2968, 9530]. Puller [8600, 10666]. Pulley [6, 3362, 4470, 6065, 5935, 7666, 7816, 8483, 11884, 12798]. Pulleys [1431, 2864, 3591, 8596, 10776, 12449]. Pullman [983, 7843]. Pulp [10197, 10086]. Pulsometer [4767]. Pulverizer [11069]. Pump [1161, 1990, 3556, 2518, 3124, 2459, 4726, 4705, 5220, 4180, 4010, 4667, 4042, 5706, 5609, 5720, 6302, 7882, 6750, 7005, 7025, 7239, 7412, 6971, 8389, 9181, 9117, 8797, 8020, 8626, 8358, 9434, 10379, 10352, 9832, 9804, 9862, 11221, 11563, 11264, 11623, 11235, 12125, 12126, 12738, 12938, 13287, 13049, 12388]. Pumping [154, 76, 4458, 4225, 5634, 5537, 5423, 8180, 8186, 11376, 11796, 11264, 11191, 12368]. Pumps [2950, 5961, 7959, 10042, 10837, 11019, 12943, 13642]. Punch [5854, 7451, 13362]. Punching [452, 2878, 3353, 6064, 9025, 8769, 9704, 10314, 11237, 13334]. Purchase [5158]. Purchasers [4397]. Pure [2529, 11991]. Purification [5472]. Purifier [4790, 7427, 8187, 11495, 11007]. Purifying [1342, 12872, 273]. Purity [12259]. Purposes [508, 832, 370, 1855, 2462, 2488, 4693, 6445, 5363, 7022, 921]. Push [1765, 10503]. Push-Pin [10503]. Pushing [13483]. Put [7316]. Putnam [9395]. Putt [3785]. Putting [9696, 31]. Puzzle [9212, 10447, 11186, 12879]. Pygmy [12860]. Pyramid [10878]. Pyramids [3952]. Pyrenean [11433]. Pyrites [5321]. Pyrometer [1141, 1398, 7307]. Pyrophone [5669]. Pyrotechnic [1103]. Pyx [12094].

Quacamaia [13083]. Quadruplex [10144]. Quality [3510, 5305]. Quandary [12360]. Quantitative [5091, 6914, 2763]. Quantitativelt [8250]. Quantitatively [8222]. Quantities [11149, 5245]. Quantity [1878]. Quantivalence [5818]. Quarries [12134]. Quarrying [4142]. Quarter [1917]. Quartz [1321, 4525]. Queensland [4031]. Queer [6216, 7405, 12003]. Queries [1082, 1107, 1129, 1151, 1202, 1230, 1907, 1925, 1950, 1979, 2003, 2025, 2049, 2074, 2092, 2135, 2157, 2176, 2207, 2230, 2254, 2278, 2302, 2329, 2354, 2403, 1255, 1307, 1336, 1363, 1385, 1416, 1447, 1470, 1502, 1524, 1555, 1579, 1604, 1629, 1654, 1678, 1705, 1731, 1756, 1781, 1812, 1834, 1859, 1882, 1751, 2431, 2456, 3053, 3076, 3101, 3121, 3147, 3173, 3204, 3232, 3257, 3290, 3323, 3350, 3380, 3411, 3438, 3465, 3519, 3553, 3579, 3615, 3640, 3672, 3701, 3737, 2484, 2512, 2538, 2560, 2582, 2611, 2637, 2662, 2680, 2704, 2732]. Queries [2754, 2779, 2805, 2831, 2854, 2875, 2896, 2913, 2933, 2959, 2980, 3003, 3029, 3798, 3824, 4591, 4627, 4655, 4714, 4740, 4765, 4788, 4818, 4871, 4900, 4921, 4948, 4974, 5030, 5056, 5080, 5102, 3878, 3910, 3939, 3967, 3996, 4025, 4054, 4089, 4113, 4140, 4167, 4196, 4223, 4255, 4280, 4312, 4337, 4365, 4392, 4419, 4448, 4481, 4539, 4560, 6594, 6617, 7252, 7278, 7298, 7323, 7353, 7377, 7400, 7425, 7449, 7477, 7500, 7523, 7547, 7573, 7599, 7629, 7653, 7682, 7712, 7743, 7776, 7805, 7836, 7866, 7894, 6640, 6667, 6690]. Queries [6711, 6735, 6763, 6789, 6814, 6836, 6863, 6888, 6905, 6932, 6959, 6989, 7013, 7041, 7063, 7090, 7121, 7146, 7171, 7202, 7229, 7922, 7950, 8672, 8700, 8729, 8759, 8789, 8821, 8857, 8889, 8917, 8948, 8978, 9009, 9041, 9073, 9107, 9133, 9167, 9198, 9231, 9265, 9298, 9331, 9361, 9391, 9422, 9454, 7978, 9490, 8009, 8039, 8065, 8070, 8124, 8151, 8178, 8209, 8239, 8268, 8294, 8319, 8343, 8379, 8411, 8442, 8469, 8501, 8532, 8557, 8584, 8614, 8643, 9953, 10217, 10817, 9522, 9547, 10274, 10302, 10335, 10367, 10400, 10429, 10459]. Queries [10493, 10521, 10556, 10590, 10622, 10653, 10679, 10709, 10737, 10763, 10791, 10847, 10872, 10897, 10923, 10950, 9574, 9604, 9630, 9661, 9688, 9715, 9740, 9764, 9794, 9819, 9847, 9873, 9902, 9931, 9977, 10001, 10025, 10056, 10078, 10107, 10135, 10162, 10188, 10245, 11053, 11077, 11180, 11252, 11278, 11656, 12165, 10974, 11002, 11769, 11793, 11821, 11852, 11876, 11899, 11919, 11946, 11965, 11986, 12016, 12039, 12062, 12085, 12111, 12140, 12191, 12218, 12243, 12273, 11024, 11101, 11127, 11154, 11202, 11225, 11307, 11335, 11358, 11387, 11414, 11444, 11469, 11492, 11520, 11546, 11576]. Queries [11603, 11630, 11677, 11700, 11720, 11745, 12571, 12869, 12900, 12926, 12953, 12977, 13003, 12299, 12326, 13193, 13216, 13243, 13268, 13295, 13323, 13351, 13375, 13396, 13422, 13452, 13477, 13502, 13526, 13550, 13576, 13601, 13627, 13656, 13683, 12351, 12379, 12402, 12429, 12456, 12480, 12511, 12547, 12600, 12636, 12662, 12692, 12724, 12755, 12784, 12811, 12842, 13032, 13061, 13088, 13115, 13140, 13167]. Queris [12571]. Querise [12165]. Question [507, 555, 2320, 2391, 1668, 2587, 3948, 5658, 10566, 11081]. Quests [11594]. Quick [6188, 11087, 11839, 2963]. Quick-Speed [11087]. Quicklime [12252]. Quill [11065]. Quinby [2623]. Quinine [13343]. Quinnipiac [2318]. Quintino [4774]. Quoins [4463]. Quries [11656].

R [3093, 3116, 3165, 3197, 3224, 3249, 3285, 3315, 3343, 3403, 3419, 3428, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4936, 4967, 5042, 4607, 4642, 5444, 10094]. R. [9282]. Rabbits [10942]. Race [9745]. Races [495, 7332, 8893]. Rack [2869, 10892, 9674]. Radial [4213, 10432, 11066, 11671]. Radiant [2707, 13129]. Radiata [9026]. Radiating [4044]. Radiation [815, 6249]. Radiator [4465]. Radiometer [7194, 9464]. Rafferty [5829]. Raffineurs [786]. Raft [222, 5661, 7437]. Rafter [2292, 3502, 7435]. Rag [4406, 7366, 8166]. Raid [11472]. Raids [10578]. Rail [3112, 3336, 2961, 7264, 7313, 8216, 7963, 8772, 8964, 11186, 12176]. Railroad [650, 909, 1092, 323, 1044, 1453, 1303, 1274, 1253, 1496, 2619, 3150, 3788, 3013, 2820, 3111, 3112, 2961, 4913, 4199, 5014, 4887, 4994, 3986, 3980, 5493, 6273, 5668, 6326, 8714, 8011, 8213, 8247, 8860, 10344, 9905, 10121, 11658, 13232, 13180, 983, 415, 1403, 4517, 9723, 12144, 11981]. Railroads [963, 1394, 3540, 9475]. Rails [323, 1572, 4199, 6470, 10504, 12225, 12816]. Railway [9, 16, 922, 872, 265, 1465, 1554, 2105, 2221, 1634, 2339, 2344, 1514, 2290, 1960, 2706, 3277, 2834, 2972, 3213, 3223, 3010, 3757, 3019, 3218, 3336, 3601, 3598, 2721, 5225, 4378, 4873, 4950, 4060, 4158, 4016, 4883, 4015, 5046, 3813, 6466, 6210, 6503, 6111, 6542, 6470, 6408, 6430, 6450, 6500, 6553, 6171, 6656, 7181, 7206, 6945, 6828, 7309, 7288, 7587, 7355, 7695, 7667, 7264, 7313, 7552, 6938, 7610, 6579, 6677, 6720, 6934, 8962, 7963, 8772, 8516, 8833, 8477, 8098, 8895, 10769, 10434, 10311, 10463, 10524, 10373, 9664]. Railway [10961, 11080, 12128, 11913, 11012, 12225, 11362, 11254, 13382, 13484, 12464, 13300, 12796, 12913, 13103, 13155, 13144, 425, 2735, 2687, 4259, 4287, 3838, 7980]. Railways [619, 1440, 1833, 1971, 1575, 3166, 3136, 3274, 4124, 4011, 4225, 4598, 6340, 6364, 5655, 6062, 5923, 5667, 9112, 9374, 8021, 10682, 10626, 11579, 13541, 13274, 13586, 13455, 773]. Rain [857, 91]. Rainfall [9961]. Rains [6216, 7453]. Raising [863, 4726, 6255, 8184, 10836, 11957, 11175, 12533, 12594]. Rake [669, 4937, 7219, 7242]. Ralston [4158]. Ram [4261, 5434, 6333, 11241, 13465]. Rambles [1458, 1484]. Rambling [9153, 9185]. Ramie [10754, 11095]. Rammer [13515]. Ramming [5088]. Rancid [12872]. Rand [1820]. Range [8874]. Ransom [168, 11218]. Ransome [1711]. Rapid [342, 3067, 4058, 3815, 4035, 6401, 7483, 6964, 7125, 6911, 7327, 8284, 7987, 9376, 8321, 12048, 13304]. Rapieff [12315]. Rare [4998, 9015]. Rasps [11841]. Rat [4266, 9780]. Ratchet [111, 12530]. Rate [7304, 9029]. Rates [7098, 6755]. Rating [1050, 1515, 2017]. Rats [9937]. Rattan [1291]. Ravages [527, 547]. Raw [4348]. Ray [5463]. Rays [3741, 4034, 3924, 3589]. Razing [1081]. Razor [2014]. Re [1975, 1801, 3340, 7751, 6778, 8762, 8127, 10471, 13027, 13056]. Re-Agent [1975]. Re-Bore [10471]. Re-Discovered [7751]. Re-Discoveries [8127]. Re-Established [13027, 13056]. Re-Inventions [8127]. Re-Sawing [6778, 8762]. Re-Vaccination [1801]. Read [674]. Reader [4029]. Readers [2586, 2596, 11060]. Reading [4288, 7874, 9635, 13276]. Ready [8503, 11450]. Real [1776, 2015, 3508, 9272, 11836]. Reaper [2309, 4105]. Reapers [6850, 11608]. Reaping [4603, 7418]. Reasons [2870, 7876, 12263, 13506]. Reat [1744]. Rebolting [6104]. Rebuilt [5108]. Receive [553]. Received [3432, 3548, 3665, 3761, 3792, 3819, 3906, 4684, 4710, 4735, 4762, 4784, 4815, 4840, 4867, 4896, 4918, 4945, 4970, 5002, 5026, 5053, 5076, 5099, 5127, 5151, 5182, 5205, 5234, 3964, 3992, 4021, 4050, 4084, 4109, 4137, 4164, 4192, 4251, 4277, 4308, 4333, 4361, 4389, 4415, 4444, 4477, 4505, 4535, 4556, 4584, 4620, 4650, 5259, 5286, 5968, 5992, 6018, 6043, 6072, 6098, 6123, 6149, 6176, 6202, 6234, 6261, 6286, 6317, 6343, 6368, 6395, 6416, 6439, 6460, 6484, 6508, 6536, 6567, 5315, 5339, 5367, 5397, 5428, 5455, 5478]. Received [5505, 5531, 5561, 5590, 5623, 5645, 5674, 5699, 5726, 5755, 5779, 5807, 5835, 5861, 5887, 5913, 5942, 6590, 6613, 7248, 7274, 7294, 7318, 7349, 7373, 7396, 7421, 7444, 7473, 7496, 7544, 7569, 7594, 7624, 7650, 7678, 7708, 7739, 7771, 7800, 7833, 7861, 7891, 6635, 7919, 6662, 6686, 6707, 6730, 6759, 6784, 6809, 6833, 6858, 6884, 6901, 6927, 6956, 6985, 7010, 7037, 7060, 7087, 7119, 7143, 7168, 7198, 7226, 7946, 7973, 8696, 8725, 8755, 8787, 8818, 8852, 8884, 8913, 8943, 8975, 9005, 9036, 9068, 9103, 9129, 9162]. Received [9192, 9226, 9261, 9293, 9326, 9356, 9386, 9418, 9450, 9487, 8005, 9519, 8035, 8061, 8092, 8119, 8146, 8175, 8205, 8234, 8264, 8291, 8316, 8340, 8375, 8408, 8438, 8467, 8498, 8529, 8553, 8581, 8611, 8639, 8669, 9545, 9572, 10272, 10299, 10333, 10365, 10397, 10427, 10457, 10490, 10519, 10553, 10587, 10620, 10651, 10676, 10706, 10735, 10761, 10788, 10814, 10868, 10893, 10919, 10945, 10971, 9601, 9627, 9659, 9685, 9712, 9737, 9762, 9792, 9817, 9844, 9870, 9899, 9929, 9950, 9975, 9999, 10022, 10053, 10075, 10104, 10132, 10185, 10215, 10242, 10997]. Received [11020, 11942, 11963, 12013, 12035, 12082, 12108, 12136, 12162, 12188, 12215, 12269, 12296, 11049, 11098, 11249, 11276, 11304, 11332, 11356, 11383, 11410, 11440, 11464, 11489, 11518, 11544, 11573, 11600, 11627, 11742, 11766, 11790, 11849, 11872, 11896, 12323, 12346, 13085, 13111, 13137, 13189, 13212, 13239, 13264, 13291, 13320, 13348, 13371, 13393, 13418, 13450, 13473, 13498, 13523, 13546, 13623, 13653, 13681, 12399, 12425, 12451, 12476, 12506, 12542, 12567, 12598, 12633, 12660, 12688, 12721, 12752, 12780, 12808, 12839, 12865, 12897, 12924, 12950, 12975, 13000, 13028, 13057]. Receiver [11436]. Receiving [10932]. Recently [11983, 12544]. Reception [2976]. Recipes [667, 1397, 1435, 1460, 1509, 3155, 3226, 3260, 3497, 5415, 7718]. Reciprocating [1322, 8279, 12938]. Reclaiming [3441, 5630, 8562]. Recoil [1986, 7909, 10908]. Recommended [1438]. Record [2337, 2397, 2423, 2520, 2436, 2446, 3779, 4288, 4002, 4352, 5296, 6718, 8095]. Recorder [783, 7515]. Recording [4236, 4149, 11130, 13443]. Records [2390, 2298, 6275, 9366, 10797]. Recover [1148]. Recreations [470]. Rectifying [5174]. Red [6899, 8430, 12008, 11272, 12803, 13303, 10781]. Redier [11379]. Reduction [562, 2486, 7660, 7733, 7968, 6391]. Reductions [3066]. Reed [10038]. Reel [2839, 9243]. Reeves [10995]. Reference [2797, 13364, 13386, 13408]. Refined [12259]. Refinements [7178]. Refineries [2217]. Refining [554, 1404, 3892, 6875, 12003]. Reflecting [13147]. Reflection [8617]. Reflections [2271]. Reflector [8691, 12821]. Reform [1940, 2918, 3065, 4512, 6446]. Refrigerating [161, 4688, 5925, 12933]. Refrigerative [388]. Refrigerator [1055, 3016, 4700, 7672, 7756, 6915, 8101, 12095]. Refrigerators [7106]. Refuge [12436]. Refuse [1649]. Regard [2129, 13134]. Regarded [11475]. Regards [861]. Regia [5499, 6338, 11378, 5167]. Regions [1797, 3629, 3311, 3442, 599]. Register [1766, 4214]. Registering [4066, 7213, 8224, 10569, 11379]. Registration [13631]. Registrations [2145, 2151]. Regnault [11514]. Regnier [11036]. Regulating [1438, 2022, 5740, 8711, 12179]. Regulations [2000]. Regulator [1043, 938, 803, 3383, 3414, 2903, 5460, 6165, 7103, 7642, 9054, 8196, 9967, 11796, 11507, 11298, 12123, 12053, 11506, 13206, 12648, 12524, 12387]. Reid [6852]. Rein [573, 5615, 6478, 8873, 11868]. Rejection [6869]. Relating [5106, 5897, 6698, 12046, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 13491]. Relation [1175, 2846, 2866, 6424, 6243, 11989, 2385]. Relations [925, 2770, 10852, 11134, 5142, 5177, 5201]. Relationships [7576, 10714]. Relative [1666, 5515, 5245, 2366]. Relays [1433]. Releasing [13023]. Reliable [11555]. Relics [1458, 1484, 9410]. Relief [967, 10111]. Reliefs [10943]. Religion [1546, 2846, 2600]. Reluctance [553]. Remaining [4906]. Remains [2614, 4314, 5710, 11595]. Remarkable [2316, 1334, 1407, 2202, 1355, 3507, 2486, 3682, 3645, 3654, 3312, 2798, 3677, 3163, 4153, 4424, 5857, 5537, 6727, 7020, 6829, 8605, 8477, 9499, 8481, 8186, 9555, 10799, 10044, 9961, 12067, 12117, 11062, 11562, 11997, 12803, 12817, 13689, 2385]. Remarks [2315]. Remedies [4028, 10404]. Remedy [9833, 2645, 2650, 6162, 6294, 7303, 9426, 12649, 12329]. Remington [2808]. Reminiscences [1918]. Removable [5550]. Removal [1623, 1658, 2588, 9121, 8098]. Removing [1173, 2734, 7049, 6963]. Rendering [4854, 8027]. Renewable [2761]. Renewal [4513]. Renewed [3340]. Rennes [12125]. Renovator [4468]. Repaired [67]. Repairing [1775]. Repairs [2324]. Repeal [6587]. Repealed [8829]. Repeat [3109]. Repeating [3214]. Repetition [2389, 10797]. Reply [2266]. Report [966, 1143, 203, 43, 180, 626, 703, 402, 49, 323, 1382, 1438, 1809, 1695, 1494, 2068, 2566, 2569, 2766, 3999, 3979, 5133, 5243, 5537, 6792, 9728, 11132, 11361, 2366, 6480]. Reporter [3368]. Reports [1688, 2759, 6353, 12172]. Representation [10310]. Representative [4791]. Representatives [12552]. Reproduction [11074]. Reptiles [90]. Republic [561]. Required [3884, 5976]. Requirements [2974, 12874]. Rescue [4906]. Research [1093, 3882, 4977]. Researches [2212, 3712, 3931, 6133, 6749, 11674].

Resemblances [662]. Reservoir [1199, 2940, 5776, 7333]. Reservoirs [10837]. Residences [7271]. Resignation [1078, 7409, 9556]. Resistance [1433, 3745, 3039, 4233, 9091]. Resistances [13652]. Resisted [3494]. Resonant [13514]. Resources [929, 1497, 6131, 8845, 8877, 8183, 9801, 9707]. Respect [4373, 4342, 5515, 7234]. Respiration [5767, 9647, 11762]. Respirator [6844]. Rest [3526, 4672, 8129]. Restaurant [2124]. Restoration [11995]. Restoring [12003]. Restrictions [281]. Result [2925, 10835]. Results [2503, 2486, 4093, 5542, 5294, 7019, 9406, 9851, 9961]. Retirement [3776]. Retouching [9913]. Retreating [7880]. Retrospect [5950]. Return [5522, 9304]. Rev. [2212]. Revealed [3773]. Revelations [6845, 7482, 5718]. Revenue [4755, 49, 3801]. Reversible [421, 2196, 1894, 3157, 5140, 4170, 7538]. Reversing [941, 4775]. Review [674, 291]. Reviewer [440]. Revival [3706]. Revolver [5266]. Revolving [392, 1790, 4530, 4169, 5920, 6752, 7241, 8094, 10940, 12336]. Reward [1613, 1622, 3974, 4513, 6162, 5378, 8042, 8073, 10821, 13530]. Rewarded [695]. Reynier [12002, 13361, 13487]. Rheostat [12105, 11706, 13282]. Rheostatic [11246]. Rheumatism [4344]. Rhinoceros [11785]. Rhyme [2787]. Ribbed [9314]. Ribbon [5446]. Rice [3006, 9858, 10987, 10480]. Rich [12894]. Richard [714]. Rider [8047, 8078, 11191]. Rifle [26, 2079, 3214, 3756, 4519, 6313, 7292, 7431, 9119, 10599, 10985, 12006]. Rifles [9343]. Riggs [8364]. Righi [12629]. Right [201, 385, 705, 2064, 7152, 9312, 8734, 10014, 13036, 12815]. Right-Handed [8734]. Rights [623, 4551, 4397, 5515, 6738, 7234, 8246, 9803, 11516, 11312, 12121, 11856, 11472]. Rigi [4883]. Rim [12449]. Ring [810, 1744, 3536, 2997, 11844]. Ringing [13283]. Rings [11584]. Rinsing [12102]. Ripping [7664]. Rise [10060, 11256]. Rivals [6085]. Rive [5324]. River [635, 1792, 1547, 1809, 1403, 1271, 1674, 1408, 2409, 3769, 3092, 3524, 3269, 4151, 3895, 5572, 5346, 7807, 6896, 7583, 7050, 7953, 7952, 9369, 9636, 10371, 10338, 9767, 9743, 11471, 11815, 11310, 12087, 12855, 13427, 13663, 13483, 12817, 12670, 156, 1394, 1410, 1799, 3195, 4106, 4491, 5023, 5049, 6006, 5281, 5309, 5335, 6503, 7602, 9238, 9810, 10628, 13091]. Rivers [11161, 11474]. Rivet [2164, 10508]. Rivet-Making [10508]. Riveter [8570]. Riveting [1821, 2989, 5770, 7300, 8354, 9616, 13462]. Rivets [9616]. Road [508, 1065, 318, 1075, 644, 2036, 3190, 3361, 3528, 2563, 4013, 3949, 5762, 7101, 6580, 8457, 10532, 11005, 12470, 12762, 12610, 12994]. Road-Building [3528]. Road-Making [1075, 8457]. Road-Ways [508]. Roads [176, 1075, 660, 1775, 11797, 11968]. Roadways [1572, 2128, 2471]. Robaugh [9804]. Robbery [3341, 12117]. Robbins [1664]. Robert [12608, 4613, 5489, 12958]. Roberts [7540, 11522]. Robertson [7879, 9920]. Rock [3195, 8225, 10122]. Rock-Drilling [8225]. Rocking [1486, 1392, 4677]. Rockport [12134]. Rocks [1517, 9121, 10807]. Rocky [8955, 9964]. Rod [2617, 2847, 5228, 4724, 5640, 7899, 9250, 8628, 8363, 9908, 9800, 11709, 13201]. Roderick [2250]. Rodman [1967]. Rods [1828, 1943, 2826, 3126, 4719, 4795, 7507, 7555, 8458, 9017, 10170, 9773, 13222].

Roebling [626]. Roger [1034]. Rogers [733]. Role [2997]. Roll [5828, 8988]. Rolled [450, 3668, 3697, 3732, 3764, 12449]. Roller [548, 3363, 4938, 4409, 13209]. Rollers [2122, 6580, 10532]. Rolling [1005, 644, 2413, 3587, 4775, 6007]. Rolls [1431]. Roman [13542]. Romance [2202, 1350]. Rome [9619]. Roof [2619]. Roofing [1597, 3185, 7051, 8769, 12208]. Roofs [2310, 1597, 11429]. Room [666, 958, 4799, 5325, 7745]. Rooms [7874]. Rooster [7462]. Root [538, 1100, 698, 4348, 10692]. Roots [7818, 9647, 11967, 12984]. Rope [534, 484, 1634, 2739, 2756, 3418, 4873, 7262, 6934, 8902, 9595]. Rose [7966]. Roses [9427, 13303, 13250]. Ross [10859]. Rosset [13407]. Rot [5022, 6189]. Rotary [546, 2667, 2738, 3186, 2564, 4629, 5140, 4510, 4437, 6400, 6080, 5797, 7266, 7643, 6642, 7005, 7366, 6700, 6628, 8704, 9055, 9457, 8345, 10796, 9832, 10345, 11563, 12097, 12768, 12889, 13021, 13049, 12388]. Rotating [307]. Rotation [2812, 2564, 11657, 11809]. Rotative [6080]. Roth [10988]. Roughing [3294]. Round [2044, 4467]. Route [4607, 4642]. Roving [11848]. Rovings [87]. Row [12467]. Rowe [9531]. Rowell [12826]. Rowing [8572, 8893]. Rowland [6226]. Rowlock [10354]. Royal [10065, 5883]. Rubber [774, 488, 1867, 2087, 1491, 3048, 2945, 4635, 6167, 5790, 6492, 6820, 8349, 9214, 10042, 9781, 11147, 11817, 11723]. Rubies [10981]. Rudder [6221, 11478]. Rudimentary [7240]. Ruhe [368]. Ruhmkorff [1104, 11110]. Ruins [6384]. Rule [1533, 4342, 6354, 6817, 12153]. Ruler [1617, 12497]. Rules [2000, 2757, 5517, 11762, 13344]. Rulofson [12307]. Rumford [151, 1660]. Runaway [7055]. Runk [1691]. Running [1692, 10315]. Ruptures [1343]. Rural [4354]. Rush [6768]. Russia [873, 1960, 3565, 6772, 12518]. Russian [2572, 3703, 7370, 9218, 12130, 11529]. Rustic [10446]. Rusting [6470]. Ruthenium [487]. Rutherfurd [255]. Ryi [6646].

S [274, 37, 353, 376, 434, 627, 690, 1832, 1842, 11171]. Saccharometer [11651, 11268]. Sack [12337]. Sad [6802, 6798, 6751, 13209]. Sad-Iron [13209]. Sadiron [2196, 7823]. Sadlier [11175]. Safe [743, 597, 2239, 3784, 5120, 5362, 9370, 9053, 8045, 8076, 11807, 11112, 12918]. Safes [1643, 2228, 2001, 2310, 4285]. Safety [720, 384, 445, 2215, 2105, 1937, 2651, 3501, 3041, 3364, 2648, 4198, 4067, 4667, 5046, 3864, 5615, 5597, 5667, 6505, 5981, 6825, 7027, 7539, 6890, 8714, 10728, 9700, 10115, 10111, 9607, 11134, 11633, 11482, 11764, 13696, 12612]. Safety-Cap [720]. Sahara [12748]. Said [160]. Sail [5716]. Sailing [4119, 4718, 10140, 11362]. Sailor [3417]. Sails [3723]. Saki [12007]. Salamander [10449, 10781]. Sale [295, 790, 1438, 2022, 2295, 3366, 5178]. Salesman [9755]. Salicylic [7381]. Saline [10601]. Salisbury [12170]. Saloon [6651, 6681]. Saloons [5411]. Salt [53, 297, 1659, 7562, 7045, 6894, 9369, 9140, 10033, 11829]. Salts [161]. Sambur [7791]. Sample [4127, 9244, 12965]. Sample-Weighing [9244]. Sampler [9916]. Samson [7179]. Samuel [5251, 6608, 2823, 12517, 1119].

San [3223, 11948]. Sanatorium [13389]. Sand [598, 3454, 3388, 4927, 5381, 9021, 11646, 13555]. Sand-Wasp [598]. Sandal [13198]. Sandpapering [12448]. Sands [11434]. Sandy [863]. Sane [8984]. Sanitary [4512, 4435, 4383, 4464, 4502, 4526, 12093, 11448, 12226, 13200, 12312]. Sanitation [12023]. Sap [9344, 11937]. Saponifier [2168]. Sapphire [5278, 5304]. Sapphires [10981]. Saratoga [531, 768, 2619]. Sardine [6704]. Sargent [10561]. Sarvin [2123]. Sash [2107, 1894, 1716, 3242, 3591, 3363, 4989, 4751, 4933, 6410, 5639, 8163, 8256, 10610]. Sashes [1872, 6879]. Satchel [12881]. Satellites [5793, 10498, 10729, 11584]. Satisfactory [12092]. Saturated [4696]. Saturator [12387]. Savants [3509]. Save [1089]. Saver [10613]. Saves [6428]. Saving [3595, 3131, 7533, 10081, 11804]. Savings [5171, 8045, 8076]. Saw [1071, 874, 293, 690, 1359, 1537, 1820, 3786, 2546, 2718, 3720, 3035, 4077, 3941, 4215, 4156, 4115, 5045, 3894, 5980, 5553, 5881, 5496, 6922, 7074, 6669, 7159, 7632, 6778, 7937, 8808, 8279, 9213, 8964, 9282, 10913, 9987, 10663, 12233, 11409, 11297, 11453, 12293, 10988, 11330, 13184, 12302, 13016]. Saw-Gumming [293]. Sawdust [13636]. Sawing [892, 834, 1790, 2769, 3295, 3035, 6155, 6388, 7714, 7336, 6778, 8762, 8050, 8081, 7936, 9076, 10937, 10806, 13078, 13409]. Sawmill [200, 9332, 9534]. Saws [1081, 1210, 4078, 5468, 9962]. Sawyer [12194, 12553]. Saxby [3601, 6890]. Say [639, 24, 8448]. Scaffold [5580]. Scaffolding [13257]. Scale [3399, 3338, 7435, 9244, 10347, 10113, 12012, 12910]. Scaler [10037]. Scales [13412]. Scarabæids [875, 833]. Scarfs [2811]. Scarlet [12789]. Scarus [13083]. Scenery [9526]. Scenes [8712]. Scertain [8628]. Schedule [4585, 4621, 4651, 6177]. Schemes [5874]. Schiller [7095]. Schofield [5548]. Scholarships [6513]. Schönbrunn [8026]. School [884, 581, 883, 179, 1401, 5325, 6602, 9336, 11120, 11586, 577, 2011]. Schooling [6891]. Schools [1598, 2348, 4395, 4092, 7067, 8937, 10317, 11311, 11836, 11994]. Schopp [1765]. Science [22, 1142, 674, 505, 96, 807, 89, 804, 1126, 489, 1192, 1191, 1987, 2337, 1250, 1194, 1721, 1668, 1350, 2397, 2423, 2348, 1227, 1251, 1932, 1912, 3600, 3631, 3212, 2846, 2600, 2824, 2692, 3426, 2436, 2446, 3779, 3307, 3569, 3866, 3890, 3926, 4171, 4206, 5064, 4146, 5062, 5124, 5136, 4002, 4352, 5171, 5768, 5571, 6169, 5897, 5296, 5802, 6909, 7330, 6891, 7553, 6718, 6694, 6742, 9093, 8325, 9048, 8095, 8953, 8415, 10808, 10278, 9553, 10571, 10606, 10615, 9745, 10715, 10317, 10560, 12236, 11391, 12226, 11774, 11449, 13307]. Science [12836, 13659, 12932, 13359, 12437, 12575, 13355, 504, 821, 835, 939, 959, 794, 3243, 3370, 3402, 3504, 4809, 4835, 4862, 5691, 6451, 6110, 6144, 6257, 6311, 5719, 7441, 7470, 9061, 13197, 12731]. Sciences [73, 5069, 5095, 7066, 9275, 9311, 8488, 8566, 8748, 8844, 9430, 10774, 9566, 9643, 9727, 9785, 9893, 9993, 10069, 10097, 10126, 11462, 12160, 10983, 11088, 11428, 12225, 13544, 12737, 12892, 12940, 13607, 13664]. Scientific [138, 1146, 594, 694, 100, 172, 1035, 1070, 1127, 1149, 232, 327, 391, 451, 473, 515, 558, 571, 616, 649, 677, 704, 716, 752, 775, 795, 820, 846, 885, 931, 994, 1014, 21, 883, 179, 1324, 1973, 1624, 1660, 1546, 1361, 1413, 2326, 2350, 2374, 2398, 1455, 1498, 1552, 1578, 1601, 1626, 1647, 1672, 1727, 1778, 1800, 1831, 1879, 1923, 1945, 2021, 2046, 2090, 2142, 2204, 2227, 2299, 1489, 1328, 3298, 2942, 3531, 2604, 2447, 2475, 3632, 3681, 2531, 2554, 3360, 3416, 3545, 3571, 3663, 3684]. Scientific [3783, 2576, 2602, 2629, 2656, 2675, 2697, 2726, 2742, 2767, 2792, 2818, 2838, 2885, 2922, 2949, 2975, 3025, 3069, 3132, 3167, 3198, 3284, 3314, 3334, 5034, 3882, 4423, 3901, 3843, 3869, 4850, 4881, 4929, 4957, 4985, 5012, 5038, 5066, 5096, 5112, 5137, 5163, 5194, 5226, 3900, 4036, 4075, 4098, 4121, 4152, 4186, 4212, 4291, 4321, 4347, 4376, 4426, 4457, 4493, 4518, 4553, 4636, 4664, 4697, 4731, 4749, 4773, 4828, 5921, 6119, 5252, 5276, 6139, 6191, 6328, 6380, 6427, 6449, 6481, 6497, 5310, 5327]. Scientific [5354, 5382, 5416, 5445, 5466, 5491, 5520, 5605, 5637, 5665, 5687, 5713, 5796, 5831, 5875, 5909, 5928, 5958, 5979, 6031, 6087, 6113, 5542, 6268, 6577, 6650, 6674, 7675, 7689, 7720, 7752, 7784, 7847, 6699, 6747, 6766, 6801, 6849, 6897, 6941, 6969, 7023, 7079, 7115, 7129, 7180, 7238, 7268, 7305, 7345, 7408, 7457, 7486, 7536, 7581, 7608, 7636, 8461, 8992, 7933, 7997, 8018, 8053, 8084, 8110, 8199, 8276, 8401, 8451, 8514, 8545, 8597, 9673, 10222, 10957, 10497, 10577, 11231, 11033, 12117, 11342, 11189, 12441, 13042]. Scientific [12308, 13007, 12364, 551, 787, 10866, 1045, 1066, 1090, 2514, 680, 1099, 122, 1121, 1150, 1176, 2419, 1201, 1229, 3131, 4588, 4624, 5568, 6376, 7845, 7876, 8327, 7958, 8271, 8382, 8414, 8472, 8536, 10530, 10845, 10895, 10947, 11497, 11663, 11776, 12280, 10999, 11022, 11051, 11384, 11411, 11441, 11466, 12263, 12453, 12478, 12507, 13630, 12543, 13690]. Scientifically [9174, 12024]. Scientists [5087, 9081, 11058]. Scissors [1935, 5660, 7664, 8834]. Scolloping [9339]. Scoop [11729, 11017]. Scope [4229]. Scorpion [7083, 572]. Scorpions [6192, 8718]. Scotland [863, 6602]. Scott [10881, 7051]. Scotton [4467]. Scourer [9116]. Scourge [8538]. Scouring [2165, 12102]. Scow [9809, 13091]. Scrap [10255]. Scraper [4013, 4407]. Scraps [1974, 5384]. Screen [6374, 8452, 13370]. Screens [13508]. Screw [783, 394, 941, 1456, 2309, 2164, 2998, 4293, 4702, 4129, 4549, 6382, 5920, 5723, 5414, 6295, 5603, 6791, 7436, 7883, 7878, 7029, 6842, 9312, 9278, 10439, 9730, 9912, 9618, 11732, 11847, 11478, 11175, 12319, 13614, 13251, 13257, 11590]. Screw-Cutting [9278, 9730, 9618]. Screw-Headed [11847]. Screwdriver [5854, 10043]. Screws [397, 1542, 9988]. Scroll [892, 1790, 2769, 8279, 7936, 9213, 10913, 9987, 10937, 10663, 11330]. Scroll-Sawing [892]. Scrubbing [1319, 11349]. Scuffle [12990]. Scums [2734]. Scura [2687]. Sea [731, 894, 390, 1702, 1588, 2294, 1637, 1670, 3835, 3805, 4852, 4433, 4962, 4667, 5739, 5791, 7558, 7603, 7361, 9141, 9183, 9409, 9140, 10883, 10096, 10713, 10356, 11259, 11780, 11029, 11302, 12079, 11369, 13580, 13425, 12612, 13495, 1029, 2261, 13703]. Sea-Horse [1637]. Seaforthia [5775]. Seal [386, 1685]. Seals [1822]. Seance [11636]. Search [4367, 7236]. Seas [2756]. Season [8794]. Seasonable [11355]. Seasonal [9215]. Seasoning [10496]. Seasons [6974]. Seat [3753, 7786, 6753, 8139, 10064, 13565]. Seating [12486]. Seats [9113]. Sèbillot [13199]. Second [1034, 3837, 4324, 4350, 12020, 2182]. Secret [5006, 7151, 10203, 11976]. Secretary [6980, 9281, 10716]. Secrets [8385, 10201]. Section [11447, 11906, 12362, 12408]. Sectional [4027, 4102, 5814, 6061, 7025, 11169, 13534]. Secure [11497]. Securing [2590, 7695, 8832]. Security [3271, 12670]. See [6697, 7452]. Seed [3179, 3715, 3927, 101, 130, 143, 189]. Seedlings [3659]. Seeds [2643, 2916, 11959, 1590, 2860]. Seeing [6578, 7961, 10222]. Seel [7368]. Seen [2081]. Seifert [1619]. Selden [4180, 9862]. Selecting [570]. Selection [9823]. Selenium [4148]. Self [545, 988, 855, 764, 334, 70, 918, 690, 1217, 1823, 2387, 1819, 2467, 3362, 3496, 2984, 5227, 5019, 4056, 3959, 4072, 4910, 5798, 5545, 6361, 6303, 5423, 7139, 6802, 6701, 8800, 8711, 8357, 8475, 8333, 9935, 10202, 10752, 9617, 10593, 10358, 9780, 10886, 12254, 11672, 13039, 13310]. Self-Acting [764, 5227, 4056, 7139, 10593]. Self-Adjusting [545, 2467]. Self-Centering [918, 10202]. Self-Clamp [12254]. Self-Closing [2387, 5798, 5545, 8357]. Self-Connecting [334]. Self-Cooling [10358]. Self-Corking [6701]. Self-Discharging [8800]. Self-Feeding [3496, 10752]. Self-Heating [988, 6802]. Self-Improvement [8475]. Self-Laying [13310]. Self-Lighter [6303]. Self-Lighting [4072, 8333]. Self-Locking [1217, 6361]. Self-Lubricating [3362, 2984]. Self-Measuring [1823]. Self-Moving [70, 1819]. Self-Oiling [690, 9617, 11672]. Self-Propelling [13039]. Self-Pumping [5423]. Self-Regulating [8711]. Self-Setting [855, 9780]. Self-Threading [855]. Self-Tightening [4910]. Self-Watering [10886]. Sell [6743]. Sella [4774]. Sellers [5513]. Selling [10375]. Semi [7658, 8345]. Semi-Centennial [7658]. Semi-Rotary [8345]. Semolina [9506]. Senate [5169, 2996, 3852, 3915, 6673, 11925, 12458, 12362, 12552]. Sensation [5521, 6426]. Sense [3089, 9460, 8830]. Senses [7256]. Sensible [8793, 8051, 8082, 12895]. Sensitive [8478, 9438, 11314]. Sensitizing [3533]. Sentinel [7220]. Separating [13052]. Separation [1970, 5635, 9892]. Separator [3447, 8569, 9021, 9506, 11910]. Sepot [744]. September [11996]. Sequard [5610]. Sequence [11287]. Series [3419]. Serious [3364, 3131]. Sermon [9138]. Serpent [13703]. Serpents [9409, 11509, 11242]. Serpulas [12079]. Servan [6007]. Service [1569, 1642, 3950, 5133, 6575, 12168, 11283, 1276, 1688]. Set [3851, 5553, 6601, 11209, 12817]. Seth [6776, 369]. Setter [7081]. Setting [855, 1326, 3921, 7787, 8739, 8571, 9343, 8567, 9780, 10318, 12373, 10259]. Settings [11329]. Settlement [12438]. Seven [13276]. Seventy [3309]. Seventy-Nine [3309]. Sewage [1993, 2033, 4726, 4526, 10113, 10820, 10138, 636]. Seward [1114, 3498]. Sewer [5093, 9057, 8164, 11371, 13108, 13073]. sewerage [4526]. Sewers [4942, 6774, 10574, 11150]. Sewing [855, 608, 679, 1998, 1402, 1323, 2349, 1331, 2691, 2887, 2821, 3493, 3133, 2997, 2822, 3651, 3265, 4797, 4991, 4925, 3886, 3859, 4064, 4676, 4893, 4863, 6010, 5708, 7754, 6716, 9136, 8302, 8413, 9085, 8715, 10576, 10501, 10060, 12078, 12469]. Shad [1101, 11327, 11616]. Shade [769, 6607]. Shadow [5716]. Shaft [274, 1011, 2243, 1492, 2984, 7603, 8740, 10918]. Shaft-Bar [274]. Shafting [450, 2748, 3668, 3697, 3732, 3764, 5448, 9564, 9775, 11810]. Shafts [3970, 6722]. Shaker [5639]. Shall [1074, 2568, 2608, 2634, 2657, 2594, 4189, 5424, 5474, 5620, 6014, 6095, 6172, 7504, 8925, 8131, 10463, 12484, 12552, 13144]. Shameless [2082]. Shampooing [2843]. Shaper [4244, 11564]. Shaping [2585, 7906, 10152]. Shapley [6444]. Sharpener [2521, 4244, 7632]. Sharpening [1118, 293, 7134, 11646]. Shaving [1795, 11843]. Shaw [1990, 7614, 7368, 11141]. Shawl [12744]. Sheaf [11885, 11930, 12910]. Sheaf-Binding [11885]. Shear [13362, 12439]. Shearing [3353, 7451, 7881, 11237, 13334, 13462]. Shears [1349, 5072, 5306, 6655, 7141, 7043, 10777, 11645, 11194, 12057]. Sheave [7111]. Shedding [10721]. Sheep [5072, 7141, 12057, 11909]. Sheet [2013, 3185, 7380, 7051, 10277, 11761]. Sheffield [883, 179, 1210, 7690, 8216]. Sheldon [7786]. Shell [13437]. Shellac [2055]. Sheller [2883, 5071, 9754]. Shells [8806]. Shephard [8517]. Sheppard [2435]. Sherman [1490]. Sherwood [10575]. Shield [1240]. Shielded [12447]. Shilling [13513]. Shingle [6991, 6978]. Shingles [438, 13080]. Ship [75, 701, 1545, 1819, 3020, 2917, 2459, 3835, 3805, 4578, 4496, 6197, 5322, 6298, 7370, 12221, 11825, 13274, 12393, 13300, 13155, 13144, 3709, 10724]. Ship-Railway [13300]. Shipbuilding [3706, 9983, 141]. Shipping [12473, 12676]. Ships [731, 5015, 5653, 11606]. Shive [3055]. Shoals [7885]. Shoe [2665, 4925, 3943, 8104, 9339, 10258, 10940, 11734, 11613, 12443]. Shoe-Scolloping [9339]. Shoeing [2213]. Shoemaker [1270, 6304, 4941, 7648]. Shoemakers [1829]. Shoes [2087, 3261, 4611, 6167, 387]. Shooting [10880]. Shop [10930, 11138, 13016]. Shops [3297, 3333, 11586]. Short [2628, 10746, 11071, 12248]. Short-Horned [11071]. Shortt [918]. Shot [1220, 6572, 10227, 10305, 12774]. Should [7876, 7981, 12071, 12263, 12718]. Shovel [532, 5172]. Show [3852, 3860, 3862, 3812, 4690, 7932, 3915]. Shower [3178, 4695, 11265]. Showing [6161, 11657, 11809]. Shown [5769]. Shrew [13050]. Shrinking [2367]. Shrub [6880]. Shrubbery [6479]. Shut [427]. Shute [11482]. Shutter [981, 11320, 12096, 12228]. Shutters [3587]. Shuttle [1089, 1159, 3263, 12078]. Shuttle-Motion [1089, 1159]. Siam [11948]. Siamese [5486, 5377, 5443]. Sicily [6391]. Sick [3835, 3805]. Side [19, 425, 4950, 4463]. Sideboard [7106]. Sided [11081]. Sidewalk [2834]. Sidewalks [125, 2274]. Siege [917, 948]. Siemen [1141]. Siemens [2361, 9438, 12625]. Sieve [10035]. Sifter [1351, 11530]. Sight [7481, 7221, 10222]. Sight-Seeing [10222]. Sigl [4408]. Sign [3023]. Signal [1569, 1276, 1688, 2815, 4913, 4378, 4094, 4072, 5133, 5667, 7897, 8741, 8516, 8833, 9475, 10444, 11553, 11893, 3950]. Signaling [13669]. Signals [872, 878, 1276, 3010, 3601, 7456, 6890]. Signature [1842]. Signor [10669].

Sile [11496]. Silence [9401]. Silent [347, 12253]. Silica [925, 1599]. Silicic [10563]. Silk [1485, 1701, 3630, 7911, 7157, 13554, 12469, 12562]. Silk-Spinning [7911]. Silkworm [4704, 4671, 5356]. Silkworms [3210, 4494]. Silky [10991]. Silver [1968, 1635, 2737, 9437, 8505, 11750, 11255, 11758, 11555, 11583, 11951, 11580, 12750, 13633, 8222, 8250, 3043, 336, 5359]. Silvering [77, 3953, 10984]. Silverware [10008]. Simian [11504]. Simla [11712]. Simons [2648]. Simonson [2317]. Simple [3331, 5107, 4405, 4201, 3955, 4663, 4909, 5333, 9089, 8191, 8511, 8684, 9508, 8902, 10544, 11688, 11642, 11718, 11842, 12005, 12504, 12885, 8393, 11958]. Simplified [4489]. Simultaneous [6026]. Sinai [5518]. Sinclair [2106]. Singeing [6522]. Single [274, 4170]. Singular [1033, 4175, 9208, 8703, 9910, 13647, 13278, 8335]. Sink [7551]. Sinking [4301]. Siphon [783, 168, 4149, 6355, 9310, 11218, 12683]. Siphon-Recorder [783]. Siphons [4774]. Sir [2250, 1817, 5500, 6130, 6608, 8706]. Siren [5670]. Siriped [595]. Sirius [3339]. Sirups [641]. Site [12759]. Six [601, 199, 1805, 1193, 2445, 2468, 6085, 7876, 8988, 13007]. Six-Roll [8988]. Six-Way [199]. Sixpenny [12939]. Sixth [2419, 3954]. Sixty [6162]. Size [843, 4262, 6873, 8508]. Skein [7081]. Sketch [361, 11772]. Sketching [3331, 5415]. Skill [1046, 9692]. Skilled [11526, 12173]. Skimmer [10533]. Skimmias [7731]. Skimming [13333]. Skipper [4699]. Skull [10667]. Sky [2841]. Skylights [6556]. Slade [9334]. Slat [7110, 7338]. Slate [129, 9203]. Slates [8769]. Slaughter [4849]. Slaughtering [4854, 6942]. Slaves [243]. Slavonski [516]. Slawson [1947]. Slayton [3758]. Sled [11557]. Sledge [7308]. Sleep [580, 804]. Sleeping [10665, 13536]. Sleeping-Car [13536]. Sleeve [918, 5935]. Sleigh [2243, 3261]. Sleighs [274]. Slicer [5043]. Slicing [3225]. Slide [421, 1426, 1744, 3921, 4672, 6166, 7789, 7337, 9113, 8630, 9239, 9970, 10259]. Slides [187]. Sliding [121, 9052, 10631]. Slim [9050]. Sling [6572]. Slipped [3810]. Slipper [10476]. Slippery [5462]. Slipping [4497, 12993]. Slope [7330]. Slotting [5536, 7924, 11368]. Small [1123, 66, 2645, 2797, 2552, 2650, 4028, 6353, 6086, 7284, 8471, 10696, 10113]. Smallest [12550]. Smell [3104]. Smith [7368, 1947, 1487, 2265, 2367, 3544, 5500, 6478, 7110, 7339, 7411, 7111, 6997, 7220, 8254, 11113]. Smoke [989, 1902, 2371, 2368, 5582, 8685, 10670, 12045, 11137, 11749, 12444]. Smoke-Burning [8685, 10670]. Smooth [3703]. Smoothing [10967]. Smut [13225]. Snake [4599, 10124, 10417, 12686]. Snake-Eating [10124]. Snap [3998]. Snapping [12624]. Snaps [7116]. Snoring [1243]. Snow [1173, 1353, 5172, 5190, 3980, 6963, 6845, 8098, 9912, 12964]. Snowy [10476]. Snowy-Lipped [10476]. Snyder [294, 11748]. Soap [1342, 4126, 6251, 11976, 12354, 10090]. Social [856, 1668, 1251, 5768, 5802, 9093, 8620, 13359, 12437]. Societies [5006]. Society [674, 2465, 2689, 2789, 2884, 251, 2621, 3531, 9580]. Socket [363, 7262, 7853]. Soda [641, 2650, 2516, 4469, 5540, 9914, 10699]. Soda-Carbonating [9914]. Sodium [2272]. Soft [11859, 12354, 12344].

Softening [12562, 12704]. Soil [1536, 4202, 7554]. Soissons [922]. Solanum [7007, 8519]. Solar [1195, 3741, 2573, 5617, 6249, 6793, 7480, 9503, 10254, 10405, 9697, 10822, 11596, 12026, 3589]. Soldering [4529, 8930]. Soldiers [876]. Sole [11734]. Soled [6434]. Solid [508, 115, 3140, 7640, 12186, 2715]. Solidified [5820]. Solids [6145, 5750]. Solomon [7179]. Soluble [940, 771, 1325]. Solution [1353]. Solutions [152, 1635, 10601]. Solve [12575]. Solvent [3005]. Solvents [6492, 13404]. Some [1047, 6360, 5924, 7747, 7019, 7482, 8212, 8244, 9236, 8657, 9049, 8795, 9271, 8875, 10936, 10775, 9890, 10436, 9634, 11882, 11374, 11541, 12263, 11355, 12618, 12997, 12589, 12408, 12677, 13448, 13506, 12565, 12620, 13591, 12133, 44, 10641, 2315]. Something [841, 298, 556, 2249, 2370, 3129, 2641, 3084, 6672, 11901, 1590, 336]. Song [7291]. Sophora [5219]. Sorel [991, 1592, 2041]. Sound [857, 1342, 1844, 3627, 2652, 4211, 6109, 5521, 5899, 1739]. Sounder [13025]. Sounding [606, 4962, 7822, 7558, 9475]. Soundings [894, 7361]. Sounds [6026, 11365]. Soup [3509]. Source [579, 4514, 5631, 7715, 7812, 8016, 8241, 13154, 4295, 4323]. Sources [4258, 5873, 8795, 13430]. South [2974, 5164, 9801, 12357, 1128, 1057, 972, 1016, 1036, 1080, 1105, 929, 3177, 2642, 2553, 6306, 5442, 7237, 8843, 10112, 11457, 11804, 12074, 13379, 12306]. Southern [79, 103, 131, 474, 838, 1253, 1667, 4472, 4580]. Southport [6390]. Sovereigns [6643]. Sowing [2643]. Space [7716]. Spade [669, 4888]. Spain [895, 4038, 3862, 11830, 12642]. Spanish [187, 8590, 11758]. Spanner [5906, 8901]. Spar [449, 624, 7160, 13156]. Spark [2743, 7765, 9648]. Sparks [3621, 7873, 7925, 11749]. Sparrows [10374]. Spawn [6776]. Spawn-Carrying [6776]. Spayth [7435]. Speaking [10110, 11870]. Spear [5494]. Special [3556, 6376, 9692, 13364, 13386, 13408, 5305, 49]. Specialistic [74]. Specialty [6696]. Species [8779]. Specific [1850, 2696, 4344, 4605, 4903, 6114, 2385, 2364]. Specification [2787]. Specifications [300, 11465]. Specimens [12948]. Spectabilis [5907]. Spectacles [2098, 12826]. Spectra [7765, 6749, 8923]. Spectroscope [1713, 1920, 3773, 3981, 4298, 4489, 8733, 8419, 12026, 9836]. Spectroscopic [5348, 6914]. Spectroscopy [13148]. Spectrum [4698, 5091, 8391, 10405]. Speculation [2641]. Speculations [320, 6577]. Speculative [11904]. Speech [1779, 8961, 10746, 10797]. Speed [1322, 4262, 5740, 7515, 8596, 8907, 10041, 10724, 11087, 11839, 13505]. Speeds [9280]. Speedy [9754]. Spelman [10354]. Spelter [12287]. Sphere [6698]. Spheres [11597]. Spherical [1073]. Spider [6359, 11890, 12235, 12775, 13303, 897, 11351]. Spiders [1690, 8693]. Spike [9805, 3788, 6526]. Spinal [10717]. Spindle [4500, 4483, 13568]. Spinnerets [6359]. Spinning [7911]. Spiral [2769, 10985]. Spirit [2698, 7431, 11498, 13176]. Spirits [1773, 2067, 5174]. Spiritual [9203]. Spiritualism [1126, 1870, 1227, 2824, 6519, 6694, 6742]. Spiritualist [9272]. Spirophore [9473]. Splendid [120, 7717]. Splices [1479]. Split [8927].

Splitter [3942]. Spoke [8571, 9678, 11481]. Spoke-Making [9678]. Spoke-Setting [8571]. Spoked [240, 1512]. Sponge [12594]. Sponges [7212, 6892, 12806, 12658]. Spontaneous [793, 3189, 4744, 6769, 9141, 8618, 10642, 11275]. Spool [1874, 3294]. Spooler [7082]. Spooling [11805]. Spoon [5744, 13151]. Spots [4825, 10597, 10822, 12358]. Spotted [7617]. Spout [9344]. Spouts [2056]. Sprague [4933]. Spread [2552]. Spreading [1071, 12372]. Spring [528, 531, 275, 2242, 1511, 3034, 2769, 2434, 2782, 3265, 4993, 3864, 6335, 7269, 7113, 7151, 11511, 11146]. Spring-Lock [528]. Springs [7008, 10531, 10152, 9889, 13069, 768, 2619, 6194]. Sprinkler [8159, 12229]. Spurts [10483]. Square [2547, 7759, 7563, 10445, 12129]. Square-Inch [7563]. Squaring [12439]. Squibbel [595]. St [2311, 2335, 9619, 5073]. St. [1143, 607, 836, 1563, 1799, 1771, 2393, 3960, 6429, 6007, 7410, 8247, 11905, 11474, 11696, 13641]. Stable [4439, 10864]. Stables [9087]. Stack [12444]. Stacking [3650]. Stacking-Apparatus [3650]. Staff [2113]. Stafford [11330]. Stained [204]. Stairways [1511]. Stalk [8600, 10666]. Stall [438]. Stalls [13023]. Stammering [814]. Stamp [3181, 4245, 4405, 7266, 10382, 10686]. Stamping [10116, 12292]. Stamps [3415, 3445, 13692]. Stancliff [2382]. Stand [1138, 2196, 1209, 2795, 3474, 8573, 10967, 11955, 12999, 12829]. Standard [1197, 5032, 5955, 6804, 10773, 10805, 12718]. Standing [5654]. Stanisłas [2041]. Stanley [7982]. Star [255, 3339, 4695, 10683, 10007, 11322]. Starch [5873]. Starling [6220]. Starr [5308]. Stars [2888, 4346, 9826, 10880]. Starting [2916]. State [559, 2022, 2016, 2114, 1941, 3608, 2587, 2759, 4490, 4859, 3974, 5035, 4225, 5435, 6738, 8541, 8246, 9693, 12314, 12383, 961, 3044, 5293]. Staterooms [482]. States [79, 103, 131, 405, 474, 432, 2168, 1688, 2996, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 5169, 4936, 4967, 5042, 4607, 4642, 4801, 5410, 7404, 8274, 8602, 10878, 4992, 5016, 538, 179, 1598, 1253, 1667, 2693, 2850, 3036, 6195, 6229, 6250, 6280, 6307, 5923, 6968, 6630, 7071, 8839, 9984, 9797, 12226, 13428, 13332, 13091, 13507]. Static [722, 10195]. Statics [671]. Station [4008]. Stationary [11371]. Stations [6273]. Statistical [3272]. Statistics [1401, 1445, 6082, 1695]. Statue [1798, 1807, 2547, 3902, 9399, 11886, 12056]. Statues [171]. Status [2800]. Stave [2785]. Stay [6156]. Steak [9169, 12143, 13561, 13515]. Stealing [7484]. Steam [17, 938, 1073, 42, 1026, 87, 228, 618, 1117, 197, 92, 418, 121, 324, 1050, 423, 469, 660, 150, 127, 70, 492, 1023, 1092, 644, 445, 47, 215, 1322, 2191, 1890, 1914, 1889, 1394, 1573, 1530, 1544, 1621, 1353, 1368, 2059, 1648, 1746, 1752, 2017, 2043, 1619, 1751, 1854, 1858, 1445, 1696, 1196, 1738, 1315, 3335, 3556, 3238, 2671, 2517, 2486, 2575, 2624, 2525, 2684, 3392, 2898, 3468, 3685, 3054, 3674, 3326, 3127, 3278, 3124, 2908, 2598, 3717, 3567, 2849, 3739, 2917, 2625, 2649, 2993, 2466]. Steam [3683, 2592, 4450, 3851, 4061, 4603, 4705, 3826, 4293, 4987, 4027, 4747, 4438, 3913, 4940, 4214, 4208, 4180, 4437, 5104, 3827, 4716, 4044, 4882, 4409, 4155, 4723, 4468, 4696, 4725, 4169, 4801, 3985, 3947, 4203, 4562, 4225, 4105, 5190, 5090, 5901, 5652, 5949, 6445, 6324, 6327, 6001, 6309, 5347, 5930, 5609, 5961, 5597, 5850, 5959, 5303, 6305, 5363, 5410, 5350, 5670, 5746, 7049, 7407, 7231, 7148, 7694, 7412, 7849, 6824, 6847, 7645, 6604, 6580, 7530, 6769, 7877, 7220, 7291, 7938, 9309, 9120, 8471, 9117, 9205, 8249, 8841]. Steam [9404, 8797, 8570, 7959, 8892, 9094, 9403, 9434, 8137, 9095, 9395, 9239, 8860, 8021, 8042, 8073, 8196, 9562, 10471, 10532, 10064, 9808, 10507, 9876, 9967, 10117, 10670, 9529, 9621, 10029, 10288, 10191, 10626, 10312, 10960, 11221, 11540, 11485, 12124, 11323, 11633, 11370, 12028, 11264, 11045, 11235, 11784, 11171, 11281, 11158, 11401, 11659, 11797, 11968, 12178, 12122, 11353, 12115, 11019, 12145, 11684, 11160, 13661, 12738, 12368, 13127, 13336, 12709, 13227, 13229, 12859, 12914, 13124, 12942, 12639, 12495, 13095, 13665, 13634, 13620, 13610, 2963, 2240, 313, 1736, 1767, 1791]. Steam [1843, 1869, 9733, 1166, 830, 1715, 11262]. Steamboat [2651, 10933, 10857, 10220, 12550]. Steamboilers [2345]. Steamer [318, 1611, 2036, 3273, 2563, 4260, 6160, 5821, 5788, 5988, 7130, 7756, 8248, 9641, 9797, 10577, 11501, 11722]. Steamers [1065, 3190, 3500, 4667, 3970, 6363]. Steaming [87, 3395, 8481]. Steamship [863, 2966, 3193, 6681, 10807, 13013]. Steamships [4398]. Stearin [5490, 8278]. Steel [723, 919, 277, 614, 152, 323, 290, 1463, 1477, 2361, 1538, 1988, 2655, 2589, 5464, 6469, 7388, 6678, 7511, 7668, 7020, 7454, 7690, 6966, 8216, 9270, 8964, 8402, 9217, 10373, 10962, 12003, 12149, 13246, 13356, 12449, 13582, 8774, 5305]. Steeling [3507]. Steep [5849, 7814]. Steering [783, 6168, 9095]. Stellar [2716, 6913]. Steller [7338, 8256]. Stench [11436]. Stencil [12650]. Stenography [7182]. Step [2988, 4500, 8906, 11711, 13701]. Stephenson [173, 8593]. Steppes [8562]. Steps [1511]. Stereometrical [2962]. Stereoscope [1068, 9367]. Sterhydraulic [2495]. Sterilization [11394]. Stern [5723]. Sternberg [803]. Stethoscopic [12011]. Stevens [26, 648, 3712, 2976, 6039, 5880, 4382, 4295, 4323, 4410, 4351]. Stevenson [10961]. Stewart [7077]. Stick [4463, 9968, 11107, 13051]. Still [7993]. Stillè [13221]. Stilt [11482]. Stilwell [4068]. Stock [699, 1743, 8016, 10438, 12067, 11645, 13539, 13530, 12886]. Stocking [3129]. Stockings [12021]. Stockley [6999]. Stocks [10635, 11425]. Stockwell [420, 8333]. Stolp [11016]. Stomach [8692]. Stone [494, 991, 614, 834, 1574, 2141, 2347, 1570, 1592, 2274, 2203, 2411, 1370, 1939, 1711, 1378, 2470, 2442, 3309, 4234, 4077, 4802, 4142, 5358, 6155, 6669, 8457, 9076, 8683, 10570, 10028, 10859, 9887, 13702, 2314, 2341, 1710]. Stone-Dressing [10028, 13702]. Stone-Sawing [9076]. Stone-Working [9887]. Stones [3745, 5978]. Stool [1162, 3057, 2901, 8485, 12947]. Stop [121, 1161, 1540, 7438, 7461, 9150, 10111, 12496]. Stop-Cocks [1161]. Stopcock [9277]. Stopper [2765, 4071, 6943, 11093, 12291, 12058, 13108]. Stopping [9918]. Stops [6005]. Storage [6921, 10225]. Stories [7811, 7018]. Storing [6876, 10254]. Storks [10942]. Storm [144, 878, 1213, 1569, 1276, 1688, 7456, 9462, 10031, 9589]. Storms [3682, 7073]. Story [2202, 7610]. Stove [1168, 900, 2105, 1217, 2685, 4644, 6209, 5582, 5962, 7077, 10612, 9561, 11195, 11506, 12446]. Stovepipe [10990]. Stovepipes [13464]. Stoves [670, 1278, 5242, 11531, 31]. Stow [1239]. Stowmarket [2062]. Stoy [10965]. Straightener [3653, 2961]. Straightening [268, 8512, 9885, 9962, 9778, 9831]. Strain [4233, 8446, 5718, 5664]. Strainer [1616]. Strands [11471]. Strange [689]. Stranger [2202]. Strap [1097, 3355]. Straps [3389]. Strasbourg [944, 6924]. Strata [8241]. Straub [12364]. Straw [3650, 4524, 5008, 13233, 12596, 12910]. Straw-Burning [4524]. Straw-Necked [12596]. Strawberries [11724]. Stray [7531]. Streaks [6242]. Stream [7779, 2715]. Streams [7240]. Street [1136, 1173, 2273, 1931, 3596, 3674, 3137, 3471, 4204, 4983, 4595, 5303, 7264, 7313, 6619, 6938, 6934, 8866, 9277, 8021, 8417, 10682, 9751, 10112, 10820, 11579, 11862]. Streets [1353, 2721, 6963, 10626]. Streit [10776]. Strength [4154, 1673, 1196, 2756, 4396, 6424, 7326, 10959, 11954, 13687, 5750]. Strengthening [2990]. Stress [5106, 5815]. Stretcher [220, 1267, 3943, 9613]. Stretching [12704]. Stricta [7763]. Strike [786, 3138, 3018, 3046, 3064, 3091, 7602, 10799, 12982]. Strikes [10344, 12515]. Striking [2794]. Stringing [9234]. Strings [11498]. Strip [4803, 4933]. Strong [7745, 1095]. Strongest [9620]. Strop [2014]. Structure [3659, 8955]. Structures [417, 3645, 6406, 7732, 10539]. Struggles [1192]. Strychnia [3209]. Stub [4471]. Student [2215, 2336]. Students [5006]. Studied [6145, 10822]. Studies [9694, 12341, 8775]. Study [348, 153, 446, 2348, 1517, 2500, 2528, 4770, 10479, 11836, 13545, 12814, 13149]. Stuffing [4380, 10285, 11709]. Stump [958, 6388]. Stupidity [3654, 2673]. Sturtevant [3721]. Style [6795, 12371]. Stylographic [11981]. Subject [13220]. Subjects [5141, 5711]. Submarine [511, 783, 3897, 4241, 3888, 4238, 6406, 11096, 11222]. Submerged [417, 8190]. Submitted [651]. Subscribers [2399, 2424, 2449, 2479, 2505, 2532, 2555, 5156, 5467, 5492, 5555, 5576, 6496, 5271, 9495, 8301, 9512, 9850, 9877, 11011, 11034, 11064]. Subscription [5682]. Subscriptions [12251]. Subsoil [12180]. Substances [2694, 2886, 2671, 2771, 3338, 3388, 4383, 1589]. Substantial [8051, 8082]. Substantiate [2571]. Substitute [1693, 4635, 4665, 8863, 11687, 12252, 13343, 3495, 12186]. Substitutes [2863]. Substituting [2625, 2649]. Subterranean [2416, 6405, 773]. Subways [5058]. Success [1174, 2202, 7096, 9144, 11005, 12070, 13506]. Successful [4575, 7099, 8244, 11057, 12875, 1166]. Successfully [13273, 13331]. Such [1438]. Sucker [11698]. Suction [6467]. Sudden [7417]. Suez [7, 14]. Sufficient [13067]. Sugar [538, 1100, 698, 513, 554, 5, 1698, 2693, 2850, 3387, 3235, 3892, 5155, 4877, 5633, 6875, 10858, 9670, 12018, 11289, 12580, 13010, 12735, 12259, 10528]. Sugar-Making [5155]. Sugar-Refining [6875]. Sugars [10747]. Suggested [2271]. Suggestion [2890, 3772, 9720, 13441]. Suggestions [1177, 1575, 2375, 2400, 2425, 3689, 2450, 6013, 7903, 8621, 7939, 9634]. Suicide [8130, 10901]. Suitable [2268]. Suits [1148, 12279]. Sulky [110, 4243]. Sulphate [5635]. Sulphite [2650]. Sulpho [8198]. Sulpho-Carbonates [8198]. Sulphozone [3495]. Sulphur [1699, 5573, 6391, 3495]. Sulphureted [12396]. Sulphuric [607]. Sulphurous [811, 3444, 8106]. Sumac [864]. Summary [25, 403, 419, 457, 517, 642, 1602, 1675, 1880, 1904, 1977, 2047, 2155, 2190, 2545, 2603]. Summer [3114, 7033, 8953]. Sun [11798, 10977, 2291, 3423, 3946, 4153, 4040, 4772, 4825, 5463, 6936, 6974, 9211, 10406, 10882, 10906, 9706, 10597, 10822, 11725, 11978, 11736, 13129, 12358, 11705, 11727, 11778, 13154, 6531, 6557, 6611]. Sunday [2905, 3598, 3813]. Sundridge [6011]. Sunken [6255]. Sunlight [10036, 4295]. Sunshine [2195, 2171, 10408]. Superheating [11171]. Superintendent [2015]. Superintendents [157]. Superior [807, 3311]. Supervising [10716]. Supplement [2928, 7845, 7958, 8271, 8382, 8414, 8472, 8536]. Supplementary [323]. Supply [2327, 1670, 3164, 5436, 8017, 9511, 10629, 10006, 10854, 13532, 13400]. Support [2037, 7141, 7905]. Supporter [8777]. Supporting [11142]. Suppressed [5162, 5601]. Supreme [432, 11007, 12684]. Surf [10613, 13365, 2063]. Surface [1010, 6198, 5707, 6422, 6411, 7469, 7282, 7048, 7746, 9509, 10694, 11989, 8719]. Surfaces [508, 7637, 6653, 8270]. Surgery [1264, 5902, 7019, 9111]. Surplus [9959]. Surprise [12611]. Survey [5816, 672, 13173]. Survivors [4906]. Susan [2061]. Suspended [1648, 7580, 1715]. Suspender [2467]. Suspending [1936]. Suspends [816]. Suspension [1047, 1566, 3276, 2983, 7631, 7177, 8476, 9909, 9911, 10961]. Sustained [9334]. Sustainer [10959]. Swage [8593]. Swan [5122]. Swans [7516]. Swartwout [1692]. Sweden [873, 4750, 6008, 11779]. Swedish [48, 8966, 12359]. Sweeper [13588]. Sweepstakes [12559]. Sweetzer [2919]. Swelled [13288]. Swim [6052]. Swimming [7585, 6583, 7222]. Swims [42]. Swindles [7258]. Swindling [4232, 5513]. Swine [13283]. Swing [421]. Swinging [9052]. Swiss [10312, 11971, 11694, 11561]. Switch [4994, 5977, 8516, 13006, 12407]. Switches [5046, 6890]. Swivel [11140]. Sydney [8706, 13401]. Synopsis [696]. Synthesis [10036]. Syphon [11019]. Syracuse [2365]. Syrian [6892]. System [144, 909, 51, 1665, 2170, 2327, 1895, 1686, 1748, 2082, 2344, 2664, 3213, 3019, 4853, 5178, 3980, 4116, 5766, 6719, 7181, 7897, 7215, 7123, 7778, 7656, 6907, 6981, 8617, 9218, 10769, 10098, 10225, 11210, 11059, 11171, 11561, 12956, 13271, 13227, 291]. Systematic [252]. Système [13199].

Table [760, 1745, 1391, 1262, 3755, 3371, 5227, 4498, 5881, 6922, 6797, 9118, 10260, 11165, 12151, 12386, 12647, 13231]. Tableau [2962]. Tables [1356]. Tack [6975]. Tackle [9469]. Tadpoles [13594]. Tag [12676]. Tail [7821].

Tailed [12030, 11404]. Tailor [11911]. Tails [6083]. Tainted [7233]. Take [1290, 3449]. Taking [7301, 12277]. Tale [4066, 7639]. Talent [880]. Talk [6715]. Talking [1241, 10928]. Talks [3068]. Tall [8985]. Talley [11217]. Tallow [4528, 2879]. Tamarind [13390]. Tamping [6059]. Tangye [2368]. Tanite [115, 4244, 6349, 9561, 4577]. Tank [3622, 4056, 3986, 6875, 8510]. Tannate [4469]. Tannin [8592]. Tannin-Testing [8592]. Tanning [980, 4911, 6979, 13659]. Tap [199, 6675, 9312]. Tapayaxin [8999]. Tape [9084]. Taper [5221, 5935]. Tapered [2413]. Tapers [10936]. Tapper [11070]. Tapping [4427, 6277, 6092]. Tar [1700, 1481, 5192, 5011, 5635, 273]. Tarantula [5797]. Target [524]. Tariff [3066]. Tarsier [13446]. Tasimeter [11609]. Taste [11502]. Tasteful [9507]. Tax [49, 3066, 10632]. Taxes [1829]. Taxidermy [6816]. Taxing [5951, 10142]. Tay [1970, 4326]. Taylor [4257, 4349, 9269]. Tea [2795, 4402, 5820, 6601, 7663, 8631, 10627, 10575, 12460, 13040, 13250]. Teach [2527]. Teachers [9014]. Teaching [1250, 11449]. Teachings [9527]. Teas [216]. Teaser [9971]. Teaspoonfuls [7578]. Technical [179, 1489, 3477, 4977, 6308, 8937, 8839, 13355]. Technology [648, 2068, 2465, 2689, 2789, 2884, 9022, 2621, 7100]. Tedder [4937, 8223]. Tedious [2349]. Teeth [1071, 2035, 5468, 5442, 6649, 6965, 9282, 11209, 11196]. Telegraph [898, 958, 1825, 2387, 2699, 3113, 2646, 2746, 2835, 2708, 3918, 5217, 5139, 4602, 4317, 5089, 4149, 4210, 5138, 6555, 5514, 6546, 6083, 5824, 6332, 6907, 8747, 8417, 10746, 12652, 12935, 12651, 12972, 2267]. Telegraphed [1842]. Telegraphic [1433, 2798, 5488]. Telegraphing [6794, 11133]. Telegraphs [1507, 5440, 11639, 5671]. Telegraphy [43, 342, 1723, 3787, 4659, 3828, 6135, 6330, 6755, 7659, 8968, 10144, 1190, 1028, 6610, 6631]. Telephone [10625, 10378, 9969, 10110, 9849, 11551, 11207, 11621, 11893, 11009, 11774, 11314, 12769, 13366, 13584, 12936, 13045, 13694, 12742, 12605, 12629, 13094, 11534, 11916, 12133, 8217]. Telephones [12005]. Telephonic [11674, 11365]. Telescope [3535, 3773, 5159, 5084, 4455, 3925, 5511, 8126]. Telescopes [3800]. Tell [4066, 7639, 12459, 5664]. Tell-Tale [7639]. Tellier [1031]. Tellurium [10602]. Telo [398]. Telo-Dynamic [398]. Telodynamic [803]. Temperance [1467]. Temperature [161, 5063, 3884, 6379, 6576, 10596, 10717, 13400]. Temperatures [1995, 3627, 2696, 2908, 5871, 6855]. Tempered [9557]. Tempering [277]. Temple [9732]. Temporary [10007]. Ten [2994, 8042, 8073]. Tenacity [10798]. Tendency [9608]. Tenderer [7031]. Tenements [12604]. Tenoning [4314]. Tensile [1673]. Tension [2014, 11488]. Tent [5044]. Tenth [3283]. Teredoes [12485]. Terminus [635]. Terms [8826]. Terneu [12393]. Terneu-Zen [12393]. Terra [2760, 10720]. Terrestrial [3624, 3267, 3337, 7382, 12945]. Terrible [2197, 2062, 3754, 5776]. Tertiary [3558]. Tessie [1970]. Test [401, 197, 155, 2058, 3677, 3881, 6966, 7563, 9273, 10467, 11980, 11555, 2714, 2740, 9892]. Testament [4177]. Tester [9243, 9176, 10450]. Testers [11089]. Testimonial [8708]. Testimony [2012, 11561]. Testing [146, 1196, 2706, 2756, 3163, 3911, 4808, 5735, 7904, 6645, 7315, 9468, 8283, 8592, 8684, 8395, 8274, 10662, 9679, 9888, 10094, 11190, 11954, 11977, 12615, 12852, 5305, 2366]. Tests [224, 282, 1769, 2720, 2690, 5065, 4663, 10115, 11061, 13579]. Texas [929, 4517, 12623, 4855]. Textile [1947, 3605, 2898, 4039, 9044, 7942, 8004, 8116, 8172, 8525, 8551, 8576, 8667, 8752, 8783, 8974, 9034, 9067, 9354, 9416, 9446, 9486, 9787, 9839, 9927, 9949, 10051, 10073, 10182, 10267, 10294, 10488, 10552, 10702]. Thallium [1516]. Thames [2384, 6197, 11515, 11569]. Thatch [1331]. Theater [2642, 9526]. Theaters [9498, 9579]. Their [1538, 2462, 2488, 3359, 4514, 4451, 6294, 5268, 5925, 7781, 8908, 8487, 9113, 8099, 8826, 9366, 8923, 9080, 9014, 9589, 12440, 5575, 56, 1896, 3081, 2771, 3017, 3139, 2822, 3834, 4768, 4848, 6219, 6039, 6192, 5298, 6326, 8861, 11894, 806, 785, 875, 1591, 5718]. Thelemark [8356]. Them [1646, 3667, 3731, 3795, 4586, 4618, 4648, 4708, 4760, 5256, 5283, 5312, 5340, 9357, 9451, 12121, 12036, 12477, 176, 1290, 3696, 3763]. Theology [3307, 3569]. Theories [4825, 7493, 8654, 9854, 3648]. Theory [808, 4097, 3836, 6269, 6050, 5818, 5387, 6982, 9045, 8647, 8043, 8074, 9505, 7996, 8924, 9212, 8013, 10825, 10405, 10195, 11006, 13038, 5142, 5177, 5201, 11690, 2493]. Therapeutic [12819]. There [1355, 4003, 10877]. Therefrom [9990]. Thermo [4579, 5956, 9078]. Thermo-Diffusion [9078]. Thermo-Electric [4579, 5956]. Thermometer [317, 4433, 6249, 7213, 10569, 11780, 11379, 13311]. Thermometers [749]. Thermometry [10823]. Thermoscope [7335]. Thermostatic [9307]. Thick [4486, 13051]. Thick-Thighed [13051]. Thiers [10560]. Thieves [3271]. Thighed [13051]. Thill [2262, 2441, 11122]. Thimble [1714]. Things [66, 5107]. Think [11364]. Thinker [4029]. Thinking [252, 2654]. Third [3865]. Thirst [11567]. Thirty [583, 12850]. Thirty-Ninth [583]. Thoma [3960]. Thomas [939, 959, 11667, 1465, 1185]. Thompson [11146]. Thomson [10567]. Thorneycroft [10005]. Thoroughness [10609]. Those [1355]. Thought [715, 413, 439]. Thoughts [8212, 9271]. Thousand [1613, 1622, 3675, 4995, 4513, 5328, 6162, 5378, 6381, 6376, 7328, 8042, 8073, 9938, 13530]. Thrasher [12910]. Thrashers [3179]. Thrashing [4603, 6562, 5385, 8109, 13158]. Thread [1741, 11641, 11972]. Threaded [1899]. Threader [1714, 6277, 5548]. Threading [855, 5470, 9020, 8160, 12967]. Threads [4549, 6791, 10014, 10439]. Three [221, 3896, 5041, 6005, 6381, 6433, 7239, 6848, 11081, 11404, 12340, 13344, 12875, 1054]. Three-Cylinder [6433, 7239]. Three-Tailed [11404]. Three-Wheeled [6848]. Threshold [4235, 5578]. Throat [6170]. Thumb [5138]. Thunderbolt [12595]. Thunderer [8897, 12827, 9016, 13252]. Thunderstorms [3312]. Thurston [3093, 3116, 3165, 3197, 3224, 3249, 3285, 3315, 3343, 3403, 3419, 3090, 3428, 3743, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 4566, 4936, 4967, 5042, 4607, 4642, 5489, 10094]. Tickets [909]. Tidal [3772]. Tide [7583, 8904, 11046, 13009]. Tides [4200, 3916, 6055, 6996]. Tidings [4878]. Tie [1044, 2435, 4294, 6656, 6803, 7587, 7411, 9314, 11347, 13180]. Ties [2811]. Tiger [4640, 8843]. Tightener [1483, 1691, 7464, 9242]. Tightening [4910]. Tile [10827]. Tile-Making [10827]. Tilghman [1589]. Tiller [12639]. Timber [4154, 570, 248, 1148, 55, 2362, 4015, 8130, 9863, 12719, 12835, 13041]. Time [352, 1077, 2247, 1197, 3131, 4516, 5769, 7724, 8588, 7928, 9371, 9936, 9594, 12168, 12207, 12053, 13306, 12879]. Time-Indicating [12207]. Timely [1177, 2375, 2400, 2425, 2952, 2450, 7903, 8896, 7939]. Times [5157, 4368, 5083, 7657, 12734, 2112]. Timpe [2122]. Tin [770, 579, 123, 2973, 4031, 10255, 11848, 11810, 266]. Tinkering [49]. Tinned [12667]. Tip [488, 2945]. Tipped [6820, 11329]. Tipping [13508]. Tirade [2825]. Tire [7613]. Tire-Upsetting [7613]. Tires [1576, 1491, 2367]. Tissues [6973]. Titicaca [5512]. Toad [10349]. Tobacco [501, 465, 2783, 4523, 9306, 9433, 9234, 9536, 11740]. Todd [5772, 5829]. Toddy [11243]. Tody [12080]. Toilet [2869]. Toller [11172]. Tolman [8745]. Tom [5138]. Tomes [2079]. Tomlinson [6950]. Tompkins [6628]. Ton [7601, 9779, 10960, 11167]. Tones [4059, 11074]. Tongs [1535, 4237, 8456, 10043, 10287, 11435, 13675]. Tongue [6850, 11196]. Too [491, 2343, 11483]. Tool [420, 1482, 1292, 1268, 1270, 3718, 2844, 2987, 3526, 4672, 4529, 5116, 5552, 5536, 5275, 5524, 6304, 7078, 8802, 9345, 9470, 9085, 9090, 10861, 11425, 10986, 11266, 12989, 12525, 13702]. Tool-Grinding [9345]. Tools [1821, 1770, 5905, 6561, 7459, 8167, 8160, 9247, 10147, 9618, 11418, 11423, 13567, 10964]. Tooth [2761, 4215, 7074, 12249, 12293]. Toothed [5045]. Top [5554]. Topeka [688]. Topical [56]. Torch [5114]. Tornado [11864]. Tornadoes [1667, 2056]. Tornelia [13159]. Torpedo [4, 3744, 4771, 4593, 6527, 6132, 5390, 7656, 7160, 10196, 10339, 10224, 10566, 9934, 10289, 9797, 10083, 11068, 11111, 11529, 12114, 11673, 12406, 12674, 12502, 13459, 13156]. Torpedoes [818, 2131, 6671, 10084]. Torrey [12009]. Torsional [4233]. Tortoise [12264, 13437, 10540]. Tortoises [7647]. Toselli [5390]. Total [11736]. Toughened [7214]. Toughening [1404]. Tow [13605]. Tow-Boat [13605]. Towage [1696, 2466, 3418, 6429]. Tower [1674, 2543, 4007, 5869, 5328]. Towers [360]. Towing [3190, 2891, 4718, 13271, 13620]. Town [12434]. Toxicological [12727]. Toy [1294, 1738, 5229, 7389, 8482, 9588, 11686]. Trace [10608, 11838, 12464]. Tracelock [270]. Traces [3804]. Tracing [13650]. Tracings [12181]. Track [6271, 6210, 7288, 9957]. Tracks [8098]. Tractile [9699]. Traction [1491, 2684, 3683, 4058, 4562, 3949, 6934, 12124, 13645]. Trade [748, 1650, 2145, 2151, 3085, 2836, 2615, 4000, 4226, 5035, 5709, 5435, 7359, 7097, 6869, 6935, 8216, 9492, 9218, 9695, 11751, 12258, 12200, 11497, 12153, 12100, 12042, 12196, 12147, 13247, 13378, 13238, 13263, 13369, 13392, 13440, 13596, 13649, 13631, 13692, 12522, 13604, 13491, 983]. Trademark [2115]. Trademarks [2723, 12360]. Trades [2129, 2114, 2825]. Traffic [13541]. Tragopan [11173]. Train [8860]. Training [876]. Trains [3598, 3813, 10311]. Trajectory [12930]. Tramway [484, 5300, 10312, 12176, 13609, 13310].

Tramways [6498]. Trance [10193]. Transatlantic [5788]. Transferring [4857]. Transfusion [6136, 9497, 11687]. Transit [507, 536, 555, 1514, 3067, 4058, 3815, 4035, 5975, 6401, 7483, 6964, 7125, 6911, 7327, 8284, 9376, 8321, 12048, 11499, 13387, 13341, 5222, 4037, 6088, 6547, 11543]. Transits [4290, 4320]. Transmission [1461, 3472, 3583, 3647, 3687, 3719, 9936, 9822]. Transmitted [10746]. Transmitter [13486]. Transmitting [2564, 6060, 11662, 13316]. Transmutation [12382]. Transpiration [984]. Transplantation [7099]. Transplanter [6528]. Transplanting [5982]. Transportation [1055]. Transporting [2739, 5826, 7280]. Trap [938, 3335, 3326, 5093, 4960, 4300, 4266, 5901, 7174, 6822, 6824, 9057, 8330, 8164, 8137, 9921, 9780, 11436, 11888, 12235, 13073]. Trapless [10890]. Traps [1536]. Traube [7485]. Travel [917, 4595, 8323]. Traveling [2834, 5090, 5743, 10719, 10957, 11134]. Treadle [1639, 3651, 7336, 13666]. Treasure [1724, 4860, 9525]. Treasures [3834]. Treat [5845]. Treated [13273]. Treatment [2797, 4721, 6000, 6170, 6268, 11163, 11762, 12148, 11208, 13431]. Treaty [12147]. Tree [4467, 4909, 4356, 5421, 5654, 5356, 7207, 6872, 6703, 9378, 10194, 11405, 11147, 11939, 11287, 12835, 13260, 12503, 46, 2879]. Trees [1089, 5982, 5845]. Trellis [9507]. Trench [5471]. Trevithick [714]. Trial [1898, 2287, 1722, 6160, 9169, 9851, 9779, 10614, 9723, 12094, 2498]. Trials [127, 4981, 5193, 4297, 6327, 8273, 11931]. Triangular [1566]. Tribe [12022]. Trichinze [11616]. Trick [6644]. Tricks [2544, 2494]. Tricycle [10473, 10865, 11047, 11301, 11458]. Tried [2645]. Triggers [9343]. Trigonometer [12075]. Trimmer [8362, 12206]. Trimmers [8959]. Trimming [6278, 8769]. Trinity [7885]. Trip [929, 415, 6160, 9723]. Tripler [1044]. Triscott [2665]. Triumph [4717]. Triumphs [1142, 1191, 1264, 12959]. Trocadero [11565]. Trolling [9724]. Trophy [11480]. Tropical [13122]. Tropics [30]. Trotting [5273]. Trouble [2550, 12276, 12557]. Troubles [6772]. Trough [12706]. Trout [9208]. Trouve [11009, 10828, 11621]. Trove [9525]. Troy [2685, 6384, 1306, 7807]. Truck [9142]. Trucks [11344]. True [808, 2202, 10905, 11972, 12053]. Truffle [9049]. Truncated [10889]. Trunk [1486, 13288]. Trunnion [12206]. Truss [2983]. Trustworthiness [7256]. Trustworthy [2720]. Truth [4096]. Trying [6560]. Trying-Up [6560]. Tub [5449]. Tube [560, 4465, 8837, 8402, 10204, 10909, 9640, 3644]. Tubes [4301, 7341, 9918, 12373]. Tubing [13621]. Tubular [15, 7853, 6652, 9607, 13281]. Tucking [4676]. Tufted [11938]. Tug [13097]. Tulip [6703]. Tun [5616, 7388]. Tungsten [312]. Tuning [13514]. Tunis [11948]. Tunnel [202, 244, 1139, 1272, 1238, 2126, 1223, 2541, 3269, 3895, 5434, 6503, 6946, 7552, 6912, 10820, 12283, 12855, 12846, 1011, 175, 1492, 2393, 3774, 5168, 5376, 5572, 6896, 10059, 13427]. Tunneling [2285, 2409]. Tunnelling [2721]. Tunnels [159, 6377, 6171, 7355, 11092, 4151]. Tunnicliffe [4639]. Tunny [10322]. Tuns [3309]. Turbine [245, 294, 155, 1640, 3328, 2720, 3156, 4832, 5119, 5685, 5583, 5692, 8629, 13579, 12592, 9209]. Turbines [428]. Turkey [11802, 11881]. Turkish [422, 4641, 4860, 10794, 10146, 11174, 12948]. Turks [1355]. Turn [760, 11901]. Turn-Table [760]. Turned [1917]. Turner [10384, 5379]. Turning [1764, 6240, 5877, 10802, 11884, 11254, 8774]. Turpentine [1773, 7078]. Turret [6197]. Turrets [307, 3163]. Turtle [12277, 12624, 12313]. Turtles [5442]. Turton [3864]. Tussock [617]. Tussock-Moths [617]. Tuttle [10925]. Tweer [7368]. Tweers [4100]. Twelve [7884]. Twenty [2419, 3703]. Twenty-Sixth [2419]. Twin [8248, 11722]. Twine [2496]. Twins [5486, 5377, 5443, 11696]. Twist [9345]. Twitchell [2903]. Two [318, 593, 3135, 5009, 6516, 6829, 7610, 7150, 8935, 8520, 8738, 9077, 10086, 11037, 11188, 10781]. Two-Feeder [593]. Two-Horse [11037]. Two-Lined [10781]. Two-Story [7610]. Tying [1007, 13676]. Tying-in [13676]. Tyler [13444]. Tyndall [5164, 1067, 256, 1902, 1844, 1932, 2707, 3506, 3529, 4006, 4034, 3889, 3924, 3802, 3837, 3811, 3954, 3865, 5463, 6184, 6212, 6518, 9333, 10768]. Type [1326, 3176, 12465]. Type-Setting [1326]. Typhoid [6518, 12589]. Typical [6516]. Typo [3902]. Typo-Statue [3902]. Tyres [7149].

U. [627]. U.S. [1276, 10658]. Uchatius [10962]. Ultra [5794, 9301]. Ultra-Darwinism [9301]. Ultramarine [700]. Umbrella [2819, 9051, 8777, 12106]. Unauthorized [5517]. Uncommon [1614]. Undergo [8131]. Underground [2128, 2344, 3540, 4259, 4287, 4598, 5655, 6408, 6430, 6450, 6500, 6553, 6915, 6579, 6677, 6720, 12935]. Undescribed [3589]. Undulata [6530]. Undulatory [8654]. Unfit [7782]. Unhealthy [2983]. Unicorn [390]. Uniform [4116]. Uninflammable [9526]. Uninvited [11594]. Union [453, 2825, 3980, 6907, 1592]. Unions [2129, 4000]. Unit [3039, 2205]. United [4801, 10878, 538, 179, 79, 103, 131, 405, 474, 432, 1598, 2168, 1688, 2693, 2850, 3036, 2996, 4574, 4674, 4701, 4727, 4752, 4778, 4800, 4831, 4856, 4886, 4912, 5169, 4936, 4967, 5042, 4607, 4642, 6195, 6229, 6250, 6280, 6307, 5923, 5410, 6968, 6630, 7071, 7404, 8839, 8274, 8602, 9984, 9797, 12226, 13428, 13332, 13091, 13507, 4992, 5016]. Units [2340]. Universal [442, 246, 740, 2040, 2186, 3302, 5198, 5842, 6385, 6713, 7510, 7910, 7582, 8329, 8732, 10445, 13045, 6659, 6683]. Universally [1938]. Universe [8922, 11828]. University [7100, 2336, 3357, 3925, 8989]. Unknown [12981]. Unlimited [5159]. Unloading [6079]. Unrecognized [4692]. Unruly [6477]. Unscientific [1932, 11058]. Unseen [11828]. Unsinkable [11606, 11684]. Unveiling [1798]. Upheaval [7176]. Upon [1463, 2822, 990, 812, 857, 788, 452, 622, 43, 2245, 2626, 2890, 4548, 6192, 5750]. Upright [1594, 3617, 6628, 9672]. Upsetting [7613]. Urchin [1029]. Urea [6331]. Urgently [2228]. Use [98, 352, 670, 463, 450, 877, 927, 50, 297, 818, 1454, 2009, 1517, 2272, 3444, 3584, 4322, 4662, 3919, 6375, 7261, 7430, 7550, 7368, 7361, 7953, 9528, 12667, 12581, 13381, 13426, 13582, 11731, 1045, 1090, 764, 534, 3892, 3887, 1066]. Used [9178, 1592, 1569]. Useful [2462, 2488, 3772, 5746, 7764, 7827, 7674, 8331, 9618]. Useless [7656]. Uses [449, 624, 1254, 1333, 1302, 1516, 1380, 1198, 2323, 1327, 1599, 3721, 2973, 3857, 4407, 5170, 5487, 5817, 5924, 7403, 8128, 9171, 8420, 8419, 11894, 12312, 13402, 241, 2879, 811, 771, 11703, 188]. Using [1634]. Utero [9609]. Utility [1913, 12983]. Utilization [1974, 2866, 5384, 5321, 8867, 10754, 10138, 9937, 10255, 10977, 12195, 11923, 11392]. Utilizations [8386, 10086, 11065]. Utilize [12361, 13636]. Utilized [324, 66]. Utilizing [1702, 1315, 10254].

V [2212, 1494]. Vaccination [1801, 13531]. Vacuum [1969, 3124, 4515, 6750, 6573, 12420]. Vagaries [1793]. Vagrancy [7094]. Validity [2248, 8447, 8853, 8885, 8914, 8944, 8976, 9006, 9037, 9069, 9104, 9130, 9163, 9194, 9227, 9262, 9294, 9327, 9358, 9387, 9419]. Valley [295, 344]. Valuable [1965, 2790, 7871, 9428, 10687, 11450]. Valuables [2239, 7745]. Valuble [4999]. Value [1037, 489, 1857, 1328, 2173, 3372, 2890, 3667, 3696, 3731, 3763, 3795, 4586, 4618, 4648, 4708, 4760, 3934, 4682, 5105, 5568, 5297, 5256, 5283, 5312, 5340, 6992, 9429, 9357, 9451, 9193, 10598, 12145, 11903, 12036, 12477, 8425, 5245, 2366]. Valve [384, 421, 1426, 1619, 1937, 3986, 4466, 4155, 3864, 6166, 7781, 7075, 7027, 7337, 7539, 7563, 7461, 8630, 9150, 10574, 10115, 10111, 11611, 11454, 11371, 11685, 11867, 11784]. Valved [3893]. Valveless [1892]. Valves [121, 1161, 803, 1214, 3921, 4067, 9113, 9239, 9307, 9970, 9700, 12260, 13100, 13170, 10259]. Valvoline [12204]. Vanadium [3881]. Vanda [10477]. Vander [3398]. Vanderbilt [9585]. Vanier [11168]. Vanne [8303]. Vapart [10350]. Vapor [1278, 3393, 3476, 2773, 2565, 3009, 2649, 3222, 4297, 5497]. Vaporizing [11274]. Vapors [2192, 2625, 2649, 2440]. Variable [1262, 2888, 4923, 6842, 11702, 11402]. Variance [2592]. Variations [5063]. Variety [10202]. Various [1995, 1300, 3114, 5938, 8779]. Variously [3247]. Varnish [720, 1768]. Varnishes [550]. Vase [8511, 12656, 12537]. Vases [7913, 12423, 13391]. Vassar [678, 12020]. Vault [12199]. Vaults [2239, 8185]. Vegetable [483, 3247, 5043, 4694, 7191, 11598, 12410, 6659, 6683]. Vegetation [859, 5963]. Vehicle [5300, 9889, 10443, 11481, 11408, 12762]. Vehicles [903, 6340, 6364, 11146]. Veil [4159]. Velocipede [1765, 9342, 11788, 13585]. Velocipedes [9587]. Velocities [3354, 3371]. Velocity [2812, 5213, 5404]. Veneered [7719]. Veneers [474]. Venerable [10853]. Venom [6192]. Vent [8, 8308]. Ventilating [2263, 1496, 2814, 4644, 5420, 5412, 7644, 7778]. Ventilation [1941, 3162, 2652, 2548, 5169, 4942, 6171, 7355, 8348, 8895, 10692, 12198]. Ventilator [1488, 2414, 2459, 4702, 11318]. Venus [5222, 4037, 6088, 6547]. Verdict [2131, 2015, 9852]. Vermicelli [1171]. Vermont [3481, 12793]. Versus [605, 2490, 2748, 6080, 7606, 9692, 2240, 5684]. Vertical [223, 393, 1289, 3011, 4465, 6058, 6163, 5688, 7931, 10568, 10594, 11535, 11585, 12613, 13383]. Very [2696, 9050]. Vesicular [3222]. Vessel [2263, 3723, 4484, 5331, 9620, 12114, 12406, 13156]. Vessels [747, 2013, 4718, 6255, 7825, 8282, 10204, 10165, 9669, 11270, 11960, 11684, 12198, 12713].

Vesuvius [4176]. VI [6063]. Viaduct [265, 5058]. Vibert [8455]. Vibration [8863, 12847]. Vibrations [6145, 8604]. Victims [1148]. Victor [11514, 5498, 9115, 12592]. Victoria [6338, 5499, 11378]. Victorious [13053]. Victory [4812, 11607]. Vienna [2420, 2297, 3268, 3724, 3746, 3626, 3675, 3740, 3499, 3743, 3594, 3652, 4798, 4425, 4837, 4318, 4793, 4797, 4821, 3860, 5021, 4545, 4573, 4601, 4343, 4794, 4834, 5090, 4860, 4399, 4001, 4150, 4456, 4893, 4874, 4863, 4846, 4978, 3812, 3915, 4690, 5294, 5247, 8385, 12321, 6659, 6683]. View [1344, 7508]. Views [7150, 11834]. Villa [7565]. Ville [5189]. Villette [786]. Vincent [1687]. Vine [7219]. Vinegar [3313, 4673, 7904]. Vines [9151, 13275]. Virgin [9369]. Virginia [1663, 5957, 7092, 12184]. Vis [278, 249]. Vise [545, 27, 1938, 3525, 2434, 3912, 5773, 7417, 7537, 8024, 9558, 12227, 11535, 12262, 12319]. Vision [10192, 413, 439, 9836]. Visit [792, 290]. Visitors [11504]. Visits [8012]. Vital [78, 5375]. Vitality [751, 5375]. Vitriol [4742]. Vitt [2125, 11172]. Viva [249]. Vivisection [6213]. Vogel [509]. Voice [10479]. Voices [6748]. Volcanic [7812]. Volcano [7482, 12250, 13319]. Volcanoes [11035]. Volpicelli [10195]. Voltaic [12405]. Volume [602, 1172, 1804, 5870, 7208, 8677, 10929, 6855]. Volumes [9511]. Vortex [8649]. Voting [7991]. Voyage [4823, 4844, 7150, 10915]. Vs [9269, 12099, 11726, 11513, 919, 774, 789, 1101, 445, 159, 3569, 9496, 9333, 8246, 9111, 11991, 11859, 12816]. Vulcan [9302, 9337, 11857]. Vulcanized [1867]. Vulcanizing [11723]. Vulture [8646].

W [674, 626, 2212, 3498, 3543]. W. [1989, 1400]. Wages [1032, 225, 8298, 11990]. Wagner [2068]. Wagon [3562, 2622, 4299, 6093, 6226, 7727, 8771, 8332, 11452, 11556, 12119, 11240, 12495]. Wagons [1692, 10315]. Waiting [11901]. Wakefield [1291]. Wales [2974, 5164, 9801, 12357]. Walk [1209, 7101]. Walker [10663]. Walking [13051]. Wall [3541, 5361, 8362, 10657]. Wallace [12392]. Wallingford [11864]. Walls [147, 2037, 4953]. Walsh [2717]. Want [2909, 7093, 12540, 837]. Wanted [6, 125, 1353, 1746, 2796, 3478, 3599, 7154, 6938, 10958, 10524, 10339, 10083, 12018, 11255, 12146, 11294, 11229, 13277, 12523, 13636, 13200, 12308]. Wanted-Torpedo [10083]. Wants [10844, 10869, 10894, 10920, 10946, 10972, 10998, 11021, 11050, 11099, 11125, 11151]. War [1100, 788, 817, 807, 1289, 7758, 7748, 8479, 9981, 9620, 486, 882]. Wardian [7394, 6981]. Wardwell [5330]. Ware [9705, 12667]. Warehouses [7525, 7848]. Warfare [766, 927, 3744, 4593, 7160, 9934, 10289, 11446, 11111]. Warmer [10751]. Warming [1496]. Warner [2167]. Warning [1199]. Warrants [2015]. Was [5602, 5824, 9302, 2288, 2202, 8793]. Wash [3785, 4131, 3806, 5422, 11454, 11371, 68]. Washbasin [13100]. Washboiler [10410]. Washer [442, 2040, 4101, 4725, 5523, 9277, 10926, 12586]. Washing [1003, 1432, 1457, 1962, 1240, 3446, 4122, 4638, 4184, 7309, 7052, 6582, 9472, 10938, 12685]. Washings [1826]. Washington [11981, 493, 1830, 3646, 12103, 11164]. Washtub [8573]. Wasp [598, 11997]. Wasps [8219]. Wastage [12197]. Waste [280, 66, 226, 1968, 1649, 1315, 1226, 3441, 5635, 6821, 8130, 11287, 11392, 12558]. Watch [8994, 10503, 10390, 10415, 10451, 10093, 13072]. Watches [13364, 13386, 13408]. Watchmaker [8304]. Watchmaking [854]. Watchman [3055, 7724]. Water [731, 770, 938, 384, 428, 148, 673, 1117, 364, 69, 121, 1012, 373, 161, 417, 641, 199, 155, 124, 224, 282, 1849, 1520, 2312, 2327, 1640, 1368, 1211, 2059, 1648, 1246, 2019, 1377, 1642, 1553, 1334, 1239, 2056, 1670, 1316, 2193, 3061, 3328, 2858, 2438, 2542, 2565, 3009, 2690, 3240, 3157, 3772, 3584, 2866, 3222, 2717, 3971, 5231, 4061, 4889, 4143, 4934, 4631, 3840, 4214, 4790, 4832, 4240, 4729, 4062, 4173, 4427, 6116, 5776, 5685, 6375, 6267, 5876, 6196, 5525, 5936, 6060, 5843, 6092, 7562, 7174, 7360]. Water [7363, 7024, 7761, 7316, 7325, 7427, 7054, 7760, 6898, 6745, 7513, 7240, 7315, 7810, 7534, 8287, 8388, 9407, 8429, 8180, 8995, 8599, 8187, 9276, 8017, 8795, 8619, 7988, 8810, 8196, 10629, 9986, 9565, 10006, 10033, 10035, 10541, 10854, 9587, 10143, 11056, 12044, 12001, 12264, 11976, 11495, 12245, 11369, 11044, 13254, 13429, 12732, 13068, 12335, 12471, 13532, 13663, 12802, 13099, 13098, 12750, 13009, 13400, 12581, 13402, 6855, 2616, 313, 11397, 7782, 9209, 10417, 1715]. Water-Proof [1246]. Waterbury [2193]. Watercloset [10890]. Watering [5014, 10886, 12706]. Waterlow [8706]. Waterproof [5388]. Waterr [7030]. Waters [2489, 8243, 8328, 11502]. Waterspouts [11510]. Watson [5546, 12126]. Watts [1262]. Wave [3710, 3221]. Waves [7104, 7105]. Way [199, 2739, 5469, 6743, 6936, 10769, 10544, 11284, 13539, 13299, 13489, 5826]. Ways [508]. Weak [6274, 7873, 7925]. Weakness [10539]. Wealth [4185, 7127]. Weapons [766]. Weather [990, 1688, 3396, 4803, 4933, 4235, 5894, 7151, 8154, 8828, 12581]. Weatherhead [986]. Weatherstrip [7588]. Weaving [3305, 3605, 3582]. Web [5240]. Webb [7585]. Webs [6359]. Webster [6142]. Weeds [11923]. Week [8827, 2928]. Weeping [5219]. Weighed [8394]. Weigher [2487]. Weighing [721, 281, 3309, 5721, 9244, 11017, 13412]. Weight [847, 2171, 3047, 3063, 6645, 10434, 13687, 6855, 2385]. Weights [2870, 8925]. Welcome [2351]. Well [3481, 4633, 4301, 5423, 6609, 7780, 8837, 8735, 8875, 13606]. Well-Boring [5423]. Wellman [6753]. Wells [1145, 1275, 2889, 2542, 4769, 5604, 7311, 47, 5736]. Wendell [2418]. Wenham [4459]. Wenzel [11887]. Werdermann [12232, 13487]. were [1355]. West [3442, 5516, 11593, 11256, 425, 3297, 11861, 9964]. Western [2469, 929, 2061, 6094, 7633, 6907, 11660, 11993, 13120, 12678, 13605]. Westfield [1972, 2012, 2015]. Weston [8996, 8363]. Wetli [9723]. Weyde [3398]. Whale [13317, 12536]. Whalebone [11121]. Whales [11121]. Whaling [11661]. Wharfdale [593]. Wharton [2498]. Wheat [753, 1611, 5821, 12700]. Wheel [245, 294, 115, 224, 282, 1512, 2363, 2123, 1640, 3107, 3328, 3562, 2690, 2991, 3363, 2717, 5121, 3948, 4316, 3840, 3863, 4832, 5119, 4170, 6116, 5693, 6267, 5583, 5876, 6196, 5935, 5797, 5692, 7255, 7030, 8107, 8190, 8927, 8629, 8599, 9876, 10443, 10631, 10150, 11481, 11672, 12119, 12829, 13099, 13689, 13383]. Wheelbarrows [12154]. Wheeled [6340, 6364, 6848]. Wheeler [2665]. Wheels [428, 240, 155, 1211, 2720, 2490, 2990, 5685, 6680, 7140, 6581, 6624, 8596, 9254, 10343, 11390, 9209]. Wheelwright [7910]. Where [123, 429, 1221, 4796, 3827, 4045, 11800, 13149]. Which [546, 1148, 553, 2870, 4316, 1715, 1648, 11526, 11967]. Whiffletree [1916, 2842, 4471, 6997, 12499]. Whiffletrees [6220]. Whip [363, 5615]. Whipping [505]. Whipple [7000]. Whirlwinds [13035]. Whisk [3088]. Whiskers [11196]. Whiskey [1549]. Whistle [13229]. Whistling [3508, 13095]. White [4367, 9251, 10698, 12287, 595, 501, 2369, 2175, 8573, 9810, 10697, 11673]. White-Flowered [9251]. Whitehead [5116]. Whitford [6241]. Whitmore [5692]. Whittling [11994]. Whitton [7461]. Whitus [2920]. Whitworth [6513, 5464]. Who [516, 1052, 772, 813, 1646, 2015, 2645, 3398, 3508, 9501, 7985, 12121, 11969, 12092, 12878]. Wholesale [9668]. Whom [2646, 4057]. Whooping [13673]. Wick [2101]. Wicked [6767]. Wicker [7387]. Wickersham [8743]. Wicket [896]. Wicks [34, 2599, 11506]. Wide [1298, 2012]. Width [3582]. Wieting [9559]. Wiggin [10111]. Wilcox [2386]. Wild [8197, 11294, 13447]. Wilhelm [4613]. Wilkinson [10151]. Will [346, 2170, 1969, 2577, 2605, 2630, 6550, 7551, 11256, 12092, 12459]. Willahs [6433]. Willes [9531]. William [6130, 10917, 270, 3270, 3544, 5786, 12307]. Williams [321, 2290, 1400]. Wilmington [3020]. Wilson [3070, 12078, 12973, 5302]. Wimbledon [2124]. Winch [1963, 6604]. Wind [351, 3723, 4119, 5876, 10140, 13511]. Winding [12858]. Windmill [6940, 6923, 11113]. Window [204, 2107, 2414, 1894, 4803, 8452, 10778, 10726, 11913, 12422, 12988, 13593, 12830]. Windows [11320]. Windsor [4242, 1989]. Wine [1693, 10640, 12294]. Winged [12971]. Wink [6188]. Winn [2843]. Wins [11607]. Winter [99, 3803, 6425, 6390, 8659, 9151, 9676, 10930, 13171]. Winter-Flowering [8659]. Winterburn [11506]. Wintering [10744]. Wire [534, 484, 28, 1634, 1267, 1754, 2739, 2756, 3418, 4873, 5827, 6498, 5826, 7558, 6803, 7361, 6934, 9743, 12238, 11089, 12557, 12700]. Wires [1507, 6083, 8832, 11133, 12935]. Wiring [3103]. Wise [11144]. Without [3035, 1161, 53, 4854, 5733, 6185, 8763, 11133, 715, 5305]. Wolf [11272]. Wolff [7053]. Woman [2309, 1344, 4551, 6185, 8415, 11340, 12699]. Women [594, 694, 3976, 8591, 11316]. Wonder [5795]. Wonderful [3442, 10797, 9525, 10093, 13354, 12503]. Wonders [4381, 4404, 4434, 4324, 4350, 4267, 4296, 5442]. Wood [595, 740, 276, 215, 293, 690, 1094, 1847, 1369, 1773, 2186, 1481, 1553, 1519, 2540, 3035, 2551, 3942, 5008, 5414, 6713, 7379, 7839, 7931, 9279, 8424, 10387, 11625, 11734, 11212, 11460, 11712, 12485, 13286, 12947, 13198, 13426, 1429, 1772, 1871, 7563, 9056, 11844]. Wood-Boring [7839, 7931, 1772, 1871].

Wood-Carving [7379]. Wood-Chuck [595]. Wood-Preserving [11212]. Wood-Sole [11734]. Woodbury [967, 7748, 6598, 7503, 8479, 12489]. Woodcock [13647, 12340]. Woodcut [7643]. Wooden [10771, 11146]. Woodruff [10497, 10577, 12441]. Woods [4233]. Woodwork [7830, 7889, 3870]. Woodworking [7706, 7797, 7918, 8933, 8987, 9364, 8001, 8032, 9066, 9100, 9128, 9160, 9188, 9223, 9260, 9291, 9322, 9353, 9414, 9447, 8114, 8145, 8173, 8202, 8232, 8262, 8313, 8371, 8405, 8434, 8465, 8526, 8550, 8578, 8609, 8635, 8666, 8723, 8816, 8849, 8883, 8912, 8970, 9032, 8158, 8102, 9884, 9569, 9658, 10548, 10586, 10647, 10674, 10701, 10758, 10786, 9684, 9709, 9759, 9791, 9816, 9867, 9897, 9926, 9948, 9996, 10018, 10049, 10131, 10159, 10184, 10210, 10240, 10268, 10295, 10330, 10362, 10394, 10425, 10456]. Wool [1186, 2165, 2588, 6582, 11429, 11013, 12102, 190]. Woolwich [6272]. Worcester [5247, 12731]. Word [2324, 5156, 4029]. Words [2970, 5519, 13336]. Work [1047, 1074, 170, 1053, 2321, 1847, 1827, 1369, 1776, 2172, 1973, 1850, 2393, 3394, 2994, 2631, 2604, 3708, 3047, 3063, 4319, 4063, 6423, 5748, 5410, 5708, 6797, 7855, 7281, 9437, 9280, 8900, 8182, 7996, 8331, 8602, 8298, 8415, 10446, 10659, 9692, 10318, 10930, 11193, 11033, 11610, 11340, 12223, 11801, 11258, 12026, 11257, 12152, 13425, 12717, 13564, 13231, 12321]. Work-Men [1827]. Worker [5851, 6713, 9116, 10226]. Workers [11832]. Working [740, 269, 2186, 2149, 1539, 3424, 4234, 7514, 7193, 7874, 7814, 9270, 8983, 8706, 8012, 7981, 10688, 9887, 9771, 11255, 11806, 11761, 11970, 11704, 11313, 11316]. Workmanship [2524, 9429]. Workmen [1646, 3524, 4144, 4284, 6423, 5515, 8861, 9080, 9014, 10463, 13272]. Works [792, 2361, 2238, 1822, 2784, 2808, 4946, 4302, 3897, 4241, 5164, 5040, 6373, 6008, 6765, 7429, 8874, 8180, 8964, 8221, 8097, 10309, 10006, 10772, 12539, 2312, 4269, 6678, 7843, 9565, 12732, 12335]. Workshop [11374]. World [1406, 2044, 4341, 4633, 4812, 4319, 4185, 4516, 6495, 7855, 7871, 9675, 9620, 12199, 12846, 12550, 12879, 1053, 9802, 12118, 12845, 12730]. World-Circuit [12879]. Worlds [621, 10749]. Worm [907, 1615, 1701, 5447, 7157, 9435, 11928]. Wormood [392]. Worms [10598, 12079]. Worry [8507]. Worse [12329]. Worswick [6302]. Would [10749, 12408]. Wounds [6378, 11163]. Woven [28, 1754, 13687]. Woven-Wire [28]. Wrapper [7369]. Wreck [4230, 7095]. Wren [10772]. Wrench [1664, 3525, 4805, 6723, 8284, 8901, 8801, 8188, 10043, 11407, 11670, 12127, 13494]. Wright [2765]. Wringer [1163]. Wrinkle [12027]. Wrinkles [7718]. Wrists [11971]. Write [10916]. Writer [466, 3176, 4029]. Writers [1912]. Writes [12465]. Writing [750, 2338, 2261, 1480, 3327, 5789, 10203, 12652, 12651, 12972]. Wrongs [9803]. Wrought [1673, 1960, 2816, 2589, 9778, 9831, 9607, 5302]. Wurtz [8605]. Wyckoff [1094].

XXI [6947]. XXII [602]. XXIII [1172]. XXX [5870]. XXXVII [10929]. Xylonite [1263].

Yacht [495, 6279, 12728]. Yachting [8894, 9784]. Yachts [8894, 13505]. Yak [12831]. Yale [883, 179, 2088, 2709]. Yama [5356]. Yama-Mai [5356]. Yard [9968]. Yarn [7082, 9243]. Yaupon [216]. Year [13, 73, 180, 1374, 2419, 1445, 1317, 2566, 3770, 4511, 3814, 6974, 9493, 9667, 13633, 5351, 7383, 12152, 13686]. Yearly [6480]. Years [1428, 6381, 7328, 12850]. Yellow [9139, 12760, 13221]. Yellowstone [6194]. Yielding [11649]. Yoke [1744, 7306]. Yokes [12494]. York [651, 402, 1520, 1803, 2112, 3368, 3240, 4273, 4876, 4491, 3895, 3829, 6523, 6542, 6062, 6408, 6430, 6450, 6500, 6553, 7502, 6755, 7125, 6579, 6677, 6720, 8566, 8748, 8844, 9430, 8793, 8298, 10176, 9566, 9643, 9727, 9785, 9893, 9993, 10069, 10097, 10126, 10820, 10854, 12144, 10983, 11088, 11428, 11258, 12225, 12173, 12737, 12892, 12940, 13607, 13664, 966, 1047, 453, 961, 512, 529, 203, 202, 929, 792, 1395, 1798, 2390, 1547, 1577, 1554, 1674, 1464, 1748, 2388, 2082, 1304, 2042, 1571, 1514, 2409, 1941, 1539, 2859, 3279]. York [2809, 3018, 2641, 2708, 3091, 3082, 3067, 3023, 3269, 4065, 4902, 4956, 4966, 5132, 4400, 3974, 4093, 3889, 3924, 3954, 3865, 3815, 4035, 5999, 6503, 6401, 6765, 6896, 6647, 7050, 6907, 7177, 6838, 6865, 6964, 7327, 9477, 8476, 7952, 9305, 8321, 9399, 9695, 9911, 9721, 10934, 10408, 10626, 10110, 11232, 12168, 12118, 11579, 11310, 11012, 13541, 13378, 12412, 12817, 12759, 12981, 12730, 13402, 13457, 551]. Yorker [1659]. Yosemite [295, 344]. Young [1695, 13630, 1766, 3628, 5403, 6092, 8621, 10942, 12733, 13336, 13610]. Youngman [6166]. Youth [12904]. Youthful [2586]. Yucatan [139]. Yucca [5144, 7763]. Yuccas [6529, 8520]. Yule [11955].

Zealand [7559, 13355]. Zee [8299]. Zen [12393]. Zerah [466]. Zimmer [2585]. Zimmerman [7156]. Zinc [747, 659, 2223, 2175, 4367, 4662, 10143, 12287, 188, 145]. Zinc-Lined [10143]. Zircon [257]. Zirconium [472, 4263]. Zodiacal [4632, 5734]. Zoological [90, 11760, 11982, 13457]. Zuyder [8299].

References

Anonymous:1870:Ca

[1] Anonymous. Correspondence. Scientific American, 22(1):11–12, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

       (electronic). URL http://www.nature.com/scientificamerican/
       journal/v22/n1/pdf/scientificamerican01011870-11.pdf.

Anonymous:1870:BPa

[2] Anonymous. Business and personal. Scientific American, 22(1):15, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

       (electronic). URL http://www.nature.com/scientificamerican/
       journal/v22/n1/pdf/scientificamerican01011870-15a.pdf.

Anonymous:1870:BNBa

[3] Anonymous. Books: New books and publications. Scientific American, 22(1):17, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
17.pdf.

Anonymous:1870:GTB

[4] Anonymous. Gunboats, the torpedo boat problem. Scientific American, 22(1):1, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
1.pdf.

Anonymous:1870:SML

[5] Anonymous. Sugar making in Louisiana. Scientific American, 22(1):1–2, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-1a.pdf.

Anonymous:1870:IHP

[6] Anonymous. An improved hoisting pulley wanted. Scientific American, 22(1):2, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
2.pdf.

Anonymous:1870:FLC

[7] Anonymous. Ferdinand de Lesseps — chief promoter of the Suez Canal. Scientific American, 22(1):2, January 1, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n1/pdf/scientificamerican01011870-2a.pdf.

Anonymous:1870:IVP

[8] Anonymous. An ingenious vent peg. Scientific American, 22(1):2–3, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-2b.pdf.

Anonymous:1870:AER

[9] Anonymous. American and English railway practice contrasted. Scientific American, 22(1):4, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
4.pdf.

Ohara:1870:BC

[10] C. M. O’hara. Boiler covering. Scientific American, 22(1):4–5, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-4a.pdf.

Anonymous:1870:PD

[11] Anonymous. Patent decision. Scientific American, 22(1):5–6, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-5.pdf.

Anonymous:1870:GPG

[12] Anonymous. Girard’s “palier glissant”. Scientific American, 22(1):6, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-6.pdf.

Anonymous:1870:HNY

[13] Anonymous. A happy new year! Scientific American, 22(1):7, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-7.pdf.

Anonymous:1870:SCY

[14] Anonymous. The Suez Canal not yet a failure. Scientific American, 22(1):7, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
7a.pdf.

Anonymous:1870:TBB

[15] Anonymous. Tubular boilers and boiler explosions. Scientific American, 22(1):7, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
7b.pdf.

Anonymous:1870:ARM

[16] Anonymous. American railway management. Scientific American, 22(1):7, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
7c.pdf.

Anonymous:1870:AIP

[17] Anonymous. The American institute prizes awarded to steam engines. Scientific American, 22(1):7, January 1, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n1/pdf/scientificamerican01011870-7d.pdf.

Anonymous:1870:PAC

[18] Anonymous. A protest against the Canadian patent law. Scientific American, 22(1):7–8, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
7e.pdf.

Anonymous:1870:BS

[19] Anonymous. The brighter side. Scientific American, 22(1):8, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-8.pdf.

Anonymous:1870:PFL

[20] Anonymous. Professor Fiske’s lectures at Harvard. Scientific American, 22(1):8, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
8a.pdf.

Anonymous:1870:SLBa

[21] Anonymous. Scientific lectures before the American institute. Scientific American, 22(1):8, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
8b.pdf.

Anonymous:1870:BFS

[22] Anonymous. The battle fields of science. Scientific American, 22(1):8–9, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-8c.pdf.

Anonymous:1870:HFBa

[23] Anonymous. How French bank notes are made. Scientific American, 22(1):9, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
9.pdf.

Anonymous:1870:WNS

[24] Anonymous. What the newspapers say. Scientific American, 22(1):9, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-9a.pdf.

Anonymous:1870:ESa

[25] Anonymous. Editorial summary. Scientific American, 22(1):9, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-9b.pdf.

Anonymous:1870:SBL

[26] Anonymous. The Stevens breech-loading rifle. Scientific American, 22(1):10, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
10.pdf.

Anonymous:1870:NFH

[27] Anonymous. A novel French hand vise. Scientific American, 22(1):10, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-10a.pdf.

Anonymous:1870:WWMa

[28] Anonymous. The woven-wire mattress. Scientific American, 22(1):10, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-10b.pdf.

Colton:1870:PI

[29] Henry E. Colton. Preservation of iron. Scientific American, 22(1):12, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-12.pdf.

Anonymous:1870:BPT

[30] Anonymous. The bananas and plantains of the tropics. Scientific American, 22(1):12–13, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
12a.pdf.

Twain:1870:PS

[31] Mark Twain. Putting up stoves. Scientific American, 22(1):13, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-13.pdf.

Anonymous:1870:MLa

[32] Anonymous. The magic lantern. Scientific American, 22(1):13, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-13a.pdf.

Anonymous:1870:PBa

[33] Anonymous. Paper for building. Scientific American, 22(1):13–15, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-13b.pdf.

Anonymous:1870:ILW

[34] Anonymous. Improvement in lamp wicks. Scientific American, 22(1):15, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-15.pdf.

Anonymous:1870:ACa

[35] Anonymous. Answers to correspondents. Scientific American, 22(1):16, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-16.pdf.

Anonymous:1870:RAFa

[36] Anonymous. Recent American and foreign patents. Scientific American, 22(1):16–17, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n1/pdf/scientificamerican01011870-
16a.pdf.

Anonymous:1870:HPOa

[37] Anonymous. H. S. Patent Office. Scientific American, 22(1):18, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-18.pdf.

Anonymous:1870:OLPa

[38] Anonymous. Official list of patents. Scientific American, 22(1):18–20, January 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n1/pdf/scientificamerican01011870-18a.pdf.

Anonymous:1870:Cb

[39] Anonymous. Correspondence. Scientific American, 22(2):27–28, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-27.pdf.

Anonymous:1870:BPb

[40] Anonymous. Business and personal. Scientific American, 22(2):34, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-34.pdf.

Anonymous:1870:BNBb

[41] Anonymous. Books: New books and publications. Scientific American, 22(2):35, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
35.pdf.

Anonymous:1870:CSE

[42] Anonymous. The Corliss steam engine, how a fish swims. Scientific American, 22(2):23–24, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
23.pdf.

Anonymous:1870:PMO

[43] Anonymous. Professor Morse’s official report upon the telegraphy of the French Exposition. Scientific American, 22(2):24–25, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-24.pdf.

Joy:1870:MSC

[44] Charles A. Joy. On manganese and some of its compounds. Scientific American, 22(2):25–26, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
25.pdf.

Cantini:1870:ALP

[45] I. Cantini. Artificial light from the pinewood chip to the gas chandelier. Scientific American, 22(2):28–29, January 8, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n2/pdf/scientificamerican01081870-28.pdf.

Gordon:1870:CT

[46] John Ramsay Gordon. The calabash tree. Scientific American, 22(2):29–30, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
29.pdf.

Anonymous:1870:WIE

[47] Anonymous. Wells’ improved eccentric for steam engines. Scientific American, 22(2):30, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
30.pdf.

Anonymous:1870:SCA

[48] Anonymous. Swedish cooking appliances. Scientific American, 22(2):30, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-30a.pdf.

Anonymous:1870:RSCa

[49] Anonymous. Report of the Special Commissioner of Revenue — tinkering the income tax. Scientific American, 22(2):31, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-31.pdf.

Anonymous:1870:UE

[50] Anonymous. The use of eyes. Scientific American, 22(2):31, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-31a.pdf.

Anonymous:1870:MSa

[51] Anonymous. The metric system. Scientific American, 22(2):31, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-31b.pdf.

Anonymous:1870:GM

[52] Anonymous. The growth of monopolies. Scientific American, 22 (2):31–32, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
31c.pdf.

Anonymous:1870:PMSa

[53] Anonymous. The preservation of meats without salt. Scientific American, 22(2):32, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
32.pdf.

Anonymous:1870:CTI

[54] Anonymous. Canal through the Isthmus of Darien. Scientific American, 22(2):32, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
32a.pdf.

Anonymous:1870:NFA

[55] Anonymous. New facts about the preservation of timber. Scientific American, 22(2):32, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
32b.pdf.

Anonymous:1870:MET

[56] Anonymous. Madder extracts and their application in topical dyeing. Scientific American, 22(2):32–33, January 8, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n2/pdf/scientificamerican01081870-32c.pdf.

Anonymous:1870:AN

[57] Anonymous. Aerial navigation. Scientific American, 22(2):33, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-33.pdf.

Anonymous:1870:ACb

[58] Anonymous. Answers to correspondents. Scientific American, 22(2):34, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-34a.pdf.

Anonymous:1870:RAFb

[59] Anonymous. Recent American and foreign patents. Scientific American, 22(2):34–35, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
34b.pdf.

Anonymous:1870:OLPb

[60] Anonymous. Official list of patents. Scientific American, 22(2):35–36, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n2/pdf/scientificamerican01081870-35a.pdf.

Anonymous:1870:AEPa

[61] Anonymous. Applications for extension of patents. Scientific American, 22(2):36, January 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n2/pdf/scientificamerican01081870-
36.pdf.

Anonymous:1870:Cc

[62] Anonymous. Correspondence. Scientific American, 22(3):43–44, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-43b.pdf.

Anonymous:1870:BPc

[63] Anonymous. Business and personal. Scientific American, 22(3):50, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-50b.pdf.

Anonymous:1870:BNBc

[64] Anonymous. Books: New books and publications. Scientific American, 22(3):51, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
51.pdf.

Anonymous:1870:MBA

[65] Anonymous. Machine for branching artificial flowers. Scientific American, 22(3):39, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
39.pdf.

Anonymous:1870:WHS

[66] Anonymous. Waste not — how small things are utilized. Scientific American, 22(3):39–40, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
39a.pdf.

Anonymous:1870:RBa

[67] Anonymous. Repaired boilers. Scientific American, 22(3):40–41, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-40.pdf.

Reimann:1870:WLG

[68] Reimann. To wash leather gloves. Scientific American, 22(3): 41–43, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
41.pdf.

Anonymous:1870:LCI

[69] Anonymous. Laying cast iron pipes under water. Scientific American, 22(3):43, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
43.pdf.

Anonymous:1870:SMAa

[70] Anonymous. Self-moving anchor for steam cultivation. Scientific American, 22(3):43, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
43a.pdf.

Joy:1870:C

[71] Charles A. Joy. On chromium. Scientific American, 22(3):45, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-45.pdf.

Anonymous:1870:MDC

[72] Anonymous. Milk — from the depot to the consumer. Scientific American, 22(3):45–46, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
45a.pdf.

Anonymous:1870:PPSa

[73] Anonymous. The progress of the physical sciences in the year 1869. Scientific American, 22(3):47, January 15, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n3/pdf/scientificamerican01151870-47.pdf.

Anonymous:1870:SJ

[74] Anonymous. Specialistic journalism. Scientific American, 22(3): 47–48, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
47a.pdf.

Anonymous:1870:AAF

[75] Anonymous. American aeronautics — a flying ship to be built. Scientific American, 22(3):48, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
48.pdf.

Anonymous:1870:PBb

[76] Anonymous. Pumping down buildings. Scientific American, 22 (3):48, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
48a.pdf.

Anonymous:1870:SMb

[77] Anonymous. Silvering mirrors. Scientific American, 22(3):48, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-48b.pdf.

Anonymous:1870:CVP

[78] Anonymous. Correlation of vital and physical forces. Scientific American, 22(3):48–50, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
48c.pdf.

Anonymous:1870:USCa

[79] Anonymous. United States Circuit Court — Southern District before Judge Blatchford. Scientific American, 22(3):50, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n3/pdf/scientificamerican01151870-50.pdf.

Anonymous:1870:ACc

[80] Anonymous. Answers to correspondents. Scientific American, 22 (3):50, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
50a.pdf.

Anonymous:1870:RAFc

[81] Anonymous. Recent American and foreign patents. Scientific American, 22(3):50–51, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
50c.pdf.

Anonymous:1870:AEPb

[82] Anonymous. Applications for extension of patents. Scientific American, 22(3):51, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
51a.pdf.

Anonymous:1870:OLPc

[83] Anonymous. Official list of patents. Scientific American, 22 (3):51–52, January 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n3/pdf/scientificamerican01151870-
51b.pdf.

Anonymous:1870:Cd

[84] Anonymous. Correspondence. Scientific American, 22(4):60–61, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-60.pdf.

Anonymous:1870:BNBd

[85] Anonymous. Books: New books and publications. Scientific American, 22(4):66, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
66a.pdf.

Anonymous:1870:BPd

[86] Anonymous. Business and personal. Scientific American, 22(4):66, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-66b.pdf.

Anonymous:1870:DSR

[87] Anonymous. Device for steaming rovings, pressure in steam boilers, and more. Scientific American, 22(4):55, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n4/pdf/scientificamerican01221870-55.pdf.

Anonymous:1870:LB

[88] Anonymous. The Leclanche battery. Scientific American, 22 (4):55–56, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
55a.pdf.

Anonymous:1870:SPH

[89] Anonymous. Science and the public health. Scientific American, 22(4):56, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
56.pdf.

Anonymous:1870:RZG

[90] Anonymous. The reptiles at the zoological gardens. Scientific American, 22(4):56–57, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
56a.pdf.

Anonymous:1870:RRD

[91] Anonymous. Rain and rain doctors. Scientific American, 22 (4):57–59, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
57.pdf.

Anonymous:1870:IPP

[92] Anonymous. Improved packing for the pistons of steam engines. Scientific American, 22(4):59, January 22, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n4/pdf/scientificamerican01221870-59.pdf.

Anonymous:1870:CIFa

[93] Anonymous. Curious ice formations. Scientific American, 22 (4):59–60, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
59a.pdf.

VanderWeyde:1870:NIM

[94] P. R. Vander Weyde. On the nature of the interplanetary medium. Scientific American, 22(4):61, January 22, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n4/pdf/scientificamerican01221870-61.pdf.

Wideumann:1870:MPN

[95] C. Wideumann. Manufacture of the points of needles and pins by electricity. Scientific American, 22(4):61–62, January 22, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n4/pdf/scientificamerican01221870-61a.pdf.

Anonymous:1870:PME

[96] Anonymous. Progress in mechanical and engineering science and invention in 1869. Scientific American, 22(4):63, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-63.pdf.

Anonymous:1870:NDP

[97] Anonymous. Nature of design patents. Scientific American, 22 (4):63, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
63a.pdf.

Anonymous:1870:CUL

[98] Anonymous. The careless use of Lucifer matches. Scientific American, 22(4):63–64, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
63b.pdf.

Anonymous:1870:MW

[99] Anonymous. The mild winter. Scientific American, 22(4):64, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-64.pdf.

Anonymous:1870:SIa

[100] Anonymous. Scientific intelligence. Scientific American, 22(4):64, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-64a.pdf.

Widemann:1870:CSCa

[101] C. Widemann and Chemist. On cotton seed and cotton seed oil. Scientific American, 22(4):64, January 22, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n4/pdf/scientificamerican01221870-64b.pdf.

Anonymous:1870:MFa

[102] Anonymous. Michael Faraday. Scientific American, 22(4):65–66, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-65.pdf.

Anonymous:1870:USCb

[103] Anonymous. United States Circuit Court — Southern District before Judge Blatchford. Scientific American, 22(4):66, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n4/pdf/scientificamerican01221870-66.pdf.

Anonymous:1870:ACd

[104] Anonymous. Answers to correspondents. Scientific American, 22 (4):66, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
66c.pdf.

Anonymous:1870:RAFd

[105] Anonymous. Recent American and foreign patents. Scientific American, 22(4):66–67, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
66d.pdf.

Anonymous:1870:OLPd

[106] Anonymous. Official list of patents. Scientific American, 22 (4):67–68, January 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n4/pdf/scientificamerican01221870-
67.pdf.

Anonymous:1870:Ce

[107] Anonymous. Correspondence. Scientific American, 22(5):76–77, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-76.pdf.

Anonymous:1870:BPe

[108] Anonymous. Business and personal. Scientific American, 22(5):83, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-83.pdf.

Anonymous:1870:BNBe

[109] Anonymous. Books: New books and publications. Scientific American, 22(5):83, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
83b.pdf.

Anonymous:1870:ISHa

[110] Anonymous. Improvement in sulky harrows, lighting mines with gas. Scientific American, 22(5):71, January 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n5/pdf/scientificamerican01291870-71.pdf.

Anonymous:1870:DRB

[111] Anonymous. Double ratchet brace. Scientific American, 22(5):71, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-71a.pdf.

Anonymous:1870:AE

[112] Anonymous. Agricultural engineering. Scientific American, 22 (5):71–72, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
71b.pdf.

Anonymous:1870:AMIa

[113] Anonymous. Ancient and modern industries of the Chinese. Scientific American, 22(5):72–73, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
72.pdf.

Anonymous:1870:MII

[114] Anonymous. Manufacture of Indian ink. Scientific American, 22(5):73, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
73.pdf.

Anonymous:1870:TSE

[115] Anonymous. The tanite solid emery wheel. Scientific American, 22(5):73–75, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
73a.pdf.

Anonymous:1870:IOF

[116] Anonymous. Improved open fire grate. Scientific American, 22 (5):75, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
75.pdf.

Cantini:1870:EF

[117] I. Cantini. The Egyptian fellah. Scientific American, 22(5):77, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-77.pdf.

Gordon:1870:B

[118] John Ramsay Gordon. The barracouta. Scientific American, 22 (5):77, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
77a.pdf.

Anonymous:1870:ILE

[119] Anonymous. Improved lever escapement. Scientific American, 22(5):78, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
78.pdf.

Anonymous:1870:OSE

[120] Anonymous. Our splendid engraving. Scientific American, 22 (5):78, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
78a.pdf.

Anonymous:1870:LSS

[121] Anonymous. Ludlow’s sliding stop valves for steam, water, and gas. Scientific American, 22(5):78, January 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n5/pdf/scientificamerican01291870-78b.pdf.

Anonymous:1870:SAa

[122] Anonymous. The “Scientific American” in 1869. Scientific American, 22(5):79, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
79.pdf.

Anonymous:1870:TWH

[123] Anonymous. Tin — where and how it is obtained. Scientific American, 22(5):79, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
79a.pdf.

Anonymous:1870:WNM

[124] Anonymous. Is water necessary to mitigate the effects of dry air from heating furnaces? Scientific American, 22(5):79–80, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-79b.pdf.

Anonymous:1870:WBS

[125] Anonymous. Wanted — better sidewalks. Scientific American, 22(5):80, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
80.pdf.

Anonymous:1870:PFA

[126] Anonymous. A peculiar feature of agricultural invention. Scientific American, 22(5):80, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
80a.pdf.

Anonymous:1870:SET

[127] Anonymous. The steam engine trials at the American Institute Exhibition. Scientific American, 22(5):80, January 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n5/pdf/scientificamerican01291870-80b.pdf.

Anonymous:1870:AME

[128] Anonymous. Ancient modes of extinguishing fires. Scientific American, 22(5):80–81, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
80c.pdf.

Anonymous:1870:SPa

[129] Anonymous. Slate pencils. Scientific American, 22(5):81, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-81.pdf.

Widemann:1870:CSCb

[130] C. Widemann and Chemist. On cotton seed and cotton seed oil. Scientific American, 22(5):81–82, January 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n5/pdf/scientificamerican01291870-81a.pdf.

Anonymous:1870:USCc

[131] Anonymous. United States Circuit Court — Southern District before Judge Blatchford. Scientific American, 22(5):82, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n5/pdf/scientificamerican01291870-82.pdf.

Anonymous:1870:RAFe

[132] Anonymous. Recent American and foreign patents. Scientific American, 22(5):82–83, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
82a.pdf.

Anonymous:1870:ACe

[133] Anonymous. Answers to correspondents. Scientific American, 22 (5):83, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
83a.pdf.

Anonymous:1870:AEPc

[134] Anonymous. Applications for extension of patents. Scientific American, 22(5):83, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
83c.pdf.

Anonymous:1870:OLPe

[135] Anonymous. Official list of patents. Scientific American, 22 (5):83–84, January 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n5/pdf/scientificamerican01291870-
83d.pdf.

Anonymous:1870:Cf

[136] Anonymous. Correspondence. Scientific American, 22(6):91–92, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-91a.pdf.

Anonymous:1870:BPf

[137] Anonymous. Business and personal. Scientific American, 22 (6):99, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
99.pdf.

Anonymous:1870:CMS

[138] Anonymous. Cigar machine, scientific nonsense, and more. Scientific American, 22(6):87, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
87.pdf.

Anonymous:1870:NY

[139] Anonymous. The Niin of Yucatan. Scientific American, 22(6):88, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-88.pdf.

Hay:1870:CGI

[140] Elzy Hay. The cotton gin and its inventors. Scientific American, 22(6):88, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
88a.pdf.

Colton:1870:ISA

[141] Henry E. Colton. Iron shipbuilding and American commerce. Scientific American, 22(6):89, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
89.pdf.

Widemann:1870:A

[142] C. Widemann. Albumen. Scientific American, 22(6):89–91, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-89a.pdf.

Widemann:1870:MCSa

[143] C. Widemann and Chemist. Manufacture of cotton seed, cotton seed cakes, and meal. Scientific American, 22(6):91, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-91.pdf.

Anonymous:1870:FSS

[144] Anonymous. The French system of storm maps. Scientific American, 22(6):92, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
92.pdf.

Joy:1870:Z

[145] Charles A. Joy. On zinc. Scientific American, 22(6):93, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-93.pdf.

Anonymous:1870:NEO

[146] Anonymous. A new English oil testing apparatus. Scientific American, 22(6):94, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
94.pdf.

Anonymous:1870:CHWa

[147] Anonymous. Construction of hollow walls. Scientific American, 22(6):94, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
94a.pdf.

Anonymous:1870:DWF

[148] Anonymous. Domestic water filters. Scientific American, 22 (6):94, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
94b.pdf.

Anonymous:1870:NAM

[149] Anonymous. New apparatus for measuring oil. Scientific American, 22(6):94, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
94c.pdf.

Anonymous:1870:SEA

[150] Anonymous. Steam engines at the American institute. Scientific American, 22(6):94, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
94d.pdf.

Anonymous:1870:RM

[151] Anonymous. The rumford medals. Scientific American, 22(6):95, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-95.pdf.

Anonymous:1870:SHS

[152] Anonymous. Solutions for hardening steel. Scientific American, 22(6):95, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
95a.pdf.

Anonymous:1870:SFP

[153] Anonymous. Study of first principles by inventors. Scientific American, 22(6):95, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
95b.pdf.

Anonymous:1870:PBF

[154] Anonymous. Pumping down buildings — fall of the Dining Hall of King’s College, London. Scientific American, 22(6):95–96, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-95c.pdf.

Anonymous:1870:TTW

[155] Anonymous. Test of turbine water wheels at Lowell, Mass. Scientific American, 22(6):96, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
96.pdf.

Anonymous:1870:BRF

[156] Anonymous. Black River flood case. Scientific American, 22 (6):96, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
96a.pdf.

Anonymous:1870:FS

[157] Anonymous. Foremen and superintendents. Scientific American, 22(6):96–97, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
96b.pdf.

Anonymous:1870:DDA

[158] Anonymous. Diving and diving apparatus. Scientific American, 22(6):97, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
97.pdf.

Anonymous:1870:TVB

[159] Anonymous. Tunnels vs bridges. Scientific American, 22(6):97, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n6/pdf/scientificamerican02051870-97a.pdf.

Anonymous:1870:WSOa

[160] Anonymous. What is said of our prize engraving. Scientific American, 22(6):97, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
97b.pdf.

Anonymous:1870:RMD

[161] Anonymous. On refrigerating mixtures and the depression of temperature produced by dissolving salts in water. Scientific American, 22(6):98, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
98.pdf.

Anonymous:1870:AEPd

[162] Anonymous. Applications for extension of patents. Scientific American, 22(6):98, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
98a.pdf.

Anonymous:1870:ACf

[163] Anonymous. Answers to correspondents. Scientific American, 22 (6):99, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
99a.pdf.

Anonymous:1870:RAFf

[164] Anonymous. Recent American and foreign patents. Scientific American, 22(6):99, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
99b.pdf.

Anonymous:1870:OLPf

[165] Anonymous. Official list of patents. Scientific American, 22(6): 100–101, February 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n6/pdf/scientificamerican02051870-
100.pdf.

Anonymous:1870:Cg

[166] Anonymous. Correspondence. Scientific American, 22(7):108– 110, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
108a.pdf.

Anonymous:1870:BPg

[167] Anonymous. Business and personal. Scientific American, 22(7): 116, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
116.pdf.

Anonymous:1870:RSCb

[168] Anonymous. The ransom siphon condenser. Scientific American, 22(7):103, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
103.pdf.

Anonymous:1870:FC

[169] Anonymous. Flooring clamp. Scientific American, 22(7):103, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n7/pdf/scientificamerican02121870-103a.pdf.

Anonymous:1870:LW

[170] Anonymous. Lapidary work. Scientific American, 22(7):104, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n7/pdf/scientificamerican02121870-104.pdf.

Anonymous:1870:HOF

[171] Anonymous. How to obtain a fine patina on bronze statues. Scientific American, 22(7):104, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
104a.pdf.

Anonymous:1870:SIb

[172] Anonymous. Scientific intelligence. Scientific American, 22(7): 104, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
104b.pdf.

Anonymous:1870:WGS

[173] Anonymous. What George Stephenson accomplished. Scientific American, 22(7):104–105, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
104c.pdf.

Anonymous:1870:MEM

[174] Anonymous. Microscopical examination of milk under certain conditions. Scientific American, 22(7):105, February 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n7/pdf/scientificamerican02121870-105.pdf.

Anonymous:1870:HT

[175] Anonymous. The Hoosac Tunnel. Scientific American, 22(7): 106–107, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
106.pdf.

Anonymous:1870:FAR

[176] Anonymous. French asphalt roads, and American attempts to imitate them. Scientific American, 22(7):107–108, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n7/pdf/scientificamerican02121870-107.pdf.

Widemann:1870:NP

[177] C. Widemann. New paint. Scientific American, 22(7):108, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n7/pdf/scientificamerican02121870-108.pdf.

Anonymous:1870:ABC

[178] Anonymous. The attainment of beauty in commonplace handiwork. Scientific American, 22(7):111, February 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n7/pdf/scientificamerican02121870-111.pdf.

Anonymous:1870:TEU

[179] Anonymous. Technical education in the United States — the Sheffield Scientific School of Yale College. Scientific American, 22(7):111, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n7/pdf/scientificamerican02121870-111a.pdf.

Anonymous:1870:RCP

[180] Anonymous. Report of the Commissioner of Patents for the year 1869. Scientific American, 22(7):111–114, February 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n7/pdf/scientificamerican02121870-111b.pdf.

Anonymous:1870:RAFg

[181] Anonymous. Recent American and foreign patents. Scientific American, 22(7):114–115, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
114.pdf.

Anonymous:1870:OLPg

[182] Anonymous. Official list of patents. Scientific American, 22(7): 115–116, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
115.pdf.

Anonymous:1870:AEPe

[183] Anonymous. Applications for extension of patents. Scientific American, 22(7):116, February 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n7/pdf/scientificamerican02121870-
116a.pdf.

Anonymous:1870:Ch (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
123.pdf.

Anonymous:1870:BNBf

[185] Anonymous. Books: New books and publications. Scientific American, 22(8):130, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
130.pdf.

Anonymous:1870:BPh

[186] Anonymous. Business and personal. Scientific American, 22(8): 131, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
131a.pdf.

Anonymous:1870:CSS

[187] Anonymous. Carriages and slides of the Spanish gunboats designed by Captain Ericsson. Scientific American, 22(8):119, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n8/pdf/scientificamerican02191870-119.pdf.

Joy:1870:UZ

[188] Charles A. Joy. On the uses of zinc. Scientific American, 22 (8):120, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
120.pdf.

Widemann:1870:MCSb

[189] C. Widemann. Manufacture of cotton seed and cotton seed oil. Scientific American, 22(8):120–121, February 19, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n8/pdf/scientificamerican02191870-120a.pdf.

Reimann:1870:AGW

[190] M. Reimann. Aniline Green for wool. Scientific American, 22 (8):121, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
121.pdf.

Widemann:1870:OG

[191] C. Widemann. On ox gall. Scientific American, 22(8):121, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n8/pdf/scientificamerican02191870-121a.pdf.

Anonymous:1870:PIb

[192] Anonymous. Preservation of ice. Scientific American, 22(8):122, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n8/pdf/scientificamerican02191870-122.pdf.

Anonymous:1870:IBM

[193] Anonymous. Improved bedstead and mattress for invalids. Scientific American, 22(8):122, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
122a.pdf.

Anonymous:1870:PMa

[194] Anonymous. The preservation of meat. Scientific American, 22 (8):122, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
122b.pdf.

Anonymous:1870:ICD

[195] Anonymous. Improved clothes dryer. Scientific American, 22(8): 122–123, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
122c.pdf.

Anonymous:1870:MAN

[196] Anonymous. More about niin. Scientific American, 22(8):125, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n8/pdf/scientificamerican02191870-125.pdf.

Anonymous:1870:HTSa

[197] Anonymous. The hydrostatic test for steam boilers. Scientific American, 22(8):125, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
125a.pdf.

Anonymous:1870:IKH

[198] Anonymous. Improvement in key-hole guards. Scientific American, 22(8):126, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
126.pdf.

Anonymous:1870:SWT

[199] Anonymous. Six-way tap for gas and water. Scientific American, 22(8):126, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
126a.pdf.

Anonymous:1870:PIG

[200] Anonymous. Penny’s improved gang Sawmill. Scientific American, 22(8):126, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
126b.pdf.

Anonymous:1870:PMPa

[201] Anonymous. Proposed modification of the patent laws — an attempt to break down the right of appeal. Scientific American, 22(8):127, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n8/pdf/scientificamerican02191870-127.pdf.

Anonymous:1870:PTUa

[202] Anonymous. The pneumatic tunnel under Broadway, New York. Scientific American, 22(8):127, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
127a.pdf.

Anonymous:1870:ORP

[203] Anonymous. Official report of Professor Chandler to the Board of Health on the New York gas companies. Scientific American, 22 (8):127–128, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
127b.pdf.

Anonymous:1870:BSO

[204] Anonymous. Bent, stained, and ornamental window glass. Scientific American, 22(8):128, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
128.pdf.

Anonymous:1870:WCa

[205] Anonymous. What is chemistry? Scientific American, 22(8): 128–129, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
128a.pdf.

Anonymous:1870:CIFb

[206] Anonymous. Curious ice formations. Scientific American, 22 (8):129, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
129.pdf.

Anonymous:1870:AOC

[207] Anonymous. Award of our cash prizes. Scientific American, 22 (8):129–130, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
129a.pdf.

Anonymous:1870:RAFh

[208] Anonymous. Recent American and foreign patents. Scientific American, 22(8):130–131, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
130a.pdf.

Anonymous:1870:ACg

[209] Anonymous. Answers to correspondents. Scientific American, 22 (8):131, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
131.pdf.

Anonymous:1870:OLPh

[210] Anonymous. Official list of patents. Scientific American, 22(8): 131–132, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
131b.pdf.

Anonymous:1870:AEPf

[211] Anonymous. Applications for extension of patents. Scientific American, 22(8):132, February 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n8/pdf/scientificamerican02191870-
132.pdf.

Anonymous:1870:Ci

[212] Anonymous. Correspondence. Scientific American, 22(9):139– 141, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
139b.pdf.

Anonymous:1870:BNBg

[213] Anonymous. Books: New books and publications. Scientific American, 22(9):145, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
145b.pdf.

Anonymous:1870:BPi

[214] Anonymous. Business and personal. Scientific American, 22(9): 146, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
146a.pdf.

Anonymous:1870:WMMa

[215] Anonymous. Wood molding machine, improvement in steam hammers. Scientific American, 22(9):135–136, February 26, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n9/pdf/scientificamerican02261870-135.pdf.

Colton:1870:YTC

[216] H. E. Colton. Yaupon teas of Carolina — maté of Paraguay. Scientific American, 22(9):136, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
136.pdf.

Widemann:1870:PLGa

[217] C. Widemann. Platinized looking-glasses. Scientific American, 22 (9):136–137, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
136a.pdf.

Cantini:1870:GIP

[218] I. Cantini. Guatemala, its inhabitants and products — advice to emigrants. Scientific American, 22(9):137, February 26, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n9/pdf/scientificamerican02261870-137.pdf.

Anonymous:1870:EAI

[219] Anonymous. English and American iron-clads compared. Scientific American, 22(9):138–139, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
138.pdf.

Anonymous:1870:ECS

[220] Anonymous. The Eagle carpet stretcher. Scientific American, 22(9):139, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
139.pdf.

Anonymous:1870:HHT

[221] Anonymous. How to hitch three horses to one plow. Scientific American, 22(9):139, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
139a.pdf.

Anonymous:1870:JR

[222] Anonymous. A journey with a raft. Scientific American, 22(9): 141–142, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
141.pdf.

Anonymous:1870:IPV

[223] Anonymous. Improved patent vertical drill. Scientific American, 22(9):142, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
142.pdf.

Anonymous:1870:WWTa

[224] Anonymous. The water wheel tests at Lowell. Scientific American, 22(9):143, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
143.pdf.

Anonymous:1870:WMb

[225] Anonymous. The wages of machinery. Scientific American, 22 (9):143, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
143a.pdf.

Anonymous:1870:WLB

[226] Anonymous. Waste of labor in building. Scientific American, 22(9):143, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
143b.pdf.

Anonymous:1870:FEL

[227] Anonymous. French experiments with liquid fuel. Scientific American, 22(9):143–144, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
143c.pdf.

Anonymous:1870:ESE

[228] Anonymous. The economy of steam engines. Scientific American, 22(9):144, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
144.pdf.

Anonymous:1870:LFa

[229] Anonymous. Liebig on fermentation. Scientific American, 22 (9):144, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
144a.pdf.

Anonymous:1870:PP

[230] Anonymous. Positive philosophy. Scientific American, 22(9):144, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n9/pdf/scientificamerican02261870-144b.pdf.

Anonymous:1870:MMb

[231] Anonymous. A mile a minute. Scientific American, 22(9):144– 145, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
144c.pdf.

Anonymous:1870:SIc

[232] Anonymous. Scientific intelligence. Scientific American, 22(9): 145, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
145.pdf.

Anonymous:1870:PODa

[233] Anonymous. Patent Office decisions. Scientific American, 22 (9):145, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
145a.pdf.

Anonymous:1870:RAFi

[234] Anonymous. Recent American and foreign patents. Scientific American, 22(9):145–146, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
145c.pdf.

Anonymous:1870:ACh

[235] Anonymous. Answers to correspondents. Scientific American, 22 (9):146, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
146.pdf.

Anonymous:1870:OLPi

[236] Anonymous. Official list of patents. Scientific American, 22(9): 147–148, February 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n9/pdf/scientificamerican02261870-
147.pdf.

Anonymous:1870:Cj

[237] Anonymous. Correspondence. Scientific American, 22(10):157–158, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-157.pdf.

Anonymous:1870:BPj

[238] Anonymous. Business and personal. Scientific American, 22(10):162, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-162b.pdf.

Anonymous:1870:BNBh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
164a.pdf.

Anonymous:1870:IMM

[240] Anonymous. Improved machine for the manufacture of spoked wheels, polishing granite, and more. Scientific American, 22(10):151–152, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-151.pdf.

Joy:1870:CPU

[241] Charles A. Joy. Cobalt — its properties and uses. Scientific American, 22(10):152, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
152.pdf.

Widemann:1870:PLGb

[242] C. Widemann. Platinized looking-glasses. Scientific American, 22(10):152, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
152a.pdf.

Anonymous:1870:PODb

[243] Anonymous. A plea for “our dumb slaves”. Scientific American, 22(10):153–154, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
153.pdf.

Anonymous:1870:PTUb

[244] Anonymous. The pneumatic tunnel under Broadway, N.Y. Scientific American, 22(10):154–156, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
154.pdf.

Anonymous:1870:HCI

[245] Anonymous. Hubbell & Capron’s improved turbine wheel. Scientific American, 22(10):156, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
156.pdf.

Anonymous:1870:UBM

[246] Anonymous. Universal balance for millstones. Scientific American, 22(10):158, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
158.pdf.

Anonymous:1870:NED

[247] Anonymous. The new explosive Dualin. Scientific American, 22 (10):159, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
159.pdf.

Anonymous:1870:CT

[248] Anonymous. The cultivation of timber. Scientific American, 22(10):159, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-159a.pdf.

Anonymous:1870:LVV

[249] Anonymous. The law of “vis viva”. Scientific American, 22(10):159, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-159b.pdf.

Anonymous:1870:WOE

[250] Anonymous. What oysters eat. Scientific American, 22(10):159, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-159c.pdf.

Anonymous:1870:PSP

[251] Anonymous. A plea for the Society for the Prevention of Cruelty to Animals. Scientific American, 22(10):159–160, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-159d.pdf.

Anonymous:1870:ST

[252] Anonymous. Systematic thinking. Scientific American, 22(10):160, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-160.pdf.

Anonymous:1870:CPO

[253] Anonymous. Concerning Patent Office matters. Scientific American, 22(10):160, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
160a.pdf.

Anonymous:1870:LM

[254] Anonymous. The lightest machine. Scientific American, 22(10):160, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-160b.pdf.

Anonymous:1870:MRS

[255] Anonymous. Mr. Rutherfurd’s star photographs. Scientific American, 22(10):160–161, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
160c.pdf.

Anonymous:1870:THD

[256] Anonymous. Tyndall on haze and dust. Scientific American, 22(10):161, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-161.pdf.

Anonymous:1870:ZL

[257] Anonymous. The zircon light. Scientific American, 22(10):162, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n10/pdf/scientificamerican03051870-162.pdf.

Anonymous:1870:ACi

[258] Anonymous. Answers to correspondents. Scientific American, 22 (10):162, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
162a.pdf.

Anonymous:1870:RAFj

[259] Anonymous. Recent American and foreign patents. Scientific American, 22(10):163, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
163.pdf.

Anonymous:1870:OLPj

[260] Anonymous. Official list of patents. Scientific American, 22 (10):163–164, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
163a.pdf.

Anonymous:1870:AEPg

[261] Anonymous. Applications for extension of patents. Scientific American, 22(10):164, March 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n10/pdf/scientificamerican03051870-
164.pdf.

Anonymous:1870:Ck

[262] Anonymous. Correspondence. Scientific American, 22(11):172–174, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-172c.pdf.

Anonymous:1870:BPk

[263] Anonymous. Business and personal. Scientific American, 22(11):177, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-177a.pdf.

Anonymous:1870:BNBi

[264] Anonymous. Books: New books and publications. Scientific American, 22(11):178–179, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
178a.pdf.

Anonymous:1870:CVR

[265] Anonymous. The cere viaduct, railway economy. Scientific American, 22(11):167–168, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
167.pdf.

Joy:1870:Ta

[266] Charles A. Joy. On tin. Scientific American, 22(11):168, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-168.pdf.

Gordon:1870:CP

[267] John Ramsay Gordon. The century plant. Scientific American, 22(11):168, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
168a.pdf.

Anonymous:1870:NMS

[268] Anonymous. On a new method of straightening high chimneys. Scientific American, 22(11):169, March 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n11/pdf/scientificamerican03121870-169.pdf.

Anonymous:1870:WMA

[269] Anonymous. Working millstones with an air blast. Scientific American, 22(11):170, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
170.pdf.

Anonymous:1870:WIT

[270] Anonymous. William’s improved tracelock. Scientific American, 22(11):170, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
170a.pdf.

Anonymous:1870:JDM

[271] Anonymous. J. D. Michael’s patent egg box. Scientific American, 22(11):170–171, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
170b.pdf.

Widemann:1870:PLGc

[272] C. Widemann. Platinized looking-glasses. Scientific American, 22(11):172, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
172.pdf.

Widemann:1870:APA

[273] C. Widemann. Apparatus for purifying the air by the evaporation of coal tar, pitch, carbolic acid, phenic acid, or any other disinfectant for apartments, or hospitals. Scientific American, 22(11):172, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-172a.pdf.

Anonymous:1870:CIS

[274] Anonymous. L. S. Clark’s improved shaft-bar for single sleighs. Scientific American, 22(11):172, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
172b.pdf.

Anonymous:1870:SBC

[275] Anonymous. Spring balance for chemicals. Scientific American, 22(11):174, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
174.pdf.

Anonymous:1870:WEa

[276] Anonymous. Wood engravings. Scientific American, 22(11):175, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-175.pdf.

Anonymous:1870:HTSb

[277] Anonymous. Hardening and tempering steel. Scientific American, 22(11):175, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
175a.pdf.

Anonymous:1870:IVI

[278] Anonymous. Inertia and vis inertiæ. Scientific American, 22(11):175, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n11/pdf/scientificamerican03121870-175b.pdf.

Anonymous:1870:IP

[279] Anonymous. An inexplicable power. Scientific American, 22 (11):175–176, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
175c.pdf.

Anonymous:1870:LWO

[280] Anonymous. Lockage waste on our canals. Scientific American, 22(11):176, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
176.pdf.

Anonymous:1870:RWC

[281] Anonymous. Restrictions on the weighing of coal. Scientific American, 22(11):176, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
176a.pdf.

Anonymous:1870:WWTb

[282] Anonymous. The water wheel tests at Lowell. Scientific American, 22(11):176–177, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
176b.pdf.

Anonymous:1870:ACj

[283] Anonymous. Answers to correspondents. Scientific American, 22(11):177, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
177.pdf.

Anonymous:1870:RAFk

[284] Anonymous. Recent American and foreign patents. Scientific American, 22(11):178, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
178.pdf.

Anonymous:1870:OLPk

[285] Anonymous. Official list of patents. Scientific American, 22 (11):179–180, March 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n11/pdf/scientificamerican03121870-
179.pdf.

Anonymous:1870:Cl

[286] Anonymous. Correspondence. Scientific American, 22(12):188–190, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n12/pdf/scientificamerican03191870-188.pdf.

Anonymous:1870:BPl

[287] Anonymous. Business and personal. Scientific American, 22(12):193, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n12/pdf/scientificamerican03191870-193b.pdf.

Anonymous:1870:BNBj

[288] Anonymous. Books: New books and publications. Scientific American, 22(12):196, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
196.pdf.

Anonymous:1870:PIF

[289] Anonymous. The pedespeed, improved fan blower. Scientific American, 22(12):183–184, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
183.pdf.

Anonymous:1870:VSM

[290] Anonymous. Visit to a steel manufactory. Scientific American, 22 (12):184–185, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
184.pdf.

Duncan:1870:IRA

[291] Samuel A. Duncan. An interesting review of the American patent system. Scientific American, 22(12):185, March 19, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n12/pdf/scientificamerican03191870-185.pdf.

Colton:1870:PPF

[292] Henry E. Colton. Products of the pine forest. Scientific American, 22(12):186–187, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
186.pdf.

Anonymous:1870:WSG

[293] Anonymous. Woods’ saw-gumming and sharpening machine. Scientific American, 22(12):187, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
187.pdf.

Anonymous:1870:SITa

[294] Anonymous. Snyder’s improved turbine wheel. Scientific American, 22(12):187–188, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
187a.pdf.

Anonymous:1870:PSY

[295] Anonymous. Proposed sale of the yosemite valley. Scientific American, 22(12):191, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
191.pdf.

Anonymous:1870:COA

[296] Anonymous. Co-operation among inventors. Scientific American, 22(12):191, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
191a.pdf.

Anonymous:1870:USAa

[297] Anonymous. The use of salt in agriculture. Scientific American, 22(12):191, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
191b.pdf.

Anonymous:1870:SAG

[298] Anonymous. Something about gases. Scientific American, 22 (12):191–192, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
191c.pdf.

Anonymous:1870:NMMa

[299] Anonymous. New mechanical movements. Scientific American, 22(12):192, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
192.pdf.

Anonymous:1870:PPP

[300] Anonymous. Proposed plan for publishing Patent Office specifications and drawings. Scientific American, 22(12):192, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n12/pdf/scientificamerican03191870-192a.pdf.

Anonymous:1870:CEPa

[301] Anonymous. Congress extending patents. Scientific American, 22 (12):192–193, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
192b.pdf.

Anonymous:1870:AEPh

[302] Anonymous. Applications for extension of patents. Scientific American, 22(12):193, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
193.pdf.

Anonymous:1870:ACk

[303] Anonymous. Answers to correspondents. Scientific American, 22(12):193, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
193a.pdf.

Anonymous:1870:RAFl

[304] Anonymous. Recent American and foreign patents. Scientific American, 22(12):193–194, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
193c.pdf.

Anonymous:1870:PAI

[305] Anonymous. Progress of American invention in Europe. Scientific American, 22(12):194, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
194.pdf.

Anonymous:1870:OLPl

[306] Anonymous. Official list of patents. Scientific American, 22 (12):194–195, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
194a.pdf.

Anonymous:1870:RFT

[307] Anonymous. Rotating and fixed turrets. Scientific American, 22 (12):195–196, March 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n12/pdf/scientificamerican03191870-
195.pdf.

Anonymous:1870:Cm

[308] Anonymous. Correspondence. Scientific American, 22(13):203–204, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-203b.pdf.

Anonymous:1870:BPm

[309] Anonymous. Business and personal. Scientific American, 22(13):209, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-209.pdf.

Anonymous:1870:BNBk

[310] Anonymous. Books: New books and publications. Scientific American, 22(13):211–212, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
211.pdf.

Anonymous:1870:IMDa

[311] Anonymous. Improvement in machines for dressing millstones. Scientific American, 22(13):199, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
199.pdf.

Joy:1870:Tb

[312] Charles A. Joy. On tungsten. Scientific American, 22(13):200, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-200.pdf.

Emery:1870:CWS

[313] Chas. E. Emery. Circulation of water in steam boilers. Scientific American, 22(13):200–201, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
200a.pdf.

Anonymous:1870:BIa

[314] Anonymous. Boiler incrustations. Scientific American, 22(13):201, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-201.pdf.

Anonymous:1870:HP

[315] Anonymous. The home of the perfumes. Scientific American, 22(13):201, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
201a.pdf.

Anonymous:1870:NGO

[316] Anonymous. The new grand opera house, Paris. Scientific American, 22(13):202, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
202.pdf.

Anonymous:1870:HMD

[317] Anonymous. A home-made differential thermometer. Scientific American, 22(13):203, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
203.pdf.

Anonymous:1870:OHR

[318] Anonymous. One-horse road steamer to carry two persons. Scientific American, 22(13):203, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
203a.pdf.

Anonymous:1870:PA

[319] Anonymous. Profits of advertising. Scientific American, 22(13):204, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-204.pdf.

Anonymous:1870:RSEa

[320] Anonymous. Recent speculations on the Earth’s origin. Scientific American, 22(13):204–206, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
204a.pdf.

Anonymous:1870:WIL

[321] Anonymous. Williams’ improvement in locking nuts. Scientific American, 22(13):206, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
206.pdf.

Anonymous:1870:ONP

[322] Anonymous. Our new paper currency. Scientific American, 22(13):207, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-207.pdf.

Anonymous:1870:SRM

[323] Anonymous. Supplementary report of the Massachusetts Railroad Commissioners on steel rails. Scientific American, 22(13):207, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-207a.pdf.

Anonymous:1870:PHU

[324] Anonymous. Proportion of heat utilized in steam engines. Scientific American, 22(13):207–208, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
207b.pdf.

Anonymous:1870:EPG

[325] Anonymous. Every pane of glass is a hole to heat. Scientific American, 22(13):208, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
208.pdf.

Anonymous:1870:EPa

[326] Anonymous. Engineering projects. Scientific American, 22(13):208, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-208a.pdf.

Anonymous:1870:SId

[327] Anonymous. Scientific intelligence. Scientific American, 22(13):208, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n13/pdf/scientificamerican03261870-208b.pdf.

Anonymous:1870:ACl

[328] Anonymous. Answers to correspondents. Scientific American, 22(13):209, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
209a.pdf.

Anonymous:1870:RAFm

[329] Anonymous. Recent American and foreign patents. Scientific American, 22(13):209–210, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
209b.pdf.

Anonymous:1870:OLPm

[330] Anonymous. Official list of patents. Scientific American, 22 (13):210–211, March 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n13/pdf/scientificamerican03261870-
210.pdf.

Anonymous:1870:Cn

[331] Anonymous. Correspondence. Scientific American, 22(14):220–222, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-220.pdf.

Anonymous:1870:BPn

[332] Anonymous. Business and personal. Scientific American, 22(14):226, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-226.pdf.

Anonymous:1870:BNBl

[333] Anonymous. Books: New books and publications. Scientific American, 22(14):227, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
227.pdf.

Anonymous:1870:IBC

[334] Anonymous. Improved bolt cutter, improved self-connecting car coupling, and more. Scientific American, 22(14):215–216, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-215.pdf.

Anonymous:1870:MFB

[335] Anonymous. The manufacture of fire bricks. Scientific American, 22(14):216–217, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
216.pdf.

Joy:1870:SAS

[336] Charles A. Joy. Something about silver. Scientific American, 22 (14):217, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
217.pdf.

Gordon:1870:G

[337] John R. Gordon. The guava. Scientific American, 22(14):217, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-217a.pdf.

Anonymous:1870:CNa

[338] Anonymous. Chemical news. Scientific American, 22(14):217–218, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-217b.pdf.

Anonymous:1870:WCb

[339] Anonymous. What are condiments? Scientific American, 22(14):218, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-218.pdf.

Orton:1870:NHG

[340] James Orton. Natural history — the great auk. Scientific American, 22(14):219, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
219.pdf.

Anonymous:1870:IID

[341] Anonymous. Improved implement for drawing nails. Scientific American, 22(14):219, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
219a.pdf.

Anonymous:1870:RT

[342] Anonymous. Rapid telegraphy. Scientific American, 22(14):222, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-222.pdf.

Anonymous:1870:ICB

[343] Anonymous. Improved coal-breaking machine. Scientific American, 22(14):222, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
222a.pdf.

Anonymous:1870:YV

[344] Anonymous. The Yosemite Valley. Scientific American, 22(14):222, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-222b.pdf.

Anonymous:1870:PLG

[345] Anonymous. Portable laughing gas. Scientific American, 22(14):222, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-222c.pdf.

Anonymous:1870:DCW

[346] Anonymous. The Darien Canal — will it pay to build it? Scientific American, 22(14):223, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
223.pdf.

Anonymous:1870:ICS

[347] Anonymous. An invisible chain and a silent conflict. Scientific American, 22(14):223–224, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
223a.pdf.

Anonymous:1870:SAb

[348] Anonymous. The study of alloys. Scientific American, 22(14):224, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-224.pdf.

Anonymous:1870:BA

[349] Anonymous. Butter and its adulterations. Scientific American, 22(14):224, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
224a.pdf.

Anonymous:1870:HFBb

[350] Anonymous. How fire bricks are made. Scientific American, 22(14):224, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-224b.pdf.

Anonymous:1870:WMc

[351] Anonymous. The wind month. Scientific American, 22(14):224, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-224c.pdf.

Anonymous:1870:EUT

[352] Anonymous. Economy in the use of time. Scientific American, 22(14):224–225, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
224d.pdf.

Anonymous:1870:HPOb

[353] Anonymous. H. S. Patent Office. Scientific American, 22(14):225–226, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n14/pdf/scientificamerican04021870-225.pdf.

Anonymous:1870:ACm

[354] Anonymous. Answers to correspondents. Scientific American, 22 (14):226, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
226a.pdf.

Anonymous:1870:RAFn

[355] Anonymous. Recent American and foreign patents. Scientific American, 22(14):226–227, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
226b.pdf.

Anonymous:1870:OLPn

[356] Anonymous. Official list of patents. Scientific American, 22 (14):227–228, April 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n14/pdf/scientificamerican04021870-
227a.pdf.

Anonymous:1870:Co

[357] Anonymous. Correspondence. Scientific American, 22(15):236–238, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-236.pdf.

Anonymous:1870:BPo

[358] Anonymous. Business and personal. Scientific American, 22(15):242, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-242a.pdf.

Anonymous:1870:MBC

[359] Anonymous. Mr. Bessemer’s conservatory. Scientific American, 22(15):231–232, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
231.pdf.

Anonymous:1870:BBT

[360] Anonymous. Bells and Bell towers. Scientific American, 22(15):232–233, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-232.pdf.

Anonymous:1870:ISD

[361] Anonymous. An interesting sketch of the distinguished aeronaut, John La Mountain. Scientific American, 22(15):233, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n15/pdf/scientificamerican04091870-233.pdf.

Brigham:1870:MD

[362] William T. Brigham. Mineral deposits. Scientific American, 22(15):233, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-233a.pdf.

Anonymous:1870:HLWa

[363] Anonymous. Hill’s locking whip socket. Scientific American, 22(15):234, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-234.pdf.

Anonymous:1870:HWE

[364] Anonymous. Hamilton’s water elevator. Scientific American, 22 (15):234, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
234a.pdf.

Anonymous:1870:DIG

[365] Anonymous. Denn’s improved gridiron and cake Baker. Scientific American, 22(15):234, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
234b.pdf.

Anonymous:1870:BFI

[366] Anonymous. The bitricycle, a French invention. Scientific American, 22(15):235, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
235.pdf.

Anonymous:1870:BPU

[367] Anonymous. The Bessemer process under pressure. Scientific American, 22(15):235–236, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
235a.pdf.

Anonymous:1870:GRI

[368] Anonymous. Galmann and Ruhe’s improved joist protector. Scientific American, 22(15):238, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
238.pdf.

Anonymous:1870:OSB

[369] Anonymous. Obituary — Seth Boyden. Scientific American, 22(15):238, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-238a.pdf.

Anonymous:1870:GBB

[370] Anonymous. Glazed bricks for building purposes. Scientific American, 22(15):239, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
239.pdf.

Anonymous:1870:DWC

[371] Anonymous. Dualin — what is claimed for it. Scientific American, 22(15):239, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
239a.pdf.

Anonymous:1870:NE

[372] Anonymous. Nature’s elevator. Scientific American, 22(15):239, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-239b.pdf.

Anonymous:1870:MWP

[373] Anonymous. The measurement of water power. Scientific American, 22(15):239, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
239c.pdf.

Anonymous:1870:MFc

[374] Anonymous. Michael Faraday. Scientific American, 22(15):240–241, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-240.pdf.

Anonymous:1870:PODc

[375] Anonymous. Patent Office decisions. Scientific American, 22(15):241, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-241.pdf.

Anonymous:1870:HPOc

[376] Anonymous. H. S. Patent Office. Scientific American, 22(15):242, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n15/pdf/scientificamerican04091870-242.pdf.

Anonymous:1870:ACn

[377] Anonymous. Answers to correspondents. Scientific American, 22 (15):242, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
242b.pdf.

Anonymous:1870:RAFo

[378] Anonymous. Recent American and foreign patents. Scientific American, 22(15):242–243, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
242c.pdf.

Anonymous:1870:AEPi

[379] Anonymous. Applications for extension of patents. Scientific American, 22(15):243, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
243.pdf.

Anonymous:1870:OLPo

[380] Anonymous. Official list of patents. Scientific American, 22 (15):243–244, April 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n15/pdf/scientificamerican04091870-
243a.pdf.

Anonymous:1870:Cp

[381] Anonymous. Correspondence. Scientific American, 22(16):251, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-251b.pdf.

Anonymous:1870:BNBm

[382] Anonymous. Books: New books and publications. Scientific American, 22(16):257, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
257a.pdf.

Anonymous:1870:BPp

[383] Anonymous. Business and personal. Scientific American, 22(16):258, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-258.pdf.

Anonymous:1870:CSV

[384] Anonymous. Combined safety valve and low-water alarm. Scientific American, 22(16):247, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
247.pdf.

Anonymous:1870:PRCa

[385] Anonymous. Patent right and copyright. Scientific American, 22 (16):247–248, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
247a.pdf.

Anonymous:1870:SFA

[386] Anonymous. Fur seal fishery in Alaska. Scientific American, 22 (16):248–249, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
248.pdf.

Child:1870:DFS

[387] Duff Child. Deformed feet and shoes. Scientific American, 22(16):249, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-249.pdf.

Anonymous:1870:CRE

[388] Anonymous. Comparison of the refrigerative effects of the carre apparatus and common ice. Scientific American, 22(16):249, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-249a.pdf.

Anonymous:1870:Bb

[389] Anonymous. Barnacles. Scientific American, 22(16):250, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-250.pdf.

Anonymous:1870:NSU

[390] Anonymous. The narwhal or sea unicorn. Scientific American, 22(16):250, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
250a.pdf.

Anonymous:1870:SIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-251.pdf.

Anonymous:1870:WRP

[392] Anonymous. Wormood’s revolving pay box. Scientific American, 22(16):251, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
251a.pdf.

Anonymous:1870:VM

[393] Anonymous. Vertical multiplier. Scientific American, 22(16):251–253, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-251c.pdf.

Anonymous:1870:SPb

[394] Anonymous. Screw piles. Scientific American, 22(16):253–254, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-253.pdf.

Anonymous:1870:CDL

[395] Anonymous. Conklin’s detachable lip for bowls, basins, etc. Scientific American, 22(16):254, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
254.pdf.

Anonymous:1870:IEL

[396] Anonymous. International Exhibition of 1871, in London. Scientific American, 22(16):255, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
255.pdf.

Anonymous:1870:PES

[397] Anonymous. Power and efficiency of screws. Scientific American, 22(16):255, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
255a.pdf.

Anonymous:1870:TDC

[398] Anonymous. Telo-dynamic cables. Scientific American, 22(16):255, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-255b.pdf.

Anonymous:1870:BEa

[399] Anonymous. Boiler explosions. Scientific American, 22(16):255–256, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-255c.pdf.

Anonymous:1870:CCa

[400] Anonymous. Charlatans and charlatanism. Scientific American, 22(16):256, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
256.pdf.

Anonymous:1870:ETL

[401] Anonymous. Experimental test of the laws of the deflection of beams. Scientific American, 22(16):256, April 16, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n16/pdf/scientificamerican04161870-256a.pdf.

Anonymous:1870:RNY

[402] Anonymous. Report of the New York Commissioners of Fisheries. Scientific American, 22(16):256–257, April 16, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n16/pdf/scientificamerican04161870-256b.pdf.

Anonymous:1870:ESb

[403] Anonymous. Editorial summary. Scientific American, 22(16):257, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-257.pdf.

Anonymous:1870:PODd

[404] Anonymous. Patent Office decision. Scientific American, 22(16):257, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n16/pdf/scientificamerican04161870-257b.pdf.

Anonymous:1870:USCd

[405] Anonymous. United States Circuit Court at New Orleans. Scientific American, 22(16):257–258, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
257c.pdf.

Anonymous:1870:ACo

[406] Anonymous. Answers to correspondents. Scientific American, 22 (16):258, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
258a.pdf.

Anonymous:1870:RAFp

[407] Anonymous. Recent American and foreign patents. Scientific American, 22(16):258–259, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
258b.pdf.

Anonymous:1870:OLPp

[408] Anonymous. Official list of patents. Scientific American, 22 (16):259–260, April 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n16/pdf/scientificamerican04161870-
259.pdf.

Anonymous:1870:Cq

[409] Anonymous. Correspondence. Scientific American, 22(17):268–269, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-268.pdf.

Anonymous:1870:BNBn

[410] Anonymous. Books: New books and publications. Scientific American, 22(17):273–274, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
273b.pdf.

Anonymous:1870:BPq

[411] Anonymous. Business and personal. Scientific American, 22(17):274, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-274a.pdf.

Anonymous:1870:BPr

[412] Anonymous. The ballard pavement. Scientific American, 22(17):263, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-263.pdf.

Knapp:1870:LVMa

[413] H. Knapp. Light and vision — more especially how light is converted into thought. Scientific American, 22(17):264, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-264.pdf.

Anonymous:1870:EHG

[414] Anonymous. Experiments on the head of a guillotined patient. Scientific American, 22(17):264–265, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
264a.pdf.

Anonymous:1870:TCP

[415] Anonymous. A trip over the Central Pacific Railroad. Scientific American, 22(17):265–266, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
265.pdf.

Anonymous:1870:BIb

[416] Anonymous. Boring insects. Scientific American, 22(17):266–267, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-266.pdf.

Anonymous:1870:SBB

[417] Anonymous. Submerged buoys as a basis for structures in deep water. Scientific American, 22(17):267, April 23, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n17/pdf/scientificamerican04231870-267.pdf.

Anonymous:1870:IPS

[418] Anonymous. Improved packing steam drying cylinders on paper machines. Scientific American, 22(17):267, April 23, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n17/pdf/scientificamerican04231870-267a.pdf.

Anonymous:1870:ESc

[419] Anonymous. Editorial summary. Scientific American, 22(17):269, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-269.pdf.

Anonymous:1870:SICa

[420] Anonymous. Stockwell’s improved combination tool. Scientific American, 22(17):270, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
270.pdf.

Anonymous:1870:KPR

[421] Anonymous. Kay’s patent reversible slide valve and swing grate. Scientific American, 22(17):270, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
270a.pdf.

Anonymous:1870:AAT

[422] Anonymous. An Anglo–American Turkish bath. Scientific American, 22(17):271, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
271.pdf.

Anonymous:1870:SBIa

[423] Anonymous. Steam boiler inspection. Scientific American, 22(17):271, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-271a.pdf.

Anonymous:1870:PL

[424] Anonymous. Principle of the lever. Scientific American, 22(17):271, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-271b.pdf.

Anonymous:1870:WSE

[425] Anonymous. The West Side Elevated Railway. Scientific American, 22(17):272, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
272.pdf.

Anonymous:1870:PO

[426] Anonymous. The Paris Observatory. Scientific American, 22(17):272, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-272a.pdf.

Anonymous:1870:SYM

[427] Anonymous. Shut your mouth. Scientific American, 22(17):272–273, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-272b.pdf.

Anonymous:1870:CTO

[428] Anonymous. Comparison of turbines with other water wheels. Scientific American, 22(17):273, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
273.pdf.

Anonymous:1870:WHC

[429] Anonymous. Where and how corks are cut. Scientific American, 22(17):273, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
273a.pdf.

Anonymous:1870:ACp

[430] Anonymous. Answers to correspondents. Scientific American, 22 (17):274, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
274.pdf.

Anonymous:1870:RAFq

[431] Anonymous. Recent American and foreign patents. Scientific American, 22(17):274–275, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
274b.pdf.

Anonymous:1870:USS

[432] Anonymous. United States Supreme Court. Scientific American, 22(17):275, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
275.pdf.

Anonymous:1870:OLPq

[433] Anonymous. Official list of patents. Scientific American, 22 (17):275–276, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n17/pdf/scientificamerican04231870-
275a.pdf.

Anonymous:1870:HPOd

[434] Anonymous. H. S. Patent Office. Scientific American, 22(17):276, April 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n17/pdf/scientificamerican04231870-276.pdf.

Anonymous:1870:Cr

[435] Anonymous. Correspondence. Scientific American, 22(18):284, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-284.pdf.

Anonymous:1870:BPs

[436] Anonymous. Business and personal. Scientific American, 22(18):290, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-290a.pdf.

Anonymous:1870:BNBo

[437] Anonymous. Books: New books and publications. Scientific American, 22(18):292, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
292.pdf.

Anonymous:1870:ISF

[438] Anonymous. Improved stall floor for horses, to prevent decay of shingles, and more. Scientific American, 22(18):279, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-279.pdf.

Knapp:1870:LVMb

[439] H. Knapp. Light and vision — more especially how light is converted into thought. Scientific American, 22(18):280–281, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-280.pdf.

Anonymous:1870:MFF

[440] Anonymous. Mammalian fossils — a facetious reviewer. Scientific American, 22(18):281, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
281.pdf.

Cantini:1870:CP

[441] Rx J. Cantini. The cocoa palm. Scientific American, 22(18):281, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-281a.pdf.

Anonymous:1870:DUC

[442] Anonymous. Dugdale’s universal clothes washer. Scientific American, 22(18):282–284, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
282.pdf.

Anonymous:1870:GGM

[443] Anonymous. Gentlemen glass-makers. Scientific American, 22(18):285, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-285.pdf.

Anonymous:1870:LIN

[444] Anonymous. Langhorne’s improved nut-lock. Scientific American, 22(18):285–286, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
285a.pdf.

Anonymous:1870:SVE

[445] Anonymous. Safety vs. economy in the construction of steam boilers. Scientific American, 22(18):287, April 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n18/pdf/scientificamerican04301870-287.pdf.

Anonymous:1870:SIf

[446] Anonymous. The study of impracticabilities. Scientific American, 22(18):287, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
287a.pdf.

Anonymous:1870:NMMb

[447] Anonymous. New mechanical movements. Scientific American, 22(18):287, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
287b.pdf.

Anonymous:1870:MH

[448] Anonymous. Metallic hydrogen. Scientific American, 22(18):287, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-287c.pdf.

Anonymous:1870:UFSa

[449] Anonymous. Uses of fluor spar. Scientific American, 22(18):287–288, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-287d.pdf.

Anonymous:1870:IUC

[450] Anonymous. The increased use of cold-rolled shafting. Scientific American, 22(18):288, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
288.pdf.

Anonymous:1870:SIg

[451] Anonymous. Scientific intelligence. Scientific American, 22(18):288, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-288a.pdf.

Anonymous:1870:MEU

[452] Anonymous. Morin’s experiments upon the punching of metals. Scientific American, 22(18):288, April 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n18/pdf/scientificamerican04301870-288b.pdf.

Anonymous:1870:GUD

[453] Anonymous. The great union depot on fourth avenue, New York. Scientific American, 22(18):288–289, April 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n18/pdf/scientificamerican04301870-288c.pdf.

Anonymous:1870:BAP

[454] Anonymous. Bill to amend the patent laws now pending before Congress. Scientific American, 22(18):289, April 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n18/pdf/scientificamerican04301870-289.pdf.

Anonymous:1870:EF

[455] Anonymous. Electric forces. Scientific American, 22(18):289, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-289a.pdf.

Anonymous:1870:IMDb

[456] Anonymous. Instrument for measuring the deflection of girders. Scientific American, 22(18):289–290, April 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n18/pdf/scientificamerican04301870-289b.pdf.

Anonymous:1870:ESd

[457] Anonymous. Editorial summary. Scientific American, 22(18):290, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n18/pdf/scientificamerican04301870-290.pdf.

Anonymous:1870:ACq

[458] Anonymous. Answers to correspondents. Scientific American, 22 (18):290, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
290b.pdf.

Anonymous:1870:RAFr

[459] Anonymous. Recent American and foreign patents. Scientific American, 22(18):290–291, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
290c.pdf.

Anonymous:1870:OLPr

[460] Anonymous. Official list of patents. Scientific American, 22 (18):291–292, April 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n18/pdf/scientificamerican04301870-
291.pdf.

Anonymous:1870:Cs

[461] Anonymous. Correspondence. Scientific American, 22(20):315–318, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-315a.pdf.

Anonymous:1870:BPv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-322.pdf.

Anonymous:1870:IPE

[463] Anonymous. Improved portable engine for agricultural use. Scientific American, 22(20):311–312, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
311.pdf.

Anonymous:1870:Ma

[464] Anonymous. Meteors. Scientific American, 22(20):312–314, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-312.pdf.

Anonymous:1870:MIT

[465] Anonymous. McNeil’s improved tobacco pipe. Scientific American, 22(20):314–315, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
314.pdf.

Anonymous:1870:OZC

[466] Anonymous. Obituary — Zerah Colburn, engineer, and leading writer of engineering papers. Scientific American, 22(20):315, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-315.pdf.

Anonymous:1870:AMN

[467] Anonymous. The American Museum of Natural History. Scientific American, 22(20):319, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
319.pdf.

Anonymous:1870:DD

[468] Anonymous. The dangers of drugs. Scientific American, 22(20):319, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-319a.pdf.

Anonymous:1870:SBIb

[469] Anonymous. Steam boiler inspection. Scientific American, 22(20):319, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-319b.pdf.

Anonymous:1870:MRa

[470] Anonymous. Mechanical recreations. Scientific American, 22 (20):319–320, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
319c.pdf.

Anonymous:1870:CAAa

[471] Anonymous. Curious associations among animals. Scientific American, 22(20):320, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
320.pdf.

Anonymous:1870:CZ

[472] Anonymous. Chemistry of zirconium. Scientific American, 22(20):320, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n20/pdf/scientificamerican05141870-320a.pdf.

Anonymous:1870:SIh

[473] Anonymous. Scientific intelligence. Scientific American, 22(20): 320–321, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
320b.pdf.

Anonymous:1870:USD

[474] Anonymous. United States District Court — Southern District before Judge Blatchford patent for making veneers. Scientific American, 22(20):321, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
321.pdf.

Anonymous:1870:AEPj

[475] Anonymous. Applications for extension of patents. Scientific American, 22(20):321, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
321a.pdf.

Anonymous:1870:RAFs

[476] Anonymous. Recent American and foreign patents. Scientific American, 22(20):321–322, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
321b.pdf.

Anonymous:1870:ACr

[477] Anonymous. Answers to correspondents. Scientific American, 22 (20):323, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
323.pdf.

Anonymous:1870:OLPs

[478] Anonymous. Official list of patents. Scientific American, 22 (20):323–324, May 14, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n20/pdf/scientificamerican05141870-
323a.pdf.

Anonymous:1870:Cu

[479] Anonymous. Correspondence. Scientific American, 22(21):332, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-332.pdf.

Anonymous:1870:BPw

[480] Anonymous. Business and personal. Scientific American, 22(21):338, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-338a.pdf.

Anonymous:1870:BNBp

[481] Anonymous. Books: New books and publications. Scientific American, 22(21):340, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
340.pdf.

Anonymous:1870:IMSa

[482] Anonymous. Improvement in marine staterooms and furniture, presence of mind, and more. Scientific American, 22(21):327, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-327.pdf.

Anonymous:1870:CFM

[483] Anonymous. Codfish as food for man, beast, and vegetable. Scientific American, 22(21):328–329, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
328.pdf.

Anonymous:1870:WRT

[484] Anonymous. The wire rope tramway at Brighton England. Scientific American, 22(21):329, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
329.pdf.

Winton:1870:FPL

[485] John G. Winton. Furrowing and pitting in locomotive boilers. Scientific American, 22(21):329–330, May 21, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n21/pdf/scientificamerican05211870-329a.pdf.

Day:1870:BWP

[486] Edward C. H. Day. A beetle on the war path. Scientific American, 22(21):331, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
331.pdf.

Joy:1870:R

[487] Charles A. Joy. On ruthenium. Scientific American, 22(21):332–334, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-332a.pdf.

Anonymous:1870:RTF

[488] Anonymous. Rubber tip for furniture legs. Scientific American, 22(21):334, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
334.pdf.

Anonymous:1870:VS

[489] Anonymous. The value of science. Scientific American, 22(21):335, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-335.pdf.

Anonymous:1870:HMa

[490] Anonymous. Is hydrogen a metal? Scientific American, 22(21):335, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-335a.pdf.

Anonymous:1870:HPL

[491] Anonymous. How people live too fast. Scientific American, 22(21):335, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-335b.pdf.

Anonymous:1870:SMc

[492] Anonymous. The steam man. Scientific American, 22(21):335–336, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-335c.pdf.

Anonymous:1870:WCP

[493] Anonymous. Washington considered as a place for an exhibition. Scientific American, 22(21):336, May 21, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n21/pdf/scientificamerican05211870-336.pdf.

Anonymous:1870:AS

[494] Anonymous. Artificial stone. Scientific American, 22(21):336, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-336a.pdf.

Anonymous:1870:YR

[495] Anonymous. The yacht races. Scientific American, 22(21):336–337, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n21/pdf/scientificamerican05211870-336b.pdf.

Anonymous:1870:ACs

[496] Anonymous. Answers to correspondents. Scientific American, 22 (21):338, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
338.pdf.

Anonymous:1870:RAFt

[497] Anonymous. Recent American and foreign patents. Scientific American, 22(21):338, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
338b.pdf.

Anonymous:1870:OLPt

[498] Anonymous. Official list of patents. Scientific American, 22 (21):339–340, May 21, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n21/pdf/scientificamerican05211870-
339.pdf.

Anonymous:1870:Cv

[499] Anonymous. Correspondence. Scientific American, 22(22):348, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-348.pdf.

Anonymous:1870:BPx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-354a.pdf.

Anonymous:1870:ITP

[501] Anonymous. Improved tobacco press, White brass, and more. Scientific American, 22(22):343, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
343.pdf.

Adams:1870:NFM

[502] W. Bridges Adams. The nature and food of man. Scientific American, 22(22):344–346, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
344.pdf.

Anonymous:1870:CMA

[503] Anonymous. The clothes-moth and its allies. Scientific American, 22(22):346–348, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
346.pdf.

Widemann:1870:RAE

[504] C. Widemann. Recent advances in electric science. Scientific American, 22(22):348, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
348a.pdf.

Anonymous:1870:HSA

[505] Anonymous. The history, science, and art of whipping. Scientific American, 22(22):349–351, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
349.pdf.

Anonymous:1870:POAa

[506] Anonymous. Patent Office affairs. Scientific American, 22(22):351, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-351.pdf.

Anonymous:1870:CTQ

[507] Anonymous. The city transit question. Scientific American, 22(22):351, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-351a.pdf.

Anonymous:1870:ACAa

[508] Anonymous. Advantages of compressed asphalte for road-ways, and the production of solid surfaces for other purposes. Scientific American, 22(22):351–352, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
351b.pdf.

Anonymous:1870:PED

[509] Anonymous. Photographic exhibition and Dr. Vogel. Scientific American, 22(22):352, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
352.pdf.

Anonymous:1870:DDP

[510] Anonymous. Death of a distinguished professor. Scientific American, 22(22):352, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
352a.pdf.

Anonymous:1870:CEN

[511] Anonymous. Captain Ericsson’s new method of submarine attack. Scientific American, 22(22):352, May 28, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n22/pdf/scientificamerican05281870-352b.pdf.

Anonymous:1870:NYLa

[512] Anonymous. New York lyceum of natural history. Scientific American, 22(22):352, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
352c.pdf.

Anonymous:1870:MGS

[513] Anonymous. Manufacture of grape sugar. Scientific American, 22(22):352, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
352d.pdf.

Anonymous:1870:NPF

[514] Anonymous. A new process for fixing images by exposure to heat. Scientific American, 22(22):352–353, May 28, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n22/pdf/scientificamerican05281870-352e.pdf.

Anonymous:1870:SIi

[515] Anonymous. Scientific intelligence. Scientific American, 22(22):353, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-353.pdf.

Anonymous:1870:CHWb

[516] Anonymous. A clever hoax — who is Dr. Ivan Slavonski? Scientific American, 22(22):353, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
353a.pdf.

Anonymous:1870:ESf

[517] Anonymous. Editorial summary. Scientific American, 22(22):353–354, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n22/pdf/scientificamerican05281870-353b.pdf.

Anonymous:1870:ACt

[518] Anonymous. Answers to correspondents. Scientific American, 22 (22):354, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
354.pdf.

Anonymous:1870:RAFu

[519] Anonymous. Recent American and foreign patents. Scientific American, 22(22):354–355, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
354b.pdf.

Anonymous:1870:OLPu

[520] Anonymous. Official list of patents. Scientific American, 22 (22):355–356, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
355.pdf.

Anonymous:1870:AEPk

[521] Anonymous. Applications for extension of patents. Scientific American, 22(22):356, May 28, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n22/pdf/scientificamerican05281870-
356.pdf.

Anonymous:1870:Cw

[522] Anonymous. Correspondence. Scientific American, 22(23):364, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-364.pdf.

Anonymous:1870:BPy

[523] Anonymous. Business and personal. Scientific American, 22(23):370, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-370.pdf.

Anonymous:1870:ATP

[524] Anonymous. Air target pistol, improved peat machine. Scientific American, 22(23):359, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
359.pdf.

Anonymous:1870:WEb

[525] Anonymous. What is energy? Scientific American, 22(23):360, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-360.pdf.

Anonymous:1870:MPM

[526] Anonymous. Meat preserving in Melbourne, Australia. Scientific American, 22(23):361, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
361.pdf.

Day:1870:CRa

[527] Edward C. H. Day. The cockchafer and its ravages. Scientific American, 22(23):362–363, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
362.pdf.

Anonymous:1870:KSL

[528] Anonymous. Kemlo’s spring-lock fishhooks. Scientific American, 22(23):363–364, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
363.pdf.

Anonymous:1870:NYLb

[529] Anonymous. New York lyceum of natural history. Scientific American, 22(23):365, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
365.pdf.

Anonymous:1870:IME

[530] Anonymous. Italian Maritime Exposition. Scientific American, 22(23):365, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
365a.pdf.

Anonymous:1870:NSS

[531] Anonymous. A new spring at Saratoga. Scientific American, 22 (23):365–366, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
365b.pdf.

Anonymous:1870:ISHb

[532] Anonymous. Improved shovel handle. Scientific American, 22(23):366, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-366.pdf.

Anonymous:1870:DCE

[533] Anonymous. Deterioration of coal by exposure to the air. Scientific American, 22(23):367, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
367.pdf.

Anonymous:1870:UWR

[534] Anonymous. The use of wire rope in civil and mechanical engineering. Scientific American, 22(23):367, June 4, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n23/pdf/scientificamerican06041870-367a.pdf.

Anonymous:1870:IB

[535] Anonymous. The impact of bodies. Scientific American, 22(23):367, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-367b.pdf.

Anonymous:1870:PCT

[536] Anonymous. The problem of city transit. Scientific American, 22(23):367–369, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
367c.pdf.

Anonymous:1870:POAb

[537] Anonymous. Patent Office affairs. Scientific American, 22(23):369, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n23/pdf/scientificamerican06041870-369.pdf.

Anonymous:1870:BRSa

[538] Anonymous. Beet root sugar in the United States. Scientific American, 22(23):369, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
369a.pdf.

Anonymous:1870:AEPl

[539] Anonymous. Applications for extension of patents. Scientific American, 22(23):369–370, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
369b.pdf.

Anonymous:1870:ACu

[540] Anonymous. Answers to correspondents. Scientific American, 22 (23):370, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
370a.pdf.

Anonymous:1870:RAFv

[541] Anonymous. Recent American and foreign patents. Scientific American, 22(23):370–371, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
370b.pdf.

Anonymous:1870:OLPv

[542] Anonymous. Official list of patents. Scientific American, 22 (23):371–372, June 4, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n23/pdf/scientificamerican06041870-
371.pdf.

Anonymous:1870:Cx

[543] Anonymous. Correspondence. Scientific American, 22(24):380, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-380.pdf.

Anonymous:1870:BPz

[544] Anonymous. Business and personal. Scientific American, 22(24):386, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-386a.pdf.

Anonymous:1870:BNI

[545] Anonymous. Beck’s new improved self-adjusting vise. Scientific American, 22(24):375–376, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
375.pdf.

Anonymous:1870:CLW

[546] Anonymous. On the conditions and limits which govern the proportions of rotary fans. Scientific American, 22(24):376–377, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-376.pdf.

Day:1870:CRb

[547] Edward C. H. Day. The cockchafer and its ravages. Scientific American, 22(24):378, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
378.pdf.

Anonymous:1870:RBb

[548] Anonymous. The roller Barrow. Scientific American, 22(24):378–380, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-378a.pdf.

Gordon:1870:CC

[549] John R. Gordon. A chapter on cockroaches. Scientific American, 22(24):381, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
381.pdf.

Wise:1870:BV

[550] John Wise. Balloon varnishes. Scientific American, 22(24):381, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-381a.pdf.

Bolton:1870:SPF

[551] H. Carrington Bolton. Scientific periodicals found in New York libraries. Scientific American, 22(24):381–382, June 11, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n24/pdf/scientificamerican06111870-381b.pdf.

Anonymous:1870:MMA

[552] Anonymous. The metropolitan museum of art. Scientific American, 22(24):383, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
383.pdf.

Anonymous:1870:ORW

[553] Anonymous. The origin of the reluctance with which mankind receive new ideas and inventions. Scientific American, 22(24):383, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-383a.pdf.

Anonymous:1870:NSR

[554] Anonymous. New sugar refining process. Scientific American, 22 (24):383, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
383b.pdf.

Anonymous:1870:QCT

[555] Anonymous. The question of city transit. Scientific American, 22(24):384, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
384.pdf.

Anonymous:1870:SAH

[556] Anonymous. Something about horseshoes. Scientific American, 22(24):384, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
384a.pdf.

Anonymous:1870:RFI

[557] Anonymous. Recent foreign improvements. Scientific American, 22(24):384, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
384b.pdf.

Anonymous:1870:SIj

[558] Anonymous. Scientific intelligence. Scientific American, 22(24): 384–385, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
384c.pdf.

Anonymous:1870:GSF

[559] Anonymous. Georgia State Fair. Scientific American, 22(24):385, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-385.pdf.

Anonymous:1870:PTF

[560] Anonymous. A pneumatic tube four hundred miles long. Scientific American, 22(24):385, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
385a.pdf.

Anonymous:1870:NEA

[561] Anonymous. National exhibition of the Argentine Republic. Scientific American, 22(24):385, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
385b.pdf.

Anonymous:1870:RPD

[562] Anonymous. Reduction of the public debt. Scientific American, 22(24):385, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
385c.pdf.

Anonymous:1870:POAc

[563] Anonymous. Patent Office affairs. Scientific American, 22(24):386, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n24/pdf/scientificamerican06111870-386.pdf.

Anonymous:1870:ACv

[564] Anonymous. Answers to correspondents. Scientific American, 22 (24):386, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
386b.pdf.

Anonymous:1870:RAFw

[565] Anonymous. Recent American and foreign patents. Scientific American, 22(24):386–387, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
386c.pdf.

Anonymous:1870:OLPw

[566] Anonymous. Official list of patents. Scientific American, 22 (24):387–388, June 11, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n24/pdf/scientificamerican06111870-
387.pdf.

Anonymous:1870:Cy

[567] Anonymous. Correspondence. Scientific American, 22(25):396–398, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-396.pdf.

Anonymous:1870:BPba

[568] Anonymous. Business and personal. Scientific American, 22(25):402, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-402.pdf.

Anonymous:1870:BNBq

[569] Anonymous. Books: New books and publications. Scientific American, 22(25):404, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
404.pdf.

Anonymous:1870:ABP

[570] Anonymous. The Alleghany Bridge at Pittsburgh, Pa, selecting durable and perishable timber. Scientific American, 22(25):391, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-391.pdf.

Anonymous:1870:SIk

[571] Anonymous. Scientific intelligence. Scientific American, 22(25): 392–393, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
392.pdf.

Day:1870:Sa

[572] Edward C. H. Day. The scorpion. Scientific American, 22(25):394–395, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-394.pdf.

Anonymous:1870:ARH

[573] Anonymous. Achenbach’s rein holder. Scientific American, 22 (25):395–396, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
395.pdf.

Anonymous:1870:HIC

[574] Anonymous. Hotte’s improved coffeepot. Scientific American, 22 (25):398, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
398.pdf.

Anonymous:1870:MMc

[575] Anonymous. Molecular motion. Scientific American, 22(25):399, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-399.pdf.

Anonymous:1870:PFb

[576] Anonymous. Potash from Feldspar. Scientific American, 22(25):399, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-399a.pdf.

Anonymous:1870:SMCa

[577] Anonymous. The School of Mines, Columbia College. Scientific American, 22(25):399, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
399b.pdf.

Anonymous:1870:CCH

[578] Anonymous. Conduction and convection of heat. Scientific American, 22(25):399–400, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
399c.pdf.

Anonymous:1870:TFC

[579] Anonymous. Are tin fruit cans a source of metallic poisoning? Scientific American, 22(25):400, June 18, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n25/pdf/scientificamerican06181870-400.pdf.

Anonymous:1870:MSb

[580] Anonymous. The mystery of sleep. Scientific American, 22(25):400, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-400a.pdf.

Anonymous:1870:PSE

[581] Anonymous. Primary school education. Scientific American, 22 (25):400, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
400b.pdf.

Anonymous:1870:NAO

[582] Anonymous. A new alkaloid in opium. Scientific American, 22 (25):400–401, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
400c.pdf.

Anonymous:1870:TNI

[583] Anonymous. Thirty-ninth industrial exhibition of the American institute. Scientific American, 22(25):401, June 18, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n25/pdf/scientificamerican06181870-401.pdf.

Anonymous:1870:CIE

[584] Anonymous. Cincinnati Industrial Exhibition. Scientific American, 22(25):401, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
401a.pdf.

Anonymous:1870:NMMc

[585] Anonymous. New mechanical movements. Scientific American, 22(25):401, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
401b.pdf.

Anonymous:1870:POAd

[586] Anonymous. Patent Office affairs. Scientific American, 22(25):401–402, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n25/pdf/scientificamerican06181870-401c.pdf.

Anonymous:1870:ACw

[587] Anonymous. Answers to correspondents. Scientific American, 22 (25):402, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
402a.pdf.

Anonymous:1870:RAFx

[588] Anonymous. Recent American and foreign patents. Scientific American, 22(25):402–403, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
402b.pdf.

Anonymous:1870:OLPx

[589] Anonymous. Official list of patents. Scientific American, 22 (25):403–404, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
403.pdf.

Anonymous:1870:AEPm

[590] Anonymous. Applications for the extension of patents. Scientific American, 22(25):404–406, June 18, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n25/pdf/scientificamerican06181870-
404a.pdf.

Anonymous:1870:BPbb

[591] Anonymous. Business and personal. Scientific American, 22(26):417, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n26/pdf/scientificamerican06251870-417.pdf.

Anonymous:1870:Ia

[592] Anonymous. Index. Scientific American, 22(26):419–420, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n26/pdf/scientificamerican06251870-419a.pdf.

Anonymous:1870:WTF

[593] Anonymous. The Wharfdale two-feeder printing press. Scientific American, 22(26):407–408, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
407.pdf.

Anonymous:1870:SEWa

[594] Anonymous. Scientific education for women. Scientific American, 22(26):408, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
408.pdf.

Anonymous:1870:HSS

[595] Anonymous. Habits of the siriped squibbel — a White woodchuck. Scientific American, 22(26):409, June 25, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n26/pdf/scientificamerican06251870-409.pdf.

Anonymous:1870:AMC

[596] Anonymous. Ancient and modern coinage. Scientific American, 22(26):409–410, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
409a.pdf.

Anonymous:1870:IMSb

[597] Anonymous. Improved match safe. Scientific American, 22(26): 410–411, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
410.pdf.

Anonymous:1870:BES

[598] Anonymous. The bee-eating sand-wasp. Scientific American, 22 (26):411–412, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
411.pdf.

Vestal:1870:MCC

[599] Tilghman B. Vestal. The Mammoth Cave and the cave regions of Kentucky. Scientific American, 22(26):412–413, June 25, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n26/pdf/scientificamerican06251870-412.pdf.

Anonymous:1870:POAe

[600] Anonymous. Patent Office affairs. Scientific American, 22(26):413–414, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n26/pdf/scientificamerican06251870-413.pdf.

Anonymous:1870:SMPa

[601] Anonymous. Six mouths of progress in mechanical and civil engineering. Scientific American, 22(26):415, June 25, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n26/pdf/scientificamerican06251870-415.pdf.

Anonymous:1870:CVXa

[602] Anonymous. The close of volume XXII. Scientific American, 22(26):415, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v22/n26/pdf/scientificamerican06251870-415a.pdf.

Anonymous:1870:DEC

[603] Anonymous. Danger of explosion and collapse in kitchen boilershow to avoid it. Scientific American, 22(26):415, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n26/pdf/scientificamerican06251870-415b.pdf.

Anonymous:1870:DII

[604] Anonymous. The dawn of an important industry. Scientific American, 22(26):415–416, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
415c.pdf.

Anonymous:1870:CVN

[605] Anonymous. Civilization versus nature. Scientific American, 22 (26):416, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
416.pdf.

Anonymous:1870:ISA

[606] Anonymous. Instruments for sounding alarms. Scientific American, 22(26):416, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
416a.pdf.

Anonymous:1870:SAN

[607] Anonymous. St. Anthony’s nose and the manufacture of sulphuric acid. Scientific American, 22(26):416, June 25, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v22/
n26/pdf/scientificamerican06251870-416b.pdf.

Anonymous:1870:ESM

[608] Anonymous. Extension of a sewing machine patent. Scientific American, 22(26):416–417, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
416c.pdf.

Anonymous:1870:RAFy

[609] Anonymous. Recent American and foreign patents. Scientific American, 22(26):417, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
417a.pdf.

Anonymous:1870:OLPy

[610] Anonymous. Official list of patents. Scientific American, 22 (26):418–419, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
418.pdf.

Anonymous:1870:ACx

[611] Anonymous. Answers to correspondents. Scientific American, 22 (26):419, June 25, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v22/n26/pdf/scientificamerican06251870-
419.pdf.

Anonymous:1870:BNBr

[612] Anonymous. Books: New books and publications. Scientific American, 23(1):10, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n1/pdf/scientificamerican07021870-
10a.pdf.

Anonymous:1870:BPbc

[613] Anonymous. Business and personal. Scientific American, 23(1):10, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-10b.pdf.

Anonymous:1870:PPMa

[614] Anonymous. The Peters process of manufacturing steel from pig iron, improved oil-stone holder, and more. Scientific American, 23(1):1–2, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-1.pdf.

Anonymous:1870:PHC

[615] Anonymous. Pacific Hotel, Chicago. Scientific American, 23(1):2, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-2.pdf.

Anonymous:1870:SIl

[616] Anonymous. Scientific intelligence. Scientific American, 23(1):2, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-2a.pdf.

Day:1870:TM

[617] Edward C. H. Day. Tussock-moths. Scientific American, 23(1):3–4, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-3.pdf.

Anonymous:1870:FSE

[618] Anonymous. The friction of steam engines. Scientific American, 23(1):4–6, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n1/pdf/scientificamerican07021870-
4.pdf.

Anonymous:1870:MRb

[619] Anonymous. The management of railways. Scientific American, 23(1): 7, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-7.pdf.

Anonymous:1870:RPC

[620] Anonymous. Recent progress in chemistry. Scientific American, 23(1): 7, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-7a.pdf.

Anonymous:1870:OWT

[621] Anonymous. Other worlds than ours. Scientific American, 23(1):7–8, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-7b.pdf.

Anonymous:1870:MIU

[622] Anonymous. Macfie the irrepressible upon patent law. Scientific American, 23(1):8, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n1/pdf/scientificamerican07021870-
8.pdf.

Anonymous:1870:PRCb

[623] Anonymous. Patent rights in Congress. Scientific American, 23(1):8, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-8a.pdf.

Anonymous:1870:UFSb

[624] Anonymous. Uses of fluor spar. Scientific American, 23(1):8, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-8b.pdf.

Anonymous:1870:POAf

[625] Anonymous. Patent Office affairs. Scientific American, 23(1):8, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-8c.pdf.

Anonymous:1870:RCW

[626] Anonymous. The report of Col. W. A. Roebling, Chief Engineer of the N.Y. Bridge Co. Scientific American, 23(1):8–10, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-8d.pdf.

Anonymous:1870:IPC

[627] Anonymous. Ice pitcher case. — in the U. S. Circuit Court, Eastern District of Pennsylvania. Scientific American, 23(1):10, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-10.pdf.

Anonymous:1870:ACy

[628] Anonymous. Answers to correspondents. Scientific American, 23(1):10, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-10c.pdf.

Anonymous:1870:RAFz

[629] Anonymous. Recent American and foreign patents. Scientific American, 23(1):10–11, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n1/pdf/scientificamerican07021870-
10d.pdf.

Anonymous:1870:OLPz

[630] Anonymous. Office list of patents. Scientific American, 23(1):11–12, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n1/pdf/scientificamerican07021870-11.pdf.

Anonymous:1870:AEPn

[631] Anonymous. Applications for the extension of patents. Scientific American, 23(1):12, July 2, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n1/pdf/scientificamerican07021870-
12.pdf.

Anonymous:1870:Cba

[632] Anonymous. Correspondence. Scientific American, 23(2):20, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-20.pdf.

Anonymous:1870:BPbd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-26.pdf.

Anonymous:1870:BNBs

[634] Anonymous. Books: New books and publications. Scientific American, 23(2):28, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
28a.pdf.

Anonymous:1870:LGC

[635] Anonymous. Launching of the great caisson for the Brooklyn Terminus of the East River Bridge. Scientific American, 23(2):15–16, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-15.pdf.

Darby:1870:DS

[636] John Darby. The disposal of sewage. Scientific American, 23(2):16–17, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-16.pdf.

Anonymous:1870:WEc

[637] Anonymous. What is energy? Scientific American, 23(2):17–18, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-17.pdf.

Day:1870:B

[638] Edward C. H. Day. The butterfly. Scientific American, 23(2):19–20, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-19.pdf.

Anonymous:1870:WIS

[639] Anonymous. What inventors say. Scientific American, 23(2):20–21, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-20a.pdf.

Colton:1870:AP

[640] Henry E. Colton. Adulteration of paints. Scientific American, 23(2):21, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-21.pdf.

Anonymous:1870:SSW

[641] Anonymous. Sirups for soda water. Scientific American, 23(2):21, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-21a.pdf.

Anonymous:1870:ESg

[642] Anonymous. Editorial summary. Scientific American, 23(2):21–22, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-21b.pdf.

Anonymous:1870:PC

[643] Anonymous. The progress of chemistry in 1870. Scientific American, 23(2):23, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
23.pdf.

Anonymous:1870:SRR

[644] Anonymous. Steam road rolling. Scientific American, 23(2):23, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-23a.pdf.

Anonymous:1870:GF

[645] Anonymous. Gas as fuel. Scientific American, 23(2):23–24, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-23b.pdf.

Anonymous:1870:HDM

[646] Anonymous. Hydrophobia and dog muzzling. Scientific American, 23(2):24, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
24.pdf.

Anonymous:1870:PIAa

[647] Anonymous. Progress of invention abroad. Scientific American, 23(2): 24, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-24a.pdf.

Anonymous:1870:SITb

[648] Anonymous. The Stevens institute of technology. Scientific American, 23(2):24, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
24b.pdf.

Anonymous:1870:SIm

[649] Anonymous. Scientific intelligence. Scientific American, 23(2):24–25, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-24c.pdf.

Anonymous:1870:CRC

[650] Anonymous. City railroad cars. Scientific American, 23(2):25, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-25.pdf.

Anonymous:1870:NYD

[651] Anonymous. The New York Dock Commission — the principal plans submitted. Scientific American, 23(2):25–26, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-25a.pdf.

Anonymous:1870:ACz

[652] Anonymous. Answers to correspondents. Scientific American, 23(2):26, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-26a.pdf.

Anonymous:1870:RAFba

[653] Anonymous. Recent American and foreign patents. Scientific American, 23(2):26–27, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
26b.pdf.

Anonymous:1870:OLPba

[654] Anonymous. Official list of patents. Scientific American, 23(2):27–28, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n2/pdf/scientificamerican07091870-27.pdf.

Anonymous:1870:AEPo

[655] Anonymous. Applications for the extension of patents. Scientific American, 23(2):28, July 9, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n2/pdf/scientificamerican07091870-
28.pdf.

Anonymous:1870:BNBt

[656] Anonymous. Books: New books and publications. Scientific American, 23(3):42, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
42.pdf.

Anonymous:1870:BPbe

[657] Anonymous. Business and personal. Scientific American, 23(3):42, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-42a.pdf.

Anonymous:1870:PNPa

[658] Anonymous. Patent notes: Patents. Scientific American, 23(3):44, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-44a.pdf.

Anonymous:1870:MZO

[659] Anonymous. Manufacture of zinc oxide and Bartlett lead. Scientific American, 23(3):31–32, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
31.pdf.

Anonymous:1870:SCR

[660] Anonymous. Steam on common roads. Scientific American, 23(3):32, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-32.pdf.

Anonymous:1870:MFAa

[661] Anonymous. Is man a feeble animal? Scientific American, 23(3):33, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-33.pdf.

Anonymous:1870:RAA

[662] Anonymous. Resemblances among animals. Scientific American, 23(3):33–34, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
33a.pdf.

Anonymous:1870:LFb

[663] Anonymous. The lantern fly. Scientific American, 23(3):35, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-35.pdf.

Anonymous:1870:CEPb

[664] Anonymous. Curculio extirpation possible. Scientific American, 23(3):35, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
35a.pdf.

Anonymous:1870:IIN

[665] Anonymous. Imported insects and native American insects. Scientific American, 23(3):36, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
36.pdf.

Anonymous:1870:CRM

[666] Anonymous. A chill room in a meat packers’ establishment. Scientific American, 23(3):37, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
37.pdf.

Anonymous:1870:FR

[667] Anonymous. Facts and recipes. Scientific American, 23(3):37, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-37a.pdf.

Anonymous:1870:KLM

[668] Anonymous. Kelly’s lasting machine boring bit. Scientific American, 23(3):38, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
38.pdf.

Anonymous:1870:FCC

[669] Anonymous. Foster’s convertible cultivator, hoe, fork, rake, and spade. Scientific American, 23(3):38, July 16, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n3/pdf/scientificamerican07161870-38a.pdf.

Anonymous:1870:GSO

[670] Anonymous. Gas stoves for ordinary domestic use. Scientific American, 23(3):39, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
39.pdf.

Anonymous:1870:SDF

[671] Anonymous. Statics and dynamics of friction. Scientific American, 23(3):39, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
39a.pdf.

Anonymous:1870:DCS

[672] Anonymous. The Darien Canal Survey. Scientific American, 23(3):39, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-39b.pdf.

Anonymous:1870:DWM

[673] Anonymous. A denunciation of water meters. Scientific American, 23(3):39–40, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
39c.pdf.

Anonymous:1870:GAS

[674] Anonymous. A glance at aeronautical science — review of a paper read at a meeting of the Aeronautical Society in London, by W. Clare. Scientific American, 23(3):40, July 16, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n3/pdf/scientificamerican07161870-40.pdf.

Anonymous:1870:HDH

[675] Anonymous. The history of a defunct horse. Scientific American, 23(3):40, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
40a.pdf.

Anonymous:1870:IMa

[676] Anonymous. Ice by machinery. Scientific American, 23(3):40, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-40b.pdf.

Anonymous:1870:SIn

[677] Anonymous. Scientific intelligence. Scientific American, 23(3):40–41, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-40c.pdf.

Anonymous:1870:VC

[678] Anonymous. Vassar College. Scientific American, 23(3):41, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-41.pdf.

Anonymous:1870:SMPb

[679] Anonymous. A sewing machine-patent extension — an irate opponent. Scientific American, 23(3):41, July 16, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n3/pdf/scientificamerican07161870-41a.pdf.

Anonymous:1870:CSA

[680] Anonymous. Correspondence of the Scientific American. Scientific American, 23(3):41–42, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
41b.pdf.

Anonymous:1870:ACba

[681] Anonymous. Answers to correspondents. Scientific American, 23(3):42, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-42b.pdf.

Anonymous:1870:RAFbb

[682] Anonymous. Recent American and foreign patents. Scientific American, 23(3):42–43, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
42c.pdf.

Anonymous:1870:OLPbb

[683] Anonymous. Official list of patents. Scientific American, 23(3):43–44, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n3/pdf/scientificamerican07161870-43.pdf.

Anonymous:1870:AEPp

[684] Anonymous. Applications for the extension of patents. Scientific American, 23(3):44, July 16, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n3/pdf/scientificamerican07161870-
44.pdf.

Anonymous:1870:Cbb

[685] Anonymous. Correspondence. Scientific American, 23(4):51–52, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-51a.pdf.

Anonymous:1870:BPbf

[686] Anonymous. Business and personal. Scientific American, 23(4):58, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-58a.pdf.

Anonymous:1870:PNPb

[687] Anonymous. Patent notes: Patents. Scientific American, 23(4):60, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-60.pdf.

Anonymous:1870:IBT

[688] Anonymous. Iron bridge at Topeka, Kansas. Scientific American, 23(4):48, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
48.pdf.

Anonymous:1870:SD

[689] Anonymous. Strange drinks. Scientific American, 23(4):49, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-49.pdf.

Anonymous:1870:WSOb

[690] Anonymous. S. a, woods’ self-oiling circular-saw arbors and adjustable cone bushing. Scientific American, 23(4):50, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-50.pdf.

Anonymous:1870:NPI

[691] Anonymous. Nickel-plating — interesting to marine engineers. Scientific American, 23(4):50, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
50a.pdf.

Anonymous:1870:MFd

[692] Anonymous. Mechanical fingers. Scientific American, 23(4):50, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-50b.pdf.

Anonymous:1870:H

[693] Anonymous. The Hesperians. Scientific American, 23(4):51, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-51.pdf.

Anonymous:1870:SEWb

[694] Anonymous. The scientific education of women. Scientific American, 23(4):53–54, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
53.pdf.

Anonymous:1870:HHI

[695] Anonymous. How the heirs of an inventor may be rewarded. Scientific American, 23(4):55, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
55.pdf.

Anonymous:1870:SAP

[696] Anonymous. Synopsis of the amended patent law. Scientific American, 23(4):55, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
55a.pdf.

Anonymous:1870:EGC

[697] Anonymous. Experiments with gun cotton. Scientific American, 23(4):55, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
55b.pdf.

Anonymous:1870:MBR

[698] Anonymous. The manufacture of beet-root sugar in England. Scientific American, 23(4):55–56, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
55c.pdf.

Anonymous:1870:PPSb

[699] Anonymous. Paper and paper stock. Scientific American, 23(4):56, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-56.pdf.

Anonymous:1870:AU

[700] Anonymous. Artificial ultramarine. Scientific American, 23(4):56, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-56a.pdf.

Anonymous:1870:PSC

[701] Anonymous. Projected ship canals. Scientific American, 23(4):56, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-56b.pdf.

Anonymous:1870:FM

[702] Anonymous. Friction matches. Scientific American, 23(4):56–57, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-56c.pdf.

Anonymous:1870:RDA

[703] Anonymous. Report of the department of agriculture for may and June. Scientific American, 23(4):57, July 23, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n4/pdf/scientificamerican07231870-57.pdf.

Anonymous:1870:SIo

[704] Anonymous. Scientific intelligence. Scientific American, 23(4):57, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-57a.pdf.

Anonymous:1870:RAM

[705] Anonymous. The right of appeal maintained. Scientific American, 23(4):57, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
57b.pdf.

Anonymous:1870:AIE

[706] Anonymous. American Institute Exposition. Scientific American, 23(4):57–58, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
57c.pdf.

Anonymous:1870:AEPq

[707] Anonymous. Applications for the extension of patents. Scientific American, 23(4):58, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
58.pdf.

Anonymous:1870:ACbb

[708] Anonymous. Answers to correspondents. Scientific American, 23(4):58, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-58b.pdf.

Anonymous:1870:RAFbc

[709] Anonymous. Recent American and foreign patents. Scientific American, 23(4):58–59, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n4/pdf/scientificamerican07231870-
58c.pdf.

Anonymous:1870:OLPbc

[710] Anonymous. Official list of patents. Scientific American, 23(4):59–60, July 23, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n4/pdf/scientificamerican07231870-59.pdf.

Anonymous:1870:Cbc

[711] Anonymous. Correspondence. Scientific American, 23(5):68–69, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-68.pdf.

Anonymous:1870:BPbg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-74b.pdf.

Anonymous:1870:IEM

[713] Anonymous. Improved engraving machine. Scientific American, 23(5):63–64, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
63.pdf.

Anonymous:1870:RTI

[714] Anonymous. Richard Trevithick inventor of the locomotive. Scientific American, 23(5):64–65, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
64.pdf.

Burleigh:1870:TL

[715] George S. Burleigh. Of thought without language. Scientific American, 23(5):65–66, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
65.pdf.

Anonymous:1870:SIp

[716] Anonymous. Scientific intelligence. Scientific American, 23(5):66, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-66.pdf.

Day:1870:H

[717] Edward C. H. Day. The hymenopters. Scientific American, 23(5):67–68, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-67.pdf.

Howig:1870:EHC

[718] F. A. Howig. Experiments with hydrate of chloral. Scientific American, 23(5):69, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
69.pdf.

Anonymous:1870:NDH

[719] Anonymous. A new dodge in hair dyes. Scientific American, 23(5):69–70, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
69a.pdf.

Anonymous:1870:ASC

[720] Anonymous. Alexander’s safety-cap for molasses, varnish, and oil cans. Scientific American, 23(5):70, July 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n5/pdf/scientificamerican07301870-70.pdf.

Anonymous:1870:DHW

[721] Anonymous. Duckham’s hydrostatic weighing machine. Scientific American, 23(5):70, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
70a.pdf.

Anonymous:1870:FSB

[722] Anonymous. Floato-static breakwater. Scientific American, 23(5):70, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-70b.pdf.

Anonymous:1870:BSA

[723] Anonymous. Bessemer steel — American improvements. Scientific American, 23(5):71, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
71.pdf.

Anonymous:1870:NMMd

[724] Anonymous. New mechanical movements. Scientific American, 23(5):71, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
71a.pdf.

Anonymous:1870:NBA

[725] Anonymous. A new bleaching agent. Scientific American, 23(5):71, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-71b.pdf.

Anonymous:1870:CMM

[726] Anonymous. Can mineral matter be assimilated in the animal economy? Scientific American, 23(5):71–72, July 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n5/pdf/scientificamerican07301870-71c.pdf.

Anonymous:1870:PIAb

[727] Anonymous. Progress of invention abroad. Scientific American, 23(5):72, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
72.pdf.

Anonymous:1870:LEM

[728] Anonymous. Liebig’s extract of meat. Scientific American, 23(5):72, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-72a.pdf.

Anonymous:1870:OC

[729] Anonymous. Our correspondence. Scientific American, 23(5):72–73, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-72b.pdf.

Anonymous:1870:HIA

[730] Anonymous. How inventors are appreciated in Congress. Scientific American, 23(5):73, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
73.pdf.

Anonymous:1870:ASW

[731] Anonymous. Action of sea water on the metallic fastenings of ships. Scientific American, 23(5):73, July 30, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n5/pdf/scientificamerican07301870-73a.pdf.

Anonymous:1870:PPF

[732] Anonymous. Public parks and floral decorations. Scientific American, 23(5):73–74, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
73b.pdf.

Anonymous:1870:RPP

[733] Anonymous. Rogers’ patent price current books. Scientific American, 23(5):74, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
74.pdf.

Anonymous:1870:AEPr

[734] Anonymous. Applications for the extension of patents. Scientific American, 23(5):74, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
74a.pdf.

Anonymous:1870:ACbc

[735] Anonymous. Answers to correspondents. Scientific American, 23(5):74, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-74c.pdf.

Anonymous:1870:RAFbd

[736] Anonymous. Recent American and foreign patents. Scientific American, 23(5):74–75, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n5/pdf/scientificamerican07301870-
74d.pdf.

Anonymous:1870:OLPbd

[737] Anonymous. Official list of patents. Scientific American, 23(5):75–76, July 30, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n5/pdf/scientificamerican07301870-75.pdf.

Anonymous:1870:Cbd

[738] Anonymous. Correspondence. Scientific American, 23(6):84–85, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-84.pdf.

Anonymous:1870:BPbh

[739] Anonymous. Business and personal. Scientific American, 23(6):91, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-91.pdf.

Anonymous:1870:UWW

[740] Anonymous. Universal wood working machine. Scientific American, 23(6):79, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
79.pdf.

Anonymous:1870:GPP

[741] Anonymous. The grinding and polishing of plate glass by machinery. Scientific American, 23(6):80–82, August 6, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n6/pdf/scientificamerican08061870-80.pdf.

Anonymous:1870:FF

[742] Anonymous. The “farmer” fountain. Scientific American, 23(6):82, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-82.pdf.

Anonymous:1870:AOS

[743] Anonymous. Anderson’s oil safe. Scientific American, 23(6):82, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-82a.pdf.

Anonymous:1870:PNG

[744] Anonymous. The Prussian needle gun and the French Chas-Sepot needle gun. Scientific American, 23(6):83, August 6, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n6/pdf/scientificamerican08061870-83.pdf.

Anonymous:1870:HMb

[745] Anonymous. Hydraulic machinery. Scientific American, 23(6):86, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-86.pdf.

Anonymous:1870:MIG

[746] Anonymous. The manufacture of illuminating gas. Scientific American, 23(6):87, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
87.pdf.

Anonymous:1870:CVZ

[747] Anonymous. Coating vessels with zinc. Scientific American, 23(6):87, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-87a.pdf.

Anonymous:1870:NPL

[748] Anonymous. The new patent laws — important changes affecting American and foreign manufacturers — free trade in patents now fully established. Scientific American, 23(6):87–88, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-87b.pdf.

Anonymous:1870:HMT

[749] Anonymous. How mercurial thermometers are made. Scientific American, 23(6):88, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
88.pdf.

Anonymous:1870:KEG

[750] Anonymous. Is the knowledge of English grammar necessary to the writing of good English? Scientific American, 23(6):88, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-88a.pdf.

Anonymous:1870:DAV

[751] Anonymous. Depreciation of American vitality. Scientific American, 23(6):88, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
88b.pdf.

Anonymous:1870:SIq

[752] Anonymous. Scientific intelligence. Scientific American, 23(6):88–89, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-88c.pdf.

Anonymous:1870:HOW

[753] Anonymous. How our wheat crops are handled. Scientific American, 23(6):89–90, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
89.pdf.

Anonymous:1870:RAFbe

[754] Anonymous. Recent Americans and foreign patents. Scientific American, 23(6):90, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
90.pdf.

Anonymous:1870:ACbd

[755] Anonymous. Answers to correspondents. Scientific American, 23(6):90–91, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n6/pdf/scientificamerican08061870-
90a.pdf.

Anonymous:1870:OLPbe

[756] Anonymous. Official list of patents. Scientific American, 23(6):91–92, August 6, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n6/pdf/scientificamerican08061870-91a.pdf.

Anonymous:1870:Cbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-100.pdf.

Anonymous:1870:BNBu

[758] Anonymous. Books: New books and publications. Scientific American, 23(7):106, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
106a.pdf.

Anonymous:1870:BPbi

[759] Anonymous. Business and personal. Scientific American, 23(7):106, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-106d.pdf.

Anonymous:1870:ITT

[760] Anonymous. Improved turn-table. Scientific American, 23(7):95, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-95.pdf.

Anonymous:1870:NB

[761] Anonymous. New bread. Scientific American, 23(7):95–96, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-95a.pdf.

Anonymous:1870:MDa

[762] Anonymous. Magnetism and diamagnetism. Scientific American, 23(7):97, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
97.pdf.

Day:1870:LCB

[763] Edward C. H. Day. The leaf-cutter bee. Scientific American, 23(7):98, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
98.pdf.

Anonymous:1870:HPS

[764] Anonymous. Harris’ patent self-acting mashing machine for the use of brewers. Scientific American, 23(7):99, August 13, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n7/pdf/scientificamerican08131870-99.pdf.

Anonymous:1870:PMPb

[765] Anonymous. Patent metallic post butt. Scientific American, 23 (7):99–100, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
99a.pdf.

Anonymous:1870:IWE

[766] Anonymous. Improved weapons in European warfare. Scientific American, 23(7):103, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
103.pdf.

Anonymous:1870:LC

[767] Anonymous. The law of compensation. Scientific American, 23 (7):103–104, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
103a.pdf.

Anonymous:1870:SSa

[768] Anonymous. Saratoga Springs. Scientific American, 23(7):104, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-104.pdf.

Anonymous:1870:NDS

[769] Anonymous. Ninety degrees in the shade. Scientific American, 23(7):104, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
104a.pdf.

Anonymous:1870:AWL

[770] Anonymous. The action of water on lead, tin, and copper. Scientific American, 23(7):104, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
104b.pdf.

Anonymous:1870:USG

[771] Anonymous. The uses of soluble glass. Scientific American, 23 (7):104, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
104c.pdf.

Anonymous:1870:WDN

[772] Anonymous. Who discovered nitro-glycerin. Scientific American, 23(7):104–105, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
104d.pdf.

Steele:1870:CAM

[773] J. Dutton Steele. Compressed air as a motor for subterranean railways. Scientific American, 23(7):105, August 13, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n7/pdf/scientificamerican08131870-105.pdf.

Anonymous:1870:CAAb

[774] Anonymous. Compressed air, ammonia, and rubber vs. mule power. Scientific American, 23(7):105, August 13, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n7/pdf/scientificamerican08131870-105a.pdf.

Anonymous:1870:SIr

[775] Anonymous. Scientific intelligence. Scientific American, 23(7):105, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-105b.pdf.

Anonymous:1870:Mb

[776] Anonymous. The mitrailleuse. Scientific American, 23(7):106, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n7/pdf/scientificamerican08131870-106.pdf.

Anonymous:1870:AEPs

[777] Anonymous. Applications for the extension of patents. Scientific American, 23(7):106, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
106b.pdf.

Anonymous:1870:ACbe

[778] Anonymous. Answers to correspondents. Scientific American, 23(7):106, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
106c.pdf.

Anonymous:1870:RAFbf

[779] Anonymous. Recent American and foreign patents. Scientific American, 23(7):107, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
107.pdf.

Anonymous:1870:OLPbf

[780] Anonymous. Official list of patents. Scientific American, 23 (7):107–108, August 13, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n7/pdf/scientificamerican08131870-
107a.pdf.

Anonymous:1870:Cbf

[781] Anonymous. Correspondence. Scientific American, 23(8):116–117, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n8/pdf/scientificamerican08201870-116.pdf.

Anonymous:1870:BPbj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n8/pdf/scientificamerican08201870-122.pdf.

Anonymous:1870:ISSa

[783] Anonymous. Improved screw steering apparatus, the “siphon-recorder” for submarine cables, and more. Scientific American, 23(8):111, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n8/pdf/scientificamerican08201870-111.pdf.

Anonymous:1870:AFB

[784] Anonymous. The artisan in France, Belgium, and Holland. Scientific American, 23(8):112–113, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
112.pdf.

Day:1870:OGT

[785] Edward C. H. Day. The orthopters — the grasshoppers and their allies. Scientific American, 23(8):114, August 20, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n8/pdf/scientificamerican08201870-114.pdf.

Anonymous:1870:SRV

[786] Anonymous. The strike of the “raffineurs” at La Villette, France. Scientific American, 23(8):115–116, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
115.pdf.

Piesse:1870:SPF

[787] Septimus Piesse. Scientific and practical facts and items. Scientific American, 23(8):117–118, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
117.pdf.

Anonymous:1870:IWU

[788] Anonymous. The influence of war upon inventions. Scientific American, 23(8):119, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
119.pdf.

Anonymous:1870:PMV

[789] Anonymous. Paper money vs. gold. Scientific American, 23(8):119, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n8/pdf/scientificamerican08201870-119a.pdf.

Anonymous:1870:SEI

[790] Anonymous. The sale of explosive illuminating fluids. Scientific American, 23(8):119, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
119b.pdf.

Anonymous:1870:IMI

[791] Anonymous. Important to manufacturers and importers. Scientific American, 23(8):119–120, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
119c.pdf.

Anonymous:1870:VDI

[792] Anonymous. Visit to the delamater iron works, New York. Scientific American, 23(8):120, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
120.pdf.

Anonymous:1870:SG

[793] Anonymous. Spontaneous generation. Scientific American, 23 (8):120–121, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
120a.pdf.

Anonymous:1870:OJH

[794] Anonymous. The Orange Judd Hall of Natural Science. Scientific American, 23(8):121, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
121.pdf.

Anonymous:1870:SIs

[795] Anonymous. Scientific intelligence. Scientific American, 23(8): 121–122, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
121a.pdf.

Anonymous:1870:ACbf

[796] Anonymous. Answers to correspondents. Scientific American, 23(8):122, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
122a.pdf.

Anonymous:1870:RAFbg

[797] Anonymous. Recent American and foreign patents. Scientific American, 23(8):123, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
123.pdf.

Anonymous:1870:AEPt

[798] Anonymous. Applications for the extension of patents. Scientific American, 23(8):123, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
123a.pdf.

Anonymous:1870:OLPbg

[799] Anonymous. Official list of patents. Scientific American, 23 (8):123–124, August 20, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n8/pdf/scientificamerican08201870-
123b.pdf.

Anonymous:1870:Cbg

[800] Anonymous. Correspondence. Scientific American, 23(9):132–133, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-132.pdf.

Anonymous:1870:BPbk

[801] Anonymous. Business and personal. Scientific American, 23(9):138, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-138a.pdf.

Anonymous:1870:BNBv

[802] Anonymous. Books: New books and publications. Scientific American, 23(9):140, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
140.pdf.

Anonymous:1870:SEM

[803] Anonymous. Sternberg’s electro-magnetic regulator for dampers and valves, Hirn’s telodynamic cable, and more. Scientific American, 23(9):127, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
127.pdf.

Anonymous:1870:SSD

[804] Anonymous. On the science of sleep and dreams. Scientific American, 23(9):128–129, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
128.pdf.

Anonymous:1870:MLb

[805] Anonymous. Mountain locomotive. Scientific American, 23(9):130, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-130.pdf.

Day:1870:MLT

[806] Edward C. H. Day. The migratory locusts and their American cousins. Scientific American, 23(9):131, August 27, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n9/pdf/scientificamerican08271870-131.pdf.

Anonymous:1870:SGW

[807] Anonymous. The science of going to war — how Prussia is superior to France. Scientific American, 23(9):131, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-131a.pdf.

Anonymous:1870:TTF

[808] Anonymous. The true theory of flying. Scientific American, 23 (9):133, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
133.pdf.

Anonymous:1870:DCA

[809] Anonymous. Diamond cutting in Amsterdam. Scientific American, 23(9):134, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
134.pdf.

Anonymous:1870:MIC

[810] Anonymous. Melendy’s improved combined key ring and door fastener with compensator. Scientific American, 23(9):134, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-134a.pdf.

Anonymous:1870:USAb

[811] Anonymous. The uses of sulphurous acid. Scientific American, 23(9):135, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
135.pdf.

Anonymous:1870:EAU

[812] Anonymous. Effect of association upon mental development. Scientific American, 23(9):135, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
135a.pdf.

Anonymous:1870:WMMb

[813] Anonymous. Who make mistakes? Scientific American, 23(9):135, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-135b.pdf.

Anonymous:1870:PS

[814] Anonymous. Pen stammering. Scientific American, 23(9):135–136, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-135c.pdf.

Anonymous:1870:RH

[815] Anonymous. Radiation of heat. Scientific American, 23(9):136, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n9/pdf/scientificamerican08271870-136.pdf.

Anonymous:1870:CSP

[816] Anonymous. The Commissioner suspends a patent agent for gross misconduct. Scientific American, 23(9):136, August 27, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n9/pdf/scientificamerican08271870-136a.pdf.

Anonymous:1870:PEW

[817] Anonymous. The present European war. Scientific American, 23(9):136, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
136b.pdf.

Anonymous:1870:UTC

[818] Anonymous. The use of torpedoes for coast defense. Scientific American, 23(9):136, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
136c.pdf.

Anonymous:1870:DPP

[819] Anonymous. Death of Professor Palmstedt. Scientific American, 23(9):136, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
136d.pdf.

Anonymous:1870:SIu

[820] Anonymous. Scientific intelligence. Scientific American, 23(9): 136–137, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
136e.pdf.

Anonymous:1870:AAAa

[821] Anonymous. American Association for the Advancement of Science. Scientific American, 23(9):137–138, August 27, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n9/pdf/scientificamerican08271870-137.pdf.

Anonymous:1870:ACbg

[822] Anonymous. Answers to correspondents. Scientific American, 23(9):138, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
138.pdf.

Anonymous:1870:RAFbh

[823] Anonymous. Recent American and foreign patents. Scientific American, 23(9):139, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
139.pdf.

Anonymous:1870:OLPbh

[824] Anonymous. Official list of patents. Scientific American, 23 (9):139–140, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
139a.pdf.

Anonymous:1870:AEPu

[825] Anonymous. Applications for the extension of patents. Scientific American, 23(9):140, August 27, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n9/pdf/scientificamerican08271870-
140a.pdf.

Anonymous:1870:Cbh

[826] Anonymous. Correspondence. Scientific American, 23(10):148, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n10/pdf/scientificamerican09031870-148.pdf.

Anonymous:1870:BNBw

[827] Anonymous. Books: New books and publications. Scientific American, 23(10):154, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
154.pdf.

Anonymous:1870:BPbl

[828] Anonymous. Business and personal. Scientific American, 23(10): 154–155, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
154b.pdf.

Anonymous:1870:IDM

[829] Anonymous. Improved dredging machine. Scientific American, 23 (10):143, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
143.pdf.

Marten:1870:SBI

[830] E. B. Marten. On steam boiler inspection. Scientific American, 23(10):144, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
144.pdf.

Anonymous:1870:APa

[831] Anonymous. The artisan in Prussia. Scientific American, 23(10): 144–145, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
144a.pdf.

Anonymous:1870:APM

[832] Anonymous. The application of photography to military purposes. Scientific American, 23(10):145–146, September 3, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n10/pdf/scientificamerican09031870-145.pdf.

Day:1870:Sb

[833] Edward C. H. Day. The scarabæids. Scientific American, 23(10): 147–148, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
147.pdf.

Anonymous:1870:SSb

[834] Anonymous. Stone sawing. Scientific American, 23(10):149, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n10/pdf/scientificamerican09031870-149.pdf.

Anonymous:1870:AAAb

[835] Anonymous. American Association for the Advancement of Science. Scientific American, 23(10):151, September 3, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n10/pdf/scientificamerican09031870-151.pdf.

Anonymous:1870:SLBb

[836] Anonymous. The St. Louis Bridge. Scientific American, 23 (10):151, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
151a.pdf.

Anonymous:1870:WLE

[837] Anonymous. A want in locomotive engineering. Scientific American, 23(10):151, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
151b.pdf.

Anonymous:1870:SDM

[838] Anonymous. The Southern demand for machinery. Scientific American, 23(10):151–152, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
151c.pdf.

Anonymous:1870:PIc

[839] Anonymous. Paraffine industry. Scientific American, 23(10): 152, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
152.pdf.

Anonymous:1870:CPA

[840] Anonymous. Craik’s practical American millwright and miller. Scientific American, 23(10):152, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
152a.pdf.

Anonymous:1870:MSL

[841] Anonymous. The millennium, or something like it. Scientific American, 23(10):152, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
152b.pdf.

Anonymous:1870:AM

[842] Anonymous. The analysis of milk. Scientific American, 23(10): 152–153, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
152c.pdf.

Anonymous:1870:DMSa

[843] Anonymous. Divisibility of matter and size of chemical atoms. Scientific American, 23(10):153, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
153.pdf.

Anonymous:1870:EJ

[844] Anonymous. Enterprising journalism. Scientific American, 23 (10):153, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
153a.pdf.

Anonymous:1870:BNP

[845] Anonymous. $20,000 bonus for a new press. Scientific American, 23(10):153, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
153b.pdf.

Anonymous:1870:SIv

[846] Anonymous. Scientific intelligence. Scientific American, 23(10): 153, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
153c.pdf.

Anonymous:1870:HW

[847] Anonymous. Hight and weight. Scientific American, 23(10): 153–154, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
153d.pdf.

Anonymous:1870:ACbh

[848] Anonymous. Answers to correspondents. Scientific American, 23 (10):154, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
154a.pdf.

Anonymous:1870:RAFbi

[849] Anonymous. Recent American and foreign patents. Scientific American, 23(10):155, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
155.pdf.

Anonymous:1870:OLPbi

[850] Anonymous. Official list of patents. Scientific American, 23(10): 155–156, September 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n10/pdf/scientificamerican09031870-
155a.pdf.

Anonymous:1870:Cbi

[851] Anonymous. Correspondence. Scientific American, 23(11):164– 166, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
164a.pdf.

Anonymous:1870:BNBx

[852] Anonymous. Books: New books and publications. Scientific American, 23(11):170, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
170.pdf.

Anonymous:1870:BPbm

[853] Anonymous. Business and personal. Scientific American, 23 (11):170, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
170a.pdf.

Anonymous:1870:WAa

[854] Anonymous. Watchmaking in America. Scientific American, 23 (11):159–164, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
159.pdf.

Anonymous:1870:CST

[855] Anonymous. The carpenter self-threading and self-setting needle for sewing machines. Scientific American, 23(11):164, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n11/pdf/scientificamerican09101870-164.pdf.

Anonymous:1870:SCb

[856] Anonymous. Social classification. Scientific American, 23(11): 167, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
167.pdf.

Anonymous:1870:ISU

[857] Anonymous. Influence of sound upon rain. Scientific American, 23(11):167, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
167a.pdf.

Anonymous:1870:CPb

[858] Anonymous. Concrete paving. Scientific American, 23(11):167– 168, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
167b.pdf.

Anonymous:1870:DAI

[859] Anonymous. Destructive action of illuminating gas on vegetation. Scientific American, 23(11):168, September 10, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n11/pdf/scientificamerican09101870-168.pdf.

Anonymous:1870:BIc

[860] Anonymous. Boiler incrustations. Scientific American, 23(11): 168, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
168a.pdf.

Anonymous:1870:RP

[861] Anonymous. As regards protoplasm. Scientific American, 23 (11):168, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
168b.pdf.

Anonymous:1870:MK

[862] Anonymous. Manufacture of kumis. Scientific American, 23 (11):168, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
168c.pdf.

Anonymous:1870:RSSa

[863] Anonymous. Raising of the steamship “Scotland” off Sandy Hook. Scientific American, 23(11):168–169, September 10, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n11/pdf/scientificamerican09101870-168d.pdf.

Colton:1870:AS

[864] H. E. Colton. American sumac. Scientific American, 23(11): 169–170, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
169.pdf.

Anonymous:1870:ACbi

[865] Anonymous. Answers to correspondents. Scientific American, 23 (11):171, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
171.pdf.

Anonymous:1870:RAFbj

[866] Anonymous. Recent American and foreign patents. Scientific American, 23(11):171, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
171a.pdf.

Anonymous:1870:OLPbj

[867] Anonymous. Official list of patents. Scientific American, 23(11): 171–172, September 10, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n11/pdf/scientificamerican09101870-
171b.pdf.

Anonymous:1870:Cbj

[868] Anonymous. Correspondence. Scientific American, 23(12):180– 181, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
180.pdf.

Anonymous:1870:BPbn

[869] Anonymous. Business and personal. Scientific American, 23 (12):186, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
186.pdf.

Anonymous:1870:ICH

[870] Anonymous. Improved cotton and hay press. Scientific American, 23(12):175, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
175.pdf.

Anonymous:1870:BBM

[871] Anonymous. Bells and Bell metal. Scientific American, 23(12): 175–176, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
175a.pdf.

Anonymous:1870:BRSb

[872] Anonymous. British railway signals. Scientific American, 23 (12):176, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
176.pdf.

Anonymous:1870:AAR

[873] Anonymous. The artisan in Austria, Russia, and Sweden. Scientific American, 23(12):177–178, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
177.pdf.

Anonymous:1870:GIB

[874] Anonymous. Gleason’s improved band saw. Scientific American, 23 (12):178–179, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
178.pdf.

Day:1870:SFB

[875] Edward C. H. Day. The scarabæids — the flower beetles and their allies. Scientific American, 23(12):179–180, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n12/pdf/scientificamerican09171870-179.pdf.

Maclaren:1870:TPF

[876] Archibald Maclaren. The training of Prussian and French soldiers. Scientific American, 23(12):181–182, September 17, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n12/pdf/scientificamerican09171870-181.pdf.

Anonymous:1870:IUG

[877] Anonymous. The increasing use of glycerin. Scientific American, 23(12):183, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
183.pdf.

Anonymous:1870:GSS

[878] Anonymous. The government storm signals. Scientific American, 23(12):183, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
183a.pdf.

Anonymous:1870:MDb

[879] Anonymous. Mechanical drawing. Scientific American, 23(12): 183–184, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
183b.pdf.

Anonymous:1870:FIT

[880] Anonymous. Female inventive talent. Scientific American, 23 (12):184, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
184.pdf.

Anonymous:1870:AML

[881] Anonymous. American magazine literature. Scientific American, 23(12):184, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
184a.pdf.

Anonymous:1870:FPW

[882] Anonymous. The Franco–Prussian War. Scientific American, 23 (12):184, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
184b.pdf.

Anonymous:1870:SSS

[883] Anonymous. The Sheffield Scientific School of Yale College. Scientific American, 23(12):185, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
185.pdf.

Anonymous:1870:HSD

[884] Anonymous. Harvard School of Dentistry. Scientific American, 23 (12):185, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
185a.pdf.

Anonymous:1870:SIw

[885] Anonymous. Scientific intelligence. Scientific American, 23(12): 185–186, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
185b.pdf.

Anonymous:1870:ACbj

[886] Anonymous. Answers to correspondents. Scientific American, 23 (12):186, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
186a.pdf.

Anonymous:1870:RAFbk

[887] Anonymous. Recent American and foreign patents. Scientific American, 23(12):186–187, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
186b.pdf.

Anonymous:1870:OLPbk

[888] Anonymous. Official list of patents. Scientific American, 23(12): 187–188, September 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n12/pdf/scientificamerican09171870-
187.pdf.

Anonymous:1870:Cbk

[889] Anonymous. Correspondence. Scientific American, 23(13):196, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n13/pdf/scientificamerican09241870-196.pdf.

Anonymous:1870:BPbo

[890] Anonymous. Business and personal. Scientific American, 23 (13):202, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
202a.pdf.

Anonymous:1870:BNBy

[891] Anonymous. Books: New books and publications. Scientific American, 23(13):202, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
202b.pdf.

Anonymous:1870:ISSb

[892] Anonymous. Improved scroll-sawing machinery, gutta-percha. Scientific American, 23(13):191, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
191.pdf.

Anonymous:1870:HFM

[893] Anonymous. The history and fortifications of Metz. Scientific American, 23(13):191–192, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
191a.pdf.

Anonymous:1870:DSEa

[894] Anonymous. Deep-sea exploration — how soundings are obtained. Scientific American, 23(13):192–193, September 24, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n13/pdf/scientificamerican09241870-192.pdf.

Anonymous:1870:ADS

[895] Anonymous. The artisan in Denmark, Spain, Portugal and Greece. Scientific American, 23(13):193–194, September 24, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n13/pdf/scientificamerican09241870-193.pdf.

Anonymous:1870:IWC

[896] Anonymous. Improved wicket for canal locks. Scientific American, 23(13):194, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
194.pdf.

Day:1870:Sc

[897] Edward C. H. Day. The spider. Scientific American, 23(13): 195, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
195.pdf.

Anonymous:1870:ET

[898] Anonymous. The electric telegraph. Scientific American, 23(13): 196–198, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
196a.pdf.

Anonymous:1870:FLN

[899] Anonymous. Freeland’s lock nuts. Scientific American, 23(13): 198, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
198.pdf.

Anonymous:1870:ISC

[900] Anonymous. Improved stove cover, dish, and kettle lifter. Scientific American, 23(13):198, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
198a.pdf.

Anonymous:1870:MC

[901] Anonymous. Modern chemistry. Scientific American, 23(13): 199, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
199.pdf.

Anonymous:1870:WOC

[902] Anonymous. What are our colleges doing? Scientific American, 23(13):199, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
199a.pdf.

Anonymous:1870:DV

[903] Anonymous. The draft of vehicles. Scientific American, 23(13): 199, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
199b.pdf.

Anonymous:1870:FDM

[904] Anonymous. The finishing and decoration of machinery. Scientific American, 23(13):199–200, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
199c.pdf.

Anonymous:1870:PFIa

[905] Anonymous. Progress of foreign invention. Scientific American, 23 (13):200, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
200.pdf.

Anonymous:1870:CEa

[906] Anonymous. Chemical experiments. Scientific American, 23(13): 200, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
200a.pdf.

Anonymous:1870:WAb

[907] Anonymous. The “worm” on the Ailanthus. Scientific American, 23 (13):200–201, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
200b.pdf.

Anonymous:1870:FAIa

[908] Anonymous. Fair of the American institute. Scientific American, 23(13):201, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
201.pdf.

Anonymous:1870:MSR

[909] Anonymous. Mileage system of railroad commutation tickets. Scientific American, 23(13):201–202, September 24, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n13/pdf/scientificamerican09241870-201a.pdf.

Anonymous:1870:AEPv

[910] Anonymous. Applications for the extension of patents. Scientific American, 23(13):202, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
202.pdf.

Anonymous:1870:ACbk

[911] Anonymous. Answers to correspondents. Scientific American, 23 (13):202, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
202c.pdf.

Anonymous:1870:RAFbl

[912] Anonymous. Recent American and foreign patents. Scientific American, 23(13):202–203, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
202d.pdf.

Anonymous:1870:OLPbl

[913] Anonymous. Official list of patents. Scientific American, 23(13): 203–204, September 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n13/pdf/scientificamerican09241870-
203.pdf.

Anonymous:1870:Cbl

[914] Anonymous. Correspondence. Scientific American, 23(14):212–214, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-212.pdf.

Anonymous:1870:BNBz

[915] Anonymous. Books: New books and publications. Scientific American, 23(14):218, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
218a.pdf.

Anonymous:1870:BPbp

[916] Anonymous. Business and personal. Scientific American, 23(14):218, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-218c.pdf.

Anonymous:1870:KPS

[917] Anonymous. Krupp’s 1,000-pounder siege gun, great circle “travel”. Scientific American, 23(14):207, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
207.pdf.

Anonymous:1870:SSC

[918] Anonymous. Shortt’s self-centering sleeve coupling. Scientific American, 23(14):207–208, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
207a.pdf.

Anonymous:1870:BVS

[919] Anonymous. Bronze vs. steel and iron for field artillery. Scientific American, 23(14):208, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
208.pdf.

Anonymous:1870:DE

[920] Anonymous. The dissipation of energy. Scientific American, 23 (14):208–209, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
208a.pdf.

Skey:1870:AHB

[921] W. Skey. On the application of the hot blast to blow-pipe purposes. Scientific American, 23(14):209, October 1, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n14/pdf/scientificamerican10011870-209.pdf.

Anonymous:1870:BLR

[922] Anonymous. The Bridge de Laisne on the railway from Soissons to Laon, in France. Scientific American, 23(14):210, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-210.pdf.

Anonymous:1870:IPO

[923] Anonymous. Important Patent Office decisions. Scientific American, 23(14):211, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
211.pdf.

Anonymous:1870:BDC

[924] Anonymous. Bussell’s disk cultivator and harrow. Scientific American, 23(14):211, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
211a.pdf.

Anonymous:1870:NRS

[925] Anonymous. New relations of silica. Scientific American, 23(14):215, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-215.pdf.

Anonymous:1870:HPP

[926] Anonymous. History and progress of photography. Scientific American, 23(14):215, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
215a.pdf.

Anonymous:1870:UBW

[927] Anonymous. The use of balloons in warfare. Scientific American, 23(14):215, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
215b.pdf.

Anonymous:1870:FAIb

[928] Anonymous. Fair of the American institute. Scientific American, 23(14):216, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
216.pdf.

Anonymous:1870:RSN

[929] Anonymous. Resources of the South — notes of a trip from New York to Western Texas. Scientific American, 23(14):216–217, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-216a.pdf.

Anonymous:1870:NMMe

[930] Anonymous. New mechanical movements. Scientific American, 23 (14):217–218, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
217.pdf.

Anonymous:1870:SIx

[931] Anonymous. Scientific intelligence. Scientific American, 23(14):218, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n14/pdf/scientificamerican10011870-218.pdf.

Anonymous:1870:ACbl

[932] Anonymous. Answers to correspondents. Scientific American, 23(14):218, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
218b.pdf.

Anonymous:1870:RAFbm

[933] Anonymous. Recent American and foreign patents. Scientific American, 23(14):219, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
219.pdf.

Anonymous:1870:OLPbm

[934] Anonymous. Official list of patents. Scientific American, 23 (14):219–220, October 1, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n14/pdf/scientificamerican10011870-
219a.pdf.

Anonymous:1870:Cbm

[935] Anonymous. Correspondence. Scientific American, 23(15):228–229, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-228.pdf.

Anonymous:1870:BNBba

[936] Anonymous. Books: New books and publications. Scientific American, 23(15):234, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
234.pdf.

Anonymous:1870:BPbq

[937] Anonymous. Business and personal. Scientific American, 23 (15):234–237, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
234a.pdf.

Anonymous:1870:BFR

[938] Anonymous. Boiler-feed regulator, low-water alarm and steam trap. Scientific American, 23(15):223, October 8, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n15/pdf/scientificamerican10081870-223.pdf.

Anonymous:1870:IAPa

[939] Anonymous. Inaugural address of the President, Thomas H. Huxley, LL.D., F.R.S., Etc., before the British Association for the Advancement of Science. Scientific American, 23(15):224–225, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-224.pdf.

Anonymous:1870:MSG

[940] Anonymous. The manufacture of soluble glass. Scientific American, 23(15):225, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
225.pdf.

Anonymous:1870:SRG

[941] Anonymous. Screw reversing gear for locomotives. Scientific American, 23(15):226, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
226.pdf.

Anonymous:1870:HPG

[942] Anonymous. Hydro-pneumatic gun carriage. Scientific American, 23(15):227, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
227.pdf.

Anonymous:1870:PIAc

[943] Anonymous. Ponsard’s improvements in apparatus for puddling iron, etc. Scientific American, 23(15):227–228, October 8, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n15/pdf/scientificamerican10081870-227a.pdf.

Anonymous:1870:FHS

[944] Anonymous. Fortifications and history of Strasbourg. Scientific American, 23(15):229–230, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
229.pdf.

Anonymous:1870:NAE

[945] Anonymous. Naval architecture and engineering. Scientific American, 23(15):231, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
231.pdf.

Anonymous:1870:NAL

[946] Anonymous. A new artificial light. Scientific American, 23(15):231, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-231a.pdf.

Anonymous:1870:APb

[947] Anonymous. Affairs in Paris. Scientific American, 23(15):231, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-231b.pdf.

Anonymous:1870:SPc

[948] Anonymous. The Siege of Paris. Scientific American, 23(15):231, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-231c.pdf.

Anonymous:1870:CPP

[949] Anonymous. The cheap production of potash. Scientific American, 23(15):231–232, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
231d.pdf.

Anonymous:1870:BC

[950] Anonymous. Balancing cylinders. Scientific American, 23(15):232, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n15/pdf/scientificamerican10081870-232.pdf.

Anonymous:1870:FAIc

[951] Anonymous. Fair of the American institute. Scientific American, 23(15):232–234, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
232a.pdf.

Anonymous:1870:ACbm

[952] Anonymous. Answers to correspondents. Scientific American, 23(15):235, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
235.pdf.

Anonymous:1870:RAFbn

[953] Anonymous. Recent American and foreign patents. Scientific American, 23(15):235, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
235a.pdf.

Anonymous:1870:OLPbn

[954] Anonymous. Official list of patents. Scientific American, 23 (15):235–236, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
235b.pdf.

Anonymous:1870:AEPw

[955] Anonymous. Applications for the extension of patents. Scientific American, 23(15):236, October 8, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n15/pdf/scientificamerican10081870-
236.pdf.

Anonymous:1870:Cbn

[956] Anonymous. Correspondence. Scientific American, 23(16):243–244, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n16/pdf/scientificamerican10151870-243.pdf.

Anonymous:1870:BPbr

[957] Anonymous. Business and personal. Scientific American, 23 (16):250, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
250a.pdf.

Anonymous:1870:ISE

[958] Anonymous. Improved stump extractor, the “instrument room” of the electric telegraph, of London, and more. Scientific American, 23(16):239, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
239.pdf.

Anonymous:1870:IAPb

[959] Anonymous. Inaugural address of the President, Thomas H. Huxley, Ll.D., F.R.S., Etc., before the British Association for the Advancement of Science. Scientific American, 23(16):240–241, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n16/pdf/scientificamerican10151870-240.pdf.

Anonymous:1870:CAb

[960] Anonymous. Costume and art. Scientific American, 23(16):241, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n16/pdf/scientificamerican10151870-241.pdf.

Anonymous:1870:NCS

[961] Anonymous. New Capitol for the State of New York. Scientific American, 23(16):242–243, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
242.pdf.

Anonymous:1870:DP

[962] Anonymous. Design patents. Scientific American, 23(16):244, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n16/pdf/scientificamerican10151870-244.pdf.

Abbott:1870:HRM

[963] Jacob Abbott. How railroads are made. Scientific American, 23 (16):245–246, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
245.pdf.

Anonymous:1870:IHD

[964] Anonymous. Improved hand-drilling machine. Scientific American, 23(16):246, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
246.pdf.

Anonymous:1870:HC

[965] Anonymous. History of chloroform. Scientific American, 23(16): 246, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
246a.pdf.

Anonymous:1870:CPN

[966] Anonymous. Central Park. New York — report of the Commissioners. Scientific American, 23(16):247, October 15, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n16/pdf/scientificamerican10151870-247.pdf.

Anonymous:1870:WPR

[967] Anonymous. Woodbury’s photo-relief process. Scientific American, 23(16):247, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
247a.pdf.

Anonymous:1870:GBG

[968] Anonymous. The Gatling battery gun in England. Scientific American, 23(16):247–248, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
247b.pdf.

Anonymous:1870:CAM

[969] Anonymous. Comparative accuracy of mercurial and aneroid barometers. Scientific American, 23(16):248, October 15, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n16/pdf/scientificamerican10151870-248.pdf.

Anonymous:1870:FAId

[970] Anonymous. Fair of the American institute. Scientific American, 23(16):248–249, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
248a.pdf.

Anonymous:1870:WCc

[971] Anonymous. What a contrast. Scientific American, 23(16):249, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n16/pdf/scientificamerican10151870-249.pdf.

Anonymous:1870:LSE

[972] Anonymous. Letters from the South, etc. Scientific American, 23(16):249, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
249a.pdf.

Anonymous:1870:ACbn

[973] Anonymous. Answers to correspondents. Scientific American, 23 (16):250, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
250.pdf.

Anonymous:1870:RAFbo

[974] Anonymous. Recent American and foreign patents. Scientific American, 23(16):250–251, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
250b.pdf.

Anonymous:1870:OLPbo

[975] Anonymous. Official list of patents. Scientific American, 23(16): 251–252, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
251.pdf.

Anonymous:1870:AEPx

[976] Anonymous. Applications for the extension of patents. Scientific American, 23(16):252, October 15, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n16/pdf/scientificamerican10151870-
252.pdf.

Anonymous:1870:Cbo

[977] Anonymous. Correspondence. Scientific American, 23(17):260–262, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n17/pdf/scientificamerican10221870-260.pdf.

Anonymous:1870:BNBbb

[978] Anonymous. Books: New books and publications. Scientific American, 23(17):265, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
265a.pdf.

Anonymous:1870:BPbs

[979] Anonymous. Business and personal. Scientific American, 23 (17):266, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
266a.pdf.

Anonymous:1870:IAT

[980] Anonymous. Improved apparatus for tanning leather, hard on Evans, and more. Scientific American, 23(17):255, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n17/pdf/scientificamerican10221870-255.pdf.

Anonymous:1870:IDS

[981] Anonymous. Improved door and shutter bolt. Scientific American, 23(17):255, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
255a.pdf.

Anonymous:1870:EI

[982] Anonymous. Electrolytic insulation. Scientific American, 23(17): 256, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
256.pdf.

Anonymous:1870:EBB

[983] Anonymous. The excursion of the Boston Board of Trade over the Pacific Railroad — the Pullman Palace cars. Scientific American, 23(17):257, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
257.pdf.

Anonymous:1870:TP

[984] Anonymous. Transpiration in plants. Scientific American, 23 (17):258, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
258.pdf.

Anonymous:1870:HHC

[985] Anonymous. Hickman’s hand corn planter. Scientific American, 23(17):258, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
258a.pdf.

Anonymous:1870:WD

[986] Anonymous. Weatherhead’s dividers. Scientific American, 23(17): 258–259, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
258b.pdf.

Anonymous:1870:Me

[987] Anonymous. The mantis. Scientific American, 23(17):259, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n17/pdf/scientificamerican10221870-259.pdf.

Anonymous:1870:CES

[988] Anonymous. The carp enter self-heating fluting machine. Scientific American, 23(17):262, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
262.pdf.

Anonymous:1870:SCc

[989] Anonymous. Smoke consumption. Scientific American, 23(17):263, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n17/pdf/scientificamerican10221870-263.pdf.

Anonymous:1870:EAD

[990] Anonymous. Effect of artillery discharges upon weather. Scientific American, 23(17):263, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
263a.pdf.

Anonymous:1870:ASS

[991] Anonymous. Artificial stone — the Sorel process. Scientific American, 23(17):263, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
263b.pdf.

Anonymous:1870:PMC

[992] Anonymous. Progress in the manufacture of colors. Scientific American, 23(17):264, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
264.pdf.

Anonymous:1870:PDH

[993] Anonymous. Portable dwelling houses. Scientific American, 23 (17):264, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
264a.pdf.

Anonymous:1870:SIy

[994] Anonymous. Scientific intelligence. Scientific American, 23(17): 264–265, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
264b.pdf.

Anonymous:1870:FAIe

[995] Anonymous. Fair of the American institute. Scientific American, 23(17):265, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
265.pdf.

Anonymous:1870:AEPy

[996] Anonymous. Applications for the extension of patents. Scientific American, 23(17):265–266, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
265b.pdf.

Anonymous:1870:ACbo

[997] Anonymous. Answers to correspondents. Scientific American, 23 (17):266, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
266.pdf.

Anonymous:1870:RAFbp

[998] Anonymous. Recent American and foreign patents. Scientific American, 23(17):266–267, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
266b.pdf.

Anonymous:1870:OLPbp

[999] Anonymous. Official list of patents. Scientific American, 23(17): 267–268, October 22, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n17/pdf/scientificamerican10221870-
267.pdf.

Anonymous:1870:Cbp

[1000] Anonymous. Correspondence. Scientific American, 23(18):276, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-276.pdf.

Anonymous:1870:BNBbc

[1001] Anonymous. Books: New books and publications. Scientific American, 23(18):281–282, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
281b.pdf.

Anonymous:1870:BPbt

[1002] Anonymous. Business and personal. Scientific American, 23 (18):282, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
282.pdf.

Anonymous:1870:IDW

[1003] Anonymous. Improved dish-washing machine, how counterfeit notes are made, and more. Scientific American, 23(18):271, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-271.pdf.

Anonymous:1870:IC

[1004] Anonymous. The Island of Capri. Scientific American, 23(18):272, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-272.pdf.

Anonymous:1870:RL

[1005] Anonymous. Rolling loads. Scientific American, 23(18):272–273, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-272a.pdf.

Deacon:1870:EFM

[1006] C. F. Deacon. On the efficiency of furnaces and mechanical firing. Scientific American, 23(18):274, October 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n18/pdf/scientificamerican10291870-274.pdf.

Anonymous:1870:ATK

[1007] Anonymous. The art of tying knots. Scientific American, 23 (18):275, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
275.pdf.

Anonymous:1870:MFCa

[1008] Anonymous. The Mechanics’ Fair at Cincinnati, Ohio. Scientific American, 23(18):276–277, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
276a.pdf.

Anonymous:1870:EPb

[1009] Anonymous. Enlarging photographs. Scientific American, 23(18): 277–278, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
277.pdf.

Anonymous:1870:CSC

[1010] Anonymous. Conical surface condenser. Scientific American, 23 (18):278, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
278.pdf.

Anonymous:1870:CSH

[1011] Anonymous. Central shaft — Hoosac Tunnel. Scientific American, 23(18):279, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
279.pdf.

Anonymous:1870:MWF

[1012] Anonymous. Measurement of water flowing through pipes — water meters. Scientific American, 23(18):279, October 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n18/pdf/scientificamerican10291870-279a.pdf.

Anonymous:1870:CFAa

[1013] Anonymous. Chemistry at the fair of the American institute. Scientific American, 23(18):279–280, October 29, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n18/pdf/scientificamerican10291870-279b.pdf.

Anonymous:1870:SIz

[1014] Anonymous. Scientific intelligence. Scientific American, 23(18):280, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-280.pdf.

Anonymous:1870:FAIf

[1015] Anonymous. Fair of the American institute. Scientific American, 23(18):280, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
280a.pdf.

Anonymous:1870:LSa

[1016] Anonymous. Letters from the South. Scientific American, 23 (18):281, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
281.pdf.

Anonymous:1870:E

[1017] Anonymous. The earthquake. Scientific American, 23(18):281, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n18/pdf/scientificamerican10291870-281a.pdf.

Anonymous:1870:RAFbq

[1018] Anonymous. Recent American and foreign patents. Scientific American, 23(18):282–283, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
282a.pdf.

Anonymous:1870:OLPbq

[1019] Anonymous. Official list of patents. Scientific American, 23(18): 283–284, October 29, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n18/pdf/scientificamerican10291870-
283.pdf.

Anonymous:1870:Cbq

[1020] Anonymous. Correspondence. Scientific American, 23(19):291– 292, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
291a.pdf.

Anonymous:1870:BPbu

[1021] Anonymous. Business and personal. Scientific American, 23(19): 298–299, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
298a.pdf.

Anonymous:1870:BNBbd

[1022] Anonymous. Books: New books and publications. Scientific American, 23(19):300, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
300.pdf.

Anonymous:1870:SPI

[1023] Anonymous. Steam plow, on improved appliances for the production of heavy forgings. Scientific American, 23(19):287–288, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n19/pdf/scientificamerican11051870-287.pdf.

Anonymous:1870:LME

[1024] Anonymous. The labor movement in England. Scientific American, 23(19):288–289, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
288.pdf.

Gamgee:1870:CNA

[1025] John Gamgee. Chloralum — a new antiseptic. Scientific American, 23(19):289–290, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
289.pdf.

Anonymous:1870:DSEb

[1026] Anonymous. A dollar steam engine. Scientific American, 23(19): 290–291, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
290.pdf.

Anonymous:1870:IE

[1027] Anonymous. Improved elevator. Scientific American, 23(19): 291, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
291.pdf.

Prescott:1870:AT

[1028] George B. Prescott. Automatic telegraphy. Scientific American, 23 (19):292–293, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
292.pdf.

Gordon:1870:SU

[1029] J. R. Gordon. The sea urchin. Scientific American, 23(19): 293–294, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
293.pdf.

Anonymous:1870:CFAb

[1030] Anonymous. Chemistry at the fair of the American institute. Scientific American, 23(19):295, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
295.pdf.

Anonymous:1870:TIM

[1031] Anonymous. The Tellier ice machine. Scientific American, 23 (19):295–296, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
295a.pdf.

Anonymous:1870:WL

[1032] Anonymous. The wages of labor. Scientific American, 23(19): 296, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
296.pdf.

Anonymous:1870:SMd

[1033] Anonymous. A singular monstrosity. Scientific American, 23 (19):296, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
296a.pdf.

Anonymous:1870:SRB

[1034] Anonymous. A second roger bacon, and a century of inventions (almost). Scientific American, 23(19):296, November 5, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n19/pdf/scientificamerican11051870-296b.pdf.

Anonymous:1870:SIba

[1035] Anonymous. Scientific intelligence. Scientific American, 23(19): 296–297, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
296c.pdf.

Anonymous:1870:LSb

[1036] Anonymous. Letters from the South. Scientific American, 23 (19):297, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
297.pdf.

Anonymous:1870:VHB

[1037] Anonymous. The value of the honey bee in agriculture. Scientific American, 23(19):297–298, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
297a.pdf.

Anonymous:1870:ACbp

[1038] Anonymous. Answers to correspondents. Scientific American, 23 (19):298, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
298.pdf.

Anonymous:1870:RAFbr

[1039] Anonymous. Recent American and foreign patents. Scientific American, 23(19):299, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
299.pdf.

Anonymous:1870:OLPbr

[1040] Anonymous. Official list of patents. Scientific American, 23(19): 299–300, November 5, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n19/pdf/scientificamerican11051870-
299a.pdf.

Anonymous:1870:Cbr

[1041] Anonymous. Correspondence. Scientific American, 23(20):307– 308, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
307a.pdf.

Anonymous:1870:BPbv

[1042] Anonymous. Business and personal. Scientific American, 23 (20):314, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
314.pdf.

Anonymous:1870:ADR

[1043] Anonymous. Automatic damper regulator, important patent decision, and more. Scientific American, 23(20):303, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n20/pdf/scientificamerican11121870-303.pdf.

Anonymous:1870:TRT

[1044] Anonymous. Tripler’s railroad tie. Scientific American, 23(20): 303, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
303a.pdf.

Tyndall:1870:SUIa

[1045] John Tyndall. Scientific use of the imagination. Scientific American, 23(20):304–305, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
304.pdf.

Anonymous:1870:PMSb

[1046] Anonymous. The power of modern skill in the mechanic arts. Scientific American, 23(20):305–306, November 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n20/pdf/scientificamerican11121870-305.pdf.

Anonymous:1870:GSB

[1047] Anonymous. The great suspension bridge between New York and Brooklyn — progress of the work — its present condition — some account of the method of procedure. Scientific American, 23(20):307, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n20/pdf/scientificamerican11121870-307.pdf.

Anonymous:1870:Bd

[1048] Anonymous. Blackboards. Scientific American, 23(20):309–310, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n20/pdf/scientificamerican11121870-309.pdf.

Anonymous:1870:HLG

[1049] Anonymous. Hutchinson’s locking grate bar. Scientific American, 23(20):310, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
310.pdf.

Anonymous:1870:RSEb

[1050] Anonymous. Rating steam engines. Scientific American, 23(20): 311, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
311.pdf.

Anonymous:1870:PPMb

[1051] Anonymous. Power of prime motors. Scientific American, 23 (20):311, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
311a.pdf.

Anonymous:1870:PWA

[1052] Anonymous. The people who attend fairs. Scientific American, 23(20):311, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
311b.pdf.

Anonymous:1870:WWMb

[1053] Anonymous. The World’s work, and its moral necessity. Scientific American, 23(20):311–312, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
311c.pdf.

Matthiessen:1870:DTD

[1054] Augustus Matthiessen. Death of three distinguished chemists. Scientific American, 23(20):312, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
312.pdf.

Anonymous:1870:TFM

[1055] Anonymous. Transportation of fresh meats and fruits, etc., through long distances — the Davis refrigerator car. Scientific American, 23(20):312, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
312a.pdf.

Anonymous:1870:CFAc

[1056] Anonymous. Chemistry at the fair of the American institute. Scientific American, 23(20):312–313, November 12, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n20/pdf/scientificamerican11121870-312b.pdf.

Anonymous:1870:DMSb

[1057] Anonymous. Diamond mining in South Africa. Scientific American, 23(20):313–314, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
313.pdf.

Anonymous:1870:ACbq

[1058] Anonymous. Answers to correspondents. Scientific American, 23 (20):314, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
314a.pdf.

Anonymous:1870:AEPz

[1059] Anonymous. Applications for the extension of patents. Scientific American, 23(20):314, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
314b.pdf.

Anonymous:1870:RAFbs

[1060] Anonymous. Recent American and foreign patents. Scientific American, 23(20):315, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
315.pdf.

Anonymous:1870:OLPbs

[1061] Anonymous. Official list of patents. Scientific American, 23(20): 315–316, November 12, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n20/pdf/scientificamerican11121870-
315a.pdf.

Anonymous:1870:Cbs

[1062] Anonymous. Correspondence. Scientific American, 23(21):324, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n21/pdf/scientificamerican11191870-324.pdf.

Anonymous:1870:BPbw

[1063] Anonymous. Business and personal. Scientific American, 23 (21):330, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
330.pdf.

Anonymous:1870:BNBbe

[1064] Anonymous. Books: New books and publications. Scientific American, 23(21):332, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
332.pdf.

Anonymous:1870:IRS

[1065] Anonymous. Improvement in road steamers. Scientific American, 23(21):319, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
319.pdf.

Tyndall:1870:SUIb

[1066] John Tyndall. Scientific use of the imagination. Scientific American, 23(21):320–321, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
320.pdf.

Anonymous:1870:PTL

[1067] Anonymous. Professor Tyndall’s lecture on electrical phenomena. Scientific American, 23(21):321–322, November 19, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n21/pdf/scientificamerican11191870-321.pdf.

Anonymous:1870:S

[1068] Anonymous. The stereoscope. Scientific American, 23(21):322– 324, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
322.pdf.

Anonymous:1870:PIAd

[1069] Anonymous. Progress of invention abroad. Scientific American, 23(21):325, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
325.pdf.

Anonymous:1870:SIbb

[1070] Anonymous. Scientific intelligence. Scientific American, 23(21): 325, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
325a.pdf.

Anonymous:1870:DSC

[1071] Anonymous. Device for spreading circular saw teeth. Scientific American, 23(21):326, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
326.pdf.

Anonymous:1870:PEL

[1072] Anonymous. Power’s endless ladder hod elevator. Scientific American, 23(21):326, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
326a.pdf.

Anonymous:1870:BSS

[1073] Anonymous. Bourne’s spherical steam engine governor. Scientific American, 23(21):326, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
326b.pdf.

Anonymous:1870:HLS

[1074] Anonymous. How long shall nitro-glycerin continue its work of destruction? — the Fairport explosion. Scientific American, 23 (21):327, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
327.pdf.

Anonymous:1870:RRM

[1075] Anonymous. Roads and road-making. Scientific American, 23 (21):327, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
327a.pdf.

Anonymous:1870:MPE

[1076] Anonymous. The motive power of expanding gases. Scientific American, 23(21):327–328, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
327b.pdf.

Anonymous:1870:TD

[1077] Anonymous. Time and distance. Scientific American, 23(21): 328, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
328.pdf.

Anonymous:1870:RCF

[1078] Anonymous. Resignation of Commissioner Fisher. Scientific American, 23(21):328, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
328a.pdf.

Anonymous:1870:AAP

[1079] Anonymous. Attempt to abolish the Patent Office. Scientific American, 23(21):328, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
328b.pdf.

Anonymous:1870:LSc

[1080] Anonymous. Letters from the South. Scientific American, 23(21): 328–329, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
328c.pdf.

Anonymous:1870:RBS

[1081] Anonymous. Razing band saws. Scientific American, 23(21): 329, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
329.pdf.

Anonymous:1870:Qa

[1082] Anonymous. Queries. Scientific American, 23(21):329, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n21/pdf/scientificamerican11191870-329a.pdf.

Anonymous:1870:ACbr

[1083] Anonymous. Answers to correspondents. Scientific American, 23 (21):329–330, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
329b.pdf.

Anonymous:1870:RAFbt

[1084] Anonymous. Recent American and foreign patents. Scientific American, 23(21):330–331, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
330a.pdf.

Anonymous:1870:OLPbt

[1085] Anonymous. Official list of patents. Scientific American, 23(21): 331–332, November 19, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n21/pdf/scientificamerican11191870-
331.pdf.

Anonymous:1870:Cbt

[1086] Anonymous. Correspondence. Scientific American, 23(22):339– 340, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
339.pdf.

Anonymous:1870:BPbx

[1087] Anonymous. Business and personal. Scientific American, 23(22): 345–346, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
345a.pdf.

Anonymous:1870:BNBbf

[1088] Anonymous. Books: New books and publications. Scientific American, 23(22):346, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
346.pdf.

Anonymous:1870:IISa

[1089] Anonymous. Improved independent shuttle-motion loom, how to save peach trees. Scientific American, 23(22):335, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n22/pdf/scientificamerican11261870-335.pdf.

Tyndall:1870:SUIc

[1090] John Tyndall. Scientific use of the imagination. Scientific American, 23(22):336–337, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
336.pdf.

Lewis:1870:NMA

[1091] Dio Lewis. Notes and maxims about health. Scientific American, 23(22):337, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
337.pdf.

Anonymous:1870:SRF

[1092] Anonymous. Steam railroad ferries between England and France. Scientific American, 23(22):338–339, November 26, 1870. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v23/
n22/pdf/scientificamerican11261870-338.pdf.

Anonymous:1870:BAM

[1093] Anonymous. The balloon as an aid to meteorological research. Scientific American, 23(22):341, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
341.pdf.

Anonymous:1870:WWP

[1094] Anonymous. The Wyckoff wood pavement. Scientific American, 23(22):342, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
342.pdf.

Anonymous:1870:SPF

[1095] Anonymous. Strong’s patent fireplace. Scientific American, 23 (22):342, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
342a.pdf.

Anonymous:1870:PCM

[1096] Anonymous. Pierced cores in metal casting. Scientific American, 23(22):342, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
342b.pdf.

Anonymous:1870:IBS

[1097] Anonymous. Improved breast-strap protector for harness. Scientific American, 23(22):342, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
342c.pdf.

Anonymous:1870:PMb

[1098] Anonymous. Perpetual motion. Scientific American, 23(22): 342, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
342d.pdf.

Anonymous:1870:SAW

[1099] Anonymous. The “Scientific American” — what it has done and what it expects to do. Scientific American, 23(22):343, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n22/pdf/scientificamerican11261870-343.pdf.

Anonymous:1870:BRSc

[1100] Anonymous. Beet-root sugar — how the war is likely to affect the manufacture in France. Scientific American, 23(22):343, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n22/pdf/scientificamerican11261870-343a.pdf.

Anonymous:1870:SVB

[1101] Anonymous. Shad vs. black Bass. Scientific American, 23(22): 343, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
343b.pdf.

Anonymous:1870:PBc

[1102] Anonymous. The principle of bracing. Scientific American, 23 (22):343–344, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
343c.pdf.

Anonymous:1870:PMd

[1103] Anonymous. Pyrotechnic mixtures. Scientific American, 23(22): 344, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
344.pdf.

Anonymous:1870:ARI

[1104] Anonymous. Application of Ruhmkorff’s induction coil to the copying of drawings. Scientific American, 23(22):344, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n22/pdf/scientificamerican11261870-344a.pdf.

Anonymous:1870:LSd

[1105] Anonymous. Letters from the South. Scientific American, 23(22): 344–345, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
344b.pdf.

Anonymous:1870:AEPba

[1106] Anonymous. Applications for the extension of patents. Scientific American, 23(22):345, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
345.pdf.

Anonymous:1870:Qb

[1107] Anonymous. Queries. Scientific American, 23(22):346, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n22/pdf/scientificamerican11261870-346a.pdf.

Anonymous:1870:ACbs

[1108] Anonymous. Answers to correspondents. Scientific American, 23 (22):346, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
346b.pdf.

Anonymous:1870:RAFbu

[1109] Anonymous. Recent American and foreign patents. Scientific American, 23(22):346–347, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
346c.pdf.

Anonymous:1870:OLPbu

[1110] Anonymous. Official list of patents. Scientific American, 23(22): 347–348, November 26, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n22/pdf/scientificamerican11261870-
347.pdf.

Anonymous:1870:Cbu

[1111] Anonymous. Correspondence. Scientific American, 23(23):356, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-356b.pdf.

Anonymous:1870:BNBbg

[1112] Anonymous. Books: New books and publications. Scientific American, 23(23):362, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
362.pdf.

Anonymous:1870:BPby

[1113] Anonymous. Business and personal. Scientific American, 23 (23):362, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
362a.pdf.

Anonymous:1870:MCP

[1114] Anonymous. Molding, carving, and paneling machine, Mount Seward in the Adirondacks, and more. Scientific American, 23(23):351, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-351.pdf.

Anonymous:1870:IMb

[1115] Anonymous. Iron and its manufacture. Scientific American, 23 (23):352–355, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
352.pdf.

Anonymous:1870:PMf

[1116] Anonymous. Perpetual motion. Scientific American, 23(23):355– 356, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
355.pdf.

Anonymous:1870:HLWb

[1117] Anonymous. High and low water alarm for steam boilers. Scientific American, 23(23):356, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
356.pdf.

Anonymous:1870:DSH

[1118] Anonymous. Device for sharpening horseshoe calks. Scientific American, 23(23):356, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
356a.pdf.

Parton:1870:SFB

[1119] James Parton. Samuel F. B. Morse. Scientific American, 23(23): 357–358, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
357.pdf.

Anonymous:1870:AMIb

[1120] Anonymous. Apparatus for making and inhaling oxygen. Scientific American, 23(23):358, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
358.pdf.

Anonymous:1870:SAc

[1121] Anonymous. Scientific American. Scientific American, 23(23):359, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-359.pdf.

Anonymous:1870:HB

[1122] Anonymous. Heating of buildings. Scientific American, 23(23):359, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-359a.pdf.

Anonymous:1870:MCG

[1123] Anonymous. To make a cheap and good small ice house. Scientific American, 23(23):359, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
359b.pdf.

Anonymous:1870:BEb

[1124] Anonymous. Boiler explosions. Scientific American, 23(23):359– 360, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
359c.pdf.

Anonymous:1870:BEc

[1125] Anonymous. A brutal exhibition. Scientific American, 23(23):360, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-360.pdf.

Anonymous:1870:SSe

[1126] Anonymous. Spiritualism and science. Scientific American, 23 (23):360–361, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
360a.pdf.

Anonymous:1870:SIbc

[1127] Anonymous. Scientific intelligence. Scientific American, 23(23):361, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-361.pdf.

Anonymous:1870:LSf

[1128] Anonymous. Letters from the south. Scientific American, 23 (23):361, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
361a.pdf.

Anonymous:1870:Qc

[1129] Anonymous. Queries. Scientific American, 23(23):362, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n23/pdf/scientificamerican12031870-362b.pdf.

Anonymous:1870:ACbt

[1130] Anonymous. Answers to correspondents. Scientific American, 23 (23):362–363, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
362c.pdf.

Anonymous:1870:RAFbv

[1131] Anonymous. Recent American and foreign patents. Scientific American, 23(23):363, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
363.pdf.

Anonymous:1870:OLPbv

[1132] Anonymous. Official list of patents. Scientific American, 23(23): 363–364, December 3, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n23/pdf/scientificamerican12031870-
363a.pdf.

Anonymous:1870:Cbv

[1133] Anonymous. Correspondence. Scientific American, 23(25):388– 390, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
388a.pdf.

Anonymous:1870:BNBbh

[1134] Anonymous. Books: New books and publications. Scientific American, 23(25):394, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
394.pdf.

Anonymous:1870:BPbz

[1135] Anonymous. Business and personal. Scientific American, 23 (25):395, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
395.pdf.

Anonymous:1870:ISLa

[1136] Anonymous. Improved street letter-box for lamp posts, history of carpeting, and more. Scientific American, 23(25):383–385, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n25/pdf/scientificamerican12171870-383.pdf.

Anonymous:1870:PMg

[1137] Anonymous. Perpetual motion. Scientific American, 23(25):385– 386, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
385.pdf.

Anonymous:1870:ASD

[1138] Anonymous. Adjustable stand for draftsmen. Scientific American, 23 (25):386–387, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
386.pdf.

Anonymous:1870:TE

[1139] Anonymous. Tunnel excavator. Scientific American, 23(25): 387, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
387.pdf.

Anonymous:1870:EA

[1140] Anonymous. Exercising apparatus. Scientific American, 23(25): 387, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
387a.pdf.

Anonymous:1870:SPd

[1141] Anonymous. Siemens’ pyrometer. Scientific American, 23(25): 387–388, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
387b.pdf.

Anonymous:1870:DDT

[1142] Anonymous. Dr. Doremus on the triumphs of science. Scientific American, 23(25):388, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
388.pdf.

Anonymous:1870:ISLb

[1143] Anonymous. Illinois and St. Louis Bridge Company — report of Capt. James B. Eads, Chief Engineer. Scientific American, 23 (25):391, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
391.pdf.

Anonymous:1870:EHP

[1144] Anonymous. Explosions from hydraulic pressure. Scientific American, 23(25):391, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
391a.pdf.

Anonymous:1870:AW

[1145] Anonymous. Artesian wells. Scientific American, 23(25):391, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n25/pdf/scientificamerican12171870-391b.pdf.

Anonymous:1870:SAd

[1146] Anonymous. Scientific administration. Scientific American, 23(25): 391–392, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
391c.pdf.

Anonymous:1870:GBP

[1147] Anonymous. The great British problem. Scientific American, 23 (25):392, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
392.pdf.

Anonymous:1870:FPP

[1148] Anonymous. The foremanizing process for preserving timber, the victims of its poisonous effects, and the suits at law which have been instituted to recover damages. Scientific American, 23(25):392, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n25/pdf/scientificamerican12171870-392a.pdf.

Anonymous:1870:SIbd

[1149] Anonymous. Scientific intelligence. Scientific American, 23(25): 392–393, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
392b.pdf.

Anonymous:1870:SAe

[1150] Anonymous. Scientific American. Scientific American, 23(25): 393, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
393.pdf.

Anonymous:1870:Qd

[1151] Anonymous. Queries. Scientific American, 23(25):394, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n25/pdf/scientificamerican12171870-394a.pdf.

Anonymous:1870:ACbu

[1152] Anonymous. Answers to correspondents. Scientific American, 23 (25):394–395, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
394b.pdf.

Anonymous:1870:RAFbw

[1153] Anonymous. Recent American and foreign patents. Scientific American, 23(25):395, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
395a.pdf.

Anonymous:1870:OLPbw

[1154] Anonymous. Official list of patents. Scientific American, 23(25): 395–396, December 17, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n25/pdf/scientificamerican12171870-
395b.pdf.

Anonymous:1870:Cbw

[1155] Anonymous. Correspondence. Scientific American, 23(26):404, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n26/pdf/scientificamerican12241870-404.pdf.

Anonymous:1870:BNBbi

[1156] Anonymous. Books: New books and publications. Scientific American, 23(26):408, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
408b.pdf.

Anonymous:1870:Id

[1157] Anonymous. Index. Scientific American, 23(26):410–411, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n26/pdf/scientificamerican12241870-410b.pdf.

Anonymous:1870:BPca

[1158] Anonymous. Business and personal. Scientific American, 23 (26):412, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
412.pdf.

Anonymous:1870:IISb

[1159] Anonymous. Improved independent shuttle-motion loom. Scientific American, 23(26):399, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
399.pdf.

Anonymous:1870:PFIb

[1160] Anonymous. Progress of foreign invention. Scientific American, 23 (26):400–401, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
400.pdf.

Anonymous:1870:MPV

[1161] Anonymous. Mercurial pump without valves or stop-cocks. Scientific American, 23(26):402, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
402.pdf.

Anonymous:1870:MSP

[1162] Anonymous. Milking-stool and pail-holder. Scientific American, 23(26):402, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
402a.pdf.

Anonymous:1870:IMW

[1163] Anonymous. Improved mop wringer. Scientific American, 23 (26):402, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
402b.pdf.

Anonymous:1870:NIP

[1164] Anonymous. A new ink for printers. Scientific American, 23 (26):402, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
402c.pdf.

Anonymous:1870:PMh

[1165] Anonymous. Perpetual motion. Scientific American, 23(26):402– 404, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
402d.pdf.

Main:1870:SAS

[1166] T. Main. Successful application of steam to canal boats. Scientific American, 23(26):405, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
405.pdf.

Anonymous:1870:EM

[1167] Anonymous. Eyesight and the microscope. Scientific American, 23 (26):405–406, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
405a.pdf.

Anonymous:1870:GS

[1168] Anonymous. Gas stove. Scientific American, 23(26):406, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v23/n26/pdf/scientificamerican12241870-406.pdf.

Anonymous:1870:ABF

[1169] Anonymous. Automatic boiler feeder. Scientific American, 23 (26):406, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
406a.pdf.

Anonymous:1870:PFc

[1170] Anonymous. Proprietors and foremen. Scientific American, 23 (26):407, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407.pdf.

Anonymous:1870:HVM

[1171] Anonymous. How vermicelli and macaroni are made. Scientific American, 23(26):407, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407a.pdf.

Anonymous:1870:CVXb

[1172] Anonymous. Close of volume XXIII. Scientific American, 23 (26):407, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407b.pdf.

Anonymous:1870:RSSb

[1173] Anonymous. Removing street snow and ice. Scientific American, 23(26):407, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407c.pdf.

Anonymous:1870:SMAb

[1174] Anonymous. Success as the measure of ability. Scientific American, 23(26):407, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407d.pdf.

Anonymous:1870:RMC

[1175] Anonymous. The relation of mineralogy to chemistry. Scientific American, 23(26):407–408, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
407e.pdf.

Anonymous:1870:SAf

[1176] Anonymous. Scientific American. Scientific American, 23(26): 408, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
408.pdf.

Anonymous:1870:TS

[1177] Anonymous. Timely suggestions. Scientific American, 23(26): 408, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
408a.pdf.

Anonymous:1870:RAFbx

[1178] Anonymous. Recent American and foreign patents. Scientific American, 23(26):408–409, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
408c.pdf.

Anonymous:1870:OLPbx

[1179] Anonymous. Official list of patents. Scientific American, 23(26): 409–410, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
409.pdf.

Anonymous:1870:AEPbb

[1180] Anonymous. Applications for extension of patents. Scientific American, 23(26):410, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
410.pdf.

Anonymous:1870:ACbv

[1181] Anonymous. Answers to correspondents. Scientific American, 23 (26):410, December 24, 1870. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v23/n26/pdf/scientificamerican12241870-
410a.pdf.

Anonymous:1871:Ca

[1182] Anonymous. Correspondence. Scientific American, 24(1):4, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-4.pdf.

Anonymous:1871:PNPa

[1183] Anonymous. Patent notes: Patents. Scientific American, 24(1):9, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-9c.pdf.

Anonymous:1871:BPa

[1184] Anonymous. Business and personal. Scientific American, 24(1):10, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-10.pdf.

Anonymous:1871:THH

[1185] Anonymous. Thomas Henry Huxley. Scientific American, 24(1):1–2, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-1.pdf.

Anonymous:1871:MOW

[1186] Anonymous. Machine for oiling wool. Scientific American, 24(1):2, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-2.pdf.

Anonymous:1871:PMa

[1187] Anonymous. Perpetual motion. Scientific American, 24(1):2, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-2a.pdf.

Anonymous:1871:ICC

[1188] Anonymous. Improved chimney cleaner. Scientific American, 24(1):2–3, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-2b.pdf.

Anonymous:1871:CHG

[1189] Anonymous. Carbureted hydrogen gas. Scientific American, 24(1):3–4, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-3.pdf.

Craig:1871:AT

[1190] D. H. Craig. Automatic telegraphy. Scientific American, 24(1):4, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-4a.pdf.

Anonymous:1871:DDT

[1191] Anonymous. Dr. Doremus on the triumphs of science. Scientific American, 24(1):5, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n1/pdf/scientificamerican01011871-
5.pdf.

Anonymous:1871:AIL

[1192] Anonymous. American institute lectures. — the struggles of science. Scientific American, 24(1):5–6, January 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n1/pdf/scientificamerican01011871-5a.pdf.

Anonymous:1871:SMPa

[1193] Anonymous. Six months of progress in engineering. Scientific American, 24(1):7, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n1/pdf/scientificamerican01011871-
7.pdf.

Anonymous:1871:PCSa

[1194] Anonymous. Progress of chemical science. Scientific American, 24(1):7, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-7a.pdf.

Anonymous:1871:SED

[1195] Anonymous. The solar eclipse of 22d December, 1870. Scientific American, 24(1):7–8, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n1/pdf/scientificamerican01011871-
7b.pdf.

Anonymous:1871:TSBa

[1196] Anonymous. Testing the strength of boilers by steam pressure. Scientific American, 24(1):8, January 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n1/pdf/scientificamerican01011871-8.pdf.

Anonymous:1871:SPT

[1197] Anonymous. Standard public time. Scientific American, 24(1):8, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-8a.pdf.

Anonymous:1871:UD

[1198] Anonymous. The uses of dissociation. Scientific American, 24(1):8, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-8b.pdf.

Anonymous:1871:FEG

[1199] Anonymous. Fatal explosion of a gasoline reservoir — a warning to mechanics. Scientific American, 24(1):8–9, January 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n1/pdf/scientificamerican01011871-8c.pdf.

Anonymous:1871:HB

[1200] Anonymous. Harvey Brown. Scientific American, 24(1):9, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-9.pdf.

Anonymous:1871:SAa

[1201] Anonymous. Scientific American 1871. Scientific American, 24(1):9, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-9a.pdf.

Anonymous:1871:Qa

[1202] Anonymous. Queries. Scientific American, 24(1):9, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-9b.pdf.

Anonymous:1871:ACa

[1203] Anonymous. Answers to correspondents. Scientific American, 24(1):9, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-9d.pdf.

Anonymous:1871:RAFa

[1204] Anonymous. Recent American and foreign patents. Scientific American, 24(1):10–11, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n1/pdf/scientificamerican01011871-
10a.pdf.

Anonymous:1871:OLPa

[1205] Anonymous. Official list of patents. Scientific American, 24(1):11–12, January 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n1/pdf/scientificamerican01011871-11.pdf.

Anonymous:1871:Cb

[1206] Anonymous. Correspondence. Scientific American, 24(2):20–21, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-20.pdf.

Anonymous:1871:BPb

[1207] Anonymous. Business and personal. Scientific American, 24(2):26, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-26.pdf.

Anonymous:1871:MCL

[1208] Anonymous. Mammoth cylinder lithographic machine. Scientific American, 24(2):15, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
15.pdf.

Anonymous:1871:HWS

[1209] Anonymous. How we stand and walk. Scientific American, 24(2):16, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-16.pdf.

Anonymous:1871:MSSa

[1210] Anonymous. Manufacture of saws in Sheffield. Scientific American, 24(2):16–18, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
16a.pdf.

Anonymous:1871:IWW

[1211] Anonymous. Improvement in water wheels. Scientific American, 24(2):18, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
18.pdf.

Anonymous:1871:PMb

[1212] Anonymous. Perpetual motion. Scientific American, 24(2):19, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-19.pdf.

Anonymous:1871:HSJ

[1213] Anonymous. The hail-storm of June 20, 1870. Scientific American, 24(2):19, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
19a.pdf.

Anonymous:1871:IGV

[1214] Anonymous. Implement for grinding valves. Scientific American, 24(2):19–20, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
19b.pdf.

Anonymous:1871:ELI

[1215] Anonymous. Eastern Long Island — menhaden oil and the fisheries. Scientific American, 24(2):21, January 7, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n2/pdf/scientificamerican01071871-21.pdf.

Winder:1871:WCA

[1216] Daniel Knode Winder. What causes Aurora Borealis? Scientific American, 24(2):21, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
21a.pdf.

Anonymous:1871:ISLa

[1217] Anonymous. Improved self-locking stove leg. Scientific American, 24(2):22, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
22.pdf.

Anonymous:1871:GC

[1218] Anonymous. Gun cotton. Scientific American, 24(2):22, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-22a.pdf.

Anonymous:1871:CPP

[1219] Anonymous. Chubbuck’s patent piston and packing. Scientific American, 24(2):22, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
22b.pdf.

Anonymous:1871:BSC

[1220] Anonymous. Booton’s shot case and distributer. Scientific American, 24(2):22, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
22c.pdf.

Anonymous:1871:WLI

[1221] Anonymous. Where is the limit of invention? Scientific American, 24(2):23, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
23.pdf.

Anonymous:1871:PMD

[1222] Anonymous. The profession of the mechanical or dynamical engineer. Scientific American, 24(2):23, January 7, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n2/pdf/scientificamerican01071871-23a.pdf.

Anonymous:1871:MCT

[1223] Anonymous. The mont cenis tunnel completed. Scientific American, 24(2):23, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
23b.pdf.

Anonymous:1871:Pa

[1224] Anonymous. Pavements. Scientific American, 24(2):23–24, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-23c.pdf.

Anonymous:1871:ALa

[1225] Anonymous. Air light. Scientific American, 24(2):24, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-24.pdf.

Anonymous:1871:WLP

[1226] Anonymous. Waste liquors of paper mills. Scientific American, 24(2):24, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-24a.pdf.

Anonymous:1871:SSa

[1227] Anonymous. Spiritualism and science. Scientific American, 24(2):24, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-24b.pdf.

Anonymous:1871:ALP

[1228] Anonymous. Alloy of lead and platinum. Scientific American, 24(2):25, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-25.pdf.

Anonymous:1871:SAb

[1229] Anonymous. Scientific American 1871. Scientific American, 24(2):25, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-25a.pdf.

Anonymous:1871:Qb

[1230] Anonymous. Queries. Scientific American, 24(2):25, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-25b.pdf.

Anonymous:1871:ACb

[1231] Anonymous. Answers to correspondents. Scientific American, 24(2):26, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-26a.pdf.

Anonymous:1871:AEPa

[1232] Anonymous. Applications for extension of patents. Scientific American, 24(2):26, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
26b.pdf.

Anonymous:1871:RAFb

[1233] Anonymous. Recent American and foreign patents. Scientific American, 24(2):26–27, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n2/pdf/scientificamerican01071871-
26c.pdf.

Anonymous:1871:OLPb

[1234] Anonymous. Official list of patents. Scientific American, 24(2):27–28, January 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n2/pdf/scientificamerican01071871-27.pdf.

Anonymous:1871:Cc

[1235] Anonymous. Correspondence. Scientific American, 24(3):35–38, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-35.pdf.

Anonymous:1871:BNBa

[1236] Anonymous. Books: New books and publications. Scientific American, 24(3):42, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
42.pdf.

Anonymous:1871:BPc

[1237] Anonymous. Business and personal. Scientific American, 24(3):42, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-42a.pdf.

Anonymous:1871:ICPa

[1238] Anonymous. Improved cotton press, boring machines at the mont cenis tunnel. Scientific American, 24(3):31, January 14, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n3/pdf/scientificamerican01141871-31.pdf.

Anonymous:1871:SNC

[1239] Anonymous. Stow’s non-corroding water and gas pipe. Scientific American, 24(3):31, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
31a.pdf.

Anonymous:1871:WSa

[1240] Anonymous. Washing shield. Scientific American, 24(3):31, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-31b.pdf.

Anonymous:1871:TM

[1241] Anonymous. Talking machines. Scientific American, 24(3):32–33, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-32.pdf.

Anonymous:1871:PMc

[1242] Anonymous. Perpetual motion. Scientific American, 24(3):34, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-34.pdf.

Anonymous:1871:ASDa

[1243] Anonymous. Anti-snoring device. Scientific American, 24(3):34, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-34a.pdf.

Anonymous:1871:IHPa

[1244] Anonymous. Improved hitching post. Scientific American, 24 (3):34, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
34b.pdf.

Anonymous:1871:PPH

[1245] Anonymous. Paint-pot holder. Scientific American, 24(3):34, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-34c.pdf.

Anonymous:1871:MIWa

[1246] Anonymous. Merriman’s improved water-proof dress and life preserver. Scientific American, 24(3):38, January 14, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n3/pdf/scientificamerican01141871-38.pdf.

Anonymous:1871:NCPa

[1247] Anonymous. The new Commissioner of Patents. Scientific American, 24(3):39, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
39.pdf.

Anonymous:1871:PFA

[1248] Anonymous. The proposed fair of the American institute in 1876. Scientific American, 24(3):39, January 14, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n3/pdf/scientificamerican01141871-39a.pdf.

Anonymous:1871:MP

[1249] Anonymous. Mechanical power. Scientific American, 24(3):39, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-39b.pdf.

Anonymous:1871:OTS

[1250] Anonymous. Object teaching and science. Scientific American, 24(3):39–40, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
39c.pdf.

Anonymous:1871:SSb

[1251] Anonymous. Social science. Scientific American, 24(3):40, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-40.pdf.

Anonymous:1871:FHO

[1252] Anonymous. The failure of the Hon. Oakes Ames. Scientific American, 24(3):40, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
40a.pdf.

Anonymous:1871:RPSa

[1253] Anonymous. Railroad progress in the Southern States. Scientific American, 24(3):40–41, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
40b.pdf.

Anonymous:1871:BU

[1254] Anonymous. On blood and its uses. Scientific American, 24(3):41, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-41.pdf.

Anonymous:1871:Qc

[1255] Anonymous. Queries. Scientific American, 24(3):41, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n3/pdf/scientificamerican01141871-41a.pdf.

Anonymous:1871:ACc

[1256] Anonymous. Answers to correspondents. Scientific American, 24 (3):42, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
42b.pdf.

Anonymous:1871:RAFc

[1257] Anonymous. Recent American and foreign patents. Scientific American, 24(3):42–43, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
42c.pdf.

Anonymous:1871:OLPc

[1258] Anonymous. Official list of patents. Scientific American, 24 (3):43–44, January 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n3/pdf/scientificamerican01141871-
43.pdf.

Anonymous:1871:Cd

[1259] Anonymous. Correspondence. Scientific American, 24(4):52, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-52.pdf.

Anonymous:1871:BPd

[1260] Anonymous. Business and personal. Scientific American, 24 (4):58–59, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
58a.pdf.

Anonymous:1871:BNBb

[1261] Anonymous. Books: New books and publications. Scientific American, 24(4):60, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
60.pdf.

Anonymous:1871:WPV

[1262] Anonymous. Watts and Phelps’ variable cut-off, French ironing table. Scientific American, 24(4):47, January 21, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n4/pdf/scientificamerican01211871-47.pdf.

Anonymous:1871:X

[1263] Anonymous. Xylonite. Scientific American, 24(4):48, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-48.pdf.

Anonymous:1871:TSa

[1264] Anonymous. The triumphs of surgery. Scientific American, 24 (4):48–49, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
48a.pdf.

Anonymous:1871:IDG

[1265] Anonymous. Improved door guard. Scientific American, 24(4):50, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-50.pdf.

Anonymous:1871:PMe

[1266] Anonymous. Perpetual motion. Scientific American, 24(4):50–51, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-50a.pdf.

Anonymous:1871:WSF

[1267] Anonymous. Wire stretcher for fence building. Scientific American, 24(4):51, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
51.pdf.

Anonymous:1871:ITa

[1268] Anonymous. Intrenching tool. Scientific American, 24(4):51, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-51a.pdf.

Anonymous:1871:PHB

[1269] Anonymous. Post-hole borer. Scientific American, 24(4):51, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-51b.pdf.

Anonymous:1871:ST

[1270] Anonymous. Shoemaker’s tool. Scientific American, 24(4):51, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-51c.pdf.

Anonymous:1871:IDF

[1271] Anonymous. An idle day at Fall River, Mass. Scientific American, 24(4):53–54, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
53.pdf.

Anonymous:1871:FFA

[1272] Anonymous. Facts and figures about the mont cenis tunnel. Scientific American, 24(4):55, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
55.pdf.

Anonymous:1871:WOM

[1273] Anonymous. What ought a mechanic to know? Scientific American, 24(4):55, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
55a.pdf.

Anonymous:1871:RM

[1274] Anonymous. Railroad monopolies. Scientific American, 24(4): 55–56, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
55b.pdf.

Anonymous:1871:AGW

[1275] Anonymous. American gas wells. Scientific American, 24(4):56, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-56.pdf.

Anonymous:1871:USSa

[1276] Anonymous. The U.S. Signal Service and storm signals. Scientific American, 24(4):56, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
56a.pdf.

Anonymous:1871:NCPb

[1277] Anonymous. New chemical products. Scientific American, 24 (4):56, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
56b.pdf.

Anonymous:1871:VLV

[1278] Anonymous. Vapor lamps and vapor stoves. Scientific American, 24(4):56, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
56c.pdf.

Anonymous:1871:PFIa

[1279] Anonymous. Progress of foreign invention. Scientific American, 24(4):57–58, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
57.pdf.

Anonymous:1871:IPDa

[1280] Anonymous. Important patent decision. Scientific American, 24 (4):58, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
58.pdf.

Anonymous:1871:ACd

[1281] Anonymous. Answers to correspondents. Scientific American, 24 (4):59, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
59.pdf.

Anonymous:1871:RAFd

[1282] Anonymous. Recent American and foreign patents. Scientific American, 24(4):59, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
59a.pdf.

Anonymous:1871:AEPb

[1283] Anonymous. Applications for extension of patents. Scientific American, 24(4):59, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n4/pdf/scientificamerican01211871-
59b.pdf.

Anonymous:1871:OLPd

[1284] Anonymous. Official list of patents. Scientific American, 24(4):60, January 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n4/pdf/scientificamerican01211871-60a.pdf.

Anonymous:1871:Ce

[1285] Anonymous. Correspondence. Scientific American, 24(5):68–70, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-68.pdf.

Anonymous:1871:BPe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-74a.pdf.

Anonymous:1871:BNBc

[1287] Anonymous. Books: New books and publications. Scientific American, 24(5):75, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
75.pdf.

Anonymous:1871:PNPb

[1288] Anonymous. Patent notes: Patents. Scientific American, 24(5):75, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-75b.pdf.

Anonymous:1871:IVP

[1289] Anonymous. Improved vertical portable engine, the place of the mitrailleur in war. Scientific American, 24(5):63, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-63.pdf.

Anonymous:1871:OEH

[1290] Anonymous. Our eyes, and how to take care of them. Scientific American, 24(5):64, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
64.pdf.

Anonymous:1871:RFW

[1291] Anonymous. The rattan factories at Wakefield Mass. Scientific American, 24(5):64–65, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
64a.pdf.

Anonymous:1871:CT

[1292] Anonymous. Combined tool. Scientific American, 24(5):66, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-66.pdf.

Anonymous:1871:CICa

[1293] Anonymous. Cast iron chimney. Scientific American, 24(5):66, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-66a.pdf.

Anonymous:1871:AT

[1294] Anonymous. Automatic toy. Scientific American, 24(5):66, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-66b.pdf.

Anonymous:1871:IPa

[1295] Anonymous. Improved pliers. Scientific American, 24(5):66–67, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-66c.pdf.

Anonymous:1871:PMf

[1296] Anonymous. Perpetual motion. Scientific American, 24(5):67–68, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-67.pdf.

Anonymous:1871:IG

[1297] Anonymous. Improved gate. Scientific American, 24(5):70, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-70.pdf.

Anonymous:1871:NWG

[1298] Anonymous. Narrow and wide gages. Scientific American, 24 (5):71, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
71.pdf.

Anonymous:1871:HMa

[1299] Anonymous. Hints to mechanics. Scientific American, 24(5):71, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-71a.pdf.

Anonymous:1871:CVF

[1300] Anonymous. The conveyance of various fluids through pipes. Scientific American, 24(5):71–72, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
71b.pdf.

Anonymous:1871:IICa

[1301] Anonymous. International industrial competition. Scientific American, 24(5):72, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
72.pdf.

Anonymous:1871:NU

[1302] Anonymous. Naphthaline and its uses. Scientific American, 24 (5):72, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
72a.pdf.

Anonymous:1871:RL

[1303] Anonymous. Railroad litigation. Scientific American, 24(5):72, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-72b.pdf.

Anonymous:1871:NYPa

[1304] Anonymous. New York photograph galleries. Scientific American, 24(5):72–73, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
72c.pdf.

Anonymous:1871:PPa

[1305] Anonymous. The present and the past. Scientific American, 24(5):73, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
73.pdf.

Anonymous:1871:OHB

[1306] Anonymous. Obituary — Henry Burden of Troy, N.Y. Scientific American, 24(5):73–74, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
73a.pdf.

Anonymous:1871:Qd

[1307] Anonymous. Queries. Scientific American, 24(5):74, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n5/pdf/scientificamerican01281871-74.pdf.

Anonymous:1871:ACe

[1308] Anonymous. Answers to correspondents. Scientific American, 24 (5):74, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
74b.pdf.

Anonymous:1871:AEPc

[1309] Anonymous. Applications for extension of patents. Scientific American, 24(5):75, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
75a.pdf.

Anonymous:1871:RAFe

[1310] Anonymous. Recent American and foreign patents. Scientific American, 24(5):75, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
75c.pdf.

Anonymous:1871:OLPe

[1311] Anonymous. Official list of patents. Scientific American, 24 (5):75–76, January 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n5/pdf/scientificamerican01281871-
75d.pdf.

Anonymous:1871:Cf

[1312] Anonymous. Correspondence. Scientific American, 24(6):84–86, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-84.pdf.

Anonymous:1871:BPf

[1313] Anonymous. Business and personal. Scientific American, 24 (6):91, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
91.pdf.

Anonymous:1871:BNBd

[1314] Anonymous. Books: New books and publications. Scientific American, 24(6):91, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
91c.pdf.

Anonymous:1871:UWH

[1315] Anonymous. Utilizing waste heat from steam engines and boilers. Scientific American, 24(6):79, February 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n6/pdf/scientificamerican02041871-79.pdf.

Anonymous:1871:WWD

[1316] Anonymous. The water we drink. Scientific American, 24(6):80, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-80.pdf.

Anonymous:1871:YD

[1317] Anonymous. The year and the day. Scientific American, 24 (6):81, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
81.pdf.

Anonymous:1871:IEA

[1318] Anonymous. Improved externally adjustable packing for pistons. Scientific American, 24(6):82, February 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n6/pdf/scientificamerican02041871-82.pdf.

Anonymous:1871:SM

[1319] Anonymous. Scrubbing machine. Scientific American, 24(6):82, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-82a.pdf.

Anonymous:1871:PMg

[1320] Anonymous. Perpetual motion. Scientific American, 24(6):82– 83, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
82b.pdf.

Anonymous:1871:QCa

[1321] Anonymous. Quartz crusher. Scientific American, 24(6):83, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-83.pdf.

Anonymous:1871:ARP

[1322] Anonymous. Action of the reciprocating parts of steam engines, and its influence on the problem of high piston-speed. Scientific American, 24(6):83–84, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
83a.pdf.

Anonymous:1871:ISMa

[1323] Anonymous. Improved sewing machine. Scientific American, 24 (6):86, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
86.pdf.

Anonymous:1871:EAS

[1324] Anonymous. English and American scientific, mechanical, and engineering journalism. Scientific American, 24(6):87, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-87.pdf.

Anonymous:1871:SG

[1325] Anonymous. Soluble glass. Scientific American, 24(6):87, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-87a.pdf.

Anonymous:1871:TSM

[1326] Anonymous. Type-setting machines. Scientific American, 24(6): 87, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
87b.pdf.

Anonymous:1871:UH

[1327] Anonymous. The uses of habit. Scientific American, 24(6):87, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-87c.pdf.

Anonymous:1871:SVC

[1328] Anonymous. The scientific value of the Central Park. Scientific American, 24(6):88, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
88.pdf.

Anonymous:1871:EEa

[1329] Anonymous. The Eclipse Expeditions. Scientific American, 24 (6):88, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
88a.pdf.

Anonymous:1871:PPb

[1330] Anonymous. The present and the past. Scientific American, 24 (6):88–89, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
88b.pdf.

Anonymous:1871:TMS

[1331] Anonymous. Thatch made by sewing machines. Scientific American, 24(6):89, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
89.pdf.

Anonymous:1871:OHD

[1332] Anonymous. Obituary. — Hon. David Lyman. Scientific American, 24(6):89, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
89a.pdf.

Anonymous:1871:CU

[1333] Anonymous. Cryolite and its uses. Scientific American, 24(6):89, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-89b.pdf.

Anonymous:1871:RBW

[1334] Anonymous. A remarkable body of water. Scientific American, 24(6):90, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
90.pdf.

Anonymous:1871:ACf

[1335] Anonymous. Answers to correspondents. Scientific American, 24 (6):90, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
90a.pdf.

Anonymous:1871:Qe

[1336] Anonymous. Queries. Scientific American, 24(6):91, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n6/pdf/scientificamerican02041871-91a.pdf.

Anonymous:1871:RAFf

[1337] Anonymous. Recent American and foreign patents. Scientific American, 24(6):91, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
91b.pdf.

Anonymous:1871:OLPf

[1338] Anonymous. Official list of patents. Scientific American, 24 (6):91–92, February 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n6/pdf/scientificamerican02041871-
91d.pdf.

Anonymous:1871:Cg

[1339] Anonymous. Correspondence. Scientific American, 24(7):99– 100, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
99b.pdf.

Anonymous:1871:BNBe

[1340] Anonymous. Books: New books and publications. Scientific American, 24(7):105–106, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
105a.pdf.

Anonymous:1871:BPg

[1341] Anonymous. Business and personal. Scientific American, 24(7): 106, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
106.pdf.

Anonymous:1871:CMS

[1342] Anonymous. Cochrane’s Mariner’s sound indicator, purifying gas and soap limes. Scientific American, 24(7):95, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n7/pdf/scientificamerican02111871-95.pdf.

Anonymous:1871:DCR

[1343] Anonymous. Device for closing ruptures in fire hose. Scientific American, 24(7):95, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
95a.pdf.

Anonymous:1871:WVP

[1344] Anonymous. A woman’s view of the Patent Office. Scientific American, 24(7):96, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
96.pdf.

Anonymous:1871:IB

[1345] Anonymous. The incrustation of boilers. Scientific American, 24(7):97, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
97.pdf.

Anonymous:1871:LFI

[1346] Anonymous. The land of fire and ice. Scientific American, 24 (7):97–98, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
97a.pdf.

Anonymous:1871:PMh

[1347] Anonymous. Perpetual motion. Scientific American, 24(7):98– 99, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
98.pdf.

Anonymous:1871:ICF

[1348] Anonymous. Improved cistern filter. Scientific American, 24(7): 99, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
99.pdf.

Anonymous:1871:PSa

[1349] Anonymous. Pruning shears. Scientific American, 24(7):99, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n7/pdf/scientificamerican02111871-99a.pdf.

Anonymous:1871:RS

[1350] Anonymous. A romance of science. Scientific American, 24(7): 101–102, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
101.pdf.

Anonymous:1871:ICA

[1351] Anonymous. Improved coal ash sifter. Scientific American, 24 (7):102, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
102.pdf.

Anonymous:1871:CLO

[1352] Anonymous. Cod-liver oil. Scientific American, 24(7):102, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n7/pdf/scientificamerican02111871-102a.pdf.

Anonymous:1871:IWC

[1353] Anonymous. An invention wanted to clean the streets of snow. solution of the problem to be found in steam. Scientific American, 24(7):103, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
103.pdf.

Anonymous:1871:AP

[1354] Anonymous. The adulteration of petroleum. Scientific American, 24(7):103, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
103a.pdf.

Anonymous:1871:TWG

[1355] Anonymous. “and there were giants in those days” — the largest inventor yet — a most remarkable family of gigantic turks. Scientific American, 24(7):103, February 11, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n7/pdf/scientificamerican02111871-103b.pdf.

Anonymous:1871:EET

[1356] Anonymous. Excavation and embankment tables. Scientific American, 24(7):103, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
103c.pdf.

Anonymous:1871:ACg

[1357] Anonymous. The alloys of copper. Scientific American, 24(7): 103–104, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
103d.pdf.

Anonymous:1871:DP

[1358] Anonymous. The downfall of Paris. Scientific American, 24 (7):104, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
104.pdf.

Anonymous:1871:BLS

[1359] Anonymous. Boynton’s lightning saw. Scientific American, 24 (7):104, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
104a.pdf.

Anonymous:1871:PPc

[1360] Anonymous. The present and the past. Scientific American, 24 (7):104–105, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
104b.pdf.

Anonymous:1871:SIa

[1361] Anonymous. Scientific intelligence. Scientific American, 24(7): 105, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
105.pdf.

Anonymous:1871:ACh

[1362] Anonymous. Answers to correspondents. Scientific American, 24 (7):106, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
106a.pdf.

Anonymous:1871:Qf

[1363] Anonymous. Queries. Scientific American, 24(7):106–107, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n7/pdf/scientificamerican02111871-106b.pdf.

Anonymous:1871:RAFg

[1364] Anonymous. Recent American and foreign patents. Scientific American, 24(7):107, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
107.pdf.

Anonymous:1871:OLPg

[1365] Anonymous. Official list of patents. Scientific American, 24(7): 107–108, February 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n7/pdf/scientificamerican02111871-
107a.pdf.

Anonymous:1871:Ch

[1366] Anonymous. Correspondence. Scientific American, 24(8):115– 118, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
115b.pdf.

Anonymous:1871:BPh

[1367] Anonymous. Business and personal. Scientific American, 24(8): 122, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
122.pdf.

Anonymous:1871:IWG

[1368] Anonymous. Improved water grate and steam generator, how eyes are made. Scientific American, 24(8):111–112, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-111.pdf.

Anonymous:1871:EWW

[1369] Anonymous. Enameling wood work. Scientific American, 24(8): 112–113, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
112.pdf.

Anonymous:1871:NSa

[1370] Anonymous. A new stone. Scientific American, 24(8):113, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-113.pdf.

Anonymous:1871:ECa

[1371] Anonymous. Exercising club. Scientific American, 24(8):114, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-114.pdf.

Anonymous:1871:PPd

[1372] Anonymous. The present and the past. Scientific American, 24 (8):114, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
114a.pdf.

Anonymous:1871:KFF

[1373] Anonymous. Knapp’s fluting and flat iron. Scientific American, 24(8):114, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
114b.pdf.

Anonymous:1871:APY

[1374] Anonymous. Agricultural patents of the year. Scientific American, 24(8):115, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
115.pdf.

Anonymous:1871:PMi

[1375] Anonymous. Perpetual motion. Scientific American, 24(8):115, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-115a.pdf.

Anonymous:1871:KM

[1376] Anonymous. Kerosene murder. Scientific American, 24(8):119, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-119.pdf.

Anonymous:1871:OMW

[1377] Anonymous. Organic matter in water. Scientific American, 24 (8):119, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
119a.pdf.

Anonymous:1871:SCM

[1378] Anonymous. Stone cutting by machinery. Scientific American, 24(8):119, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
119b.pdf.

Anonymous:1871:EDC

[1379] Anonymous. The electro deposition of copper and brass. Scientific American, 24(8):119–120, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
119c.pdf.

Anonymous:1871:UA

[1380] Anonymous. The uses of apatite. Scientific American, 24(8):120, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-120.pdf.

Anonymous:1871:ACDa

[1381] Anonymous. The average city dwelling house. Scientific American, 24(8):120–121, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
120a.pdf.

Anonymous:1871:ARC

[1382] Anonymous. Annual report of Commissioner of Patents. Scientific American, 24(8):121, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
121.pdf.

Anonymous:1871:NPLa

[1383] Anonymous. New patent law of 1870. Scientific American, 24 (8):121–122, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
121a.pdf.

Anonymous:1871:ACi

[1384] Anonymous. Answers to correspondents. Scientific American, 24 (8):122, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
122a.pdf.

Anonymous:1871:Qg

[1385] Anonymous. Queries. Scientific American, 24(8):122, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n8/pdf/scientificamerican02181871-122b.pdf.

Anonymous:1871:RAFh

[1386] Anonymous. Recent American and foreign patents. Scientific American, 24(8):123, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
123.pdf.

Anonymous:1871:AEPd

[1387] Anonymous. Applications for extension of patents. Scientific American, 24(8):123, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
123a.pdf.

Anonymous:1871:OLPh

[1388] Anonymous. Official list of patents. Scientific American, 24 (8):123, February 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n8/pdf/scientificamerican02181871-
123b.pdf.

Anonymous:1871:Ci

[1389] Anonymous. Correspondence. Scientific American, 24(9):132– 133, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
132.pdf.

Anonymous:1871:BPi

[1390] Anonymous. Business and personal. Scientific American, 24(9): 138, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
138a.pdf.

Anonymous:1871:IET

[1391] Anonymous. Improved extension table. Scientific American, 24 (9):127, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
127.pdf.

Anonymous:1871:EMR

[1392] Anonymous. Elastic metallic rocking chair. Scientific American, 24(9):127, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
127a.pdf.

Anonymous:1871:PPe

[1393] Anonymous. Poetry and patents. Scientific American, 24(9):128, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-128.pdf.

Anonymous:1871:FAR

[1394] Anonymous. Facts about the River Amazon. — steam navigation and railroads. Scientific American, 24(9):128, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-128a.pdf.

Anonymous:1871:ABN

[1395] Anonymous. The ancient brewers of New York. Scientific American, 24(9):129, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
129.pdf.

Anonymous:1871:OG

[1396] Anonymous. Oak graining. Scientific American, 24(9):129, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-129a.pdf.

Anonymous:1871:DRa

[1397] Anonymous. Dyers’ recipes. Scientific American, 24(9):129–130, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-129b.pdf.

Anonymous:1871:Pb

[1398] Anonymous. Pyrometer. Scientific American, 24(9):130, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-130.pdf.

Anonymous:1871:PMj

[1399] Anonymous. Perpetual motion. Scientific American, 24(9):131, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-131.pdf.

Anonymous:1871:WMW

[1400] Anonymous. W. Mattieu Williams on the Bessemer process. Scientific American, 24(9):131–132, February 25, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n9/pdf/scientificamerican02251871-131a.pdf.

Anonymous:1871:ISS

[1401] Anonymous. Interesting school statistics. Scientific American, 24 (9):133–134, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
133.pdf.

Anonymous:1871:HSMa

[1402] Anonymous. Hand sewing machine. Scientific American, 24(9): 134, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
134.pdf.

Anonymous:1871:HRR

[1403] Anonymous. The Hudson River Railroad disaster. Scientific American, 24(9):135, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
135.pdf.

Anonymous:1871:TRG

[1404] Anonymous. Toughening and refining gold. Scientific American, 24(9):135, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
135a.pdf.

Anonymous:1871:EMa

[1405] Anonymous. Electro-metallurgy. Scientific American, 24(9):135, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-135b.pdf.

Anonymous:1871:LGW

[1406] Anonymous. The largest gun in the world. Scientific American, 24(9):135, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
135c.pdf.

Anonymous:1871:RGD

[1407] Anonymous. A remarkable geological discovery. Scientific American, 24(9):136, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
136.pdf.

Anonymous:1871:PPEa

[1408] Anonymous. The pecuniary prospects of the East River Bridge. Scientific American, 24(9):136, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
136a.pdf.

Anonymous:1871:ILA

[1409] Anonymous. Intelligent legislation about medical prescriptions. Scientific American, 24(9):136, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
136b.pdf.

Anonymous:1871:HIB

[1410] Anonymous. How the ice bridge is formed in the East River. Scientific American, 24(9):136, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
136c.pdf.

Anonymous:1871:PFIb

[1411] Anonymous. Progress of foreign invention. Scientific American, 24 (9):136–137, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
136d.pdf.

Anonymous:1871:PPf

[1412] Anonymous. The present and the past. Scientific American, 24 (9):137, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
137.pdf.

Anonymous:1871:SIb

[1413] Anonymous. Scientific intelligence. Scientific American, 24(9): 137, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
137a.pdf.

Anonymous:1871:RDP

[1414] Anonymous. Recent decisions at the Patent Office. Scientific American, 24(9):137–138, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
137b.pdf.

Anonymous:1871:ACj

[1415] Anonymous. Answers to correspondents. Scientific American, 24 (9):138, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
138.pdf.

Anonymous:1871:Qh

[1416] Anonymous. Queries. Scientific American, 24(9):138, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n9/pdf/scientificamerican02251871-138b.pdf.

Anonymous:1871:NBPa

[1417] Anonymous. New books and publications. Scientific American, 24(9):138, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
138c.pdf.

Anonymous:1871:AEPe

[1418] Anonymous. Applications for extension of patents. Scientific American, 24(9):139, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
139.pdf.

Anonymous:1871:NPLb

[1419] Anonymous. New patent law of 1870. Scientific American, 24 (9):139, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
139a.pdf.

Anonymous:1871:RAFi

[1420] Anonymous. Recent American and foreign patents. Scientific American, 24(9):139, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
139b.pdf.

Anonymous:1871:OLPi

[1421] Anonymous. Official list of patents. Scientific American, 24(9): 139–140, February 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n9/pdf/scientificamerican02251871-
139c.pdf.

Anonymous:1871:Cj

[1422] Anonymous. Correspondence. Scientific American, 24(10):147–148, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-147.pdf.

Anonymous:1871:BPj

[1423] Anonymous. Business and personal. Scientific American, 24(10):154, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-154.pdf.

Anonymous:1871:BNBf 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
155.pdf.

Anonymous:1871:PNPc

[1425] Anonymous. Patent notes: Patents. Scientific American, 24(10):155, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-155a.pdf.

Anonymous:1871:IBS

[1426] Anonymous. Improved balanced slide valve, curious case of asymmetry. Scientific American, 24(10):143, March 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n10/pdf/scientificamerican03041871-143.pdf.

Anonymous:1871:MFa

[1427] Anonymous. Measuring faucet. Scientific American, 24(10):143, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-143b.pdf.

Anonymous:1871:OHY

[1428] Anonymous. One hundred years ago. Scientific American, 24(10):144, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-144.pdf.

Crace:1871:WIC

[1429] J. D. Crace. Wood and ivory carving. Scientific American, 24 (10):144–145, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
144a.pdf.

Anonymous:1871:PMk

[1430] Anonymous. Perpetual motion. Scientific American, 24(10):146, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-146.pdf.

Anonymous:1871:ICPb

[1431] Anonymous. Improvements in casting pulleys and grooved rolls. Scientific American, 24(10):146, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
146a.pdf.

Anonymous:1871:IPW

[1432] Anonymous. Infantile power washing machine. Scientific American, 24(10):146, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
146b.pdf.

Anonymous:1871:PRT

[1433] Anonymous. On the proper resistance of telegraphic relays. Scientific American, 24(10):148, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
148.pdf.

Colton:1871:PPO

[1434] H. E. Colton. Peanuts and peanut oil. Scientific American, 24(10):149, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-149.pdf.

Anonymous:1871:DRb

[1435] Anonymous. Dyers’ recipes. Scientific American, 24(10):149–150, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-149a.pdf.

Anonymous:1871:IHPb

[1436] Anonymous. Improved horse-power hayfork. Scientific American, 24(10):150, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
150.pdf.

Anonymous:1871:MEa

[1437] Anonymous. The mechanical equation. Scientific American, 24(10):151, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-151.pdf.

Anonymous:1871:DPO

[1438] Anonymous. Dangers of petroleum oils. report of the Health Board of New Orleans. A law regulating the sale of such oils recommended. Scientific American, 24(10):151, March 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n10/pdf/scientificamerican03041871-151a.pdf.

Anonymous:1871:GBM

[1439] Anonymous. A Government Bureau of Mines. Scientific American, 24(10):151, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
151b.pdf.

Anonymous:1871:NGRa

[1440] Anonymous. Narrow gage railways. Scientific American, 24(10): 151–152, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
151c.pdf.

Anonymous:1871:AEE

[1441] Anonymous. The American Eclipse Expedition. Scientific American, 24(10):152, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
152.pdf.

Anonymous:1871:PFL

[1442] Anonymous. Peat fuel for locomotives. Scientific American, 24(10):152, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-152a.pdf.

Anonymous:1871:LC

[1443] Anonymous. The lifting cure. Scientific American, 24(10):152, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-152b.pdf.

Anonymous:1871:ACP

[1444] Anonymous. Alaska. — climate and products. Scientific American, 24(10):152, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
152c.pdf.

Anonymous:1871:SSBa

[1445] Anonymous. Statistics of steam boiler explosions for the year 1870. Scientific American, 24(10):153–154, March 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n10/pdf/scientificamerican03041871-153.pdf.

Anonymous:1871:ACk

[1446] Anonymous. Answers to correspondents. Scientific American, 24 (10):154, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
154a.pdf.

Anonymous:1871:Qi

[1447] Anonymous. Queries. Scientific American, 24(10):154, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n10/pdf/scientificamerican03041871-154b.pdf.

Anonymous:1871:RAFj

[1448] Anonymous. Recent American and foreign patents. Scientific American, 24(10):154–155, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
154c.pdf.

Anonymous:1871:OLPj

[1449] Anonymous. Official list of patents. Scientific American, 24 (10):155–156, March 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n10/pdf/scientificamerican03041871-
155b.pdf.

Anonymous:1871:Ck

[1450] Anonymous. Correspondence. Scientific American, 24(11):164, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-164.pdf.

Anonymous:1871:BPk

[1451] Anonymous. Business and personal. Scientific American, 24(11):170, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-170.pdf.

Anonymous:1871:ICH

[1452] Anonymous. Improved corn-husking machine. Scientific American, 24(11):159, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
159.pdf.

Anonymous:1871:RJ

[1453] Anonymous. Railroad jack. Scientific American, 24(11):159–160, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-159a.pdf.

Anonymous:1871:AAG

[1454] Anonymous. Adulteration of articles in general use. Scientific American, 24(11):160–161, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
160.pdf.

Anonymous:1871:SIc

[1455] Anonymous. Scientific intelligence. Scientific American, 24(11): 161–162, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
161.pdf.

Anonymous:1871:PSb

[1456] Anonymous. Planking screw. Scientific American, 24(11):162–163, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-162.pdf.

Anonymous:1871:IWM

[1457] Anonymous. Improved washing machine. Scientific American, 24(11):163, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
163.pdf.

Anonymous:1871:RRa

[1458] Anonymous. Rambles for relics. Scientific American, 24(11):164–165, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-164a.pdf.

Anonymous:1871:PFIc

[1459] Anonymous. Progress of foreign invention. Scientific American, 24(11):165, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
165.pdf.

Anonymous:1871:DRc

[1460] Anonymous. Dyers’ recipes. Scientific American, 24(11):165–166, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-165a.pdf.

Anonymous:1871:TPC

[1461] Anonymous. The transmission of power by compressed air. Scientific American, 24(11):167, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
167.pdf.

Anonymous:1871:PEMa

[1462] Anonymous. Payne’s electro-magnetic motor. Scientific American, 24(11):167, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
167a.pdf.

Anonymous:1871:ECU

[1463] Anonymous. Effect of cold upon iron and steel. Scientific American, 24(11):167–168, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
167b.pdf.

Anonymous:1871:MCR

[1464] Anonymous. Mr. Cooper’s recent gift to the mechanics of New York. Scientific American, 24(11):168, March 11, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n11/pdf/scientificamerican03111871-168.pdf.

Anonymous:1871:DTB

[1465] Anonymous. Death of Thomas Brassey, the great English railway contractor. Scientific American, 24(11):168, March 11, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n11/pdf/scientificamerican03111871-168a.pdf.

Anonymous:1871:PPg

[1466] Anonymous. The present and the past. Scientific American, 24 (11):168–169, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
168b.pdf.

Anonymous:1871:PTM

[1467] Anonymous. A philanthropic temperance man. Scientific American, 24(11):169, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
169.pdf.

Anonymous:1871:NPLc

[1468] Anonymous. New patent law of 1870. Scientific American, 24(11):169, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-169a.pdf.

Anonymous:1871:NBPb

[1469] Anonymous. New books and publications. Scientific American, 24(11):169, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
169b.pdf.

Anonymous:1871:Qj

[1470] Anonymous. Queries. Scientific American, 24(11):169, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n11/pdf/scientificamerican03111871-169c.pdf.

Anonymous:1871:ACl

[1471] Anonymous. Answers to correspondents. Scientific American, 24(11):170, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
170a.pdf.

Anonymous:1871:RAFk

[1472] Anonymous. Recent American and foreign patents. Scientific American, 24(11):170–171, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
170b.pdf.

Anonymous:1871:OLPk

[1473] Anonymous. Official list of patents. Scientific American, 24 (11):171–174, March 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n11/pdf/scientificamerican03111871-
171.pdf.

Anonymous:1871:Cl

[1474] Anonymous. Correspondence. Scientific American, 24(12):179–180, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-179b.pdf.

Anonymous:1871:BNBg 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
185b.pdf.

Anonymous:1871:BPl

[1476] Anonymous. Business and personal. Scientific American, 24(12):186, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-186.pdf.

Anonymous:1871:IICb

[1477] Anonymous. The influence of intense cold on steel and iron. Scientific American, 24(12):176, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
176.pdf.

Anonymous:1871:OGO

[1478] Anonymous. Oak graining in oil colors. Scientific American, 24 (12):176–177, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
176a.pdf.

Anonymous:1871:KS

[1479] Anonymous. Knots and splices. Scientific American, 24(12):177–178, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-177.pdf.

Anonymous:1871:PWC

[1480] Anonymous. Portable writing and copying case. Scientific American, 24(12):178, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
178.pdf.

Anonymous:1871:IWT

[1481] Anonymous. Impregnating wood with tar or other preserving material. Scientific American, 24(12):178, March 18, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n12/pdf/scientificamerican03181871-178a.pdf.

Anonymous:1871:BCT

[1482] Anonymous. Boardman’s combined tool. Scientific American, 24(12):179, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
179.pdf.

Anonymous:1871:BTa

[1483] Anonymous. Belt tightener. Scientific American, 24(12):179, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-179a.pdf.

Anonymous:1871:RRb

[1484] Anonymous. Rambles for relics. Scientific American, 24(12):180, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-180.pdf.

Andrews:1871:SC

[1485] W. V. Andrews. Silk culture. Scientific American, 24(12):181–182, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-181.pdf.

Anonymous:1871:CTR

[1486] Anonymous. Combined trunk and rocking chair. Scientific American, 24(12):182, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
182.pdf.

Anonymous:1871:SID

[1487] Anonymous. Smith’s infant dining chair. Scientific American, 24(12):182, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
182a.pdf.

Anonymous:1871:BVM

[1488] Anonymous. Barnes’ ventilator for mattresses, etc. Scientific American, 24(12):182, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
182b.pdf.

Anonymous:1871:STA

[1489] Anonymous. A scientific and technical awakening. Scientific American, 24(12):183, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
183.pdf.

Anonymous:1871:SP

[1490] Anonymous. The Sherman process. Scientific American, 24(12):183, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-183a.pdf.

Anonymous:1871:RTT

[1491] Anonymous. Rubber tires for traction engines. Scientific American, 24(12):183–184, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
183b.pdf.

Anonymous:1871:CSH

[1492] Anonymous. Central shaft of the Hoosac Tunnel. Scientific American, 24(12):184, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
184.pdf.

Anonymous:1871:MAN

[1493] Anonymous. A museum of art and natural history. Scientific American, 24(12):184, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
184a.pdf.

Anonymous:1871:RJG

[1494] Anonymous. Report of the judges of group 1, department V. of the Exhibition of the American Institute for 1870. the Allen engine. Scientific American, 24(12):184, March 18, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n12/pdf/scientificamerican03181871-184b.pdf.

Anonymous:1871:LNHa

[1495] Anonymous. Lyceum of natural history. Scientific American, 24(12):184, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
184c.pdf.

Anonymous:1871:WVR

[1496] Anonymous. Warming and ventilating railroad cars. Scientific American, 24(12):184–185, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
184d.pdf.

Anonymous:1871:MRM

[1497] Anonymous. The mineral resources of Missouri. Scientific American, 24(12):185, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
185.pdf.

Anonymous:1871:SIe

[1498] Anonymous. Scientific intelligence. Scientific American, 24(12):185, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-185a.pdf.

Anonymous:1871:ACm

[1499] Anonymous. Answers to correspondents. Scientific American, 24(12):186, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
186a.pdf.

Anonymous:1871:AEPf

[1500] Anonymous. Applications for extension of patents. Scientific American, 24(12):186, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
186b.pdf.

Anonymous:1871:RAFl

[1501] Anonymous. Recent American and foreign patents. Scientific American, 24(12):187, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
187.pdf.

Anonymous:1871:Qk

[1502] Anonymous. Queries. Scientific American, 24(12):187, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n12/pdf/scientificamerican03181871-187a.pdf.

Anonymous:1871:OLPl

[1503] Anonymous. Official list of patents. Scientific American, 24 (12):187–188, March 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n12/pdf/scientificamerican03181871-
187b.pdf.

Anonymous:1871:Cm

[1504] Anonymous. Correspondence. Scientific American, 24(13):195–198, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-195c.pdf.

Anonymous:1871:BNBh

[1505] Anonymous. Books: New books and publications. Scientific American, 24(13):201, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
201a.pdf.

Anonymous:1871:BPm

[1506] Anonymous. Business and personal. Scientific American, 24(13):201, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-201b.pdf.

Anonymous:1871:IHO

[1507] Anonymous. Improvement in heavy ordnance, conducting wires for military telegraphs. Scientific American, 24(13):191, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-191.pdf.

Anonymous:1871:CMM

[1508] Anonymous. The condensed milk manufacture. Scientific American, 24(13):192, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
192.pdf.

Anonymous:1871:DRd

[1509] Anonymous. Dyers’ recipes. Scientific American, 24(13):193, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-193.pdf.

Davidson:1871:FB

[1510] R. O. Davidson. The flight of birds. Scientific American, 24(13):194, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-194.pdf.

Anonymous:1871:SSSa

[1511] Anonymous. Spring steps for stairways. Scientific American, 24(13):195, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
195.pdf.

Anonymous:1871:AIS

[1512] Anonymous. Archibald’s improved spoked wheel. Scientific American, 24(13):195, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
195a.pdf.

Anonymous:1871:IEb

[1513] Anonymous. Improvement in envelopes. Scientific American, 24 (13):195, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
195b.pdf.

Anonymous:1871:RTN

[1514] Anonymous. Railway transit for New York. Scientific American, 24(13):199, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
199.pdf.

Anonymous:1871:RB

[1515] Anonymous. Rating of boilers. Scientific American, 24(13):199, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-199a.pdf.

Anonymous:1871:TPU

[1516] Anonymous. Thallium, its properties and uses. Scientific American, 24(13):199–200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
199b.pdf.

Anonymous:1871:UMS

[1517] Anonymous. The use of the microscope in the study of rocks. Scientific American, 24(13):200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
200.pdf.

Anonymous:1871:HMHa

[1518] Anonymous. The human mind and the human hand. Scientific American, 24(13):200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
200a.pdf.

Anonymous:1871:WC

[1519] Anonymous. Wood carpeting. Scientific American, 24(13):200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-200b.pdf.

Anonymous:1871:CWN

[1520] Anonymous. Cost of water in New York City. Scientific American, 24(13):200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
200c.pdf.

Anonymous:1871:NMC

[1521] Anonymous. On the nature and manufacture of charcoal. Scientific American, 24(13):200, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
200d.pdf.

Anonymous:1871:RPDa

[1522] Anonymous. Recent patent decision. Scientific American, 24(13):201, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-201.pdf.

Anonymous:1871:ACn

[1523] Anonymous. Answers to correspondents. Scientific American, 24(13):201, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
201c.pdf.

Anonymous:1871:Ql

[1524] Anonymous. Queries. Scientific American, 24(13):201, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n13/pdf/scientificamerican03251871-201d.pdf.

Anonymous:1871:RAFm

[1525] Anonymous. Recent American and foreign patents. Scientific American, 24(13):202, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
202.pdf.

Anonymous:1871:AEPg

[1526] Anonymous. Applications for extension of patents. Scientific American, 24(13):203, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
203.pdf.

Anonymous:1871:OLPm

[1527] Anonymous. Official list of patents. Scientific American, 24 (13):203–204, March 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n13/pdf/scientificamerican03251871-
203a.pdf.

Anonymous:1871:Cn

[1528] Anonymous. Correspondence. Scientific American, 24(14):212, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-212.pdf.

Anonymous:1871:BPn

[1529] Anonymous. Business and personal. Scientific American, 24(14):218, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-218a.pdf.

Anonymous:1871:IHS

[1530] Anonymous. Improved horizontal steam engine, potato diggers. Scientific American, 24(14):207, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
207.pdf.

Anonymous:1871:AJ

[1531] Anonymous. Artificial jewelry. Scientific American, 24(14):208, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-208.pdf.

Barnard:1871:AAC

[1532] F. A. P. Barnard. Annual assay of the coinage. Scientific American, 24(14):209, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
209.pdf.

Anonymous:1871:HDY

[1533] Anonymous. How do you prove your plumb rule? Scientific American, 24(14):209, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
209a.pdf.

Day:1871:LM

[1534] E. C. H. Day. The “luna” moth. Scientific American, 24(14):210–211, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-210.pdf.

Anonymous:1871:ET

[1535] Anonymous. Egg tongs. Scientific American, 24(14):211, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-211.pdf.

Anonymous:1871:DCT

[1536] Anonymous. Device for cleaning traps in soil pipes. Scientific American, 24(14):211, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
211a.pdf.

Anonymous:1871:CPH

[1537] Anonymous. Combined pruning hook and saw. Scientific American, 24(14):211, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
211b.pdf.

Anonymous:1871:WBA

[1538] Anonymous. What becomes of all the steel pens? — their manufacture. Scientific American, 24(14):212–213, April 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n14/pdf/scientificamerican04011871-212a.pdf.

Anonymous:1871:WNY

[1539] Anonymous. The working of the New York fire department. Scientific American, 24(14):213–214, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
213.pdf.

Anonymous:1871:IDS

[1540] Anonymous. Improved door stop. Scientific American, 24(14):214, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-214.pdf.

Anonymous:1871:JIA

[1541] Anonymous. Johnson’s improved adjusting plumb and level. Scientific American, 24(14):214, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
214a.pdf.

Anonymous:1871:IMA

[1542] Anonymous. Improved method of attaching knobs to screws. Scientific American, 24(14):214–215, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
214b.pdf.

Anonymous:1871:PIM

[1543] Anonymous. Practical instruction in mechanics and physics. Scientific American, 24(14):215, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
215.pdf.

Anonymous:1871:ISBa

[1544] Anonymous. Inspection of steam boilers in Ohio. Scientific American, 24(14):215, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
215a.pdf.

Anonymous:1871:PDS

[1545] Anonymous. Progress of the Darien Ship Canal. Scientific American, 24(14):215, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
215b.pdf.

Anonymous:1871:SER

[1546] Anonymous. Scientific education and religion. Scientific American, 24(14):216, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
216.pdf.

Anonymous:1871:CNY

[1547] Anonymous. Caisson for the New York end of the East River Bridge. Scientific American, 24(14):216, April 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n14/pdf/scientificamerican04011871-216a.pdf.

Anonymous:1871:MLa

[1548] Anonymous. The moral lubricator. Scientific American, 24(14):216, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-216b.pdf.

Anonymous:1871:WNO

[1549] Anonymous. Whiskey, new and old. Scientific American, 24(14):216, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-216c.pdf.

Anonymous:1871:DD

[1550] Anonymous. Dangerous dentistry. Scientific American, 24(14):216, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-216d.pdf.

Anonymous:1871:ON

[1551] Anonymous. Obituary notices. Scientific American, 24(14):217, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-217.pdf.

Anonymous:1871:SIf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-217a.pdf.

Anonymous:1871:PWU

[1553] Anonymous. Preservation of wood under water. Scientific American, 24(14):217, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
217b.pdf.

Anonymous:1871:EPBa

[1554] Anonymous. Endless platform belt railway for New York. Scientific American, 24(14):217, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
217c.pdf.

Anonymous:1871:Qm

[1555] Anonymous. Queries. Scientific American, 24(14):217, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n14/pdf/scientificamerican04011871-217d.pdf.

Anonymous:1871:ACo

[1556] Anonymous. Answers to correspondents. Scientific American, 24 (14):218, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
218.pdf.

Anonymous:1871:AEPh

[1557] Anonymous. Applications for extension of patents. Scientific American, 24(14):218, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
218b.pdf.

Anonymous:1871:RAFn

[1558] Anonymous. Recent American and foreign patents. Scientific American, 24(14):219, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
219.pdf.

Anonymous:1871:OLPn

[1559] Anonymous. Official list of patents. Scientific American, 24 (14):219–220, April 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n14/pdf/scientificamerican04011871-
219a.pdf.

Anonymous:1871:Co

[1560] Anonymous. Correspondence. Scientific American, 24(16):244–245, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-244.pdf.

Anonymous:1871:BPo

[1561] Anonymous. Business and personal. Scientific American, 24(16):250, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-250.pdf.

Anonymous:1871:BNBi

[1562] Anonymous. Books: New books and publications. Scientific American, 24(16):252, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
252a.pdf.

Anonymous:1871:CEPa

[1563] Anonymous. The caisson for the east pier of the St. Louis Bridge. Scientific American, 24(16):239–240, April 15, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n16/pdf/scientificamerican04151871-239.pdf.

Anonymous:1871:PL

[1564] Anonymous. The patent laws. Scientific American, 24(16):241, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-241.pdf.

Anonymous:1871:IPb

[1565] Anonymous. The ivory plant. Scientific American, 24(16):242, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-242.pdf.

Anonymous:1871:TSBb

[1566] Anonymous. Triangular suspension bridge. Scientific American, 24(16):242–243, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
242a.pdf.

Anonymous:1871:IPH

[1567] Anonymous. Improved power hoisting machine. Scientific American, 24(16):243, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
243.pdf.

Anonymous:1871:BO

[1568] Anonymous. Boring for oil. Scientific American, 24(16):243–244, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-243a.pdf.

Anonymous:1871:IUS

[1569] Anonymous. Instruments used in storm signal service. Scientific American, 24(16):245–246, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
245.pdf.

Anonymous:1871:FHS

[1570] Anonymous. Furrowing hammer for stone dressing. Scientific American, 24(16):246, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
246.pdf.

Anonymous:1871:PNY

[1571] Anonymous. Pisciculture in New York and New Jersey. Scientific American, 24(16):247, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
247.pdf.

Anonymous:1871:IRC

[1572] Anonymous. Iron rails for common roadways. Scientific American, 24(16):247, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
247a.pdf.

Anonymous:1871:FSB

[1573] Anonymous. Firing of steam boilers. Scientific American, 24 (16):247–248, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
247b.pdf.

Anonymous:1871:AS

[1574] Anonymous. Artificial stone. Scientific American, 24(16):248, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-248.pdf.

Anonymous:1871:SIR

[1575] Anonymous. Suggestions for the improvement of railways. Scientific American, 24(16):248, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
248a.pdf.

Anonymous:1871:BTb

[1576] Anonymous. The battle of the tires. Scientific American, 24(16):248, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-248b.pdf.

Anonymous:1871:EFN

[1577] Anonymous. Employment of females in New York. Scientific American, 24(16):248–249, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
248c.pdf.

Anonymous:1871:SIg

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-249.pdf.

Anonymous:1871:Qn

[1579] Anonymous. Queries. Scientific American, 24(16):249–250, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n16/pdf/scientificamerican04151871-249a.pdf.

Anonymous:1871:ACq

[1580] Anonymous. Answers to correspondents. Scientific American, 24 (16):250, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
250a.pdf.

Anonymous:1871:RAFo

[1581] Anonymous. Recent American and foreign patents. Scientific American, 24(16):250–251, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
250b.pdf.

Anonymous:1871:OLPo

[1582] Anonymous. Official list of patents. Scientific American, 24 (16):251–252, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
251.pdf.

Anonymous:1871:AEPi

[1583] Anonymous. Applications for extension of patents. Scientific American, 24(16):252, April 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n16/pdf/scientificamerican04151871-
252.pdf.

Anonymous:1871:Cp

[1584] Anonymous. Correspondence. Scientific American, 24(18):276–278, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-276.pdf.

Anonymous:1871:BPp

[1585] Anonymous. Business and personal. Scientific American, 24(18):282, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-282.pdf.

Anonymous:1871:BNBj

[1586] Anonymous. Books: New books and publications. Scientific American, 24(18):282, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
282b.pdf.

Anonymous:1871:IEc

[1587] Anonymous. Improved excavator. Scientific American, 24(18):271, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-271.pdf.

Anonymous:1871:CSa

[1588] Anonymous. Colors of the sea. Scientific American, 24(18):271–272, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-271a.pdf.

Sellers:1871:TPC

[1589] Coleman Sellers. Tilghman’s process of cutting hard substances. Scientific American, 24(18):272, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
272.pdf.

Bailey:1871:SAS

[1590] W. W. Bailey. Something about seeds. Scientific American, 24(18):273, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-273.pdf.

Day:1871:CIT

[1591] E. C. H. Day. Cochineal insects and their allies. Scientific American, 24(18):274, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
274.pdf.

Anonymous:1871:HMM

[1592] Anonymous. Holding and mixing machines, used by the Union Stone Company — the Sorel process of manufacturing artificial stone. Scientific American, 24(18):275, April 29, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n18/pdf/scientificamerican04291871-275.pdf.

Anonymous:1871:MLb

[1593] Anonymous. Mortar light. Scientific American, 24(18):275, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-275a.pdf.

Anonymous:1871:CMU

[1594] Anonymous. Cope & Maxwell’s upright boiler feeder. Scientific American, 24(18):278–279, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
278.pdf.

Anonymous:1871:PCB

[1595] Anonymous. Proposed changes in the British patent laws. Scientific American, 24(18):279, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
279.pdf.

Anonymous:1871:CNa

[1596] Anonymous. Canal navigation. Scientific American, 24(18):279, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-279a.pdf.

Anonymous:1871:RRc

[1597] Anonymous. Roofs and roofing. Scientific American, 24(18):279–280, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-279b.pdf.

Anonymous:1871:MSU

[1598] Anonymous. Mining schools in the United States. Scientific American, 24(18):280, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
280.pdf.

Anonymous:1871:UIS

[1599] Anonymous. Uses of infusorial silica. Scientific American, 24(18):280, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-280a.pdf.

Anonymous:1871:CAa

[1600] Anonymous. Carbolic acid. Scientific American, 24(18):280, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-280b.pdf.

Anonymous:1871:SIh

[1601] Anonymous. Scientific intelligence. Scientific American, 24(18): 280–281, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
280c.pdf.

Anonymous:1871:ESa

[1602] Anonymous. Editorial summary. Scientific American, 24(18):281, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-281.pdf.

Anonymous:1871:ACr

[1603] Anonymous. Answers to correspondents. Scientific American, 24 (18):282, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
282a.pdf.

Anonymous:1871:Qo

[1604] Anonymous. Queries. Scientific American, 24(18):282, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n18/pdf/scientificamerican04291871-282c.pdf.

Anonymous:1871:RAFp

[1605] Anonymous. Recent American and foreign patents. Scientific American, 24(18):282–283, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
282d.pdf.

Anonymous:1871:AEPj

[1606] Anonymous. Applications for extension of patents. Scientific American, 24(18):283, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
283.pdf.

Anonymous:1871:OLPp

[1607] Anonymous. Official lists of patents. Scientific American, 24 (18):283–284, April 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n18/pdf/scientificamerican04291871-
283a.pdf.

Anonymous:1871:Cq

[1608] Anonymous. Correspondence. Scientific American, 24(19):291–294, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-291c.pdf.

Anonymous:1871:BPq

[1609] Anonymous. Business and personal. Scientific American, 24(19):298, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-298.pdf.

Anonymous:1871:BNBk

[1610] Anonymous. Books: New books and publications. Scientific American, 24(19):298, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
298b.pdf.

Anonymous:1871:IWS

[1611] Anonymous. Improved wheat steamer and dryer. Scientific American, 24(19):287, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
287.pdf.

Anonymous:1871:BBE

[1612] Anonymous. Bartlett and Boomer’s extension hemmer. Scientific American, 24(19):287–288, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
287a.pdf.

Anonymous:1871:OHTa

[1613] Anonymous. One hundred thousand dollars reward for a new plan for canal propulsion. Scientific American, 24(19):288, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-288.pdf.

Anonymous:1871:UF

[1614] Anonymous. Uncommon food. Scientific American, 24(19):289–290, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-289.pdf.

Anonymous:1871:CW

[1615] Anonymous. The cotton worm. Scientific American, 24(19):290, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-290.pdf.

Anonymous:1871:CMP

[1616] Anonymous. Combined milk pail and strainer. Scientific American, 24(19):291, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
291.pdf.

Anonymous:1871:CBR

[1617] Anonymous. Combined blotter, ruler, and paper cutter. Scientific American, 24(19):291, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
291a.pdf.

Lockwood:1871:MC

[1618] S. Lockwood. Mussel climbing. Scientific American, 24(19):291, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-291b.pdf.

Anonymous:1871:SBS

[1619] Anonymous. Seifert’s balanced steam valve. Scientific American, 24(19):294–295, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
294.pdf.

Anonymous:1871:CEa

[1620] Anonymous. Compound engines. Scientific American, 24(19):295, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-295.pdf.

Anonymous:1871:ISBb

[1621] Anonymous. Inspection of steam boilers. Scientific American, 24(19):295, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
295a.pdf.

Anonymous:1871:OHTb

[1622] Anonymous. One hundred thousand dollars reward for a new invention for propelling canal boats. Scientific American, 24(19):295, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-295b.pdf.

Anonymous:1871:RHG

[1623] Anonymous. The removal of the Hell Gate obstructions. Scientific American, 24(19):295–296, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
295c.pdf.

Anonymous:1871:MHT

[1624] Anonymous. More humorous than scientific. Scientific American, 24(19):296, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
296.pdf.

Anonymous:1871:NEI

[1625] Anonymous. Are noisome effluvia injurious to health? Scientific American, 24(19):296, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
296a.pdf.

Anonymous:1871:SIi

[1626] Anonymous. Scientific intelligence. Scientific American, 24(19): 296–298, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
296b.pdf.

Anonymous:1871:ACs

[1627] Anonymous. Answers to correspondence. Scientific American, 24 (19):298, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
298a.pdf.

Anonymous:1871:RAFq

[1628] Anonymous. Recent American and foreign patents. Scientific American, 24(19):299, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
299.pdf.

Anonymous:1871:Qp

[1629] Anonymous. Queries. Scientific American, 24(19):299, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n19/pdf/scientificamerican05061871-299a.pdf.

Anonymous:1871:OLPq

[1630] Anonymous. Official list of patents. Scientific American, 24 (19):299–300, May 6, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n19/pdf/scientificamerican05061871-
299b.pdf.

Anonymous:1871:Cr

[1631] Anonymous. Correspondence. Scientific American, 24(20):308–310, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-308.pdf.

Anonymous:1871:BPr

[1632] Anonymous. Business and personal. Scientific American, 24(20):314, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-314.pdf.

Anonymous:1871:BNBl

[1633] Anonymous. Books: New books and publications. Scientific American, 24(20):314, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
314b.pdf.

Anonymous:1871:HPM

[1634] Anonymous. Hannahs’ patent metropolitan railway, precautions in using wire rope. Scientific American, 24(20):303–304, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-303.pdf.

Anonymous:1871:SSP

[1635] Anonymous. Solutions for silver plating. Scientific American, 24 (20):305, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
305.pdf.

Anonymous:1871:CIA

[1636] Anonymous. The cochineal insect and its allies. Scientific American, 24(20):306, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
306.pdf.

Anonymous:1871:SH

[1637] Anonymous. The sea-horse. Scientific American, 24(20):306, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-306a.pdf.

Anonymous:1871:MKI

[1638] Anonymous. Mitchell & Kessinger’s improved flour bolt. Scientific American, 24(20):307, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
307.pdf.

Anonymous:1871:GKP

[1639] Anonymous. G. K. Proctor’s improved treadle motion. Scientific American, 24(20):307, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
307a.pdf.

Anonymous:1871:ITW

[1640] Anonymous. Improved turbine water wheel. Scientific American, 24(20):310, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
310.pdf.

Anonymous:1871:ECb

[1641] Anonymous. The Erie Canal. Scientific American, 24(20):311, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-311.pdf.

Anonymous:1871:PDW

[1642] Anonymous. Pipes for domestic water service. Scientific American, 24(20):311, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
311a.pdf.

Anonymous:1871:BAD

[1643] Anonymous. The blowpipe as an aid to the drill in opening safes. Scientific American, 24(20):311–312, May 13, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n20/pdf/scientificamerican05131871-311b.pdf.

Anonymous:1871:IDCa

[1644] Anonymous. Important decision by the Commissioner of Patents. Scientific American, 24(20):312, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
312.pdf.

Anonymous:1871:HOA

[1645] Anonymous. Health in old age. Scientific American, 24(20):312, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-312a.pdf.

Anonymous:1871:IMWa

[1646] Anonymous. Inventions made by workmen — who owns them? Scientific American, 24(20):312, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
312b.pdf.

Anonymous:1871:SIj

[1647] Anonymous. Scientific intelligence. Scientific American, 24(20):312, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-312c.pdf.

Anonymous:1871:MDP

[1648] Anonymous. On a method of determining the percentage of water mechanically suspended in steam delivered from boilers which prime. Scientific American, 24(20):312–313, May 13, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n20/pdf/scientificamerican05131871-312d.pdf.

Anonymous:1871:RW

[1649] Anonymous. Refuse and waste. Scientific American, 24(20):313, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-313.pdf.

Anonymous:1871:IDCb

[1650] Anonymous. Important decision of the Commissioner of Patents — trade mark patents. Scientific American, 24(20):313–314, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-313a.pdf.

Anonymous:1871:ACt

[1651] Anonymous. Answers to correspondents. Scientific American, 24 (20):314, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
314a.pdf.

Anonymous:1871:RAFr

[1652] Anonymous. Recent American and foreign patents. Scientific American, 24(20):314–315, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
314c.pdf.

Anonymous:1871:AEPk

[1653] Anonymous. Applications for extension of patents. Scientific American, 24(20):315, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
315.pdf.

Anonymous:1871:Qq

[1654] Anonymous. Queries. Scientific American, 24(20):315, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n20/pdf/scientificamerican05131871-315a.pdf.

Anonymous:1871:OLPr

[1655] Anonymous. Official list of patents. Scientific American, 24 (20):315–316, May 13, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n20/pdf/scientificamerican05131871-
315b.pdf.

Anonymous:1871:Cv

[1656] Anonymous. Correspondence. Scientific American, 24(21):324, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-324.pdf.

Anonymous:1871:BPs

[1657] Anonymous. Business and personal. Scientific American, 24(21):330, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-330.pdf.

Anonymous:1871:ROH

[1658] Anonymous. Removal of the obstructions at Hell Gate. Scientific American, 24(21):319–320, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
319.pdf.

Anonymous:1871:NYGa

[1659] Anonymous. A New Yorker in the Great Salt Mine of Cracow. Scientific American, 24(21):320, May 20, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n21/pdf/scientificamerican05201871-320.pdf.

Anonymous:1871:RSD

[1660] Anonymous. Rumford’s scientific discoveries. Scientific American, 24(21):321–322, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
321.pdf.

Anonymous:1871:HBH

[1661] Anonymous. Horse block and hitching post. Scientific American, 24(21):322–323, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
322.pdf.

Anonymous:1871:EPL

[1662] Anonymous. Eye pointed lasting awl. Scientific American, 24(21):323, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-323.pdf.

Anonymous:1871:EV

[1663] Anonymous. Eastern Virginia. Scientific American, 24(21):324–326, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-324a.pdf.

Anonymous:1871:RIPa

[1664] Anonymous. Robbins’ improved pipe wrench. Scientific American, 24(21):326, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
326.pdf.

Anonymous:1871:ECSa

[1665] Anonymous. The Earth closet system — an improved Earth closet. Scientific American, 24(21):326, May 20, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n21/pdf/scientificamerican05201871-326a.pdf.

Anonymous:1871:FRE

[1666] Anonymous. Facts relative to the Erie Canal. Scientific American, 24(21):327, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
327.pdf.

Anonymous:1871:TSS

[1667] Anonymous. Tornadoes of the Southern States. Scientific American, 24(21):327, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
327a.pdf.

Anonymous:1871:QSS

[1668] Anonymous. A question in social science. Scientific American, 24(21):327–328, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
327b.pdf.

Anonymous:1871:DO

[1669] Anonymous. From Duluth to the ocean. Scientific American, 24 (21):328, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
328.pdf.

Anonymous:1871:WSL

[1670] Anonymous. The water supply of London — sea baths in cities. Scientific American, 24(21):328, May 20, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n21/pdf/scientificamerican05201871-328a.pdf.

Anonymous:1871:DCL

[1671] Anonymous. The drug clerk law. Scientific American, 24(21):328, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-328b.pdf.

Anonymous:1871:SIk

[1672] Anonymous. Scientific intelligence. Scientific American, 24(21): 328–329, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
328c.pdf.

Anonymous:1871:ETS

[1673] Anonymous. Elasticity and tensile strength of wrought iron. Scientific American, 24(21):329, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
329.pdf.

Anonymous:1871:LCN

[1674] Anonymous. Launching of the caisson for the New York tower of the East River Bridge. Scientific American, 24(21):329, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-329a.pdf.

Anonymous:1871:ESb

[1675] Anonymous. Editorial summary. Scientific American, 24(21):329–330, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-329b.pdf.

Anonymous:1871:ACu

[1676] Anonymous. Answers to correspondents. Scientific American, 24 (21):330, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
330a.pdf.

Anonymous:1871:RAFs

[1677] Anonymous. Recent American and foreign patents. Scientific American, 24(21):330–331, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
330b.pdf.

Anonymous:1871:Qr

[1678] Anonymous. Queries. Scientific American, 24(21):331, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n21/pdf/scientificamerican05201871-331.pdf.

Anonymous:1871:OLPs

[1679] Anonymous. Official list of patents. Scientific American, 24 (21):331–332, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
331a.pdf.

Anonymous:1871:AEPl

[1680] Anonymous. Applications for extension of patents. Scientific American, 24(21):332, May 20, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n21/pdf/scientificamerican05201871-
332.pdf.

Anonymous:1871:Cx

[1681] Anonymous. Correspondence. Scientific American, 24(22):340, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-340.pdf.

Anonymous:1871:BPt

[1682] Anonymous. Business and personal. Scientific American, 24(22):346, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-346.pdf.

Anonymous:1871:BNBm

[1683] Anonymous. Books: New books and publications. Scientific American, 24(22):346, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
346b.pdf.

Anonymous:1871:OM

[1684] Anonymous. The old molder. Scientific American, 24(22):335, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-335.pdf.

Anonymous:1871:SF

[1685] Anonymous. Seal fishing. Scientific American, 24(22):336, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-336.pdf.

Anonymous:1871:MS

[1686] Anonymous. The metric system. Scientific American, 24(22):336, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-336a.pdf.

Anonymous:1871:VPPa

[1687] Anonymous. Vincent’s process of preparing linseed oil. Scientific American, 24(22):337, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
337.pdf.

Anonymous:1871:USSb

[1688] Anonymous. The United States Storm Signal Service weather charts and reports. Scientific American, 24(22):338–339, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-338.pdf.

Anonymous:1871:GIM

[1689] Anonymous. Gecmen’s improved malt kiln. Scientific American, 24(22):339, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
339.pdf.

Anonymous:1871:FS

[1690] Anonymous. Flying spiders. Scientific American, 24(22):341, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-341.pdf.

Anonymous:1871:RBT

[1691] Anonymous. Runk’s belt tightener. Scientific American, 24(22):342, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-342.pdf.

Anonymous:1871:SFR

[1692] Anonymous. Swartwout’s front running gear for wagons. Scientific American, 24(22):342, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
342a.pdf.

Anonymous:1871:SWC

[1693] Anonymous. Substitute for wine cellars. Scientific American, 24 (22):342, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
342b.pdf.

Anonymous:1871:PPBa

[1694] Anonymous. Patent packing box. Scientific American, 24(22):342–343, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-342c.pdf.

Anonymous:1871:IRE

[1695] Anonymous. Immigration — report of Edward Young, Chief of the Bureau of Statistics. Scientific American, 24(22):343, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-343.pdf.

Anonymous:1871:STE

[1696] Anonymous. Steam towage for the Erie Canal. Scientific American, 24(22):343, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
343a.pdf.

Anonymous:1871:NL

[1697] Anonymous. The necessities of life. Scientific American, 24(22):343, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-343b.pdf.

Anonymous:1871:GSa

[1698] Anonymous. Grape sugar. Scientific American, 24(22):343–344, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-343c.pdf.

Anonymous:1871:SLa

[1699] Anonymous. Sulphur in Louisiana. Scientific American, 24(22):344, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-344.pdf.

Anonymous:1871:CTP

[1700] Anonymous. Coal tar products. Scientific American, 24(22):344, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-344a.pdf.

Anonymous:1871:MSSb

[1701] Anonymous. Manufacture of silk — silk worm eggs. Scientific American, 24(22):344, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
344b.pdf.

Anonymous:1871:BOI

[1702] Anonymous. The bathometer, and other instruments for exploring, measuring, and utilizing the sea. Scientific American, 24(22):345–346, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-345.pdf.

Anonymous:1871:ACv

[1703] Anonymous. Answers to correspondents. Scientific American, 24 (22):346, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
346a.pdf.

Anonymous:1871:RAFt

[1704] Anonymous. Recent American and foreign patents. Scientific American, 24(22):347, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
347.pdf.

Anonymous:1871:Qs

[1705] Anonymous. Queries. Scientific American, 24(22):347, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n22/pdf/scientificamerican05271871-347a.pdf.

Anonymous:1871:AEPm

[1706] Anonymous. Applications for extension of patents. Scientific American, 24(22):347, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
347b.pdf.

Anonymous:1871:OLPt

[1707] Anonymous. Official list of patents. Scientific American, 24 (22):347–348, May 27, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n22/pdf/scientificamerican05271871-
347c.pdf.

Anonymous:1871:Cy

[1708] Anonymous. Correspondence. Scientific American, 24(23):356, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-356.pdf.

Anonymous:1871:BPu

[1709] Anonymous. Business and personal. Scientific American, 24(23):362, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-362.pdf.

Anonymous:1871:MLF

[1710] Anonymous. Machine for labeling fruit and other cans, the Moabite Stone, and more. Scientific American, 24(23):351, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n23/pdf/scientificamerican06031871-351.pdf.

Anonymous:1871:RAS

[1711] Anonymous. Ransome artificial stone. Scientific American, 24 (23):352–355, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
352.pdf.

Anonymous:1871:ADH

[1712] Anonymous. Apparatus for dyeing hair. Scientific American, 24 (23):355, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
355.pdf.

Anonymous:1871:NSb

[1713] Anonymous. A new spectroscope. Scientific American, 24(23):355, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-355a.pdf.

Anonymous:1871:CTN

[1714] Anonymous. Combined thimble and needle threader. Scientific American, 24(23):355–356, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
355b.pdf.

VanderWeyde:1871:FCM

[1715] P. R. Vander Weyde. A few corrections of the method of determining the percentage of water mechanically suspended in steam delivered from boilers which prime. Scientific American, 24(23):357–358, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-357.pdf.

Anonymous:1871:SLb

[1716] Anonymous. Sash lock. Scientific American, 24(23):358, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-358.pdf.

Anonymous:1871:CEb

[1717] Anonymous. Cork extractor. Scientific American, 24(23):358–359, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-358a.pdf.

Anonymous:1871:ECNa

[1718] Anonymous. The Erie Canal navigation problem. Scientific American, 24(23):359, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
359.pdf.

Anonymous:1871:HLE

[1719] Anonymous. How labor enslaves itself to capital. Scientific American, 24(23):359, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
359a.pdf.

Anonymous:1871:CEC

[1720] Anonymous. Charles E. Emery on compound engines. Scientific American, 24(23):359–360, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
359b.pdf.

Anonymous:1871:PSc

[1721] Anonymous. Popularizing science. Scientific American, 24(23):360, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-360.pdf.

Anonymous:1871:TOE

[1722] Anonymous. Trial of Olmstead’s electric car brake. Scientific American, 24(23):360, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
360a.pdf.

Anonymous:1871:OT

[1723] Anonymous. Ocean telegraphy. Scientific American, 24(23):360, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-360b.pdf.

Anonymous:1871:JLH

[1724] Anonymous. Jean Lafitte and his treasure. Scientific American, 24(23):360, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
360c.pdf.

Anonymous:1871:CAM

[1725] Anonymous. Canals ancient and modern. Scientific American, 24(23):360–361, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
360d.pdf.

Anonymous:1871:DM

[1726] Anonymous. The descent of man. Scientific American, 24(23):361, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-361.pdf.

Anonymous:1871:SIl

[1727] Anonymous. Scientific intelligence. Scientific American, 24(23): 361–362, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
361a.pdf.

Anonymous:1871:ACw

[1728] Anonymous. Answers to correspondents. Scientific American, 24 (23):362, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
362a.pdf.

Anonymous:1871:AEPn

[1729] Anonymous. Applications for extension of patents. Scientific American, 24(23):362, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
362b.pdf.

Anonymous:1871:RAFu

[1730] Anonymous. Recent American and foreign patents. Scientific American, 24(23):363, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
363.pdf.

Anonymous:1871:Qt

[1731] Anonymous. Queries. Scientific American, 24(23):363, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n23/pdf/scientificamerican06031871-363a.pdf.

Anonymous:1871:OLPu

[1732] Anonymous. Official list of patents. Scientific American, 24 (23):363–364, June 3, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n23/pdf/scientificamerican06031871-
363b.pdf.

Anonymous:1871:Cz

[1733] Anonymous. Correspondence. Scientific American, 24(24):371–372, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-371a.pdf.

Anonymous:1871:BPv

[1734] Anonymous. Business and personal. Scientific American, 24(24):378, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-378.pdf.

Anonymous:1871:BFK

[1735] Anonymous. The Bickford family knitting machine. Scientific American, 24(24):367–368, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
367.pdf.

Harding:1871:ASCa

[1736] George Edward Harding. The application of steam to canals — no. I. Scientific American, 24(24):369, June 10, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n24/pdf/scientificamerican06101871-369.pdf.

Anonymous:1871:Cba

[1737] Anonymous. Caravans. Scientific American, 24(24):369–371, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-369a.pdf.

Anonymous:1871:TSE

[1738] Anonymous. Toy steam engine. Scientific American, 24(24):371, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-371.pdf.

Tydall:1871:S

[1739] Tydall. On sound. Scientific American, 24(24):373, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-373.pdf.

Anonymous:1871:PMCa

[1740] Anonymous. Papier maché and carton pierre. Scientific American, 24(24):373, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
373a.pdf.

Anonymous:1871:TC

[1741] Anonymous. Thread cutter. Scientific American, 24(24):374, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-374.pdf.

Anonymous:1871:ACHa

[1742] Anonymous. Apparatus for carbureting hydrogen gas. Scientific American, 24(24):374, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
374a.pdf.

Anonymous:1871:ISC

[1743] Anonymous. Improvement in stock cars. Scientific American, 24 (24):374, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
374b.pdf.

Anonymous:1871:RID

[1744] Anonymous. Reat’s improved design for neck-yoke ring and slide. Scientific American, 24(24):374, June 10, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n24/pdf/scientificamerican06101871-374c.pdf.

Anonymous:1871:HPT

[1745] Anonymous. Hire’s patent table hinge. Scientific American, 24(24):374, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-374d.pdf.

Anonymous:1871:PCC

[1746] Anonymous. Plowing and cultivation of crops by steam improvements wanted. Scientific American, 24(24):374–375, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n24/pdf/scientificamerican06101871-374e.pdf.

Anonymous:1871:PEMb

[1747] Anonymous. Paine’s electro-motor. Scientific American, 24(24):375, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-375.pdf.

Anonymous:1871:NSP

[1748] Anonymous. The new system of piers for New York. Scientific American, 24(24):375, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
375a.pdf.

Anonymous:1871:DI

[1749] Anonymous. Doctoring iron. Scientific American, 24(24):375, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-375b.pdf.

Anonymous:1871:MEb

[1750] Anonymous. Mental emaciation. Scientific American, 24(24):375–376, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-375c.pdf.

Anonymous:1871:SECa

[1751] Anonymous. Steam on the Erie Canal — answers to queries. Scientific American, 24(24):376, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
376.pdf.

Anonymous:1871:PCSb

[1752] Anonymous. Plowing and cultivating by steam. Scientific American, 24(24):376–377, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
376a.pdf.

Houston:1871:NCI

[1753] Edwin J. Houston. On a new connection for the induction coil. Scientific American, 24(24):377, June 10, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n24/pdf/scientificamerican06101871-377.pdf.

Anonymous:1871:WWMa

[1754] Anonymous. Woven wire mattresses. Scientific American, 24(24):377, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-377a.pdf.

Anonymous:1871:PODa

[1755] Anonymous. Patent Office decision. Scientific American, 24(24): 377–378, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
377b.pdf.

Anonymous:1871:Qu

[1756] Anonymous. Queries. Scientific American, 24(24):378, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n24/pdf/scientificamerican06101871-378a.pdf.

Anonymous:1871:ACx

[1757] Anonymous. Answers to correspondents. Scientific American, 24 (24):378–379, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
378b.pdf.

Anonymous:1871:RAFv

[1758] Anonymous. Recent American and foreign patents. Scientific American, 24(24):379, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
379.pdf.

Anonymous:1871:OLPv

[1759] Anonymous. Official list of patents. Scientific American, 24 (24):379–380, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
379a.pdf.

Anonymous:1871:AEPo

[1760] Anonymous. Applications for extension of patents. Scientific American, 24(24):380, June 10, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n24/pdf/scientificamerican06101871-
380.pdf.

Anonymous:1871:Cbb

[1761] Anonymous. Correspondence. Scientific American, 24(25):388–390, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-388.pdf.

Anonymous:1871:BPw

[1762] Anonymous. Business and personal. Scientific American, 24(25):394, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-394.pdf.

Anonymous:1871:BNBn

[1763] Anonymous. Books: New books and publications. Scientific American, 24(25):394, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
394b.pdf.

Anonymous:1871:MTE

[1764] Anonymous. Machine for turning the ends of axletrees. Scientific American, 24(25):383, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
383.pdf.

Anonymous:1871:SPC

[1765] Anonymous. Schopp’s push chair or ice velocipede. Scientific American, 24(25):383, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
383a.pdf.

Anonymous:1871:YIH

[1766] Anonymous. Young’s improved hot air register. Scientific American, 24(25):383–384, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
383b.pdf.

Harding:1871:ASCb

[1767] George Edward Harding. The application of steam to canals — no. 2. Scientific American, 24(25):384, June 17, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n25/pdf/scientificamerican06171871-384.pdf.

Anonymous:1871:MVE

[1768] Anonymous. Manufacture of varnish in England. Scientific American, 24(25):384, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
384a.pdf.

Anonymous:1871:MT

[1769] Anonymous. Mechanical tests. Scientific American, 24(25):385, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-385.pdf.

Anonymous:1871:TCM

[1770] Anonymous. Tools for cutting metals. Scientific American, 24(25):385, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-385a.pdf.

Anonymous:1871:ISLb

[1771] Anonymous. The Illinois and St. Louis Bridge. Scientific American, 24(25):386, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
386.pdf.

Day:1871:WBBa

[1772] E. C. H. Day. Wood-boring beetles. Scientific American, 24(25):387, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-387.pdf.

Anonymous:1871:IES

[1773] Anonymous. Improvement in extracting spirits of turpentine from pine wood. Scientific American, 24(25):390, June 17, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n25/pdf/scientificamerican06171871-390.pdf.

Anonymous:1871:PCE

[1774] Anonymous. The proposed Centennial Exhibition in Philadelphia, in 1876. Scientific American, 24(25):391, June 17, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n25/pdf/scientificamerican06171871-391.pdf.

Anonymous:1871:RRd

[1775] Anonymous. Repairing roads. Scientific American, 24(25):391, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-391a.pdf.

Anonymous:1871:GWR

[1776] Anonymous. Guess work and real work. Scientific American, 24 (25):391, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
391b.pdf.

Anonymous:1871:EL

[1777] Anonymous. The electric light. Scientific American, 24(25):391–392, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-391c.pdf.

Anonymous:1871:SIm

[1778] Anonymous. Scientific intelligence. Scientific American, 24(25):392, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-392.pdf.

Anonymous:1871:GSb

[1779] Anonymous. A great speech. Scientific American, 24(25):392–393, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-392a.pdf.

Anonymous:1871:KAI

[1780] Anonymous. Klinkerfues’ apparatus for igniting gas and other lights. Scientific American, 24(25):393–394, June 17, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n25/pdf/scientificamerican06171871-393.pdf.

Anonymous:1871:Qv

[1781] Anonymous. Queries. Scientific American, 24(25):394, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n25/pdf/scientificamerican06171871-394a.pdf.

Anonymous:1871:ACy

[1782] Anonymous. Answer to correspondence. Scientific American, 24 (25):394, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
394c.pdf.

Anonymous:1871:RAFw

[1783] Anonymous. Recent American and foreign patents. Scientific American, 24(25):395, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
395.pdf.

Anonymous:1871:AEPp

[1784] Anonymous. Applications for extension of patents. Scientific American, 24(25):395, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
395a.pdf.

Anonymous:1871:OLPw

[1785] Anonymous. Official list of patents. Scientific American, 24 (25):395–396, June 17, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n25/pdf/scientificamerican06171871-
395b.pdf.

Anonymous:1871:Cbc

[1786] Anonymous. Correspondence. Scientific American, 24(26):404, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-404.pdf.

Anonymous:1871:BPx

[1787] Anonymous. Business and personal. Scientific American, 24(26):408, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-408a.pdf.

Anonymous:1871:BNBo

[1788] Anonymous. Books: New books and publications. Scientific American, 24(26):409, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
409b.pdf.

Anonymous:1871:Ia

[1789] Anonymous. Index. Scientific American, 24(26):410–412, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-410.pdf.

Anonymous:1871:IRB

[1790] Anonymous. Improved revolving blade scroll sawing machine, improved carriage hub, and more. Scientific American, 24(26):399–400, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-399.pdf.

Harding:1871:ASCc

[1791] George Edward Harding. The application of steam to canals — no. 3. Scientific American, 24(26):400–401, June 24, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n26/pdf/scientificamerican06241871-400.pdf.

Anonymous:1871:BEC

[1792] Anonymous. Blowing out of the east caisson of the East River Bridge. Scientific American, 24(26):402, June 24, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n26/pdf/scientificamerican06241871-402.pdf.

Anonymous:1871:VPO

[1793] Anonymous. Vagaries of the Patent Office. Scientific American, 24(26):402, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
402a.pdf.

Anonymous:1871:FMa

[1794] Anonymous. Flying machine. Scientific American, 24(26):402–403, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-402b.pdf.

Anonymous:1871:ISMb

[1795] Anonymous. Ice shaving machine. Scientific American, 24(26):403, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-403.pdf.

Anonymous:1871:DEM

[1796] Anonymous. Dredging and excavating machine. Scientific American, 24(26):403, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
403a.pdf.

Anonymous:1871:NEA

[1797] Anonymous. New expeditions to the Arctic regions. Scientific American, 24(26):403–404, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
403b.pdf.

Anonymous:1871:APM

[1798] Anonymous. Address of Professor Morse, at the ceremony of unveiling the statue in Central Park New York. Scientific American, 24(26):405, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-405.pdf.

Anonymous:1871:IMR

[1799] Anonymous. Improvement of the Missouri River at St. Joseph, MO. Scientific American, 24(26):405, June 24, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n26/pdf/scientificamerican06241871-405a.pdf.

Anonymous:1871:SIn

[1800] Anonymous. Scientific intelligence. Scientific American, 24(26): 405–406, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
405b.pdf.

Anonymous:1871:RVG

[1801] Anonymous. Re-vaccination — glycerin lymph. Scientific American, 24(26):406, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
406.pdf.

Anonymous:1871:BMM

[1802] Anonymous. Boring and mortising machine. Scientific American, 24(26):406, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
406a.pdf.

Anonymous:1871:GNY

[1803] Anonymous. The government of New York City. Scientific American, 24(26):406–407, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
406b.pdf.

Anonymous:1871:CAV

[1804] Anonymous. Close of another volume. Scientific American, 24(26):407, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-407.pdf.

Anonymous:1871:LSM

[1805] Anonymous. The last six months of chemistry. Scientific American, 24(26):407, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
407a.pdf.

Anonymous:1871:PEMc

[1806] Anonymous. Paine’s electro-motor. Scientific American, 24(26):407, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-407b.pdf.

Anonymous:1871:IMS

[1807] Anonymous. The inauguration of the Morse statue. Scientific American, 24(26):407, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
407c.pdf.

Anonymous:1871:FMb

[1808] Anonymous. Flying machine. Scientific American, 24(26):407–408, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-407d.pdf.

Anonymous:1871:ERB

[1809] Anonymous. The East River Bridge — report of the Chief Engineer. Scientific American, 24(26):408, June 24, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v24/
n26/pdf/scientificamerican06241871-408.pdf.

Anonymous:1871:ACz

[1810] Anonymous. Answer to correspondents. Scientific American, 24 (26):408, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
408b.pdf.

Anonymous:1871:RAFx

[1811] Anonymous. Recent American and foreign patents. Scientific American, 24(26):409, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v24/n26/pdf/scientificamerican06241871-
409.pdf.

Anonymous:1871:Qw

[1812] Anonymous. Queries. Scientific American, 24(26):409, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-409a.pdf.

Anonymous:1871:OLPx

[1813] Anonymous. Official list of patents. Scientific American, 24(26):412, June 24, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v24/n26/pdf/scientificamerican06241871-412.pdf.

Anonymous:1871:Cbd

[1814] Anonymous. Correspondence. Scientific American, 25(1):4–5, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-4.pdf.

Anonymous:1871:BNBp

[1815] Anonymous. Books: New books and publications. Scientific American, 25(1):9, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
9a.pdf.

Anonymous:1871:BPy

[1816] Anonymous. Business and personal. Scientific American, 25(1):10, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-10a.pdf.

Anonymous:1871:SJH

[1817] Anonymous. Sir John Herschel. Scientific American, 25(1):1–2, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-1.pdf.

Anonymous:1871:IPL

[1818] Anonymous. Instrument for parting ladies’ hair. Scientific American, 25(1):2–3, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
2.pdf.

Anonymous:1871:SMS

[1819] Anonymous. A self-moving ship. Scientific American, 25(1):3–4, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-3.pdf.

Anonymous:1871:RIS

[1820] Anonymous. Rand’s improved saw arbor. Scientific American, 25(1):6, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-6.pdf.

Anonymous:1871:IRT

[1821] Anonymous. Improvement in riveting tools. Scientific American, 25(1): 6, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-6a.pdf.

Anonymous:1871:OIS

[1822] Anonymous. Odiorne’s improvement in seals for hydraulic mains of gas works. Scientific American, 25(1):6, July 1, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n1/pdf/scientificamerican07011871-6b.pdf.

Anonymous:1871:ISMc

[1823] Anonymous. Improved self-measuring faucet. Scientific American, 25(1):6, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
6c.pdf.

Anonymous:1871:PID

[1824] Anonymous. Philippi’s improved drill chuck. Scientific American, 25(1): 6, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-6d.pdf.

Anonymous:1871:JHE

[1825] Anonymous. Joseph Henry and the electric telegraph. Scientific American, 25(1):7, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
7.pdf.

Anonymous:1871:EGW

[1826] Anonymous. Extracting gold from washings and poor ores. Scientific American, 25(1):7, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
7a.pdf.

Anonymous:1871:ECE

[1827] Anonymous. Effect of corporation employment on work-men. Scientific American, 25(1):7, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
7b.pdf.

Anonymous:1871:LRa

[1828] Anonymous. Lightning rods. Scientific American, 25(1):7–8, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-7c.pdf.

Anonymous:1871:MST

[1829] Anonymous. Massachusetts shoemakers on the taxes. Scientific American, 25(1):8, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
8.pdf.

Anonymous:1871:CBBa

[1830] Anonymous. Colossal bronze bust of Washington Irving. Scientific American, 25(1):8, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
8a.pdf.

Anonymous:1871:SIo

[1831] Anonymous. Scientific intelligence. Scientific American, 25(1):8, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-8b.pdf.

Anonymous:1871:KPE

[1832] Anonymous. The Kelly Patent Extension Case. — address of Hon. S. S. Fisher. — competitive examinations. — female applicants for clerkships in the Patent Office. Scientific American, 25(1):8–9, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-8c.pdf.

Anonymous:1871:NGRb

[1833] Anonymous. Narrow gages for railways. Scientific American, 25(1):9, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-9.pdf.

Anonymous:1871:Qx

[1834] Anonymous. Queries. Scientific American, 25(1):9, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n1/pdf/scientificamerican07011871-9b.pdf.

Anonymous:1871:ACba

[1835] Anonymous. Answers to correspondence. Scientific American, 25(1):10, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-10.pdf.

Anonymous:1871:AEPq

[1836] Anonymous. Applications for extension of patents. Scientific American, 25(1):10, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
10b.pdf.

Anonymous:1871:RAFy

[1837] Anonymous. Recent American and foreign patents. Scientific American, 25(1):11, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n1/pdf/scientificamerican07011871-
11.pdf.

Anonymous:1871:OLPy

[1838] Anonymous. Official list of patents. Scientific American, 25(1):11–12, July 1, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n1/pdf/scientificamerican07011871-11a.pdf.

Anonymous:1871:Cbe

[1839] Anonymous. Correspondence. Scientific American, 25(2):20–22, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-20.pdf.

Anonymous:1871:BPz

[1840] Anonymous. Business and personal. Scientific American, 25(2):26, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-26.pdf.

Anonymous:1871:BNBq

[1841] Anonymous. Books: New books and publications. Scientific American, 25(2):26, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
26b.pdf.

Anonymous:1871:PMAa

[1842] Anonymous. Professor Morse affixing his signature to the greeting telegraphed by Miss S. E. Cornwell at the Morse celebration. Scientific American, 25(2):15, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
15.pdf.

Harding:1871:ASCd

[1843] George Edward Harding. The application of steam to canals. — no.4. Scientific American, 25(2):16, July 8, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n2/pdf/scientificamerican07081871-16.pdf.

Anonymous:1871:PTS

[1844] Anonymous. Professor Tyndall on “sound”. Scientific American, 25(2):16–17, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
16a.pdf.

Anonymous:1871:FP

[1845] Anonymous. Fowler’s propeller. Scientific American, 25(2):17–18, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-17.pdf.

Anonymous:1871:IHL

[1846] Anonymous. Improved hand and light power drill. Scientific American, 25(2):18–19, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
18.pdf.

Anonymous:1871:DCC

[1847] Anonymous. Dorn’s circular cutter for wood work. Scientific American, 25(2):19, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
19.pdf.

Anonymous:1871:EBOa

[1848] Anonymous. Experiences of a Bureau officer. Scientific American, 25(2):19–20, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
19a.pdf.

Anonymous:1871:CLW

[1849] Anonymous. Coates and Lascell’s water engine. Scientific American, 25(2):22, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
22.pdf.

Anonymous:1871:MPS

[1850] Anonymous. Mechanical power and specific work. Scientific American, 25(2):23, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
23.pdf.

Anonymous:1871:EFC

[1851] Anonymous. Extinguishing fires by carbonic acid. Scientific American, 25(2):23, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
23a.pdf.

Anonymous:1871:HBM

[1852] Anonymous. How bodies may be frozen by heat. Scientific American, 25(2):23, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
23b.pdf.

Anonymous:1871:COC

[1853] Anonymous. Co-operative colonies. Scientific American, 25(2):23–24, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-23c.pdf.

Anonymous:1871:SECb

[1854] Anonymous. Steam on the Erie Canal. Scientific American, 25(2):24, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-24.pdf.

Anonymous:1871:BMP

[1855] Anonymous. Buildings for manufacturing purposes. Scientific American, 25(2):24, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
24a.pdf.

Anonymous:1871:ECF

[1856] Anonymous. Ex-Commissioner Fisher’s experience in the Patent Office. Scientific American, 25(2):24, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
24b.pdf.

Anonymous:1871:BHP

[1857] Anonymous. Burden horseshoe patent extended — value of the invention. Scientific American, 25(2):24, July 8, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n2/pdf/scientificamerican07081871-24c.pdf.

Anonymous:1871:SFE

[1858] Anonymous. Steam fire engines. Scientific American, 25(2):24–25, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-24d.pdf.

Anonymous:1871:Qy

[1859] Anonymous. Queries. Scientific American, 25(2):25–26, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-25.pdf.

Anonymous:1871:ACbb

[1860] Anonymous. Answers to correspondents. Scientific American, 25(2):26, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-26a.pdf.

Anonymous:1871:RAFz

[1861] Anonymous. Recent American and foreign patents. Scientific American, 25(2):26–27, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
26c.pdf.

Anonymous:1871:AEPr

[1862] Anonymous. Applications for extension of patents. Scientific American, 25(2):27, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n2/pdf/scientificamerican07081871-
27.pdf.

Anonymous:1871:OLPz

[1863] Anonymous. Official list of patents. Scientific American, 25(2):27–28, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-27a.pdf.

Anonymous:1871:PHIa

[1864] Anonymous. Practical hints to inventors. Scientific American, 25(2):28, July 8, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n2/pdf/scientificamerican07081871-28.pdf.

Anonymous:1871:Cbf

[1865] Anonymous. Correspondence. Scientific American, 25(3):36–38, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-36.pdf.

Anonymous:1871:BPba

[1866] Anonymous. Business and personal. Scientific American, 25(3):42, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-42.pdf.

Anonymous:1871:IGH

[1867] Anonymous. Improved grain hulling machine, applications of vulcanized rubber. Scientific American, 25(3):31–32, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3/pdf/scientificamerican07151871-31.pdf.

Anonymous:1871:EBOb

[1868] Anonymous. Experiences of a Bureau officer. Scientific American, 25(3):32, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
32.pdf.

Harding:1871:ASCe

[1869] George Edward Harding. The application of steam to canals. — no. 5. Scientific American, 25(3):32–33, July 15, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3/pdf/scientificamerican07151871-32a.pdf.

Anonymous:1871:PSLa

[1870] Anonymous. Phrenology and spiritualism. — a lively discussion. Scientific American, 25(3):33–34, July 15, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n3/pdf/scientificamerican07151871-33.pdf.

Day:1871:WBBb

[1871] E. C. H. Day and Professor. Wood-boring beetles. — the “girdlers”. Scientific American, 25(3):35, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
35.pdf.

Anonymous:1871:MCHa

[1872] Anonymous. Magrath’s chain for hanging sashes. Scientific American, 25(3):38, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
38.pdf.

Anonymous:1871:ALb

[1873] Anonymous. Artificial leech. Scientific American, 25(3):38, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-38a.pdf.

Anonymous:1871:MLS

[1874] Anonymous. Metallic lined spool head. Scientific American, 25(3):38, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-38b.pdf.

Anonymous:1871:CCa

[1875] Anonymous. College commencements. Scientific American, 25(3):39, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-39.pdf.

Anonymous:1871:CDa

[1876] Anonymous. The causes of disease. Scientific American, 25(3):39, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-39a.pdf.

Anonymous:1871:PAM

[1877] Anonymous. Prejudices, ancient and modern. Scientific American, 25(3):39, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
39b.pdf.

Anonymous:1871:INH

[1878] Anonymous. Ice a necessity. — how it is obtained in quantity. Scientific American, 25(3):40, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
40.pdf.

Anonymous:1871:SIp

[1879] Anonymous. Scientific intelligence. Scientific American, 25(3):40–41, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-40a.pdf.

Anonymous:1871:ESc

[1880] Anonymous. Editorial summary. Scientific American, 25(3):41–42, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-41.pdf.

Anonymous:1871:ACbc

[1881] Anonymous. Answers to correspondents. Scientific American, 25(3):42, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-42a.pdf.

Anonymous:1871:Qz

[1882] Anonymous. Queries. Scientific American, 25(3):43, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-43.pdf.

Anonymous:1871:RAFba

[1883] Anonymous. Recent American and foreign patents. Scientific American, 25(3):43, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
43a.pdf.

Anonymous:1871:OLPba

[1884] Anonymous. Official list of patents. Scientific American, 25(3):43–44, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-43b.pdf.

Anonymous:1871:AEPs

[1885] Anonymous. Applications for extension of patents. Scientific American, 25(3):44, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n3/pdf/scientificamerican07151871-
44.pdf.

Anonymous:1871:PHIb

[1886] Anonymous. Practical hints to inventors. Scientific American, 25(3):44, July 15, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n3/pdf/scientificamerican07151871-44a.pdf.

Anonymous:1871:Cbg

[1887] Anonymous. Correspondence. Scientific American, 25(4):51–54, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-51b.pdf.

Anonymous:1871:BPbb

[1888] Anonymous. Business and personal. Scientific American, 25(4):59, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-59a.pdf.

Anonymous:1871:ESE

[1889] Anonymous. The Estes steam engine, the general oceanic circulation. Scientific American, 25(4):47–48, July 22, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n4/pdf/scientificamerican07221871-47.pdf.

Anonymous:1871:DMHa

[1890] Anonymous. Discussion on Mr. Harding’s paper, entitled “The Application of Steam to Canals,” concluded in our last issue. — no. 1. Scientific American, 25(4):48–49, July 22, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n4/pdf/scientificamerican07221871-48.pdf.

Anonymous:1871:EBOc

[1891] Anonymous. Experiences of a Bureau officer. Scientific American, 25(4):49–50, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
49.pdf.

Anonymous:1871:KVE

[1892] Anonymous. King’s “valveless” engine. Scientific American, 25 (4):50–51, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
50.pdf.

Anonymous:1871:FIM

[1893] Anonymous. Fowler’s improved method of adjusting eccentrics. Scientific American, 25(4):51, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
51.pdf.

Anonymous:1871:NRWa

[1894] Anonymous. Nelson’s reversible window sash. Scientific American, 25(4):51, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
51a.pdf.

Anonymous:1871:MOS

[1895] Anonymous. Mental order and system. Scientific American, 25(4):55, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-55.pdf.

Anonymous:1871:EFT

[1896] Anonymous. Extinguishing fires at their commencement. Scientific American, 25(4):55, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
55a.pdf.

Anonymous:1871:GPB

[1897] Anonymous. Graphite, plumbago, black lead. Scientific American, 25(4):55, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
55b.pdf.

Anonymous:1871:ITb

[1898] Anonymous. Infringement trial. Scientific American, 25(4):55–56, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-55c.pdf.

Anonymous:1871:TEP

[1899] Anonymous. The threaded envelope patents. Scientific American, 25(4):56, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
56.pdf.

Anonymous:1871:WDP

[1900] Anonymous. What is doing at the Patent Office — great number of applicants for extensions. Scientific American, 25(4):56, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-56a.pdf.

Anonymous:1871:PN

[1901] Anonymous. Photographic notes. Scientific American, 25(4):56, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-56b.pdf.

Anonymous:1871:PTD

[1902] Anonymous. Professor Tyndall on “dust and smoke”. Scientific American, 25(4):57–58, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
57.pdf.

Anonymous:1871:NLP

[1903] Anonymous. No list of patentees. Scientific American, 25(4):58, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-58.pdf.

Anonymous:1871:ESd

[1904] Anonymous. Editorial summary. Scientific American, 25(4):58–59, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-58a.pdf.

Anonymous:1871:ACbd

[1905] Anonymous. Answers to correspondents. Scientific American, 25(4):59, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-59.pdf.

Anonymous:1871:AEPt

[1906] Anonymous. Applications for extension of patents. Scientific American, 25(4):59, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
59b.pdf.

Anonymous:1871:Qba

[1907] Anonymous. Queries. Scientific American, 25(4):60, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n4/pdf/scientificamerican07221871-60.pdf.

Anonymous:1871:RAFbb

[1908] Anonymous. Recent American and foreign patents. Scientific American, 25(4):60, July 22, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n4/pdf/scientificamerican07221871-
60a.pdf.

Anonymous:1871:Cbh

[1909] Anonymous. Correspondence. Scientific American, 25(5):68–70, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-68.pdf.

Anonymous:1871:BPbc

[1910] Anonymous. Business and personal. Scientific American, 25(5):72, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-72a.pdf.

Anonymous:1871:FFB

[1911] Anonymous. The Frere Fountain, Bombay, India. Scientific American, 25(5):63, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
63.pdf.

Anonymous:1871:WSb

[1912] Anonymous. Writers on science. Scientific American, 25(5):63–64, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-63a.pdf.

Anonymous:1871:UPL

[1913] Anonymous. Utility of patent laws. Scientific American, 25(5):64–65, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-64.pdf.

Anonymous:1871:DMHb

[1914] Anonymous. Discussion on Mr. Harding’s paper, entitled “The Application of Steam to Canals” concluded in our issue of July 15

— no. 2. Scientific American, 25(5):65–66, July 29, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n5/pdf/scientificamerican07291871-65.pdf.

Anonymous:1871:NMM

[1915] Anonymous. New mechanical movement. Scientific American, 25(5):66, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-66.pdf.

Anonymous:1871:CWC

[1916] Anonymous. Crocker’s whiffletree coupling. Scientific American, 25(5):67, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
67.pdf.

Anonymous:1871:BHQ

[1917] Anonymous. Belt holes — quarter turned belts. Scientific American, 25(5):67, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
67a.pdf.

Anonymous:1871:RI

[1918] Anonymous. Reminiscences of inventors. Scientific American, 25 (5):67–68, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
67b.pdf.

Anonymous:1871:BDF

[1919] Anonymous. Baer’s door fastener. Scientific American, 25(5):70, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-70.pdf.

Anonymous:1871:SAc

[1920] Anonymous. The spectroscope and its applications. Scientific American, 25(5):71, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
71.pdf.

Anonymous:1871:ECNb

[1921] Anonymous. The Erie Canal navigation problem. Scientific American, 25(5):71, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
71a.pdf.

Anonymous:1871:OC

[1922] Anonymous. Our correspondence. Scientific American, 25(5):71, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-71b.pdf.

Anonymous:1871:SIq

[1923] Anonymous. Scientific intelligence. Scientific American, 25(5):72, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-72.pdf.

Anonymous:1871:ACbe

[1924] Anonymous. Answers to correspondents. Scientific American, 25 (5):72–73, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
72b.pdf.

Anonymous:1871:Qbb

[1925] Anonymous. Queries. Scientific American, 25(5):73–74, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-73.pdf.

Anonymous:1871:RAFbc 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
74.pdf.

Anonymous:1871:OLPbb

[1927] Anonymous. Official list of patents. Scientific American, 25(5):75–76, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n5/pdf/scientificamerican07291871-75.pdf.

Anonymous:1871:AEPu

[1928] Anonymous. Applications for extension of patents. Scientific American, 25(5):76, July 29, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n5/pdf/scientificamerican07291871-
76.pdf.

Anonymous:1871:Cbi

[1929] Anonymous. Correspondence. Scientific American, 25(6):83–86, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-83b.pdf.

Anonymous:1871:BPbd

[1930] Anonymous. Business and personal. Scientific American, 25(6):90, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-90.pdf.

Anonymous:1871:SALa

[1931] Anonymous. Street architecture in London. Scientific American, 25(6):79–80, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
79.pdf.

Anonymous:1871:TFS

[1932] Anonymous. Tyndall’s fragments of science for unscientific people. Scientific American, 25(6):80, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
80.pdf.

Anonymous:1871:EBOd

[1933] Anonymous. Experiences of a Bureau officer. Scientific American, 25(6):81, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
81.pdf.

Day:1871:N

[1934] E. C. H. Day. The neuropters. Scientific American, 25(6):82–83, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-82.pdf.

Anonymous:1871:CSBa

[1935] Anonymous. Combined scissors and button-hole cutter. Scientific American, 25(6):83, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
83.pdf.

Anonymous:1871:DSA

[1936] Anonymous. Device for suspending and attaching articles. Scientific American, 25(6):83, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
83a.pdf.

Anonymous:1871:SVL

[1937] Anonymous. Safety valve lever. Scientific American, 25(6):86, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-86.pdf.

Anonymous:1871:HUA

[1938] Anonymous. Hildreth’s universally adjustable vise. Scientific American, 25(6):86, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
86a.pdf.

Anonymous:1871:OSC

[1939] Anonymous. Ore and stone crushing machine. Scientific American, 25(6):86, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
86b.pdf.

Anonymous:1871:PLR

[1940] Anonymous. Patent law reform in England. Scientific American, 25(6):87, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
87.pdf.

Anonymous:1871:VNY

[1941] Anonymous. Ventilation of the New York state lunatic asylum. Scientific American, 25(6):87, August 5, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n6/pdf/scientificamerican08051871-87a.pdf.

Anonymous:1871:FIF

[1942] Anonymous. Folly and ignorance — followed by death and destruction. Scientific American, 25(6):87, August 5, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n6/pdf/scientificamerican08051871-87b.pdf.

Anonymous:1871:LRb

[1943] Anonymous. Lightning rods. Scientific American, 25(6):87, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-87c.pdf.

Anonymous:1871:EMb

[1944] Anonymous. Electromagnetic motors. Scientific American, 25(6):88, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-88.pdf.

Anonymous:1871:SIs

[1945] Anonymous. Scientific intelligence. Scientific American, 25(6):88, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-88a.pdf.

Anonymous:1871:PPL

[1946] Anonymous. Patent proceedings in London. Scientific American, 25(6):88, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
88b.pdf.

Anonymous:1871:ECSb

[1947] Anonymous. Extension cases — Smith’s plow patent — Slawson’s car fare box — textile manufactures in India. Scientific American, 25(6):88–89, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
88c.pdf.

Anonymous:1871:CCb

[1948] Anonymous. Cancer cure. Scientific American, 25(6):89, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-89.pdf.

Anonymous:1871:ACbf

[1949] Anonymous. Answers to correspondents. Scientific American, 25(6):89–90, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
89a.pdf.

Anonymous:1871:Qbc

[1950] Anonymous. Queries. Scientific American, 25(6):90, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-90a.pdf.

Anonymous:1871:RAFbd

[1951] Anonymous. Recent American and foreign patents. Scientific American, 25(6):90–91, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
90b.pdf.

Anonymous:1871:OLPbc

[1952] Anonymous. Official list of patents. Scientific American, 25(6):91–92, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n6/pdf/scientificamerican08051871-91.pdf.

Anonymous:1871:AEPv

[1953] Anonymous. Applications for extension of patents. Scientific American, 25(6):92, August 5, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n6/pdf/scientificamerican08051871-
92.pdf.

Anonymous:1871:Cbj

[1954] Anonymous. Correspondence. Scientific American, 25(7):100–102, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-100.pdf.

Anonymous:1871:BNBr

[1955] Anonymous. Books: New books and publications. Scientific American, 25(7):105–106, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
105d.pdf.

Anonymous:1871:BPbe

[1956] Anonymous. Business and personal. Scientific American, 25(7):106, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-106.pdf.

Anonymous:1871:NPO

[1957] Anonymous. The new Post Office, Melbourne, Australia. Scientific American, 25(7):95, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
95.pdf.

Anonymous:1871:EBOe

[1958] Anonymous. Experiences of a Bureau officer. Scientific American, 25(7):96, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
96.pdf.

Morris:1871:CAG

[1959] Charles Morris. A chapter on American geology. Scientific American, 25(7):97, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
97.pdf.

Anonymous:1871:WIB

[1960] Anonymous. Wrought iron bridges for the Nicolai Railway of Russia. Scientific American, 25(7):98, August 12, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n7/pdf/scientificamerican08121871-98.pdf.

Crookes:1871:EIN

[1961] William Crookes. Experimental investigation of a new force. Scientific American, 25(7):99–100, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
99.pdf.

Anonymous:1871:MMC

[1962] Anonymous. Mechanical movement for churns, washing machines, etc. Scientific American, 25(7):102, August 12, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n7/pdf/scientificamerican08121871-102.pdf.

Anonymous:1871:PGW

[1963] Anonymous. Paget’s gipsy winch. Scientific American, 25(7):102, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-102a.pdf.

Anonymous:1871:EIM

[1964] Anonymous. Engravings of inventions and machinery. Scientific American, 25(7):102, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
102b.pdf.

Anonymous:1871:VB

[1965] Anonymous. A valuable book. Scientific American, 25(7):102, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-102c.pdf.

Anonymous:1871:DCPa

[1966] Anonymous. Dr. Crookes on psychic force. Scientific American, 25(7):103, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
103.pdf.

Anonymous:1871:LGR

[1967] Anonymous. The late General Rodman. Scientific American, 25(7):103, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
103a.pdf.

Anonymous:1871:GSW

[1968] Anonymous. Gold and silver waste. Scientific American, 25 (7):103–104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
103b.pdf.

Anonymous:1871:WGE

[1969] Anonymous. Will gunpowder explode in a vacuum? Scientific American, 25(7):104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
104.pdf.

Anonymous:1871:NAO

[1970] Anonymous. New application of the oxyhydrogen light to the separation of metals, by Tessie Du Mo-Tay of Paris, France. Scientific American, 25(7):104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
104a.pdf.

Anonymous:1871:NGRc

[1971] Anonymous. Narrow gage railways. Scientific American, 25(7):104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-104b.pdf.

Anonymous:1871:WE

[1972] Anonymous. The westfield explosion. Scientific American, 25 (7):104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
104c.pdf.

Anonymous:1871:HSMb

[1973] Anonymous. How scientific men work. Scientific American, 25 (7):104, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
104d.pdf.

Anonymous:1871:ALU

[1974] Anonymous. Artificial leather — utilization of leather scraps and cuttings. Scientific American, 25(7):105, August 12, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n7/pdf/scientificamerican08121871-105.pdf.

Anonymous:1871:RAD

[1975] Anonymous. Re-agent to detect nitric acid. Scientific American, 25(7):105, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
105a.pdf.

Anonymous:1871:OE

[1976] Anonymous. Ozonized ether. Scientific American, 25(7):105, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-105b.pdf.

Anonymous:1871:ESf

[1977] Anonymous. Editorial summary. Scientific American, 25(7):105, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-105c.pdf.

Anonymous:1871:ACbg

[1978] Anonymous. Answers to correspondents. Scientific American, 25 (7):106–107, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
106a.pdf.

Anonymous:1871:Qbd

[1979] Anonymous. Queries. Scientific American, 25(7):107, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n7/pdf/scientificamerican08121871-107.pdf.

Anonymous:1871:RAFbe

[1980] Anonymous. Recent American and foreign patents. Scientific American, 25(7):107, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
107a.pdf.

Anonymous:1871:OLPbd

[1981] Anonymous. Official lists of patents. Scientific American, 25 (7):107–108, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
107b.pdf.

Anonymous:1871:AEPw

[1982] Anonymous. Applications for extension of patents. Scientific American, 25(7):108, August 12, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n7/pdf/scientificamerican08121871-
108.pdf.

Anonymous:1871:Cbk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-116.pdf.

Anonymous:1871:BPbf

[1984] Anonymous. Business and personal. Scientific American, 25(8):122, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-122.pdf.

Anonymous:1871:FCa

[1985] Anonymous. File cutting. Scientific American, 25(8):111–113, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-111.pdf.

Anonymous:1871:HBC

[1986] Anonymous. Hydraulic buffer for checking the recoil of heavy guns. Scientific American, 25(8):114, August 19, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n8/pdf/scientificamerican08191871-114.pdf.

Anonymous:1871:ELM

[1987] Anonymous. The East London Museum of Science and Art. Scientific American, 25(8):115, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
115.pdf.

Anonymous:1871:WSc

[1988] Anonymous. What is steel? Scientific American, 25(8):115–116, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-115a.pdf.

Anonymous:1871:WHW

[1989] Anonymous. W. H. Windsor’s printer’s chase. Scientific American, 25(8):118, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
118.pdf.

Anonymous:1871:SCP

[1990] Anonymous. Shaw’s compound propeller pump. Scientific American, 25(8):118, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
118a.pdf.

Anonymous:1871:EPN

[1991] Anonymous. Emile Prevost’s new battery. Scientific American, 25(8):118, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
118b.pdf.

Anonymous:1871:BEa

[1992] Anonymous. Boiler explosions. Scientific American, 25(8):119, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-119.pdf.

Anonymous:1871:DS

[1993] Anonymous. The disposal of sewage. Scientific American, 25 (8):119, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
119a.pdf.

Anonymous:1871:HBB

[1994] Anonymous. How to become bankrupt by manufacturing. Scientific American, 25(8):119, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
119b.pdf.

Anonymous:1871:CDV

[1995] Anonymous. Combustibility of diamonds — various effects produced by high temperatures. Scientific American, 25(8):119–120, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-119c.pdf.

Anonymous:1871:AFA

[1996] Anonymous. The approaching fair of the American institute. Scientific American, 25(8):120, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
120.pdf.

Anonymous:1871:PAB

[1997] Anonymous. The patent agency business. Scientific American, 25(8):120, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
120a.pdf.

Anonymous:1871:EOB

[1998] Anonymous. Electrical organ blowers and sewing machine motors. Scientific American, 25(8):120, August 19, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n8/pdf/scientificamerican08191871-120b.pdf.

Anonymous:1871:NBPc

[1999] Anonymous. A new bleaching process. Scientific American, 25 (8):120, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
120c.pdf.

Anonymous:1871:IECa

[2000] Anonymous. Interesting extension cases — new rules and regulations. Scientific American, 25(8):121, August 19, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n8/pdf/scientificamerican08191871-121.pdf.

Anonymous:1871:NIS

[2001] Anonymous. A new idea in safes. Scientific American, 25(8):121, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-121a.pdf.

Anonymous:1871:ACbh

[2002] Anonymous. Answers to correspondents. Scientific American, 25 (8):121–122, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
121b.pdf.

Anonymous:1871:Qbe

[2003] Anonymous. Queries. Scientific American, 25(8):122–123, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n8/pdf/scientificamerican08191871-122a.pdf.

Anonymous:1871:RAFbf

[2004] Anonymous. Recent American and foreign patents. Scientific American, 25(8):123, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
123.pdf.

Anonymous:1871:OLPbe

[2005] Anonymous. Official list of patents. Scientific American, 25 (8):123–124, August 19, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n8/pdf/scientificamerican08191871-
123a.pdf.

Anonymous:1871:Cbl

[2006] Anonymous. Correspondence. Scientific American, 25(9):131–132, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-131a.pdf.

Anonymous:1871:BPbg

[2007] Anonymous. Business and personal. Scientific American, 25(9): 137–138, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
137b.pdf.

Anonymous:1871:BNBs 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
139a.pdf.

Anonymous:1871:MBL

[2009] Anonymous. Magazine breech loading gun, new light for the use of photographers, and more. Scientific American, 25(9):127–129, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-127.pdf.

Day:1871:CC

[2010] E. C. H. Day. Chrysalids and cocoons. Scientific American, 25(9):130, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
130.pdf.

Anonymous:1871:NSM

[2011] Anonymous. The new School of Mechanical Engineering at Hoboken, N.J. Scientific American, 25(9):131, August 26, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n9/pdf/scientificamerican08261871-131.pdf.

Anonymous:1871:WEH

[2012] Anonymous. The Westfield Explosion — how boilers ought to be inspected — interesting testimony by a wide awake man. Scientific American, 25(9):133, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
133.pdf.

Anonymous:1871:IMD

[2013] Anonymous. Improved method of drawing patterns for flaring vessels of sheet metal. Scientific American, 25(9):134, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-134.pdf.

Anonymous:1871:ATR

[2014] Anonymous. Adjustable tension razor strop. Scientific American, 25(9):134, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
134a.pdf.

Anonymous:1871:WEV

[2015] Anonymous. The Westfield Explosion: The verdict — warrants for the arrest of the President, Superintendent, and Engineer. who are the real culprits? Scientific American, 25(9):135, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-135.pdf.

Anonymous:1871:SIPa

[2016] Anonymous. State interference with patents. Scientific American, 25(9):135, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
135a.pdf.

Anonymous:1871:RSB

[2017] Anonymous. Rating steam boilers. Scientific American, 25(9):135, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-135b.pdf.

Anonymous:1871:OD

[2018] Anonymous. Objections to Darwinism. Scientific American, 25 (9):135–136, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
135c.pdf.

Anonymous:1871:NWM

[2019] Anonymous. A new water meter. Scientific American, 25(9):136, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-136.pdf.

Anonymous:1871:IMP

[2020] Anonymous. Improvements in marbleizing paper and oil cloth. Scientific American, 25(9):136, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
136a.pdf.

Anonymous:1871:SIt

[2021] Anonymous. Scientific intelligence. Scientific American, 25(9):136, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-136b.pdf.

Anonymous:1871:ISLc

[2022] Anonymous. Invalidity of state laws regulating the sale of patents. Scientific American, 25(9):137, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
137.pdf.

Anonymous:1871:PODb

[2023] Anonymous. Patent Office decisions. Scientific American, 25 (9):137, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
137a.pdf.

Anonymous:1871:ACbi

[2024] Anonymous. Answers to correspondents. Scientific American, 25(9):138, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
138.pdf.

Anonymous:1871:Qbf

[2025] Anonymous. Queries. Scientific American, 25(9):138, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n9/pdf/scientificamerican08261871-138a.pdf.

Anonymous:1871:RAFbg

[2026] Anonymous. Recent American and foreign patents. Scientific American, 25(9):139, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
139.pdf.

Anonymous:1871:OLPbf

[2027] Anonymous. Official list of patents. Scientific American, 25 (9):139–140, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
139b.pdf.

Anonymous:1871:AEPx

[2028] Anonymous. Applications for extension of patents. Scientific American, 25(9):140, August 26, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n9/pdf/scientificamerican08261871-
140.pdf.

Anonymous:1871:Cbm

[2029] Anonymous. Correspondence. Scientific American, 25(10):148, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n10/pdf/scientificamerican09021871-148.pdf.

Anonymous:1871:BNBt

[2030] Anonymous. Books: New books and publications. Scientific American, 25(10):153, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
153a.pdf.

Anonymous:1871:BPbh

[2031] Anonymous. Business and personal. Scientific American, 25(10): 153–154, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
153b.pdf.

Anonymous:1871:BBH

[2032] Anonymous. The burdict bolt heading machine. Scientific American, 25(10):143, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
143.pdf.

Anonymous:1871:DSI

[2033] Anonymous. The dangers of sewage irrigation. Scientific American, 25(10):143–144, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
143a.pdf.

Anonymous:1871:PPM

[2034] Anonymous. Plaster of Paris manufacture. Scientific American, 25(10):144, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
144.pdf.

Anonymous:1871:T

[2035] Anonymous. Teeth. Scientific American, 25(10):145, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n10/pdf/scientificamerican09021871-145.pdf.

Anonymous:1871:RSO

[2036] Anonymous. The road steamer and omnibus. Scientific American, 25(10):146, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
146.pdf.

Anonymous:1871:ABS

[2037] Anonymous. Anchor beam support for brick walls. Scientific American, 25(10):146, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
146a.pdf.

Anonymous:1871:NHC

[2038] Anonymous. The new Hotel Cairo Egypt. Scientific American, 25 (10):147–148, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
147.pdf.

H:1871:ONa

[2039] G. E. H. Occasional notes. Scientific American, 25(10):149– 150, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
149.pdf.

Anonymous:1871:DUC

[2040] Anonymous. Dugdale’s universal clothes washer. Scientific American, 25(10):150, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
150.pdf.

Anonymous:1871:SSc

[2041] Anonymous. Stanisłas Sorel. Scientific American, 25(10):151, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n10/pdf/scientificamerican09021871-151.pdf.

Anonymous:1871:NYPb

[2042] Anonymous. New York police inspection of boilers. Scientific American, 25(10):151, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
151a.pdf.

Anonymous:1871:SBL

[2043] Anonymous. Steam boiler legislation in England. Scientific American, 25(10):151, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
151b.pdf.

Anonymous:1871:WRF

[2044] Anonymous. Is the world round or flat? Scientific American, 25 (10):151–152, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
151c.pdf.

Anonymous:1871:GAP

[2045] Anonymous. The Grahamite asphalt pavement. Scientific American, 25(10):152, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
152.pdf.

Anonymous:1871:SIu

[2046] Anonymous. Scientific intelligence. Scientific American, 25(10): 152–153, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
152a.pdf.

Anonymous:1871:ESg

[2047] Anonymous. Editorial summary. Scientific American, 25(10): 153, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
153.pdf.

Anonymous:1871:ACbj

[2048] Anonymous. Answers to correspondents. Scientific American, 25 (10):154, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
154.pdf.

Anonymous:1871:Qbg

[2049] Anonymous. Queries. Scientific American, 25(10):154–155, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n10/pdf/scientificamerican09021871-154a.pdf.

Anonymous:1871:RAFbh

[2050] Anonymous. Recent American and foreign patents. Scientific American, 25(10):155, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
155.pdf.

Anonymous:1871:AEPy

[2051] Anonymous. Applications for extension of patents. Scientific American, 25(10):155, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
155a.pdf.

Anonymous:1871:OLPbg

[2052] Anonymous. Official list of patents. Scientific American, 25(10): 155–156, September 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n10/pdf/scientificamerican09021871-
155b.pdf.

Anonymous:1871:Cbn

[2053] Anonymous. Correspondence. Scientific American, 25(11):164, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-164.pdf.

Anonymous:1871:BPbi

[2054] Anonymous. Business and personal. Scientific American, 25(11): 169–170, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
169b.pdf.

Anonymous:1871:PGG

[2055] Anonymous. Portable gas generator, bleaching shellac, and more. Scientific American, 25(11):159–160, September 9, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n11/pdf/scientificamerican09091871-159.pdf.

Anonymous:1871:TWS

[2056] Anonymous. Tornadoes and water spouts. Scientific American, 25(11):161, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
161.pdf.

Anonymous:1871:LAF

[2057] Anonymous. Lemman’s anti-friction ball gate hinge. Scientific American, 25(11):162, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
162.pdf.

Anonymous:1871:NTK

[2058] Anonymous. New test for kerosene oils. Scientific American, 25 (11):162, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
162a.pdf.

Anonymous:1871:MDC

[2059] Anonymous. Massey’s double column high pressure and low water alarm for steam boilers. Scientific American, 25(11):162–163, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-162b.pdf.

Joy:1871:ILB

[2060] Charles A. Joy. An incident in the life of Baron Liebig. Scientific American, 25(11):163, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
163.pdf.

Anonymous:1871:PSLb

[2061] Anonymous. Plucky Susan, or luck at last — a leaf from the life of a Western inventor. Scientific American, 25(11):163, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-163a.pdf.

Anonymous:1871:TGC

[2062] Anonymous. Terrible gun cotton explosion at Stowmarket, England. Scientific American, 25(11):165, September 9, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n11/pdf/scientificamerican09091871-165.pdf.

Morris:1871:S

[2063] Charles Morris. The surf. Scientific American, 25(11):165–166, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-165a.pdf.

Anonymous:1871:RPI

[2064] Anonymous. Right of property in inventions an abstract right. Scientific American, 25(11):167, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
167.pdf.

Anonymous:1871:OL

[2065] Anonymous. The origin of life. Scientific American, 25(11):167, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-167a.pdf.

Anonymous:1871:JFI

[2066] Anonymous. Japan a field for inventions and machinery. Scientific American, 25(11):167, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
167b.pdf.

Anonymous:1871:SD

[2067] Anonymous. Spirits at a discount. Scientific American, 25(11): 168, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
168.pdf.

Anonymous:1871:WRT

[2068] Anonymous. Wagner’s report on technology. Scientific American, 25(11):168, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
168a.pdf.

Anonymous:1871:LPP

[2069] Anonymous. The late Professor Payen. Scientific American, 25 (11):168, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
168b.pdf.

Anonymous:1871:PFb

[2070] Anonymous. Psychic force. Scientific American, 25(11):168, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-168c.pdf.

Anonymous:1871:IECb

[2071] Anonymous. Interesting extension cases. Scientific American, 25 (11):169, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
169.pdf.

Anonymous:1871:AIF

[2072] Anonymous. American Institute Fair. Scientific American, 25 (11):169, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
169a.pdf.

Anonymous:1871:ACbk

[2073] Anonymous. Answers to correspondents. Scientific American, 25 (11):170, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
170.pdf.

Anonymous:1871:Qbh

[2074] Anonymous. Queries. Scientific American, 25(11):170, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n11/pdf/scientificamerican09091871-170a.pdf.

Anonymous:1871:RAFbi

[2075] Anonymous. Recent American and foreign patents. Scientific American, 25(11):170–171, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
170b.pdf.

Anonymous:1871:OLPbh

[2076] Anonymous. Official list of patents. Scientific American, 25 (11):171, September 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n11/pdf/scientificamerican09091871-
171.pdf.

Anonymous:1871:Cbo

[2077] Anonymous. Correspondence. Scientific American, 25(12):179– 182, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
179.pdf.

Anonymous:1871:BPbj

[2078] Anonymous. Business and personal. Scientific American, 25 (12):186, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
186.pdf.

Anonymous:1871:JTB

[2079] Anonymous. The Joslyn Tomes breech loading rifle, does the internal administration of drugs ever defeat disease and death? Scientific American, 25(12):175–176, September 16, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n12/pdf/scientificamerican09161871-175.pdf.

Anonymous:1871:JDJ

[2080] Anonymous. Jonathan Dennis, Jr’s appeal. Scientific American, 25 (12):176–177, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
176.pdf.

Anonymous:1871:FBS

[2081] Anonymous. The fog bow as seen from the Matterhorn. Scientific American, 25(12):178–179, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
178.pdf.

Anonymous:1871:NYI

[2082] Anonymous. The New York iron founders grievance — a system of shameless extortion. Scientific American, 25(12):183, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n12/pdf/scientificamerican09161871-183.pdf.

Anonymous:1871:NGN

[2083] Anonymous. Narrow gage nonsense. Scientific American, 25 (12):183, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
183a.pdf.

Anonymous:1871:BEb

[2084] Anonymous. Boiler explosions. Scientific American, 25(12):183– 184, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
183b.pdf.

Anonymous:1871:CHC

[2085] Anonymous. On certain habits considered as necessities of modern civilization. Scientific American, 25(12):184, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n12/pdf/scientificamerican09161871-184.pdf.

Anonymous:1871:PPi

[2086] Anonymous. Pharmacy in Prussia. Scientific American, 25(12): 184, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
184a.pdf.

Anonymous:1871:IIR

[2087] Anonymous. The introduction of India rubber shoes. Scientific American, 25(12):184, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
184b.pdf.

Anonymous:1871:LPC

[2088] Anonymous. The labor problem — class oration for 1871 at Yale. Scientific American, 25(12):184–185, September 16, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n12/pdf/scientificamerican09161871-184c.pdf.

H:1871:ONb

[2089] G. E. H. Occasional notes. Scientific American, 25(12):185, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n12/pdf/scientificamerican09161871-185.pdf.

Anonymous:1871:SIv

[2090] Anonymous. Scientific intelligence. Scientific American, 25(12): 185–186, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
185a.pdf.

Anonymous:1871:ACbl

[2091] Anonymous. Answers to correspondents. Scientific American, 25 (12):186, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
186a.pdf.

Anonymous:1871:Qbi

[2092] Anonymous. Queries. Scientific American, 25(12):187, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n12/pdf/scientificamerican09161871-187.pdf.

Anonymous:1871:RAFbj

[2093] Anonymous. Recent American and foreign patents. Scientific American, 25(12):187–188, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
187a.pdf.

Anonymous:1871:OLPbi

[2094] Anonymous. Official list of patents. Scientific American, 25 (12):188, September 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n12/pdf/scientificamerican09161871-
188.pdf.

Anonymous:1871:Cbp

[2095] Anonymous. Correspondence. Scientific American, 25(13):196– 197, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
196.pdf.

Anonymous:1871:BPbk

[2096] Anonymous. Business and personal. Scientific American, 25 (13):201, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
201.pdf.

Anonymous:1871:FEA

[2097] Anonymous. Fire extinguishing apparatus. Scientific American, 25 (13):191–192, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
191.pdf.

Anonymous:1871:LMS

[2098] Anonymous. The loss of my spectacles. Scientific American, 25 (13):192, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
192.pdf.

Anonymous:1871:IPDb

[2099] Anonymous. Important patent decision. Scientific American, 25 (13):193, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
193.pdf.

Anonymous:1871:APA

[2100] Anonymous. The aged pilgrims’ asylum. Scientific American, 25 (13):194–195, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
194.pdf.

Anonymous:1871:FILa

[2101] Anonymous. Farrell’s improved lamp Wick. Scientific American, 25(13):195, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
195.pdf.

Anonymous:1871:GCP

[2102] Anonymous. George’s clamp pen holder. Scientific American, 25 (13):195, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
195a.pdf.

Anonymous:1871:PFc

[2103] Anonymous. On the psychic force. Scientific American, 25(13): 197, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
197.pdf.

Anonymous:1871:FAIa

[2104] Anonymous. Fair of the American institute. Scientific American, 25(13):197, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
197a.pdf.

Anonymous:1871:ESS

[2105] Anonymous. Evens’ safety stove for railway cars. Scientific American, 25(13):198, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
198.pdf.

Anonymous:1871:SMPb

[2106] Anonymous. Sinclair’s mill pick holder. Scientific American, 25 (13):198, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
198a.pdf.

Anonymous:1871:BAW

[2107] Anonymous. Butler’s adjustable window sash. Scientific American, 25(13):198, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
198b.pdf.

Anonymous:1871:PGC

[2108] Anonymous. The perils of gun cotton. Scientific American, 25 (13):199, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
199.pdf.

Anonymous:1871:AMP

[2109] Anonymous. Ammonia motive power for organs. Scientific American, 25(13):199, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
199a.pdf.

Anonymous:1871:WCPa

[2110] Anonymous. What constitutes patentable novelty? Scientific American, 25(13):199, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
199b.pdf.

Anonymous:1871:DME

[2111] Anonymous. A defect in marine engineering. Scientific American, 25(13):199, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
199c.pdf.

Anonymous:1871:MHN

[2112] Anonymous. Mr. Highton and the New York Times. Scientific American, 25(13):199–200, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
199d.pdf.

Anonymous:1871:SLBa

[2113] Anonymous. The staff of life — the breadmakers of the American Metropolis. Scientific American, 25(13):200, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n13/pdf/scientificamerican09231871-200.pdf.

Anonymous:1871:SMT

[2114] Anonymous. The state of the manufacturing trades. Scientific American, 25(13):200, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
200a.pdf.

Anonymous:1871:ITDa

[2115] Anonymous. Important trademark decision. Scientific American, 25 (13):200–201, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
200b.pdf.

Anonymous:1871:RAFbk

[2116] Anonymous. Recent American and foreign patents. Scientific American, 25(13):201–203, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
201a.pdf.

Anonymous:1871:OLPbj

[2117] Anonymous. Official list of patents. Scientific American, 25(13): 203–204, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
203.pdf.

Anonymous:1871:AEPz

[2118] Anonymous. Applications for extension of patents. Scientific American, 25(13):204, September 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n13/pdf/scientificamerican09231871-
204.pdf.

Anonymous:1871:Cbq

[2119] Anonymous. Correspondence. Scientific American, 25(14):212– 213, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
212.pdf.

Anonymous:1871:BPbl

[2120] Anonymous. Business and personal. Scientific American, 25(14): 217–218, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
217.pdf.

Anonymous:1871:BNBu

[2121] Anonymous. Books: New books and publications. Scientific American, 25(14):220, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
220.pdf.

Anonymous:1871:TIE

[2122] Anonymous. Timpe’s improved elevators, mechanism for adjusting the rollers of carding machines, and more. Scientific American, 25 (14):207, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
207.pdf.

Anonymous:1871:ESCa

[2123] Anonymous. Extension of Sarvin’s carriage wheel patent. Scientific American, 25(14):208–209, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
208.pdf.

Anonymous:1871:NRWb

[2124] Anonymous. The new restaurant at Wimbledon Camp. Scientific American, 25(14):210–211, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
210.pdf.

Anonymous:1871:VIE

[2125] Anonymous. Vitt’s improved elevator cup. Scientific American, 25 (14):211, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
211.pdf.

Anonymous:1871:IMC

[2126] Anonymous. Inauguration of the mont cenis tunnel. Scientific American, 25(14):213, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
213.pdf.

Anonymous:1871:DPD

[2127] Anonymous. Death of Professor D. H. Mahan. Scientific American, 25(14):213–214, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
213a.pdf.

Anonymous:1871:PURa

[2128] Anonymous. Progress of underground roadways. Scientific American, 25(14):215, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
215.pdf.

Anonymous:1871:ATU

[2129] Anonymous. The action of trades unions in regard to apprenticeship. Scientific American, 25(14):215, September 30, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n14/pdf/scientificamerican09301871-215a.pdf.

Anonymous:1871:FCb

[2130] Anonymous. The fear of cholera. Scientific American, 25(14): 215, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
215b.pdf.

Anonymous:1871:FET

[2131] Anonymous. Fearful explosion of torpedoes — Coroner’s verdict. Scientific American, 25(14):215–216, September 30, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n14/pdf/scientificamerican09301871-215c.pdf.

Anonymous:1871:IECc

[2132] Anonymous. Interesting extension cases. Scientific American, 25 (14):216, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
216.pdf.

Anonymous:1871:FAIb

[2133] Anonymous. Fair of the American institute. Scientific American, 25 (14):216–217, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
216a.pdf.

Anonymous:1871:ACbm

[2134] Anonymous. Answers to correspondents. Scientific American, 25 (14):218, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
218.pdf.

Anonymous:1871:Qbj

[2135] Anonymous. Queries. Scientific American, 25(14):218, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n14/pdf/scientificamerican09301871-218a.pdf.

Anonymous:1871:AEPba

[2136] Anonymous. Applications for extension of patents. Scientific American, 25(14):218–219, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
218b.pdf.

Anonymous:1871:RAFbl

[2137] Anonymous. Recent American and foreign patents. Scientific American, 25(14):219, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
219.pdf.

Anonymous:1871:OLPbk

[2138] Anonymous. Official list of patents. Scientific American, 25(14): 219–220, September 30, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n14/pdf/scientificamerican09301871-
219a.pdf.

Anonymous:1871:Cbs

[2139] Anonymous. Correspondence. Scientific American, 25(15):227–229, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-227b.pdf.

Anonymous:1871:BPbm

[2140] Anonymous. Business and personal. Scientific American, 25(15):233, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-233a.pdf.

Anonymous:1871:ASDb

[2141] Anonymous. Anderson’s stone dresser. Scientific American, 25(15):223, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-223.pdf.

Anonymous:1871:SIw

[2142] Anonymous. Scientific intelligence. Scientific American, 25(15): 224–226, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
224.pdf.

Anonymous:1871:CDb

[2143] Anonymous. Churn dasher. Scientific American, 25(15):227, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-227.pdf.

Anonymous:1871:CICb

[2144] Anonymous. Copson’s improved corn broom. Scientific American, 25(15):227, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
227a.pdf.

Anonymous:1871:TMRa

[2145] Anonymous. Trade mark registrations. — interesting decision by the Hon. M. D. Leggett, Commissioner of Patents. Scientific American, 25(15):229, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
229.pdf.

Anonymous:1871:WCPb

[2146] Anonymous. What constitutes patentability in an article of manufacture. — important decision by the Commissioner of Patents. Scientific American, 25(15):229, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
229a.pdf.

Anonymous:1871:ECM

[2147] Anonymous. Extension of centering machine patent. Scientific American, 25(15):229–230, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
229b.pdf.

Anonymous:1871:PIL

[2148] Anonymous. Parish’s improved land mark. Scientific American, 25(15):230, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
230.pdf.

Anonymous:1871:LEI

[2149] Anonymous. Labor in england, and the international working Men’s association. Scientific American, 25(15):231, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-231.pdf.

Anonymous:1871:IDCc

[2150] Anonymous. Important decisions by the Commissioner of Patents. Scientific American, 25(15):231, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
231a.pdf.

Anonymous:1871:TMRb

[2151] Anonymous. Trade mark registrations. Scientific American, 25(15):231, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-231b.pdf.

Anonymous:1871:CMa

[2152] Anonymous. Chemical manures. Scientific American, 25(15):231–232, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-231c.pdf.

Anonymous:1871:ECc

[2153] Anonymous. The ether controversy. Scientific American, 25(15):232, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-232.pdf.

Anonymous:1871:FAIc

[2154] Anonymous. Fair of the American institute. Scientific American, 25(15):232–233, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
232a.pdf.

Anonymous:1871:ESh

[2155] Anonymous. Editorial summary. Scientific American, 25(15):233, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-233.pdf.

Anonymous:1871:ACbn

[2156] Anonymous. Answers to correspondents. Scientific American, 25 (15):233–234, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
233b.pdf.

Anonymous:1871:Qbk

[2157] Anonymous. Queries. Scientific American, 25(15):234, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-234.pdf.

Anonymous:1871:RAFbm

[2158] Anonymous. Recent American and foreign patents. Scientific American, 25(15):234–235, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n15/pdf/scientificamerican10071871-
234a.pdf.

Anonymous:1871:OLPbl

[2159] Anonymous. Official list of patents. Scientific American, 25(15):235, October 7, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n15/pdf/scientificamerican10071871-235.pdf.

Anonymous:1871:Cbt

[2160] Anonymous. Correspondence. Scientific American, 25(16):243–245, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n16/pdf/scientificamerican10141871-243a.pdf.

Anonymous:1871:BPbn

[2161] Anonymous. Business and personal. Scientific American, 25 (16):249, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
249b.pdf.

Anonymous:1871:BNBv

[2162] Anonymous. Books: New books and publications. Scientific American, 25(16):250, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
250a.pdf.

Anonymous:1871:BHO

[2163] Anonymous. Bessemer’s heavy ordnance. Scientific American, 25 (16):239–240, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
239.pdf.

Anonymous:1871:SRM

[2164] Anonymous. Screw and rivet manufacture. Scientific American, 25(16):240, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
240.pdf.

Anonymous:1871:NAW

[2165] Anonymous. New appliances for wool scouring and dyeing. Scientific American, 25(16):241–242, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
241.pdf.

Anonymous:1871:BPBo

[2166] Anonymous. Bain’s portable buildings. Scientific American, 25 (16):242–243, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
242.pdf.

Anonymous:1871:WPD

[2167] Anonymous. Warner and Payne’s door lock. Scientific American, 25(16):243, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
243.pdf.

Anonymous:1871:RIPb

[2168] Anonymous. Recent important patent decisions before the United States Circuit Court for the Eastern District of Pennsylvania. — Mc’Kennan, Judge Saponifier Patent. Scientific American, 25(16): 245–246, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
245.pdf.

Anonymous:1871:HMb

[2169] Anonymous. Homes for mechanics. Scientific American, 25(16):247, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n16/pdf/scientificamerican10141871-247.pdf.

Anonymous:1871:EPS

[2170] Anonymous. The English patent system. — what will they do with it. Scientific American, 25(16):247, October 14, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n16/pdf/scientificamerican10141871-247a.pdf.

Anonymous:1871:GWS

[2171] Anonymous. The gold weight of sunshine. Scientific American, 25(16):247, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
247b.pdf.

Anonymous:1871:HMHb

[2172] Anonymous. How many hours ought to constitute a Day’s work? Scientific American, 25(16):247–248, October 14, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n16/pdf/scientificamerican10141871-247c.pdf.

Anonymous:1871:VPPb

[2173] Anonymous. The value of property in patents illustrated. recent important patent decisions. Scientific American, 25(16):248, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n16/pdf/scientificamerican10141871-248.pdf.

Anonymous:1871:FAId

[2174] Anonymous. Fair of the American institute. Scientific American, 25(16):248–249, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
248a.pdf.

Anonymous:1871:ZWP

[2175] Anonymous. The zinc White patent decision. Scientific American, 25(16):249, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
249.pdf.

Anonymous:1871:Qbl

[2176] Anonymous. Queries. Scientific American, 25(16):249, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n16/pdf/scientificamerican10141871-249a.pdf.

Anonymous:1871:ACbo

[2177] Anonymous. Answer to correspondents. Scientific American, 25 (16):250, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
250.pdf.

Anonymous:1871:AEPbb

[2178] Anonymous. Applications for extension of patents. Scientific American, 25(16):250, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
250b.pdf.

Anonymous:1871:PHIc

[2179] Anonymous. Practical hints to inventors. Scientific American, 25(16):250, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
250c.pdf.

Anonymous:1871:RAFbn

[2180] Anonymous. Recent American and foreign patents. Scientific American, 25(16):251, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
251.pdf.

Anonymous:1871:OLPbm

[2181] Anonymous. Official list of patents. Scientific American, 25(16): 251–252, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n16/pdf/scientificamerican10141871-
251a.pdf.

Anonymous:1871:SAFa

[2182] Anonymous. The Second Annual Fair — of the — Alabama Agricultural — and — Mechanical Association. Scientific American, 25(16):252, October 14, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n16/pdf/scientificamerican10141871-252.pdf.

Anonymous:1871:Cbu

[2183] Anonymous. Correspondence. Scientific American, 25(17):260–261, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-260.pdf.

Anonymous:1871:BPRa

[2184] Anonymous. Business and personal: Recent American and foreign patents. Scientific American, 25(17):266, October 21, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n17/pdf/scientificamerican10211871-266a.pdf.

Anonymous:1871:BPbp

[2185] Anonymous. Business and personal. Scientific American, 25(17): 266–267, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
266b.pdf.

Anonymous:1871:IUW

[2186] Anonymous. Improved universal wood working machine. Scientific American, 25(17):255, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
255.pdf.

Anonymous:1871:CCc

[2187] Anonymous. Coloring cold. Scientific American, 25(17):255, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-255a.pdf.

Anonymous:1871:BCI

[2188] Anonymous. Breadstuff’s and cotton imported into great Britain. Scientific American, 25(17):255, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
255b.pdf.

Anonymous:1871:COa

[2189] Anonymous. Colorado ores. Scientific American, 25(17):256, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-256.pdf.

Anonymous:1871:ESi

[2190] Anonymous. Editorial summary. Scientific American, 25(17): 256–258, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
256a.pdf.

Anonymous:1871:BPSa

[2191] Anonymous. Baxter’s portable steam engine. Scientific American, 25(17):258, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
258.pdf.

Anonymous:1871:AEPbc

[2192] Anonymous. Apparatus for exhibiting the properties of vapors. Scientific American, 25(17):259, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
259.pdf.

Anonymous:1871:WWMb

[2193] Anonymous. The Waterbury water meter. Scientific American, 25(17):259, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
259a.pdf.

Anonymous:1871:MCP

[2194] Anonymous. M’Gee’s collar pin. Scientific American, 25(17):259, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-259b.pdf.

Anonymous:1871:CPS

[2195] Anonymous. The Cerulean Pleasonton’s sunshine patent. Scientific American, 25(17):261–262, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
261.pdf.

Anonymous:1871:CSS

[2196] Anonymous. Combined sadiron and stand, and combined reversible sadiron and polishing iron. Scientific American, 25(17):262, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-262.pdf.

Anonymous:1871:GFC

[2197] Anonymous. The great fire at Chicago. a national calamity and a terrible lesson. Scientific American, 25(17):263, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-263.pdf.

Anonymous:1871:EMc

[2198] Anonymous. An example to manufacturers. Scientific American, 25(17):263, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
263a.pdf.

Anonymous:1871:RPM

[2199] Anonymous. Recent progress in metallurgy. Scientific American, 25 (17):263–264, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
263b.pdf.

Anonymous:1871:MA

[2200] Anonymous. Master and apprentice. Scientific American, 25 (17):264, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
264.pdf.

Anonymous:1871:FAIe

[2201] Anonymous. Fair of the American institute. Scientific American, 25(17):264, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
264a.pdf.

Anonymous:1871:RHT

[2202] Anonymous. A remarkable history — a true story that is stranger than romance — how misfortune was crowned by success. Scientific American, 25(17):264–265, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
264b.pdf.

Anonymous:1871:MASa

[2203] Anonymous. Marquard’s artificial stone. Scientific American, 25 (17):265, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
265.pdf.

Anonymous:1871:SIx

[2204] Anonymous. Scientific intelligence. Scientific American, 25(17):265, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-265a.pdf.

Draper:1871:LU

[2205] John C. Draper. A light unit. Scientific American, 25(17):266, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-266.pdf.

Anonymous:1871:ACbp

[2206] Anonymous. Answers to correspondents. Scientific American, 25 (17):267, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
267.pdf.

Anonymous:1871:Qbm

[2207] Anonymous. Queries. Scientific American, 25(17):267, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n17/pdf/scientificamerican10211871-267a.pdf.

Anonymous:1871:OLPbn

[2208] Anonymous. Official list of patents. Scientific American, 25(17): 267–268, October 21, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n17/pdf/scientificamerican10211871-
267b.pdf.

Anonymous:1871:Cbv

[2209] Anonymous. Correspondence. Scientific American, 25(18):276, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-276.pdf.

Anonymous:1871:BPbq

[2210] Anonymous. Business and personal. Scientific American, 25 (18):281, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
281a.pdf.

Anonymous:1871:PMAb

[2211] Anonymous. Proposed monument in Athens to commemorate Greek independence. Scientific American, 25(18):271, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n18/pdf/scientificamerican10281871-271.pdf.

Anonymous:1871:LRW

[2212] Anonymous. The late Rev. W. V. Harcourt’s researches on glass. Scientific American, 25(18):272, October 28, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n18/pdf/scientificamerican10281871-272.pdf.

Fleming:1871:HS

[2213] G. Fleming. Horse shoeing. Scientific American, 25(18):272, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-272a.pdf.

Anonymous:1871:PMJa

[2214] Anonymous. Paper making in Japan. Scientific American, 25 (18):273–274, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
273.pdf.

Anonymous:1871:ASS

[2215] Anonymous. The American safety student lamp. Scientific American, 25(18):274, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
274.pdf.

Anonymous:1871:LBE

[2216] Anonymous. Lowe’s bucket elevator. Scientific American, 25 (18):275, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
275.pdf.

Draper:1871:FRM

[2217] John C. Draper. Filtration in refineries and manufactories. Scientific American, 25(18):275, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
275a.pdf.

Anonymous:1871:FILb

[2218] Anonymous. Fairlie’s improvements in locomotive engines. Scientific American, 25(18):277–278, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
277.pdf.

Anonymous:1871:HPF

[2219] Anonymous. Herr’s paper file. Scientific American, 25(18):278, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-278.pdf.

Anonymous:1871:FBa

[2220] Anonymous. Fireproof building. Scientific American, 25(18):279, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-279.pdf.

Anonymous:1871:GRP

[2221] Anonymous. A gigantic railway project. Scientific American, 25 (18):279, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
279a.pdf.

Anonymous:1871:GFN

[2222] Anonymous. The great fires in the Northwest. Scientific American, 25(18):279, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
279b.pdf.

Anonymous:1871:NJZ

[2223] Anonymous. The New Jersey Zinc Company. Scientific American, 25(18):279, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
279c.pdf.

Anonymous:1871:Cbw

[2224] Anonymous. Comets. Scientific American, 25(18):279–280, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-279d.pdf.

Anonymous:1871:EMd

[2225] Anonymous. Electricity in medicine. Scientific American, 25 (18):280, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
280.pdf.

Anonymous:1871:FAIf

[2226] Anonymous. Fair of the American institute. Scientific American, 25(18):280, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
280a.pdf.

Anonymous:1871:SIy

[2227] Anonymous. Scientific intelligence. Scientific American, 25(18): 280–281, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
280b.pdf.

Anonymous:1871:FSI

[2228] Anonymous. Fireproof safes — improvements urgently called for. Scientific American, 25(18):281, October 28, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n18/pdf/scientificamerican10281871-281.pdf.

Anonymous:1871:ACbq

[2229] Anonymous. Answers to correspondents. Scientific American, 25 (18):281, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
281b.pdf.

Anonymous:1871:Qbn

[2230] Anonymous. Queries. Scientific American, 25(18):282, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n18/pdf/scientificamerican10281871-282.pdf.

Anonymous:1871:RAFbo

[2231] Anonymous. Recent American and foreign patents. Scientific American, 25(18):282–283, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
282a.pdf.

Anonymous:1871:AEPbd

[2232] Anonymous. Applications for extension of patents. Scientific American, 25(18):283, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
283.pdf.

Anonymous:1871:OLPbo

[2233] Anonymous. Official list of patents. Scientific American, 25(18): 283–284, October 28, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n18/pdf/scientificamerican10281871-
283a.pdf.

Anonymous:1871:Cbx

[2234] Anonymous. Correspondence. Scientific American, 25(19):292– 293, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
292.pdf.

Anonymous:1871:BNBw

[2235] Anonymous. Books: New books and publications. Scientific American, 25(19):297, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
297a.pdf.

Anonymous:1871:BPbr

[2236] Anonymous. Business and personal. Scientific American, 25 (19):298, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
298.pdf.

Anonymous:1871:OBA

[2237] Anonymous. Organ blowing apparatus. Scientific American, 25 (19):287–288, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
287.pdf.

Anonymous:1871:MIWb

[2238] Anonymous. Model iron works. Scientific American, 25(19): 289–291, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
289.pdf.

Anonymous:1871:IVS

[2239] Anonymous. Ireland’s vaults for the safe keeping of valuables. Scientific American, 25(19):291, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
291.pdf.

Draper:1871:SVD

[2240] John O. Draper. Steam versus disease. Scientific American, 25 (19):291, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
291a.pdf.

Anonymous:1871:IECd

[2241] Anonymous. Improvement in Earth closets. Scientific American, 25(19):293, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
293.pdf.

Anonymous:1871:CSBb

[2242] Anonymous. The champion Spring bed. Scientific American, 25 (19):294, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
294.pdf.

Anonymous:1871:BIO

[2243] Anonymous. Bock’s improved one horse sleigh shaft. Scientific American, 25(19):294, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
294a.pdf.

Anonymous:1871:CMB

[2244] Anonymous. Coogan’s machine for boarding pebbling, and glossing leather. Scientific American, 25(19):294, November 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n19/pdf/scientificamerican11041871-294b.pdf.

Anonymous:1871:DFM

[2245] Anonymous. The dependence of future mechanical progress upon the discovery of new materials. Scientific American, 25(19):295, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n19/pdf/scientificamerican11041871-295.pdf.

Anonymous:1871:FBb

[2246] Anonymous. Fireproof building. Scientific American, 25(19): 295, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
295a.pdf.

Anonymous:1871:LTB

[2247] Anonymous. The leisure time of Boys. Scientific American, 25 (19):295, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
295b.pdf.

Anonymous:1871:VPI

[2248] Anonymous. Validity of patents issued under the new law. Scientific American, 25(19):295–296, November 4, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n19/pdf/scientificamerican11041871-295c.pdf.

Anonymous:1871:SAFb

[2249] Anonymous. Something about faces. Scientific American, 25 (19):296, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
296.pdf.

Anonymous:1871:DSR

[2250] Anonymous. Death of Sir Roderick Impey Murchison. Scientific American, 25(19):296, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
296a.pdf.

Anonymous:1871:FAIg

[2251] Anonymous. Fair of the American institute. — additional objects of interest. Scientific American, 25(19):296–297, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n19/pdf/scientificamerican11041871-296b.pdf.

Anonymous:1871:LNP

[2252] Anonymous. Large number of patents extended. Scientific American, 25(19):297, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
297.pdf.

Anonymous:1871:ACbr

[2253] Anonymous. Answers to correspondents. Scientific American, 25 (19):298, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
298a.pdf.

Anonymous:1871:Qbo

[2254] Anonymous. Queries. Scientific American, 25(19):298, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n19/pdf/scientificamerican11041871-298b.pdf.

Anonymous:1871:RAFbp

[2255] Anonymous. Recent American and foreign patents. Scientific American, 25(19):298–299, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
298c.pdf.

Anonymous:1871:OLPbp

[2256] Anonymous. Official list of patents. Scientific American, 25(19): 299–300, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
299.pdf.

Anonymous:1871:AEPbe

[2257] Anonymous. Applications for extension of patents. Scientific American, 25(19):300, November 4, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n19/pdf/scientificamerican11041871-
300.pdf.

Anonymous:1871:Cby

[2258] Anonymous. Correspondence. Scientific American, 25(20):308– 309, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
308.pdf.

Anonymous:1871:BNBx

[2259] Anonymous. Books: New books and publications. Scientific American, 25(20):313, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
313b.pdf.

Anonymous:1871:BPbs

[2260] Anonymous. Business and personal. Scientific American, 25 (20):314, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
314.pdf.

Anonymous:1871:EPP

[2261] Anonymous. Electro-pneumatic protector, an Italian diving Bell — the inventor writing letter at the bottom of the sea, and more. Scientific American, 25(20):303–304, November 11, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n20/pdf/scientificamerican11111871-303.pdf.

Anonymous:1871:FTC

[2262] Anonymous. Foster’s thill coupling. Scientific American, 25(20): 306, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
306.pdf.

Anonymous:1871:VGV

[2263] Anonymous. Ventilating grain vessel and grain car. Scientific American, 25(20):306, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
306a.pdf.

Anonymous:1871:IPO

[2264] Anonymous. Improved pocket oiler. Scientific American, 25(20): 306, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
306b.pdf.

Anonymous:1871:SIHa

[2265] Anonymous. Smith’s improved hollow Auger. Scientific American, 25(20):306, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
306c.pdf.

Anonymous:1871:PFF

[2266] Anonymous. Psychic force. — further experiments by Dr Crookes. — reply to his opponents. Scientific American, 25(20):307, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n20/pdf/scientificamerican11111871-307.pdf.

VanderWeyde:1871:MTC

[2267] P. H. Vander Weyde. On musical telegraph companies. Scientific American, 25(20):309–310, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
309.pdf.

Anonymous:1871:SBM

[2268] Anonymous. Suitable building material for cities. Scientific American, 25(20):311, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
311.pdf.

Anonymous:1871:DCPb

[2269] Anonymous. Dr. Crookes and psychic force. Scientific American, 25(20):311, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
311a.pdf.

Anonymous:1871:AF

[2270] Anonymous. American forests. Scientific American, 25(20):311– 312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
311b.pdf.

Anonymous:1871:GRS

[2271] Anonymous. General reflections suggested: by the fair of the American institute. Scientific American, 25(20):312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n20/pdf/scientificamerican11111871-312.pdf.

Anonymous:1871:USB

[2272] Anonymous. Use of sodium for blasting. Scientific American, 25 (20):312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
312a.pdf.

Anonymous:1871:PSCa

[2273] Anonymous. Propulsion of street cars. Scientific American, 25 (20):312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
312b.pdf.

Anonymous:1871:IPAa

[2274] Anonymous. Improvement in pavements — artificial stone flagging for sidewalks. Scientific American, 25(20):312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n20/pdf/scientificamerican11111871-312c.pdf.

Anonymous:1871:HBL

[2275] Anonymous. The health of Baron Liebig. Scientific American, 25 (20):312, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
312d.pdf.

Anonymous:1871:OCE

[2276] Anonymous. Our Chicago exchanges. Scientific American, 25 (20):313, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
313.pdf.

Anonymous:1871:RPDb

[2277] Anonymous. Recent patent decisions. Scientific American, 25 (20):313, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
313a.pdf.

Anonymous:1871:Qbp

[2278] Anonymous. Queries. Scientific American, 25(20):313, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n20/pdf/scientificamerican11111871-313c.pdf.

Anonymous:1871:ACbs

[2279] Anonymous. Answers to correspondents. Scientific American, 25 (20):314, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
314a.pdf.

Anonymous:1871:AEPbf

[2280] Anonymous. Applications for extension of patents. Scientific American, 25(20):314, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
314b.pdf.

Anonymous:1871:RAFbq

[2281] Anonymous. Recent American and foreign patents. Scientific American, 25(20):314–315, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
314c.pdf.

Anonymous:1871:OLPbq

[2282] Anonymous. Official list of patents. Scientific American, 25(20): 315–316, November 11, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n20/pdf/scientificamerican11111871-
315.pdf.

Anonymous:1871:Cbz

[2283] Anonymous. Correspondence. Scientific American, 25(21):324– 326, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
324.pdf.

Anonymous:1871:BPbt

[2284] Anonymous. Business and personal. Scientific American, 25 (21):330, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
330.pdf.

Anonymous:1871:ITDb

[2285] Anonymous. Improved tunneling drill. Scientific American, 25(21): 319–320, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
319.pdf.

Anonymous:1871:EAI

[2286] Anonymous. Electricity at the American Institute Exhibition. Scientific American, 25(21):320, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
320.pdf.

Anonymous:1871:TFA

[2287] Anonymous. Trial of fire arms. Scientific American, 25(21):320, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n21/pdf/scientificamerican11181871-320a.pdf.

Anonymous:1871:MCHb

[2288] Anonymous. Mammoth cave and how it was made. Scientific American, 25(21):321, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
321.pdf.

Anonymous:1871:HMc

[2289] Anonymous. The Humboldt Monument. Scientific American, 25 (21):322, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
322.pdf.

Anonymous:1871:WIA

[2290] Anonymous. Williams’ improved axle box for railway cars. Scientific American, 25(21):323, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
323.pdf.

Anonymous:1871:ESj

[2291] Anonymous. An explosion on the Sun. Scientific American, 25 (21):323, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
323a.pdf.

Anonymous:1871:BRH

[2292] Anonymous. Bebout’s rafter hook. Scientific American, 25(21): 326, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
326.pdf.

Anonymous:1871:EPBb

[2293] Anonymous. Evaporative power of boilers. Scientific American, 25(21):327, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
327.pdf.

Anonymous:1871:OPS

[2294] Anonymous. The open polar sea. Scientific American, 25(21): 327, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
327a.pdf.

Anonymous:1871:SPA

[2295] Anonymous. The sale of patented articles. Scientific American, 25(21):327, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
327b.pdf.

Anonymous:1871:HPH

[2296] Anonymous. How to prevent and how to extinguish fires. Scientific American, 25(21):327–328, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
327c.pdf.

Anonymous:1871:PEV

[2297] Anonymous. The proposed exhibition of vienna, in 1873. Scientific American, 25(21):328, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
328.pdf.

Anonymous:1871:FBP

[2298] Anonymous. Fireproof buildings for preserving public records. Scientific American, 25(21):328, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
328a.pdf.

Anonymous:1871:SIz

[2299] Anonymous. Scientific intelligence. Scientific American, 25(21): 328–329, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
328b.pdf.

Anonymous:1871:RPDc

[2300] Anonymous. Recent patent decisions. Scientific American, 25 (21):329, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
329.pdf.

Anonymous:1871:ACbt

[2301] Anonymous. Answers to correspondents. Scientific American, 25 (21):329–330, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
329a.pdf.

Anonymous:1871:Qbq

[2302] Anonymous. Queries. Scientific American, 25(21):330, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n21/pdf/scientificamerican11181871-330a.pdf.

Anonymous:1871:AEPbg

[2303] Anonymous. Applications for extension of patents. Scientific American, 25(21):330, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
330b.pdf.

Anonymous:1871:RAFbr

[2304] Anonymous. Recent American and foreign patents. Scientific American, 25(21):331, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
331.pdf.

Anonymous:1871:OLPbr

[2305] Anonymous. Official list of patents. Scientific American, 25(21): 331–332, November 18, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n21/pdf/scientificamerican11181871-
331a.pdf.

Anonymous:1871:Ccd

[2306] Anonymous. Correspondence. Scientific American, 25(22):340, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-340.pdf.

Anonymous:1871:BPbu

[2307] Anonymous. Business and personal. Scientific American, 25(22): 345–346, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
345a.pdf.

Anonymous:1871:BNBy 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
347.pdf.

Anonymous:1871:SMR

[2309] Anonymous. The screw mower and reaper, a Burmese hairy woman. Scientific American, 25(22):335, November 25, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n22/pdf/scientificamerican11251871-335.pdf.

Anonymous:1871:RPSb

[2310] Anonymous. Roofs, pavements, and safes, under fire at Chicago. Scientific American, 25(22):335–336, November 25, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n22/pdf/scientificamerican11251871-335a.pdf.

Anonymous:1871:CEPb

[2311] Anonymous. Down in the caisson of the East Pier of the St. Louis Bridge. Scientific American, 25(22):336–337, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-336.pdf.

Anonymous:1871:CWW

[2312] Anonymous. The Chicago Water Works as they appeared before the late fire. Scientific American, 25(22):339, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-339.pdf.

Anonymous:1871:NCPc

[2313] Anonymous. The new City Post Office. Scientific American, 25 (22):339–340, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
339a.pdf.

Draper:1871:AMB

[2314] John C. Draper. Absorption of moisture by brick and stone. Scientific American, 25(22):340–341, November 25, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n22/pdf/scientificamerican11251871-340a.pdf.

VanderWeyde:1871:SRP

[2315] P. H. Vander Weyde. Some remarks on Professor Crookes’ latest communications. Scientific American, 25(22):341–342, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-341.pdf.

Anonymous:1871:RBM

[2316] Anonymous. A remarkable boy mechanic. Scientific American, 25 (22):342, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
342.pdf.

Anonymous:1871:SIFa

[2317] Anonymous. Simonson’s improved filtering cistern. Scientific American, 25(22):342, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
342a.pdf.

Anonymous:1871:QDa

[2318] Anonymous. Quinnipiac dam. Scientific American, 25(22):342, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-342b.pdf.

Anonymous:1871:HI

[2319] Anonymous. Hints to inventors. Scientific American, 25(22): 343, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
343.pdf.

Anonymous:1871:MAL

[2320] Anonymous. More about the labor question. Scientific American, 25(22):343, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
343a.pdf.

Anonymous:1871:AGP

[2321] Anonymous. Another great public work. Scientific American, 25 (22):343–344, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
343b.pdf.

Anonymous:1871:HOE

[2322] Anonymous. The Hassler oceanic exploration. Scientific American, 25(22):344, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
344.pdf.

Anonymous:1871:UGA

[2323] Anonymous. Uses of glass in architectural decoration etc. Scientific American, 25(22):344, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
344a.pdf.

Anonymous:1871:WAR

[2324] Anonymous. A word about repairs. Scientific American, 25 (22):344, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
344b.pdf.

Anonymous:1871:BM

[2325] Anonymous. Books for mechanics. Scientific American, 25(22): 344, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
344c.pdf.

Anonymous:1871:SIba

[2326] Anonymous. Scientific intelligence. Scientific American, 25(22): 344–345, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
344d.pdf.

Anonymous:1871:HSF

[2327] Anonymous. Holly’s system of fire protection and water supply. Scientific American, 25(22):345, November 25, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n22/pdf/scientificamerican11251871-345.pdf.

Anonymous:1871:ACbu

[2328] Anonymous. Answers to correspondents. Scientific American, 25 (22):346, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
346.pdf.

Anonymous:1871:Qbr

[2329] Anonymous. Queries. Scientific American, 25(22):346, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n22/pdf/scientificamerican11251871-346a.pdf.

Anonymous:1871:RAFbs

[2330] Anonymous. Recent American and foreign patents. Scientific American, 25(22):346–347, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
346b.pdf.

Anonymous:1871:OLPbs

[2331] Anonymous. Official list of patents. Scientific American, 25(22): 347–348, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
347a.pdf.

Anonymous:1871:AEPbh

[2332] Anonymous. Applications for extension of patents. Scientific American, 25(22):348, November 25, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n22/pdf/scientificamerican11251871-
348.pdf.

Anonymous:1871:Cce

[2333] Anonymous. Correspondence. Scientific American, 25(23):356, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n23/pdf/scientificamerican12021871-356.pdf.

Anonymous:1871:BNBz

[2334] Anonymous. Books: New books and publications. Scientific American, 25(23):361, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
361b.pdf.

Anonymous:1871:CSP

[2335] Anonymous. Church of St. Peter’s, Bournemouth England. Scientific American, 25(23):351–352, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
351.pdf.

Anonymous:1871:CUS

[2336] Anonymous. Cornell University — a Student’s labor for knowledge. Scientific American, 25(23):352–353, December 2, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n23/pdf/scientificamerican12021871-352.pdf.

Anonymous:1871:NPS

[2337] Anonymous. New publication — the science record, for 1872. Scientific American, 25(23):353–354, December 2, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n23/pdf/scientificamerican12021871-353.pdf.

Anonymous:1871:BWA

[2338] Anonymous. Briesen’s writing apparatus for the blind etc. Scientific American, 25(23):354–355, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
354.pdf.

Anonymous:1871:IAR

[2339] Anonymous. Improved axle for railway cars and locomotives. Scientific American, 25(23):355, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
355.pdf.

Anonymous:1871:UHF

[2340] Anonymous. Units of heat and of force. Scientific American, 25(23):356, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
356a.pdf.

Draper:1871:DSB

[2341] John C. Draper. Decay of stone and brick. Scientific American, 25(23):357, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
357.pdf.

H:1871:ONc

[2342] G. E. H. Occasional notes. Scientific American, 25(23):357–358, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n23/pdf/scientificamerican12021871-357a.pdf.

Anonymous:1871:TMG

[2343] Anonymous. Too much gas; not enough light. Scientific American, 25(23):359, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
359.pdf.

Anonymous:1871:PURb

[2344] Anonymous. Progress of the underground railway system. Scientific American, 25(23):359, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
359a.pdf.

Anonymous:1871:EMDa

[2345] Anonymous. An exact method of determining the evaporative efficiency of steamboilers. Scientific American, 25(23):359, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n23/pdf/scientificamerican12021871-359b.pdf.

Anonymous:1871:CBbb

[2346] Anonymous. Concrete building. Scientific American, 25(23):359– 360, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
359c.pdf.

Anonymous:1871:DDE

[2347] Anonymous. Dr. Draper’s experiments with brick and stone. Scientific American, 25(23):360, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
360.pdf.

Anonymous:1871:SSSb

[2348] Anonymous. The study of science in schools. Scientific American, 25(23):360, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
360a.pdf.

Anonymous:1871:NET

[2349] Anonymous. A number of extensions — tedious sewing machine case. Scientific American, 25(23):360–361, December 2, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n23/pdf/scientificamerican12021871-360b.pdf.

Anonymous:1871:SIbb

[2350] Anonymous. Scientific intelligence. Scientific American, 25(23):361, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n23/pdf/scientificamerican12021871-361.pdf.

Anonymous:1871:PAH

[2351] Anonymous. Prince Alexis — his welcome to America. Scientific American, 25(23):361, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
361a.pdf.

Anonymous:1871:BPbv

[2352] Anonymous. Business and personal. Scientific American, 25 (23):361, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
361c.pdf.

Anonymous:1871:ACbv

[2353] Anonymous. Answers to correspondents. Scientific American, 25 (23):362, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
362.pdf.

Anonymous:1871:Qbs

[2354] Anonymous. Queries. Scientific American, 25(23):362, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n23/pdf/scientificamerican12021871-362a.pdf.

Anonymous:1871:RAFbt

[2355] Anonymous. Recent American and foreign patents. Scientific American, 25(23):362–363, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
362b.pdf.

Anonymous:1871:OLPbt

[2356] Anonymous. Official list of patents. Scientific American, 25(23): 363–364, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
363.pdf.

Anonymous:1871:AEPbi

[2357] Anonymous. Applications for extension of patents. Scientific American, 25(23):364, December 2, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n23/pdf/scientificamerican12021871-
364.pdf.

Anonymous:1871:Ccf

[2358] Anonymous. Correspondence. Scientific American, 25(24):372, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n24/pdf/scientificamerican12091871-372.pdf.

Anonymous:1871:BNBba

[2359] Anonymous. Books: New books and publications. Scientific American, 25(24):377, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
377a.pdf.

Anonymous:1871:BPbw

[2360] Anonymous. Business and personal. Scientific American, 25 (24):377, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
377b.pdf.

Anonymous:1871:LSS

[2361] Anonymous. The Landore Siemens Steel Works England. Scientific American, 25(24):367–368, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
367.pdf.

Anonymous:1871:DT

[2362] Anonymous. Durability of timber. Scientific American, 25(24): 368–370, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
368.pdf.

Anonymous:1871:CHA

[2363] Anonymous. Calles’ hydro-aero-dynamic wheel. Scientific American, 25(24):370–372, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
370.pdf.

VanderWeyde:1871:SH

[2364] P. H. Vander Weyde. On specific heat. Scientific American, 25 (24):372, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
372a.pdf.

Anonymous:1871:SMI

[2365] Anonymous. Syracuse — its mechanical industries. Scientific American, 25(24):372–373, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
372b.pdf.

VanCleve:1871:RET

[2366] A. H. Van Cleve. Report of experiments for testing the relative value of lubricating oils. Scientific American, 25(24):373–374, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n24/pdf/scientificamerican12091871-373.pdf.

Anonymous:1871:SPF

[2367] Anonymous. Smith’s portable furnace for shrinking on tires. Scientific American, 25(24):374, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
374.pdf.

Anonymous:1871:TSH

[2368] Anonymous. Tangye’s smoke and heat precipitator. Scientific American, 25(24):374, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
374a.pdf.

Anonymous:1871:IMWb

[2369] Anonymous. Improvement in the manufacture of White lead. Scientific American, 25(24):374, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
374b.pdf.

Anonymous:1871:SALb

[2370] Anonymous. Something about lubricators. Scientific American, 25(24):375, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
375.pdf.

Anonymous:1871:SSd

[2371] Anonymous. To smoke or not to smoke. Scientific American, 25(24):375, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
375a.pdf.

Anonymous:1871:CML

[2372] Anonymous. The condition of mechanics and laborers in England. Scientific American, 25(24):375–376, December 9, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n24/pdf/scientificamerican12091871-375b.pdf.

Anonymous:1871:PAI

[2373] Anonymous. Protection to American industry. Scientific American, 25(24):376, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
376.pdf.

Anonymous:1871:SIbc

[2374] Anonymous. Scientific intelligence. Scientific American, 25(24):376, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n24/pdf/scientificamerican12091871-376a.pdf.

Anonymous:1871:TSc

[2375] Anonymous. Timely suggestions. Scientific American, 25(24):377, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n24/pdf/scientificamerican12091871-377.pdf.

Anonymous:1871:RAFbu

[2376] Anonymous. Recent American and foreign patents. Scientific American, 25(24):378–379, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
378a.pdf.

Anonymous:1871:AEPbj

[2377] Anonymous. Applications for extension of patents. Scientific American, 25(24):379, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
379.pdf.

Anonymous:1871:OLPbu

[2378] Anonymous. Official list of patents. Scientific American, 25(24): 379–380, December 9, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n24/pdf/scientificamerican12091871-
379a.pdf.

Anonymous:1871:Ccg

[2379] Anonymous. Correspondence. Scientific American, 25(25):388– 389, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
388.pdf.

Anonymous:1871:BNBbb

[2380] Anonymous. Books: New books and publications. Scientific American, 25(25):394, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
394.pdf.

Anonymous:1871:BPbx

[2381] Anonymous. Business and personal. Scientific American, 25 (25):394, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
394b.pdf.

Anonymous:1871:GSAa

[2382] Anonymous. Green and Stancliff’s apparatus for hoisting and conveying coal. Scientific American, 25(25):383, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n25/pdf/scientificamerican12161871-383.pdf.

Anonymous:1871:NFFa

[2383] Anonymous. Notes on flying and flying machines. Scientific American, 25(25):384–386, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
384.pdf.

Anonymous:1871:TE

[2384] Anonymous. The Thames Embankment. Scientific American, 25 (25):387–388, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
387.pdf.

VanderWeyde:1871:RRB

[2385] P. H. Vander Weyde. Remarkable relation between the specific heat and the atomic weight. Scientific American, 25(25):389–390, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n25/pdf/scientificamerican12161871-389.pdf.

Anonymous:1871:WFH

[2386] Anonymous. Wilcox’s folding hencoop. Scientific American, 25 (25):390, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
390.pdf.

Anonymous:1871:SCT

[2387] Anonymous. Self-closing telegraph key. Scientific American, 25 (25):390, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
390a.pdf.

Anonymous:1871:NYGb

[2388] Anonymous. The New York gas monopolies. Scientific American, 25(25):391, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
391.pdf.

Anonymous:1871:RE

[2389] Anonymous. The repetition of experiments. Scientific American, 25(25):391, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
391a.pdf.

Anonymous:1871:CHP

[2390] Anonymous. The condemnation of the Hall of public records for the city and county of New York. Scientific American, 25(25):391, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n25/pdf/scientificamerican12161871-391b.pdf.

Anonymous:1871:MNM

[2391] Anonymous. A Mare’s nest — more about the gas question. Scientific American, 25(25):391–392, December 16, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n25/pdf/scientificamerican12161871-391c.pdf.

Anonymous:1871:EEb

[2392] Anonymous. Education of the eye. Scientific American, 25(25): 392, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
392.pdf.

Anonymous:1871:SGT

[2393] Anonymous. The St. Gothard Tunnel — another grand engineering work. Scientific American, 25(25):392, December 16, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n25/pdf/scientificamerican12161871-392a.pdf.

Anonymous:1871:MIM

[2394] Anonymous. A most ingenious machine. Scientific American, 25 (25):392, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
392b.pdf.

Anonymous:1871:MB

[2395] Anonymous. Mechanical bigots. Scientific American, 25(25): 392, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
392c.pdf.

Anonymous:1871:IIG

[2396] Anonymous. Important improvement in glass and plate engraving. Scientific American, 25(25):392, December 16, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n25/pdf/scientificamerican12161871-392d.pdf.

Anonymous:1871:SRa

[2397] Anonymous. Science record for 1872. Scientific American, 25 (25):393, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
393.pdf.

Anonymous:1871:SIbd

[2398] Anonymous. Scientific intelligence. Scientific American, 25(25): 393, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
393a.pdf.

Anonymous:1871:CSc

[2399] Anonymous. To city subscribers. Scientific American, 25(25): 393, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
393b.pdf.

Anonymous:1871:TSd

[2400] Anonymous. Timely suggestions. Scientific American, 25(25): 393–394, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
393c.pdf.

Anonymous:1871:AEPbk

[2401] Anonymous. Applications for extension of patents. Scientific American, 25(25):394, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
394a.pdf.

Anonymous:1871:ACbw

[2402] Anonymous. Answers to correspondents. Scientific American, 25 (25):394–395, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
394c.pdf.

Anonymous:1871:Qbt

[2403] Anonymous. Queries. Scientific American, 25(25):395, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n25/pdf/scientificamerican12161871-395.pdf.

Anonymous:1871:RAFbv

[2404] Anonymous. Recent American and foreign patents. Scientific American, 25(25):395, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
395a.pdf.

Anonymous:1871:OLPbv

[2405] Anonymous. Official list of patents. Scientific American, 25(25): 395–396, December 16, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n25/pdf/scientificamerican12161871-
395b.pdf.

Anonymous:1871:Cch

[2406] Anonymous. Correspondence. Scientific American, 25(26):404– 405, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
404.pdf.

Anonymous:1871:BPby

[2407] Anonymous. Business and personal. Scientific American, 25 (26):408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
408d.pdf.

Anonymous:1871:Ic

[2408] Anonymous. Index. Scientific American, 25(26):410–412, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n26/pdf/scientificamerican12231871-410.pdf.

Anonymous:1871:TBO

[2409] Anonymous. The tunneling and blasting operations at Hell Gate, East River New York. Scientific American, 25(26):399–400, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n26/pdf/scientificamerican12231871-399.pdf.

Anonymous:1871:NFFb

[2410] Anonymous. Notes on flying and flying machines. Scientific American, 25(26):400–401, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
400.pdf.

Anonymous:1871:MASb

[2411] Anonymous. Manufacture of artificial stone for the Carthagena breakwater. Scientific American, 25(26):402, December 23, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n26/pdf/scientificamerican12231871-402.pdf.

Anonymous:1871:CBbc

[2412] Anonymous. Charles Babbage. Scientific American, 25(26):402– 403, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
402a.pdf.

Anonymous:1871:MRT

[2413] Anonymous. Machine for rolling tapered bars. Scientific American, 25(26):403, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
403.pdf.

Anonymous:1871:MPW

[2414] Anonymous. Maine’s portable window ventilator. Scientific American, 25(26):403, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
403a.pdf.

Anonymous:1871:AMM

[2415] Anonymous. The application of mechanics and mechanical principles in agricultural operations. Scientific American, 25(26):403–404, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n26/pdf/scientificamerican12231871-403b.pdf.

Anonymous:1871:SE

[2416] Anonymous. Subterranean explorations. Scientific American, 25 (26):405, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
405.pdf.

VanderWeyde:1871:LHL

[2417] P. H. Vander Weyde. Latent heat of liquids. Scientific American, 25 (26):405–406, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
405a.pdf.

Anonymous:1871:WCD

[2418] Anonymous. Wendell’s combined door buffer and door fastener. Scientific American, 25(26):406, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
406.pdf.

Anonymous:1871:CTS

[2419] Anonymous. Close of the twenty-sixth year of the Scientific American. Scientific American, 25(26):407, December 23, 1871. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v25/
n26/pdf/scientificamerican12231871-407.pdf.

Anonymous:1871:GEV

[2420] Anonymous. The great exhibition at vienna, 1873. Scientific American, 25(26):407, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
407a.pdf.

Anonymous:1871:LNHb

[2421] Anonymous. Lyceum of Natural History — discussion on milk and disinfectants. Scientific American, 25(26):407, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v25/n26/pdf/scientificamerican12231871-407b.pdf.

Anonymous:1871:MFb

[2422] Anonymous. The mysteries of flight. Scientific American, 25(26): 407–408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
407c.pdf.

Anonymous:1871:SRb

[2423] Anonymous. Science record for 1872. Scientific American, 25 (26):408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
408.pdf.

Anonymous:1871:CSd

[2424] Anonymous. To city subscribers. Scientific American, 25(26): 408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
408a.pdf.

Anonymous:1871:TSf

[2425] Anonymous. Timely suggestions. Scientific American, 25(26): 408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
408b.pdf.

Anonymous:1871:AEPbl

[2426] Anonymous. Applications for extension of patents. Scientific American, 25(26):408, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
408c.pdf.

Anonymous:1871:ACbx

[2427] Anonymous. Answers to correspondents. Scientific American, 25 (26):409, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
409.pdf.

Anonymous:1871:OLPbw

[2428] Anonymous. Official list of patents. Scientific American, 25 (26):409, December 23, 1871. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v25/n26/pdf/scientificamerican12231871-
409a.pdf.

Anonymous:1872:Ca

[2429] Anonymous. Correspondence. Scientific American, 26(1):4–5, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-4.pdf.

Anonymous:1872:BPa

[2430] Anonymous. Business and personal. Scientific American, 26(1):9, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-9.pdf.

Anonymous:1872:NQa

[2431] Anonymous. Notes and queries. Scientific American, 26(1):9, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-9a.pdf.

Anonymous:1872:PNPa

[2432] Anonymous. Patent notes: Patents. Scientific American, 26(1):12–14, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-12b.pdf.

Anonymous:1872:PLA

[2433] Anonymous. Professor Louis Agassiz. Scientific American, 26(1):1–2, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-1.pdf.

Anonymous:1872:ESV

[2434] Anonymous. Eccentric Spring vise. Scientific American, 26(1):2, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-2.pdf.

Anonymous:1872:SCT

[2435] Anonymous. Sheppard’s cotton tie. Scientific American, 26(1):2, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-2a.pdf.

Anonymous:1872:SRa

[2436] Anonymous. Science record for 1872. Scientific American, 26(1):2, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-2b.pdf.

Anonymous:1872:GIE

[2437] Anonymous. Goux’s improvements in Earth closets. Scientific American, 26(1):3, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
3.pdf.

Anonymous:1872:HWH

[2438] Anonymous. Hot water heater for greenhouses. Scientific American, 26(1):3, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
3a.pdf.

Young:1872:EC

[2439] O. A. Young. Encke’s comet. Scientific American, 26(1):3–4, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-3b.pdf.

Weyde:1872:LHV

[2440] P. H. Vandee Weyde. Latent heat of vapors and gases. Scientific American, 26(1):5–6, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
5.pdf.

Anonymous:1872:BTC

[2441] Anonymous. Bridgman’s thill coupling. Scientific American, 26(1):6, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-6.pdf.

Anonymous:1872:ASF

[2442] Anonymous. Artificial stone and fireproof buildings. Scientific American, 26(1):7, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
7.pdf.

Anonymous:1872:EE

[2443] Anonymous. Education of the ear. Scientific American, 26(1):7, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-7a.pdf.

Anonymous:1872:MCF

[2444] Anonymous. Meteorology, its capabilities and its future. Scientific American, 26(1):7, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
7b.pdf.

Anonymous:1872:LSM

[2445] Anonymous. The last six months of chemistry. Scientific American, 26(1):8, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
8.pdf.

Anonymous:1872:SRb

[2446] Anonymous. Science record for 1872. Scientific American, 26(1):8, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-8a.pdf.

Anonymous:1872:SIa

[2447] Anonymous. Scientific intelligence. Scientific American, 26(1):8, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-8b.pdf.

Anonymous:1872:PDa

[2448] Anonymous. Patent decision. Scientific American, 26(1):8, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-8c.pdf.

Anonymous:1872:CSa

[2449] Anonymous. To city subscribers. Scientific American, 26(1):8, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-8d.pdf.

Anonymous:1872:TS

[2450] Anonymous. Timely suggestions. Scientific American, 26(1):8–9, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-8e.pdf.

Anonymous:1872:ACa

[2451] Anonymous. Answer to correspondents. Scientific American, 26 (1):9–10, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
9b.pdf.

Anonymous:1872:RAFa

[2452] Anonymous. Recent American and foreign patents. Scientific American, 26(1):10–12, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
10.pdf.

Anonymous:1872:OLPa

[2453] Anonymous. Official list of patents. Scientific American, 26(1):12, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n1/pdf/scientificamerican01011872-12.pdf.

Anonymous:1872:AEPa

[2454] Anonymous. Applications for extension of patents. Scientific American, 26(1):12, January 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n1/pdf/scientificamerican01011872-
12a.pdf.

Anonymous:1872:Cb

[2455] Anonymous. Correspondence. Scientific American, 26(2):20, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-20.pdf.

Anonymous:1872:NQb

[2456] Anonymous. Notes and queries. Scientific American, 26(2):26, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-26a.pdf.

Anonymous:1872:BPb

[2457] Anonymous. Business and personal. Scientific American, 26(2):26, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-26b.pdf.

Anonymous:1872:BNBa

[2458] Anonymous. Books: New books and publications. Scientific American, 26(2):27, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
27a.pdf.

Anonymous:1872:SVF

[2459] Anonymous. Ship ventilator, fog alarm, and bilge pump, celestial engineering. Scientific American, 26(2):15–16, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-15.pdf.

Anonymous:1872:CMa

[2460] Anonymous. On the colors of metals. Scientific American, 26(2):16, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-16.pdf.

Anonymous:1872:KIN

[2461] Anonymous. Keith’s improvement in nickel plating. Scientific American, 26(2):16, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
16a.pdf.

Anonymous:1872:HCTa

[2462] Anonymous. Hydraulic cements — their adaptability for useful and ornamental purposes. Scientific American, 26(2):17, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-17.pdf.

Anonymous:1872:PIL

[2463] Anonymous. The Philosophical Institution and Library, Bristol, England. Scientific American, 26(2):18–19, January 6, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n2/pdf/scientificamerican01061872-18.pdf.

VanderWeyde:1872:LHD

[2464] P. H. Vander Weyde. Latent heat of dissociation. Scientific American, 26(2):21, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
21.pdf.

Anonymous:1872:SAMa

[2465] Anonymous. Society of arts of the Massachusetts institute of technology. Scientific American, 26(2):21–22, January 6, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n2/pdf/scientificamerican01061872-21a.pdf.

Anonymous:1872:STC

[2466] Anonymous. Steam towage on canals. Scientific American, 26(2):22, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-22.pdf.

Anonymous:1872:BSA

[2467] Anonymous. Bartholomew’s self-adjusting suspender. Scientific American, 26(2):22, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
22a.pdf.

Anonymous:1872:SME

[2468] Anonymous. Six months of engineering progress. Scientific American, 26(2):23, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
23.pdf.

Anonymous:1872:PAGa

[2469] Anonymous. Proposed Atlantic and Great Western Canal. Scientific American, 26(2):23, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
23a.pdf.

Anonymous:1872:AS

[2470] Anonymous. Artificial stone. Scientific American, 26(2):23, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-23b.pdf.

Anonymous:1872:GAR

[2471] Anonymous. Granite and asphalte roadways. Scientific American, 26(2):24, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
24.pdf.

Anonymous:1872:MJG

[2472] Anonymous. Mr. Justice Grove, F.R.S. Scientific American, 26(2):24, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-24a.pdf.

Anonymous:1872:FLA

[2473] Anonymous. The first locomotives in America. Scientific American, 26(2):24–25, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
24b.pdf.

Anonymous:1872:MGP

[2474] Anonymous. Mr. Grove on patent law. Scientific American, 26(2):25, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-25.pdf.

Anonymous:1872:SIb

[2475] Anonymous. Scientific intelligence. Scientific American, 26(2):25–26, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-25a.pdf.

Anonymous:1872:AEPb

[2476] Anonymous. Applications for extension of patents. Scientific American, 26(2):26, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
26.pdf.

Anonymous:1872:ACb

[2477] Anonymous. Answers to correspondents. Scientific American, 26(2):27, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-27.pdf.

Anonymous:1872:RAFb

[2478] Anonymous. Recent American and foreign patents. Scientific American, 26(2):27, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n2/pdf/scientificamerican01061872-
27b.pdf.

Anonymous:1872:CSb

[2479] Anonymous. To city subscribers. Scientific American, 26(2):27, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-27c.pdf.

Anonymous:1872:OLPb

[2480] Anonymous. Official list of patents. Scientific American, 26(2):28, January 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n2/pdf/scientificamerican01061872-28.pdf.

Anonymous:1872:Cc

[2481] Anonymous. Correspondence. Scientific American, 26(3):36, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-36.pdf.

Anonymous:1872:BNBb

[2482] Anonymous. Books: New books and publications. Scientific American, 26(3):41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
41b.pdf.

Anonymous:1872:BPc

[2483] Anonymous. Business and personal. Scientific American, 26(3):41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-41d.pdf.

Anonymous:1872:NQc

[2484] Anonymous. Notes and queries. Scientific American, 26(3):42, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-42.pdf.

Anonymous:1872:IIa

[2485] Anonymous. Index of inventions. Scientific American, 26(3):44, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-44.pdf.

Anonymous:1872:GRCa

[2486] Anonymous. Great reduction in cost of steam power — remarkable results. Scientific American, 26(3):31, January 13, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n3/pdf/scientificamerican01131872-31.pdf.

Anonymous:1872:PAGb

[2487] Anonymous. Pooley’s automatic grain weigher. Scientific American, 26(3):31–32, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
31a.pdf.

Anonymous:1872:HCTb

[2488] Anonymous. Hydraulic cements — their adaptability for useful and ornamental purposes. Scientific American, 26(3):32–33, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-32.pdf.

Anonymous:1872:MW

[2489] Anonymous. Mineral waters. Scientific American, 26(3):33, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-33.pdf.

Anonymous:1872:FGV

[2490] Anonymous. Friction gearing versus belts and cog wheels. Scientific American, 26(3):33–35, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
33a.pdf.

Draper:1872:AM

[2491] John C. Draper. The attraction of the Moon. Scientific American, 26(3):35, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
35.pdf.

Anonymous:1872:AO

[2492] Anonymous. The Allegheny Observatory. Scientific American, 26 (3):35–36, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
35a.pdf.

VanderWeyde:1872:MTH

[2493] P. H. Vander Weyde. The modern theory of heat applied to the phenomena of latent heat, combustion, etc. Scientific American, 26(3):37, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-37.pdf.

Weyde:1872:ETD

[2494] P. H. Vandee Weyde. Expose of the tricks of the Davenport brothers. Scientific American, 26(3):37, January 13, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n3/pdf/scientificamerican01131872-37a.pdf.

Anonymous:1872:SAa

[2495] Anonymous. Sterhydraulic apparatus. Scientific American, 26 (3):38, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
38.pdf.

Anonymous:1872:LTC

[2496] Anonymous. Lewis’ twine cutter. Scientific American, 26(3):38, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-38a.pdf.

Anonymous:1872:WM

[2497] Anonymous. What is a machine? Scientific American, 26(3):39, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-39.pdf.

Anonymous:1872:PAW

[2498] Anonymous. Poisoning by antimony — the Wharton Trial. Scientific American, 26(3):39, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
39a.pdf.

Anonymous:1872:HDa

[2499] Anonymous. Health and disease. Scientific American, 26(3):39–40, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-39b.pdf.

Anonymous:1872:SAb

[2500] Anonymous. The study of alloys. Scientific American, 26(3):40, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-40.pdf.

Anonymous:1872:AFa

[2501] Anonymous. Artificial fuel. Scientific American, 26(3):40, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-40a.pdf.

Anonymous:1872:AMM

[2502] Anonymous. The American master mechanics’ association. Scientific American, 26(3):40, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
40b.pdf.

Anonymous:1872:AEE

[2503] Anonymous. The Agassiz Exploring Expedition — interesting results already obtained. Scientific American, 26(3):40–41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-40c.pdf.

Anonymous:1872:LCF

[2504] Anonymous. Loiseau’s compressed fuel. Scientific American, 26 (3):41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
41.pdf.

Anonymous:1872:CSc

[2505] Anonymous. To city subscribers. Scientific American, 26(3):41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n3/pdf/scientificamerican01131872-41a.pdf.

Anonymous:1872:AEPc

[2506] Anonymous. Applications for extension of patents. Scientific American, 26(3):41, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
41c.pdf.

Anonymous:1872:ACc

[2507] Anonymous. Answers to correspondents. Scientific American, 26 (3):42, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
42a.pdf.

Anonymous:1872:RAFc

[2508] Anonymous. Recent American and foreign patents. Scientific American, 26(3):42–43, January 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n3/pdf/scientificamerican01131872-
42b.pdf.

Anonymous:1872:Cd

[2509] Anonymous. Correspondence. Scientific American, 26(4):52–53, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-52.pdf.

Anonymous:1872:BNBc

[2510] Anonymous. Books: New books and publications. Scientific American, 26(4):57, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
57.pdf.

Anonymous:1872:BPd

[2511] Anonymous. Business and personal. Scientific American, 26(4):57, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-57a.pdf.

Anonymous:1872:NQd

[2512] Anonymous. Notes and queries. Scientific American, 26(4):58, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-58.pdf.

Anonymous:1872:IIb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-60.pdf.

Anonymous:1872:SCJ

[2514] Anonymous. Scientific Commission of Japan. Scientific American, 26(4):47, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
47.pdf.

Anonymous:1872:HIM

[2515] Anonymous. History of ice-making machinery. Scientific American, 26(4):47–49, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
47a.pdf.

Anonymous:1872:Sd

[2516] Anonymous. Soda. Scientific American, 26(4):49, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-49.pdf.

Anonymous:1872:EES

[2517] Anonymous. Extraordinary experiments on steam boilers. Scientific American, 26(4):50–51, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
50.pdf.

Anonymous:1872:FPA

[2518] Anonymous. Free piston air pump. Scientific American, 26(4):51, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-51.pdf.

Anonymous:1872:PCAa

[2519] Anonymous. Preparation and composition of alloys. Scientific American, 26(4):53, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
53.pdf.

Anonymous:1872:AHR

[2520] Anonymous. The American historical record. Scientific American, 26(4):53–54, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
53a.pdf.

Anonymous:1872:MNS

[2521] Anonymous. Mechanical needle sharpener. Scientific American, 26(4):54, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
54.pdf.

Anonymous:1872:HIE

[2522] Anonymous. Haupt’s improved easy chair. Scientific American, 26(4):54, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
54a.pdf.

Anonymous:1872:KCP

[2523] Anonymous. Knife cleaner and polisher. Scientific American, 26(4):54, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
54b.pdf.

Anonymous:1872:AW

[2524] Anonymous. Accurate workmanship. Scientific American, 26 (4):55, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
55.pdf.

Anonymous:1872:HLS

[2525] Anonymous. High and low steam. Scientific American, 26(4):55, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-55a.pdf.

Anonymous:1872:LJF

[2526] Anonymous. The late James Fisk, Jr. Scientific American, 26 (4):55, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
55b.pdf.

Anonymous:1872:TCD

[2527] Anonymous. Teach the children to draw. Scientific American, 26(4):55–56, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
55c.pdf.

Anonymous:1872:SB

[2528] Anonymous. The study of botany. Scientific American, 26(4):56, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-56.pdf.

Anonymous:1872:PAa

[2529] Anonymous. Pure air. Scientific American, 26(4):56, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-56a.pdf.

Anonymous:1872:PFE

[2530] Anonymous. Portable fire extinguisher. Scientific American, 26 (4):56, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
56b.pdf.

Anonymous:1872:SPIa

[2531] Anonymous. Scientific and practical information. Scientific American, 26(4):56–57, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
56c.pdf.

Anonymous:1872:CSd

[2532] Anonymous. To city subscribers. Scientific American, 26(4):57, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n4/pdf/scientificamerican01201872-57b.pdf.

Anonymous:1872:ACd

[2533] Anonymous. Answers to correspondents. Scientific American, 26 (4):58, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
58a.pdf.

Anonymous:1872:RAFd

[2534] Anonymous. Recent American and foreign patents. Scientific American, 26(4):58–59, January 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n4/pdf/scientificamerican01201872-
58b.pdf.

Anonymous:1872:Ce

[2535] Anonymous. Correspondence. Scientific American, 26(5):67–68, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-67.pdf.

Anonymous:1872:BNBd

[2536] Anonymous. Books: New books and publications. Scientific American, 26(5):73, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
73b.pdf.

Anonymous:1872:BPe

[2537] Anonymous. Business and personal. Scientific American, 26(5):74, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-74.pdf.

Anonymous:1872:NQe

[2538] Anonymous. Notes and queries. Scientific American, 26(5):74, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-74a.pdf.

Anonymous:1872:IIc

[2539] Anonymous. Index of inventions. Scientific American, 26(5):76, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-76.pdf.

Anonymous:1872:IWP

[2540] Anonymous. Improved wood pavement. Scientific American, 26 (5):63, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
63.pdf.

Anonymous:1872:ACM

[2541] Anonymous. The air compressors at the mont cenis tunnel. Scientific American, 26(5):63–64, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
63a.pdf.

Anonymous:1872:IGN

[2542] Anonymous. Illuminating gas — natural gas wells — coal gas — water gas fallacies. Scientific American, 26(5):64–65, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-64.pdf.

Anonymous:1872:KTH

[2543] Anonymous. Kenwood tower, highgate, near London. Scientific American, 26(5):66, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
66.pdf.

VanderWeyde:1872:ETD

[2544] P. H. Vander Weyde. Expose of the tricks of the Davenport Brothers. Scientific American, 26(5):68–69, January 27, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n5/pdf/scientificamerican01271872-68.pdf.

Anonymous:1872:ESa

[2545] Anonymous. Editorial summary. Scientific American, 26(5):69–70, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-69.pdf.

Anonymous:1872:BSCa

[2546] Anonymous. Baldwin’s saw clamp. Scientific American, 26(5):70, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-70.pdf.

Anonymous:1872:FSP

[2547] Anonymous. The Franklin statue in printing house square. Scientific American, 26(5):70, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
70a.pdf.

Anonymous:1872:HVB

[2548] Anonymous. Heating and ventilation of buildings. Scientific American, 26(5):71, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
71.pdf.

Anonymous:1872:AIM

[2549] Anonymous. The antiquity of the iron manufacture. Scientific American, 26(5):71, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
71a.pdf.

Anonymous:1872:WDW

[2550] Anonymous. What to do when you are in trouble. Scientific American, 26(5):71, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
71b.pdf.

Anonymous:1872:PW

[2551] Anonymous. Polishing wood. Scientific American, 26(5):72, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-72.pdf.

Anonymous:1872:SSP

[2552] Anonymous. The spread of small pox. Scientific American, 26 (5):72, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
72a.pdf.

Anonymous:1872:NFS

[2553] Anonymous. The natural features of South America — new and interesting discoveries. Scientific American, 26(5):72, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-72b.pdf.

Anonymous:1872:SPIb

[2554] Anonymous. Scientific and practical information. Scientific American, 26(5):73, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
73.pdf.

Anonymous:1872:CSe

[2555] Anonymous. To city subscribers. Scientific American, 26(5):73, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n5/pdf/scientificamerican01271872-73a.pdf.

Anonymous:1872:ACe

[2556] Anonymous. Answers to correspondents. Scientific American, 26 (5):74, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
74b.pdf.

Anonymous:1872:RAFe

[2557] Anonymous. Recent American and foreign patents. Scientific American, 26(5):74–76, January 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n5/pdf/scientificamerican01271872-
74c.pdf.

Anonymous:1872:Cf

[2558] Anonymous. Correspondence. Scientific American, 26(6):84, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-84.pdf.

Anonymous:1872:BPf

[2559] Anonymous. Business and personal. Scientific American, 26 (6):90, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
90.pdf.

Anonymous:1872:NQf

[2560] Anonymous. Notes and queries. Scientific American, 26(6):90, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-90a.pdf.

Anonymous:1872:IId

[2561] Anonymous. Index of inventions. Scientific American, 26(6):92, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-92.pdf.

Anonymous:1872:BNBe

[2562] Anonymous. Books: New books and publications. Scientific American, 26(6):92, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
92a.pdf.

Anonymous:1872:RSAa

[2563] Anonymous. The road steamer “America”. Scientific American, 26(6):79–82, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
79.pdf.

Anonymous:1872:TRM

[2564] Anonymous. Transmitting rotary motion at angles from the plane of rotation. Scientific American, 26(6):82, February 3, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n6/pdf/scientificamerican02031872-82.pdf.

Anonymous:1872:IWVa

[2565] Anonymous. On ice, water, vapor, and air. Scientific American, 26(6):83–84, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
83.pdf.

Anonymous:1872:AAR

[2566] Anonymous. Abridgement of the annual report of the Commissioner of Patents for the year 1871. Scientific American, 26(6):84–85, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-84a.pdf.

Anonymous:1872:IBH

[2567] Anonymous. Ice boating on the Hudson. Scientific American, 26(6):86, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
86.pdf.

Anonymous:1872:HSIa

[2568] Anonymous. How shall I introduce my invention? Scientific American, 26(6):87, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
87.pdf.

Anonymous:1872:ARC

[2569] Anonymous. Annual report of the Commissioner of Patents. Scientific American, 26(6):87, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
87a.pdf.

Anonymous:1872:HGB

[2570] Anonymous. How general business is influenced by patents. Scientific American, 26(6):87–88, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
87b.pdf.

Anonymous:1872:FFI

[2571] Anonymous. Facts and figures — importance of the latter to substantiate the former. Scientific American, 26(6):88, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-88.pdf.

Anonymous:1872:RIE

[2572] Anonymous. Russian International Exhibition. Scientific American, 26(6):88, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
88a.pdf.

Anonymous:1872:SED

[2573] Anonymous. The solar eclipse, December, 1871. Scientific American, 26(6):88, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
88b.pdf.

Anonymous:1872:HBI

[2574] Anonymous. How to build an ice boat. Scientific American, 26(6):88, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
88c.pdf.

Anonymous:1872:HBSa

[2575] Anonymous. The heating of buildings by steam. Scientific American, 26(6):88–89, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
88d.pdf.

Anonymous:1872:SPIc

[2576] Anonymous. Scientific and practical information. Scientific American, 26(6):89, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
89.pdf.

Anonymous:1872:WYFa

[2577] Anonymous. Will you favor us? Scientific American, 26(6):89, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n6/pdf/scientificamerican02031872-89a.pdf.

Anonymous:1872:ACf

[2578] Anonymous. Answers to correspondents. Scientific American, 26 (6):90–91, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
90b.pdf.

Anonymous:1872:RAFf

[2579] Anonymous. Recent American and foreign patents. Scientific American, 26(6):91–92, February 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n6/pdf/scientificamerican02031872-
91.pdf.

Anonymous:1872:Cg

[2580] Anonymous. Correspondence. Scientific American, 26(7):100, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-100.pdf.

Anonymous:1872:BPg

[2581] Anonymous. Business and personal. Scientific American, 26(7): 106, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
106b.pdf.

Anonymous:1872:NQg

[2582] Anonymous. Notes and queries. Scientific American, 26(7):106, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-106c.pdf.

Anonymous:1872:IIe

[2583] Anonymous. Index of inventions. Scientific American, 26(7):108, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-108.pdf.

Anonymous:1872:BNBf

[2584] Anonymous. Books: New books and publications. Scientific American, 26(7):108, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
108a.pdf.

Anonymous:1872:PZD

[2585] Anonymous. The Pennell and Zimmer dovetail and shaping machine. Scientific American, 26(7):95, February 10, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n7/pdf/scientificamerican02101872-95.pdf.

Anonymous:1872:CEY

[2586] Anonymous. Chemical experiments for youthful readers. Scientific American, 26(7):95–96, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
95a.pdf.

Anonymous:1872:SEL

[2587] Anonymous. State education and the labor question. Scientific American, 26(7):97, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
97.pdf.

Anonymous:1872:ROW

[2588] Anonymous. Removal of oil from wool. Scientific American, 26(7):98, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
98.pdf.

Anonymous:1872:WIS

[2589] Anonymous. Wrought iron and steel posts. Scientific American, 26(7):99, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
99.pdf.

Anonymous:1872:AMS

[2590] Anonymous. Albert’s method of securing flag halyards. Scientific American, 26(7):99, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
99a.pdf.

Anonymous:1872:PLb

[2591] Anonymous. The phosphoric light. Scientific American, 26(7): 99–100, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
99b.pdf.

Anonymous:1872:VBH

[2592] Anonymous. Variance between hydrostatic and steam pressure in boilers. Scientific American, 26(7):100–101, February 10, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n7/pdf/scientificamerican02101872-100a.pdf.

Anonymous:1872:IHH

[2593] Anonymous. Ice harvesting on the Hudson. Scientific American, 26(7):102, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
102.pdf.

Anonymous:1872:WSD

[2594] Anonymous. What shall be done with the models at the Patent Office? Scientific American, 26(7):103, February 10, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n7/pdf/scientificamerican02101872-103.pdf.

Anonymous:1872:DCb

[2595] Anonymous. Drying by cold. Scientific American, 26(7):103, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-103a.pdf.

Anonymous:1872:OR

[2596] Anonymous. To our readers. Scientific American, 26(7):103, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-103b.pdf.

Anonymous:1872:EDDa

[2597] Anonymous. The education of the deaf and dumb. Scientific American, 26(7):103, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
103c.pdf.

Anonymous:1872:PSB

[2598] Anonymous. Production of steam in boilers. Scientific American, 26(7):103–104, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
103d.pdf.

Anonymous:1872:ILW

[2599] Anonymous. Imperfection in lamps, wicks, fluids. Scientific American, 26(7):104, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
104.pdf.

Anonymous:1872:RSb

[2600] Anonymous. Religion and science. Scientific American, 26(7): 104, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
104a.pdf.

Anonymous:1872:NF

[2601] Anonymous. The Navy of the future. Scientific American, 26 (7):104, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
104b.pdf.

Anonymous:1872:SPId

[2602] Anonymous. Scientific and practical information. Scientific American, 26(7):105, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
105.pdf.

Anonymous:1872:ESb

[2603] Anonymous. Editorial summary. Scientific American, 26(7):105, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-105a.pdf.

Anonymous:1872:NSW

[2604] Anonymous. A new scientific work. Scientific American, 26(7): 106, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
106.pdf.

Anonymous:1872:WYFb

[2605] Anonymous. Will you favor us? Scientific American, 26(7):106, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n7/pdf/scientificamerican02101872-106a.pdf.

Anonymous:1872:ACg

[2606] Anonymous. Answers to correspondents. Scientific American, 26 (7):106–107, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
106d.pdf.

Anonymous:1872:RAFg

[2607] Anonymous. Recent American and foreign patents. Scientific American, 26(7):107, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
107.pdf.

Anonymous:1872:HSIb

[2608] Anonymous. How shall I introduce my invention? Scientific American, 26(7):108, February 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n7/pdf/scientificamerican02101872-
108b.pdf.

Anonymous:1872:Ch

[2609] Anonymous. Correspondence. Scientific American, 26(8):115– 116, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
115b.pdf.

Anonymous:1872:BPh

[2610] Anonymous. Business and personal. Scientific American, 26(8): 121–122, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
121c.pdf.

Anonymous:1872:NQh

[2611] Anonymous. Notes and queries. Scientific American, 26(8):122, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-122.pdf.

Anonymous:1872:IIf

[2612] Anonymous. Index of inventions. Scientific American, 26(8):124, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-124.pdf.

Anonymous:1872:BNBg

[2613] Anonymous. Books: New books and publications. Scientific American, 26(8):124, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
124b.pdf.

Anonymous:1872:IHPa

[2614] Anonymous. Improved horse power, mastodon remains, and more. Scientific American, 26(8):111–112, February 17, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n8/pdf/scientificamerican02171872-111.pdf.

Anonymous:1872:LTM

[2615] Anonymous. The law of trade marks in England. Scientific American, 26(8):112, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
112.pdf.

Chandler:1872:EAE

[2616] Charles F. Chandler. Effect of animal excreta in water. Scientific American, 26(8):113, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
113.pdf.

Anonymous:1872:APR

[2617] Anonymous. Asbestos piston rod packing. Scientific American, 26 (8):113–114, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
113a.pdf.

Anonymous:1872:HIH

[2618] Anonymous. Harvesting ice on the Hudson. Scientific American, 26(8):115, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
115.pdf.

Anonymous:1872:FRR

[2619] Anonymous. Fall of the roof of the railroad depot at Saratoga Springs. Scientific American, 26(8):115, February 17, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n8/pdf/scientificamerican02171872-115a.pdf.

Anonymous:1872:J

[2620] Anonymous. Japan. Scientific American, 26(8):116, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-116.pdf.

Anonymous:1872:MSA

[2621] Anonymous. Meeting of the Society of Arts of the Massachusetts Institute of Technology. Scientific American, 26(8):116–118, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-116a.pdf.

Anonymous:1872:GIW

[2622] Anonymous. Gerhart’s improved wagon brake. Scientific American, 26(8):118, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
118.pdf.

Anonymous:1872:QGA

[2623] Anonymous. Quinby’s geometrical arching bricks. Scientific American, 26(8):118, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
118a.pdf.

Anonymous:1872:HBSb

[2624] Anonymous. The heating of buildings by steam. Scientific American, 26(8):119, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
119.pdf.

Anonymous:1872:SOVa

[2625] Anonymous. Substituting other vapors for steam — ether and bisulphide of carbon. Scientific American, 26(8):119, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-119a.pdf.

Anonymous:1872:III

[2626] Anonymous. The indirect influence of invention upon manufactures, arts, and commerce. Scientific American, 26(8):119–120, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-119b.pdf.

Anonymous:1872:ACCa

[2627] Anonymous. Agricultural chemistry and chemical manures. Scientific American, 26(8):120, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
120.pdf.

Anonymous:1872:SEF

[2628] Anonymous. Short extracts from a few letters. Scientific American, 26(8):120–121, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
120a.pdf.

Anonymous:1872:SPIe

[2629] Anonymous. Scientific and practical information. Scientific American, 26(8):121, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
121.pdf.

Anonymous:1872:WYFc

[2630] Anonymous. Will you favor us? Scientific American, 26(8):121, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n8/pdf/scientificamerican02171872-121a.pdf.

Anonymous:1872:GMW

[2631] Anonymous. A good Month’s work. Scientific American, 26 (8):121, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
121b.pdf.

Anonymous:1872:ACh

[2632] Anonymous. Answers to correspondents. Scientific American, 26 (8):122, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
122a.pdf.

Anonymous:1872:RAFh

[2633] Anonymous. Recent American and foreign patents. Scientific American, 26(8):123, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
123.pdf.

Anonymous:1872:HSIc

[2634] Anonymous. How shall I introduce my invention? Scientific American, 26(8):124, February 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n8/pdf/scientificamerican02171872-
124a.pdf.

Anonymous:1872:Ci

[2635] Anonymous. Correspondence. Scientific American, 26(9):132– 133, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
132.pdf.

Anonymous:1872:BPi

[2636] Anonymous. Business and personal. Scientific American, 26(9): 137, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
137.pdf.

Anonymous:1872:NQi

[2637] Anonymous. Notes and queries. Scientific American, 26(9):138, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n9/pdf/scientificamerican02241872-138.pdf.

Anonymous:1872:IIg

[2638] Anonymous. Index of inventions. Scientific American, 26(9):140, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n9/pdf/scientificamerican02241872-140.pdf.

Anonymous:1872:BNBh

[2639] Anonymous. Books: New books and publications. Scientific American, 26(9):140, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
140a.pdf.

Anonymous:1872:IHPb

[2640] Anonymous. Improved high pressure alarm. Scientific American, 26(9):127, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
127.pdf.

Anonymous:1872:MSS

[2641] Anonymous. Mining speculation — something as applicable to New York as to London. Scientific American, 26(9):128–129, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n9/pdf/scientificamerican02241872-128.pdf.

Anonymous:1872:LTB

[2642] Anonymous. Lecture theater buildings at South Kensington. Scientific American, 26(9):130, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
130.pdf.

Anonymous:1872:CSSa

[2643] Anonymous. Coloring seeds for sowing. Scientific American, 26 (9):131, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
131.pdf.

Anonymous:1872:PCC

[2644] Anonymous. Patent crocheted caps for infants. Scientific American, 26(9):131, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
131a.pdf.

Anonymous:1872:RSP

[2645] Anonymous. A Remedy for small pox by one who has tried it. Scientific American, 26(9):133, February 24, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n9/pdf/scientificamerican02241872-133.pdf.

Anonymous:1872:EMTa

[2646] Anonymous. The electro-magnetic telegraph — honor to whom honor is due. Scientific American, 26(9):133, February 24, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n9/pdf/scientificamerican02241872-133a.pdf.

Anonymous:1872:PODa

[2647] Anonymous. Patent Office decision. Scientific American, 26(9): 133–134, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
133b.pdf.

Anonymous:1872:SSH

[2648] Anonymous. Simons’ safety hold-back for carriages. Scientific American, 26(9):134, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
134.pdf.

Anonymous:1872:SOVb

[2649] Anonymous. Substituting other vapors for steam — vapor of liquefied carbonic acid. Scientific American, 26(9):135, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n9/pdf/scientificamerican02241872-135.pdf.

Anonymous:1872:SSR

[2650] Anonymous. Sulphite of soda as a Remedy for small pox. Scientific American, 26(9):135, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
135a.pdf.

Anonymous:1872:CSG

[2651] Anonymous. Compulsory safety gages — the steamboat owners’ protest. Scientific American, 26(9):135, February 24, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n9/pdf/scientificamerican02241872-135b.pdf.

Anonymous:1872:ESBa

[2652] Anonymous. The effect of sound in buildings as influenced by ventilation. Scientific American, 26(9):135–136, February 24, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n9/pdf/scientificamerican02241872-135c.pdf.

Anonymous:1872:DBB

[2653] Anonymous. The difference between benzole and benzine. Scientific American, 26(9):136, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
136.pdf.

Anonymous:1872:AT

[2654] Anonymous. The art of thinking. Scientific American, 26(9):136, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n9/pdf/scientificamerican02241872-136a.pdf.

Anonymous:1872:JNS

[2655] Anonymous. Japanese native steel. Scientific American, 26(9): 136, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
136b.pdf.

Anonymous:1872:SPIf

[2656] Anonymous. Scientific and practical information. Scientific American, 26(9):136–137, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
136c.pdf.

Anonymous:1872:HSId

[2657] Anonymous. How shall I introduce my invention? Scientific American, 26(9):137, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
137a.pdf.

Anonymous:1872:ACi

[2658] Anonymous. Answers to correspondents. Scientific American, 26 (9):138, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
138a.pdf.

Anonymous:1872:RAFi

[2659] Anonymous. Recent American and foreign patents. Scientific American, 26(9):139, February 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n9/pdf/scientificamerican02241872-
139.pdf.

Anonymous:1872:Cj

[2660] Anonymous. Correspondence. Scientific American, 26(10):148–149, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-148.pdf.

Anonymous:1872:BPj

[2661] Anonymous. Business and personal. Scientific American, 26(10):154, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-154.pdf.

Anonymous:1872:NQj

[2662] Anonymous. Notes and queries. Scientific American, 26(10):154–155, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-154a.pdf.

Anonymous:1872:IIh

[2663] Anonymous. Index of inventions. Scientific American, 26(10):156, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-156.pdf.

Anonymous:1872:GSFa

[2664] Anonymous. The Gatling system of firearms. Scientific American, 26(10):143–146, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
143.pdf.

Anonymous:1872:TWS

[2665] Anonymous. Triscott and Wheeler’s shoe fastenings. Scientific American, 26(10):146, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
146.pdf.

Anonymous:1872:PIMa

[2666] Anonymous. Patton’s improved meat chopper. Scientific American, 26(10):147, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
147.pdf.

Anonymous:1872:REa

[2667] Anonymous. Rotary engines. Scientific American, 26(10):147–148, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-147a.pdf.

Anonymous:1872:MPC

[2668] Anonymous. Middle Park, Colorado. Scientific American, 26 (10):149–150, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
149.pdf.

Anonymous:1872:PP

[2669] Anonymous. Publication of the patents. Scientific American, 26 (10):151, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
151.pdf.

Anonymous:1872:CPE

[2670] Anonymous. Congress and patent extension cases. Scientific American, 26(10):151, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
151a.pdf.

Anonymous:1872:DSS

[2671] Anonymous. Drying substances by steam. Scientific American, 26 (10):151–152, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
151b.pdf.

Anonymous:1872:AHa

[2672] Anonymous. American homesteads. Scientific American, 26(10):152, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n10/pdf/scientificamerican03021872-152.pdf.

Anonymous:1872:SIc

[2673] Anonymous. The stupidity of ignorance. Scientific American, 26(10):152, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
152a.pdf.

Anonymous:1872:PML

[2674] Anonymous. The positive motion loom grand medal of honor. Scientific American, 26(10):152, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
152b.pdf.

Anonymous:1872:SPIg

[2675] Anonymous. Scientific and practical information. Scientific American, 26(10):152, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
152c.pdf.

Anonymous:1872:ECB

[2676] Anonymous. Extension cases before Congress. Scientific American, 26(10):153, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
153.pdf.

Anonymous:1872:ACj

[2677] Anonymous. Answers to correspondents. Scientific American, 26 (10):155, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
155.pdf.

Anonymous:1872:RAFj

[2678] Anonymous. Recent American and foreign patents. Scientific American, 26(10):155–156, March 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n10/pdf/scientificamerican03021872-
155a.pdf.

Anonymous:1872:Ck

[2679] Anonymous. Correspondence. Scientific American, 26(11):163–164, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-163.pdf.

Anonymous:1872:NQk

[2680] Anonymous. Notes and queries. Scientific American, 26(11):169–170, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-169b.pdf.

Anonymous:1872:BPk

[2681] Anonymous. Business and personal. Scientific American, 26(11):170, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-170.pdf.

Anonymous:1872:IIj

[2682] Anonymous. Index of inventions. Scientific American, 26(11):172, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-172.pdf.

Anonymous:1872:BNBi

[2683] Anonymous. Books: New books and publications. Scientific American, 26(11):172, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
172a.pdf.

Anonymous:1872:IPS

[2684] Anonymous. Improved portable steam brick machine, traction engine. Scientific American, 26(11):159–160, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-159.pdf.

Anonymous:1872:SMTa

[2685] Anonymous. Stove manufactories in troy. Scientific American, 26(11):160, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
160.pdf.

Anonymous:1872:IPCa

[2686] Anonymous. Inventions and patents connected with building ing. Scientific American, 26(11):161–162, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-161.pdf.

Anonymous:1872:CSBa

[2687] Anonymous. The Comba Scura Bridge — Mont Cenis Railway. Scientific American, 26(11):162–163, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-162.pdf.

Anonymous:1872:MCa

[2688] Anonymous. Mining in Colorado. Scientific American, 26(11):162–163, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-162a.pdf.

Anonymous:1872:SAMb

[2689] Anonymous. Society of arts of the Massachusetts institute of technology. Scientific American, 26(11):164, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-164.pdf.

Anonymous:1872:LWW

[2690] Anonymous. The Lowell water wheel tests. Scientific American, 26(11):165–166, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
165.pdf.

Anonymous:1872:BIS

[2691] Anonymous. The Beckwith improved sewing machine. Scientific American, 26(11):166, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
166.pdf.

Anonymous:1872:SCa

[2692] Anonymous. Science in the courts. Scientific American, 26(11):167, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-167.pdf.

Anonymous:1872:BSUa

[2693] Anonymous. Beet sugar in the United States. Scientific American, 26(11):167, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
167a.pdf.

Anonymous:1872:DSH

[2694] Anonymous. Drying substances by, hot air. Scientific American, 26(11):167–168, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
167b.pdf.

Anonymous:1872:AOP

[2695] Anonymous. The amalgamation of ores of the precious metals. Scientific American, 26(11):168, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-168.pdf.

Anonymous:1872:KSH

[2696] Anonymous. The knowledge of specific heat applied to the determination of very high temperatures. Scientific American, 26(11):168, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n11/pdf/scientificamerican03091872-168a.pdf.

Anonymous:1872:SPIh

[2697] Anonymous. Scientific and practical information. Scientific American, 26(11):168, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
168b.pdf.

Anonymous:1872:PFS

[2698] Anonymous. Psychic force — spirit faces. Scientific American, 26(11):169, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
169.pdf.

Anonymous:1872:ADT

[2699] Anonymous. American district telegraph company — a new enterprise. Scientific American, 26(11):169, March 9, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n11/pdf/scientificamerican03091872-169a.pdf.

Anonymous:1872:ACk

[2700] Anonymous. Answers to correspondents. Scientific American, 26 (11):170, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
170a.pdf.

Anonymous:1872:RAFk

[2701] Anonymous. Recent American and foreign patents. Scientific American, 26(11):171, March 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n11/pdf/scientificamerican03091872-
171.pdf.

Anonymous:1872:Cl

[2702] Anonymous. Correspondence. Scientific American, 26(12):179–180, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-179b.pdf.

Anonymous:1872:BPl

[2703] Anonymous. Business and personal. Scientific American, 26(12):185, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-185.pdf.

Anonymous:1872:NQl

[2704] Anonymous. Notes and queries. Scientific American, 26(12):185, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-185a.pdf.

Anonymous:1872:IIk

[2705] Anonymous. Index of inventions. Scientific American, 26(12):187, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-187.pdf.

Anonymous:1872:CAT

[2706] Anonymous. Colt’s armory testing machine, railway power brakes. Scientific American, 26(12):175–176, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
175.pdf.

Anonymous:1872:ILP

[2707] Anonymous. Interesting lecture by Professor Tyndall — the identity of light and radiant heat. Scientific American, 26(12):176–177, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-176.pdf.

Anonymous:1872:NYF

[2708] Anonymous. The New York fire alarm telegraph. Scientific American, 26(12):177, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
177.pdf.

Anonymous:1872:YEE

[2709] Anonymous. The Yale Exploring Expedition of 1871. Scientific American, 26(12):177, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
177a.pdf.

Anonymous:1872:PPD

[2710] Anonymous. Printing the patent drawings. Scientific American, 26(12):177, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
177b.pdf.

Anonymous:1872:IG

[2711] Anonymous. Interior of a greenhouse. Scientific American, 26(12):178, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-178.pdf.

Anonymous:1872:JIP

[2712] Anonymous. Jones’ improved projectile for firearms. Scientific American, 26(12):178–179, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
178a.pdf.

Anonymous:1872:MEL

[2713] Anonymous. Millener’s extension ladder. Scientific American, 26(12):179, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
179.pdf.

Draper:1872:MMT

[2714] John C. Draper. Magnesium in Marsh’s test for arsenic. Scientific American, 26(12):179, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
179a.pdf.

Walker:1872:CGS

[2715] William L. Walker. The color of the Gulf Stream, and its excess of solid matter accounted for. Scientific American, 26(12):180, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-180.pdf.

Anonymous:1872:PSC

[2716] Anonymous. The progress of stellar chemistry. Scientific American, 26(12):181–182, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
181.pdf.

Anonymous:1872:WWW

[2717] Anonymous. Walsh’s water wheel governor. Scientific American, 26(12):182, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
182.pdf.

Anonymous:1872:CIS

[2718] Anonymous. Chapman’s improved saw filing machine. Scientific American, 26(12):182, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
182a.pdf.

Anonymous:1872:AAI

[2719] Anonymous. American art industry. Scientific American, 26(12):183, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-183.pdf.

Anonymous:1872:ETT

[2720] Anonymous. Are experimental tests of turbine wheels trustworthy? Scientific American, 26(12):183, March 16, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n12/pdf/scientificamerican03161872-183a.pdf.

Anonymous:1872:TUC

[2721] Anonymous. Tunnelling under city streets — the atmospheric railway. Scientific American, 26(12):183, March 16, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n12/pdf/scientificamerican03161872-183b.pdf.

Anonymous:1872:MCb

[2722] Anonymous. A Munchausen comet. Scientific American, 26(12):183, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n12/pdf/scientificamerican03161872-183c.pdf.

Anonymous:1872:CLT

[2723] Anonymous. The common law of trademarks. Scientific American, 26(12):183–184, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
183d.pdf.

Anonymous:1872:LPC

[2724] Anonymous. Labor and personal cleanliness. Scientific American, 26(12):184, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
184.pdf.

Anonymous:1872:MHC

[2725] Anonymous. Measuring the heat of combustion. Scientific American, 26(12):184, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
184a.pdf.

Anonymous:1872:SPIi

[2726] Anonymous. Scientific and practical information. Scientific American, 26(12):184–185, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
184b.pdf.

Anonymous:1872:ACl

[2727] Anonymous. Answers to correspondents. Scientific American, 26 (12):185–186, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
185b.pdf.

Anonymous:1872:RAFl

[2728] Anonymous. Recent American and foreign patents. Scientific American, 26(12):186–187, March 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n12/pdf/scientificamerican03161872-
186.pdf.

Anonymous:1872:BNBj

[2729] Anonymous. Books: New books and publications. Scientific American, 26(13):192, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
192a.pdf.

Anonymous:1872:Cn

[2730] Anonymous. Correspondence. Scientific American, 26(13):195–197, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-195c.pdf.

Anonymous:1872:BPm

[2731] Anonymous. Business and personal. Scientific American, 26(13):200, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-200.pdf.

Anonymous:1872:NQm

[2732] Anonymous. Notes and queries. Scientific American, 26(13):200, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-200a.pdf.

Anonymous:1872:IIl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-203.pdf.

Anonymous:1872:CRS

[2734] Anonymous. Collecting and removing scums from boilers. Scientific American, 26(13):191, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
191.pdf.

Anonymous:1872:BMB

[2735] Anonymous. Bridge over the Mississippi at Brainerd, Minn — Northern Pacific Railway. Scientific American, 26(13):191, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-191a.pdf.

Anonymous:1872:NBa

[2736] Anonymous. A navigable balloon. Scientific American, 26(13):192, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-192.pdf.

Anonymous:1872:PSE

[2737] Anonymous. The Parkes silver extracting process in Germany. Scientific American, 26(13):193–194, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
193.pdf.

Anonymous:1872:REb

[2738] Anonymous. Rotary engines. Scientific American, 26(13):194, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-194.pdf.

Anonymous:1872:HWR

[2739] Anonymous. Hallidie’s wire rope way for transporting ores. Scientific American, 26(13):195, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
195.pdf.

Draper:1872:PAH

[2740] John C. Draper. Precipitation of arsenic by heat in Marsh’s test. Scientific American, 26(13):195, March 23, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n13/pdf/scientificamerican03231872-195a.pdf.

Anonymous:1872:DHD

[2741] Anonymous. Determination of high degrees of heat by the melting of ice. Scientific American, 26(13):195, March 23, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n13/pdf/scientificamerican03231872-195b.pdf.

Anonymous:1872:SPIj

[2742] Anonymous. Scientific and practical information. Scientific American, 26(13):197, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
197.pdf.

Anonymous:1872:BSCb

[2743] Anonymous. Beach’s spark catcher and consumer. Scientific American, 26(13):198, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
198.pdf.

Anonymous:1872:HEB

[2744] Anonymous. Himmer’s electric battery. Scientific American, 26(13):198, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-198a.pdf.

Anonymous:1872:KPC

[2745] Anonymous. Ketchum’s Pitman connection. Scientific American, 26(13):198, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
198b.pdf.

Anonymous:1872:EMTb

[2746] Anonymous. The electro magnetic telegraph. Scientific American, 26(13):199, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
199.pdf.

Anonymous:1872:PFa

[2747] Anonymous. The problem of flight. Scientific American, 26(13):199, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n13/pdf/scientificamerican03231872-199a.pdf.

Anonymous:1872:LVH

[2748] Anonymous. Light versus heavy shafting. Scientific American, 26 (13):199–200, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
199b.pdf.

Anonymous:1872:ACn

[2749] Anonymous. Answers to correspondents. Scientific American, 26 (13):200–201, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
200b.pdf.

Anonymous:1872:RAFm

[2750] Anonymous. Recent American and foreign patents. Scientific American, 26(13):201–203, March 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n13/pdf/scientificamerican03231872-
201.pdf.

Anonymous:1872:Co

[2751] Anonymous. Correspondence. Scientific American, 26(14):212, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-212.pdf.

Anonymous:1872:BNBk 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
216a.pdf.

Anonymous:1872:BPn

[2753] Anonymous. Business and personal. Scientific American, 26(14):216, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-216b.pdf.

Anonymous:1872:NQn

[2754] Anonymous. Notes and queries. Scientific American, 26(14):217, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-217.pdf.

Anonymous:1872:IIm

[2755] Anonymous. Index of inventions. Scientific American, 26(14):219, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-219.pdf.

Anonymous:1872:MTS

[2756] Anonymous. Machine for testing the strength of metals, wire, rope, chain, cast iron columns, bridge bolts, and boiler plates, junction of the Black and Caspian Seas, and more. Scientific American, 26(14):207, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-207.pdf.

Anonymous:1872:MBPa

[2757] Anonymous. Machine for bending printers’ leads and rules. Scientific American, 26(14):207–208, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
207a.pdf.

Anonymous:1872:GM

[2758] Anonymous. Gunmaking by machinery. Scientific American, 26 (14):208, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
208.pdf.

Anonymous:1872:SRc

[2759] Anonymous. State reports. Scientific American, 26(14):208, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-208a.pdf.

Anonymous:1872:TC

[2760] Anonymous. Terra cotta. Scientific American, 26(14):209–210, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-209.pdf.

Anonymous:1872:JRT

[2761] Anonymous. Johnson’s renewable tooth comb. Scientific American, 26(14):210, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
210.pdf.

Anonymous:1872:MG

[2762] Anonymous. Moustache guard. Scientific American, 26(14):210, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-210a.pdf.

Draper:1872:QDA

[2763] John C. Draper. The quantitative determination of arsenic. Scientific American, 26(14):211, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
211.pdf.

Anonymous:1872:PJa

[2764] Anonymous. The Planet Jupiter. Scientific American, 26(14):211, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-211a.pdf.

Anonymous:1872:WBS

[2765] Anonymous. Wright’s bottle stopper. Scientific American, 26(14):211, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-211b.pdf.

Anonymous:1872:GRN

[2766] Anonymous. Geological report for New Jersey. Scientific American, 26(14):212–213, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
212a.pdf.

Anonymous:1872:SPIk

[2767] Anonymous. Scientific and practical information. Scientific American, 26(14):213–214, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
213.pdf.

Anonymous:1872:MDF

[2768] Anonymous. Melendy’s door fastener. Scientific American, 26(14):214, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-214.pdf.

Anonymous:1872:BSS

[2769] Anonymous. Beach’s spiral Spring scroll sawing machine. Scientific American, 26(14):214, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
214a.pdf.

Anonymous:1872:RAI

[2770] Anonymous. The relations of art to industry — museums of the arts of design. Scientific American, 26(14):215, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n14/pdf/scientificamerican03301872-215.pdf.

Anonymous:1872:MOS

[2771] Anonymous. Making organic substances from their inorganic elements. Scientific American, 26(14):215, March 30, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n14/pdf/scientificamerican03301872-215a.pdf.

Anonymous:1872:MIa

[2772] Anonymous. Mechanical impossibilities. Scientific American, 26 (14):215, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
215b.pdf.

Anonymous:1872:HVM

[2773] Anonymous. Hydrogen the vapor of the metal hydrium. Scientific American, 26(14):215–216, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
215c.pdf.

Anonymous:1872:ABP

[2774] Anonymous. Another big patent job before Congress. Scientific American, 26(14):216, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
216.pdf.

Anonymous:1872:ACo

[2775] Anonymous. Answer to correspondents. Scientific American, 26 (14):217–218, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
217a.pdf.

Anonymous:1872:RAFn

[2776] Anonymous. Recent American and foreign patents. Scientific American, 26(14):218–219, March 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n14/pdf/scientificamerican03301872-
218.pdf.

Anonymous:1872:Cp

[2777] Anonymous. Correspondence. Scientific American, 26(15):228, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-228.pdf.

Anonymous:1872:BPo

[2778] Anonymous. Business and personal. Scientific American, 26 (15):232–231, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
232a.pdf.

Anonymous:1872:NQo

[2779] Anonymous. Notes and queries. Scientific American, 26(15):233, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-233.pdf.

Anonymous:1872:IIn

[2780] Anonymous. Index of inventions. Scientific American, 26(15):235, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-235.pdf.

Anonymous:1872:BNBl

[2781] Anonymous. Books: New books and publications. Scientific American, 26(15):236, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
236.pdf.

Anonymous:1872:PSG

[2782] Anonymous. Portable Spring grist mill, Price’s composition pavement, and more. Scientific American, 26(15):223–224, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-223.pdf.

Anonymous:1872:TMB

[2783] Anonymous. The tobacco manufactories in Brooklyn. Scientific American, 26(15):224, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
224.pdf.

Anonymous:1872:GWH

[2784] Anonymous. The government works at Hell Gate. Scientific American, 26(15):224–226, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
224a.pdf.

Anonymous:1872:PSJ

[2785] Anonymous. Palmer’s stave jointing machine. Scientific American, 26(15):226, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
226.pdf.

Leach:1872:DHCa

[2786] J. H. Leach. The discovery and history of the chromosphere. Scientific American, 26(15):227, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
227.pdf.

Anonymous:1872:POC

[2787] Anonymous. A Patent Office curiosity, and a specification in rhyme. Scientific American, 26(15):227, April 6, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n15/pdf/scientificamerican04061872-227a.pdf.

Wicklin:1872:FGa

[2788] E. S. Wicklin. Frictional gearing. Scientific American, 26(15):227–228, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-227b.pdf.

Anonymous:1872:SAMc

[2789] Anonymous. Society of arts of the Massachusetts institute of technology. Scientific American, 26(15):229, April 6, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n15/pdf/scientificamerican04061872-229.pdf.

Anonymous:1872:VPI

[2790] Anonymous. Valuable photographic improvement. Scientific American, 26(15):229, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
229a.pdf.

Anonymous:1872:HEM

[2791] Anonymous. The Holtz electrical machine. Scientific American, 26(15):229, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
229b.pdf.

Anonymous:1872:SPIl

[2792] Anonymous. Scientific and practical information. Scientific American, 26(15):229, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
229c.pdf.

Anonymous:1872:FG

[2793] Anonymous. Friction gearing. Scientific American, 26(15):229–230, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n15/pdf/scientificamerican04061872-229d.pdf.

Anonymous:1872:KIS

[2794] Anonymous. Kehoe’s improved striking bag. Scientific American, 26(15):230, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
230.pdf.

Anonymous:1872:CTC

[2795] Anonymous. Currier’s tea and coffee pot stand. Scientific American, 26(15):230, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
230a.pdf.

Anonymous:1872:ADI

[2796] Anonymous. African diamonds — an invention wanted. Scientific American, 26(15):231, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
231.pdf.

Anonymous:1872:SPP

[2797] Anonymous. Small pox — its prevention — the treatment of patients with reference to the protection of the public. Scientific American, 26(15):231, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
231a.pdf.

Anonymous:1872:RTI

[2798] Anonymous. A remarkable telegraphic improvement. Scientific American, 26(15):231, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
231b.pdf.

Anonymous:1872:EDDb

[2799] Anonymous. Epidemics of disease, disaster, and crime. Scientific American, 26(15):231–232, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
231c.pdf.

Anonymous:1872:PSO

[2800] Anonymous. The present status of organic chemistry. Scientific American, 26(15):232, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
232.pdf.

Anonymous:1872:ACp

[2801] Anonymous. Answers to correspondents. Scientific American, 26 (15):233, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
233a.pdf.

Anonymous:1872:RAFo

[2802] Anonymous. Recent American and foreign patents. Scientific American, 26(15):234–235, April 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n15/pdf/scientificamerican04061872-
234.pdf.

Anonymous:1872:BNBm

[2803] Anonymous. Books: New books and publications. Scientific American, 26(16):248, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
248c.pdf.

Anonymous:1872:BPp

[2804] Anonymous. Business and personal. Scientific American, 26(16):249, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-249.pdf.

Anonymous:1872:NQp

[2805] Anonymous. Notes and queries. Scientific American, 26(16):249, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-249a.pdf.

Anonymous:1872:IIo

[2806] Anonymous. Index of inventions. Scientific American, 26(16):251, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-251.pdf.

Anonymous:1872:IMP

[2807] Anonymous. Improved method of propelling canal boats. Scientific American, 26(16):239–240, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
239.pdf.

Anonymous:1872:RWI

[2808] Anonymous. The Remington Works Ilion, N.Y. Scientific American, 26(16):240, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
240.pdf.

Anonymous:1872:DCN

[2809] Anonymous. Diamond cutting in New York. Scientific American, 26(16):241–242, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
241.pdf.

Anonymous:1872:BBI

[2810] Anonymous. The Baltimore Boiler Inspection Law. Scientific American, 26(16):242, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
242.pdf.

Anonymous:1872:FNT

[2811] Anonymous. Fastenings for neck ties, scarfs, etc. Scientific American, 26(16):242, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
242a.pdf.

Anonymous:1872:MVR

[2812] Anonymous. Measuring the velocity of rotation. Scientific American, 26(16):242, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
242b.pdf.

Leach:1872:DHCb

[2813] J. H. Leach. The discovery and history of the chromosphere. Scientific American, 26(16):243, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
243.pdf.

Anonymous:1872:HV

[2814] Anonymous. Heard’s ventilating hat. Scientific American, 26(16):243, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-243a.pdf.

Anonymous:1872:ISL

[2815] Anonymous. Inextinguishable signal light. Scientific American, 26(16):243, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
243b.pdf.

Anonymous:1872:CCI

[2816] Anonymous. Combined cast iron and wrought iron arch girders. Scientific American, 26(16):243–244, April 13, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n16/pdf/scientificamerican04131872-243c.pdf.

Wicklin:1872:FGb

[2817] E. S. Wicklin. Frictional gearing. Scientific American, 26(16):244, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-244.pdf.

Anonymous:1872:SPIm

[2818] Anonymous. Scientific and practical information. Scientific American, 26(16):245, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
245.pdf.

Anonymous:1872:UF

[2819] Anonymous. Umbrella frames. Scientific American, 26(16):245, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-245a.pdf.

Anonymous:1872:PCRa

[2820] Anonymous. Proposed city railroad. Scientific American, 26(16):246, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-246.pdf.

Anonymous:1872:GSM

[2821] Anonymous. The great sewing machine job before Congress. Scientific American, 26(16):247, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
247.pdf.

Anonymous:1872:SMTb

[2822] Anonymous. Sewing machines and their effect upon health. Scientific American, 26(16):247, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
247a.pdf.

Anonymous:1872:DSF

[2823] Anonymous. Death of Samuel F. B. Morse. Scientific American, 26(16):247–248, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
247b.pdf.

Anonymous:1872:SAS

[2824] Anonymous. Spiritualism answered by science. Scientific American, 26(16):248, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
248.pdf.

Anonymous:1872:TUT

[2825] Anonymous. A trades union tirade. Scientific American, 26(16):248, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-248a.pdf.

Anonymous:1872:LRa

[2826] Anonymous. Lightning rods. Scientific American, 26(16):248, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n16/pdf/scientificamerican04131872-248b.pdf.

Anonymous:1872:ACq

[2827] Anonymous. Answers to correspondents. Scientific American, 26 (16):249, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
249b.pdf.

Anonymous:1872:RAFp

[2828] Anonymous. Recent American and foreign patents. Scientific American, 26(16):250, April 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n16/pdf/scientificamerican04131872-
250.pdf.

Anonymous:1872:Cq

[2829] Anonymous. Correspondence. Scientific American, 26(17):260–262, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-260.pdf.

Anonymous:1872:BPq

[2830] Anonymous. Business and personal. Scientific American, 26(17):265, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-265.pdf.

Anonymous:1872:NQq

[2831] Anonymous. Notes and queries. Scientific American, 26(17):265–266, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-265a.pdf.

Anonymous:1872:BNBn

[2832] Anonymous. Books: New books and publications. Scientific American, 26(17):267, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
267.pdf.

Anonymous:1872:IIp

[2833] Anonymous. Index of inventions. Scientific American, 26(17):268, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-268.pdf.

Anonymous:1872:ETR

[2834] Anonymous. Endless traveling or railway sidewalk, the crystallization of metals, and more. Scientific American, 26(17):255–257, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-255.pdf.

Anonymous:1872:EMTc

[2835] Anonymous. The electro-magnetic telegraph. Scientific American, 26(17):257, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
257.pdf.

Anonymous:1872:ITM

[2836] Anonymous. Important trade mark decision. Scientific American, 26(17):257, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
257a.pdf.

Anonymous:1872:PMa

[2837] Anonymous. Professor Morse. Scientific American, 26(17):258, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-258.pdf.

Anonymous:1872:SPIn

[2838] Anonymous. Scientific and practical information. Scientific American, 26(17):259, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
259.pdf.

Anonymous:1872:CLR

[2839] Anonymous. Clothes line reel. Scientific American, 26(17):259, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-259a.pdf.

Wicklin:1872:FGc

[2840] E. S. Wicklin. Frictional gearing. Scientific American, 26(17):259, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-259b.pdf.

Anonymous:1872:MCS

[2841] Anonymous. Measuring the clearness of the sky. Scientific American, 26(17):259–260, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
259c.pdf.

Anonymous:1872:GWa

[2842] Anonymous. Gibbs’ whiffletree. Scientific American, 26(17):262, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-262.pdf.

Anonymous:1872:WSA

[2843] Anonymous. Winn’s shampooing apparatus. Scientific American, 26(17):262, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
262a.pdf.

Anonymous:1872:IGT

[2844] Anonymous. An improved grafting tool. Scientific American, 26 (17):262, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
262b.pdf.

Anonymous:1872:DDA

[2845] Anonymous. Drying by the direct application of heat-distillation. Scientific American, 26(17):263, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
263.pdf.

Anonymous:1872:RSR

[2846] Anonymous. The relation of science to religion. Scientific American, 26(17):263, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
263a.pdf.

Anonymous:1872:PRP

[2847] Anonymous. Piston rod packing. Scientific American, 26(17):263–264, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-263b.pdf.

Anonymous:1872:API

[2848] Anonymous. The advance in the price of iron. Scientific American, 26(17):264, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
264.pdf.

Anonymous:1872:SEC

[2849] Anonymous. Steam on the Erie Canal. Scientific American, 26(17):264, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n17/pdf/scientificamerican04201872-264a.pdf.

Anonymous:1872:BSUb

[2850] Anonymous. Beet sugar in the United States. Scientific American, 26(17):264–265, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
264b.pdf.

Anonymous:1872:ACr

[2851] Anonymous. Answers to correspondents. Scientific American, 26 (17):266, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
266.pdf.

Anonymous:1872:RAFq

[2852] Anonymous. Recent American and foreign patents. Scientific American, 26(17):266–267, April 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n17/pdf/scientificamerican04201872-
266a.pdf.

Anonymous:1872:BPr

[2853] Anonymous. Business and personal. Scientific American, 26 (18):280–281, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
280c.pdf.

Anonymous:1872:NQr

[2854] Anonymous. Notes and queries. Scientific American, 26(18):281, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-281.pdf.

Anonymous:1872:BNBo

[2855] Anonymous. Books: New books and publications. Scientific American, 26(18):281, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
281b.pdf.

Anonymous:1872:IIq

[2856] Anonymous. Index of inventions. Scientific American, 26(18):283, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-283.pdf.

Anonymous:1872:IBK

[2857] Anonymous. Improved brick kiln. Scientific American, 26(18):271–272, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-271.pdf.

Anonymous:1872:HWB

[2858] Anonymous. On the heating of water and boilers — curious explosive phenomena. Scientific American, 26(18):272, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-272.pdf.

Anonymous:1872:CKM

[2859] Anonymous. Carpets and knitted manufactures of Amsterdam, New York. Scientific American, 26(18):272–273, April 27, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n18/pdf/scientificamerican04271872-272a.pdf.

Fraser:1872:CS

[2860] Horatio N. Fraser. On cotton seeds. Scientific American, 26 (18):273–275, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
273.pdf.

Wioklin:1872:FG

[2861] E. S. Wioklin. Frictional gearing. Scientific American, 26(18):275, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-275.pdf.

Ott:1872:PC

[2862] Adolph Ott. Portland cement. Scientific American, 26(18):276, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-276.pdf.

Anonymous:1872:HHS

[2863] Anonymous. Human hair and its substitutes. Scientific American, 26(18):276, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
276a.pdf.

Anonymous:1872:BPs

[2864] Anonymous. Belting and pulleys. Scientific American, 26(18):277–278, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-277.pdf.

Anonymous:1872:DPa

[2865] Anonymous. Double plows. Scientific American, 26(18):279, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-279.pdf.

Anonymous:1872:UWR

[2866] Anonymous. The utilization of water in its relation to public health. Scientific American, 26(18):279, April 27, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n18/pdf/scientificamerican04271872-279a.pdf.

Anonymous:1872:AFb

[2867] Anonymous. The air filter. Scientific American, 26(18):279, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-279b.pdf.

Anonymous:1872:FCL

[2868] Anonymous. Folding chair, lounge, and bed. Scientific American, 26(18):280, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
280.pdf.

Anonymous:1872:PTR

[2869] Anonymous. Photographic toilet and hat rack. Scientific American, 26(18):280, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
280a.pdf.

Anonymous:1872:RWC

[2870] Anonymous. On the reasons which have caused the adoption of the new atomic weights. Scientific American, 26(18):280, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n18/pdf/scientificamerican04271872-280b.pdf.

Anonymous:1872:ACs

[2871] Anonymous. Answers to correspondents. Scientific American, 26 (18):281, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
281a.pdf.

Anonymous:1872:RAFr

[2872] Anonymous. Recent American and foreign patents. Scientific American, 26(18):282–283, April 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n18/pdf/scientificamerican04271872-
282.pdf.

Anonymous:1872:Cr

[2873] Anonymous. Correspondence. Scientific American, 26(19):292, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-292.pdf.

Anonymous:1872:BPt

[2874] Anonymous. Business and personal. Scientific American, 26(19):297, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-297.pdf.

Anonymous:1872:NQs

[2875] Anonymous. Notes and queries. Scientific American, 26(19):297–298, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-297a.pdf.

Anonymous:1872:BNBp

[2876] Anonymous. Books: New books and publications. Scientific American, 26(19):298, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
298a.pdf.

Anonymous:1872:IIr

[2877] Anonymous. Index of inventions. Scientific American, 26(19):300, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-300.pdf.

Anonymous:1872:IPD

[2878] Anonymous. Improved punching and drop press, the grain elevators at Liverpool, and more. Scientific American, 26(19):287–288, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-287.pdf.

Macgowan:1872:TTU

[2879] D. J. Macgowan. The tallow tree and its uses. Scientific American, 26(19):289, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
289.pdf.

Anonymous:1872:FGI

[2880] Anonymous. First Grand Industrial Exposition of Louisiana. Scientific American, 26(19):290, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
290.pdf.

Anonymous:1872:GFE

[2881] Anonymous. Gardner’s fire extinguisher. Scientific American, 26 (19):291, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
291.pdf.

Anonymous:1872:BC

[2882] Anonymous. Butter from California. Scientific American, 26(19):291, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-291a.pdf.

Anonymous:1872:PCSa

[2883] Anonymous. Patch’s corn sheller. Scientific American, 26(19):291, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-291b.pdf.

Anonymous:1872:SAMd

[2884] Anonymous. Society of arts of the Massachusetts institute of technology. Scientific American, 26(19):292–293, May 4, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n19/pdf/scientificamerican05041872-292a.pdf.

Anonymous:1872:SPIo

[2885] Anonymous. Scientific and practical information. Scientific American, 26(19):293–294, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
293.pdf.

Anonymous:1872:DSM

[2886] Anonymous. Drying substances by mechanical action. Scientific American, 26(19):295, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
295.pdf.

Anonymous:1872:CSMa

[2887] Anonymous. The Congressional sewing machine job. Scientific American, 26(19):295, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
295a.pdf.

Anonymous:1872:PAV

[2888] Anonymous. Progress of astronomy — the variable stars. Scientific American, 26(19):295–296, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
295b.pdf.

Anonymous:1872:AWM

[2889] Anonymous. Artesian wells and malarial fevers. Scientific American, 26(19):296, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
296.pdf.

Anonymous:1872:SIV

[2890] Anonymous. A suggestion to inventors. — value of patented improvements upon patented inventions. Scientific American, 26(19):296, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n19/pdf/scientificamerican05041872-296a.pdf.

Anonymous:1872:CTE

[2891] Anonymous. Cable towing on the Erie Canal. Scientific American, 26(19):296–297, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
296b.pdf.

Anonymous:1872:ACt

[2892] Anonymous. Answers to correspondents. Scientific American, 26 (19):298, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
298.pdf.

Anonymous:1872:RAFs

[2893] Anonymous. Recent American and foreign patents. Scientific American, 26(19):298–299, May 4, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n19/pdf/scientificamerican05041872-
298b.pdf.

Anonymous:1872:Ct

[2894] Anonymous. Correspondence. Scientific American, 26(20):308, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-308.pdf.

Anonymous:1872:BPu

[2895] Anonymous. Business and personal. Scientific American, 26 (20):312–313, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
312b.pdf.

Anonymous:1872:NQt

[2896] Anonymous. Notes and queries. Scientific American, 26(20):313, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-313.pdf.

Anonymous:1872:IIs

[2897] Anonymous. Index of inventions. Scientific American, 26(20):315, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-315.pdf.

Anonymous:1872:ISEa

[2898] Anonymous. Improved steam engine, Japanese textile art, and more. Scientific American, 26(20):303–304, May 11, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n20/pdf/scientificamerican05111872-303.pdf.

Anonymous:1872:BGM

[2899] Anonymous. Birmingham Gun Makers’ and Inventors’ Club. Scientific American, 26(20):304–306, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
304.pdf.

Anonymous:1872:PB

[2900] Anonymous. Pocket bootjack. Scientific American, 26(20):306, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-306.pdf.

Anonymous:1872:GS

[2901] Anonymous. Gardener’s stool. Scientific American, 26(20):306–307, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-306a.pdf.

Anonymous:1872:BM

[2902] Anonymous. Biscuit manufacture. Scientific American, 26(20):309–310, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-309.pdf.

Anonymous:1872:TDR

[2903] Anonymous. Twitchell’s draft regulator and heat economizer. Scientific American, 26(20):310, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
310.pdf.

Anonymous:1872:CAL

[2904] Anonymous. Compressed air locomotives. Scientific American, 26 (20):310, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
310a.pdf.

Anonymous:1872:OPL

[2905] Anonymous. Opening the public libraries on Sunday. Scientific American, 26(20):311, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
311.pdf.

Anonymous:1872:DCA

[2906] Anonymous. Drying by chemical action. Scientific American, 26 (20):311, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
311a.pdf.

Anonymous:1872:FBL

[2907] Anonymous. Fireproof building — the lesson of Chicago. Scientific American, 26(20):311–312, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
311b.pdf.

Anonymous:1872:PPSa

[2908] Anonymous. Power produced by steam, under different temperatures and pressures. Scientific American, 26(20):312, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-312.pdf.

Anonymous:1872:LFW

[2909] Anonymous. A long felt want. Scientific American, 26(20):312, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n20/pdf/scientificamerican05111872-312a.pdf.

Anonymous:1872:ACu

[2910] Anonymous. Answers to correspondents. Scientific American, 26 (20):313, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
313a.pdf.

Anonymous:1872:RAFt

[2911] Anonymous. Recent American and foreign patents. Scientific American, 26(20):313–315, May 11, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n20/pdf/scientificamerican05111872-
313b.pdf.

Anonymous:1872:BPv

[2912] Anonymous. Business and personal. Scientific American, 26(21):329, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-329.pdf.

Anonymous:1872:NQu

[2913] Anonymous. Notes and queries. Scientific American, 26(21):330, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-330.pdf.

Anonymous:1872:IIt

[2914] Anonymous. Index of inventions. Scientific American, 26(21):331, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-331.pdf.

Anonymous:1872:BNBq

[2915] Anonymous. Books: New books and publications. Scientific American, 26(21):332, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
332.pdf.

Anonymous:1872:MBPb

[2916] Anonymous. Machine for boarding pebbling and glossing leather, starting flower seeds. Scientific American, 26(21):319, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-319.pdf.

Anonymous:1872:SFS

[2917] Anonymous. Steam flying ship. Scientific American, 26(21):319–320, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-319a.pdf.

Anonymous:1872:PLR

[2918] Anonymous. Patent law reform in England. Scientific American, 26(21):320–322, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
320.pdf.

Anonymous:1872:SBB

[2919] Anonymous. Sweetzer’s blacking and brush holder. Scientific American, 26(21):322, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
322.pdf.

Anonymous:1872:WAB

[2920] Anonymous. Whitus’ attachment for bit braces. Scientific American, 26(21):322, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
322a.pdf.

Anonymous:1872:G

[2921] Anonymous. Guano. Scientific American, 26(21):323–324, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n21/pdf/scientificamerican05181872-323.pdf.

Anonymous:1872:SPIp

[2922] Anonymous. Scientific and practical information. Scientific American, 26(21):325–326, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
325.pdf.

Anonymous:1872:OOA

[2923] Anonymous. Our obligations to the Arabian alchemists. Scientific American, 26(21):327, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
327.pdf.

Anonymous:1872:PIMb

[2924] Anonymous. Puddling iron by machinery — an important improvement. Scientific American, 26(21):327, May 18, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n21/pdf/scientificamerican05181872-327a.pdf.

Anonymous:1872:RIN

[2925] Anonymous. Result of illustrating a new invention. Scientific American, 26(21):327–328, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
327b.pdf.

Anonymous:1872:APP

[2926] Anonymous. On atmospheric and pneumatic propulsion. Scientific American, 26(21):328, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
328.pdf.

Anonymous:1872:PBP

[2927] Anonymous. Proposed book on patents. Scientific American, 26 (21):328, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
328a.pdf.

Anonymous:1872:SWE

[2928] Anonymous. Supplement to this Week’s edition. Scientific American, 26(21):328–329, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
328b.pdf.

Anonymous:1872:ACv

[2929] Anonymous. Answers to correspondents. Scientific American, 26 (21):330, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
330a.pdf.

Anonymous:1872:RAFu

[2930] Anonymous. Recent American and foreign patents. Scientific American, 26(21):330–331, May 18, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n21/pdf/scientificamerican05181872-
330b.pdf.

Anonymous:1872:Cu

[2931] Anonymous. Correspondence. Scientific American, 26(22):348–350, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-348.pdf.

Anonymous:1872:BPw

[2932] Anonymous. Business and personal. Scientific American, 26(22):353, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-353b.pdf.

Anonymous:1872:NQv

[2933] Anonymous. Notes and queries. Scientific American, 26(22):354, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-354.pdf.

Anonymous:1872:IIu

[2934] Anonymous. Index of inventions. Scientific American, 26(22):356, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-356.pdf.

Anonymous:1872:BNBr

[2935] Anonymous. Books: New books and publications. Scientific American, 26(22):356, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
356a.pdf.

Anonymous:1872:PDM

[2936] Anonymous. Portable drilling machines, bone dust as a fertilizer, and more. Scientific American, 26(22):343–344, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-343.pdf.

Anonymous:1872:PICa

[2937] Anonymous. Patent infringement cases. Scientific American, 26 (22):344–345, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
344.pdf.

Anonymous:1872:NH

[2938] Anonymous. The nebular hypothesis. Scientific American, 26(22):345, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-345.pdf.

Anonymous:1872:CLL

[2939] Anonymous. City of London Library and Museum. Scientific American, 26(22):346–347, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
346.pdf.

Anonymous:1872:RP

[2940] Anonymous. Reservoir palette. Scientific American, 26(22):347, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-347.pdf.

Anonymous:1872:ECC

[2941] Anonymous. Electro chemical copying press. Scientific American, 26(22):347, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
347a.pdf.

Anonymous:1872:HCN

[2942] Anonymous. How to conduct new scientific investigations. Scientific American, 26(22):351, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
351.pdf.

Anonymous:1872:EB

[2943] Anonymous. Engineering in Brussels. Scientific American, 26(22):351, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-351a.pdf.

Anonymous:1872:CPB

[2944] Anonymous. The Colorado potato beetle. Scientific American, 26(22):351, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
351b.pdf.

Anonymous:1872:RTPa

[2945] Anonymous. The rubber tip patent case. Scientific American, 26(22):351, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
351c.pdf.

Anonymous:1872:NA

[2946] Anonymous. Nature and art. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-352.pdf.

Anonymous:1872:CPa

[2947] Anonymous. Canadian patents. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-352a.pdf.

Anonymous:1872:EII

[2948] Anonymous. East Indian iron. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-352b.pdf.

Anonymous:1872:SPIq

[2949] Anonymous. Scientific and practical information. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
352c.pdf.

Anonymous:1872:IPa

[2950] Anonymous. Improvement in pumps. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-352d.pdf.

Anonymous:1872:BH

[2951] Anonymous. Builders’ hardware. Scientific American, 26(22):352, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n22/pdf/scientificamerican05251872-352e.pdf.

Anonymous:1872:THP

[2952] Anonymous. Timely hints for Patent Office examiners. Scientific American, 26(22):353, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
353.pdf.

Anonymous:1872:DIN

[2953] Anonymous. Death of an Indian nobleman. Scientific American, 26(22):353, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
353a.pdf.

Anonymous:1872:ACw

[2954] Anonymous. Answers to correspondents. Scientific American, 26 (22):354, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
354a.pdf.

Anonymous:1872:RAFv

[2955] Anonymous. Recent American and foreign patents. Scientific American, 26(22):354–355, May 25, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n22/pdf/scientificamerican05251872-
354b.pdf.

Anonymous:1872:Cv

[2956] Anonymous. Correspondence. Scientific American, 26(23):364–365, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-364a.pdf.

Anonymous:1872:BNBs

[2957] Anonymous. Books: New books and publications. Scientific American, 26(23):369, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
369a.pdf.

Anonymous:1872:BPx

[2958] Anonymous. Business and personal. Scientific American, 26(23):369, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-369b.pdf.

Anonymous:1872:NQw

[2959] Anonymous. Notes and queries. Scientific American, 26(23):370, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-370.pdf.

Anonymous:1872:IIv

[2960] Anonymous. Index of inventions. Scientific American, 26(23):372, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-372.pdf.

Anonymous:1872:RRS

[2961] Anonymous. Railroad rail straightener, the effect of cold on iron. Scientific American, 26(23):359, June 1, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n23/pdf/scientificamerican06011872-359.pdf.

Anonymous:1872:BST

[2962] Anonymous. Baillairge’s stereometrical tableau. Scientific American, 26(23):359–360, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
359a.pdf.

Bramwell:1872:QSL

[2963] F. J. Bramwell. Quick steam launches. Scientific American, 26(23):360, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-360.pdf.

Anonymous:1872:B

[2964] Anonymous. Beetles. Scientific American, 26(23):361–362, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-361.pdf.

Anonymous:1872:HLa

[2965] Anonymous. Heat and light. Scientific American, 26(23):363–364, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-363.pdf.

Anonymous:1872:SEa

[2966] Anonymous. The steamship “Egypt”. Scientific American, 26(23):364, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-364.pdf.

Anonymous:1872:IEa

[2967] Anonymous. The international exhibition of 1872. Scientific American, 26(23):365–366, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
365.pdf.

Anonymous:1872:BPBa

[2968] Anonymous. Bell pull and Bell. Scientific American, 26(23):366, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-366.pdf.

Anonymous:1872:PFb

[2969] Anonymous. Projectile for firearms. Scientific American, 26(23):366, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-366a.pdf.

Anonymous:1872:FWA

[2970] Anonymous. A few words of advice and encouragement to investigators and inventors. Scientific American, 26(23):367, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-367.pdf.

Anonymous:1872:GMJ

[2971] Anonymous. The great musical jubilee. Scientific American, 26(23):367, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-367a.pdf.

Anonymous:1872:GRCb

[2972] Anonymous. A gigantic railway car. Scientific American, 26(23):367, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n23/pdf/scientificamerican06011872-367b.pdf.

Anonymous:1872:TFU

[2973] Anonymous. Tin foil — its uses and manufacture. Scientific American, 26(23):367, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
367c.pdf.

Anonymous:1872:IPRa

[2974] Anonymous. The industrial progress and requirements of New South Wales. Scientific American, 26(23):368, June 1, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n23/pdf/scientificamerican06011872-368.pdf.

Anonymous:1872:SPIr

[2975] Anonymous. Scientific and practical information. Scientific American, 26(23):368, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
368a.pdf.

Anonymous:1872:RME

[2976] Anonymous. The reception of the mining engineers at the Stevens institute. Scientific American, 26(23):368, June 1, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n23/pdf/scientificamerican06011872-368b.pdf.

Anonymous:1872:PICb

[2977] Anonymous. Patent infringement cases. Scientific American, 26 (23):369, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
369.pdf.

Anonymous:1872:ACx

[2978] Anonymous. Answers to correspondents. Scientific American, 26 (23):370, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
370a.pdf.

Anonymous:1872:RAFw

[2979] Anonymous. Recent American and foreign patents. Scientific American, 26(23):370–371, June 1, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n23/pdf/scientificamerican06011872-
370b.pdf.

Anonymous:1872:NQx

[2980] Anonymous. Notes and queries. Scientific American, 26(24):385–386, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-385.pdf.

Anonymous:1872:BPy

[2981] Anonymous. Business and personal. Scientific American, 26(24):386, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-386a.pdf.

Anonymous:1872:BNBt 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
388.pdf.

Anonymous:1872:STB

[2983] Anonymous. Suspension truss bridge, powdered coal for unhealthy plants, and more. Scientific American, 26(24):375, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n24/pdf/scientificamerican06081872-375.pdf.

Anonymous:1872:SLB

[2984] Anonymous. Self-lubricating box and shaft bearing. Scientific American, 26(24):375–376, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
375a.pdf.

Anonymous:1872:ABG

[2985] Anonymous. Alloys for bells, gongs, cymbals, etc. Scientific American, 26(24):377, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
377.pdf.

Anonymous:1872:HLb

[2986] Anonymous. Heat and light. Scientific American, 26(24):378–379, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-378.pdf.

Anonymous:1872:IPT

[2987] Anonymous. Improved puddling tool. Scientific American, 26 (24):379–381, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
379.pdf.

Anonymous:1872:MS

[2988] Anonymous. Mill step. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-382.pdf.

Anonymous:1872:RD

[2989] Anonymous. Riveting device. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-382a.pdf.

Anonymous:1872:SEW

[2990] Anonymous. Strengthening emery wheels. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
382b.pdf.

Anonymous:1872:PPPa

[2991] Anonymous. Peart’s patent pit car wheel. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
382c.pdf.

Anonymous:1872:DCc

[2992] Anonymous. Discharging coal. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-382d.pdf.

Anonymous:1872:SPC

[2993] Anonymous. Steam propulsion on canals. Scientific American, 26(24):382, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
382e.pdf.

Anonymous:1872:EHW

[2994] Anonymous. Eight hours’ work and ten hours’ pay. Scientific American, 26(24):383, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
383.pdf.

Anonymous:1872:IGC

[2995] Anonymous. Information on gum copal. Scientific American, 26 (24):383, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
383a.pdf.

Anonymous:1872:USS

[2996] Anonymous. The United States Senate on compensation to inventors. Scientific American, 26(24):383–384, June 8, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v26/
n24/pdf/scientificamerican06081872-383b.pdf.

Anonymous:1872:SMR

[2997] Anonymous. The sewing machine ring in a new role. Scientific American, 26(24):384, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
384.pdf.

Anonymous:1872:SEb

[2998] Anonymous. Screw elevators. Scientific American, 26(24):384–385, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n24/pdf/scientificamerican06081872-384a.pdf.

Anonymous:1872:ACy

[2999] Anonymous. Answers to correspondents. Scientific American, 26 (24):386, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
386.pdf.

Anonymous:1872:RAFx

[3000] Anonymous. Recent American and foreign patents. Scientific American, 26(24):387, June 8, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n24/pdf/scientificamerican06081872-
387.pdf.

Anonymous:1872:Cw

[3001] Anonymous. Correspondence. Scientific American, 26(25):396, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-396.pdf.

Anonymous:1872:BPz

[3002] Anonymous. Business and personal. Scientific American, 26(25):401, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-401a.pdf.

Anonymous:1872:NQy

[3003] Anonymous. Notes and queries. Scientific American, 26(25):401, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-401b.pdf.

Anonymous:1872:IIw

[3004] Anonymous. Index of inventions. Scientific American, 26(25):403–404, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-403.pdf.

Anonymous:1872:ICG

[3005] Anonymous. Improved coffee and grain dryer, solvent for coralline. Scientific American, 26(25):391, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
391.pdf.

Anonymous:1872:GPH

[3006] Anonymous. Guardiola’s pestle for hulling coffee, rice, etc. Scientific American, 26(25):391–392, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
391a.pdf.

Anonymous:1872:APa

[3007] Anonymous. From the Atlantic to the Pacific. Scientific American, 26(25):392–393, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
392.pdf.

Anonymous:1872:GPG

[3008] Anonymous. Gunpowder pressure gages. Scientific American, 26 (25):394, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
394.pdf.

Anonymous:1872:IWVb

[3009] Anonymous. On ice, water, vapor, and air. Scientific American, 26(25):394–395, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
394a.pdf.

Anonymous:1872:NDR

[3010] Anonymous. New device for railway signals. Scientific American, 26(25):395, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
395.pdf.

Anonymous:1872:VBG

[3011] Anonymous. Vertical balance galvanometer. Scientific American, 26(25):395–396, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
395a.pdf.

Anonymous:1872:CCCa

[3012] Anonymous. Corks and cork cutting. Scientific American, 26(25):397, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-397.pdf.

Anonymous:1872:NRB

[3013] Anonymous. The new railroad bridge at Albany. Scientific American, 26(25):397–398, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
397a.pdf.

Anonymous:1872:BMC

[3014] Anonymous. Bunker’s metal cutter. Scientific American, 26(25):398, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-398.pdf.

Anonymous:1872:DB

[3015] Anonymous. Durand’s burette. Scientific American, 26(25):398, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-398a.pdf.

Anonymous:1872:DR

[3016] Anonymous. Dynamic refrigerator. Scientific American, 26(25):398, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-398b.pdf.

Anonymous:1872:MTO

[3017] Anonymous. Meteorites, their origin. Scientific American, 26(25):399, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-399.pdf.

Anonymous:1872:EHS

[3018] Anonymous. The eight hour strike in New York. Scientific American, 26(25):399, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
399a.pdf.

Anonymous:1872:PRB

[3019] Anonymous. The pneumatic railway brake and other applications of the pneumatic system. Scientific American, 26(25):399, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n25/pdf/scientificamerican06151872-399b.pdf.

Anonymous:1872:ISBa

[3020] Anonymous. Iron ship building in Wilmington. Scientific American, 26(25):399–400, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
399c.pdf.

Anonymous:1872:HPCa

[3021] Anonymous. How paper collars are made. Scientific American, 26(25):400, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
400.pdf.

Anonymous:1872:AIEa

[3022] Anonymous. American inventions in Europe. Scientific American, 26(25):400, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
400a.pdf.

Anonymous:1872:SBN

[3023] Anonymous. The sign boards of New York. Scientific American, 26(25):400, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
400b.pdf.

Anonymous:1872:DJG

[3024] Anonymous. Death of James Gordon Bennett. Scientific American, 26(25):400, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
400c.pdf.

Anonymous:1872:SPIs

[3025] Anonymous. Scientific and practical information. Scientific American, 26(25):401, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
401.pdf.

Anonymous:1872:ACz

[3026] Anonymous. Answers to correspondents. Scientific American, 26 (25):401–402, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
401c.pdf.

Anonymous:1872:RAFy

[3027] Anonymous. Recent American and foreign patents. Scientific American, 26(25):402, June 15, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n25/pdf/scientificamerican06151872-
402.pdf.

Anonymous:1872:Cx

[3028] Anonymous. Correspondence. Scientific American, 26(26):412–413, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-412.pdf.

Anonymous:1872:NQz

[3029] Anonymous. Notes and queries. Scientific American, 26(26):416, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-416.pdf.

Anonymous:1872:BPbb

[3030] Anonymous. Business and personal. Scientific American, 26(26):416, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-416a.pdf.

Anonymous:1872:IIx

[3031] Anonymous. Index of inventions. Scientific American, 26(26):417, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-417.pdf.

Anonymous:1872:PNPb

[3032] Anonymous. Patent notes: Patents. Scientific American, 26(26):417, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-417a.pdf.

Anonymous:1872:Ia

[3033] Anonymous. Index. Scientific American, 26(26):418–420, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-418.pdf.

Anonymous:1872:ASP

[3034] Anonymous. Air Spring printing press. Scientific American, 26(26):407, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-407.pdf.

Anonymous:1872:NMSa

[3035] Anonymous. Novel mode of sawing wood without a saw. Scientific American, 26(26):407–408, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
407a.pdf.

Anonymous:1872:KUS

[3036] Anonymous. The kaolin of the United States. Scientific American, 26(26):408, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
408.pdf.

Anonymous:1872:PFc

[3037] Anonymous. Portable frescoes. Scientific American, 26(26):409, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-409.pdf.

Anonymous:1872:CEa

[3038] Anonymous. Cincinnati Exposition. Scientific American, 26(26):409, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-409a.pdf.

Anonymous:1872:UME

[3039] Anonymous. The unit measure of electrical resistance. Scientific American, 26(26):409, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
409b.pdf.

Anonymous:1872:NME

[3040] Anonymous. New magneto-electric machine. Scientific American, 26(26):410, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
410.pdf.

Anonymous:1872:SKL

[3041] Anonymous. Safety kerosene lamp. Scientific American, 26(26):410, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-410a.pdf.

Anonymous:1872:PLc

[3042] Anonymous. Pencil leads. Scientific American, 26(26):410–411, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-410b.pdf.

Gladstone:1872:CSG

[3043] John Hall Gladstone, F. R. S., and F. C. S. Crystallization of silver, gold, and other metals. Scientific American, 26(26):411–412, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v26/n26/pdf/scientificamerican06221872-411.pdf.

Anonymous:1872:NSC

[3044] Anonymous. The new State Capitol at Albany, N.Y. Scientific American, 26(26):413–414, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
413.pdf.

Anonymous:1872:NPLa

[3045] Anonymous. A new patent law in Canada. Scientific American, 26(26):415, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
415.pdf.

Anonymous:1872:PEHa

[3046] Anonymous. Progress of the eight hour strike. Scientific American, 26(26):415, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
415a.pdf.

Anonymous:1872:WPFa

[3047] Anonymous. Weight, pressure, force, power, work. Scientific American, 26(26):415, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
415b.pdf.

Anonymous:1872:RGP

[3048] Anonymous. Rubber graphite paint. Scientific American, 26 (26):415–416, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
415c.pdf.

Anonymous:1872:RAFz

[3049] Anonymous. Recent American and foreign patents. Scientific American, 26(26):416, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
416b.pdf.

Anonymous:1872:NPLb

[3050] Anonymous. New patent law in Canada. Scientific American, 26(26):416, June 22, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v26/n26/pdf/scientificamerican06221872-
416c.pdf.

Anonymous:1872:Cy

[3051] Anonymous. Correspondence. Scientific American, 27(1):4, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-4.pdf.

Anonymous:1872:BPbc

[3052] Anonymous. Business and personal. Scientific American, 27(1):9, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-9a.pdf.

Anonymous:1872:NQba

[3053] Anonymous. Notes and queries. Scientific American, 27(1):10, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-10a.pdf.

Anonymous:1872:ISH

[3054] Anonymous. Improved steam heater. Scientific American, 27(1):1, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-1.pdf.

Anonymous:1872:SWC

[3055] Anonymous. Shive’s Watchman’s clock. Scientific American, 27(1):1, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-1a.pdf.

Anonymous:1872:LBE

[3056] Anonymous. Locomotive boiler explosion. Scientific American, 27(1): 2, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-2.pdf.

Anonymous:1872:DS

[3057] Anonymous. The ducking stool. Scientific American, 27(1):2–3, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-2a.pdf.

Anonymous:1872:MMI

[3058] Anonymous. On the measurement of musical intervals. Scientific American, 27(1):3, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n1/pdf/scientificamerican07061872-
3.pdf.

Anonymous:1872:HH

[3059] Anonymous. Horse hair. Scientific American, 27(1):4, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-4a.pdf.

Anonymous:1872:NOL

[3060] Anonymous. Nitrous oxide or laughing gas. Scientific American, 27(1): 5, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-5.pdf.

Anonymous:1872:BFW

[3061] Anonymous. Berryman’s feed water heater. Scientific American, 27(1): 6, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-6.pdf.

Anonymous:1872:IPb

[3062] Anonymous. Ice preserver. Scientific American, 27(1):6, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-6a.pdf.

Anonymous:1872:WPFb

[3063] Anonymous. Weight, pressure, force, power, energy, work. Scientific American, 27(1):7, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n1/pdf/scientificamerican07061872-
7.pdf.

Anonymous:1872:PEHb

[3064] Anonymous. Progress of the eight hour strike. Scientific American, 27(1):7–8, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n1/pdf/scientificamerican07061872-
7a.pdf.

Anonymous:1872:PR

[3065] Anonymous. Postal reform. Scientific American, 27(1):8, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-8.pdf.

Anonymous:1872:TTR

[3066] Anonymous. The tariff and tax reductions. Scientific American, 27(1): 8, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-8a.pdf.

Anonymous:1872:RTN

[3067] Anonymous. Rapid transit in New York. Scientific American, 27(1):8, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-8b.pdf.

Anonymous:1872:MT

[3068] Anonymous. A machine that talks. Scientific American, 27(1):8–9, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-8c.pdf.

Anonymous:1872:SPIt

[3069] Anonymous. Scientific and practical information. Scientific American, 27(1):9, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n1/pdf/scientificamerican07061872-
9.pdf.

Anonymous:1872:GWb

[3070] Anonymous. Grant and Wilson. Scientific American, 27(1):10, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-10.pdf.

Anonymous:1872:ACba

[3071] Anonymous. Answers to correspondents. Scientific American, 27(1):10, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-10b.pdf.

Anonymous:1872:NPLc

[3072] Anonymous. New patent law in Canada. Scientific American, 27(1):10, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n1/pdf/scientificamerican07061872-10c.pdf.

Anonymous:1872:RAFba

[3073] Anonymous. Recent American and foreign patents. Scientific American, 27(1):10–11, July 6, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n1/pdf/scientificamerican07061872-
10d.pdf.

Anonymous:1872:Cz

[3074] Anonymous. Correspondence. Scientific American, 27(2):20, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-20.pdf.

Anonymous:1872:BPbd

[3075] Anonymous. Business and personal. Scientific American, 27(2):25, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-25a.pdf.

Anonymous:1872:NQbb

[3076] Anonymous. Notes and queries. Scientific American, 27(2):26, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-26.pdf.

Anonymous:1872:IIy

[3077] Anonymous. Index of inventions. Scientific American, 27(2):28, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-28.pdf.

Anonymous:1872:BNBu

[3078] Anonymous. Books: New books and publications. Scientific American, 27(2):28, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
28a.pdf.

Anonymous:1872:ICB

[3079] Anonymous. Improved canal boat. Scientific American, 27(2):15–16, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-15.pdf.

Anonymous:1872:PMI

[3080] Anonymous. Papier mache for interior decoration. Scientific American, 27(2):16, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
16.pdf.

Anonymous:1872:CTM

[3081] Anonymous. Crackers and their manufacture. Scientific American, 27(2):16, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
16a.pdf.

Anonymous:1872:PFCa

[3082] Anonymous. The progress of fish culture in New York. Scientific American, 27(2):17, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
17.pdf.

Anonymous:1872:MIb

[3083] Anonymous. The mosquito illustrated. Scientific American, 27 (2):18–20, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
18.pdf.

Anonymous:1872:SAP

[3084] Anonymous. Something about pipes. Scientific American, 27(2):20–21, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-20a.pdf.

Anonymous:1872:EBT

[3085] Anonymous. The English building trade lock-out. Scientific American, 27(2):21, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
21.pdf.

Anonymous:1872:IEb

[3086] Anonymous. Improved ellipsograph. Scientific American, 27(2):22, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-22.pdf.

Anonymous:1872:CHPa

[3087] Anonymous. Corn husker and picker. Scientific American, 27(2):22, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-22a.pdf.

Anonymous:1872:CFW

[3088] Anonymous. Combined flask and whisk. Scientific American, 27(2):22, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-22b.pdf.

Anonymous:1872:PCSb

[3089] Anonymous. A plea for common sense. Scientific American, 27(2):23, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-23.pdf.

Anonymous:1872:LPT

[3090] Anonymous. Letters from Professor Thurston. Scientific American, 27(2):23, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
23a.pdf.

Anonymous:1872:PEHc

[3091] Anonymous. Progress of the eight hour strike in New York. Scientific American, 27(2):23–24, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
23b.pdf.

Anonymous:1872:ERB

[3092] Anonymous. The East River Bridge. Scientific American, 27(2):24, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-24.pdf.

Anonymous:1872:LPRa

[3093] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(2):24, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
24a.pdf.

Anonymous:1872:DPF

[3094] Anonymous. Design patents to foreigners. Scientific American, 27(2):24–25, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
24b.pdf.

Anonymous:1872:NIEa

[3095] Anonymous. The Newark Industrial Exhibition. Scientific American, 27(2):25, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
25.pdf.

Anonymous:1872:NPLd

[3096] Anonymous. New patent law in Canada. Scientific American, 27(2):25, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-25b.pdf.

Anonymous:1872:ACbb

[3097] Anonymous. Answer to correspondents. Scientific American, 27(2):26, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n2/pdf/scientificamerican07131872-26a.pdf.

Anonymous:1872:RAFbb

[3098] Anonymous. Recent American and foreign patents. Scientific American, 27(2):26–27, July 13, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n2/pdf/scientificamerican07131872-
26b.pdf.

Anonymous:1872:Cba

[3099] Anonymous. Correspondence. Scientific American, 27(3):36, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-36.pdf.

Anonymous:1872:BPbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-41.pdf.

Anonymous:1872:NQbc

[3101] Anonymous. Notes and queries. Scientific American, 27(3):41, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-41a.pdf.

Anonymous:1872:IIz

[3102] Anonymous. Index of inventions. Scientific American, 27(3):43–44, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-43.pdf.

Anonymous:1872:MBW

[3103] Anonymous. Machine for bunching, wiring, and inserting bristles in brush backs. Scientific American, 27(3):31–32, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-31.pdf.

Anonymous:1872:Sf

[3104] Anonymous. Smell. Scientific American, 27(3):32–33, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-32.pdf.

Anonymous:1872:PAH

[3105] Anonymous. Phenomena associated with a hydrogen flame. Scientific American, 27(3):33, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
33.pdf.

Anonymous:1872:Cbb

[3106] Anonymous. Cincinnati. Scientific American, 27(3):34, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-34.pdf.

Anonymous:1872:CWa

[3107] Anonymous. Carriage wheel. Scientific American, 27(3):35, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-35.pdf.

Anonymous:1872:BME

[3108] Anonymous. Brick and mortar elevator and distributor. Scientific American, 27(3):35, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
35a.pdf.

Anonymous:1872:GGD

[3109] Anonymous. The Gatling gun does history repeat itself? Scientific American, 27(3):36, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
36a.pdf.

Anonymous:1872:PFd

[3110] Anonymous. Pressure on foundations. Scientific American, 27(3):37, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-37.pdf.

Anonymous:1872:RCB

[3111] Anonymous. Railroad car brake. Scientific American, 27(3):38, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-38.pdf.

Anonymous:1872:RR

[3112] Anonymous. Railroad rail. Scientific American, 27(3):38, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-38a.pdf.

Anonymous:1872:DT

[3113] Anonymous. The domestic telegraph. Scientific American, 27(3):39, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-39.pdf.

Anonymous:1872:SHV

[3114] Anonymous. Summer heats in various countries. Scientific American, 27(3):39, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
39a.pdf.

Anonymous:1872:NPLe

[3115] Anonymous. The new patent law of Canada. Scientific American, 27(3):39–40, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
39b.pdf.

Anonymous:1872:LPRb

[3116] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(3):40–41, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
40.pdf.

Anonymous:1872:ACbc

[3117] Anonymous. Answers to correspondents. Scientific American, 27(3):41, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n3/pdf/scientificamerican07201872-41b.pdf.

Anonymous:1872:RAFbc

[3118] Anonymous. Recent American and foreign patents. Scientific American, 27(3):41–42, July 20, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n3/pdf/scientificamerican07201872-
41c.pdf.

Anonymous:1872:Cbc

[3119] Anonymous. Correspondence. Scientific American, 27(4):52, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-52.pdf.

Anonymous:1872:BPbf

[3120] Anonymous. Business and personal. Scientific American, 27(4):57, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-57a.pdf.

Anonymous:1872:NQbd

[3121] Anonymous. Notes and queries. Scientific American, 27(4):58, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-58.pdf.

Anonymous:1872:BNBv

[3122] Anonymous. Books: New books and publications. Scientific American, 27(4):59, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
59.pdf.

Anonymous:1872:IIba

[3123] Anonymous. Index of inventions. Scientific American, 27(4):60, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-60.pdf.

Anonymous:1872:PAS

[3124] Anonymous. The Prall automatic steam vacuum pump. Scientific American, 27(4):47–48, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
47.pdf.

Anonymous:1872:HN

[3125] Anonymous. The hands of the nation. Scientific American, 27(4):48, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-48.pdf.

Anonymous:1872:LRb

[3126] Anonymous. Lightning rods. Scientific American, 27(4):49, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-49.pdf.

Anonymous:1872:NSD

[3127] Anonymous. New steam dredging boats. Scientific American, 27(4):50–51, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
50.pdf.

Anonymous:1872:IBa

[3128] Anonymous. Improved bung. Scientific American, 27(4):51, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-51.pdf.

Anonymous:1872:BBS

[3129] Anonymous. A bridle bit that is something more, and an improved stocking for horses. Scientific American, 27(4):51, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-51a.pdf.

Anonymous:1872:ACbd

[3130] Anonymous. Armored can. Scientific American, 27(4):51, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-51b.pdf.

Anonymous:1872:STS

[3131] Anonymous. Saving of time — a serious charge against the Scientific American. Scientific American, 27(4):52–53, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n4/pdf/scientificamerican07271872-52a.pdf.

Anonymous:1872:SPIu

[3132] Anonymous. Scientific and practical information. Scientific American, 27(4):53, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
53.pdf.

Anonymous:1872:SM

[3133] Anonymous. Sewing machine. Scientific American, 27(4):54, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-54.pdf.

Anonymous:1872:HC

[3134] Anonymous. Hose coupling. Scientific American, 27(4):54, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-54a.pdf.

Anonymous:1872:TO

[3135] Anonymous. The two oceans. Scientific American, 27(4):55, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-55.pdf.

Anonymous:1872:NCR

[3136] Anonymous. Narrow and cheap railways. Scientific American, 27(4):55, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-55a.pdf.

Anonymous:1872:OSL

[3137] Anonymous. Oxyhydrogen street lamps. Scientific American, 27 (4):55–56, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
55b.pdf.

Anonymous:1872:EEH

[3138] Anonymous. End of the eight hour strike. Scientific American, 27(4):56, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-56.pdf.

Anonymous:1872:PHTa

[3139] Anonymous. Paper hangings and their manufacture. Scientific American, 27(4):56, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
56a.pdf.

Anonymous:1872:SEG

[3140] Anonymous. Solid emery grinder. Scientific American, 27(4):56–57, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-56b.pdf.

Anonymous:1872:NC

[3141] Anonymous. The nature of comets. Scientific American, 27(4):57, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-57.pdf.

Anonymous:1872:ACbe

[3142] Anonymous. Answers to correspondents. Scientific American, 27(4):58, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-58a.pdf.

Anonymous:1872:RAFbd

[3143] Anonymous. Recent American and foreign patents. Scientific American, 27(4):58–59, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n4/pdf/scientificamerican07271872-
58b.pdf.

Anonymous:1872:NPLf

[3144] Anonymous. New patent law in Canada. Scientific American, 27(4):60, July 27, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n4/pdf/scientificamerican07271872-60a.pdf.

Anonymous:1872:Cbd

[3145] Anonymous. Correspondence. Scientific American, 27(5):68, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-68.pdf.

Anonymous:1872:BPbg

[3146] Anonymous. Business and personal. Scientific American, 27(5):73, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-73a.pdf.

Anonymous:1872:NQbe

[3147] Anonymous. Notes and queries. Scientific American, 27(5):73–74, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-73b.pdf.

Anonymous:1872:IIbb

[3148] Anonymous. Index of inventions. Scientific American, 27(5):76, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-76.pdf.

Anonymous:1872:BNBw

[3149] Anonymous. Books: New books and publications. Scientific American, 27(5):76, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
76b.pdf.

Anonymous:1872:IFC

[3150] Anonymous. Improved form of coal railroad. Scientific American, 27(5):63, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
63.pdf.

Anonymous:1872:DIC

[3151] Anonymous. Doane’s improved canal construction. Scientific American, 27(5):63, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
63a.pdf.

Anonymous:1872:DM

[3152] Anonymous. Domestic motors. Scientific American, 27(5):64–65, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-64.pdf.

Anonymous:1872:CGH

[3153] Anonymous. Combined gin house, horse power, and cotton press. Scientific American, 27(5):66, August 3, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n5/pdf/scientificamerican08031872-66.pdf.

Riley:1872:NHC

[3154] C. V. Riley. The natural history of the curculio. Scientific American, 27(5):67, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
67.pdf.

Anonymous:1872:REc

[3155] Anonymous. Recipes and experiments. Scientific American, 27 (5):68–69, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
68a.pdf.

Anonymous:1872:MCT

[3156] Anonymous. Mastersun’s Canadian turbine. Scientific American, 27(5):70, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
70.pdf.

Anonymous:1872:RWF

[3157] Anonymous. Reversible water filter. Scientific American, 27(5):70, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-70a.pdf.

Anonymous:1872:CCa

[3158] Anonymous. Changes of climate. Scientific American, 27(5):71, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-71.pdf.

Anonymous:1872:CPc

[3159] Anonymous. The Chironectes Pictus. Scientific American, 27(5):71, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-71a.pdf.

Anonymous:1872:DPb

[3160] Anonymous. Drawings of the patents. Scientific American, 27(5):71, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-71b.pdf.

Anonymous:1872:DCPa

[3161] Anonymous. A decision by the Commissioner of Patents. Scientific American, 27(5):71, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
71c.pdf.

Anonymous:1872:AV

[3162] Anonymous. Aids to ventilation. Scientific American, 27(5):71, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-71d.pdf.

Anonymous:1872:TAT

[3163] Anonymous. Testing armored turrets — a remarkable experiment. Scientific American, 27(5):71–72, August 3, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n5/pdf/scientificamerican08031872-71e.pdf.

Anonymous:1872:DOF

[3164] Anonymous. The decrease of our fish supply. Scientific American, 27(5):72, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
72.pdf.

Anonymous:1872:LPRc

[3165] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(5):72, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
72a.pdf.

Anonymous:1872:CRa

[3166] Anonymous. Cheap railways. Scientific American, 27(5):72, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-72b.pdf.

Anonymous:1872:SPIv

[3167] Anonymous. Scientific and practical information. Scientific American, 27(5):73, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
73.pdf.

Anonymous:1872:ACbf

[3168] Anonymous. Answers to correspondents. Scientific American, 27(5):74, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-74.pdf.

Anonymous:1872:RAFbe

[3169] Anonymous. Recent American and foreign patents. Scientific American, 27(5):74–75, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n5/pdf/scientificamerican08031872-
74a.pdf.

Anonymous:1872:NPLg

[3170] Anonymous. New patent law in Canada. Scientific American, 27(5):76, August 3, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n5/pdf/scientificamerican08031872-76a.pdf.

Anonymous:1872:Cbe

[3171] Anonymous. Correspondence. Scientific American, 27(6):84–85, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-84.pdf.

Anonymous:1872:BPbh

[3172] Anonymous. Business and personal. Scientific American, 27(6): 89–90, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
89a.pdf.

Anonymous:1872:NQbf

[3173] Anonymous. Notes and queries. Scientific American, 27(6):90, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-90.pdf.

Anonymous:1872:IIbc

[3174] Anonymous. Index of inventions. Scientific American, 27(6):92, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-92.pdf.

Anonymous:1872:BNBx

[3175] Anonymous. Books: New books and publications. Scientific American, 27(6):92, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
92b.pdf.

Anonymous:1872:TWa

[3176] Anonymous. The type writer. Scientific American, 27(6):79, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-79.pdf.

Anonymous:1872:GAS

[3177] Anonymous. Glacial action in South America. Scientific American, 27(6):80, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
80.pdf.

Anonymous:1872:AAS

[3178] Anonymous. The annual August shower of meteors. Scientific American, 27(6):81, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
81.pdf.

Anonymous:1872:ISCa

[3179] Anonymous. Improved seed cleaner for thrashers. Scientific American, 27(6):82, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
82.pdf.

Anonymous:1872:MBa

[3180] Anonymous. A musical barometer. Scientific American, 27(6):82, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-82a.pdf.

Anonymous:1872:PSH

[3181] Anonymous. Postage stamp holder. Scientific American, 27(6):82, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-82b.pdf.

Anonymous:1872:POG

[3182] Anonymous. The Patent Office Gazette. Scientific American, 27 (6):82, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
82c.pdf.

Anonymous:1872:AFc

[3183] Anonymous. Automatic fan. Scientific American, 27(6):82, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-82d.pdf.

Anonymous:1872:BMG

[3184] Anonymous. The British monitor “Glatton”. Scientific American, 27(6):83, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
83.pdf.

Anonymous:1872:ISI

[3185] Anonymous. Improved sheet iron roofing. Scientific American, 27(6):86, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
86.pdf.

Anonymous:1872:RPB

[3186] Anonymous. Rotary pressure blower. Scientific American, 27(6):86, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-86a.pdf.

Anonymous:1872:CCCb

[3187] Anonymous. Cosmical causes of changes of climate. Scientific American, 27(6):87, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87.pdf.

Anonymous:1872:NMM

[3188] Anonymous. The new Manhattan market. Scientific American, 27(6):87, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87a.pdf.

Anonymous:1872:SCb

[3189] Anonymous. Spontaneous combustion. Scientific American, 27 (6):87, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87b.pdf.

Anonymous:1872:CBT

[3190] Anonymous. Canal boat towing by road steamers. Scientific American, 27(6):87, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87c.pdf.

Anonymous:1872:LM

[3191] Anonymous. Look out for the meteors. Scientific American, 27 (6):87, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87d.pdf.

Anonymous:1872:BL

[3192] Anonymous. Because I am so lazy. Scientific American, 27 (6):87–88, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
87e.pdf.

Anonymous:1872:SN

[3193] Anonymous. Steamship notes. Scientific American, 27(6):88, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-88.pdf.

Anonymous:1872:BD

[3194] Anonymous. Bright’s disease. Scientific American, 27(6):88, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-88a.pdf.

Anonymous:1872:BRB

[3195] Anonymous. The Black Rock Bridge over the Niagara River. Scientific American, 27(6):88, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
88b.pdf.

Anonymous:1872:Aa

[3196] Anonymous. The Amanians. Scientific American, 27(6):88, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n6/pdf/scientificamerican08101872-88c.pdf.

Anonymous:1872:LPRd

[3197] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(6):88–89, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
88d.pdf.

Anonymous:1872:SPIw

[3198] Anonymous. Scientific and practical information. Scientific American, 27(6):89, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
89.pdf.

Anonymous:1872:ACbg

[3199] Anonymous. Answer to correspondents. Scientific American, 27 (6):90, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
90a.pdf.

Anonymous:1872:RAFbf

[3200] Anonymous. Recent American and foreign patents. Scientific American, 27(6):90–91, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
90b.pdf.

Anonymous:1872:NPLh

[3201] Anonymous. New patent law in Canada. Scientific American, 27(6):92, August 10, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n6/pdf/scientificamerican08101872-
92a.pdf.

Anonymous:1872:Cbf

[3202] Anonymous. Correspondence. Scientific American, 27(7):100, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-100.pdf.

Anonymous:1872:BPbi

[3203] Anonymous. Business and personal. Scientific American, 27(7): 105–106, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
105b.pdf.

Anonymous:1872:NQbg

[3204] Anonymous. Notes and queries. Scientific American, 27(7):106, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-106.pdf.

Anonymous:1872:BNBy

[3205] Anonymous. Books: New books and publications. Scientific American, 27(7):107, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
107.pdf.

Anonymous:1872:IIbd

[3206] Anonymous. Index of inventions. Scientific American, 27(7):108, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-108.pdf.

Anonymous:1872:EIM

[3207] Anonymous. Ether ice machine. Scientific American, 27(7):95, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-95.pdf.

Anonymous:1872:HKO

[3208] Anonymous. How kerosene oil is made. Scientific American, 27 (7):96, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
96.pdf.

Anonymous:1872:SH

[3209] Anonymous. Strychnia and hemlock. Scientific American, 27(7):97, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-97.pdf.

Anonymous:1872:Sg

[3210] Anonymous. Silkworms. Scientific American, 27(7):98, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-98.pdf.

Anonymous:1872:MBb

[3211] Anonymous. The manora breakwater. Scientific American, 27 (7):99, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
99.pdf.

Anonymous:1872:NES

[3212] Anonymous. Notes of English science. Scientific American, 27 (7):100–101, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
100a.pdf.

Anonymous:1872:IRSa

[3213] Anonymous. Improved railway system. Scientific American, 27 (7):102, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
102.pdf.

Anonymous:1872:IRR

[3214] Anonymous. Improved repeating rifle. Scientific American, 27 (7):102, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
102a.pdf.

Anonymous:1872:AK

[3215] Anonymous. The appreciation of knowledge. Scientific American, 27(7):103, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
103.pdf.

Anonymous:1872:PLd

[3216] Anonymous. Protection from lightning. Scientific American, 27 (7):103, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
103a.pdf.

Anonymous:1872:PFf

[3217] Anonymous. A petroleum fire. Scientific American, 27(7):103, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-103b.pdf.

Anonymous:1872:RAE

[3218] Anonymous. A railway accident explained. Scientific American, 27(7):103, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
103c.pdf.

Anonymous:1872:IPCb

[3219] Anonymous. The international prison Congress. Scientific American, 27(7):103–104, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
103d.pdf.

Anonymous:1872:NBb

[3220] Anonymous. Newspaper benevolence. Scientific American, 27 (7):104, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
104.pdf.

Anonymous:1872:WPP

[3221] Anonymous. Wave power propellers. Scientific American, 27 (7):104, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
104a.pdf.

Anonymous:1872:WVV

[3222] Anonymous. Water vapor not vesicular. Scientific American, 27(7):104, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
104b.pdf.

Anonymous:1872:IRSb

[3223] Anonymous. Inclined railway in San Francisco. Scientific American, 27(7):104, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
104c.pdf.

Anonymous:1872:LPRe

[3224] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(7):104–105, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
104d.pdf.

Anonymous:1872:SAc

[3225] Anonymous. Slicing apples. Scientific American, 27(7):105, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n7/pdf/scientificamerican08171872-105.pdf.

Anonymous:1872:REd

[3226] Anonymous. Recipes and experiments. Scientific American, 27 (7):105, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
105a.pdf.

Anonymous:1872:ACbh

[3227] Anonymous. Answers to correspondents. Scientific American, 27(7):106, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
106a.pdf.

Anonymous:1872:RAFbg

[3228] Anonymous. Recent American and foreign patents. Scientific American, 27(7):106–107, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
106b.pdf.

Anonymous:1872:NPLi

[3229] Anonymous. New patent law in Canada. Scientific American, 27(7):107, August 17, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n7/pdf/scientificamerican08171872-
107a.pdf.

Anonymous:1872:Cbg

[3230] Anonymous. Correspondence. Scientific American, 27(8):116–117, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-116a.pdf.

Anonymous:1872:BPbj

[3231] Anonymous. Business and personal. Scientific American, 27(8):122, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-122.pdf.

Anonymous:1872:NQbh

[3232] Anonymous. Notes and queries. Scientific American, 27(8):122, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-122a.pdf.

Anonymous:1872:IIbe

[3233] Anonymous. Index of inventions. Scientific American, 27(8):124, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-124.pdf.

Anonymous:1872:BNBz

[3234] Anonymous. Books: New books and publications. Scientific American, 27(8):124, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
124a.pdf.

Anonymous:1872:ISEb

[3235] Anonymous. Improved sugar evaporating apparatus. Scientific American, 27(8):111–112, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
111.pdf.

Anonymous:1872:HFH

[3236] Anonymous. How felt hats are made. Scientific American, 27 (8):113, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
113.pdf.

Anonymous:1872:ACbi

[3237] Anonymous. Adulteration of chemicals. Scientific American, 27 (8):113, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
113a.pdf.

Anonymous:1872:DAS

[3238] Anonymous. Direct-acting steam engine. Scientific American, 27(8):114, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
114.pdf.

Anonymous:1872:BBa

[3239] Anonymous. Blacksmith’s butteris. Scientific American, 27(8):114, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-114a.pdf.

Anonymous:1872:NWF

[3240] Anonymous. The new water front, New York City. Scientific American, 27(8):116, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
116.pdf.

Anonymous:1872:FLb

[3241] Anonymous. Fireless locomotive. Scientific American, 27(8):118, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-118.pdf.

Anonymous:1872:SCM

[3242] Anonymous. Sash coupling machine. Scientific American, 27 (8):118, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
118a.pdf.

Anonymous:1872:AAAa

[3243] Anonymous. The American Association for the Advancement of Science. Scientific American, 27(8):119, August 24, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n8/pdf/scientificamerican08241872-119.pdf.

Anonymous:1872:CPd

[3244] Anonymous. Canadian patents. Scientific American, 27(8):119, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-119a.pdf.

Anonymous:1872:Ab

[3245] Anonymous. The Aurora. Scientific American, 27(8):119–120, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-119b.pdf.

Anonymous:1872:CCb

[3246] Anonymous. Country churchyards. Scientific American, 27(8):120, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-120.pdf.

Anonymous:1872:IVC

[3247] Anonymous. Influence of variously colored light animal and vegetable growth. Scientific American, 27(8):120, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n8/pdf/scientificamerican08241872-120a.pdf.

Anonymous:1872:PC

[3248] Anonymous. Plaster casts. Scientific American, 27(8):120, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-120b.pdf.

Anonymous:1872:LPRf

[3249] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(8):121, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
121.pdf.

Anonymous:1872:MIc

[3250] Anonymous. Miscellaneous items. Scientific American, 27(8):121, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n8/pdf/scientificamerican08241872-121a.pdf.

Anonymous:1872:ACbj

[3251] Anonymous. Answers to correspondents. Scientific American, 27(8):122, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
122b.pdf.

Anonymous:1872:RAFbh

[3252] Anonymous. Recent American and foreign patents. Scientific American, 27(8):122–123, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
122c.pdf.

Anonymous:1872:NPLj

[3253] Anonymous. New patent law in Canada. Scientific American, 27(8):124, August 24, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n8/pdf/scientificamerican08241872-
124b.pdf.

Anonymous:1872:Cbh

[3254] Anonymous. Correspondence. Scientific American, 27(9):131–133, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-131.pdf.

Anonymous:1872:BNBba

[3255] Anonymous. Books: New books and publications. Scientific American, 27(9):137, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
137b.pdf.

Anonymous:1872:BPbk

[3256] Anonymous. Business and personal. Scientific American, 27(9):138, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-138.pdf.

Anonymous:1872:NQbi

[3257] Anonymous. Notes and queries. Scientific American, 27(9):138, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-138a.pdf.

Anonymous:1872:IIbf

[3258] Anonymous. Index of inventions. Scientific American, 27(9):140, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-140.pdf.

Anonymous:1872:INM

[3259] Anonymous. Improved napping machine. Scientific American, 27 (9):127–128, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
127.pdf.

Anonymous:1872:REe

[3260] Anonymous. Recipes and experiments. Scientific American, 27 (9):129, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
129.pdf.

Anonymous:1872:MFC

[3261] Anonymous. Mold fob casting and chilling sleigh shoes. Scientific American, 27(9):130, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
130.pdf.

Anonymous:1872:HMC

[3262] Anonymous. How to make a cheap hydrogen lamp. Scientific American, 27(9):130, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
130a.pdf.

Anonymous:1872:LSB

[3263] Anonymous. Law’s Shuttle binder. Scientific American, 27(9): 130–131, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
130b.pdf.

Anonymous:1872:MP

[3264] Anonymous. Manufacture of pins. Scientific American, 27(9):133, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-133.pdf.

Anonymous:1872:SPS

[3265] Anonymous. Spring power for sewing machines. Scientific American, 27(9):134, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
134.pdf.

Anonymous:1872:GAL

[3266] Anonymous. Goodridge’s automatic lamp extinguisher. Scientific American, 27(9):134, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
134a.pdf.

Anonymous:1872:TMa

[3267] Anonymous. Terrestrial magnetism. Scientific American, 27(9):135, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-135.pdf.

Anonymous:1872:IEV

[3268] Anonymous. The International Exposition at Vienna, 1873. Scientific American, 27(9):135, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
135a.pdf.

Anonymous:1872:TUH

[3269] Anonymous. Tunnel under the Harlem River, New York. Scientific American, 27(9):135, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
135b.pdf.

Anonymous:1872:WBA

[3270] Anonymous. William Bridges Adams. Scientific American, 27 (9):135, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
135c.pdf.

Anonymous:1872:SAT

[3271] Anonymous. Security against thieves. Scientific American, 27 (9):135, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
135d.pdf.

Anonymous:1872:ISCb

[3272] Anonymous. The international statistical Congress. Scientific American, 27(9):135–136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
135e.pdf.

Anonymous:1872:NCS

[3273] Anonymous. A new canal steamer. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-136.pdf.

Anonymous:1872:NRA

[3274] Anonymous. The new railways across the continent. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136a.pdf.

Anonymous:1872:IMN

[3275] Anonymous. Improvements that are much needed. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136b.pdf.

Anonymous:1872:NSB

[3276] Anonymous. A new suspension bridge. Scientific American, 27 (9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136c.pdf.

Anonymous:1872:ERA

[3277] Anonymous. The electrical railway alarm. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136d.pdf.

Anonymous:1872:ONS

[3278] Anonymous. Old and new steam engines. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136e.pdf.

Anonymous:1872:CPN

[3279] Anonymous. The coolest place in New York. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136f.pdf.

Anonymous:1872:PPO

[3280] Anonymous. Promotions at the Patent Office. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136g.pdf.

Anonymous:1872:DA

[3281] Anonymous. Diamonds in Arizona. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n9/pdf/scientificamerican08311872-136h.pdf.

Anonymous:1872:WIB

[3282] Anonymous. Why it has been so hot. Scientific American, 27 (9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136i.pdf.

Anonymous:1872:MAT

[3283] Anonymous. The meteors of August Tenth. Scientific American, 27(9):136, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
136j.pdf.

Anonymous:1872:SPIx

[3284] Anonymous. Scientific and practical information. Scientific American, 27(9):137, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
137.pdf.

Anonymous:1872:LPRg

[3285] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(9):137, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
137a.pdf.

Anonymous:1872:ACbk

[3286] Anonymous. Answers to correspondents. Scientific American, 27(9):138, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
138b.pdf.

Anonymous:1872:RAFbi

[3287] Anonymous. Recent American and foreign patents. Scientific American, 27(9):139, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
139.pdf.

Anonymous:1872:NPLk

[3288] Anonymous. New patent law in Canada. Scientific American, 27(9):140, August 31, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n9/pdf/scientificamerican08311872-
140a.pdf.

Anonymous:1872:Cbi

[3289] Anonymous. Correspondence. Scientific American, 27(10):148, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n10/pdf/scientificamerican09071872-148.pdf.

Anonymous:1872:NQbj

[3290] Anonymous. Notes and queries. Scientific American, 27(10): 153, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
153a.pdf.

Anonymous:1872:BPbl

[3291] Anonymous. Business and personal. Scientific American, 27 (10):154, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
154.pdf.

Anonymous:1872:IIbg

[3292] Anonymous. Index of inventions. Scientific American, 27(10): 156, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
156.pdf.

Anonymous:1872:BNBbb

[3293] Anonymous. Books: New books and publications. Scientific American, 27(10):156, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
156c.pdf.

Anonymous:1872:SBR

[3294] Anonymous. Spool blank roughing machine. Scientific American, 27(10):143, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
143.pdf.

Anonymous:1872:CHF

[3295] Anonymous. Combined hand and foot circular sawing machine. Scientific American, 27(10):143, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
143a.pdf.

Anonymous:1872:HAM

[3296] Anonymous. How axes are made. Scientific American, 27(10): 144, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
144.pdf.

Anonymous:1872:CSW

[3297] Anonymous. The car shops of West Albany. Scientific American, 27(10):144, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
144a.pdf.

Anonymous:1872:ASE

[3298] Anonymous. The advantages of scientific education. Scientific American, 27(10):145–146, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
145.pdf.

Anonymous:1872:DIM

[3299] Anonymous. Detection of impurities in milk. Scientific American, 27(10):146, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
146.pdf.

Anonymous:1872:HBB

[3300] Anonymous. Handles for barrels and boxes. Scientific American, 27(10):146, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
146a.pdf.

Anonymous:1872:BBb

[3301] Anonymous. Bed bottom. Scientific American, 27(10):147, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n10/pdf/scientificamerican09071872-147.pdf.

Anonymous:1872:GUP

[3302] Anonymous. Giffard’s universal piston. Scientific American, 27 (10):147, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
147a.pdf.

Anonymous:1872:AHb

[3303] Anonymous.  American horses.  Scientific American, 27(10):

147, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
147b.pdf.

Anonymous:1872:MId

[3304] Anonymous. Miscellaneous items. Scientific American, 27(10): 148–149, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
148a.pdf.

Anonymous:1872:FWL

[3305] Anonymous. Fancy weaving loom. Scientific American, 27(10): 150, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
150.pdf.

Anonymous:1872:GDP

[3306] Anonymous. Griswold’s dental plugging instrument. Scientific American, 27(10):150, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
150a.pdf.

Anonymous:1872:STa

[3307] Anonymous. Science and theology. Scientific American, 27(10): 151, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
151.pdf.

Anonymous:1872:PAIa

[3308] Anonymous. Progress of American industry — inauguration of five Grand Industrial Expositions. Scientific American, 27(10):151, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n10/pdf/scientificamerican09071872-151a.pdf.

Anonymous:1872:PND

[3309] Anonymous. Progress of the new docks in this city. — artificial stone blocks weighing seventy-nine tuns. Scientific American, 27 (10):151, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
151b.pdf.

Anonymous:1872:IEc

[3310] Anonymous. Iridescent engravings. Scientific American, 27(10): 151–152, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
151c.pdf.

Anonymous:1872:LSC

[3311] Anonymous. The Lake Superior copper regions. Scientific American, 27(10):152, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
152.pdf.

Anonymous:1872:RT

[3312] Anonymous. Remarkable thunderstorms. Scientific American, 27 (10):152, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
152a.pdf.

Anonymous:1872:MV

[3313] Anonymous. Manufacture of vinegar. Scientific American, 27 (10):152, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
152b.pdf.

Anonymous:1872:SPIy

[3314] Anonymous. Scientific and practical information. Scientific American, 27(10):152, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
152c.pdf.

Anonymous:1872:LPRh

[3315] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(10):153, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
153.pdf.

Anonymous:1872:ACbl

[3316] Anonymous. Answers to correspondents. Scientific American, 27 (10):154, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
154a.pdf.

Anonymous:1872:RAFbj

[3317] Anonymous. Recent American and foreign patents. Scientific American, 27(10):154–155, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
154b.pdf.

Anonymous:1872:PHIa

[3318] Anonymous. Practical hints to inventors. Scientific American, 27 (10):155, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
155.pdf.

Anonymous:1872:FPHa

[3319] Anonymous. Foreign patents — a hint to patentees. Scientific American, 27(10):156, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
156a.pdf.

Anonymous:1872:NPLl

[3320] Anonymous. New patent law in Canada. Scientific American, 27 (10):156, September 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n10/pdf/scientificamerican09071872-
156b.pdf.

Anonymous:1872:Cbj

[3321] Anonymous. Correspondence. Scientific American, 27(11):164– 165, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
164.pdf.

Anonymous:1872:BPbm

[3322] Anonymous. Business and personal. Scientific American, 27(11): 169–170, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
169.pdf.

Anonymous:1872:NQbk

[3323] Anonymous. Notes and queries. Scientific American, 27(11): 170, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
170.pdf.

Anonymous:1872:IIbh

[3324] Anonymous. Index of inventions. Scientific American, 27(11): 171–172, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
171.pdf.

Anonymous:1872:BNBbc

[3325] Anonymous. Books: New books and publications. Scientific American, 27(11):172, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
172b.pdf.

Anonymous:1872:IST

[3326] Anonymous. Improved steam trap. Scientific American, 27(11): 159, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
159.pdf.

Anonymous:1872:WF

[3327] Anonymous. Writing fluids. Scientific American, 27(11):160– 161, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
160.pdf.

Anonymous:1872:DTW

[3328] Anonymous. Double turbine water wheel. Scientific American, 27 (11):162, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
162.pdf.

Anonymous:1872:APb

[3329] Anonymous. Adjustable plow. Scientific American, 27(11):162, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n11/pdf/scientificamerican09141872-162a.pdf.

Anonymous:1872:CBk

[3330] Anonymous. The carpenter bee. Scientific American, 27(11): 163, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
163.pdf.

Anonymous:1872:SAD

[3331] Anonymous. Simple apparatus for out-door sketching. Scientific American, 27(11):163, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
163a.pdf.

Anonymous:1872:PBH

[3332] Anonymous. Paper boats — how made. Scientific American, 27 (11):163, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
163b.pdf.

Anonymous:1872:ISa

[3333] Anonymous. Improved shops. Scientific American, 27(11):163, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n11/pdf/scientificamerican09141872-163c.pdf.

Anonymous:1872:SPIz

[3334] Anonymous. Scientific and practical information. Scientific American, 27(11):165, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
165.pdf.

Anonymous:1872:AST

[3335] Anonymous. Albany steam trap. Scientific American, 27(11): 166, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
166.pdf.

Anonymous:1872:RRJ

[3336] Anonymous. Railway rail joint. Scientific American, 27(11): 166, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
166a.pdf.

Anonymous:1872:TMc

[3337] Anonymous. Terrestrial magnetism. Scientific American, 27(11): 167, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
167.pdf.

Anonymous:1872:SAMe

[3338] Anonymous. Substances acting mechanically to prevent the formation of scale in boilers. Scientific American, 27(11):167, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n11/pdf/scientificamerican09141872-167a.pdf.

Anonymous:1872:SSa

[3339] Anonymous. The star Sirius. Scientific American, 27(11):167– 168, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
167b.pdf.

Anonymous:1872:RAP

[3340] Anonymous. Renewed activity at the Patent Officer — re organization and new classification. — appointment and promotions. Scientific American, 27(11):168, September 14, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n11/pdf/scientificamerican09141872-168.pdf.

Anonymous:1872:AR

[3341] Anonymous. An astronomical robbery. Scientific American, 27 (11):168, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
168a.pdf.

Anonymous:1872:NIEb

[3342] Anonymous. The Newark Industrial Exposition. Scientific American, 27(11):168, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
168b.pdf.

Anonymous:1872:LPRi

[3343] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(11):168–169, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
168c.pdf.

Anonymous:1872:ACbm

[3344] Anonymous. Answers to correspondents. Scientific American, 27 (11):170, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
170a.pdf.

Anonymous:1872:RAFbk

[3345] Anonymous. Recent American and foreign patents. Scientific American, 27(11):170–171, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
170b.pdf.

Anonymous:1872:FPHb

[3346] Anonymous. Foreign patents — a hint to patentees. Scientific American, 27(11):172, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
172.pdf.

Anonymous:1872:NPLm

[3347] Anonymous. New patent law in Canada. Scientific American, 27 (11):172, September 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n11/pdf/scientificamerican09141872-
172a.pdf.

Anonymous:1872:Cbl

[3348] Anonymous. Correspondence. Scientific American, 27(12):180– 181, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
180.pdf.

Anonymous:1872:BPbn

[3349] Anonymous. Business and personal. Scientific American, 27 (12):185, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
185a.pdf.

Anonymous:1872:NQbl

[3350] Anonymous. Notes and queries. Scientific American, 27(12): 185, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
185b.pdf.

Anonymous:1872:IIbi

[3351] Anonymous. Index of inventions. Scientific American, 27(12): 187–188, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
187c.pdf.

Anonymous:1872:MMF

[3352] Anonymous. Molecular motion the fundamental problem of nature. Scientific American, 27(12):175, September 21, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n12/pdf/scientificamerican09211872-175.pdf.

Anonymous:1872:PPSb

[3353] Anonymous. Patent punching and shearing machine. Scientific American, 27(12):175–176, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
175a.pdf.

Anonymous:1872:LV

[3354] Anonymous. A list of velocities. Scientific American, 27(12): 177, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
177.pdf.

Anonymous:1872:IFS

[3355] Anonymous. Improved flush strap hinge. Scientific American, 27 (12):178, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
178.pdf.

Anonymous:1872:GPM

[3356] Anonymous. Gold pen making. Scientific American, 27(12): 178, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
178a.pdf.

Anonymous:1872:GUB

[3357] Anonymous. Glasgow University buildings. Scientific American, 27 (12):179, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
179.pdf.

Anonymous:1872:PH

[3358] Anonymous. Photographic hints. Scientific American, 27(12): 179, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
179a.pdf.

Anonymous:1872:KGT

[3359] Anonymous. Knitted goods and their manufacture. Scientific American, 27(12):181, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
181.pdf.

Anonymous:1872:SPIba

[3360] Anonymous. Scientific and practical information. Scientific American, 27(12):181, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
181a.pdf.

Anonymous:1872:IRF

[3361] Anonymous. Improved road or farm gate. Scientific American, 27 (12):182, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
182.pdf.

Anonymous:1872:GSL

[3362] Anonymous. Grosvenor’s self-lubricating loose pulley. Scientific American, 27(12):182, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
182a.pdf.

Anonymous:1872:WRS

[3363] Anonymous. Wheel and roller sash catch. Scientific American, 27 (12):182, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
182b.pdf.

Anonymous:1872:SMD

[3364] Anonymous. Serious marine disasters — improved safety appliances greatly needed. Scientific American, 27(12):183, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n12/pdf/scientificamerican09211872-183.pdf.

Anonymous:1872:CCDa

[3365] Anonymous. Car coupling dangers. Scientific American, 27(12): 183, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
183a.pdf.

Anonymous:1872:ASM

[3366] Anonymous. An auction sale of models at the Patent Office. Scientific American, 27(12):183, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
183b.pdf.

Anonymous:1872:CIE

[3367] Anonymous. The Cincinnati Industrial Exposition. Scientific American, 27(12):183, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
183c.pdf.

Anonymous:1872:MBAa

[3368] Anonymous. Meeting of the British Association. — address by a New York Herald reporter. — Livingstone in Africa. Scientific American, 27(12):183–184, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
183d.pdf.

Anonymous:1872:OAI

[3369] Anonymous. The opening of the American Institute Fair. Scientific American, 27(12):184, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
184.pdf.

Anonymous:1872:AAAb

[3370] Anonymous. The American Association for the Advancement of Science. Scientific American, 27(12):184, September 21, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n12/pdf/scientificamerican09211872-184a.pdf.

Anonymous:1872:TV

[3371] Anonymous. Table of velocities. Scientific American, 27(12): 184–185, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
184b.pdf.

Anonymous:1872:EPV

[3372] Anonymous. Extension of patents. value of the invention. Scientific American, 27(12):185, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
185.pdf.

Anonymous:1872:ACbn

[3373] Anonymous. Answers to correspondents. Scientific American, 27 (12):185, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
185c.pdf.

Anonymous:1872:RAFbl

[3374] Anonymous. Recent American and foreign patents. Scientific American, 27(12):186–187, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
186.pdf.

Anonymous:1872:NPLn

[3375] Anonymous. New patent law in Canada. Scientific American, 27 (12):187, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
187.pdf.

Anonymous:1872:FPHc

[3376] Anonymous. Foreign patents — a hint to patentees. Scientific American, 27(12):187, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
187a.pdf.

Anonymous:1872:PHIb

[3377] Anonymous. Practical hints to inventors. Scientific American, 27 (12):187, September 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n12/pdf/scientificamerican09211872-
187b.pdf.

Anonymous:1872:Cbm

[3378] Anonymous. Correspondence. Scientific American, 27(13):196, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n13/pdf/scientificamerican09281872-196.pdf.

Anonymous:1872:BPbo

[3379] Anonymous. Business and personal. Scientific American, 27 (13):202, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
202.pdf.

Anonymous:1872:NQbm

[3380] Anonymous. Notes and queries. Scientific American, 27(13): 202, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
202a.pdf.

Anonymous:1872:IIbj

[3381] Anonymous. Index of inventions. Scientific American, 27(13): 203, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
203.pdf.

Anonymous:1872:BNBbd

[3382] Anonymous. Books: New books and publications. Scientific American, 27(13):204, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
204c.pdf.

Anonymous:1872:CRAa

[3383] Anonymous. Combined regulator and alarm. Scientific American, 27(13):191, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
191.pdf.

Anonymous:1872:MOO

[3384] Anonymous. The manufacture of olive oil in California. Scientific American, 27(13):192, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
192.pdf.

Anonymous:1872:IMC

[3385] Anonymous. Iron as a material of construction. Scientific American, 27(13):192–193, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
192a.pdf.

Anonymous:1872:PDb

[3386] Anonymous. The pneumatic despatch. Scientific American, 27 (13):193, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
193.pdf.

Anonymous:1872:CSMb

[3387] Anonymous. The Cuban sugar manufacture — modern machinery. Scientific American, 27(13):195, September 28, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n13/pdf/scientificamerican09281872-195.pdf.

Anonymous:1872:SBP

[3388] Anonymous. The sand blast process for cutting hard substances, engraving on glass, etc. Scientific American, 27(13):195, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n13/pdf/scientificamerican09281872-195a.pdf.

Anonymous:1872:IHBa

[3389] Anonymous. Improvement in harness breast straps. Scientific American, 27(13):195, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
195b.pdf.

Rains:1872:PMB

[3390] George W. Rains and Professor. Portable medical batteries. Scientific American, 27(13):196–197, September 28, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n13/pdf/scientificamerican09281872-196a.pdf.

Anonymous:1872:CHPb

[3391] Anonymous. Coffee hulling and polishing machine. Scientific American, 27(13):198, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
198.pdf.

Anonymous:1872:ISBb

[3392] Anonymous. Improved steam boiler. Scientific American, 27 (13):198, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
198a.pdf.

Anonymous:1872:BVEa

[3393] Anonymous. Binary vapor engines. Scientific American, 27(13): 199, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
199.pdf.

Anonymous:1872:AGE

[3394] Anonymous. Another great engineering work. Scientific American, 27(13):199, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
199a.pdf.

Anonymous:1872:FCS

[3395] Anonymous. Fast canal steaming. Scientific American, 27(13): 199, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
199b.pdf.

Anonymous:1872:CEC

[3396] Anonymous. The causes of extraordinary conditions of weather. Scientific American, 27(13):199, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
199c.pdf.

Anonymous:1872:GEB

[3397] Anonymous. The government examination of breech loading arms. Scientific American, 27(13):199, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
199d.pdf.

Anonymous:1872:WDV

[3398] Anonymous. Who is Dr. Vander Weyde? Scientific American, 27(13):200, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
200.pdf.

Anonymous:1872:CMP

[3399] Anonymous. Chemical means of preventing scale in boilers. Scientific American, 27(13):200, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
200a.pdf.

Anonymous:1872:MFH

[3400] Anonymous. The marriage of father hyacinthe. Scientific American, 27(13):200, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
200b.pdf.

Anonymous:1872:POIa

[3401] Anonymous. Patent Office items. Scientific American, 27(13): 200, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
200c.pdf.

Anonymous:1872:AAAc

[3402] Anonymous. The American Association for the Advancement of Science. Scientific American, 27(13):200–201, September 28, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n13/pdf/scientificamerican09281872-200d.pdf.

Anonymous:1872:LPRj

[3403] Anonymous. Letter from Professor R. H. Thurston. Scientific American, 27(13):201–202, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
201.pdf.

Anonymous:1872:ACbo

[3404] Anonymous. Answers to correspondents. Scientific American, 27 (13):202, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
202b.pdf.

Anonymous:1872:RAFbm

[3405] Anonymous. Recent American and foreign patents. Scientific American, 27(13):202–203, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
202c.pdf.

Anonymous:1872:PHIc

[3406] Anonymous. Practical hints to inventors. Scientific American, 27 (13):204, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
204.pdf.

Anonymous:1872:NPLo

[3407] Anonymous. New patent law in Canada. Scientific American, 27 (13):204, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
204a.pdf.

Anonymous:1872:FPHd

[3408] Anonymous. Foreign patents — a hint to patentees. Scientific American, 27(13):204, September 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n13/pdf/scientificamerican09281872-
204b.pdf.

Anonymous:1872:Cbn

[3409] Anonymous. Correspondence. Scientific American, 27(14):212, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-212.pdf.

Anonymous:1872:BPbp

[3410] Anonymous. Business and personal. Scientific American, 27 (14):216–217, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
216d.pdf.

Anonymous:1872:NQbn

[3411] Anonymous. Notes and queries. Scientific American, 27(14):217, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-217.pdf.

Anonymous:1872:IIbk

[3412] Anonymous. Index of inventions. Scientific American, 27(14):219–220, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-219.pdf.

Anonymous:1872:CFA

[3413] Anonymous. Cod fishing in Alaska. Scientific American, 27(14):207, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-207.pdf.

Anonymous:1872:IFR

[3414] Anonymous. Improved feed regulator for boilers. Scientific American, 27(14):207–208, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
207a.pdf.

Anonymous:1872:GPSa

[3415] Anonymous. “greenbacks” and postage stamps. Scientific American, 27(14):208, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
208.pdf.

Anonymous:1872:SPIbb

[3416] Anonymous. Scientific and practical information. Scientific American, 27(14):208–209, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
208a.pdf.

Anonymous:1872:SHB

[3417] Anonymous. Sailors’ Home, Bombay, India. Scientific American, 27(14):210, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
210.pdf.

Anonymous:1872:WRT

[3418] Anonymous. Wire rope towage on the Danube. Scientific American, 27(14):211–212, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
211.pdf.

Anonymous:1872:LPRk

[3419] Anonymous. Letter from Professor R. H. Thurston. — last letter of the series. Scientific American, 27(14):213, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-213.pdf.

Anonymous:1872:ICP

[3420] Anonymous. Improved cotton press. Scientific American, 27(14):214, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-214.pdf.

Anonymous:1872:IHC

[3421] Anonymous. Improved hose coupling. Scientific American, 27(14):214, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-214a.pdf.

Anonymous:1872:DCF

[3422] Anonymous. The disintegrating or centrifugal flour mill. Scientific American, 27(14):214, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
214b.pdf.

Anonymous:1872:LDN

[3423] Anonymous. The latest discoveries in the nature of the Sun. Scientific American, 27(14):215, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
215.pdf.

Anonymous:1872:WMC

[3424] Anonymous. A working Man’s city. Scientific American, 27(14):215, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-215a.pdf.

Anonymous:1872:NCT

[3425] Anonymous. New canal through Florida. Scientific American, 27(14):215, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
215b.pdf.

Anonymous:1872:SCc

[3426] Anonymous. Science in court. Scientific American, 27(14):215–216, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-215c.pdf.

Anonymous:1872:FAIa

[3427] Anonymous. The fair of the American institute. Scientific American, 27(14):216, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
216.pdf.

Anonymous:1872:PRH

[3428] Anonymous. Professor R. H. Thurston’s letters. Scientific American, 27(14):216, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
216a.pdf.

Anonymous:1872:POIb

[3429] Anonymous. Patent Office items. Scientific American, 27(14):216, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-216b.pdf.

Anonymous:1872:CCDb

[3430] Anonymous. Car coupling dangers. Scientific American, 27(14):216, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-216c.pdf.

Anonymous:1872:ACbp

[3431] Anonymous. Answer to correspondents. Scientific American, 27 (14):217, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
217a.pdf.

Anonymous:1872:CRb

[3432] Anonymous. Communications received. Scientific American, 27 (14):217, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
217b.pdf.

Anonymous:1872:RAFbn

[3433] Anonymous. Recent American and foreign patents. Scientific American, 27(14):217–219, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
217c.pdf.

Anonymous:1872:PHId

[3434] Anonymous. Practical hints to inventors. Scientific American, 27(14):220, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n14/pdf/scientificamerican10051872-
220.pdf.

Anonymous:1872:Ic

[3435] Anonymous. To investors. Scientific American, 27(14):220, October 5, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n14/pdf/scientificamerican10051872-220a.pdf.

Anonymous:1872:Cbo

[3436] Anonymous. Correspondence. Scientific American, 27(15):228–229, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-228.pdf.

Anonymous:1872:BPbq

[3437] Anonymous. Business and personal. Scientific American, 27 (15):234, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
234.pdf.

Anonymous:1872:NQbo

[3438] Anonymous. Notes and queries. Scientific American, 27(15):234, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-234a.pdf.

Anonymous:1872:BNBbe

[3439] Anonymous. Books: New books and publications. Scientific American, 27(15):235, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
235a.pdf.

Anonymous:1872:IIbl 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
235b.pdf.

Anonymous:1872:NMD

[3441] Anonymous. New method of docking and excavating canals and of reclaiming waste land. Scientific American, 27(15):223–224, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-223.pdf.

Anonymous:1872:WRW

[3442] Anonymous. The wonderful regions of the west. Scientific American, 27(15):224–225, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
224.pdf.

Anonymous:1872:PMB

[3443] Anonymous. Pneumatic machinery for bridges. Scientific American, 27(15):226–227, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
226.pdf.

Anonymous:1872:USA

[3444] Anonymous. The use of sulphurous acid in distilleries. Scientific American, 27(15):227, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
227.pdf.

Anonymous:1872:GPSb

[3445] Anonymous. “greenbacks” and postage stamps. Scientific American, 27(15):227–228, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
227a.pdf.

Anonymous:1872:CWM

[3446] Anonymous. Coffee washing machine. Scientific American, 27 (15):230, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
230.pdf.

Anonymous:1872:IOS

[3447] Anonymous. Improved ore separator. Scientific American, 27 (15):230, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
230a.pdf.

Anonymous:1872:AIA

[3448] Anonymous. American inventors in Austria. Scientific American, 27(15):231, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
231.pdf.

Anonymous:1872:HDW

[3449] Anonymous. How do we take cold? Scientific American, 27(15):231, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-231a.pdf.

Anonymous:1872:ICa

[3450] Anonymous. Interoceanic canal. Scientific American, 27(15):231, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-231b.pdf.

Anonymous:1872:EGP

[3451] Anonymous. Efforts in Germany to prevent emigration. Scientific American, 27(15):231, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
231c.pdf.

Anonymous:1872:DYI

[3452] Anonymous. Develop your inventions. Scientific American, 27 (15):232, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
232.pdf.

Anonymous:1872:D

[3453] Anonymous. Disinfectants. Scientific American, 27(15):232, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-232a.pdf.

Anonymous:1872:MSI

[3454] Anonymous. Molding sand for iron founderies. Scientific American, 27(15):232, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
232b.pdf.

Anonymous:1872:AHD

[3455] Anonymous. Animal heat and disease. Scientific American, 27 (15):232–233, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
232c.pdf.

Anonymous:1872:EHK

[3456] Anonymous. Edmund Henry Knight. Scientific American, 27 (15):233, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
233.pdf.

Anonymous:1872:FAIb

[3457] Anonymous. The fair of the American institute. Scientific American, 27(15):233, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
233a.pdf.

Anonymous:1872:CPf

[3458] Anonymous. Copies of patents. Scientific American, 27(15):233, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-233b.pdf.

Anonymous:1872:ACbq

[3459] Anonymous. Answers to correspondents. Scientific American, 27 (15):234–235, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
234b.pdf.

Anonymous:1872:RAFbo

[3460] Anonymous. Recent American and foreign patents. Scientific American, 27(15):235, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
235.pdf.

Anonymous:1872:PHIe

[3461] Anonymous. Practical hints to inventors. Scientific American, 27(15):236, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n15/pdf/scientificamerican10121872-
236.pdf.

Anonymous:1872:Id

[3462] Anonymous. To investors. Scientific American, 27(15):236, October 12, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n15/pdf/scientificamerican10121872-236a.pdf.

Anonymous:1872:Cbp

[3463] Anonymous. Correspondence. Scientific American, 27(16):244–245, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-244.pdf.

Anonymous:1872:BPbr

[3464] Anonymous. Business and personal. Scientific American, 27 (16):249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
249c.pdf.

Anonymous:1872:NQbp

[3465] Anonymous. Notes and queries. Scientific American, 27(16):249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-249d.pdf.

Anonymous:1872:IIbm

[3466] Anonymous. Index of inventions. Scientific American, 27(16):251, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-251.pdf.

Anonymous:1872:BNBbf

[3467] Anonymous. Books: New books and publication. Scientific American, 27(16):252, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
252a.pdf.

Anonymous:1872:ISEc

[3468] Anonymous. Improved steam engine. Scientific American, 27 (16):239–240, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
239.pdf.

Anonymous:1872:EDE

[3469] Anonymous. Exhibition of Domestic Economy in Paris. Scientific American, 27(16):240–241, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
240.pdf.

Anonymous:1872:FB

[3470] Anonymous. Floating breakwater. Scientific American, 27(16):242, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-242.pdf.

Anonymous:1872:SPa

[3471] Anonymous. Street pavements. Scientific American, 27(16):242, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-242a.pdf.

Anonymous:1872:TMd

[3472] Anonymous. Transmission of motion. Scientific American, 27 (16):243–244, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
243.pdf.

Anonymous:1872:NMP

[3473] Anonymous. New method of propelling canal boats. Scientific American, 27(16):246, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
246.pdf.

Anonymous:1872:PMS

[3474] Anonymous. Portable music stand. Scientific American, 27(16):246, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-246a.pdf.

Anonymous:1872:LIEa

[3475] Anonymous. The London International Exhibition of 1873. Scientific American, 27(16):247, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
247.pdf.

Anonymous:1872:BVEb

[3476] Anonymous. Binary vapor engines. Scientific American, 27(16):247, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-247a.pdf.

Anonymous:1872:ATE

[3477] Anonymous. Advantages of a technical education. Scientific American, 27(16):247, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
247b.pdf.

Anonymous:1872:IWa

[3478] Anonymous. An invention wanted. Scientific American, 27(16):247, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-247c.pdf.

Anonymous:1872:LNA

[3479] Anonymous. The latest news about the Moon. Scientific American, 27(16):247–248, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
247d.pdf.

Anonymous:1872:FAIc

[3480] Anonymous. The fair of the American institute. Scientific American, 27(16):248, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
248.pdf.

Anonymous:1872:FWB

[3481] Anonymous. The frozen well of Brandon, Vermont. Scientific American, 27(16):248–249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
248a.pdf.

Anonymous:1872:MIE

[3482] Anonymous. The Moscow Industrial Exposition. Scientific American, 27(16):249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
249.pdf.

Anonymous:1872:EAC

[3483] Anonymous. Expected advance in crude petroleum. Scientific American, 27(16):249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
249a.pdf.

Anonymous:1872:FLc

[3484] Anonymous. Francis Lieber. Scientific American, 27(16):249, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-249b.pdf.

Anonymous:1872:ACbr

[3485] Anonymous. Answers to correspondents. Scientific American, 27 (16):249–250, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
249e.pdf.

Anonymous:1872:CPg

[3486] Anonymous. Copies of patents. Scientific American, 27(16):250, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-250.pdf.

Anonymous:1872:RAFbp

[3487] Anonymous. Recent American and foreign patents. Scientific American, 27(16):250–251, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
250a.pdf.

Anonymous:1872:PHIf

[3488] Anonymous. Practical hints to inventors. Scientific American, 27(16):252, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n16/pdf/scientificamerican10191872-
252.pdf.

Anonymous:1872:If

[3489] Anonymous. To investors. Scientific American, 27(16):252, October 19, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n16/pdf/scientificamerican10191872-252b.pdf.

Anonymous:1872:Cbq

[3490] Anonymous. Correspondence. Scientific American, 27(17):259–260, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-259b.pdf.

Anonymous:1872:BPbs

[3491] Anonymous. Business and personal. Scientific American, 27 (17):266, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
266.pdf.

Anonymous:1872:IIbn

[3492] Anonymous. Index of inventions. Scientific American, 27(17):267, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-267.pdf.

Anonymous:1872:ISM

[3493] Anonymous. Improved sewing machine. Scientific American, 27 (17):255, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
255.pdf.

Anonymous:1872:PPA

[3494] Anonymous. The Page patent. — the attempts to enforce it to be resisted. Scientific American, 27(17):256, October 26, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n17/pdf/scientificamerican10261872-256.pdf.

Roberts:1872:SSS

[3495] Charles Roberts, F. R. C. S., and Etc. Sulphozone, a substitute for sulphur. Scientific American, 27(17):257, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-257.pdf.

Anonymous:1872:SFD

[3496] Anonymous. Self-feeding drill. Scientific American, 27(17):258, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-258.pdf.

Anonymous:1872:REf

[3497] Anonymous. Recipes and experiments. Scientific American, 27 (17):259, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
259.pdf.

Anonymous:1872:DMW

[3498] Anonymous. Death of Mr. W. H. Seward. Scientific American, 27(17):259, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
259a.pdf.

Anonymous:1872:VEB

[3499] Anonymous. The Vienna Exposition buildings. Scientific American, 27(17):260, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
260.pdf.

Anonymous:1872:NLA

[3500] Anonymous. New line of Atlantic steamers. Scientific American, 27 (17):260–261, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
260a.pdf.

Anonymous:1872:ISG

[3501] Anonymous. Improved safety gun lock. Scientific American, 27 (17):262, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
262.pdf.

Anonymous:1872:RHF

[3502] Anonymous. Rafter hook and fire escape. Scientific American, 27(17):262, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
262a.pdf.

Anonymous:1872:ML

[3503] Anonymous. More light. Scientific American, 27(17):263, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-263.pdf.

Anonymous:1872:FAA

[3504] Anonymous. The French Association for the Advancement of Science. Scientific American, 27(17):263, October 26, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n17/pdf/scientificamerican10261872-263a.pdf.

Anonymous:1872:NNB

[3505] Anonymous. A novel news bulletin. Scientific American, 27 (17):263, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
263b.pdf.

Anonymous:1872:PJT

[3506] Anonymous. Professor John Tyndall. Scientific American, 27 (17):263, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
263c.pdf.

Anonymous:1872:ASN

[3507] Anonymous. American steeling — a new and remarkable alloy. Scientific American, 27(17):263–264, October 26, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n17/pdf/scientificamerican10261872-263d.pdf.

Anonymous:1872:WLW

[3508] Anonymous. The whistling lantern — who is the real inventor? Scientific American, 27(17):264, October 26, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n17/pdf/scientificamerican10261872-264.pdf.

Anonymous:1872:SSb

[3509] Anonymous. Soup and savants. Scientific American, 27(17):264, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-264a.pdf.

Anonymous:1872:LBF

[3510] Anonymous. Leather belting — facts for the determination of the quality of the leather. Scientific American, 27(17):264, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-264b.pdf.

Anonymous:1872:ME

[3511] Anonymous. A mechanical eye. Scientific American, 27(17):265, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-265.pdf.

Anonymous:1872:FAId

[3512] Anonymous. The fair of the American institute. Scientific American, 27(17):265, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
265a.pdf.

Anonymous:1872:ACbs

[3513] Anonymous. Answers to correspondents. Scientific American, 27 (17):266, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
266a.pdf.

Anonymous:1872:RAFbq

[3514] Anonymous. Recent American and foreign patents. Scientific American, 27(17):266–267, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
266b.pdf.

Anonymous:1872:CPh

[3515] Anonymous. Copies of patents. Scientific American, 27(17):267, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n17/pdf/scientificamerican10261872-267a.pdf.

Anonymous:1872:PHIg

[3516] Anonymous. Practical hints to inventors. Scientific American, 27(17):268, October 26, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n17/pdf/scientificamerican10261872-
268.pdf.

Anonymous:1872:Cbr

[3517] Anonymous. Correspondence. Scientific American, 27(18):276– 277, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
276.pdf.

Anonymous:1872:BPbt

[3518] Anonymous. Business and personal. Scientific American, 27 (18):281, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
281.pdf.

Anonymous:1872:NQbq

[3519] Anonymous. Notes and queries. Scientific American, 27(18): 281, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
281b.pdf.

Anonymous:1872:IIbo

[3520] Anonymous. Index of inventions. Scientific American, 27(18): 283, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
283.pdf.

Anonymous:1872:BNBbg

[3521] Anonymous. Books: New books and publications. Scientific American, 27(18):284, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
284a.pdf.

Anonymous:1872:PE

[3522] Anonymous. Pneumatic elevator. Scientific American, 27(18): 271, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
271.pdf.

Anonymous:1872:PAb

[3523] Anonymous. Plating with aluminum. Scientific American, 27 (18):272, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
272.pdf.

Anonymous:1872:HWE

[3524] Anonymous. The health of the workmen of the East River Bridge. Scientific American, 27(18):272–273, November 2, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n18/pdf/scientificamerican11021872-272a.pdf.

Anonymous:1872:CPW

[3525] Anonymous. Combined pipe wrench and vise. Scientific American, 27(18):274, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
274.pdf.

Anonymous:1872:ITR

[3526] Anonymous. Improved tool rest. Scientific American, 27(18): 274, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
274a.pdf.

Anonymous:1872:CNC

[3527] Anonymous. Corundum of North Carolina and Georgia. Scientific American, 27(18):274, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
274b.pdf.

Anonymous:1872:RBC

[3528] Anonymous. Road-building in the Caucasus. Scientific American, 27(18):275, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
275.pdf.

Anonymous:1872:PTF

[3529] Anonymous. Professor Tyndall’s first lecture in America. Scientific American, 27(18):275, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
275a.pdf.

Anonymous:1872:HPCb

[3530] Anonymous. How patterns in carpets are made. Scientific American, 27(18):275, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
275b.pdf.

Anonymous:1872:MSM

[3531] Anonymous. The Manchester Scientific and Mechanical Society. Scientific American, 27(18):275, November 2, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n18/pdf/scientificamerican11021872-275c.pdf.

Anonymous:1872:MIf

[3532] Anonymous. Malleable iron. Scientific American, 27(18):277, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n18/pdf/scientificamerican11021872-277.pdf.

Anonymous:1872:NMSb

[3533] Anonymous. New method for sensitizing collodion. Scientific American, 27(18):277, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
277a.pdf.

Anonymous:1872:PCAb

[3534] Anonymous. Publishers’s card to advertisers. Scientific American, 27(18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279.pdf.

Anonymous:1872:MDT

[3535] Anonymous. A million dollar telescope. Scientific American, 27 (18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279a.pdf.

Anonymous:1872:DPC

[3536] Anonymous. Downfall of the paper collar ring. Scientific American, 27(18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279b.pdf.

Anonymous:1872:DPE

[3537] Anonymous. Discouraging prospects in England. Scientific American, 27(18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279c.pdf.

Anonymous:1872:IPE

[3538] Anonymous. An important patent extension denied. Scientific American, 27(18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279d.pdf.

Anonymous:1872:PJHa

[3539] Anonymous. Professor J. H. Pepper. Scientific American, 27 (18):279, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279e.pdf.

Anonymous:1872:UR

[3540] Anonymous. Underground railroads. Scientific American, 27(18): 279–280, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
279f.pdf.

Anonymous:1872:HWP

[3541] Anonymous. Hanging wall paper. Scientific American, 27(18): 280, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
280.pdf.

Anonymous:1872:EMC

[3542] Anonymous. Edwin Marcus Chaffee. Scientific American, 27 (18):280, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
280a.pdf.

Anonymous:1872:PJW

[3543] Anonymous. Professor John W. Frazer. Scientific American, 27 (18):280, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
280b.pdf.

Anonymous:1872:WPS

[3544] Anonymous. William Prescott Smith. Scientific American, 27 (18):280, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
280c.pdf.

Anonymous:1872:SPIbc

[3545] Anonymous. Scientific and practical information. Scientific American, 27(18):280–281, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
280d.pdf.

Anonymous:1872:CPi

[3546] Anonymous. Copies of patents. Scientific American, 27(18): 281, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
281a.pdf.

Anonymous:1872:ACbt

[3547] Anonymous. Answer to correspondence. Scientific American, 27 (18):281–282, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
281c.pdf.

Anonymous:1872:CRc

[3548] Anonymous. Communications received. Scientific American, 27 (18):282, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
282.pdf.

Anonymous:1872:RAFbr

[3549] Anonymous. Recent American and foreign patents. Scientific American, 27(18):282–283, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
282a.pdf.

Anonymous:1872:PHIh

[3550] Anonymous. Practical hints to inventors. Scientific American, 27 (18):284, November 2, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n18/pdf/scientificamerican11021872-
284.pdf.

Anonymous:1872:Cbs

[3551] Anonymous. Correspondence. Scientific American, 27(19):292– 293, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
292.pdf.

Anonymous:1872:BPbu

[3552] Anonymous. Business and personal. Scientific American, 27 (19):297, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
297b.pdf.

Anonymous:1872:NQbr

[3553] Anonymous. Notes and queries. Scientific American, 27(19): 298, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
298.pdf.

Anonymous:1872:IIbp

[3554] Anonymous. Index of inventions. Scientific American, 27(19): 299–300, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
299.pdf.

Anonymous:1872:BNBbh

[3555] Anonymous. Books: New books and publications. Scientific American, 27(19):300, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
300a.pdf.

Anonymous:1872:CSSb

[3556] Anonymous. The Cameron special steam pump. Scientific American, 27(19):287, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
287.pdf.

Parton:1872:JP

[3557] James Parton. Joseph Priestley. Scientific American, 27(19): 288, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
288.pdf.

Anonymous:1872:NTP

[3558] Anonymous. New tertiary and post tertiary birds. Scientific American, 27(19):289–290, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
289.pdf.

Anonymous:1872:EHF

[3559] Anonymous. Earthenware hotbed frames. Scientific American, 27 (19):290, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
290.pdf.

Anonymous:1872:CCH

[3560] Anonymous. Cheap country homes. Scientific American, 27(19): 291, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
291.pdf.

Anonymous:1872:IGB

[3561] Anonymous. Improved grate bars and bearer. Scientific American, 27(19):294, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
294.pdf.

Anonymous:1872:FIW

[3562] Anonymous. Fetta’s improved wagon wheel. Scientific American, 27(19):294, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
294a.pdf.

Anonymous:1872:PCAc

[3563] Anonymous. Publishers’ card to advertisers. Scientific American, 27(19):295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295.pdf.

Anonymous:1872:MPE

[3564] Anonymous. The molecular phenomena of evaporation. Scientific American, 27(19):295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295a.pdf.

Anonymous:1872:IPRb

[3565] Anonymous. Industrial progress in Russia. Scientific American, 27(19):295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295b.pdf.

Anonymous:1872:ABa

[3566] Anonymous. The Aurora Borealis. Scientific American, 27(19): 295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295c.pdf.

Anonymous:1872:SCd

[3567] Anonymous. Steam on the canals. Scientific American, 27(19): 295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295d.pdf.

Anonymous:1872:FHE

[3568] Anonymous. A fearful horse epidemic. Scientific American, 27 (19):295, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295e.pdf.

Anonymous:1872:TVS

[3569] Anonymous. Theology vs. science. Scientific American, 27(19): 295–296, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
295f.pdf.

Anonymous:1872:EPA

[3570] Anonymous. Extension of patents — the accounts. Scientific American, 27(19):296, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
296.pdf.

Anonymous:1872:SPIbd

[3571] Anonymous. Scientific and practical information. Scientific American, 27(19):296–297, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
296a.pdf.

Anonymous:1872:Ih

[3572] Anonymous. To investors. Scientific American, 27(19):297, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n19/pdf/scientificamerican11091872-297.pdf.

Anonymous:1872:CPj

[3573] Anonymous. Copies of patents. Scientific American, 27(19): 297, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
297a.pdf.

Anonymous:1872:ACbu

[3574] Anonymous. Answer to correspondents. Scientific American, 27 (19):298, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
298a.pdf.

Anonymous:1872:RAFbs

[3575] Anonymous. Recent American and foreign patents. Scientific American, 27(19):298–299, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
298b.pdf.

Anonymous:1872:PHIi

[3576] Anonymous. Practical hints to inventors. Scientific American, 27 (19):300, November 9, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n19/pdf/scientificamerican11091872-
300.pdf.

Anonymous:1872:Cbu

[3577] Anonymous. Correspondence. Scientific American, 27(20):308– 309, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
308.pdf.

Anonymous:1872:BPbv

[3578] Anonymous. Business and personal. Scientific American, 27 (20):313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313d.pdf.

Anonymous:1872:NQbs

[3579] Anonymous. Notes and queries. Scientific American, 27(20): 314, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
314.pdf.

Anonymous:1872:BNBbi

[3580] Anonymous. Books: New books and publications. Scientific American, 27(20):315, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
315.pdf.

Anonymous:1872:IIbq

[3581] Anonymous. Index of inventions. Scientific American, 27(20): 315–316, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
315a.pdf.

Anonymous:1872:LWF

[3582] Anonymous. Loom for weaving fabrics of any width. Scientific American, 27(20):303–304, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
303.pdf.

Anonymous:1872:TMe

[3583] Anonymous. Transmission of motion. Scientific American, 27 (20):304, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
304.pdf.

Anonymous:1872:UWG

[3584] Anonymous. The use of water glass in building. Scientific American, 27(20):305, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
305.pdf.

Anonymous:1872:OCC

[3585] Anonymous. Has our climate changed! Scientific American, 27 (20):305–306, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
305a.pdf.

Anonymous:1872:ICc

[3586] Anonymous. Improved conservatory. Scientific American, 27(20): 306, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
306.pdf.

Anonymous:1872:RSc

[3587] Anonymous. Rolling shutters. Scientific American, 27(20):307, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n20/pdf/scientificamerican11161872-307.pdf.

Anonymous:1872:PCM

[3588] Anonymous. Plating and coating metals. Scientific American, 27 (20):307–308, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
307a.pdf.

Robinson:1872:CUP

[3589] George Robinson. On certain undescribed properties of the concentrated solar rays. Scientific American, 27(20):309, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n20/pdf/scientificamerican11161872-309.pdf.

Anonymous:1872:FPC

[3590] Anonymous. Flexible pipe coupling. Scientific American, 27 (20):310, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
310.pdf.

Anonymous:1872:SPb

[3591] Anonymous. Sash pulleys. Scientific American, 27(20):310, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n20/pdf/scientificamerican11161872-310a.pdf.

Anonymous:1872:MIg

[3592] Anonymous. Manganiferous iron. Scientific American, 27(20): 310, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
310b.pdf.

Anonymous:1872:AM

[3593] Anonymous. The antiquity of man. Scientific American, 27 (20):311, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
311.pdf.

Anonymous:1872:VEa

[3594] Anonymous. The Vienna Exposition. Scientific American, 27 (20):311, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
311a.pdf.

Anonymous:1872:LSI

[3595] Anonymous. Life-saving inventions needed. Scientific American, 27(20):311, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
311b.pdf.

Anonymous:1872:HDS

[3596] Anonymous. The horse disease and street cars. Scientific American, 27(20):311, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
311c.pdf.

Anonymous:1872:BEa

[3597] Anonymous. Binary engines. Scientific American, 27(20):311, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n20/pdf/scientificamerican11161872-311d.pdf.

Anonymous:1872:SRT

[3598] Anonymous. Sunday railway trains. Scientific American, 27(20): 312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312.pdf.

Anonymous:1872:IWb

[3599] Anonymous. An improvement wanted. Scientific American, 27 (20):312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312a.pdf.

Anonymous:1872:CSf

[3600] Anonymous. The contradictions of science. Scientific American, 27(20):312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312b.pdf.

Anonymous:1872:SFR

[3601] Anonymous. The Saxby and Farmer railway signals. Scientific American, 27(20):312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312c.pdf.

Anonymous:1872:CPk

[3602] Anonymous. Corundum in Pennsylvania. Scientific American, 27 (20):312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312d.pdf.

Anonymous:1872:PMc

[3603] Anonymous. Petroleum in Massachusetts. Scientific American, 27 (20):312, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312e.pdf.

Anonymous:1872:DHJ

[3604] Anonymous. Death of Hon. John A. Griswold. Scientific American, 27(20):312–313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
312f.pdf.

Anonymous:1872:IITa

[3605] Anonymous. Important improvement in textile weaving. Scientific American, 27(20):313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313.pdf.

Anonymous:1872:BEb

[3606] Anonymous. The burning of the Esourial. Scientific American, 27(20):313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313a.pdf.

Anonymous:1872:IBb

[3607] Anonymous. Incrustation of boilers. Scientific American, 27 (20):313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313b.pdf.

Anonymous:1872:GSFb

[3608] Anonymous. The Georgia state fair. Scientific American, 27 (20):313, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313c.pdf.

Anonymous:1872:CPl

[3609] Anonymous. Copies of patents. Scientific American, 27(20): 313–314, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
313e.pdf.

Anonymous:1872:ACbv

[3610] Anonymous. Answers to correspondents. Scientific American, 27 (20):314, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
314a.pdf.

Anonymous:1872:RAFbt

[3611] Anonymous. Recent American and foreign patents. Scientific American, 27(20):314–315, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
314b.pdf.

Anonymous:1872:PHIj

[3612] Anonymous. Practical hints to inventors. Scientific American, 27 (20):316, November 16, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n20/pdf/scientificamerican11161872-
316.pdf.

Anonymous:1872:Cbv

[3613] Anonymous. Correspondence. Scientific American, 27(21):321– 323, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
321.pdf.

Anonymous:1872:BPbw

[3614] Anonymous. Business and personal. Scientific American, 27 (21):330, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
330.pdf.

Anonymous:1872:NQbt

[3615] Anonymous. Notes and queries. Scientific American, 27(21): 330, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
330a.pdf.

Anonymous:1872:IIbr

[3616] Anonymous. Index of inventions. Scientific American, 27(21): 331–332, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
331a.pdf.

Anonymous:1872:UBM

[3617] Anonymous. Upright boring machine. Scientific American, 27 (21):319, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
319.pdf.

Anonymous:1872:Ij

[3618] Anonymous. On instinct. Scientific American, 27(21):320, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n21/pdf/scientificamerican11231872-320.pdf.

Anonymous:1872:HML

[3619] Anonymous. Horsenail making in London. Scientific American, 27 (21):320–321, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
320a.pdf.

Anonymous:1872:IMH

[3620] Anonymous. Improved mop, head. Scientific American, 27(21): 323, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
323.pdf.

Anonymous:1872:AES

[3621] Anonymous. Apparatus for extinguishing sparks. Scientific American, 27(21):323–325, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
323a.pdf.

Anonymous:1872:ITL

[3622] Anonymous. Improved tank locomotive. Scientific American, 27 (21):326, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
326.pdf.

Anonymous:1872:ABD

[3623] Anonymous. Automatic boat detaching apparatus. Scientific American, 27(21):326, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
326a.pdf.

Anonymous:1872:NDC

[3624] Anonymous. New discoveries concerning terrestrial heat. Scientific American, 27(21):327, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
327.pdf.

Anonymous:1872:DPO

[3625] Anonymous. Drawings for the Patent Office. Scientific American, 27(21):327, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
327a.pdf.

Anonymous:1872:LIEb

[3626] Anonymous. Lake of interest in England for the Vienna Exposition. Scientific American, 27(21):327, November 23, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n21/pdf/scientificamerican11231872-327b.pdf.

Anonymous:1872:DHT

[3627] Anonymous. The determination of high temperatures by sound. Scientific American, 27(21):327–328, November 23, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n21/pdf/scientificamerican11231872-327c.pdf.

Anonymous:1872:COY

[3628] Anonymous. Choosing an occupation for a Young man. Scientific American, 27(21):328, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
328.pdf.

Anonymous:1872:LDP

[3629] Anonymous. The latest discoveries in the polar regions. Scientific American, 27(21):328, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
328a.pdf.

Anonymous:1872:IMS

[3630] Anonymous. Improvements in the manufacture of silk. Scientific American, 27(21):328–329, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
328b.pdf.

Anonymous:1872:ISd

[3631] Anonymous. Intolerance in science. Scientific American, 27(21): 329, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
329.pdf.

Anonymous:1872:SMPa

[3632] Anonymous. Scientific and mechanical possibilities. Scientific American, 27(21):329, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
329a.pdf.

Anonymous:1872:CBw

[3633] Anonymous. The conflagration in Boston. Scientific American, 27 (21):329–330, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
329b.pdf.

Anonymous:1872:ACbw

[3634] Anonymous. Answers to correspondents. Scientific American, 27 (21):330, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
330b.pdf.

Anonymous:1872:RAFbu

[3635] Anonymous. Recent American and foreign patents. Scientific American, 27(21):331, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
331.pdf.

Anonymous:1872:PHIk

[3636] Anonymous. Practical hints to inventors. Scientific American, 27 (21):332, November 23, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n21/pdf/scientificamerican11231872-
332.pdf.

Anonymous:1872:Cbx

[3637] Anonymous. Correspondence. Scientific American, 27(22):340, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n22/pdf/scientificamerican11301872-340.pdf.

Anonymous:1872:BNBbj

[3638] Anonymous. Books: New books and publications. Scientific American, 27(22):345, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
345a.pdf.

Anonymous:1872:BPbx

[3639] Anonymous. Business and personal. Scientific American, 27 (22):345, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
345b.pdf.

Anonymous:1872:NQbu

[3640] Anonymous. Notes and queries. Scientific American, 27(22): 345, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
345c.pdf.

Anonymous:1872:IIbs

[3641] Anonymous. Index of inventions. Scientific American, 27(22): 347, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
347.pdf.

Anonymous:1872:CMb

[3642] Anonymous. Cigar making. Scientific American, 27(22):335, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n22/pdf/scientificamerican11301872-335.pdf.

Anonymous:1872:ICE

[3643] Anonymous. The institution of civil engineers. Scientific American, 27(22):336, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
336.pdf.

Pile:1872:NAT

[3644] Wilson H. Pile. A new application of tube hydrometers. Scientific American, 27(22):336, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
336a.pdf.

Anonymous:1872:RSAb

[3645] Anonymous. Remarkable structures of the ancient Americans. Scientific American, 27(22):337, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
337.pdf.

Anonymous:1872:ADW

[3646] Anonymous. The agricultural department at Washington. Scientific American, 27(22):338, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
338.pdf.

Anonymous:1872:TMf

[3647] Anonymous. Transmission of motion. Scientific American, 27 (22):339, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
339.pdf.

Robinson:1872:ACT

[3648] W. T. Robinson and A. M. Absurd cosmical theories. Scientific American, 27(22):340–341, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
340a.pdf.

Anonymous:1872:F

[3649] Anonymous. Fire. Scientific American, 27(22):341, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n22/pdf/scientificamerican11301872-341.pdf.

Anonymous:1872:HSS

[3650] Anonymous. Hay and straw stacking-apparatus. Scientific American, 27(22):342, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
342.pdf.

Anonymous:1872:SMTc

[3651] Anonymous. Sewing machine treadle and casters. Scientific American, 27(22):342, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
342a.pdf.

Anonymous:1872:VEc

[3652] Anonymous. The Vienna Exposition. Scientific American, 27 (22):343, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
343.pdf.

Anonymous:1872:NS

[3653] Anonymous. The nose straightener. Scientific American, 27(22): 343, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
343a.pdf.

Anonymous:1872:RSd

[3654] Anonymous. Remarkable stupidity. Scientific American, 27(22): 343, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
343b.pdf.

Anonymous:1872:HDb

[3655] Anonymous. The horse disease. Scientific American, 27(22): 343, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
343c.pdf.

Anonymous:1872:OI

[3656] Anonymous. Oxyhydric illumination. Scientific American, 27(22): 343–344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
343d.pdf.

Anonymous:1872:OCD

[3657] Anonymous. Our concrete docks. Scientific American, 27(22): 344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
344.pdf.

Anonymous:1872:AGL

[3658] Anonymous. Air gas light in England. Scientific American, 27 (22):344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
344a.pdf.

Anonymous:1872:GES

[3659] Anonymous. The growth or evolution of structure in seedlings. Scientific American, 27(22):344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
344b.pdf.

Anonymous:1872:PAIb

[3660] Anonymous. Progress of American iron industry. Scientific American, 27(22):344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
344c.pdf.

Anonymous:1872:PDc

[3661] Anonymous. Patent decision. Scientific American, 27(22):344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n22/pdf/scientificamerican11301872-344d.pdf.

Anonymous:1872:PJHb

[3662] Anonymous. Professor James Hadley. Scientific American, 27 (22):344, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
344e.pdf.

Anonymous:1872:SPIbe

[3663] Anonymous. Scientific and practical information. Scientific American, 27(22):345, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
345.pdf.

Anonymous:1872:ACbx

[3664] Anonymous. Answers to correspondents. Scientific American, 27 (22):345–346, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
345d.pdf.

Anonymous:1872:CRd

[3665] Anonymous. Communications received. Scientific American, 27 (22):346, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
346.pdf.

Anonymous:1872:RAFbv

[3666] Anonymous. Recent American and foreign patents. Scientific American, 27(22):346–347, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
346a.pdf.

Anonymous:1872:VPHa

[3667] Anonymous. Value of patents, and how to obtain them. Scientific American, 27(22):348, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
348.pdf.

Anonymous:1872:PCRb

[3668] Anonymous. Patent cold rolled shafting. Scientific American, 27 (22):350, November 30, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n22/pdf/scientificamerican11301872-
350.pdf.

Anonymous:1872:Cby

[3669] Anonymous. Correspondence. Scientific American, 27(23):355–356, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-355a.pdf.

Anonymous:1872:BNBbk

[3670] Anonymous. Books: New books and publications. Scientific American, 27(23):361, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
361.pdf.

Anonymous:1872:BPby

[3671] Anonymous. Business and personal. Scientific American, 27 (23):361, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
361a.pdf.

Anonymous:1872:NQbv

[3672] Anonymous. Notes and queries. Scientific American, 27(23):362, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-362.pdf.

Anonymous:1872:IIbt

[3673] Anonymous. Index of inventions. Scientific American, 27(23): 363–364, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
363.pdf.

Anonymous:1872:ISS

[3674] Anonymous. Improved steam street cars. Scientific American, 27 (23):351, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
351.pdf.

Anonymous:1872:OFH

[3675] Anonymous. Our five hundred thousand dollar Commissioner to Vienna. Scientific American, 27(23):352, December 7, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n23/pdf/scientificamerican12071872-352.pdf.

Anonymous:1872:ANa

[3676] Anonymous. Astronomical notes. Scientific American, 27(23):352, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-352a.pdf.

Anonymous:1872:RTPb

[3677] Anonymous. A remarkable test plate. Scientific American, 27 (23):352, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
352b.pdf.

Anonymous:1872:IIPa

[3678] Anonymous. Italian industrial progress. Scientific American, 27 (23):352, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
352c.pdf.

Anonymous:1872:FAIe

[3679] Anonymous. The fair of the American institute. Scientific American, 27(23):352–353, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
352d.pdf.

Anonymous:1872:NPLp

[3680] Anonymous. New plan for a local fire department. Scientific American, 27(23):353, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
353.pdf.

Anonymous:1872:SMPb

[3681] Anonymous. Scientific and mechanical possibilities. Scientific American, 27(23):353, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
353a.pdf.

Anonymous:1872:RMS

[3682] Anonymous. Remarkable magnetic storms and auroras in Europe. Scientific American, 27(23):353, December 7, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n23/pdf/scientificamerican12071872-353b.pdf.

Anonymous:1872:STd

[3683] Anonymous. Steam traction. Scientific American, 27(23):353, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-353c.pdf.

Anonymous:1872:SPIbf

[3684] Anonymous. Scientific and practical information. Scientific American, 27(23):353, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
353d.pdf.

Anonymous:1872:ISEd

[3685] Anonymous. Improved steam engine. Scientific American, 27 (23):354, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
354.pdf.

Anonymous:1872:LCG

[3686] Anonymous. Lathe center grinder. Scientific American, 27(23):354, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-354a.pdf.

Anonymous:1872:TMg

[3687] Anonymous. Transmission of motion. Scientific American, 27 (23):355, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
355.pdf.

Anonymous:1872:ANb

[3688] Anonymous. Astronomical notes. Scientific American, 27(23):356, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n23/pdf/scientificamerican12071872-356.pdf.

Anonymous:1872:ISPa

[3689] Anonymous. Inventions and suggestions for the prevention of fires. Scientific American, 27(23):357, December 7, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n23/pdf/scientificamerican12071872-357.pdf.

Anonymous:1872:CDa

[3690] Anonymous. The consciousness of dogs. Scientific American, 27 (23):357, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
357a.pdf.

Anonymous:1872:HMMa

[3691] Anonymous. How money is made. Scientific American, 27(23): 358–360, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
358.pdf.

Anonymous:1872:DCPb

[3692] Anonymous. Decisions by the Commissioner of Patents. Scientific American, 27(23):360, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
360.pdf.

Anonymous:1872:DCd

[3693] Anonymous. Decisions of the courts. Scientific American, 27(23): 360–361, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
360a.pdf.

Anonymous:1872:ACby

[3694] Anonymous. Answers to correspondents. Scientific American, 27 (23):362, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
362a.pdf.

Anonymous:1872:RAFbw

[3695] Anonymous. Recent American and foreign patents. Scientific American, 27(23):362–363, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
362b.pdf.

Anonymous:1872:VPHb

[3696] Anonymous. Value of patents, and how to obtain them. Scientific American, 27(23):364, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
364.pdf.

Anonymous:1872:PCRc

[3697] Anonymous. Patent cold rolled shafting. Scientific American, 27 (23):366, December 7, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n23/pdf/scientificamerican12071872-
366.pdf.

Anonymous:1872:Cbz

[3698] Anonymous. Correspondence. Scientific American, 27(24):372, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n24/pdf/scientificamerican12141872-372.pdf.

Anonymous:1872:PNPc

[3699] Anonymous. Patent notes: Patents. Scientific American, 27 (24):373, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
373.pdf.

Anonymous:1872:BPbz

[3700] Anonymous. Business and personal. Scientific American, 27 (24):378, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
378.pdf.

Anonymous:1872:NQbw

[3701] Anonymous. Notes and queries. Scientific American, 27(24): 378, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
378a.pdf.

Anonymous:1872:IIbu

[3702] Anonymous. Index of inventions. Scientific American, 27(24): 379–380, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
379.pdf.

Anonymous:1872:TIS

[3703] Anonymous. Twenty inch smooth bore gun for the Russian Government. Scientific American, 27(24):367, December 14, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n24/pdf/scientificamerican12141872-367.pdf.

Anonymous:1872:MCc

[3704] Anonymous. Marine casualties. Scientific American, 27(24): 367, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
367a.pdf.

Anonymous:1872:PHM

[3705] Anonymous. The proposed half million grab. Scientific American, 27(24):368, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
368.pdf.

Anonymous:1872:RAS

[3706] Anonymous. Revival of American shipbuilding. Scientific American, 27(24):368, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
368a.pdf.

Anonymous:1872:CDB

[3707] Anonymous. The California diamond bubble. Scientific American, 27(24):368, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
368b.pdf.

Anonymous:1872:OPA

[3708] Anonymous. Official publication of the American patents — a new and important work. Scientific American, 27(24):368–369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n24/pdf/scientificamerican12141872-368c.pdf.

Anonymous:1872:ISCc

[3709] Anonymous. Improved Ship’s compass. Scientific American, 27 (24):369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
369.pdf.

Anonymous:1872:NAW

[3710] Anonymous. The November atmospheric wave. Scientific American, 27(24):369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
369a.pdf.

Anonymous:1872:DHG

[3711] Anonymous. Death of Horace Greeley. Scientific American, 27 (24):369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
369b.pdf.

Anonymous:1872:ESI

[3712] Anonymous. Electricity at the Stevens Institute. — novel researches by Professor Morton concerning the induced current. Scientific American, 27(24):369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
369c.pdf.

Anonymous:1872:CLa

[3713] Anonymous. The Christian leader. Scientific American, 27(24): 369, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
369d.pdf.

Anonymous:1872:NIH

[3714] Anonymous. New ice harvesting invention. Scientific American, 27(24):370, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
370.pdf.

Anonymous:1872:ISPb

[3715] Anonymous. Improved seed planter. Scientific American, 27 (24):370, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
370a.pdf.

Anonymous:1872:HMP

[3716] Anonymous. How to make paper caps. Scientific American, 27 (24):370, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
370b.pdf.

Anonymous:1872:SBE

[3717] Anonymous. Steam boiler explosions. Scientific American, 27 (24):370, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
370c.pdf.

Anonymous:1872:CTa

[3718] Anonymous. Combined tool. Scientific American, 27(24):370, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n24/pdf/scientificamerican12141872-370d.pdf.

Anonymous:1872:TMh

[3719] Anonymous. Transmission of motion. Scientific American, 27 (24):371, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
371.pdf.

Anonymous:1872:HBSc

[3720] Anonymous. Head block for saw mills. Scientific American, 27 (24):371, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
371a.pdf.

Anonymous:1872:SBU

[3721] Anonymous. The Sturtevant blower and its uses. Scientific American, 27(24):372, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
372a.pdf.

Anonymous:1872:IHBb

[3722] Anonymous. Improved hot blast oven. Scientific American, 27 (24):375, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
375.pdf.

Anonymous:1872:HVS

[3723] Anonymous. How a vessel sails against the wind. Scientific American, 27(24):375–376, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
375a.pdf.

Anonymous:1872:BVEc

[3724] Anonymous. The buildings of the Vienna Exposition. Scientific American, 27(24):376, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
376.pdf.

Anonymous:1872:HMMb

[3725] Anonymous. How to make money by patents. Scientific American, 27(24):376–377, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
376a.pdf.

Anonymous:1872:NFC

[3726] Anonymous. New fireproof construction. Scientific American, 27 (24):377, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
377.pdf.

Anonymous:1872:PODb

[3727] Anonymous. Patent Office decisions. Scientific American, 27 (24):377, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
377a.pdf.

Anonymous:1872:DCe

[3728] Anonymous. Decisions of the courts. Scientific American, 27(24): 377–378, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
377b.pdf.

Anonymous:1872:ACbz

[3729] Anonymous. Answers to correspondents. Scientific American, 27 (24):378, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
378b.pdf.

Anonymous:1872:RAFbx

[3730] Anonymous. Recent American and foreign patents. Scientific American, 27(24):378–379, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
378c.pdf.

Anonymous:1872:VPHc

[3731] Anonymous. Value of patents, and how to obtain them. Scientific American, 27(24):380, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
380.pdf.

Anonymous:1872:PCRd

[3732] Anonymous. Patent cold rolled shafting. Scientific American, 27 (24):382, December 14, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n24/pdf/scientificamerican12141872-
382.pdf.

Anonymous:1872:Cce

[3733] Anonymous. Correspondence. Scientific American, 27(25):388–389, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:PNPd

[3734] Anonymous. Patent notes: Patents. Scientific American, 27(25):392, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:BNBbl

[3735] Anonymous. Books: New books and publications. Scientific American, 27(25):393, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:BPca

[3736] Anonymous. Business and personal. Scientific American, 27(25):393, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:NQbx

[3737] Anonymous. Notes and queries. Scientific American, 27(25):393–394, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:IIbv

[3738] Anonymous. Index of inventions. Scientific American, 27(25): 395–396, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
395.pdf.

Anonymous:1872:SEE

[3739] Anonymous. Steam engine economy and steam boiler economics. Scientific American, 27(25):384, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:PMV

[3740] Anonymous. The President’s message and the Vienna Exposition. Scientific American, 27(25):384, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:CSR

[3741] Anonymous. Concentrated solar rays. Scientific American, 27(25):384, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:BCB

[3742] Anonymous. A badly constructed boiler. Scientific American, 27(25): 384, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:VEP

[3743] Anonymous. The Vienna Exposition — Professor Thurston explains. Scientific American, 27(25):384–385, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:TWb

[3744] Anonymous. Torpedo warfare. Scientific American, 27(25):385, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:RBS

[3745] Anonymous. On the resistance of buildings stones to fire. Scientific American, 27(25):385, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:GVE

[3746] Anonymous. Germany in the Vienna Exposition. Scientific American, 27(25):385, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:HG

[3747] Anonymous. Horace Greeley. Scientific American, 27(25):385, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:NBH

[3748] Anonymous. New Burlington House, Piccadilly, London. Scientific American, 27(25):386, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:NMB

[3749] Anonymous. New method of brewing. Scientific American, 27(25): 387, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:BHC

[3750] Anonymous. Button hole cutter. Scientific American, 27(25):387, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:PJb

[3751] Anonymous. Patents in Japan. Scientific American, 27(25):389, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:PPPb

[3752] Anonymous. A patent poisonous pavement. Scientific American, 27 (25):389, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:KSC

[3753] Anonymous. Key seat cutting machine. Scientific American, 27(25): 390, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:ATB

[3754] Anonymous. Another terrible boiler explosion in Ohio. Scientific American, 27(25):390, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:IDT

[3755] Anonymous. Improved drawing table. Scientific American, 27(25):390, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:NPR

[3756] Anonymous. The new Prussian rifle. Scientific American, 27(25):391, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:NRF

[3757] Anonymous. A novel railway ferry boat. Scientific American, 27(25): 391, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:PS

[3758] Anonymous. Phineas L. Slayton. Scientific American, 27(25):392, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:DCg

[3759] Anonymous. Decisions of the courts. Scientific American, 27(25):393, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1872:ACcb

[3760] Anonymous. Answers to correspondents. Scientific American, 27 (25):394, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
394.pdf.

Anonymous:1872:CRe

[3761] Anonymous. Communications received. Scientific American, 27 (25):394, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
394a.pdf.

Anonymous:1872:RAFby

[3762] Anonymous. Recent American and foreign patents. Scientific American, 27(25):394–395, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
394b.pdf.

Anonymous:1872:VPHd

[3763] Anonymous. Value of patents, and how to obtain them. Scientific American, 27(25):396, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
396.pdf.

Anonymous:1872:PCRe

[3764] Anonymous. Patent cold rolled shafting. Scientific American, 27 (25):398, December 21, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n25/pdf/scientificamerican12211872-
398.pdf.

Anonymous:1872:Ccf

[3765] Anonymous. Correspondence. Scientific American, 27(26):404, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n26/pdf/scientificamerican12281872-404.pdf.

Anonymous:1872:BPcb

[3766] Anonymous. Business and personal. Scientific American, 27 (26):408, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
408c.pdf.

Anonymous:1872:IIbw

[3767] Anonymous. Index of inventions. Scientific American, 27(26): 409–410, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
409a.pdf.

Anonymous:1872:Ik

[3768] Anonymous. Index. Scientific American, 27(26):410–412, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n26/pdf/scientificamerican12281872-410a.pdf.

Anonymous:1872:BPEa

[3769] Anonymous. The Brooklyn Pier of the East River Bridge. Scientific American, 27(26):399, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
399.pdf.

Anonymous:1872:EAY

[3770] Anonymous. The end of another year. Scientific American, 27 (26):400, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
400.pdf.

Anonymous:1872:PCP

[3771] Anonymous. Proposed change in the patent law. Scientific American, 27(26):400, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
400a.pdf.

Anonymous:1872:TWP

[3772] Anonymous. Tidal water power; a new and useful suggestion. Scientific American, 27(26):400, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
400b.pdf.

Anonymous:1872:PJR

[3773] Anonymous. The Planet Jupiter as revealed by the modern spectroscope, photometer, and telescope. Scientific American, 27(26): 400–401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
400c.pdf.

Anonymous:1872:PHTb

[3774] Anonymous. Progress of the Hoosic Tunnel. Scientific American, 27(26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401.pdf.

Anonymous:1872:FFA

[3775] Anonymous. The fire at the fifth avenue hotel. Scientific American, 27(26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401a.pdf.

Anonymous:1872:RJN

[3776] Anonymous. Retirement of judge Nelson. Scientific American, 27 (26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401b.pdf.

Anonymous:1872:Ccg

[3777] Anonymous. The Centennial of 1876. Scientific American, 27 (26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401c.pdf.

Anonymous:1872:AIEb

[3778] Anonymous. American inventions in Europe. Scientific American, 27(26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401d.pdf.

Anonymous:1872:SRd

[3779] Anonymous. Science record for 1873. Scientific American, 27 (26):401, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
401e.pdf.

Anonymous:1872:MMO

[3780] Anonymous. Mallet’s method for oxygen. Scientific American, 27 (26):402, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
402.pdf.

Anonymous:1872:ACFa

[3781] Anonymous. The application of chemical fertilizers to horticulture. Scientific American, 27(26):402, December 28, 1872. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v27/
n26/pdf/scientificamerican12281872-402a.pdf.

Anonymous:1872:E

[3782] Anonymous. Electricity. Scientific American, 27(26):403, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v27/n26/pdf/scientificamerican12281872-403.pdf.

Anonymous:1872:SPIbg

[3783] Anonymous. Scientific and practical information. Scientific American, 27(26):405, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
405.pdf.

Anonymous:1872:ISPc

[3784] Anonymous. Improved safe protector. Scientific American, 27 (26):406, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
406.pdf.

Anonymous:1872:PWB

[3785] Anonymous. Putt’s wash boiler. Scientific American, 27(26): 406, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
406a.pdf.

Anonymous:1872:AJS

[3786] Anonymous. Amateur’s jig saw. Scientific American, 27(26): 407, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
407.pdf.

Anonymous:1872:GT

[3787] Anonymous. Government telegraphy. Scientific American, 27 (26):407, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
407a.pdf.

Anonymous:1872:IRSc

[3788] Anonymous. Improved railroad Spike. Scientific American, 27 (26):407, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
407b.pdf.

Anonymous:1872:DGC

[3789] Anonymous. Dayton’s gas carburetter. Scientific American, 27 (26):407, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
407c.pdf.

Anonymous:1872:PODc

[3790] Anonymous. Patent Office decisions. Scientific American, 27 (26):408, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
408.pdf.

Anonymous:1872:DCh

[3791] Anonymous. Decisions of the courts. Scientific American, 27 (26):408, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
408a.pdf.

Anonymous:1872:CRf

[3792] Anonymous. Communications received. Scientific American, 27 (26):408, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
408b.pdf.

Anonymous:1872:ACcc

[3793] Anonymous. Answers to correspondents. Scientific American, 27 (26):408, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
408d.pdf.

Anonymous:1872:RAFbz

[3794] Anonymous. Recent American and foreign patents. Scientific American, 27(26):409, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
409.pdf.

Anonymous:1872:VPHe

[3795] Anonymous. Value of patents and how to obtain them. Scientific American, 27(26):410, December 28, 1872. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v27/n26/pdf/scientificamerican12281872-
410.pdf.

Anonymous:1873:Ca

[3796] Anonymous. Correspondence. Scientific American, 28(1):4, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-4.pdf.

Anonymous:1873:BPa

[3797] Anonymous. Business and personal. Scientific American, 28(1):9, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-9b.pdf.

Anonymous:1873:NQa

[3798] Anonymous. Notes and queries. Scientific American, 28(1):10, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-10.pdf.

Anonymous:1873:IIa

[3799] Anonymous. Index of inventions. Scientific American, 28(1):11–12, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-11.pdf.

Anonymous:1873:GTa

[3800] Anonymous. Great telescopes. Scientific American, 28(1):1, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-1.pdf.

Anonymous:1873:NBM

[3801] Anonymous. New buildings in Madras for the Revenue Board offices. Scientific American, 28(1):2, January 4, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n1/pdf/scientificamerican01041873-2.pdf.

Anonymous:1873:PTLa

[3802] Anonymous. Professor Tyndall on light. Scientific American, 28(1):3, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-3.pdf.

Anonymous:1873:FGW

[3803] Anonymous. Fresh grapes in winter. Scientific American, 28(1):3, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-3a.pdf.

Anonymous:1873:ETM

[3804] Anonymous. Early traces of men. Scientific American, 28(1):5, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-5.pdf.

Anonymous:1873:BAS

[3805] Anonymous. Bessemer’s anti-sea-sick ship. Scientific American, 28(1):5, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-5a.pdf.

Anonymous:1873:PWB

[3806] Anonymous. Prime’s wash boiler. Scientific American, 28(1):6, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-6.pdf.

Anonymous:1873:BIB

[3807] Anonymous. Brown’s improved beehive. Scientific American, 28(1):6, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-6a.pdf.

Anonymous:1873:ON

[3808] Anonymous. Our Navy. Scientific American, 28(1):7, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-7.pdf.

Anonymous:1873:ME

[3809] Anonymous. The Moscow Exposition. Scientific American, 28(1):7, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-7a.pdf.

Anonymous:1873:SEW

[3810] Anonymous. A slipped eccentric and what came of it. Scientific American, 28(1):7, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n1/pdf/scientificamerican01041873-
7b.pdf.

Anonymous:1873:PTM

[3811] Anonymous. Professor Tyndall as a money maker. Scientific American, 28(1):7, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n1/pdf/scientificamerican01041873-
7c.pdf.

Anonymous:1873:VSC

[3812] Anonymous. The Vienna show in Congress. Scientific American, 28(1):7–8, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n1/pdf/scientificamerican01041873-
7d.pdf.

Anonymous:1873:SRTa

[3813] Anonymous. Sunday railway trains. Scientific American, 28(1):8, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-8.pdf.

Anonymous:1873:NY

[3814] Anonymous. The new year. Scientific American, 28(1):8, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-8a.pdf.

Anonymous:1873:RTNa

[3815] Anonymous. Rapid transit in New York. Scientific American, 28(1):8–9, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-8b.pdf.

Anonymous:1873:PODa

[3816] Anonymous. Patent Office decisions. Scientific American, 28(1):9, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-9.pdf.

Anonymous:1873:DCa

[3817] Anonymous. Decisions of the courts. Scientific American, 28(1):9, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-9a.pdf.

Anonymous:1873:ACa

[3818] Anonymous. Answers to correspondents. Scientific American, 28(1):10, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-10a.pdf.

Anonymous:1873:CRa

[3819] Anonymous. Communications received. Scientific American, 28(1):10, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n1/pdf/scientificamerican01041873-10b.pdf.

Anonymous:1873:RAFa

[3820] Anonymous. Recent American and foreign patents. Scientific American, 28(1):10–11, January 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n1/pdf/scientificamerican01041873-
10c.pdf.

Anonymous:1873:Cb

[3821] Anonymous. Correspondence. Scientific American, 28(2):20–21, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-20.pdf.

Anonymous:1873:BNBa

[3822] Anonymous. Books: New books and publications. Scientific American, 28(2):25, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
25b.pdf.

Anonymous:1873:BPb

[3823] Anonymous. Business and personal. Scientific American, 28(2):27, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-27.pdf.

Anonymous:1873:NQb

[3824] Anonymous. Notes and queries. Scientific American, 28(2):27, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-27a.pdf.

Anonymous:1873:IIb

[3825] Anonymous. Index of inventions. Scientific American, 28(2):27–28, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-27c.pdf.

Anonymous:1873:DAS

[3826] Anonymous. Direct acting steam and hydrostatic press. Scientific American, 28(2):15, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
15.pdf.

Anonymous:1873:LPS

[3827] Anonymous. Losses of power in the steam engine. where it may be improved and to what extent. Scientific American, 28(2):16, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-16.pdf.

Anonymous:1873:GTb

[3828] Anonymous. Government telegraphy. Scientific American, 28 (2):16, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
16a.pdf.

Anonymous:1873:NYI

[3829] Anonymous. The New York Industrial Exhibition. Scientific American, 28(2):16, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
16b.pdf.

Anonymous:1873:CF

[3830] Anonymous. Corn as fuel. Scientific American, 28(2):16, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-16c.pdf.

Anonymous:1873:HIP

[3831] Anonymous. The Henderson iron process. Scientific American, 28(2):16–17, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
16d.pdf.

Anonymous:1873:FJ

[3832] Anonymous. Friction of journals. Scientific American, 28(2):17, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-17.pdf.

Anonymous:1873:PBC

[3833] Anonymous. Patent business in Congress. Scientific American, 28(2):17, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
17a.pdf.

Anonymous:1873:ATC

[3834] Anonymous. Art treasures from Cyprus — English criticism on their destination. Scientific American, 28(2):17, January 11, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n2/pdf/scientificamerican01111873-17b.pdf.

Anonymous:1873:ASS

[3835] Anonymous. The anti-sea-sick ship. Scientific American, 28(2):18, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-18.pdf.

Anonymous:1873:NTA

[3836] Anonymous. New theory of the Aurora. Scientific American, 28(2):18, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
18a.pdf.

Anonymous:1873:PTLb

[3837] Anonymous. Professor Tyndall on light — second lecture. Scientific American, 28(2):19–20, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
19.pdf.

Anonymous:1873:LEE

[3838] Anonymous. The Lyons Exposition Elevated Railway. Scientific American, 28(2):22, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
22.pdf.

Anonymous:1873:IBF

[3839] Anonymous. Improved blast furnace. Scientific American, 28 (2):22, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
22a.pdf.

Anonymous:1873:GLW

[3840] Anonymous. The great laxey water wheel. Scientific American, 28(2):23, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
23.pdf.

Anonymous:1873:PCb

[3841] Anonymous. The progress of chemistry in 1872. Scientific American, 28(2):24, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
24.pdf.

Anonymous:1873:GFE

[3842] Anonymous. A gigantic fire engine. Scientific American, 28(2):24, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n2/pdf/scientificamerican01111873-24a.pdf.

Anonymous:1873:SPIa

[3843] Anonymous. Scientific and practical information. Scientific American, 28(2):25, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
25.pdf.

Anonymous:1873:RPD

[3844] Anonymous. Recent patent decisions. Scientific American, 28 (2):25, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
25a.pdf.

Anonymous:1873:RAFb

[3845] Anonymous. Recent American and foreign patents. Scientific American, 28(2):25–26, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
25c.pdf.

Anonymous:1873:ACb

[3846] Anonymous. Answers to correspondents. Scientific American, 28 (2):27, January 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n2/pdf/scientificamerican01111873-
27b.pdf.

Anonymous:1873:Cc

[3847] Anonymous. Correspondence. Scientific American, 28(3):36, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-36.pdf.

Anonymous:1873:BPc

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-43.pdf.

Anonymous:1873:IIc

[3849] Anonymous. Index of inventions. Scientific American, 28(3):43–44, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-43b.pdf.

Anonymous:1873:IBMa

[3850] Anonymous. Improved brick machine. Scientific American, 28 (3):31, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
31.pdf.

Anonymous:1873:CBS

[3851] Anonymous. Can buildings be set on fire by steam pipes? Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
32.pdf.

Anonymous:1873:ASB

[3852] Anonymous. The Austrian show before the Senate. Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
32a.pdf.

Anonymous:1873:CBF

[3853] Anonymous. Collapsing boiler flues. Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-32b.pdf.

Anonymous:1873:Pa

[3854] Anonymous. Prospects for 1873. Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-32c.pdf.

Anonymous:1873:DMP

[3855] Anonymous. Du Motay’s process for oxyhydric illumination. Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
32d.pdf.

Anonymous:1873:OA

[3856] Anonymous. Oxidizing agents. Scientific American, 28(3):32, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-32e.pdf.

Anonymous:1873:PUK

[3857] Anonymous. Properties and uses of kieserite. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
33.pdf.

Anonymous:1873:BMa

[3858] Anonymous. Balancing machinery. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-33a.pdf.

Anonymous:1873:SMMa

[3859] Anonymous. The sewing machine monopoly. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
33b.pdf.

Anonymous:1873:IEV

[3860] Anonymous. To intending exhibitors at the Vienna show. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
33c.pdf.

Anonymous:1873:PAEa

[3861] Anonymous. Practical advice to exhibitors in Austria. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
33d.pdf.

Anonymous:1873:ISSa

[3862] Anonymous. An international show in Spain. Scientific American, 28(3):33, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
33e.pdf.

Anonymous:1873:IFP

[3863] Anonymous. Improved feathering paddle wheel. Scientific American, 28(3):34, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
34.pdf.

Anonymous:1873:TCSa

[3864] Anonymous. Turton’s compensated Spring safety valve. Scientific American, 28(3):34, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
34a.pdf.

Anonymous:1873:PTT

[3865] Anonymous. Professor Tyndall’s third lecture in New York crystals and light. Scientific American, 28(3):35, January 18, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n3/pdf/scientificamerican01181873-35.pdf.

Anonymous:1873:GSAa

[3866] Anonymous. Gleanings in science and art. Scientific American, 28(3):37, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
37.pdf.

Anonymous:1873:NHPa

[3867] Anonymous. Nine horse power compound boat engine. Scientific American, 28(3):38, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
38.pdf.

Anonymous:1873:DBG

[3868] Anonymous. Design for ball ground pavilion. Scientific American, 28(3):39, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
39.pdf.

Anonymous:1873:SPIb

[3869] Anonymous. Scientific and practical information. Scientific American, 28(3):40, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
40.pdf.

Porter:1873:BWC

[3870] Henry F. Porter. Bent woodwork in carriage making. Scientific American, 28(3):40–41, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
40a.pdf.

Anonymous:1873:PODb

[3871] Anonymous. Patent Office decisions. Scientific American, 28(3):41, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n3/pdf/scientificamerican01181873-41.pdf.

Anonymous:1873:DCb

[3872] Anonymous. Decisions of the courts. Scientific American, 28 (3):41–42, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
41a.pdf.

Anonymous:1873:RAFc

[3873] Anonymous. Recent American and foreign patents. Scientific American, 28(3):42–43, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
42.pdf.

Anonymous:1873:ACc

[3874] Anonymous. Answers correspondents. Scientific American, 28 (3):43, January 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n3/pdf/scientificamerican01181873-
43a.pdf.

Anonymous:1873:Cd

[3875] Anonymous. Correspondence. Scientific American, 28(4):52–53, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-52.pdf.

Anonymous:1873:BNBb

[3876] Anonymous. Books: New books and publications. Scientific American, 28(4):58, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
58a.pdf.

Anonymous:1873:BPd

[3877] Anonymous. Business and personal. Scientific American, 28(4):59, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-59.pdf.

Anonymous:1873:NQc

[3878] Anonymous. Notes and queries. Scientific American, 28(4):59, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-59a.pdf.

Anonymous:1873:IId

[3879] Anonymous. Index of inventions. Scientific American, 28(4):60, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-60a.pdf.

Anonymous:1873:CMC

[3880] Anonymous. Compound metallic columns and connections. Scientific American, 28(4):47, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
47.pdf.

Anonymous:1873:HTV

[3881] Anonymous. How to test for vanadium, and what we know of the metal. Scientific American, 28(4):47–48, January 25, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n4/pdf/scientificamerican01251873-47a.pdf.

Anonymous:1873:PSRa

[3882] Anonymous. The promotion of scientific research. Scientific American, 28(4):48, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
48.pdf.

Anonymous:1873:GPA

[3883] Anonymous. Government pap for Austria. Scientific American, 28(4):48, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
48a.pdf.

Anonymous:1873:TRI

[3884] Anonymous. The temperature required for ignition. Scientific American, 28(4):48, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
48b.pdf.

Anonymous:1873:MI

[3885] Anonymous. The mystery of the injector. Scientific American, 28(4):48–49, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
48c.pdf.

Anonymous:1873:GSM

[3886] Anonymous. The great sewing machine job before Congress. Scientific American, 28(4):49, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
49.pdf.

Anonymous:1873:UAC

[3887] Anonymous. The use of arsenic colors. Scientific American, 28 (4):49, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
49a.pdf.

Anonymous:1873:SD

[3888] Anonymous. Submarine diving. Scientific American, 28(4):50–51, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-50.pdf.

Anonymous:1873:PTFa

[3889] Anonymous. Professor Tyndall’s fourth lecture in New York — polarized light. Scientific American, 28(4):51, January 25, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n4/pdf/scientificamerican01251873-51.pdf.

Anonymous:1873:GSAb

[3890] Anonymous. Gleanings in science and art. Scientific American, 28(4):53, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
53.pdf.

Anonymous:1873:B

[3891] Anonymous. Bridle. Scientific American, 28(4):54, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-54.pdf.

Anonymous:1873:ACU

[3892] Anonymous. Animal charcoal and its use in the refining of sugar. Scientific American, 28(4):54, January 25, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n4/pdf/scientificamerican01251873-54a.pdf.

Anonymous:1873:VOC

[3893] Anonymous. Valved oil can nozzle. Scientific American, 28(4):54, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-54b.pdf.

Anonymous:1873:SCa

[3894] Anonymous. Saw clamp. Scientific American, 28(4):54, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-54c.pdf.

Anonymous:1873:NTU

[3895] Anonymous. New tunnel under the Harlem River, New York City. Scientific American, 28(4):55, January 25, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n4/pdf/scientificamerican01251873-55.pdf.

Anonymous:1873:TCEa

[3896] Anonymous. Three cylinder engine. Scientific American, 28(4):55, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-55a.pdf.

Anonymous:1873:GSWa

[3897] Anonymous. The government submarine works at Hell Gate. Scientific American, 28(4):56, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
56.pdf.

Anonymous:1873:ENa

[3898] Anonymous. Engineering notes. Scientific American, 28(4):56, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n4/pdf/scientificamerican01251873-56a.pdf.

Anonymous:1873:NEF

[3899] Anonymous. Nascent elements as fertilizers. Scientific American, 28(4):56, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
56b.pdf.

Anonymous:1873:SPIc

[3900] Anonymous. Scientific and practical information. Scientific American, 28(4):57, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
57.pdf.

Anonymous:1873:SMP

[3901] Anonymous. Scientific and mechanical possibilities. Scientific American, 28(4):57, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
57a.pdf.

Anonymous:1873:TSH

[3902] Anonymous. A typo-statue to Horace Greeley. Scientific American, 28(4):57, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
57b.pdf.

Anonymous:1873:PODc

[3903] Anonymous. Patent Office decisions. Scientific American, 28 (4):57–58, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
57c.pdf.

Anonymous:1873:RAFd

[3904] Anonymous. Recent American and foreign patents. Scientific American, 28(4):58, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
58.pdf.

Anonymous:1873:ACd

[3905] Anonymous. Answers correspondents. Scientific American, 28 (4):59–60, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
59b.pdf.

Anonymous:1873:CRb

[3906] Anonymous. Communications received. Scientific American, 28 (4):60, January 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n4/pdf/scientificamerican01251873-
60.pdf.

Anonymous:1873:Ce

[3907] Anonymous. Correspondence. Scientific American, 28(5):68, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-68.pdf.

Anonymous:1873:BNBc

[3908] Anonymous. Books: New books and publications. Scientific American, 28(5):73, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
73a.pdf.

Anonymous:1873:BPe

[3909] Anonymous. Business and personal. Scientific American, 28 (5):74, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
74.pdf.

Anonymous:1873:NQd

[3910] Anonymous. Notes and queries. Scientific American, 28(5):74, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-74a.pdf.

Anonymous:1873:IHT

[3911] Anonymous. Improved horizontal testing machine. Scientific American, 28(5):63, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
63.pdf.

Anonymous:1873:LV

[3912] Anonymous. Lathe vise. Scientific American, 28(5):63, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-63a.pdf.

Anonymous:1873:GLS

[3913] Anonymous. Gas light and steam heat. Scientific American, 28(5):64, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
64.pdf.

Anonymous:1873:ECa

[3914] Anonymous. Entomology in Congress. Scientific American, 28 (5):64, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
64a.pdf.

Anonymous:1873:VSS

[3915] Anonymous. The Vienna Show in the Senate. Scientific American, 28(5):64, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
64b.pdf.

Anonymous:1873:OTP

[3916] Anonymous. The ocean tides as a prime motor. Scientific American, 28(5):64–65, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
64c.pdf.

Anonymous:1873:BMb

[3917] Anonymous. Balancing machinery. Scientific American, 28(5): 65, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
65.pdf.

Anonymous:1873:DTIa

[3918] Anonymous. The duplex telegraph instrument in England. Scientific American, 28(5):65, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
65a.pdf.

Anonymous:1873:WU

[3919] Anonymous. What is the use? Scientific American, 28(5):65, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-65b.pdf.

Anonymous:1873:EM

[3920] Anonymous. Early metallurgy. Scientific American, 28(5):66, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-66.pdf.

Anonymous:1873:GSS

[3921] Anonymous. Guide for setting slide valves. Scientific American, 28(5):66, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
66a.pdf.

Anonymous:1873:GCD

[3922] Anonymous. Gears for changing direction of motion. Scientific American, 28(5):66, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
66b.pdf.

Anonymous:1873:PPa

[3923] Anonymous. The proportions of pipes. Scientific American, 28 (5):66, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
66c.pdf.

Anonymous:1873:PTFb

[3924] Anonymous. Professor Tyndall’s fifth lecture in New York — the invisible rays. Scientific American, 28(5):67, February 1, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n5/pdf/scientificamerican02011873-67.pdf.

Anonymous:1873:TCU

[3925] Anonymous. The telescope in the Cambridge University. Scientific American, 28(5):68, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
68a.pdf.

Anonymous:1873:GSAc

[3926] Anonymous. Gleanings in science and art. Scientific American, 28(5):69, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
69.pdf.

Anonymous:1873:ISPa

[3927] Anonymous. Improved seed planter. Scientific American, 28 (5):70, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
70.pdf.

Anonymous:1873:COH

[3928] Anonymous. Cooling our houses. Scientific American, 28(5):70, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-70a.pdf.

Anonymous:1873:NBN

[3929] Anonymous. New bridge over the Nile. Scientific American, 28 (5):71, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
71.pdf.

Anonymous:1873:LD

[3930] Anonymous. Learn to draw. Scientific American, 28(5):72, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-72.pdf.

Anonymous:1873:DDR

[3931] Anonymous. Dr. Draper’s researches in actino-chemistry. Scientific American, 28(5):72, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
72a.pdf.

Anonymous:1873:RAFe

[3932] Anonymous. Recent American and foreign patents. Scientific American, 28(5):73, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
73.pdf.

Anonymous:1873:DCc

[3933] Anonymous. Decisions of the courts. Scientific American, 28 (5):73, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
73b.pdf.

Anonymous:1873:VPa

[3934] Anonymous. Value of patents. Scientific American, 28(5):73, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n5/pdf/scientificamerican02011873-73c.pdf.

Anonymous:1873:ACe

[3935] Anonymous. Answers to correspondents. Scientific American, 28 (5):74, February 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n5/pdf/scientificamerican02011873-
74b.pdf.

Anonymous:1873:Cg

[3936] Anonymous. Correspondence. Scientific American, 28(6):83–84, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-83a.pdf.

Anonymous:1873:BNBd 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
89a.pdf.

Anonymous:1873:BPf

[3938] Anonymous. Business and personal. Scientific American, 28 (6):90, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
90.pdf.

Anonymous:1873:NQe

[3939] Anonymous. Notes and queries. Scientific American, 28(6):90, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-90a.pdf.

Anonymous:1873:IIe

[3940] Anonymous. Index of inventions. Scientific American, 28(6): 91–92, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
91a.pdf.

Anonymous:1873:FPB

[3941] Anonymous. Foot power band saw. Scientific American, 28 (6):79, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
79.pdf.

Anonymous:1873:KWS

[3942] Anonymous. Kindling wood splitter. Scientific American, 28 (6):79, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
79a.pdf.

Anonymous:1873:SSa

[3943] Anonymous. Shoe stretcher. Scientific American, 28(6):79, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-79b.pdf.

Anonymous:1873:DAa

[3944] Anonymous. The doctors and the apothecaries. Scientific American, 28(6):80, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
80.pdf.

Anonymous:1873:PPI

[3945] Anonymous. Photograph portraits — an improvement. Scientific American, 28(6):80, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
80a.pdf.

Anonymous:1873:ESa

[3946] Anonymous. The eye and the Sun. Scientific American, 28(6):80, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-80b.pdf.

Anonymous:1873:SFE

[3947] Anonymous. Steam as a fire extinguisher. Scientific American, 28(6):80–81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
80c.pdf.

Anonymous:1873:BWQ

[3948] Anonymous. The balance wheel question. Scientific American, 28(6):81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
81.pdf.

Anonymous:1873:TER

[3949] Anonymous. Traction engines or road locomotives. Scientific American, 28(6):81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
81a.pdf.

Anonymous:1873:SSB

[3950] Anonymous. The Signal Service Bureau. Scientific American, 28(6):81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
81b.pdf.

Anonymous:1873:AF

[3951] Anonymous. Artificial fertilization. Scientific American, 28(6): 81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
81c.pdf.

Anonymous:1873:RDP

[3952] Anonymous. Recent discoveries in the pyramids. Scientific American, 28(6):81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
81d.pdf.

Anonymous:1873:SGa

[3953] Anonymous. Silvering glass. Scientific American, 28(6):81, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-81e.pdf.

Anonymous:1873:PTSa

[3954] Anonymous. Professor Tyndall’s sixth and concluding lecture in New York. Scientific American, 28(6):82–83, February 8, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n6/pdf/scientificamerican02081873-82.pdf.

Anonymous:1873:SPa

[3955] Anonymous. A simple photometer. Scientific American, 28(6): 83, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
83.pdf.

Anonymous:1873:ANa

[3956] Anonymous. Astronomical notes. Scientific American, 28(6):84, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n6/pdf/scientificamerican02081873-84.pdf.

Anonymous:1873:PBL

[3957] Anonymous. Plumbago, black lead, graphite. Scientific American, 28(6):84–85, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
84a.pdf.

Anonymous:1873:CGM

[3958] Anonymous. Combination grinding machine. Scientific American, 28(6):86, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
86.pdf.

Anonymous:1873:ISCa

[3959] Anonymous. Improved self car coupler. Scientific American, 28 (6):86, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
86a.pdf.

Anonymous:1873:NST

[3960] Anonymous. The new St. Thomas’ Hospital. Scientific American, 28(6):87, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
87.pdf.

Anonymous:1873:EAI

[3961] Anonymous. English agricultural implements. Scientific American, 28(6):88, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
88.pdf.

Anonymous:1873:RAFf

[3962] Anonymous. Recent American and foreign patents. Scientific American, 28(6):89, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
89.pdf.

Anonymous:1873:ACf

[3963] Anonymous. Answers to correspondents. Scientific American, 28 (6):90–91, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
90b.pdf.

Anonymous:1873:CRc

[3964] Anonymous. Communications received. Scientific American, 28 (6):91, February 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n6/pdf/scientificamerican02081873-
91.pdf.

Anonymous:1873:Ch

[3965] Anonymous. Correspondence. Scientific American, 28(7):100, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-100.pdf.

Anonymous:1873:BPg

[3966] Anonymous. Business and personal. Scientific American, 28(7): 106, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
106.pdf.

Anonymous:1873:NQf

[3967] Anonymous. Notes and queries. Scientific American, 28(7):106, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-106a.pdf.

Anonymous:1873:IIf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-107a.pdf.

Anonymous:1873:BMc

[3969] Anonymous. Boring machine. Scientific American, 28(7):95, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-95.pdf.

Anonymous:1873:PSP

[3970] Anonymous. Paddle shafts for Pacific mail steamers. Scientific American, 28(7):95, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
95a.pdf.

Anonymous:1873:BWF

[3971] Anonymous. Burning water as fuel. Scientific American, 28(7): 96, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
96.pdf.

Anonymous:1873:MMA

[3972] Anonymous. The metropolitan museum of art. Scientific American, 28(7):96, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
96a.pdf.

Anonymous:1873:PIa

[3973] Anonymous. Are the planets inhabited? Scientific American, 28(7):96, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
96b.pdf.

Anonymous:1873:NYS

[3974] Anonymous. The New York state reward for improvements in canal navigation. Scientific American, 28(7):96–97, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-96c.pdf.

Anonymous:1873:ANP

[3975] Anonymous. Army and Navy patents. Scientific American, 28 (7):97, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
97.pdf.

Anonymous:1873:EWa

[3976] Anonymous. The employment of women. Scientific American, 28 (7):97–98, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
97a.pdf.

Anonymous:1873:ACg

[3977] Anonymous. Arsenic colors. Scientific American, 28(7):98, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-98.pdf.

Anonymous:1873:FMP

[3978] Anonymous. Free mail privileges ended. Scientific American, 28(7):98, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
98a.pdf.

Anonymous:1873:BFC

[3979] Anonymous. The Boston Fire Commissioners’ report. Scientific American, 28(7):98, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
98b.pdf.

Anonymous:1873:SSP

[3980] Anonymous. The system of snow protection on the union Pacific railroad. Scientific American, 28(7):98, February 15, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n7/pdf/scientificamerican02151873-98c.pdf.

Anonymous:1873:CDS

[3981] Anonymous. The chemical discoveries by the spectroscope. Scientific American, 28(7):99, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
99.pdf.

Anonymous:1873:MCC

[3982] Anonymous. Mixing cylinder for concrete. Scientific American, 28(7):99, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
99a.pdf.

Anonymous:1873:Pb

[3983] Anonymous. The pantagraph. Scientific American, 28(7):99, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-99b.pdf.

Anonymous:1873:LA

[3984] Anonymous. The lost arts. Scientific American, 28(7):101, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n7/pdf/scientificamerican02151873-101.pdf.

Anonymous:1873:SFC

[3985] Anonymous. Steam filter and condenser. Scientific American, 28(7):102, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
102.pdf.

Anonymous:1873:LAF

[3986] Anonymous. Lynn’s anti-freezing railroad tank valve and feeder. Scientific American, 28(7):102–103, February 15, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n7/pdf/scientificamerican02151873-102a.pdf.

Anonymous:1873:PGC

[3987] Anonymous. Pneumatic gun carriage and elevator. Scientific American, 28(7):103–105, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
103.pdf.

Anonymous:1873:PODd

[3988] Anonymous. Patent Office decisions. Scientific American, 28 (7):105, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
105.pdf.

Anonymous:1873:DCd

[3989] Anonymous. Decisions of the courts. Scientific American, 28 (7):105, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
105a.pdf.

Anonymous:1873:RAFg

[3990] Anonymous. Recent American and foreign patents. Scientific American, 28(7):105, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
105b.pdf.

Anonymous:1873:ACh

[3991] Anonymous. Answers to correspondents. Scientific American, 28 (7):106–107, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
106b.pdf.

Anonymous:1873:CRd

[3992] Anonymous. Communications received. Scientific American, 28 (7):107, February 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n7/pdf/scientificamerican02151873-
107.pdf.

Anonymous:1873:Ci (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
115b.pdf.

Anonymous:1873:BNBe

[3994] Anonymous. Books: New books and publications. Scientific American, 28(8):121, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
121b.pdf.

Anonymous:1873:BPh

[3995] Anonymous. Business and personal. Scientific American, 28(8): 122, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
122.pdf.

Anonymous:1873:NQg

[3996] Anonymous. Notes and queries. Scientific American, 28(8):122, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-122a.pdf.

Anonymous:1873:IIg

[3997] Anonymous. Index of inventions. Scientific American, 28(8): 123–124, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
123a.pdf.

Anonymous:1873:SFa

[3998] Anonymous. The snap flask. Scientific American, 28(8):111, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-111.pdf.

Anonymous:1873:ARC

[3999] Anonymous. Annual report of the Commissioner of Patents. Scientific American, 28(8):112, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
112.pdf.

Anonymous:1873:BTU

[4000] Anonymous. A blow to the trades’ unions. Scientific American, 28(8):112, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
112a.pdf.

Anonymous:1873:VEa

[4001] Anonymous. The Vienna Exposition. Scientific American, 28 (8):112, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
112b.pdf.

Anonymous:1873:SR

[4002] Anonymous. The science record for 1873. Scientific American, 28(8):112, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
112c.pdf.

Anonymous:1873:TFB

[4003] Anonymous. There fearful boiler explosions in one day. Scientific American, 28(8):112–113, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
112d.pdf.

Anonymous:1873:MFM

[4004] Anonymous. Matthew F. Maury. Scientific American, 28(8):113, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-113.pdf.

Anonymous:1873:OFG

[4005] Anonymous. An old friend gone. Scientific American, 28(8):113, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-113a.pdf.

Anonymous:1873:DPT

[4006] Anonymous. Departure of Professor Tyndall. Scientific American, 28(8):113, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
113b.pdf.

Anonymous:1873:ET

[4007] Anonymous. An electrical tower. Scientific American, 28(8):113, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-113c.pdf.

Anonymous:1873:FSF

[4008] Anonymous. Floating station for floating fire engines. Scientific American, 28(8):114, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
114.pdf.

Anonymous:1873:IE

[4009] Anonymous. Improved evaporator. Scientific American, 28(8): 114, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
114a.pdf.

Anonymous:1873:IPP

[4010] Anonymous. Improved pail pump. Scientific American, 28(8): 115, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
115.pdf.

Anonymous:1873:NCR

[4011] Anonymous. Novel and cheap railways in Portugal. Scientific American, 28(8):115, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
115a.pdf.

Anonymous:1873:ENb

[4012] Anonymous. Engineering notes. Scientific American, 28(8):117, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-117.pdf.

Anonymous:1873:IRSa

[4013] Anonymous. Improved road scraper. Scientific American, 28 (8):118, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
118.pdf.

Anonymous:1873:ACCa

[4014] Anonymous. Automatic car coupling. Scientific American, 28 (8):118, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
118a.pdf.

Anonymous:1873:RTB

[4015] Anonymous. Railway timber bridge. Scientific American, 28 (8):119, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
119.pdf.

Anonymous:1873:RP

[4016] Anonymous. Railway progress in 1872. Scientific American, 28 (8):119–120, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
119a.pdf.

Anonymous:1873:Cj

[4017] Anonymous. Coffee. Scientific American, 28(8):120, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n8/pdf/scientificamerican02221873-120.pdf.

Anonymous:1873:PODe

[4018] Anonymous. Patent Office decisions. Scientific American, 28 (8):121, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
121.pdf.

Anonymous:1873:RAFh

[4019] Anonymous. Recent American and foreign patents. Scientific American, 28(8):121, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
121a.pdf.

Anonymous:1873:ACi

[4020] Anonymous. Answers to correspondents. Scientific American, 28 (8):122–123, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
122b.pdf.

Anonymous:1873:CRe

[4021] Anonymous. Communications received. Scientific American, 28 (8):123, February 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n8/pdf/scientificamerican02221873-
123.pdf.

Anonymous:1873:Ck

[4022] Anonymous. Correspondence. Scientific American, 28(9):132–133, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-132.pdf.

Anonymous:1873:BNBf

[4023] Anonymous. Books: New books and publications. Scientific American, 28(9):136, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
136c.pdf.

Anonymous:1873:BPi

[4024] Anonymous. Business and personal. Scientific American, 28(9):138, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-138.pdf.

Anonymous:1873:NQh

[4025] Anonymous. Notes and queries. Scientific American, 28(9):138, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-138a.pdf.

Anonymous:1873:IIh

[4026] Anonymous. Index of inventions. Scientific American, 28(9):139–140, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-139a.pdf.

Anonymous:1873:ESS

[4027] Anonymous. The eclipse sectional steam boiler. Scientific American, 28(9):127, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
127.pdf.

Anonymous:1873:SPO

[4028] Anonymous. Small pox — old and new remedies. Scientific American, 28(9):128, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
128.pdf.

Anonymous:1873:WRT

[4029] Anonymous. A word with the reader, thinker, and writer. Scientific American, 28(9):128, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
128a.pdf.

Anonymous:1873:PAA

[4030] Anonymous. The photometer applied to astronomy. Scientific American, 28(9):128, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
128b.pdf.

Anonymous:1873:DTQ

[4031] Anonymous. Discoveries of tin in Queensland. Scientific American, 28(9):128, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
128c.pdf.

Anonymous:1873:FGH

[4032] Anonymous. A few gratuitous hints to inventors. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
129.pdf.

Anonymous:1873:PAO

[4033] Anonymous. Professor Agassiz on our colleges. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
129a.pdf.

Anonymous:1873:DTIb

[4034] Anonymous. Draper and Tyndall on the invisible rays. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
129b.pdf.

Anonymous:1873:RTNb

[4035] Anonymous. Rapid transit in New York. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-129c.pdf.

Anonymous:1873:SPId

[4036] Anonymous. Scientific and practical information. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
129d.pdf.

Anonymous:1873:TV

[4037] Anonymous. The Transit of Venus. Scientific American, 28(9):129, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-129e.pdf.

Anonymous:1873:CCSa

[4038] Anonymous. Cork cutting in Spain. Scientific American, 28(9):130–131, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-130.pdf.

Anonymous:1873:TFU

[4039] Anonymous. Textile fibers under microscope. Scientific American, 28(9):131, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
131.pdf.

Anonymous:1873:OPK

[4040] Anonymous. Our present knowledge of the Sun. Scientific American, 28(9):131–132, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
131a.pdf.

Anonymous:1873:CBa

[4041] Anonymous. Culinary boiler. Scientific American, 28(9):134, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-134.pdf.

Anonymous:1873:OP

[4042] Anonymous. Oscillating pump. Scientific American, 28(9):134, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-134a.pdf.

Anonymous:1873:NGG

[4043] Anonymous. New gyroscope governor. Scientific American, 28(9):134, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-134b.pdf.

Anonymous:1873:NRS

[4044] Anonymous. Non-radiating steam engine. Scientific American, 28 (9):135–136, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
135.pdf.

Anonymous:1873:WHAa

[4045] Anonymous. Where and how to advertise. Scientific American, 28(9):136, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
136.pdf.

Anonymous:1873:PODf

[4046] Anonymous. Patent Office decisions. Scientific American, 28(9):136, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-136a.pdf.

Anonymous:1873:DCe

[4047] Anonymous. Decisions of the courts. Scientific American, 28(9):136, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-136b.pdf.

Anonymous:1873:RAFi

[4048] Anonymous. Recent American and foreign patents. Scientific American, 28(9):137, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
137.pdf.

Anonymous:1873:ACj

[4049] Anonymous. Answers to correspondents. Scientific American, 28 (9):138–139, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n9/pdf/scientificamerican03011873-
138b.pdf.

Anonymous:1873:CRf

[4050] Anonymous. Communications received. Scientific American, 28(9):139, March 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n9/pdf/scientificamerican03011873-139.pdf.

Anonymous:1873:Cl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-148.pdf.

Anonymous:1873:BNBg

[4052] Anonymous. Books: New books and publications. Scientific American, 28(10):152, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
152b.pdf.

Anonymous:1873:BPj

[4053] Anonymous. Business and personal. Scientific American, 28(10):154, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-154.pdf.

Anonymous:1873:NQi

[4054] Anonymous. Notes and queries. Scientific American, 28(10):154, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-154a.pdf.

Anonymous:1873:IIi

[4055] Anonymous. Index of inventions. Scientific American, 28(10):155, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-155a.pdf.

Anonymous:1873:ISA

[4056] Anonymous. Improved self-acting fire tank. Scientific American, 28(10):143–144, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
143.pdf.

Anonymous:1873:HWH

[4057] Anonymous. Honor to whom honor is due. Scientific American, 28(10):144, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
144.pdf.

Anonymous:1873:NTEa

[4058] Anonymous. A novel traction engine for rapid transit. Scientific American, 28(10):144, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
144a.pdf.

Anonymous:1873:EPH

[4059] Anonymous. Explosions produced by high musical tones. Scientific American, 28(10):144, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
144b.pdf.

Anonymous:1873:PRB

[4060] Anonymous. Power railway brakes. Scientific American, 28(10): 144–145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
144c.pdf.

Anonymous:1873:CLW

[4061] Anonymous. Collapse from low water in steam boilers. Scientific American, 28(10):145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
145.pdf.

Anonymous:1873:PIW

[4062] Anonymous. The proposed introduction of water meters in cities. Scientific American, 28(10):145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
145a.pdf.

Anonymous:1873:IMC

[4063] Anonymous. The International Metric Commission and its work. Scientific American, 28(10):145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
145b.pdf.

Anonymous:1873:SMMb

[4064] Anonymous. The sewing machine monopoly. Scientific American, 28(10):145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
145c.pdf.

Anonymous:1873:ABN

[4065] Anonymous. The annexation of Brooklyn to New York. Scientific American, 28(10):145, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
145d.pdf.

Anonymous:1873:RTT

[4066] Anonymous. Registering tell tale clock. Scientific American, 28(10):146, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-146.pdf.

Anonymous:1873:ISV

[4067] Anonymous. Improvement in safety valves. Scientific American, 28(10):146, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
146a.pdf.

Anonymous:1873:SLE

[4068] Anonymous. Stilwell’s lime extracting heater and filter. Scientific American, 28(10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
147.pdf.

Anonymous:1873:LBI

[4069] Anonymous. Local boiler inspection. Scientific American, 28(10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-147a.pdf.

Anonymous:1873:NFK

[4070] Anonymous. A novel form of key. Scientific American, 28(10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-147b.pdf.

Anonymous:1873:NBS

[4071] Anonymous. A new bottle stopper. Scientific American, 28(10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-147c.pdf.

Anonymous:1873:SLS

[4072] Anonymous. Self-lighting signal lantern. Scientific American, 28 (10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
147d.pdf.

Anonymous:1873:CBB

[4073] Anonymous. Combined blacking brush, box, and holder. Scientific American, 28(10):147, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
147e.pdf.

Anonymous:1873:ANb

[4074] Anonymous. Astronomical notes. Scientific American, 28(10):148–149, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-148a.pdf.

Anonymous:1873:SPIe

[4075] Anonymous. Scientific and practical information. Scientific American, 28(10):149, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
149.pdf.

Anonymous:1873:IMF

[4076] Anonymous. Instrument for measuring fibers. Scientific American, 28(10):150, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
150.pdf.

Anonymous:1873:EDS

[4077] Anonymous. Emerson’s diamond stone saw. Scientific American, 28(10):150, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
150a.pdf.

Anonymous:1873:NMA

[4078] Anonymous. New method of attaching handles to saws. Scientific American, 28(10):150, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
150b.pdf.

Anonymous:1873:NPH

[4079] Anonymous. The new Parliament house in Berlin. Scientific American, 28(10):151–152, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
151.pdf.

Anonymous:1873:PODg

[4080] Anonymous. Patent Office decisions. Scientific American, 28(10):152, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-152.pdf.

Anonymous:1873:DCf

[4081] Anonymous. Decisions of the courts. Scientific American, 28(10):152, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-152a.pdf.

Anonymous:1873:RAFj

[4082] Anonymous. Recent American and foreign patents. Scientific American, 28(10):152–153, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
152c.pdf.

Anonymous:1873:ACk

[4083] Anonymous. Answers to correspondents. Scientific American, 28 (10):154–155, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
154b.pdf.

Anonymous:1873:CRg

[4084] Anonymous. Communications received. Scientific American, 28 (10):155, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n10/pdf/scientificamerican03081873-
155.pdf.

Anonymous:1873:CPa

[4085] Anonymous. Canadian patents. Scientific American, 28(10):155, March 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n10/pdf/scientificamerican03081873-155b.pdf.

Anonymous:1873:Cm

[4086] Anonymous. Correspondence. Scientific American, 28(11):164–165, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-164.pdf.

Anonymous:1873:BNBh

[4087] Anonymous. Books: New books and publications. Scientific American, 28(11):168, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
168.pdf.

Anonymous:1873:BPk

[4088] Anonymous. Business and personal. Scientific American, 28(11):170, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-170.pdf.

Anonymous:1873:NQj

[4089] Anonymous. Notes and queries. Scientific American, 28(11):170, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-170a.pdf.

Anonymous:1873:IIj

[4090] Anonymous. Index of inventions. Scientific American, 28(11):171–172, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-171a.pdf.

Anonymous:1873:IHP

[4091] Anonymous. Improved hot-pressing machine. Scientific American, 28(11):159, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
159.pdf.

Anonymous:1873:PAS

[4092] Anonymous. Poisonous air in schools and factories. Scientific American, 28(11):160, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
160.pdf.

Anonymous:1873:PRE

[4093] Anonymous. Practical results of the eight hour plan in New York. Scientific American, 28(11):160, March 15, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n11/pdf/scientificamerican03151873-160a.pdf.

Anonymous:1873:FAS

[4094] Anonymous. Foul air signal. Scientific American, 28(11):160, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-160b.pdf.

Anonymous:1873:IM

[4095] Anonymous. Invertebrate men. Scientific American, 28(11):160, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-160c.pdf.

Anonymous:1873:KT

[4096] Anonymous. The knowledge of the truth. Scientific American, 28 (11):160–161, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
160d.pdf.

Anonymous:1873:CTB

[4097] Anonymous. Colburn’s theory and the boiler explosion at Pittsburgh. Scientific American, 28(11):161, March 15, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n11/pdf/scientificamerican03151873-161.pdf.

Anonymous:1873:SPIf

[4098] Anonymous. Scientific and practical information. Scientific American, 28(11):161, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
161a.pdf.

Anonymous:1873:KPA

[4099] Anonymous. The king penguin and the Australian bustard. Scientific American, 28(11):162–163, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
162.pdf.

Anonymous:1873:PBT

[4100] Anonymous. Phosphor-bronze tweers. Scientific American, 28(11):163, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-163.pdf.

Anonymous:1873:EWb

[4101] Anonymous. Elastic washer. Scientific American, 28(11):163, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-163a.pdf.

Anonymous:1873:SLB

[4102] Anonymous. Sectional life boat. Scientific American, 28(11):163, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-163b.pdf.

Anonymous:1873:OMP

[4103] Anonymous. Odd materials for paper making. Scientific American, 28(11):165, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
165.pdf.

Anonymous:1873:IGC

[4104] Anonymous. Improved grain cleaner. Scientific American, 28(11):166, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n11/pdf/scientificamerican03151873-166.pdf.

Anonymous:1873:SRM

[4105] Anonymous. Steam reaper and mower. Scientific American, 28 (11):166–167, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
166a.pdf.

Anonymous:1873:BMR

[4106] Anonymous. Bridge over the Missouri River, near Leavenworth, Kansas. Scientific American, 28(11):167–168, March 15, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n11/pdf/scientificamerican03151873-167.pdf.

Anonymous:1873:RAFk

[4107] Anonymous. Recent American and foreign patents. Scientific American, 28(11):168–169, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
168a.pdf.

Anonymous:1873:ACl

[4108] Anonymous. Answers to correspondents. Scientific American, 28 (11):170–171, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
170b.pdf.

Anonymous:1873:CRh

[4109] Anonymous. Communications received. Scientific American, 28 (11):171, March 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n11/pdf/scientificamerican03151873-
171.pdf.

Anonymous:1873:Cn

[4110] Anonymous. Correspondence. Scientific American, 28(12):179–181, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-179d.pdf.

Anonymous:1873:BNBi

[4111] Anonymous. Books: New books and publications. Scientific American, 28(12):184, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
184a.pdf.

Anonymous:1873:BPl

[4112] Anonymous. Business and personal. Scientific American, 28(12):186, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-186.pdf.

Anonymous:1873:NQk

[4113] Anonymous. Notes and queries. Scientific American, 28(12):186, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-186a.pdf.

Anonymous:1873:IIk

[4114] Anonymous. Index of inventions. Scientific American, 28(12):187, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-187a.pdf.

Anonymous:1873:MBS

[4115] Anonymous. A mammoth band saw. Scientific American, 28(12):175, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-175.pdf.

Anonymous:1873:USC

[4116] Anonymous. A uniform system of colors. Scientific American, 28(12):176, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
176.pdf.

Anonymous:1873:PDC

[4117] Anonymous. Papin’s digester for cooking. Scientific American, 28(12):176, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
176a.pdf.

Anonymous:1873:DMO

[4118] Anonymous. Dangerous mineral oils. Scientific American, 28(12):176, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-176b.pdf.

Anonymous:1873:SFT

[4119] Anonymous. Sailing faster than the wind. Scientific American, 28 (12):176–177, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
176c.pdf.

Anonymous:1873:ILJ

[4120] Anonymous. The invention of the Leyden Jar. Scientific American, 28(12):177, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
177.pdf.

Anonymous:1873:SPIg

[4121] Anonymous. Scientific and practical information. Scientific American, 28(12):177, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
177a.pdf.

Anonymous:1873:EWM

[4122] Anonymous. Elastic washing mitts. Scientific American, 28(12):178, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-178.pdf.

Anonymous:1873:BHBa

[4123] Anonymous. Barber’s hair blowing apparatus. Scientific American, 28(12):178, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
178a.pdf.

Anonymous:1873:EPR

[4124] Anonymous. Early plans for railways. Scientific American, 28(12):178, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-178b.pdf.

Anonymous:1873:IFCa

[4125] Anonymous. Improved folding chair. Scientific American, 28(12):178, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-178c.pdf.

Anonymous:1873:SHa

[4126] Anonymous. Soap holder. Scientific American, 28(12):179, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-179.pdf.

Anonymous:1873:GSC

[4127] Anonymous. Grocer’s sample case. Scientific American, 28(12):179, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-179a.pdf.

Anonymous:1873:Pd

[4128] Anonymous. The pantagraph. Scientific American, 28(12):179, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-179b.pdf.

Anonymous:1873:SDA

[4129] Anonymous. Screw driver attachment. Scientific American, 28(12):179, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-179c.pdf.

Anonymous:1873:BCH

[4130] Anonymous. Bradley’s cushioned hammer. Scientific American, 28(12):182, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
182.pdf.

Anonymous:1873:IWBa

[4131] Anonymous. Improved wash boiler attachment. Scientific American, 28(12):182, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
182a.pdf.

Anonymous:1873:BCMa

[4132] Anonymous. Blasting in a coal mine. Scientific American, 28(12):184, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-184.pdf.

Anonymous:1873:PODh

[4133] Anonymous. Patent Office decisions. Scientific American, 28(12):184, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n12/pdf/scientificamerican03221873-184b.pdf.

Anonymous:1873:DCg

[4134] Anonymous. Decisions of the courts. Scientific American, 28 (12):184–185, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
184c.pdf.

Anonymous:1873:RAFl

[4135] Anonymous. Recent American and foreign patents. Scientific American, 28(12):185, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
185.pdf.

Anonymous:1873:ACm

[4136] Anonymous. Answers to correspondents. Scientific American, 28 (12):186–187, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
186b.pdf.

Anonymous:1873:CRi

[4137] Anonymous. Communications received. Scientific American, 28 (12):187, March 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n12/pdf/scientificamerican03221873-
187.pdf.

Anonymous:1873:Co

[4138] Anonymous. Correspondence. Scientific American, 28(13):195–196, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-195.pdf.

Anonymous:1873:BPm

[4139] Anonymous. Business and personal. Scientific American, 28(13):202, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-202.pdf.

Anonymous:1873:NQl

[4140] Anonymous. Notes and queries. Scientific American, 28(13):202, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-202a.pdf.

Anonymous:1873:IIl

[4141] Anonymous. Index of inventions. Scientific American, 28(13):203–204, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-203a.pdf.

Anonymous:1873:ISQ

[4142] Anonymous. Improved stone quarrying machine. Scientific American, 28(13):191, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
191.pdf.

Anonymous:1873:EDL

[4143] Anonymous. Explosions due to low water. Scientific American, 28(13):192, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
192.pdf.

Anonymous:1873:IWa

[4144] Anonymous. Injustice to workmen. Scientific American, 28(13):192, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-192a.pdf.

Anonymous:1873:PNP

[4145] Anonymous. Proposed new patent law in England. Scientific American, 28(13):192, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
192b.pdf.

Anonymous:1873:PSb

[4146] Anonymous. Pseudo science. Scientific American, 28(13):192–193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-192c.pdf.

Anonymous:1873:LF

[4147] Anonymous. Lens fires. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-193.pdf.

Anonymous:1873:CEL

[4148] Anonymous. Curious effect of light on selenium. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
193a.pdf.

Anonymous:1873:SRTb

[4149] Anonymous. The siphon recording telegraph instrument. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
193b.pdf.

Anonymous:1873:VEb

[4150] Anonymous. The Vienna Exhibition. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-193c.pdf.

Anonymous:1873:DRT

[4151] Anonymous. The Detroit River Tunnels. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
193d.pdf.

Anonymous:1873:SPIh

[4152] Anonymous. Scientific and practical information. Scientific American, 28(13):193, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
193e.pdf.

Anonymous:1873:KRO

[4153] Anonymous. Kircher’s remarkable observations concerning the Sun. Scientific American, 28(13):193–194, March 29, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n13/pdf/scientificamerican03291873-193f.pdf.

Anderson:1873:ST

[4154] John Anderson. The strength of timber. Scientific American, 28(13):197, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
197.pdf.

Anonymous:1873:NVG

[4155] Anonymous. New valve gearing for steam engines. Scientific American, 28(13):198, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
198.pdf.

Anonymous:1873:GSF

[4156] Anonymous. Gin saw filing machine. Scientific American, 28(13):198, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-198a.pdf.

Anonymous:1873:CGA

[4157] Anonymous. The coffee gardens of Arabia. Scientific American, 28(13):199, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
199.pdf.

Anonymous:1873:RIR

[4158] Anonymous. The Ralston inclined railway. Scientific American, 28(13):199, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
199a.pdf.

Anonymous:1873:BVH

[4159] Anonymous. The Bridal Veil, Havana Glen N.Y. Scientific American, 28(13):199, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
199b.pdf.

Anonymous:1873:HB

[4160] Anonymous. The Hixson beehive. Scientific American, 28(13):199–200, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-199c.pdf.

Anonymous:1873:DCh

[4161] Anonymous. Decisions of the courts. Scientific American, 28(13):200, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n13/pdf/scientificamerican03291873-200.pdf.

Anonymous:1873:RAFm

[4162] Anonymous. Recent American and foreign patents. Scientific American, 28(13):200–201, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
200a.pdf.

Anonymous:1873:ACn

[4163] Anonymous. Answers to correspondents. Scientific American, 28 (13):202–203, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
202b.pdf.

Anonymous:1873:CRj

[4164] Anonymous. Communications received. Scientific American, 28 (13):203, March 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n13/pdf/scientificamerican03291873-
203.pdf.

Anonymous:1873:BNBj

[4165] Anonymous. Books: New books and publications. Scientific American, 28(14):216, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
216a.pdf.

Anonymous:1873:BPn

[4166] Anonymous. Business and personal. Scientific American, 28(14):218, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-218.pdf.

Anonymous:1873:NQm

[4167] Anonymous. Notes and queries. Scientific American, 28(14):218, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-218a.pdf.

Anonymous:1873:IIm

[4168] Anonymous. Index of inventions. Scientific American, 28(14):219–220, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-219a.pdf.

Anonymous:1873:RSE

[4169] Anonymous. Revolving steam engine. Scientific American, 28(14):207, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-207.pdf.

Anonymous:1873:RBS

[4170] Anonymous. Reversible best single wheel grinder. Scientific American, 28(14):207–208, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
207a.pdf.

Anonymous:1873:IES

[4171] Anonymous. The intellectual enjoyments of science. Scientific American, 28(14):208, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
208.pdf.

Anonymous:1873:PDL

[4172] Anonymous. Pile driving and the laws of impact. Scientific American, 28(14):208, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
208a.pdf.

Anonymous:1873:WF

[4173] Anonymous. Water as fuel. Scientific American, 28(14):208, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-208b.pdf.

Anonymous:1873:CLa

[4174] Anonymous. Chemistry in Leipsic. Scientific American, 28(14):208–209, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-208c.pdf.

Anonymous:1873:SCF

[4175] Anonymous. Singular cause: of fire. Scientific American, 28(14):209, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-209.pdf.

Anonymous:1873:V

[4176] Anonymous. Vesuvius. Scientific American, 28(14):209, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-209a.pdf.

Anonymous:1873:GNT

[4177] Anonymous. The Greek new testament. Scientific American, 28 (14):209, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
209b.pdf.

Anonymous:1873:CID

[4178] Anonymous. Contagious and infectious diseases. Scientific American, 28(14):209, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
209c.pdf.

Anonymous:1873:CPM

[4179] Anonymous. Corn plow and marker. Scientific American, 28(14):210, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-210.pdf.

Anonymous:1873:IFS

[4180] Anonymous. An improved form of the selden steam pump. Scientific American, 28(14):210, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
210a.pdf.

Anonymous:1873:ANc

[4181] Anonymous. Astronomical notes. Scientific American, 28(14):211, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-211.pdf.

Anonymous:1873:Cq

[4182] Anonymous. Child’s hat. Scientific American, 28(14):211, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-211a.pdf.

Anonymous:1873:NDH

[4183] Anonymous. Nail driving hatchet. Scientific American, 28(14):211, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-211b.pdf.

Anonymous:1873:WM

[4184] Anonymous. Washing machine. Scientific American, 28(14):211, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-211c.pdf.

Anonymous:1873:IWWa

[4185] Anonymous. The increasing wealth of the world. Scientific American, 28(14):212, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
212.pdf.

Anonymous:1873:SPIi

[4186] Anonymous. Scientific and practical information. Scientific American, 28(14):212–213, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
212a.pdf.

Anonymous:1873:HNM

[4187] Anonymous. Horseshoe nail machinery. Scientific American, 28 (14):214–215, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
214.pdf.

Anonymous:1873:DCi

[4188] Anonymous. Decisions of the courts. Scientific American, 28(14):216, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-216.pdf.

Anonymous:1873:HSI

[4189] Anonymous. How shall I introduce my invention? Scientific American, 28(14):216, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
216b.pdf.

Anonymous:1873:RAFn

[4190] Anonymous. Recent American and foreign patents. Scientific American, 28(14):216–217, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
216c.pdf.

Anonymous:1873:ACo

[4191] Anonymous. Answers to correspondents. Scientific American, 28 (14):218–219, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
218b.pdf.

Anonymous:1873:CRk

[4192] Anonymous. Communications received. Scientific American, 28 (14):219, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n14/pdf/scientificamerican04051873-
219.pdf.

Anonymous:1873:CPb

[4193] Anonymous. Canadian patents. Scientific American, 28(14):220, April 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n14/pdf/scientificamerican04051873-220.pdf.

Anonymous:1873:Cs

[4194] Anonymous. Correspondence. Scientific American, 28(15):227–228, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-227.pdf.

Anonymous:1873:BPo

[4195] Anonymous. Business and personal. Scientific American, 28(15):234, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-234.pdf.

Anonymous:1873:NQn

[4196] Anonymous. Notes and queries. Scientific American, 28(15):234, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-234b.pdf.

Anonymous:1873:IIn

[4197] Anonymous. Index of inventions. Scientific American, 28(15):235–236, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-235.pdf.

Anonymous:1873:IDS

[4198] Anonymous. Improved duplex safety gun lock. Scientific American, 28(15):223, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
223.pdf.

Anonymous:1873:CCC

[4199] Anonymous. Compound chair for connecting railroad rails. Scientific American, 28(15):223, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
223a.pdf.

Anonymous:1873:AEO

[4200] Anonymous. Astronomical effects of the ocean tides. Scientific American, 28(15):224, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
224.pdf.

Anonymous:1873:SMI

[4201] Anonymous. A simple method of improving the health of cities. Scientific American, 28(15):224, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
224a.pdf.

Anonymous:1873:CAS

[4202] Anonymous. Carbonic acid in the soil. Scientific American, 28(15):224, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-224b.pdf.

Anonymous:1873:SFb

[4203] Anonymous. Steam on the farm. Scientific American, 28(15):224, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-224c.pdf.

Anonymous:1873:ESC

[4204] Anonymous. European street car improvements. Scientific American, 28(15):224, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
224d.pdf.

Anonymous:1873:NPL

[4205] Anonymous. Nebraska patent laws. Scientific American, 28(15): 224–225, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
224e.pdf.

Anonymous:1873:MGS

[4206] Anonymous. A munificent gift to science. Scientific American, 28(15):225, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
225.pdf.

Anonymous:1873:EI

[4207] Anonymous. Eccentric inventors. Scientific American, 28(15):225, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-225a.pdf.

Anonymous:1873:ICO

[4208] Anonymous. The ignition of combustibles by overheated steam. Scientific American, 28(15):225, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
225b.pdf.

Anonymous:1873:NMb

[4209] Anonymous. A new mitrailleuse. Scientific American, 28(15):225, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-225c.pdf.

Anonymous:1873:TPa

[4210] Anonymous. Telegraph poles. Scientific American, 28(15):226, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-226.pdf.

Anonymous:1873:SHb

[4211] Anonymous. Sound and hearing. Scientific American, 28(15):226–227, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-226a.pdf.

Anonymous:1873:SPIj

[4212] Anonymous. Scientific and practical information. Scientific American, 28(15):229, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
229.pdf.

Anonymous:1873:IRD

[4213] Anonymous. Improved radial drilling machine. Scientific American, 28(15):230, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
230.pdf.

Anonymous:1873:HSG

[4214] Anonymous. High steam gage and low water register. Scientific American, 28(15):230, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
230a.pdf.

Anonymous:1873:FST

[4215] Anonymous. Flexible saw and tooth filing apparatus. Scientific American, 28(15):231–232, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
231.pdf.

Anonymous:1873:DCj

[4216] Anonymous. Decisions of the courts. Scientific American, 28(15):232, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n15/pdf/scientificamerican04121873-232.pdf.

Anonymous:1873:RAFo

[4217] Anonymous. Recent American and foreign patents. Scientific American, 28(15):232–233, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
232a.pdf.

Anonymous:1873:BMIa

[4218] Anonymous. Best mode of introducing new inventions. Scientific American, 28(15):234, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
234a.pdf.

Anonymous:1873:ACp

[4219] Anonymous. Answers to correspondents. Scientific American, 28 (15):234–235, April 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n15/pdf/scientificamerican04121873-
234c.pdf.

Anonymous:1873:Ct

[4220] Anonymous. Correspondence. Scientific American, 28(16):243–245, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-243b.pdf.

Anonymous:1873:BNBk

[4221] Anonymous. Books: New books and publications. Scientific American, 28(16):248, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
248b.pdf.

Anonymous:1873:BPp

[4222] Anonymous. Business and personal. Scientific American, 28(16):250, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-250.pdf.

Anonymous:1873:NQo

[4223] Anonymous. Notes and queries. Scientific American, 28(16):250, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-250b.pdf.

Anonymous:1873:IIo

[4224] Anonymous. Index of inventions. Scientific American, 28(16):251–252, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-251a.pdf.

Anonymous:1873:SPE

[4225] Anonymous. Steam pumping engines for India state railways. Scientific American, 28(16):239, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
239.pdf.

Anonymous:1873:ETO

[4226] Anonymous. The evil of trade organizations. Scientific American, 28(16):240, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
240.pdf.

Anonymous:1873:DPL

[4227] Anonymous. Driving power of leather belts. Scientific American, 28(16):240, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
240a.pdf.

Anonymous:1873:Cu

[4228] Anonymous. Copyrights. Scientific American, 28(16):240, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-240b.pdf.

Anonymous:1873:SI

[4229] Anonymous. The Scope of the infinitesimal. Scientific American, 28(16):240–241, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
240c.pdf.

Anonymous:1873:WA

[4230] Anonymous. The wreck of the Atlantic. Scientific American, 28(16):241, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-241.pdf.

Anonymous:1873:PCA

[4231] Anonymous. The present condition of Arctic explorations. Scientific American, 28(16):241, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
241a.pdf.

Anonymous:1873:SO

[4232] Anonymous. Swindling operations. Scientific American, 28(16):241, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-241b.pdf.

Anonymous:1873:RWT

[4233] Anonymous. Resistance of woods to torsional strain. Scientific American, 28(16):241, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
241c.pdf.

Anonymous:1873:DAD

[4234] Anonymous. Diamonds applied to dressing and working moldings in stone. Scientific American, 28(16):241–242, April 19, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n16/pdf/scientificamerican04191873-241d.pdf.

Anonymous:1873:WT

[4235] Anonymous. Weather threshold. Scientific American, 28(16):242, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-242.pdf.

Anonymous:1873:RB

[4236] Anonymous. Recording barometer. Scientific American, 28(16):242, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-242a.pdf.

Anonymous:1873:PT

[4237] Anonymous. Pipe tongs. Scientific American, 28(16):242, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-242b.pdf.

Anonymous:1873:SOP

[4238] Anonymous. A submarine observatory and photographic gallery. Scientific American, 28(16):242, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
242c.pdf.

Anonymous:1873:MC

[4239] Anonymous. Manufacture of combs. Scientific American, 28(16):243, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-243.pdf.

Anonymous:1873:IWF

[4240] Anonymous. Improved water fire-grate. Scientific American, 28 (16):243, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
243a.pdf.

Anonymous:1873:GSWb

[4241] Anonymous. The government submarine works at Hell Gate. Scientific American, 28(16):245, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
245.pdf.

Anonymous:1873:LAW

[4242] Anonymous. Lunatic Asylum, Windsor, England. Scientific American, 28(16):246, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
246.pdf.

Anonymous:1873:SPb

[4243] Anonymous. Sulky plow. Scientific American, 28(16):247, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-247.pdf.

Anonymous:1873:TCSb

[4244] Anonymous. The tanite Company’s shaper and sharpener. Scientific American, 28(16):247, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
247a.pdf.

Anonymous:1873:PSA

[4245] Anonymous. Postage stamp affixer. Scientific American, 28(16): 247–248, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
247b.pdf.

Anonymous:1873:PODi

[4246] Anonymous. Patent Office decisions. Scientific American, 28(16):248, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-248.pdf.

Anonymous:1873:DCk

[4247] Anonymous. Decisions of the courts. Scientific American, 28(16):248, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n16/pdf/scientificamerican04191873-248a.pdf.

Anonymous:1873:RAFp

[4248] Anonymous. Recent American and foreign patents. Scientific American, 28(16):248–249, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
248c.pdf.

Anonymous:1873:BMIb

[4249] Anonymous. Best mode of introducing new inventions. Scientific American, 28(16):250, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
250a.pdf.

Anonymous:1873:ACq

[4250] Anonymous. Answers to correspondents. Scientific American, 28 (16):250–251, April 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
250c.pdf.

Anonymous:1873:CRl (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n16/pdf/scientificamerican04191873-
251.pdf.

Anonymous:1873:Cv

[4252] Anonymous. Correspondence. Scientific American, 28(17):260–261, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-260.pdf.

Anonymous:1873:BNBl

[4253] Anonymous. Books: New books and publications. Scientific American, 28(17):264, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
264.pdf.

Anonymous:1873:BPq

[4254] Anonymous. Business and personal. Scientific American, 28(17):266, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-266.pdf.

Anonymous:1873:NQp

[4255] Anonymous. Notes and queries. Scientific American, 28(17):266, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-266b.pdf.

Anonymous:1873:IIp

[4256] Anonymous. Index of inventions. Scientific American, 28(17):267–268, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-267a.pdf.

Anonymous:1873:TAF

[4257] Anonymous. Taylor’s American field piece. Scientific American, 28(17):255–256, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
255.pdf.

Anonymous:1873:SOM

[4258] Anonymous. The sources of our modern knowledge. Scientific American, 28(17):256, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
256.pdf.

Anonymous:1873:BURa

[4259] Anonymous. The Broadway Underground Railway. Scientific American, 28(17):256, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
256a.pdf.

Anonymous:1873:FBS

[4260] Anonymous. The fire on board the steamer Alaska. Scientific American, 28(17):256–257, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
256b.pdf.

Anonymous:1873:HR

[4261] Anonymous. The hydraulic ram. Scientific American, 28(17):257, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-257.pdf.

Anonymous:1873:EPS

[4262] Anonymous. Estimating power by size and speed of belts. Scientific American, 28(17):257, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
257a.pdf.

Anonymous:1873:Z

[4263] Anonymous. Zirconium. Scientific American, 28(17):257, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-257b.pdf.

Anonymous:1873:PMa

[4264] Anonymous. Patent medicines. Scientific American, 28(17):257, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-257c.pdf.

Anonymous:1873:NMM

[4265] Anonymous. A novel mode of marine propulsion. Scientific American, 28(17):258, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
258.pdf.

Anonymous:1873:NRT

[4266] Anonymous. Novel rat trap. Scientific American, 28(17):258–259, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-258a.pdf.

Anonymous:1873:WEa

[4267] Anonymous. The wonders of the egg. Scientific American, 28(17):259, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-259.pdf.

Anonymous:1873:CDB

[4268] Anonymous. The configuration and dimensions of blast furnaces. Scientific American, 28(17):259–260, April 26, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n17/pdf/scientificamerican04261873-259a.pdf.

Anonymous:1873:ALB

[4269] Anonymous. American locomotives — the Baldwin Works. Scientific American, 28(17):261, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
261.pdf.

Anonymous:1873:CLb

[4270] Anonymous. Cart loader. Scientific American, 28(17):262, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-262.pdf.

Anonymous:1873:CPc

[4271] Anonymous. Combination padlock. Scientific American, 28(17):262, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-262a.pdf.

Anonymous:1873:CBc

[4272] Anonymous. Culinary boiler. Scientific American, 28(17):262, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n17/pdf/scientificamerican04261873-262b.pdf.

Anonymous:1873:CPN

[4273] Anonymous. Central Park, New York City. Scientific American, 28(17):263–264, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
263.pdf.

Anonymous:1873:RAFq

[4274] Anonymous. Recent American and foreign patents. Scientific American, 28(17):264–265, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
264a.pdf.

Anonymous:1873:BMIc

[4275] Anonymous. Best mode of introducing new inventions. Scientific American, 28(17):266, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
266a.pdf.

Anonymous:1873:ACr

[4276] Anonymous. Answers to correspondents. Scientific American, 28 (17):266–267, April 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
266c.pdf.

Anonymous:1873:CRm (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n17/pdf/scientificamerican04261873-
267.pdf.

Anonymous:1873:Cw

[4278] Anonymous. Correspondence. Scientific American, 28(18):277, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-277a.pdf.

Anonymous:1873:BPr

[4279] Anonymous. Business and personal. Scientific American, 28(18):282, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-282.pdf.

Anonymous:1873:NQq

[4280] Anonymous. Notes and queries. Scientific American, 28(18):282, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-282b.pdf.

Anonymous:1873:IIq

[4281] Anonymous. Index of inventions. Scientific American, 28(18):283–284, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-283a.pdf.

Anonymous:1873:IPB

[4282] Anonymous. Improved planer bar. Scientific American, 28(18):271, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-271.pdf.

Anonymous:1873:CBd

[4283] Anonymous. The circulation of the blood. Scientific American, 28(18):272, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
272.pdf.

Anonymous:1873:IWc

[4284] Anonymous. Injustice to workmen. Scientific American, 28(18):272, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-272a.pdf.

Anonymous:1873:FS

[4285] Anonymous. Fireproof safes. Scientific American, 28(18):272, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-272b.pdf.

Anonymous:1873:CPDa

[4286] Anonymous. The case of Phelps, Dodge & Co. Scientific American, 28(18):272, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
272c.pdf.

Anonymous:1873:BURb

[4287] Anonymous. The Broadway Underground Railway. Scientific American, 28(18):272, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
272d.pdf.

Anonymous:1873:NRO

[4288] Anonymous. A new reading of an old record. Scientific American, 28(18):272–273, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
272e.pdf.

Anonymous:1873:Cx

[4289] Anonymous. Codeine. Scientific American, 28(18):273, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-273.pdf.

Anonymous:1873:PCTa

[4290] Anonymous. The problem of the coming transits. Scientific American, 28(18):273, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
273a.pdf.

Anonymous:1873:SPIk

[4291] Anonymous. Scientific and practical information. Scientific American, 28(18):273, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
273b.pdf.

Anonymous:1873:DBL

[4292] Anonymous. Death of Baron von Liebig. Scientific American, 28(18):273–274, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
273c.pdf.

Anonymous:1873:EHS

[4293] Anonymous. Early history of steam navigation — the first screw propeller. Scientific American, 28(18):274, May 3, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n18/pdf/scientificamerican05031873-274.pdf.

Anonymous:1873:BTF

[4294] Anonymous. Bale tie fastener. Scientific American, 28(18):274–275, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-274a.pdf.

Morton:1873:SILa

[4295] Henry Morton. Stevens institute lectures — sunlight and its source. Scientific American, 28(18):275, May 3, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n18/pdf/scientificamerican05031873-275.pdf.

Anonymous:1873:WEb

[4296] Anonymous. The wonders of the egg. Scientific American, 28 (18):275–276, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
275a.pdf.

Anonymous:1873:TEV

[4297] Anonymous. Trials of the Ellis vapor engine. Scientific American, 28(18):276, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
276.pdf.

Anonymous:1873:IS

[4298] Anonymous. Improved spectroscope. Scientific American, 28(18):277, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-277.pdf.

Anonymous:1873:IWBb

[4299] Anonymous. Improved wagon brake. Scientific American, 28(18):278, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-278.pdf.

Anonymous:1873:IFT

[4300] Anonymous. Improved fly trap. Scientific American, 28(18):278–279, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-278a.pdf.

Anonymous:1873:NMS

[4301] Anonymous. New machine for sinking well tubes. Scientific American, 28(18):279, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
279.pdf.

Anonymous:1873:AMW

[4302] Anonymous. The agricultural machine works of Chicago. Scientific American, 28(18):279–280, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
279a.pdf.

Anonymous:1873:CNa

[4303] Anonymous. Chemical notes. Scientific American, 28(18):280–281, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-280.pdf.

Anonymous:1873:DCl

[4304] Anonymous. Decisions of the courts. Scientific American, 28(18):281, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n18/pdf/scientificamerican05031873-281.pdf.

Anonymous:1873:RAFr

[4305] Anonymous. Recent American and foreign patents. Scientific American, 28(18):281, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
281a.pdf.

Anonymous:1873:BMId

[4306] Anonymous. Best mode of introducing new inventions. Scientific American, 28(18):282, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
282a.pdf.

Anonymous:1873:ACs

[4307] Anonymous. Answers to correspondents. Scientific American, 28 (18):282–283, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
282c.pdf.

Anonymous:1873:CRn

[4308] Anonymous. Communications received. Scientific American, 28 (18):283, May 3, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n18/pdf/scientificamerican05031873-
283.pdf.

Anonymous:1873:Cy

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-292.pdf.

Anonymous:1873:BNBm

[4310] Anonymous. Books: New books and publications. Scientific American, 28(19):297, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
297.pdf.

Anonymous:1873:BPs

[4311] Anonymous. Business and personal. Scientific American, 28(19):298, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-298.pdf.

Anonymous:1873:NQr

[4312] Anonymous. Notes and queries. Scientific American, 28(19):298, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-298b.pdf.

Anonymous:1873:IIr

[4313] Anonymous. Index of inventions. Scientific American, 28(19):299–300, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-299a.pdf.

Anonymous:1873:DTMa

[4314] Anonymous. Double tenoning machine, mammoth remains. Scientific American, 28(19):287, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
287.pdf.

Anonymous:1873:PCPa

[4315] Anonymous. The physical condition of the planets. Scientific American, 28(19):288, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
288.pdf.

Anonymous:1873:FWE

[4316] Anonymous. The fly wheel. — exact formulas by which to proportion it. Scientific American, 28(19):288, May 10, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n19/pdf/scientificamerican05101873-288a.pdf.

Anonymous:1873:NOT

[4317] Anonymous. The new ocean telegraph cable. Scientific American, 28(19):288, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
288b.pdf.

Anonymous:1873:APVa

[4318] Anonymous. Another phase of the Vienna Exposition. Scientific American, 28(19):288, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
288c.pdf.

Anonymous:1873:GWW

[4319] Anonymous. The geographical work of the world in 1872. Scientific American, 28(19):288–289, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
288d.pdf.

Anonymous:1873:PCTb

[4320] Anonymous. The problem of the coming transits. Scientific American, 28(19):289, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
289.pdf.

Anonymous:1873:SPIl

[4321] Anonymous. Scientific and practical information. Scientific American, 28(19):289, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
289a.pdf.

Anonymous:1873:OUM

[4322] Anonymous. The ornamental use of metal. Scientific American, 28(19):290, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
290.pdf.

Morton:1873:SILb

[4323] Henry Morton. Stevens institute lectures. — moonlight and its source. Scientific American, 28(19):291, May 10, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n19/pdf/scientificamerican05101873-291.pdf.

Anonymous:1873:WESa

[4324] Anonymous. The wonders of the egg. — second lecture. Scientific American, 28(19):291–292, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
291a.pdf.

Anonymous:1873:CNb

[4325] Anonymous. Chemical notes. Scientific American, 28(19):293, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-293.pdf.

Anonymous:1873:TB

[4326] Anonymous. The Tay Bridge. Scientific American, 28(19):294–295, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-294.pdf.

Anonymous:1873:PCPb

[4327] Anonymous. A patent cannon plow. Scientific American, 28(19):295, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-295.pdf.

Anonymous:1873:ANd

[4328] Anonymous. Astronomical notes. Scientific American, 28(19):296, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n19/pdf/scientificamerican05101873-296.pdf.

Anonymous:1873:DCm

[4329] Anonymous. Decisions of the courts. Scientific American, 28 (19):296–297, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
296a.pdf.

Anonymous:1873:RAFs

[4330] Anonymous. Recent American and foreign patents. Scientific American, 28(19):297, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
297a.pdf.

Anonymous:1873:BMIe

[4331] Anonymous. Best mode of introducing new inventions. Scientific American, 28(19):298, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
298a.pdf.

Anonymous:1873:ACt

[4332] Anonymous. Answers to correspondents. Scientific American, 28 (19):298–299, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
298c.pdf.

Anonymous:1873:CRo

[4333] Anonymous. Communications received. Scientific American, 28 (19):299, May 10, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n19/pdf/scientificamerican05101873-
299.pdf.

Anonymous:1873:Cz

[4334] Anonymous. Correspondence. Scientific American, 28(20):309, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-309.pdf.

Anonymous:1873:BNBn

[4335] Anonymous. Books: New books and publications. Scientific American, 28(20):312, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
312a.pdf.

Anonymous:1873:BPt

[4336] Anonymous. Business and personal. Scientific American, 28(20):314, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-314.pdf.

Anonymous:1873:NQs

[4337] Anonymous. Notes and queries. Scientific American, 28(20):314, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-314a.pdf.

Anonymous:1873:IIs

[4338] Anonymous. Index of inventions. Scientific American, 28(20):315–316, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-315a.pdf.

Anonymous:1873:IBH

[4339] Anonymous. Improved bolt heading machine. Scientific American, 28(20):303, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
303.pdf.

Anonymous:1873:IPa

[4340] Anonymous. Improved planer. Scientific American, 28(20):303, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-303a.pdf.

Anonymous:1873:EOI

[4341] Anonymous. Is the Earth the only inhabited world? Scientific American, 28(20):304, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
304.pdf.

Anonymous:1873:NRR

[4342] Anonymous. A new rule in respect to caveats. Scientific American, 28(20):304, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
304a.pdf.

Anonymous:1873:OVE

[4343] Anonymous. Opening of the Vienna Exposition. Scientific American, 28(20):304, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
304b.pdf.

Anonymous:1873:NSRa

[4344] Anonymous. A new specific for rheumatism. Scientific American, 28(20):304–305, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
304c.pdf.

Anonymous:1873:NAM

[4345] Anonymous. Nature and art — the mystery of the magneto-electric machine. Scientific American, 28(20):305, May 17, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n20/pdf/scientificamerican05171873-305.pdf.

Anonymous:1873:MS

[4346] Anonymous. The movements of the stars. Scientific American, 28(20):305, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
305a.pdf.

Anonymous:1873:SPIm

[4347] Anonymous. Scientific and practical information. Scientific American, 28(20):305, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
305b.pdf.

Anonymous:1873:RPJ

[4348] Anonymous. Root power jig raw. Scientific American, 28(20):306, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-306.pdf.

Anonymous:1873:TPF

[4349] Anonymous. Taylor’s patent firkin head. Scientific American, 28 (20):306, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
306a.pdf.

Anonymous:1873:WESb

[4350] Anonymous. The wonders of the egg, — second lecture. Scientific American, 28(20):306–307, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
306b.pdf.

Silliman:1873:SIL

[4351] Silliman. Stevens institute lectures. — gold mining in California. Scientific American, 28(20):307, May 17, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n20/pdf/scientificamerican05171873-307.pdf.

Anonymous:1873:SRN

[4352] Anonymous. Science record. — new edition. Scientific American, 28(20):307, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
307a.pdf.

Anonymous:1873:PMb

[4353] Anonymous. Paper making. Scientific American, 28(20):308–309, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-308.pdf.

Anonymous:1873:CHRa

[4354] Anonymous. Country homes — rural cottages. Scientific American, 28(20):310–311, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
310.pdf.

Anonymous:1873:FFa

[4355] Anonymous. Fruit fork. Scientific American, 28(20):311, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-311.pdf.

Anonymous:1873:TPb

[4356] Anonymous. Tree protector. Scientific American, 28(20):311–312, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-311a.pdf.

Anonymous:1873:RMM

[4357] Anonymous. A.r.m.m.a. Scientific American, 28(20):312, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-312.pdf.

Anonymous:1873:DCn

[4358] Anonymous. Decisions of the courts. Scientific American, 28(20):312, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n20/pdf/scientificamerican05171873-312b.pdf.

Anonymous:1873:RAFt

[4359] Anonymous. Recent American and foreign patents. Scientific American, 28(20):312–313, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
312c.pdf.

Anonymous:1873:ACv

[4360] Anonymous. Answers to correspondents. Scientific American, 28 (20):314–315, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
314b.pdf.

Anonymous:1873:CRp

[4361] Anonymous. Communications received. Scientific American, 28 (20):315, May 17, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n20/pdf/scientificamerican05171873-
315.pdf.

Anonymous:1873:Cbe

[4362] Anonymous. Correspondence. Scientific American, 28(21):325, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-325.pdf.

Anonymous:1873:BNBo

[4363] Anonymous. Books: New books and publications. Scientific American, 28(21):328, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
328b.pdf.

Anonymous:1873:BPu

[4364] Anonymous. Business and personal. Scientific American, 28(21):330, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-330.pdf.

Anonymous:1873:NQt

[4365] Anonymous. Notes and queries. Scientific American, 28(21):330, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-330a.pdf.

Anonymous:1873:IIt

[4366] Anonymous. Index of inventions. Scientific American, 28(21):331–332, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-331a.pdf.

Anonymous:1873:MZZ

[4367] Anonymous. The manufacture of zinc and zinc white, how to search for metals, and more. Scientific American, 28(21):319, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-319.pdf.

Anonymous:1873:MPE

[4368] Anonymous. Medical practice in early times. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
320.pdf.

Anonymous:1873:HPP

[4369] Anonymous. The horse power and the power of the horse. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
320a.pdf.

Anonymous:1873:CHI

[4370] Anonymous. Chemical heat indicators. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-320b.pdf.

Anonymous:1873:LM

[4371] Anonymous. Leaping by machinery. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-320c.pdf.

Anonymous:1873:LP

[4372] Anonymous. Loss of the Polaris. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-320d.pdf.

Anonymous:1873:IDR

[4373] Anonymous. Important decision in respect to assignments. Scientific American, 28(21):320, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
320e.pdf.

Anonymous:1873:DL

[4374] Anonymous. The death of Liebig. Scientific American, 28(21):320–321, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-320f.pdf.

Anonymous:1873:FDB

[4375] Anonymous. The falling of the Dixon bridge. Scientific American, 28(21):321, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
321.pdf.

Anonymous:1873:SPIn

[4376] Anonymous. Scientific and practical information. Scientific American, 28(21):321, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
321a.pdf.

Anonymous:1873:JPL

[4377] Anonymous. Judson’s patent lathe chuck. Scientific American, 28(21):322, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
322.pdf.

Anonymous:1873:ERS

[4378] Anonymous. Electric railway signal. Scientific American, 28(21):322, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-322a.pdf.

Anonymous:1873:CHa

[4379] Anonymous. Hat and coat hook. Scientific American, 28(21):322, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-322b.pdf.

Anonymous:1873:IMS

[4380] Anonymous. Improved metallic stuffing box. Scientific American, 28(21):322, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
322c.pdf.

Anonymous:1873:WEIa

[4381] Anonymous. The wonders of the egg — III. Scientific American, 28(21):323, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
323.pdf.

Chandler:1873:SIL

[4382] Charles F. Chandler. Stevens institute lectures — dyeing and calico printing. Scientific American, 28(21):324, May 24, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n21/pdf/scientificamerican05241873-324.pdf.

Anonymous:1873:SNFa

[4383] Anonymous. Sanitary notes — flavoring substances for food and drink. Scientific American, 28(21):324, May 24, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n21/pdf/scientificamerican05241873-324a.pdf.

Anonymous:1873:IBCa

[4384] Anonymous. Iron bridge construction. Scientific American, 28 (21):326–328, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
326.pdf.

Anonymous:1873:PODj

[4385] Anonymous. Patent Office decisions. Scientific American, 28(21):328, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-328.pdf.

Anonymous:1873:DCo

[4386] Anonymous. Decisions of the courts. Scientific American, 28(21):328, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n21/pdf/scientificamerican05241873-328a.pdf.

Anonymous:1873:RAFu

[4387] Anonymous. Recent American and foreign patents. Scientific American, 28(21):328–329, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
328c.pdf.

Anonymous:1873:ACw

[4388] Anonymous. Answers to correspondents. Scientific American, 28 (21):330–331, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
330b.pdf.

Anonymous:1873:CRq

[4389] Anonymous. Communications received. Scientific American, 28 (21):331, May 24, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n21/pdf/scientificamerican05241873-
331.pdf.

Anonymous:1873:Cbg

[4390] Anonymous. Correspondence. Scientific American, 28(22):341, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-341.pdf.

Anonymous:1873:BPv

[4391] Anonymous. Business and personal. Scientific American, 28(22):346, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-346.pdf.

Anonymous:1873:NQu

[4392] Anonymous. Notes and queries. Scientific American, 28(22):346, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-346b.pdf.

Anonymous:1873:IIu

[4393] Anonymous. Index of inventions. Scientific American, 28(22):347–348, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-347a.pdf.

Anonymous:1873:GBA

[4394] Anonymous. Grain binder attachment to harvesters. Scientific American, 28(22):335, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
335.pdf.

Anonymous:1873:MSCa

[4395] Anonymous. Mining schools and colleges. Scientific American, 28(22):336, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
336.pdf.

Anonymous:1873:ESM

[4396] Anonymous. Experiments on the strength of materials. Scientific American, 28(22):336, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
336a.pdf.

Anonymous:1873:RPP

[4397] Anonymous. Rights of purchasers of patented articles. Scientific American, 28(22):336, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
336b.pdf.

Anonymous:1873:NPSa

[4398] Anonymous. The new Philadelphia steamships. Scientific American, 28(22):336, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
336c.pdf.

Anonymous:1873:VEAa

[4399] Anonymous. The Vienna Exposition — appointment of a new Commissioner. Scientific American, 28(22):336, May 31, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n22/pdf/scientificamerican05311873-336d.pdf.

Anonymous:1873:NYP

[4400] Anonymous. The New York post office. Scientific American, 28 (22):336–337, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
336e.pdf.

Anonymous:1873:MPM

[4401] Anonymous. Modern printing machinery and appliances — how a daily newspaper is made. Scientific American, 28(22):337, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-337.pdf.

Anonymous:1873:AT

[4402] Anonymous. About tea. Scientific American, 28(22):338, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-338.pdf.

Anonymous:1873:IBCb

[4403] Anonymous. Iron bridge construction. Scientific American, 28 (22):338–340, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
338a.pdf.

Anonymous:1873:WEIb

[4404] Anonymous. The wonders of the egg — IV. Scientific American, 28(22):340–341, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
340.pdf.

Anonymous:1873:SHS

[4405] Anonymous. Simple hand stamp. Scientific American, 28(22):342, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-342.pdf.

Anonymous:1873:RCE

[4406] Anonymous. Rag cutting engine. Scientific American, 28(22):342, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-342a.pdf.

Anonymous:1873:SHU

[4407] Anonymous. Scraper for household uses. Scientific American, 28 (22):342, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
342b.pdf.

Anonymous:1873:SMa

[4408] Anonymous. The Sigl mitrailleuse. Scientific American, 28(22):342, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-342c.pdf.

Anonymous:1873:NSRb

[4409] Anonymous. New steam roller. Scientific American, 28(22):343, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-343.pdf.

Morton:1873:SILc

[4410] Henry Morton. Stevens institute lectures — fluorescence. Scientific American, 28(22):343–344, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
343a.pdf.

Anonymous:1873:DCp

[4411] Anonymous. Decisions of the courts. Scientific American, 28(22):344, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-344.pdf.

Anonymous:1873:RAFv

[4412] Anonymous. Recent American and foreign patents. Scientific American, 28(22):344–345, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
344a.pdf.

Anonymous:1873:Cbh

[4413] Anonymous. Copyrights. Scientific American, 28(22):346, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n22/pdf/scientificamerican05311873-346a.pdf.

Anonymous:1873:ACx

[4414] Anonymous. Answers to correspondents. Scientific American, 28 (22):346–347, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
346c.pdf.

Anonymous:1873:CRr

[4415] Anonymous. Communications received. Scientific American, 28 (22):347, May 31, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n22/pdf/scientificamerican05311873-
347.pdf.

Anonymous:1873:Cbi

[4416] Anonymous. Correspondence. Scientific American, 28(23):357, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-357.pdf.

Anonymous:1873:BNBp

[4417] Anonymous. Books: New books and publications. Scientific American, 28(23):360, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
360.pdf.

Anonymous:1873:BPw

[4418] Anonymous. Business and personal. Scientific American, 28(23):362, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-362.pdf.

Anonymous:1873:NQv

[4419] Anonymous. Notes and queries. Scientific American, 28(23):362, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-362a.pdf.

Anonymous:1873:IIv

[4420] Anonymous. Index of inventions. Scientific American, 28(23):363–364, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-363a.pdf.

Anonymous:1873:CMM

[4421] Anonymous. Chain making machine. Scientific American, 28(23):351, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-351.pdf.

Anonymous:1873:LBa

[4422] Anonymous. Leather belting. Scientific American, 28(23):352, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-352.pdf.

Anonymous:1873:SE

[4423] Anonymous. Scientific executions. Scientific American, 28(23):352, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-352a.pdf.

Anonymous:1873:LAE

[4424] Anonymous. The latest Arctic explorations — the remarkable escape of the Polaris party. Scientific American, 28(23):352–353, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-352b.pdf.

Anonymous:1873:AEV

[4425] Anonymous. American exhibits at the Vienna Exposition. Scientific American, 28(23):353, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
353.pdf.

Anonymous:1873:SPIo

[4426] Anonymous. Scientific and practical information. Scientific American, 28(23):353, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
353a.pdf.

Anonymous:1873:WMT

[4427] Anonymous. Water main tapping machine. Scientific American, 28(23):354, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
354.pdf.

Anonymous:1873:ICN

[4428] Anonymous. Improved cutting nippers. Scientific American, 28(23):354, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-354a.pdf.

Anonymous:1873:MDB

[4429] Anonymous. Miner’s drill bar and bits. Scientific American, 28(23):354, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-354b.pdf.

Anonymous:1873:IPH

[4430] Anonymous. Improved plow handle. Scientific American, 28(23):354, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-354c.pdf.

Anonymous:1873:CLBa

[4431] Anonymous. A collapsing life boat. Scientific American, 28(23):355, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-355.pdf.

Anonymous:1873:CPDb

[4432] Anonymous. Chandler’s patent dip pipe. Scientific American, 28(23):355, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
355a.pdf.

Anonymous:1873:NDSa

[4433] Anonymous. A new deep sea thermometer. Scientific American, 28(23):355, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
355b.pdf.

Anonymous:1873:WEIc

[4434] Anonymous. The wonders of the egg. — IV. Scientific American, 28(23):356, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
356.pdf.

Anonymous:1873:SND

[4435] Anonymous. Sanitary notes — drainage for health. Scientific American, 28(23):356–357, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
356a.pdf.

Anonymous:1873:ANe

[4436] Anonymous. Astronomical notes. Scientific American, 28(23):357, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-357a.pdf.

Anonymous:1873:IRSb

[4437] Anonymous. Improved rotary steam engine. Scientific American, 28(23):358, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
358.pdf.

Anonymous:1873:FFS

[4438] Anonymous. Feed and filter for steam boilers. Scientific American, 28(23):358, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
358a.pdf.

Anonymous:1873:CCSb

[4439] Anonymous. Country cottage and stable. Scientific American, 28(23):359–360, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
359.pdf.

Anonymous:1873:PODk

[4440] Anonymous. Patent Office decisions. Scientific American, 28(23):360, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-360a.pdf.

Anonymous:1873:RAFw

[4441] Anonymous. Recent American and foreign patents. Scientific American, 28(23):360–361, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
360b.pdf.

Anonymous:1873:Cbj

[4442] Anonymous. Copyrights. Scientific American, 28(23):362, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n23/pdf/scientificamerican06071873-362b.pdf.

Anonymous:1873:ACy

[4443] Anonymous. Answers to correspondents. Scientific American, 28 (23):362–363, June 7, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
362c.pdf.

Anonymous:1873:CRs (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n23/pdf/scientificamerican06071873-
363.pdf.

Anonymous:1873:Cbk

[4445] Anonymous. Correspondence. Scientific American, 28(24):373, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-373.pdf.

Anonymous:1873:BNBq

[4446] Anonymous. Books: New books and publications. Scientific American, 28(24):376, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
376b.pdf.

Anonymous:1873:BPx

[4447] Anonymous. Business and personal. Scientific American, 28(24):378, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-378.pdf.

Anonymous:1873:NQw

[4448] Anonymous. Notes and queries. Scientific American, 28(24):378, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-378b.pdf.

Anonymous:1873:IIw

[4449] Anonymous. Index of inventions. Scientific American, 28(24):379–380, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-379b.pdf.

Anonymous:1873:ASE

[4450] Anonymous. Automatic steam engine. Scientific American, 28 (24):367–368, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
367.pdf.

Anonymous:1873:OCD

[4451] Anonymous. Our coal deposits — their formation and extent. Scientific American, 28(24):368, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
368.pdf.

Anonymous:1873:EFA

[4452] Anonymous. Electrical fire arms. Scientific American, 28(24):368, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-368a.pdf.

Anonymous:1873:BC

[4453] Anonymous. Bisulphide of carbon. Scientific American, 28(24):368, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-368b.pdf.

Anonymous:1873:GPB

[4454] Anonymous. A gas-propelled boat. Scientific American, 28(24):368, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-368c.pdf.

Anonymous:1873:OLT

[4455] Anonymous. An objection to the large telescope answered. Scientific American, 28(24):368–369, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
368d.pdf.

Anonymous:1873:VEc

[4456] Anonymous. The Vienna Exposition. Scientific American, 28(24):369, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-369.pdf.

Anonymous:1873:SPIp

[4457] Anonymous. Scientific and practical information. Scientific American, 28(24):369, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
369a.pdf.

Anonymous:1873:PE

[4458] Anonymous. Pumping engine. Scientific American, 28(24):370, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-370.pdf.

Anonymous:1873:WHAb

[4459] Anonymous. Wenham’s heated air engine. Scientific American, 28(24):370, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
370a.pdf.

Anonymous:1873:IGH

[4460] Anonymous. Improved garden hoe. Scientific American, 28(24):371, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-371.pdf.

Anonymous:1873:FFb

[4461] Anonymous. Fireproof floor. Scientific American, 28(24):371, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-371a.pdf.

Anonymous:1873:BD

[4462] Anonymous. The Belgian dray. Scientific American, 28(24):371, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-371b.pdf.

Anonymous:1873:MPS

[4463] Anonymous. M’Grath’s patent side stick and quoins. Scientific American, 28(24):371, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
371c.pdf.

Anonymous:1873:SNFb

[4464] Anonymous. Sanitary notes — food and its preparation. Scientific American, 28(24):372, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
372.pdf.

Anonymous:1873:PCVa

[4465] Anonymous. Positive circulating vertical tube radiator. Scientific American, 28(24):374, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
374.pdf.

Anonymous:1873:NVA

[4466] Anonymous. The needle valve automatic oil feeder. Scientific American, 28(24):374, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
374a.pdf.

Anonymous:1873:SPR

[4467] Anonymous. Scotton’s patent round tree box. Scientific American, 28(24):374, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
374b.pdf.

Anonymous:1873:PSRb

[4468] Anonymous. Portable steam renovator. Scientific American, 28(24):375, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-375.pdf.

Anonymous:1873:TS

[4469] Anonymous. Tannate of soda. Scientific American, 28(24):375, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-375a.pdf.

Anonymous:1873:APBa

[4470] Anonymous. Adjustable pulley bolt. Scientific American, 28(24):375, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-375b.pdf.

Anonymous:1873:WS

[4471] Anonymous. Whiffletree stub. Scientific American, 28(24):375, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-375c.pdf.

Anonymous:1873:SCb

[4472] Anonymous. The Southern Canal. Scientific American, 28(24):376, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-376.pdf.

Anonymous:1873:DCq

[4473] Anonymous. Decisions of the courts. Scientific American, 28(24):376, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-376a.pdf.

Anonymous:1873:RAFx

[4474] Anonymous. Recent American and foreign patents. Scientific American, 28(24):377, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
377.pdf.

Anonymous:1873:Cbl

[4475] Anonymous. Copyrights. Scientific American, 28(24):378, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n24/pdf/scientificamerican06141873-378a.pdf.

Anonymous:1873:ACz

[4476] Anonymous. Answers to correspondents. Scientific American, 28 (24):378–379, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
378c.pdf.

Anonymous:1873:CRt

[4477] Anonymous. Communications received. Scientific American, 28 (24):379, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
379.pdf.

Anonymous:1873:BMIf

[4478] Anonymous. Best mode of introducing new inventions. Scientific American, 28(24):379, June 14, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n24/pdf/scientificamerican06141873-
379a.pdf.

Anonymous:1873:Cbm

[4479] Anonymous. Correspondence. Scientific American, 28(25):388–389, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-388.pdf.

Anonymous:1873:BPy

[4480] Anonymous. Business and personal. Scientific American, 28(25):395, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-395.pdf.

Anonymous:1873:NQx

[4481] Anonymous. Notes and queries. Scientific American, 28(25):395, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-395a.pdf.

Anonymous:1873:IIx

[4482] Anonymous. Index of inventions. Scientific American, 28(25):396, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-396a.pdf.

Anonymous:1873:NDSb

[4483] Anonymous. New double spindle drill. Scientific American, 28(25):383, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-383.pdf.

Anonymous:1873:ICV

[4484] Anonymous. Improved cooking vessel. Scientific American, 28 (25):383–384, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
383a.pdf.

Anonymous:1873:DAb

[4485] Anonymous. The digestive apparatus. Scientific American, 28(25):384, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-384.pdf.

Anonymous:1873:TC

[4486] Anonymous. Thick cylinders. Scientific American, 28(25):384, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-384a.pdf.

Anonymous:1873:PF

[4487] Anonymous. Pneumatic foundations. Scientific American, 28(25):384, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-384b.pdf.

Anonymous:1873:PG

[4488] Anonymous. The preparation of gelatin. Scientific American, 28 (25):384, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
384c.pdf.

Anonymous:1873:SSc

[4489] Anonymous. The spectroscope simplified. Scientific American, 28 (25):384–385, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
384d.pdf.

Anonymous:1873:CIS

[4490] Anonymous. The Chicago Inter-State Industrial Exposition. Scientific American, 28(25):385, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
385.pdf.

Anonymous:1873:FBN

[4491] Anonymous. Freight business of New York City across the North River. Scientific American, 28(25):385, June 21, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n25/pdf/scientificamerican06211873-385a.pdf.

Anonymous:1873:AIF

[4492] Anonymous. The American Institute Fair. Scientific American, 28(25):385, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
385b.pdf.

Anonymous:1873:SPIq

[4493] Anonymous. Scientific and practical information. Scientific American, 28(25):385, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
385c.pdf.

Anonymous:1873:S

[4494] Anonymous. Silkworms. Scientific American, 28(25):386, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-386.pdf.

Anonymous:1873:IPc

[4495] Anonymous. Improved plow. Scientific American, 28(25):387, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-387.pdf.

Anonymous:1873:ISCb

[4496] Anonymous. Iron ship construction. Scientific American, 28 (25):387–388, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
387a.pdf.

Anonymous:1873:NPSb

[4497] Anonymous. A new preventive for slipping belts. Scientific American, 28(25):389, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
389.pdf.

Anonymous:1873:EGA

[4498] Anonymous. Emery grinder with adjustable table. Scientific American, 28(25):390, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
390.pdf.

Anonymous:1873:LE

[4499] Anonymous. Lock envelope. Scientific American, 28(25):390, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-390a.pdf.

Anonymous:1873:ISSb

[4500] Anonymous. Improved spindle step. Scientific American, 28(25):390, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n25/pdf/scientificamerican06211873-390b.pdf.

Anonymous:1873:CPEa

[4501] Anonymous. Car and passenger elevator. Scientific American, 28 (25):391–392, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
391.pdf.

Anonymous:1873:SNM

[4502] Anonymous. Sanitary notes — milk and its adulterations. Scientific American, 28(25):392–393, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
392.pdf.

Anonymous:1873:RAFy

[4503] Anonymous. Recent American and foreign patents. Scientific American, 28(25):394, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
394.pdf.

Anonymous:1873:ACba

[4504] Anonymous. Answers correspondents. Scientific American, 28 (25):395–396, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
395b.pdf.

Anonymous:1873:CRu

[4505] Anonymous. Communications received. Scientific American, 28 (25):396, June 21, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n25/pdf/scientificamerican06211873-
396.pdf.

Anonymous:1873:Cbn

[4506] Anonymous. Correspondence. Scientific American, 28(26):404, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-404.pdf.

Anonymous:1873:BPz

[4507] Anonymous. Business and personal. Scientific American, 28(26):409, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-409.pdf.

Anonymous:1873:IIy

[4508] Anonymous. Index of inventions. Scientific American, 28(26):410, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-410a.pdf.

Anonymous:1873:Ia

[4509] Anonymous. Index. Scientific American, 28(26):411–412, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-411.pdf.

Anonymous:1873:IFBa

[4510] Anonymous. Improved force blast rotary blower. Scientific American, 28(26):399–400, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
399.pdf.

Anonymous:1873:EAH

[4511] Anonymous. End of another half year. Scientific American, 28(26):400, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-400.pdf.

Anonymous:1873:NSRc

[4512] Anonymous. Necessity for sanitary reform. Scientific American, 28(26):400, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
400a.pdf.

Anonymous:1873:RRO

[4513] Anonymous. Renewal of the reward of one hundred thousand dollars. Scientific American, 28(26):400, June 28, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v28/
n26/pdf/scientificamerican06281873-400b.pdf.

Anonymous:1873:CTC

[4514] Anonymous. Comets — their character and source. Scientific American, 28(26):400, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
400c.pdf.

Anonymous:1873:PV

[4515] Anonymous. A perfect vacuum. Scientific American, 28(26):400–401, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-400d.pdf.

Anonymous:1873:TAW

[4516] Anonymous. Time around the world. Scientific American, 28(26):401, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-401.pdf.

Anonymous:1873:TPR

[4517] Anonymous. The Texas Pacific Railroad. Scientific American, 28(26):401, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
401a.pdf.

Anonymous:1873:SPIr

[4518] Anonymous. Scientific and practical information. Scientific American, 28(26):401, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
401b.pdf.

Anonymous:1873:NRA

[4519] Anonymous. New rifle attachment for fowling pieces. Scientific American, 28(26):402, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
402.pdf.

Anonymous:1873:NMC

[4520] Anonymous. New mode of chair construction. Scientific American, 28(26):402, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
402a.pdf.

Anonymous:1873:BB

[4521] Anonymous. Blacking box. Scientific American, 28(26):402, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-402b.pdf.

Anonymous:1873:CEMa

[4522] Anonymous. Carre’s electrical machine. Scientific American, 28 (26):402, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
402c.pdf.

Anonymous:1873:TH

[4523] Anonymous. Tobacco hanger. Scientific American, 28(26):402, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-402d.pdf.

Anonymous:1873:SBP

[4524] Anonymous. Straw-burning portable engine. Scientific American, 28(26):403–404, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
403.pdf.

Anonymous:1873:AFQ

[4525] Anonymous. Alcohol from flint and quartz. Scientific American, 28(26):404–405, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
404a.pdf.

Anonymous:1873:SNS

[4526] Anonymous. Sanitary notes — sewerage and sewage. Scientific American, 28(26):405, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
405.pdf.

Anonymous:1873:BPba

[4527] Anonymous. Bee protector. Scientific American, 28(26):406, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-406.pdf.

Anonymous:1873:BPCa

[4528] Anonymous. Brewster’s patent combination tallow cup. Scientific American, 28(26):406, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
406a.pdf.

Anonymous:1873:STa

[4529] Anonymous. Soldering tool. Scientific American, 28(26):406, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-406b.pdf.

Anonymous:1873:BRE

[4530] Anonymous. Boynton’s revolving evaporator. Scientific American, 28(26):406, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
406c.pdf.

Anonymous:1873:PAF

[4531] Anonymous. Parlor aquarium and fountain. Scientific American, 28(26):407, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
407.pdf.

Anonymous:1873:RAFz

[4532] Anonymous. Recent American and foreign patents. Scientific American, 28(26):408, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
408.pdf.

Anonymous:1873:DCr

[4533] Anonymous. Decisions of the courts. Scientific American, 28(26):408, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v28/n26/pdf/scientificamerican06281873-408a.pdf.

Anonymous:1873:ACbb

[4534] Anonymous. Answers to correspondents. Scientific American, 28 (26):409–410, June 28, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
409a.pdf.

Anonymous:1873:CRv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v28/n26/pdf/scientificamerican06281873-
410.pdf.

Anonymous:1873:Cbo

[4536] Anonymous. Correspondence. Scientific American, 29(1):4, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-4.pdf.

Anonymous:1873:BNBr

[4537] Anonymous. Books: New books and publications. Scientific American, 29(1):9, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
9.pdf.

Anonymous:1873:BPbb

[4538] Anonymous. Business and personal. Scientific American, 29(1):11, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-11.pdf.

Anonymous:1873:NQy

[4539] Anonymous. Notes and queries. Scientific American, 29(1):11, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-11a.pdf.

Anonymous:1873:IIz

[4540] Anonymous. Index of inventions. Scientific American, 29(1):12, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-12a.pdf.

Anonymous:1873:PPFa

[4541] Anonymous. The preparation of peat for fuel. Scientific American, 29(1):1–2, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
1.pdf.

Anonymous:1873:HLJ

[4542] Anonymous. The Herculean lifting jack. Scientific American, 29(1):2, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-2.pdf.

Anonymous:1873:LBb

[4543] Anonymous. Life in the beehive. Scientific American, 29(1):3–4, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-3.pdf.

Anonymous:1873:ANf

[4544] Anonymous. Astronomical notes. Scientific American, 29(1):5, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-5.pdf.

Anonymous:1873:NVEa

[4545] Anonymous. Notes from the Vienna Exposition. Scientific American, 29(1):5, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
5a.pdf.

Anonymous:1873:IHL

[4546] Anonymous. Improved hub lathe. Scientific American, 29(1):6, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-6.pdf.

Anonymous:1873:IFG

[4547] Anonymous. Improved farm gate. Scientific American, 29(1):6, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-6a.pdf.

Anonymous:1873:IUP

[4548] Anonymous. An imposition upon the public. Scientific American, 29(1): 7, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-7.pdf.

Anonymous:1873:STE

[4549] Anonymous. Screw threads — English and American proportions. Scientific American, 29(1):7, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
7a.pdf.

Anonymous:1873:FLa

[4550] Anonymous. Fire ladders. Scientific American, 29(1):7, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-7b.pdf.

Anonymous:1873:PWR

[4551] Anonymous. Progress of woman’s rights. Scientific American, 29(1):7, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-7c.pdf.

Anonymous:1873:PCc

[4552] Anonymous. A patent Congress. Scientific American, 29(1):7–8, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-7d.pdf.

Anonymous:1873:SPIs

[4553] Anonymous. Scientific and practical information. Scientific American, 29(1):8, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
8.pdf.

Anonymous:1873:RAFba

[4554] Anonymous. Recent American and foreign patents. Scientific American, 29(1):9–10, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n1/pdf/scientificamerican07051873-
9a.pdf.

Anonymous:1873:ACbc

[4555] Anonymous. Answers to correspondents. Scientific American, 29(1):11– 12, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-11b.pdf.

Anonymous:1873:CRw

[4556] Anonymous. Communications received. Scientific American, 29(1):12, July 5, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n1/pdf/scientificamerican07051873-12.pdf.

Anonymous:1873:Cbp

[4557] Anonymous. Correspondence. Scientific American, 29(2):21, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-21.pdf.

Anonymous:1873:BNBs

[4558] Anonymous. Books: New books and publications. Scientific American, 29(2):24, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
24b.pdf.

Anonymous:1873:BPbc

[4559] Anonymous. Business and personal. Scientific American, 29(2):27, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-27.pdf.

Anonymous:1873:NQz

[4560] Anonymous. Notes and queries. Scientific American, 29(2):27, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-27a.pdf.

Anonymous:1873:IIba

[4561] Anonymous. Index of inventions. Scientific American, 29(2):28, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-28.pdf.

Anonymous:1873:SPD

[4562] Anonymous. Steam plowing by direct traction. Scientific American, 29(2):15, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
15.pdf.

Anonymous:1873:PNa

[4563] Anonymous. Publishers’ notice. Scientific American, 29(2):16, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-16.pdf.

Anonymous:1873:MPa

[4564] Anonymous. The measurement of power. Scientific American, 29(2):16, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-16a.pdf.

Anonymous:1873:CPf

[4565] Anonymous. The coming pavement. Scientific American, 29(2):16, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-16b.pdf.

Anonymous:1873:LCT

[4566] Anonymous. Letters from Commissioner Thurston. Scientific American, 29(2):16, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
16c.pdf.

Anonymous:1873:DAc

[4567] Anonymous. The digestive apparatus. Scientific American, 29(2):16–17, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-16d.pdf.

Anonymous:1873:MAC

[4568] Anonymous. Mordants for aniline colors. Scientific American, 29(2):17, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-17.pdf.

Anonymous:1873:FCF

[4569] Anonymous. Fish culture by farmers. Scientific American, 29(2):17, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-17a.pdf.

Anonymous:1873:RAE

[4570] Anonymous. The recent Arctic Expedition. Scientific American, 29(2):17–18, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
17b.pdf.

Anonymous:1873:MPb

[4571] Anonymous. The mechanical pigeon. Scientific American, 29(2):18, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-18.pdf.

Anonymous:1873:CBq

[4572] Anonymous. The common beetle. Scientific American, 29(2):18, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-18a.pdf.

Anonymous:1873:NVEb

[4573] Anonymous. Notes from the Vienna Exposition. Scientific American, 29(2):19, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
19.pdf.

Anonymous:1873:GELa

[4574] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(2):20, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-20.pdf.

Anonymous:1873:SLF

[4575] Anonymous. Successful laying of the fifth Atlantic cable. Scientific American, 29(2):21, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
21a.pdf.

Anonymous:1873:IPD

[4576] Anonymous. Improved portable dumping car. Scientific American, 29(2):22, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
22.pdf.

Anonymous:1873:TCEb

[4577] Anonymous. The Tanite Company’s emery grinder “e.”. Scientific American, 29(2):22, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
22a.pdf.

Anonymous:1873:CSCa

[4578] Anonymous. Composite ship construction. Scientific American, 29(2):23–24, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
23.pdf.

Anonymous:1873:NTEb

[4579] Anonymous. A new thermo-electric battery. Scientific American, 29(2):24, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
24.pdf.

Anonymous:1873:SCc

[4580] Anonymous. The Southern Canal. Scientific American, 29(2):24, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-24a.pdf.

Anonymous:1873:DCs

[4581] Anonymous. Decisions of the courts. Scientific American, 29(2):24, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-24c.pdf.

Anonymous:1873:RAFbb

[4582] Anonymous. Recent American and foreign patents. Scientific American, 29(2):24–26, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
24d.pdf.

Anonymous:1873:ACbd

[4583] Anonymous. Answers to correspondents. Scientific American, 29(2):27, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-27b.pdf.

Anonymous:1873:CRx

[4584] Anonymous. Communications received. Scientific American, 29(2):27, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-27c.pdf.

Anonymous:1873:SPFa

[4585] Anonymous. Schedule of patent fees. Scientific American, 29(2):28, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-28a.pdf.

Anonymous:1873:VPHa

[4586] Anonymous. Value of patents and how to obtain them. Scientific American, 29(2):28, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n2/pdf/scientificamerican07121873-
28b.pdf.

Anonymous:1873:Ib

[4587] Anonymous. To investors. Scientific American, 29(2):28, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-28c.pdf.

Anonymous:1873:SAa

[4588] Anonymous. Scientific American. Scientific American, 29(2):30, July 12, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n2/pdf/scientificamerican07121873-30.pdf.

Anonymous:1873:Cbr

[4589] Anonymous. Correspondence. Scientific American, 29(3):37, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-37.pdf.

Anonymous:1873:BPbd

[4590] Anonymous. Business and personal. Scientific American, 29(3):43, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-43.pdf.

Anonymous:1873:NQba

[4591] Anonymous. Notes and queries. Scientific American, 29(3):43, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-43a.pdf.

Anonymous:1873:IIbb

[4592] Anonymous. Index of inventions. Scientific American, 29(3):44, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-44a.pdf.

Anonymous:1873:RIT

[4593] Anonymous. Recent improvements in torpedo warfare. Scientific American, 29(3):31–32, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
31.pdf.

Anonymous:1873:PNb

[4594] Anonymous. Publishers’ notice. Scientific American, 29(3):32, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-32.pdf.

Anonymous:1873:STb

[4595] Anonymous. Street travel. Scientific American, 29(3):32, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-32a.pdf.

Anonymous:1873:IMN

[4596] Anonymous. Invention the mother of necessity. Scientific American, 29(3):32, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
32b.pdf.

Anonymous:1873:BBO

[4597] Anonymous. Boilers and boiler owners. Scientific American, 29(3):32, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-32c.pdf.

Anonymous:1873:URA

[4598] Anonymous. Underground railways in American cities. Scientific American, 29(3):32–33, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
32d.pdf.

Anonymous:1873:SPc

[4599] Anonymous. Snake poisons. Scientific American, 29(3):33, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-33.pdf.

Anonymous:1873:EBa

[4600] Anonymous. To Europe in a balloon. Scientific American, 29(3):33, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-33a.pdf.

Anonymous:1873:NVEc

[4601] Anonymous. Notes from the Vienna Exposition. Scientific American, 29(3):34, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
34.pdf.

Anonymous:1873:NFM

[4602] Anonymous. New form for metallic telegraph poles. Scientific American, 29(3):35, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
35.pdf.

Anonymous:1873:CSM

[4603] Anonymous. Cummings’ steam mowing, reaping, and thrashing machine. Scientific American, 29(3):35, July 19, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n3/pdf/scientificamerican07191873-35a.pdf.

Anonymous:1873:EGL

[4604] Anonymous. Electric gas lighter. Scientific American, 29(3):35, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-35b.pdf.

Anonymous:1873:SGI

[4605] Anonymous. Specific gravity indicator. Scientific American, 29 (3):35–36, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
35c.pdf.

Anonymous:1873:GLM

[4606] Anonymous. The great leopard moth. Scientific American, 29(3):36, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-36.pdf.

Anonymous:1873:RGEa

[4607] Anonymous. En route to the Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(3):36, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
36a.pdf.

Anonymous:1873:ICa

[4608] Anonymous. Improved cultivator. Scientific American, 29(3):38, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-38.pdf.

Anonymous:1873:IHC

[4609] Anonymous. An improved horse collar brace. Scientific American, 29(3):38, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
38a.pdf.

Anonymous:1873:NCH

[4610] Anonymous. Novel claw hatchet. Scientific American, 29(3):38, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-38b.pdf.

Anonymous:1873:NHPb

[4611] Anonymous. New heel plate for boots and shoes. Scientific American, 29(3):38, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
38c.pdf.

Anonymous:1873:NFE

[4612] Anonymous. A new fire escape. Scientific American, 29(3):38–39, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-38d.pdf.

Anonymous:1873:RWB

[4613] Anonymous. Robert Wilhelm Bunsen. Scientific American, 29 (3):39–40, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
39.pdf.

Anonymous:1873:NEL

[4614] Anonymous. The new electric light in London. Scientific American, 29(3):40, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
40.pdf.

Anonymous:1873:LL

[4615] Anonymous. Light and life. Scientific American, 29(3):40, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-40a.pdf.

Anonymous:1873:EBb

[4616] Anonymous. To Europe in a balloon. Scientific American, 29(3):41, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-41.pdf.

Anonymous:1873:RAFbc

[4617] Anonymous. Recent American and foreign patents. Scientific American, 29(3):41–42, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
41a.pdf.

Anonymous:1873:VPHb

[4618] Anonymous. Value of patents, and how to obtain them practical hints to inventors. Scientific American, 29(3):42, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-42.pdf.

Anonymous:1873:ACbe

[4619] Anonymous. Answers to correspondents. Scientific American, 29 (3):43–44, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n3/pdf/scientificamerican07191873-
43b.pdf.

Anonymous:1873:CRy

[4620] Anonymous. Communications received. Scientific American, 29(3):44, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-44.pdf.

Anonymous:1873:SPFb

[4621] Anonymous. Schedule of patent fees. Scientific American, 29(3):44, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-44b.pdf.

Anonymous:1873:Id

[4622] Anonymous. To investors. Scientific American, 29(3):45, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-45.pdf.

Anonymous:1873:If

[4623] Anonymous. Inventors. Scientific American, 29(3):45, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-45a.pdf.

Anonymous:1873:SAb

[4624] Anonymous. Scientific American. Scientific American, 29(3):46, July 19, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n3/pdf/scientificamerican07191873-46.pdf.

Anonymous:1873:Cbt

[4625] Anonymous. Correspondence. Scientific American, 29(4):51–52, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-51a.pdf.

Anonymous:1873:BPbe

[4626] Anonymous. Business and personal. Scientific American, 29(4):59, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-59.pdf.

Anonymous:1873:NQbb

[4627] Anonymous. Notes and queries. Scientific American, 29(4):59, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-59a.pdf.

Anonymous:1873:IIbc

[4628] Anonymous. Index of inventions. Scientific American, 29(4):60, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-60a.pdf.

Anonymous:1873:BCR

[4629] Anonymous. The Baltimore concentric rotary engine. Scientific American, 29(4):47–48, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
47.pdf.

Anonymous:1873:PNc

[4630] Anonymous. Publishers’ notice. Scientific American, 29(4):48, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-48.pdf.

Anonymous:1873:FWP

[4631] Anonymous. Friction of water in pipes. Scientific American, 29 (4):48–49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
48a.pdf.

Anonymous:1873:ZL

[4632] Anonymous. The zodiacal light. Scientific American, 29(4):49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-49.pdf.

Anonymous:1873:GGW

[4633] Anonymous. The greatest gas well in the world. Scientific American, 29(4):49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
49a.pdf.

Anonymous:1873:EBc

[4634] Anonymous. To Europe by balloon. Scientific American, 29(4):49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-49b.pdf.

Anonymous:1873:NSRd

[4635] Anonymous. A new substitute for rubber. Scientific American, 29(4):49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-49c.pdf.

Anonymous:1873:SPIt

[4636] Anonymous. Scientific and practical information. Scientific American, 29(4):49, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
49d.pdf.

Anonymous:1873:GG

[4637] Anonymous. The Gardner governor. Scientific American, 29(4):50, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-50.pdf.

Anonymous:1873:IWMa

[4638] Anonymous. Improved washing machine. Scientific American, 29(4):50, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-50a.pdf.

Anonymous:1873:TFA

[4639] Anonymous. The Tunnicliffe fire alarm. Scientific American, 29(4):50, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-50b.pdf.

Anonymous:1873:ITMa

[4640] Anonymous. The Isabella tiger moth. Scientific American, 29(4):51, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-51.pdf.

Anonymous:1873:NTB

[4641] Anonymous. New Turkish bath. Scientific American, 29(4):52, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-52.pdf.

Anonymous:1873:RGEb

[4642] Anonymous. En route to the Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(4):52, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
52a.pdf.

Anonymous:1873:GFB

[4643] Anonymous. On girders and floor beams. — the effect of crossbridging. Scientific American, 29(4):53, July 26, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n4/pdf/scientificamerican07261873-53.pdf.

Anonymous:1873:PVH

[4644] Anonymous. Patent ventilating and heating stove. Scientific American, 29(4):54, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
54.pdf.

Anonymous:1873:ACCb

[4645] Anonymous. Automatic car coupler. Scientific American, 29(4):54, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-54a.pdf.

Anonymous:1873:GB

[4646] Anonymous. Great balloons. Scientific American, 29(4):55–57, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-55.pdf.

Anonymous:1873:RAFbd

[4647] Anonymous. Recent American and foreign patents. Scientific American, 29(4):57–58, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
57.pdf.

Anonymous:1873:VPHc

[4648] Anonymous. Value of patents and how to obtain them practical hints to inventors. Scientific American, 29(4):58, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-58.pdf.

Anonymous:1873:ACbf

[4649] Anonymous. Answers to correspondence. Scientific American, 29(4):59–60, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n4/pdf/scientificamerican07261873-
59b.pdf.

Anonymous:1873:CRz

[4650] Anonymous. Communications received. Scientific American, 29(4):60, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-60.pdf.

Anonymous:1873:SPFc

[4651] Anonymous. Schedule of patent fees. Scientific American, 29(4):60, July 26, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n4/pdf/scientificamerican07261873-60b.pdf.

Anonymous:1873:Cbu

[4652] Anonymous. Correspondence. Scientific American, 29(5):68, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-68a.pdf.

Anonymous:1873:BNBt

[4653] Anonymous. Books: New books and publications. Scientific American, 29(5):73, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
73a.pdf.

Anonymous:1873:BPbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-75.pdf.

Anonymous:1873:NQbc

[4655] Anonymous. Notes and queries. Scientific American, 29(5):75, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-75a.pdf.

Anonymous:1873:IIbd

[4656] Anonymous. Index of inventions. Scientific American, 29(5):76, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-76a.pdf.

Anonymous:1873:PDF

[4657] Anonymous. A public drinking fountain. Scientific American, 29(5):63–64, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
63.pdf.

Anonymous:1873:LH

[4658] Anonymous. Latent heat. Scientific American, 29(5):64, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-64.pdf.

Anonymous:1873:FTE

[4659] Anonymous. French telegraphy at the Exposition. Scientific American, 29(5):64, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
64a.pdf.

Anonymous:1873:AAE

[4660] Anonymous. Another Arctic Expedition begun. Scientific American, 29(5):64, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
64b.pdf.

Anonymous:1873:AFE

[4661] Anonymous. An Austrian farmer’s entertainment. Scientific American, 29(5):64–65, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
64c.pdf.

Anonymous:1873:UZ

[4662] Anonymous. The use of zinc. Scientific American, 29(5):65, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-65.pdf.

Anonymous:1873:STM

[4663] Anonymous. Simple tests for minerals. Scientific American, 29(5):65, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-65a.pdf.

Anonymous:1873:SPIu

[4664] Anonymous. Scientific and practical information. Scientific American, 29(5):65, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
65b.pdf.

Anonymous:1873:SBC

[4665] Anonymous. A substitute for the Bell cord. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
66.pdf.

Anonymous:1873:IBa

[4666] Anonymous. Improvement in bricks. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-66a.pdf.

Anonymous:1873:NSP

[4667] Anonymous. New safety pump for sea going steamers. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
66b.pdf.

Anonymous:1873:IPe

[4668] Anonymous. Improved propeller. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-66c.pdf.

Anonymous:1873:IOB

[4669] Anonymous. Improved ox bow. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-66d.pdf.

Anonymous:1873:PPFb

[4670] Anonymous. Patent pig feeder. Scientific American, 29(5):66, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-66e.pdf.

Anonymous:1873:PSc

[4671] Anonymous. The Pernyi silkworm. Scientific American, 29(5):67, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-67.pdf.

Anonymous:1873:NTH

[4672] Anonymous. New tool holder for the slide rest. Scientific American, 29(5):67, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
67a.pdf.

Anonymous:1873:MV

[4673] Anonymous. On the manufacture of vinegar. Scientific American, 29(5):68, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
68.pdf.

Anonymous:1873:GELb

[4674] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(5):69, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-69.pdf.

Anonymous:1873:ICPa

[4675] Anonymous. Improved chain propeller. Scientific American, 29(5):70, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-70.pdf.

Anonymous:1873:TAS

[4676] Anonymous. Tucking attachment for sewing machines. Scientific American, 29(5):70, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
70a.pdf.

Anonymous:1873:CCR

[4677] Anonymous. Combined cradle and rocking chair. Scientific American, 29(5):70, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
70b.pdf.

Anonymous:1873:BL

[4678] Anonymous. Baron Von Liebig. Scientific American, 29(5):71, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-71.pdf.

Anonymous:1873:ENc

[4679] Anonymous. Engineering notes. Scientific American, 29(5):72–73, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-72.pdf.

Anonymous:1873:DCt

[4680] Anonymous. Decisions of the courts. Scientific American, 29(5):73, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-73.pdf.

Anonymous:1873:RAFbe

[4681] Anonymous. Recent American and foreign patents. Scientific American, 29(5):73–74, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
73b.pdf.

Anonymous:1873:VPb

[4682] Anonymous. Value of patents. Scientific American, 29(5):74, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-74.pdf.

Anonymous:1873:ACbg

[4683] Anonymous. Answers to correspondents. Scientific American, 29(5):75–76, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n5/pdf/scientificamerican08021873-
75b.pdf.

Anonymous:1873:CRba

[4684] Anonymous. Communications received. Scientific American, 29(5):76, August 2, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n5/pdf/scientificamerican08021873-76.pdf.

Anonymous:1873:Cbv

[4685] Anonymous. Correspondence. Scientific American, 29(6):84, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-84.pdf.

Anonymous:1873:BPbg

[4686] Anonymous. Business and personal. Scientific American, 29(6):91, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-91.pdf.

Anonymous:1873:IIbe

[4687] Anonymous. Index of inventions. Scientific American, 29(6):92, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-92a.pdf.

Anonymous:1873:IIRa

[4688] Anonymous. Improved ice and refrigerating machine. Scientific American, 29(6):79–80, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
79.pdf.

Anonymous:1873:IDO

[4689] Anonymous. Is it a dastardly outrage? Scientific American, 29(6):80, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-80.pdf.

Anonymous:1873:VSa

[4690] Anonymous. The Vienna show. Scientific American, 29(6):80, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-80a.pdf.

Anonymous:1873:PCd

[4691] Anonymous. Progress of cabs. Scientific American, 29(6):80, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-80b.pdf.

Anonymous:1873:GUI

[4692] Anonymous. A great unrecognized inventor. Scientific American, 29(6):80, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
80c.pdf.

Anonymous:1873:CBPa

[4693] Anonymous. Concrete for building purposes. Scientific American, 29(6):80–81, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
80d.pdf.

Anonymous:1873:MVC

[4694] Anonymous. Madder, vegetable and chemical. Scientific American, 29(6):81, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
81.pdf.

Anonymous:1873:SSAa

[4695] Anonymous. The star shower of August 10. Scientific American, 29(6):81, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
81a.pdf.

Anonymous:1873:PSSa

[4696] Anonymous. Properties of Saturated steam. Scientific American, 29(6):81, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
81b.pdf.

Anonymous:1873:SPIv

[4697] Anonymous. Scientific and practical information. Scientific American, 29(6):81, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
81c.pdf.

Anonymous:1873:NAS

[4698] Anonymous. A new application of spectrum analysis. Scientific American, 29(6):82–83, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
82.pdf.

Anonymous:1873:CSEa

[4699] Anonymous. Cheese skipper exterminator. Scientific American, 29(6):83, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
83.pdf.

Anonymous:1873:IMR

[4700] Anonymous. Improved milk refrigerator. Scientific American, 29(6):83, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-83a.pdf.

Anonymous:1873:GELc

[4701] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(6):85, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-85.pdf.

Anonymous:1873:PSV

[4702] Anonymous. Pneumatic screw ventilator for cars. Scientific American, 29(6):86, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
86.pdf.

Anonymous:1873:APBb

[4703] Anonymous. Abbe’s patent bolt forging machine. Scientific American, 29(6):86, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
86a.pdf.

Anonymous:1873:CSa

[4704] Anonymous. The Cecropia silkworm. Scientific American, 29(6):87–88, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-87.pdf.

Anonymous:1873:DABa

[4705] Anonymous. Double-acting bucket plunger steam pump. Scientific American, 29(6):88–89, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
88.pdf.

Anonymous:1873:DCu

[4706] Anonymous. Decisions of the courts. Scientific American, 29(6):89, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-89.pdf.

Anonymous:1873:RAFbf

[4707] Anonymous. Recent American and foreign patents. Scientific American, 29(6):89–90, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
89a.pdf.

Anonymous:1873:VPHd

[4708] Anonymous. Value of patents, and how to obtain them. Scientific American, 29(6):90, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
90.pdf.

Anonymous:1873:ACbh

[4709] Anonymous. Answers to correspondents. Scientific American, 29(6):91–92, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n6/pdf/scientificamerican08091873-
91a.pdf.

Anonymous:1873:CRbb

[4710] Anonymous. Communications received. Scientific American, 29(6):92, August 9, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n6/pdf/scientificamerican08091873-92.pdf.

Anonymous:1873:Cbw

[4711] Anonymous. Correspondence. Scientific American, 29(7):100–101, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-100a.pdf.

Anonymous:1873:BNBu 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
104a.pdf.

Anonymous:1873:BPbh

[4713] Anonymous. Business and personal. Scientific American, 29(7):107, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-107.pdf.

Anonymous:1873:NQbd

[4714] Anonymous. Notes and queries. Scientific American, 29(7):107, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-107a.pdf.

Anonymous:1873:IIbf

[4715] Anonymous. Index of inventions. Scientific American, 29(7):108, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-108a.pdf.

Anonymous:1873:LSC

[4716] Anonymous. A large steam crane. Scientific American, 29(7):95, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-95.pdf.

Anonymous:1873:TAH

[4717] Anonymous. Triumph of American harvesters in Austria — the English exhibitors back out of the competition. Scientific American, 29(7):96, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
96.pdf.

Anonymous:1873:TSV

[4718] Anonymous. Towing sailing vessels. Scientific American, 29(7):96, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-96a.pdf.

Anonymous:1873:LR

[4719] Anonymous. Lightning rods. Scientific American, 29(7):96–97, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-96b.pdf.

Anonymous:1873:RMI

[4720] Anonymous. Recent meteoric investigations. Scientific American, 29(7):97, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
97.pdf.

Anonymous:1873:LPT

[4721] Anonymous. Lead poisoning and its treatment. Scientific American, 29(7):97, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
97a.pdf.

Anonymous:1873:PFM

[4722] Anonymous. Paper feeding machine. Scientific American, 29(7):98, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-98.pdf.

Anonymous:1873:PSCa

[4723] Anonymous. Progress of steam cultivation. Scientific American, 29(7):98, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
98a.pdf.

Anonymous:1873:IFR

[4724] Anonymous. Improved fishing rod. Scientific American, 29(7):98, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-98b.pdf.

Anonymous:1873:PSW

[4725] Anonymous. “The People’s” steam washer. Scientific American, 29(7):98, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
98c.pdf.

Anonymous:1873:APR

[4726] Anonymous. Air pump for raising sewage. Scientific American, 29(7):99, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
99.pdf.

Anonymous:1873:GELd

[4727] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(7):100, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-100.pdf.

Anonymous:1873:ICb

[4728] Anonymous. Improved churn. Scientific American, 29(7):102, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-102.pdf.

Anonymous:1873:OWC

[4729] Anonymous. Odorless water closet attachment. Scientific American, 29(7):102, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
102a.pdf.

Anonymous:1873:JB

[4730] Anonymous. Japanese birds. Scientific American, 29(7):103, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n7/pdf/scientificamerican08161873-103.pdf.

Anonymous:1873:SPIw

[4731] Anonymous. Scientific and practical information. Scientific American, 29(7):104, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
104.pdf.

Anonymous:1873:DCv

[4732] Anonymous. Decisions of the courts. Scientific American, 29 (7):105, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
105.pdf.

Anonymous:1873:RAFbg

[4733] Anonymous. Recent American and foreign patents. Scientific American, 29(7):105–106, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
105a.pdf.

Anonymous:1873:ACbi

[4734] Anonymous. Answers to correspondents. Scientific American, 29 (7):107–108, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
107b.pdf.

Anonymous:1873:CRbc

[4735] Anonymous. Communications received. Scientific American, 29 (7):108, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
108.pdf.

Anonymous:1873:CPg

[4736] Anonymous. Canadian patents. Scientific American, 29(7):108– 109, August 16, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n7/pdf/scientificamerican08161873-
108b.pdf.

Anonymous:1873:Cbx

[4737] Anonymous. Correspondence. Scientific American, 29(8):116–117, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-116a.pdf.

Anonymous:1873:BNBv

[4738] Anonymous. Books: New books and publications. Scientific American, 29(8):121, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
121a.pdf.

Anonymous:1873:BPbi

[4739] Anonymous. Business and personal. Scientific American, 29(8):123, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-123.pdf.

Anonymous:1873:NQbe

[4740] Anonymous. Notes and queries. Scientific American, 29(8):123, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-123a.pdf.

Anonymous:1873:IIbg

[4741] Anonymous. Index of inventions. Scientific American, 29(8):124, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-124a.pdf.

Anonymous:1873:MOV

[4742] Anonymous. Manufacture of oil of vitriol. Scientific American, 29(8):111, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
111.pdf.

Anonymous:1873:DOA

[4743] Anonymous. That “dastardly outrage” again. Scientific American, 29(8):112, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
112.pdf.

Anonymous:1873:SGb

[4744] Anonymous. Spontaneous generation. Scientific American, 29 (8):112, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
112a.pdf.

Anonymous:1873:RGI

[4745] Anonymous. Recent geological investigations. Scientific American, 29(8):112, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
112b.pdf.

Anonymous:1873:BLA

[4746] Anonymous. The base line of astronomy. Scientific American, 29 (8):112–113, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
112c.pdf.

Anonymous:1873:ESb

[4747] Anonymous. Efflux of steam. Scientific American, 29(8):113, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-113.pdf.

Anonymous:1873:EPa

[4748] Anonymous. Ezekiel Page. Scientific American, 29(8):113, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-113a.pdf.

Anonymous:1873:SPIx

[4749] Anonymous. Scientific and practical information. Scientific American, 29(8):113, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
113b.pdf.

Anonymous:1873:NS

[4750] Anonymous. Norway and Sweden. Scientific American, 29(8):115, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-115.pdf.

Anonymous:1873:HHP

[4751] Anonymous. Harris & Hewitt’s patent sash cord fastener. Scientific American, 29(8):115, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
115a.pdf.

Anonymous:1873:GELe

[4752] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(8):116, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-116.pdf.

Anonymous:1873:NDM

[4753] Anonymous. A new domestic motor. Scientific American, 29 (8):118, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
118.pdf.

Anonymous:1873:PCe

[4754] Anonymous. Perambulating cot. Scientific American, 29(8):118, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n8/pdf/scientificamerican08231873-118a.pdf.

Anonymous:1873:PRC

[4755] Anonymous. The patent revenue cigar box. Scientific American, 29(8):118, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
118b.pdf.

Anonymous:1873:CPEb

[4756] Anonymous. The Cooper portable engine. Scientific American, 29(8):119, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
119.pdf.

Anonymous:1873:FAF

[4757] Anonymous. The first automatic fan. Scientific American, 29 (8):119–120, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
119a.pdf.

Anonymous:1873:DCw

[4758] Anonymous. Decisions of the courts. Scientific American, 29 (8):121, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
121.pdf.

Anonymous:1873:RAFbh

[4759] Anonymous. Recent American and foreign patents. Scientific American, 29(8):121–122, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
121b.pdf.

Anonymous:1873:VPHe

[4760] Anonymous. Value of patents and how to obtain them. Scientific American, 29(8):122, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
122.pdf.

Anonymous:1873:ACbj

[4761] Anonymous. Answers to correspondents. Scientific American, 29 (8):123–124, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
123b.pdf.

Anonymous:1873:CRbd

[4762] Anonymous. Communications received. Scientific American, 29 (8):124, August 23, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n8/pdf/scientificamerican08231873-
124.pdf.

Anonymous:1873:Cby

[4763] Anonymous. Correspondence. Scientific American, 29(9):131–132, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-131b.pdf.

Anonymous:1873:BPbj

[4764] Anonymous. Business and personal. Scientific American, 29(9):138, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-138.pdf.

Anonymous:1873:NQbf

[4765] Anonymous. Notes and queries. Scientific American, 29(9):138, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-138a.pdf.

Anonymous:1873:IIbh

[4766] Anonymous. Index of inventions. Scientific American, 29(9):140, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-140a.pdf.

Anonymous:1873:Ph

[4767] Anonymous. The pulsometer. Scientific American, 29(9):127–128, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-127.pdf.

Anonymous:1873:GTO

[4768] Anonymous. The “Granges” and their object. Scientific American, 29(9):128, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
128.pdf.

Anonymous:1873:FOW

[4769] Anonymous. The flowing oil wells of Pennsylvania — great decline in the price of oil. Scientific American, 29(9):128, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-128a.pdf.

Anonymous:1873:SMb

[4770] Anonymous. The study of mathematics. Scientific American, 29 (9):128–129, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
128b.pdf.

Anonymous:1873:EAP

[4771] Anonymous. Engines of Admiral Porter’s torpedo boat. Scientific American, 29(9):129, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
129.pdf.

Anonymous:1873:SDH

[4772] Anonymous. The Sun’s distance and how it is measured. Scientific American, 29(9):129, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
129a.pdf.

Anonymous:1873:SPIy

[4773] Anonymous. Scientific and practical information. Scientific American, 29(9):129, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
129b.pdf.

Anonymous:1873:CSCb

[4774] Anonymous. Concrete siphons on the Canal Quintino Sella, Lomellina, Italy. Scientific American, 29(9):130, August 30, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n9/pdf/scientificamerican08301873-130.pdf.

Anonymous:1873:RGR

[4775] Anonymous. Reversing gear for rolling mills. Scientific American, 29(9):130, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
130a.pdf.

Anonymous:1873:NA

[4776] Anonymous. Nature and art. Scientific American, 29(9):131, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-131.pdf.

Anonymous:1873:DF

[4777] Anonymous. The devil fish. Scientific American, 29(9):131, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-131a.pdf.

Anonymous:1873:GELf

[4778] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(9):133, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n9/pdf/scientificamerican08301873-133.pdf.

Anonymous:1873:BCMb

[4779] Anonymous. Brick compressing machine. Scientific American, 29(9):134, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
134.pdf.

Anonymous:1873:DCCa

[4780] Anonymous. The diamond cotton chopper and cultivator. Scientific American, 29(9):134, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
134a.pdf.

Anonymous:1873:FFAa

[4781] Anonymous. French four axled articulated locomotive. Scientific American, 29(9):135–137, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
135.pdf.

Anonymous:1873:RAFbi

[4782] Anonymous. Recent American and foreign patents. Scientific American, 29(9):137, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
137.pdf.

Anonymous:1873:ACbk

[4783] Anonymous. Answers to correspondents. Scientific American, 29 (9):138–139, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
138b.pdf.

Anonymous:1873:CRbe

[4784] Anonymous. Communications received. Scientific American, 29 (9):140, August 30, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n9/pdf/scientificamerican08301873-
140.pdf.

Anonymous:1873:Cbz

[4785] Anonymous. Correspondence. Scientific American, 29(10):149, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:BNBw

[4786] Anonymous. Books: New books and publications. Scientific American, 29(10):154, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:BPbk

[4787] Anonymous. Business and personal. Scientific American, 29(10):155, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:NQbg

[4788] Anonymous. Notes and queries. Scientific American, 29(10):155, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IIbi

[4789] Anonymous. Index of inventions. Scientific American, 29(10):156, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IFW

[4790] Anonymous. Improved feed water heater and purifier. Scientific American, 29(10):143–144, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:PMR

[4791] Anonymous. The poor Man’s representative. Scientific American, 29 (10):144, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:BCA

[4792] Anonymous. The bisulphide of carbon auxiliary. Scientific American, 29(10):144, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:APVb

[4793] Anonymous. Award of prizes at the Vienna Exposition. Scientific American, 29(10):144, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:PCVb

[4794] Anonymous. The patent Congress at Vienna. Scientific American, 29(10):144, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:LRF

[4795] Anonymous. Lightning rods — a few practical hints. Scientific American, 29(10):144–145, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:CWI

[4796] Anonymous. Caoutchouc — where it comes from, and how it is prepared. Scientific American, 29(10):145, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:ASMa

[4797] Anonymous. American sewing machine awards at Vienna. Scientific American, 29(10):145, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:ABV

[4798] Anonymous. The architect and the builder of the Vienna Exposition Palace. Scientific American, 29(10):147, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:MRG

[4799] Anonymous. The machine room of the Great Exposition. Scientific American, 29(10):147, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:GELg

[4800] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(10):148, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:SBG

[4801] Anonymous. Steam boilers at the Great Exposition — letter from the Engineer of the United States Department. Scientific American, 29 (10):148–149, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IHS

[4802] Anonymous. The ironclad hulling stone. Scientific American, 29(10): 150, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IWWb

[4803] Anonymous. Improved window weather strip. Scientific American, 29(10):150, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IFA

[4804] Anonymous. Improved friction attachment. Scientific American, 29(10): 150, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:IWd

[4805] Anonymous. Improved wrench. Scientific American, 29(10):150, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:KCM

[4806] Anonymous. Kansas City, Mo., and its coming Exposition. Scientific American, 29(10):151, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:CLBb

[4807] Anonymous. Counterbalanced lifting bridge. Scientific American, 29 (10):151–152, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:ITMb

[4808] Anonymous. Improved testing machine. Scientific American, 29(10): 152, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:AAAa

[4809] Anonymous. American Association for the Advancement of Science. Scientific American, 29(10):153, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:FAA

[4810] Anonymous. Foreign appreciation of American inventions. Scientific American, 29(10):153–154, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:DCx

[4811] Anonymous. Decisions of the courts. Scientific American, 29(10):154, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:GVE

[4812] Anonymous. A grand victory over every competitor in the world. Scientific American, 29(10):154, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:RAFbj

[4813] Anonymous. Recent American and foreign patents. Scientific American, 29(10):154, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:ACbl

[4814] Anonymous. Answers to correspondents. Scientific American, 29(10): 155–156, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:CRbf

[4815] Anonymous. Communications received. Scientific American, 29(10): 156, September 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1873:Cca

[4816] Anonymous. Correspondence. Scientific American, 29(11):164, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n11/pdf/scientificamerican09131873-164.pdf.

Anonymous:1873:BPbl

[4817] Anonymous. Business and personal. Scientific American, 29 (11):171, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
171.pdf.

Anonymous:1873:NQbh

[4818] Anonymous. Notes and queries. Scientific American, 29(11): 171, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
171a.pdf.

Anonymous:1873:IIbj

[4819] Anonymous. Index of inventions. Scientific American, 29(11): 172, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
172a.pdf.

Anonymous:1873:BAa

[4820] Anonymous. The balloon and the aeronauts. Scientific American, 29(11):159, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
159.pdf.

Anonymous:1873:EVP

[4821] Anonymous. End of the Vienna patent Congress. Scientific American, 29(11):160, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
160.pdf.

Anonymous:1873:HNP

[4822] Anonymous. Hoe’s new printing press. Scientific American, 29 (11):160, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
160a.pdf.

Anonymous:1873:BVE

[4823] Anonymous. The balloon voyage to Europe. Scientific American, 29(11):160, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
160b.pdf.

Anonymous:1873:MM

[4824] Anonymous.  Modern miracles.  Scientific American, 29(11):

160, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
160c.pdf.

Anonymous:1873:TSS

[4825] Anonymous. Theories of the Sun spots. Scientific American, 29 (11):160–161, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
160d.pdf.

Anonymous:1873:ANDa

[4826] Anonymous. Accident to the new docks. Scientific American, 29 (11):161, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
161.pdf.

Anonymous:1873:HF

[4827] Anonymous. Hay fever. Scientific American, 29(11):161, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n11/pdf/scientificamerican09131873-161a.pdf.

Anonymous:1873:SPIz

[4828] Anonymous. Scientific and practical information. Scientific American, 29(11):161, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
161b.pdf.

Anonymous:1873:LFD

[4829] Anonymous. The large floating Derrick. Scientific American, 29 (11):162, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
162.pdf.

Morton:1873:MLMa

[4830] Henry Morton. The magic lantern as a means of demonstration. Scientific American, 29(11):163–164, September 13, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n11/pdf/scientificamerican09131873-163.pdf.

Anonymous:1873:GELh

[4831] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29 (11):165, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
165.pdf.

Anonymous:1873:ITWa

[4832] Anonymous. Improved turbine water wheel. Scientific American, 29(11):166, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
166.pdf.

Anonymous:1873:HPN

[4833] Anonymous. The hydraulic propeller — a novel mode of marine propulsion. Scientific American, 29(11):166–167, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n11/pdf/scientificamerican09131873-166a.pdf.

Anonymous:1873:PPV

[4834] Anonymous. The Persian pavilion at Vienna. Scientific American, 29(11):167, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
167.pdf.

Anonymous:1873:AAAb

[4835] Anonymous. The American Association for the Advancement of Science. Scientific American, 29(11):168–169, September 13, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n11/pdf/scientificamerican09131873-168.pdf.

Anonymous:1873:GMC

[4836] Anonymous. A grand medal for Cleveland. Scientific American, 29 (11):169–170, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
169.pdf.

Anonymous:1873:AMV

[4837] Anonymous. American manufactures at Vienna. Scientific American, 29(11):170, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
170.pdf.

Anonymous:1873:RAFbk

[4838] Anonymous. Recent American and foreign patents. Scientific American, 29(11):170, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
170a.pdf.

Anonymous:1873:ACbm

[4839] Anonymous. Answers to correspondents. Scientific American, 29 (11):171–172, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
171b.pdf.

Anonymous:1873:CRbg

[4840] Anonymous. Communications received. Scientific American, 29 (11):172, September 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n11/pdf/scientificamerican09131873-
172.pdf.

Anonymous:1873:Ccb

[4841] Anonymous. Correspondence. Scientific American, 29(12):180, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n12/pdf/scientificamerican09201873-180a.pdf.

Anonymous:1873:BPbm

[4842] Anonymous. Business and personal. Scientific American, 29 (12):187, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
187.pdf.

Anonymous:1873:IIbk

[4843] Anonymous. Index of inventions. Scientific American, 29(12): 188, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
188a.pdf.

Anonymous:1873:VGB

[4844] Anonymous. The voyage of the graphic balloon. Scientific American, 29(12):176, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
176.pdf.

Anonymous:1873:HG

[4845] Anonymous. Hell Gate. Scientific American, 29(12):176, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n12/pdf/scientificamerican09201873-176a.pdf.

Anonymous:1873:VPc

[4846] Anonymous. The Vienna prizes. Scientific American, 29(12): 176, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
176b.pdf.

Anonymous:1873:PIM

[4847] Anonymous. The patent ice machine. Scientific American, 29 (12):177, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
177.pdf.

Anonymous:1873:PTFc

[4848] Anonymous. Pens and their failings. Scientific American, 29 (12):177, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
177a.pdf.

Anonymous:1873:DOS

[4849] Anonymous. Deodorizing the offal from slaughter houses. Scientific American, 29(12):177, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
177b.pdf.

Anonymous:1873:SPIba

[4850] Anonymous. Scientific and practical information. Scientific American, 29(12):177, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
177c.pdf.

Anonymous:1873:HCM

[4851] Anonymous. How to catch moles. Scientific American, 29(12): 178, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
178.pdf.

Anonymous:1873:DSD

[4852] Anonymous. Deep sea dredging apparatus. Scientific American, 29(12):178, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
178a.pdf.

Anonymous:1873:NSC

[4853] Anonymous. A new system of constructing blast furnaces. Scientific American, 29(12):179, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
179.pdf.

Anonymous:1873:HCSa

[4854] Anonymous. How to conduct slaughtering packing and rendering without offense. Scientific American, 29(12):179–180, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n12/pdf/scientificamerican09201873-179a.pdf.

Edwards:1873:HMA

[4855] Henry Edwards. On the honey-making ant of Texas and new Mexico. Scientific American, 29(12):180, September 20, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n12/pdf/scientificamerican09201873-180.pdf.

Anonymous:1873:GELi

[4856] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29 (12):181, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
181.pdf.

Anonymous:1873:DTMb

[4857] Anonymous. Device for transferring motion. Scientific American, 29(12):182, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
182.pdf.

Anonymous:1873:ABD

[4858] Anonymous. Automatic boat detaching apparatus. Scientific American, 29(12):182, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
182a.pdf.

Anonymous:1873:ISE

[4859] Anonymous. The Indiana State Exposition. Scientific American, 29 (12):182–183, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
182b.pdf.

Anonymous:1873:TTP

[4860] Anonymous. The Turkish treasure pavilion at Vienna. Scientific American, 29(12):183, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
183.pdf.

Morton:1873:MLMb

[4861] Henry Morton. The magic lantern as a means of demonstration. Scientific American, 29(12):184, September 20, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n12/pdf/scientificamerican09201873-184.pdf.

Anonymous:1873:AAAc

[4862] Anonymous. The American Association for the Advancement of Science. Scientific American, 29(12):185, September 20, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n12/pdf/scientificamerican09201873-185.pdf.

Anonymous:1873:VPS

[4863] Anonymous. Vienna premiums and sewing machines. Scientific American, 29(12):185–186, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
185a.pdf.

Anonymous:1873:DCy

[4864] Anonymous. Decisions of the courts. Scientific American, 29 (12):186, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
186.pdf.

Anonymous:1873:RAFbl

[4865] Anonymous. Recent American and foreign patents. Scientific American, 29(12):186, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
186a.pdf.

Anonymous:1873:ACbn

[4866] Anonymous. Answers to correspondents. Scientific American, 29 (12):187–188, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
187a.pdf.

Anonymous:1873:CRbh

[4867] Anonymous. Communications received. Scientific American, 29 (12):188, September 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n12/pdf/scientificamerican09201873-
188.pdf.

Anonymous:1873:Ccd

[4868] Anonymous. Correspondence. Scientific American, 29(13):196, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n13/pdf/scientificamerican09271873-196.pdf.

Anonymous:1873:BNBx

[4869] Anonymous. Books: New books and publications. Scientific American, 29(13):201, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
201a.pdf.

Anonymous:1873:BPbn

[4870] Anonymous. Business and personal. Scientific American, 29 (13):203, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
203.pdf.

Anonymous:1873:NQbi

[4871] Anonymous. Notes and queries. Scientific American, 29(13): 203, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
203a.pdf.

Anonymous:1873:IIbl

[4872] Anonymous. Index of inventions. Scientific American, 29(13): 204, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
204.pdf.

Anonymous:1873:HAW

[4873] Anonymous. Havens’ automatic wire rope railway. Scientific American, 29(13):191–192, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
191.pdf.

Anonymous:1873:VPCa

[4874] Anonymous. The Vienna patent Congress once more. Scientific American, 29(13):192, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
192.pdf.

Anonymous:1873:MIG

[4875] Anonymous. Manufacture of illuminating gas from crude petroleum. Scientific American, 29(13):192, September 27, 1873. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
192a.pdf.

Anonymous:1873:DPN

[4876] Anonymous. The Department of Parks, New York City. Scientific American, 29(13):192, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
192b.pdf.

Anonymous:1873:CSb

[4877] Anonymous. Chemical sugar. Scientific American, 29(13):192– 193, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
192c.pdf.

Anonymous:1873:TMA

[4878] Anonymous. Tidings from the missing Arctic explorers. Scientific American, 29(13):193, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
193.pdf.

Anonymous:1873:FBE

[4879] Anonymous. Failure of the balloon to Europe. Scientific American, 29(13):193, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
193a.pdf.

Anonymous:1873:GF

[4880] Anonymous. Gas as fuel. Scientific American, 29(13):193, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n13/pdf/scientificamerican09271873-193b.pdf.

Anonymous:1873:SPIbb

[4881] Anonymous. Scientific and practical information. Scientific American, 29(13):193, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
193c.pdf.

Anonymous:1873:NSM

[4882] Anonymous. New steam motor. Scientific American, 29(13): 194, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
194.pdf.

Anonymous:1873:RR

[4883] Anonymous. The rigi railway. Scientific American, 29(13):194, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n13/pdf/scientificamerican09271873-194a.pdf.

Anonymous:1873:PMIa

[4884] Anonymous. The Planet Mars — is it inhabited? Scientific American, 29(13):195, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
195.pdf.

Anonymous:1873:E

[4885] Anonymous. Earwigs. Scientific American, 29(13):195–196, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n13/pdf/scientificamerican09271873-195a.pdf.

Anonymous:1873:GELj

[4886] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29 (13):197, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
197.pdf.

Anonymous:1873:IRH

[4887] Anonymous. Improved railroad hand brake. Scientific American, 29(13):198, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
198.pdf.

Anonymous:1873:CSF

[4888] Anonymous. Combined spade and fork. Scientific American, 29 (13):198, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
198a.pdf.

Anonymous:1873:CWP

[4889] Anonymous. Combined water and pressure gage. Scientific American, 29(13):198, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
198b.pdf.

Anonymous:1873:ICD

[4890] Anonymous. Improved clutch drill. Scientific American, 29(13): 199, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
199.pdf.

Anonymous:1873:HNH

[4891] Anonymous. Down Hall, near Harlow, Essex, England. Scientific American, 29(13):199, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
199a.pdf.

Morton:1873:MLMc

[4892] Henry Morton. The magic lantern as a means of demonstration. Scientific American, 29(13):200–201, September 27, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n13/pdf/scientificamerican09271873-200.pdf.

Anonymous:1873:VPAa

[4893] Anonymous. Vienna premiums again and the American sewing machines. Scientific American, 29(13):201, September 27, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n13/pdf/scientificamerican09271873-201.pdf.

Anonymous:1873:RAFbm

[4894] Anonymous. Recent American and foreign patents. Scientific American, 29(13):201–202, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
201b.pdf.

Anonymous:1873:ACbo

[4895] Anonymous. Answer to correspondents. Scientific American, 29 (13):203, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
203b.pdf.

Anonymous:1873:CRbi

[4896] Anonymous. Communications received. Scientific American, 29 (13):203, September 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n13/pdf/scientificamerican09271873-
203c.pdf.

Anonymous:1873:Cce

[4897] Anonymous. Correspondence. Scientific American, 29(14):212, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-212.pdf.

Anonymous:1873:BNBy

[4898] Anonymous. Books: New books and publications. Scientific American, 29(14):217, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
217.pdf.

Anonymous:1873:BPbo

[4899] Anonymous. Business and personal. Scientific American, 29(14):219, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-219.pdf.

Anonymous:1873:NQbj

[4900] Anonymous. Notes and queries. Scientific American, 29(14):219, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-219a.pdf.

Anonymous:1873:IIbm

[4901] Anonymous. Index of inventions. Scientific American, 29(14):220, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-220a.pdf.

Anonymous:1873:DCNa

[4902] Anonymous. Diamond cutting — a new industry in New York. Scientific American, 29(14):207, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
207.pdf.

Anonymous:1873:SHc

[4903] Anonymous. Specific heat. Scientific American, 29(14):208, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-208.pdf.

Anonymous:1873:Ccf

[4904] Anonymous. Camphor. Scientific American, 29(14):208, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-208a.pdf.

Anonymous:1873:CPFa

[4905] Anonymous. Crude petroleum for fuel and for illuminating gas. Scientific American, 29(14):208–209, October 4, 1873. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
208b.pdf.

Anonymous:1873:RRS

[4906] Anonymous. Rescue of the remaining survivors of the Polaris. Scientific American, 29(14):209, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
209.pdf.

Anonymous:1873:FAIa

[4907] Anonymous. The fair of the American institute. Scientific American, 29(14):209, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
209a.pdf.

Anonymous:1873:PMIb

[4908] Anonymous. The Planet Mars — is it inhabited? Scientific American, 29(14):210, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
210.pdf.

Anonymous:1873:STP

[4909] Anonymous. A simple tree protector. Scientific American, 29 (14):210–211, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
210a.pdf.

Anonymous:1873:STD

[4910] Anonymous. Self-tightening drill chuck. Scientific American, 29 (14):211, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
211.pdf.

Anonymous:1873:CTE

[4911] Anonymous. Calfskin tanning in Europe. Scientific American, 29(14):211, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
211a.pdf.

Anonymous:1873:GELk

[4912] Anonymous. The Great Exposition — letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(14):213, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-213.pdf.

Anonymous:1873:ARG

[4913] Anonymous. Automatic railroad gate and signal. Scientific American, 29(14):214, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
214.pdf.

Anonymous:1873:DDC

[4914] Anonymous. The Danbury drill chuck. Scientific American, 29(14):214, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n14/pdf/scientificamerican10041873-214a.pdf.

Anonymous:1873:IPAa

[4915] Anonymous. Improved plow attachment. Scientific American, 29(14):214, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
214b.pdf.

Anonymous:1873:RAFbn

[4916] Anonymous. Recent American and foreign patents. Scientific American, 29(14):217–218, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
217a.pdf.

Anonymous:1873:ACbp

[4917] Anonymous. Answer to correspondents. Scientific American, 29 (14):219–220, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
219b.pdf.

Anonymous:1873:CRbj

[4918] Anonymous. Communications received. Scientific American, 29 (14):220, October 4, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n14/pdf/scientificamerican10041873-
220.pdf.

Anonymous:1873:Ccg

[4919] Anonymous. Correspondence. Scientific American, 29(15):228, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-228.pdf.

Anonymous:1873:BPbp

[4920] Anonymous. Business and personal. Scientific American, 29 (15):234, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
234.pdf.

Anonymous:1873:NQbk

[4921] Anonymous. Notes and queries. Scientific American, 29(15):234, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-234a.pdf.

Anonymous:1873:IIbn

[4922] Anonymous. Index of inventions. Scientific American, 29(15): 235–236, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
235a.pdf.

Anonymous:1873:IVC

[4923] Anonymous. Improved variable cut-off. Scientific American, 29 (15):223, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
223.pdf.

Anonymous:1873:FC

[4924] Anonymous. The financial crisis. Scientific American, 29(15):224, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-224.pdf.

Anonymous:1873:DOB

[4925] Anonymous. Discovery of an old boot and shoe sewing machine patent. Scientific American, 29(15):224, October 11, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n15/pdf/scientificamerican10111873-224a.pdf.

Anonymous:1873:HPI

[4926] Anonymous. How pianos are injured. Scientific American, 29 (15):224, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
224b.pdf.

Anonymous:1873:SMBa

[4927] Anonymous. Sand and mud baths. Scientific American, 29(15):224, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-224c.pdf.

Anonymous:1873:MB

[4928] Anonymous. The mechanics of the brain. Scientific American, 29 (15):224–225, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
224d.pdf.

Anonymous:1873:SPIbc

[4929] Anonymous. Scientific and practical information. Scientific American, 29(15):225, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
225.pdf.

Anonymous:1873:FAIb

[4930] Anonymous. The fair of the American institute. Scientific American, 29(15):225, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
225a.pdf.

Anonymous:1873:IBMb

[4931] Anonymous. Improved boring machine. Scientific American, 29 (15):226, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
226.pdf.

Anonymous:1873:CLP

[4932] Anonymous. Checker lozenge package. Scientific American, 29 (15):226, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
226a.pdf.

Anonymous:1873:SCS

[4933] Anonymous. Sprague’s combined sash lock, sash holder, and weather strip. Scientific American, 29(15):226, October 11, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n15/pdf/scientificamerican10111873-226b.pdf.

Anonymous:1873:EWP

[4934] Anonymous. The evaporation of water by plants. Scientific American, 29(15):227–228, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
227.pdf.

Anonymous:1873:Aa

[4935] Anonymous. Up the Amazons. Scientific American, 29(15):228, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-228a.pdf.

Anonymous:1873:LUSa

[4936] Anonymous. Letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(15):229, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-229.pdf.

Anonymous:1873:CHRb

[4937] Anonymous. Combined hay rake and tedder. Scientific American, 29(15):230, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
230.pdf.

Anonymous:1873:ICR

[4938] Anonymous. Improved combined roller and irrigator. Scientific American, 29(15):230, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
230a.pdf.

Anonymous:1873:IMP

[4939] Anonymous. Improved match plane. Scientific American, 29 (15):231, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
231.pdf.

Anonymous:1873:GSG

[4940] Anonymous. The globe steam gage. Scientific American, 29(15): 231–232, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
231a.pdf.

Anonymous:1873:ISPb

[4941] Anonymous. Improved Shoemaker’s pinchers. Scientific American, 29(15):232, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
232.pdf.

Anonymous:1873:VSb

[4942] Anonymous. Ventilation of sewers. Scientific American, 29(15):232, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-232a.pdf.

Anonymous:1873:RAFbo

[4943] Anonymous. Recent American and foreign patents. Scientific American, 29(15):232–233, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
232b.pdf.

Anonymous:1873:ACbq

[4944] Anonymous. Answers to correspondents. Scientific American, 29 (15):234–235, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
234b.pdf.

Anonymous:1873:CRbk

[4945] Anonymous. Communications received. Scientific American, 29 (15):235, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n15/pdf/scientificamerican10111873-
235.pdf.

Anonymous:1873:AIW

[4946] Anonymous. American iron works. Scientific American, 29(15):236, October 11, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n15/pdf/scientificamerican10111873-236.pdf.

Anonymous:1873:BPbq

[4947] Anonymous. Business and personal. Scientific American, 29 (16):250, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
250.pdf.

Anonymous:1873:NQbl

[4948] Anonymous. Notes and queries. Scientific American, 29(16):250, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-250a.pdf.

Anonymous:1873:IIbo

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-252.pdf.

Anonymous:1873:HSR

[4950] Anonymous. Hill side railway. Scientific American, 29(16):239–240, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-239.pdf.

Anonymous:1873:PPM

[4951] Anonymous. Patent planing machine litigation. Scientific American, 29(16):240, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
240.pdf.

Anonymous:1873:Pj

[4952] Anonymous. The panic over. Scientific American, 29(16):240, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-240a.pdf.

Anonymous:1873:PW

[4953] Anonymous. Puddle walls. Scientific American, 29(16):240, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-240b.pdf.

Anonymous:1873:FLb

[4954] Anonymous. The fireless locomotive. Scientific American, 29 (16):240, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
240c.pdf.

Anonymous:1873:AI

[4955] Anonymous. The art of inventing. Scientific American, 29(16):241, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-241.pdf.

Anonymous:1873:MABa

[4956] Anonymous. Manufacture of artificial butter in New York. Scientific American, 29(16):241, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
241a.pdf.

Anonymous:1873:SPIbd

[4957] Anonymous. Scientific and practical information. Scientific American, 29(16):241, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
241b.pdf.

Anonymous:1873:FAIc

[4958] Anonymous. The fair of the American institute. Scientific American, 29(16):241, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
241c.pdf.

Anonymous:1873:ICBa

[4959] Anonymous. Improved car brake. Scientific American, 29(16):242, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-242.pdf.

Anonymous:1873:IAT

[4960] Anonymous. Improved animal trap. Scientific American, 29 (16):242, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
242a.pdf.

Anonymous:1873:FMM

[4961] Anonymous. The fossil man of mentone. Scientific American, 29(16):243, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
243.pdf.

Anonymous:1873:NDSc

[4962] Anonymous. A new deep sea sounding instrument. Scientific American, 29(16):243, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
243a.pdf.

Anonymous:1873:FFI

[4963] Anonymous. The fecundation of flowers by insects. Scientific American, 29(16):243–244, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
243b.pdf.

Anonymous:1873:Ab

[4964] Anonymous. Up the Amazons. Scientific American, 29(16):244– 245, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
244.pdf.

Plante:1873:NED

[4965] Gaston Plante and Alf. Maudet-Breguet. Note on an electrodynamic experiment. Scientific American, 29(16):245, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-245.pdf.

Anonymous:1873:MABb

[4966] Anonymous. Manufacture of artificial butter in New York. Scientific American, 29(16):246–248, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
246.pdf.

Anonymous:1873:LUSb

[4967] Anonymous. Letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(16):248–249, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n16/pdf/scientificamerican10181873-248.pdf.

Anonymous:1873:RAFbp

[4968] Anonymous. Recent American and foreign patents. Scientific American, 29(16):249, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
249.pdf.

Anonymous:1873:ACbr

[4969] Anonymous. Answers to correspondents. Scientific American, 29 (16):250–251, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
250b.pdf.

Anonymous:1873:CRbl

[4970] Anonymous. Communications received. Scientific American, 29 (16):251, October 18, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n16/pdf/scientificamerican10181873-
251.pdf.

Anonymous:1873:Cch

[4971] Anonymous. Correspondence. Scientific American, 29(17):260–261, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-260a.pdf.

Anonymous:1873:BNBz

[4972] Anonymous. Books: New books and publications. Scientific American, 29(17):265, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
265.pdf.

Anonymous:1873:BPbr

[4973] Anonymous. Business and personal. Scientific American, 29 (17):266, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
266.pdf.

Anonymous:1873:NQbm

[4974] Anonymous. Notes and queries. Scientific American, 29(17):266, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-266a.pdf.

Anonymous:1873:IIbp

[4975] Anonymous. Index of inventions. Scientific American, 29(17):267, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-267.pdf.

Anonymous:1873:AEC

[4976] Anonymous. Apparatus for emptying cesspools. Scientific American, 29(17):255–256, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
255.pdf.

Anonymous:1873:TRE

[4977] Anonymous. Technical research and education. Scientific American, 29(17):256, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
256.pdf.

Anonymous:1873:VPd

[4978] Anonymous. The Vienna prizes. Scientific American, 29(17):256, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-256a.pdf.

Anonymous:1873:FNB

[4979] Anonymous. Failure no. 2 of the balloon to Europe. Scientific American, 29(17):256, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
256b.pdf.

Anonymous:1873:CCP

[4980] Anonymous. Caoutchouc from common plants. Scientific American, 29(17):256–257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
256c.pdf.

Anonymous:1873:RGT

[4981] Anonymous. Recent gunnery trials. Scientific American, 29(17): 257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
257.pdf.

Anonymous:1873:CPFb

[4982] Anonymous. Crude petroleum for fuel. Scientific American, 29 (17):257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
257a.pdf.

Anonymous:1873:SPg

[4983] Anonymous. Street pavements. Scientific American, 29(17):257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-257b.pdf.

Anonymous:1873:CBCa

[4984] Anonymous. Curiosities of butter and churning. Scientific American, 29(17):257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
257c.pdf.

Anonymous:1873:SPIbe

[4985] Anonymous. Scientific and practical information. Scientific American, 29(17):257, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
257d.pdf.

Anonymous:1873:IFLa

[4986] Anonymous. Inextinguishable fire for life buoys. Scientific American, 29(17):258, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
258.pdf.

Anonymous:1873:ESL

[4987] Anonymous. The English steam launch firefly. Scientific American, 29(17):258–259, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
258a.pdf.

Anonymous:1873:PIb

[4988] Anonymous. Peripolar induction. Scientific American, 29(17):259, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n17/pdf/scientificamerican10251873-259.pdf.

Anonymous:1873:DGS

[4989] Anonymous. Demming’s grown sash fastener. Scientific American, 29(17):259, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
259a.pdf.

Anonymous:1873:IPF

[4990] Anonymous. Improved press for fruit, lard etc. Scientific American, 29(17):259–260, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
259b.pdf.

Anonymous:1873:ASMb

[4991] Anonymous. That ancient sewing machine. Scientific American, 29(17):260, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
260.pdf.

Thurston:1873:LUSa

[4992] R. H. Thurston. Letter from United States Commissioner. Scientific American, 29(17):261, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
261.pdf.

Anonymous:1873:PPS

[4993] Anonymous. Palmer’s power Spring hammer. Scientific American, 29(17):262, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
262.pdf.

Anonymous:1873:IRSc

[4994] Anonymous. Improved railroad switch. Scientific American, 29 (17):262, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
262a.pdf.

Anonymous:1873:FTD

[4995] Anonymous. A forty thousand dollar cow. Scientific American, 29(17):263, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
263.pdf.

Anonymous:1873:HCC

[4996] Anonymous. Helicoidal concavo–convex propeller. Scientific American, 29(17):263, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
263a.pdf.

Anonymous:1873:MBA

[4997] Anonymous. Meeting of the British Association. Scientific American, 29(17):264, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
264.pdf.

Anonymous:1873:RCA

[4998] Anonymous. Rare chance to advertisers. Scientific American, 29 (17):264, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
264a.pdf.

Anonymous:1873:VDI

[4999] Anonymous. A valuble drafting instrument. Scientific American, 29(17):264, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
264b.pdf.

Anonymous:1873:RAFbq

[5000] Anonymous. Recent American and foreign patents. Scientific American, 29(17):265, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
265a.pdf.

Anonymous:1873:ACbs

[5001] Anonymous. Answers to correspondents. Scientific American, 29 (17):266, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
266b.pdf.

Anonymous:1873:CRbm

[5002] Anonymous. Communications received. Scientific American, 29 (17):266, October 25, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n17/pdf/scientificamerican10251873-
266c.pdf.

Anonymous:1873:Cci

[5003] Anonymous. Correspondence. Scientific American, 29(18):275– 277, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
275a.pdf.

Anonymous:1873:BPbs

[5004] Anonymous. Business and personal. Scientific American, 29 (18):282, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
282.pdf.

Anonymous:1873:JG

[5005] Anonymous. Joseph Gillott. Scientific American, 29(18):271– 272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
271.pdf.

Anonymous:1873:SSAb

[5006] Anonymous. Secret societies among college students. Scientific American, 29(18):272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272.pdf.

Anonymous:1873:PPb

[5007] Anonymous. Progress of patents. Scientific American, 29(18): 272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272a.pdf.

Anonymous:1873:WSP

[5008] Anonymous. Wood and straw paper making in France. Scientific American, 29(18):272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272b.pdf.

Anonymous:1873:TID

[5009] Anonymous. Two interesting discoveries. Scientific American, 29 (18):272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272c.pdf.

Anonymous:1873:NOC

[5010] Anonymous. New order by the Commissioner of Patents. Scientific American, 29(18):272, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272d.pdf.

Anonymous:1873:PPT

[5011] Anonymous. Petroleum and pine tar gases. Scientific American, 29 (18):272–273, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
272e.pdf.

Anonymous:1873:SPIbf

[5012] Anonymous. Scientific and practical information. Scientific American, 29(18):273, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
273.pdf.

Anonymous:1873:BAb

[5013] Anonymous. The British Association. Scientific American, 29 (18):273, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
273a.pdf.

Anonymous:1873:FWC

[5014] Anonymous. Feeding and watering cattle in railroad cars. Scientific American, 29(18):274, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
274.pdf.

Anonymous:1873:CSd

[5015] Anonymous. Circular ships. Scientific American, 29(18):275, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n18/pdf/scientificamerican11011873-275.pdf.

Thurston:1873:LUSb

[5016] R. H. Thurston. Letter from United States Commissioner. Scientific American, 29(18):277, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
277.pdf.

Anonymous:1873:ACCc

[5017] Anonymous. Automatic car coupling. Scientific American, 29 (18):278, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
278.pdf.

Anonymous:1873:CIM

[5018] Anonymous. Cummings’ improved mill pick. Scientific American, 29(18):278, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
278a.pdf.

Anonymous:1873:HCSb

[5019] Anonymous. Hamilton’s continuous self feeding permeator. Scientific American, 29(18):278, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
278b.pdf.

Anonymous:1873:IL

[5020] Anonymous. Improved lubricator. Scientific American, 29(18): 279, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
279.pdf.

Anonymous:1873:JDV

[5021] Anonymous. The Japanese department at Vienna. Scientific American, 29(18):279, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
279a.pdf.

Anonymous:1873:PBR

[5022] Anonymous. Potato blight and rot. Scientific American, 29(18): 279–280, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
279b.pdf.

Anonymous:1873:NEAa

[5023] Anonymous. The new exploration of the Amazon River, by Professor Orton, — up the Amazons. Scientific American, 29(18): 280–281, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
280.pdf.

Anonymous:1873:RAFbr

[5024] Anonymous. Recent American and foreign patents. Scientific American, 29(18):281, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
281.pdf.

Anonymous:1873:ACbt

[5025] Anonymous. Answers to correspondents. Scientific American, 29 (18):282–283, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
282a.pdf.

Anonymous:1873:CRbn

[5026] Anonymous. Communications received. Scientific American, 29 (18):283, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
283.pdf.

Anonymous:1873:IIbq

[5027] Anonymous. Index of inventions. Scientific American, 29(18): 283–284, November 1, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n18/pdf/scientificamerican11011873-
283a.pdf.

Anonymous:1873:Ccj

[5028] Anonymous. Correspondence. Scientific American, 29(19):292, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n19/pdf/scientificamerican11081873-292.pdf.

Anonymous:1873:BPbt

[5029] Anonymous. Business and personal. Scientific American, 29 (19):299, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
299.pdf.

Anonymous:1873:NQbn

[5030] Anonymous. Notes and queries. Scientific American, 29(19): 299, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
299a.pdf.

Anonymous:1873:IIbr

[5031] Anonymous. Index of inventions. Scientific American, 29(19): 300, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
300a.pdf.

Anonymous:1873:ISCc

[5032] Anonymous. Improved standard compress. Scientific American, 29 (19):287–288, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
287.pdf.

Anonymous:1873:ECb

[5033] Anonymous. The electrical condenser. Scientific American, 29 (19):288, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
288.pdf.

Anonymous:1873:PSE

[5034] Anonymous. Progress of scientific education. Scientific American, 29(19):288, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
288a.pdf.

Anonymous:1873:SIT

[5035] Anonymous. The state of the iron trade. Scientific American, 29(19):288, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
288b.pdf.

Anonymous:1873:EPb

[5036] Anonymous. Elastic profanity. Scientific American, 29(19):288– 289, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
288c.pdf.

Anonymous:1873:LDN

[5037] Anonymous. The late Dr Nelaton. Scientific American, 29(19): 289, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
289.pdf.

Anonymous:1873:SPIbg

[5038] Anonymous. Scientific and practical information. Scientific American, 29(19):289, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
289a.pdf.

Anonymous:1873:BAc

[5039] Anonymous. The British Association. Scientific American, 29 (19):289–290, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
289b.pdf.

Anonymous:1873:PAH

[5040] Anonymous. Proposed addition to the harbor works at Dover England. Scientific American, 29(19):291, November 8, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n19/pdf/scientificamerican11081873-291.pdf.

Anonymous:1873:TCEc

[5041] Anonymous. The three cylinder engine. Scientific American, 29 (19):291–292, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
291a.pdf.

Anonymous:1873:LUSc

[5042] Anonymous. Letter from United States Commissioner Professor R. H. Thurston. Scientific American, 29(19):293, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n19/pdf/scientificamerican11081873-293.pdf.

Anonymous:1873:FVS

[5043] Anonymous. Fruit and vegetable slicer. Scientific American, 29 (19):294, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
294.pdf.

Anonymous:1873:TAL

[5044] Anonymous. Tent attachment for life boats. Scientific American, 29(19):294, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
294a.pdf.

Anonymous:1873:PTSb

[5045] Anonymous. Planer-toothed saw and adjustable iswage. Scientific American, 29(19):294, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
294b.pdf.

Anonymous:1873:SRS

[5046] Anonymous. Safety railway switches. Scientific American, 29 (19):294, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
294c.pdf.

Anonymous:1873:IFBb

[5047] Anonymous. Improved furnace for buildings. Scientific American, 29(19):295, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
295.pdf.

Anonymous:1873:LCC

[5048] Anonymous. Lloyd’s car coupling. Scientific American, 29(19): 295, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
295a.pdf.

Anonymous:1873:NEAb

[5049] Anonymous. The new exploration of the Amazon River, by Professor Orton. — up the Amazons. Scientific American, 29(19):296, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n19/pdf/scientificamerican11081873-296.pdf.

Anonymous:1873:ANg

[5050] Anonymous. Astronomical notes. Scientific American, 29(19): 297, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
297.pdf.

Anonymous:1873:RAFbs

[5051] Anonymous. Recent American and foreign patents. Scientific American, 29(19):297–298, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
297a.pdf.

Anonymous:1873:ACbu

[5052] Anonymous. Answers to correspondents. Scientific American, 29 (19):299–300, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
299b.pdf.

Anonymous:1873:CRbo

[5053] Anonymous. Communications received. Scientific American, 29 (19):300, November 8, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n19/pdf/scientificamerican11081873-
300.pdf.

Anonymous:1873:Cck

[5054] Anonymous. Correspondence. Scientific American, 29(20):308, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n20/pdf/scientificamerican11151873-308.pdf.

Anonymous:1873:BPbu

[5055] Anonymous. Business and personal. Scientific American, 29 (20):315, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
315.pdf.

Anonymous:1873:NQbo

[5056] Anonymous. Notes and queries. Scientific American, 29(20): 315, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
315a.pdf.

Anonymous:1873:IIbs

[5057] Anonymous. Index of inventions. Scientific American, 29(20): 316, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
316a.pdf.

Anonymous:1873:HVS

[5058] Anonymous. Holborn viaduct subways. Scientific American, 29 (20):303, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
303.pdf.

Anonymous:1873:PLC

[5059] Anonymous. The panic and the laboring classes. Scientific American, 29(20):304, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
304.pdf.

Anonymous:1873:PHG

[5060] Anonymous. The progress of the Hell Gate excavations. Scientific American, 29(20):304, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
304a.pdf.

Anonymous:1873:RIG

[5061] Anonymous. A recent improvement in gas making. Scientific American, 29(20):304, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
304b.pdf.

Anonymous:1873:PSd

[5062] Anonymous. Postal science. Scientific American, 29(20):304, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n20/pdf/scientificamerican11151873-304c.pdf.

Anonymous:1873:MCP

[5063] Anonymous. On the molecular changes produced by variations of temperature. Scientific American, 29(20):304–305, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n20/pdf/scientificamerican11151873-304d.pdf.

Anonymous:1873:MSCb

[5064] Anonymous. Medical science in court. Scientific American, 29 (20):305, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
305.pdf.

Anonymous:1873:GBT

[5065] Anonymous. The government boiler tests. Scientific American, 29 (20):305, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
305a.pdf.

Anonymous:1873:SPIbh

[5066] Anonymous. Scientific and practical information. Scientific American, 29(20):305, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
305b.pdf.

Anonymous:1873:AL

[5067] Anonymous. American lighthouses. Scientific American, 29(20): 306, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
306.pdf.

Anonymous:1873:NIG

[5068] Anonymous. A new improvement in gas manufacture. Scientific American, 29(20):307–308, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
307.pdf.

Anonymous:1873:AASa

[5069] Anonymous. American Academy of Sciences. Scientific American, 29(20):309, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
309.pdf.

Anonymous:1873:IFLb

[5070] Anonymous. Improved frost log dog. Scientific American, 29 (20):310, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
310.pdf.

Anonymous:1873:CCSc

[5071] Anonymous. Combination corn sheller boot jack etc. Scientific American, 29(20):310, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
310a.pdf.

Anonymous:1873:LSS

[5072] Anonymous. Lightning sheep shears. Scientific American, 29 (20):310, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
310b.pdf.

Anonymous:1873:GBS

[5073] Anonymous. The great bridge at St Louis. Scientific American, 29 (20):311–313, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
311.pdf.

Anonymous:1873:RAFbt

[5074] Anonymous. Recent American and foreign patents. Scientific American, 29(20):313–314, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
313.pdf.

Anonymous:1873:ACbv

[5075] Anonymous. Answers to correspondents. Scientific American, 29 (20):315–316, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
315b.pdf.

Anonymous:1873:CRbp

[5076] Anonymous. Communications received. Scientific American, 29 (20):316, November 15, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n20/pdf/scientificamerican11151873-
316.pdf.

Anonymous:1873:Ccl

[5077] Anonymous. Correspondence. Scientific American, 29(21):324, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n21/pdf/scientificamerican11221873-324.pdf.

Anonymous:1873:BNBba

[5078] Anonymous. Books: New books and publications. Scientific American, 29(21):330, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
330.pdf.

Anonymous:1873:BPbv

[5079] Anonymous. Business and personal. Scientific American, 29 (21):331, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
331.pdf.

Anonymous:1873:NQbp

[5080] Anonymous. Notes and queries. Scientific American, 29(21): 331, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
331a.pdf.

Anonymous:1873:IIbt

[5081] Anonymous. Index of inventions. Scientific American, 29(21): 332, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
332a.pdf.

Anonymous:1873:IMA

[5082] Anonymous. Improved manufacture of artificial fuel. Scientific American, 29(21):319–320, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
319.pdf.

Anonymous:1873:WDH

[5083] Anonymous. What to do in hard times. Scientific American, 29 (21):320, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
320.pdf.

Anonymous:1873:GLA

[5084] Anonymous. A great “Lick” in astronomy — the million dollar telescope provided for. Scientific American, 29(21):320–321, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n21/pdf/scientificamerican11221873-320a.pdf.

Anonymous:1873:EPP

[5085] Anonymous. English printing presses in America. Scientific American, 29(21):321, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
321.pdf.

Anonymous:1873:NMd

[5086] Anonymous. The November meteors. Scientific American, 29 (21):321, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
321a.pdf.

Anonymous:1873:DS

[5087] Anonymous. Decorated scientists. Scientific American, 29(21): 321, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
321b.pdf.

Anonymous:1873:RM

[5088] Anonymous. Ramming the mold. Scientific American, 29(21): 322, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
322.pdf.

Anonymous:1873:PTC

[5089] Anonymous. Postal telegraph carriage. Scientific American, 29 (21):323, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
323.pdf.

Anonymous:1873:STC

[5090] Anonymous. Steam traveling cranes at the Vienna Exposition. Scientific American, 29(21):323, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
323a.pdf.

Anonymous:1873:QSA

[5091] Anonymous. Quantitative spectrum analysis. Scientific American, 29(21):325, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
325.pdf.

Anonymous:1873:IFCb

[5092] Anonymous. Improved fireproof construction. Scientific American, 29(21):326, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
326.pdf.

Anonymous:1873:ASG

[5093] Anonymous. Anthon’s sewer gas trap. Scientific American, 29 (21):326, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
326a.pdf.

Anonymous:1873:MMC

[5094] Anonymous. A mansion and museum combined. Scientific American, 29(21):327–328, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
327.pdf.

Anonymous:1873:AASb

[5095] Anonymous. American Academy of Sciences. Scientific American, 29 (21):328–329, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
328.pdf.

Anonymous:1873:SPIbi

[5096] Anonymous. Scientific and practical information. Scientific American, 29(21):329, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
329.pdf.

Anonymous:1873:RAFbu

[5097] Anonymous. Recent American and foreign patents. Scientific American, 29(21):329–330, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
329a.pdf.

Anonymous:1873:ACbw

[5098] Anonymous. Answers to correspondents. Scientific American, 29 (21):331–332, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
331b.pdf.

Anonymous:1873:CRbq

[5099] Anonymous. Communications received. Scientific American, 29 (21):332, November 22, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n21/pdf/scientificamerican11221873-
332.pdf.

Anonymous:1873:Ccm

[5100] Anonymous. Correspondence. Scientific American, 29(22):340, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n22/pdf/scientificamerican11291873-340.pdf.

Anonymous:1873:BPbw

[5101] Anonymous. Business and personal. Scientific American, 29 (22):347, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
347.pdf.

Anonymous:1873:NQbq

[5102] Anonymous. Notes and queries. Scientific American, 29(22): 347, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
347a.pdf.

Anonymous:1873:IIbu

[5103] Anonymous. Index of inventions. Scientific American, 29(22): 348–349, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
348a.pdf.

Anonymous:1873:KMS

[5104] Anonymous. The K. M. I. steam engine. Scientific American, 29 (22):335, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
335.pdf.

Anonymous:1873:VPe

[5105] Anonymous. The value of patents. Scientific American, 29(22): 336, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
336.pdf.

Anonymous:1873:NID

[5106] Anonymous. A new and important discovery relating to the behavior of metals under stress. Scientific American, 29(22):336, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n22/pdf/scientificamerican11291873-336a.pdf.

Anonymous:1873:PST

[5107] Anonymous. Patents for simple things. Scientific American, 29 (22):336, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
336b.pdf.

Anonymous:1873:BR

[5108] Anonymous. Boston rebuilt. Scientific American, 29(22):336– 337, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
336c.pdf.

Anonymous:1873:PMC

[5109] Anonymous. Physical and mental over-culture. Scientific American, 29(22):337, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
337.pdf.

Anonymous:1873:BE

[5110] Anonymous. Boiler explosion. Scientific American, 29(22):337, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n22/pdf/scientificamerican11291873-337a.pdf.

Anonymous:1873:PBa

[5111] Anonymous. Phosphor-bronze. Scientific American, 29(22):337, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n22/pdf/scientificamerican11291873-337b.pdf.

Anonymous:1873:SPIbj

[5112] Anonymous. Scientific and practical information. Scientific American, 29(22):337, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
337c.pdf.

Anonymous:1873:ICAa

[5113] Anonymous. Improved churning apparatus. Scientific American, 29(22):338, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
338.pdf.

Anonymous:1873:BIT

[5114] Anonymous. Bartholomew’s improved torch. Scientific American, 29 (22):338–339, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
338a.pdf.

Anonymous:1873:PHF

[5115] Anonymous. Pardee Hall and its founder. Scientific American, 29(22):339, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
339.pdf.

Anonymous:1873:WCT

[5116] Anonymous. Whitehead’s combination tool. Scientific American, 29 (22):339–340, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
339a.pdf.

Sharples:1873:LN

[5117] S. P. Sharples. Laboratory notes. Scientific American, 29(22): 340, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
340a.pdf.

Anonymous:1873:PIc

[5118] Anonymous. Prizes for improvements. Scientific American, 29 (22):341, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
341.pdf.

Anonymous:1873:ITWb

[5119] Anonymous. Improved turbine wheel governor. Scientific American, 29(22):342, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
342.pdf.

Anonymous:1873:GBM

[5120] Anonymous. Gas bracket match safe. Scientific American, 29 (22):342, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
342a.pdf.

Anonymous:1873:BMW

[5121] Anonymous. Blacksmith’s measuring wheel. Scientific American, 29(22):343, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
343.pdf.

Anonymous:1873:BNS

[5122] Anonymous. The black necked swan. Scientific American, 29(22): 343–344, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
343a.pdf.

Anonymous:1873:DIC

[5123] Anonymous. The deep injection of chloroform. Scientific American, 29(22):344, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
344.pdf.

Anonymous:1873:SHd

[5124] Anonymous. Science and health. Scientific American, 29(22): 345, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
345.pdf.

Anonymous:1873:RAFbv

[5125] Anonymous. Recent American and foreign patents. Scientific American, 29(22):345–346, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
345a.pdf.

Anonymous:1873:ACbx

[5126] Anonymous. Answers to correspondents. Scientific American, 29 (22):347–348, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
347b.pdf.

Anonymous:1873:CRbr

[5127] Anonymous. Communications received. Scientific American, 29 (22):348, November 29, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n22/pdf/scientificamerican11291873-
348.pdf.

Anonymous:1873:Ccn

[5128] Anonymous. Correspondence. Scientific American, 29(23):356, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n23/pdf/scientificamerican12061873-356.pdf.

Anonymous:1873:BPbx

[5129] Anonymous. Business and personal. Scientific American, 29 (23):363, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
363.pdf.

Anonymous:1873:IIbv

[5130] Anonymous. Index of inventions. Scientific American, 29(23): 364–365, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
364a.pdf.

Anonymous:1873:RPE

[5131] Anonymous. Recent progress in electromagnetism. Scientific American, 29(23):351–352, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
351.pdf.

Anonymous:1873:NYC

[5132] Anonymous. New York and the Centennial. Scientific American, 29(23):352, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
352.pdf.

Anonymous:1873:SSR

[5133] Anonymous. The signal service report. Scientific American, 29 (23):352, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
352a.pdf.

Anonymous:1873:RED

[5134] Anonymous. Recent experiments with diamonds. Scientific American, 29(23):353, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
353.pdf.

Anonymous:1873:PId

[5135] Anonymous. Photography of the invisible. Scientific American, 29(23):353, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
353a.pdf.

Anonymous:1873:SK

[5136] Anonymous. Science in the kitchen. Scientific American, 29 (23):353, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
353b.pdf.

Anonymous:1873:SPIbk

[5137] Anonymous. Scientific and practical information. Scientific American, 29(23):353–354, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
353c.pdf.

Anonymous:1873:TTT

[5138] Anonymous. The Tom Thumb telegraph. Scientific American, 29(23):354, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
354.pdf.

Anonymous:1873:MT

[5139] Anonymous. The miniature telegraph. Scientific American, 29 (23):354, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
354a.pdf.

Anonymous:1873:CRRa

[5140] Anonymous. The challenge reversible rotary hoisting machine. Scientific American, 29(23):355, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
355.pdf.

Anonymous:1873:SEP

[5141] Anonymous. Subjects for engineering papers. Scientific American, 29(23):356, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
356a.pdf.

Barnard:1873:GTRa

[5142] F. A. P. Barnard. The germ theory and its relations to hygiene. Scientific American, 29(23):357, December 6, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n23/pdf/scientificamerican12061873-357.pdf.

Anonymous:1873:IBCc

[5143] Anonymous. Improved box corner grooving machine. Scientific American, 29(23):358, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
358.pdf.

Anonymous:1873:YP

[5144] Anonymous. The yucca pendula. Scientific American, 29(23): 358–359, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
358a.pdf.

Anonymous:1873:PPBa

[5145] Anonymous. The pashiuba palm of Brazil. Scientific American, 29(23):359, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
359.pdf.

Anonymous:1873:ANh

[5146] Anonymous. Astronomical notes. Scientific American, 29(23):360, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n23/pdf/scientificamerican12061873-360.pdf.

Young:1873:N

[5147] C. A. Young. The nebulæ. Scientific American, 29(23):360, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n23/pdf/scientificamerican12061873-360a.pdf.

Anonymous:1873:DCz

[5148] Anonymous. Decisions of the courts. Scientific American, 29 (23):361, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
361.pdf.

Anonymous:1873:RAFbw

[5149] Anonymous. Recent American and foreign patents. Scientific American, 29(23):361–362, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
361a.pdf.

Anonymous:1873:ACby

[5150] Anonymous. Answers to correspondence. Scientific American, 29 (23):363–364, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
363a.pdf.

Anonymous:1873:CRbs

[5151] Anonymous. Communications received. Scientific American, 29 (23):364, December 6, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n23/pdf/scientificamerican12061873-
364.pdf.

Anonymous:1873:Cco

[5152] Anonymous. Correspondence. Scientific American, 29(24):372, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n24/pdf/scientificamerican12131873-372.pdf.

Anonymous:1873:BPby

[5153] Anonymous. Business and personal. Scientific American, 29 (24):379, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
379.pdf.

Anonymous:1873:IIbw

[5154] Anonymous. Index of inventions. Scientific American, 29(24): 380, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
380a.pdf.

Anonymous:1873:AIS

[5155] Anonymous. Aschenbrenner’s improved sugar-making process. Scientific American, 29(24):367, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
367.pdf.

Anonymous:1873:WOS

[5156] Anonymous. A word to our subscribers. Scientific American, 29 (24):368, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
368.pdf.

Anonymous:1873:BT

[5157] Anonymous. Better times. Scientific American, 29(24):368, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n24/pdf/scientificamerican12131873-368a.pdf.

Anonymous:1873:PPc

[5158] Anonymous. On the purchase of patents. Scientific American, 29 (24):368, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
368b.pdf.

Anonymous:1873:CTU

[5159] Anonymous. Concerning a telescope of unlimited power. Scientific American, 29(24):368–369, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
368c.pdf.

Anonymous:1873:PPAa

[5160] Anonymous. The protection of plants by artificial clouds. Scientific American, 29(24):369, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
369.pdf.

Anonymous:1873:EPF

[5161] Anonymous. Eloquence and patent fuel. Scientific American, 29 (24):369, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
369a.pdf.

Anonymous:1873:SMc

[5162] Anonymous. The suppressed member. Scientific American, 29 (24):369, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
369b.pdf.

Anonymous:1873:SPIbl

[5163] Anonymous. Scientific and practical information. Scientific American, 29(24):369, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
369c.pdf.

Anonymous:1873:IWMb

[5164] Anonymous. The iron works at Merthyr Tyndall, South Wales. Scientific American, 29(24):370, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
370.pdf.

Anonymous:1873:BHBb

[5165] Anonymous. The baby hippopotamus’ birthday. Scientific American, 29(24):370, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
370a.pdf.

Anonymous:1873:FP

[5166] Anonymous. Flax and its preparation. Scientific American, 29 (24):371, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
371.pdf.

Anonymous:1873:ORP

[5167] Anonymous. The Oreodoxa Regia palm. Scientific American, 29 (24):371, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
371a.pdf.

Anonymous:1873:CHT

[5168] Anonymous. Completion of the Hoosac Tunnel. Scientific American, 29(24):372, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
372a.pdf.

Anonymous:1873:IVU

[5169] Anonymous. Improvements in the ventilation of the United States Senate Chamber. Scientific American, 29(24):372, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n24/pdf/scientificamerican12131873-372b.pdf.

Anonymous:1873:UN

[5170] Anonymous. The uses of nickel. Scientific American, 29(24): 372, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
372c.pdf.

Anonymous:1873:SSd

[5171] Anonymous. The savings of science. Scientific American, 29 (24):373, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
373.pdf.

Anonymous:1873:CSEb

[5172] Anonymous. Combined snow and Earth shovel and plow. Scientific American, 29(24):374, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
374.pdf.

Anonymous:1873:DIP

[5173] Anonymous. Dunning’s improved pipe coupling. Scientific American, 29(24):374, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
374a.pdf.

Anonymous:1873:AFR

[5174] Anonymous. Apparatus for filtering and rectifying spirits. Scientific American, 29(24):374, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
374b.pdf.

Anonymous:1873:Ccq

[5175] Anonymous. The Cocuyo. Scientific American, 29(24):375, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n24/pdf/scientificamerican12131873-375.pdf.

Anonymous:1873:BAd

[5176] Anonymous. The Brighton Aquarium. Scientific American, 29 (24):375, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
375a.pdf.

Barnard:1873:GTRb

[5177] F. A. P. Barnard. The germ theory and its relations to hygiene. Scientific American, 29(24):376–377, December 13, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n24/pdf/scientificamerican12131873-376.pdf.

Anonymous:1873:PSSb

[5178] Anonymous. A practical system for the sale of patents. Scientific American, 29(24):377, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
377.pdf.

Anonymous:1873:DCba

[5179] Anonymous. Decisions of the courts. Scientific American, 29 (24):378, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
378.pdf.

Anonymous:1873:RAFbx

[5180] Anonymous. Recent American and foreign patents. Scientific American, 29(24):378, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
378a.pdf.

Anonymous:1873:ACbz

[5181] Anonymous. Answers to correspondents. Scientific American, 29 (24):379–380, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
379a.pdf.

Anonymous:1873:CRbt

[5182] Anonymous. Communications received. Scientific American, 29 (24):380, December 13, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n24/pdf/scientificamerican12131873-
380.pdf.

Anonymous:1873:Cct

[5183] Anonymous. Correspondence. Scientific American, 29(25):388– 389, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
388.pdf.

Anonymous:1873:BNBbb

[5184] Anonymous. Books: New books and publications. Scientific American, 29(25):394, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
394.pdf.

Anonymous:1873:BPbz

[5185] Anonymous. Business and personal. Scientific American, 29 (25):395, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
395.pdf.

Anonymous:1873:IIbx

[5186] Anonymous. Index of inventions. Scientific American, 29(25): 395–396, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
395c.pdf.

Anonymous:1873:JM

[5187] Anonymous. The jamin magnet. Scientific American, 29(25): 383–384, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
383.pdf.

Anonymous:1873:PCf

[5188] Anonymous. Publishers’ card. Scientific American, 29(25):384, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n25/pdf/scientificamerican12201873-384.pdf.

Anonymous:1873:LVD

[5189] Anonymous. The loss of the Ville Du Havre. Scientific American, 29(25):384, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
384a.pdf.

Anonymous:1873:SSM

[5190] Anonymous. A steam snow melter. Scientific American, 29 (25):384, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
384b.pdf.

Anonymous:1873:NFC

[5191] Anonymous. A new fuel — carbonite. Scientific American, 29 (25):384, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
384c.pdf.

Anonymous:1873:CTI

[5192] Anonymous. The coal tar interest. Scientific American, 29(25): 384–385, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
384d.pdf.

Anonymous:1873:ROT

[5193] Anonymous. Recent ordnance trials. Scientific American, 29 (25):385, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
385.pdf.

Anonymous:1873:SPIbm

[5194] Anonymous. Scientific and practical information. Scientific American, 29(25):385, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
385a.pdf.

Anonymous:1873:AAB

[5195] Anonymous. Albinism among birds. Scientific American, 29 (25):386, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
386.pdf.

Anonymous:1873:ICPb

[5196] Anonymous. Improved coffee pot. Scientific American, 29(25): 386, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
386a.pdf.

Anonymous:1873:HDC

[5197] Anonymous. The hull dock Company’s new offices. Scientific American, 29(25):387–388, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
387.pdf.

Anonymous:1873:UBM

[5198] Anonymous. The universal boring machine. Scientific American, 29(25):390, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
390.pdf.

Anonymous:1873:CCHa

[5199] Anonymous. Combined cooking, heating, and drying apparatus. Scientific American, 29(25):390, December 20, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n25/pdf/scientificamerican12201873-390a.pdf.

Anonymous:1873:DCbb

[5200] Anonymous. Design for a conservatory. Scientific American, 29 (25):391, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
391.pdf.

Barnard:1873:GTRc

[5201] F. A. P. Barnard. The germ theory and its relations to hygiene. Scientific American, 29(25):392–393, December 20, 1873. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v29/
n25/pdf/scientificamerican12201873-392.pdf.

Anonymous:1873:DCbc

[5202] Anonymous. Decisions of the courts. Scientific American, 29 (25):393, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
393.pdf.

Anonymous:1873:RAFby

[5203] Anonymous. Recent American and foreign patents. Scientific American, 29(25):393–394, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
393a.pdf.

Anonymous:1873:ACcd

[5204] Anonymous. Answers to correspondents. Scientific American, 29 (25):395, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
395a.pdf.

Anonymous:1873:CRbu

[5205] Anonymous. Communications received. Scientific American, 29 (25):395, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
395b.pdf.

Anonymous:1873:CPh

[5206] Anonymous. Canadian patents. Scientific American, 29(25): 396, December 20, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n25/pdf/scientificamerican12201873-
396.pdf.

Anonymous:1873:Ccu

[5207] Anonymous. Correspondence. Scientific American, 29(26):404, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n26/pdf/scientificamerican12271873-404.pdf.

Anonymous:1873:BPcb

[5208] Anonymous. Business and personal. Scientific American, 29 (26):409, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
409.pdf.

Anonymous:1873:IIby

[5209] Anonymous. Index of inventions. Scientific American, 29(26): 410, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
410a.pdf.

Anonymous:1873:Ig

[5210] Anonymous. Index. Scientific American, 29(26):411–412, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n26/pdf/scientificamerican12271873-411.pdf.

Gesner:1873:MPP

[5211] J. F. Gesner. The manufacture of plaster of Paris. Scientific American, 29(26):399–400, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
399.pdf.

Anonymous:1873:PCg

[5212] Anonymous. Publishers’ card. Scientific American, 29(26):400, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n26/pdf/scientificamerican12271873-400.pdf.

Anonymous:1873:MIV

[5213] Anonymous. Measurement of the initial velocity of projectiles. Scientific American, 29(26):400, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
400a.pdf.

Anonymous:1873:PAEb

[5214] Anonymous. The practical aspect of the eight hour law. Scientific American, 29(26):400–401, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
400b.pdf.

Anonymous:1873:JBA

[5215] Anonymous. Jonah and the Belgian Academy. Scientific American, 29(26):401, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
401.pdf.

Anonymous:1873:HCF

[5216] Anonymous. Hints for cheap floral decoration. Scientific American, 29(26):401, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
401a.pdf.

Anonymous:1873:ETC

[5217] Anonymous. An easy telegraph code. Scientific American, 29 (26):401, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
401b.pdf.

Anonymous:1873:DPA

[5218] Anonymous. Death of Professor Agassiz. Scientific American, 29 (26):401, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
401c.pdf.

Anonymous:1873:WSJ

[5219] Anonymous. The weeping sophora of Japan. Scientific American, 29(26):402, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
402.pdf.

Anonymous:1873:FGC

[5220] Anonymous. Friction geared centrifugal pump. Scientific American, 29(26):402, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
402a.pdf.

Anonymous:1873:ATL

[5221] Anonymous. The automatic taper lamp. Scientific American, 29 (26):402, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
402b.pdf.

Anonymous:1873:ATV

[5222] Anonymous. An automaton Transit of Venus. Scientific American, 29(26):403, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
403.pdf.

Anonymous:1873:RGN

[5223] Anonymous. Recent geographical notes. Scientific American, 29 (26):403, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
403a.pdf.

Anonymous:1873:CPi

[5224] Anonymous. Careless plumbing. Scientific American, 29(26): 403, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
403b.pdf.

Anonymous:1873:ARL

[5225] Anonymous. The atmospheric railway in London. Scientific American, 29(26):404–405, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
404a.pdf.

Anonymous:1873:SPIbn

[5226] Anonymous. Scientific and practical information. Scientific American, 29(26):405, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
405.pdf.

Anonymous:1873:CSAa

[5227] Anonymous. The crystal self-acting table fountain. Scientific American, 29(26):406, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
406.pdf.

Anonymous:1873:CEMb

[5228] Anonymous. Combined extension measuring rod and divider. Scientific American, 29(26):406, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
406a.pdf.

Anonymous:1873:TG

[5229] Anonymous. Toy gymnast. Scientific American, 29(26):406, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v29/n26/pdf/scientificamerican12271873-406b.pdf.

Anonymous:1873:MDL

[5230] Anonymous. Moistening device for lame animals. Scientific American, 29(26):406, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
406c.pdf.

Anonymous:1873:CFW

[5231] Anonymous. Chinese fresh water fish. Scientific American, 29(26): 407–408, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
407.pdf.

Anonymous:1873:RAFbz

[5232] Anonymous. Recent American and foreign patents. Scientific American, 29(26):408, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
408.pdf.

Anonymous:1873:ACce

[5233] Anonymous. Answers to correspondents. Scientific American, 29 (26):409, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
409a.pdf.

Anonymous:1873:CRbv

[5234] Anonymous. Communications received. Scientific American, 29 (26):410, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
410.pdf.

Anonymous:1873:CPj

[5235] Anonymous. Canadian patents. Scientific American, 29(26):410– 411, December 27, 1873. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v29/n26/pdf/scientificamerican12271873-
410b.pdf.

Anonymous:1874:Ca

[5236] Anonymous. Correspondence. Scientific American, 30(1):7, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-7.pdf.

Anonymous:1874:BNBa

[5237] Anonymous. Books: New books and publications. Scientific American, 30(1):9, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
9.pdf.

Anonymous:1874:BPa

[5238] Anonymous. Business and personal. Scientific American, 30(1):11, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-11.pdf.

Anonymous:1874:IIa

[5239] Anonymous. Index of inventions. Scientific American, 30(1):12, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-12a.pdf.

Anonymous:1874:AWP

[5240] Anonymous. The Augsburg Web printing press. Scientific American, 30(1):1, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
1.pdf.

Anonymous:1874:LAa

[5241] Anonymous. Louis Agassiz. Scientific American, 30(1):2, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-2.pdf.

Anonymous:1874:LYS

[5242] Anonymous. Look to your stoves. Scientific American, 30(1):2, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-2a.pdf.

Anonymous:1874:RCE

[5243] Anonymous. The report of the Chief Engineer of the Navy. Scientific American, 30(1):2–3, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
2b.pdf.

Anonymous:1874:ME

[5244] Anonymous. The manufacture of earthenware. Scientific American, 30(1):3, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
3.pdf.

Thurston:1874:NRE

[5245] R. H. Thurston. Note relative to the estimation of the commercial value of coals containing large quantities of ash. Scientific American, 30(1):3, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-3a.pdf.

Anonymous:1874:AIB

[5246] Anonymous. The Albert iron bridge. Scientific American, 30(1):4–5, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-4.pdf.

Anonymous:1874:WJP

[5247] Anonymous. Worcester Japanese porcelain at Vienna. Scientific American, 30(1):5, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
5.pdf.

Anonymous:1874:IFC

[5248] Anonymous. Improved furniture casters. Scientific American, 30(1):6, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-6.pdf.

Anonymous:1874:ABa

[5249] Anonymous. Automatic brake. Scientific American, 30(1):6, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-6a.pdf.

Anonymous:1874:IP

[5250] Anonymous. Improved plow. Scientific American, 30(1):6, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-6b.pdf.

Anonymous:1874:SN

[5251] Anonymous. Samuel Nelson. Scientific American, 30(1):7, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-7a.pdf.

Anonymous:1874:SPIa

[5252] Anonymous. Scientific and practical information. Scientific American, 30(1):7, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
7b.pdf.

Anonymous:1874:ANa

[5253] Anonymous. Astronomical notes. Scientific American, 30(1):8, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-8.pdf.

Anonymous:1874:DCa

[5254] Anonymous. Decisions of the courts. Scientific American, 30(1):9, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-9a.pdf.

Anonymous:1874:CDa

[5255] Anonymous. Commissioner’s decisions. Scientific American, 30(1):9, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-9b.pdf.

Anonymous:1874:VPHa

[5256] Anonymous. Value of patents, and how to obtain them. Scientific American, 30(1):9, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
9c.pdf.

Anonymous:1874:RAFa

[5257] Anonymous. Recent American and foreign patents. Scientific American, 30(1):10, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
10.pdf.

Anonymous:1874:ACa

[5258] Anonymous. Answers to correspondents. Scientific American, 30(1):11, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-11a.pdf.

Anonymous:1874:CRa

[5259] Anonymous. Communications received. Scientific American, 30(1):12, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-12.pdf.

Anonymous:1874:CPa

[5260] Anonymous. Canadian patents. Scientific American, 30(1):12, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n1/pdf/scientificamerican01031874-12b.pdf.

Anonymous:1874:HOPa

[5261] Anonymous. How to obtain patents and caveats in Canada. Scientific American, 30(1):12, January 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n1/pdf/scientificamerican01031874-
12c.pdf.

Anonymous:1874:Cb

[5262] Anonymous. Correspondence. Scientific American, 30(2):20, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-20.pdf.

Anonymous:1874:BNBb

[5263] Anonymous. Books: New books and publications. Scientific American, 30(2):25, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
25a.pdf.

Anonymous:1874:BPb

[5264] Anonymous. Business and personal. Scientific American, 30(2):27, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-27.pdf.

Anonymous:1874:IIb

[5265] Anonymous. Index of inventions. Scientific American, 30(2):28, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-28a.pdf.

Anonymous:1874:HCR

[5266] Anonymous. The Hotchkiss cannon revolver. Scientific American, 30(2):15–16, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
15.pdf.

Anonymous:1874:NPC

[5267] Anonymous. New patent conventions. Scientific American, 30 (2):16, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
16.pdf.

Anonymous:1874:PMTa

[5268] Anonymous. The prime movers and their recent progress. Scientific American, 30(2):16–17, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
16a.pdf.

Anonymous:1874:PCa

[5269] Anonymous. Progress of the Centennial. Scientific American, 30(2):17, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
17.pdf.

Anonymous:1874:EEa

[5270] Anonymous. Expert engineering. Scientific American, 30(2):17, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-17a.pdf.

Anonymous:1874:OS

[5271] Anonymous. To our subscribers. Scientific American, 30(2):17, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-17b.pdf.

Anonymous:1874:IPFa

[5272] Anonymous. Improved portable fence. Scientific American, 30 (2):18, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
18.pdf.

Anonymous:1874:ITG

[5273] Anonymous. Improved trotting gear. Scientific American, 30 (2):18, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
18a.pdf.

Anonymous:1874:IAa

[5274] Anonymous. Improved atomizer. Scientific American, 30(2):18, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-18b.pdf.

Anonymous:1874:ITH

[5275] Anonymous. Improved tool holder. Scientific American, 30(2):18, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-18c.pdf.

Anonymous:1874:SPIb

[5276] Anonymous. Scientific and practical information. Scientific American, 30(2):19, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
19.pdf.

Anonymous:1874:M

[5277] Anonymous. Metaline. Scientific American, 30(2):19, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-19a.pdf.

Anonymous:1874:ACSa

[5278] Anonymous. Alumina, from the clay to the Sapphire. Scientific American, 30(2):20–21, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
20a.pdf.

Anonymous:1874:ICO

[5279] Anonymous. Inside a Church organ. Scientific American, 30(2):21, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-21.pdf.

Anonymous:1874:BFL

[5280] Anonymous. Barrow in Furness, Lancashire, England. Scientific American, 30(2):22–23, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
22.pdf.

Anonymous:1874:NEAa

[5281] Anonymous. The new exploration of the Amazon River, by Professor Orton. — over the Andes. Scientific American, 30(2):24, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-24.pdf.

Anonymous:1874:DCb

[5282] Anonymous. Decisions of the courts. Scientific American, 30(2):25, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-25.pdf.

Anonymous:1874:VPHb

[5283] Anonymous. Value of patents, and how to obtain them. Scientific American, 30(2):25, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
25b.pdf.

Anonymous:1874:RAFb

[5284] Anonymous. Recent American and foreign patents. Scientific American, 30(2):25–26, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
25c.pdf.

Anonymous:1874:ACb

[5285] Anonymous. Answers to correspondents. Scientific American, 30 (2):27–28, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
27a.pdf.

Anonymous:1874:CRb

[5286] Anonymous. Communications received. Scientific American, 30 (2):28, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
28.pdf.

Anonymous:1874:CPb

[5287] Anonymous. Canadian patents. Scientific American, 30(2):28, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n2/pdf/scientificamerican01101874-28b.pdf.

Anonymous:1874:HOPb

[5288] Anonymous. How to obtain patents and caveats in Canada. Scientific American, 30(2):28–29, January 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n2/pdf/scientificamerican01101874-
28c.pdf.

Anonymous:1874:Cc

[5289] Anonymous. Correspondence. Scientific American, 30(3):36, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-36.pdf.

Anonymous:1874:BNBc

[5290] Anonymous. Books: New books and publications. Scientific American, 30(3):41, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
41b.pdf.

Anonymous:1874:BPc

[5291] Anonymous. Business and personal. Scientific American, 30(3):43, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-43.pdf.

Anonymous:1874:IIc

[5292] Anonymous. Index of inventions. Scientific American, 30(3):44, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-44.pdf.

Anonymous:1874:NSC

[5293] Anonymous. The new State Capitol at Hartford, Conn. Scientific American, 30(3):31, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
31.pdf.

Anonymous:1874:VER

[5294] Anonymous. The Vienna Exhibition and its results. Scientific American, 30(3):32, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
32.pdf.

Anonymous:1874:POP

[5295] Anonymous. Patent Office printing. Scientific American, 30(3):32, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-32a.pdf.

Anonymous:1874:SRa

[5296] Anonymous. Science record for 1874. Scientific American, 30 (3):32, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
32b.pdf.

Anonymous:1874:VKC

[5297] Anonymous. The value of a kernel of corn. Scientific American, 30(3):32–33, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
32c.pdf.

Anonymous:1874:PMTb

[5298] Anonymous. The prime movers and their recent progress. Scientific American, 30(3):33, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
33.pdf.

Anonymous:1874:Oa

[5299] Anonymous. Obituary. Scientific American, 30(3):33, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-33a.pdf.

Anonymous:1874:ITV

[5300] Anonymous. Improved tramway and vehicle. Scientific American, 30(3):34, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
34.pdf.

Anonymous:1874:PHC

[5301] Anonymous. Prosper Henry’s comet. Scientific American, 30 (3):34, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
34a.pdf.

Grant:1874:WPM

[5302] Edward M. Grant. The Wilson process for making wrought iron direct from the ore. Scientific American, 30(3):35, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-35.pdf.

Anonymous:1874:NSO

[5303] Anonymous. A new steam omnibus and street car. Scientific American, 30(3):35–36, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
35a.pdf.

Anonymous:1874:ACSb

[5304] Anonymous. Alumina, from the clay to the Sapphire. Scientific American, 30(3):36–37, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
36a.pdf.

Thurston:1874:TQI

[5305] Thurston. Testing the quality of iron, steel, and other metals, without special apparatus. Scientific American, 30(3):38–39, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-38.pdf.

Anonymous:1874:IPSa

[5306] Anonymous. Improved pruning shears. Scientific American, 30 (3):39, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
39.pdf.

Anonymous:1874:INL

[5307] Anonymous. Improved nut lock. Scientific American, 30(3):39, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-39a.pdf.

Anonymous:1874:SCB

[5308] Anonymous. Starr’s cotton bale fastener. Scientific American, 30(3):39, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
39b.pdf.

Anonymous:1874:NEAb

[5309] Anonymous. The new exploration of the Amazon River, by Professor Orton. — over the Andes. Scientific American, 30(3):40, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-40.pdf.

Anonymous:1874:SPIc

[5310] Anonymous. Scientific and practical information. Scientific American, 30(3):41, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
41.pdf.

Anonymous:1874:DCc

[5311] Anonymous. Decisions of the courts. Scientific American, 30(3):41, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-41a.pdf.

Anonymous:1874:VPHc

[5312] Anonymous. Value of patents, and how to obtain them. Scientific American, 30(3):41, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
41c.pdf.

Anonymous:1874:RAFc

[5313] Anonymous. Recent American and foreign patents. Scientific American, 30(3):41–42, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
41d.pdf.

Anonymous:1874:ACc

[5314] Anonymous. Answers to correspondents. Scientific American, 30 (3):43, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
43a.pdf.

Anonymous:1874:CRc

[5315] Anonymous. Communications received. Scientific American, 30 (3):43, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
43b.pdf.

Anonymous:1874:CPc

[5316] Anonymous. Canadian patents. Scientific American, 30(3):44, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n3/pdf/scientificamerican01171874-44a.pdf.

Anonymous:1874:HOPc

[5317] Anonymous. How to obtain patents and caveats in Canada. Scientific American, 30(3):44, January 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n3/pdf/scientificamerican01171874-
44b.pdf.

Anonymous:1874:Ce

[5318] Anonymous. Correspondence. Scientific American, 30(4):52–53, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-52.pdf.

Anonymous:1874:BPd

[5319] Anonymous. Business and personal. Scientific American, 30(4):58, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-58.pdf.

Anonymous:1874:IId

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-60.pdf.

Anonymous:1874:UIP

[5321] Anonymous. The utilization of iron pyrites. Scientific American, 30(4):48, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
48.pdf.

Anonymous:1874:ICM

[5322] Anonymous. The influence of cargoes of machinery and hardware on ships’ compasses. Scientific American, 30(4):48, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-48a.pdf.

Anonymous:1874:NIA

[5323] Anonymous. New iron alloys. Scientific American, 30(4):48–49, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-48b.pdf.

Anonymous:1874:AR

[5324] Anonymous. Auguste De La Rive. Scientific American, 30(4):49, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-49.pdf.

Anonymous:1874:DSR

[5325] Anonymous. Death in the school room. Scientific American, 30(4):49, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
49a.pdf.

Anonymous:1874:OF

[5326] Anonymous. To our friends. Scientific American, 30(4):49, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-49b.pdf.

Anonymous:1874:SPId

[5327] Anonymous. Scientific and practical information. Scientific American, 30(4):49, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
49c.pdf.

Anonymous:1874:CTOa

[5328] Anonymous. The Centennial Tower one thousand feet high. Scientific American, 30(4):50, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
50.pdf.

Anonymous:1874:RIB

[5329] Anonymous. Recent iron bridge construction. Scientific American, 30(4):51, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
51.pdf.

Anonymous:1874:WAB

[5330] Anonymous. Wardwell’s adjustable bench jack. Scientific American, 30(4):54, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
54.pdf.

Anonymous:1874:ICV

[5331] Anonymous. Improved cooking vessel. Scientific American, 30 (4):54, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
54a.pdf.

Anonymous:1874:IM

[5332] Anonymous. Improved mallet. Scientific American, 30(4):54, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-54b.pdf.

Anonymous:1874:SFM

[5333] Anonymous. A simple freezing machine. Scientific American, 30(4):54, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
54c.pdf.

Anonymous:1874:MOH

[5334] Anonymous. A modern orchid house. Scientific American, 30 (4):55, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
55.pdf.

Anonymous:1874:NEAc

[5335] Anonymous. The new exploration of the Amazon River, by Professor Orton. — over the Andes. Scientific American, 30(4):56, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-56.pdf.

Anonymous:1874:DCd

[5336] Anonymous. Decisions of the courts. Scientific American, 30(4):56, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n4/pdf/scientificamerican01241874-56a.pdf.

Anonymous:1874:RAFd

[5337] Anonymous. Recent American and foreign patents. Scientific American, 30(4):56–57, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
56b.pdf.

Anonymous:1874:ACd

[5338] Anonymous. Answers to correspondents. Scientific American, 30 (4):58–59, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
58a.pdf.

Anonymous:1874:CRd

[5339] Anonymous. Communications received. Scientific American, 30 (4):59, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
59.pdf.

Anonymous:1874:VPHd

[5340] Anonymous. Value of patents and how to obtain them. Scientific American, 30(4):60, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
60a.pdf.

Anonymous:1874:HOPd

[5341] Anonymous. How to obtain patents and caveats in Canada. Scientific American, 30(4):60, January 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n4/pdf/scientificamerican01241874-
60b.pdf.

Anonymous:1874:Cf

[5342] Anonymous. Correspondence. Scientific American, 30(5):68–69, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-68.pdf.

Anonymous:1874:BNBd

[5343] Anonymous. Books: New books and publications. Scientific American, 30(5):74, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
74.pdf.

Anonymous:1874:BPe

[5344] Anonymous. Business and personal. Scientific American, 30(5):75, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-75.pdf.

Anonymous:1874:IIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-76a.pdf.

Anonymous:1874:ERB

[5346] Anonymous. The East River Bridge. Scientific American, 30 (5):63–64, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
63.pdf.

Anonymous:1874:ISEa

[5347] Anonymous. Indicating steam engines. Scientific American, 30 (5):64, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
64.pdf.

Anonymous:1874:DHD

[5348] Anonymous. Dr. Henry Draper’s recent discoveries in spectroscopic analysis. Scientific American, 30(5):64, January 31, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n5/pdf/scientificamerican01311874-64a.pdf.

Anonymous:1874:GB

[5349] Anonymous. Gail borden. Scientific American, 30(5):64, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-64b.pdf.

Anonymous:1874:SCa

[5350] Anonymous. Steam on canals. Scientific American, 30(5):64–65, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-64c.pdf.

Anonymous:1874:YIP

[5351] Anonymous. A Year’s inventive progress. Scientific American, 30(5):65, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
65.pdf.

Anonymous:1874:PPCa

[5352] Anonymous. Proceedings of the patent convention. Scientific American, 30(5):65, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
65a.pdf.

Anonymous:1874:PFa

[5353] Anonymous. Psychic force. Scientific American, 30(5):65, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-65b.pdf.

Anonymous:1874:SPIe

[5354] Anonymous. Scientific and practical information. Scientific American, 30(5):65, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
65c.pdf.

Anonymous:1874:EPB

[5355] Anonymous. Electro-plating balance. Scientific American, 30 (5):66, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
66.pdf.

Anonymous:1874:YMO

[5356] Anonymous. The yama-mai, or oak tree silkworm. Scientific American, 30(5):66, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
66a.pdf.

Anonymous:1874:HEE

[5357] Anonymous. Hydraulic engineering extraordinary. Scientific American, 30(5):67, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
67.pdf.

Anonymous:1874:CSa

[5358] Anonymous. The Canopus stone. Scientific American, 30(5):67–68, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-67a.pdf.

Orton:1874:SMP

[5359] Tames Orton. The silver mines of Peru. Scientific American, 30(5):69, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
69.pdf.

Anonymous:1874:IPCa

[5360] Anonymous. Improved patent candy cutter. Scientific American, 30(5):70, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
70.pdf.

Anonymous:1874:IFB

[5361] Anonymous. Improved furnace bridge wall. Scientific American, 30(5):70, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
70a.pdf.

Anonymous:1874:BDP

[5362] Anonymous. The baffle drill-proof safe. Scientific American, 30 (5):70, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
70b.pdf.

Anonymous:1874:OSC

[5363] Anonymous. Overhead steam cranes for dock purposes. Scientific American, 30(5):71–72, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
71.pdf.

Anonymous:1874:DCe

[5364] Anonymous. Decisions of the courts. Scientific American, 30(5):72, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-72.pdf.

Anonymous:1874:RAFe

[5365] Anonymous. Recent American foreign patents. Scientific American, 30(5):73–74, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
73.pdf.

Anonymous:1874:ACe

[5366] Anonymous. Answers correspondents. Scientific American, 30 (5):75–76, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
75a.pdf.

Anonymous:1874:CRe

[5367] Anonymous. Communications received. Scientific American, 30 (5):76, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n5/pdf/scientificamerican01311874-
76.pdf.

Anonymous:1874:CPd

[5368] Anonymous. Canadian patents. Scientific American, 30(5):76, January 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n5/pdf/scientificamerican01311874-76b.pdf.

Anonymous:1874:Cg

[5369] Anonymous. Correspondence. Scientific American, 30(6):84, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-84.pdf.

Anonymous:1874:BNBe 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
88b.pdf.

Anonymous:1874:BPf

[5371] Anonymous. Business and personal. Scientific American, 30 (6):90, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
90.pdf.

Anonymous:1874:IIf

[5372] Anonymous. Index of inventions. Scientific American, 30(6): 91–92, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
91a.pdf.

Anonymous:1874:GG

[5373] Anonymous. The Gatling gun. Scientific American, 30(6):79–80, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-79.pdf.

Anonymous:1874:IAb

[5374] Anonymous. Importance of advertising. Scientific American, 30 (6):80, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
80.pdf.

Anonymous:1874:VV

[5375] Anonymous. Is vitality vital? Scientific American, 30(6):80, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-80a.pdf.

Anonymous:1874:DCT

[5376] Anonymous. The Dover and Calais Tunnel. Scientific American, 30(6):80, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
80b.pdf.

Anonymous:1874:STa

[5377] Anonymous. The Siamese twins. Scientific American, 30(6): 80–81, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
80c.pdf.

Anonymous:1874:OHT

[5378] Anonymous. The one hundred thousand dollar canal reward. Scientific American, 30(6):81, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
81.pdf.

Anonymous:1874:TCB

[5379] Anonymous. The Turner car brake — application for an extension. Scientific American, 30(6):81, February 7, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n6/pdf/scientificamerican02071874-81a.pdf.

Anonymous:1874:NEAd

[5380] Anonymous. The new England Association of inventors and patent owners. Scientific American, 30(6):81, February 7, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n6/pdf/scientificamerican02071874-81b.pdf.

Anonymous:1874:PSB

[5381] Anonymous. The philosophy of the sand blast. Scientific American, 30(6):81, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
81c.pdf.

Anonymous:1874:SPIf

[5382] Anonymous. Scientific and practical information. Scientific American, 30(6):81, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
81d.pdf.

Anonymous:1874:PCb

[5383] Anonymous. Professor Coffin. Scientific American, 30(6):82, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-82.pdf.

Anonymous:1874:UCS

[5384] Anonymous. Utilization of copper scraps in electro plating. Scientific American, 30(6):82, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
82a.pdf.

Anonymous:1874:NET

[5385] Anonymous. New English thrashing machines. Scientific American, 30(6):83–84, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
83.pdf.

Anonymous:1874:ANb

[5386] Anonymous. Astronomical notes. Scientific American, 30(6):84, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-84a.pdf.

Anonymous:1874:PAD

[5387] Anonymous. Professor Agassiz on the Darwinian theory — interesting facsimile letter from the great naturalist. Scientific American, 30(6):85, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
85.pdf.

Anonymous:1874:FWC

[5388] Anonymous. Fire and waterproof chimney guard. Scientific American, 30(6):86, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
86.pdf.

Anonymous:1874:PCc

[5389] Anonymous. Physician’s cane. Scientific American, 30(6):86, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-86a.pdf.

Anonymous:1874:TTa

[5390] Anonymous. The Toselli torpedo. Scientific American, 30(6):86, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-86b.pdf.

Anonymous:1874:PBH

[5391] Anonymous. Patent barrel head. Scientific American, 30(6):86, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-86c.pdf.

Anonymous:1874:PPA

[5392] Anonymous. Professor Proctor’s astronomical’ lectures. Scientific American, 30(6):87–88, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
87.pdf.

Anonymous:1874:PODa

[5393] Anonymous. Patent Office decisions. Scientific American, 30 (6):88, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
88.pdf.

Anonymous:1874:DCf

[5394] Anonymous. Decisions of the courts. Scientific American, 30 (6):88, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
88a.pdf.

Anonymous:1874:RAFf

[5395] Anonymous. Recent American and foreign patents. Scientific American, 30(6):88–89, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
88c.pdf.

Anonymous:1874:ACf

[5396] Anonymous. Answers correspondents. Scientific American, 30 (6):90–91, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
90a.pdf.

Anonymous:1874:CRf

[5397] Anonymous. Communications received. Scientific American, 30 (6):91, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n6/pdf/scientificamerican02071874-
91.pdf.

Anonymous:1874:CPe

[5398] Anonymous. Canadian patents. Scientific American, 30(6):92, February 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n6/pdf/scientificamerican02071874-92.pdf.

Anonymous:1874:Ch

[5399] Anonymous. Correspondence. Scientific American, 30(7):100, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n7/pdf/scientificamerican02141874-100.pdf.

Anonymous:1874:BNBf

[5400] Anonymous. Books: New books and publications. Scientific American, 30(7):104, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
104b.pdf.

Anonymous:1874:BPg

[5401] Anonymous. Business and personal. Scientific American, 30(7): 106, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
106.pdf.

Anonymous:1874:IIg

[5402] Anonymous. Index of inventions. Scientific American, 30(7): 107–108, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
107a.pdf.

Anonymous:1874:JYF

[5403] Anonymous. James Young founder of the paraffin oil industry. Scientific American, 30(7):95–96, February 14, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n7/pdf/scientificamerican02141874-95.pdf.

Anonymous:1874:VNI

[5404] Anonymous. Velocity of nervous impulses. Scientific American, 30(7):96, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
96.pdf.

Anonymous:1874:APA

[5405] Anonymous. The autopsy of Professor Agassiz. Scientific American, 30(7):96, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
96a.pdf.

Anonymous:1874:MCP

[5406] Anonymous. Microscopic crystals in plants. Scientific American, 30(7):96, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
96b.pdf.

Anonymous:1874:FL

[5407] Anonymous. The fireless locomotive. Scientific American, 30 (7):96–97, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
96c.pdf.

Anonymous:1874:OD

[5408] Anonymous. The origin of the diamond. Scientific American, 30(7):97, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
97.pdf.

Anonymous:1874:DDL

[5409] Anonymous. The death of Dr. Livingstone. Scientific American, 30(7):97, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
97a.pdf.

Anonymous:1874:SBE

[5410] Anonymous. Steam boiler explosions. — the work accomplished by the United States Commission. Scientific American, 30(7):97, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n7/pdf/scientificamerican02141874-97b.pdf.

Anonymous:1874:AFB

[5411] Anonymous. Allan’s floating batteries and saloons. Scientific American, 30(7):98, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
98.pdf.

Anonymous:1874:CGV

[5412] Anonymous. Captain Galton’s ventilating fireplace. Scientific American, 30(7):99, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
99.pdf.

Anonymous:1874:HGC

[5413] Anonymous. How greenbacks are cancelled. Scientific American, 30(7):99, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
99a.pdf.

Anonymous:1874:IWSa

[5414] Anonymous. Improved wood screw. Scientific American, 30(7): 99, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
99b.pdf.

Anonymous:1874:DSP

[5415] Anonymous. Drawing and sketching — practical hints and recipes. Scientific American, 30(7):99–100, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
99c.pdf.

Anonymous:1874:SPIg

[5416] Anonymous. Scientific and practical information. Scientific American, 30(7):101, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
101.pdf.

Anonymous:1874:AO

[5417] Anonymous. About ourselves. Scientific American, 30(7):101, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n7/pdf/scientificamerican02141874-101a.pdf.

Anonymous:1874:IGM

[5418] Anonymous. Improved gaining machine. Scientific American, 30 (7):102, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
102.pdf.

Anonymous:1874:MIA

[5419] Anonymous. Monson’s improved Auger. Scientific American, 30 (7):102, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
102a.pdf.

Anonymous:1874:BHV

[5420] Anonymous. Barker’s heating and ventilating apparatus. Scientific American, 30(7):102, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
102b.pdf.

Anonymous:1874:AFT

[5421] Anonymous. The Australian fever tree. Scientific American, 30 (7):103, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
103.pdf.

Anonymous:1874:IWB

[5422] Anonymous. Improved wash boiler. Scientific American, 30(7): 103, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
103a.pdf.

Anonymous:1874:SPW

[5423] Anonymous. Self-pumping well-boring drill. Scientific American, 30 (7):103–104, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
103b.pdf.

Anonymous:1874:HSIa

[5424] Anonymous. How shall I introduce my invention? Scientific American, 30(7):104, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
104.pdf.

Anonymous:1874:DCg

[5425] Anonymous. Decisions of the courts. Scientific American, 30 (7):104, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
104a.pdf.

Anonymous:1874:RAFg

[5426] Anonymous. Recent American and foreign patents. Scientific American, 30(7):104–105, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
104c.pdf.

Anonymous:1874:ACg

[5427] Anonymous. Answers correspondents. Scientific American, 30(7): 106–107, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
106a.pdf.

Anonymous:1874:CRg

[5428] Anonymous. Communications received. Scientific American, 30 (7):107, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n7/pdf/scientificamerican02141874-
107.pdf.

Anonymous:1874:CPf

[5429] Anonymous. Canadian patents. Scientific American, 30(7):108, February 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n7/pdf/scientificamerican02141874-108.pdf.

Anonymous:1874:Ci

[5430] Anonymous. Correspondence. Scientific American, 30(8):116– 117, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
116.pdf.

Anonymous:1874:BNBg

[5431] Anonymous. Books: New books and publications. Scientific American, 30(8):120, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
120b.pdf.

Anonymous:1874:BPh

[5432] Anonymous. Business and personal. Scientific American, 30(8): 122, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
122.pdf.

Anonymous:1874:IIh

[5433] Anonymous. Index of inventions. Scientific American, 30(8): 123–124, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
123a.pdf.

Anonymous:1874:BRT

[5434] Anonymous. The battering ram as a tunnel borer. Scientific American, 30(8):111, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
111.pdf.

Anonymous:1874:SIT

[5435] Anonymous. The state of the iron trade. Scientific American, 30(8):112, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
112.pdf.

Anonymous:1874:ONC

[5436] Anonymous. Ozone — a new and correct method of supply. Scientific American, 30(8):112, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
112a.pdf.

Anonymous:1874:EBK

[5437] Anonymous. Education and book knowledge. Scientific American, 30(8):112, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
112b.pdf.

Anonymous:1874:IOM

[5438] Anonymous. The iron ores of Missouri. Scientific American, 30 (8):112–113, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
112c.pdf.

Anonymous:1874:MPC

[5439] Anonymous. Modern progress of chemical industries in Europe. Scientific American, 30(8):113, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
113.pdf.

Anonymous:1874:ET

[5440] Anonymous. The English telegraphs. Scientific American, 30 (8):114, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
114.pdf.

Anonymous:1874:HB

[5441] Anonymous. The Heberlein brake. Scientific American, 30(8): 114, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
114a.pdf.

Anonymous:1874:SAWa

[5442] Anonymous. South African wonders — explosive diamonds and turtles with teeth. Scientific American, 30(8):114, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n8/pdf/scientificamerican02211874-114b.pdf.

Anonymous:1874:STb

[5443] Anonymous. The Siamese twins. Scientific American, 30(8):115, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n8/pdf/scientificamerican02211874-115.pdf.

Anonymous:1874:PRP

[5444] Anonymous. Professor R. A. Proctor. Scientific American, 30 (8):115–116, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
115a.pdf.

Anonymous:1874:SPIh

[5445] Anonymous. Scientific and practical information. Scientific American, 30(8):117, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
117.pdf.

Anonymous:1874:RP

[5446] Anonymous. The ribbon post. Scientific American, 30(8):117, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n8/pdf/scientificamerican02211874-117a.pdf.

Anonymous:1874:ICWa

[5447] Anonymous. Improved cotton worm destroyer. Scientific American, 30(8):118, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
118.pdf.

Anonymous:1874:IBS

[5448] Anonymous. Improved bearing for shafting. Scientific American, 30(8):118, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
118a.pdf.

Anonymous:1874:CLB

[5449] Anonymous. Combined lounge and bath tub. Scientific American, 30(8):118, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
118b.pdf.

Anonymous:1874:AMB

[5450] Anonymous. The artificial manufacture of butter. Scientific American, 30(8):119–120, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
119.pdf.

Anonymous:1874:IAc

[5451] Anonymous. Importance of advertising. Scientific American, 30 (8):120, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
120.pdf.

Anonymous:1874:DCh

[5452] Anonymous. Decisions of the courts. Scientific American, 30 (8):120, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
120a.pdf.

Anonymous:1874:RAFh

[5453] Anonymous. Recent American and foreign patents. Scientific American, 30(8):120–121, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
120c.pdf.

Anonymous:1874:ACh

[5454] Anonymous. Answers correspondents. Scientific American, 30(8): 122–123, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
122a.pdf.

Anonymous:1874:CRh

[5455] Anonymous. Communications received. Scientific American, 30 (8):123, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n8/pdf/scientificamerican02211874-
123.pdf.

Anonymous:1874:CPg

[5456] Anonymous. Canadian patents. Scientific American, 30(8):124, February 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n8/pdf/scientificamerican02211874-124.pdf.

Anonymous:1874:Cj

[5457] Anonymous. Correspondence. Scientific American, 30(9):132– 133, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
132.pdf.

Anonymous:1874:BPi

[5458] Anonymous. Business and personal. Scientific American, 30(9): 138, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
138.pdf.

Anonymous:1874:IIi

[5459] Anonymous. Index of inventions. Scientific American, 30(9):139, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n9/pdf/scientificamerican02281874-139a.pdf.

Anonymous:1874:CPR

[5460] Anonymous. Casement’s pressure regulator for fluids. Scientific American, 30(9):127, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
127.pdf.

Anonymous:1874:PMC

[5461] Anonymous. Patent monopolies in Congress. Scientific American, 30(9):128, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
128.pdf.

Anonymous:1874:SP

[5462] Anonymous. Slippery pavements. Scientific American, 30(9): 128, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
128a.pdf.

Anonymous:1874:HTD

[5463] Anonymous. Herschel, Tyndall, and Draper on the Sun ray. Scientific American, 30(9):128–129, February 28, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n9/pdf/scientificamerican02281874-128b.pdf.

Anonymous:1874:WSa

[5464] Anonymous. Whitworth steel. Scientific American, 30(9):129, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n9/pdf/scientificamerican02281874-129.pdf.

Anonymous:1874:PGD

[5465] Anonymous. A paper and glass debate. Scientific American, 30 (9):129, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
129a.pdf.

Anonymous:1874:SPIi

[5466] Anonymous. Scientific and practical information. Scientific American, 30(9):129, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
129b.pdf.

Anonymous:1874:NSa

[5467] Anonymous. To new subscribers. Scientific American, 30(9):129, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n9/pdf/scientificamerican02281874-129c.pdf.

Anonymous:1874:IMTa

[5468] Anonymous. Improved movable teeth for saws. Scientific American, 30(9):130, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
130.pdf.

Anonymous:1874:HIP

[5469] Anonymous. Hilf’s iron permanent way. Scientific American, 30 (9):130, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
130a.pdf.

Anonymous:1874:PCTa

[5470] Anonymous. Pipe cutting and threading machine. Scientific American, 30(9):131, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
131.pdf.

Anonymous:1874:ITB

[5471] Anonymous. Improved trench braces. Scientific American, 30 (9):131, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
131a.pdf.

Leeds:1874:PM

[5472] Albert R. Leeds. On the purification of mercury. Scientific American, 30(9):131–132, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
131b.pdf.

Anonymous:1874:CIM

[5473] Anonymous. Casement’s improved method of burning fuel. Scientific American, 30(9):135–136, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
135.pdf.

Anonymous:1874:HSIb

[5474] Anonymous. How shall I introduce my invention? Scientific American, 30(9):136, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
136.pdf.

Anonymous:1874:DCi

[5475] Anonymous. Decisions of the courts. Scientific American, 30 (9):136, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
136a.pdf.

Anonymous:1874:RAFi

[5476] Anonymous. Recent American and foreign patents. Scientific American, 30(9):137, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
137.pdf.

Anonymous:1874:ACi

[5477] Anonymous. Answers correspondents. Scientific American, 30(9): 138–139, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
138a.pdf.

Anonymous:1874:CRi

[5478] Anonymous. Communications received. Scientific American, 30 (9):139, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n9/pdf/scientificamerican02281874-
139.pdf.

Anonymous:1874:CPh

[5479] Anonymous. Canadian patents. Scientific American, 30(9):140, February 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n9/pdf/scientificamerican02281874-140.pdf.

Anonymous:1874:Ck

[5480] Anonymous. Correspondence. Scientific American, 30(10):148–149, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-148.pdf.

Anonymous:1874:BNBh 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
152b.pdf.

Anonymous:1874:BPj

[5482] Anonymous. Business and personal. Scientific American, 30(10):154, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-154.pdf.

Anonymous:1874:IIj

[5483] Anonymous. Index of inventions. Scientific American, 30(10):155, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-155a.pdf.

Anonymous:1874:IGCa

[5484] Anonymous. Improved grinder for castings. Scientific American, 30(10):143, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
143.pdf.

Anonymous:1874:PLC

[5485] Anonymous. Proposed law for castings. Scientific American, 30(10):144, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-144.pdf.

Anonymous:1874:AST

[5486] Anonymous. The autopsy on the Siamese twins. Scientific American, 30(10):144, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
144a.pdf.

Anonymous:1874:IUB

[5487] Anonymous. The industrial uses of bisulphide of carbon. Scientific American, 30(10):144, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
144b.pdf.

Anonymous:1874:TP

[5488] Anonymous. Telegraphic progress in 1873. Scientific American, 30(10):144–145, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
144c.pdf.

Anonymous:1874:RTa

[5489] Anonymous. Robert L. Thurston. Scientific American, 30(10):145, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-145.pdf.

Anonymous:1874:SCc

[5490] Anonymous. Stearin candles. Scientific American, 30(10):145, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-145a.pdf.

Anonymous:1874:SPIj

[5491] Anonymous. Scientific and practical information. Scientific American, 30(10):145, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
145b.pdf.

Anonymous:1874:NSb

[5492] Anonymous. To new subscribers. Scientific American, 30(10):145, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-145c.pdf.

Anonymous:1874:ARB

[5493] Anonymous. The Albany Railroad Bridge. Scientific American, 30(10):147, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
147.pdf.

Anonymous:1874:IFG

[5494] Anonymous. Improved fish grappling spear. Scientific American, 30(10):147, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
147a.pdf.

Anonymous:1874:ANc

[5495] Anonymous. Astronomical notes. Scientific American, 30(10):147, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-147b.pdf.

Anonymous:1874:ISG

[5496] Anonymous. Improved saw gummer. Scientific American, 30(10):150, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-150.pdf.

Anonymous:1874:IEV

[5497] Anonymous. Improved electric and vapor chair. Scientific American, 30(10):150, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
150a.pdf.

Anonymous:1874:VCD

[5498] Anonymous. The Victor churn dasher. Scientific American, 30(10):150, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-150b.pdf.

Anonymous:1874:VRH

[5499] Anonymous. The Victoria Regia House at Chatsworth, Eng. Scientific American, 30(10):151, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
151.pdf.

Anonymous:1874:SFP

[5500] Anonymous. Sir Francis Pettit Smith. Scientific American, 30 (10):151–152, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
151a.pdf.

Anonymous:1874:IAd

[5501] Anonymous. Importance of advertising. Scientific American, 30 (10):152, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
152.pdf.

Anonymous:1874:PODb

[5502] Anonymous. Patent Office decisions. Scientific American, 30(10):152, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-152a.pdf.

Anonymous:1874:RAFj

[5503] Anonymous. Recent American and foreign patents. Scientific American, 30(10):152–153, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
152c.pdf.

Anonymous:1874:ACj

[5504] Anonymous. Answers to correspondents. Scientific American, 30 (10):154–155, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
154a.pdf.

Anonymous:1874:CRj

[5505] Anonymous. Communications received. Scientific American, 30 (10):155, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n10/pdf/scientificamerican03071874-
155.pdf.

Anonymous:1874:CPi

[5506] Anonymous. Canadian patents. Scientific American, 30(10):155–156, March 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n10/pdf/scientificamerican03071874-155b.pdf.

Anonymous:1874:Cl

[5507] Anonymous. Correspondence. Scientific American, 30(11):164–165, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-164.pdf.

Anonymous:1874:BNBi

[5508] Anonymous. Books: New books and publications. Scientific American, 30(11):168, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
168.pdf.

Anonymous:1874:BPk

[5509] Anonymous. Business and personal. Scientific American, 30(11):170, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-170.pdf.

Anonymous:1874:IIk

[5510] Anonymous. Index of inventions. Scientific American, 30(11):171–172, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-171a.pdf.

Anonymous:1874:NGT

[5511] Anonymous. New and gigantic telescope. Scientific American, 30 (11):159–160, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
159.pdf.

Anonymous:1874:LT

[5512] Anonymous. Lake Titicaca. Scientific American, 30(11):160, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-160.pdf.

Anonymous:1874:SPSa

[5513] Anonymous. Swindling patent sellers. Scientific American, 30(11):160, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-160a.pdf.

Anonymous:1874:FNT

[5514] Anonymous. Failure of the new telegraph line between Europe and America. Scientific American, 30(11):160, March 14, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n11/pdf/scientificamerican03141874-160b.pdf.

Anonymous:1874:RRE

[5515] Anonymous. Relative rights of employers and workmen in respect to inventions. Scientific American, 30(11):160, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-160c.pdf.

Anonymous:1874:ODG

[5516] Anonymous. The oil deposits of the great west. Scientific American, 30(11):160–161, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
160d.pdf.

Anonymous:1874:EUC

[5517] Anonymous. Enforcement of unauthorized caveat rules. Scientific American, 30(11):161, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
161.pdf.

Anonymous:1874:MSa

[5518] Anonymous. Mount Sinai. Scientific American, 30(11):161, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-161a.pdf.

Anonymous:1874:PW

[5519] Anonymous. Pleasant words. Scientific American, 30(11):161, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-161b.pdf.

Anonymous:1874:SPIk

[5520] Anonymous. Scientific and practical information. Scientific American, 30(11):161, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
161c.pdf.

Anonymous:1874:MSSa

[5521] Anonymous. On the motion and sensation of sound. Scientific American, 30(11):162–163, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
162.pdf.

Anonymous:1874:GHR

[5522] Anonymous. Gilberds & Harris’ return butter and oyster pail. Scientific American, 30(11):163–164, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
163.pdf.

Anonymous:1874:OCW

[5523] Anonymous. Oakley’s currant washer. Scientific American, 30(11):166, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-166.pdf.

Anonymous:1874:NTH

[5524] Anonymous. A new tool holder. Scientific American, 30(11):166, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-166a.pdf.

Anonymous:1874:NWM

[5525] Anonymous. A new water meter. Scientific American, 30(11):167, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-167.pdf.

Anonymous:1874:PODc

[5526] Anonymous. Patent Office decisions. Scientific American, 30(11):168, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-168a.pdf.

Anonymous:1874:DCj

[5527] Anonymous. Decisions of the courts. Scientific American, 30(11):168, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n11/pdf/scientificamerican03141874-168b.pdf.

Anonymous:1874:IAe

[5528] Anonymous. Importance of advertising. Scientific American, 30 (11):168, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
168c.pdf.

Anonymous:1874:RAFk

[5529] Anonymous. Recent American and foreign patents. Scientific American, 30(11):168–169, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
168d.pdf.

Anonymous:1874:ACk

[5530] Anonymous. Answer to correspondents. Scientific American, 30 (11):170–171, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
170a.pdf.

Anonymous:1874:CRk

[5531] Anonymous. Communications received. Scientific American, 30 (11):171, March 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n11/pdf/scientificamerican03141874-
171.pdf.

Anonymous:1874:Cm

[5532] Anonymous. Correspondence. Scientific American, 30(12):181–182, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-181a.pdf.

Anonymous:1874:BNBj 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
184a.pdf.

Anonymous:1874:BPl

[5534] Anonymous. Business and personal. Scientific American, 30(12):186, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-186.pdf.

Anonymous:1874:IIl

[5535] Anonymous. Index of inventions. Scientific American, 30(12):187, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-187a.pdf.

Anonymous:1874:ISMa

[5536] Anonymous. Improved slotting machine and slotting and planing tool. Scientific American, 30(12):175, March 21, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n12/pdf/scientificamerican03211874-175.pdf.

Anonymous:1874:RRA

[5537] Anonymous. A remarkable report about remarkable pumping engines. Scientific American, 30(12):176, March 21, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n12/pdf/scientificamerican03211874-176.pdf.

Anonymous:1874:CPDa

[5538] Anonymous. A Congressional patent discussion. Scientific American, 30(12):176, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
176a.pdf.

Anonymous:1874:DNE

[5539] Anonymous. Dangers of nursery explosives. Scientific American, 30(12):176, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
176b.pdf.

Anonymous:1874:ASa

[5540] Anonymous. American soda. Scientific American, 30(12):176, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-176c.pdf.

Anonymous:1874:GLF

[5541] Anonymous. The great Lava flood of Oregon. Scientific American, 30(12):176–177, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
176d.pdf.

Anonymous:1874:SRP

[5542] Anonymous. Scientific results of the Polaris Expedition. Scientific American, 30(12):177, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
177.pdf.

Anonymous:1874:NF

[5543] Anonymous. Nerve force. Scientific American, 30(12):177, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-177a.pdf.

Anonymous:1874:ADM

[5544] Anonymous. Acoustic darkness and mental light. Scientific American, 30(12):177, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
177b.pdf.

Anonymous:1874:ISCa

[5545] Anonymous. Improved self-closing faucet attachment. Scientific American, 30(12):178, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
178.pdf.

Anonymous:1874:WBF

[5546] Anonymous. Watson’s butter former. Scientific American, 30(12):178, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-178a.pdf.

Anonymous:1874:FPC

[5547] Anonymous. Font pen with a capillary feeder. Scientific American, 30(12):178, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
178b.pdf.

Anonymous:1874:SIN

[5548] Anonymous. Schofield’s improved needle threader. Scientific American, 30(12):178, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
178c.pdf.

Anonymous:1874:NDF

[5549] Anonymous. The Newfoundland devil fish. Scientific American, 30(12):179, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
179.pdf.

Anonymous:1874:BRH

[5550] Anonymous. Barnum’s removable horseshoe calk. Scientific American, 30(12):179–181, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
179a.pdf.

Anonymous:1874:CN

[5551] Anonymous. Chemical notes. Scientific American, 30(12):181, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-181.pdf.

Anonymous:1874:IATa

[5552] Anonymous. Improved automatic tool grinding machine. Scientific American, 30(12):182, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
182.pdf.

Anonymous:1874:ICBa

[5553] Anonymous. Improved circular and band saw set. Scientific American, 30(12):182, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
182a.pdf.

Anonymous:1874:HTG

[5554] Anonymous. House top gardening. Scientific American, 30(12):183–184, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-183.pdf.

Anonymous:1874:NSd

[5555] Anonymous. To new subscribers. Scientific American, 30(12):184, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-184.pdf.

Anonymous:1874:PODd

[5556] Anonymous. Patent Office decisions. Scientific American, 30(12):184, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-184b.pdf.

Anonymous:1874:DCk

[5557] Anonymous. Decisions of the courts. Scientific American, 30(12):184, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-184c.pdf.

Anonymous:1874:IAf

[5558] Anonymous. Importance of advertising. Scientific American, 30 (12):184, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
184d.pdf.

Anonymous:1874:RAFl

[5559] Anonymous. Recent American and foreign patents. Scientific American, 30(12):184–185, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
184e.pdf.

Anonymous:1874:ACl

[5560] Anonymous. Answers to correspondents. Scientific American, 30 (12):186–187, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
186a.pdf.

Anonymous:1874:CRl

[5561] Anonymous. Communications received. Scientific American, 30 (12):187, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n12/pdf/scientificamerican03211874-
187.pdf.

Anonymous:1874:CPj

[5562] Anonymous. Canadian patents. Scientific American, 30(12):187–188, March 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n12/pdf/scientificamerican03211874-187b.pdf.

Anonymous:1874:Co

[5563] Anonymous. Correspondence. Scientific American, 30(14):212–213, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-212.pdf.

Anonymous:1874:BPm

[5564] Anonymous. Business and personal. Scientific American, 30(14):218, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-218.pdf.

Anonymous:1874:IIm

[5565] Anonymous. Index of inventions. Scientific American, 30(14):219–220, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-219a.pdf.

Anonymous:1874:PNPa

[5566] Anonymous. Patent notes: Patents. Scientific American, 30(14):220, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-220a.pdf.

Anonymous:1874:FEA

[5567] Anonymous. Fire extinguishing apparatus for buildings. Scientific American, 30(14):207, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
207.pdf.

Anonymous:1874:SAP

[5568] Anonymous. The Scientific American as a preventive of boiler explosions — value of its correspondence. Scientific American, 30(14):208, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-208.pdf.

Anonymous:1874:BLE

[5569] Anonymous. The Brotherhood of Locomotive Engineers. Scientific American, 30(14):208, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
208a.pdf.

Anonymous:1874:MMa

[5570] Anonymous. The marvels of mechanics. Scientific American, 30 (14):208, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
208b.pdf.

Anonymous:1874:MSb

[5571] Anonymous. Metrological science. Scientific American, 30(14):208, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-208c.pdf.

Anonymous:1874:DRT

[5572] Anonymous. The Detroit River Tunnel. Scientific American, 30 (14):208–209, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
208d.pdf.

Anonymous:1874:HS

[5573] Anonymous. A Hill of sulphur. Scientific American, 30(14):209, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-209.pdf.

Anonymous:1874:LCP

[5574] Anonymous. The launch of the city of peking. Scientific American, 30(14):209, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
209a.pdf.

Thurston:1874:CME

[5575] R. H. Thurston. The character of metals as exhibited by their fracture. Scientific American, 30(14):209, April 4, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n14/pdf/scientificamerican04041874-209b.pdf.

Anonymous:1874:NSe

[5576] Anonymous. To new subscribers. Scientific American, 30(14):209, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-209c.pdf.

Anonymous:1874:PMa

[5577] Anonymous. The Palæotherium Magnum. Scientific American, 30(14):210, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
210.pdf.

Anonymous:1874:IATb

[5578] Anonymous. Improved adjustable threshold. Scientific American, 30(14):210, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
210a.pdf.

Anonymous:1874:PBa

[5579] Anonymous. Patent bottle. Scientific American, 30(14):210, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-210b.pdf.

Anonymous:1874:IASa

[5580] Anonymous. Improved adjustable scaffold. Scientific American, 30(14):211, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
211.pdf.

Anonymous:1874:NME

[5581] Anonymous. A new magneto-electrical instrument. Scientific American, 30(14):211, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
211a.pdf.

Anonymous:1874:ISSa

[5582] Anonymous. Improved stove for smoke houses. Scientific American, 30(14):214, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
214.pdf.

Anonymous:1874:ITW

[5583] Anonymous. Improved turbine wheel. Scientific American, 30(14):214, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-214a.pdf.

Anonymous:1874:NBPa

[5584] Anonymous. A new bird of paradise. Scientific American, 30(14):215, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-215.pdf.

Anonymous:1874:ANd

[5585] Anonymous. Astronomical notes. Scientific American, 30(14):216, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-216.pdf.

Anonymous:1874:IAg

[5586] Anonymous. Importance of advertising. Scientific American, 30 (14):216, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
216a.pdf.

Anonymous:1874:DCl

[5587] Anonymous. Decisions of the courts. Scientific American, 30(14):216, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-216b.pdf.

Anonymous:1874:RAFm

[5588] Anonymous. Recent American and foreign patents. Scientific American, 30(14):217, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
217.pdf.

Anonymous:1874:ACm

[5589] Anonymous. Answers to correspondents. Scientific American, 30 (14):218–219, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
218a.pdf.

Anonymous:1874:CRm

[5590] Anonymous. Communications received. Scientific American, 30 (14):219, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n14/pdf/scientificamerican04041874-
219.pdf.

Anonymous:1874:CPk

[5591] Anonymous. Canadian patents. Scientific American, 30(14):220, April 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n14/pdf/scientificamerican04041874-220.pdf.

Anonymous:1874:Cq

[5592] Anonymous. Correspondence. Scientific American, 30(16):245, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-245.pdf.

Anonymous:1874:BNBk

[5593] Anonymous. Books: New books and publications. Scientific American, 30(16):248, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
248c.pdf.

Anonymous:1874:BPn

[5594] Anonymous. Business and personal. Scientific American, 30(16):250, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-250.pdf.

Anonymous:1874:IIn

[5595] Anonymous. Index of inventions. Scientific American, 30(16):251–252, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-251a.pdf.

Anonymous:1874:PNPb

[5596] Anonymous. Patent notes: Patents. Scientific American, 30(16):252, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-252a.pdf.

Anonymous:1874:MSSb

[5597] Anonymous. The Mitchell safety steam generator. Scientific American, 30(16):239, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
239.pdf.

Anonymous:1874:DIP

[5598] Anonymous. Decision of an important patent litigation. Scientific American, 30(16):240, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
240.pdf.

Anonymous:1874:AM

[5599] Anonymous. The Agassiz Memorial. Scientific American, 30(16):240, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-240a.pdf.

Anonymous:1874:GH

[5600] Anonymous. The genesis of the horse. Scientific American, 30(16):240, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-240b.pdf.

Anonymous:1874:SMA

[5601] Anonymous. The suppressed member again. Scientific American, 30(16):240–241, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
240c.pdf.

Anonymous:1874:HGD

[5602] Anonymous. How a great discovery was made. Scientific American, 30(16):241, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
241.pdf.

Anonymous:1874:OSP

[5603] Anonymous. The origination of screw propulsion. Scientific American, 30(16):241, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
241a.pdf.

Anonymous:1874:AW

[5604] Anonymous. Artesian wells. Scientific American, 30(16):241, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-241b.pdf.

Anonymous:1874:SPIl

[5605] Anonymous. Scientific and practical information. Scientific American, 30(16):241, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
241c.pdf.

Anonymous:1874:NED

[5606] Anonymous. New electric disengaging gear for knitting machines. Scientific American, 30(16):242, April 18, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n16/pdf/scientificamerican04181874-242.pdf.

Anonymous:1874:HICa

[5607] Anonymous. Horton’s improved chuck jaw. Scientific American, 30(16):242, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
242a.pdf.

Anonymous:1874:AIM

[5608] Anonymous. The ammonia ice-making process. Scientific American, 30(16):243, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
243.pdf.

Anonymous:1874:ISPa

[5609] Anonymous. Improved steam pump. Scientific American, 30(16):243, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-243a.pdf.

Anonymous:1874:DBS

[5610] Anonymous. Dr. Brown Sequard on nerve force. Scientific American, 30(16):243–244, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
243b.pdf.

Anonymous:1874:LAb

[5611] Anonymous. In the laboratory with Agassiz. Scientific American, 30(16):244–245, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
244.pdf.

Anonymous:1874:IPGa

[5612] Anonymous. Improved plane guide. Scientific American, 30(16):246, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-246.pdf.

Anonymous:1874:LII

[5613] Anonymous. Lucetti’s Italian ice cream freezer. Scientific American, 30(16):246, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
246a.pdf.

Anonymous:1874:MIC

[5614] Anonymous. Magnifying insect case. Scientific American, 30(16):246, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-246b.pdf.

Anonymous:1874:MSC

[5615] Anonymous. Matson’s safety combined whip and rein holder. Scientific American, 30(16):246, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
246c.pdf.

Anonymous:1874:CFH

[5616] Anonymous. Casting a five hundred tun anvil. Scientific American, 30(16):247, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
247.pdf.

Anonymous:1874:SCd

[5617] Anonymous. The solar counter. Scientific American, 30(16):247–248, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-247a.pdf.

Anonymous:1874:ANe

[5618] Anonymous. Astronomical notes. Scientific American, 30(16):248, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-248.pdf.

Anonymous:1874:DCm

[5619] Anonymous. Decisions of the courts. Scientific American, 30(16):248, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-248a.pdf.

Anonymous:1874:HSIc

[5620] Anonymous. How shall I introduce my invention? Scientific American, 30(16):248, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
248b.pdf.

Anonymous:1874:RAFn

[5621] Anonymous. Recent American and foreign patents. Scientific American, 30(16):248–249, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
248d.pdf.

Anonymous:1874:ACn

[5622] Anonymous. Answers to correspondents. Scientific American, 30 (16):250–251, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
250a.pdf.

Anonymous:1874:CRn

[5623] Anonymous. Communications received. Scientific American, 30 (16):251, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n16/pdf/scientificamerican04181874-
251.pdf.

Anonymous:1874:CPl

[5624] Anonymous. Canadian patents. Scientific American, 30(16):252, April 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n16/pdf/scientificamerican04181874-252.pdf.

Anonymous:1874:BNBl

[5625] Anonymous. Books: New books and publications. Scientific American, 30(17):264, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
264.pdf.

Anonymous:1874:BPo

[5626] Anonymous. Business and personal. Scientific American, 30(17):266, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-266.pdf.

Anonymous:1874:IIo

[5627] Anonymous. Index of inventions. Scientific American, 30(17):267–268, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-267a.pdf.

Anonymous:1874:PNPc

[5628] Anonymous. Patent notes: Patents. Scientific American, 30(17):268, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-268a.pdf.

Anonymous:1874:IPE

[5629] Anonymous. Improved portable elevator and conveyer. Scientific American, 30(17):255, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
255.pdf.

Anonymous:1874:RDP

[5630] Anonymous. Reclaiming the desert places. Scientific American, 30(17):256, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
256.pdf.

Anonymous:1874:NSF

[5631] Anonymous. A neglected source of food. Scientific American, 30(17):256, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
256a.pdf.

Anonymous:1874:ACLa

[5632] Anonymous. The astronomical conditions of life. Scientific American, 30(17):256–257, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
256b.pdf.

Anonymous:1874:ISMb

[5633] Anonymous. Improvements in sugar making. Scientific American, 30(17):257, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
257.pdf.

Anonymous:1874:PEE

[5634] Anonymous. Pumping engine economy. Scientific American, 30 (17):257–258, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
257a.pdf.

Anonymous:1874:STM

[5635] Anonymous. The separation of tar and the manufacture of sulphate of ammonia from waste products. Scientific American, 30(17):258–260, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-258.pdf.

Anonymous:1874:RBE

[5636] Anonymous. Recent boiler explosions. Scientific American, 30(17):260, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-260.pdf.

Anonymous:1874:SPIm

[5637] Anonymous. Scientific and practical information. Scientific American, 30(17):261, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
261.pdf.

Anonymous:1874:IGCb

[5638] Anonymous. Improved grate clearing attachment. Scientific American, 30(17):262, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
262.pdf.

Anonymous:1874:SSB

[5639] Anonymous. The shaker sash balance. Scientific American, 30(17):262, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-262a.pdf.

Anonymous:1874:MILa

[5640] Anonymous. Munson’s improved lightning rod. Scientific American, 30(17):262, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
262b.pdf.

Anonymous:1874:MAL

[5641] Anonymous. Modern American lighthouses. Scientific American, 30(17):263–264, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
263.pdf.

Anonymous:1874:IAh

[5642] Anonymous. Importance of advertising. Scientific American, 30 (17):264, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
264a.pdf.

Anonymous:1874:RAFo

[5643] Anonymous. Recent American and foreign patents. Scientific American, 30(17):264–265, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
264b.pdf.

Anonymous:1874:ACo

[5644] Anonymous. Answers to correspondents. Scientific American, 30 (17):266–267, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
266a.pdf.

Anonymous:1874:CRo

[5645] Anonymous. Communications received. Scientific American, 30 (17):267, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n17/pdf/scientificamerican04251874-
267.pdf.

Anonymous:1874:CPm

[5646] Anonymous. Canadian patents. Scientific American, 30(17):268, April 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n17/pdf/scientificamerican04251874-268.pdf.

Anonymous:1874:Ct

[5647] Anonymous. Correspondence. Scientific American, 30(18):276, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-276b.pdf.

Anonymous:1874:BPp

[5648] Anonymous. Business and personal. Scientific American, 30(18):282, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-282.pdf.

Anonymous:1874:IIp

[5649] Anonymous. Index of inventions. Scientific American, 30(18):283–284, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-283a.pdf.

Anonymous:1874:PNPd

[5650] Anonymous. Patent notes: Patents. Scientific American, 30(18):286, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-286.pdf.

Anonymous:1874:IG

[5651] Anonymous. Improved governor. Scientific American, 30(18):271–272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-271.pdf.

Anonymous:1874:EPS

[5652] Anonymous. The effective power of steam engines. Scientific American, 30(18):272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
272.pdf.

Anonymous:1874:CSb

[5653] Anonymous. Compartment ships. Scientific American, 30(18):272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-272a.pdf.

Anonymous:1874:TKS

[5654] Anonymous. A tree that keeps a standing army. Scientific American, 30(18):272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
272b.pdf.

Anonymous:1874:PUR

[5655] Anonymous. Progress of underground railways in London. Scientific American, 30(18):272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
272c.pdf.

Anonymous:1874:NC

[5656] Anonymous. A new comet. Scientific American, 30(18):272, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-272d.pdf.

Anonymous:1874:ANM

[5657] Anonymous. Another new motor. Scientific American, 30(18):273, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-273.pdf.

Anonymous:1874:OQ

[5658] Anonymous. The oremation question. Scientific American, 30 (18):273–274, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
273a.pdf.

Anonymous:1874:IPO

[5659] Anonymous. Improved portable oiler. Scientific American, 30(18):274, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-274.pdf.

Anonymous:1874:ICSa

[5660] Anonymous. Improved combination scissors. Scientific American, 30(18):274, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
274a.pdf.

Anonymous:1874:ILP

[5661] Anonymous. An instantaneous life preserving raft. Scientific American, 30(18):275–276, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
275.pdf.

Anonymous:1874:JB

[5662] Anonymous. James Bogardus. Scientific American, 30(18):276, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-276.pdf.

Anonymous:1874:PPCb

[5663] Anonymous. Prevention in place of cure. Scientific American, 30(18):276, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
276a.pdf.

Thurston:1874:SDW

[5664] R. H. Thurston. Strain diagrams — what they are and what they tell us. Scientific American, 30(18):277, May 2, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n18/pdf/scientificamerican05021874-277.pdf.

Anonymous:1874:SPIn

[5665] Anonymous. Scientific and practical information. Scientific American, 30(18):277, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
277a.pdf.

Anonymous:1874:IMMa

[5666] Anonymous. Improved mitering machine. Scientific American, 30(18):278, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
278.pdf.

Anonymous:1874:SBS

[5667] Anonymous. Safety Bell signal for railways. Scientific American, 30(18):278, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
278a.pdf.

Anonymous:1874:CRDa

[5668] Anonymous. Christie’s railroad detachable claw bar. Scientific American, 30(18):278, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
278b.pdf.

Anonymous:1874:Pa

[5669] Anonymous. The pyrophone. Scientific American, 30(18):279, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-279.pdf.

Anonymous:1874:SSF

[5670] Anonymous. The steam siren or fog Horn. Scientific American, 30(18):279, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
279a.pdf.

Prescott:1874:ET

[5671] B. Prescott. The English telegraphs. Scientific American, 30(18):280, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-280.pdf.

Anonymous:1874:RAFp

[5672] Anonymous. Recent American and foreign patents. Scientific American, 30(18):280–281, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
280a.pdf.

Anonymous:1874:ACp

[5673] Anonymous. Answers to correspondents. Scientific American, 30 (18):282–283, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
282a.pdf.

Anonymous:1874:CRp

[5674] Anonymous. Communications received. Scientific American, 30 (18):283, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n18/pdf/scientificamerican05021874-
283.pdf.

Anonymous:1874:CPn

[5675] Anonymous. Canadian patents. Scientific American, 30(18):284, May 2, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n18/pdf/scientificamerican05021874-284.pdf.

Anonymous:1874:Cu

[5676] Anonymous. Correspondence. Scientific American, 30(19):291, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-291.pdf.

Anonymous:1874:BNBm

[5677] Anonymous. Books: New books and publications. Scientific American, 30(19):297, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
297a.pdf.

Anonymous:1874:BPq

[5678] Anonymous. Business and personal. Scientific American, 30(19):298, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-298.pdf.

Anonymous:1874:IIq

[5679] Anonymous. Index of inventions. Scientific American, 30(19):299–300, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-299a.pdf.

Anonymous:1874:PNPe

[5680] Anonymous. Patent notes: Patents. Scientific American, 30(19):300, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-300a.pdf.

Anonymous:1874:NBCa

[5681] Anonymous. A novelty in bolt cutters. Scientific American, 30(19):287, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-287.pdf.

Anonymous:1874:CSc

[5682] Anonymous. The Centennial subscription. Scientific American, 30(19):288, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
288.pdf.

Anonymous:1874:FB

[5683] Anonymous. The famine in Bengal. Scientific American, 30(19):288, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-288a.pdf.

Leeds:1874:EVB

[5684] Albert R. Leeds. Eremacausis versus burial and cremation. Scientific American, 30(19):288–289, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
288b.pdf.

Anonymous:1874:EPTa

[5685] Anonymous. The effective power of turbine water wheels. Scientific American, 30(19):289, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
289.pdf.

Anonymous:1874:AAb

[5686] Anonymous. Army ants. Scientific American, 30(19):289, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-289a.pdf.

Anonymous:1874:SPIo

[5687] Anonymous. Scientific and practical information. Scientific American, 30(19):289, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
289b.pdf.

Anonymous:1874:VMB

[5688] Anonymous. The vertical multiplier boring machine. Scientific American, 30(19):290, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
290.pdf.

Anonymous:1874:IFE

[5689] Anonymous. Improved fuel economizer. Scientific American, 30 (19):290–291, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
290a.pdf.

Anonymous:1874:ANf

[5690] Anonymous. Astronomical notes. Scientific American, 30(19):292, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-292.pdf.

Anonymous:1874:MNA

[5691] Anonymous. The meeting of the National Academy of Science. Scientific American, 30(19):293, May 9, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n19/pdf/scientificamerican05091874-293.pdf.

Anonymous:1874:WTW

[5692] Anonymous. The Whitmore turbine wheel. Scientific American, 30(19):294, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
294.pdf.

Anonymous:1874:GIW

[5693] Anonymous. Gardner’s improved wheel. Scientific American, 30 (19):294, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
294a.pdf.

Anonymous:1874:CFa

[5694] Anonymous. Cremation furnace. Scientific American, 30(19):295, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-295.pdf.

Anonymous:1874:CEa

[5695] Anonymous. The Centennial Exhibition. Scientific American, 30 (19):296, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
296.pdf.

Anonymous:1874:DCn

[5696] Anonymous. Decisions of the courts. Scientific American, 30(19):297, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-297.pdf.

Anonymous:1874:RAFq

[5697] Anonymous. Recent American and foreign patents. Scientific American, 30(19):297, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
297b.pdf.

Anonymous:1874:ACq

[5698] Anonymous. Answers to correspondents. Scientific American, 30 (19):298–299, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
298a.pdf.

Anonymous:1874:CRq

[5699] Anonymous. Communications received. Scientific American, 30 (19):299, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n19/pdf/scientificamerican05091874-
299.pdf.

Anonymous:1874:CPo

[5700] Anonymous. Canadian patents. Scientific American, 30(19):300, May 9, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n19/pdf/scientificamerican05091874-300.pdf.

Anonymous:1874:Cv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-308a.pdf.

Anonymous:1874:BNBn

[5702] Anonymous. Books: New books and publications. Scientific American, 30(20):312, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
312.pdf.

Anonymous:1874:BPr

[5703] Anonymous. Business and personal. Scientific American, 30(20):314, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-314.pdf.

Anonymous:1874:IIr

[5704] Anonymous. Index of inventions. Scientific American, 30(20):315–316, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-315a.pdf.

Anonymous:1874:PNPf

[5705] Anonymous. Patent notes: Patents. Scientific American, 30(20):316, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-316a.pdf.

Anonymous:1874:DCPa

[5706] Anonymous. The Dayton cam pump. Scientific American, 30(20):303, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-303.pdf.

Anonymous:1874:HSPa

[5707] Anonymous. The Hamilton surface planer. Scientific American, 30(20):303–304, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
303a.pdf.

Anonymous:1874:SMM

[5708] Anonymous. The sewing machine monopolists again at work. Scientific American, 30(20):304, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
304.pdf.

Anonymous:1874:DIT

[5709] Anonymous. The depression in the iron trade. Scientific American, 30(20):304, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
304a.pdf.

Anonymous:1874:GRC

[5710] Anonymous. Glacial remains in Central America. Scientific American, 30(20):304–305, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
304b.pdf.

Anonymous:1874:DSL

[5711] Anonymous. Dead subjects and a live discussion. Scientific American, 30(20):305, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
305.pdf.

Anonymous:1874:Ob

[5712] Anonymous. Oleomargarin. Scientific American, 30(20):305, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-305a.pdf.

Anonymous:1874:SPIp

[5713] Anonymous. Scientific and practical information. Scientific American, 30(20):305, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
305b.pdf.

Anonymous:1874:IIs

[5714] Anonymous. An inexhaustible inkstand. Scientific American, 30 (20):306, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
306.pdf.

Anonymous:1874:CPBa

[5715] Anonymous. The Colorado potato bug. Scientific American, 30(20):306, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-306a.pdf.

Anonymous:1874:SSa

[5716] Anonymous. The shadow sail. Scientific American, 30(20):306, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-306b.pdf.

Anonymous:1874:ILB

[5717] Anonymous. Improved lard and butter cutter. Scientific American, 30(20):306, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
306c.pdf.

Thurston:1874:SDT

[5718] R. H. Thurston. Strain diagrams and their revelations. Scientific American, 30(20):307, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
307.pdf.

Anonymous:1874:PNA

[5719] Anonymous. Proceedings of the National Academy of Science. Scientific American, 30(20):308, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
308.pdf.

Anonymous:1874:NDD

[5720] Anonymous. The Niagara direct and double acting pump. Scientific American, 30(20):310, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
310.pdf.

Anonymous:1874:IAP

[5721] Anonymous. Improved automatic packing and weighing machine. Scientific American, 30(20):310, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
310a.pdf.

Anonymous:1874:MEG

[5722] Anonymous. Modern Egyptian gardens. Scientific American, 30 (20):311, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
311.pdf.

Anonymous:1874:BSS

[5723] Anonymous. The bow and Stern screw propeller. Scientific American, 30(20):311–312, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
311a.pdf.

Anonymous:1874:RAFr

[5724] Anonymous. Recent American and foreign patents. Scientific American, 30(20):312–313, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
312a.pdf.

Anonymous:1874:ACr

[5725] Anonymous. Answers to correspondents. Scientific American, 30 (20):314–315, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
314a.pdf.

Anonymous:1874:CRr (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n20/pdf/scientificamerican05161874-
315.pdf.

Anonymous:1874:CPp

[5727] Anonymous. Canadian patents. Scientific American, 30(20):316, May 16, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n20/pdf/scientificamerican05161874-316.pdf.

Anonymous:1874:Cw

[5728] Anonymous. Correspondence. Scientific American, 30(21):324, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-324a.pdf.

Anonymous:1874:BPs

[5729] Anonymous. Business and personal. Scientific American, 30(21):330, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-330.pdf.

Anonymous:1874:IIt

[5730] Anonymous. Index of inventions. Scientific American, 30(21):331–332, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-331a.pdf.

Anonymous:1874:PNPg

[5731] Anonymous. Patent notes: Patents. Scientific American, 30(21):332, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-332a.pdf.

Anonymous:1874:ABb

[5732] Anonymous. The Albion boiler. Scientific American, 30(21):319, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-319.pdf.

Anonymous:1874:MML

[5733] Anonymous. Muscular movement without life. Scientific American, 30(21):320, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
320.pdf.

Anonymous:1874:DCZ

[5734] Anonymous. Discovery of the cause of the zodiacal light. Scientific American, 30(21):320, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
320a.pdf.

Anonymous:1874:PTL

[5735] Anonymous. A proposed testing laboratory. Scientific American, 30(21):320, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
320b.pdf.

Anonymous:1874:HDW

[5736] Anonymous. Hon. David A. Wells and the French Institute. Scientific American, 30(21):320–321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
320c.pdf.

Anonymous:1874:CCa

[5737] Anonymous. From chaos to coral. Scientific American, 30(21):321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-321.pdf.

Anonymous:1874:AAA

[5738] Anonymous. Art among the Ashantees. Scientific American, 30(21):321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-321a.pdf.

Anonymous:1874:BSa

[5739] Anonymous. Burial in the sea. Scientific American, 30(21):321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-321b.pdf.

Anonymous:1874:RSEa

[5740] Anonymous. Regulating the speed of an engine. Scientific American, 30(21):321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
321c.pdf.

Anonymous:1874:MEH

[5741] Anonymous. The magnetic equivalent of heat. Scientific American, 30(21):321, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
321d.pdf.

Anonymous:1874:DPH

[5742] Anonymous. Device for preventing horses from cribbing. Scientific American, 30(21):322, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
322.pdf.

Anonymous:1874:OTC

[5743] Anonymous. Overhead traveling crane. Scientific American, 30(21):322, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-322a.pdf.

Anonymous:1874:IMS

[5744] Anonymous. Improved mixing spoon. Scientific American, 30(21):322, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-322b.pdf.

Anonymous:1874:KHB

[5745] Anonymous. The Keokuk and Hamilton Bridge. Scientific American, 30(21):324, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
324.pdf.

Anonymous:1874:UIS

[5746] Anonymous. Useful information on steam power. Scientific American, 30(21):325, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
325.pdf.

Anonymous:1874:IPFb

[5747] Anonymous. Improved picket fence. Scientific American, 30(21):326, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-326.pdf.

Anonymous:1874:IWH

[5748] Anonymous. Improved work holder for lathes. Scientific American, 30(21):326, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
326a.pdf.

Anonymous:1874:CPq

[5749] Anonymous. Curious plants. Scientific American, 30(21):327–328, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-327.pdf.

Thurston:1874:FSE

[5750] R. H. Thurston. The flow of solids and its effect upon the strength of materials. Scientific American, 30(21):328, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-328.pdf.

Anonymous:1874:DCPb

[5751] Anonymous. Decisions of the Commissioner of Patents. Scientific American, 30(21):328, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
328a.pdf.

Anonymous:1874:DCo

[5752] Anonymous. Decisions of the courts. Scientific American, 30 (21):328–329, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
328b.pdf.

Anonymous:1874:RAFs

[5753] Anonymous. Recent American and foreign patents. Scientific American, 30(21):329, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
329.pdf.

Anonymous:1874:ACt

[5754] Anonymous. Answer correspondents. Scientific American, 30 (21):330–331, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
330a.pdf.

Anonymous:1874:CRs

[5755] Anonymous. Communications received. Scientific American, 30 (21):331, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n21/pdf/scientificamerican05231874-
331.pdf.

Anonymous:1874:CPs

[5756] Anonymous. Canadian patents. Scientific American, 30(21):332, May 23, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n21/pdf/scientificamerican05231874-332.pdf.

Anonymous:1874:Cx

[5757] Anonymous. Correspondence. Scientific American, 30(22):340–341, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-340.pdf.

Anonymous:1874:BNBo

[5758] Anonymous. Books: New books and publications. Scientific American, 30(22):345, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
345.pdf.

Anonymous:1874:BPt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-346.pdf.

Anonymous:1874:IIu

[5760] Anonymous. Index of inventions. Scientific American, 30(22):348, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-348.pdf.

Anonymous:1874:PNPh

[5761] Anonymous. Patent notes: Patents. Scientific American, 30(22):349, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-349.pdf.

Anonymous:1874:IRF

[5762] Anonymous. Improved road and farm locomotive. Scientific American, 30(22):335–336, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
335.pdf.

Anonymous:1874:PMP

[5763] Anonymous. Pressure not a motive power. Scientific American, 30(22):336, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
336.pdf.

Anonymous:1874:GPM

[5764] Anonymous. Governors for prime movers. Scientific American, 30(22):336, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
336a.pdf.

Anonymous:1874:WDP

[5765] Anonymous. Why do plants absorb oxygen during the night? Scientific American, 30(22):336–337, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
336b.pdf.

Anonymous:1874:MDE

[5766] Anonymous. Milk as a diet and its effect on the system. Scientific American, 30(22):337, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
337.pdf.

Anonymous:1874:RO

[5767] Anonymous. The respiration of oxygen. Scientific American, 30 (22):337, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
337a.pdf.

Anonymous:1874:ASS

[5768] Anonymous. The American social science Congress. Scientific American, 30(22):337–338, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
337b.pdf.

Anonymous:1874:PCTb

[5769] Anonymous. The polar clock — the time of day shown by colors. Scientific American, 30(22):338, May 30, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n22/pdf/scientificamerican05301874-338.pdf.

Anonymous:1874:HRM

[5770] Anonymous. Hydraulic riveting machine. Scientific American, 30 (22):339, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
339.pdf.

Anonymous:1874:IDa

[5771] Anonymous. Improved doubletree. Scientific American, 30(22):342, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-342.pdf.

Anonymous:1874:TIH

[5772] Anonymous. Todd’s improved horseshoe. Scientific American, 30 (22):342, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
342a.pdf.

Anonymous:1874:IMMb

[5773] Anonymous. Improved miter machine and frame vise. Scientific American, 30(22):342, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
342b.pdf.

Anonymous:1874:JPC

[5774] Anonymous. Jenkins’ patent compound hatchet. Scientific American, 30(22):342, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
342c.pdf.

Anonymous:1874:SEa

[5775] Anonymous. The Seaforthia Elegans. Scientific American, 30(22):343, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-343.pdf.

Anonymous:1874:BMW

[5776] Anonymous. Bursting of a mountain water reservoir in Massachusetts — a terrible calamity. Scientific American, 30(22):344, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-344.pdf.

Anonymous:1874:RAFt

[5777] Anonymous. Recent American and foreign patents. Scientific American, 30(22):344–345, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
344a.pdf.

Anonymous:1874:ACu

[5778] Anonymous. Answers correspondents. Scientific American, 30 (22):346–347, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
346a.pdf.

Anonymous:1874:CRt

[5779] Anonymous. Communications received. Scientific American, 30 (22):347, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n22/pdf/scientificamerican05301874-
347.pdf.

Anonymous:1874:CPt

[5780] Anonymous. Canadian patents. Scientific American, 30(22):348, May 30, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n22/pdf/scientificamerican05301874-348a.pdf.

Anonymous:1874:Cy

[5781] Anonymous. Correspondence. Scientific American, 30(23):355–356, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-355.pdf.

Anonymous:1874:BNBp

[5782] Anonymous. Books: New books and publications. Scientific American, 30(23):361, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
361a.pdf.

Anonymous:1874:BPu

[5783] Anonymous. Business and personal. Scientific American, 30(23):362, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-362.pdf.

Anonymous:1874:IIv

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-363a.pdf.

Anonymous:1874:PNPi

[5785] Anonymous. Patent notes: Patents. Scientific American, 30(23):365, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-365.pdf.

Anonymous:1874:WBH

[5786] Anonymous. William Baxter and his inventions. Scientific American, 30(23):351–352, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
351.pdf.

Anonymous:1874:FFF

[5787] Anonymous. Formic fungus farmers. Scientific American, 30(23):352, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-352.pdf.

Anonymous:1874:MTS

[5788] Anonymous. A model transatlantic steamer. Scientific American, 30(23):353, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
353.pdf.

Anonymous:1874:LHW

[5789] Anonymous. Left hand writing. Scientific American, 30(23):353, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-353a.pdf.

Anonymous:1874:RDA

[5790] Anonymous. Rubber as a defensive armor. Scientific American, 30(23):353, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
353b.pdf.

Anonymous:1874:PBS

[5791] Anonymous. Photography at the bottom of the sea. Scientific American, 30(23):353, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
353c.pdf.

Anonymous:1874:FCM

[5792] Anonymous. Fattening chickens by machinery. Scientific American, 30(23):354, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
354.pdf.

Anonymous:1874:JS

[5793] Anonymous. Jupiter’s satellites. Scientific American, 30(23):354, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-354a.pdf.

Anonymous:1874:NUN

[5794] Anonymous. The neplus ultra newspaper and magazine file. Scientific American, 30(23):354, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
354b.pdf.

Anonymous:1874:WC

[5795] Anonymous. The wonder camera. Scientific American, 30(23):354, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-354c.pdf.

Anonymous:1874:SPIq

[5796] Anonymous. Scientific and practical information. Scientific American, 30(23):357, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
357.pdf.

Anonymous:1874:TWR

[5797] Anonymous. The tarantula wheel rotary harrow. Scientific American, 30(23):358, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
358.pdf.

Anonymous:1874:HPSa

[5798] Anonymous. Hotz’s patent self-closing faucet. Scientific American, 30(23):358, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
358a.pdf.

Anonymous:1874:CFM

[5799] Anonymous. Caron’s fountain masking brush. Scientific American, 30(23):358, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
358b.pdf.

Anonymous:1874:ECC

[5800] Anonymous. Emanuel Church, Clifton, England. Scientific American, 30(23):359, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
359.pdf.

Anonymous:1874:ANg

[5801] Anonymous. Astronomical notes. Scientific American, 30(23):360, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-360.pdf.

Anonymous:1874:SSc

[5802] Anonymous. Social science. Scientific American, 30(23):360, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-360a.pdf.

Anonymous:1874:CDb

[5803] Anonymous. Commissioner’s decisions. Scientific American, 30(23):360, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-360b.pdf.

Anonymous:1874:DCq

[5804] Anonymous. Decisions of the courts. Scientific American, 30(23):360, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-360c.pdf.

Anonymous:1874:RAFu

[5805] Anonymous. Recent American and foreign patents. Scientific American, 30(23):361, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
361.pdf.

Anonymous:1874:ACv

[5806] Anonymous. Answers of correspondents. Scientific American, 30 (23):362–363, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
362a.pdf.

Anonymous:1874:CRu

[5807] Anonymous. Communications received. Scientific American, 30 (23):363, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n23/pdf/scientificamerican06061874-
363.pdf.

Anonymous:1874:CPu

[5808] Anonymous. Canadian patents. Scientific American, 30(23):364, June 6, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n23/pdf/scientificamerican06061874-364.pdf.

Anonymous:1874:Cz

[5809] Anonymous. Correspondence. Scientific American, 30(24):371–372, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-371c.pdf.

Anonymous:1874:BNBq 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
377a.pdf.

Anonymous:1874:BPv

[5811] Anonymous. Business and personal. Scientific American, 30(24):378, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-378.pdf.

Anonymous:1874:IIw

[5812] Anonymous. Index of inventions. Scientific American, 30(24):380, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-380.pdf.

Anonymous:1874:PNPj

[5813] Anonymous. Patent notes: Patents. Scientific American, 30(24):381, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-381.pdf.

Anonymous:1874:IPSb

[5814] Anonymous. Improved patent sectional boiler. Scientific American, 30(24):367, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
367.pdf.

Anonymous:1874:EEL

[5815] Anonymous. Elevation of the elastic limit by stress. Scientific American, 30(24):368, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
368.pdf.

Anonymous:1874:GSJ

[5816] Anonymous. The geological survey of Japan. Scientific American, 30(24):368, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
368a.pdf.

Anonymous:1874:IUM

[5817] Anonymous. The increased uses of the members. Scientific American, 30(24):368, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
368b.pdf.

Anonymous:1874:NTQ

[5818] Anonymous. The new theory of quantivalence. Scientific American, 30(24):368–369, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
368c.pdf.

Anonymous:1874:FHa

[5819] Anonymous. Farmers’ health. Scientific American, 30(24):369, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-369.pdf.

Anonymous:1874:STc

[5820] Anonymous. Solidified tea. Scientific American, 30(24):370, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-370.pdf.

Anonymous:1874:IWSb

[5821] Anonymous. Improved wheat steamer and dryer. Scientific American, 30(24):370, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
370a.pdf.

Anonymous:1874:NPM

[5822] Anonymous. A new parallel motion. Scientific American, 30(24):371, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-371.pdf.

Anonymous:1874:ICBb

[5823] Anonymous. Improved car bumper. Scientific American, 30(24):371, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-371a.pdf.

Anonymous:1874:NAT

[5824] Anonymous. The new Atlantic telegraph cable — how it was made. Scientific American, 30(24):371, June 13, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n24/pdf/scientificamerican06131874-371b.pdf.

Rose:1874:PMNa

[5825] Joshua Rose. Practical mechanism number I. Scientific American, 30(24):373, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
373.pdf.

Anonymous:1874:WWT

[5826] Anonymous. Wire way for transporting ores, etc. Scientific American, 30(24):374, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
374.pdf.

Anonymous:1874:WCB

[5827] Anonymous. A wire cloth boot. Scientific American, 30(24):374, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-374a.pdf.

Anonymous:1874:HFR

[5828] Anonymous. The Hotchkiss or friction roll drop hammer. Scientific American, 30(24):374–375, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
374b.pdf.

Anonymous:1874:TRH

[5829] Anonymous. The Todd and Rafferty hoisting engine. Scientific American, 30(24):375, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
375.pdf.

Anonymous:1874:CAa

[5830] Anonymous. The coral æchmea. Scientific American, 30(24):375, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-375a.pdf.

Anonymous:1874:SPIr

[5831] Anonymous. Scientific and practical information. Scientific American, 30(24):376, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
376.pdf.

Anonymous:1874:DCr

[5832] Anonymous. Decisions of the courts. Scientific American, 30(24):376, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-376a.pdf.

Anonymous:1874:RAFv

[5833] Anonymous. Recent American and foreign patents. Scientific American, 30(24):377, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
377.pdf.

Anonymous:1874:ACw

[5834] Anonymous. Answers correspondents. Scientific American, 30 (24):378–379, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
378a.pdf.

Anonymous:1874:CRv

[5835] Anonymous. Communications received. Scientific American, 30 (24):379, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n24/pdf/scientificamerican06131874-
379.pdf.

Anonymous:1874:CPv

[5836] Anonymous. Canadian patents. Scientific American, 30(24):380, June 13, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n24/pdf/scientificamerican06131874-380a.pdf.

Anonymous:1874:Cba

[5837] Anonymous. Correspondence. Scientific American, 30(25):388, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-388.pdf.

Anonymous:1874:BNBr

[5838] Anonymous. Books: New books and publications. Scientific American, 30(25):392, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
392a.pdf.

Anonymous:1874:BPw

[5839] Anonymous. Business and personal. Scientific American, 30(25):394, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-394.pdf.

Anonymous:1874:IIx

[5840] Anonymous. Index of inventions. Scientific American, 30(25):395–396, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-395a.pdf.

Anonymous:1874:PNPk

[5841] Anonymous. Patent notes: Patents. Scientific American, 30(25):397, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-397.pdf.

Anonymous:1874:PUJ

[5842] Anonymous. Patent universal jointer. Scientific American, 30(25):383, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-383.pdf.

Anonymous:1874:WF

[5843] Anonymous. Water as fuel. Scientific American, 30(25):384, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-384.pdf.

Anonymous:1874:HMa

[5844] Anonymous. The horse in market. Scientific American, 30(25):384, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-384a.pdf.

Anonymous:1874:HTF

[5845] Anonymous. How to treat fruit trees. Scientific American, 30(25):384, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-384b.pdf.

Anonymous:1874:CGa

[5846] Anonymous. Coffee grounds. Scientific American, 30(25):385, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-385.pdf.

Anonymous:1874:ECa

[5847] Anonymous. Everybody’s Centennial. Scientific American, 30(25):385, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-385a.pdf.

Anonymous:1874:AMN

[5848] Anonymous. The American Museum of Natural History. Scientific American, 30(25):385, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
385b.pdf.

Anonymous:1874:SG

[5849] Anonymous. Steep gradients. Scientific American, 30(25):385, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-385c.pdf.

Anonymous:1874:NDS

[5850] Anonymous. A new domestic steam engine. Scientific American, 30(25):386, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
386.pdf.

Anonymous:1874:MPB

[5851] Anonymous. May’s patent butter worker. Scientific American, 30(25):387, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
387.pdf.

Anonymous:1874:DMC

[5852] Anonymous. Dr. Marey’s chronograph. Scientific American, 30(25):387, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-387a.pdf.

Rose:1874:PMa

[5853] Joshua Rose. Practical mechanism. Scientific American, 30(25):389, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-389.pdf.

Anonymous:1874:IPSc

[5854] Anonymous. Improved punch and screwdriver. Scientific American, 30(25):390, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
390.pdf.

Anonymous:1874:IHa

[5855] Anonymous. Improved hammer. Scientific American, 30(25):390, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-390a.pdf.

Anonymous:1874:MA

[5856] Anonymous. A microscopic aquarium. Scientific American, 30(25):391, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-391.pdf.

Anonymous:1874:RBA

[5857] Anonymous. Remarkable balloon ascent. Scientific American, 30 (25):391–392, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
391a.pdf.

Anonymous:1874:DCs

[5858] Anonymous. Decisions of the courts. Scientific American, 30(25):392, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-392.pdf.

Anonymous:1874:RAFw

[5859] Anonymous. Recent American and foreign patents. Scientific American, 30(25):392–393, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
392b.pdf.

Anonymous:1874:ACx

[5860] Anonymous. Answers correspondents. Scientific American, 30 (25):394–395, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
394a.pdf.

Anonymous:1874:CRw

[5861] Anonymous. Communications received. Scientific American, 30 (25):395, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n25/pdf/scientificamerican06201874-
395.pdf.

Anonymous:1874:CPw

[5862] Anonymous. Canadian patents. Scientific American, 30(25):396–397, June 20, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n25/pdf/scientificamerican06201874-396.pdf.

Anonymous:1874:Cbb

[5863] Anonymous. Correspondence. Scientific American, 30(26):404, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-404.pdf.

Anonymous:1874:BNBs

[5864] Anonymous. Books: New books and publications. Scientific American, 30(26):409–410, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
409a.pdf.

Anonymous:1874:BPx

[5865] Anonymous. Business and personal. Scientific American, 30(26):410, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-410.pdf.

Anonymous:1874:IIy

[5866] Anonymous. Index of inventions. Scientific American, 30(26):410, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-410a.pdf.

Anonymous:1874:Ib

[5867] Anonymous. Index. Scientific American, 30(26):411–412, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-411.pdf.

Anonymous:1874:PNPl

[5868] Anonymous. Patent notes: Patents. Scientific American, 30(26):414, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-414.pdf.

Anonymous:1874:CTD

[5869] Anonymous. The clock tower at Delhi, India. Scientific American, 30(26):399–400, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
399.pdf.

Anonymous:1874:EVX

[5870] Anonymous. The end of volume XXX. Scientific American, 30(26):400, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-400.pdf.

Anonymous:1874:HAH

[5871] Anonymous. How to attain high temperatures. Scientific American, 30(26):400, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
400a.pdf.

Anonymous:1874:CPx

[5872] Anonymous. A curious problem. Scientific American, 30(26):400–401, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-400b.pdf.

Anonymous:1874:SESa

[5873] Anonymous. Sources of edible starch. Scientific American, 30(26):401, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-401.pdf.

Anonymous:1874:FPE

[5874] Anonymous. Failure of patent extension schemes. Scientific American, 30(26):401, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
401a.pdf.

Anonymous:1874:SPIs

[5875] Anonymous. Scientific and practical information. Scientific American, 30(26):401, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
401b.pdf.

Anonymous:1874:IWWa

[5876] Anonymous. Improved wind wheel and water elevator. Scientific American, 30(26):402, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
402.pdf.

Anonymous:1874:MTC

[5877] Anonymous. Machine for turning Crank pins and journals of locomotives. Scientific American, 30(26):402, June 27, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v30/
n26/pdf/scientificamerican06271874-402a.pdf.

Anonymous:1874:EAPa

[5878] Anonymous. Effects of air pressure on animal life. Scientific American, 30(26):403–404, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
403.pdf.

Rose:1874:PMb

[5879] Joshua Rose. Practical mechanism. Scientific American, 30(26):405, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-405.pdf.

Anonymous:1874:SIH

[5880] Anonymous. Stevens’ improved hinge. Scientific American, 30(26):406, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-406.pdf.

Anonymous:1874:IPGb

[5881] Anonymous. Improved patent gang saw table. Scientific American, 30(26):406, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
406a.pdf.

Anonymous:1874:DIHa

[5882] Anonymous. Davis’ improved hydrant. Scientific American, 30(26):406, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-406b.pdf.

Anonymous:1874:RGC

[5883] Anonymous. The Royal Gardens at Caserta, Italy. Scientific American, 30(26):407, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
407.pdf.

Anonymous:1874:MNb

[5884] Anonymous. Medical notes. Scientific American, 30(26):407–408, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v30/n26/pdf/scientificamerican06271874-407a.pdf.

Anonymous:1874:CCEa

[5885] Anonymous. The convention of the civil engineers. Scientific American, 30(26):408–409, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
408.pdf.

Anonymous:1874:RAFx

[5886] Anonymous. Recent American and foreign patents. Scientific American, 30(26):409, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
409.pdf.

Anonymous:1874:CRx

[5887] Anonymous. Communications received. Scientific American, 30 (26):410, June 27, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v30/n26/pdf/scientificamerican06271874-
410b.pdf.

Anonymous:1874:Cbc

[5888] Anonymous. Correspondence. Scientific American, 31(1):5–6, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-5b.pdf.

Anonymous:1874:BNBt

[5889] Anonymous. Books: New books and publications. Scientific American, 31(1):9, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n1/pdf/scientificamerican07041874-
9b.pdf.

Anonymous:1874:BPy

[5890] Anonymous. Business and personal. Scientific American, 31(1):11, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-11.pdf.

Anonymous:1874:IIz

[5891] Anonymous. Index of inventions. Scientific American, 31(1):12, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-12.pdf.

Anonymous:1874:PNPm

[5892] Anonymous. Patent notes: Patents. Scientific American, 31(1):13, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-13.pdf.

Anonymous:1874:IDM

[5893] Anonymous. Improved dovetailing machine. Scientific American, 31(1): 1, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-1.pdf.

Anonymous:1874:HWH

[5894] Anonymous. Hot weather and hydrophobia. Scientific American, 31(1): 2, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-2.pdf.

Anonymous:1874:NCC

[5895] Anonymous. The new comet. Coggia. Scientific American, 31(1):2–3, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-2a.pdf.

Anonymous:1874:PFb

[5896] Anonymous. Protection from fire. Scientific American, 31(1):3, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-3.pdf.

Anonymous:1874:SBO

[5897] Anonymous. Science beads an oracle relating to the metallurgy of the future. Scientific American, 31(1):3, July 4, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n1/pdf/scientificamerican07041874-3a.pdf.

Anonymous:1874:AL

[5898] Anonymous. The Astor Library. Scientific American, 31(1):3, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-3b.pdf.

Anonymous:1874:PS

[5899] Anonymous. Philosophers on sound. Scientific American, 31(1):3, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-3c.pdf.

Anonymous:1874:EPL

[5900] Anonymous. Electric phenomena of the leaves of the fly catcher. Scientific American, 31(1):4, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n1/pdf/scientificamerican07041874-
4.pdf.

Anonymous:1874:BPSa

[5901] Anonymous. Bonser’s patent steam trap. Scientific American, 31(1):4, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-4a.pdf.

Anonymous:1874:BSb

[5902] Anonymous. Bloodless surgery. Scientific American, 31(1):5, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-5.pdf.

Anonymous:1874:FAH

[5903] Anonymous. Feathering arrow heads. Scientific American, 31(1):5, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-5a.pdf.

Anonymous:1874:NMM

[5904] Anonymous. A novel mechanical movement. Scientific American, 31(1): 6, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-6.pdf.

Anonymous:1874:IGT

[5905] Anonymous. Improved grappling tools. Scientific American, 31(1):6, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-6a.pdf.

Anonymous:1874:PPS

[5906] Anonymous. Pond’s patent spanner. Scientific American, 31(1):6, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-6b.pdf.

Anonymous:1874:JSP

[5907] Anonymous. The Jubea Spectabilis palm. Scientific American, 31(1):7, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-7.pdf.

Anonymous:1874:CCEb

[5908] Anonymous. The convention of the civil engineers. Scientific American, 31(1):8–9, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n1/pdf/scientificamerican07041874-
8.pdf.

Anonymous:1874:SPIt

[5909] Anonymous. Scientific and practical information. Scientific American, 31(1):9, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n1/pdf/scientificamerican07041874-
9.pdf.

Anonymous:1874:DCu

[5910] Anonymous. Decisions of to courts. Scientific American, 31(1):9, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-9a.pdf.

Anonymous:1874:RAFy

[5911] Anonymous. Recent American and foreign patents. Scientific American, 31(1):9–10, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n1/pdf/scientificamerican07041874-
9c.pdf.

Anonymous:1874:ACy

[5912] Anonymous. Answers to correspondents. Scientific American, 31(1):11, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-11a.pdf.

Anonymous:1874:CRy

[5913] Anonymous. Communication received. Scientific American, 31(1):11, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-11b.pdf.

Anonymous:1874:CPy

[5914] Anonymous. Canadian patents. Scientific American, 31(1):12, July 4, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n1/pdf/scientificamerican07041874-12a.pdf.

Anonymous:1874:Cbd

[5915] Anonymous. Correspondence. Scientific American, 31(2):20, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-20.pdf.

Anonymous:1874:BNBu

[5916] Anonymous. Books: New books and publications. Scientific American, 31(2):24, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
24a.pdf.

Anonymous:1874:BPz

[5917] Anonymous. Business and personal. Scientific American, 31(2):26, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-26.pdf.

Anonymous:1874:IIba

[5918] Anonymous. Index of inventions. Scientific American, 31(2):27–28, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-27a.pdf.

Anonymous:1874:PNPn

[5919] Anonymous. Patent notes: Patents. Scientific American, 31(2):29, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-29.pdf.

Anonymous:1874:BSR

[5920] Anonymous. Balanced screw and revolving cotton press. Scientific American, 31(2):15, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
15.pdf.

Anonymous:1874:GAS

[5921] Anonymous. Government aid to scientific investigation. Scientific American, 31(2):16, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
16.pdf.

Anonymous:1874:FFM

[5922] Anonymous. Four footed motion. Scientific American, 31(2):16, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-16a.pdf.

Anonymous:1874:RUS

[5923] Anonymous. The railways of the United States. Scientific American, 31(2):16, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
16b.pdf.

Anonymous:1874:SUP

[5924] Anonymous. Some of the uses of paraffin. Scientific American, 31(2):16–17, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
16c.pdf.

Anonymous:1874:RMT

[5925] Anonymous. Refrigerating mixtures and their physiological effects. Scientific American, 31(2):17, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
17.pdf.

Anonymous:1874:PFL

[5926] Anonymous. Progress of the fireless locomotive. Scientific American, 31(2):17, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
17a.pdf.

Anonymous:1874:NLC

[5927] Anonymous. New law concerning copyrights for labels. Scientific American, 31(2):17, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
17b.pdf.

Anonymous:1874:SPIu

[5928] Anonymous. Scientific and practical information. Scientific American, 31(2):17, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
17c.pdf.

Anonymous:1874:EAPb

[5929] Anonymous. The effect of air pressure on animal life. Scientific American, 31(2):18, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
18.pdf.

Anonymous:1874:ISEb

[5930] Anonymous. Improved steam engine. Scientific American, 31(2):19, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-19.pdf.

Anonymous:1874:GC

[5931] Anonymous. The gravitation compass. Scientific American, 31(2):19, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-19a.pdf.

Anonymous:1874:NDB

[5932] Anonymous. A navigable diving Bell. Scientific American, 31(2):19–20, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-19b.pdf.

Anonymous:1874:ANh

[5933] Anonymous. Astronomical notes. Scientific American, 31(2):20, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-20a.pdf.

Rose:1874:PMc

[5934] Joshua Rose. Practical mechanism. Scientific American, 31(2):21, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-21.pdf.

Anonymous:1874:TSP

[5935] Anonymous. Taper sleeve pulley and wheel fastener. Scientific American, 31(2):22, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
22.pdf.

Anonymous:1874:NWP

[5936] Anonymous. A new water pitcher. Scientific American, 31(2):22, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-22a.pdf.

Anonymous:1874:NNa

[5937] Anonymous. A novel needle. Scientific American, 31(2):22, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-22b.pdf.

Anonymous:1874:VMC

[5938] Anonymous. Various methods of cooling air. Scientific American, 31(2):23–24, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
23.pdf.

Anonymous:1874:MNc

[5939] Anonymous. Medical notes. Scientific American, 31(2):24, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-24.pdf.

Anonymous:1874:RAFz

[5940] Anonymous. Recent American and foreign patents. Scientific American, 31(2):24–25, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
24b.pdf.

Anonymous:1874:ACz

[5941] Anonymous. Answers and correspondents. Scientific American, 31(2):26–27, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n2/pdf/scientificamerican07111874-
26a.pdf.

Anonymous:1874:CRz

[5942] Anonymous. Communication received. Scientific American, 31(2):27, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-27.pdf.

Anonymous:1874:CPz

[5943] Anonymous. Canadian patents. Scientific American, 31(2):28, July 11, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n2/pdf/scientificamerican07111874-28.pdf.

Anonymous:1874:Cbe

[5944] Anonymous. Correspondence. Scientific American, 31(3):36, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-36a.pdf.

Anonymous:1874:BNBv

[5945] Anonymous. Books: New books and publications. Scientific American, 31(3):42, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
42.pdf.

Anonymous:1874:BPba

[5946] Anonymous. Business and personal. Scientific American, 31(3):43, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-43.pdf.

Anonymous:1874:IIbb

[5947] Anonymous. Index of inventions. Scientific American, 31(3):44, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-44.pdf.

Anonymous:1874:PNPo

[5948] Anonymous. Patent notes: Patents. Scientific American, 31(3):45, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-45.pdf.

Anonymous:1874:ESG

[5949] Anonymous. The eclipse steam generator. Scientific American, 31(3):31, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
31.pdf.

Anonymous:1874:CRba

[5950] Anonymous. A cometary retrospect. Scientific American, 31(3):32, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-32.pdf.

Anonymous:1874:TPL

[5951] Anonymous. Taxing power of the legislature. Scientific American, 31(3):32, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
32a.pdf.

Anonymous:1874:RPP

[5952] Anonymous. Recent printing press improvements. Scientific American, 31(3):32–33, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
32b.pdf.

Anonymous:1874:PCI

[5953] Anonymous. Patented car improvements. Scientific American, 31(3):33, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-33.pdf.

Anonymous:1874:MMI

[5954] Anonymous. The mines and minerals of India. Scientific American, 31(3):33, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
33a.pdf.

Anonymous:1874:CSM

[5955] Anonymous. Casting the standard meter. Scientific American, 31(3):34, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-34.pdf.

Anonymous:1874:CTE

[5956] Anonymous. The clamond thermo-electric battery. Scientific American, 31(3):34–31, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
34a.pdf.

Anonymous:1874:FCV

[5957] Anonymous. Fountain Cave, Virginia. Scientific American, 31(3):36, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-36.pdf.

Anonymous:1874:SPIv

[5958] Anonymous. Scientific and practical information. Scientific American, 31(3):37, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
37.pdf.

Anonymous:1874:NIS

[5959] Anonymous. Novel improvements in steam boilers. Scientific American, 31(3):37, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
37a.pdf.

Anonymous:1874:IHCa

[5960] Anonymous. Improved hay and cotton press. Scientific American, 31(3):38, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
38.pdf.

Anonymous:1874:ISPb

[5961] Anonymous. An improvement in steam pumps. Scientific American, 31(3):38, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
38a.pdf.

Anonymous:1874:NCSa

[5962] Anonymous. A novel cooking stove. Scientific American, 31(3):38, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-38b.pdf.

Anonymous:1874:MCV

[5963] Anonymous. More curious vegetation. Scientific American, 31(3):39, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-39.pdf.

Anonymous:1874:MNd

[5964] Anonymous. Medical notes. Scientific American, 31(3):40, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-40.pdf.

Anonymous:1874:PMb

[5965] Anonymous. A petroleum motor. Scientific American, 31(3):41–42, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-41.pdf.

Anonymous:1874:RAFba

[5966] Anonymous. Recent American and foreign patents. Scientific American, 31(3):42, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
42a.pdf.

Anonymous:1874:ACba

[5967] Anonymous. Answers and correspondents. Scientific American, 31(3):43, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
43a.pdf.

Anonymous:1874:CRbb

[5968] Anonymous. Communications received. Scientific American, 31 (3):43–44, July 18, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n3/pdf/scientificamerican07181874-
43b.pdf.

Anonymous:1874:CPbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n3/pdf/scientificamerican07181874-44a.pdf.

Anonymous:1874:Cbf

[5970] Anonymous. Correspondence. Scientific American, 31(4):52–53, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-52.pdf.

Anonymous:1874:BPbb

[5971] Anonymous. Business and personal. Scientific American, 31(4):58, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-58.pdf.

Anonymous:1874:IIbc

[5972] Anonymous. Index of inventions. Scientific American, 31(4):59–60, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-59a.pdf.

Anonymous:1874:IHM

[5973] Anonymous. Improved hoisting machine. Scientific American, 31(4):47–48, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
47.pdf.

Anonymous:1874:MMC

[5974] Anonymous. Morbid mental conditions. Scientific American, 31(4):48, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-48.pdf.

Anonymous:1874:CTOb

[5975] Anonymous. Cheap transit for oil. Scientific American, 31(4):48, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-48a.pdf.

Anonymous:1874:PRD

[5976] Anonymous. Power required to drive cotton machinery. Scientific American, 31(4):48–49, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
48b.pdf.

Anonymous:1874:SA

[5977] Anonymous. A switch accident. Scientific American, 31(4):49, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-49.pdf.

Anonymous:1874:MGS

[5978] Anonymous. The manufacture of gem stones. Scientific American, 31(4):49, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
49a.pdf.

Anonymous:1874:SPIw

[5979] Anonymous. Scientific and practical information. Scientific American, 31(4):49–50, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
49b.pdf.

Anonymous:1874:GSI

[5980] Anonymous. Gang saw improvements. Scientific American, 31(4):50, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-50.pdf.

Anonymous:1874:SF

[5981] Anonymous. Safety furnace. Scientific American, 31(4):50, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-50a.pdf.

Anonymous:1874:ATT

[5982] Anonymous. Apparatus for transplanting trees. Scientific American, 31(4):51, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
51.pdf.

Anonymous:1874:PE

[5983] Anonymous. Practical education. Scientific American, 31(4):53, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-53.pdf.

Anonymous:1874:NIM

[5984] Anonymous. New ice making process. Scientific American, 31(4):54, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-54.pdf.

Anonymous:1874:EII

[5985] Anonymous. Earle’s improved ice creeper. Scientific American, 31(4):54, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
54a.pdf.

Anonymous:1874:IFM

[5986] Anonymous. Improved furnace for melting brass. Scientific American, 31(4):54, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
54b.pdf.

Anonymous:1874:NBA

[5987] Anonymous. A novel balloon ascent. Scientific American, 31(4):55–56, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-55.pdf.

Anonymous:1874:NSBa

[5988] Anonymous. The new steamer Britannic — a new propeller improvement. Scientific American, 31(4):56, July 25, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n4/pdf/scientificamerican07251874-56.pdf.

Anonymous:1874:DCv

[5989] Anonymous. Decisions of the courts. Scientific American, 31(4):56, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-56a.pdf.

Anonymous:1874:RAFbb

[5990] Anonymous. Recent American and foreign patents. Scientific American, 31(4):57, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
57.pdf.

Anonymous:1874:ACbb

[5991] Anonymous. Answers and correspondents. Scientific American, 31(4):58–59, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n4/pdf/scientificamerican07251874-
58a.pdf.

Anonymous:1874:CRbc

[5992] Anonymous. Communications received. Scientific American, 31(4):59, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-59.pdf.

Anonymous:1874:CPbc

[5993] Anonymous. Canadian patents. Scientific American, 31(4):60, July 25, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n4/pdf/scientificamerican07251874-60.pdf.

Anonymous:1874:Cbg

[5994] Anonymous. Correspondence. Scientific American, 31(5):68, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-68.pdf.

Anonymous:1874:BPbc

[5995] Anonymous. Business and personal. Scientific American, 31(5):74, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-74.pdf.

Anonymous:1874:IIbd

[5996] Anonymous. Index of inventions. Scientific American, 31(5):75–76, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-75a.pdf.

Anonymous:1874:PNPp

[5997] Anonymous. Patent notes: Patents. Scientific American, 31(5):76, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-76a.pdf.

Anonymous:1874:IGCc

[5998] Anonymous. Improved gear cutting attachment for lathes. Scientific American, 31(5):63, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
63.pdf.

Anonymous:1874:AGF

[5999] Anonymous. Another great fire in Chicago, and the materials accumulating for one in New York. Scientific American, 31(5):64, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-64.pdf.

Anonymous:1874:CTa

[6000] Anonymous. Cholera and its treatment. Scientific American, 31(5):64, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-64a.pdf.

Anonymous:1874:IISa

[6001] Anonymous. The inspection and insurance of steam boilers. Scientific American, 31(5):64–65, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
64b.pdf.

Anonymous:1874:POJ

[6002] Anonymous. Patent Office justice. Scientific American, 31(5):65, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-65.pdf.

Anonymous:1874:PBb

[6003] Anonymous. Politics in the beehive. Scientific American, 31(5):65, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-65a.pdf.

Anonymous:1874:LMa

[6004] Anonymous. The locust in Minnesota. Scientific American, 31(5):65, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-65b.pdf.

Anonymous:1874:FHF

[6005] Anonymous. Four hundred and forty-four miles, at over forty miles per hour, and three stops. Scientific American, 31(5):65, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-65c.pdf.

Anonymous:1874:MCB

[6006] Anonymous. The Mezzana-Corti Bridge over the River Po. Scientific American, 31(5):66–67, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
66.pdf.

Anonymous:1874:RBB

[6007] Anonymous. Rolling bridge between St. Servan and St. Malo. Scientific American, 31(5):67, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
67.pdf.

Anonymous:1874:DIW

[6008] Anonymous. The Degerfors iron works, Sweden. Scientific American, 31(5):67–68, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
67a.pdf.

Rose:1874:PMd

[6009] Joshua Rose. Practical mechanism. Scientific American, 31(5):69, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-69.pdf.

Anonymous:1874:ISMc

[6010] Anonymous. Improved sewing machine motor. Scientific American, 31(5):70, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
70.pdf.

Anonymous:1874:SPC

[6011] Anonymous. The Sundridge Park Conservatory, England. Scientific American, 31(5):71, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
71.pdf.

Anonymous:1874:BIE

[6012] Anonymous. The British International Exhibitions. Scientific American, 31(5):71, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
71a.pdf.

Anonymous:1874:HPSb

[6013] Anonymous. Hydrophobia-practical suggestions for:its prevention and cure. Scientific American, 31(5):72, August 1, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n5/pdf/scientificamerican08011874-72.pdf.

Anonymous:1874:HSId

[6014] Anonymous. How shall I introduce my invention? Scientific American, 31(5):72, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
72a.pdf.

Anonymous:1874:DCw

[6015] Anonymous. Decisions of the courts. Scientific American, 31(5):72, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-72b.pdf.

Anonymous:1874:RAFbc

[6016] Anonymous. Recent American and foreign patents. Scientific American, 31(5):72–73, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
72c.pdf.

Anonymous:1874:ACbc

[6017] Anonymous. Answers and correspondents. Scientific American, 31(5):74–75, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n5/pdf/scientificamerican08011874-
74a.pdf.

Anonymous:1874:CRbd

[6018] Anonymous. Communications received. Scientific American, 31(5):75, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-75.pdf.

Anonymous:1874:CPbd

[6019] Anonymous. Canadian patents. Scientific American, 31(5):76, August 1, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n5/pdf/scientificamerican08011874-76.pdf.

Anonymous:1874:Cbh

[6020] Anonymous. Correspondence. Scientific American, 31(6):84, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-84.pdf.

Anonymous:1874:BNBw

[6021] Anonymous. Books: New books and publications. Scientific American, 31(6):88, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
88a.pdf.

Anonymous:1874:BPbd

[6022] Anonymous. Business and personal. Scientific American, 31(6):90, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-90.pdf.

Anonymous:1874:IIbe

[6023] Anonymous. Index of inventions. Scientific American, 31(6):92, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-92a.pdf.

Anonymous:1874:PNPq

[6024] Anonymous. Patent notes: Patents. Scientific American, 31(6):93, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-93.pdf.

Anonymous:1874:NGB

[6025] Anonymous. New granary at Bristol, England. Scientific American, 31(6):79–80, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
79.pdf.

Anonymous:1874:SPSb

[6026] Anonymous. The simultaneous perception of sounds. Scientific American, 31(6):80, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
80.pdf.

Anonymous:1874:A

[6027] Anonymous. The aquarium. Scientific American, 31(6):80, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-80a.pdf.

Anonymous:1874:PMd

[6028] Anonymous. The practical man. Scientific American, 31(6):80–81, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-80b.pdf.

Anonymous:1874:LIA

[6029] Anonymous. Language of insects and animals. Scientific American, 31(6):81, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
81.pdf.

Anonymous:1874:DCx

[6030] Anonymous. Dead cities. Scientific American, 31(6):81, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-81a.pdf.

Anonymous:1874:SPIx

[6031] Anonymous. Scientific and practical information. Scientific American, 31(6):81–82, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
81b.pdf.

Anonymous:1874:IO

[6032] Anonymous. An Iris and an orchid. Scientific American, 31(6):82, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-82.pdf.

Anonymous:1874:OAG

[6033] Anonymous. Oak apples or galls. Scientific American, 31(6):82, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-82a.pdf.

Anonymous:1874:FMF

[6034] Anonymous. The flower markets of Florence, Italy. Scientific American, 31(6):83, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
83.pdf.

Anonymous:1874:RAE

[6035] Anonymous. Recent archeological explorations. Scientific American, 31(6):83–84, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
83a.pdf.

Anonymous:1874:MNe

[6036] Anonymous. Medical notes. Scientific American, 31(6):85, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-85.pdf.

Anonymous:1874:OCG

[6037] Anonymous. Ord’s cut-off governor. Scientific American, 31(6):86, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-86.pdf.

Anonymous:1874:IPBa

[6038] Anonymous. Improved paint brush. Scientific American, 31(6):86, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-86a.pdf.

Anonymous:1874:MSH

[6039] Anonymous. The Messrs. Stevens of Hoboken, and their ironclads. Scientific American, 31(6):87–88, August 8, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n6/pdf/scientificamerican08081874-87.pdf.

Anonymous:1874:DCy

[6040] Anonymous. Decisions of the courts. Scientific American, 31(6):88, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-88.pdf.

Anonymous:1874:RAFbd

[6041] Anonymous. Recent American and foreign patents. Scientific American, 31(6):88–89, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
88b.pdf.

Anonymous:1874:ACbd

[6042] Anonymous. Answers to correspondents. Scientific American, 31(6):90–92, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n6/pdf/scientificamerican08081874-
90a.pdf.

Anonymous:1874:CRbe

[6043] Anonymous. Communications received. Scientific American, 31(6):92, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-92.pdf.

Anonymous:1874:CPbe

[6044] Anonymous. Canadian patents. Scientific American, 31(6):92, August 8, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n6/pdf/scientificamerican08081874-92b.pdf.

Anonymous:1874:Cbi

[6045] Anonymous. Correspondence. Scientific American, 31(7):100, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-100a.pdf.

Anonymous:1874:BPbe

[6046] Anonymous. Business and personal. Scientific American, 31(7):106, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-106.pdf.

Anonymous:1874:IIbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-108a.pdf.

Anonymous:1874:PNPr

[6048] Anonymous. Patent notes: Patents. Scientific American, 31(7):109, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-109.pdf.

Anonymous:1874:GII

[6049] Anonymous. The Gartsherrie iron industry. Scientific American, 31(7):95, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
95.pdf.

Anonymous:1874:NTF

[6050] Anonymous. A new theory of the formation of diamonds. Scientific American, 31(7):96, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
96.pdf.

Anonymous:1874:ELD

[6051] Anonymous. The exploration of the Libyan Desert. Scientific American, 31(7):96, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
96a.pdf.

Anonymous:1874:CYS

[6052] Anonymous. Can you swim. Scientific American, 31(7):96–97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-96b.pdf.

Anonymous:1874:PHH

[6053] Anonymous. Professor Huxley and Harvard. Scientific American, 31(7):97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
97.pdf.

Anonymous:1874:DF

[6054] Anonymous. Disastrous floods. Scientific American, 31(7):97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-97a.pdf.

Anonymous:1874:TGM

[6055] Anonymous. Tides in the Gulf of Mexico. Scientific American, 31(7):97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
97b.pdf.

Anonymous:1874:FLC

[6056] Anonymous. Fracture by long-continued jarring. Scientific American, 31(7):97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
97c.pdf.

Anonymous:1874:EFA

[6057] Anonymous. English food adulteration. Scientific American, 31 (7):97, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
97d.pdf.

Anonymous:1874:IVB

[6058] Anonymous. Improved vertical boiler. Scientific American, 31 (7):98, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
98.pdf.

Anonymous:1874:ICTa

[6059] Anonymous. Improved crow and tamping bar. Scientific American, 31(7):98, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
98a.pdf.

Anonymous:1874:WBT

[6060] Anonymous. A water belt for transmitting motion. Scientific American, 31(7):99, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
99.pdf.

Anonymous:1874:NSBb

[6061] Anonymous. A new sectional balloon. Scientific American, 31 (7):99–100, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
99a.pdf.

Anonymous:1874:RNY

[6062] Anonymous. Railways in New York City. Scientific American, 31(7):100, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
100.pdf.

Rose:1874:PMNb

[6063] Joshua Rose. Practical mechanism. number VI. Scientific American, 31(7):101, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
101.pdf.

Anonymous:1874:IMC

[6064] Anonymous. Improved metal cutting and punching machine. Scientific American, 31(7):102, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
102.pdf.

Anonymous:1874:HIL

[6065] Anonymous. Holden’s improved loose pulley. Scientific American, 31(7):102, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
102a.pdf.

Anonymous:1874:DIHb

[6066] Anonymous. Davis’ improved hub. Scientific American, 31(7): 102–103, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
102b.pdf.

Anonymous:1874:CEb

[6067] Anonymous. The Chilian Exposition. Scientific American, 31 (7):103–104, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
103.pdf.

Anonymous:1874:IAi

[6068] Anonymous. Importance of advertising. Scientific American, 31 (7):104, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
104.pdf.

Anonymous:1874:DCba

[6069] Anonymous. Decisions of the courts. Scientific American, 31 (7):104, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
104a.pdf.

Anonymous:1874:RAFbe

[6070] Anonymous. Recent American and foreign patents. Scientific American, 31(7):105, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
105.pdf.

Anonymous:1874:ACbe

[6071] Anonymous. Answers to correspondents. Scientific American, 31 (7):106–108, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
106a.pdf.

Anonymous:1874:CRbf

[6072] Anonymous. Communications received. Scientific American, 31 (7):108, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n7/pdf/scientificamerican08151874-
108.pdf.

Anonymous:1874:CPbf

[6073] Anonymous. Canadian patents. Scientific American, 31(7):108, August 15, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n7/pdf/scientificamerican08151874-108b.pdf.

Anonymous:1874:Cbj

[6074] Anonymous. Correspondence. Scientific American, 31(8):116–117, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-116.pdf.

Anonymous:1874:BNBx

[6075] Anonymous. Books: New books and publications. Scientific American, 31(8):121, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
121a.pdf.

Anonymous:1874:BPbf

[6076] Anonymous. Business and personal. Scientific American, 31(8):123, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-123.pdf.

Anonymous:1874:IIbg

[6077] Anonymous. Index of inventions. Scientific American, 31(8):124, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-124a.pdf.

Anonymous:1874:PNPs

[6078] Anonymous. Patent notes: Patents. Scientific American, 31(8):126, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-126.pdf.

Anonymous:1874:IMU

[6079] Anonymous. Improved machine for unloading grain. Scientific American, 31(8):111, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
111.pdf.

Anonymous:1874:RVR

[6080] Anonymous. Rotative versus rotary engines. Scientific American, 31(8):112, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
112.pdf.

Anonymous:1874:FA

[6081] Anonymous. The fabrication of antiques. Scientific American, 31(8):112, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
112a.pdf.

Anonymous:1874:MMS

[6082] Anonymous. Meteorological and medical statistics. Scientific American, 31(8):112, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
112b.pdf.

Anonymous:1874:KTT

[6083] Anonymous. Kite tails and telegraph wires. Scientific American, 31(8):113, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
113.pdf.

Anonymous:1874:IEF

[6084] Anonymous. Influence of the Earth’s figure on geologic changes. Scientific American, 31(8):113, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
113a.pdf.

Anonymous:1874:OSL

[6085] Anonymous. Our six legged rivals. Scientific American, 31(8):113, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-113b.pdf.

Anonymous:1874:SEB

[6086] Anonymous. Small engines and boilers. Scientific American, 31 (8):113, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
113c.pdf.

Anonymous:1874:SPIy

[6087] Anonymous. Scientific and practical information. Scientific American, 31(8):113, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
113d.pdf.

Anonymous:1874:TVa

[6088] Anonymous. The Transit of Venus. Scientific American, 31(8):114, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-114.pdf.

Anonymous:1874:FAD

[6089] Anonymous. A four-antlered Deer’s head. Scientific American, 31(8):114, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
114a.pdf.

Anonymous:1874:MGA

[6090] Anonymous. Modern German artillery. Scientific American, 31 (8):115, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
115.pdf.

Anonymous:1874:NGD

[6091] Anonymous. New geological discovery. Scientific American, 31 (8):115–116, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
115a.pdf.

Anonymous:1874:YWM

[6092] Anonymous. Young’s water main tapping machine. Scientific American, 31(8):118, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
118.pdf.

Anonymous:1874:IDW

[6093] Anonymous. Improved dumping wagon. Scientific American, 31 (8):118, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
118a.pdf.

Anonymous:1874:WLP

[6094] Anonymous. The Western locust plague. Scientific American, 31 (8):119–121, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
119.pdf.

Anonymous:1874:HSIe

[6095] Anonymous. How shall I introduce my invention! Scientific American, 31(8):121, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
121.pdf.

Anonymous:1874:RAFbf

[6096] Anonymous. Recent American and foreign patents. Scientific American, 31(8):121–122, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
121b.pdf.

Anonymous:1874:ACbf

[6097] Anonymous. Answers to correspondents. Scientific American, 31 (8):123–124, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
123a.pdf.

Anonymous:1874:CRbg

[6098] Anonymous. Communications received. Scientific American, 31 (8):124, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n8/pdf/scientificamerican08221874-
124.pdf.

Anonymous:1874:CPbg

[6099] Anonymous. Canadian patents. Scientific American, 31(8):124, August 22, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n8/pdf/scientificamerican08221874-124b.pdf.

Anonymous:1874:Cbk

[6100] Anonymous. Correspondence. Scientific American, 31(9):132, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-132.pdf.

Anonymous:1874:BNBy

[6101] Anonymous. Books: New books and publications. Scientific American, 31(9):136, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
136b.pdf.

Anonymous:1874:BPbg

[6102] Anonymous. Business and personal. Scientific American, 31(9):138, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-138.pdf.

Anonymous:1874:IIbh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-140a.pdf.

Anonymous:1874:IRM

[6104] Anonymous. Improved rebolting machine. Scientific American, 31(9):127, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
127.pdf.

Anonymous:1874:CFL

[6105] Anonymous. The combination foot lathe. Scientific American, 31 (9):127–128, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
127a.pdf.

Anonymous:1874:LCD

[6106] Anonymous. The light of coming days. Scientific American, 31 (9):128, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
128.pdf.

Anonymous:1874:CIa

[6107] Anonymous. A chance for inventors. Scientific American, 31(9):128, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-128a.pdf.

Anonymous:1874:CBO

[6108] Anonymous. A city built by one man. Scientific American, 31 (9):128, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
128b.pdf.

Anonymous:1874:FDE

[6109] Anonymous. Fire damp explosions produced by sound. Scientific American, 31(9):129, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
129.pdf.

Anonymous:1874:PAAa

[6110] Anonymous. Proceedings of the American Association for the Advancement of Science. Scientific American, 31(9):129, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-129a.pdf.

Anonymous:1874:PRa

[6111] Anonymous. A portable railway. Scientific American, 31(9): 130–131, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
130.pdf.

Anonymous:1874:FPM

[6112] Anonymous. Fabric printing machine. Scientific American, 31 (9):131, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
131.pdf.

Anonymous:1874:SPIz

[6113] Anonymous. Scientific and practical information. Scientific American, 31(9):131, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
131a.pdf.

Buel:1874:SH

[6114] Richard H. Buel. Specific heat. Scientific American, 31(9):132–133, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-132a.pdf.

Rose:1874:PMe

[6115] Joshua Rose. Practical mechanism. Scientific American, 31(9):133, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-133.pdf.

Anonymous:1874:BEW

[6116] Anonymous. The Blackstone and Elmer water wheel. Scientific American, 31(9):134, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
134.pdf.

Anonymous:1874:IHCb

[6117] Anonymous. Improved hand Crank drill. Scientific American, 31(9):134, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
134a.pdf.

Anonymous:1874:BFG

[6118] Anonymous. Bouquets from field and garden. Scientific American, 31(9):135–136, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
135.pdf.

Anonymous:1874:NSM

[6119] Anonymous. A new scientific museum. Scientific American, 31 (9):136, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
136.pdf.

Anonymous:1874:DCbb

[6120] Anonymous. Decisions of the courts. Scientific American, 31 (9):136, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
136a.pdf.

Anonymous:1874:RAFbg

[6121] Anonymous. Recent American and foreign patents. Scientific American, 31(9):136–137, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
136c.pdf.

Anonymous:1874:ACbg

[6122] Anonymous. Answers to correspondents. Scientific American, 31 (9):138–140, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
138a.pdf.

Anonymous:1874:CRbh

[6123] Anonymous. Communications received. Scientific American, 31 (9):140, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n9/pdf/scientificamerican08291874-
140.pdf.

Anonymous:1874:CPbh

[6124] Anonymous. Canadian patents. Scientific American, 31(9):140, August 29, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n9/pdf/scientificamerican08291874-140b.pdf.

Anonymous:1874:Cbl

[6125] Anonymous. Correspondence. Scientific American, 31(10):148, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n10/pdf/scientificamerican09051874-148.pdf.

Anonymous:1874:BNBz

[6126] Anonymous. Books: New books and publications. Scientific American, 31(10):153, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
153b.pdf.

Anonymous:1874:BPbh

[6127] Anonymous. Business and personal. Scientific American, 31 (10):155, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
155.pdf.

Anonymous:1874:IIbi

[6128] Anonymous. Index of inventions. Scientific American, 31(10): 156–157, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
156a.pdf.

Anonymous:1874:PNPt

[6129] Anonymous. Patent notes: Patents. Scientific American, 31 (10):158, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
158.pdf.

Anonymous:1874:SWF

[6130] Anonymous. Sir William Fairbairn. Scientific American, 31(10): 143–144, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
143.pdf.

Anonymous:1874:NIR

[6131] Anonymous. New industrial resources of French colonies. Scientific American, 31(10):144, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
144.pdf.

Anonymous:1874:TPN

[6132] Anonymous. The torpedo practice at Newport. Scientific American, 31(10):144, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
144a.pdf.

Anonymous:1874:RMR

[6133] Anonymous. Recent metallurgical researches. Scientific American, 31(10):144, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
144b.pdf.

Anonymous:1874:PL

[6134] Anonymous. Protection from lightning. Scientific American, 31 (10):144–145, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
144c.pdf.

Anonymous:1874:END

[6135] Anonymous. The electromotograph...a new discovery in telegraphy. Scientific American, 31(10):145, September 5, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n10/pdf/scientificamerican09051874-145.pdf.

Anonymous:1874:TB

[6136] Anonymous. The transfusion of blood. Scientific American, 31 (10):146, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
146.pdf.

Anonymous:1874:PFc

[6137] Anonymous. Popular fallacies. Scientific American, 31(10):148, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n10/pdf/scientificamerican09051874-148a.pdf.

Anonymous:1874:MNf

[6138] Anonymous. Medical notes. Scientific American, 31(10):149, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n10/pdf/scientificamerican09051874-149.pdf.

Anonymous:1874:SPIba

[6139] Anonymous. Scientific and practical information. Scientific American, 31(10):149, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
149a.pdf.

Anonymous:1874:NM

[6140] Anonymous. A new motor. Scientific American, 31(10):150, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n10/pdf/scientificamerican09051874-150.pdf.

Anonymous:1874:FIL

[6141] Anonymous. Folsom’s improved locking latch. Scientific American, 31(10):150, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
150a.pdf.

Anonymous:1874:WIF

[6142] Anonymous. Webster’s improved faucet. Scientific American, 31 (10):150, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
150b.pdf.

Anonymous:1874:MMB

[6143] Anonymous. Manufacture of mortar bricks. Scientific American, 31(10):151, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
151.pdf.

Anonymous:1874:PAAb

[6144] Anonymous. Proceedings of the American Association for the Advancement of Science. Scientific American, 31(10):152–153, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n10/pdf/scientificamerican09051874-152.pdf.

Anonymous:1874:VSO

[6145] Anonymous. The vibrations of solids optically studied. Scientific American, 31(10):153, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
153.pdf.

Anonymous:1874:IAj

[6146] Anonymous. Importance of advertising. Scientific American, 31 (10):153, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
153a.pdf.

Anonymous:1874:RAFbh

[6147] Anonymous. Recent American and foreign patents. Scientific American, 31(10):153–154, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
153c.pdf.

Anonymous:1874:ACbh

[6148] Anonymous. Answers to correspondents. Scientific American, 31 (10):155–156, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
155a.pdf.

Anonymous:1874:CRbi

[6149] Anonymous. Communications received. Scientific American, 31 (10):156, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
156.pdf.

Anonymous:1874:CPbi

[6150] Anonymous. Canadian patents. Scientific American, 31(10): 157, September 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n10/pdf/scientificamerican09051874-
157.pdf.

Anonymous:1874:Cbm

[6151] Anonymous. Correspondence. Scientific American, 31(11):165, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n11/pdf/scientificamerican09121874-165.pdf.

Anonymous:1874:BPbi

[6152] Anonymous. Business and personal. Scientific American, 31 (11):171, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
171.pdf.

Anonymous:1874:IIbj

[6153] Anonymous. Index of inventions. Scientific American, 31(11): 172, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
172a.pdf.

Anonymous:1874:PNPu

[6154] Anonymous. Patent notes: Patents. Scientific American, 31 (11):174, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
174.pdf.

Anonymous:1874:IDS

[6155] Anonymous. Improved diamond stone sawing machine. Scientific American, 31(11):159, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
159.pdf.

Anonymous:1874:CWS

[6156] Anonymous. Can we stay? Scientific American, 31(11):160, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n11/pdf/scientificamerican09121874-160.pdf.

Anonymous:1874:PCd

[6157] Anonymous. Prepare for the Centennial. Scientific American, 31 (11):160, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
160a.pdf.

Anonymous:1874:Pd

[6158] Anonymous. Penholders. Scientific American, 31(11):160, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n11/pdf/scientificamerican09121874-160b.pdf.

Anonymous:1874:MCD

[6159] Anonymous. The microscope as a criminal detective. Scientific American, 31(11):160–161, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
160c.pdf.

Anonymous:1874:ETT

[6160] Anonymous. The Engineer’s trial trip of the Pacific mail steamer city of peking. Scientific American, 31(11):161, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n11/pdf/scientificamerican09121874-161.pdf.

Anonymous:1874:DSG

[6161] Anonymous. Diagram showing the general arrangement of the machinery and coal bunkers of the city of piking. Scientific American, 31(11):161, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
161a.pdf.

Anonymous:1874:GPS

[6162] Anonymous. The grape phylloxera — sixty thousand dollars reward offered for a Remedy for the grape disease. Scientific American, 31 (11):162–163, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
162.pdf.

Anonymous:1874:IVP

[6163] Anonymous. Improved vertical planing machine. Scientific American, 31(11):163, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
163.pdf.

Rose:1874:PMf

[6164] Joshua Rose. Practical mechanism. Scientific American, 31(11): 164–165, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
164.pdf.

Anonymous:1874:IGR

[6165] Anonymous. Improved gas regulator. Scientific American, 31 (11):166, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
166.pdf.

Anonymous:1874:YIS

[6166] Anonymous. Youngman’s improved slide valve. Scientific American, 31(11):166, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
166a.pdf.

Anonymous:1874:IRS

[6167] Anonymous. India rubber shoes for horses. Scientific American, 31(11):166, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
166b.pdf.

Anonymous:1874:BSD

[6168] Anonymous. A balloon steering device. Scientific American, 31 (11):167–168, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
167.pdf.

Anonymous:1874:PAAc

[6169] Anonymous. Proceedings of the American Association for the advancement of science. Scientific American, 31(11):168, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n11/pdf/scientificamerican09121874-168.pdf.

Anonymous:1874:NTT

[6170] Anonymous. New treatment for throat and nose diseases. Scientific American, 31(11):168, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
168a.pdf.

Anonymous:1874:VRT

[6171] Anonymous. Ventilation of railway tunnels. Scientific American, 31(11):169, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
169.pdf.

Anonymous:1874:HSIf

[6172] Anonymous. How shall I introduce my invention? Scientific American, 31(11):169, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
169a.pdf.

Anonymous:1874:NBPb

[6173] Anonymous. New books and publications. Scientific American, 31 (11):169, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
169b.pdf.

Anonymous:1874:RAFbi

[6174] Anonymous. Recent American and foreign patents. Scientific American, 31(11):170, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
170.pdf.

Anonymous:1874:ACbi

[6175] Anonymous. Answers to correspondents. Scientific American, 31 (11):171–172, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
171a.pdf.

Anonymous:1874:CRbj

[6176] Anonymous. Communications received. Scientific American, 31 (11):172, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
172.pdf.

Anonymous:1874:SPF

[6177] Anonymous. Schedule of patent fees. Scientific American, 31 (11):172, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
172b.pdf.

Anonymous:1874:CPbj

[6178] Anonymous. Canadian patents. Scientific American, 31(11): 172, September 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n11/pdf/scientificamerican09121874-
172c.pdf.

Anonymous:1874:Cbn

[6179] Anonymous. Correspondence. Scientific American, 31(12):180, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n12/pdf/scientificamerican09191874-180.pdf.

Anonymous:1874:BPbj

[6180] Anonymous. Business and personal. Scientific American, 31 (12):187, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
187.pdf.

Anonymous:1874:IIbk

[6181] Anonymous. Index of inventions. Scientific American, 31(12): 188, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
188a.pdf.

Anonymous:1874:PNPv

[6182] Anonymous. Patent notes: Patents. Scientific American, 31 (12):190, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
190.pdf.

Anonymous:1874:CII

[6183] Anonymous. Cleveland Iron Industry and its founders. Scientific American, 31(12):175, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
175.pdf.

Anonymous:1874:PTAa

[6184] Anonymous. Professor Tyndall’s address before the British Association. Scientific American, 31(12):176, September 19, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n12/pdf/scientificamerican09191874-176.pdf.

Anonymous:1874:WB

[6185] Anonymous. A woman without bones. Scientific American, 31 (12):176, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
176a.pdf.

Anonymous:1874:LB

[6186] Anonymous. Living barometers. Scientific American, 31(12): 176, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
176b.pdf.

Anonymous:1874:PCe

[6187] Anonymous. Plumbers’ carelessness. Scientific American, 31(12): 176–177, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
176c.pdf.

Anonymous:1874:QW

[6188] Anonymous. Quick as wink. Scientific American, 31(12):177, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n12/pdf/scientificamerican09191874-177.pdf.

Anonymous:1874:WMA

[6189] Anonymous. What makes the apples rot? Scientific American, 31(12):177, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
177a.pdf.

Anonymous:1874:FFP

[6190] Anonymous. Fat in forage plants. Scientific American, 31(12): 177, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
177b.pdf.

Anonymous:1874:SPIbb

[6191] Anonymous. Scientific and practical information. Scientific American, 31(12):177, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
177c.pdf.

Anonymous:1874:NIU

[6192] Anonymous. New investigations upon scorpions and their venom. Scientific American, 31(12):178, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
178.pdf.

Anonymous:1874:HCC

[6193] Anonymous. A hand coal cutter. Scientific American, 31(12): 178, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
178a.pdf.

Anonymous:1874:YHS

[6194] Anonymous. The Yellowstone Hot Springs. Scientific American, 31 (12):179–180, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
179.pdf.

Anonymous:1874:DUSa

[6195] Anonymous. Dentistry in the United States. Scientific American, 31(12):181, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
181.pdf.

Anonymous:1874:IWWb

[6196] Anonymous. Improved water wheel governor. Scientific American, 31(12):182, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
182.pdf.

Anonymous:1874:DLT

[6197] Anonymous. Disaster in launching a turret ship on the Thames. Scientific American, 31(12):183, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
183.pdf.

Anonymous:1874:BSC

[6198] Anonymous. Burnishing surface colored papers. Scientific American, 31(12):183–185, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
183a.pdf.

Anonymous:1874:DCbc

[6199] Anonymous. Decisions of the courts. Scientific American, 31 (12):185, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
185.pdf.

Anonymous:1874:RAFbj

[6200] Anonymous. Recent American and foreign patents. Scientific American, 31(12):185–186, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
185a.pdf.

Anonymous:1874:ACbj

[6201] Anonymous. Answers to correspondents. Scientific American, 31 (12):187–188, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
187a.pdf.

Anonymous:1874:CRbk

[6202] Anonymous. Communications received. Scientific American, 31 (12):188, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
188.pdf.

Anonymous:1874:CPbk

[6203] Anonymous. Canadian patents. Scientific American, 31(12): 188, September 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n12/pdf/scientificamerican09191874-
188b.pdf.

Anonymous:1874:Cbp

[6204] Anonymous. Correspondence. Scientific American, 31(13):196, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n13/pdf/scientificamerican09261874-196.pdf.

Anonymous:1874:BNBba

[6205] Anonymous. Books: New books and publications. Scientific American, 31(13):200, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
200c.pdf.

Anonymous:1874:BPbk

[6206] Anonymous. Business and personal. Scientific American, 31 (13):202, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
202.pdf.

Anonymous:1874:IIbl

[6207] Anonymous. Index of inventions. Scientific American, 31(13): 204, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
204a.pdf.

Anonymous:1874:PNPw

[6208] Anonymous. Patent notes: Patents. Scientific American, 31 (13):206, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
206.pdf.

Anonymous:1874:ASP

[6209] Anonymous. The American stove plate dressing machine. Scientific American, 31(13):191, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
191.pdf.

Anonymous:1874:KRT

[6210] Anonymous. Kinzel’s railway track lifter. Scientific American, 31 (13):191, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
191a.pdf.

Anonymous:1874:AAc

[6211] Anonymous. Animals are automata. Scientific American, 31 (13):192, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
192.pdf.

Anonymous:1874:PTAb

[6212] Anonymous. Professor Tyndall’s address. Scientific American, 31 (13):192, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
192a.pdf.

Anonymous:1874:V

[6213] Anonymous. Vivisection. Scientific American, 31(13):192–193, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n13/pdf/scientificamerican09261874-192b.pdf.

Anonymous:1874:ABL

[6214] Anonymous. American beacon lights. Scientific American, 31 (13):193, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
193.pdf.

Anonymous:1874:AI

[6215] Anonymous. Alaska and the Icelanders. Scientific American, 31 (13):193, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
193a.pdf.

Anonymous:1874:QR

[6216] Anonymous. Queer rains. Scientific American, 31(13):193, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n13/pdf/scientificamerican09261874-193b.pdf.

Anonymous:1874:FAIa

[6217] Anonymous. The fair of the American institute. Scientific American, 31(13):193, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
193c.pdf.

Anonymous:1874:IACa

[6218] Anonymous. Improved adjustable clamp. Scientific American, 31 (13):194, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
194.pdf.

Anonymous:1874:FTC

[6219] Anonymous. Furnaces, their construction and management. Scientific American, 31(13):194, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
194a.pdf.

Anonymous:1874:SFH

[6220] Anonymous. Starling’s ferrule and hook for whiffletrees. Scientific American, 31(13):194, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
194b.pdf.

Anonymous:1874:IR

[6221] Anonymous.  Improved rudder.  Scientific American, 31(13):

195, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
195.pdf.

Anonymous:1874:L

[6222] Anonymous.  On lifeboats.  Scientific American, 31(13):195–

196, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
195a.pdf.

Rose:1874:PMg

[6223] Joshua Rose. Practical mechanism. Scientific American, 31(13): 197–198, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
197.pdf.

Anonymous:1874:LIL

[6224] Anonymous. Lordon’s improved lantern. Scientific American, 31 (13):198, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
198.pdf.

Anonymous:1874:FHb

[6225] Anonymous. The fire on the hearth. Scientific American, 31 (13):198, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
198a.pdf.

Anonymous:1874:RWJ

[6226] Anonymous. Rowland’s wagon jack. Scientific American, 31 (13):198, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
198b.pdf.

Anonymous:1874:IBP

[6227] Anonymous. Improved brick; pressing machine. Scientific American, 31(13):199, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
199.pdf.

Anonymous:1874:NEB

[6228] Anonymous. Novel engine and boiler. Scientific American, 31 (13):199, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
199a.pdf.

Anonymous:1874:DUSb

[6229] Anonymous. Dentistry in the United States. Scientific American, 31(13):200, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
200.pdf.

Anonymous:1874:MNg

[6230] Anonymous. Medical notes. Scientific American, 31(13):200, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n13/pdf/scientificamerican09261874-200a.pdf.

Anonymous:1874:DCbd

[6231] Anonymous. Decisions of the courts. Scientific American, 31 (13):200, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
200b.pdf.

Anonymous:1874:RAFbk

[6232] Anonymous. Recent American and foreign patents. Scientific American, 31(13):201, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
201.pdf.

Anonymous:1874:ACbk

[6233] Anonymous. Answers to correspondents. Scientific American, 31 (13):202–204, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
202a.pdf.

Anonymous:1874:CRbl

[6234] Anonymous. Communications received. Scientific American, 31 (13):204, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
204.pdf.

Anonymous:1874:CPbl

[6235] Anonymous. Canadian patents. Scientific American, 31(13): 204, September 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n13/pdf/scientificamerican09261874-
204b.pdf.

Anonymous:1874:Cbq

[6236] Anonymous. Correspondence. Scientific American, 31(14):212–213, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-212a.pdf.

Anonymous:1874:BPbl

[6237] Anonymous. Business and personal. Scientific American, 31(14):218, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-218.pdf.

Anonymous:1874:IIbm

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-220a.pdf.

Anonymous:1874:PNPx

[6239] Anonymous. Patent notes: Patents. Scientific American, 31(14):222, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-222.pdf.

Anonymous:1874:IMTb

[6240] Anonymous. Improved machine for turning carriage axles. Scientific American, 31(14):207, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
207.pdf.

Anonymous:1874:WIPa

[6241] Anonymous. Whitford’s improved potato coverer. Scientific American, 31(14):207, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
207a.pdf.

Anonymous:1874:CFS

[6242] Anonymous. Candle flames and streaks of cloud. Scientific American, 31(14):208, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
208.pdf.

Anonymous:1874:RME

[6243] Anonymous. The relation of mechanical engineering to industrial operations. Scientific American, 31(14):208, October 3, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n14/pdf/scientificamerican10031874-208a.pdf.

Anonymous:1874:COG

[6244] Anonymous. Co-operation in great Britain. Scientific American, 31 (14):208–209, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
208b.pdf.

Anonymous:1874:HAB

[6245] Anonymous. Hints to architects and builders. Scientific American, 31(14):209, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
209.pdf.

Anonymous:1874:PBc

[6246] Anonymous. Pneumatic burial. Scientific American, 31(14):209, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-209a.pdf.

Anonymous:1874:CAM

[6247] Anonymous. Compressed air machinery. Scientific American, 31 (14):209, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
209b.pdf.

Anonymous:1874:HCE

[6248] Anonymous. Horizontal compound engine. Scientific American, 31(14):211, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
211.pdf.

Anonymous:1874:SRT

[6249] Anonymous. Solar radiation thermometer. Scientific American, 31(14):211, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
211a.pdf.

Anonymous:1874:DUSc

[6250] Anonymous. Dentistry: In the United States. Scientific American, 31(14):212, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
212.pdf.

Anonymous:1874:ISCb

[6251] Anonymous. Improved soap crutching machine. Scientific American, 31(14):214, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
214.pdf.

Anonymous:1874:DIF

[6252] Anonymous. Davidson’s improved floor clamp. Scientific American, 31(14):214, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
214a.pdf.

Anonymous:1874:NG

[6253] Anonymous. A new grapnel. Scientific American, 31(14):214, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-214b.pdf.

Anonymous:1874:DBL

[6254] Anonymous. Double bogie locomotive. Scientific American, 31(14):215, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-215.pdf.

Anonymous:1874:ARSa

[6255] Anonymous. Apparatus for raising sunken vessels. Scientific American, 31(14):215, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
215a.pdf.

Anonymous:1874:HMb

[6256] Anonymous. Hydraulic mining. Scientific American, 31(14):216, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-216.pdf.

Anonymous:1874:PBAa

[6257] Anonymous. The Proceedings of the British Association for the Advancement of Science. Scientific American, 31(14):216, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-216a.pdf.

Anonymous:1874:DCbe

[6258] Anonymous. Decisions of the courts. Scientific American, 31(14):216, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-216b.pdf.

Anonymous:1874:RAFbl

[6259] Anonymous. Recent American and foreign patents. Scientific American, 31(14):217, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
217.pdf.

Anonymous:1874:ACbl

[6260] Anonymous. Answers to correspondents. Scientific American, 31 (14):218–220, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
218a.pdf.

Anonymous:1874:CRbm

[6261] Anonymous. Communications received. Scientific American, 31 (14):220, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n14/pdf/scientificamerican10031874-
220.pdf.

Anonymous:1874:CPbm

[6262] Anonymous. Canadian patents. Scientific American, 31(14):220, October 3, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n14/pdf/scientificamerican10031874-220b.pdf.

Anonymous:1874:Cbr

[6263] Anonymous. Correspondence. Scientific American, 31(15):228, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-228a.pdf.

Anonymous:1874:BPbm

[6264] Anonymous. Business and personal. Scientific American, 31 (15):235, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
235.pdf.

Anonymous:1874:IIbn

[6265] Anonymous. Index of inventions. Scientific American, 31(15):236, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-236.pdf.

Anonymous:1874:PNPy

[6266] Anonymous. Patent notes: Patents. Scientific American, 31 (15):237, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
237.pdf.

Anonymous:1874:ICWb

[6267] Anonymous. Improved current water wheel. Scientific American, 31(15):223, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
223.pdf.

Anonymous:1874:STCa

[6268] Anonymous. The scientific treatment of criminals. Scientific American, 31(15):224, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
224.pdf.

Anonymous:1874:HTM

[6269] Anonymous. Huxley’s theory of man. Scientific American, 31 (15):224–225, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
224a.pdf.

Anonymous:1874:AFF

[6270] Anonymous. Alcohol, food, and force. Scientific American, 31 (15):225, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
225.pdf.

Anonymous:1874:BT

[6271] Anonymous. Bulls on the track. Scientific American, 31(15):225, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-225a.pdf.

Anonymous:1874:FLH

[6272] Anonymous. Furnace for the large hammer at Woolwich England. Scientific American, 31(15):226, October 10, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n15/pdf/scientificamerican10101874-226.pdf.

Anonymous:1874:ARSb

[6273] Anonymous. American railroad stations. Scientific American, 31 (15):227, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
227.pdf.

Anonymous:1874:WPI

[6274] Anonymous. The weak points in iron castings. Scientific American, 31(15):227–228, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
227a.pdf.

Anonymous:1874:GRL

[6275] Anonymous. Geological records of life. Scientific American, 31 (15):228, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
228.pdf.

Rose:1874:PMh

[6276] Joshua Rose. Practical mechanism. Scientific American, 31(15):229, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-229.pdf.

Anonymous:1874:IPBb

[6277] Anonymous. Improved pipe and bolt threader, cutter, and nut tapping machine. Scientific American, 31(15):230, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-230.pdf.

Anonymous:1874:JHT

[6278] Anonymous. Jones’ heel trimming machine. Scientific American, 31(15):230, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
230a.pdf.

Anonymous:1874:NYK

[6279] Anonymous. New yacht for the Khedive of Egypt. Scientific American, 31(15):231, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
231.pdf.

Anonymous:1874:DUSd

[6280] Anonymous. Dentistry in the United States. Scientific American, 31(15):232, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
232.pdf.

Anonymous:1874:FAIb

[6281] Anonymous. The fair of the American institute. Scientific American, 31(15):233, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
233.pdf.

Anonymous:1874:MNh

[6282] Anonymous. Medical notes. Scientific American, 31(15):233, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-233a.pdf.

Anonymous:1874:DCbf

[6283] Anonymous. Decisions of the courts. Scientific American, 31 (15):233, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
233b.pdf.

Anonymous:1874:RAFbm

[6284] Anonymous. Recent American and foreign patents. Scientific American, 31(15):234, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
234.pdf.

Anonymous:1874:ACbm

[6285] Anonymous. Answers to correspondents. Scientific American, 31 (15):235, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
235a.pdf.

Anonymous:1874:CRbn

[6286] Anonymous. Communications received. Scientific American, 31 (15):235, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n15/pdf/scientificamerican10101874-
235b.pdf.

Anonymous:1874:CPbn

[6287] Anonymous. Canadian patents. Scientific American, 31(15):236, October 10, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n15/pdf/scientificamerican10101874-236a.pdf.

Anonymous:1874:Cbs

[6288] Anonymous. Correspondence. Scientific American, 31(16):244, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-244.pdf.

Anonymous:1874:BNBbb

[6289] Anonymous. Books: New books and publications. Scientific American, 31(16):248, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
248c.pdf.

Anonymous:1874:BPbn

[6290] Anonymous. Business and personal. Scientific American, 31 (16):250, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
250.pdf.

Anonymous:1874:IIbo

[6291] Anonymous. Index of inventions. Scientific American, 31(16):252, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-252.pdf.

Anonymous:1874:PNPz

[6292] Anonymous. Patent notes: Patents. Scientific American, 31 (16):254, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
254.pdf.

Anonymous:1874:MMM

[6293] Anonymous. Machine for molding metal castings. Scientific American, 31(16):239–240, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
239.pdf.

Anonymous:1874:OFN

[6294] Anonymous. Our fever nests and their Remedy. Scientific American, 31(16):240, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
240.pdf.

Anonymous:1874:MSP

[6295] Anonymous. Measurement of a screw propeller. Scientific American, 31(16):240–241, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
240a.pdf.

Anonymous:1874:CEc

[6296] Anonymous. Crime epidemics. Scientific American, 31(16):241, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-241.pdf.

Anonymous:1874:CPbo

[6297] Anonymous. The Commissionership of Patents. Scientific American, 31(16):241, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
241a.pdf.

Anonymous:1874:IPCb

[6298] Anonymous. Influence of the price of coal on ship building. Scientific American, 31(16):241, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
241b.pdf.

Anonymous:1874:PPb

[6299] Anonymous. Peculiar people. Scientific American, 31(16):241, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-241c.pdf.

Anonymous:1874:ACMa

[6300] Anonymous. The automatic cow milker. Scientific American, 31 (16):242, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
242.pdf.

Anonymous:1874:EEP

[6301] Anonymous. Enameled and embossed photographs. Scientific American, 31(16):242, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
242a.pdf.

Anonymous:1874:WIPb

[6302] Anonymous. Worswick’s improved pump mechanism. Scientific American, 31(16):242, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
242b.pdf.

Anonymous:1874:MNEa

[6303] Anonymous. Muller’s non-explosive self-lighter. Scientific American, 31(16):242, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
242c.pdf.

Anonymous:1874:SCT

[6304] Anonymous. Shoemakers’ combination tool. Scientific American, 31(16):242, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
242d.pdf.

Anonymous:1874:NSP

[6305] Anonymous. New steam plows. Scientific American, 31(16): 243–244, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
243.pdf.

Anonymous:1874:CBLa

[6306] Anonymous. Coal burning locomotives in the South. Scientific American, 31(16):244, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
244a.pdf.

Anonymous:1874:DUSe

[6307] Anonymous. Dentistry in the United States. Scientific American, 31(16):245, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
245.pdf.

Anonymous:1874:AAT

[6308] Anonymous. Arsenic in agricultural and technical products. Scientific American, 31(16):245, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
245a.pdf.

Anonymous:1874:ISBa

[6309] Anonymous. Improved steam boiler. Scientific American, 31 (16):246, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
246.pdf.

Anonymous:1874:GBC

[6310] Anonymous. The Guernsey breed of cattle. Scientific American, 31(16):247, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
247.pdf.

Anonymous:1874:PBAb

[6311] Anonymous. The Proceedings of the British Association for the Advancement of Science. Scientific American, 31(16):247, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-247a.pdf.

Anonymous:1874:FAIc

[6312] Anonymous. The fair of the American institute. Scientific American, 31(16):248, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
248.pdf.

Anonymous:1874:IRC

[6313] Anonymous. The international rifle contest. Scientific American, 31(16):248, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
248a.pdf.

Anonymous:1874:DCbg

[6314] Anonymous. Decisions of the courts. Scientific American, 31 (16):248, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
248b.pdf.

Anonymous:1874:RAFbn

[6315] Anonymous. Recent American and foreign patents. Scientific American, 31(16):249, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
249.pdf.

Anonymous:1874:ACbn

[6316] Anonymous. Answers to correspondents. Scientific American, 31 (16):250–251, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
250a.pdf.

Anonymous:1874:CRbo

[6317] Anonymous. Communications received. Scientific American, 31 (16):251–252, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n16/pdf/scientificamerican10171874-
251.pdf.

Anonymous:1874:CPbp

[6318] Anonymous. Canadian patents. Scientific American, 31(16):252, October 17, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n16/pdf/scientificamerican10171874-252a.pdf.

Anonymous:1874:Cbt

[6319] Anonymous. Correspondence. Scientific American, 31(17):260, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n17/pdf/scientificamerican10241874-260.pdf.

Anonymous:1874:BPbo

[6320] Anonymous. Business and personal. Scientific American, 31 (17):266, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
266.pdf.

Anonymous:1874:IIbp

[6321] Anonymous. Index of inventions. Scientific American, 31(17): 267–268, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
267a.pdf.

Anonymous:1874:PNPba

[6322] Anonymous. Patent notes: Patents. Scientific American, 31 (17):268, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
268a.pdf.

Anonymous:1874:CIF

[6323] Anonymous. Combination ice and freezing house. Scientific American, 31(17):255, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
255.pdf.

Anonymous:1874:GAC

[6324] Anonymous. Gain from the application of condensers to steam engines. Scientific American, 31(17):256, October 24, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n17/pdf/scientificamerican10241874-256.pdf.

Anonymous:1874:IIbq

[6325] Anonymous. Impracticable inventors. Scientific American, 31 (17):256, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
256a.pdf.

Anonymous:1874:RET

[6326] Anonymous. Railroad employees and their pay. Scientific American, 31(17):256–257, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
256b.pdf.

Anonymous:1874:IASb

[6327] Anonymous. Improvised apparatus for steam boiler trials. Scientific American, 31(17):257, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
257.pdf.

Anonymous:1874:SPIbc

[6328] Anonymous. Scientific and practical information. Scientific American, 31(17):257, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
257a.pdf.

Anonymous:1874:NPE

[6329] Anonymous. New portable engine. Scientific American, 31(17):258, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n17/pdf/scientificamerican10241874-258.pdf.

Anonymous:1874:ICTb

[6330] Anonymous. Improvement in cable telegraphy. Scientific American, 31(17):258, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
258a.pdf.

Anonymous:1874:NIE

[6331] Anonymous. New instrument for estimating urea. Scientific American, 31(17):258, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
258b.pdf.

Anonymous:1874:POT

[6332] Anonymous. The Pacific Ocean telegraph. Scientific American, 31(17):259, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
259.pdf.

Anonymous:1874:HR

[6333] Anonymous. The hydraulic ram. Scientific American, 31(17): 259–260, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
259a.pdf.

Rose:1874:PMi

[6334] Joshua Rose. Practical mechanism. Scientific American, 31(17):261, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n17/pdf/scientificamerican10241874-261.pdf.

Anonymous:1874:IASc

[6335] Anonymous. Improved air Spring forge hammer. Scientific American, 31(17):262, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
262.pdf.

Anonymous:1874:NBCb

[6336] Anonymous. A new boiler covering. Scientific American, 31 (17):262, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
262a.pdf.

Anonymous:1874:DKC

[6337] Anonymous. Dough kneader and cutter. Scientific American, 31 (17):262, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
262b.pdf.

Anonymous:1874:CVR

[6338] Anonymous. The cultivation of the Victoria Regia. Scientific American, 31(17):263, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
263.pdf.

Anonymous:1874:LMb

[6339] Anonymous. Long moss. Scientific American, 31(17):263, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n17/pdf/scientificamerican10241874-263a.pdf.

Anonymous:1874:EHWa

[6340] Anonymous. The early history of wheeled vehicles and railways. Scientific American, 31(17):264, October 24, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n17/pdf/scientificamerican10241874-264.pdf.

Anonymous:1874:RAFbo

[6341] Anonymous. Recent American and foreign patents. Scientific American, 31(17):264–265, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
264a.pdf.

Anonymous:1874:ACbo

[6342] Anonymous. Answers to correspondents. Scientific American, 31 (17):266–267, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
266a.pdf.

Anonymous:1874:CRbp

[6343] Anonymous. Communications received. Scientific American, 31 (17):267, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n17/pdf/scientificamerican10241874-
267.pdf.

Anonymous:1874:CPbq

[6344] Anonymous. Canadian patents. Scientific American, 31(17):268, October 24, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n17/pdf/scientificamerican10241874-268.pdf.

Anonymous:1874:Cbu

[6345] Anonymous. Correspondence. Scientific American, 31(18):276–277, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n18/pdf/scientificamerican10311874-276a.pdf.

Anonymous:1874:BNBbc

[6346] Anonymous. Books: New books and publications. Scientific American, 31(18):280, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
280a.pdf.

Anonymous:1874:BPbp

[6347] Anonymous. Business and personal. Scientific American, 31 (18):282, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
282.pdf.

Anonymous:1874:IIbr

[6348] Anonymous. Index of inventions. Scientific American, 31(18):284, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n18/pdf/scientificamerican10311874-284a.pdf.

Anonymous:1874:TCN

[6349] Anonymous. The tanite Company’s new milling machine. Scientific American, 31(18):271, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
271.pdf.

Anonymous:1874:PPO

[6350] Anonymous. A panic at the Patent Office. Scientific American, 31(18):272, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
272.pdf.

Anonymous:1874:EEb

[6351] Anonymous. Economy in eating. Scientific American, 31(18): 272–273, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
272a.pdf.

Anonymous:1874:DP

[6352] Anonymous. Demoniacal possessions. Scientific American, 31 (18):273, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
273.pdf.

Anonymous:1874:RSEb

[6353] Anonymous. Reports on small engines. Scientific American, 31 (18):273, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
273a.pdf.

Anonymous:1874:IDR

[6354] Anonymous. Improved Draftsman’s rule. Scientific American, 31 (18):274, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
274.pdf.

Anonymous:1874:SPLa

[6355] Anonymous. A siphon for poisonous liquids and acids. Scientific American, 31(18):274, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
274a.pdf.

Anonymous:1874:FPB

[6356] Anonymous. The Fairmount Park bear pits. Scientific American, 31(18):275, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
275.pdf.

Anonymous:1874:MCC

[6357] Anonymous. Milk cooling can. Scientific American, 31(18):275, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n18/pdf/scientificamerican10311874-275a.pdf.

Anonymous:1874:LG

[6358] Anonymous. Lights for greenhouses. Scientific American, 31(18):275– 276, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1874:SWS

[6359] Anonymous. Spiders’ Webs and spinnerets. Scientific American, 31(18):276, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
276.pdf.

Anonymous:1874:SNG

[6360] Anonymous. Some new galvanic batteries. Scientific American, 31(18):277, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
277.pdf.

Anonymous:1874:ISL

[6361] Anonymous. Improved self-locking padlock. Scientific American, 31(18):278, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
278.pdf.

Anonymous:1874:IDb

[6362] Anonymous. Improved disintegrator. Scientific American, 31 (18):278, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
278a.pdf.

Anonymous:1874:ECS

[6363] Anonymous. The English channel steamers. Scientific American, 31(18):279, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
279.pdf.

Anonymous:1874:EHWb

[6364] Anonymous. The early history of wheeled vehicles and railways. Scientific American, 31(18):279–280, October 31, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n18/pdf/scientificamerican10311874-279a.pdf.

Anonymous:1874:DCbh

[6365] Anonymous. Decisions of the courts. Scientific American, 31 (18):280, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
280.pdf.

Anonymous:1874:RAFbp

[6366] Anonymous. Recent American and foreign patents. Scientific American, 31(18):281, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
281.pdf.

Anonymous:1874:ACbp

[6367] Anonymous. Answers to correspondents. Scientific American, 31 (18):282–283, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
282a.pdf.

Anonymous:1874:CRbq

[6368] Anonymous. Communications received. Scientific American, 31 (18):284, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n18/pdf/scientificamerican10311874-
284.pdf.

Anonymous:1874:CPbr

[6369] Anonymous. Canadian patents. Scientific American, 31(18):284, October 31, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n18/pdf/scientificamerican10311874-284b.pdf.

Anonymous:1874:Cbv

[6370] Anonymous. Correspondence. Scientific American, 31(19):292, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n19/pdf/scientificamerican11071874-292a.pdf.

Anonymous:1874:BPbq

[6371] Anonymous. Business and personal. Scientific American, 31 (19):298, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
298.pdf.

Anonymous:1874:IIbs

[6372] Anonymous. Index of inventions. Scientific American, 31(19): 300, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
300.pdf.

Anonymous:1874:AFG

[6373] Anonymous. Automatic fire governor for gas works. Scientific American, 31(19):287, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
287.pdf.

Anonymous:1874:ICSb

[6374] Anonymous. Improved coal screen. Scientific American, 31(19): 287–288, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
287a.pdf.

Anonymous:1874:GUF

[6375] Anonymous. Gain from the use of feed water heaters. Scientific American, 31(19):288, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
288.pdf.

Anonymous:1874:SESb

[6376] Anonymous. A special edition of the Scientific American — one hundred thousand copies. Scientific American, 31(19):288, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n19/pdf/scientificamerican11071874-288a.pdf.

Anonymous:1874:CTb

[6377] Anonymous. Cost of tunnels. Scientific American, 31(19):288– 289, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
288b.pdf.

Anonymous:1874:EWN

[6378] Anonymous. Explosive wounds from non-explosive bullets. Scientific American, 31(19):289, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
289.pdf.

Anonymous:1874:WTK

[6379] Anonymous. What temperature kills? Scientific American, 31 (19):289, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
289a.pdf.

Anonymous:1874:SPIbd

[6380] Anonymous. Scientific and practical information. Scientific American, 31(19):289, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
289b.pdf.

Anonymous:1874:PTT

[6381] Anonymous. Peas three thousand years old. Scientific American, 31(19):290, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
290.pdf.

Anonymous:1874:ASb

[6382] Anonymous. The aerial screw. Scientific American, 31(19):290, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n19/pdf/scientificamerican11071874-290a.pdf.

Anonymous:1874:MGP

[6383] Anonymous. M. De Groof’s parachute. Scientific American, 31 (19):290, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
290b.pdf.

Anonymous:1874:RTb

[6384] Anonymous. The ruins of troy. Scientific American, 31(19):291, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n19/pdf/scientificamerican11071874-291.pdf.

Anonymous:1874:UAB

[6385] Anonymous. Universal angle bricks. Scientific American, 31 (19):291, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
291a.pdf.

Anonymous:1874:ANi

[6386] Anonymous. Astronomical notes. Scientific American, 31(19): 292, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
292.pdf.

Rose:1874:PMj

[6387] Joshua Rose. Practical mechanism. Scientific American, 31(19): 293, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
293.pdf.

Anonymous:1874:ISSb

[6388] Anonymous. Improved stump sawing machine. Scientific American, 31(19):294, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
294.pdf.

Anonymous:1874:MILb

[6389] Anonymous. Moat’s improved lock. Scientific American, 31(19): 294, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
294a.pdf.

Anonymous:1874:SAWb

[6390] Anonymous. The Southport Aquarium and Winter Garden. Scientific American, 31(19):295, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
295.pdf.

Rath:1874:SSR

[6391] G. Von Rath. Sulphur in Sicily, and its reduction from the ore. Scientific American, 31(19):295–296, November 7, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n19/pdf/scientificamerican11071874-295a.pdf.

Anonymous:1874:FAId

[6392] Anonymous. The fair of the American institute. Scientific American, 31(19):296, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
296.pdf.

Anonymous:1874:RAFbq

[6393] Anonymous. Recent American and foreign patents. Scientific American, 31(19):297, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
297.pdf.

Anonymous:1874:ACbq

[6394] Anonymous. Answers to correspondents. Scientific American, 31 (19):298–299, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
298a.pdf.

Anonymous:1874:CRbr

[6395] Anonymous. Communications received. Scientific American, 31 (19):299, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
299.pdf.

Anonymous:1874:CPbs

[6396] Anonymous. Canadian patents. Scientific American, 31(19): 300, November 7, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n19/pdf/scientificamerican11071874-
300a.pdf.

Anonymous:1874:Cbw

[6397] Anonymous. Correspondence. Scientific American, 31(20):308– 309, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
308.pdf.

Anonymous:1874:BPbr

[6398] Anonymous. Business and personal. Scientific American, 31 (20):314, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
314.pdf.

Anonymous:1874:IIbt

[6399] Anonymous. Index of inventions. Scientific American, 31(20): 315–316, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
315a.pdf.

Anonymous:1874:MRE

[6400] Anonymous. The Myers rotary engine. Scientific American, 31 (20):303–304, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
303.pdf.

Anonymous:1874:RTC

[6401] Anonymous. Rapid transit in the city of New York. Scientific American, 31(20):304, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
304.pdf.

Anonymous:1874:HICb

[6402] Anonymous. Hints to inventors and capitalists. Scientific American, 31(20):304–305, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
304a.pdf.

Anonymous:1874:CDc

[6403] Anonymous. Chloroform dangers. Scientific American, 31(20): 305, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
305.pdf.

Anonymous:1874:MMd

[6404] Anonymous. Masters and men. Scientific American, 31(20): 305, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
305a.pdf.

Anonymous:1874:DSH

[6405] Anonymous. The development of subterranean heat. Scientific American, 31(20):305, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
305b.pdf.

Anonymous:1874:NPB

[6406] Anonymous. Novel plan for building submarine structures. Scientific American, 31(20):306, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
306.pdf.

Anonymous:1874:CPbt

[6407] Anonymous. Carrier pigeons. Scientific American, 31(20):306, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n20/pdf/scientificamerican11141874-306a.pdf.

Anonymous:1874:URNa

[6408] Anonymous. The underground railway, New York City. Scientific American, 31(20):307–308, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
307.pdf.

Anonymous:1874:Cbx

[6409] Anonymous. Crucibles. Scientific American, 31(20):309, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n20/pdf/scientificamerican11141874-309.pdf.

Anonymous:1874:ISBb

[6410] Anonymous. Improved sash boring and grooving machine. Scientific American, 31(20):310, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
310.pdf.

Anonymous:1874:ISBc

[6411] Anonymous. Improved surface blow-off. Scientific American, 31 (20):310, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
310a.pdf.

Anonymous:1874:BFE

[6412] Anonymous. Barrow-in-furness (england) and its founder. Scientific American, 31(20):311–312, November 14, 1874. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
311.pdf.

Anonymous:1874:FIEa

[6413] Anonymous. The Franklin Institute Exhibition. Scientific American, 31(20):312, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
312.pdf.

Anonymous:1874:RAFbr

[6414] Anonymous. Recent American and foreign patents. Scientific American, 31(20):313, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
313.pdf.

Anonymous:1874:ACbr

[6415] Anonymous. Answers to correspondents. Scientific American, 31 (20):314–315, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
314a.pdf.

Anonymous:1874:CRbs

[6416] Anonymous. Communications received. Scientific American, 31 (20):315, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
315.pdf.

Anonymous:1874:CPbu

[6417] Anonymous. Canadian patents. Scientific American, 31(20): 316, November 14, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n20/pdf/scientificamerican11141874-
316.pdf.

Anonymous:1874:Cby (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
324.pdf.

Anonymous:1874:BNBbd

[6419] Anonymous. Books: New books and publications. Scientific American, 31(21):330, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
330a.pdf.

Anonymous:1874:BPbs

[6420] Anonymous. Business and personal. Scientific American, 31 (21):331, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
331.pdf.

Anonymous:1874:IIbu

[6421] Anonymous. Index of inventions. Scientific American, 31(21): 332, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
332a.pdf.

Anonymous:1874:HSPb

[6422] Anonymous. The Hamilton surface planer. Scientific American, 31(21):319, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
319.pdf.

Anonymous:1874:CWM

[6423] Anonymous. Cheap workmen make dear work. Scientific American, 31(21):320, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
320.pdf.

Anonymous:1874:RAP

[6424] Anonymous. The relation of alcohol to physical strength. Scientific American, 31(21):320, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
320a.pdf.

Anonymous:1874:LPW

[6425] Anonymous. The labor prospects for the winter. Scientific American, 31(21):320–321, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
320b.pdf.

Anonymous:1874:SPLb

[6426] Anonymous. The sensation of pain in the lower animals. Scientific American, 31(21):321, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
321.pdf.

Anonymous:1874:SPIbe

[6427] Anonymous. Scientific and practical information. Scientific American, 31(21):321, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
321a.pdf.

Anonymous:1874:AGS

[6428] Anonymous. The automatic gas saves. Scientific American, 31 (21):322, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
322.pdf.

Anonymous:1874:CTS

[6429] Anonymous. Chain towage on the St. Lawrence. Scientific American, 31(21):322, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
322a.pdf.

Anonymous:1874:URNb

[6430] Anonymous. The underground railway, New York City. Scientific American, 31(21):323, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
323.pdf.

Rose:1874:PMk

[6431] Joshua Rose. Practical mechanism. Scientific American, 31(21): 325–326, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
325.pdf.

Anonymous:1874:NDH

[6432] Anonymous. A new detachable horseshoe. Scientific American, 31 (21):326, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
326.pdf.

Anonymous:1874:WTC

[6433] Anonymous. Willahs’ three-cylinder engine. Scientific American, 31(21):326, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
326a.pdf.

Anonymous:1874:ICSc

[6434] Anonymous. Improved cork-soled boots. Scientific American, 31 (21):326, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
326b.pdf.

Anonymous:1874:BCE

[6435] Anonymous. The buildings for the Centennial Exhibition. Scientific American, 31(21):327–328, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
327.pdf.

Anonymous:1874:FIEb

[6436] Anonymous. The Franklin Institute Exhibition. Scientific American, 31(21):328–329, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
328.pdf.

Anonymous:1874:RAFbs

[6437] Anonymous. Recent American and foreign patents. Scientific American, 31(21):330, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
330.pdf.

Anonymous:1874:ACbs

[6438] Anonymous. Answers to correspondents. Scientific American, 31 (21):331, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
331a.pdf.

Anonymous:1874:CRbt

[6439] Anonymous. Communications received. Scientific American, 31 (21):332, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
332.pdf.

Anonymous:1874:CPbv

[6440] Anonymous. Canadian patents. Scientific American, 31(21): 332, November 21, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n21/pdf/scientificamerican11211874-
332b.pdf.

Anonymous:1874:Cbz

[6441] Anonymous. Correspondence. Scientific American, 31(22):340, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n22/pdf/scientificamerican11281874-340.pdf.

Anonymous:1874:BPbt

[6442] Anonymous. Business and personal. Scientific American, 31 (22):346, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
346.pdf.

Anonymous:1874:IIbv

[6443] Anonymous. Index of inventions. Scientific American, 31(22): 348, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
348.pdf.

Anonymous:1874:SEc

[6444] Anonymous. The Shapley engine. Scientific American, 31(22): 335, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
335.pdf.

Anonymous:1874:ESH

[6445] Anonymous. Exhaust steam for heating purposes. Scientific American, 31(22):336, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
336.pdf.

Anonymous:1874:PRb

[6446] Anonymous. Prison reform. Scientific American, 31(22):336, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n22/pdf/scientificamerican11281874-336a.pdf.

Anonymous:1874:PIH

[6447] Anonymous. A possible improvement in house heating. Scientific American, 31(22):336–337, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
336b.pdf.

Anonymous:1874:MAP

[6448] Anonymous. Mechanical aesthetics and practical men. Scientific American, 31(22):337, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
337.pdf.

Anonymous:1874:SPIbf

[6449] Anonymous. Scientific and practical information. Scientific American, 31(22):337, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
337a.pdf.

Anonymous:1874:URNc

[6450] Anonymous. To underground railway, New York City. Scientific American, 31(22):338–340, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
338.pdf.

Anonymous:1874:NAS

[6451] Anonymous. The National Academy of Science. Scientific American, 31(22):341, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
341.pdf.

Anonymous:1874:IPI

[6452] Anonymous. Improved pinking iron. Scientific American, 31 (22):342, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
342.pdf.

Anonymous:1874:BHJ

[6453] Anonymous. Biddle’s hydraulic jack. Scientific American, 31 (22):342, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
342a.pdf.

Anonymous:1874:NMCa

[6454] Anonymous. A novel milk can. Scientific American, 31(22):342, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n22/pdf/scientificamerican11281874-342b.pdf.

Anonymous:1874:IHb

[6455] Anonymous. Improved hatchet. Scientific American, 31(22): 342, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
342c.pdf.

Anonymous:1874:LJL

[6456] Anonymous. The late John Laird. Scientific American, 31(22): 343, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
343.pdf.

Anonymous:1874:FAIe

[6457] Anonymous. The fair of the American institute. Scientific American, 31(22):344, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
344.pdf.

Anonymous:1874:RAFbt

[6458] Anonymous. Recent American and foreign patents. Scientific American, 31(22):344–345, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
344a.pdf.

Anonymous:1874:ACbt

[6459] Anonymous. Answers to correspondents. Scientific American, 31 (22):346–347, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
346a.pdf.

Anonymous:1874:CRbu

[6460] Anonymous. Communication received. Scientific American, 31(22): 347–348, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
347.pdf.

Anonymous:1874:CPbw

[6461] Anonymous. Canadian patents. Scientific American, 31(22): 348, November 28, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n22/pdf/scientificamerican11281874-
348a.pdf.

Anonymous:1874:Ccb

[6462] Anonymous. Correspondence. Scientific American, 31(23):356–357, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-356b.pdf.

Anonymous:1874:BNBbe

[6463] Anonymous. Books: New books and publications. Scientific American, 31(23):361, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
361a.pdf.

Anonymous:1874:BPbu

[6464] Anonymous. Business and personal. Scientific American, 31 (23):362, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
362.pdf.

Anonymous:1874:IIbw

[6465] Anonymous. Index of inventions. Scientific American, 31(23):364, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-364a.pdf.

Anonymous:1874:BMC

[6466] Anonymous. Bridge on the mont cenis railway. Scientific American, 31(23):351, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
351.pdf.

Anonymous:1874:S

[6467] Anonymous. Suction. Scientific American, 31(23):352, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-352.pdf.

Anonymous:1874:OPO

[6468] Anonymous. Operations of the Patent Office. Scientific American, 31(23):352, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
352a.pdf.

Anonymous:1874:WSb

[6469] Anonymous. What is steel! Scientific American, 31(23):352, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-352b.pdf.

Anonymous:1874:RRR

[6470] Anonymous. Rusting of railway rails. Scientific American, 31 (23):352–353, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
352c.pdf.

Anonymous:1874:MF

[6471] Anonymous. The mechanic of the future. Scientific American, 31(23):353, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
353.pdf.

Anonymous:1874:CBKa

[6472] Anonymous. Characteristics of a business king. Scientific American, 31(23):353–354, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
353a.pdf.

Anonymous:1874:DPE

[6473] Anonymous. The docks at Portsmouth, England. Scientific American, 31(23):354, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
354.pdf.

Anonymous:1874:PFd

[6474] Anonymous. Peat fuel. Scientific American, 31(23):356, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-356.pdf.

Anonymous:1874:ANj

[6475] Anonymous. Astronomical notes. Scientific American, 31(23):356, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-356a.pdf.

Anonymous:1874:NFA

[6476] Anonymous. The new French Army gun. Scientific American, 31(23):358, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
358.pdf.

Anonymous:1874:BBU

[6477] Anonymous. A bridle and a blinder for unruly horses. Scientific American, 31(23):358, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
358a.pdf.

Anonymous:1874:SIR

[6478] Anonymous. Smith’s improved rein holder. Scientific American, 31(23):358, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
358b.pdf.

Anonymous:1874:IGS

[6479] Anonymous. The indoor garden and the shrubbery. Scientific American, 31(23):359–360, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
359.pdf.

Anonymous:1874:POY

[6480] Anonymous. Patent Office Yearly Report. Scientific American, 31(23):360, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
360.pdf.

Anonymous:1874:SPIbg

[6481] Anonymous. Scientific and practical information. Scientific American, 31(23):360–361, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
360a.pdf.

Anonymous:1874:RAFbu

[6482] Anonymous. Recent American and foreign patents. Scientific American, 31(23):361, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
361.pdf.

Anonymous:1874:ACbu

[6483] Anonymous. Answers to correspondents. Scientific American, 31 (23):362–363, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
362a.pdf.

Anonymous:1874:CRbv

[6484] Anonymous. Communications received. Scientific American, 31 (23):364, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n23/pdf/scientificamerican12051874-
364.pdf.

Anonymous:1874:CPbx

[6485] Anonymous. Canadian patents. Scientific American, 31(23):364, December 5, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n23/pdf/scientificamerican12051874-364b.pdf.

Anonymous:1874:Ccc

[6486] Anonymous. Correspondence. Scientific American, 31(24):372, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n24/pdf/scientificamerican12121874-372.pdf.

Anonymous:1874:BPbv

[6487] Anonymous. Business and personal. Scientific American, 31 (24):379, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
379.pdf.

Anonymous:1874:IIbx

[6488] Anonymous. Index of inventions. Scientific American, 31(24): 380, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
380a.pdf.

Anonymous:1874:KIB

[6489] Anonymous. The Keokuk (Iowa) Bridge. Scientific American, 31 (24):367, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
367.pdf.

Anonymous:1874:PO

[6490] Anonymous. The Patent Office. Scientific American, 31(24): 368, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
368.pdf.

Anonymous:1874:IAGa

[6491] Anonymous. The illumination of art galleries. Scientific American, 31(24):368, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
368a.pdf.

Anonymous:1874:SRb

[6492] Anonymous. Solvents for rubber. Scientific American, 31(24): 368, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
368b.pdf.

Anonymous:1874:PPd

[6493] Anonymous. Polychrome printing. Scientific American, 31(24): 368, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
368c.pdf.

Anonymous:1874:AOC

[6494] Anonymous. American oyster culture. Scientific American, 31(24): 368–369, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
368d.pdf.

Anonymous:1874:CIP

[6495] Anonymous. The coal and iron products of the world. Scientific American, 31(24):369, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
369.pdf.

Anonymous:1874:OPF

[6496] Anonymous. To our present and future subscribers. Scientific American, 31(24):369, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
369a.pdf.

Anonymous:1874:SPIbh

[6497] Anonymous. Scientific and practical information. Scientific American, 31(24):369, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
369b.pdf.

Anonymous:1874:WT

[6498] Anonymous. Wire tramways. Scientific American, 31(24):370, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n24/pdf/scientificamerican12121874-370.pdf.

Anonymous:1874:DPP

[6499] Anonymous. A device for preventing priming. Scientific American, 31(24):370, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
370a.pdf.

Anonymous:1874:URNd

[6500] Anonymous. The underground railway, New York City. Scientific American, 31(24):371–372, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
371.pdf.

Rose:1874:PMl

[6501] Joshua Rose. Practical mechanism. Scientific American, 31(24): 373, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
373.pdf.

Anonymous:1874:CPIa

[6502] Anonymous. The City of Pittsburgh and its industries. Scientific American, 31(24):376, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
376.pdf.

Anonymous:1874:NRT

[6503] Anonymous. New railway tunnel under the Hudson River, between New York and Jersey City. Scientific American, 31(24):376, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n24/pdf/scientificamerican12121874-376a.pdf.

Anonymous:1874:NHA

[6504] Anonymous. A new hot air balloon. Scientific American, 31 (24):377, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
377.pdf.

Anonymous:1874:SCCa

[6505] Anonymous. Safety catch for cranes. Scientific American, 31(24): 377–378, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
377a.pdf.

Anonymous:1874:RAFbv

[6506] Anonymous. Recent American and foreign patents. Scientific American, 31(24):378, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
378.pdf.

Anonymous:1874:ACbv

[6507] Anonymous. Answers to correspondents. Scientific American, 31 (24):379–380, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
379a.pdf.

Anonymous:1874:CRbw

[6508] Anonymous. Communications received. Scientific American, 31 (24):380, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
380.pdf.

Anonymous:1874:CPby

[6509] Anonymous. Canadian patents. Scientific American, 31(24): 380, December 12, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n24/pdf/scientificamerican12121874-
380b.pdf.

Anonymous:1874:Ccd

[6510] Anonymous. Correspondence. Scientific American, 31(25):388, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n25/pdf/scientificamerican12191874-388a.pdf.

Anonymous:1874:BPbw

[6511] Anonymous. Business and personal. Scientific American, 31 (25):395, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
395.pdf.

Anonymous:1874:IIby

[6512] Anonymous. Index of inventions. Scientific American, 31(25): 395–396, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
395c.pdf.

Anonymous:1874:WIS

[6513] Anonymous. The Whitworth Industrial Scholarships. Scientific American, 31(25):383, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
383.pdf.

Anonymous:1874:PCf

[6514] Anonymous. Publishers’ card. Scientific American, 31(25):384, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n25/pdf/scientificamerican12191874-384.pdf.

Anonymous:1874:POA

[6515] Anonymous. The Patent Office again. Scientific American, 31 (25):384, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
384a.pdf.

Anonymous:1874:TTE

[6516] Anonymous. Two typical experiments. Scientific American, 31 (25):384–385, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
384b.pdf.

Anonymous:1874:MMN

[6517] Anonymous. Mica mines in North Carolina. Scientific American, 31(25):385, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
385.pdf.

Anonymous:1874:TTb

[6518] Anonymous. Tyndall on typhoid. Scientific American, 31(25): 385, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
385a.pdf.

Anonymous:1874:HIS

[6519] Anonymous. How to investigate spiritualism. Scientific American, 31(25):385, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
385b.pdf.

Anonymous:1874:FFE

[6520] Anonymous. Floating fire engines. Scientific American, 31(25): 386, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
386.pdf.

Anonymous:1874:CGb

[6521] Anonymous. The crystallization of glass. Scientific American, 31 (25):386, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
386a.pdf.

Anonymous:1874:NCSb

[6522] Anonymous. New cloth-singeing machine. Scientific American, 31 (25):386, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
386b.pdf.

Anonymous:1874:NCN

[6523] Anonymous. The Normal College, New York City. Scientific American, 31(25):387, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
387.pdf.

Anonymous:1874:AEP

[6524] Anonymous. Apid estimation of phosphoric acid, magnesia, and lime. Scientific American, 31(25):388, December 19, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n25/pdf/scientificamerican12191874-388.pdf.

Anonymous:1874:PGa

[6525] Anonymous. Pendulum governors. Scientific American, 31(25): 389, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
389.pdf.

Anonymous:1874:ISEc

[6526] Anonymous. Improved Spike extractor. Scientific American, 31 (25):390, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
390.pdf.

Anonymous:1874:EPTb

[6527] Anonymous. The Ericsson pneumatic torpedo. Scientific American, 31(25):390, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
390a.pdf.

Anonymous:1874:CIT

[6528] Anonymous. Copper’s improved transplanter. Scientific American, 31(25):390, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
390b.pdf.

Anonymous:1874:Y

[6529] Anonymous. The yuccas. Scientific American, 31(25):391, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n25/pdf/scientificamerican12191874-391.pdf.

Anonymous:1874:HU

[6530] Anonymous. The Hymenocalyx Undulata. Scientific American, 31 (25):391–392, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
391a.pdf.

Young:1874:CSa

[6531] C. A. Young. The constitution of the Sun. Scientific American, 31(25):393, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
393.pdf.

Sprague:1874:EDM

[6532] J. T. Sprague. Electro-deposition of metals. Scientific American, 31 (25):393–394, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
393a.pdf.

Anonymous:1874:DCbi

[6533] Anonymous. Decisions of the courts. Scientific American, 31 (25):394, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
394.pdf.

Anonymous:1874:RAFbw

[6534] Anonymous. Recent American and foreign patents. Scientific American, 31(25):394, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
394a.pdf.

Anonymous:1874:ACbw

[6535] Anonymous. Answers to correspondents. Scientific American, 31 (25):395, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
395a.pdf.

Anonymous:1874:CRbx

[6536] Anonymous. Communications received. Scientific American, 31 (25):395, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
395b.pdf.

Anonymous:1874:CPbz

[6537] Anonymous. Canadian patents. Scientific American, 31(25): 396, December 19, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n25/pdf/scientificamerican12191874-
396.pdf.

Anonymous:1874:Ccf

[6538] Anonymous. Correspondence. Scientific American, 31(26):404, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-404b.pdf.

Anonymous:1874:BPbx

[6539] Anonymous. Business and personal. Scientific American, 31 (26):409, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
409.pdf.

Anonymous:1874:IIbz

[6540] Anonymous. Index of inventions. Scientific American, 31(26): 410, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
410a.pdf.

Anonymous:1874:Ic

[6541] Anonymous. Index. Scientific American, 31(26):410–412, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-410b.pdf.

Anonymous:1874:RBN

[6542] Anonymous. Railway bridge, near york, England. Scientific American, 31(26):399, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
399.pdf.

Anonymous:1874:LY

[6543] Anonymous. The largest yet. Scientific American, 31(26):400, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-400.pdf.

Anonymous:1874:PCg

[6544] Anonymous. Publishers’ card. Scientific American, 31(26):400, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-400a.pdf.

Anonymous:1874:ANAa

[6545] Anonymous. Another needed amendment of the patent law. Scientific American, 31(26):400, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
400b.pdf.

Anonymous:1874:GMT

[6546] Anonymous. Government monopoly of the telegraph lines. Scientific American, 31(26):400, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
400c.pdf.

Anonymous:1874:TVb

[6547] Anonymous. The Transit of Venus. Scientific American, 31(26): 400–401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
400d.pdf.

Anonymous:1874:POC

[6548] Anonymous. The Patent Office clerks. Scientific American, 31 (26):401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
401.pdf.

Anonymous:1874:ECb

[6549] Anonymous. Ezra Cornell. Scientific American, 31(26):401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-401a.pdf.

Anonymous:1874:WDI

[6550] Anonymous. Will do it after a funeral. Scientific American, 31 (26):401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
401b.pdf.

Anonymous:1874:LPE

[6551] Anonymous. The latest polar expedition. Scientific American, 31 (26):401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
401c.pdf.

Anonymous:1874:CFb

[6552] Anonymous. Cheap freights. Scientific American, 31(26):401, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-401d.pdf.

Anonymous:1874:URNe

[6553] Anonymous. The underground railway, New York City. Scientific American, 31(26):402, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
402.pdf.

Anonymous:1874:NNb

[6554] Anonymous. A new nail. Scientific American, 31(26):403, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v31/n26/pdf/scientificamerican12261874-403.pdf.

Anonymous:1874:CTc

[6555] Anonymous. Conception of the telegraph. Scientific American, 31 (26):403, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
403a.pdf.

Anonymous:1874:FPS

[6556] Anonymous. French photo skylights. Scientific American, 31 (26):403, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
403b.pdf.

Young:1874:CSb

[6557] C. A. Young. The constitution of the Sun. Scientific American, 31(26):404, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
404.pdf.

Anonymous:1874:PGb

[6558] Anonymous. Pendulum governors. Scientific American, 31(26): 404, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
404a.pdf.

Rose:1874:PMm

[6559] Joshua Rose. Practical mechanism. Scientific American, 31(26): 405, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
405.pdf.

Anonymous:1874:TFC

[6560] Anonymous. Trying-up and four-cutter planing and molding machine. Scientific American, 31(26):406, December 26, 1874. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v31/
n26/pdf/scientificamerican12261874-406.pdf.

Anonymous:1874:NMCb

[6561] Anonymous. New mechanical construction for compound tools. Scientific American, 31(26):406, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
406a.pdf.

Anonymous:1874:IFT

[6562] Anonymous. Iron-framed thrashing machines. Scientific American, 31(26):407, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
407.pdf.

Anonymous:1874:CAb

[6563] Anonymous. Cheerful for advertisers. Scientific American, 31 (26):408, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
408.pdf.

Anonymous:1874:DCbj

[6564] Anonymous. Decisions of the courts. Scientific American, 31 (26):408, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
408a.pdf.

Anonymous:1874:RAFbx

[6565] Anonymous. Recent American and foreign patents. Scientific American, 31(26):408, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
408b.pdf.

Anonymous:1874:ACbx

[6566] Anonymous. Answers to correspondents. Scientific American, 31 (26):409–410, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
409a.pdf.

Anonymous:1874:CRby

[6567] Anonymous. Communications received. Scientific American, 31 (26):410, December 26, 1874. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v31/n26/pdf/scientificamerican12261874-
410.pdf.

Anonymous:1875:BNBa

[6568] Anonymous. Books: New books and publications. Scientific American, 32(1):8, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
8b.pdf.

Anonymous:1875:BPa

[6569] Anonymous. Business and personal. Scientific American, 32(1):10, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-10.pdf.

Anonymous:1875:IIa

[6570] Anonymous. Index of inventions. Scientific American, 32(1):11–12, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-11a.pdf.

Anonymous:1875:IPM

[6571] Anonymous. Improved planing and matching machine. Scientific American, 32(1):1, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
1.pdf.

Anonymous:1875:ISPa

[6572] Anonymous. Improved shot pouch sling. Scientific American, 32(1):1, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-1a.pdf.

Anonymous:1875:VAN

[6573] Anonymous. The vacuum an absolute non-conductor of electricity. Scientific American, 32(1):2, January 2, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n1/pdf/scientificamerican01021875-2.pdf.

Anonymous:1875:DKB

[6574] Anonymous. Don’t kiss the baby! Scientific American, 32(1):2, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-2a.pdf.

Anonymous:1875:PDSa

[6575] Anonymous. Postal detective service. Scientific American, 32(1):2, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-2b.pdf.

Anonymous:1875:TICa

[6576] Anonymous. Temperature of ignition of charcoal. Scientific American, 32(1):2–3, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
2c.pdf.

Anonymous:1875:SFSa

[6577] Anonymous. Scientific facts and speculations. Scientific American, 32(1):3, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
3.pdf.

Anonymous:1875:STC

[6578] Anonymous. Seeing through colored glasses. Scientific American, 32(1):3, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
3a.pdf.

Anonymous:1875:URNa

[6579] Anonymous. The underground railway, New York City. Scientific American, 32(1):4, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
4.pdf.

Anonymous:1875:RFL

[6580] Anonymous. Road and farm locomotives and steam road rollers. Scientific American, 32(1):5, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
5.pdf.

Anonymous:1875:FLWa

[6581] Anonymous. Finishing locomotive wheels. Scientific American, 32(1):6, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-6.pdf.

Anonymous:1875:WWM

[6582] Anonymous. Wool washing machinery. Scientific American, 32(1):6–7, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-6a.pdf.

Anonymous:1875:NSBa

[6583] Anonymous. New swimming bath in London. Scientific American, 32(1):7, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
7.pdf.

Anonymous:1875:ANa

[6584] Anonymous. Astronomical notes. Scientific American, 32(1):8, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-8.pdf.

Anonymous:1875:Ga

[6585] Anonymous. Glue. Scientific American, 32(1):8, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n1/pdf/scientificamerican01021875-8a.pdf.

Anonymous:1875:DCa

[6586] Anonymous. Decisions of the courts. Scientific American, 32(1):8–9, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-8c.pdf.

Anonymous:1875:RPO

[6587] Anonymous. Repeal of patents — opinion of the attorney general. Scientific American, 32(1):9, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
9.pdf.

Anonymous:1875:RAFa

[6588] Anonymous. Recent American and foreign patents. Scientific American, 32(1):9, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
9a.pdf.

Anonymous:1875:AC

[6589] Anonymous. Answers to correspondents. Scientific American, 32 (1):10–11, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n1/pdf/scientificamerican01021875-
10a.pdf.

Anonymous:1875:CRa

[6590] Anonymous. Communications received. Scientific American, 32(1):11, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-11.pdf.

Anonymous:1875:CPa

[6591] Anonymous. Canadian patents. Scientific American, 32(1):12, January 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n1/pdf/scientificamerican01021875-12.pdf.

Anonymous:1875:Ca

[6592] Anonymous. Correspondence. Scientific American, 32(2):20–21, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-20.pdf.

Anonymous:1875:BPb

[6593] Anonymous. Business and personal. Scientific American, 32(2):26, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-26.pdf.

Anonymous:1875:NQa

[6594] Anonymous. Notes and queries. Scientific American, 32(2):26–27, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-26a.pdf.

Anonymous:1875:IIb

[6595] Anonymous. Index of inventions. Scientific American, 32(2):28, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-28.pdf.

Anonymous:1875:BAEa

[6596] Anonymous. The buckeye automatic engine. Scientific American, 32(2):15, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
15.pdf.

Anonymous:1875:EMa

[6597] Anonymous. Is the ether matter? Scientific American, 32(2):16, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-16.pdf.

Anonymous:1875:WPa

[6598] Anonymous. The Woodbury Patent. Scientific American, 32(2):16, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-16a.pdf.

Anonymous:1875:PFM

[6599] Anonymous. The physical forces are modes of ether pressure. Scientific American, 32(2):16, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
16b.pdf.

Anonymous:1875:COa

[6600] Anonymous. The cultivation of oysters. Scientific American, 32(2):17, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-17.pdf.

Anonymous:1875:CPP

[6601] Anonymous. The Commissioner of Patents on the Patent Office tea set. Scientific American, 32(2):17, January 9, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n2/pdf/scientificamerican01091875-17a.pdf.

Anonymous:1875:MHS

[6602] Anonymous. The Morgan Hospital School, Dundee, Scotland. Scientific American, 32(2):18, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
18.pdf.

Anonymous:1875:BNG

[6603] Anonymous. British Naval guns. Scientific American, 32(2):19, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-19.pdf.

Anonymous:1875:PSW

[6604] Anonymous. Portable steam winch. Scientific American, 32(2):19, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-19a.pdf.

Anonymous:1875:IGMa

[6605] Anonymous. Improved grubbing machine. Scientific American, 32(2):22, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
22.pdf.

Anonymous:1875:IHD

[6606] Anonymous. Improved horse detacher. Scientific American, 32(2):22, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-22a.pdf.

Anonymous:1875:SAPa

[6607] Anonymous. Shade attachment for plows. Scientific American, 32(2):22, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-22b.pdf.

Anonymous:1875:SSB

[6608] Anonymous. Sir Samuel Baker’s expedition. Scientific American, 32(2):23, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
23.pdf.

Anonymous:1875:GWP

[6609] Anonymous. The great well at Prospect Park, Brooklyn. Scientific American, 32(2):23, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
23a.pdf.

Prescott:1875:OTa

[6610] George B. Prescott. Ocean telegraphy. Scientific American, 32(2):24, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-24.pdf.

Young:1875:CS

[6611] C. A. Young. The constitution of the Sun. Scientific American, 32(2):24–25, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
24a.pdf.

Anonymous:1875:RAFb

[6612] Anonymous. Recent American and foreign patents. Scientific American, 32(2):25–26, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n2/pdf/scientificamerican01091875-
25.pdf.

Anonymous:1875:CRb

[6613] Anonymous. Communications received. Scientific American, 32(2):27, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-27.pdf.

Anonymous:1875:CPb

[6614] Anonymous. Canadian patents. Scientific American, 32(2):28, January 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n2/pdf/scientificamerican01091875-28a.pdf.

Anonymous:1875:Cb

[6615] Anonymous. Correspondence. Scientific American, 32(3):36, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-36.pdf.

Anonymous:1875:BPc

[6616] Anonymous. Business and personal. Scientific American, 32(3):43, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-43.pdf.

Anonymous:1875:NQb

[6617] Anonymous. Notes and queries. Scientific American, 32(3):43–44, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-43a.pdf.

Anonymous:1875:IIc

[6618] Anonymous. Index of inventions. Scientific American, 32(3):44, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-44a.pdf.

Anonymous:1875:PCS

[6619] Anonymous. Patent chariot street car. Scientific American, 32 (3):31, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
31.pdf.

Anonymous:1875:ADA

[6620] Anonymous. Anstie and Dupré on the action of alcohol. Scientific American, 32(3):32, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
32.pdf.

Anonymous:1875:ONE

[6621] Anonymous. Our Naval efficiency. Scientific American, 32(3): 32–33, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
32a.pdf.

Anonymous:1875:PPa

[6622] Anonymous. Printing the patents. Scientific American, 32(3):33, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-33.pdf.

Anonymous:1875:COb

[6623] Anonymous. The cultivation of oysters. Scientific American, 32 (3):33, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
33a.pdf.

Anonymous:1875:FLWb

[6624] Anonymous. Finishing locomotive wheels. Scientific American, 32(3):34, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
34.pdf.

Anonymous:1875:HMCa

[6625] Anonymous. Hillside mining in Colorado. Scientific American, 32(3):35, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
35.pdf.

Rose:1875:PMa

[6626] Joshua Rose. Practical mechanism. Scientific American, 32(3):37, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-37.pdf.

Anonymous:1875:NGC

[6627] Anonymous. A new gas carburetter. Scientific American, 32(3):38, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-38.pdf.

Anonymous:1875:TUR

[6628] Anonymous. The Tompkins upright rotary knitting machine. Scientific American, 32(3):38, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
38a.pdf.

Anonymous:1875:NEM

[6629] Anonymous. A new electro-magnetic motor. Scientific American, 32(3):38, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
38b.pdf.

Anonymous:1875:LUS

[6630] Anonymous. The lighthouses of the United States. Scientific American, 32(3):39–40, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
39.pdf.

Prescott:1875:OTb

[6631] George H. Prescott. Ocean telegraphy. Scientific American, 32 (3):40–41, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
40.pdf.

Anonymous:1875:AOM

[6632] Anonymous. The arithmetical operations of multiplication and division. Scientific American, 32(3):41–42, January 16, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n3/pdf/scientificamerican01161875-41.pdf.

Anonymous:1875:RAFc

[6633] Anonymous. Recent American and foreign patents. Scientific American, 32(3):42, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
42.pdf.

Anonymous:1875:DCb

[6634] Anonymous. Decisions of the courts. Scientific American, 32 (3):42–43, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
42a.pdf.

Anonymous:1875:CRc

[6635] Anonymous. Communications received. Scientific American, 32 (3):44, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n3/pdf/scientificamerican01161875-
44.pdf.

Anonymous:1875:CPc

[6636] Anonymous. Canadian patents. Scientific American, 32(3):44–45, January 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n3/pdf/scientificamerican01161875-44b.pdf.

Anonymous:1875:Cc

[6637] Anonymous. Correspondence. Scientific American, 32(4):52, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-52.pdf.

Anonymous:1875:BNBb

[6638] Anonymous. Books: New books and publications. Scientific American, 32(4):58, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
58a.pdf.

Anonymous:1875:BPd

[6639] Anonymous. Business and personal. Scientific American, 32(4):59, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-59.pdf.

Anonymous:1875:NQc

[6640] Anonymous. Notes and queries. Scientific American, 32(4):59–60, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-59a.pdf.

Anonymous:1875:IId

[6641] Anonymous. Index of inventions. Scientific American, 32(4):60, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-60a.pdf.

Anonymous:1875:IRPa

[6642] Anonymous. Improved rotary pressure blower. Scientific American, 32(4):47, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
47.pdf.

Anonymous:1875:SIa

[6643] Anonymous. The sovereigns of industry. Scientific American, 32(4):48, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
48.pdf.

Anonymous:1875:BTB

[6644] Anonymous. A box trick to beat Hartz’. Scientific American, 32(4):48, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
48a.pdf.

Anonymous:1875:TCF

[6645] Anonymous. Testing the correctness of figures by weight. Scientific American, 32(4):48, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
48b.pdf.

Anonymous:1875:PR

[6646] Anonymous. P’i-Ryi. Scientific American, 32(4):48–49, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-48c.pdf.

Anonymous:1875:LLN

[6647] Anonymous. Loss of a large New York mail. Scientific American, 32(4):49, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
49.pdf.

Anonymous:1875:ICN

[6648] Anonymous. Improvement in canal navigation. Scientific American, 32(4):49, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
49a.pdf.

Anonymous:1875:AFTa

[6649] Anonymous. Amalgam fillings for the teeth. Scientific American, 32(4):49, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
49b.pdf.

Anonymous:1875:SPIa

[6650] Anonymous. Scientific and practical information. Scientific American, 32(4):49, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
49c.pdf.

Anonymous:1875:BOS

[6651] Anonymous. The Bessemer oscillating saloon. Scientific American, 32(4):50, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
50.pdf.

Anonymous:1875:TFD

[6652] Anonymous. Tubular floating docks. Scientific American, 32 (4):51–52, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
51.pdf.

Anonymous:1875:NMM

[6653] Anonymous. A new method of measuring surfaces, applied to the circle. Scientific American, 32(4):52–53, January 23, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n4/pdf/scientificamerican01231875-52a.pdf.

Anonymous:1875:IPCa

[6654] Anonymous. Improved potato cutter and planter. Scientific American, 32(4):54, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
54.pdf.

Anonymous:1875:BEP

[6655] Anonymous. Broadbooks’ Excelsior pruning shears. Scientific American, 32(4):54, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
54a.pdf.

Anonymous:1875:BIR

[6656] Anonymous. Blaisdell’s improved railway tie. Scientific American, 32(4):54, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
54b.pdf.

Anonymous:1875:CAP

[6657] Anonymous. Combined aquarium and plant case. Scientific American, 32(4):55–56, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
55.pdf.

Anonymous:1875:PPG

[6658] Anonymous. The physical phenomena of germination. Scientific American, 32(4):57, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
57.pdf.

Wiesner:1875:VFUa

[6659] Julius Wiesner. The vegetable fibers at the Universal Exposition, Vienna. Scientific American, 32(4):57, January 23, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n4/pdf/scientificamerican01231875-57a.pdf.

Anonymous:1875:DCc

[6660] Anonymous. Decisions of the courts. Scientific American, 32(4):58, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-58.pdf.

Anonymous:1875:RAFd

[6661] Anonymous. Recent American and foreign patents. Scientific American, 32(4):58, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
58b.pdf.

Anonymous:1875:CRd

[6662] Anonymous. Communications received. Scientific American, 32 (4):60, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n4/pdf/scientificamerican01231875-
60.pdf.

Anonymous:1875:CPd

[6663] Anonymous. Canadian patents. Scientific American, 32(4):60–61, January 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n4/pdf/scientificamerican01231875-60b.pdf.

Anonymous:1875:Cd

[6664] Anonymous. Correspondence. Scientific American, 32(5):68, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-68.pdf.

Anonymous:1875:BNBc

[6665] Anonymous. Books: New books and publications. Scientific American, 32(5):73, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
73.pdf.

Anonymous:1875:BPe

[6666] Anonymous. Business and personal. Scientific American, 32(5):74, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-74.pdf.

Anonymous:1875:NQd

[6667] Anonymous. Notes and queries. Scientific American, 32(5):74–76, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-74a.pdf.

Anonymous:1875:IIe

[6668] Anonymous. Index of inventions. Scientific American, 32(5):76, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-76a.pdf.

Anonymous:1875:IDS

[6669] Anonymous. Improved diamond stone saw. Scientific American, 32(5):63, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
63.pdf.

Anonymous:1875:EEB

[6670] Anonymous. Electric Earth batteries. Scientific American, 32 (5):64, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
64.pdf.

Anonymous:1875:THD

[6671] Anonymous. Torpedoes for harbor defense. Scientific American, 32(5):64, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
64a.pdf.

Anonymous:1875:SAB

[6672] Anonymous. Something about balloons. Scientific American, 32 (5):64–65, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
64b.pdf.

Anonymous:1875:SCa

[6673] Anonymous. Senate confirmations. Scientific American, 32(5):65, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-65.pdf.

Anonymous:1875:SPIb

[6674] Anonymous. Scientific and practical information. Scientific American, 32(5):65, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
65a.pdf.

Anonymous:1875:IT

[6675] Anonymous. Improved die tap. Scientific American, 32(5):66, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-66.pdf.

Anonymous:1875:FUP

[6676] Anonymous. Filtration under pressure. Scientific American, 32 (5):66, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
66a.pdf.

Anonymous:1875:URNb

[6677] Anonymous. The underground railway, New York City. Scientific American, 32(5):67–68, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
67.pdf.

Anonymous:1875:CSWa

[6678] Anonymous. The Crescent Steel Works at Pittsburgh, PA. Scientific American, 32(5):68, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
68a.pdf.

Rose:1875:PMb

[6679] Joshua Rose. Practical mechanism. Scientific American, 32(5):69, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-69.pdf.

Anonymous:1875:CGW

[6680] Anonymous. Cutting gear wheels. Scientific American, 32(5):70, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-70.pdf.

Anonymous:1875:BSS

[6681] Anonymous. The Bessemer saloon steamship. Scientific American, 32(5):70, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
70a.pdf.

Anonymous:1875:HLC

[6682] Anonymous. Highland and lowland cattle. Scientific American, 32(5):71, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
71.pdf.

Wiesner:1875:VFUb

[6683] Julius Wiesner. The vegetable fibers at the Universal Exposition Vienna. Scientific American, 32(5):72, January 30, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n5/pdf/scientificamerican01301875-72.pdf.

Anonymous:1875:DCd

[6684] Anonymous. Decisions of the courts. Scientific American, 32 (5):72–73, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
72a.pdf.

Anonymous:1875:RAFe

[6685] Anonymous. Recent American and foreign patents. Scientific American, 32(5):73, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
73a.pdf.

Anonymous:1875:CRe

[6686] Anonymous. Communications received. Scientific American, 32 (5):76, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n5/pdf/scientificamerican01301875-
76.pdf.

Anonymous:1875:CPe

[6687] Anonymous. Canadian patents. Scientific American, 32(5):76, January 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n5/pdf/scientificamerican01301875-76b.pdf.

Anonymous:1875:Ce

[6688] Anonymous. Correspondence. Scientific American, 32(6):83–85, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-83a.pdf.

Anonymous:1875:BPf

[6689] Anonymous. Business and personal. Scientific American, 32 (6):90, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
90.pdf.

Anonymous:1875:NQe

[6690] Anonymous. Notes and queries. Scientific American, 32(6):90– 92, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
90a.pdf.

Anonymous:1875:IIf

[6691] Anonymous. Index of inventions. Scientific American, 32(6):92, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-92a.pdf.

Anonymous:1875:HNB

[6692] Anonymous. Holmes’ new barrel machinery. Scientific American, 32(6):79, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
79.pdf.

Anonymous:1875:EF

[6693] Anonymous. The example of France. Scientific American, 32 (6):80, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
80.pdf.

Anonymous:1875:SSa

[6694] Anonymous. The “science” of spiritualism. Scientific American, 32(6):80, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
80a.pdf.

Anonymous:1875:OF

[6695] Anonymous. To our friends. Scientific American, 32(6):80, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-80b.pdf.

Anonymous:1875:Sa

[6696] Anonymous. Have a specialty. Scientific American, 32(6):80–81, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-80c.pdf.

Anonymous:1875:OSU

[6697] Anonymous. “ourselves, as others see us”. Scientific American, 32(6):81, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
81.pdf.

Anonymous:1875:PRSa

[6698] Anonymous. A problem relating to the sphere. Scientific American, 32(6):81, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
81a.pdf.

Anonymous:1875:SPIc

[6699] Anonymous. Scientific and practical information. Scientific American, 32(6):81, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
81b.pdf.

Anonymous:1875:PRP

[6700] Anonymous. The Pernot rotary puddler. Scientific American, 32(6):82, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
82.pdf.

Anonymous:1875:SCB

[6701] Anonymous. A self-corking bottle. Scientific American, 32(6):82, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-82a.pdf.

Anonymous:1875:CKP

[6702] Anonymous. Combined knife and pepper box. Scientific American, 32(6):82, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
82b.pdf.

Anonymous:1875:TT

[6703] Anonymous. The tulip tree. Scientific American, 32(6):83, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-83.pdf.

Anonymous:1875:SIb

[6704] Anonymous. The sardine industry. Scientific American, 32(6): 87–88, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
87.pdf.

Anonymous:1875:ANb

[6705] Anonymous. Astronomical notes. Scientific American, 32(6):88, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-88.pdf.

Anonymous:1875:RAFf

[6706] Anonymous. Recent American and foreign patents. Scientific American, 32(6):89, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
89.pdf.

Anonymous:1875:CRf

[6707] Anonymous. Communications received. Scientific American, 32 (6):92, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n6/pdf/scientificamerican02061875-
92.pdf.

Anonymous:1875:CPf

[6708] Anonymous. Canadian patents. Scientific American, 32(6):92, February 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n6/pdf/scientificamerican02061875-92b.pdf.

Anonymous:1875:Cf

[6709] Anonymous. Correspondence. Scientific American, 32(7):99– 101, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
99c.pdf.

Anonymous:1875:BPg

[6710] Anonymous. Business and personal. Scientific American, 32(7): 106, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
106.pdf.

Anonymous:1875:NQf

[6711] Anonymous. Notes and queries. Scientific American, 32(7):106– 107, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
106a.pdf.

Anonymous:1875:IIg

[6712] Anonymous. Index of inventions. Scientific American, 32(7): 107–108, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
107a.pdf.

Anonymous:1875:EUW

[6713] Anonymous. The excelsior universal wood worker. Scientific American, 32(7):95–96, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
95.pdf.

Anonymous:1875:PLAa

[6714] Anonymous. Patent law amendments. Scientific American, 32 (7):96–97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
96.pdf.

Anonymous:1875:CAT

[6715] Anonymous. Can ants talk? Scientific American, 32(7):97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n7/pdf/scientificamerican02131875-97.pdf.

Anonymous:1875:SMa

[6716] Anonymous. The sewing machine. Scientific American, 32(7):97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n7/pdf/scientificamerican02131875-97a.pdf.

Anonymous:1875:Pb

[6717] Anonymous. Prospects for 1875. Scientific American, 32(7):97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n7/pdf/scientificamerican02131875-97b.pdf.

Anonymous:1875:SRa

[6718] Anonymous. Science record for 1875. Scientific American, 32 (7):97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
97c.pdf.

Anonymous:1875:AMS

[6719] Anonymous. An ancient metric system. Scientific American, 32(7):97, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
97d.pdf.

Anonymous:1875:URNc

[6720] Anonymous. The underground railway, New York City. Scientific American, 32(7):98, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
98.pdf.

Anonymous:1875:BIB

[6721] Anonymous. Baker’s improved bootjack. Scientific American, 32(7):99, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
99.pdf.

Anonymous:1875:BML

[6722] Anonymous. The Belgian mode of lowering miners in shafts. Scientific American, 32(7):99, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
99a.pdf.

Anonymous:1875:CWB

[6723] Anonymous. Combined wrench and bolt cutter. Scientific American, 32(7):99, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
99b.pdf.

Rose:1875:PMc

[6724] Joshua Rose. Practical mechanism. Scientific American, 32(7): 101, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
101.pdf.

Anonymous:1875:ICL

[6725] Anonymous. Improved camp lounge. Scientific American, 32 (7):102, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
102.pdf.

Anonymous:1875:IGBa

[6726] Anonymous. Improved grate bar. Scientific American, 32(7):102, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n7/pdf/scientificamerican02131875-102a.pdf.

Anonymous:1875:RB

[6727] Anonymous. Remarkable birds. Scientific American, 32(7):103– 104, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
103.pdf.

Anonymous:1875:DCe

[6728] Anonymous. Decisions of the courts. Scientific American, 32 (7):105, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
105.pdf.

Anonymous:1875:RAFg

[6729] Anonymous. Recent American and foreign patents. Scientific American, 32(7):105, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
105a.pdf.

Anonymous:1875:CRg

[6730] Anonymous. Communications received. Scientific American, 32 (7):107, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n7/pdf/scientificamerican02131875-
107.pdf.

Anonymous:1875:CPg

[6731] Anonymous. Canadian patents. Scientific American, 32(7):108, February 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n7/pdf/scientificamerican02131875-108.pdf.

Anonymous:1875:Cg

[6732] Anonymous. Correspondence. Scientific American, 32(8):117, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n8/pdf/scientificamerican02201875-117.pdf.

Anonymous:1875:BNBd

[6733] Anonymous. Books: New books and publications. Scientific American, 32(8):121–122, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
121b.pdf.

Anonymous:1875:BPh

[6734] Anonymous. Business and personal. Scientific American, 32(8): 123, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
123.pdf.

Anonymous:1875:NQg

[6735] Anonymous. Notes and queries. Scientific American, 32(8):123– 124, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
123a.pdf.

Anonymous:1875:IIh

[6736] Anonymous. Index of inventions. Scientific American, 32(8):124, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n8/pdf/scientificamerican02201875-124a.pdf.

Anonymous:1875:PNB

[6737] Anonymous. The proposed Niagara Bridge at Lewiston, N.Y. Scientific American, 32(8):111–112, February 20, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n8/pdf/scientificamerican02201875-111.pdf.

Anonymous:1875:BSL

[6738] Anonymous. Bogus state laws concerning patent rights. Scientific American, 32(8):112, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
112.pdf.

Anonymous:1875:MAI

[6739] Anonymous. Metaline at the American institute. Scientific American, 32(8):112, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
112a.pdf.

Anonymous:1875:AA

[6740] Anonymous. Accurate alignments. Scientific American, 32(8): 112, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
112b.pdf.

Anonymous:1875:FM

[6741] Anonymous. Flying machines. Scientific American, 32(8):112– 113, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
112c.pdf.

Anonymous:1875:SSc

[6742] Anonymous. The “science” of spiritualism. Scientific American, 32(8):113, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
113.pdf.

Anonymous:1875:BWS

[6743] Anonymous. Not the best way to sell a patent. Scientific American, 32(8):113–114, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
113a.pdf.

Anonymous:1875:ICa

[6744] Anonymous. Irrigation in Colorado. Scientific American, 32(8): 114, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
114.pdf.

Anonymous:1875:IWMa

[6745] Anonymous. Improved water motor. Scientific American, 32 (8):115, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
115.pdf.

Anonymous:1875:MMa

[6746] Anonymous. Measuring mountains. Scientific American, 32(8): 115, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
115a.pdf.

Anonymous:1875:SPId

[6747] Anonymous. Scientific and practical information. Scientific American, 32(8):115–117, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
115b.pdf.

Orton:1875:VA

[6748] James Orton. The voices of animals. Scientific American, 32 (8):117, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
117a.pdf.

Anonymous:1875:RRS

[6749] Anonymous. Recent researches in the spectra of the planets. Scientific American, 32(8):117, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
117b.pdf.

Anonymous:1875:IPV

[6750] Anonymous. Improved plenum and vacuum pump. Scientific American, 32(8):118, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
118.pdf.

Anonymous:1875:NSIa

[6751] Anonymous. A new sad iron. Scientific American, 32(8):118, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n8/pdf/scientificamerican02201875-118a.pdf.

Anonymous:1875:HRC

[6752] Anonymous. Hayden’s revolving cartridge belt. Scientific American, 32(8):118, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
118b.pdf.

Anonymous:1875:WICa

[6753] Anonymous. Wellman’s improved car seat. Scientific American, 32(8):118, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
118c.pdf.

Anonymous:1875:HFFa

[6754] Anonymous. A hairy faced family. Scientific American, 32(8): 119–120, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
119.pdf.

Anonymous:1875:OTF

[6755] Anonymous. Ocean telegraphy. — the foreign connections of New York City and the existing rates of charges. Scientific American, 32(8):120–121, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
120.pdf.

Anonymous:1875:POC

[6756] Anonymous. The post office a carrier of merchandise. Scientific American, 32(8):121, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
121.pdf.

Anonymous:1875:DPOa

[6757] Anonymous. Decisions of the Patent Office. Scientific American, 32(8):121, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
121a.pdf.

Anonymous:1875:RAFh

[6758] Anonymous. Recent American and foreign patents. Scientific American, 32(8):122, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
122.pdf.

Anonymous:1875:CRh

[6759] Anonymous. Communications received. Scientific American, 32 (8):124, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n8/pdf/scientificamerican02201875-
124.pdf.

Anonymous:1875:CPh

[6760] Anonymous. Canadian patents. Scientific American, 32(8):124, February 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n8/pdf/scientificamerican02201875-124b.pdf.

Anonymous:1875:Ch

[6761] Anonymous. Correspondence. Scientific American, 32(9):132, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n9/pdf/scientificamerican02271875-132.pdf.

Anonymous:1875:BPi

[6762] Anonymous. Business and personal. Scientific American, 32(9): 138, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
138.pdf.

Anonymous:1875:NQh

[6763] Anonymous. Notes and queries. Scientific American, 32(9):138– 140, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
138a.pdf.

Anonymous:1875:IIi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n9/pdf/scientificamerican02271875-140a.pdf.

Anonymous:1875:EWT

[6765] Anonymous. The engineering: Works through, under, and over the Bergen Hills, opposite New York. Scientific American, 32(9): 127–128, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
127.pdf.

Anonymous:1875:SPIe

[6766] Anonymous. Scientific and practical information. Scientific American, 32(9):128, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
128.pdf.

Anonymous:1875:GW

[6767] Anonymous. The generation of the wicked. Scientific American, 32 (9):128–129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
128a.pdf.

Anonymous:1875:HRP

[6768] Anonymous. How to rush a patent through. Scientific American, 32(9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
129.pdf.

Anonymous:1875:SCSa

[6769] Anonymous. Spontaneous combustion at steam heat. Scientific American, 32(9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
129a.pdf.

Anonymous:1875:IPO

[6770] Anonymous. An important Patent Office decision. Scientific American, 32(9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
129b.pdf.

Anonymous:1875:CDa

[6771] Anonymous. Children and dogs. Scientific American, 32(9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n9/pdf/scientificamerican02271875-129c.pdf.

Anonymous:1875:LTR

[6772] Anonymous. Labor troubles in Russia. Scientific American, 32 (9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
129d.pdf.

Anonymous:1875:HFFb

[6773] Anonymous. Horse flesh as food in France. Scientific American, 32(9):129, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
129e.pdf.

Anonymous:1875:PS

[6774] Anonymous. The proportions of sewers. Scientific American, 32 (9):130, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
130.pdf.

Anonymous:1875:NCA

[6775] Anonymous. New Chilian armor clads. Scientific American, 32 (9):131, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
131.pdf.

Anonymous:1875:MSG

[6776] Anonymous. Mr. Seth Green’s new spawn-carrying device. Scientific American, 32(9):131, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
131a.pdf.

Rose:1875:PMd

[6777] Joshua Rose. Practical mechanism. Scientific American, 32(9): 133, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
133.pdf.

Anonymous:1875:RSB

[6778] Anonymous. Re-sawing band saw. Scientific American, 32(9): 134, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
134.pdf.

Anonymous:1875:NPM

[6779] Anonymous. A new pie marker. Scientific American, 32(9):134, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n9/pdf/scientificamerican02271875-134a.pdf.

Anonymous:1875:IA

[6780] Anonymous. Improved awning. Scientific American, 32(9):134– 135, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
134b.pdf.

Anonymous:1875:PBB

[6781] Anonymous. Public buildings in Brussels. Scientific American, 32 (9):135–137, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
135.pdf.

Anonymous:1875:DPOb

[6782] Anonymous. Decisions of the Patent Office. Scientific American, 32(9):137, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
137.pdf.

Anonymous:1875:RAFi

[6783] Anonymous. Recent American and foreign patents. Scientific American, 32(9):137, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
137a.pdf.

Anonymous:1875:CRi

[6784] Anonymous. Communications received. Scientific American, 32 (9):140, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n9/pdf/scientificamerican02271875-
140.pdf.

Anonymous:1875:CPi

[6785] Anonymous. Canadian patents. Scientific American, 32(9):140, February 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n9/pdf/scientificamerican02271875-140b.pdf.

Anonymous:1875:Ci

[6786] Anonymous. Correspondence. Scientific American, 32(10):148–149, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-148.pdf.

Anonymous:1875:BNBe

[6787] Anonymous. Books: New books and publications. Scientific American, 32(10):152, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
152.pdf.

Anonymous:1875:BPj

[6788] Anonymous. Business and personal. Scientific American, 32(10):154, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-154.pdf.

Anonymous:1875:NQi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-154a.pdf.

Anonymous:1875:IIj

[6790] Anonymous. Index of inventions. Scientific American, 32(10):156, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-156a.pdf.

Anonymous:1875:CST

[6791] Anonymous. Chasing screw threads. Scientific American, 32(10):143, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-143.pdf.

Anonymous:1875:RCP

[6792] Anonymous. Report of the Commissioner of Patents. Scientific American, 32(10):144, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
144.pdf.

Anonymous:1875:PPS

[6793] Anonymous. Progress and prospects of solar chemistry. Scientific American, 32(10):144, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
144a.pdf.

Anonymous:1875:ITE

[6794] Anonymous. An immense telegraphing establishment. Scientific American, 32(10):144–145, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
144b.pdf.

Anonymous:1875:NSBb

[6795] Anonymous. A new style of book-making needed. Scientific American, 32(10):145, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
145.pdf.

Anonymous:1875:WB

[6796] Anonymous. What are bacteria? Scientific American, 32(10):145, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-145a.pdf.

Anonymous:1875:ILWa

[6797] Anonymous. Improved Lady’s work table. Scientific American, 32(10):146, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
146.pdf.

Anonymous:1875:ISI

[6798] Anonymous. Improved sad iron. Scientific American, 32(10):146, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-146a.pdf.

Anonymous:1875:MH

[6799] Anonymous. A machine horse. Scientific American, 32(10):147–148, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-147.pdf.

Anonymous:1875:Ab

[6800] Anonymous. Ants. Scientific American, 32(10):149, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-149.pdf.

Anonymous:1875:SPIf

[6801] Anonymous. Scientific and practical information. Scientific American, 32(10):149, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
149a.pdf.

Anonymous:1875:ISHa

[6802] Anonymous. Improved self-heating fluting and sad iron. Scientific American, 32(10):150, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
150.pdf.

Anonymous:1875:IBT

[6803] Anonymous. Improved bale tie for wire bands. Scientific American, 32(10):150, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
150a.pdf.

Anonymous:1875:CSG

[6804] Anonymous. Cylindrical standard gages. Scientific American, 32 (10):150, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
150b.pdf.

Anonymous:1875:LPM

[6805] Anonymous. Life-preserving mattress. Scientific American, 32 (10):150–151, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
150c.pdf.

Anonymous:1875:BCa

[6806] Anonymous. The Breton cattle. Scientific American, 32(10):151–152, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-151.pdf.

Anonymous:1875:DPOc

[6807] Anonymous. Decisions of the Patent Office. Scientific American, 32(10):152, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
152a.pdf.

Anonymous:1875:RAFj

[6808] Anonymous. Recent American and foreign patents. Scientific American, 32(10):153, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
153.pdf.

Anonymous:1875:CRj

[6809] Anonymous. Communications received. Scientific American, 32 (10):156, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n10/pdf/scientificamerican03061875-
156.pdf.

Anonymous:1875:CPj

[6810] Anonymous. Canadian patents. Scientific American, 32(10):156, March 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n10/pdf/scientificamerican03061875-156b.pdf.

Anonymous:1875:Cj

[6811] Anonymous. Correspondence. Scientific American, 32(11):165, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-165a.pdf.

Anonymous:1875:BNBf

[6812] Anonymous. Books: New books and publications. Scientific American, 32(11):168, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
168a.pdf.

Anonymous:1875:BPk

[6813] Anonymous. Business and personal. Scientific American, 32(11):170, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-170.pdf.

Anonymous:1875:NQj

[6814] Anonymous. Notes and queries. Scientific American, 32(11):170–171, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-170a.pdf.

Anonymous:1875:IIk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-172a.pdf.

Anonymous:1875:AT

[6816] Anonymous. Amateur taxidermy. Scientific American, 32(11):159–160, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-159.pdf.

Anonymous:1875:NRP

[6817] Anonymous. The new rule of the Patent Office. Scientific American, 32(11):160–161, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
160.pdf.

Anonymous:1875:PPb

[6818] Anonymous. Piscicultural progress. Scientific American, 32(11):161, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-161.pdf.

Anonymous:1875:NE

[6819] Anonymous. Nature’s efforts. Scientific American, 32(11):161, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-161a.pdf.

Anonymous:1875:RTP

[6820] Anonymous. Rubber tipped pencils. Scientific American, 32(11):161, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-161b.pdf.

Anonymous:1875:WLF

[6821] Anonymous. Waste land and forest culture. Scientific American, 32(11):161, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
161c.pdf.

Anonymous:1875:IAT

[6822] Anonymous. Improved animal trap. Scientific American, 32(11):162, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-162.pdf.

Anonymous:1875:MPa

[6823] Anonymous. The manipulation of pruning. Scientific American, 32(11):162, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
162a.pdf.

Anonymous:1875:NFF

[6824] Anonymous. Non-freezing float steam trap. Scientific American, 32(11):163, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
163.pdf.

Anonymous:1875:ISGa

[6825] Anonymous. Improved safety governor. Scientific American, 32 (11):163–164, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
163a.pdf.

Rose:1875:PMe

[6826] Joshua Rose. Practical mechanism. Scientific American, 32(11): 164–165, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
164.pdf.

Anonymous:1875:ANc

[6827] Anonymous. Astronomical notes. Scientific American, 32(11):165, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-165.pdf.

Anonymous:1875:IRA

[6828] Anonymous. Improved railway axle box. Scientific American, 32(11):166, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
166.pdf.

Anonymous:1875:TRP

[6829] Anonymous. Two remarkable plants. Scientific American, 32 (11):167–168, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
167.pdf.

Anonymous:1875:ECMa

[6830] Anonymous. Engineers, contractors, machinists, and manufacturers. Scientific American, 32(11):168, March 13, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n11/pdf/scientificamerican03131875-168.pdf.

Anonymous:1875:DCf

[6831] Anonymous. Decisions of the courts. Scientific American, 32(11):168, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-168b.pdf.

Anonymous:1875:RAFk

[6832] Anonymous. Recent American and foreign patents. Scientific American, 32(11):169, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
169.pdf.

Anonymous:1875:CRk

[6833] Anonymous. Communications received. Scientific American, 32 (11):172, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n11/pdf/scientificamerican03131875-
172.pdf.

Anonymous:1875:CPk

[6834] Anonymous. Canadian patents. Scientific American, 32(11):172, March 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n11/pdf/scientificamerican03131875-172b.pdf.

Anonymous:1875:BPl

[6835] Anonymous. Business and personal. Scientific American, 32(12):186, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-186.pdf.

Anonymous:1875:NQk

[6836] Anonymous. Notes and queries. Scientific American, 32(12):186–188, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-186a.pdf.

Anonymous:1875:IIl

[6837] Anonymous. Index of inventions. Scientific American, 32(12):188, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-188a.pdf.

Anonymous:1875:PMNa

[6838] Anonymous. Porcelain manufacture in New York. Scientific American, 32(12):175–176, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
175.pdf.

Anonymous:1875:CBE

[6839] Anonymous. Crossing the boundary of the experimental evidence. Scientific American, 32(12):176, March 20, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n12/pdf/scientificamerican03201875-176.pdf.

Anonymous:1875:FEC

[6840] Anonymous. Foreign exhibitors at the Centennial. Scientific American, 32(12):176–177, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
176a.pdf.

Anonymous:1875:APa

[6841] Anonymous. American porcelain. Scientific American, 32(12):177, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-177.pdf.

Anonymous:1875:VS

[6842] Anonymous. A variable screw. Scientific American, 32(12):177, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-177a.pdf.

Anonymous:1875:MI

[6843] Anonymous. The metallurgy of iridium. Scientific American, 32(12):177, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
177b.pdf.

Anonymous:1875:FR

[6844] Anonymous. The Fireman’s respirator. Scientific American, 32(12):178, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-178.pdf.

Anonymous:1875:NRS

[6845] Anonymous. The new revelations of a snow flake. Scientific American, 32(12):178, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
178a.pdf.

Anonymous:1875:IL

[6846] Anonymous. Improved Lewis. Scientific American, 32(12):179, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-179.pdf.

Anonymous:1875:NSL

[6847] Anonymous. A new steam launch. Scientific American, 32(12):179, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-179a.pdf.

Anonymous:1875:TWO

[6848] Anonymous. A three-wheeled omnibus. Scientific American, 32 (12):179–180, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
179b.pdf.

Anonymous:1875:SPIg

[6849] Anonymous. Scientific and practical information. Scientific American, 32(12):181, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
181.pdf.

Anonymous:1875:ICDa

[6850] Anonymous. Improved center draft tongue for reapers and mowers. Scientific American, 32(12):182, March 20, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n12/pdf/scientificamerican03201875-182.pdf.

Anonymous:1875:IMD

[6851] Anonymous. Improved mining dial. Scientific American, 32(12):182, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-182a.pdf.

Anonymous:1875:RAH

[6852] Anonymous. Reid’s automatic hatchway. Scientific American, 32(12):182, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
182b.pdf.

Anonymous:1875:IPJ

[6853] Anonymous. Improved pipe joint. Scientific American, 32(12):182, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-182c.pdf.

Anonymous:1875:Ck

[6854] Anonymous. The crane. Scientific American, 32(12):183–184, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-183.pdf.

Buel:1875:VWD

[6855] Richard H. Buel. The volume and weight of distilled water at different temperatures. Scientific American, 32(12):184, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-184.pdf.

Anonymous:1875:DCg

[6856] Anonymous. Decisions of the courts. Scientific American, 32(12):185, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-185.pdf.

Anonymous:1875:RAFl

[6857] Anonymous. Recent American and foreign patents. Scientific American, 32(12):185–186, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
185a.pdf.

Anonymous:1875:CRl

[6858] Anonymous. Communications received. Scientific American, 32 (12):188, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n12/pdf/scientificamerican03201875-
188.pdf.

Anonymous:1875:CPl

[6859] Anonymous. Canadian patents. Scientific American, 32(12):188, March 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n12/pdf/scientificamerican03201875-188b.pdf.

Anonymous:1875:Cl

[6860] Anonymous. Correspondence. Scientific American, 32(13):196–197, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-196.pdf.

Anonymous:1875:BNBg

[6861] Anonymous. Books: New books and publications. Scientific American, 32(13):200, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
200a.pdf.

Anonymous:1875:BPm

[6862] Anonymous. Business and personal. Scientific American, 32(13):202, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-202.pdf.

Anonymous:1875:NQl

[6863] Anonymous. Notes and queries. Scientific American, 32(13):202–204, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-202a.pdf.

Anonymous:1875:IIm

[6864] Anonymous. Index of inventions. Scientific American, 32(13):204, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-204a.pdf.

Anonymous:1875:PMNb

[6865] Anonymous. Porcelain manufacture in New York. Scientific American, 32(13):191–192, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
191.pdf.

Anonymous:1875:FPH

[6866] Anonymous. Frozen pipes and hydrants. Scientific American, 32(13):192, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
192.pdf.

Anonymous:1875:IPCb

[6867] Anonymous. Increase of postal charges. Scientific American, 32(13):192, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-192a.pdf.

Anonymous:1875:MMb

[6868] Anonymous. Models by mail. Scientific American, 32(13):192, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-192b.pdf.

Anonymous:1875:TMR

[6869] Anonymous. A trade mark rejection. Scientific American, 32(13):192, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-192c.pdf.

Anonymous:1875:PAP

[6870] Anonymous. Partnerships of ants and plants. Scientific American, 32(13):192–193, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
192d.pdf.

Thurston:1875:MCM

[6871] R. H. Thurston. Molecular changes in metals. Scientific American, 32(13):193, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
193.pdf.

Anonymous:1875:ITPa

[6872] Anonymous. Improved tree protector. Scientific American, 32(13):194, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-194.pdf.

Anonymous:1875:MOC

[6873] Anonymous. To make ovals and circles of any size. Scientific American, 32(13):194, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
194a.pdf.

Anonymous:1875:MBM

[6874] Anonymous. Measuring building materials. Scientific American, 32(13):194, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
194b.pdf.

Anonymous:1875:ISR

[6875] Anonymous. Improved sugar-refining tank. Scientific American, 32(13):194, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
194c.pdf.

Anonymous:1875:GSIa

[6876] Anonymous. Gathering and storing ice. Scientific American, 32 (13):195–196, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
195.pdf.

Rose:1875:PMf

[6877] Joshua Rose. Practical mechanism. Scientific American, 32(13):197, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-197.pdf.

Anonymous:1875:IHMa

[6878] Anonymous. Improved hydraulic motor. Scientific American, 32 (13):198, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
198.pdf.

Anonymous:1875:IDL

[6879] Anonymous. Improved device for laying out sashes. Scientific American, 32(13):198, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
198a.pdf.

Anonymous:1875:EIO

[6880] Anonymous. An east Indian ornamental shrub. Scientific American, 32(13):199, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
199.pdf.

Anonymous:1875:CLB

[6881] Anonymous. The cut-leaved begonia. Scientific American, 32 (13):199–200, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
199a.pdf.

Anonymous:1875:DCh

[6882] Anonymous. Decisions of the courts. Scientific American, 32(13):200, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-200.pdf.

Anonymous:1875:RAFm

[6883] Anonymous. Recent American and foreign patents. Scientific American, 32(13):200–202, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
200b.pdf.

Anonymous:1875:CRm

[6884] Anonymous. Communications received. Scientific American, 32 (13):204, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n13/pdf/scientificamerican03271875-
204.pdf.

Anonymous:1875:CPm

[6885] Anonymous. Canadian patents. Scientific American, 32(13):204, March 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n13/pdf/scientificamerican03271875-204b.pdf.

Anonymous:1875:Cm

[6886] Anonymous. Correspondence. Scientific American, 32(14):212–213, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-212.pdf.

Anonymous:1875:BPn

[6887] Anonymous. Business and personal. Scientific American, 32(14):218, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-218.pdf.

Anonymous:1875:NQm

[6888] Anonymous. Notes and queries. Scientific American, 32(14):218–220, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-218a.pdf.

Anonymous:1875:IIn

[6889] Anonymous. Index of inventions. Scientific American, 32(14):220, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-220a.pdf.

Anonymous:1875:SFSb

[6890] Anonymous. The Saxby and Farmer safety switches and signals. Scientific American, 32(14):207–208, April 3, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n14/pdf/scientificamerican04031875-207.pdf.

Anonymous:1875:SCSb

[6891] Anonymous. Science in common schooling. Scientific American, 32(14):208, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
208.pdf.

Anonymous:1875:SSd

[6892] Anonymous. Syrian sponges. Scientific American, 32(14):208–209, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-208a.pdf.

Anonymous:1875:IB

[6893] Anonymous. Inspection of boilers. Scientific American, 32(14):209, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-209.pdf.

Anonymous:1875:MSI

[6894] Anonymous. Michigan’s salt interests. Scientific American, 32(14):209, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-209a.pdf.

Anonymous:1875:MF

[6895] Anonymous. The mission of the fly. Scientific American, 32(14):209, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-209b.pdf.

Anonymous:1875:HRT

[6896] Anonymous. The Hudson River Tunnel at New York. Scientific American, 32(14):211, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
211.pdf.

Anonymous:1875:SPIh

[6897] Anonymous. Scientific and practical information. Scientific American, 32(14):213, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
213.pdf.

Anonymous:1875:IWIa

[6898] Anonymous. Improved water injectors. Scientific American, 32 (14):214–215, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
214.pdf.

Anonymous:1875:RD

[6899] Anonymous. Red deer. Scientific American, 32(14):215–216, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-215.pdf.

Anonymous:1875:RAFn

[6900] Anonymous. Recent American and foreign patents. Scientific American, 32(14):216–218, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
216.pdf.

Anonymous:1875:CRn

[6901] Anonymous. Communications received. Scientific American, 32 (14):220, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n14/pdf/scientificamerican04031875-
220.pdf.

Anonymous:1875:CPn

[6902] Anonymous. Canadian patents. Scientific American, 32(14):221, April 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n14/pdf/scientificamerican04031875-221.pdf.

Anonymous:1875:Cn

[6903] Anonymous. Correspondence. Scientific American, 32(15):228, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-228.pdf.

Anonymous:1875:BPo

[6904] Anonymous. Business and personal. Scientific American, 32(15):234, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-234.pdf.

Anonymous:1875:NQn

[6905] Anonymous. Notes and queries. Scientific American, 32(15):234–236, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-234a.pdf.

Anonymous:1875:IIo

[6906] Anonymous. Index of inventions. Scientific American, 32(15):236, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-236a.pdf.

Anonymous:1875:PDSb

[6907] Anonymous. The pneumatic despatch system in the Western Union Telegraph Building, New York. Scientific American, 32(15):223–224, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-223.pdf.

Anonymous:1875:ICGa

[6908] Anonymous. The industrial condition of Germany. Scientific American, 32(15):224, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
224.pdf.

Anonymous:1875:HSAa

[6909] Anonymous. How science is annihilated. Scientific American, 32 (15):224–225, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
224a.pdf.

Anonymous:1875:HPC

[6910] Anonymous. How a piece of coal Burns. Scientific American, 32(15):225, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-225.pdf.

Anonymous:1875:RT

[6911] Anonymous. Up or down in rapid transit? Scientific American, 32(15):225, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
225a.pdf.

Anonymous:1875:PTU

[6912] Anonymous. Proposed tunnel under Newark Bay. Scientific American, 32(15):225, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
225b.pdf.

Anonymous:1875:NSIb

[6913] Anonymous. A new stellar indicator. Scientific American, 32(15):226, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-226.pdf.

Anonymous:1875:SQA

[6914] Anonymous. Spectroscopic quantitative analysis. Scientific American, 32(15):226, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
226a.pdf.

Anonymous:1875:URB

[6915] Anonymous. Underground refrigerator for butchers. Scientific American, 32(15):226, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
226b.pdf.

Anonymous:1875:EI

[6916] Anonymous. The earthquake indicator. Scientific American, 32(15):226, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-226c.pdf.

Anonymous:1875:NDF

[6917] Anonymous. A novel decoy fish. Scientific American, 32(15):226, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-226d.pdf.

Anonymous:1875:MMF

[6918] Anonymous. Machine for making fence pickets. Scientific American, 32(15):227, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
227.pdf.

Anonymous:1875:IHMb

[6919] Anonymous. Improved hoisting machine. Scientific American, 32 (15):227–228, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
227a.pdf.

Anonymous:1875:ANd

[6920] Anonymous. Astronomical notes. Scientific American, 32(15):228–229, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-228a.pdf.

Anonymous:1875:SHF

[6921] Anonymous. The storage and handling of fluids. Scientific American, 32(15):229, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
229.pdf.

Anonymous:1875:ICSa

[6922] Anonymous. Improved circular saw table. Scientific American, 32(15):230, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
230.pdf.

Anonymous:1875:IW

[6923] Anonymous. Improved windmill. Scientific American, 32(15):230, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-230a.pdf.

Anonymous:1875:SCC

[6924] Anonymous. Strasbourg Cathedral and its clock. Scientific American, 32(15):231–232, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
231.pdf.

Anonymous:1875:DCi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-232.pdf.

Anonymous:1875:RAFo

[6926] Anonymous. Recent American and foreign patents. Scientific American, 32(15):232–233, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
232a.pdf.

Anonymous:1875:CRo

[6927] Anonymous. Communications received. Scientific American, 32 (15):236, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n15/pdf/scientificamerican04101875-
236.pdf.

Anonymous:1875:CPo

[6928] Anonymous. Canadian patents. Scientific American, 32(15):236, April 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n15/pdf/scientificamerican04101875-236b.pdf.

Anonymous:1875:Cq

[6929] Anonymous. Correspondence. Scientific American, 32(16):244, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-244.pdf.

Anonymous:1875:BNBh

[6930] Anonymous. Books: New books and publications. Scientific American, 32(16):248, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
248b.pdf.

Anonymous:1875:BPp

[6931] Anonymous. Business and personal. Scientific American, 32(16):250, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-250.pdf.

Anonymous:1875:NQo

[6932] Anonymous. Notes and queries. Scientific American, 32(16):250–252, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-250a.pdf.

Anonymous:1875:IIp

[6933] Anonymous. Index of inventions. Scientific American, 32(16):252, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-252a.pdf.

Anonymous:1875:WRT

[6934] Anonymous. Wire rope traction street railway. Scientific American, 32(16):239, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
239.pdf.

Anonymous:1875:TMa

[6935] Anonymous. Trade marks. Scientific American, 32(16):240, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-240.pdf.

Anonymous:1875:HNW

[6936] Anonymous. Home news by way of the Sun. Scientific American, 32(16):240, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
240a.pdf.

Anonymous:1875:GPL

[6937] Anonymous. German patent law. Scientific American, 32(16):240–241, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-240b.pdf.

Anonymous:1875:SRI

[6938] Anonymous. A street railway improvement wanted. Scientific American, 32(16):241, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
241.pdf.

Anonymous:1875:ND

[6939] Anonymous. A new dodge. Scientific American, 32(16):241, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-241a.pdf.

Anonymous:1875:HBW

[6940] Anonymous. How to build a windmill. Scientific American, 32(16):241, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-241b.pdf.

Anonymous:1875:SPIi

[6941] Anonymous. Scientific and practical information. Scientific American, 32(16):241, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
241c.pdf.

Anonymous:1875:NMS

[6942] Anonymous. New mode of slaughtering cattle. Scientific American, 32(16):242, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
242.pdf.

Anonymous:1875:IBS

[6943] Anonymous. Improved bottle stopper. Scientific American, 32(16):242, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-242a.pdf.

Anonymous:1875:EPa

[6944] Anonymous. Extermination of the phylloxera. Scientific American, 32(16):242, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
242b.pdf.

Anonymous:1875:GRC

[6945] Anonymous. Giffard’s railway car. Scientific American, 32(16):242, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-242c.pdf.

Anonymous:1875:NDB

[6946] Anonymous. A novel design for a British channel tunnel. Scientific American, 32(16):243–244, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
243.pdf.

Rose:1875:PMN

[6947] Joshua Rose. Practical mechanism number XXI. Scientific American, 32(16):245, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
245.pdf.

Anonymous:1875:IBCa

[6948] Anonymous. Improved bolt cutter. Scientific American, 32(16):246, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-246.pdf.

Anonymous:1875:IDPa

[6949] Anonymous. Improved dinner pail. Scientific American, 32(16):246, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-246a.pdf.

Anonymous:1875:TICb

[6950] Anonymous. Tomlinson’s improved chimney cowl. Scientific American, 32(16):246, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
246b.pdf.

Anonymous:1875:Gb

[6951] Anonymous. The grayhound. Scientific American, 32(16):247, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-247.pdf.

Anonymous:1875:K

[6952] Anonymous. The Kagu. Scientific American, 32(16):247–248, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-247a.pdf.

Anonymous:1875:PPF

[6953] Anonymous. A prize plan for a fireproof house. Scientific American, 32(16):248, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
248.pdf.

Anonymous:1875:DCj

[6954] Anonymous. Decisions of the courts. Scientific American, 32(16):248, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n16/pdf/scientificamerican04171875-248a.pdf.

Anonymous:1875:RAFp

[6955] Anonymous. Recent American and foreign patents. Scientific American, 32(16):249, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
249.pdf.

Anonymous:1875:CRp

[6956] Anonymous. Communications received. Scientific American, 32 (16):252, April 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n16/pdf/scientificamerican04171875-
252.pdf.

Anonymous:1875:Cs

[6957] Anonymous. Correspondence. Scientific American, 32(17):260, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-260.pdf.

Anonymous:1875:BPq

[6958] Anonymous. Business and personal. Scientific American, 32(17):267, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-267.pdf.

Anonymous:1875:NQp

[6959] Anonymous. Notes and queries. Scientific American, 32(17):267, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-267a.pdf.

Anonymous:1875:IIq

[6960] Anonymous. Index of inventions. Scientific American, 32(17):267–268, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-267c.pdf.

Anonymous:1875:IBCb

[6961] Anonymous. Improved bridge construction. Scientific American, 32(17):255, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
255.pdf.

Anonymous:1875:ICPa

[6962] Anonymous. Important changes proposed in the British patent laws. Scientific American, 32(17):256, April 24, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n17/pdf/scientificamerican04241875-256.pdf.

Anonymous:1875:NMR

[6963] Anonymous. A new machine for removing snow from streets. Scientific American, 32(17):256, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
256a.pdf.

Anonymous:1875:PRTa

[6964] Anonymous. Progress of rapid transit in New York. Scientific American, 32(17):256–257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
256b.pdf.

Anonymous:1875:AFTb

[6965] Anonymous. Amalgams for filling the teeth. Scientific American, 32(17):257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
257.pdf.

Anonymous:1875:TAI

[6966] Anonymous. Test of American iron and steel. Scientific American, 32(17):257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
257a.pdf.

Anonymous:1875:AEI

[6967] Anonymous. Another earthquake indicator. Scientific American, 32(17):257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
257b.pdf.

Anonymous:1875:KUS

[6968] Anonymous. Kaolin in the United States. Scientific American, 32(17):257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
257c.pdf.

Anonymous:1875:SPIj

[6969] Anonymous. Scientific and practical information. Scientific American, 32(17):257, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
257d.pdf.

Anonymous:1875:BIL

[6970] Anonymous. Beckwith’s improved lumber dog. Scientific American, 32(17):258, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
258.pdf.

Anonymous:1875:NMA

[6971] Anonymous. A new Mercury air pump. Scientific American, 32(17):258, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-258a.pdf.

Anonymous:1875:DIL

[6972] Anonymous. Dittman’s improved locking nut. Scientific American, 32(17):258, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
258b.pdf.

Anonymous:1875:ADI

[6973] Anonymous. Apparatus for determining the impermeability of tissues. Scientific American, 32(17):258, April 24, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n17/pdf/scientificamerican04241875-258c.pdf.

Anonymous:1875:SPD

[6974] Anonymous. The Sun’s position at different seasons of the year. Scientific American, 32(17):259, April 24, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n17/pdf/scientificamerican04241875-259.pdf.

Anonymous:1875:NTH

[6975] Anonymous. A new tack hammer. Scientific American, 32(17):259, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-259a.pdf.

Anonymous:1875:Ad

[6976] Anonymous. The Anatifa. Scientific American, 32(17):259, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-259b.pdf.

Anonymous:1875:CPq

[6977] Anonymous. A Chinese patentee. Scientific American, 32(17):261, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-261.pdf.

Anonymous:1875:ISM

[6978] Anonymous. Improved shingle machine. Scientific American, 32 (17):262, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
262.pdf.

Anonymous:1875:ITPb

[6979] Anonymous. Improved tanning process. Scientific American, 32 (17):262, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
262a.pdf.

Anonymous:1875:SBb

[6980] Anonymous. The secretary bird. Scientific American, 32(17):263, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-263.pdf.

Anonymous:1875:WSP

[6981] Anonymous. The Wardian system of plant cases. Scientific American, 32(17):263, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
263a.pdf.

Anonymous:1875:MTC

[6982] Anonymous. The modern theory of color. Scientific American, 32(17):264–265, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
264.pdf.

Anonymous:1875:DCk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n17/pdf/scientificamerican04241875-266.pdf.

Anonymous:1875:RAFq

[6984] Anonymous. Recent American and foreign patents. Scientific American, 32(17):266, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
266a.pdf.

Anonymous:1875:CRq

[6985] Anonymous. Communications received. Scientific American, 32 (17):267, April 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n17/pdf/scientificamerican04241875-
267b.pdf.

Anonymous:1875:Ct

[6986] Anonymous. Correspondence. Scientific American, 32(18):276, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-276.pdf.

Anonymous:1875:BNBi

[6987] Anonymous. Books: New books and publications. Scientific American, 32(18):281, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
281.pdf.

Anonymous:1875:BPr

[6988] Anonymous. Business and personal. Scientific American, 32(18):282, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-282.pdf.

Anonymous:1875:NQq

[6989] Anonymous. Notes and queries. Scientific American, 32(18):282–284, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-282a.pdf.

Anonymous:1875:IIr

[6990] Anonymous. Index of inventions. Scientific American, 32(18):284, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-284a.pdf.

Anonymous:1875:ISHb

[6991] Anonymous. Improved shingle and heading machinery. Scientific American, 32(18):271, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
271.pdf.

Anonymous:1875:MVE

[6992] Anonymous. The money value of education. Scientific American, 32(18):272, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
272.pdf.

Anonymous:1875:PMP

[6993] Anonymous. Photo-mechanical printing. Scientific American, 32 (18):272–273, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
272a.pdf.

Anonymous:1875:IMM

[6994] Anonymous. Improvements at the mouth of the Mississippi. Scientific American, 32(18):273, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
273.pdf.

Anonymous:1875:MMBa

[6995] Anonymous. Motion on a moving body. Scientific American, 32(18):273, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-273a.pdf.

Anonymous:1875:WCT

[6996] Anonymous. What is the cause of tides? Scientific American, 32(18):273–274, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
273b.pdf.

Anonymous:1875:SIW

[6997] Anonymous. Smith’s improved whiffletree hook. Scientific American, 32(18):274, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
274.pdf.

Anonymous:1875:COI

[6998] Anonymous. A curious ocular illusion. Scientific American, 32(18):274, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-274a.pdf.

Anonymous:1875:SIAa

[6999] Anonymous. Stockley’s improved anti-primer. Scientific American, 32(18):274, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
274b.pdf.

Anonymous:1875:WICb

[7000] Anonymous. Whipple’s improved clothes dryer. Scientific American, 32(18):275, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
275.pdf.

Anonymous:1875:IEL

[7001] Anonymous. Improved electric lamp lighter. Scientific American, 32(18):275, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
275a.pdf.

Anonymous:1875:NLB

[7002] Anonymous. A new locking bolt. Scientific American, 32(18):275–276, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-275b.pdf.

Rose:1875:PMg

[7003] Joshua Rose. Practical mechanism. Scientific American, 32(18):277, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-277.pdf.

Anonymous:1875:MFB

[7004] Anonymous. Movable fronts for buildings. Scientific American, 32(18):278, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
278.pdf.

Anonymous:1875:IRPb

[7005] Anonymous. Improved rotary pump. Scientific American, 32(18):278, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-278a.pdf.

Anonymous:1875:GO

[7006] Anonymous. Growing orchids. Scientific American, 32(18):279, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n18/pdf/scientificamerican05011875-279.pdf.

Anonymous:1875:SCd

[7007] Anonymous. The Solanum Crinitum. Scientific American, 32 (18):279–280, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
279a.pdf.

Anonymous:1875:CCS

[7008] Anonymous. Car and carriage springs. Scientific American, 32 (18):280–281, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
280.pdf.

Anonymous:1875:RAFr

[7009] Anonymous. Recent American and foreign patents. Scientific American, 32(18):281, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
281a.pdf.

Anonymous:1875:CRr

[7010] Anonymous. Communications received. Scientific American, 32 (18):284, May 1, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n18/pdf/scientificamerican05011875-
284.pdf.

Anonymous:1875:Cu

[7011] Anonymous. Correspondence. Scientific American, 32(19):292, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-292.pdf.

Anonymous:1875:BPs

[7012] Anonymous. Business and personal. Scientific American, 32(19):298, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-298.pdf.

Anonymous:1875:NQr

[7013] Anonymous. Notes and queries. Scientific American, 32(19):298–299, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-298a.pdf.

Anonymous:1875:IIs

[7014] Anonymous. Index of inventions. Scientific American, 32(19):299–300, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-299a.pdf.

Anonymous:1875:ICDb

[7015] Anonymous. Improved coffer dam. Scientific American, 32(19):287–288, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-287.pdf.

Anonymous:1875:IP

[7016] Anonymous. An inland penikese. Scientific American, 32(19):288, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://en.wikipedia.org/wiki/Penikese;
 http://www.nature.com/scientificamerican/journal/v32/n19/
pdf/scientificamerican05081875-288.pdf.

Anonymous:1875:CCE

[7017] Anonymous. The chain of criminal entailment. Scientific American, 32(19):288, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
288a.pdf.

Anonymous:1875:PBT

[7018] Anonymous. Pictures better than stories. Scientific American, 32(19):288, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
288b.pdf.

Anonymous:1875:SCR

[7019] Anonymous. Some curious results of experimental surgery. Scientific American, 32(19):288–289, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
288c.pdf.

Anonymous:1875:RRP

[7020] Anonymous. The recent remarkable progress in the steel industry. Scientific American, 32(19):289, May 8, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n19/pdf/scientificamerican05081875-289.pdf.

Anonymous:1875:CLa

[7021] Anonymous. Cimex Lectularius. Scientific American, 32(19):289, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-289a.pdf.

Anonymous:1875:IGBb

[7022] Anonymous. Improved gas burner for heating purposes. Scientific American, 32(19):290, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
290.pdf.

Anonymous:1875:SPIk

[7023] Anonymous. Scientific and practical information. Scientific American, 32(19):290, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
290a.pdf.

Anonymous:1875:HBH

[7024] Anonymous. Heating buildings with hot water. Scientific American, 32(19):290, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
290b.pdf.

Anonymous:1875:ISPb

[7025] Anonymous. Improved sectional pump bucket. Scientific American, 32(19):290, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
290c.pdf.

Anonymous:1875:NGF

[7026] Anonymous. The new gold fields. Scientific American, 32(19):291, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-291.pdf.

Anonymous:1875:ISVa

[7027] Anonymous. Improved safety valve. Scientific American, 32(19): 291–292, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
291a.pdf.

Orton:1875:PA

[7028] James Orton. The palms of the Amazons. Scientific American, 32(19):292–293, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
292a.pdf.

Anonymous:1875:SPA

[7029] Anonymous. The screw pine of the Amazons. Scientific American, 32(19):293, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
293.pdf.

Anonymous:1875:IWW

[7030] Anonymous. Improved waterr wheel governor. Scientific American, 32(19):294, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
294.pdf.

Anonymous:1875:IMTa

[7031] Anonymous. Improved meat tenderer. Scientific American, 32(19):294, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-294a.pdf.

Anonymous:1875:NDC

[7032] Anonymous. New design for a conservatory. Scientific American, 32(19):295, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
295.pdf.

Anonymous:1875:SH

[7033] Anonymous. Summer houses. Scientific American, 32(19):295–296, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-295a.pdf.

Anonymous:1875:ANe

[7034] Anonymous. Astronomical notes. Scientific American, 32(19):296, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-296.pdf.

Anonymous:1875:DCl

[7035] Anonymous. Decisions of the courts. Scientific American, 32(19):296, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n19/pdf/scientificamerican05081875-296a.pdf.

Anonymous:1875:RAFs

[7036] Anonymous. Recent American and foreign patents. Scientific American, 32(19):296–297, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
296b.pdf.

Anonymous:1875:CRs

[7037] Anonymous. Communications received. Scientific American, 32 (19):299, May 8, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n19/pdf/scientificamerican05081875-
299.pdf.

Anonymous:1875:Cv

[7038] Anonymous. Correspondence. Scientific American, 32(20):308, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-308.pdf.

Anonymous:1875:BNBj

[7039] Anonymous. Books: New books and publications. Scientific American, 32(20):312–313, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
312b.pdf.

Anonymous:1875:BPt

[7040] Anonymous. Business and personal. Scientific American, 32(20):314, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-314.pdf.

Anonymous:1875:NQs

[7041] Anonymous. Notes and queries. Scientific American, 32(20):314–315, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-314a.pdf.

Anonymous:1875:IIt

[7042] Anonymous. Index of inventions. Scientific American, 32(20):315–316, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-315a.pdf.

Anonymous:1875:ISCa

[7043] Anonymous. Improved shears for cutting bar iron. Scientific American, 32(20):303, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
303.pdf.

Anonymous:1875:GI

[7044] Anonymous. The genesis of inventors. Scientific American, 32(20):304, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-304.pdf.

Anonymous:1875:DSC

[7045] Anonymous. Death in the salt cellar. Scientific American, 32(20):304, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-304a.pdf.

Anonymous:1875:ADC

[7046] Anonymous. The agricultural display at the Centennial. Scientific American, 32(20):304–305, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
304b.pdf.

Anonymous:1875:MBE

[7047] Anonymous. The management of boilers and engines. Scientific American, 32(20):305, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
305.pdf.

Anonymous:1875:ISPc

[7048] Anonymous. Improved surface planer. Scientific American, 32(20):306, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-306.pdf.

Anonymous:1875:ARC

[7049] Anonymous. Apparatus for removing condensed steam from pipes. Scientific American, 32(20):306, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
306a.pdf.

Anonymous:1875:NYR

[7050] Anonymous. The New York river front improvements. Scientific American, 32(20):307–308, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
307.pdf.

Anonymous:1875:SPS

[7051] Anonymous. Scott’s patent sheet iron roofing. Scientific American, 32(20):309, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
309.pdf.

Anonymous:1875:IWMb

[7052] Anonymous. Improved washing machine. Scientific American, 32 (20):310, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
310.pdf.

Anonymous:1875:WBH

[7053] Anonymous. Wolff’s button hole cutter. Scientific American, 32 (20):310, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
310a.pdf.

Anonymous:1875:ILWb

[7054] Anonymous. Improved low water indicator. Scientific American, 32(20):310, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
310b.pdf.

Anonymous:1875:NDP

[7055] Anonymous. Novel device for preventing runaway horses. Scientific American, 32(20):311, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
311.pdf.

Anonymous:1875:LDD

[7056] Anonymous. The little dodo or dodlet. Scientific American, 32 (20):311–312, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
311a.pdf.

Anonymous:1875:MB

[7057] Anonymous. The Meidinger battery. Scientific American, 32(20):312, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-312.pdf.

Anonymous:1875:DCm

[7058] Anonymous. Decisions of the courts. Scientific American, 32(20):312, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n20/pdf/scientificamerican05151875-312a.pdf.

Anonymous:1875:RAFt

[7059] Anonymous. Recent American and foreign patents. Scientific American, 32(20):313, May 15, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
313.pdf.

Anonymous:1875:CRt (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n20/pdf/scientificamerican05151875-
315.pdf.

Anonymous:1875:Cw

[7061] Anonymous. Correspondence. Scientific American, 32(21):324, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-324.pdf.

Anonymous:1875:BPu

[7062] Anonymous. Business and personal. Scientific American, 32(21):330, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-330.pdf.

Anonymous:1875:NQt

[7063] Anonymous. Notes and queries. Scientific American, 32(21):330–332, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-330a.pdf.

Anonymous:1875:IIu

[7064] Anonymous. Index of inventions. Scientific American, 32(21):332, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-332a.pdf.

Anonymous:1875:IFMa

[7065] Anonymous. Improved fulling mill. Scientific American, 32(21):319, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-319.pdf.

Anonymous:1875:PNA

[7066] Anonymous. Proceedings of the National Academy of Sciences. Scientific American, 32(21):319, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
319a.pdf.

Anonymous:1875:SOa

[7067] Anonymous. Schools of observation. Scientific American, 32(21):320, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-320.pdf.

Anonymous:1875:AL

[7068] Anonymous. The antiquity of life. Scientific American, 32(21):320, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-320a.pdf.

Anonymous:1875:DHa

[7069] Anonymous. Dangerous houses. Scientific American, 32(21):320, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-320b.pdf.

Anonymous:1875:CAE

[7070] Anonymous. The coming Arctic Expeditions. Scientific American, 32(21):320–321, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
320c.pdf.

Anonymous:1875:PLE

[7071] Anonymous. Patent litigation in England and the United States. Scientific American, 32(21):321, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
321.pdf.

Anonymous:1875:DB

[7072] Anonymous. Deadly ballooning. Scientific American, 32(21):321, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-321a.pdf.

Anonymous:1875:LSb

[7073] Anonymous. The laws of storms. Scientific American, 32(21):321, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-321b.pdf.

Anonymous:1875:ICT

[7074] Anonymous. Improved chisel tooth saw. Scientific American, 32 (21):322, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
322.pdf.

Anonymous:1875:EVM

[7075] Anonymous. Engine valve motion. Scientific American, 32(21):322, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-322a.pdf.

Anonymous:1875:IAI

[7076] Anonymous. Improved attachment for injectors. Scientific American, 32(21):323, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
323.pdf.

Anonymous:1875:SISa

[7077] Anonymous. Stewart’s improved stove damper. Scientific American, 32(21):323, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
323a.pdf.

Anonymous:1875:ITT

[7078] Anonymous. Improved turpentine tool. Scientific American, 32 (21):323–324, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
323b.pdf.

Anonymous:1875:SPIl

[7079] Anonymous. Scientific and practical information. Scientific American, 32(21):324, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
324a.pdf.

Rose:1875:PMh

[7080] Joshua Rose. Practical mechanism. Scientific American, 32(21):325, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-325.pdf.

Anonymous:1875:ISS

[7081] Anonymous. Improved Skein setter. Scientific American, 32(21):326, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-326.pdf.

Anonymous:1875:IYS

[7082] Anonymous. Improved yarn spooler. Scientific American, 32(21):326, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-326a.pdf.

Anonymous:1875:SF

[7083] Anonymous. The scorpion fly. Scientific American, 32(21):327, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n21/pdf/scientificamerican05221875-327.pdf.

Anonymous:1875:DHM

[7084] Anonymous. The Death’s head moth. Scientific American, 32 (21):327–328, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
327a.pdf.

Anonymous:1875:CAC

[7085] Anonymous. Car axles and couplings. Scientific American, 32 (21):328–329, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
328.pdf.

Anonymous:1875:RAFu

[7086] Anonymous. Recent American and foreign patents. Scientific American, 32(21):329, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
329.pdf.

Anonymous:1875:CRu

[7087] Anonymous. Communications received. Scientific American, 32 (21):332, May 22, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n21/pdf/scientificamerican05221875-
332.pdf.

Anonymous:1875:Cx

[7088] Anonymous. Correspondence. Scientific American, 32(22):340, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-340.pdf.

Anonymous:1875:BPv

[7089] Anonymous. Business and personal. Scientific American, 32(22):347, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-347.pdf.

Anonymous:1875:NQu

[7090] Anonymous. Notes and queries. Scientific American, 32(22):347–348, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-347a.pdf.

Anonymous:1875:IIv

[7091] Anonymous. Index of inventions. Scientific American, 32(22):348, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-348a.pdf.

Anonymous:1875:VD

[7092] Anonymous. The Virginia ditcher. Scientific American, 32(22):335, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-335.pdf.

Anonymous:1875:W

[7093] Anonymous. A want. Scientific American, 32(22):336, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-336.pdf.

Anonymous:1875:PCV

[7094] Anonymous. A practical cure for vagrancy. Scientific American, 32(22):336, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
336a.pdf.

Anonymous:1875:WSa

[7095] Anonymous. The wreck of the Schiller. Scientific American, 32 (22):336–337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
336b.pdf.

Anonymous:1875:SFC

[7096] Anonymous. A success for the Fishery Commission. Scientific American, 32(22):337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
337.pdf.

Anonymous:1875:DPTa

[7097] Anonymous. Design patents and trade marks. Scientific American, 32(22):337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
337a.pdf.

Anonymous:1875:NIP

[7098] Anonymous. The new international postal rates. Scientific American, 32(22):337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
337b.pdf.

Anonymous:1875:STB

[7099] Anonymous. Successful transplantation of bone. Scientific American, 32(22):337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
337c.pdf.

Anonymous:1875:ONU

[7100] Anonymous. Our National University of Technology. Scientific American, 32(22):337, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
337d.pdf.

Anonymous:1875:IWR

[7101] Anonymous. Improved walk and road cleaner. Scientific American, 32(22):338, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
338.pdf.

Anonymous:1875:MLP

[7102] Anonymous. The Merriman life-preserving dress. Scientific American, 32(22):338, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
338a.pdf.

Anonymous:1875:IBFa

[7103] Anonymous. Improved boiler feed regulator. Scientific American, 32(22):339, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
339.pdf.

Anonymous:1875:HW

[7104] Anonymous. Hight of waves. Scientific American, 32(22):339, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-339a.pdf.

Anonymous:1875:MPW

[7105] Anonymous. Motive power from waves. Scientific American, 32 (22):339–340, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
339b.pdf.

Anonymous:1875:SRb

[7106] Anonymous. Sideboard refrigerators. Scientific American, 32(22):341, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-341.pdf.

Anonymous:1875:AGB

[7107] Anonymous. Arrangement of galvanic batteries. Scientific American, 32(22):341, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
341a.pdf.

Anonymous:1875:PE

[7108] Anonymous. Psycho exposed. Scientific American, 32(22):341, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-341b.pdf.

Anonymous:1875:IHB

[7109] Anonymous. Improved hub borer. Scientific American, 32(22):342, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-342.pdf.

Anonymous:1875:SBSa

[7110] Anonymous. Smith’s blind slat holder. Scientific American, 32(22):342, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-342a.pdf.

Anonymous:1875:SISb

[7111] Anonymous. Smith’s improved sheave. Scientific American, 32(22):342, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-342b.pdf.

Anonymous:1875:OG

[7112] Anonymous. An ornamental gourd. Scientific American, 32(22):343, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-343.pdf.

Anonymous:1875:SBH

[7113] Anonymous. A Spring button hole bouquet. Scientific American, 32(22):343, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
343a.pdf.

Anonymous:1875:CHa

[7114] Anonymous. Cottage horticulture. Scientific American, 32(22):343–344, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-343b.pdf.

Anonymous:1875:SPIm

[7115] Anonymous. Scientific and practical information. Scientific American, 32(22):344, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
344.pdf.

Anonymous:1875:OCS

[7116] Anonymous. The origin of cold snaps. Scientific American, 32(22):345, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n22/pdf/scientificamerican05291875-345.pdf.

Anonymous:1875:DCn

[7117] Anonymous. Decisions of the courts. Scientific American, 32 (22):345–346, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
345a.pdf.

Anonymous:1875:RAFv

[7118] Anonymous. Recent American and foreign patents. Scientific American, 32(22):346, May 29, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
346.pdf.

Anonymous:1875:CRv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n22/pdf/scientificamerican05291875-
348.pdf.

Anonymous:1875:BPw

[7120] Anonymous. Business and personal. Scientific American, 32(23):362, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-362.pdf.

Anonymous:1875:NQv

[7121] Anonymous. Notes and queries. Scientific American, 32(23):362–363, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-362a.pdf.

Anonymous:1875:IIw

[7122] Anonymous. Index of inventions. Scientific American, 32(23):364, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-364.pdf.

Anonymous:1875:NSG

[7123] Anonymous. A new system of gas lighting. Scientific American, 32(23):351, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
351.pdf.

Anonymous:1875:OOE

[7124] Anonymous. The organic origin of the Earth’s crust. Scientific American, 32(23):352, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
352.pdf.

Anonymous:1875:PRTb

[7125] Anonymous. Progress of rapid transit in New York City. Scientific American, 32(23):352, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
352a.pdf.

Anonymous:1875:GG

[7126] Anonymous. Great guns. Scientific American, 32(23):352–353, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-352b.pdf.

Anonymous:1875:DW

[7127] Anonymous. The distribution of wealth. Scientific American, 32 (23):353, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
353.pdf.

Anonymous:1875:DG

[7128] Anonymous. The durability of glass. Scientific American, 32(23):353, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-353a.pdf.

Anonymous:1875:SPIn

[7129] Anonymous. Scientific and practical information. Scientific American, 32(23):353, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
353b.pdf.

Anonymous:1875:BCS

[7130] Anonymous. The Bessemer channel steamer. Scientific American, 32(23):354–356, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
354.pdf.

Anonymous:1875:ANf

[7131] Anonymous. Astronomical notes. Scientific American, 32(23):356–357, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-356.pdf.

Rose:1875:PMi

[7132] Joshua Rose. Practical mechanism. Scientific American, 32(23):357, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-357.pdf.

Anonymous:1875:IBG

[7133] Anonymous. Improved belt-geared cotton press. Scientific American, 32(23):358, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
358.pdf.

Anonymous:1875:NSIc

[7134] Anonymous. New sharpening instrument. Scientific American, 32(23):358, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
358a.pdf.

Anonymous:1875:IBE

[7135] Anonymous. Improved bucket ear. Scientific American, 32(23):358, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-358b.pdf.

Anonymous:1875:IHa

[7136] Anonymous. Improved hydrant. Scientific American, 32(23):358, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-358c.pdf.

Anonymous:1875:FBA

[7137] Anonymous. The fatal balloon ascension. Scientific American, 32(23):359, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
359.pdf.

Anonymous:1875:HAB

[7138] Anonymous. Hood’s adjustable brush handle. Scientific American, 32(23):360, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
360.pdf.

Anonymous:1875:ESA

[7139] Anonymous. Eldridge’s self-acting key fastener. Scientific American, 32(23):360, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
360a.pdf.

Anonymous:1875:CWa

[7140] Anonymous. Car wheels. Scientific American, 32(23):360, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n23/pdf/scientificamerican06051875-360b.pdf.

Anonymous:1875:HSS

[7141] Anonymous. Hand support for sheep shears. Scientific American, 32(23):360–361, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
360c.pdf.

Anonymous:1875:RAFw

[7142] Anonymous. Recent American and foreign patents. Scientific American, 32(23):361, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
361.pdf.

Anonymous:1875:CRw

[7143] Anonymous. Communications received. Scientific American, 32 (23):363, June 5, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n23/pdf/scientificamerican06051875-
363.pdf.

Anonymous:1875:Cy

[7144] Anonymous. Correspondence. Scientific American, 32(24):372, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-372.pdf.

Anonymous:1875:BPx

[7145] Anonymous. Business and personal. Scientific American, 32(24):378, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-378.pdf.

Anonymous:1875:NQw

[7146] Anonymous. Notes and queries. Scientific American, 32(24):378–379, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-378a.pdf.

Anonymous:1875:IIx

[7147] Anonymous. Index of inventions. Scientific American, 32(24):380, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-380.pdf.

Anonymous:1875:ICSb

[7148] Anonymous. Improved compound steam boiler. Scientific American, 32(24):367, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
367.pdf.

Anonymous:1875:DTH

[7149] Anonymous. Driving on tyres by hydraulic pressure. Scientific American, 32(24):367, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
367a.pdf.

Anonymous:1875:TVV

[7150] Anonymous. Two views of the voyage of life. Scientific American, 32(24):368, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
368.pdf.

Anonymous:1875:SSW

[7151] Anonymous. The secret of Spring weather. Scientific American, 32(24):368, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
368a.pdf.

Anonymous:1875:RI

[7152] Anonymous. The right to interfere. Scientific American, 32 (24):368–369, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
368b.pdf.

Anonymous:1875:IHc

[7153] Anonymous. The iron horse. Scientific American, 32(24):369, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-369.pdf.

Anonymous:1875:GIW

[7154] Anonymous. Grasshopper inventions wanted. Scientific American, 32(24):369, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
369a.pdf.

Anonymous:1875:CIA

[7155] Anonymous. Climate of the ice age. Scientific American, 32(24):369, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-369b.pdf.

Anonymous:1875:ZHE

[7156] Anonymous. Zimmerman’s hydro-electric lamp and extinguisher. Scientific American, 32(24):370, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
370.pdf.

Anonymous:1875:DSW

[7157] Anonymous. The diseases of the silk worm. Scientific American, 32(24):370, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
370a.pdf.

Anonymous:1875:CEI

[7158] Anonymous. Curious experiment in instantaneous crystallization. Scientific American, 32(24):370, June 12, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v32/
n24/pdf/scientificamerican06121875-370b.pdf.

Anonymous:1875:ISGb

[7159] Anonymous. Improved saw gummer. Scientific American, 32(24):371, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-371.pdf.

Anonymous:1875:STW

[7160] Anonymous. Spar torpedo warfare. Scientific American, 32(24): 371–372, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
371a.pdf.

Anonymous:1875:HHa

[7161] Anonymous. Household hints. — I. Scientific American, 32 (24):372–373, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
372a.pdf.

Anonymous:1875:ANg

[7162] Anonymous. Arctic notes. Scientific American, 32(24):373, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-373.pdf.

Anonymous:1875:IEa

[7163] Anonymous. Improved elevator. Scientific American, 32(24):374, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-374.pdf.

Anonymous:1875:IBFb

[7164] Anonymous. Improved bill file. Scientific American, 32(24):374, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-374a.pdf.

Anonymous:1875:EG

[7165] Anonymous. English gardens. Scientific American, 32(24):375, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-375.pdf.

Anonymous:1875:BBa

[7166] Anonymous. The birds of Brazil. Scientific American, 32(24):375–376, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n24/pdf/scientificamerican06121875-375a.pdf.

Anonymous:1875:RAFx

[7167] Anonymous. Recent American and foreign patents. Scientific American, 32(24):376–377, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
376.pdf.

Anonymous:1875:CRx

[7168] Anonymous. Communications received. Scientific American, 32 (24):379–380, June 12, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n24/pdf/scientificamerican06121875-
379.pdf.

Anonymous:1875:Cz

[7169] Anonymous. Correspondence. Scientific American, 32(25):388, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-388.pdf.

Anonymous:1875:BPy

[7170] Anonymous. Business and personal. Scientific American, 32(25):395, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-395.pdf.

Anonymous:1875:NQx

[7171] Anonymous. Notes and queries. Scientific American, 32(25):395, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-395a.pdf.

Anonymous:1875:IIy

[7172] Anonymous. Index of inventions. Scientific American, 32(25):396, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-396.pdf.

Anonymous:1875:ICGb

[7173] Anonymous. Improved cotton gin feeder. Scientific American, 32(25):383, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
383.pdf.

Anonymous:1875:CWT

[7174] Anonymous. Condensed water trap. Scientific American, 32(25):383, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-383a.pdf.

Anonymous:1875:PPO

[7175] Anonymous. Patent politics in Ohio. Scientific American, 32(25):384, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-384.pdf.

Anonymous:1875:CUa

[7176] Anonymous. Civilization by upheaval. Scientific American, 32(25):384, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-384a.pdf.

Anonymous:1875:PGS

[7177] Anonymous. Progress of the great suspension bridge between New York and Brooklyn. Scientific American, 32(25):384, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-384b.pdf.

Anonymous:1875:RMA

[7178] Anonymous. Refinements in modern astronomical observations. Scientific American, 32(25):384–385, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
384c.pdf.

Anonymous:1875:SSe

[7179] Anonymous. Samson or Solomon? Scientific American, 32(25):385, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-385.pdf.

Anonymous:1875:SPIo

[7180] Anonymous. Scientific and practical information. Scientific American, 32(25):385, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
385a.pdf.

Anonymous:1875:FRSa

[7181] Anonymous. The fireless railway system in England. Scientific American, 32(25):386, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
386.pdf.

Anonymous:1875:SMb

[7182] Anonymous. Stenography by machinery. Scientific American, 32 (25):387, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
387.pdf.

Anonymous:1875:CGa

[7183] Anonymous. Cooking by gas. Scientific American, 32(25):387, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-387a.pdf.

Anonymous:1875:Ae

[7184] Anonymous. The Alexandra. Scientific American, 32(25):387, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-387b.pdf.

Anonymous:1875:B

[7185] Anonymous. Brachiospongia. Scientific American, 32(25):387–388, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-387c.pdf.

Anonymous:1875:NGA

[7186] Anonymous. Novel gas apparatus at the Paris grand opera. Scientific American, 32(25):388, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
388a.pdf.

Anonymous:1875:ICPb

[7187] Anonymous. Improvement in the construction of prisons. Scientific American, 32(25):388, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
388b.pdf.

Rose:1875:PMj

[7188] Joshua Rose. Practical mechanism. Scientific American, 32(25):389, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-389.pdf.

Anonymous:1875:IPCc

[7189] Anonymous. Improved power couplings. Scientific American, 32 (25):390, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
390.pdf.

Anonymous:1875:IBH

[7190] Anonymous. Improved barrel hoop. Scientific American, 32(25):390, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-390a.pdf.

Anonymous:1875:ADV

[7191] Anonymous. Abbe’s dish and vegetable drainer. Scientific American, 32(25):390, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
390b.pdf.

Anonymous:1875:FCP

[7192] Anonymous. Fuller’s combination packing. Scientific American, 32(25):390, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
390c.pdf.

Anonymous:1875:HWC

[7193] Anonymous. The homes of the working classes. Scientific American, 32(25):391–392, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
391.pdf.

Anonymous:1875:R

[7194] Anonymous. The radiometer. Scientific American, 32(25):392–393, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-392.pdf.

Anonymous:1875:HHI

[7195] Anonymous. Household hints — II. Scientific American, 32(25):393, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-393.pdf.

Anonymous:1875:DCo

[7196] Anonymous. Decisions of the courts. Scientific American, 32(25):393, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n25/pdf/scientificamerican06191875-393a.pdf.

Anonymous:1875:RAFy

[7197] Anonymous. Recent American and foreign patents. Scientific American, 32(25):393–394, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
393b.pdf.

Anonymous:1875:CRy

[7198] Anonymous. Communications received. Scientific American, 32 (25):395, June 19, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n25/pdf/scientificamerican06191875-
395b.pdf.

Anonymous:1875:Cba

[7199] Anonymous. Correspondence. Scientific American, 32(26):404, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-404.pdf.

Anonymous:1875:BNBk

[7200] Anonymous. Books: New books and publications. Scientific American, 32(26):408, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
408a.pdf.

Anonymous:1875:BPz

[7201] Anonymous. Business and personal. Scientific American, 32(26):409, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-409.pdf.

Anonymous:1875:NQy

[7202] Anonymous. Notes and queries. Scientific American, 32(26):409–410, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-409a.pdf.

Anonymous:1875:IIz

[7203] Anonymous. Index of inventions. Scientific American, 32(26):410–411, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-410a.pdf.

Anonymous:1875:PNP

[7204] Anonymous. Patent notes: Patents. Scientific American, 32(26):411, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-411.pdf.

Anonymous:1875:Ig

[7205] Anonymous. Index. Scientific American, 32(26):411–412, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-411a.pdf.

Anonymous:1875:GCR

[7206] Anonymous. The grand central railway depot. Scientific American, 32(26):399, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
399.pdf.

Anonymous:1875:BPTa

[7207] Anonymous. A beautiful palm tree. Scientific American, 32(26):399, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-399a.pdf.

Anonymous:1875:EAV

[7208] Anonymous. The end of another volume. Scientific American, 32(26):400, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
400.pdf.

Anonymous:1875:KMDa

[7209] Anonymous. The Keely motor deception. Scientific American, 32 (26):400–401, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
400a.pdf.

Anonymous:1875:BPGa

[7210] Anonymous. The Bastie patent glass. Scientific American, 32(26):401, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-401.pdf.

Anonymous:1875:NFP

[7211] Anonymous. A new form of patent litigation. Scientific American, 32(26):401, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
401a.pdf.

Anonymous:1875:CSa

[7212] Anonymous. Commercial sponges. Scientific American, 32(26):401, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-401b.pdf.

Anonymous:1875:NRT

[7213] Anonymous. A new registering thermometer. Scientific American, 32(26):402, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
402.pdf.

Anonymous:1875:BTG

[7214] Anonymous. The Bastie toughened glass. Scientific American, 32(26):402, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
402a.pdf.

Anonymous:1875:GFS

[7215] Anonymous. A gigantic ferry system. Scientific American, 32(26):403, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-403.pdf.

Anonymous:1875:MFL

[7216] Anonymous. The motive force of light. Scientific American, 32 (26):403–404, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
403a.pdf.

Anonymous:1875:EMb

[7217] Anonymous. Electro-metallurgy. Scientific American, 32(26):405, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-405.pdf.

Anonymous:1875:IHCa

[7218] Anonymous. Improved hand cart. Scientific American, 32(26):406, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-406.pdf.

Anonymous:1875:DIV

[7219] Anonymous. Dunn’s improved vine rake. Scientific American, 32(26):406, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
406a.pdf.

Anonymous:1875:SSS

[7220] Anonymous. Smith’s steam sentinel. Scientific American, 32(26):406, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-406b.pdf.

Anonymous:1875:HIG

[7221] Anonymous. Harrington’s improved gun sight. Scientific American, 32(26):406, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
406c.pdf.

Anonymous:1875:Sd

[7222] Anonymous. Swimming. Scientific American, 32(26):407, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-407.pdf.

Anonymous:1875:GF

[7223] Anonymous. The grouse family. Scientific American, 32(26):407–408, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-407a.pdf.

Anonymous:1875:DCp

[7224] Anonymous. Decisions of the courts. Scientific American, 32(26):408, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v32/n26/pdf/scientificamerican06261875-408.pdf.

Anonymous:1875:RAFz

[7225] Anonymous. Recent American and foreign patents. Scientific American, 32(26):408, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
408b.pdf.

Anonymous:1875:CRz

[7226] Anonymous. Communications received. Scientific American, 32 (26):410, June 26, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v32/n26/pdf/scientificamerican06261875-
410.pdf.

Anonymous:1875:Cbb

[7227] Anonymous. Correspondence. Scientific American, 33(1):8–9, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-8a.pdf.

Anonymous:1875:BPba

[7228] Anonymous. Business and personal. Scientific American, 33(1):11, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-11.pdf.

Anonymous:1875:NQz

[7229] Anonymous. Notes and queries. Scientific American, 33(1):11–12, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-11a.pdf.

Anonymous:1875:IIba

[7230] Anonymous. Index of inventions. Scientific American, 33(1):12, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-12a.pdf.

Anonymous:1875:HDS

[7231] Anonymous. Hot drafts for steam boiler furnaces. Scientific American, 33(1):1, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n1/pdf/scientificamerican07031875-
1.pdf.

Anonymous:1875:CHb

[7232] Anonymous. Cremation in the household. Scientific American, 33(1): 2, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-2.pdf.

Anonymous:1875:TMb

[7233] Anonymous. Tainted meat. Scientific American, 33(1):2, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-2a.pdf.

Anonymous:1875:REE

[7234] Anonymous. Rights of employer and employed in respect to a new invention. Scientific American, 33(1):2, July 3, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n1/pdf/scientificamerican07031875-2b.pdf.

Anonymous:1875:KMDb

[7235] Anonymous. The Keely motor deception. Scientific American, 33(1):2, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-2c.pdf.

Anonymous:1875:SPb

[7236] Anonymous. The search for the pole. Scientific American, 33(1):2–3, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-2d.pdf.

Anonymous:1875:HSAb

[7237] Anonymous. The hydrology of South Africa. Scientific American, 33(1): 3, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-3.pdf.

Anonymous:1875:SPIp

[7238] Anonymous. Scientific and practical information. Scientific American, 33(1):3, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n1/pdf/scientificamerican07031875-
3a.pdf.

Anonymous:1875:ITC

[7239] Anonymous. Improved three-cylinder pump. Scientific American, 33(1): 4, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-4.pdf.

Anonymous:1875:REF

[7240] Anonymous. Rudimentary existence in fresh water streams. Scientific American, 33(1):4, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n1/pdf/scientificamerican07031875-
4a.pdf.

Anonymous:1875:RFB

[7241] Anonymous. Revolving fire bars. Scientific American, 33(1):5, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-5.pdf.

Anonymous:1875:IHH

[7242] Anonymous. Improved horse hay rake. Scientific American, 33(1):6, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-6.pdf.

Anonymous:1875:INB

[7243] Anonymous. Improved napper and brusher. Scientific American, 33(1):6–7, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n1/pdf/scientificamerican07031875-
6a.pdf.

Anonymous:1875:APb

[7244] Anonymous. The animalism of plants. Scientific American, 33(1):7–8, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-7.pdf.

Anonymous:1875:NMEa

[7245] Anonymous. A new magneto-electric engine. Scientific American, 33(1): 8, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-8.pdf.

Anonymous:1875:DCq

[7246] Anonymous. Decisions of the courts. Scientific American, 33(1):9, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-9.pdf.

Anonymous:1875:RAFba

[7247] Anonymous. Recent American and foreign patents. Scientific American, 33(1):10, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n1/pdf/scientificamerican07031875-
10.pdf.

Anonymous:1875:CRba

[7248] Anonymous. Communications received. Scientific American, 33(1):12, July 3, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-12.pdf.

Anonymous:1875:CPr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n1/pdf/scientificamerican07031875-12b.pdf.

Anonymous:1875:Cbc

[7250] Anonymous. Correspondence. Scientific American, 33(2):20–21, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-20.pdf.

Anonymous:1875:BPbb

[7251] Anonymous. Business and personal. Scientific American, 33(2):26, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-26.pdf.

Anonymous:1875:NQba

[7252] Anonymous. Notes and queries. Scientific American, 33(2):26–27, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-26a.pdf.

Anonymous:1875:IIbb

[7253] Anonymous. Index of inventions. Scientific American, 33(2):28, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-28.pdf.

Anonymous:1875:IFMb

[7254] Anonymous. Improved flanging machine. Scientific American, 33(2):15, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-15.pdf.

Anonymous:1875:DWL

[7255] Anonymous. Duplex wheel lathe. Scientific American, 33(2):15, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-15a.pdf.

Anonymous:1875:TSa

[7256] Anonymous. The trustworthiness of the senses. Scientific American, 33(2):16, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
16.pdf.

Anonymous:1875:KMDc

[7257] Anonymous. The Keely motor deception. Scientific American, 33(2):16–17, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
16a.pdf.

Anonymous:1875:NHS

[7258] Anonymous. The natural history of swindles. Scientific American, 33(2):17, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
17.pdf.

Anonymous:1875:RIG

[7259] Anonymous. Recent improvements in glass manufacture. Scientific American, 33(2):17, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
17a.pdf.

Anonymous:1875:PA

[7260] Anonymous. The planetary atmospheres. Scientific American, 33(2):17, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-17b.pdf.

Anonymous:1875:EBM

[7261] Anonymous. Engine for blooming mill use. Scientific American, 33(2):18, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
18.pdf.

Anonymous:1875:IRS

[7262] Anonymous. Improved rope socket. Scientific American, 33(2):18–19, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-18a.pdf.

Anonymous:1875:NGB

[7263] Anonymous. A new gas blowpipe. Scientific American, 33(2):19, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-19.pdf.

Anonymous:1875:NSRa

[7264] Anonymous. Nessle’s street railway rail. Scientific American, 33(2):19, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-19a.pdf.

Anonymous:1875:ANh

[7265] Anonymous. Astronomical notes. Scientific American, 33(2):19, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-19b.pdf.

Anonymous:1875:BRH

[7266] Anonymous. Baldwin’s rotary hand stamp. Scientific American, 33(2):19–20, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
19c.pdf.

Rose:1875:PMk

[7267] Joshua Rose. Practical mechanism. Scientific American, 33(2):21, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-21.pdf.

Anonymous:1875:SPIq

[7268] Anonymous. Scientific and practical information. Scientific American, 33(2):21, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
21a.pdf.

Anonymous:1875:IGP

[7269] Anonymous. Improved gang plow Spring. Scientific American, 33(2):22, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-22.pdf.

Anonymous:1875:ID

[7270] Anonymous. Improved dashboard. Scientific American, 33(2):22, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-22a.pdf.

Anonymous:1875:MR

[7271] Anonymous. Model residences. Scientific American, 33(2):23, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-23.pdf.

Anonymous:1875:BCP

[7272] Anonymous. Bullets, cannon projectiles, and cartridges. Scientific American, 33(2):24, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
24.pdf.

Anonymous:1875:RAFbb

[7273] Anonymous. Recent American and foreign patents. Scientific American, 33(2):25, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n2/pdf/scientificamerican07101875-
25.pdf.

Anonymous:1875:CRbb

[7274] Anonymous. Communications received. Scientific American, 33(2):27, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-27.pdf.

Anonymous:1875:CPs

[7275] Anonymous. Canadian patents. Scientific American, 33(2):28, July 10, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n2/pdf/scientificamerican07101875-28a.pdf.

Anonymous:1875:Cbd

[7276] Anonymous. Correspondence. Scientific American, 33(3):36–37, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-36.pdf.

Anonymous:1875:BPbc

[7277] Anonymous. Business and personal. Scientific American, 33(3):42, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-42.pdf.

Anonymous:1875:NQbb

[7278] Anonymous. Notes and queries. Scientific American, 33(3):42–44, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-42a.pdf.

Anonymous:1875:IIbc

[7279] Anonymous. Index of inventions. Scientific American, 33(3):44, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-44a.pdf.

Anonymous:1875:IMTb

[7280] Anonymous. Improved method of transporting bricks during manufacture. Scientific American, 33(3):31, July 17, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n3/pdf/scientificamerican07171875-31.pdf.

Anonymous:1875:WAE

[7281] Anonymous. Work for Arctic explorers. Scientific American, 33(3):32, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-32.pdf.

Anonymous:1875:IES

[7282] Anonymous. Instability of the Earth’s surface. Scientific American, 33(3):32, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
32a.pdf.

Anonymous:1875:CPBa

[7283] Anonymous. The Colorado potato bug. Scientific American, 33 (3):32–33, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
32b.pdf.

Anonymous:1875:PSEa

[7284] Anonymous. The power of small engines. Scientific American, 33(3):33, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-33.pdf.

Anonymous:1875:KMDd

[7285] Anonymous. The Keely motor deception. Scientific American, 33(3):33, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-33a.pdf.

Anonymous:1875:IEM

[7286] Anonymous. Improved engraving machines. Scientific American, 33(3):34, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
34.pdf.

Anonymous:1875:EL

[7287] Anonymous. The electric light. Scientific American, 33(3):35–36, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-35.pdf.

Anonymous:1875:IRTa

[7288] Anonymous. Improved railway track. Scientific American, 33(3):38, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-38.pdf.

Anonymous:1875:IFB

[7289] Anonymous. Improved fire box for locomotive furnaces. Scientific American, 33(3):38, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
38a.pdf.

Anonymous:1875:KPH

[7290] Anonymous. Knebworth Park, Hertfordshire, England. Scientific American, 33(3):39, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
39.pdf.

Anonymous:1875:SSf

[7291] Anonymous. The song of steam. Scientific American, 33(3):40, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-40.pdf.

Anonymous:1875:IARa

[7292] Anonymous. The Irish–American rifle contest. Scientific American, 33(3):40, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
40a.pdf.

Anonymous:1875:RAFbc

[7293] Anonymous. Recent American and foreign patents. Scientific American, 33(3):40–41, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n3/pdf/scientificamerican07171875-
40b.pdf.

Anonymous:1875:CRbc

[7294] Anonymous. Communications received. Scientific American, 33(3):44, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-44.pdf.

Anonymous:1875:CPt

[7295] Anonymous. Canadian patents. Scientific American, 33(3):44, July 17, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n3/pdf/scientificamerican07171875-44b.pdf.

Anonymous:1875:Cbf

[7296] Anonymous. Correspondence. Scientific American, 33(4):52–53, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-52.pdf.

Anonymous:1875:BPbd

[7297] Anonymous. Business and personal. Scientific American, 33(4):59, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-59.pdf.

Anonymous:1875:NQbc

[7298] Anonymous. Notes and queries. Scientific American, 33(4):59–60, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-59a.pdf.

Anonymous:1875:IIbd

[7299] Anonymous. Index of inventions. Scientific American, 33(4):60, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-60a.pdf.

Anonymous:1875:RHP

[7300] Anonymous. Riveting by hydraulic power. Scientific American, 33(4):47, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
47.pdf.

Anonymous:1875:TC

[7301] Anonymous. Taking a chance. Scientific American, 33(4):48, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-48.pdf.

Anonymous:1875:KMDe

[7302] Anonymous. The Keely motor deception. Scientific American, 33(4):48, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-48a.pdf.

Anonymous:1875:NPR

[7303] Anonymous. The new phylloxera Remedy. Scientific American, 33(4):48–49, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
48b.pdf.

Anonymous:1875:RCG

[7304] Anonymous. Rate and cause of glacial motion. Scientific American, 33(4):49, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
49.pdf.

Anonymous:1875:SPIr

[7305] Anonymous. Scientific and practical information. Scientific American, 33(4):49, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
49a.pdf.

Anonymous:1875:IHY

[7306] Anonymous. Improved horse yoke. Scientific American, 33(4):50, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-50.pdf.

Anonymous:1875:NPa

[7307] Anonymous. A new pyrometer. Scientific American, 33(4):50, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-50a.pdf.

Anonymous:1875:NAS

[7308] Anonymous. The new Arctic sledge. Scientific American, 33(4):51, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-51.pdf.

Anonymous:1875:IRCa

[7309] Anonymous. Improved railway car-washing machine. Scientific American, 33(4):51, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
51a.pdf.

Rose:1875:PMl

[7310] Joshua Rose. Practical mechanism. Scientific American, 33(4):53, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-53.pdf.

Anonymous:1875:AW

[7311] Anonymous. Artesian wells. Scientific American, 33(4):54–55, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-54.pdf.

Anonymous:1875:Fa

[7312] Anonymous. The finches. Scientific American, 33(4):55, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-55.pdf.

Anonymous:1875:NSRb

[7313] Anonymous. Nessle’s street railway rail. Scientific American, 33(4):55, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-55a.pdf.

Anonymous:1875:MN

[7314] Anonymous. Medical notes. Scientific American, 33(4):56, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-56.pdf.

Anonymous:1875:TCW

[7315] Anonymous. Testing the color of water. Scientific American, 33(4):57, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-57.pdf.

Anonymous:1875:HPW

[7316] Anonymous. How to put up a water closet. Scientific American, 33(4):57, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
57a.pdf.

Anonymous:1875:RAFbd

[7317] Anonymous. Recent American and foreign patents. Scientific American, 33(4):57–58, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n4/pdf/scientificamerican07241875-
57b.pdf.

Anonymous:1875:CRbd

[7318] Anonymous. Communications received. Scientific American, 33(4):60, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-60.pdf.

Anonymous:1875:CPu

[7319] Anonymous. Canadian patents. Scientific American, 33(4):60, July 24, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n4/pdf/scientificamerican07241875-60b.pdf.

Anonymous:1875:Cbg

[7320] Anonymous. Correspondence. Scientific American, 33(5):68–69, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-68.pdf.

Anonymous:1875:BNBl

[7321] Anonymous. Books: New books and publications. Scientific American, 33(5):73, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
73a.pdf.

Anonymous:1875:BPbe

[7322] Anonymous. Business and personal. Scientific American, 33(5):75, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-75.pdf.

Anonymous:1875:NQbd

[7323] Anonymous. Notes and queries. Scientific American, 33(5):75–76, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-75a.pdf.

Anonymous:1875:IIbe

[7324] Anonymous. Index of inventions. Scientific American, 33(5):76, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-76a.pdf.

Anonymous:1875:IBFc

[7325] Anonymous. Improved boiler and feed water heater. Scientific American, 33(5):63, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
63.pdf.

Anonymous:1875:SIc

[7326] Anonymous. The strength of insects. Scientific American, 33(5):64, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-64.pdf.

Anonymous:1875:RTN

[7327] Anonymous. Rapid transit in New York. Scientific American, 33(5):64, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-64a.pdf.

Anonymous:1875:COH

[7328] Anonymous. A city one hundred and eighty thousand years old. Scientific American, 33(5):64–65, July 31, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n5/pdf/scientificamerican07311875-64b.pdf.

Anonymous:1875:Af

[7329] Anonymous. Apparitions. Scientific American, 33(5):65, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-65.pdf.

Anonymous:1875:PSP

[7330] Anonymous. Prospects of science on the Pacific slope. Scientific American, 33(5):65, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
65a.pdf.

Anonymous:1875:DCDa

[7331] Anonymous. Dangers of chloral drinking. Scientific American, 33(5):65, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-65b.pdf.

Anonymous:1875:CRbe

[7332] Anonymous. Collegiate races. Scientific American, 33(5):65, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-65c.pdf.

Anonymous:1875:IARb

[7333] Anonymous. Improved air reservoir. Scientific American, 33(5):66, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-66.pdf.

Anonymous:1875:DPH

[7334] Anonymous. Device for protecting horses. Scientific American, 33(5):66, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
66a.pdf.

Anonymous:1875:NCT

[7335] Anonymous. New clinical thermoscope. Scientific American, 33(5):66, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-66b.pdf.

Anonymous:1875:ITS

[7336] Anonymous. Improved treadle sawing machine. Scientific American, 33(5):66, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
66c.pdf.

Anonymous:1875:ISVb

[7337] Anonymous. Improved slide valve calculator. Scientific American, 33(5):67–68, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
67.pdf.

Anonymous:1875:SBSb

[7338] Anonymous. Steller’s blind slat holder. Scientific American, 33(5):70, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-70.pdf.

Anonymous:1875:SIAb

[7339] Anonymous. Smith’s improved awl. Scientific American, 33(5):70, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-70a.pdf.

Anonymous:1875:ABF

[7340] Anonymous. Alley’s bearing feeler. Scientific American, 33(5):70, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-70b.pdf.

Anonymous:1875:IMTc

[7341] Anonymous. Improved milking tubes. Scientific American, 33(5):70, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-70c.pdf.

Anonymous:1875:CDC

[7342] Anonymous. Cavender and Dallas’ corn husking implement. Scientific American, 33(5):70, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
70d.pdf.

Anonymous:1875:Cbh

[7343] Anonymous. The cycads. Scientific American, 33(5):71, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-71.pdf.

Anonymous:1875:Fb

[7344] Anonymous. The ficaria. Scientific American, 33(5):71, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-71a.pdf.

Anonymous:1875:SPIs

[7345] Anonymous. Scientific and practical information. Scientific American, 33(5):72, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
72.pdf.

Anonymous:1875:JM

[7346] Anonymous. The jamin magnet. Scientific American, 33(5):72–73, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-72a.pdf.

Anonymous:1875:DCr

[7347] Anonymous. Decisions of the courts. Scientific American, 33(5):73, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-73.pdf.

Anonymous:1875:RAFbe

[7348] Anonymous. Recent American and foreign patents. Scientific American, 33(5):74, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n5/pdf/scientificamerican07311875-
74.pdf.

Anonymous:1875:CRbf

[7349] Anonymous. Communications received. Scientific American, 33(5):76, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-76.pdf.

Anonymous:1875:CPv

[7350] Anonymous. Canadian patents. Scientific American, 33(5):76, July 31, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n5/pdf/scientificamerican07311875-76b.pdf.

Anonymous:1875:Cbi

[7351] Anonymous. Correspondence. Scientific American, 33(6):84, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-84.pdf.

Anonymous:1875:BPbf

[7352] Anonymous. Business and personal. Scientific American, 33(6):91, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-91.pdf.

Anonymous:1875:NQbe

[7353] Anonymous. Notes and queries. Scientific American, 33(6):91–92, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-91a.pdf.

Anonymous:1875:IIbf

[7354] Anonymous. Index of inventions. Scientific American, 33(6):92, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-92a.pdf.

Anonymous:1875:IVR

[7355] Anonymous. Improvement in the ventilation of railway tunnels. Scientific American, 33(6):79, August 7, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n6/pdf/scientificamerican08071875-79.pdf.

Anonymous:1875:PLAb

[7356] Anonymous. Physical life after death. Scientific American, 33(6):80, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-80.pdf.

Anonymous:1875:KMDf

[7357] Anonymous. The Keely motor deception. Scientific American, 33(6):80, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-80a.pdf.

Anonymous:1875:EPL

[7358] Anonymous. English patent law discussion. Scientific American, 33(6):80–81, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
80b.pdf.

Anonymous:1875:CTM

[7359] Anonymous. Centennial trade marks. Scientific American, 33(6):81, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-81.pdf.

Anonymous:1875:DBW

[7360] Anonymous. The difference between water and nitro-glycerin. Scientific American, 33(6):81, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
81a.pdf.

Anonymous:1875:UWD

[7361] Anonymous. The use of wire in deep sea soundings. Scientific American, 33(6):81, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
81b.pdf.

Anonymous:1875:IEG

[7362] Anonymous. Improved elevator for granular matters. Scientific American, 33(6):82, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
82.pdf.

Anonymous:1875:FWB

[7363] Anonymous. Frozen water in bottles. Scientific American, 33(6):82, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-82a.pdf.

Anonymous:1875:Fc

[7364] Anonymous. Fastenings. Scientific American, 33(6):83–84, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-83.pdf.

Rose:1875:PMm

[7365] Joshua Rose. Practical mechanism. Scientific American, 33(6):84–85, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-84a.pdf.

Anonymous:1875:IRR

[7366] Anonymous. Improved rotary rag boiler. Scientific American, 33(6):86, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-86.pdf.

Anonymous:1875:CKH

[7367] Anonymous. Coleman’s key holder. Scientific American, 33(6):86, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-86a.pdf.

Anonymous:1875:SST

[7368] Anonymous. Seel and Shaw’s tweer for smith’s use. Scientific American, 33(6):86, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
86b.pdf.

Anonymous:1875:BCW

[7369] Anonymous. Brechbiel’s cigar wrapper knife. Scientific American, 33(6):86, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
86c.pdf.

Anonymous:1875:RCS

[7370] Anonymous. The Russian circular ship. Scientific American, 33 (6):87–89, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
87.pdf.

Anonymous:1875:DCs

[7371] Anonymous. Decisions of the courts. Scientific American, 33(6):89, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-89.pdf.

Anonymous:1875:RAFbf

[7372] Anonymous. Recent American and foreign patents. Scientific American, 33(6):90, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n6/pdf/scientificamerican08071875-
90.pdf.

Anonymous:1875:CRbg

[7373] Anonymous. Communications received. Scientific American, 33(6):92, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-92.pdf.

Anonymous:1875:CPw

[7374] Anonymous. Canadian patents. Scientific American, 33(6):92, August 7, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n6/pdf/scientificamerican08071875-92b.pdf.

Anonymous:1875:Cbj

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-100.pdf.

Anonymous:1875:BPbg

[7376] Anonymous. Business and personal. Scientific American, 33(7):107, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-107.pdf.

Anonymous:1875:NQbf

[7377] Anonymous. Notes and queries. Scientific American, 33(7):107– 108, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
107a.pdf.

Anonymous:1875:IIbg

[7378] Anonymous. Index of inventions. Scientific American, 33(7):108, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-108a.pdf.

Anonymous:1875:IWC

[7379] Anonymous. Improved wood-carving machine. Scientific American, 33(7):95, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
95.pdf.

Anonymous:1875:GIS

[7380] Anonymous. The great ice sheet. Scientific American, 33(7):96, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-96.pdf.

Anonymous:1875:SAa

[7381] Anonymous. Salicylic acid. Scientific American, 33(7):96, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-96a.pdf.

Anonymous:1875:TMc

[7382] Anonymous. Terrestrial magnetism. Scientific American, 33(7):97, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-97.pdf.

Anonymous:1875:YEB

[7383] Anonymous. A Year’s experience in boiler inspection. Scientific American, 33(7):97, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
97a.pdf.

Anonymous:1875:EPCa

[7384] Anonymous. An enormous peach crop. Scientific American, 33 (7):97, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
97b.pdf.

Anonymous:1875:Pc

[7385] Anonymous. The podometer. Scientific American, 33(7):98, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-98.pdf.

Anonymous:1875:PCa

[7386] Anonymous. A powerful crane. Scientific American, 33(7):98, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-98a.pdf.

Anonymous:1875:WC

[7387] Anonymous. Wicker coffins. Scientific American, 33(7):99, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-99.pdf.

Anonymous:1875:CSO

[7388] Anonymous. Casting steel in one tun ingots. Scientific American, 33(7):99–100, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
99a.pdf.

Anonymous:1875:ITG

[7389] Anonymous. Improved toy gun. Scientific American, 33(7):102, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-102.pdf.

Anonymous:1875:IFCa

[7390] Anonymous. Improved floor clamp. Scientific American, 33(7):102, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-102a.pdf.

Anonymous:1875:IPB

[7391] Anonymous. Improved pressure blocks for planing machines. Scientific American, 33(7):102–103, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
102b.pdf.

Anonymous:1875:HA

[7392] Anonymous. Hardy azaleas. Scientific American, 33(7):103, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-103.pdf.

Anonymous:1875:CAa

[7393] Anonymous. Cottage architecture. Scientific American, 33(7):103, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-103a.pdf.

Anonymous:1875:CPAa

[7394] Anonymous. Combined portable aquarium and Wardian case. Scientific American, 33(7):103–105, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
103b.pdf.

Anonymous:1875:RAFbg

[7395] Anonymous. Recent American and foreign patents. Scientific American, 33(7):105–106, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
105.pdf.

Anonymous:1875:CRbh

[7396] Anonymous. Communications received. Scientific American, 33 (7):108, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n7/pdf/scientificamerican08141875-
108.pdf.

Anonymous:1875:CPx

[7397] Anonymous. Canadian patents. Scientific American, 33(7):108, August 14, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n7/pdf/scientificamerican08141875-108b.pdf.

Anonymous:1875:Cbk

[7398] Anonymous. Correspondence. Scientific American, 33(8):116, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-116.pdf.

Anonymous:1875:BPbh

[7399] Anonymous. Business and personal. Scientific American, 33(8):123, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-123.pdf.

Anonymous:1875:NQbg

[7400] Anonymous. Notes and queries. Scientific American, 33(8):123– 124, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
123a.pdf.

Anonymous:1875:IIbh

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-124a.pdf.

Anonymous:1875:HGI

[7402] Anonymous. Hell Gate improvement. Scientific American, 33 (8):111–112, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
111.pdf.

Anonymous:1875:UN

[7403] Anonymous. The uses of nature. Scientific American, 33(8):112, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-112.pdf.

Anonymous:1875:USC

[7404] Anonymous. The United States Commission on Boiler Explosions. Scientific American, 33(8):112, August 21, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n8/pdf/scientificamerican08211875-112a.pdf.

Anonymous:1875:QC

[7405] Anonymous. Queer cattle. Scientific American, 33(8):112–113, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-112b.pdf.

Anonymous:1875:COL

[7406] Anonymous. Constancy of the ocean level. Scientific American, 33(8):113, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
113.pdf.

Anonymous:1875:ESB

[7407] Anonymous. Experimental steam boiler explosions. Scientific American, 33(8):113, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
113a.pdf.

Anonymous:1875:SPIt

[7408] Anonymous. Scientific and practical information. Scientific American, 33(8):113, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
113b.pdf.

Anonymous:1875:ERC

[7409] Anonymous. Expected resignation of the Commissioner of Patents. Scientific American, 33(8):113, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
113c.pdf.

Anonymous:1875:SLB

[7410] Anonymous. The St. Louis Bridge. Scientific American, 33(8):114, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-114.pdf.

Anonymous:1875:SIBa

[7411] Anonymous. Smith’s improved bale tie. Scientific American, 33 (8):115, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
115.pdf.

Anonymous:1875:KSJ

[7412] Anonymous. Korting’s steam jet bilge pump. Scientific American, 33(8):115, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
115a.pdf.

Anonymous:1875:JCB

[7413] Anonymous. Joining up cells of batteries. Scientific American, 33(8):115, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
115b.pdf.

Anonymous:1875:EIC

[7414] Anonymous. The electrostatic or induction coil. Scientific American, 33(8):115–116, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
115c.pdf.

Rose:1875:PMo

[7415] Joshua Rose. Practical mechanism. Scientific American, 33(8):117, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-117.pdf.

Anonymous:1875:IFCb

[7416] Anonymous. Improved floor clamp. Scientific American, 33(8):118, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-118.pdf.

Anonymous:1875:ISGc

[7417] Anonymous. Improved sudden grip vise. Scientific American, 33(8):118, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
118a.pdf.

Anonymous:1875:NRM

[7418] Anonymous. New reaping and mowing machinery. Scientific American, 33(8):119–121, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
119.pdf.

Wright:1875:CFF

[7419] C. W. Wright. The chemical fire-fly. Scientific American, 33(8):121, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-121.pdf.

Anonymous:1875:RAFbh

[7420] Anonymous. Recent American and foreign patents. Scientific American, 33(8):121–122, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
121a.pdf.

Anonymous:1875:CRbi

[7421] Anonymous. Communications received. Scientific American, 33 (8):124, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n8/pdf/scientificamerican08211875-
124.pdf.

Anonymous:1875:CPy

[7422] Anonymous. Canadian patents. Scientific American, 33(8):124, August 21, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n8/pdf/scientificamerican08211875-124b.pdf.

Anonymous:1875:Cbl

[7423] Anonymous. Correspondence. Scientific American, 33(9):132, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-132.pdf.

Anonymous:1875:BPbi

[7424] Anonymous. Business and personal. Scientific American, 33(9):138, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-138.pdf.

Anonymous:1875:NQbh

[7425] Anonymous. Notes and queries. Scientific American, 33(9):138– 140, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
138a.pdf.

Anonymous:1875:IIbi

[7426] Anonymous. Index of inventions. Scientific American, 33(9):140, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-140a.pdf.

Anonymous:1875:IFWa

[7427] Anonymous. Improved feed water heater and purifier. Scientific American, 33(9):127–128, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
127.pdf.

Anonymous:1875:EPB

[7428] Anonymous. The English patent bill. Scientific American, 33 (9):128, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
128.pdf.

Anonymous:1875:IWIb

[7429] Anonymous. The irrigation works of India. Scientific American, 33(9):128, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
128a.pdf.

Anonymous:1875:ECU

[7430] Anonymous. Is education for culture or for use? Scientific American, 33(9):128–129, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
128b.pdf.

Anonymous:1875:SRP

[7431] Anonymous. Spirit rifle practice. Scientific American, 33(9):129, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-129.pdf.

Anonymous:1875:KMDg

[7432] Anonymous. The Keely motor deception. Scientific American, 33(9):129, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
129a.pdf.

Anonymous:1875:BAEb

[7433] Anonymous. The British Arctic Expedition. Scientific American, 33(9):130–132, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
130.pdf.

Anonymous:1875:Ag

[7434] Anonymous. Augers. Scientific American, 33(9):133, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-133.pdf.

Anonymous:1875:SRS

[7435] Anonymous. Spayth’s rafter scale and bevel gage. Scientific American, 33(9):134, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
134.pdf.

Anonymous:1875:HSP

[7436] Anonymous. The Hercules screw propeller. Scientific American, 33(9):134, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
134a.pdf.

Anonymous:1875:NLR

[7437] Anonymous. New life raft. Scientific American, 33(9):134, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-134b.pdf.

Anonymous:1875:JIB

[7438] Anonymous. Jones’ improved blind stop. Scientific American, 33(9):134, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
134c.pdf.

Anonymous:1875:AEF

[7439] Anonymous. The ant-eater family. Scientific American, 33(9):135, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-135.pdf.

Anonymous:1875:CAM

[7440] Anonymous. The cobweb Apple moth. Scientific American, 33 (9):135–136, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
135a.pdf.

Anonymous:1875:AAAa

[7441] Anonymous. The American Association for the Advancement of Science. Scientific American, 33(9):136, August 28, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n9/pdf/scientificamerican08281875-136.pdf.

Anonymous:1875:DCt

[7442] Anonymous. Decisions of the courts. Scientific American, 33 (9):136, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
136a.pdf.

Anonymous:1875:RAFbi

[7443] Anonymous. Recent American and foreign patents. Scientific American, 33(9):137, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
137.pdf.

Anonymous:1875:CRbj

[7444] Anonymous. Communications received. Scientific American, 33 (9):140, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n9/pdf/scientificamerican08281875-
140.pdf.

Anonymous:1875:CPz

[7445] Anonymous. Canadian patents. Scientific American, 33(9):140, August 28, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n9/pdf/scientificamerican08281875-140b.pdf.

Anonymous:1875:Cbm

[7446] Anonymous. Correspondence. Scientific American, 33(10):148, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n10/pdf/scientificamerican09041875-148.pdf.

Anonymous:1875:BNBm

[7447] Anonymous. Books: New books and publications. Scientific American, 33(10):153, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
153a.pdf.

Anonymous:1875:BPbj

[7448] Anonymous. Business and personal. Scientific American, 33 (10):155, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
155.pdf.

Anonymous:1875:NQbi

[7449] Anonymous. Notes and queries. Scientific American, 33(10): 155–156, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
155a.pdf.

Anonymous:1875:IIbj

[7450] Anonymous. Index of inventions. Scientific American, 33(10): 156, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
156a.pdf.

Anonymous:1875:IPP

[7451] Anonymous. Improved power punch and shearing machine. Scientific American, 33(10):143, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
143.pdf.

Anonymous:1875:WDW

[7452] Anonymous. What do we see? Scientific American, 33(10):144, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n10/pdf/scientificamerican09041875-144.pdf.

Anonymous:1875:CPRa

[7453] Anonymous. The cause of profuse rains. Scientific American, 33 (10):144, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
144a.pdf.

Anonymous:1875:SBA

[7454] Anonymous. Steel bronze an American invention. Scientific American, 33(10):144–145, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
144b.pdf.

Anonymous:1875:DNB

[7455] Anonymous. Death in the nursing bottle. Scientific American, 33(10):145, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
145.pdf.

Anonymous:1875:CSS

[7456] Anonymous. Clouds as storm signals. Scientific American, 33 (10):145, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
145a.pdf.

Anonymous:1875:SPIu

[7457] Anonymous. Scientific and practical information. Scientific American, 33(10):145, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
145b.pdf.

Anonymous:1875:NG

[7458] Anonymous. A new gunboat. Scientific American, 33(10):146, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n10/pdf/scientificamerican09041875-146.pdf.

Anonymous:1875:DPTb

[7459] Anonymous. Driving portable tools. Scientific American, 33(10): 146–147, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
146a.pdf.

Anonymous:1875:NCC

[7460] Anonymous. A new circular compass. Scientific American, 33 (10):147, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
147.pdf.

Anonymous:1875:WSV

[7461] Anonymous. Whitton’s stop valve. Scientific American, 33(10): 147, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
147a.pdf.

Anonymous:1875:CRbk

[7462] Anonymous. The cataleptic rooster. Scientific American, 33(10): 147–148, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
147b.pdf.

Rose:1875:PMp

[7463] Joshua Rose. Practical mechanism. Scientific American, 33(10): 148–149, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
148a.pdf.

Anonymous:1875:CBT

[7464] Anonymous. Combined belt tightener and countershaft hanger. Scientific American, 33(10):150, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
150.pdf.

Anonymous:1875:NMB

[7465] Anonymous. New method of bending pipes. Scientific American, 33(10):150, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
150a.pdf.

Anonymous:1875:INE

[7466] Anonymous. Improved nail extractor. Scientific American, 33 (10):150, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
150b.pdf.

Anonymous:1875:BBC

[7467] Anonymous. Bohart’s beer cooler. Scientific American, 33(10): 150, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
150c.pdf.

Anonymous:1875:BH

[7468] Anonymous. El Baron De Hughes. Scientific American, 33(10): 151–152, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
151.pdf.

Anonymous:1875:HSBa

[7469] Anonymous. The heating surface of boilers. Scientific American, 33(10):152, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
152.pdf.

Anonymous:1875:AAAb

[7470] Anonymous. The American Association for the Advancement of Science. Scientific American, 33(10):152–153, September 4, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n10/pdf/scientificamerican09041875-152a.pdf.

Anonymous:1875:DCu

[7471] Anonymous. Decisions of the courts. Scientific American, 33 (10):153, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
153.pdf.

Anonymous:1875:RAFbj

[7472] Anonymous. Recent American and foreign patents. Scientific American, 33(10):154–155, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
154.pdf.

Anonymous:1875:CRbl

[7473] Anonymous. Communications received. Scientific American, 33 (10):156, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
156.pdf.

Anonymous:1875:CPbb

[7474] Anonymous. Canadian patents. Scientific American, 33(10): 156, September 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n10/pdf/scientificamerican09041875-
156b.pdf.

Anonymous:1875:Cbn

[7475] Anonymous. Correspondence. Scientific American, 33(11):164– 165, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
164.pdf.

Anonymous:1875:BPbk

[7476] Anonymous. Business and personal. Scientific American, 33 (11):171, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
171.pdf.

Anonymous:1875:NQbj

[7477] Anonymous. Notes and queries. Scientific American, 33(11): 171–172, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
171a.pdf.

Anonymous:1875:IIbk

[7478] Anonymous. Index of inventions. Scientific American, 33(11): 172, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
172a.pdf.

Anonymous:1875:CBo

[7479] Anonymous. The Centennial buildings. Scientific American, 33 (11):159, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
159.pdf.

Anonymous:1875:SCe

[7480] Anonymous. A solar chronometer. Scientific American, 33(11): 159–160, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
159a.pdf.

Anonymous:1875:ESc

[7481] Anonymous. The education of sight. Scientific American, 33 (11):160, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
160.pdf.

Anonymous:1875:SNV

[7482] Anonymous. Some new volcano revelations. Scientific American, 33(11):160, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
160a.pdf.

Anonymous:1875:PRC

[7483] Anonymous. Progress of rapid city transit. Scientific American, 33 (11):160–161, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
160b.pdf.

Anonymous:1875:SBc

[7484] Anonymous. Stealing brains. Scientific American, 33(11):161, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n11/pdf/scientificamerican09111875-161.pdf.

Anonymous:1875:TAC

[7485] Anonymous. Traube’s artificial cells. Scientific American, 33 (11):161, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
161a.pdf.

Anonymous:1875:SPIv

[7486] Anonymous. Scientific and practical information. Scientific American, 33(11):161, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
161b.pdf.

Anonymous:1875:FFE

[7487] Anonymous. Floating fire engines. Scientific American, 33(11): 162–163, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
162.pdf.

Anonymous:1875:BBb

[7488] Anonymous. Boring bits. Scientific American, 33(11):165, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n11/pdf/scientificamerican09111875-165.pdf.

Anonymous:1875:ICC

[7489] Anonymous. Improved cooking chamber. Scientific American, 33 (11):166, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
166.pdf.

Anonymous:1875:IOM

[7490] Anonymous. Improved ornamental molding machine. Scientific American, 33(11):166, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
166a.pdf.

Anonymous:1875:CJa

[7491] Anonymous. Cochliostema Jacobianum. Scientific American, 33 (11):167, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
167.pdf.

Anonymous:1875:Ma

[7492] Anonymous. Mesembryanthemums. Scientific American, 33(11): 167–168, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
167a.pdf.

Anonymous:1875:TEI

[7493] Anonymous. Theories of the Earth’s interior. Scientific American, 33 (11):168–169, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
168.pdf.

Anonymous:1875:DCv

[7494] Anonymous. Decisions of the courts. Scientific American, 33 (11):169, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
169.pdf.

Anonymous:1875:RAFbk

[7495] Anonymous. Recent American and foreign patents. Scientific American, 33(11):169–170, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
169a.pdf.

Anonymous:1875:CRbm

[7496] Anonymous. Communications received. Scientific American, 33 (11):172, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
172.pdf.

Anonymous:1875:CPbc

[7497] Anonymous. Canadian patents. Scientific American, 33(11):172– 173, September 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n11/pdf/scientificamerican09111875-
172b.pdf.

Anonymous:1875:Cbp

[7498] Anonymous. Correspondence. Scientific American, 33(12):180, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n12/pdf/scientificamerican09181875-180.pdf.

Anonymous:1875:BPbl

[7499] Anonymous. Business and personal. Scientific American, 33 (12):187, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
187.pdf.

Anonymous:1875:NQbk

[7500] Anonymous. Notes and queries. Scientific American, 33(12): 187–188, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
187a.pdf.

Anonymous:1875:IIbl

[7501] Anonymous. Index of inventions. Scientific American, 33(12): 188, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
188.pdf.

Anonymous:1875:NPO

[7502] Anonymous. The new post office, New York City. Scientific American, 33(12):175, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
175.pdf.

Anonymous:1875:WPb

[7503] Anonymous. The Woodbury Patent. Scientific American, 33 (12):176, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
176.pdf.

Anonymous:1875:SWE

[7504] Anonymous. Shall we eat the horse? Scientific American, 33 (12):176, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
176a.pdf.

Anonymous:1875:MDI

[7505] Anonymous. Me. Dawson’s idea of evolution. Scientific American, 33 (12):176–177, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
176b.pdf.

Anonymous:1875:NDA

[7506] Anonymous. The new department of aesthetics. Scientific American, 33(12):177, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
177.pdf.

Anonymous:1875:LRa

[7507] Anonymous. Lightning rods. Scientific American, 33(12):177, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n12/pdf/scientificamerican09181875-177a.pdf.

Anonymous:1875:AVM

[7508] Anonymous. Another view of Mr. Crookes’ last discovery. Scientific American, 33(12):177, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
177b.pdf.

Anonymous:1875:EGM

[7509] Anonymous. Explosions in gunpowder mills by electricity. Scientific American, 33(12):177–178, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
177c.pdf.

Anonymous:1875:IUG

[7510] Anonymous. Improved universal grinder. Scientific American, 33 (12):179, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
179.pdf.

Anonymous:1875:CSb

[7511] Anonymous. Compresses steel. Scientific American, 33(12):179– 180, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
179a.pdf.

Rose:1875:PMq

[7512] Joshua Rose. Practical mechanism. Scientific American, 33(12): 180–181, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
180a.pdf.

Anonymous:1875:NLF

[7513] Anonymous. New locomotive feed water heater. Scientific American, 33(12):182, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
182.pdf.

Anonymous:1875:FWP

[7514] Anonymous. Furnace working with petroleum. Scientific American, 33(12):182, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
182a.pdf.

Anonymous:1875:ESR

[7515] Anonymous. Electrical speed recorder. Scientific American, 33 (12):182, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
182b.pdf.

Anonymous:1875:Se

[7516] Anonymous. Swans. Scientific American, 33(12):183, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n12/pdf/scientificamerican09181875-183.pdf.

Anonymous:1875:Mc

[7517] Anonymous. Magnetism. Scientific American, 33(12):184–185, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n12/pdf/scientificamerican09181875-184.pdf.

Anonymous:1875:RAFbl

[7518] Anonymous. Recent American and foreign patents. Scientific American, 33(12):186, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
186.pdf.

Anonymous:1875:CPbd

[7519] Anonymous. Canadian patents. Scientific American, 33(12): 187, September 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n12/pdf/scientificamerican09181875-
187b.pdf.

Anonymous:1875:Cbq

[7520] Anonymous. Correspondence. Scientific American, 33(13):196– 197, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
196.pdf.

Anonymous:1875:BNBn

[7521] Anonymous. Books: New books and publications. Scientific American, 33(13):202, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
202.pdf.

Anonymous:1875:BPbm

[7522] Anonymous. Business and personal. Scientific American, 33 (13):203, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
203.pdf.

Anonymous:1875:NQbl

[7523] Anonymous. Notes and queries. Scientific American, 33(13): 203–204, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
203a.pdf.

Anonymous:1875:IIbm

[7524] Anonymous. Index of inventions. Scientific American, 33(13): 204, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
204a.pdf.

Anonymous:1875:IFWb

[7525] Anonymous. Improved floodway for warehouses. Scientific American, 33(13):191, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
191.pdf.

Anonymous:1875:IEP

[7526] Anonymous. Insect-eating plants. Scientific American, 33(13): 192, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
192.pdf.

Anonymous:1875:CHG

[7527] Anonymous. Completion of the Hell Gate excavations. Scientific American, 33(13):192, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
192a.pdf.

Anonymous:1875:DHb

[7528] Anonymous. Drugging horses. Scientific American, 33(13):192– 193, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
192b.pdf.

Anonymous:1875:Cbr

[7529] Anonymous. Celluloid. Scientific American, 33(13):193, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n13/pdf/scientificamerican09251875-193.pdf.

Anonymous:1875:SBP

[7530] Anonymous. Steam boiler phenomena. Scientific American, 33 (13):193, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
193a.pdf.

Anonymous:1875:SId

[7531] Anonymous. “stray ideas”. Scientific American, 33(13):193, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n13/pdf/scientificamerican09251875-193b.pdf.

Anonymous:1875:NCP

[7532] Anonymous. The new Commissioner of Patents. Scientific American, 33(13):193, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
193c.pdf.

Anonymous:1875:LSA

[7533] Anonymous. Life-saving apparatus. Scientific American, 33(13): 194, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
194.pdf.

Anonymous:1875:WPE

[7534] Anonymous. Water pressure engine. Scientific American, 33(13): 195–196, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
195.pdf.

Anonymous:1875:EA

[7535] Anonymous. An experimental Aurora. Scientific American, 33 (13):197, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
197.pdf.

Anonymous:1875:SPIw

[7536] Anonymous. Scientific and practical information. Scientific American, 33(13):197, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
197a.pdf.

Anonymous:1875:IVa

[7537] Anonymous. Improved vise. Scientific American, 33(13):198, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n13/pdf/scientificamerican09251875-198.pdf.

Anonymous:1875:FRC

[7538] Anonymous. Fox’s reversible cap. Scientific American, 33(13): 198, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
198a.pdf.

Anonymous:1875:MSV

[7539] Anonymous. A mercurial safety valve. Scientific American, 33 (13):198, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
198b.pdf.

Anonymous:1875:RCL

[7540] Anonymous. Roberts’ combined lamp and oil can. Scientific American, 33(13):198–199, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
198c.pdf.

Anonymous:1875:CBs

[7541] Anonymous. The Centennial buildings. Scientific American, 33 (13):199–201, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
199.pdf.

Anonymous:1875:DCw

[7542] Anonymous. Decisions of the courts. Scientific American, 33(13): 201–202, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
201.pdf.

Anonymous:1875:RAFbm

[7543] Anonymous. Recent American and foreign patents. Scientific American, 33(13):202, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
202a.pdf.

Anonymous:1875:CRbn

[7544] Anonymous. Communications received. Scientific American, 33 (13):204, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
204.pdf.

Anonymous:1875:CPbe

[7545] Anonymous. Canadian patents. Scientific American, 33(13): 204, September 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n13/pdf/scientificamerican09251875-
204b.pdf.

Anonymous:1875:BPbn

[7546] Anonymous. Business and personal. Scientific American, 33(14):219, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-219.pdf.

Anonymous:1875:NQbm

[7547] Anonymous. Notes and queries. Scientific American, 33(14):219–220, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-219a.pdf.

Anonymous:1875:IIbn

[7548] Anonymous. Index of inventions. Scientific American, 33(14):220–221, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-220a.pdf.

Anonymous:1875:ILM

[7549] Anonymous. Improved liquid mixer. Scientific American, 33 (14):207–208, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
207.pdf.

Anonymous:1875:NUC

[7550] Anonymous. A new use for criminals. Scientific American, 33(14):208, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-208.pdf.

Anonymous:1875:HFW

[7551] Anonymous. How far will bodies sink in the ocean? Scientific American, 33(14):208, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
208a.pdf.

Anonymous:1875:PRTc

[7552] Anonymous. The proposed railway tunnel under the English channel. Scientific American, 33(14):209, October 2, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n14/pdf/scientificamerican10021875-209.pdf.

Anonymous:1875:SPc

[7553] Anonymous. Science pugnacious. Scientific American, 33(14):209, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-209a.pdf.

Anonymous:1875:GNS

[7554] Anonymous. Gas from night soil and dead animals. Scientific American, 33(14):209, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
209b.pdf.

Anonymous:1875:LRb

[7555] Anonymous. Lightning rods. Scientific American, 33(14):209, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-209c.pdf.

Anonymous:1875:FAIa

[7556] Anonymous. The fair of the American institute. Scientific American, 33(14):209, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
209d.pdf.

Anonymous:1875:CMa

[7557] Anonymous. Carillon machines. Scientific American, 33(14):210, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-210.pdf.

Anonymous:1875:DSS

[7558] Anonymous. Deep sea sounding by pianoforte wire. Scientific American, 33(14):211, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
211.pdf.

Anonymous:1875:ENZ

[7559] Anonymous. Engineering in New Zealand. Scientific American, 33 (14):211–212, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
211a.pdf.

Rose:1875:PMr

[7560] Joshua Rose. Practical mechanism. Scientific American, 33(14): 212–213, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
212.pdf.

Anonymous:1875:IAP

[7561] Anonymous. Improved adding pencil. Scientific American, 33(14):214, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-214.pdf.

Anonymous:1875:CSWb

[7562] Anonymous. Changing salt water to fresh. Scientific American, 33(14):214, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
214a.pdf.

Anonymous:1875:WPS

[7563] Anonymous. Wood’s patent square-inch test valve. Scientific American, 33(14):214–215, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
214b.pdf.

Anonymous:1875:KIH

[7564] Anonymous. Knight’s improved health lift. Scientific American, 33(14):215, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
215.pdf.

Anonymous:1875:MV

[7565] Anonymous. A model villa. Scientific American, 33(14):215–216, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-215a.pdf.

Anonymous:1875:BAN

[7566] Anonymous. British Association notes. Scientific American, 33 (14):216–217, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
216.pdf.

Anonymous:1875:DCx

[7567] Anonymous. Decisions of the courts. Scientific American, 33(14):217, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-217.pdf.

Anonymous:1875:RAFbn

[7568] Anonymous. Recent American and foreign patents. Scientific American, 33(14):217–218, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
217a.pdf.

Anonymous:1875:CRbo

[7569] Anonymous. Communications received. Scientific American, 33 (14):220, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n14/pdf/scientificamerican10021875-
220.pdf.

Anonymous:1875:CPbf

[7570] Anonymous. Canadian patents. Scientific American, 33(14):221, October 2, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n14/pdf/scientificamerican10021875-221.pdf.

Anonymous:1875:Cbu

[7571] Anonymous. Correspondence. Scientific American, 33(15):228–229, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-228.pdf.

Anonymous:1875:BPbo

[7572] Anonymous. Business and personal. Scientific American, 33(15):235, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-235.pdf.

Anonymous:1875:NQbn

[7573] Anonymous. Notes and queries. Scientific American, 33(15):235–236, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-235a.pdf.

Anonymous:1875:IIbo

[7574] Anonymous. Index of inventions. Scientific American, 33(15):236–237, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-236a.pdf.

Anonymous:1875:IDE

[7575] Anonymous. Improved deodorizing excavating apparatus. Scientific American, 33(15):223–224, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
223.pdf.

Anonymous:1875:RNA

[7576] Anonymous. The relationships of native Americans. Scientific American, 33(15):224, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
224.pdf.

Anonymous:1875:AEa

[7577] Anonymous. Amateur engineering. Scientific American, 33(15):224, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-224a.pdf.

Anonymous:1875:T

[7578] Anonymous. Teaspoonfuls. Scientific American, 33(15):224, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-224b.pdf.

Anonymous:1875:PBa

[7579] Anonymous. Proportions of boilers. Scientific American, 33(15):225, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-225.pdf.

Anonymous:1875:SAPb

[7580] Anonymous. Suspended animation as a preserving agent. Scientific American, 33(15):225, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
225a.pdf.

Anonymous:1875:SPIx

[7581] Anonymous. Scientific and practical information. Scientific American, 33(15):225, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
225b.pdf.

Anonymous:1875:UFI

[7582] Anonymous. Universal field instrument. Scientific American, 33 (15):226, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
226.pdf.

Anonymous:1875:NTR

[7583] Anonymous. New tide and river gages. Scientific American, 33(15):227, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-227.pdf.

Anonymous:1875:IFMc

[7584] Anonymous. Improved friction meter. Scientific American, 33(15):227, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-227a.pdf.

Anonymous:1875:CWG

[7585] Anonymous. Captain Webb’s great swimming feat. Scientific American, 33(15):227–228, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
227b.pdf.

Anonymous:1875:IEb

[7586] Anonymous. Improved ellipsograph. Scientific American, 33(15):230, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-230.pdf.

Anonymous:1875:IRTb

[7587] Anonymous. Improved railway tie. Scientific American, 33(15):230, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-230a.pdf.

Anonymous:1875:FEW

[7588] Anonymous. Fleury’s Eureka weatherstrip. Scientific American, 33(15):230, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
230b.pdf.

Anonymous:1875:CBKa

[7589] Anonymous. The crested black kite. Scientific American, 33(15):231, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-231.pdf.

Anonymous:1875:Fd

[7590] Anonymous. The florakin. Scientific American, 33(15):231–232, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-231a.pdf.

Anonymous:1875:FAIb

[7591] Anonymous. The fair of the American institute. Scientific American, 33(15):233, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
233.pdf.

Anonymous:1875:DCy

[7592] Anonymous. Decisions of the courts. Scientific American, 33(15):234, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-234.pdf.

Anonymous:1875:RAFbo

[7593] Anonymous. Recent American and foreign patents. Scientific American, 33(15):234, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
234a.pdf.

Anonymous:1875:CRbp

[7594] Anonymous. Communications received. Scientific American, 33 (15):236, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n15/pdf/scientificamerican10091875-
236.pdf.

Anonymous:1875:CPbg

[7595] Anonymous. Canadian patents. Scientific American, 33(15):237, October 9, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n15/pdf/scientificamerican10091875-237.pdf.

Anonymous:1875:Cbv

[7596] Anonymous. Correspondence. Scientific American, 33(16):244–245, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-244.pdf.

Anonymous:1875:BNBo

[7597] Anonymous. Books: New books and publications. Scientific American, 33(16):249–250, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
249a.pdf.

Anonymous:1875:BPbp

[7598] Anonymous. Business and personal. Scientific American, 33 (16):251, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
251.pdf.

Anonymous:1875:NQbo

[7599] Anonymous. Notes and queries. Scientific American, 33(16): 251–252, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
251a.pdf.

Anonymous:1875:IIbp

[7600] Anonymous. Index of inventions. Scientific American, 33(16):252, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-252a.pdf.

Anonymous:1875:ETG

[7601] Anonymous. The eighty-ton guns. Scientific American, 33(16): 239–240, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
239.pdf.

Anonymous:1875:FRSb

[7602] Anonymous. The Fall River strike. Scientific American, 33(16):240, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-240.pdf.

Anonymous:1875:MFI

[7603] Anonymous. Mending a fifteen-inch shaft at sea. Scientific American, 33(16):240, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
240a.pdf.

Anonymous:1875:DCz

[7604] Anonymous. A disastrous Cyclone. Scientific American, 33(16):240, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-240b.pdf.

Anonymous:1875:PLAc

[7605] Anonymous. The postal laws — amendments needed. Scientific American, 33(16):240–241, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
240c.pdf.

Anonymous:1875:CVP

[7606] Anonymous. Conservatism versus progress. Scientific American, 33(16):241, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
241.pdf.

Anonymous:1875:FAIc

[7607] Anonymous. The fair of the American institute. Scientific American, 33(16):241, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
241a.pdf.

Anonymous:1875:SPIy

[7608] Anonymous. Scientific and practical information. Scientific American, 33(16):241, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
241b.pdf.

Anonymous:1875:IIMa

[7609] Anonymous. Improved ice-making machine. Scientific American, 33(16):242, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
242.pdf.

Anonymous:1875:TSR

[7610] Anonymous. Two-story railway cars. Scientific American, 33 (16):242, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
242a.pdf.

Anonymous:1875:LCA

[7611] Anonymous. Liquid carbonic acid as a motor. Scientific American, 33(16):243–244, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
243.pdf.

Rose:1875:PMs

[7612] Joshua Rose. Practical mechanism. Scientific American, 33(16):245, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-245.pdf.

Anonymous:1875:ITU

[7613] Anonymous. Improved tire-upsetting machine. Scientific American, 33(16):246, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
246.pdf.

Anonymous:1875:HSBb

[7614] Anonymous. Hennaman and Shaw’s boiling pot. Scientific American, 33(16):246, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
246a.pdf.

Anonymous:1875:IMC

[7615] Anonymous. Improved metallic column. Scientific American, 33 (16):246, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
246b.pdf.

Anonymous:1875:GFM

[7616] Anonymous. The grange farm mill. Scientific American, 33(16):246, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-246c.pdf.

Anonymous:1875:GSIb

[7617] Anonymous. The great spotted Iris. Scientific American, 33 (16):247, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
247.pdf.

Anonymous:1875:CNa

[7618] Anonymous. Cupressus Nutkaensis. Scientific American, 33(16): 247, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
247a.pdf.

Anonymous:1875:CCa

[7619] Anonymous. Caladium culture. Scientific American, 33(16):247, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-247b.pdf.

Anonymous:1875:FPa

[7620] Anonymous. The Flamingo plant. Scientific American, 33(16):247, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-247c.pdf.

Anonymous:1875:AM

[7621] Anonymous. Atmospheric machinery. Scientific American, 33(16): 248–249, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
248.pdf.

Anonymous:1875:DCba

[7622] Anonymous. Decisions of the courts. Scientific American, 33 (16):249, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
249.pdf.

Anonymous:1875:RAFbp

[7623] Anonymous. Recent American and foreign patents. Scientific American, 33(16):250, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
250.pdf.

Anonymous:1875:CRbq

[7624] Anonymous. Communications received. Scientific American, 33 (16):252, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n16/pdf/scientificamerican10161875-
252.pdf.

Anonymous:1875:CPbh

[7625] Anonymous. Canadian patents. Scientific American, 33(16):252, October 16, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n16/pdf/scientificamerican10161875-252b.pdf.

Anonymous:1875:Cbw

[7626] Anonymous. Correspondence. Scientific American, 33(17):260–261, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-260.pdf.

Anonymous:1875:BNBp

[7627] Anonymous. Books: New books and publications. Scientific American, 33(17):266, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
266.pdf.

Anonymous:1875:BPbq

[7628] Anonymous. Business and personal. Scientific American, 33 (17):267, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
267.pdf.

Anonymous:1875:NQbp

[7629] Anonymous. Notes and queries. Scientific American, 33(17): 267–268, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
267a.pdf.

Anonymous:1875:IIbq

[7630] Anonymous. Index of inventions. Scientific American, 33(17):268, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-268a.pdf.

Anonymous:1875:MSB

[7631] Anonymous. Modern suspension bridges. Scientific American, 33 (17):255, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
255.pdf.

Anonymous:1875:JJC

[7632] Anonymous. J. and J. A. Crook’s saw sharpener. Scientific American, 33(17):255, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
255a.pdf.

Anonymous:1875:GIO

[7633] Anonymous. The geological importance of our Western explorations. Scientific American, 33(17):256, October 23, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n17/pdf/scientificamerican10231875-256.pdf.

Anonymous:1875:CPbi

[7634] Anonymous. Carnivorous plants. Scientific American, 33(17):256, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-256a.pdf.

Anonymous:1875:FAId

[7635] Anonymous. The fair of the American institute. Scientific American, 33(17):256–257, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
256b.pdf.

Anonymous:1875:SPIz

[7636] Anonymous. Scientific and practical information. Scientific American, 33(17):257, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
257.pdf.

Anonymous:1875:FS

[7637] Anonymous. Flat surfaces. Scientific American, 33(17):257, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-257a.pdf.

Anonymous:1875:MGP

[7638] Anonymous. Making gas from petroleum. Scientific American, 33(17):258, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
258.pdf.

Anonymous:1875:NEC

[7639] Anonymous. Novel engine counter and tell-tale. Scientific American, 33(17):258, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
258a.pdf.

Anonymous:1875:SFB

[7640] Anonymous. A solid flame Bunsen burner. Scientific American, 33(17):258, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
258b.pdf.

Anonymous:1875:LFA

[7641] Anonymous. Lloyd’s floating apparatus. Scientific American, 33 (17):258, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
258c.pdf.

Anonymous:1875:NGRa

[7642] Anonymous. New gas regulator. Scientific American, 33(17):259, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-259.pdf.

Anonymous:1875:CBPa

[7643] Anonymous. Cottrell and Babcock’s perfecting press for woodcut printing-firm’s rotary attachment. Scientific American, 33(17): 259–260, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
259a.pdf.

Anonymous:1875:IVH

[7644] Anonymous. Improved ventilating heater. Scientific American, 33(17):262, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
262.pdf.

Anonymous:1875:PSEb

[7645] Anonymous. Portable steam engines. Scientific American, 33 (17):262, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
262a.pdf.

Anonymous:1875:MCF

[7646] Anonymous. The manatee or cow fish. Scientific American, 33 (17):263, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
263.pdf.

Anonymous:1875:LT

[7647] Anonymous. Land tortoises. Scientific American, 33(17):263, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n17/pdf/scientificamerican10231875-263a.pdf.

Anonymous:1875:SOb

[7648] Anonymous. The Shoemaker’s opportunity. Scientific American, 33 (17):264–265, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
264.pdf.

Anonymous:1875:RAFbq

[7649] Anonymous. Recent American and foreign patents. Scientific American, 33(17):266, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
266a.pdf.

Anonymous:1875:CRbr

[7650] Anonymous. Communications received. Scientific American, 33 (17):268, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
268.pdf.

Anonymous:1875:CPbj

[7651] Anonymous. Canadian patents. Scientific American, 33(17):268– 269, October 23, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n17/pdf/scientificamerican10231875-
268b.pdf.

Anonymous:1875:BPbr

[7652] Anonymous. Business and personal. Scientific American, 33 (18):283, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
283.pdf.

Anonymous:1875:NQbq

[7653] Anonymous. Notes and queries. Scientific American, 33(18): 283–284, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
283a.pdf.

Anonymous:1875:IIbr

[7654] Anonymous. Index of inventions. Scientific American, 33(18): 284–285, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
284a.pdf.

Anonymous:1875:MBL

[7655] Anonymous. Modern British locomotives. Scientific American, 33(18):271, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
271.pdf.

Anonymous:1875:OTS

[7656] Anonymous. Is our torpedo system useless? Scientific American, 33(18):272, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
272.pdf.

Anonymous:1875:CHT

[7657] Anonymous. The colleges and the hard times. Scientific American, 33(18):272, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
272a.pdf.

Anonymous:1875:LSC

[7658] Anonymous. The locomotive semi-centennial. Scientific American, 33(18):272, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
272b.pdf.

Anonymous:1875:PAT

[7659] Anonymous. Progress of American telegraphy. Scientific American, 33(18):273, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
273.pdf.

Anonymous:1875:FPR

[7660] Anonymous. Foreign patents-reduction of costs. Scientific American, 33(18):273, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
273a.pdf.

Anonymous:1875:KMDh

[7661] Anonymous. The Keely motor deception. Scientific American, 33(18):273, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
273b.pdf.

Anonymous:1875:FAIe

[7662] Anonymous. The fair of the American institute. Scientific American, 33(18):273, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
273c.pdf.

Anonymous:1875:ITK

[7663] Anonymous. Improved tea kettle. Scientific American, 33(18):274, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n18/pdf/scientificamerican10301875-274.pdf.

Anonymous:1875:CSR

[7664] Anonymous. Combined scissors and ripping knife. Scientific American, 33(18):274, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
274a.pdf.

Anonymous:1875:RHC

[7665] Anonymous. Recent hospital construction. Scientific American, 33(18):274, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
274b.pdf.

Anonymous:1875:IFCc

[7666] Anonymous. Improved friction clutch pulley. Scientific American, 33(18):274, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
274c.pdf.

Anonymous:1875:NGRb

[7667] Anonymous. The new Giffard railway carriages. Scientific American, 33(18):275, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
275.pdf.

Anonymous:1875:NMC

[7668] Anonymous. New method of casting steel ingots. Scientific American, 33(18):275, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
275a.pdf.

Anonymous:1875:HMCb

[7669] Anonymous. A home-made compound microscope. Scientific American, 33(18):275–276, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
275b.pdf.

Rose:1875:PMt

[7670] Joshua Rose. Practical mechanism. Scientific American, 33(18):277, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n18/pdf/scientificamerican10301875-277.pdf.

Anonymous:1875:IGMb

[7671] Anonymous. Improved grinding mills. Scientific American, 33 (18):278, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
278.pdf.

Anonymous:1875:IRCb

[7672] Anonymous. Improved refrigerator and counter. Scientific American, 33(18):278, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
278a.pdf.

Anonymous:1875:OH

[7673] Anonymous. Oriental horticulture. Scientific American, 33(18):279, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n18/pdf/scientificamerican10301875-279.pdf.

Anonymous:1875:UCI

[7674] Anonymous. Useful and curious inventions. Scientific American, 33 (18):280–281, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
280.pdf.

Anonymous:1875:SPIba

[7675] Anonymous. Scientific and practical information. Scientific American, 33(18):281, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
281.pdf.

Anonymous:1875:RAFbr

[7676] Anonymous. Recent American and foreign patents. Scientific American, 33(18):282, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
282.pdf.

Anonymous:1875:RAFbs

[7677] Anonymous. Recent American and foreign patents. Scientific American, 33(18):282, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
282a.pdf.

Anonymous:1875:CRbs

[7678] Anonymous. Communications received. Scientific American, 33 (18):284, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n18/pdf/scientificamerican10301875-
284.pdf.

Anonymous:1875:CPbk

[7679] Anonymous. Canadian patents. Scientific American, 33(18):285, October 30, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n18/pdf/scientificamerican10301875-285.pdf.

Anonymous:1875:Cbx

[7680] Anonymous. Correspondence. Scientific American, 33(19):292– 293, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
292.pdf.

Anonymous:1875:BPbs

[7681] Anonymous. Business and personal. Scientific American, 33 (19):298, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
298.pdf.

Anonymous:1875:NQbr

[7682] Anonymous. Notes and queries. Scientific American, 33(19): 298–300, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
298a.pdf.

Anonymous:1875:IIbs

[7683] Anonymous. Index of inventions. Scientific American, 33(19): 300, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
300a.pdf.

Anonymous:1875:PCb

[7684] Anonymous. A powerful crane. Scientific American, 33(19):287, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n19/pdf/scientificamerican11061875-287.pdf.

Anonymous:1875:GE

[7685] Anonymous. The goal of evolution. Scientific American, 33 (19):288, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
288.pdf.

Anonymous:1875:ICb

[7686] Anonymous. Instinctive calculation. Scientific American, 33(19): 288, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
288a.pdf.

Anonymous:1875:PBF

[7687] Anonymous. Pineapple and banana fibers. Scientific American, 33 (19):288–289, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
288b.pdf.

Anonymous:1875:OM

[7688] Anonymous. The origin of mountains. Scientific American, 33 (19):289, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
289.pdf.

Anonymous:1875:SPIbb

[7689] Anonymous. Scientific and practical information. Scientific American, 33(19):289, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
289a.pdf.

Anonymous:1875:SSM

[7690] Anonymous. The Sheffield steel manufactures. Scientific American, 33(19):291, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
291.pdf.

Anonymous:1875:DBG

[7691] Anonymous. Diving Bell and grapnel. Scientific American, 33 (19):291, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
291a.pdf.

Anonymous:1875:JFC

[7692] Anonymous. J. A. Fay & Co.’s mortising and boring machine. Scientific American, 33(19):291, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
291b.pdf.

Anonymous:1875:PM

[7693] Anonymous. The psycho mystery. Scientific American, 33(19): 291–292, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
291c.pdf.

Anonymous:1875:ISHc

[7694] Anonymous. Improved steam heater for dwellings. Scientific American, 33(19):294, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
294.pdf.

Anonymous:1875:MSF

[7695] Anonymous. Method of securing fish plates to railway joints. Scientific American, 33(19):294, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
294a.pdf.

Anonymous:1875:BCb

[7696] Anonymous. Black cockatoos. Scientific American, 33(19):295, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n19/pdf/scientificamerican11061875-295.pdf.

Anonymous:1875:IF

[7697] Anonymous. The iguana family. Scientific American, 33(19): 295, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
295a.pdf.

Anonymous:1875:ANi

[7698] Anonymous. Astronomical notes. Scientific American, 33(19): 296, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
296.pdf.

Anonymous:1875:EPd

[7699] Anonymous. Earth pits. Scientific American, 33(19):296, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n19/pdf/scientificamerican11061875-296a.pdf.

Anonymous:1875:FPb

[7700] Anonymous. Flower pots. Scientific American, 33(19):296, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n19/pdf/scientificamerican11061875-296b.pdf.

Anonymous:1875:DCbb

[7701] Anonymous. Decisions of the courts. Scientific American, 33 (19):296, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
296c.pdf.

Anonymous:1875:RAFbt

[7702] Anonymous. Recent American and foreign patents. Scientific American, 33(19):297, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
297.pdf.

Anonymous:1875:NAIa

[7703] Anonymous. New agricultural inventions. Scientific American, 33 (19):297, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
297a.pdf.

Anonymous:1875:NMEb

[7704] Anonymous. New mechanical and engineering inventions. Scientific American, 33(19):297, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
297b.pdf.

Anonymous:1875:NHAa

[7705] Anonymous. New household articles. Scientific American, 33 (19):297, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
297c.pdf.

Anonymous:1875:NWHa

[7706] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 33(19):297, November 6, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n19/pdf/scientificamerican11061875-297d.pdf.

Anonymous:1875:NCMa

[7707] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(19):297, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
297e.pdf.

Anonymous:1875:CRbt

[7708] Anonymous. Communications received. Scientific American, 33 (19):300, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
300.pdf.

Anonymous:1875:CPbl

[7709] Anonymous. Canadian patents. Scientific American, 33(19): 301, November 6, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n19/pdf/scientificamerican11061875-
301.pdf.

Anonymous:1875:BNBq

[7710] Anonymous. Books: New books and publications. Scientific American, 33(20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314.pdf.

Anonymous:1875:BPbt

[7711] Anonymous. Business and personal. Scientific American, 33 (20):315, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
315.pdf.

Anonymous:1875:NQbs

[7712] Anonymous. Notes and queries. Scientific American, 33(20): 315–316, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
315a.pdf.

Anonymous:1875:IIbt (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
316a.pdf.

Anonymous:1875:ISGd

[7714] Anonymous. Improved sawing and grooving machine. Scientific American, 33(20):303, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
303.pdf.

Anonymous:1875:FSP

[7715] Anonymous. Filth as the source of preventable disease. Scientific American, 33(20):304, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
304.pdf.

Anonymous:1875:WFI

[7716] Anonymous. What fills the interstellar space. Scientific American, 33 (20):304–305, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
304a.pdf.

Anonymous:1875:FPS

[7717] Anonymous. Foreign patents-splendid opportunities for Americans. Scientific American, 33(20):305, November 13, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n20/pdf/scientificamerican11131875-305.pdf.

Anonymous:1875:WR

[7718] Anonymous. “wrinkles and recipes”. Scientific American, 33 (20):305, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
305a.pdf.

Anonymous:1875:VDA

[7719] Anonymous. Veneered diamonds again. Scientific American, 33 (20):305, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
305b.pdf.

Anonymous:1875:SPIbc

[7720] Anonymous. Scientific and practical information. Scientific American, 33(20):305, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
305c.pdf.

Anonymous:1875:PCF

[7721] Anonymous. Pogpisil’s curtain fixture. Scientific American, 33 (20):306, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
306.pdf.

Anonymous:1875:JI

[7722] Anonymous. A juvenile inventor. Scientific American, 33(20): 306, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
306a.pdf.

Anonymous:1875:PG

[7723] Anonymous. Preserving grapes. Scientific American, 33(20): 306, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
306b.pdf.

Anonymous:1875:IWT

[7724] Anonymous. Improved Watchman’s time check. Scientific American, 33(20):306, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
306c.pdf.

Anonymous:1875:L

[7725] Anonymous. The locomotives of 1825 and of 1875. Scientific American, 33(20):307–308, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
307.pdf.

Rose:1875:PMu

[7726] Joshua Rose. Practical mechanism. Scientific American, 33(20): 309, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
309.pdf.

Anonymous:1875:IWB

[7727] Anonymous. Improved wagon brake. Scientific American, 33 (20):310, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
310.pdf.

Anonymous:1875:IDPb

[7728] Anonymous. Improved drawing press. Scientific American, 33 (20):310, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
310a.pdf.

Anonymous:1875:ACP

[7729] Anonymous. The artificial coloring of plants. Scientific American, 33(20):311, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
311.pdf.

Anonymous:1875:MCa

[7730] Anonymous. Melon cucumbers. Scientific American, 33(20): 311, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
311a.pdf.

Anonymous:1875:Sg

[7731] Anonymous. The skimmias. Scientific American, 33(20):311– 312, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
311b.pdf.

Anonymous:1875:ESd

[7732] Anonymous. Engineering structures. Scientific American, 33(20): 312–313, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
312.pdf.

Anonymous:1875:RCF

[7733] Anonymous. Reduction in the cost of foreign patents. Scientific American, 33(20):313, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
313.pdf.

Anonymous:1875:RAFbu

[7734] Anonymous. Recent American and foreign patents. Scientific American, 33(20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314a.pdf.

Anonymous:1875:NAIb

[7735] Anonymous. New agricultural inventions. Scientific American, 33 (20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314b.pdf.

Anonymous:1875:NCMb

[7736] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314c.pdf.

Anonymous:1875:NHAb

[7737] Anonymous. New household articles. Scientific American, 33 (20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314d.pdf.

Anonymous:1875:NMEc

[7738] Anonymous. New mechanical and engineering inventions. Scientific American, 33(20):314, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
314e.pdf.

Anonymous:1875:CRbu

[7739] Anonymous. Communications received. Scientific American, 33 (20):316, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
316.pdf.

Anonymous:1875:CPbm

[7740] Anonymous. Canadian patents. Scientific American, 33(20): 316, November 13, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n20/pdf/scientificamerican11131875-
316b.pdf.

Anonymous:1875:BNBr

[7741] Anonymous. Books: New books and publications. Scientific American, 33(21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330e.pdf.

Anonymous:1875:BPbu

[7742] Anonymous. Business and personal. Scientific American, 33 (21):331, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
331.pdf.

Anonymous:1875:NQbt

[7743] Anonymous. Notes and queries. Scientific American, 33(21): 331, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
331a.pdf.

Anonymous:1875:IIbu

[7744] Anonymous. Index of inventions. Scientific American, 33(21): 332, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
332a.pdf.

Anonymous:1875:SRV

[7745] Anonymous. A strong room for valuables. Scientific American, 33(21):319, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
319.pdf.

Anonymous:1875:SAc

[7746] Anonymous. Surface adhesion. Scientific American, 33(21):320, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n21/pdf/scientificamerican11201875-320.pdf.

Anonymous:1875:HSM

[7747] Anonymous. How some mountain gaps have been formed. Scientific American, 33(21):320, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
320a.pdf.

Anonymous:1875:WPW

[7748] Anonymous. The Woodbury planer war. Scientific American, 33 (21):320–321, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
320b.pdf.

Anonymous:1875:ECMb

[7749] Anonymous. Expansion and contraction by moisture. Scientific American, 33(21):321, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
321.pdf.

Anonymous:1875:MAa

[7750] Anonymous. Man as an automaton. Scientific American, 33 (21):321, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
321a.pdf.

Anonymous:1875:AIR

[7751] Anonymous. American inventions re-discovered in Europe. Scientific American, 33(21):321, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
321b.pdf.

Anonymous:1875:SPIbd

[7752] Anonymous. Scientific and practical information. Scientific American, 33(21):321, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
321c.pdf.

Anonymous:1875:NPb

[7753] Anonymous. A new propeller. Scientific American, 33(21):322, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n21/pdf/scientificamerican11201875-322.pdf.

Anonymous:1875:LDS

[7754] Anonymous. Lubricating device for sewing and other light machines. Scientific American, 33(21):322, November 20, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n21/pdf/scientificamerican11201875-322a.pdf.

Anonymous:1875:NPF

[7755] Anonymous. A new patent file. Scientific American, 33(21):322, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n21/pdf/scientificamerican11201875-322b.pdf.

Anonymous:1875:PRSb

[7756] Anonymous. The proposed refrigerator steamer. Scientific American, 33(21):322–323, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
322c.pdf.

Anonymous:1875:FLC

[7757] Anonymous. Flasks for liquid carbonic acid. Scientific American, 33(21):323, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
323.pdf.

Anonymous:1875:AWE

[7758] Anonymous. Ancient war engines. Scientific American, 33(21): 323–325, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
323a.pdf.

Anonymous:1875:LSd

[7759] Anonymous. Laying off a square. Scientific American, 33(21): 325, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
325.pdf.

Anonymous:1875:IOW

[7760] Anonymous. Improved odorless water closet. Scientific American, 33(21):326, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
326.pdf.

Anonymous:1875:HDW

[7761] Anonymous. Hale’s duplex water elevator. Scientific American, 33(21):326, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
326a.pdf.

Anonymous:1875:PHL

[7762] Anonymous. Production of heat and light. Scientific American, 33(21):327, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
327.pdf.

Anonymous:1875:YS

[7763] Anonymous. The Yucca Stricta. Scientific American, 33(21): 327, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
327a.pdf.

Anonymous:1875:MUIa

[7764] Anonymous. Miscellaneous useful inventions. Scientific American, 33(21):328, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
328.pdf.

Anonymous:1875:PBN

[7765] Anonymous. Professor Bunsen’s new apparatus and battery for spark spectra. Scientific American, 33(21):329–330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n21/pdf/scientificamerican11201875-329.pdf.

Anonymous:1875:RAFbv

[7766] Anonymous. Recent American and foreign patents. Scientific American, 33(21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330.pdf.

Anonymous:1875:NHAc

[7767] Anonymous. New household articles. Scientific American, 33 (21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330a.pdf.

Anonymous:1875:NAIc

[7768] Anonymous. New agricultural inventions. Scientific American, 33 (21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330b.pdf.

Anonymous:1875:NCMc

[7769] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330c.pdf.

Anonymous:1875:NMEd

[7770] Anonymous. New mechanical and engineering inventions. Scientific American, 33(21):330, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
330d.pdf.

Anonymous:1875:CRbv

[7771] Anonymous. Communications received. Scientific American, 33 (21):332, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
332.pdf.

Anonymous:1875:CPbn

[7772] Anonymous. Canadian patents. Scientific American, 33(21): 332, November 20, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n21/pdf/scientificamerican11201875-
332b.pdf.

Anonymous:1875:Cby

[7773] Anonymous. Correspondence. Scientific American, 33(22):340, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n22/pdf/scientificamerican11271875-340.pdf.

Anonymous:1875:BNBs

[7774] Anonymous. Books: New books and publications. Scientific American, 33(22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345f.pdf.

Anonymous:1875:BPbv

[7775] Anonymous. Business and personal. Scientific American, 33 (22):346, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
346.pdf.

Anonymous:1875:NQbu

[7776] Anonymous. Notes and queries. Scientific American, 33(22): 346–347, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
346a.pdf.

Anonymous:1875:IIbv

[7777] Anonymous. Index of inventions. Scientific American, 33(22): 348, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
348.pdf.

Anonymous:1875:NSV

[7778] Anonymous. New system for ventilating and cleansing drain pipes in buildings. Scientific American, 33(22):335, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n22/pdf/scientificamerican11271875-335.pdf.

Anonymous:1875:GSH

[7779] Anonymous. The Gulf Stream as a heat carrier. Scientific American, 33(22):336, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
336.pdf.

Anonymous:1875:PDW

[7780] Anonymous. The patent drive well. Scientific American, 33 (22):336, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
336a.pdf.

Anonymous:1875:CTC

[7781] Anonymous. Cereals and their chemical valve. Scientific American, 33(22):336–337, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
336b.pdf.

Leeds:1875:WWU

[7782] Albert R. Leeds. When is water unfit to drink? Scientific American, 33(22):337, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
337.pdf.

Anonymous:1875:FAIf

[7783] Anonymous. The fair of the American institute. Scientific American, 33(22):337, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
337a.pdf.

Anonymous:1875:SPIbe

[7784] Anonymous. Scientific and practical information. Scientific American, 33(22):337, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
337b.pdf.

Anonymous:1875:HIM

[7785] Anonymous. The hematite iron mines of England. Scientific American, 33(22):338, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
338.pdf.

Anonymous:1875:SACa

[7786] Anonymous. Sheldon’s auxiliary car seat. Scientific American, 33 (22):338, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
338a.pdf.

Anonymous:1875:SBd

[7787] Anonymous. Setting boilers. Scientific American, 33(22):339– 340, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
339.pdf.

Rose:1875:PMv

[7788] Joshua Rose. Practical mechanism. Scientific American, 33(22): 341, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
341.pdf.

Anonymous:1875:CHS

[7789] Anonymous. Combined hand and slide lathe. Scientific American, 33(22):342, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
342.pdf.

Anonymous:1875:Ah

[7790] Anonymous. The antelopes. Scientific American, 33(22):343, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n22/pdf/scientificamerican11271875-343.pdf.

Anonymous:1875:Sj

[7791] Anonymous. The sambur. Scientific American, 33(22):343, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n22/pdf/scientificamerican11271875-343a.pdf.

Morton:1875:NC

[7792] Henry Morton. A new chromatrope. Scientific American, 33 (22):344, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
344.pdf.

Anonymous:1875:NIC

[7793] Anonymous. A new induction coil. Scientific American, 33(22): 344–345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
344a.pdf.

Anonymous:1875:DCbc

[7794] Anonymous. Decisions of the courts. Scientific American, 33 (22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345.pdf.

Anonymous:1875:RAFbw

[7795] Anonymous. Recent American and foreign patents. Scientific American, 33(22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345a.pdf.

Anonymous:1875:NCMd

[7796] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345b.pdf.

Anonymous:1875:NWHb

[7797] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 33(22):345, November 27, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n22/pdf/scientificamerican11271875-345c.pdf.

Anonymous:1875:NMEe

[7798] Anonymous. New mechanical and engineering inventions. Scientific American, 33(22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345d.pdf.

Anonymous:1875:NHAd

[7799] Anonymous. New household articles. Scientific American, 33 (22):345, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
345e.pdf.

Anonymous:1875:CRbw

[7800] Anonymous. Communications received. Scientific American, 33 (22):347–348, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
347.pdf.

Anonymous:1875:CPbo

[7801] Anonymous. Canadian patents. Scientific American, 33(22): 348, November 27, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n22/pdf/scientificamerican11271875-
348a.pdf.

Anonymous:1875:Cbz

[7802] Anonymous. Correspondence. Scientific American, 33(23):356–357, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-356.pdf.

Anonymous:1875:BNBt

[7803] Anonymous. Books: New books and publications. Scientific American, 33(23):362, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
362.pdf.

Anonymous:1875:BPbw

[7804] Anonymous. Business and personal. Scientific American, 33 (23):363, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
363.pdf.

Anonymous:1875:NQbv

[7805] Anonymous. Notes and queries. Scientific American, 33(23):363, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-363a.pdf.

Anonymous:1875:IIbw

[7806] Anonymous. Index of inventions. Scientific American, 33(23): 363–364, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
363c.pdf.

Anonymous:1875:HRB

[7807] Anonymous. The Hudson River Bridge at Troy, N.Y. Scientific American, 33(23):351, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
351.pdf.

Anonymous:1875:PCc

[7808] Anonymous. Publishers’ card. Scientific American, 33(23):352, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-352.pdf.

Anonymous:1875:GIE

[7809] Anonymous. The Great International Exposition. Scientific American, 33(23):352, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
352a.pdf.

Anonymous:1875:WGD

[7810] Anonymous. The water gas delusion again. Scientific American, 33(23):352, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
352b.pdf.

Anonymous:1875:CPSa

[7811] Anonymous. Curious poison stories. Scientific American, 33 (23):353, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
353.pdf.

Anonymous:1875:SVA

[7812] Anonymous. The source of volcanic action. Scientific American, 33(23):353, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
353a.pdf.

Anonymous:1875:MPH

[7813] Anonymous. The meeting of the public health association. Scientific American, 33(23):353, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
353b.pdf.

Anonymous:1875:WSG

[7814] Anonymous. Working steep gradients by locomotives. Scientific American, 33(23):354, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
354.pdf.

Anonymous:1875:MPG

[7815] Anonymous. Mein’s patent governor. Scientific American, 33 (23):354, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
354a.pdf.

Anonymous:1875:MPB

[7816] Anonymous. Moore’s pulley block. Scientific American, 33(23):354, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-354b.pdf.

Anonymous:1875:PIB

[7817] Anonymous. Powell’s improved burglar alarm. Scientific American, 33(23):355, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
355.pdf.

Anonymous:1875:ECA

[7818] Anonymous. The emission of carbonic acid from roots. Scientific American, 33(23):355, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
355a.pdf.

Anonymous:1875:HPB

[7819] Anonymous. The HolbrooK patent blind hinge. Scientific American, 33(23):355, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
355b.pdf.

Anonymous:1875:MID

[7820] Anonymous. Meigs’ improved dioptric light. Scientific American, 33(23):355, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
355c.pdf.

Anonymous:1875:PIC

[7821] Anonymous. Pedden’s improved cow tail holder. Scientific American, 33(23):355, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
355d.pdf.

Anonymous:1875:CSAa

[7822] Anonymous. Le Coentre’s sounding apparatus. Scientific American, 33(23):357, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
357.pdf.

Anonymous:1875:IS

[7823] Anonymous. Improved sadiron. Scientific American, 33(23):358, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-358.pdf.

Anonymous:1875:CCC

[7824] Anonymous. The Centennial combination culinary apparatus. Scientific American, 33(23):358, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
358a.pdf.

Anonymous:1875:IVb

[7825] Anonymous. Ironclad vessels. Scientific American, 33(23):359–360, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-359.pdf.

Anonymous:1875:ANj

[7826] Anonymous. Astronomical notes. Scientific American, 33(23):360, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n23/pdf/scientificamerican12041875-360.pdf.

Anonymous:1875:MUIb

[7827] Anonymous. Miscellaneous useful inventions. Scientific American, 33(23):361, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
361.pdf.

Anonymous:1875:RAFbx

[7828] Anonymous. Recent American and foreign inventions. Scientific American, 33(23):362, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
362a.pdf.

Anonymous:1875:NMEf

[7829] Anonymous. New mechanical and engineering inventions. Scientific American, 33(23):362, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
362b.pdf.

Anonymous:1875:NWHc

[7830] Anonymous. New woodwork and house and carriage building inventions. Scientific American, 33(23):362, December 4, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n23/pdf/scientificamerican12041875-362c.pdf.

Anonymous:1875:NHAe

[7831] Anonymous. New household articles. Scientific American, 33 (23):362, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
362d.pdf.

Anonymous:1875:NCMe

[7832] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(23):362, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
362e.pdf.

Anonymous:1875:CRbx

[7833] Anonymous. Communications received. Scientific American, 33 (23):363, December 4, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n23/pdf/scientificamerican12041875-
363b.pdf.

Anonymous:1875:Ccb

[7834] Anonymous. Correspondence. Scientific American, 33(24):372, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n24/pdf/scientificamerican12111875-372.pdf.

Anonymous:1875:BPbx

[7835] Anonymous. Business and personal. Scientific American, 33 (24):378, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
378.pdf.

Anonymous:1875:NQbw

[7836] Anonymous. Notes and queries. Scientific American, 33(24): 378–379, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
378a.pdf.

Anonymous:1875:IIbx

[7837] Anonymous. Index of inventions. Scientific American, 33(24): 379–380, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
379a.pdf.

Anonymous:1875:BM

[7838] Anonymous. Ballasting machinery. Scientific American, 33(24): 367, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
367.pdf.

Anonymous:1875:WBM

[7839] Anonymous. Wood-boring machinery. Scientific American, 33 (24):367, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
367a.pdf.

Anonymous:1875:PCd

[7840] Anonymous. Publishers’ card. Scientific American, 33(24):368, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n24/pdf/scientificamerican12111875-368.pdf.

Anonymous:1875:WHM

[7841] Anonymous. What is hard money? Scientific American, 33 (24):368, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
368a.pdf.

Anonymous:1875:MD

[7842] Anonymous. Mechanical drawing. Scientific American, 33(24): 368, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
368b.pdf.

Anonymous:1875:EPCb

[7843] Anonymous. Explosion at the Pullman Car Works. Scientific American, 33(24):368–369, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
368c.pdf.

Anonymous:1875:PH

[7844] Anonymous. The preservation of hops. Scientific American, 33 (24):369, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
369.pdf.

Anonymous:1875:SAS

[7845] Anonymous. The Scientific American supplement. Scientific American, 33(24):369, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
369a.pdf.

Anonymous:1875:HHb

[7846] Anonymous. A hint for the holidays. Scientific American, 33 (24):369, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
369b.pdf.

Anonymous:1875:SPIbf

[7847] Anonymous. Scientific and practical information. Scientific American, 33(24):369, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
369c.pdf.

Anonymous:1875:IFWc

[7848] Anonymous. Improved floodway for warehouses. Scientific American, 33(24):370, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
370.pdf.

Anonymous:1875:LGS

[7849] Anonymous. The little giant steam engine. Scientific American, 33(24):371, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
371.pdf.

Anonymous:1875:MPc

[7850] Anonymous. A mechanical Phoenix. Scientific American, 33(24): 371–372, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
371a.pdf.

Rose:1875:PMw

[7851] Joshua Rose. Practical mechanism. Scientific American, 33(24): 373, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
373.pdf.

Anonymous:1875:ILP

[7852] Anonymous. Improved lever power. Scientific American, 33(24): 374, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
374.pdf.

Anonymous:1875:ITB

[7853] Anonymous. Improved tubular bow socket. Scientific American, 33(24):374, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
374a.pdf.

Anonymous:1875:HCP

[7854] Anonymous. Hampton Court Palace and Gardens. Scientific American, 33(24):375, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
375.pdf.

Anonymous:1875:OMW

[7855] Anonymous. The oldest medical work in the world. Scientific American, 33(24):376, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
376.pdf.

Anonymous:1875:RAFby

[7856] Anonymous. Recent American and foreign patents. Scientific American, 33(24):377, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
377.pdf.

Anonymous:1875:NHAf

[7857] Anonymous. New household articles. Scientific American, 33 (24):377, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
377a.pdf.

Anonymous:1875:NAId

[7858] Anonymous. New agricultural inventions. Scientific American, 33 (24):377, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
377b.pdf.

Anonymous:1875:NMEg

[7859] Anonymous. New mechanical and engineering inventions. Scientific American, 33(24):377, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
377c.pdf.

Anonymous:1875:NCMf

[7860] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(24):377, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
377d.pdf.

Anonymous:1875:CRby

[7861] Anonymous. Communications received. Scientific American, 33 (24):379, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
379.pdf.

Anonymous:1875:CPbp

[7862] Anonymous. Canadian patents. Scientific American, 33(24): 380, December 11, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n24/pdf/scientificamerican12111875-
380.pdf.

Anonymous:1875:Ccd

[7863] Anonymous. Correspondence. Scientific American, 33(25):387– 389, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
387b.pdf.

Anonymous:1875:BNBu

[7864] Anonymous. Books: New books and publications. Scientific American, 33(25):393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
393.pdf.

Anonymous:1875:BPby

[7865] Anonymous. Business and personal. Scientific American, 33 (25):394, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
394.pdf.

Anonymous:1875:NQbx

[7866] Anonymous. Notes and queries. Scientific American, 33(25): 394–395, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
394a.pdf.

Anonymous:1875:IIby

[7867] Anonymous. Index of inventions. Scientific American, 33(25): 395–396, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
395a.pdf.

Anonymous:1875:NGE

[7868] Anonymous. A new grain elevator. Scientific American, 33(25): 383–384, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
383.pdf.

Anonymous:1875:PCe

[7869] Anonymous. Publishers’ card. Scientific American, 33(25):384, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n25/pdf/scientificamerican12181875-384.pdf.

Anonymous:1875:D

[7870] Anonymous. Diphtheria. Scientific American, 33(25):384, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n25/pdf/scientificamerican12181875-384a.pdf.

Anonymous:1875:VGC

[7871] Anonymous. A valuable gift by chemistry to the world. Scientific American, 33(25):384, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
384b.pdf.

Anonymous:1875:ME

[7872] Anonymous. Making excuses. Scientific American, 33(25):384– 385, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
384c.pdf.

Anonymous:1875:WSb

[7873] Anonymous. Weak sparks. Scientific American, 33(25):385, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n25/pdf/scientificamerican12181875-385.pdf.

Anonymous:1875:WMR

[7874] Anonymous. Working Men’s reading rooms. Scientific American, 33(25):385, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
385a.pdf.

Anonymous:1875:OPL

[7875] Anonymous. That obnoxious postal law. Scientific American, 33 (25):385, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
385b.pdf.

Anonymous:1875:SGR

[7876] Anonymous. Six good reasons why every manufacturer, mechanic, inventor, and artisan should become a patron of the Scientific American. Scientific American, 33(25):385, December 18, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n25/pdf/scientificamerican12181875-385c.pdf.

Anonymous:1875:SCf

[7877] Anonymous. Steam cultivation. Scientific American, 33(25): 386, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
386.pdf.

Anonymous:1875:RIS

[7878] Anonymous. Recent improvements in the screw propeller. Scientific American, 33(25):386–387, December 18, 1875. CODEN SCAMAC.

ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
386a.pdf.

Anonymous:1875:RIF

[7879] Anonymous. Robertson’s iron fence. Scientific American, 33 (25):387, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
387.pdf.

Morton:1875:NRB

[7880] Henry Morton. A non-retreating Bunsen burner. Scientific American, 33(25):387, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
387a.pdf.

Anonymous:1875:PSM

[7881] Anonymous. Plate shearing machinery. Scientific American, 33 (25):390, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
390.pdf.

Anonymous:1875:DP

[7882] Anonymous. Donkey pump. Scientific American, 33(25):390, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n25/pdf/scientificamerican12181875-390a.pdf.

Anonymous:1875:ISCb

[7883] Anonymous. Improved screw cutting die and holder. Scientific American, 33(25):390, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
390b.pdf.

Anonymous:1875:KTH

[7884] Anonymous. The Krupp twelve hundred pounder gun. Scientific American, 33(25):391, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
391.pdf.

Anonymous:1875:TSL

[7885] Anonymous. The trinity shoals lighthouse. Scientific American, 33 (25):391–393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
391a.pdf.

Anonymous:1875:RAFbz

[7886] Anonymous. Recent American and foreign patents. Scientific American, 33(25):393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
393a.pdf.

Anonymous:1875:NMEh

[7887] Anonymous. New mechanical and engineering inventions. Scientific American, 33(25):393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
393b.pdf.

Anonymous:1875:NHAg

[7888] Anonymous. New household articles. Scientific American, 33 (25):393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
393c.pdf.

Anonymous:1875:NWHd

[7889] Anonymous. New woodwork and house and carriage building inventions. Scientific American, 33(25):393, December 18, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n25/pdf/scientificamerican12181875-393d.pdf.

Anonymous:1875:NCMg

[7890] Anonymous. New chemical and miscellaneous inventions. Scientific American, 33(25):393, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
393e.pdf.

Anonymous:1875:CRbz

[7891] Anonymous. Communications received. Scientific American, 33 (25):395, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
395.pdf.

Anonymous:1875:CPbq

[7892] Anonymous. Canadian patents. Scientific American, 33(25): 396, December 18, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n25/pdf/scientificamerican12181875-
396.pdf.

Anonymous:1875:BPbz

[7893] Anonymous. Business and personal. Scientific American, 33 (26):409, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
409.pdf.

Anonymous:1875:NQby

[7894] Anonymous. Notes and queries. Scientific American, 33(26): 409, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
409a.pdf.

Anonymous:1875:IIbz

[7895] Anonymous. Index of inventions. Scientific American, 33(26): 409–410, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
409c.pdf.

Anonymous:1875:Ij

[7896] Anonymous. Index. Scientific American, 33(26):411–412, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n26/pdf/scientificamerican12251875-411.pdf.

Anonymous:1875:GCD

[7897] Anonymous. Grand central depot signal system. Scientific American, 33(26):399, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
399.pdf.

Anonymous:1875:OL

[7898] Anonymous. The origin of life. Scientific American, 33(26):400, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v33/n26/pdf/scientificamerican12251875-400.pdf.

Anonymous:1875:HLL

[7899] Anonymous. Heavy loss from lightning rod ignorance. Scientific American, 33(26):400, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
400a.pdf.

Anonymous:1875:LCM

[7900] Anonymous. Light in coal mines. Scientific American, 33(26): 400, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
400b.pdf.

Anonymous:1875:DAF

[7901] Anonymous. The discovery of another form of electricity. Scientific American, 33(26):400–401, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
400c.pdf.

Anonymous:1875:NPE

[7902] Anonymous. The new phase of electric force. Scientific American, 33(26):401, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
401.pdf.

Anonymous:1875:TSb

[7903] Anonymous. Timely suggestions. Scientific American, 33(26): 401, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
401a.pdf.

Anonymous:1875:AAT

[7904] Anonymous. Acidemetry applied to the testing of vinegar. Scientific American, 33(26):402, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
402.pdf.

Anonymous:1875:KIS

[7905] Anonymous. King’s improved support for printers’ galleys. Scientific American, 33(26):402, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
402a.pdf.

Anonymous:1875:NDS

[7906] Anonymous. New double shaping machine. Scientific American, 33(26):403, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
403.pdf.

Anonymous:1875:FLM

[7907] Anonymous. Filters and liquid meters. Scientific American, 33 (26):404–405, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
404.pdf.

Rose:1875:PMx

[7908] Joshua Rose. Practical mechanism. Scientific American, 33(26): 405, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
405.pdf.

Anonymous:1875:MRC

[7909] Anonymous. Miller’s recoil check for guns. Scientific American, 33(26):406, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
406.pdf.

Anonymous:1875:IUW

[7910] Anonymous. Improved universal Wheelwright’s machine. Scientific American, 33(26):406, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
406a.pdf.

Anonymous:1875:ASS

[7911] Anonymous. American silk-spinning machinery. Scientific American, 33(26):406, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
406b.pdf.

Anonymous:1875:GH

[7912] Anonymous. Greenhouses and hothouses. Scientific American, 33 (26):407, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
407.pdf.

Anonymous:1875:PVI

[7913] Anonymous. Plant vases for indoor decoration. Scientific American, 33(26):407–408, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
407a.pdf.

Anonymous:1875:RAFca

[7914] Anonymous. Recent American and foreign patents. Scientific American, 33(26):408, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
408.pdf.

Anonymous:1875:NHAh

[7915] Anonymous. New household articles. Scientific American, 33 (26):408, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
408a.pdf.

Anonymous:1875:NAIe

[7916] Anonymous. New agricultural inventions. Scientific American, 33 (26):408, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
408b.pdf.

Anonymous:1875:NMEi

[7917] Anonymous. New mechanical and engineering inventions. Scientific American, 33(26):408, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
408c.pdf.

Anonymous:1875:NWHe

[7918] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 33(26):408, December 25, 1875. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v33/
n26/pdf/scientificamerican12251875-408d.pdf.

Anonymous:1875:CRca

[7919] Anonymous. Communications received. Scientific American, 33 (26):409, December 25, 1875. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v33/n26/pdf/scientificamerican12251875-
409b.pdf.

Anonymous:1876:BNBa

[7920] Anonymous. Books: New books and publications. Scientific American, 34(1):8, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
8a.pdf.

Anonymous:1876:BPa

[7921] Anonymous. Business and personal. Scientific American, 34(1):10, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-10.pdf.

Anonymous:1876:NQa

[7922] Anonymous. Notes and queries. Scientific American, 34(1):10–11, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-10a.pdf.

Anonymous:1876:IIa

[7923] Anonymous. Index of inventions. Scientific American, 34(1):11–12, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-11a.pdf.

Anonymous:1876:SM

[7924] Anonymous. Slotting machines. Scientific American, 34(1):1, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-1.pdf.

Anonymous:1876:EFW

[7925] Anonymous. Etheric force and weak electric sparks. Scientific American, 34(1):2, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
2.pdf.

Anonymous:1876:E

[7926] Anonymous. To engineers. Scientific American, 34(1):2, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-2a.pdf.

Anonymous:1876:PE

[7927] Anonymous. Powerful explosives. Scientific American, 34(1):2–3, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-2b.pdf.

Anonymous:1876:PHT

[7928] Anonymous. The Plumber’s harvest time. Scientific American, 34(1):3, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-3.pdf.

Anonymous:1876:ACM

[7929] Anonymous. The analysis of Cow’s milk. Scientific American, 34(1):3, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-3a.pdf.

Anonymous:1876:DP

[7930] Anonymous. Dangerous playthings. Scientific American, 34(1):3, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-3b.pdf.

Anonymous:1876:IVW

[7931] Anonymous. Improved vertical wood-boring machine. Scientific American, 34(1):4, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
4.pdf.

Anonymous:1876:ISCa

[7932] Anonymous. Improved show case. Scientific American, 34(1):4–5, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-4a.pdf.

Anonymous:1876:SPIa

[7933] Anonymous. Scientific and practical information. Scientific American, 34(1):5, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
5.pdf.

Anonymous:1876:GP

[7934] Anonymous. A geometrical problem. Scientific American, 34(1):5, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-5a.pdf.

Morton:1876:ANC

[7935] Henry Morton. An addition to the new chromatrope. Scientific American, 34(1):5, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
5b.pdf.

Anonymous:1876:ISSa

[7936] Anonymous. Improved scroll sawing machine. Scientific American, 34(1):6, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
6.pdf.

Anonymous:1876:ISFa

[7937] Anonymous. Improved saw file guide. Scientific American, 34(1):6, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-6a.pdf.

Anonymous:1876:CSE

[7938] Anonymous. Compounding steam engines. Scientific American, 34(1):7–8, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
7.pdf.

Anonymous:1876:TS

[7939] Anonymous. Timely suggestions. Scientific American, 34(1):8, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-8.pdf.

Anonymous:1876:DCa

[7940] Anonymous. Decisions of the courts. Scientific American, 34(1):8–9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-8b.pdf.

Anonymous:1876:RAFa

[7941] Anonymous. Recent American and foreign patents. Scientific American, 34(1):9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
9.pdf.

Anonymous:1876:NTMa

[7942] Anonymous. New textile machinery. Scientific American, 34(1):9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-9a.pdf.

Anonymous:1876:NHAa

[7943] Anonymous. New household articles. Scientific American, 34(1):9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-9b.pdf.

Anonymous:1876:NAIa

[7944] Anonymous. New agricultural inventions. Scientific American, 34(1):9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-9c.pdf.

Anonymous:1876:NMEa

[7945] Anonymous. New mechanical and engineering inventions. Scientific American, 34(1):9, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n1/pdf/scientificamerican01011876-
9d.pdf.

Anonymous:1876:CRa

[7946] Anonymous. Communications received. Scientific American, 34(1):11, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-11.pdf.

Anonymous:1876:CPa

[7947] Anonymous. Canadian patents. Scientific American, 34(1):12, January 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n1/pdf/scientificamerican01011876-12.pdf.

Anonymous:1876:Ca

[7948] Anonymous. Correspondence. Scientific American, 34(2):20, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-20a.pdf.

Anonymous:1876:BPb

[7949] Anonymous. Business and personal. Scientific American, 34(2):26, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-26.pdf.

Anonymous:1876:NQb

[7950] Anonymous. Notes and queries. Scientific American, 34(2):26–27, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-26a.pdf.

Anonymous:1876:IIb

[7951] Anonymous. Index of inventions. Scientific American, 34(2):27–28, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-27a.pdf.

Anonymous:1876:NYAa

[7952] Anonymous. The New York anchorage of the East River Bridge. Scientific American, 34(2):15–16, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
15.pdf.

Anonymous:1876:NUE

[7953] Anonymous. A new use for the East River Bridge. Scientific American, 34(2):16, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
16.pdf.

Anonymous:1876:EEa

[7954] Anonymous. Electricity as an executioner. Scientific American, 34(2):16, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
16a.pdf.

Anonymous:1876:IBc

[7955] Anonymous. Is insurance a benefit? Scientific American, 34(2):17, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-17.pdf.

Anonymous:1876:MEE

[7956] Anonymous. Mr. Edison’s electric discovery. Scientific American, 34(2):17, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
17a.pdf.

Anonymous:1876:MDa

[7957] Anonymous. Mechanical drawing. Scientific American, 34(2):17, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-17b.pdf.

Anonymous:1876:SASa

[7958] Anonymous. The Scientific American supplement. Scientific American, 34(2):17, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
17c.pdf.

Anonymous:1876:KSP

[7959] Anonymous. The Knowles steam pumps. Scientific American, 34(2):18–20, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
18.pdf.

Anonymous:1876:ANa

[7960] Anonymous. Astronomical notes. Scientific American, 34(2):20, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-20.pdf.

Anonymous:1876:SDO

[7961] Anonymous. Seeing distant objects from elevations. Scientific American, 34(2):20, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
20b.pdf.

Rose:1876:PMa

[7962] Joshua Rose. Practical mechanism. Scientific American, 34(2):21, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-21.pdf.

Anonymous:1876:IRRa

[7963] Anonymous. Improved railway rail joint. Scientific American, 34(2):22, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-22.pdf.

Anonymous:1876:IFGa

[7964] Anonymous. Improved furnace grate bar. Scientific American, 34(2):22, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-22a.pdf.

Anonymous:1876:Ab

[7965] Anonymous. The alocasias. Scientific American, 34(2):23, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-23.pdf.

Anonymous:1876:GR

[7966] Anonymous. The Guelder rose. Scientific American, 34(2):23, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-23a.pdf.

Anonymous:1876:BBH

[7967] Anonymous. Bridles, bits, hubs, and horseshoes. Scientific American, 34(2):24–25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
24.pdf.

Anonymous:1876:RCF

[7968] Anonymous. Reduction in the cost of foreign patents. Scientific American, 34(2):25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
25.pdf.

Anonymous:1876:RAFb

[7969] Anonymous. Recent American and foreign patents. Scientific American, 34(2):25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
25a.pdf.

Anonymous:1876:NMEb

[7970] Anonymous. New mechanical and engineering inventions. Scientific American, 34(2):25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n2/pdf/scientificamerican01081876-
25b.pdf.

Anonymous:1876:NHAb

[7971] Anonymous. New household articles. Scientific American, 34(2):25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-25c.pdf.

Anonymous:1876:NAIb

[7972] Anonymous. New agricultural inventions. Scientific American, 34(2):25, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-25d.pdf.

Anonymous:1876:CRb

[7973] Anonymous. Communications received. Scientific American, 34(2):27, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-27.pdf.

Anonymous:1876:CPb

[7974] Anonymous. Canadian patents. Scientific American, 34(2):28, January 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n2/pdf/scientificamerican01081876-28.pdf.

Anonymous:1876:Cb

[7975] Anonymous. Correspondence. Scientific American, 34(3):36–37, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-36.pdf.

Anonymous:1876:BNBb

[7976] Anonymous. Books: New books and publications. Scientific American, 34(3):41, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
41a.pdf.

Anonymous:1876:BPc

[7977] Anonymous. Business and personal. Scientific American, 34(3):43, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-43.pdf.

Anonymous:1876:NQc

[7978] Anonymous. Notes and queries. Scientific American, 34(3):43–44, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-43a.pdf.

Anonymous:1876:IIc

[7979] Anonymous. Index of inventions. Scientific American, 34(3):44, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-44a.pdf.

Anonymous:1876:MR

[7980] Anonymous. The Mexican Railway. Scientific American, 34(3):31, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-31.pdf.

Anonymous:1876:WWM

[7981] Anonymous. What working men should exhibit at the Centennial. Scientific American, 34(3):32, January 15, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n3/pdf/scientificamerican01151876-32.pdf.

Anonymous:1876:SCN

[7982] Anonymous. Stanley, Cameron, and Nordenskjold. Scientific American, 34(3):32, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
32a.pdf.

Anonymous:1876:PIa

[7983] Anonymous. Photographing the invisible. Scientific American, 34 (3):32–33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
32b.pdf.

Anonymous:1876:OME

[7984] Anonymous. One of Mr. Edison’s curious experiments. Scientific American, 34(3):33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
33.pdf.

Anonymous:1876:WIB

[7985] Anonymous. Who invented the barrel? Scientific American, 34(3):33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
33a.pdf.

Anonymous:1876:LCa

[7986] Anonymous. Lower California. Scientific American, 34(3):33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-33b.pdf.

Anonymous:1876:RBB

[7987] Anonymous. Rapid bridge building. Scientific American, 34(3):33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-33c.pdf.

Anonymous:1876:WPa

[7988] Anonymous. Water in the pipes. Scientific American, 34(3):33, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-33d.pdf.

Anonymous:1876:IMCa

[7989] Anonymous. Improvements in the manufacture of Coke. Scientific American, 34(3):34, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
34.pdf.

Anonymous:1876:IAE

[7990] Anonymous. Improved apparatus for emptying aquaria. Scientific American, 34(3):34, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
34a.pdf.

Anonymous:1876:VM

[7991] Anonymous. Voting by machinery. Scientific American, 34(3):35, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-35.pdf.

Anonymous:1876:IMa

[7992] Anonymous. Improved mill. Scientific American, 34(3):35–36, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-35a.pdf.

Anonymous:1876:ISCb

[7993] Anonymous. Improved still column. Scientific American, 34(3):38, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-38.pdf.

Anonymous:1876:IBF

[7994] Anonymous. Improved barrel faucet bushing. Scientific American, 34(3):38, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
38a.pdf.

Anonymous:1876:J

[7995] Anonymous. The jaguars. Scientific American, 34(3):39–40, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-39.pdf.

Anonymous:1876:NNT

[7996] Anonymous. The new nebular theory — possible work for the astronomers. Scientific American, 34(3):40, January 15, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n3/pdf/scientificamerican01151876-40.pdf.

Anonymous:1876:SPIb

[7997] Anonymous. Scientific and practical information. Scientific American, 34(3):41, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
41.pdf.

Anonymous:1876:DCb

[7998] Anonymous. Decisions of the courts. Scientific American, 34(3):41, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-41b.pdf.

Anonymous:1876:RAFc

[7999] Anonymous. Recent American and foreign patents. Scientific American, 34(3):41–42, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
41c.pdf.

Anonymous:1876:NMEc

[8000] Anonymous. New mechanical and engineering inventions. Scientific American, 34(3):42, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
42.pdf.

Anonymous:1876:NWHa

[8001] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(3):42, January 15, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n3/pdf/scientificamerican01151876-42a.pdf.

Anonymous:1876:NAIc

[8002] Anonymous. New agricultural inventions. Scientific American, 34(3):42, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
42b.pdf.

Anonymous:1876:NHAc

[8003] Anonymous. New household articles. Scientific American, 34 (3):42, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
42c.pdf.

Anonymous:1876:NTMb

[8004] Anonymous. New textile machinery. Scientific American, 34(3):42, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-42d.pdf.

Anonymous:1876:CRc

[8005] Anonymous. Communications received. Scientific American, 34 (3):44, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n3/pdf/scientificamerican01151876-
44.pdf.

Anonymous:1876:CPc

[8006] Anonymous. Canadian patents. Scientific American, 34(3):44, January 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n3/pdf/scientificamerican01151876-44b.pdf.

Anonymous:1876:BNBc

[8007] Anonymous. Books: New books and publications. Scientific American, 34(4):57, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
57b.pdf.

Anonymous:1876:BPd

[8008] Anonymous. Business and personal. Scientific American, 34(4):59, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-59.pdf.

Anonymous:1876:NQd

[8009] Anonymous. Notes and queries. Scientific American, 34(4):59, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-59a.pdf.

Anonymous:1876:IId

[8010] Anonymous. Index of inventions. Scientific American, 34(4):59–60, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-59c.pdf.

Anonymous:1876:NRDa

[8011] Anonymous. New railroad depot. Scientific American, 34(4):47, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-47.pdf.

Anonymous:1876:WMV

[8012] Anonymous. Working Men’s visits to the Centennial. Scientific American, 34(4):48, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
48.pdf.

Anonymous:1876:TP

[8013] Anonymous. Theory and practice. Scientific American, 34(4):48, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-48a.pdf.

Anonymous:1876:DI

[8014] Anonymous. Drawbridge inconsistency. Scientific American, 34 (4):48–49, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
48b.pdf.

Anonymous:1876:CEa

[8015] Anonymous. The Centennial Exhibition. Scientific American, 34 (4):49, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
49.pdf.

Anonymous:1876:BSP

[8016] Anonymous. Bamboo as a source of paper stock. Scientific American, 34(4):49, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
49a.pdf.

Anonymous:1876:OWS

[8017] Anonymous. Our water supply. Scientific American, 34(4):49, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-49b.pdf.

Anonymous:1876:SPIc

[8018] Anonymous. Scientific and practical information. Scientific American, 34(4):49, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
49c.pdf.

Anonymous:1876:IMMa

[8019] Anonymous. Improved hat-making machinery. Scientific American, 34(4):50, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
50.pdf.

Anonymous:1876:KHM

[8020] Anonymous. The Knowles horizontal mining pump. Scientific American, 34(4):51, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
51.pdf.

Anonymous:1876:SHSa

[8021] Anonymous. Steam horse for street railways. Scientific American, 34(4):51–52, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
51a.pdf.

Rose:1876:PMb

[8022] Joshua Rose. Practical mechanism. Scientific American, 34(4):53, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-53.pdf.

Anonymous:1876:ICPa

[8023] Anonymous. Improved cotton planter. Scientific American, 34 (4):54, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
54.pdf.

Anonymous:1876:IV

[8024] Anonymous. Improved vise. Scientific American, 34(4):54, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-54a.pdf.

Anonymous:1876:PGE

[8025] Anonymous. Potin’s galvano-electric bath. Scientific American, 34(4):54, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
54b.pdf.

Anonymous:1876:GSa

[8026] Anonymous. The gardens at Schönbrunn. Scientific American, 34(4):55, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
55.pdf.

Anonymous:1876:CCM

[8027] Anonymous. Cheese and cider making and lard rendering. Scientific American, 34(4):56, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
56.pdf.

Beard:1876:NND

[8028] George M. Beard. The nature of the newly discovered force. Scientific American, 34(4):57, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
57.pdf.

Anonymous:1876:DCc

[8029] Anonymous. Decisions of the courts. Scientific American, 34(4):57, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-57a.pdf.

Anonymous:1876:RAFd

[8030] Anonymous. Recent American and foreign patents. Scientific American, 34(4):58, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
58.pdf.

Anonymous:1876:NMEd

[8031] Anonymous. New mechanical and engineering inventions. Scientific American, 34(4):58, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
58a.pdf.

Anonymous:1876:NWHb

[8032] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(4):58, January 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n4/pdf/scientificamerican01221876-58b.pdf.

Anonymous:1876:NCMa

[8033] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(4):58, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
58c.pdf.

Anonymous:1876:NAId

[8034] Anonymous. New agricultural inventions. Scientific American, 34(4):58, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
58d.pdf.

Anonymous:1876:CRd

[8035] Anonymous. Communications received. Scientific American, 34 (4):59, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n4/pdf/scientificamerican01221876-
59b.pdf.

Anonymous:1876:CPd

[8036] Anonymous. Canadian patents. Scientific American, 34(4):60, January 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n4/pdf/scientificamerican01221876-60.pdf.

Anonymous:1876:Cc

[8037] Anonymous. Correspondence. Scientific American, 34(5):69, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-69.pdf.

Anonymous:1876:BPe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-74.pdf.

Anonymous:1876:NQe

[8039] Anonymous. Notes and queries. Scientific American, 34(5):74–75, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-74a.pdf.

Anonymous:1876:IIe

[8040] Anonymous. Index of inventions. Scientific American, 34(5):76, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-76.pdf.

Anonymous:1876:IGAa

[8041] Anonymous. Improved gas apparatus. Scientific American, 34 (5):63, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
63.pdf.

Anonymous:1876:SHTa

[8042] Anonymous. Steam on the highways — ten thousand dollars reward. Scientific American, 34(5):64, January 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n5/pdf/scientificamerican01291876-64.pdf.

Anonymous:1876:GNTa

[8043] Anonymous. Galton’s new theory of heredity. Scientific American, 34(5):64–65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
64a.pdf.

Anonymous:1876:IMb

[8044] Anonymous. Infernal machines. Scientific American, 34(5):65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-65.pdf.

Anonymous:1876:SSIa

[8045] Anonymous. Safe savings. — an improvement needed. Scientific American, 34(5):65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
65a.pdf.

Anonymous:1876:IFPa

[8046] Anonymous. Improved fire plug. Scientific American, 34(5):66, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-66.pdf.

Anonymous:1876:RCEa

[8047] Anonymous. The rider compression engine. Scientific American, 34(5):66, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
66a.pdf.

Anonymous:1876:Cd

[8048] Anonymous. Coining. Scientific American, 34(5):68, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-68.pdf.

Anonymous:1876:BCa

[8049] Anonymous. The Buddha crab. Scientific American, 34(5):68, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-68a.pdf.

Anonymous:1876:IBSa

[8050] Anonymous. Improved band sawing machine. Scientific American, 34(5):70, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
70.pdf.

Anonymous:1876:SSGa

[8051] Anonymous. A sensible and substantial grinding machine. Scientific American, 34(5):70–71, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
70a.pdf.

Anonymous:1876:PFa

[8052] Anonymous. The penguin family. Scientific American, 34(5):71, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-71.pdf.

Anonymous:1876:SPId

[8053] Anonymous. Scientific and practical information. Scientific American, 34(5):72, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
72.pdf.

Anonymous:1876:Pa

[8054] Anonymous. The patents of 1875. Scientific American, 34(5): 72–73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
72a.pdf.

Anonymous:1876:DCd

[8055] Anonymous. Decisions of the courts. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-73.pdf.

Anonymous:1876:RAFe

[8056] Anonymous. Recent American and foreign patents. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
73a.pdf.

Anonymous:1876:NCMb

[8057] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
73b.pdf.

Anonymous:1876:NAIe

[8058] Anonymous. New agricultural inventions. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
73c.pdf.

Anonymous:1876:NMEe

[8059] Anonymous. New mechanical and engineering inventions. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
73d.pdf.

Anonymous:1876:NHAd

[8060] Anonymous. New household articles. Scientific American, 34 (5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
73e.pdf.

Anonymous:1876:CRe

[8061] Anonymous. Communications received. Scientific American, 34 (5):75, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n5/pdf/scientificamerican01291876-
75.pdf.

Anonymous:1876:CPe

[8062] Anonymous. Canadian patents. Scientific American, 34(5):76, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n5/pdf/scientificamerican01291876-76a.pdf.

Anonymous:1876:Cf

[8063] Anonymous. Correspondence. Scientific American, 34(5):69, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:BPf

[8064] Anonymous. Business and personal. Scientific American, 34(5):74, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:NQf

[8065] Anonymous. Notes and queries. Scientific American, 34(5):74–75, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IIf

[8066] Anonymous. Index of inventions. Scientific American, 34(5):76, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:Cg

[8067] Anonymous. Correspondence. Scientific American, 34(5):84, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-84.pdf.

Anonymous:1876:BNBd

[8068] Anonymous. Books: New books and publications. Scientific American, 34(5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
90.pdf.

Anonymous:1876:BPg

[8069] Anonymous. Business and personal. Scientific American, 34(5):91, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-91.pdf.

Anonymous:1876:NQg

[8070] Anonymous. Notes and queries. Scientific American, 34(5):91, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-91a.pdf.

Anonymous:1876:IIg

[8071] Anonymous. Index of inventions. Scientific American, 34(5):91–92, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-91c.pdf.

Anonymous:1876:IGAb

[8072] Anonymous. Improved gas apparatus. Scientific American, 34(5):63, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:SHTb

[8073] Anonymous. Steam on the highways — ten thousand dollars reward. Scientific American, 34(5):64, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:GNTb

[8074] Anonymous. Galton’s new theory of heredity. Scientific American, 34(5):64–65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IMd

[8075] Anonymous. Infernal machines. Scientific American, 34(5):65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:SSIb

[8076] Anonymous. Safe savings. — an improvement needed. Scientific American, 34(5):65, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IFPb

[8077] Anonymous. Improved fire plug. Scientific American, 34(5):66, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:RCEb

[8078] Anonymous. The rider compression engine. Scientific American, 34 (5):66, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:Ch

[8079] Anonymous. Coining. Scientific American, 34(5):68, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:BCb

[8080] Anonymous. The Buddha crab. Scientific American, 34(5):68, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IBSb

[8081] Anonymous. Improved band sawing machine. Scientific American, 34 (5):70, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:SSGb

[8082] Anonymous. A sensible and substantial grinding machine. Scientific American, 34(5):70–71, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:PFb

[8083] Anonymous. The penguin family. Scientific American, 34(5):71, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:SPIe

[8084] Anonymous. Scientific and practical information. Scientific American, 34(5):72, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:Pb

[8085] Anonymous. The patents of 1875. Scientific American, 34(5):72–73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:DCe

[8086] Anonymous. Decisions of the courts. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:RAFf

[8087] Anonymous. Recent American and foreign patents. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:NCMc

[8088] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:NAIf

[8089] Anonymous. New agricultural inventions. Scientific American, 34(5): 73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:NMEf

[8090] Anonymous. New mechanical and engineering inventions. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:NHAe

[8091] Anonymous. New household articles. Scientific American, 34(5):73, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:CRf

[8092] Anonymous. Communications received. Scientific American, 34(5):75, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:CPf

[8093] Anonymous. Canadian patents. Scientific American, 34(5):76, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IRF

[8094] Anonymous. Improved revolving furnace. Scientific American, 34(5):79, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
79.pdf.

Anonymous:1876:SR

[8095] Anonymous. “science record” for 1876. Scientific American, 34 (5):80, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
80.pdf.

Anonymous:1876:MM

[8096] Anonymous. Machine mercy. Scientific American, 34(5):80, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-80a.pdf.

Anonymous:1876:PGJ

[8097] Anonymous. Progress of the great jetty works at the mouth of the Mississippi. Scientific American, 34(5):80, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-80b.pdf.

Anonymous:1876:RSC

[8098] Anonymous. Removal of snow from city railway tracks. Scientific American, 34(5):80–81, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
80c.pdf.

Anonymous:1876:HGBa

[8099] Anonymous. High gas bills and their prevention. Scientific American, 34(5):81, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
81.pdf.

Anonymous:1876:NIB

[8100] Anonymous. A neglected industry — bee culture. Scientific American, 34(5):81, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
81a.pdf.

Anonymous:1876:NAR

[8101] Anonymous. A new air refrigerator. Scientific American, 34(5):82, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-82.pdf.

Anonymous:1876:WME

[8102] Anonymous. Woodworking machinery in England. Scientific American, 34(5):82, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
82a.pdf.

Anonymous:1876:MDb

[8103] Anonymous. The manufacture of dynamite. Scientific American, 34(5):83, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
83.pdf.

Anonymous:1876:BSA

[8104] Anonymous. Boot and shoe apparatus. Scientific American, 34 (5):83–84, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
83a.pdf.

Rose:1876:PMc

[8105] Joshua Rose. Practical mechanism. Scientific American, 34(5): 84–85, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
84a.pdf.

Anonymous:1876:ISA

[8106] Anonymous. Improved sulphurous acid apparatus for bleaching cane juice. Scientific American, 34(5):86, January 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n6/pdf/scientificamerican02051876-86.pdf.

Anonymous:1876:HCE

[8107] Anonymous. Harding’s cushion emery wheel. Scientific American, 34(5):86, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
86a.pdf.

Anonymous:1876:NIC

[8108] Anonymous. Noel’s improved corn planter. Scientific American, 34(5):86, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
86b.pdf.

Anonymous:1876:ITM

[8109] Anonymous. Improved thrashing machinery. Scientific American, 34(5):87–88, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
87.pdf.

Anonymous:1876:SPIf

[8110] Anonymous. Scientific and practical information. Scientific American, 34(5):88, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
88.pdf.

Anonymous:1876:ANb

[8111] Anonymous. Astronomical notes. Scientific American, 34(5):88, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-88a.pdf.

VanderWeyde:1876:NPD

[8112] P. H. Vander Weyde. The nature of the phenomena discovered by Mr. Edison. Scientific American, 34(5):89, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-89.pdf.

Anonymous:1876:RAFg

[8113] Anonymous. Recent American and foreign patents. Scientific American, 34(5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
90a.pdf.

Anonymous:1876:NWHc

[8114] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(5):90, January 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n6/pdf/scientificamerican02051876-90b.pdf.

Anonymous:1876:NHAf

[8115] Anonymous. New household articles. Scientific American, 34 (5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
90c.pdf.

Anonymous:1876:NTMc

[8116] Anonymous. New textile machinery. Scientific American, 34(5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-90d.pdf.

Anonymous:1876:NAIg

[8117] Anonymous. New agricultural inventions. Scientific American, 34(5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
90e.pdf.

Anonymous:1876:NCMd

[8118] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(5):90, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
90f.pdf.

Anonymous:1876:CRg

[8119] Anonymous. Communications received. Scientific American, 34 (5):91, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n6/pdf/scientificamerican02051876-
91b.pdf.

Anonymous:1876:CPg

[8120] Anonymous. Canadian patents. Scientific American, 34(5):92, January 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n6/pdf/scientificamerican02051876-92.pdf.

Anonymous:1876:Ci

[8121] Anonymous. Correspondence. Scientific American, 34(7):100– 101, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
100.pdf.

Anonymous:1876:BNBe

[8122] Anonymous. Books: New books and publications. Scientific American, 34(7):105, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
105.pdf.

Anonymous:1876:BPh

[8123] Anonymous. Business and personal. Scientific American, 34(7): 106, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
106.pdf.

Anonymous:1876:NQh

[8124] Anonymous. Notes and queries. Scientific American, 34(7):106– 107, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
106a.pdf.

Anonymous:1876:IIh

[8125] Anonymous. Index of inventions. Scientific American, 34(7): 107–108, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
107a.pdf.

Anonymous:1876:NTP

[8126] Anonymous. The new telescope in the Paris Observatory. Scientific American, 34(7):95–96, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
95.pdf.

Anonymous:1876:RDR

[8127] Anonymous. Re-discoveries and re-inventions. Scientific American, 34(7):96, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
96.pdf.

Anonymous:1876:FU

[8128] Anonymous. Felting and its uses. Scientific American, 34(7): 96–97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
96a.pdf.

Anonymous:1876:NR

[8129] Anonymous. No rest. Scientific American, 34(7):97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-97.pdf.

Anonymous:1876:TWN

[8130] Anonymous. Timber waste a national suicide. Scientific American, 34(7):97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
97a.pdf.

Anonymous:1876:SWU

[8131] Anonymous. Shall we undergo dissection? Scientific American, 34(7):97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
97b.pdf.

Anonymous:1876:PMB

[8132] Anonymous. Patent matters before Congress. Scientific American, 34(7):97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
97c.pdf.

Anonymous:1876:AFa

[8133] Anonymous. The Atlantic ferry. Scientific American, 34(7):97, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-97d.pdf.

Anonymous:1876:RIL

[8134] Anonymous. Recent improvements in locomotive engines. Scientific American, 34(7):98, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
98.pdf.

Anonymous:1876:IACa

[8135] Anonymous. Improved automatic cradle. Scientific American, 34(7):99, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
99.pdf.

Anonymous:1876:BIC

[8136] Anonymous. Brooks’ improved clothes dryer. Scientific American, 34(7):99, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
99a.pdf.

Anonymous:1876:NSTa

[8137] Anonymous. A new steam trap. Scientific American, 34(7):99, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-99b.pdf.

Anonymous:1876:IAAa

[8138] Anonymous. Improved adjustable almanac. Scientific American, 34(7):102, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
102.pdf.

Anonymous:1876:IJS

[8139] Anonymous. Improved jump seat for carriages. Scientific American, 34(7):102, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
102a.pdf.

Anonymous:1876:FH

[8140] Anonymous. The fish hawks. Scientific American, 34(7):103, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-103.pdf.

Anonymous:1876:PL

[8141] Anonymous. Power looms. Scientific American, 34(7):104, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-104.pdf.

Anonymous:1876:RAFh

[8142] Anonymous. Recent American and foreign patents. Scientific American, 34(7):105, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
105a.pdf.

Anonymous:1876:NCMe

[8143] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(7):105, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
105b.pdf.

Anonymous:1876:NAIh

[8144] Anonymous. New agricultural inventions. Scientific American, 34(7):105, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
105c.pdf.

Anonymous:1876:NWHd

[8145] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(7):105, February 12, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n7/pdf/scientificamerican02121876-105d.pdf.

Anonymous:1876:CRh

[8146] Anonymous. Communications received. Scientific American, 34 (7):107, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n7/pdf/scientificamerican02121876-
107.pdf.

Anonymous:1876:CPh

[8147] Anonymous. Canadian patents. Scientific American, 34(7):108, February 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n7/pdf/scientificamerican02121876-108.pdf.

Anonymous:1876:Cj

[8148] Anonymous. Correspondence. Scientific American, 34(8):116, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n8/pdf/scientificamerican02191876-116.pdf.

Anonymous:1876:BNBf

[8149] Anonymous. Books: New books and publications. Scientific American, 34(8):120, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
120a.pdf.

Anonymous:1876:BPi

[8150] Anonymous. Business and personal. Scientific American, 34(8): 122, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
122.pdf.

Anonymous:1876:NQi

[8151] Anonymous. Notes and queries. Scientific American, 34(8):122– 123, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
122a.pdf.

Anonymous:1876:IIi

[8152] Anonymous. Index of inventions. Scientific American, 34(8): 123–124, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
123a.pdf.

Anonymous:1876:IDC

[8153] Anonymous. Improved dumping cart. Scientific American, 34(8):112– 113, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:KWP

[8154] Anonymous. Kentucky weather prophets. Scientific American, 34(8): 113, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:OD

[8155] Anonymous. Opium dangers. Scientific American, 34(8):113, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:AA

[8156] Anonymous. The aristocracy of all. Scientific American, 34(8):113, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Anonymous:1876:IO

[8157] Anonymous. Improvements in the organ. Scientific American, 34(8):113, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
113c.pdf.

Anonymous:1876:NWM

[8158] Anonymous. New woodworking machinery. Scientific American, 34(8):114, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
114.pdf.

Anonymous:1876:ASE

[8159] Anonymous. Automatic sprinkler for extinguishing fires. Scientific American, 34(8):114, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
114a.pdf.

Anonymous:1876:ITT

[8160] Anonymous. Improved tools for threading bolts and nuts. Scientific American, 34(8):115, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
115.pdf.

Anonymous:1876:NBa

[8161] Anonymous. A new blowpipe. Scientific American, 34(8):115, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n8/pdf/scientificamerican02191876-115a.pdf.

Anonymous:1876:NAE

[8162] Anonymous. New applications of electricity to organ building. Scientific American, 34(8):117, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
117.pdf.

Anonymous:1876:ISFb

[8163] Anonymous. Improved sash fastener. Scientific American, 34 (8):118, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
118.pdf.

Anonymous:1876:ISGa

[8164] Anonymous. Improved sewer gas trap. Scientific American, 34 (8):118, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
118a.pdf.

Anonymous:1876:IBBa

[8165] Anonymous. Improved brush binder. Scientific American, 34 (8):118, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
118b.pdf.

Anonymous:1876:ICRa

[8166] Anonymous. Improved carpet rag looper. Scientific American, 34(8):118, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
118c.pdf.

Anonymous:1876:ICBa

[8167] Anonymous. Improved countersink for boring tools. Scientific American, 34(8):119, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
119.pdf.

Anonymous:1876:HDG

[8168] Anonymous. Household: Devices, gates, and hinges. Scientific American, 34(8):119–120, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
119a.pdf.

Anonymous:1876:DCf

[8169] Anonymous. Decisions of the courts. Scientific American, 34 (8):120, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
120.pdf.

Anonymous:1876:RAFi

[8170] Anonymous. Recent American and foreign patents. Scientific American, 34(8):120–121, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
120b.pdf.

Anonymous:1876:NAIi

[8171] Anonymous. New agricultural inventions. Scientific American, 34(8):121, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
121.pdf.

Anonymous:1876:NTMd

[8172] Anonymous. New textile machinery. Scientific American, 34(8): 121, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
121a.pdf.

Anonymous:1876:NWHe

[8173] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(8):121, February 19, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n8/pdf/scientificamerican02191876-121b.pdf.

Anonymous:1876:NCMf

[8174] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(8):121, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
121c.pdf.

Anonymous:1876:CRi

[8175] Anonymous. Communications received. Scientific American, 34 (8):123, February 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n8/pdf/scientificamerican02191876-
123.pdf.

Anonymous:1876:Ck

[8176] Anonymous. Correspondence. Scientific American, 34(9):132, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n9/pdf/scientificamerican02261876-132.pdf.

Anonymous:1876:BPj

[8177] Anonymous. Business and personal. Scientific American, 34(9): 138, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
138.pdf.

Anonymous:1876:NQj

[8178] Anonymous. Notes and queries. Scientific American, 34(9):138– 140, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
138a.pdf.

Anonymous:1876:IIj

[8179] Anonymous. Index of inventions. Scientific American, 34(9):140, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n9/pdf/scientificamerican02261876-140a.pdf.

Anonymous:1876:IPE

[8180] Anonymous. Improved pumping engine for water works. Scientific American, 34(9):127, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
127.pdf.

Anonymous:1876:WCM

[8181] Anonymous. What the coming man may be. Scientific American, 34(9):128, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
128.pdf.

Anonymous:1876:MW

[8182] Anonymous. A Man’s work. Scientific American, 34(9):128, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n9/pdf/scientificamerican02261876-128a.pdf.

Anonymous:1876:ORA

[8183] Anonymous. The oil resources of Africa. Scientific American, 34 (9):128–129, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
128b.pdf.

Anonymous:1876:ANI

[8184] Anonymous. Another neglected industry — mushroom raising. Scientific American, 34(9):129, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
129.pdf.

Anonymous:1876:CWP

[8185] Anonymous. Can we protect our bank vaults? Scientific American, 34(9):129, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
129a.pdf.

Anonymous:1876:RPE

[8186] Anonymous. Remarkable pumping engines. Scientific American, 34(9):129, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
129b.pdf.

Anonymous:1876:NAF

[8187] Anonymous. New automatic feed water purifier. Scientific American, 34(9):130, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
130.pdf.

Anonymous:1876:PIW

[8188] Anonymous. Port’s improved wrench. Scientific American, 34 (9):131, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
131.pdf.

Anonymous:1876:ILM

[8189] Anonymous. Improved lathing machine. Scientific American, 34 (9):131, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
131a.pdf.

Anonymous:1876:ISCc

[8190] Anonymous. Improved submerged current wheel. Scientific American, 34(9):131, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
131b.pdf.

Anonymous:1876:SFB

[8191] Anonymous. A simple form of blowpipe. Scientific American, 34(9):131, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
131c.pdf.

Anonymous:1876:IBA

[8192] Anonymous. Improved barbers’ apparatus. Scientific American, 34 (9):131–132, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
131d.pdf.

Anonymous:1876:NM

[8193] Anonymous. A new medicament. Scientific American, 34(9): 132–133, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
132a.pdf.

Rose:1876:PMd

[8194] Joshua Rose. Practical mechanism. Scientific American, 34(9): 133, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
133.pdf.

Anonymous:1876:ILD

[8195] Anonymous. Improved leather-dressing machine. Scientific American, 34(9):134, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
134.pdf.

Anonymous:1876:WRI

[8196] Anonymous. Water regulator and indicator for steam boilers. Scientific American, 34(9):134, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
134a.pdf.

Anonymous:1876:CWG

[8197] Anonymous. A Caucasian wild geranium. Scientific American, 34(9):135, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
135.pdf.

Anonymous:1876:EPA

[8198] Anonymous. The extermination of the phylloxera by alkaline sulphocarbonates. Scientific American, 34(9):135–136, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n9/pdf/scientificamerican02261876-135a.pdf.

Anonymous:1876:SPIg

[8199] Anonymous. Scientific and practical information. Scientific American, 34(9):136–137, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
136.pdf.

Anonymous:1876:DCg

[8200] Anonymous. Decisions of the courts. Scientific American, 34 (9):137, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
137.pdf.

Anonymous:1876:RAFj

[8201] Anonymous. Recent American and foreign patents. Scientific American, 34(9):137, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
137a.pdf.

Anonymous:1876:NWHf

[8202] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(9):137, February 26, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n9/pdf/scientificamerican02261876-137b.pdf.

Anonymous:1876:NCMg

[8203] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(9):137, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
137c.pdf.

Anonymous:1876:NAIj

[8204] Anonymous. New agricultural inventions. Scientific American, 34(9):137, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
137d.pdf.

Anonymous:1876:CRj

[8205] Anonymous. Communications received. Scientific American, 34 (9):140, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n9/pdf/scientificamerican02261876-
140.pdf.

Anonymous:1876:CPi

[8206] Anonymous. Canadian patents. Scientific American, 34(9):140, February 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n9/pdf/scientificamerican02261876-140b.pdf.

Anonymous:1876:Cl

[8207] Anonymous. Correspondence. Scientific American, 34(10):148, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-148.pdf.

Anonymous:1876:BPk

[8208] Anonymous. Business and personal. Scientific American, 34(10):154, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-154.pdf.

Anonymous:1876:NQk

[8209] Anonymous. Notes and queries. Scientific American, 34(10):154–155, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-154a.pdf.

Anonymous:1876:IIk

[8210] Anonymous. Index of inventions. Scientific American, 34(10):155–156, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-155a.pdf.

Anonymous:1876:ACDa

[8211] Anonymous. Australian coast defences. Scientific American, 34 (10):143–144, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
143.pdf.

Anonymous:1876:IASa

[8212] Anonymous. Industrial art. — some thoughts for the Centennial. Scientific American, 34(10):144, March 4, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n10/pdf/scientificamerican03041876-144.pdf.

Anonymous:1876:RAE

[8213] Anonymous. A railroad across the Eastern Continent. Scientific American, 34(10):144, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
144a.pdf.

Anonymous:1876:EC

[8214] Anonymous. The education of crows. Scientific American, 34 (10):144–145, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
144b.pdf.

Anonymous:1876:CBPa

[8215] Anonymous. The Centennial Bill passed. Scientific American, 34(10):145, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
145.pdf.

Anonymous:1876:CSA

[8216] Anonymous. The collapse of the Sheffield American steel rail trade. Scientific American, 34(10):145, March 4, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n10/pdf/scientificamerican03041876-145a.pdf.

VanderWeyde:1876:IT

[8217] P. H. Vander Weyde. The invention of the telephone. Scientific American, 34(10):145, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
145b.pdf.

Anonymous:1876:ICCa

[8218] Anonymous. Improved construction of conservatories. Scientific American, 34(10):146, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
146.pdf.

Anonymous:1876:W

[8219] Anonymous. Wasps. Scientific American, 34(10):146, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-146a.pdf.

Anonymous:1876:NCF

[8220] Anonymous. A new centrifugal filter. Scientific American, 34(10):147, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-147.pdf.

Anonymous:1876:PCL

[8221] Anonymous. Portable crane for locomotive works. Scientific American, 34(10):147–148, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
147a.pdf.

Falke:1876:NPSa

[8222] William Falke. New phosphides of silver, and a method of estimating silver quantitatively by means of phosphorus. Scientific American, 34(10):148–149, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
148a.pdf.

Anonymous:1876:IHT

[8223] Anonymous. Improved hay tedder. Scientific American, 34(10):150, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-150.pdf.

Anonymous:1876:NRP

[8224] Anonymous. A new registering pluviometer. Scientific American, 34(10):150, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
150a.pdf.

Anonymous:1876:RDH

[8225] Anonymous. Rock-drilling by hand power. Scientific American, 34(10):150, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
150b.pdf.

Anonymous:1876:CPj

[8226] Anonymous. The Chromis Paterfamilias. Scientific American, 34 (10):151–152, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
151.pdf.

Anonymous:1876:MPI

[8227] Anonymous. Medical and philosophical instruments. Scientific American, 34(10):152–153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
152.pdf.

Anonymous:1876:DCh

[8228] Anonymous. Decisions of the courts. Scientific American, 34(10):153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-153.pdf.

Anonymous:1876:RAFk

[8229] Anonymous. Recent American and foreign patents. Scientific American, 34(10):153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
153a.pdf.

Anonymous:1876:NAIk

[8230] Anonymous. New agricultural inventions. Scientific American, 34(10):153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
153b.pdf.

Anonymous:1876:NCMh

[8231] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(10):153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
153c.pdf.

Anonymous:1876:NWHg

[8232] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(10):153, March 4, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n10/pdf/scientificamerican03041876-153d.pdf.

Anonymous:1876:NHAg

[8233] Anonymous. New household articles. Scientific American, 34(10):153, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-153e.pdf.

Anonymous:1876:CRk

[8234] Anonymous. Communications received. Scientific American, 34 (10):155, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n10/pdf/scientificamerican03041876-
155.pdf.

Anonymous:1876:CPk

[8235] Anonymous. Canadian patents. Scientific American, 34(10):156, March 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n10/pdf/scientificamerican03041876-156.pdf.

Anonymous:1876:Cm

[8236] Anonymous. Correspondence. Scientific American, 34(11):164, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-164a.pdf.

Anonymous:1876:BNBg

[8237] Anonymous. Books: New books and publications. Scientific American, 34(11):168–169, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
168a.pdf.

Anonymous:1876:BPl

[8238] Anonymous. Business and personal. Scientific American, 34(11):170, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-170.pdf.

Anonymous:1876:NQl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-170a.pdf.

Anonymous:1876:IIl

[8240] Anonymous. Index of inventions. Scientific American, 34(11):171–172, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-171a.pdf.

Anonymous:1876:NSPa

[8241] Anonymous. A new source of power — absorbent strata. Scientific American, 34(11):159, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
159.pdf.

Anonymous:1876:CEb

[8242] Anonymous. The coming Emperor. Scientific American, 34(11):160, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-160.pdf.

Anonymous:1876:FWa

[8243] Anonymous. The father of waters. Scientific American, 34(11):160, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-160a.pdf.

Anonymous:1876:SASb

[8244] Anonymous. Some annals of a successful invention. Scientific American, 34(11):160–161, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
160b.pdf.

Anonymous:1876:CCa

[8245] Anonymous. Cylinder condensation. Scientific American, 34(11):161, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-161.pdf.

Anonymous:1876:PRV

[8246] Anonymous. Patent rights vs. state rights. Scientific American, 34(11):161, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
161a.pdf.

Anonymous:1876:SCR

[8247] Anonymous. The St. Charles Railroad Bridge, Mo. Scientific American, 34(11):162, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
162.pdf.

Anonymous:1876:TCS

[8248] Anonymous. The twin channel steamer Castalia. Scientific American, 34(11):163, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
163.pdf.

Anonymous:1876:IMCb

[8249] Anonymous. Improved method of covering steam boilers, etc. Scientific American, 34(11):163, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
163a.pdf.

Falke:1876:NPSb

[8250] William Falke. New phosphides of silver, and a method of estimating silver quantitativelt by means of phosphorus. Scientific American, 34(11):164, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
164.pdf.

Anonymous:1876:ANc

[8251] Anonymous. Astronomical notes. Scientific American, 34(11):164–165, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-164b.pdf.

Rose:1876:PMe

[8252] Joshua Rose. Practical mechanism. Scientific American, 34(11):165, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-165.pdf.

Anonymous:1876:PGC

[8253] Anonymous. Patent gear-cutting attachment for lathes. Scientific American, 34(11):166, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
166.pdf.

Anonymous:1876:SFM

[8254] Anonymous. Smith’s flour and meal chest. Scientific American, 34(11):166, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
166a.pdf.

Anonymous:1876:CIL

[8255] Anonymous. Cornell’s improved lifting jack. Scientific American, 34(11):166, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
166b.pdf.

Anonymous:1876:SIS

[8256] Anonymous. Steller’s improved sash holder. Scientific American, 34(11):166, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
166c.pdf.

Anonymous:1876:HHP

[8257] Anonymous. Haeckel on the human pedigree. Scientific American, 34(11):167–168, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
167.pdf.

Anonymous:1876:DCi

[8258] Anonymous. Decisions of the courts. Scientific American, 34(11):168, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-168.pdf.

Anonymous:1876:RAFl

[8259] Anonymous. Recent American and foreign patent. Scientific American, 34(11):169, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
169.pdf.

Anonymous:1876:NCMi

[8260] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(11):169, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
169a.pdf.

Anonymous:1876:NHAh

[8261] Anonymous. New household articles. Scientific American, 34(11):169, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n11/pdf/scientificamerican03111876-169b.pdf.

Anonymous:1876:NWHh

[8262] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(11):169, March 11, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n11/pdf/scientificamerican03111876-169c.pdf.

Anonymous:1876:NMEg

[8263] Anonymous. New mechanical and engineering inventions. Scientific American, 34(11):169, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
169d.pdf.

Anonymous:1876:CRl

[8264] Anonymous. Communications received. Scientific American, 34 (11):171, March 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n11/pdf/scientificamerican03111876-
171.pdf.

Anonymous:1876:Cn

[8265] Anonymous. Correspondence. Scientific American, 34(12):180, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-180.pdf.

Anonymous:1876:BNBh

[8266] Anonymous. Books: New books and publications. Scientific American, 34(12):185, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
185.pdf.

Anonymous:1876:BPm

[8267] Anonymous. Business and personal. Scientific American, 34(12):186, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-186.pdf.

Anonymous:1876:NQm

[8268] Anonymous. Notes and queries. Scientific American, 34(12):186–187, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-186a.pdf.

Anonymous:1876:IIm

[8269] Anonymous. Index of inventions. Scientific American, 34(12):187–188, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-187a.pdf.

Anonymous:1876:IMPa

[8270] Anonymous. Improved machine for polishing moldings and surfaces. Scientific American, 34(12):175, March 18, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n12/pdf/scientificamerican03181876-175.pdf.

Anonymous:1876:SASc

[8271] Anonymous. The Scientific American supplement. Scientific American, 34(12):176, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
176.pdf.

Anonymous:1876:MPa

[8272] Anonymous. Mind in plants. Scientific American, 34(12):176, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-176a.pdf.

Anonymous:1876:FET

[8273] Anonymous. Fire extinguisher trials. Scientific American, 34(12):176, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-176b.pdf.

Anonymous:1876:USM

[8274] Anonymous. The United States Metal-Testing Board. Scientific American, 34(12):176–177, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
176c.pdf.

VanderWeyde:1876:AIM

[8275] P. H. Vander Weyde. Artificial ice manufacture. Scientific American, 34(12):177, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
177.pdf.

Anonymous:1876:SPIh

[8276] Anonymous. Scientific and practical information. Scientific American, 34(12):177, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
177a.pdf.

Anonymous:1876:ICFa

[8277] Anonymous. Improved cloth-folding machine. Scientific American, 34(12):178, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
178.pdf.

Anonymous:1876:DSa

[8278] Anonymous. Distilling stearin. Scientific American, 34(12):179, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-179.pdf.

Anonymous:1876:ISRa

[8279] Anonymous. Improved scroll or reciprocating saw. Scientific American, 34(12):179–180, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
179a.pdf.

Anonymous:1876:IHM

[8280] Anonymous. Improved head muff. Scientific American, 34(12):180–181, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-180a.pdf.

Anonymous:1876:PLM

[8281] Anonymous. Production of light from mechanical force. Scientific American, 34(12):181, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
181.pdf.

Anonymous:1876:NMD

[8282] Anonymous. New mode of docking vessels. Scientific American, 34(12):182, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
182.pdf.

Anonymous:1876:IATa

[8283] Anonymous. Improved apparatus for testing: Illuminating oils. Scientific American, 34(12):182, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
182a.pdf.

Anonymous:1876:BRT

[8284] Anonymous. Bedell’s rapid transit wrench. Scientific American, 34 (12):182–183, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
182b.pdf.

Anonymous:1876:OPa

[8285] Anonymous. The owl parrot. Scientific American, 34(12):183, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-183.pdf.

Anonymous:1876:OPb

[8286] Anonymous. The opium poppy. Scientific American, 34(12):183, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n12/pdf/scientificamerican03181876-183a.pdf.

Anonymous:1876:FFL

[8287] Anonymous. Files, fuses, and low water alarms. Scientific American, 34(12):184, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
184.pdf.

Anonymous:1876:DCj

[8288] Anonymous. Decisions of the courts. Scientific American, 34 (12):184–185, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
184a.pdf.

Anonymous:1876:RAFm

[8289] Anonymous. Recent American and foreign patents. Scientific American, 34(12):185, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
185a.pdf.

Anonymous:1876:NAIl

[8290] Anonymous. New agricultural inventions. Scientific American, 34(12):185, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
185b.pdf.

Anonymous:1876:CRm

[8291] Anonymous. Communications received. Scientific American, 34 (12):187, March 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n12/pdf/scientificamerican03181876-
187.pdf.

Anonymous:1876:Co

[8292] Anonymous. Correspondence. Scientific American, 34(13):196, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-196.pdf.

Anonymous:1876:BPn

[8293] Anonymous. Business and personal. Scientific American, 34(13):202, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-202.pdf.

Anonymous:1876:NQn

[8294] Anonymous. Notes and queries. Scientific American, 34(13):202–203, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-202a.pdf.

Anonymous:1876:IIn

[8295] Anonymous. Index of inventions. Scientific American, 34(13):204, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-204.pdf.

Anonymous:1876:MDM

[8296] Anonymous. Multiple drilling machine. Scientific American, 34(13):191, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-191.pdf.

Anonymous:1876:OLL

[8297] Anonymous. The Oera Linda letters and figures. Scientific American, 34 (13):192, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic). URL http://en.wikipedia.org/wiki/Oera_
Linda;   http://www.nature.com/scientificamerican/journal/
v34/n13/pdf/scientificamerican03251876-192.pdf.

Anonymous:1876:WWN

[8298] Anonymous. Work and wages in New York City. Scientific American, 34(13):192–193, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
192a.pdf.

Anonymous:1876:DZZ

[8299] Anonymous. The draining of the Zuyder Zee. Scientific American, 34(13):193, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
193.pdf.

Anonymous:1876:AOP

[8300] Anonymous. Another obnoxious postal law. Scientific American, 34(13):193, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
193a.pdf.

Anonymous:1876:OSa

[8301] Anonymous. To our subscribers. Scientific American, 34(13):193, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-193b.pdf.

Anonymous:1876:DSM

[8302] Anonymous. Defeat of the sewing machine monopoly in Congress. Scientific American, 34(13):193, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
193c.pdf.

Anonymous:1876:AV

[8303] Anonymous. The aqueduct of La vanne. Scientific American, 34(13):194, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
194.pdf.

Anonymous:1876:IWL

[8304] Anonymous. Improved Watchmaker’s lathe. Scientific American, 34(13):194, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
194a.pdf.

Anonymous:1876:MEM

[8305] Anonymous. Magneto-electric machines. Scientific American, 34 (13):195–196, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
195.pdf.

Rose:1876:PMf

[8306] Joshua Rose. Practical mechanism. Scientific American, 34(13):197, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-197.pdf.

Anonymous:1876:IMe

[8307] Anonymous. Improved millstone. Scientific American, 34(13):198, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-198.pdf.

Anonymous:1876:IFV

[8308] Anonymous. Improved faucet and vent. Scientific American, 34 (13):198, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
198a.pdf.

Anonymous:1876:CNa

[8309] Anonymous. Cleopatra’s needles. Scientific American, 34(13):199–200, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-199.pdf.

Anonymous:1876:DCk

[8310] Anonymous. Decisions of the courts. Scientific American, 34(13):200, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-200.pdf.

Anonymous:1876:RAFn

[8311] Anonymous. Recent American and foreign patents. Scientific American, 34(13):200–201, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
200a.pdf.

Anonymous:1876:NAIm

[8312] Anonymous. New agricultural inventions. Scientific American, 34(13):201, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
201.pdf.

Anonymous:1876:NWHi

[8313] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(13):201, March 25, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n13/pdf/scientificamerican03251876-201a.pdf.

Anonymous:1876:NCMj

[8314] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(13):201, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
201b.pdf.

Anonymous:1876:NHAi

[8315] Anonymous. New household articles. Scientific American, 34(13):201, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n13/pdf/scientificamerican03251876-201c.pdf.

Anonymous:1876:CRn

[8316] Anonymous. Communications received. Scientific American, 34 (13):203–204, March 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n13/pdf/scientificamerican03251876-
203.pdf.

Anonymous:1876:Cq

[8317] Anonymous. Correspondence. Scientific American, 34(14):212, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-212.pdf.

Anonymous:1876:BPo

[8318] Anonymous. Business and personal. Scientific American, 34(14):218, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-218.pdf.

Anonymous:1876:NQo

[8319] Anonymous. Notes and queries. Scientific American, 34(14):218–219, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-218a.pdf.

Anonymous:1876:IIo

[8320] Anonymous. Index of inventions. Scientific American, 34(14):219–220, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-219a.pdf.

Anonymous:1876:RTN

[8321] Anonymous. Rapid transit in New York. Scientific American, 34(14):207, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
207.pdf.

Anonymous:1876:MCb

[8322] Anonymous. Milking cows. Scientific American, 34(14):208, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-208.pdf.

Anonymous:1876:BAT

[8323] Anonymous. A bit of ancient travel. Scientific American, 34(14):208, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-208a.pdf.

Anonymous:1876:NEC

[8324] Anonymous. The Naval Engineer Corps. Scientific American, 34 (14):208, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
208b.pdf.

Anonymous:1876:SCa

[8325] Anonymous. The science of currency. Scientific American, 34(14):208, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-208c.pdf.

Anonymous:1876:MAa

[8326] Anonymous. Milk and its adulteration. Scientific American, 34 (14):208–209, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
208d.pdf.

Anonymous:1876:FPS

[8327] Anonymous. To the friends and patrons of the Scientific American. Scientific American, 34(14):209, April 1, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n14/pdf/scientificamerican04011876-209.pdf.

Anonymous:1876:FWb

[8328] Anonymous. The father of waters. Scientific American, 34(14):209, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-209a.pdf.

Anonymous:1876:IUE

[8329] Anonymous. Improved universal emery grinder. Scientific American, 34(14):210, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
210.pdf.

Anonymous:1876:IATb

[8330] Anonymous. Improved animal trap. Scientific American, 34(14):210, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-210a.pdf.

Anonymous:1876:NUA

[8331] Anonymous. A new and useful architectural work. Scientific American, 34(14):211, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
211.pdf.

Anonymous:1876:MWB

[8332] Anonymous. Markrud’s wagon brake. Scientific American, 34(14):211, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-211a.pdf.

Anonymous:1876:SSL

[8333] Anonymous. The Stockwell self-lighting gas burner. Scientific American, 34(14):211, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
211b.pdf.

Ricco:1876:C

[8334] A. Ricco. The chromostroboscope. Scientific American, 34(14):212, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-212a.pdf.

Ricco:1876:SEE

[8335] A. Ricco. Singular electrical experiments. Scientific American, 34(14):212–213, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
212b.pdf.

Anonymous:1876:IIp

[8336] Anonymous. Improved incubator. Scientific American, 34(14):213, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n14/pdf/scientificamerican04011876-213.pdf.

Anonymous:1876:DCl

[8337] Anonymous. Decisions of the courts. Scientific American, 34 (14):216–217, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
216.pdf.

Anonymous:1876:RAFo

[8338] Anonymous. Recent American and foreign patents. Scientific American, 34(14):217, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
217.pdf.

Anonymous:1876:NCMk

[8339] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(14):217, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
217a.pdf.

Anonymous:1876:CRo

[8340] Anonymous. Communications received. Scientific American, 34 (14):219, April 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n14/pdf/scientificamerican04011876-
219.pdf.

Anonymous:1876:Cs

[8341] Anonymous. Correspondence. Scientific American, 34(15):228, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-228.pdf.

Anonymous:1876:BPp

[8342] Anonymous. Business and personal. Scientific American, 34(15):234, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-234.pdf.

Anonymous:1876:NQp

[8343] Anonymous. Notes and queries. Scientific American, 34(15):234–235, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-234a.pdf.

Anonymous:1876:IIq

[8344] Anonymous. Index of inventions. Scientific American, 34(15):235–236, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-235a.pdf.

Anonymous:1876:SRE

[8345] Anonymous. A semi-rotary engine. Scientific American, 34(15):223, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-223.pdf.

Anonymous:1876:ICN

[8346] Anonymous. Improved controlling nozzle. Scientific American, 34 (15):223, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
223a.pdf.

Anonymous:1876:CEE

[8347] Anonymous. The Centennial Exposition as an element in the development of our manufacturing interests. Scientific American, 34(15):224, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-224.pdf.

Anonymous:1876:BVC

[8348] Anonymous. Bad ventilation at the Capitol. Scientific American, 34(15):224, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
224a.pdf.

Anonymous:1876:AR

[8349] Anonymous. African rubber. Scientific American, 34(15):224, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-224b.pdf.

Anonymous:1876:CBB

[8350] Anonymous. A copper-bearing bird. Scientific American, 34 (15):224–225, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
224c.pdf.

Anonymous:1876:CMM

[8351] Anonymous. The cost of moths and mildew. Scientific American, 34(15):225, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
225.pdf.

Anonymous:1876:PBa

[8352] Anonymous. The power of boilers. Scientific American, 34(15):225, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-225a.pdf.

Anonymous:1876:PPB

[8353] Anonymous. The pay of the Patent Bureau. Scientific American, 34(15):225, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
225b.pdf.

Anonymous:1876:PHR

[8354] Anonymous. A portable hydraulic riveting machine. Scientific American, 34(15):226, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
226.pdf.

Anonymous:1876:MAb

[8355] Anonymous. The manufacture of anthracene. Scientific American, 34(15):226, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
226a.pdf.

Anonymous:1876:CTN

[8356] Anonymous. The cattle of Thelemark, Norway. Scientific American, 34(15):227, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
227.pdf.

Anonymous:1876:SCCa

[8357] Anonymous. Self-closing closet lid and anti-nuisance closet. Scientific American, 34(15):227, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
227a.pdf.

Anonymous:1876:NGP

[8358] Anonymous. A new garden pump. Scientific American, 34(15):227, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-227b.pdf.

VanderWeyde:1876:AI

[8359] P. H. Vander Weyde. Artificial ice. Scientific American, 34(15):228–229, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-228a.pdf.

Rose:1876:PMg

[8360] Joshua Rose. Practical mechanism. Scientific American, 34(15):229, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-229.pdf.

Anonymous:1876:PCEa

[8361] Anonymous. Progress of the Centennial Exhibition. Scientific American, 34(15):229, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
229a.pdf.

Anonymous:1876:IWP

[8362] Anonymous. Improved wall paper trimmer. Scientific American, 34(15):230, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
230.pdf.

Anonymous:1876:WLR

[8363] Anonymous. Weston’s lightning rod discharging point. Scientific American, 34(15):230, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
230a.pdf.

Anonymous:1876:RMM

[8364] Anonymous. Riggs’ method of mending cracked bells. Scientific American, 34(15):230, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
230b.pdf.

Anonymous:1876:NAG

[8365] Anonymous. New automatic grease cup. Scientific American, 34 (15):230, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
230c.pdf.

Anonymous:1876:AFb

[8366] Anonymous. Animal ferocity. Scientific American, 34(15):231, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-231.pdf.

Anonymous:1876:ANd

[8367] Anonymous. Astronomical notes. Scientific American, 34(15):232, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-232.pdf.

Anonymous:1876:FAF

[8368] Anonymous. Fire apparatus fog alarms and cork machinery. Scientific American, 34(15):232, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
232a.pdf.

Anonymous:1876:DCm

[8369] Anonymous. Decisions of the courts. Scientific American, 34 (15):232–233, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
232b.pdf.

Anonymous:1876:RAFp

[8370] Anonymous. Recent American and foreign patents. Scientific American, 34(15):233, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
233.pdf.

Anonymous:1876:NWHj

[8371] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(15):233, April 8, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n15/pdf/scientificamerican04081876-233a.pdf.

Anonymous:1876:NCMl

[8372] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(15):233, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
233b.pdf.

Anonymous:1876:NAIn

[8373] Anonymous. New agricultural inventions. Scientific American, 34 (15):233, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
233c.pdf.

Anonymous:1876:NHAj

[8374] Anonymous. New household articles. Scientific American, 34(15):233, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n15/pdf/scientificamerican04081876-233d.pdf.

Anonymous:1876:CRp

[8375] Anonymous. Communications received. Scientific American, 34 (15):235, April 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n15/pdf/scientificamerican04081876-
235.pdf.

Anonymous:1876:Ct

[8376] Anonymous. Correspondence. Scientific American, 34(16):244, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-244c.pdf.

Anonymous:1876:BNBi

[8377] Anonymous. Books: New books and publications. Scientific American, 34(16):248, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
248b.pdf.

Anonymous:1876:BPq

[8378] Anonymous. Business and personal. Scientific American, 34(16):250, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-250.pdf.

Anonymous:1876:NQq

[8379] Anonymous. Notes and queries. Scientific American, 34(16):250–251, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-250a.pdf.

Anonymous:1876:IIr

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-251a.pdf.

Anonymous:1876:LME

[8381] Anonymous. A locomotive magneto-electric induction engine. Scientific American, 34(16):239, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
239.pdf.

Anonymous:1876:SASd

[8382] Anonymous. The Scientific American supplement. Scientific American, 34(16):240, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
240.pdf.

Anonymous:1876:FAA

[8383] Anonymous. Fluosilicic acid in the arts. Scientific American, 34(16):240, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
240a.pdf.

Anonymous:1876:BFP

[8384] Anonymous. The baobab as a fiber plant. Scientific American, 34(16):240, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
240b.pdf.

Anonymous:1876:SMV

[8385] Anonymous. The secrets of making Vienna bread. Scientific American, 34(16):240–241, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
240c.pdf.

Anonymous:1876:UM

[8386] Anonymous. The utilizations of mica. Scientific American, 34(16):241, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-241.pdf.

Anonymous:1876:IEP

[8387] Anonymous. The international exhibition of 1876 — progress of the machinery department. Scientific American, 34(16):241, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-241a.pdf.

Anonymous:1876:IFWa

[8388] Anonymous. Improved feed water heater. Scientific American, 34(16):242, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
242.pdf.

Anonymous:1876:HPP

[8389] Anonymous. Horse power pump. Scientific American, 34(16):243, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-243.pdf.

Anonymous:1876:IKF

[8390] Anonymous. Improved kitchen furniture. Scientific American, 34 (16):243, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
243a.pdf.

Anonymous:1876:MSa

[8391] Anonymous. The magnetic spectrum. Scientific American, 34(16):244, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-244.pdf.

Anonymous:1876:NLE

[8392] Anonymous. A new lecture experiment. Scientific American, 34 (16):244, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
244a.pdf.

Deane:1876:SFG

[8393] John. C. Deane. Simple form of the gyroscope. Scientific American, 34(16):244, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
244b.pdf.

Anonymous:1876:HEW

[8394] Anonymous. How the Earth is weighed and measured. Scientific American, 34(16):245, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
245.pdf.

Anonymous:1876:TBa

[8395] Anonymous. Testing boilers. Scientific American, 34(16):246, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-246.pdf.

Anonymous:1876:AIDa

[8396] Anonymous. Atkinson’s improved desk. Scientific American, 34 (16):246, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
246a.pdf.

Anonymous:1876:IMBa

[8397] Anonymous. Improved mucilage bottle. Scientific American, 34 (16):246, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
246b.pdf.

Anonymous:1876:NFBa

[8398] Anonymous. New fusible boiler plug. Scientific American, 34(16):246, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-246c.pdf.

Anonymous:1876:OG

[8399] Anonymous. An ornamental gourd. Scientific American, 34(16):247, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-247.pdf.

Anonymous:1876:AMa

[8400] Anonymous. Acalypha Marginata. Scientific American, 34(16):247–248, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-247a.pdf.

Anonymous:1876:SPIi

[8401] Anonymous. Scientific and practical information. Scientific American, 34(16):248, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
248.pdf.

Anonymous:1876:NSTb

[8402] Anonymous. The National steel tube cleaner. Scientific American, 34(16):248, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
248a.pdf.

Anonymous:1876:RAFq

[8403] Anonymous. Recent American and foreign patents. Scientific American, 34(16):248–249, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
248c.pdf.

Anonymous:1876:NMEh

[8404] Anonymous. New mechanical and engineering inventions. Scientific American, 34(16):249, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
249.pdf.

Anonymous:1876:NWHk

[8405] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(16):249, April 15, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n16/pdf/scientificamerican04151876-249a.pdf.

Anonymous:1876:NAIo

[8406] Anonymous. New agricultural inventions. Scientific American, 34(16):249, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
249b.pdf.

Anonymous:1876:NHAk

[8407] Anonymous. New household articles. Scientific American, 34(16):249, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n16/pdf/scientificamerican04151876-249c.pdf.

Anonymous:1876:CRq

[8408] Anonymous. Communications received. Scientific American, 34 (16):251, April 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n16/pdf/scientificamerican04151876-
251.pdf.

Anonymous:1876:Cu

[8409] Anonymous. Correspondence. Scientific American, 34(17):260, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-260.pdf.

Anonymous:1876:BPr

[8410] Anonymous. Business and personal. Scientific American, 34(17):266, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-266.pdf.

Anonymous:1876:NQr

[8411] Anonymous. Notes and queries. Scientific American, 34(17):266–267, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-266a.pdf.

Anonymous:1876:IIs

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-267a.pdf.

Anonymous:1876:LMS

[8413] Anonymous. The little monitor sewing machine. Scientific American, 34(17):255, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
255.pdf.

Anonymous:1876:SASe

[8414] Anonymous. The Scientific American supplement. Scientific American, 34(17):256, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
256.pdf.

Anonymous:1876:WWS

[8415] Anonymous. A woman’s work for science. Scientific American, 34(17):256, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
256a.pdf.

Anonymous:1876:POA

[8416] Anonymous. The Patent Office appropriation. Scientific American, 34(17):256, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
256b.pdf.

Anonymous:1876:STLa

[8417] Anonymous. Street telegraph lines. Scientific American, 34(17): 256–257, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
256c.pdf.

Anonymous:1876:FPC

[8418] Anonymous. Free passes at the Centennial. Scientific American, 34(17):257, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
257.pdf.

Anonymous:1876:PUS

[8419] Anonymous. The practical uses of the spectroscope. Scientific American, 34(17):257, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
257a.pdf.

Anonymous:1876:IUA

[8420] Anonymous. The industrial uses of asbestos. Scientific American, 34(17):258, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
258.pdf.

Anonymous:1876:ILCa

[8421] Anonymous. Improved leather and cloth enameling machine. Scientific American, 34(17):259, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
259.pdf.

Anonymous:1876:IPCa

[8422] Anonymous. Improved process of capsuling medicaments. Scientific American, 34(17):259, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
259a.pdf.

Anonymous:1876:IEMa

[8423] Anonymous. Improved electro-magnetic dental engine. Scientific American, 34(17):259, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
259b.pdf.

Anonymous:1876:NPP

[8424] Anonymous. New process of preserving wood. Scientific American, 34(17):259–260, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
259c.pdf.

Bartlett:1876:EVB

[8425] F. L. Bartlett. Estimating the value of bone black. Scientific American, 34(17):260, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
260a.pdf.

Rose:1876:PMh

[8426] Joshua Rose. Practical mechanism. Scientific American, 34(17):261, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-261.pdf.

Anonymous:1876:IGC

[8427] Anonymous. Improved grain cleaner. Scientific American, 34(17):262, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-262.pdf.

Anonymous:1876:MIH

[8428] Anonymous. Miller’s improved hose carriage. Scientific American, 34(17):262, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
262a.pdf.

Anonymous:1876:IMO

[8429] Anonymous. Improved method of obtaining water. Scientific American, 34(17):262, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
262b.pdf.

Anonymous:1876:RFF

[8430] Anonymous. The red fire fish. Scientific American, 34(17):263, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-263.pdf.

Anonymous:1876:FG

[8431] Anonymous. The filamentous: Gurnard. Scientific American, 34 (17):263–264, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
263a.pdf.

Anonymous:1876:HHC

[8432] Anonymous. The horticultural Hall at the Centennial Exhibition. Scientific American, 34(17):264, April 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n17/pdf/scientificamerican04221876-264.pdf.

Anonymous:1876:RAFr

[8433] Anonymous. Recent American and foreign patents. Scientific American, 34(17):264–265, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
264a.pdf.

Anonymous:1876:NWHl

[8434] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(17):265, April 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n17/pdf/scientificamerican04221876-265.pdf.

Anonymous:1876:NAIp

[8435] Anonymous. New agricultural inventions. Scientific American, 34(17):265, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
265a.pdf.

Anonymous:1876:NHAl

[8436] Anonymous. New household articles. Scientific American, 34(17):265, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n17/pdf/scientificamerican04221876-265b.pdf.

Anonymous:1876:NMEi

[8437] Anonymous. New mechanical and engineering inventions. Scientific American, 34(17):265, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
265c.pdf.

Anonymous:1876:CRr

[8438] Anonymous. Communications received. Scientific American, 34 (17):267, April 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n17/pdf/scientificamerican04221876-
267.pdf.

Anonymous:1876:Cv

[8439] Anonymous. Correspondence. Scientific American, 34(18):276, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-276.pdf.

Anonymous:1876:BNBj

[8440] Anonymous. Books: New books and publications. Scientific American, 34(18):280, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
280a.pdf.

Anonymous:1876:BPs

[8441] Anonymous. Business and personal. Scientific American, 34(18):282, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-282.pdf.

Anonymous:1876:NQs

[8442] Anonymous. Notes and queries. Scientific American, 34(18):282–283, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-282a.pdf.

Anonymous:1876:IIt

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-283a.pdf.

Anonymous:1876:NBPa

[8444] Anonymous. A new bridge in Philadelphia. Scientific American, 34(18):271, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
271.pdf.

Anonymous:1876:LB

[8445] Anonymous. Laundry blue. Scientific American, 34(18):272, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-272.pdf.

Anonymous:1876:PEP

[8446] Anonymous. The practical effects of physical strain. Scientific American, 34(18):272, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
272a.pdf.

Anonymous:1876:VPa

[8447] Anonymous. The validity of patents. Scientific American, 34(18):272, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-272b.pdf.

Anonymous:1876:WTS

[8448] Anonymous. What they say about us. Scientific American, 34 (18):272–273, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
272c.pdf.

Anonymous:1876:DMM

[8449] Anonymous. Death of the millionaire merchant. Scientific American, 34(18):273, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
273.pdf.

Anonymous:1876:APa

[8450] Anonymous. Animal parasites. Scientific American, 34(18):273, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-273a.pdf.

Anonymous:1876:SPIj

[8451] Anonymous. Scientific and practical information. Scientific American, 34(18):273, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
273b.pdf.

Anonymous:1876:IWS

[8452] Anonymous. Improved window screen. Scientific American, 34(18):274, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-274.pdf.

Anonymous:1876:IOC

[8453] Anonymous. Improved oil can nozzle. Scientific American, 34(18):274, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-274a.pdf.

Anonymous:1876:AEC

[8454] Anonymous. Atchison’s engraving and chasing machine. Scientific American, 34(18):274, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
274b.pdf.

Anonymous:1876:VIL

[8455] Anonymous. Vibert’s improved lifting jack. Scientific American, 34(18):274, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
274c.pdf.

Anonymous:1876:IAPa

[8456] Anonymous. Improved adjustable pipe tongs. Scientific American, 34(18):274, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
274d.pdf.

Anonymous:1876:BSR

[8457] Anonymous. Broken stone road-making — Blake’s crusher. Scientific American, 34(18):275, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
275.pdf.

Anonymous:1876:ELR

[8458] Anonymous. Engines and lightning rods. Scientific American, 34(18):277, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
277.pdf.

Anonymous:1876:IMPb

[8459] Anonymous. Improved method of propelling canal boats. Scientific American, 34(18):278, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
278.pdf.

Anonymous:1876:IFGb

[8460] Anonymous. Improved flood gate. Scientific American, 34(18):278, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-278a.pdf.

Anonymous:1876:CSJ

[8461] Anonymous. A Chinese scientific journal. Scientific American, 34(18):279, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
279.pdf.

Anonymous:1876:DCn

[8462] Anonymous. Decisions of the courts. Scientific American, 34(18):280, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n18/pdf/scientificamerican04291876-280.pdf.

Anonymous:1876:RAFs

[8463] Anonymous. Recent American and foreign patents. Scientific American, 34(18):280–281, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
280b.pdf.

Anonymous:1876:NAIq

[8464] Anonymous. New agricultural inventions. Scientific American, 34(18):281, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
281.pdf.

Anonymous:1876:NWHm

[8465] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(18):281, April 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n18/pdf/scientificamerican04291876-281a.pdf.

Anonymous:1876:NMEj

[8466] Anonymous. New mechanical and engineering inventions. Scientific American, 34(18):281, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
281b.pdf.

Anonymous:1876:CRs

[8467] Anonymous. Communications received. Scientific American, 34 (18):283, April 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n18/pdf/scientificamerican04291876-
283.pdf.

Anonymous:1876:BPt

[8468] Anonymous. Business and personal. Scientific American, 34(19):298, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-298.pdf.

Anonymous:1876:NQt

[8469] Anonymous. Notes and queries. Scientific American, 34(19):298–299, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-298a.pdf.

Anonymous:1876:IIu

[8470] Anonymous. Index of inventions. Scientific American, 34(19):300, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-300.pdf.

Anonymous:1876:IASb

[8471] Anonymous. Improved automatic small steam engines. Scientific American, 34(19):287, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
287.pdf.

Anonymous:1876:SASf

[8472] Anonymous. The Scientific American supplement. Scientific American, 34(19):288, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
288.pdf.

Anonymous:1876:DCMa

[8473] Anonymous. Darwin on consanguineous marriages. Scientific American, 34(19):288, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
288a.pdf.

Anonymous:1876:EEb

[8474] Anonymous. An energetic Emperor. Scientific American, 34 (19):288–289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
288b.pdf.

Anonymous:1876:SI

[8475] Anonymous. Self-improvement. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-289.pdf.

Anonymous:1876:GSB

[8476] Anonymous. The great suspension bridge between New York and Brooklyn. Scientific American, 34(19):289, May 6, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n19/pdf/scientificamerican05061876-289a.pdf.

Anonymous:1876:ORR

[8477] Anonymous. Opening of a remarkable railway in London. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
289b.pdf.

Anonymous:1876:AEM

[8478] Anonymous. Artificial eyes made sensitive to light. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
289c.pdf.

Anonymous:1876:WPW

[8479] Anonymous. The Woodbury planer war. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
289d.pdf.

Anonymous:1876:LAB

[8480] Anonymous. The Loughbridge air brake. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
289e.pdf.

Anonymous:1876:ROS

[8481] Anonymous. Remarkable ocean steaming. Scientific American, 34(19):289, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
289f.pdf.

Anonymous:1876:ITP

[8482] Anonymous. Improved toy pistol — mechanical marble playing. Scientific American, 34(19):291, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
291.pdf.

Anonymous:1876:HDP

[8483] Anonymous. Heer’s driving pulley. Scientific American, 34(19):291, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-291a.pdf.

Anonymous:1876:HDF

[8484] Anonymous. Haptonstall’s diamond file. Scientific American, 34 (19):291, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
291b.pdf.

Anonymous:1876:NLP

[8485] Anonymous. Nash’s life-preserving stool. Scientific American, 34 (19):291, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
291c.pdf.

Anonymous:1876:BIO

[8486] Anonymous. Backus’ improved oil can. Scientific American, 34(19):291, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-291d.pdf.

Anonymous:1876:CIT

[8487] Anonymous. The coral islands. — their nature, growth, and geographical distribution. Scientific American, 34(19):292, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-292.pdf.

Anonymous:1876:NAS

[8488] Anonymous. The National Academy of Sciences. Scientific American, 34(19):292, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
292a.pdf.

Anonymous:1876:ANe

[8489] Anonymous. Astronomical notes. Scientific American, 34(19):292, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-292b.pdf.

Rose:1876:PMi

[8490] Joshua Rose. Practical mechanism. Scientific American, 34(19):293, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-293.pdf.

Anonymous:1876:IKBa

[8491] Anonymous. Improved keg and barrel machinery. Scientific American, 34(19):294–296, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
294.pdf.

Anonymous:1876:DCo

[8492] Anonymous. Decisions of the courts. Scientific American, 34(19):296, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-296.pdf.

Anonymous:1876:RAFt

[8493] Anonymous. Recent American and foreign patents. Scientific American, 34(19):297, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
297.pdf.

Anonymous:1876:NMEk 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
297a.pdf.

Anonymous:1876:NHAm

[8495] Anonymous. New household articles. Scientific American, 34(19):297, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n19/pdf/scientificamerican05061876-297b.pdf.

Anonymous:1876:NAIr

[8496] Anonymous. New agricultural inventions. Scientific American, 34(19):297, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
297c.pdf.

Anonymous:1876:NCMm

[8497] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(19):297, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
297d.pdf.

Anonymous:1876:CRt

[8498] Anonymous. Communications received. Scientific American, 34 (19):299, May 6, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n19/pdf/scientificamerican05061876-
299.pdf.

Anonymous:1876:Cw

[8499] Anonymous. Correspondence. Scientific American, 34(20):308, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-308.pdf.

Anonymous:1876:BPu

[8500] Anonymous. Business and personal. Scientific American, 34(20):315, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-315.pdf.

Anonymous:1876:NQu

[8501] Anonymous. Notes and queries. Scientific American, 34(20):315, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-315a.pdf.

Anonymous:1876:IIv

[8502] Anonymous. Index of inventions. Scientific American, 34(20):315–316, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-315c.pdf.

Anonymous:1876:BRM

[8503] Anonymous. The Brayton ready motor or hydrocarbon engine. Scientific American, 34(20):303, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
303.pdf.

Anonymous:1876:CNb

[8504] Anonymous. A Centennial number. Scientific American, 34(20):304, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-304.pdf.

Anonymous:1876:SGM

[8505] Anonymous. Silver and gold money. Scientific American, 34(20):304, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-304a.pdf.

Anonymous:1876:GC

[8506] Anonymous. La Grande Chartreuse. Scientific American, 34(20):304, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-304b.pdf.

Anonymous:1876:WPE

[8507] Anonymous. Worry and its physical effects. Scientific American, 34(20):304–305, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
304c.pdf.

Anonymous:1876:ASMa

[8508] Anonymous. The apparent size of the Moon at the horizon. Scientific American, 34(20):305, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
305.pdf.

Anonymous:1876:PCa

[8509] Anonymous. Progress of the Centennial. Scientific American, 34 (20):305, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
305a.pdf.

Anonymous:1876:IATc

[8510] Anonymous. Improved automatic tank filler. Scientific American, 34(20):306, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
306.pdf.

Anonymous:1876:SFV

[8511] Anonymous. A simple flower vase. Scientific American, 34(20):306, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-306a.pdf.

Anonymous:1876:IMSa

[8512] Anonymous. Improved machine for straightening metal bars. Scientific American, 34(20):306, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
306b.pdf.

Anonymous:1876:IKBb

[8513] Anonymous. Improved keg and barrel machinery. Scientific American, 34(20):307, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
307.pdf.

Anonymous:1876:SPIk

[8514] Anonymous. Scientific and practical information. Scientific American, 34(20):307–308, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
307a.pdf.

Anonymous:1876:M

[8515] Anonymous. The Moon. Scientific American, 34(20):308–309, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-308a.pdf.

Anonymous:1876:IRS

[8516] Anonymous. Improved railway switch signal. Scientific American, 34(20):310, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
310.pdf.

Anonymous:1876:SIC

[8517] Anonymous. Shephard’s improved churn. Scientific American, 34 (20):310, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
310a.pdf.

Anonymous:1876:BPCa

[8518] Anonymous. The Benton patent copper float. Scientific American, 34(20):310, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
310b.pdf.

Anonymous:1876:CBS

[8519] Anonymous. The cone-berried solanum. Scientific American, 34 (20):311, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
311.pdf.

Anonymous:1876:TBY

[8520] Anonymous. Two beautiful yuccas. Scientific American, 34(20):311, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-311a.pdf.

Anonymous:1876:NFFa

[8521] Anonymous. A new form of fernery. Scientific American, 34 (20):311–312, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
311b.pdf.

Anonymous:1876:AMb

[8522] Anonymous. Agricultural machinery. Scientific American, 34 (20):312–313, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
312.pdf.

Anonymous:1876:RAFu

[8523] Anonymous. Recent American and foreign patents. Scientific American, 34(20):313, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
313.pdf.

Anonymous:1876:NHAn

[8524] Anonymous. New household articles. Scientific American, 34(20):313, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-313a.pdf.

Anonymous:1876:NTMe

[8525] Anonymous. New textile machinery. Scientific American, 34(20):313, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n20/pdf/scientificamerican05131876-313b.pdf.

Anonymous:1876:NWHn

[8526] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(20):314, May 13, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n20/pdf/scientificamerican05131876-314.pdf.

Anonymous:1876:NMEl

[8527] Anonymous. New mechanical and engineering inventions. Scientific American, 34(20):314, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
314a.pdf.

Anonymous:1876:NAIs

[8528] Anonymous. New agricultural inventions. Scientific American, 34 (20):314, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
314b.pdf.

Anonymous:1876:CRu

[8529] Anonymous. Communications received. Scientific American, 34 (20):315, May 13, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n20/pdf/scientificamerican05131876-
315b.pdf.

Anonymous:1876:Cx

[8530] Anonymous. Correspondence. Scientific American, 34(21):324–325, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-324.pdf.

Anonymous:1876:BPv

[8531] Anonymous. Business and personal. Scientific American, 34(21):331, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-331.pdf.

Anonymous:1876:NQv

[8532] Anonymous. Notes and queries. Scientific American, 34(21):331–332, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-331a.pdf.

Anonymous:1876:IIw

[8533] Anonymous. Index of inventions. Scientific American, 34(21):332, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-332a.pdf.

Anonymous:1876:PNPa

[8534] Anonymous. Patent notes: Patents. Scientific American, 34(21):333, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-333.pdf.

Anonymous:1876:CIEa

[8535] Anonymous. The Centennial International Exhibition. Scientific American, 34(21):319, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
319.pdf.

Anonymous:1876:SASg

[8536] Anonymous. The Scientific American supplement. Scientific American, 34(21):320, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
320.pdf.

Anonymous:1876:APb

[8537] Anonymous. American progress — I — from 1776 to 1820. Scientific American, 34(21):320–321, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
320a.pdf.

Anonymous:1876:GSc

[8538] Anonymous. The Grasshopper scourge of 1876. Scientific American, 34(21):321, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
321.pdf.

Anonymous:1876:FI

[8539] Anonymous. Fire insurance. Scientific American, 34(21):321, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-321a.pdf.

Anonymous:1876:PCb

[8540] Anonymous. Progress at the Centennial. Scientific American, 34 (21):321, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
321b.pdf.

Anonymous:1876:CES

[8541] Anonymous. The Centennial Exhibition — the state buildings. Scientific American, 34(21):322–324, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
322.pdf.

Rose:1876:PMj

[8542] Joshua Rose. Practical mechanism. Scientific American, 34(21):325, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-325.pdf.

Anonymous:1876:CBI

[8543] Anonymous. The Centennial buildings — the industrial pavilions. Scientific American, 34(21):326–328, May 20, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n21/pdf/scientificamerican05201876-326.pdf.

Anonymous:1876:CCb

[8544] Anonymous. The Di Cesnola Collection. Scientific American, 34 (21):328–329, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
328.pdf.

Anonymous:1876:SPIl

[8545] Anonymous. Scientific and practical information. Scientific American, 34(21):329, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
329.pdf.

Anonymous:1876:DCp

[8546] Anonymous. Decisions of the courts. Scientific American, 34(21):329, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-329a.pdf.

Anonymous:1876:RAFv

[8547] Anonymous. Recent American and foreign patents. Scientific American, 34(21):330, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
330.pdf.

Anonymous:1876:NCMn

[8548] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(21):330, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
330a.pdf.

Anonymous:1876:NHAo

[8549] Anonymous. New household articles. Scientific American, 34(21):330, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-330b.pdf.

Anonymous:1876:NWHo

[8550] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(21):330, May 20, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n21/pdf/scientificamerican05201876-330c.pdf.

Anonymous:1876:NTMf

[8551] Anonymous. New textile machinery. Scientific American, 34(21):330, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n21/pdf/scientificamerican05201876-330d.pdf.

Anonymous:1876:NMEm 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
330e.pdf.

Anonymous:1876:CRv

[8553] Anonymous. Communications received. Scientific American, 34 (21):332, May 20, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n21/pdf/scientificamerican05201876-
332.pdf.

Anonymous:1876:Cy

[8554] Anonymous. Correspondence. Scientific American, 34(22):340, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-340.pdf.

Anonymous:1876:BNBk

[8555] Anonymous. Books: New books and publications. Scientific American, 34(22):344, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
344a.pdf.

Anonymous:1876:BPw

[8556] Anonymous. Business and personal. Scientific American, 34(22):346, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-346.pdf.

Anonymous:1876:NQw

[8557] Anonymous. Notes and queries. Scientific American, 34(22):346–348, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-346a.pdf.

Anonymous:1876:IIx

[8558] Anonymous. Index of inventions. Scientific American, 34(22):348, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-348a.pdf.

Anonymous:1876:PNPb

[8559] Anonymous. Patent notes: Patents. Scientific American, 34(22):349, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-349.pdf.

Anonymous:1876:AGa

[8560] Anonymous. The Allen governor. Scientific American, 34(22):335, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-335.pdf.

Anonymous:1876:APIa

[8561] Anonymous. American progress — II — from 1820 to 1840. Scientific American, 34(22):336–337, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
336.pdf.

Anonymous:1876:RS

[8562] Anonymous. Reclaiming the steppes. Scientific American, 34(22):337, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-337.pdf.

Anonymous:1876:FAC

[8563] Anonymous. French artisans at the Centennial. Scientific American, 34(22):337, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
337a.pdf.

Anonymous:1876:OC

[8564] Anonymous. The opening of the Centennial. Scientific American, 34(22):337, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
337b.pdf.

Anonymous:1876:CIEb

[8565] Anonymous. The Centennial International Exhibition. Scientific American, 34(22):338–339, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
338.pdf.

Anonymous:1876:NYAb

[8566] Anonymous. New York Academy of Sciences. Scientific American, 34(22):339, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
339.pdf.

Anonymous:1876:SEG

[8567] Anonymous. Setting engine guide bars. Scientific American, 34(22):340, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-340a.pdf.

Anonymous:1876:N

[8568] Anonymous. Nitroglycerin. Scientific American, 34(22):341, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-341.pdf.

Anonymous:1876:IMSb

[8569] Anonymous. Improved middlings separator. Scientific American, 34(22):342, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
342.pdf.

Anonymous:1876:ISRb

[8570] Anonymous. Improved steam riveter. Scientific American, 34(22):342, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-342a.pdf.

Anonymous:1876:ISSb

[8571] Anonymous. Improved spoke-setting machine. Scientific American, 34(22):343, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
343.pdf.

Anonymous:1876:IRG

[8572] Anonymous. Improved rowing gear. Scientific American, 34(22):343, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-343a.pdf.

Anonymous:1876:WIW

[8573] Anonymous. White’s improved washtub stand. Scientific American, 34(22):343–344, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
343b.pdf.

Anonymous:1876:DCq

[8574] Anonymous. Decisions of the courts. Scientific American, 34(22):344, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-344.pdf.

Anonymous:1876:RAFw

[8575] Anonymous. Recent American and foreign patents. Scientific American, 34(22):345, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
345.pdf.

Anonymous:1876:NTMg

[8576] Anonymous. New textile machinery. Scientific American, 34(22):345, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-345a.pdf.

Anonymous:1876:NCMo

[8577] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(22):345, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
345b.pdf.

Anonymous:1876:NWHp

[8578] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(22):345, May 27, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n22/pdf/scientificamerican05271876-345c.pdf.

Anonymous:1876:NHAp

[8579] Anonymous. New household articles. Scientific American, 34(22):345, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n22/pdf/scientificamerican05271876-345d.pdf.

Anonymous:1876:NAIt

[8580] Anonymous. New agricultural inventions. Scientific American, 34 (22):345, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
345e.pdf.

Anonymous:1876:CRw

[8581] Anonymous. Communications received. Scientific American, 34 (22):348, May 27, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n22/pdf/scientificamerican05271876-
348.pdf.

Anonymous:1876:BNBl

[8582] Anonymous. Books: New books and publications. Scientific American, 34(23):362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
362.pdf.

Anonymous:1876:BPx

[8583] Anonymous. Business and personal. Scientific American, 34(23):363, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-363.pdf.

Anonymous:1876:NQx

[8584] Anonymous. Notes and queries. Scientific American, 34(23):363–364, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-363a.pdf.

Anonymous:1876:IIy

[8585] Anonymous. Index of inventions. Scientific American, 34(23):364, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-364a.pdf.

Anonymous:1876:PNPc

[8586] Anonymous. Patent notes: Patents. Scientific American, 34(23):366, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-366.pdf.

Anonymous:1876:CEG

[8587] Anonymous. The Centennial Exhibition. — the great Corliss engines. Scientific American, 34(23):351, June 3, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n23/pdf/scientificamerican06031876-351.pdf.

Anonymous:1876:APIb

[8588] Anonymous. American progress — III. — from 1840 to the present time. Scientific American, 34(23):352–353, June 3, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n23/pdf/scientificamerican06031876-352.pdf.

Anonymous:1876:PCC

[8589] Anonymous. The present condition of the Centennial. Scientific American, 34(23):353, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
353.pdf.

Anonymous:1876:SCB

[8590] Anonymous. The Spanish Centennial building. Scientific American, 34(23):354, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
354.pdf.

Anonymous:1876:CBN

[8591] Anonymous. The Centennial buildings. — the New Jersey buildings and the Women’s pavilion. Scientific American, 34(23):354, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-354a.pdf.

Anonymous:1876:MTT

[8592] Anonymous. Muntz’s tannin-testing apparatus. Scientific American, 34(23):355, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
355.pdf.

Anonymous:1876:SHSb

[8593] Anonymous. Stephenson’s horseshoe and swage. Scientific American, 34(23):355, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
355a.pdf.

Anonymous:1876:CDCa

[8594] Anonymous. Collins’ dough cutter. Scientific American, 34(23):355, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-355b.pdf.

Anonymous:1876:CDCb

[8595] Anonymous. Collins’ door check. Scientific American, 34(23):355, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-355c.pdf.

Anonymous:1876:CSW

[8596] Anonymous. Calculating the speed of wheels, pulleys, etc. Scientific American, 34(23):356, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
356.pdf.

Anonymous:1876:SPIm

[8597] Anonymous. Scientific and practical information. Scientific American, 34(23):356, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
356a.pdf.

Rose:1876:PMk

[8598] Joshua Rose. Practical mechanism. Scientific American, 34(23):357, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-357.pdf.

Anonymous:1876:IWW

[8599] Anonymous. Improved water wheel. Scientific American, 34(23):358, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-358.pdf.

Anonymous:1876:ICSa

[8600] Anonymous. Improved cotton stalk puller. Scientific American, 34(23):358, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
358a.pdf.

Anonymous:1876:FFa

[8601] Anonymous. A fish that fishes. Scientific American, 34(23):359, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-359.pdf.

Anonymous:1876:WUS

[8602] Anonymous. The work of the United States Fish Commission. Scientific American, 34(23):359, June 3, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n23/pdf/scientificamerican06031876-359a.pdf.

Anonymous:1876:NCa

[8603] Anonymous. Notes on chemistry. Scientific American, 34(23):360, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-360.pdf.

Anonymous:1876:MV

[8604] Anonymous. Musical vibrations. Scientific American, 34(23):360, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-360a.pdf.

Anonymous:1876:GCR

[8605] Anonymous. Geometrical chemistry. — remarkable discovery by Professor Henry Wurtz. Scientific American, 34(23):361–362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-361.pdf.

Anonymous:1876:RAFx

[8606] Anonymous. Recent American and foreign patents. Scientific American, 34(23):362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
362a.pdf.

Anonymous:1876:NCMp

[8607] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(23):362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
362b.pdf.

Anonymous:1876:NHAq

[8608] Anonymous. New household articles. Scientific American, 34(23):362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n23/pdf/scientificamerican06031876-362c.pdf.

Anonymous:1876:NWHq

[8609] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(23):362, June 3, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n23/pdf/scientificamerican06031876-362d.pdf.

Anonymous:1876:NAIu

[8610] Anonymous. New agricultural inventions. Scientific American, 34(23):362, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
362e.pdf.

Anonymous:1876:CRx

[8611] Anonymous. Communications received. Scientific American, 34 (23):364, June 3, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n23/pdf/scientificamerican06031876-
364.pdf.

Anonymous:1876:Cz

[8612] Anonymous. Correspondence. Scientific American, 34(24):372, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-372.pdf.

Anonymous:1876:BPy

[8613] Anonymous. Business and personal. Scientific American, 34(24):379, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-379.pdf.

Anonymous:1876:NQy

[8614] Anonymous. Notes and queries. Scientific American, 34(24):379–380, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-379a.pdf.

Anonymous:1876:IIz

[8615] Anonymous. Index of inventions. Scientific American, 34(24):380, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-380a.pdf.

Anonymous:1876:PNPd

[8616] Anonymous. Patent notes: Patents. Scientific American, 34(24):382, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-382.pdf.

Anonymous:1876:NSI

[8617] Anonymous. New system of illumination by reflection. Scientific American, 34(24):367, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
367.pdf.

Anonymous:1876:SCd

[8618] Anonymous. Spontaneous combustion. Scientific American, 34(24):368, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-368.pdf.

Anonymous:1876:WBW

[8619] Anonymous. What is bad water? Scientific American, 34(24):368, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-368a.pdf.

Anonymous:1876:SP

[8620] Anonymous. A social peril. Scientific American, 34(24):368–369, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-368b.pdf.

Anonymous:1876:SYM

[8621] Anonymous. Suggestions for Young master mechanics. Scientific American, 34(24):369, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
369.pdf.

Anonymous:1876:DH

[8622] Anonymous. Drugging horses. Scientific American, 34(24):369, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-369a.pdf.

Anonymous:1876:KMD

[8623] Anonymous. The Keely motor deception again. Scientific American, 34(24):369, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
369b.pdf.

Anonymous:1876:CEc (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
371.pdf.

Anonymous:1876:CCG

[8625] Anonymous. Clark’s combined gas and oil chandelier. Scientific American, 34(24):371, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
371a.pdf.

Anonymous:1876:MAI

[8626] Anonymous. Moll & Altheide’s improved air pump. Scientific American, 34(24):371, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
371b.pdf.

Anonymous:1876:ANf

[8627] Anonymous. Astronomical notes. Scientific American, 34(24):372, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-372a.pdf.

Anonymous:1876:SPL

[8628] Anonymous. The scertain the proper length of a connecting rod. Scientific American, 34(24):373, June 10, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n24/pdf/scientificamerican06101876-373.pdf.

Anonymous:1876:ITWa

[8629] Anonymous. Improved turbine wheel. Scientific American, 34(24):374, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-374.pdf.

Anonymous:1876:IBSc

[8630] Anonymous. Improved balanced slide valve. Scientific American, 34(24):374, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
374a.pdf.

Anonymous:1876:TTC

[8631] Anonymous. Tea and tea culture. Scientific American, 34(24):375–376, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-375.pdf.

Iles:1876:NMD

[8632] Malvern W. Iles. New method for the detection of nickel in the presence of cobalt. Scientific American, 34(24):376–377, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-376.pdf.

Anonymous:1876:OPc

[8633] Anonymous. An old problem. Scientific American, 34(24):377, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n24/pdf/scientificamerican06101876-377.pdf.

Anonymous:1876:RAFy

[8634] Anonymous. Recent American and foreign patents. Scientific American, 34(24):378, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
378.pdf.

Anonymous:1876:NWHr

[8635] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(24):378, June 10, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n24/pdf/scientificamerican06101876-378a.pdf.

Anonymous:1876:NHIa

[8636] Anonymous. New household inventions. Scientific American, 34 (24):378, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
378b.pdf.

Anonymous:1876:NAIv (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
378c.pdf.

Anonymous:1876:NMEn

[8638] Anonymous. New mechanical and engineering inventions. Scientific American, 34(24):378, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
378d.pdf.

Anonymous:1876:CRy

[8639] Anonymous. Communications received. Scientific American, 34 (24):380, June 10, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n24/pdf/scientificamerican06101876-
380.pdf.

Anonymous:1876:Cba

[8640] Anonymous. Correspondence. Scientific American, 34(25):388, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-388.pdf.

Anonymous:1876:BNBm

[8641] Anonymous. Books: New books and publications. Scientific American, 34(25):393, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
393a.pdf.

Anonymous:1876:BPz

[8642] Anonymous. Business and personal. Scientific American, 34(25):395, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-395.pdf.

Anonymous:1876:NQz

[8643] Anonymous. Notes and queries. Scientific American, 34(25):395, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-395a.pdf.

Anonymous:1876:IIba

[8644] Anonymous. Index of inventions. Scientific American, 34(25):396, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-396.pdf.

Anonymous:1876:PNPe

[8645] Anonymous. Patent notes: Patents. Scientific American, 34(25):398, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-398.pdf.

Anonymous:1876:LBV

[8646] Anonymous. The lammergeyer or bearded vulture. Scientific American, 34(25):383, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
383.pdf.

Anonymous:1876:DTP

[8647] Anonymous. Darwin’s theory of pangenesis. Scientific American, 34(25):384, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
384.pdf.

Anonymous:1876:ADA

[8648] Anonymous. Another dastardly attack? Scientific American, 34 (25):384, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
384a.pdf.

Anonymous:1876:VAa

[8649] Anonymous. Vortex atoms. Scientific American, 34(25):385, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-385.pdf.

Anonymous:1876:CEd (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
385a.pdf.

Anonymous:1876:IACb

[8651] Anonymous. Industrial applications of carbonic acid. Scientific American, 34(25):386, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
386.pdf.

Anonymous:1876:ICH

[8652] Anonymous. Improved country homes. Scientific American, 34(25):387, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-387.pdf.

Anonymous:1876:NCMq

[8653] Anonymous. New cut-off mechanism. Scientific American, 34 (25):387–388, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
387a.pdf.

Anonymous:1876:CUT

[8654] Anonymous. The corpuscular and undulatory theories of light — the experiments of Crookes and Pleasonton. Scientific American, 34(25):388, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-388a.pdf.

Rose:1876:PMl

[8655] Joshua Rose. Practical mechanism. Scientific American, 34(25): 388–389, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
388b.pdf.

Anonymous:1876:ICPb

[8656] Anonymous. Improved carving and paneling machine. Scientific American, 34(25):390, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
390.pdf.

Anonymous:1876:SNA

[8657] Anonymous. Some new applications of metaline. Scientific American, 34(25):390, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
390a.pdf.

Anonymous:1876:EFa

[8658] Anonymous. An economic fountain. Scientific American, 34(25):390, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-390b.pdf.

Anonymous:1876:NWF

[8659] Anonymous. New winter-flowering Iris. Scientific American, 34(25):391, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-391.pdf.

Anonymous:1876:PMCa

[8660] Anonymous. Printing machinery at the Centennial. Scientific American, 34(25):391, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
391a.pdf.

Anonymous:1876:AMc

[8661] Anonymous. Agricultural machinery. Scientific American, 34 (25):392–393, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
392.pdf.

Anonymous:1876:DCr

[8662] Anonymous. Decisions of the courts. Scientific American, 34(25):393, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-393.pdf.

Anonymous:1876:RAFz

[8663] Anonymous. Recent American and foreign patents. Scientific American, 34(25):393, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
393b.pdf.

Anonymous:1876:NCMr

[8664] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(25):393–394, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
393c.pdf.

Anonymous:1876:NAIw

[8665] Anonymous. New agricultural inventions. Scientific American, 34 (25):394, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
394.pdf.

Anonymous:1876:NWHs

[8666] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 34(25):394, June 17, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v34/
n25/pdf/scientificamerican06171876-394a.pdf.

Anonymous:1876:NTMh

[8667] Anonymous. New textile machinery. Scientific American, 34(25):394, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-394b.pdf.

Anonymous:1876:NHAr

[8668] Anonymous. New household articles. Scientific American, 34(25):394, June 17, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n25/pdf/scientificamerican06171876-394c.pdf.

Anonymous:1876:CRz (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n25/pdf/scientificamerican06171876-
395b.pdf.

Anonymous:1876:Cbc

[8670] Anonymous. Correspondence. Scientific American, 34(26):404, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-404a.pdf.

Anonymous:1876:BPba

[8671] Anonymous. Business and personal. Scientific American, 34(26):409, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-409.pdf.

Anonymous:1876:NQba

[8672] Anonymous. Notes and queries. Scientific American, 34(26):409, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-409a.pdf.

Anonymous:1876:IIbb

[8673] Anonymous. Index of inventions. Scientific American, 34(26):410, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-410.pdf.

Anonymous:1876:PNPf

[8674] Anonymous. Patent notes: Patents. Scientific American, 34(26):410, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-410a.pdf.

Anonymous:1876:Ia

[8675] Anonymous. Index. Scientific American, 34(26):411–412, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-411.pdf.

Anonymous:1876:COa

[8676] Anonymous. Crystallized osmium. Scientific American, 34(26):400, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-400.pdf.

Anonymous:1876:EV

[8677] Anonymous. The end of the volume. Scientific American, 34(26):400, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-400a.pdf.

Anonymous:1876:IMf

[8678] Anonymous. Inventors misjudged. Scientific American, 34(26):400–401, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-400b.pdf.

Anonymous:1876:DN

[8679] Anonymous. A dangerous neighbor. Scientific American, 34(26):401, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-401.pdf.

Anonymous:1876:CEf

[8680] Anonymous. The Centennial Exposition. Scientific American, 34 (26):401, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
401a.pdf.

Anonymous:1876:MHG

[8681] Anonymous. Mechanical horse grooming. Scientific American, 34 (26):402, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
402.pdf.

Anonymous:1876:IBD

[8682] Anonymous. Improved barn door fastening. Scientific American, 34(26):402, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
402a.pdf.

Anonymous:1876:NSL

[8683] Anonymous. A new stone lifter. Scientific American, 34(26):402, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-402b.pdf.

Anonymous:1876:SKT

[8684] Anonymous. A simple kerosene-testing device. Scientific American, 34(26):402, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
402c.pdf.

Anonymous:1876:NSB

[8685] Anonymous. New smoke-burning grate. Scientific American, 34 (26):403, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
403.pdf.

Anonymous:1876:Pd

[8686] Anonymous. The plethysmograph. Scientific American, 34(26): 403–404, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
403a.pdf.

Anonymous:1876:CHB

[8687] Anonymous. The Centennial Horticultural Building. Scientific American, 34(26):404, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
404.pdf.

Anonymous:1876:HMP

[8688] Anonymous. Home-made probabilities. Scientific American, 34(26):404, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-404b.pdf.

Anonymous:1876:PP

[8689] Anonymous. Porcelain and pottery. Scientific American, 34(26):405, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v34/n26/pdf/scientificamerican06241876-405.pdf.

Anonymous:1876:IPPa

[8690] Anonymous. Improved portable photographic apparatus. Scientific American, 34(26):406, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
406.pdf.

Anonymous:1876:AIB

[8691] Anonymous. Another inventor of the Balestrieri reflector. Scientific American, 34(26):406, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
406a.pdf.

Anonymous:1876:CCSa

[8692] Anonymous. The contents of a Cod’s stomach. Scientific American, 34(26):407, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
407.pdf.

Anonymous:1876:MDBa

[8693] Anonymous. The mygales — door building spiders. Scientific American, 34(26):407–408, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
407a.pdf.

Anonymous:1876:RAFba

[8694] Anonymous. Recent American and foreign patent. Scientific American, 34(26):408, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
408.pdf.

Anonymous:1876:NCMs

[8695] Anonymous. New chemical and miscellaneous inventions. Scientific American, 34(26):408, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
408a.pdf.

Anonymous:1876:CRba

[8696] Anonymous. Communications received. Scientific American, 34 (26):409, June 24, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v34/n26/pdf/scientificamerican06241876-
409b.pdf.

Anonymous:1876:Cbd

[8697] Anonymous. Correspondence. Scientific American, 35(1):9, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-9.pdf.

Anonymous:1876:BNBn

[8698] Anonymous. Books: New books and publications. Scientific American, 35(1):9, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
9a.pdf.

Anonymous:1876:BPbb

[8699] Anonymous. Business and personal. Scientific American, 35(1):11, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-11.pdf.

Anonymous:1876:NQbb

[8700] Anonymous. Notes and queries. Scientific American, 35(1):11–12, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-11a.pdf.

Anonymous:1876:IIbc

[8701] Anonymous. Index of inventions. Scientific American, 35(1):12, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-12a.pdf.

Anonymous:1876:PNPg

[8702] Anonymous. Patent notes: Patents. Scientific American, 35(1):14, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-14.pdf.

Anonymous:1876:SCe

[8703] Anonymous. A singular collision. Scientific American, 35(1):1, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-1.pdf.

Anonymous:1876:GRE

[8704] Anonymous. The Gallahue rotary engine. Scientific American, 35(1):1, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-1a.pdf.

Anonymous:1876:NIE

[8705] Anonymous. New investigations on the Earth’s heat. Scientific American, 35(1):2, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
2.pdf.

Anonymous:1876:WMH

[8706] Anonymous. Working men’s homes. — Sir Sydney Waterlow’s London enterprise. Scientific American, 35(1):2–3, July 1, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n1/pdf/scientificamerican07011876-2a.pdf.

Anonymous:1876:FFb

[8707] Anonymous. The farmer’s foes. Scientific American, 35(1):3, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-3.pdf.

Anonymous:1876:MT

[8708] Anonymous. A misapplied testimonial. Scientific American, 35(1):3, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-3a.pdf.

Anonymous:1876:CEh

[8709] Anonymous. The Centennial Exposition. Scientific American, 35(1):3, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-3b.pdf.

Anonymous:1876:IHB

[8710] Anonymous. Improved hot blast boiler furnace. Scientific American, 35(1):4, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
4.pdf.

Anonymous:1876:MSR

[8711] Anonymous. McMillan’s self-regulating gas burner. Scientific American, 35(1):4, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
4a.pdf.

Anonymous:1876:SCf

[8712] Anonymous. Scenes at the Centennial. Scientific American, 35(1):4, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-4b.pdf.

Anonymous:1876:NCb

[8713] Anonymous. A novel cultivator. Scientific American, 35(1):4, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-4c.pdf.

Anonymous:1876:ISB

[8714] Anonymous. Improved safety bridge for railroad cars. Scientific American, 35(1):6, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
6.pdf.

Anonymous:1876:NDS

[8715] Anonymous. A novel device for sewing on buttons. Scientific American, 35(1):6, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
6a.pdf.

Anonymous:1876:AIDb

[8716] Anonymous. Almond’s improved drill chuck. Scientific American, 35(1): 6, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-6b.pdf.

Anonymous:1876:NBF

[8717] Anonymous. Nest-building fishes. Scientific American, 35(1):7, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-7.pdf.

Anonymous:1876:SE

[8718] Anonymous. The scorpions of Egypt. Scientific American, 35(1):7, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-7a.pdf.

Rose:1876:NMM

[8719] Joshua Rose. A new method of making surface plates. Scientific American, 35(1):8–9, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
8.pdf.

Anonymous:1876:DCs

[8720] Anonymous. Decisions of the court. Scientific American, 35(1):9–10, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-9b.pdf.

Anonymous:1876:RAFbb

[8721] Anonymous. Recent American and foreign patents. Scientific American, 35(1):10, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
10.pdf.

Anonymous:1876:NHIb

[8722] Anonymous. New household inventions. Scientific American, 35(1):10, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-10a.pdf.

Anonymous:1876:NWHt

[8723] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(1):10, July 1, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n1/pdf/scientificamerican07011876-10b.pdf.

Anonymous:1876:NCMt

[8724] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(1):10, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n1/pdf/scientificamerican07011876-
10c.pdf.

Anonymous:1876:CRbb

[8725] Anonymous. Communications received. Scientific American, 35(1):12, July 1, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n1/pdf/scientificamerican07011876-12.pdf.

Anonymous:1876:Cbe

[8726] Anonymous. Correspondence. Scientific American, 35(2):20, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-20.pdf.

Anonymous:1876:BNBo

[8727] Anonymous. Books: New books and publications. Scientific American, 35(2):25, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
25.pdf.

Anonymous:1876:BPbc

[8728] Anonymous. Business and personal. Scientific American, 35(2):27, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-27.pdf.

Anonymous:1876:NQbc

[8729] Anonymous. Notes and queries. Scientific American, 35(2):27–28, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-27a.pdf.

Anonymous:1876:IIbd

[8730] Anonymous. Index of inventions. Scientific American, 35(2):28, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-28a.pdf.

Anonymous:1876:PNPh

[8731] Anonymous. Patent notes: Patents. Scientific American, 35(2):30, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-30.pdf.

Anonymous:1876:IUM

[8732] Anonymous. Improves universal milling machine. Scientific American, 35(2):15–16, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
15.pdf.

Anonymous:1876:PAS

[8733] Anonymous. Practical applications of the spectroscope. Scientific American, 35(2):16, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
16.pdf.

Anonymous:1876:WWR

[8734] Anonymous. Why are we right-handed? Scientific American, 35(2):16– 17, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-16a.pdf.

Anonymous:1876:PDW

[8735] Anonymous. The patent drive well. Scientific American, 35(2):17, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-17.pdf.

Anonymous:1876:DA

[8736] Anonymous. Dangerous architecture. Scientific American, 35(2):17, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-17a.pdf.

Anonymous:1876:CEi

[8737] Anonymous. The Centennial Exposition. Scientific American, 35(2):17, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-17b.pdf.

Anonymous:1876:TND

[8738] Anonymous. Two new dental inventions. Scientific American, 35(2):18, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-18.pdf.

Anonymous:1876:ASMb

[8739] Anonymous. Apparatus for setting metal axles. Scientific American, 35(2):18, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
18a.pdf.

Anonymous:1876:ISCd

[8740] Anonymous. Improved shaft coupling. Scientific American, 35(2):19, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-19.pdf.

Anonymous:1876:IASc

[8741] Anonymous. Improved automatic signal buoy. Scientific American, 35(2):19–20, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
19a.pdf.

Rose:1876:PMm

[8742] Joshua Rose. Practical mechanism. Scientific American, 35(2):21, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-21.pdf.

Anonymous:1876:WFP

[8743] Anonymous. Wickersham’s fire place arch bar. Scientific American, 35(2):22, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
22.pdf.

Anonymous:1876:NACa

[8744] Anonymous. A new automatic car brake. Scientific American, 35(2):22, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-22a.pdf.

Anonymous:1876:TIL

[8745] Anonymous. Tolman’s improved lamp chimney. Scientific American, 35(2):22, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
22b.pdf.

Anonymous:1876:EAA

[8746] Anonymous. Elliott’s apparatus for applying Croton oil. Scientific American, 35(2):22, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
22c.pdf.

Anonymous:1876:IMT

[8747] Anonymous. Improved military telegraph. Scientific American, 35(2): 23, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-23.pdf.

Anonymous:1876:NYAc

[8748] Anonymous. New York Academy of Sciences. Scientific American, 35(2):24, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
24.pdf.

Anonymous:1876:ANg

[8749] Anonymous. Astronomical notes. Scientific American, 35(2):24–25, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-24a.pdf.

Anonymous:1876:DCt

[8750] Anonymous. Decisions of the courts. Scientific American, 35(2):25, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-25a.pdf.

Anonymous:1876:RAFbc

[8751] Anonymous. Recent American and foreign patents. Scientific American, 35(2):25–26, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
25b.pdf.

Anonymous:1876:NTMi

[8752] Anonymous. New textile machinery. Scientific American, 35(2):26, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-26.pdf.

Anonymous:1876:NAIx

[8753] Anonymous. New agricultural inventions. Scientific American, 35(2): 26, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-26a.pdf.

Anonymous:1876:NMEo

[8754] Anonymous. New mechanical and engineering inventions. Scientific American, 35(2):26, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n2/pdf/scientificamerican07081876-
26b.pdf.

Anonymous:1876:CRbc

[8755] Anonymous. Communications received. Scientific American, 35(2):28, July 8, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n2/pdf/scientificamerican07081876-28.pdf.

Anonymous:1876:Cbf

[8756] Anonymous. Correspondence. Scientific American, 35(3):36, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-36.pdf.

Anonymous:1876:BNBp

[8757] Anonymous. Books: New books and publications. Scientific American, 35(3):41, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
41.pdf.

Anonymous:1876:BPbd

[8758] Anonymous. Business and personal. Scientific American, 35(3):43, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-43.pdf.

Anonymous:1876:NQbd

[8759] Anonymous. Notes and queries. Scientific American, 35(3):43–44, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-43a.pdf.

Anonymous:1876:IIbe

[8760] Anonymous. Index of inventions. Scientific American, 35(3):44–45, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-44a.pdf.

Anonymous:1876:PNPi

[8761] Anonymous. Patent notes: Patents. Scientific American, 35(3):46, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-46.pdf.

Anonymous:1876:BRS

[8762] Anonymous. Band re-sawing machine. Scientific American, 35(3):31, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-31.pdf.

Anonymous:1876:LL

[8763] Anonymous. Life without light. Scientific American, 35(3):32, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-32.pdf.

Anonymous:1876:OAG

[8764] Anonymous. The oracles of ancient Greece. Scientific American, 35(3):32–33, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
32a.pdf.

Anonymous:1876:Cbg

[8765] Anonymous. Cams. Scientific American, 35(3):33, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n3/pdf/scientificamerican07151876-33.pdf.

Anonymous:1876:BO

[8766] Anonymous. Barometric observations. Scientific American, 35(3):33, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-33a.pdf.

Anonymous:1876:CEj

[8767] Anonymous. The Centennial Exposition. Scientific American, 35(3):33, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-33b.pdf.

Anonymous:1876:NMMa

[8768] Anonymous. A new mechanical money box. Scientific American, 35(3):34, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
34.pdf.

Anonymous:1876:TPR

[8769] Anonymous. Trimming and punching roofing slates. Scientific American, 35(3):34, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
34a.pdf.

Dunmore:1876:PC

[8770] Edward Dunmore. Photo-chemicals. Scientific American, 35(3):34, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-34b.pdf.

Anonymous:1876:IWB

[8771] Anonymous. Improved wagon brake. Scientific American, 35(3):35, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-35.pdf.

Anonymous:1876:IRRb

[8772] Anonymous. Improved railway rail. Scientific American, 35(3):35–36, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-35a.pdf.

Anonymous:1876:AMd

[8773] Anonymous. Animal mechanism. Scientific American, 35(3):36, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-36a.pdf.

Rose:1876:THS

[8774] Joshua Rose. Turning hard steel with the aid of petroleum as a lubricant. Scientific American, 35(3):36–37, July 15, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n3/pdf/scientificamerican07151876-36b.pdf.

Brackett:1876:HDO

[8775] C. F. Brackett. Harmony and discord with optical studies. Scientific American, 35(3):37, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
37.pdf.

Anonymous:1876:IPPb

[8776] Anonymous. Improved pony planer. Scientific American, 35(3):38, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-38.pdf.

Anonymous:1876:US

[8777] Anonymous. Umbrella supporter. Scientific American, 35(3):38, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-38a.pdf.

Anonymous:1876:CIH

[8778] Anonymous. Chaffin’s improved hay fork. Scientific American, 35(3):38, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-38b.pdf.

Anonymous:1876:VSA

[8779] Anonymous. Various species of ant Lions. Scientific American, 35(3):39, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-39.pdf.

Anonymous:1876:IMg

[8780] Anonymous. Ice machinery. Scientific American, 35(3):40–41, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-40.pdf.

Anonymous:1876:RAFbd

[8781] Anonymous. Recent American and foreign patents. Scientific American, 35(3):41, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
41a.pdf.

Anonymous:1876:NAIy

[8782] Anonymous. New agricultural inventions. Scientific American, 35(3):41, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-41b.pdf.

Anonymous:1876:NTMj

[8783] Anonymous. New textile machinery. Scientific American, 35(3):41, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-41c.pdf.

Anonymous:1876:NCMu

[8784] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(3):41–42, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
41d.pdf.

Anonymous:1876:NHIc

[8785] Anonymous. New household inventions. Scientific American, 35(3):42, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-42.pdf.

Anonymous:1876:NMEp

[8786] Anonymous. New mechanical and engineering inventions. Scientific American, 35(3):42, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n3/pdf/scientificamerican07151876-
42a.pdf.

Anonymous:1876:CRbd

[8787] Anonymous. Communications received. Scientific American, 35(3):44, July 15, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n3/pdf/scientificamerican07151876-44.pdf.

Anonymous:1876:BPbe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-59.pdf.

Anonymous:1876:NQbe

[8789] Anonymous. Notes and queries. Scientific American, 35(4):59–60, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-59a.pdf.

Anonymous:1876:IIbf

[8790] Anonymous. Index of inventions. Scientific American, 35(4):60, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-60a.pdf.

Anonymous:1876:PNPj

[8791] Anonymous. Patent notes: Patent. Scientific American, 35(4):62, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-62.pdf.

Anonymous:1876:CECa

[8792] Anonymous. The Centennial Exhibition. — the Campbell printing press. Scientific American, 35(4):47, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-47.pdf.

Anonymous:1876:SCW

[8793] Anonymous. A sensible celebration — what was done in New York City. Scientific American, 35(4):48, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-48.pdf.

Anonymous:1876:DGC

[8794] Anonymous. Do not get cool — health hints for the season. Scientific American, 35(4):48–49, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-48a.pdf.

Anonymous:1876:SSB

[8795] Anonymous. Some sources of bad water. Scientific American, 35(4):49, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-49.pdf.

Anonymous:1876:CEk

[8796] Anonymous. The Centennial Exposition. Scientific American, 35(4):49, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-49a.pdf.

Anonymous:1876:ISJ

[8797] Anonymous. Improved steam jet pump. Scientific American, 35(4):50, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-50.pdf.

Anonymous:1876:CPDa

[8798] Anonymous. Combined pocket and draw knife. Scientific American, 35(4):50, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
50a.pdf.

Anonymous:1876:NMMb

[8799] Anonymous. New mechanical movement. Scientific American, 35(4):50, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-50b.pdf.

Anonymous:1876:ISD

[8800] Anonymous. Improved self-discharging coal hod. Scientific American, 35(4):51, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
51.pdf.

Anonymous:1876:IW

[8801] Anonymous. Improved wrench. Scientific American, 35(4):51, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-51a.pdf.

Anonymous:1876:ITB

[8802] Anonymous. Improved tool for bending horseshoes. Scientific American, 35(4):51, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
51b.pdf.

Anonymous:1876:NMMc

[8803] Anonymous. New method of molding cylinders. Scientific American, 35(4):51, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
51c.pdf.

Anonymous:1876:FFC

[8804] Anonymous. Facts and figures concerning the Earth and its atmosphere. Scientific American, 35(4):52, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-52.pdf.

Rose:1876:PMn

[8805] Joshua Rose. Practical mechanism. Scientific American, 35(4):53, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-53.pdf.

Anonymous:1876:IAPb

[8806] Anonymous. Improved anvil for paper cartridge shells. Scientific American, 35(4):53, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
53a.pdf.

Anonymous:1876:ICBb

[8807] Anonymous. Improved car box grinder. Scientific American, 35(4):54, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-54.pdf.

Anonymous:1876:ISGb

[8808] Anonymous. Improved saw gumming machine. Scientific American, 35(4):54, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
54a.pdf.

Anonymous:1876:ICFb

[8809] Anonymous. Influence of chemical fertilizers on potatoes and grapes. Scientific American, 35(4):55, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-55.pdf.

Anonymous:1876:WPb

[8810] Anonymous. Water plants. Scientific American, 35(4):55, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-55a.pdf.

Anonymous:1876:MPb

[8811] Anonymous. Multiplying plants. Scientific American, 35(4):55, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-55b.pdf.

Anonymous:1876:OPd

[8812] Anonymous. The optics of photography. Scientific American, 35(4):56–57, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
56.pdf.

Anonymous:1876:DCu

[8813] Anonymous. Decisions of the courts. Scientific American, 35(4):58, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-58.pdf.

Anonymous:1876:RAFbe 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
58a.pdf.

Anonymous:1876:NCMv

[8815] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(4):58, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n4/pdf/scientificamerican07221876-
58b.pdf.

Anonymous:1876:NWHu

[8816] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(4):58, July 22, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n4/pdf/scientificamerican07221876-58c.pdf.

Anonymous:1876:NAIz

[8817] Anonymous. New agricultural inventions. Scientific American, 35(4):58, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-58d.pdf.

Anonymous:1876:CRbe

[8818] Anonymous. Communication received. Scientific American, 35(4):60, July 22, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n4/pdf/scientificamerican07221876-60.pdf.

Anonymous:1876:Cbh

[8819] Anonymous. Correspondence. Scientific American, 35(5):68, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-68.pdf.

Anonymous:1876:BPbf

[8820] Anonymous. Business and personal. Scientific American, 35(5):76, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-76.pdf.

Anonymous:1876:NQbf

[8821] Anonymous. Notes and queries. Scientific American, 35(5):76, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-76a.pdf.

Anonymous:1876:IIbg

[8822] Anonymous. Index of inventions. Scientific American, 35(5):76–77, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-76c.pdf.

Anonymous:1876:PNPk

[8823] Anonymous. Patent notes: Patents. Scientific American, 35(5):78, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-78.pdf.

Anonymous:1876:Ac

[8824] Anonymous. Aerotherapy. Scientific American, 35(5):63, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-63.pdf.

Anonymous:1876:EFb

[8825] Anonymous. The Emperor’s farewell. Scientific American, 35(5):64, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-64.pdf.

Anonymous:1876:HTT

[8826] Anonymous. Heated terms — their causes and dangers. Scientific American, 35(5):64, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
64a.pdf.

Anonymous:1876:LFD

[8827] Anonymous. Living on fifteen dollars a week. Scientific American, 35(5):64–65, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
64b.pdf.

Anonymous:1876:PGN

[8828] Anonymous. Poisonous garments — a new hot weather peril. Scientific American, 35(5):65, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
65.pdf.

Anonymous:1876:OPL

[8829] Anonymous. The obnoxious postal law partially repealed. Scientific American, 35(5):65, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
65a.pdf.

Anonymous:1876:OIE

[8830] Anonymous. Optical inversion — an external sense — perception. Scientific American, 35(5):65, July 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n5/pdf/scientificamerican07291876-65b.pdf.

Anonymous:1876:ICBc

[8831] Anonymous. Improved car brake. Scientific American, 35(5):66, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-66.pdf.

Anonymous:1876:NMSa

[8832] Anonymous. A new machine for securing bale wires. Scientific American, 35(5):66, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
66a.pdf.

Anonymous:1876:NER

[8833] Anonymous. A new electric railway signal. Scientific American, 35(5):66, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
66b.pdf.

Anonymous:1876:NSGa

[8834] Anonymous. A new scissors gage. Scientific American, 35(5):66, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-66c.pdf.

Anonymous:1876:IHE

[8835] Anonymous. Improved hay elevator and carrier. Scientific American, 35(5):67, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
67.pdf.

Anonymous:1876:HDA

[8836] Anonymous. Hill’s direct-acting drop hammer. Scientific American, 35(5):67, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
67a.pdf.

Anonymous:1876:ITWb

[8837] Anonymous. Improved tube well. Scientific American, 35(5):67, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-67b.pdf.

Anonymous:1876:NEB

[8838] Anonymous. A new electric battery. Scientific American, 35(5):67, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-67c.pdf.

Anonymous:1876:TEU

[8839] Anonymous. Technical education in the United States as illustrated at the Centennial. Scientific American, 35(5):68–69, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-68a.pdf.

Anonymous:1876:ICCb

[8840] Anonymous. Improved chimney cowl. Scientific American, 35(5):69, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-69.pdf.

Anonymous:1876:ISGc

[8841] Anonymous. Improved steam governor. Scientific American, 35(5):70, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-70.pdf.

Anonymous:1876:IHA

[8842] Anonymous. Improved hydraulic anchor lift. Scientific American, 35(5):70, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
70a.pdf.

Anonymous:1876:JSA

[8843] Anonymous. The jaguar or South American tiger. Scientific American, 35(5):71–72, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
71.pdf.

Anonymous:1876:NYAd

[8844] Anonymous. New York Academy of Sciences. Scientific American, 35(5):72, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
72.pdf.

Anonymous:1876:IRJa

[8845] Anonymous. The industries and resources of Japan. Scientific American, 35(5):73, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
73.pdf.

Anonymous:1876:DCv

[8846] Anonymous. Decisions of the courts. Scientific American, 35(5):73–74, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-73a.pdf.

Anonymous:1876:RAFbf

[8847] Anonymous. Recent American and foreign patents. Scientific American, 35(5):74, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
74.pdf.

Anonymous:1876:NCMw

[8848] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(5):74, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
74a.pdf.

Anonymous:1876:NWHv

[8849] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(5):74, July 29, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n5/pdf/scientificamerican07291876-74b.pdf.

Anonymous:1876:NHId

[8850] Anonymous. New household inventions. Scientific American, 35 (5):74–75, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n5/pdf/scientificamerican07291876-
74c.pdf.

Anonymous:1876:NAIba

[8851] Anonymous. New agricultural inventions. Scientific American, 35(5):75, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-75.pdf.

Anonymous:1876:CRbf

[8852] Anonymous. Communications received. Scientific American, 35(5):76, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-76b.pdf.

Anonymous:1876:VPb

[8853] Anonymous. The validity of patents. Scientific American, 35(5):77, July 29, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n5/pdf/scientificamerican07291876-77.pdf.

Anonymous:1876:Cbj

[8854] Anonymous. Correspondence. Scientific American, 35(6):84–85, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-84.pdf.

Anonymous:1876:BNBq

[8855] Anonymous. Books: New books and publications. Scientific American, 35(6):91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
91.pdf.

Anonymous:1876:BPbg

[8856] Anonymous. Business and personal. Scientific American, 35(6):92, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-92.pdf.

Anonymous:1876:NQbg

[8857] Anonymous. Notes and queries. Scientific American, 35(6):92, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-92a.pdf.

Anonymous:1876:IIbh

[8858] Anonymous. Index of inventions. Scientific American, 35(6):92, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-92c.pdf.

Anonymous:1876:PNPl

[8859] Anonymous. Patent notes: Patents. Scientific American, 35(6):94, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-94.pdf.

Anonymous:1876:SHC

[8860] Anonymous. Steam hand car — miniature railroad train. Scientific American, 35(6):79, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
79.pdf.

Anonymous:1876:WTIa

[8861] Anonymous. Workmen and their instructors. Scientific American, 35(6):80, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
80.pdf.

Anonymous:1876:AGG

[8862] Anonymous. Ancient Grecian glass. Scientific American, 35(6):80–81, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-80a.pdf.

Anonymous:1876:MVS

[8863] Anonymous. Mechanical vibration as a substitute for anæsthetics. Scientific American, 35(6):81, August 5, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n6/pdf/scientificamerican08051876-81.pdf.

Anonymous:1876:CMP

[8864] Anonymous. Canned meats poisoned with mercury. Scientific American, 35(6):81, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
81a.pdf.

Anonymous:1876:CEl

[8865] Anonymous. The Centennial Exposition. Scientific American, 35(6):81, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-81b.pdf.

Anonymous:1876:CAS

[8866] Anonymous. Compressed air as a street car motor. Scientific American, 35(6):82, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
82.pdf.

Anonymous:1876:UPN

[8867] Anonymous. The utilization of petroleum and natural gas in blast furnaces. Scientific American, 35(6):83, August 5, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n6/pdf/scientificamerican08051876-83.pdf.

Anonymous:1876:NCEa

[8868] Anonymous. A new compound engine jack. Scientific American, 35(6):83, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
83a.pdf.

Anonymous:1876:BOM

[8869] Anonymous. Bartlett’s ozone machine. Scientific American, 35(6):83, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-83b.pdf.

Rose:1876:PMo

[8870] Joshua Rose. Practical mechanism. Scientific American, 35(6):85, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-85.pdf.

Anonymous:1876:IAPc

[8871] Anonymous. Improved adding pencil. Scientific American, 35(6):86, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-86.pdf.

Anonymous:1876:FIL

[8872] Anonymous. Ficht’s improved lifting machine. Scientific American, 35(6):86, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
86a.pdf.

Anonymous:1876:ICRb

[8873] Anonymous. Improved check rein. Scientific American, 35(6):86, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-86b.pdf.

Anonymous:1876:GWK

[8874] Anonymous. Gas works on the kitchen range. Scientific American, 35(6):86, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
86c.pdf.

Anonymous:1876:SWK

[8875] Anonymous. Some well known British moths. Scientific American, 35(6):87–88, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
87.pdf.

Anonymous:1876:CFI

[8876] Anonymous. Chucks, forms of iron, and lubricators. Scientific American, 35(6):88–89, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
88.pdf.

Anonymous:1876:IRJb

[8877] Anonymous. The industries and resources of Japan. Scientific American, 35(6):89–90, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
89.pdf.

Anonymous:1876:OCC

[8878] Anonymous. The oceanic circulation controversy. Scientific American, 35(6):90–91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
90.pdf.

Anonymous:1876:RAFbg

[8879] Anonymous. Recent American and foreign patents. Scientific American, 35(6):91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
91a.pdf.

Anonymous:1876:NCMx

[8880] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(6):91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n6/pdf/scientificamerican08051876-
91b.pdf.

Anonymous:1876:NAIbb

[8881] Anonymous. New agricultural inventions. Scientific American, 35(6):91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-91c.pdf.

Anonymous:1876:NHIe

[8882] Anonymous. New household inventions. Scientific American, 35(6):91, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-91d.pdf.

Anonymous:1876:NWHw

[8883] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(6):91, August 5, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n6/pdf/scientificamerican08051876-91e.pdf.

Anonymous:1876:CRbg

[8884] Anonymous. Communications received. Scientific American, 35(6):92, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-92b.pdf.

Anonymous:1876:VPc

[8885] Anonymous. The validity of patents. Scientific American, 35(6):93, August 5, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n6/pdf/scientificamerican08051876-93.pdf.

Anonymous:1876:Cbk

[8886] Anonymous. Correspondence. Scientific American, 35(7):101, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-101.pdf.

Anonymous:1876:BNBr

[8887] Anonymous. Books: New books and publications. Scientific American, 35(7):105–106, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
105.pdf.

Anonymous:1876:BPbh

[8888] Anonymous. Business and personal. Scientific American, 35(7):107, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-107.pdf.

Anonymous:1876:NQbh

[8889] Anonymous. Notes and queries. Scientific American, 35(7):107– 108, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
107a.pdf.

Anonymous:1876:IIbi

[8890] Anonymous. Index of inventions. Scientific American, 35(7): 108–109, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
108a.pdf.

Anonymous:1876:PNPm

[8891] Anonymous. Patent notes: Patents. Scientific American, 35(7):110, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-110.pdf.

Anonymous:1876:NIH

[8892] Anonymous. New improvements in the Harris–Corliss steam engine. Scientific American, 35(7):95, August 12, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n7/pdf/scientificamerican08121876-95.pdf.

Anonymous:1876:MRR

[8893] Anonymous. Modern rowing races. Scientific American, 35(7):96, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-96.pdf.

Anonymous:1876:OYY

[8894] Anonymous. Our yachts and yachting. Scientific American, 35 (7):96, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
96a.pdf.

Anonymous:1876:VRC

[8895] Anonymous. The ventilation of railway cars. Scientific American, 35(7):96–97, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
96b.pdf.

Anonymous:1876:TKD

[8896] Anonymous. Timely kerosene dangers. Scientific American, 35 (7):97, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
97.pdf.

Anonymous:1876:TBE

[8897] Anonymous. The thunderer boiler explosion. Scientific American, 35(7):97, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
97a.pdf.

Anonymous:1876:CEm

[8898] Anonymous. The Centennial Exposition. Scientific American, 35 (7):97, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
97b.pdf.

Anonymous:1876:IFPc

[8899] Anonymous. Improved fountain pen. Scientific American, 35(7):98, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-98.pdf.

Anonymous:1876:MIW

[8900] Anonymous. Mediæval iron work. Scientific American, 35(7):99, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-99.pdf.

Anonymous:1876:ISWa

[8901] Anonymous. Improved spanner wrench. Scientific American, 35 (7):99, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
99a.pdf.

Anonymous:1876:SRC

[8902] Anonymous. A simple rope clamp. Scientific American, 35(7):99, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n7/pdf/scientificamerican08121876-99b.pdf.

Anonymous:1876:NIP

[8903] Anonymous. A new insect powder gun. Scientific American, 35 (7):99, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
99c.pdf.

Anonymous:1876:ETC

[8904] Anonymous. Effects of tide currents on harbors. Scientific American, 35(7):100–101, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
100.pdf.

Anonymous:1876:IDH

[8905] Anonymous. Improved device for heating air for furnaces. Scientific American, 35(7):102, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
102.pdf.

Anonymous:1876:ISE

[8906] Anonymous. Improved step and extension ladder. Scientific American, 35(7):102, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
102a.pdf.

Anonymous:1876:ISGd

[8907] Anonymous. Improved speed governor. Scientific American, 35 (7):102, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
102b.pdf.

Anonymous:1876:BTI

[8908] Anonymous. Bees and their institutions. Scientific American, 35 (7):103–105, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
103.pdf.

Anonymous:1876:DCw

[8909] Anonymous. Decisions of the courts. Scientific American, 35 (7):106, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
106.pdf.

Anonymous:1876:RAFbh

[8910] Anonymous. Recent American and foreign patents. Scientific American, 35(7):106, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
106a.pdf.

Anonymous:1876:NCMy

[8911] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(7):106, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
106b.pdf.

Anonymous:1876:NWHx

[8912] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(7):106, August 12, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n7/pdf/scientificamerican08121876-106c.pdf.

Anonymous:1876:CRbh

[8913] Anonymous. Communications received. Scientific American, 35 (7):108, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
108.pdf.

Anonymous:1876:VPd

[8914] Anonymous. The validity of patents. Scientific American, 35 (7):109, August 12, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n7/pdf/scientificamerican08121876-
109.pdf.

Anonymous:1876:Cbl

[8915] Anonymous. Correspondence. Scientific American, 35(8):116, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-116.pdf.

Anonymous:1876:BPbi

[8916] Anonymous. Business and personal. Scientific American, 35(8):123, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-123.pdf.

Anonymous:1876:NQbi

[8917] Anonymous. Notes and queries. Scientific American, 35(8):123– 124, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
123a.pdf.

Anonymous:1876:IIbj

[8918] Anonymous. Index of inventions. Scientific American, 35(8): 124–125, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
124a.pdf.

Anonymous:1876:PNPn

[8919] Anonymous. Patent notes: Patents. Scientific American, 35(8):126, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-126.pdf.

Anonymous:1876:HAC

[8920] Anonymous. The hydraulic annexe at the Centennial Exposition. Scientific American, 35(8):111, August 19, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n8/pdf/scientificamerican08191876-111.pdf.

Anonymous:1876:BNE

[8921] Anonymous. Berthelot’s new electro-chemical discoveries. Scientific American, 35(8):112, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
112.pdf.

Anonymous:1876:UCE

[8922] Anonymous. Is the universe composed entirely of hydrogen! Scientific American, 35(8):112, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
112a.pdf.

Anonymous:1876:OCS

[8923] Anonymous. Ocular color spectra and their causation. Scientific American, 35(8):112–113, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
112b.pdf.

Anonymous:1876:NTH

[8924] Anonymous. A new theory of hay fever. Scientific American, 35(8):113, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
113.pdf.

Anonymous:1876:SWC

[8925] Anonymous. Shall we change our weights and measures? Scientific American, 35(8):113, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
113a.pdf.

Anonymous:1876:NGG

[8926] Anonymous. New Gatling gun. Scientific American, 35(8):114, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-114.pdf.

Anonymous:1876:ISWb

[8927] Anonymous. Improved split wheel. Scientific American, 35(8):114, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-114a.pdf.

Anonymous:1876:IMMb

[8928] Anonymous. Improved mortising machine. Scientific American, 35(8):114, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
114b.pdf.

Anonymous:1876:IFE

[8929] Anonymous. Improvements in fire engines. Scientific American, 35(8):115, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
115.pdf.

Anonymous:1876:NSM

[8930] Anonymous. A new soldering machine. Scientific American, 35 (8):115, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
115a.pdf.

Riley:1876:PPP

[8931] Charles V. Riley. “potato pest poison”. Scientific American, 35 (8):116–117, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
116a.pdf.

Rose:1876:PMp

[8932] Joshua Rose. Practical mechanism. Scientific American, 35(8):117, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-117.pdf.

Anonymous:1876:IWMa

[8933] Anonymous. Improved woodworking machinery. Scientific American, 35(8):118, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
118.pdf.

Anonymous:1876:LPCa

[8934] Anonymous. Leland’s pipe coupler. Scientific American, 35(8):118, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-118a.pdf.

Anonymous:1876:TBPa

[8935] Anonymous. Two beautiful palms. Scientific American, 35(8):119, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-119.pdf.

Young:1876:M

[8936] C. A. Young. Meteorites. Scientific American, 35(8):119–120, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n8/pdf/scientificamerican08191876-119a.pdf.

Anonymous:1876:EFT

[8937] Anonymous. Exhibits of foreign technical schools at the Centennial. Scientific American, 35(8):121, August 19, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n8/pdf/scientificamerican08191876-121.pdf.

Anonymous:1876:RAFbi

[8938] Anonymous. Recent American and foreign patents. Scientific American, 35(8):121, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
121a.pdf.

Anonymous:1876:NHIf

[8939] Anonymous. New household inventions. Scientific American, 35 (8):121–122, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
121b.pdf.

Anonymous:1876:NMEq

[8940] Anonymous. New mechanical and engineering inventions. Scientific American, 35(8):122, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
122.pdf.

Anonymous:1876:NAIbc

[8941] Anonymous. New agricultural inventions. Scientific American, 35(8):122, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
122a.pdf.

Anonymous:1876:NCMz

[8942] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(8):122, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
122b.pdf.

Anonymous:1876:CRbi

[8943] Anonymous. Communications received. Scientific American, 35 (8):124, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
124.pdf.

Anonymous:1876:VPe

[8944] Anonymous. The validity of patents. Scientific American, 35 (8):125, August 19, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n8/pdf/scientificamerican08191876-
125.pdf.

Anonymous:1876:Cbm

[8945] Anonymous. Correspondence. Scientific American, 35(9):132, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-132.pdf.

Anonymous:1876:BNBs

[8946] Anonymous. Books: New books and publications. Scientific American, 35(9):137, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
137.pdf.

Anonymous:1876:BPbj

[8947] Anonymous. Business and personal. Scientific American, 35(9):139, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-139.pdf.

Anonymous:1876:NQbj

[8948] Anonymous. Notes and queries. Scientific American, 35(9):139– 140, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
139a.pdf.

Anonymous:1876:IIbk

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-140a.pdf.

Anonymous:1876:PNPo

[8950] Anonymous. Patent notes: Patents. Scientific American, 35(9):142, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-142.pdf.

Anonymous:1876:IHF

[8951] Anonymous. Improved hand fire engine and hose carriage. Scientific American, 35(9):127, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
127.pdf.

Anonymous:1876:PMA

[8952] Anonymous. Physical man in America. Scientific American, 35 (9):128, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
128.pdf.

Anonymous:1876:SSa

[8953] Anonymous. Summer science. Scientific American, 35(9):128, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-128a.pdf.

Anonymous:1876:AGb

[8954] Anonymous. Artificial guano. Scientific American, 35(9):128–129, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-128b.pdf.

Anonymous:1876:SAR

[8955] Anonymous. The structure and age of the Rocky Mountains. Scientific American, 35(9):129, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
129.pdf.

Anonymous:1876:ILa

[8956] Anonymous. Incendiary locomotives. Scientific American, 35(9): 129, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
129a.pdf.

Anonymous:1876:BLFa

[8957] Anonymous. Breech-loading firearms. Scientific American, 35 (9):130, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
130.pdf.

Anonymous:1876:IAAb

[8958] Anonymous. Improved amalgamating apparatus. Scientific American, 35(9):131, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
131.pdf.

Anonymous:1876:NLT

[8959] Anonymous. New lampwick trimmers. Scientific American, 35 (9):131, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
131a.pdf.

Anonymous:1876:NEH

[8960] Anonymous. A new egg holder. Scientific American, 35(9):131–132, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n9/pdf/scientificamerican08261876-131b.pdf.

Anonymous:1876:DSb

[8961] Anonymous. The development of speech. Scientific American, 35(9):132, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
132a.pdf.

Anonymous:1876:FCR

[8962] Anonymous. The first Chinese railway. Scientific American, 35 (9):132, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
132b.pdf.

Anonymous:1876:ILb

[8963] Anonymous. The impact of light. Scientific American, 35(9): 132–133, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
132c.pdf.

Anonymous:1876:NRS

[8964] Anonymous. New rail saw at the landore steel works. Scientific American, 35(9):134, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
134.pdf.

Anonymous:1876:JFB

[8965] Anonymous. A Japanese flower basket. Scientific American, 35 (9):134, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
134a.pdf.

Anonymous:1876:SCEa

[8966] Anonymous. A Swedish Centennial Exhibit. Scientific American, 35(9):134, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
134b.pdf.

Anonymous:1876:MAc

[8967] Anonymous. A marine aquarium. Scientific American, 35(9): 135–136, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
135.pdf.

Anonymous:1876:MTP

[8968] Anonymous. Musical telegraphy in Paris. Scientific American, 35 (9):136–137, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
136.pdf.

Anonymous:1876:RAFbj

[8969] Anonymous. Recent American and foreign patents. Scientific American, 35(9):137, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
137a.pdf.

Anonymous:1876:NWHy

[8970] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(9):137, August 26, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n9/pdf/scientificamerican08261876-137b.pdf.

Anonymous:1876:NAIbd

[8971] Anonymous. New agricultural inventions. Scientific American, 35(9):137, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
137c.pdf.

Anonymous:1876:NMEr

[8972] Anonymous. New mechanical and engineering inventions. Scientific American, 35(9):138, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
138.pdf.

Anonymous:1876:NCMba

[8973] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(9):138, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
138a.pdf.

Anonymous:1876:NTMk

[8974] Anonymous. New textile machinery. Scientific American, 35 (9):138, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
138b.pdf.

Anonymous:1876:CRbj

[8975] Anonymous. Communications received. Scientific American, 35 (9):140, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
140.pdf.

Anonymous:1876:VPf

[8976] Anonymous. The validity of patents. Scientific American, 35 (9):140, August 26, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n9/pdf/scientificamerican08261876-
140b.pdf.

Anonymous:1876:BPbk

[8977] Anonymous. Business and personal. Scientific American, 35 (10):156, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
156.pdf.

Anonymous:1876:NQbk

[8978] Anonymous. Notes and queries. Scientific American, 35(10): 156, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
156a.pdf.

Anonymous:1876:IIbl

[8979] Anonymous. Index of inventions. Scientific American, 35(10): 156–157, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
156c.pdf.

Anonymous:1876:PNPp

[8980] Anonymous. Patent notes: Patents. Scientific American, 35 (10):158, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
158.pdf.

Anonymous:1876:NBPb

[8981] Anonymous. New bridge at Pittsburgh, Pa. Scientific American, 35(10):143, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
143.pdf.

Anonymous:1876:MAd

[8982] Anonymous. The Moon’s atmosphere. Scientific American, 35 (10):144, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
144.pdf.

Anonymous:1876:RWM

[8983] Anonymous. The recent working Men’s demonstrations. Scientific American, 35(10):144, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
144a.pdf.

Anonymous:1876:AS

[8984] Anonymous. Is anybody sane? Scientific American, 35(10): 145, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
145.pdf.

Anonymous:1876:HTW

[8985] Anonymous. How tall are we? Scientific American, 35(10):145, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n10/pdf/scientificamerican09021876-145a.pdf.

Anonymous:1876:IDMa

[8986] Anonymous. Improved dredging machinery. Scientific American, 35(10):146, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
146.pdf.

Anonymous:1876:IWMb

[8987] Anonymous. Improved woodworking machinery. Scientific American, 35(10):147, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
147.pdf.

Anonymous:1876:ISRc

[8988] Anonymous. Improved six-roll planer, matcher, and bender. Scientific American, 35(10):147–148, September 2, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n10/pdf/scientificamerican09021876-147a.pdf.

Anonymous:1876:ECU

[8989] Anonymous. The exhibit of Cornell University at the Centennial. Scientific American, 35(10):148, September 2, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n10/pdf/scientificamerican09021876-148.pdf.

Anonymous:1876:FLG

[8990] Anonymous. The facts of the laws of gravitation. Scientific American, 35(10):148–149, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
148a.pdf.

Anonymous:1876:NP

[8991] Anonymous. Nickel and its preparation. Scientific American, 35 (10):149, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
149.pdf.

Anonymous:1876:SA

[8992] Anonymous. Scientific apparatus. Scientific American, 35(10): 149, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
149a.pdf.

Anonymous:1876:IBL

[8993] Anonymous. Improved boat-lowering apparatus. Scientific American, 35(10):150, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
150.pdf.

Anonymous:1876:DIWa

[8994] Anonymous. Dueber’s improved watch case. Scientific American, 35(10):150, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
150a.pdf.

Anonymous:1876:IWMc

[8995] Anonymous. Improved water motor. Scientific American, 35 (10):150, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
150b.pdf.

Anonymous:1876:WDE

[8996] Anonymous. The Weston dynamo-electric machine. Scientific American, 35(10):150, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
150c.pdf.

Anonymous:1876:OB

[8997] Anonymous. Oceanic birds. Scientific American, 35(10):151, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n10/pdf/scientificamerican09021876-151.pdf.

Anonymous:1876:BLFb

[8998] Anonymous. Breech-loading firearms. Scientific American, 35 (10):152, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
152.pdf.

Anonymous:1876:T

[8999] Anonymous. The tapayaxin. Scientific American, 35(10):153, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n10/pdf/scientificamerican09021876-153.pdf.

Anonymous:1876:IDK

[9000] Anonymous. Improved door knob. Scientific American, 35(10): 153–154, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
153a.pdf.

Anonymous:1876:RAFbk

[9001] Anonymous. Recent American and foreign patents. Scientific American, 35(10):154–155, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
154.pdf.

Anonymous:1876:NHIg

[9002] Anonymous. New household inventions. Scientific American, 35 (10):155, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
155.pdf.

Anonymous:1876:NAIbe

[9003] Anonymous. New agricultural inventions. Scientific American, 35 (10):155, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
155a.pdf.

Anonymous:1876:NMEs

[9004] Anonymous. New mechanical and engineering inventions. Scientific American, 35(10):155, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
155b.pdf.

Anonymous:1876:CRbk

[9005] Anonymous. Communications received. Scientific American, 35 (10):156, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
156b.pdf.

Anonymous:1876:VPg

[9006] Anonymous. The validity of patents. Scientific American, 35 (10):158, September 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n10/pdf/scientificamerican09021876-
158a.pdf.

Anonymous:1876:Cbo

[9007] Anonymous. Correspondence. Scientific American, 35(11):164, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n11/pdf/scientificamerican09091876-164.pdf.

Anonymous:1876:BPbl

[9008] Anonymous. Business and personal. Scientific American, 35 (11):171, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
171.pdf.

Anonymous:1876:NQbl

[9009] Anonymous. Notes and queries. Scientific American, 35(11): 171–172, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
171a.pdf.

Anonymous:1876:IIbm

[9010] Anonymous. Index of inventions. Scientific American, 35(11): 172, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
172a.pdf.

Anonymous:1876:PNPq

[9011] Anonymous. Patent notes: Patents. Scientific American, 35 (11):174, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
174.pdf.

Anonymous:1876:IDMb

[9012] Anonymous. Improved ditching machine. Scientific American, 35 (11):159, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
159.pdf.

Anonymous:1876:GFM

[9013] Anonymous. The Green fields of the Moon. Scientific American, 35(11):160, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
160.pdf.

Anonymous:1876:WTT

[9014] Anonymous. Workmen and their teachers. Scientific American, 35 (11):160–161, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
160a.pdf.

Anonymous:1876:RBD

[9015] Anonymous. Is rare beef dangerous? Scientific American, 35 (11):161, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
161.pdf.

Anonymous:1876:TE

[9016] Anonymous. The “Thunderer” explosion. Scientific American, 35 (11):161, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
161a.pdf.

Anonymous:1876:LRa

[9017] Anonymous. Lightning rods. Scientific American, 35(11):161, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n11/pdf/scientificamerican09091876-161b.pdf.

Anonymous:1876:BLFc

[9018] Anonymous. Breech-loading firearms. Scientific American, 35 (11):162, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
162.pdf.

Anonymous:1876:EN

[9019] Anonymous.  Exhibition notes.  Scientific American, 35(11):

162, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
162a.pdf.

Anonymous:1876:IPCb

[9020] Anonymous. Improved pipe cutting and threading machine. Scientific American, 35(11):163, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
163.pdf.

Anonymous:1876:ISGe

[9021] Anonymous. Improved sand and gravel separator. Scientific American, 35(11):163–164, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
163a.pdf.

Anonymous:1876:CMI

[9022] Anonymous. The calorimeter of the Massachusetts institute of technology. Scientific American, 35(11):164, September 9, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n11/pdf/scientificamerican09091876-164a.pdf.

Rose:1876:PMq

[9023] Joshua Rose. Practical mechanism. Scientific American, 35(11): 165, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
165.pdf.

Anonymous:1876:ICLa

[9024] Anonymous. Improved clothes line fastener. Scientific American, 35(11):166, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
166.pdf.

Anonymous:1876:IMPc

[9025] Anonymous. Improved metal-punching machine. Scientific American, 35(11):166, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
166a.pdf.

Anonymous:1876:R

[9026] Anonymous. The radiata. Scientific American, 35(11):167, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n11/pdf/scientificamerican09091876-167.pdf.

Anonymous:1876:CCc

[9027] Anonymous. Chemicals at the Centennial. Scientific American, 35(11):168, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
168.pdf.

Anonymous:1876:BNa

[9028] Anonymous.  Boiler notes.  Scientific American, 35(11):168–

169, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
168a.pdf.

Anonymous:1876:ODR

[9029] Anonymous. Occupation and the death rate in England. Scientific American, 35(11):169, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
169.pdf.

Anonymous:1876:RAFbl

[9030] Anonymous. Recent American and foreign patents. Scientific American, 35(11):169–170, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
169a.pdf.

Anonymous:1876:NCMbb

[9031] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(11):170, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
170.pdf.

Anonymous:1876:NWHz

[9032] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(11):170, September 9, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n11/pdf/scientificamerican09091876-170a.pdf.

Anonymous:1876:NHIh

[9033] Anonymous. New household inventions. Scientific American, 35 (11):170, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
170b.pdf.

Anonymous:1876:NTMl

[9034] Anonymous. New textile machinery. Scientific American, 35 (11):170, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
170c.pdf.

Anonymous:1876:NAIbf

[9035] Anonymous. New agricultural inventions. Scientific American, 35 (11):170, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
170d.pdf.

Anonymous:1876:CRbl

[9036] Anonymous. Communications received. Scientific American, 35 (11):172, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
172.pdf.

Anonymous:1876:VPh

[9037] Anonymous. The validity of patents. Scientific American, 35 (11):172, September 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n11/pdf/scientificamerican09091876-
172b.pdf.

Anonymous:1876:Cbq

[9038] Anonymous. Correspondence. Scientific American, 35(12):181, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-181.pdf.

Anonymous:1876:BNBt

[9039] Anonymous. Books: New books and publications. Scientific American, 35(12):185, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
185a.pdf.

Anonymous:1876:BPbm

[9040] Anonymous. Business and personal. Scientific American, 35 (12):187, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
187.pdf.

Anonymous:1876:NQbm

[9041] Anonymous. Notes and queries. Scientific American, 35(12): 187–188, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
187a.pdf.

Anonymous:1876:IIbn

[9042] Anonymous. Index of inventions. Scientific American, 35(12): 188, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
188a.pdf.

Anonymous:1876:PNPr

[9043] Anonymous. Patent notes: Patents. Scientific American, 35 (12):190, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
190.pdf.

Anonymous:1876:GTI

[9044] Anonymous. The great textile invention at the Centennial Exposition — the Lyall positive motion loom. Scientific American, 35 (12):175, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
175.pdf.

Anonymous:1876:ACDb

[9045] Anonymous. American contributions to the development theory. Scientific American, 35(12):176, September 16, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n12/pdf/scientificamerican09161876-176.pdf.

Anonymous:1876:CEHa

[9046] Anonymous. The coming explosion at Hell Gate. Scientific American, 35(12):176, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
176a.pdf.

Anonymous:1876:CCd

[9047] Anonymous. Cheap coal. Scientific American, 35(12):176–177, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-176b.pdf.

Anonymous:1876:SMM

[9048] Anonymous. The science of mechanical movements. Scientific American, 35(12):177, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
177.pdf.

Anonymous:1876:SNL

[9049] Anonymous. Some new light on the origin of the truffle. Scientific American, 35(12):177, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
177a.pdf.

Anonymous:1876:BJV

[9050] Anonymous. Is Brother Jonathan so very slim? Scientific American, 35(12):177, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
177b.pdf.

Anonymous:1876:DHP

[9051] Anonymous. The detachable handle parasol and umbrella. Scientific American, 35(12):178, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
178.pdf.

Anonymous:1876:HSS

[9052] Anonymous. Hangings for sliding and swinging gates. Scientific American, 35(12):179, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
179.pdf.

Anonymous:1876:MSb

[9053] Anonymous. A marine safe. Scientific American, 35(12):179, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-179a.pdf.

Anonymous:1876:AAD

[9054] Anonymous. The Acme automatic draft regulator. Scientific American, 35(12):179, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
179b.pdf.

Anonymous:1876:LRE

[9055] Anonymous. The Lidgerwood rotary engine at the Exhibition. Scientific American, 35(12):182, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
182.pdf.

Anonymous:1876:WIP

[9056] Anonymous. Wood’s improved piston. Scientific American, 35 (12):182, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
182a.pdf.

Anonymous:1876:BSG

[9057] Anonymous. Bedell’s sewer gas trap. Scientific American, 35 (12):182, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
182b.pdf.

Anonymous:1876:HAF

[9058] Anonymous. Hose attachment for faucets. Scientific American, 35 (12):182–183, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
182c.pdf.

Anonymous:1876:PAa

[9059] Anonymous. The pearly aloe. Scientific American, 35(12):183, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-183.pdf.

Anonymous:1876:NCEb

[9060] Anonymous. New and curious electrical experiments. Scientific American, 35(12):184, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
184.pdf.

Anonymous:1876:PAAa

[9061] Anonymous. Proceedings of the American Association for the Advancement of Science. Scientific American, 35(12):185, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-185.pdf.

Anonymous:1876:RAFbm

[9062] Anonymous. Recent American and foreign patents. Scientific American, 35(12):185–186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
185b.pdf.

Anonymous:1876:NCMbc

[9063] Anonymous. New chemical and miscellaneous inventions. Scientific American, 35(12):186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
186.pdf.

Anonymous:1876:NHIi

[9064] Anonymous. New household inventions. Scientific American, 35 (12):186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
186a.pdf.

Anonymous:1876:NAIbg

[9065] Anonymous. New agricultural inventions. Scientific American, 35 (12):186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
186b.pdf.

Anonymous:1876:NWHba

[9066] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(12):186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n12/pdf/scientificamerican09161876-186c.pdf.

Anonymous:1876:NTMm

[9067] Anonymous. New textile machinery. Scientific American, 35 (12):186, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
186d.pdf.

Anonymous:1876:CRbm

[9068] Anonymous. Communications received. Scientific American, 35 (12):188, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
188.pdf.

Anonymous:1876:VPi

[9069] Anonymous. The validity of patents. Scientific American, 35 (12):188, September 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n12/pdf/scientificamerican09161876-
188b.pdf.

Anonymous:1876:Cbr

[9070] Anonymous. Correspondence. Scientific American, 35(13):196, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n13/pdf/scientificamerican09231876-196.pdf.

Anonymous:1876:BNBu

[9071] Anonymous. Books: New books and publications. Scientific American, 35(13):201, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
201.pdf.

Anonymous:1876:BPbn

[9072] Anonymous. Business and personal. Scientific American, 35 (13):203, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
203.pdf.

Anonymous:1876:NQbn

[9073] Anonymous. Notes and queries. Scientific American, 35(13): 203–204, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
203a.pdf.

Anonymous:1876:IIbo

[9074] Anonymous. Index of inventions. Scientific American, 35(13): 204, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
204a.pdf.

Anonymous:1876:PNPs

[9075] Anonymous. Patent notes: Patents. Scientific American, 35 (13):206, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
206.pdf.

Anonymous:1876:ISSc

[9076] Anonymous. Improved stone-sawing machinery. Scientific American, 35(13):191, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
191.pdf.

Anonymous:1876:TPO

[9077] Anonymous. Two personalities in one person. Scientific American, 35(13):192, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
192.pdf.

Anonymous:1876:TDN

[9078] Anonymous. Thermo-diffusion — a new physical phenomenon. Scientific American, 35(13):192–193, September 23, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n13/pdf/scientificamerican09231876-192a.pdf.

Anonymous:1876:PIP

[9079] Anonymous. Practical information for practical men. Scientific American, 35(13):193, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
193.pdf.

Anonymous:1876:WTIb

[9080] Anonymous. Workmen and their instructors. Scientific American, 35(13):193, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
193a.pdf.

Anonymous:1876:BS

[9081] Anonymous. Born scientists. Scientific American, 35(13):193, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n13/pdf/scientificamerican09231876-193b.pdf.

Anonymous:1876:EFc

[9082] Anonymous. Educated farmers. Scientific American, 35(13): 193, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
193c.pdf.

Anonymous:1876:IAL

[9083] Anonymous. Improved apparatus for lining insoles of boots. Scientific American, 35(13):194, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
194.pdf.

Anonymous:1876:CKT

[9084] Anonymous. Combined knife and tape line. Scientific American, 35(13):194, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
194a.pdf.

Anonymous:1876:NCTa

[9085] Anonymous. New combination tool for sewing machines. Scientific American, 35(13):194, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
194b.pdf.

Anonymous:1876:NOC

[9086] Anonymous. A new oil cup. Scientific American, 35(13):194, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n13/pdf/scientificamerican09231876-194c.pdf.

Anonymous:1876:ICSb

[9087] Anonymous. Improved construction of stables. Scientific American, 35(13):195, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
195.pdf.

Anonymous:1876:PEC

[9088] Anonymous. A paper egg cup. Scientific American, 35(13):195, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n13/pdf/scientificamerican09231876-195a.pdf.

Anonymous:1876:SDM

[9089] Anonymous. A simple dividing machine. Scientific American, 35 (13):195, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
195b.pdf.

Anonymous:1876:NCTb

[9090] Anonymous. New combination tool. Scientific American, 35 (13):195, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
195c.pdf.

Anonymous:1876:NRM

[9091] Anonymous. Notes on the resistance of materials. Scientific American, 35(13):196, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
196a.pdf.

Anonymous:1876:CCe

[9092] Anonymous. Chemicals at the Centennial. Scientific American, 35 (13):197, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
197.pdf.

Anonymous:1876:ASS

[9093] Anonymous. The American social science association. Scientific American, 35(13):197, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
197a.pdf.

Anonymous:1876:NSE

[9094] Anonymous. A new steam engine. Scientific American, 35(13): 198, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
198.pdf.

Anonymous:1876:OSS

[9095] Anonymous. The original steam steering apparatus. Scientific American, 35(13):198–199, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
198a.pdf.

Anonymous:1876:HBa

[9096] Anonymous. The honey buzzard. Scientific American, 35(13): 199–200, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
199.pdf.

Anonymous:1876:CGF

[9097] Anonymous. Chain gear and fastenings. Scientific American, 35 (13):200–201, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
200.pdf.

Anonymous:1876:RAFbn

[9098] Anonymous. Recent American and foreign patents. Scientific American, 35(13):201, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
201a.pdf.

Anonymous:1876:NMEt

[9099] Anonymous. New mechanical and engineering inventions. Scientific American, 35(13):201, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
201b.pdf.

Anonymous:1876:NWHbb

[9100] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(13):202, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n13/pdf/scientificamerican09231876-202.pdf.

Anonymous:1876:NAIbh

[9101] Anonymous. New agricultural inventions. Scientific American, 35 (13):202, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
202a.pdf.

Anonymous:1876:NHIj

[9102] Anonymous. New household inventions. Scientific American, 35 (13):202, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
202b.pdf.

Anonymous:1876:CRbn

[9103] Anonymous. Communications received. Scientific American, 35 (13):204, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
204.pdf.

Anonymous:1876:VPj

[9104] Anonymous. The validity of patents. Scientific American, 35 (13):204, September 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n13/pdf/scientificamerican09231876-
204b.pdf.

Anonymous:1876:Cbt

[9105] Anonymous. Correspondence. Scientific American, 35(14):212, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n14/pdf/scientificamerican09301876-212.pdf.

Anonymous:1876:BPbo

[9106] Anonymous. Business and personal. Scientific American, 35 (14):219, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
219.pdf.

Anonymous:1876:NQbo

[9107] Anonymous. Notes and queries. Scientific American, 35(14): 219–220, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
219a.pdf.

Anonymous:1876:IIbp

[9108] Anonymous. Index of inventions. Scientific American, 35(14): 220, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
220a.pdf.

Anonymous:1876:PNPt

[9109] Anonymous. Patent notes: Patents. Scientific American, 35 (14):222, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
222.pdf.

Anonymous:1876:CECb

[9110] Anonymous. The Centennial Exposition. — the center of the main building. Scientific American, 35(14):207, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n14/pdf/scientificamerican09301876-207.pdf.

Anonymous:1876:SVP

[9111] Anonymous. Surgery vs phrenology. Scientific American, 35(14): 208–209, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
208.pdf.

Anonymous:1876:AER

[9112] Anonymous. American and English railways. Scientific American, 35(14):209, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
209.pdf.

Anonymous:1876:FSV

[9113] Anonymous. The friction of slide valves on their seats. Scientific American, 35(14):209, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
209a.pdf.

Anonymous:1876:OI

[9114] Anonymous. An opportunity for inventors. Scientific American, 35(14):209, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
209b.pdf.

Anonymous:1876:VDC

[9115] Anonymous. The Victor drill chuck. Scientific American, 35 (14):210, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
210.pdf.

Anonymous:1876:ALS

[9116] Anonymous. Automatic leather scourer and hide worker. Scientific American, 35(14):210, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
210a.pdf.

Anonymous:1876:ICSc

[9117] Anonymous. Improved compound steam pump. Scientific American, 35(14):211, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
211.pdf.

Anonymous:1876:MIB

[9118] Anonymous. May’s improved billiard table leveler. Scientific American, 35(14):211–212, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
211a.pdf.

Anonymous:1876:CIR

[9119] Anonymous. The Centennial International Rifle Match. Scientific American, 35(14):213, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
213.pdf.

Anonymous:1876:HS

[9120] Anonymous. Heating by steam. Scientific American, 35(14): 213, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
213a.pdf.

Anonymous:1876:RHG

[9121] Anonymous. Removal of the Hell Gate rocks. Scientific American, 35 (14):214–216, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
214.pdf.

Anonymous:1876:CCf

[9122] Anonymous. Chemicals at the Centennial. Scientific American, 35 (14):216–217, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
216.pdf.

Anonymous:1876:CNc

[9123] Anonymous.  Centennial notes.  Scientific American, 35(14):

217, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
217.pdf.

Anonymous:1876:RAFbo

[9124] Anonymous. Recent American and foreign patents. Scientific American, 35(14):217–218, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
217a.pdf.

Anonymous:1876:NAIbi

[9125] Anonymous. New agricultural inventions. Scientific American, 35 (14):218, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
218.pdf.

Anonymous:1876:NMIa

[9126] Anonymous. New miscellaneous inventions. Scientific American, 35(14):218, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
218a.pdf.

Anonymous:1876:NHIk

[9127] Anonymous. New household inventions. Scientific American, 35 (14):218, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
218b.pdf.

Anonymous:1876:NWHbc

[9128] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(14):218, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n14/pdf/scientificamerican09301876-218c.pdf.

Anonymous:1876:CRbo

[9129] Anonymous. Communications received. Scientific American, 35 (14):220, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
220.pdf.

Anonymous:1876:VPk

[9130] Anonymous. The validity of patents. Scientific American, 35 (14):220, September 30, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n14/pdf/scientificamerican09301876-
220b.pdf.

Anonymous:1876:Cbu

[9131] Anonymous. Correspondence. Scientific American, 35(15):228, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-228.pdf.

Anonymous:1876:BPbp

[9132] Anonymous. Business and personal. Scientific American, 35(15):235, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-235.pdf.

Anonymous:1876:NQbp

[9133] Anonymous. Notes and queries. Scientific American, 35(15):235–236, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-235a.pdf.

Anonymous:1876:IIbq

[9134] Anonymous. Index of inventions. Scientific American, 35(15):236–237, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-236a.pdf.

Anonymous:1876:PNPu

[9135] Anonymous. Patent notes: Patents. Scientific American, 35(15):238, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-238.pdf.

Anonymous:1876:BSM

[9136] Anonymous. A Bookbinder’s sewing machine. Scientific American, 35(15):223, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
223.pdf.

Anonymous:1876:PHL

[9137] Anonymous. Professor Huxley’s lectures in America. Scientific American, 35(15):224–225, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
224.pdf.

Anonymous:1876:SPM

[9138] Anonymous. A sermon preached by the microscope. Scientific American, 35(15):225, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
225.pdf.

Anonymous:1876:YFE

[9139] Anonymous. The yellow fever epidemic. Scientific American, 35(15):225, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
225c.pdf.

Anonymous:1876:WSS

[9140] Anonymous. Why is the sea salt? Scientific American, 35(15):225, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-225a.pdf.

Anonymous:1876:SCCb

[9141] Anonymous. The spontaneous combustion of coal at sea. Scientific American, 35(15):225, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
225b.pdf.

Anonymous:1876:CTF

[9142] Anonymous. Car truck frame drilling machine. Scientific American, 35(15):226, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
226.pdf.

Anonymous:1876:NG

[9143] Anonymous. A new gearing. Scientific American, 35(15):226, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-226a.pdf.

Anonymous:1876:SHG

[9144] Anonymous. Success of the Hell Gate explosion. Scientific American, 35(15):226, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
226b.pdf.

Anonymous:1876:MI

[9145] Anonymous. Mediæval ironwork. Scientific American, 35(15):227, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-227.pdf.

Anonymous:1876:OO

[9146] Anonymous. An ornithological ornament. Scientific American, 35 (15):227–228, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
227a.pdf.

Rose:1876:PMr

[9147] Joshua Rose. Practical mechanism. Scientific American, 35(15): 228–229, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
228a.pdf.

Anonymous:1876:NMMd

[9148] Anonymous. A new mechanical movement. Scientific American, 35(15):230, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
230.pdf.

Anonymous:1876:ILBa

[9149] Anonymous. Improved link block for locomotives. Scientific American, 35(15):230, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
230a.pdf.

Anonymous:1876:ISV

[9150] Anonymous. Improved stop valve. Scientific American, 35(15):230, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-230b.pdf.

Anonymous:1876:VWD

[9151] Anonymous. Vines for winter decoration. Scientific American, 35(15):231, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
231.pdf.

Anonymous:1876:CNd

[9152] Anonymous. Centennial notes. Scientific American, 35(15):232, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-232.pdf.

Anonymous:1876:RNa

[9153] Anonymous. Rambling notes. Scientific American, 35(15):232–233, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-232a.pdf.

Anonymous:1876:FAI

[9154] Anonymous. The fair of the American institute. Scientific American, 35(15):233, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
233.pdf.

Anonymous:1876:DCx

[9155] Anonymous. Decisions of the courts. Scientific American, 35(15):233, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-233a.pdf.

Anonymous:1876:DCPa

[9156] Anonymous. Decisions of the Commissioner of Patents. Scientific American, 35(15):233, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
233b.pdf.

Anonymous:1876:RAFbp

[9157] Anonymous. Recent American and foreign patents. Scientific American, 35(15):233–234, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
233c.pdf.

Anonymous:1876:NHIl

[9158] Anonymous. New household inventions. Scientific American, 35 (15):234, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
234.pdf.

Anonymous:1876:NAIbj

[9159] Anonymous. New agricultural inventions. Scientific American, 35(15):234, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
234a.pdf.

Anonymous:1876:NWHbd

[9160] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(15):234, October 7, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n15/pdf/scientificamerican10071876-234b.pdf.

Anonymous:1876:NMIb

[9161] Anonymous. New miscellaneous inventions. Scientific American, 35(15):234, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
234c.pdf.

Anonymous:1876:CRbp

[9162] Anonymous. Communications received. Scientific American, 35 (15):236, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n15/pdf/scientificamerican10071876-
236.pdf.

Anonymous:1876:VPl

[9163] Anonymous. The validity of patents. Scientific American, 35(15):237, October 7, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n15/pdf/scientificamerican10071876-237.pdf.

Anonymous:1876:Cbv

[9164] Anonymous. Correspondence. Scientific American, 35(16):244, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-244.pdf.

Anonymous:1876:BNBv

[9165] Anonymous. Books: New books and publications. Scientific American, 35(16):250, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
250.pdf.

Anonymous:1876:BPbq

[9166] Anonymous. Business and personal. Scientific American, 35 (16):251, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
251.pdf.

Anonymous:1876:NQbq

[9167] Anonymous. Notes and queries. Scientific American, 35(16):251, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-251a.pdf.

Anonymous:1876:IIbr 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
251c.pdf.

Anonymous:1876:TSF

[9169] Anonymous. Trial of steak fire engines at the Centennial Exhibition. Scientific American, 35(16):239, October 14, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n16/pdf/scientificamerican10141876-239.pdf.

Anonymous:1876:WM

[9170] Anonymous. What is materialism? Scientific American, 35(16):240, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-240.pdf.

Anonymous:1876:IEU

[9171] Anonymous. Infusorial Earth and its uses. Scientific American, 35(16):240, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
240a.pdf.

Anonymous:1876:HGBb

[9172] Anonymous. The Hell Gate blast. Scientific American, 35(16): 240–241, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
240b.pdf.

Anonymous:1876:CAa

[9173] Anonymous. The Centennial Awards. Scientific American, 35 (16):241, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
241.pdf.

Anonymous:1876:CSD

[9174] Anonymous. A cigar scientifically dissected. Scientific American, 35(16):241, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
241a.pdf.

Anonymous:1876:IBBb

[9175] Anonymous. Improved building blocks. Scientific American, 35 (16):242, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
242.pdf.

Anonymous:1876:PLT

[9176] Anonymous. A pocket lung tester. Scientific American, 35(16):242, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-242a.pdf.

Anonymous:1876:LPCb

[9177] Anonymous. A life-preserving cap and cape. Scientific American, 35(16):242, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
242b.pdf.

Anonymous:1876:AUI

[9178] Anonymous. Apparatus used in industrial processes. Scientific American, 35(16):243–244, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
243.pdf.

Rose:1876:PMs

[9179] Joshua Rose. Practical mechanism. Scientific American, 35(16): 244–245, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
244a.pdf.

Anonymous:1876:ICLb

[9180] Anonymous. Improved crosshead for locomotives. Scientific American, 35(16):246, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
246.pdf.

Anonymous:1876:IAPd

[9181] Anonymous. Improved air pump. Scientific American, 35(16):246, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-246a.pdf.

Anonymous:1876:IBE

[9182] Anonymous. Improved bucket ear. Scientific American, 35(16):246, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-246b.pdf.

Anonymous:1876:SG

[9183] Anonymous. The sea gull. Scientific American, 35(16):247, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-247.pdf.

Anonymous:1876:CNe

[9184] Anonymous. Centennial notes. Scientific American, 35(16):248, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-248.pdf.

Anonymous:1876:RNb

[9185] Anonymous. Rambling notes. Scientific American, 35(16):248–249, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-248a.pdf.

Anonymous:1876:G

[9186] Anonymous. Glycerin. Scientific American, 35(16):249, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-249.pdf.

Anonymous:1876:RAFbq

[9187] Anonymous. Recent American and foreign patents. Scientific American, 35(16):250, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
250a.pdf.

Anonymous:1876:NWHbe

[9188] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(16):250, October 14, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n16/pdf/scientificamerican10141876-250b.pdf.

Anonymous:1876:NHIm

[9189] Anonymous. New household inventions. Scientific American, 35 (16):250, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
250c.pdf.

Anonymous:1876:NMIc

[9190] Anonymous. New miscellaneous inventions. Scientific American, 35(16):250, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
250d.pdf.

Anonymous:1876:NAIbk

[9191] Anonymous. New agricultural inventions. Scientific American, 35(16):250, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
250e.pdf.

Anonymous:1876:CRbq

[9192] Anonymous. Communications received. Scientific American, 35 (16):251, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
251b.pdf.

Anonymous:1876:VPm

[9193] Anonymous. Value of patents. Scientific American, 35(16):252, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n16/pdf/scientificamerican10141876-252.pdf.

Anonymous:1876:VPn

[9194] Anonymous. The validity of patents. Scientific American, 35 (16):252, October 14, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n16/pdf/scientificamerican10141876-
252a.pdf.

Anonymous:1876:Cbw

[9195] Anonymous. Correspondence. Scientific American, 35(17):259–260, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n17/pdf/scientificamerican10211876-259a.pdf.

Anonymous:1876:BNBw

[9196] Anonymous. Books: New books and publications. Scientific American, 35(17):266–267, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
266b.pdf.

Anonymous:1876:BPbr

[9197] Anonymous. Business and personal. Scientific American, 35 (17):268, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
268.pdf.

Anonymous:1876:NQbr

[9198] Anonymous. Notes and queries. Scientific American, 35(17):268, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n17/pdf/scientificamerican10211876-268a.pdf.

Anonymous:1876:IIbs

[9199] Anonymous. Index of inventions. Scientific American, 35(17): 268–269, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
268c.pdf.

Anonymous:1876:NLD

[9200] Anonymous. A new light draft propeller. Scientific American, 35(17):255, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
255.pdf.

Anonymous:1876:PHM

[9201] Anonymous. Professor Huxley’s mistake. Scientific American, 35 (17):256, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
256.pdf.

Anonymous:1876:ACP

[9202] Anonymous. The advantages of cheap patents. Scientific American, 35(17):256–257, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
256a.pdf.

Anonymous:1876:SSc

[9203] Anonymous. The spiritual slate. Scientific American, 35(17):257, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n17/pdf/scientificamerican10211876-257.pdf.

Anonymous:1876:IMPd

[9204] Anonymous. The inter-Mercurial planet. Scientific American, 35 (17):257, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
257a.pdf.

Anonymous:1876:ILCb

[9205] Anonymous. Improved locomotive and car steam brake. Scientific American, 35(17):258, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
258.pdf.

Anonymous:1876:INE

[9206] Anonymous. Improved nail extractor. Scientific American, 35 (17):258, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
258a.pdf.

Anonymous:1876:NFFb

[9207] Anonymous. A new fodder fork. Scientific American, 35(17):258, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n17/pdf/scientificamerican10211876-258b.pdf.

Anonymous:1876:SAT

[9208] Anonymous. Singular accident to a trout. Scientific American, 35(17):259, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
259.pdf.

Robinson:1876:TWW

[9209] S. W. Robinson. Turbine water wheels. Scientific American, 35 (17):260, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
260.pdf.

Anonymous:1876:BK

[9210] Anonymous. The black knot. Scientific American, 35(17):260–261, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n17/pdf/scientificamerican10211876-260a.pdf.

Anonymous:1876:LNS

[9211] Anonymous. The latest news from the Sun. Scientific American, 35(17):261, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
261.pdf.

Anonymous:1876:TBPb

[9212] Anonymous. A theory of the ball puzzle. Scientific American, 35(17):262, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
262.pdf.

Anonymous:1876:MSS

[9213] Anonymous. The “model” scroll saw. Scientific American, 35 (17):262, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
262a.pdf.

Anonymous:1876:ROM

[9214] Anonymous. Rubber overshoe making at the Centennial. Scientific American, 35(17):262–263, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
262b.pdf.

Anonymous:1876:SBP

[9215] Anonymous. Seasonal behavior of plants. Scientific American, 35(17):264, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
264.pdf.

Anonymous:1876:BG

[9216] Anonymous. The blessings of gunpowder. Scientific American, 35(17):264, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
264a.pdf.

Anonymous:1876:PSM

[9217] Anonymous. The processes of steel manufacture. Scientific American, 35(17):265, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
265.pdf.

Anonymous:1876:RST

[9218] Anonymous. The Russian system of trade education. Scientific American, 35(17):265–266, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
265a.pdf.

Anonymous:1876:PIb

[9219] Anonymous. The profitableness of irrigation. Scientific American, 35(17):266, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
266.pdf.

Anonymous:1876:CCI

[9220] Anonymous. Comparative cost of illumination. Scientific American, 35(17):266, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
266a.pdf.

Anonymous:1876:RAFbr

[9221] Anonymous. Recent American and foreign patents. Scientific American, 35(17):267, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
267.pdf.

Anonymous:1876:NMEu

[9222] Anonymous. New mechanical and engineering inventions. Scientific American, 35(17):267, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
267a.pdf.

Anonymous:1876:NWHbf

[9223] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(17):267, October 21, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n17/pdf/scientificamerican10211876-267b.pdf.

Anonymous:1876:NHIn

[9224] Anonymous. New household inventions. Scientific American, 35 (17):267, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
267c.pdf.

Anonymous:1876:NMId

[9225] Anonymous. New miscellaneous inventions. Scientific American, 35(17):267, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
267d.pdf.

Anonymous:1876:CRbr

[9226] Anonymous. Communications received. Scientific American, 35 (17):268, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
268b.pdf.

Anonymous:1876:VPo

[9227] Anonymous. The validity of patents. Scientific American, 35 (17):269, October 21, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n17/pdf/scientificamerican10211876-
269.pdf.

Anonymous:1876:Cbx

[9228] Anonymous. Correspondence. Scientific American, 35(18):276, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-276.pdf.

Anonymous:1876:BNBx

[9229] Anonymous. Books: New books and publications. Scientific American, 35(18):281, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
281a.pdf.

Anonymous:1876:BPbs

[9230] Anonymous. Business and personal. Scientific American, 35 (18):283, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
283.pdf.

Anonymous:1876:NQbs

[9231] Anonymous. Notes and queries. Scientific American, 35(18): 283–284, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
283a.pdf.

Anonymous:1876:IIbt

[9232] Anonymous. Index of inventions. Scientific American, 35(18):284, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-284a.pdf.

Anonymous:1876:PNPv

[9233] Anonymous. Patent notes: Patents. Scientific American, 35 (18):286, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
286.pdf.

Anonymous:1876:IMSc

[9234] Anonymous. Improved machine for stringing tobacco leaves. Scientific American, 35(18):271, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
271.pdf.

Anonymous:1876:IMAa

[9235] Anonymous. Improved milling attachment for lathes. Scientific American, 35(18):271, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
271a.pdf.

Anonymous:1876:SEA

[9236] Anonymous. Some extinct American animals. Scientific American, 35(18):272, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
272.pdf.

Anonymous:1876:PNG

[9237] Anonymous. Prospect of new German patent laws. Scientific American, 35(18):272–273, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
272a.pdf.

Anonymous:1876:PMR

[9238] Anonymous. Progress of the Mississippi River jetties. Scientific American, 35(18):273, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
273.pdf.

Anonymous:1876:SES

[9239] Anonymous. Steam engine slide valves. Scientific American, 35 (18):273, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
273a.pdf.

Anonymous:1876:MCA

[9240] Anonymous. More Centennial Awards. Scientific American, 35 (18):273, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
273b.pdf.

Anonymous:1876:MBBa

[9241] Anonymous. Musical building blocks. Scientific American, 35 (18):274, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
274.pdf.

Anonymous:1876:IBH

[9242] Anonymous. Improved bale hoop tightener. Scientific American, 35(18):274, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
274a.pdf.

Anonymous:1876:IYR

[9243] Anonymous. Improved yarn reel and tester. Scientific American, 35(18):275, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
275.pdf.

Anonymous:1876:SWS

[9244] Anonymous. Sample-weighing scale. Scientific American, 35(18): 275, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
275a.pdf.

Anonymous:1876:PGH

[9245] Anonymous. Patent gas heater. Scientific American, 35(18):275, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-275b.pdf.

Anonymous:1876:FJM

[9246] Anonymous. Fishhooks, jetties, and miscellaneous devices. Scientific American, 35(18):275–276, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
275c.pdf.

Anonymous:1876:LMT

[9247] Anonymous. A lesson from the machine tools at the Centennial. Scientific American, 35(18):276–277, October 28, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n18/pdf/scientificamerican10281876-276a.pdf.

Rose:1876:PMt

[9248] Joshua Rose. Practical mechanism. Scientific American, 35(18):277, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-277.pdf.

Anonymous:1876:IEI

[9249] Anonymous. Improved engine indicator. Scientific American, 35 (18):278, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
278.pdf.

Anonymous:1876:ILR

[9250] Anonymous. Improved leveling rod. Scientific American, 35(18): 278, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
278a.pdf.

Anonymous:1876:WFB

[9251] Anonymous. A white-flowered bramble. Scientific American, 35 (18):279, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
279.pdf.

Anonymous:1876:CPl

[9252] Anonymous. The coca plant. Scientific American, 35(18):279–280, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-279a.pdf.

Anonymous:1876:CNf

[9253] Anonymous. Centennial notes. Scientific American, 35(18):280, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n18/pdf/scientificamerican10281876-280.pdf.

Anonymous:1876:CWC

[9254] Anonymous. Carriage wheels at the Centennial Exhibition. Scientific American, 35(18):280–281, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
280a.pdf.

Anonymous:1876:NAIbl

[9255] Anonymous. Notes on the American Institute Fair. Scientific American, 35(18):281, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
281.pdf.

Anonymous:1876:RAFbs

[9256] Anonymous. Recent American and foreign patents. Scientific American, 35(18):281–282, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
281b.pdf.

Anonymous:1876:NAIbm

[9257] Anonymous. New agricultural inventions. Scientific American, 35(18):282, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
282.pdf.

Anonymous:1876:NMIe

[9258] Anonymous. New miscellaneous inventions. Scientific American, 35(18):282, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
282a.pdf.

Anonymous:1876:NHIo

[9259] Anonymous. New household inventions. Scientific American, 35 (18):282, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
282b.pdf.

Anonymous:1876:NWHbg

[9260] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(18):282, October 28, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n18/pdf/scientificamerican10281876-282c.pdf.

Anonymous:1876:CRbs

[9261] Anonymous. Communications received. Scientific American, 35 (18):284, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
284.pdf.

Anonymous:1876:VPp

[9262] Anonymous. The validity of patents. Scientific American, 35 (18):284, October 28, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n18/pdf/scientificamerican10281876-
284b.pdf.

Anonymous:1876:Cby

[9263] Anonymous. Correspondence. Scientific American, 35(19):292, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n19/pdf/scientificamerican11041876-292.pdf.

Anonymous:1876:BPbt

[9264] Anonymous. Business and personal. Scientific American, 35 (19):299, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
299.pdf.

Anonymous:1876:NQbt

[9265] Anonymous. Notes and queries. Scientific American, 35(19): 299–300, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
299a.pdf.

Anonymous:1876:IIbu

[9266] Anonymous. Index of inventions. Scientific American, 35(19): 300, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
300a.pdf.

Anonymous:1876:PNPw

[9267] Anonymous. Patent notes: Patents. Scientific American, 35 (19):301, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
301.pdf.

Anonymous:1876:CBCa

[9268] Anonymous. Curious boats at the Centennial Exposition. Scientific American, 35(19):287, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
287.pdf.

Anonymous:1876:DTV

[9269] Anonymous. Dr. Taylor vs. Professor Huxley. Scientific American, 35(19):288, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
288.pdf.

Anonymous:1876:ISWc

[9270] Anonymous. Iron and steel working improvements. Scientific American, 35(19):288, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
288a.pdf.

Anonymous:1876:STLb

[9271] Anonymous. Some thoughts on labor. Scientific American, 35 (19):288–289, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
288b.pdf.

Anonymous:1876:SFR

[9272] Anonymous. Spiritualists’ facts and real facts. Scientific American, 35(19):289, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
289.pdf.

Anonymous:1876:NCTc

[9273] Anonymous. New chemical test for alcohol. Scientific American, 35(19):289, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
289a.pdf.

Anonymous:1876:PCEb

[9274] Anonymous. Petroleum changed into electricity and light. Scientific American, 35(19):289, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
289b.pdf.

Anonymous:1876:MNAa

[9275] Anonymous. The meeting of the National Academy of Sciences. Scientific American, 35(19):289, November 4, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n19/pdf/scientificamerican11041876-289c.pdf.

Anonymous:1876:NWE

[9276] Anonymous. A new water engine. Scientific American, 35(19): 290, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
290.pdf.

Anonymous:1876:ISSd

[9277] Anonymous. Improved stopcock and street washer box. Scientific American, 35(19):290, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
290a.pdf.

Anonymous:1876:NSC

[9278] Anonymous. New screw-cutting lathe. Scientific American, 35 (19):291, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
291.pdf.

Anonymous:1876:IWC

[9279] Anonymous. Improved wood carpeting and inlaid floors. Scientific American, 35(19):291, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
291a.pdf.

Anonymous:1876:CSL

[9280] Anonymous. Cutting speeds for lathe work. Scientific American, 35 (19):292–293, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
292a.pdf.

Anonymous:1876:CCSb

[9281] Anonymous. Combination chair and secretary. Scientific American, 35(19):294, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
294.pdf.

Anonymous:1876:RHC

[9282] Anonymous. R. Hoe & Co.’s improved inserted saw teeth. Scientific American, 35(19):294, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
294a.pdf.

Anonymous:1876:HBb

[9283] Anonymous. The hemiptera or bugs. Scientific American, 35 (19):295, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
295.pdf.

Anonymous:1876:B

[9284] Anonymous. The burbot. Scientific American, 35(19):295, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n19/pdf/scientificamerican11041876-295a.pdf.

Anonymous:1876:MCc

[9285] Anonymous. Minerals at the Centennial. Scientific American, 35 (19):296–297, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
296.pdf.

Anonymous:1876:CNg

[9286] Anonymous. Centennial notes. Scientific American, 35(19):297, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n19/pdf/scientificamerican11041876-297.pdf.

Anonymous:1876:ANh

[9287] Anonymous. Astronomical notes. Scientific American, 35(19): 297–298, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
297a.pdf.

Anonymous:1876:RAFbt

[9288] Anonymous. Recent American and foreign patents. Scientific American, 35(19):298, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
298.pdf.

Anonymous:1876:NMIf

[9289] Anonymous. New miscellaneous inventions. Scientific American, 35(19):298, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
298a.pdf.

Anonymous:1876:NHIp

[9290] Anonymous. New household inventions. Scientific American, 35 (19):298, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
298b.pdf.

Anonymous:1876:NWHbh

[9291] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(19):298, November 4, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n19/pdf/scientificamerican11041876-298c.pdf.

Anonymous:1876:NAIbn

[9292] Anonymous. New agricultural inventions. Scientific American, 35 (19):298–299, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
298d.pdf.

Anonymous:1876:CRbt

[9293] Anonymous. Communications received. Scientific American, 35 (19):300, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
300.pdf.

Anonymous:1876:VPq

[9294] Anonymous. The validity of patents. Scientific American, 35 (19):300, November 4, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n19/pdf/scientificamerican11041876-
300b.pdf.

Anonymous:1876:Cbz

[9295] Anonymous. Correspondence. Scientific American, 35(20):308, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n20/pdf/scientificamerican11111876-308a.pdf.

Anonymous:1876:BNBy

[9296] Anonymous. Books: New books and publications. Scientific American, 35(20):313, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
313b.pdf.

Anonymous:1876:BPbu

[9297] Anonymous. Business and personal. Scientific American, 35 (20):315, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
315.pdf.

Anonymous:1876:NQbu

[9298] Anonymous. Notes and queries. Scientific American, 35(20): 315–316, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
315a.pdf.

Anonymous:1876:IIbv

[9299] Anonymous. Index of inventions. Scientific American, 35(20): 316, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
316a.pdf.

Anonymous:1876:GMC

[9300] Anonymous. Glass making at the Centennial Exposition. Scientific American, 35(20):303, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
303.pdf.

Anonymous:1876:UD

[9301] Anonymous. Ultra-Darwinism. Scientific American, 35(20):304, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n20/pdf/scientificamerican11111876-304.pdf.

Anonymous:1876:WIV

[9302] Anonymous. Was it vulcan? Scientific American, 35(20):304– 305, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
304a.pdf.

Anonymous:1876:CAb

[9303] Anonymous. The Centennial Awards. Scientific American, 35 (20):305, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
305.pdf.

Anonymous:1876:RBA

[9304] Anonymous. The return of the British Arctic Expedition. Scientific American, 35(20):305, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
305a.pdf.

Anonymous:1876:NYAe

[9305] Anonymous. The New York aquarium. Scientific American, 35 (20):305, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
305b.pdf.

Anonymous:1876:DT

[9306] Anonymous. Detoxicated tobacco. Scientific American, 35(20): 305, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
305c.pdf.

Anonymous:1876:TAC

[9307] Anonymous. Thermostatic apparatus for controlling valves. Scientific American, 35(20):306, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
306.pdf.

Anonymous:1876:NFBb

[9308] Anonymous. A new floue bolt. Scientific American, 35(20):306, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n20/pdf/scientificamerican11111876-306a.pdf.

Anonymous:1876:CSG

[9309] Anonymous. Compensating steam generating apparatus. Scientific American, 35(20):307, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
307.pdf.

Anonymous:1876:ISH

[9310] Anonymous. Improved siphon head for bottles. etc. Scientific American, 35(20):307, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
307a.pdf.

Anonymous:1876:MNAb

[9311] Anonymous. The meeting of the National Academy of Sciences. Scientific American, 35(20):308, November 11, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n20/pdf/scientificamerican11111876-308.pdf.

Anonymous:1876:CLH

[9312] Anonymous. Cutting a left hand screw with a right hand tap. Scientific American, 35(20):308–309, November 11, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n20/pdf/scientificamerican11111876-308b.pdf.

Rose:1876:PMu

[9313] Joshua Rose. Practical mechanism. Scientific American, 35(20): 309, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
309.pdf.

Anonymous:1876:IRB

[9314] Anonymous. Improved ribbed bale tie. Scientific American, 35 (20):310, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
310.pdf.

Anonymous:1876:CEMa

[9315] Anonymous. The Camacho electro-magnetic engine. Scientific American, 35(20):310, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
310a.pdf.

Anonymous:1876:GCC

[9316] Anonymous. The grindstone column at the Centennial. Scientific American, 35(20):311, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
311.pdf.

Anonymous:1876:FL

[9317] Anonymous. Flying lizards. Scientific American, 35(20):311, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n20/pdf/scientificamerican11111876-311a.pdf.

Anonymous:1876:HA

[9318] Anonymous. Hydrostatic apparatus. Scientific American, 35(20): 312, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
312.pdf.

Anonymous:1876:CNh

[9319] Anonymous. Centennial notes. Scientific American, 35(20):313, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n20/pdf/scientificamerican11111876-313.pdf.

Anonymous:1876:DCy

[9320] Anonymous. Decisions of the courts. Scientific American, 35 (20):313, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
313a.pdf.

Anonymous:1876:RAFbu

[9321] Anonymous. Recent American and foreign patents. Scientific American, 35(20):314, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
314.pdf.

Anonymous:1876:NWHbi

[9322] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(20):314, November 11, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n20/pdf/scientificamerican11111876-314a.pdf.

Anonymous:1876:NHIq

[9323] Anonymous. New household inventions. Scientific American, 35 (20):314, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
314b.pdf.

Anonymous:1876:NAIbo

[9324] Anonymous. New agricultural inventions. Scientific American, 35 (20):314, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
314c.pdf.

Anonymous:1876:NMEv

[9325] Anonymous. New mechanical and engineering inventions. Scientific American, 35(20):314, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
314d.pdf.

Anonymous:1876:CRbu

[9326] Anonymous. Communications received. Scientific American, 35 (20):316, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
316.pdf.

Anonymous:1876:VPr

[9327] Anonymous. The validity of patents. Scientific American, 35 (20):316, November 11, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n20/pdf/scientificamerican11111876-
316b.pdf.

Anonymous:1876:Cch

[9328] Anonymous. Correspondence. Scientific American, 35(21):324, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n21/pdf/scientificamerican11181876-324.pdf.

Anonymous:1876:BNBz

[9329] Anonymous. Books: New books and publications. Scientific American, 35(21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330.pdf.

Anonymous:1876:BPbv

[9330] Anonymous. Business and personal. Scientific American, 35 (21):331, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
331.pdf.

Anonymous:1876:NQbv

[9331] Anonymous. Notes and queries. Scientific American, 35(21): 331–332, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
331a.pdf.

Anonymous:1876:IPS

[9332] Anonymous. Improved portable Sawmill. Scientific American, 35 (21):319, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
319.pdf.

Anonymous:1876:OEH

[9333] Anonymous. The ocean echo — Henry vs Tyndall again. Scientific American, 35(21):320, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
320.pdf.

Anonymous:1876:SSd

[9334] Anonymous. Slade sustained. Scientific American, 35(21):320, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n21/pdf/scientificamerican11181876-320a.pdf.

Anonymous:1876:LCb

[9335] Anonymous. The Leverrier of chemistry. Scientific American, 35 (21):320–321, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
320b.pdf.

Anonymous:1876:PSH

[9336] Anonymous. Panics in school houses. Scientific American, 35 (21):321, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
321.pdf.

Anonymous:1876:VAb

[9337] Anonymous. Vulcan again. Scientific American, 35(21):321, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n21/pdf/scientificamerican11181876-321a.pdf.

Anonymous:1876:IEMb

[9338] Anonymous. Irregularity of the Earth’s motion. Scientific American, 35(21):321, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
321b.pdf.

Anonymous:1876:ISSe

[9339] Anonymous. Improved shoe-scolloping machine. Scientific American, 35(21):322, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
322.pdf.

Anonymous:1876:CCj

[9340] Anonymous. Curiosities of the Centennial — I. Scientific American, 35(21):323–324, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
323.pdf.

Anonymous:1876:MDL

[9341] Anonymous. The merits and demerits of linnæus. Scientific American, 35(21):324–325, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
324a.pdf.

Anonymous:1876:NIV

[9342] Anonymous. New ice velocipede. Scientific American, 35(21): 325, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
325.pdf.

Anonymous:1876:NMSb

[9343] Anonymous. New method of setting hair triggers of rifles. Scientific American, 35(21):325, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
325a.pdf.

Anonymous:1876:ISSf

[9344] Anonymous. Improved sap spout. Scientific American, 35(21): 325, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
325b.pdf.

Anonymous:1876:ITD

[9345] Anonymous. Improved twist drill and tool-grinding machine. Scientific American, 35(21):326, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
326.pdf.

Anonymous:1876:ICD

[9346] Anonymous. Improved combination desk and book case. Scientific American, 35(21):326, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
326a.pdf.

Anonymous:1876:MBBb

[9347] Anonymous. A museum of British birds. Scientific American, 35 (21):327–328, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
327.pdf.

Morton:1876:GM

[9348] Henry Morton. The gas microscope. Scientific American, 35 (21):328, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
328.pdf.

Anonymous:1876:CNi

[9349] Anonymous. Centennial notes. Scientific American, 35(21):329– 330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
329.pdf.

Anonymous:1876:RAFbv

[9350] Anonymous. Recent American and foreign patents. Scientific American, 35(21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330a.pdf.

Anonymous:1876:NHIr

[9351] Anonymous. New household inventions. Scientific American, 35 (21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330b.pdf.

Anonymous:1876:NAIbp

[9352] Anonymous. New agricultural inventions. Scientific American, 35 (21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330c.pdf.

Anonymous:1876:NWHbj

[9353] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(21):330, November 18, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n21/pdf/scientificamerican11181876-330d.pdf.

Anonymous:1876:NTMn

[9354] Anonymous. New textile machinery. Scientific American, 35 (21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330e.pdf.

Anonymous:1876:NMIg

[9355] Anonymous. New miscellaneous inventions. Scientific American, 35(21):330, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
330f.pdf.

Anonymous:1876:CRbv

[9356] Anonymous. Communications received. Scientific American, 35 (21):332, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
332.pdf.

Anonymous:1876:VPHa

[9357] Anonymous. Value of patents and how to obtain them. Scientific American, 35(21):332, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
332a.pdf.

Anonymous:1876:VPs

[9358] Anonymous. The validity of patents. Scientific American, 35 (21):332, November 18, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n21/pdf/scientificamerican11181876-
332b.pdf.

Anonymous:1876:Cck

[9359] Anonymous. Correspondence. Scientific American, 35(22):340– 341, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
340.pdf.

Anonymous:1876:BPbw

[9360] Anonymous. Business and personal. Scientific American, 35 (22):347, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
347.pdf.

Anonymous:1876:NQbw

[9361] Anonymous. Notes and queries. Scientific American, 35(22): 347, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
347a.pdf.

Anonymous:1876:IIbw

[9362] Anonymous. Index of inventions. Scientific American, 35(22): 347–348, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
347c.pdf.

Anonymous:1876:PNPx

[9363] Anonymous. Patent notes: Patents. Scientific American, 35 (22):350, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
350.pdf.

Anonymous:1876:JFC

[9364] Anonymous. J. A. Fay & Co.’s exhibit of woodworking machinery at the Centennial. Scientific American, 35(22):335, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n22/pdf/scientificamerican11251876-335.pdf.

Anonymous:1876:CCEa

[9365] Anonymous. The close of the Centennial Exposition. Scientific American, 35(22):336, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
336.pdf.

Anonymous:1876:IRT

[9366] Anonymous. Intermittent records and their interpretation. Scientific American, 35(22):336–337, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
336a.pdf.

Anonymous:1876:Sb

[9367] Anonymous. The stereoscope. Scientific American, 35(22):337, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n22/pdf/scientificamerican11251876-337.pdf.

Anonymous:1876:ABa

[9368] Anonymous. Artificial butter. Scientific American, 35(22):337, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n22/pdf/scientificamerican11251876-337a.pdf.

Anonymous:1876:SBVa

[9369] Anonymous. The salt bluffs of Virgin River, Nevada. Scientific American, 35(22):337, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
337b.pdf.

Anonymous:1876:IFB

[9370] Anonymous. Improved fireproof and burglarproof safe. Scientific American, 35(22):338, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
338.pdf.

Anonymous:1876:TDA

[9371] Anonymous. Time drop attachment for alarm clocks. Scientific American, 35(22):338, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
338a.pdf.

Anonymous:1876:IBg

[9372] Anonymous. Improved bevel. Scientific American, 35(22):338, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n22/pdf/scientificamerican11251876-338b.pdf.

Anonymous:1876:NI

[9373] Anonymous. A new irrigator. Scientific American, 35(22):338, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n22/pdf/scientificamerican11251876-338c.pdf.

Anonymous:1876:LRb

[9374] Anonymous. Locomotives and railways. Scientific American, 35 (22):339–340, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
339.pdf.

Rose:1876:PMv

[9375] Joshua Rose. Practical mechanism. Scientific American, 35(22): 341, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
341.pdf.

Anonymous:1876:RTL

[9376] Anonymous. Rapid transit locomotive. Scientific American, 35 (22):342, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
342.pdf.

Anonymous:1876:CGL

[9377] Anonymous. Counter gear for lathes. Scientific American, 35 (22):342, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
342a.pdf.

Anonymous:1876:COP

[9378] Anonymous. The castor oil plant as a tree. Scientific American, 35(22):343, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
343.pdf.

Anonymous:1876:BOA

[9379] Anonymous. Bixa Orellana — Annotta. Scientific American, 35 (22):343, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
343a.pdf.

Anonymous:1876:CNj

[9380] Anonymous. Centennial notes. Scientific American, 35(22):345– 346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
345.pdf.

Anonymous:1876:RAFbw

[9381] Anonymous. Recent American and foreign patents. Scientific American, 35(22):346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
346.pdf.

Anonymous:1876:NAIbq

[9382] Anonymous. New agricultural inventions. Scientific American, 35 (22):346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
346a.pdf.

Anonymous:1876:NMEw

[9383] Anonymous. New mechanical and engineering inventions. Scientific American, 35(22):346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
346b.pdf.

Anonymous:1876:NMIh

[9384] Anonymous. New miscellaneous inventions. Scientific American, 35(22):346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
346c.pdf.

Anonymous:1876:NHIs

[9385] Anonymous. New household inventions. Scientific American, 35 (22):346, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
346d.pdf.

Anonymous:1876:CRbw

[9386] Anonymous. Communications received. Scientific American, 35 (22):347, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
347b.pdf.

Anonymous:1876:VPt

[9387] Anonymous. The validity of patents. Scientific American, 35 (22):348, November 25, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n22/pdf/scientificamerican11251876-
348.pdf.

Anonymous:1876:Ccl

[9388] Anonymous. Correspondence. Scientific American, 35(23):356, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n23/pdf/scientificamerican12021876-356.pdf.

Anonymous:1876:BNBba

[9389] Anonymous. Books: New books and publications. Scientific American, 35(23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361.pdf.

Anonymous:1876:BPbx

[9390] Anonymous. Business and personal. Scientific American, 35 (23):362, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
362.pdf.

Anonymous:1876:NQbx

[9391] Anonymous. Notes and queries. Scientific American, 35(23): 362–363, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
362a.pdf.

Anonymous:1876:IIbx

[9392] Anonymous. Index of inventions. Scientific American, 35(23): 363–364, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
363a.pdf.

Anonymous:1876:PNPy

[9393] Anonymous. Patent notes: Patents. Scientific American, 35 (23):366, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
366.pdf.

Anonymous:1876:IBCa

[9394] Anonymous. Improved boiler cleaner and mud extractor. Scientific American, 35(23):351, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
351.pdf.

Anonymous:1876:PMCb

[9395] Anonymous. The Putnam machine Company’s steam engine. Scientific American, 35(23):351, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
351a.pdf.

Anonymous:1876:PCd

[9396] Anonymous. Publishers’ card. Scientific American, 35(23):352, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n23/pdf/scientificamerican12021876-352.pdf.

Anonymous:1876:PAb

[9397] Anonymous. Prospects of aeronautics. Scientific American, 35 (23):352, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
352a.pdf.

Anonymous:1876:EDF

[9398] Anonymous. English dealings with food adulterations. Scientific American, 35(23):352–353, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
352b.pdf.

Anonymous:1876:WNY

[9399] Anonymous. What New York might do with the great French statue. Scientific American, 35(23):353, December 2, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n23/pdf/scientificamerican12021876-353.pdf.

Anonymous:1876:Ad

[9400] Anonymous. Ammoniumnitrodiphenylamin. Scientific American, 35(23):353, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
353a.pdf.

Anonymous:1876:SHD

[9401] Anonymous. Silence and haste at dinner. Scientific American, 35(23):353, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
353b.pdf.

Anonymous:1876:NMP

[9402] Anonymous. New machine for printing colored paper. Scientific American, 35(23):354, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
354.pdf.

Anonymous:1876:NSGb

[9403] Anonymous. A new steam governor. Scientific American, 35 (23):354, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
354a.pdf.

Anonymous:1876:ISGf

[9404] Anonymous. Improved steam gage. Scientific American, 35(23):354, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n23/pdf/scientificamerican12021876-354b.pdf.

Anonymous:1876:CCn

[9405] Anonymous. Curiosities of the Centennial. Scientific American, 35 (23):355–356, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
355.pdf.

Anonymous:1876:PRE

[9406] Anonymous. The practical results of the English Arctic exploring expedition. Scientific American, 35(23):357, December 2, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n23/pdf/scientificamerican12021876-357.pdf.

Anonymous:1876:IFWb

[9407] Anonymous. Improved feed water heater. Scientific American, 35(23):358, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
358.pdf.

Anonymous:1876:IFF

[9408] Anonymous. Improved furnace feeder. Scientific American, 35 (23):358, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
358a.pdf.

Anonymous:1876:SSe

[9409] Anonymous. Sea serpents. Scientific American, 35(23):359, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n23/pdf/scientificamerican12021876-359.pdf.

Anonymous:1876:PRA

[9410] Anonymous. Prehistoric relics in Arizona. Scientific American, 35(23):359, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
359a.pdf.

Anonymous:1876:CNk

[9411] Anonymous. Centennial notes. Scientific American, 35(23):360– 361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
360.pdf.

Anonymous:1876:RAFbx

[9412] Anonymous. Recent American and foreign patents. Scientific American, 35(23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361a.pdf.

Anonymous:1876:NHIt

[9413] Anonymous. New household inventions. Scientific American, 35 (23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361b.pdf.

Anonymous:1876:NWHbk

[9414] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(23):361, December 2, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n23/pdf/scientificamerican12021876-361c.pdf.

Anonymous:1876:NAIbr

[9415] Anonymous. New agricultural inventions. Scientific American, 35 (23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361d.pdf.

Anonymous:1876:NTMo

[9416] Anonymous. New textile machinery. Scientific American, 35 (23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361e.pdf.

Anonymous:1876:NMIi

[9417] Anonymous. New miscellaneous inventions. Scientific American, 35(23):361, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
361f.pdf.

Anonymous:1876:CRbx

[9418] Anonymous. Communications received. Scientific American, 35 (23):363, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
363.pdf.

Anonymous:1876:VPu

[9419] Anonymous. The validity of patents. Scientific American, 35 (23):364, December 2, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n23/pdf/scientificamerican12021876-
364.pdf.

Anonymous:1876:Cco

[9420] Anonymous. Correspondence. Scientific American, 35(24):372–373, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-372a.pdf.

Anonymous:1876:BPby

[9421] Anonymous. Business and personal. Scientific American, 35 (24):379, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
379.pdf.

Anonymous:1876:NQby

[9422] Anonymous. Notes and queries. Scientific American, 35(24): 379–380, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
379a.pdf.

Anonymous:1876:PNPz

[9423] Anonymous. Patent notes: Patents. Scientific American, 35 (24):382, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
382.pdf.

Anonymous:1876:IKG

[9424] Anonymous. Improved knife-grinding machine. Scientific American, 35(24):367, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
367.pdf.

Anonymous:1876:PCf

[9425] Anonymous. Publishers’ card. Scientific American, 35(24):368, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-368.pdf.

Anonymous:1876:CBRa

[9426] Anonymous. Color blindness and its Remedy. Scientific American, 35(24):368, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
368a.pdf.

Anonymous:1876:OR

[9427] Anonymous. Otto of roses. Scientific American, 35(24):368–369, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-368b.pdf.

Anonymous:1876:VDE

[9428] Anonymous. A valuable discovery in electric illumination. Scientific American, 35(24):369, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
369.pdf.

Anonymous:1876:VFW

[9429] Anonymous. The value of fine workmanship. Scientific American, 35(24):369, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
369a.pdf.

Anonymous:1876:NYAf

[9430] Anonymous. New York Academy of Sciences. Scientific American, 35(24):369, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
369b.pdf.

Anonymous:1876:NE

[9431] Anonymous. The Niles engine. Scientific American, 35(24):370, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-370.pdf.

Anonymous:1876:CCp

[9432] Anonymous. A curious clock. Scientific American, 35(24):371, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-371.pdf.

Anonymous:1876:IACc

[9433] Anonymous. Improved apparatus for curing tobacco. Scientific American, 35(24):371, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
371a.pdf.

Anonymous:1876:NSPb

[9434] Anonymous. A new steam pump. Scientific American, 35(24): 371–372, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
371b.pdf.

Riley:1876:AWN

[9435] C. V. Riley. The army worm: Its natural history complete. Scientific American, 35(24):372, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
372.pdf.

Rose:1876:PMw

[9436] Joshua Rose. Practical mechanism. Scientific American, 35(24):373, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-373.pdf.

Anonymous:1876:AGS

[9437] Anonymous. Antique gold and silver work. Scientific American, 35(24):374, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
374.pdf.

Anonymous:1876:SSAa

[9438] Anonymous. Siemens’ sensitive artificial eye. Scientific American, 35(24):374, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
374a.pdf.

Anonymous:1876:NBb

[9439] Anonymous. A new blowpipe. Scientific American, 35(24):374, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-374b.pdf.

Anonymous:1876:APc

[9440] Anonymous. Æchmea Paniculata. Scientific American, 35(24):375, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-375.pdf.

Anonymous:1876:HBc

[9441] Anonymous. House building. Scientific American, 35(24):376–377, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-376.pdf.

Anonymous:1876:CNl

[9442] Anonymous. Centennial notes. Scientific American, 35(24):377, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n24/pdf/scientificamerican12091876-377.pdf.

Anonymous:1876:ANj

[9443] Anonymous. Astronomical notes. Scientific American, 35(24): 377–378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
377a.pdf.

Anonymous:1876:RAFby

[9444] Anonymous. Recent American and foreign patents. Scientific American, 35(24):378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
378.pdf.

Anonymous:1876:NHIu

[9445] Anonymous. New household inventions. Scientific American, 35 (24):378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
378a.pdf.

Anonymous:1876:NTMp

[9446] Anonymous. New textile machinery. Scientific American, 35 (24):378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
378b.pdf.

Anonymous:1876:NWHbl

[9447] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 35(24):378, December 9, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n24/pdf/scientificamerican12091876-378c.pdf.

Anonymous:1876:NMEx

[9448] Anonymous. New mechanical and engineering inventions. Scientific American, 35(24):378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
378d.pdf.

Anonymous:1876:NMIj

[9449] Anonymous. New miscellaneous inventions. Scientific American, 35(24):378, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
378e.pdf.

Anonymous:1876:CRby

[9450] Anonymous. Communications received. Scientific American, 35 (24):380, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
380.pdf.

Anonymous:1876:VPHb

[9451] Anonymous. Value of patents, and how to obtain them. Scientific American, 35(24):380, December 9, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n24/pdf/scientificamerican12091876-
380a.pdf.

Anonymous:1876:Ccq

[9452] Anonymous. Correspondence. Scientific American, 35(25):388– 389, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
388.pdf.

Anonymous:1876:BPbz

[9453] Anonymous. Business and personal. Scientific American, 35 (25):395, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
395.pdf.

Anonymous:1876:NQbz

[9454] Anonymous. Notes and queries. Scientific American, 35(25): 395, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
395a.pdf.

Anonymous:1876:IIby

[9455] Anonymous. Index of inventions. Scientific American, 35(25): 395–396, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
395c.pdf.

Anonymous:1876:PNPba

[9456] Anonymous. Patent notes: Patents. Scientific American, 35 (25):398, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
398.pdf.

Anonymous:1876:NRDb

[9457] Anonymous. New rotary diamond millstone dresser. Scientific American, 35(25):383, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
383.pdf.

Anonymous:1876:PIH

[9458] Anonymous. Price’s improved hose coupling. Scientific American, 35(25):383, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
383a.pdf.

Anonymous:1876:PCg

[9459] Anonymous. Publishers’ card. Scientific American, 35(25):384, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n25/pdf/scientificamerican12161876-384.pdf.

Anonymous:1876:ES

[9460] Anonymous. The extension of sense. Scientific American, 35 (25):384, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
384a.pdf.

Anonymous:1876:GCB

[9461] Anonymous. The great cyclone in Bengal. Scientific American, 35 (25):384–385, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
384b.pdf.

Anonymous:1876:IDS

[9462] Anonymous. An Indian dust storm. Scientific American, 35 (25):385, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
385.pdf.

Anonymous:1876:LFB

[9463] Anonymous. Low flying balloons. Scientific American, 35(25): 385, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
385a.pdf.

Anonymous:1876:RLM

[9464] Anonymous. The radiometer not a light motor. Scientific American, 35(25):385, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
385b.pdf.

Anonymous:1876:LPCc

[9465] Anonymous. The limits to physical culture. Scientific American, 35(25):385, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
385c.pdf.

Anonymous:1876:JBL

[9466] Anonymous. The “John Bull” locomotive at the Centennial. Scientific American, 35(25):386, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
386.pdf.

Anonymous:1876:ILe

[9467] Anonymous. Improved lubricator. Scientific American, 35(25): 386, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
386a.pdf.

Anonymous:1876:GTL

[9468] Anonymous. Galvanometer for testing lightning conductors. Scientific American, 35(25):386, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
386b.pdf.

Anonymous:1876:NT

[9469] Anonymous. A new tackle. Scientific American, 35(25):386, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n25/pdf/scientificamerican12161876-386c.pdf.

Anonymous:1876:ITH

[9470] Anonymous. Improved tool holders and cutters. Scientific American, 35(25):387, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
387.pdf.

Corder:1876:PA

[9471] Octavius Corder. Pharmaceutical apparatus. Scientific American, 35 (25):387–388, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
387a.pdf.

Anonymous:1876:DIWb

[9472] Anonymous. Denney’s improved washing machine. Scientific American, 35(25):390, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
390.pdf.

Anonymous:1876:Sd

[9473] Anonymous. The spirophore. Scientific American, 35(25):390, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n25/pdf/scientificamerican12161876-390a.pdf.

Anonymous:1876:NACb

[9474] Anonymous. A novel air compressor. Scientific American, 35 (25):390, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
390b.pdf.

Anonymous:1876:SHSc

[9475] Anonymous. A sounding hand signal for railroads. Scientific American, 35(25):390, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
390c.pdf.

Anonymous:1876:NCD

[9476] Anonymous. New Chinese deer. Scientific American, 35(25): 391, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
391.pdf.

Anonymous:1876:CFN

[9477] Anonymous. Curious fish at the New York aquarium. Scientific American, 35(25):391, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
391a.pdf.

Riley:1876:LFE

[9478] C. V. Riley. Locust flights east of the Mississippi. Scientific American, 35(25):392, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
392.pdf.

Anonymous:1876:CNm

[9479] Anonymous. Centennial notes. Scientific American, 35(25):393, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n25/pdf/scientificamerican12161876-393.pdf.

Anonymous:1876:OHM

[9480] Anonymous. The occupations and health of the mercantile classes. Scientific American, 35(25):393–394, December 16, 1876. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n25/pdf/scientificamerican12161876-393a.pdf.

Anonymous:1876:RAFbz

[9481] Anonymous. Recent American and foreign patents. Scientific American, 35(25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394.pdf.

Anonymous:1876:NAIbs

[9482] Anonymous. New agricultural inventions. Scientific American, 35 (25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394a.pdf.

Anonymous:1876:NMEy

[9483] Anonymous. New mechanical and engineering inventions. Scientific American, 35(25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394b.pdf.

Anonymous:1876:NMIk

[9484] Anonymous. New miscellaneous inventions. Scientific American, 35(25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394c.pdf.

Anonymous:1876:NHIv

[9485] Anonymous. New household inventions. Scientific American, 35 (25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394d.pdf.

Anonymous:1876:NTMq

[9486] Anonymous. New textile machinery. Scientific American, 35 (25):394, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
394e.pdf.

Anonymous:1876:CRbz

[9487] Anonymous. Communications received. Scientific American, 35 (25):395, December 16, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n25/pdf/scientificamerican12161876-
395b.pdf.

Anonymous:1876:Ccr

[9488] Anonymous. Correspondence. Scientific American, 35(26):404, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n26/pdf/scientificamerican12231876-404a.pdf.

Anonymous:1876:BPcb

[9489] Anonymous. Business and personal. Scientific American, 35 (26):409, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
409.pdf.

Anonymous:1876:NQca

[9490] Anonymous. Notes and queries. Scientific American, 35(26): 409, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
409a.pdf.

Anonymous:1876:Ic

[9491] Anonymous. Index. Scientific American, 35(26):411–412, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n26/pdf/scientificamerican12231876-411.pdf.

Anonymous:1876:LTP

[9492] Anonymous. The lumber trade in Pennsylvania. Scientific American, 35(26):399, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
399.pdf.

Anonymous:1876:EY

[9493] Anonymous. The end of the year. Scientific American, 35(26): 400, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
400.pdf.

Anonymous:1876:PCh

[9494] Anonymous. Publishers’ card. Scientific American, 35(26):400, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n26/pdf/scientificamerican12231876-400a.pdf.

Anonymous:1876:CSb

[9495] Anonymous. City subscribers. Scientific American, 35(26):400, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v35/n26/pdf/scientificamerican12231876-400b.pdf.

Anonymous:1876:AVA

[9496] Anonymous. The artist vs. the artisan. Scientific American, 35 (26):400, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
400c.pdf.

Anonymous:1876:TBb

[9497] Anonymous. Transfusion of blood. Scientific American, 35(26): 400–401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
400d.pdf.

Anonymous:1876:PFT

[9498] Anonymous. The prevention of fires in theaters. Scientific American, 35(26):401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
401.pdf.

Anonymous:1876:RC

[9499] Anonymous. Remarkable clocks. Scientific American, 35(26): 401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
401a.pdf.

Anonymous:1876:CBPb

[9500] Anonymous. The cremation of Baron De Palm. Scientific American, 35(26):401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
401b.pdf.

Anonymous:1876:WDC

[9501] Anonymous. Who discovered the circulation of the blood? Scientific American, 35(26):401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
401c.pdf.

Anonymous:1876:BNc

[9502] Anonymous. Binding of numbers. Scientific American, 35(26): 401, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
401d.pdf.

Anonymous:1876:ISP

[9503] Anonymous. Improved solar photographic apparatus. Scientific American, 35(26):402, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
402.pdf.

Anonymous:1876:ECM

[9504] Anonymous. English cheese — its manufacture. Scientific American, 35(26):403–404, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
403.pdf.

Anonymous:1876:MPT

[9505] Anonymous. M. Plante’s theory of the formation of hail. Scientific American, 35(26):404, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
404.pdf.

Anonymous:1876:ISSg

[9506] Anonymous. Improved semolina separator. Scientific American, 35(26):405, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
405.pdf.

Anonymous:1876:TFT

[9507] Anonymous. A tasteful flower Trellis. Scientific American, 35 (26):405, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
405a.pdf.

Anonymous:1876:SLA

[9508] Anonymous. Simple laboratory apparatus. Scientific American, 35 (26):405, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
405b.pdf.

Anonymous:1876:NCH

[9509] Anonymous. New combination hand and power feed surface planing machine. Scientific American, 35(26):406, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v35/
n26/pdf/scientificamerican12231876-406.pdf.

Anonymous:1876:BAE

[9510] Anonymous. The British Arctic Expedition. Scientific American, 35 (26):407–408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
407.pdf.

Anonymous:1876:SBVb

[9511] Anonymous. Supply of bound volumes. Scientific American, 35 (26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408.pdf.

Anonymous:1876:OSb

[9512] Anonymous. To our subscribers. Scientific American, 35(26): 408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408a.pdf.

Anonymous:1876:DCz

[9513] Anonymous. Decisions of the courts. Scientific American, 35 (26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408b.pdf.

Anonymous:1876:RAFca

[9514] Anonymous. Recent American and foreign patents. Scientific American, 35(26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408c.pdf.

Anonymous:1876:NMIl

[9515] Anonymous. New miscellaneous inventions. Scientific American, 35(26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408d.pdf.

Anonymous:1876:NMEz

[9516] Anonymous. New mechanical and engineering inventions. Scientific American, 35(26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408e.pdf.

Anonymous:1876:NHIw

[9517] Anonymous. New household inventions. Scientific American, 35 (26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408f.pdf.

Anonymous:1876:NAIbt

[9518] Anonymous. New agricultural inventions. Scientific American, 35 (26):408, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
408g.pdf.

Anonymous:1876:CRca

[9519] Anonymous. Communications received. Scientific American, 35 (26):409–410, December 23, 1876. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v35/n26/pdf/scientificamerican12231876-
409b.pdf.

Anonymous:1877:BNBa

[9520] Anonymous. Books: New books and publications. Scientific American, 36(1):10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
10.pdf.

Anonymous:1877:BPa

[9521] Anonymous. Business and personal. Scientific American, 36(1):11, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-11.pdf.

Anonymous:1877:NQa

[9522] Anonymous. Notes and queries. Scientific American, 36(1):11, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-11a.pdf.

Anonymous:1877:IIa

[9523] Anonymous. Index of inventions. Scientific American, 36(1):11–12, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-11c.pdf.

Anonymous:1877:BAC

[9524] Anonymous. The Brown automatic cut-off engine. Scientific American, 36(1):1, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
1.pdf.

Anonymous:1877:WTT

[9525] Anonymous. A wonderful treasure trove. Scientific American, 36(1):2, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-2.pdf.

Anonymous:1877:UTSa

[9526] Anonymous. Uninflammable theater scenery. Scientific American, 36(1):2, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
2a.pdf.

Anonymous:1877:DGT

[9527] Anonymous. Defective geometrical teachings. Scientific American, 36(1):2–3, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
2b.pdf.

Anonymous:1877:FUC

[9528] Anonymous. The form and use of calipers. Scientific American, 36(1):3, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-3.pdf.

Anonymous:1877:NSCa

[9529] Anonymous. A new steam capstan. Scientific American, 36(1):4, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-4.pdf.

Anonymous:1877:NEF

[9530] Anonymous. A new electric fire alarm and Bell pull. Scientific American, 36(1):4, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
4a.pdf.

Anonymous:1877:WRL

[9531] Anonymous. The Willes and Rowe lightning dumper. Scientific American, 36(1):4, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
4b.pdf.

Anonymous:1877:PBM

[9532] Anonymous. Paper box making. Scientific American, 36(1):5, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-5.pdf.

Anonymous:1877:IIC

[9533] Anonymous. Improved ice cream freezer. Scientific American, 36(1):6, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-6.pdf.

Anonymous:1877:IPGa

[9534] Anonymous. Improved portable gang Sawmill. Scientific American, 36(1):6, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
6a.pdf.

Anonymous:1877:CAa

[9535] Anonymous. A colossal aquarium. Scientific American, 36(1):7, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-7.pdf.

Anonymous:1877:TPF

[9536] Anonymous. The tobacco pipe fish. Scientific American, 36(1):7, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-7a.pdf.

Rose:1877:PMa

[9537] Joshua Rose. Practical mechanism. Scientific American, 36(1):8–9, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-8.pdf.

Anonymous:1877:MCa

[9538] Anonymous. A micrometer caliper. Scientific American, 36(1):9, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-9.pdf.

Anonymous:1877:ANa

[9539] Anonymous. Astronomical notes. Scientific American, 36(1):9, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-9a.pdf.

Riley:1877:B

[9540] C. V. Riley. Bots. Scientific American, 36(1):9–10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-9b.pdf.

Anonymous:1877:DCa

[9541] Anonymous. Decisions of the courts. Scientific American, 36(1):10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-10a.pdf.

Anonymous:1877:RAFa

[9542] Anonymous. Recent American and foreign patents. Scientific American, 36(1):10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
10b.pdf.

Anonymous:1877:NMEa

[9543] Anonymous. New mechanical and engineering inventions. Scientific American, 36(1):10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
10c.pdf.

Anonymous:1877:NMIa

[9544] Anonymous. New miscellaneous inventions. Scientific American, 36(1):10, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n1/pdf/scientificamerican01061877-
10d.pdf.

Anonymous:1877:CRa

[9545] Anonymous. Communications received. Scientific American, 36(1):11, January 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n1/pdf/scientificamerican01061877-11b.pdf.

Anonymous:1877:BPb

[9546] Anonymous. Business and personal. Scientific American, 36(2):27, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-27.pdf.

Anonymous:1877:NQb

[9547] Anonymous. Notes and queries. Scientific American, 36(2):27–28, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-27a.pdf.

Anonymous:1877:IIb

[9548] Anonymous. Index of inventions. Scientific American, 36(2):28–29, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-28a.pdf.

Anonymous:1877:PNPa

[9549] Anonymous. Patent notes: Patents. Scientific American, 36(2):29, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-29.pdf.

Anonymous:1877:PMa

[9550] Anonymous. Puddling machinery. Scientific American, 36(2):15, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-15.pdf.

Anonymous:1877:BHM

[9551] Anonymous. Brayton’s hydrocarbon motor. Scientific American, 36(2):15, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
15a.pdf.

Anonymous:1877:PP

[9552] Anonymous. Physicians as patentees. Scientific American, 36 (2):16, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
16.pdf.

Anonymous:1877:OSG

[9553] Anonymous. The obligations of science to Goethe. Scientific American, 36(2):16, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
16a.pdf.

Anonymous:1877:CMa

[9554] Anonymous. Color music. Scientific American, 36(2):16–17, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-16b.pdf.

Anonymous:1877:MRCa

[9555] Anonymous. More remarkable clocks. Scientific American, 36 (2):17, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
17.pdf.

Anonymous:1877:RCP

[9556] Anonymous. The resignation of the Commissioner of Patents. Scientific American, 36(2):17, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
17a.pdf.

Anonymous:1877:TG

[9557] Anonymous. Tempered glass. Scientific American, 36(2):18–19, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-18.pdf.

Anonymous:1877:CPV

[9558] Anonymous. The Centennial pipe vise. Scientific American, 36 (2):19, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
19.pdf.

Anonymous:1877:WIF

[9559] Anonymous. Wieting’s improved frame clamp. Scientific American, 36(2):19, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
19a.pdf.

Anonymous:1877:Cb

[9560] Anonymous. Communications. Scientific American, 36(2):20–22, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-20.pdf.

Anonymous:1877:TCN

[9561] Anonymous. The tanite Company’s new planer knife and stove plate grinders. Scientific American, 36(2):22, January 13, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n2/pdf/scientificamerican01131877-22.pdf.

Anonymous:1877:DSE

[9562] Anonymous. The domestic steam engine. Scientific American, 36(2):22, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
22a.pdf.

Anonymous:1877:HHF

[9563] Anonymous. The hare and her foes. Scientific American, 36(2):23, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n2/pdf/scientificamerican01131877-23.pdf.

Anonymous:1877:HAL

[9564] Anonymous. How to adjust line shafting. Scientific American, 36(2):24, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
24.pdf.

Anonymous:1877:LWW

[9565] Anonymous. The Louisville Water Works. Scientific American, 36(2):25, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
25.pdf.

Anonymous:1877:NYAa

[9566] Anonymous. New York Academy of Sciences. Scientific American, 36(2):25–26, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
25a.pdf.

Anonymous:1877:RAFb

[9567] Anonymous. Recent American and foreign patents. Scientific American, 36(2):26, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
26.pdf.

Anonymous:1877:NMIb

[9568] Anonymous. New miscellaneous inventions. Scientific American, 36(2):26, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
26a.pdf.

Anonymous:1877:NWHa

[9569] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(2):26, January 13, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n2/pdf/scientificamerican01131877-26b.pdf.

Anonymous:1877:NAIa

[9570] Anonymous. New agricultural inventions. Scientific American, 36(2):26, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
26c.pdf.

Anonymous:1877:NHIa

[9571] Anonymous. New household inventions. Scientific American, 36 (2):26–27, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
26d.pdf.

Anonymous:1877:CRb

[9572] Anonymous. Communications received. Scientific American, 36 (2):28, January 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n2/pdf/scientificamerican01131877-
28.pdf.

Anonymous:1877:BPc

[9573] Anonymous. Business and personal. Scientific American, 36(3):44, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-44.pdf.

Anonymous:1877:NQc

[9574] Anonymous. Notes and queries. Scientific American, 36(3):44–45, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-44a.pdf.

Anonymous:1877:PNPb

[9575] Anonymous. Patent notes: Patents. Scientific American, 36(3):45, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-45a.pdf.

Anonymous:1877:BFE

[9576] Anonymous. The buildings for the French Exposition. Scientific American, 36(3):31, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
31.pdf.

Anonymous:1877:PCP

[9577] Anonymous. The prevention of criminal propagation. Scientific American, 36(3):32, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
32.pdf.

Anonymous:1877:GDA

[9578] Anonymous. The geographical distribution of animals. Scientific American, 36(3):32–33, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
32a.pdf.

Anonymous:1877:MHTa

[9579] Anonymous. Modern halls and theaters. Scientific American, 36(3):33, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
33.pdf.

Anonymous:1877:MDD

[9580] Anonymous. A Mutual Dissection and Decapitation Society. Scientific American, 36(3):33, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
33a.pdf.

Anonymous:1877:CPL

[9581] Anonymous. Colored photo-lithographs by color-blind cameras. Scientific American, 36(3):33, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
33b.pdf.

Anonymous:1877:RAE

[9582] Anonymous. The recent Arctic Expedition. Scientific American, 36(3):34–35, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
34.pdf.

Hachenberg:1877:NH

[9583] G. P. Hachenberg. A new hygrometer. Scientific American, 36 (3):35, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
35.pdf.

Anonymous:1877:Cc

[9584] Anonymous. Communications. Scientific American, 36(3):36, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-36.pdf.

Anonymous:1877:DCV

[9585] Anonymous. The death of Commodore Vanderbilt. Scientific American, 36(3):36–37, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
36a.pdf.

Anonymous:1877:CPa

[9586] Anonymous. Chemical progress in 1876. Scientific American, 36(3):37, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
37.pdf.

Anonymous:1877:WV

[9587] Anonymous. Water velocipedes. Scientific American, 36(3):38, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-38.pdf.

Anonymous:1877:TFM

[9588] Anonymous. A toy flying machine. Scientific American, 36(3):38, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-38a.pdf.

Anonymous:1877:SGT

[9589] Anonymous. Storm glasses and their indications. Scientific American, 36(3):38, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
38b.pdf.

Anonymous:1877:DCb

[9590] Anonymous. Dairy cattle. Scientific American, 36(3):39, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-39.pdf.

Rose:1877:PMb

[9591] Joshua Rose. Practical mechanism. Scientific American, 36(3): 40–41, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
40.pdf.

Anonymous:1877:DCc

[9592] Anonymous. The duplication of the cube. Scientific American, 36(3):41, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
41.pdf.

Anonymous:1877:GOM

[9593] Anonymous. Gynomania and other manias. Scientific American, 36(3):41, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
41a.pdf.

Anonymous:1877:TLa

[9594] Anonymous. Time locks. Scientific American, 36(3):42, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-42.pdf.

Anonymous:1877:RG

[9595] Anonymous. Rope gearing. Scientific American, 36(3):42, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n3/pdf/scientificamerican01201877-42a.pdf.

Anonymous:1877:RAFc

[9596] Anonymous. Recent American and foreign patents. Scientific American, 36(3):43, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
43.pdf.

Anonymous:1877:NAIb

[9597] Anonymous. New agricultural inventions. Scientific American, 36(3):43, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
43a.pdf.

Anonymous:1877:NMIc

[9598] Anonymous. New miscellaneous inventions. Scientific American, 36(3):43, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
43b.pdf.

Anonymous:1877:NHIb

[9599] Anonymous. New household inventions. Scientific American, 36 (3):43, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
43c.pdf.

Anonymous:1877:NMEb

[9600] Anonymous. New mechanical and engineering inventions. Scientific American, 36(3):43–44, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
43d.pdf.

Anonymous:1877:CRc

[9601] Anonymous. Communications received. Scientific American, 36 (3):45, January 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n3/pdf/scientificamerican01201877-
45.pdf.

Anonymous:1877:BNBb

[9602] Anonymous. Books: New books and publications. Scientific American, 36(4):58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
58a.pdf.

Anonymous:1877:BPd

[9603] Anonymous. Business and personal. Scientific American, 36(4):59, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-59.pdf.

Anonymous:1877:NQd

[9604] Anonymous. Notes and queries. Scientific American, 36(4):59, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-59a.pdf.

Anonymous:1877:IId

[9605] Anonymous. Index of inventions. Scientific American, 36(4):59–60, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-59c.pdf.

Anonymous:1877:PNPc

[9606] Anonymous. Patent notes: Patents. Scientific American, 36(4):60, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-61.pdf.

Anonymous:1877:WIT

[9607] Anonymous. Wrought iron tubular safety boiler. Scientific American, 36(4):47, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
47.pdf.

Anonymous:1877:MTM

[9608] Anonymous. The modern tendency of the medical art. Scientific American, 36(4):48, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
48.pdf.

Anonymous:1877:IU

[9609] Anonymous. Ideation in utero. Scientific American, 36(4):48, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-48a.pdf.

Anonymous:1877:IDa

[9610] Anonymous. Inebriety as a disease. Scientific American, 36 (4):48–49, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
48b.pdf.

Anonymous:1877:NHG

[9611] Anonymous. A new hunting ground. Scientific American, 36(4):49, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-49.pdf.

Anonymous:1877:NDM

[9612] Anonymous. A new dredging machine. Scientific American, 36 (4):50, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
50.pdf.

Anonymous:1877:ICS

[9613] Anonymous. Improved carpet stretcher. Scientific American, 36 (4):50, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
50a.pdf.

Anonymous:1877:PMI

[9614] Anonymous. Professor Mabey’s investigations on the action of the heart. Scientific American, 36(4):51–52, January 27, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n4/pdf/scientificamerican01271877-51.pdf.

Anonymous:1877:Cd

[9615] Anonymous. Communications. Scientific American, 36(4):52, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-52.pdf.

Anonymous:1877:RRa

[9616] Anonymous. Rivets and riveting. Scientific American, 36(4):53, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-53.pdf.

Anonymous:1877:ISO

[9617] Anonymous. Improved self-oiling axle box. Scientific American, 36(4):54, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
54.pdf.

Anonymous:1877:UKS

[9618] Anonymous. Useful kit of screw-cutting tools. Scientific American, 36(4):54, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
54a.pdf.

Anonymous:1877:BSP

[9619] Anonymous. The Basilica of St. Peter’s Rome. Scientific American, 36(4):55–56, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
55.pdf.

Anonymous:1877:SWV

[9620] Anonymous. The strongest war vessel in the world. Scientific American, 36(4):57, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
57.pdf.

Anonymous:1877:NSCb

[9621] Anonymous. A new steam carriage. Scientific American, 36 (4):57–58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
57a.pdf.

Anonymous:1877:DCd

[9622] Anonymous. Decisions of the courts. Scientific American, 36(4):58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n4/pdf/scientificamerican01271877-58.pdf.

Anonymous:1877:RAFd

[9623] Anonymous. Recent American and foreign patents. Scientific American, 36(4):58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
58b.pdf.

Anonymous:1877:NAIc

[9624] Anonymous. New agricultural inventions. Scientific American, 36(4):58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
58c.pdf.

Anonymous:1877:NMId

[9625] Anonymous. New miscellaneous inventions. Scientific American, 36(4):58, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
58d.pdf.

Anonymous:1877:NMEc

[9626] Anonymous. New mechanical and engineering inventions. Scientific American, 36(4):58–59, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
58e.pdf.

Anonymous:1877:CRd

[9627] Anonymous. Communications received. Scientific American, 36 (4):59, January 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n4/pdf/scientificamerican01271877-
59b.pdf.

Anonymous:1877:BNBc

[9628] Anonymous. Books: New books and publications. Scientific American, 36(5):74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
74.pdf.

Anonymous:1877:BPe

[9629] Anonymous. Business and personal. Scientific American, 36 (5):75, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
75.pdf.

Anonymous:1877:NQe

[9630] Anonymous. Notes and queries. Scientific American, 36(5):75, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-75a.pdf.

Anonymous:1877:IIe

[9631] Anonymous. Index of inventions. Scientific American, 36(5): 75–76, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
75c.pdf.

Anonymous:1877:PNPd

[9632] Anonymous. Patent notes: Patents. Scientific American, 36 (5):77, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
77.pdf.

Anonymous:1877:Ga

[9633] Anonymous. The gorilla. Scientific American, 36(5):63, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-63.pdf.

Anonymous:1877:SSFa

[9634] Anonymous. Some suggestions for future polar expeditions. Scientific American, 36(5):64, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
64.pdf.

Anonymous:1877:MRCb

[9635] Anonymous. Mind reading and conjurors. Scientific American, 36(5):64, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
64a.pdf.

Anonymous:1877:CER

[9636] Anonymous. The cost of the East River Bridge. Scientific American, 36(5):64–65, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
64b.pdf.

Anonymous:1877:MDA

[9637] Anonymous. The migrations and dispersal of animals. Scientific American, 36(5):65, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
65.pdf.

Anonymous:1877:CDI

[9638] Anonymous. The cause of the delay in issuing the patents. Scientific American, 36(5):65, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
65a.pdf.

Anonymous:1877:ICE

[9639] Anonymous. Improved compound engine. Scientific American, 36(5):66, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
66.pdf.

Anonymous:1877:NTE

[9640] Anonymous. A new tube expander. Scientific American, 36 (5):66, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
66a.pdf.

Anonymous:1877:NFS

[9641] Anonymous. A new food steamer. Scientific American, 36(5):67, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-67.pdf.

Anonymous:1877:NH

[9642] Anonymous. Natural hygroscopes. Scientific American, 36(5):67, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-67a.pdf.

Anonymous:1877:NYAb

[9643] Anonymous. New York Academy of Sciences. Scientific American, 36(5):67, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
67b.pdf.

Anonymous:1877:HDa

[9644] Anonymous. Hippopotamus dentistry. Scientific American, 36 (5):67–68, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
67c.pdf.

Anonymous:1877:Ce

[9645] Anonymous. Communications. Scientific American, 36(5):68, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-68.pdf.

Rose:1877:PMc

[9646] Joshua Rose. Practical mechanism. Scientific American, 36(5): 69, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
69.pdf.

Anonymous:1877:RRb

[9647] Anonymous. The respiration of roots. Scientific American, 36 (5):70, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
70.pdf.

Anonymous:1877:ILS

[9648] Anonymous. Improved locomotive spark arrester. Scientific American, 36(5):70, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
70a.pdf.

Anonymous:1877:PCa

[9649] Anonymous. The prickly comfrey. Scientific American, 36(5):71, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-71.pdf.

Anonymous:1877:CMb

[9650] Anonymous. Chemical magic. Scientific American, 36(5):71–72, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-71a.pdf.

Anonymous:1877:CPb

[9651] Anonymous. Chemical progress in 1876. Scientific American, 36 (5):72–73, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
72.pdf.

Anonymous:1877:OG

[9652] Anonymous. Œdema Glottidis. Scientific American, 36(5):73, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n5/pdf/scientificamerican02031877-73.pdf.

Anonymous:1877:ANb

[9653] Anonymous. Astronomical notes. Scientific American, 36(5): 73–74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
73a.pdf.

Anonymous:1877:RAFe

[9654] Anonymous. Recent American and foreign patents. Scientific American, 36(5):74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
74a.pdf.

Anonymous:1877:NAId

[9655] Anonymous. New agricultural inventions. Scientific American, 36(5):74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
74b.pdf.

Anonymous:1877:NMIe

[9656] Anonymous. New miscellaneous inventions. Scientific American, 36(5):74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
74c.pdf.

Anonymous:1877:NHIc

[9657] Anonymous. New household inventions. Scientific American, 36 (5):74, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
74d.pdf.

Anonymous:1877:NWHb

[9658] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(5):74, February 3, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n5/pdf/scientificamerican02031877-74e.pdf.

Anonymous:1877:CRe

[9659] Anonymous. Communications received. Scientific American, 36 (5):75, February 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n5/pdf/scientificamerican02031877-
75b.pdf.

Anonymous:1877:BPf

[9660] Anonymous. Business and personal. Scientific American, 36(6): 90, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
90.pdf.

Anonymous:1877:NQf

[9661] Anonymous. Notes and queries. Scientific American, 36(6):90– 92, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
90a.pdf.

Anonymous:1877:IIf

[9662] Anonymous. Index of inventions. Scientific American, 36(6):92, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-92a.pdf.

Anonymous:1877:PNPe

[9663] Anonymous. Patent notes: Patents. Scientific American, 36(6): 93, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
93.pdf.

Anonymous:1877:NHR

[9664] Anonymous. Novel hydraulic railway locomotive. Scientific American, 36(6):79, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
79.pdf.

Anonymous:1877:FAa

[9665] Anonymous. Force analyzed. Scientific American, 36(6):80, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-80.pdf.

Anonymous:1877:EI

[9666] Anonymous. The encouragement of insanity. Scientific American, 36(6):80, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
80a.pdf.

Anonymous:1877:MPP

[9667] Anonymous. Medical progress of the past year. Scientific American, 36(6):80–81, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
80b.pdf.

Anonymous:1877:WHa

[9668] Anonymous. Wholesale heating. Scientific American, 36(6):81, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-81.pdf.

Anonymous:1877:PV

[9669] Anonymous. Propelling vessels. Scientific American, 36(6):81, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-81a.pdf.

Anonymous:1877:GSC

[9670] Anonymous. Guardiola’s sugar and coffee machinery. Scientific American, 36(6):82–83, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
82.pdf.

Anonymous:1877:Cf

[9671] Anonymous. Communications. Scientific American, 36(6):84– 85, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
84.pdf.

Anonymous:1877:IUM

[9672] Anonymous. Improved upright moulding machine. Scientific American, 36(6):86, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
86.pdf.

Anonymous:1877:SGM

[9673] Anonymous. The scientific grain mill. Scientific American, 36 (6):86, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
86a.pdf.

Anonymous:1877:ICR

[9674] Anonymous. Improved hat and coat rack. Scientific American, 36(6):86, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
86b.pdf.

Anonymous:1877:LFW

[9675] Anonymous. The largest flower in the world. Scientific American, 36(6):87, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
87.pdf.

Anonymous:1877:AW

[9676] Anonymous. Asparagus in winter. Scientific American, 36(6):87, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-87a.pdf.

Anonymous:1877:Pa

[9677] Anonymous. The pinks. Scientific American, 36(6):87, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n6/pdf/scientificamerican02101877-87b.pdf.

Anonymous:1877:SMM

[9678] Anonymous. Spoke-making machinery. Scientific American, 36 (6):88, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
88.pdf.

Anonymous:1877:ILTa

[9679] Anonymous. Improved lubricant-testing apparatus. Scientific American, 36(6):89, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
89.pdf.

Anonymous:1877:RAFf

[9680] Anonymous. Recent American and foreign patents. Scientific American, 36(6):89, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
89a.pdf.

Anonymous:1877:NHId

[9681] Anonymous. New household inventions. Scientific American, 36 (6):89, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
89b.pdf.

Anonymous:1877:NMIf

[9682] Anonymous. New miscellaneous inventions. Scientific American, 36(6):89, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
89c.pdf.

Anonymous:1877:NAIe

[9683] Anonymous. New agricultural inventions. Scientific American, 36(6):89, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
89d.pdf.

Anonymous:1877:NWHc

[9684] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(6):89, February 10, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n6/pdf/scientificamerican02101877-89e.pdf.

Anonymous:1877:CRf

[9685] Anonymous. Communications received. Scientific American, 36 (6):92, February 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n6/pdf/scientificamerican02101877-
92.pdf.

Anonymous:1877:BNBd

[9686] Anonymous. Books: New books and publications. Scientific American, 36(7):106, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
106.pdf.

Anonymous:1877:BPg

[9687] Anonymous. Business and personal. Scientific American, 36(7): 107, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
107.pdf.

Anonymous:1877:NQg

[9688] Anonymous. Notes and queries. Scientific American, 36(7):107– 108, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
107a.pdf.

Anonymous:1877:PNPf

[9689] Anonymous. Patent notes: Patents. Scientific American, 36 (7):109, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
109.pdf.

Anonymous:1877:EAE

[9690] Anonymous. The eclipse agricultural engine. Scientific American, 36(7):95, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
95.pdf.

Anonymous:1877:MMM

[9691] Anonymous. Matter as a mode of motion. Scientific American, 36(7):96, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
96.pdf.

Anonymous:1877:SMW

[9692] Anonymous. Special machine work versus manipulative skill. Scientific American, 36(7):96–97, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
96a.pdf.

Anonymous:1877:SPL

[9693] Anonymous. State patent laws. Scientific American, 36(7):97, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n7/pdf/scientificamerican02171877-97.pdf.

Anonymous:1877:RSL

[9694] Anonymous. Recent studies of luminous flames. Scientific American, 36(7):97, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
97a.pdf.

Anonymous:1877:FTN

[9695] Anonymous. The flower trade of New York. Scientific American, 36(7):97, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
97b.pdf.

Anonymous:1877:PCb

[9696] Anonymous. Putting in coal. Scientific American, 36(7):97, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n7/pdf/scientificamerican02171877-97c.pdf.

Anonymous:1877:SPb

[9697] Anonymous. Solar photographs. Scientific American, 36(7):98, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n7/pdf/scientificamerican02171877-98.pdf.

Anonymous:1877:OD

[9698] Anonymous. Ornithological deformities. Scientific American, 36 (7):98, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
98a.pdf.

Anonymous:1877:TPE

[9699] Anonymous. A tractile power economizer. Scientific American, 36(7):99, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
99.pdf.

Anonymous:1877:ISV

[9700] Anonymous. Improved safety valves. Scientific American, 36(7): 99, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
99a.pdf.

Anonymous:1877:Cg

[9701] Anonymous. Communications. Scientific American, 36(7):100, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n7/pdf/scientificamerican02171877-100.pdf.

Rose:1877:PMd

[9702] Joshua Rose. Practical mechanism. Scientific American, 36(7): 101, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
101.pdf.

Anonymous:1877:ICMa

[9703] Anonymous. Improved cow milker. Scientific American, 36(7): 102, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
102.pdf.

Anonymous:1877:HPMa

[9704] Anonymous. Hydraulic punching machine. Scientific American, 36(7):102, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
102a.pdf.

Anonymous:1877:FW

[9705] Anonymous. Faience ware. Scientific American, 36(7):103, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n7/pdf/scientificamerican02171877-103.pdf.

Anonymous:1877:NSD

[9706] Anonymous. The new Sun and its disappearance. Scientific American, 36(7):104, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
104.pdf.

Anonymous:1877:RPC

[9707] Anonymous. The resources of the Pacific coast. Scientific American, 36(7):105, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
105.pdf.

Anonymous:1877:RAFg

[9708] Anonymous. Recent American and foreign patents. Scientific American, 36(7):106, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
106a.pdf.

Anonymous:1877:NWHd

[9709] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(7):106, February 17, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n7/pdf/scientificamerican02171877-106b.pdf.

Anonymous:1877:NHIe

[9710] Anonymous. New household inventions. Scientific American, 36 (7):106, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
106c.pdf.

Anonymous:1877:NAIf

[9711] Anonymous. New agricultural inventions. Scientific American, 36(7):106, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
106d.pdf.

Anonymous:1877:CRg

[9712] Anonymous. Communications received. Scientific American, 36 (7):108, February 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n7/pdf/scientificamerican02171877-
108.pdf.

Anonymous:1877:BNBe

[9713] Anonymous. Books: New books and publications. Scientific American, 36(8):122, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
122.pdf.

Anonymous:1877:BPh

[9714] Anonymous. Business and personal. Scientific American, 36(8): 123, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
123.pdf.

Anonymous:1877:NQh

[9715] Anonymous. Notes and queries. Scientific American, 36(8):123, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n8/pdf/scientificamerican02241877-123a.pdf.

Anonymous:1877:IIg

[9716] Anonymous. Index of inventions. Scientific American, 36(8):124, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n8/pdf/scientificamerican02241877-124.pdf.

Anonymous:1877:PNPg

[9717] Anonymous. Patent notes: Patents. Scientific American, 36 (8):126, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
126.pdf.

Anonymous:1877:DDP

[9718] Anonymous. Dates and the date palm. Scientific American, 36 (8):111, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
111.pdf.

Anonymous:1877:PNa

[9719] Anonymous. Publishers’ notice. Scientific American, 36(8):112, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n8/pdf/scientificamerican02241877-112.pdf.

Anonymous:1877:PS

[9720] Anonymous. For posterity — a suggestion. Scientific American, 36 (8):112–113, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
112a.pdf.

Anonymous:1877:LANa

[9721] Anonymous. The lost arts in New York. Scientific American, 36(8):113, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
113.pdf.

Anonymous:1877:BGDa

[9722] Anonymous. The blue glass deception. Scientific American, 36 (8):113, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
113a.pdf.

Anonymous:1877:WMR

[9723] Anonymous. The Wetli Mountain Railroad and its disastrous trial trip. Scientific American, 36(8):114, February 24, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n8/pdf/scientificamerican02241877-114.pdf.

Anonymous:1877:ITH

[9724] Anonymous. Improved trolling hook. Scientific American, 36 (8):114, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
114a.pdf.

Anonymous:1877:HP

[9725] Anonymous. The hemi-plunger. Scientific American, 36(8):115– 116, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
115.pdf.

Anonymous:1877:Ch

[9726] Anonymous. Communications. Scientific American, 36(8):116– 117, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
116.pdf.

Anonymous:1877:NYAc

[9727] Anonymous. New York Academy of Sciences. Scientific American, 36(8):117, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
117.pdf.

Anonymous:1877:ARP

[9728] Anonymous. Annual Report of the Patent Office. Scientific American, 36(8):117, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
117a.pdf.

Anonymous:1877:ILC

[9729] Anonymous. Improved lathe chuck. Scientific American, 36(8): 118, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
118.pdf.

Anonymous:1877:SCL

[9730] Anonymous. Screw-cutting lathes. Scientific American, 36(8): 118, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
118a.pdf.

Anonymous:1877:IHCa

[9731] Anonymous. Improved harness cockeye. Scientific American, 36 (8):118, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
118b.pdf.

Anonymous:1877:PCT

[9732] Anonymous. Proposed cremation Temple. Scientific American, 36(8):119, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
119.pdf.

Joy:1877:PSE

[9733] Charles A. Joy. Papin’s steam engine. Scientific American, 36 (8):120–122, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
120.pdf.

Anonymous:1877:RAFh

[9734] Anonymous. Recent American and foreign patents. Scientific American, 36(8):122, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
122a.pdf.

Anonymous:1877:NMIg

[9735] Anonymous. New miscellaneous inventions. Scientific American, 36(8):122, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
122b.pdf.

Anonymous:1877:NMEd

[9736] Anonymous. New mechanical and engineering inventions. Scientific American, 36(8):122, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
122c.pdf.

Anonymous:1877:CRh

[9737] Anonymous. Communications received. Scientific American, 36 (8):123–124, February 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n8/pdf/scientificamerican02241877-
123b.pdf.

Anonymous:1877:BNBf

[9738] Anonymous. Books: New books and publications. Scientific American, 36(9):136, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
136a.pdf.

Anonymous:1877:BPi

[9739] Anonymous. Business and personal. Scientific American, 36(9):138, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-138.pdf.

Anonymous:1877:NQi

[9740] Anonymous. Notes and queries. Scientific American, 36(9):138–139, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-138a.pdf.

Anonymous:1877:IIh

[9741] Anonymous. Index of inventions. Scientific American, 36(9):140, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-140.pdf.

Anonymous:1877:PNPh

[9742] Anonymous. Patent notes: Patents. Scientific American, 36(9):142, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-142.pdf.

Anonymous:1877:WER

[9743] Anonymous. The wire for the East River Bridge. Scientific American, 36(9):127, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
127.pdf.

Anonymous:1877:PNb

[9744] Anonymous. Publishers’ notice. Scientific American, 36(9):128, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-128.pdf.

Anonymous:1877:SPF

[9745] Anonymous. Science prophesies the future of the race. Scientific American, 36(9):128, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
128a.pdf.

Anonymous:1877:GPa

[9746] Anonymous. A good policy. Scientific American, 36(9):128, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-128b.pdf.

Anonymous:1877:PBL

[9747] Anonymous. Protection of buildings from lightning. Scientific American, 36(9):128–129, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
128c.pdf.

Anonymous:1877:BGDb

[9748] Anonymous. The blue glass deception. Scientific American, 36(9):129, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-129.pdf.

Anonymous:1877:HWA

[9749] Anonymous. How we are able to do it. Scientific American, 36(9):129, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-129a.pdf.

Anonymous:1877:NBC

[9750] Anonymous. A new boiler cleaner. Scientific American, 36(9):130, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-130.pdf.

Anonymous:1877:SCa

[9751] Anonymous. Street cars. Scientific American, 36(9):131–132, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-131.pdf.

Anonymous:1877:Ci

[9752] Anonymous. Communications. Scientific American, 36(9):132, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-132.pdf.

Rose:1877:PMe

[9753] Joshua Rose. Practical mechanism. Scientific American, 36(9):132–133, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-132a.pdf.

Anonymous:1877:LSC

[9754] Anonymous. The “Little Speedy” corn sheller. Scientific American, 36(9):134, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
134.pdf.

Anonymous:1877:MSa

[9755] Anonymous. A machine salesman. Scientific American, 36(9):134, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-134a.pdf.

Anonymous:1877:HE

[9756] Anonymous. The harpy eagle. Scientific American, 36(9):135–136, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-135.pdf.

Anonymous:1877:DCe

[9757] Anonymous. Decisions of the courts. Scientific American, 36(9):136, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n9/pdf/scientificamerican03031877-136.pdf.

Anonymous:1877:RAFi

[9758] Anonymous. Recent American and foreign patents. Scientific American, 36(9):136, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
136b.pdf.

Anonymous:1877:NWHe

[9759] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(9):136, March 3, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n9/pdf/scientificamerican03031877-136c.pdf.

Anonymous:1877:NMEe

[9760] Anonymous. New mechanical and engineering inventions. Scientific American, 36(9):136–137, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
136d.pdf.

Anonymous:1877:NMIh

[9761] Anonymous. New miscellaneous inventions. Scientific American, 36(9):137, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
137.pdf.

Anonymous:1877:CRi

[9762] Anonymous. Communications received. Scientific American, 36 (9):139–140, March 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n9/pdf/scientificamerican03031877-
139.pdf.

Anonymous:1877:BPj

[9763] Anonymous. Business and personal. Scientific American, 36(10):155, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-155.pdf.

Anonymous:1877:NQj

[9764] Anonymous. Notes and queries. Scientific American, 36(10):155–156, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-155a.pdf.

Anonymous:1877:IIi

[9765] Anonymous. Index of inventions. Scientific American, 36(10):156, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-156a.pdf.

Anonymous:1877:PNPi

[9766] Anonymous. Patent notes: Patents. Scientific American, 36(10):158, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-158.pdf.

Anonymous:1877:ERB

[9767] Anonymous. The East River Bridge. Scientific American, 36(10):143, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-143.pdf.

Anonymous:1877:PNc

[9768] Anonymous. Publishers’ notice. Scientific American, 36(10):144, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-144.pdf.

Anonymous:1877:LLP

[9769] Anonymous. Lemuria, the lost paradise. Scientific American, 36(10):144, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
144a.pdf.

Anonymous:1877:CAb

[9770] Anonymous. City architecture. Scientific American, 36(10):144–145, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-144b.pdf.

Anonymous:1877:WPCa

[9771] Anonymous. The working of patents in Canada. Scientific American, 36(10):145, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
145.pdf.

Anonymous:1877:BGDc

[9772] Anonymous. The blue glass deception. Scientific American, 36(10):145, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-145a.pdf.

Anonymous:1877:PLR

[9773] Anonymous. Pointed lightning rods. Scientific American, 36(10):145, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-145b.pdf.

Anonymous:1877:CGa

[9774] Anonymous. Curiosities of glassmaking. Scientific American, 36 (10):146–147, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
146.pdf.

Anonymous:1877:NFC

[9775] Anonymous. A new friction clutch for shafting. Scientific American, 36(10):147, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
147.pdf.

Anonymous:1877:Cj

[9776] Anonymous. Communications. Scientific American, 36(10):148, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-148.pdf.

Anonymous:1877:ANc

[9777] Anonymous. Astronomical notes. Scientific American, 36(10):148, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-148a.pdf.

Anonymous:1877:SWMa

[9778] Anonymous. Straightening wrought metal plates. Scientific American, 36(10):149, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
149.pdf.

Anonymous:1877:THT

[9779] Anonymous. The trial of the hundred-ton gun. Scientific American, 36(10):150, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
150.pdf.

Anonymous:1877:SSRa

[9780] Anonymous. A self-setting rat trap. Scientific American, 36(10):150, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-150a.pdf.

Anonymous:1877:NMR

[9781] Anonymous. New manufacture of rubber articles. Scientific American, 36(10):150, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
150b.pdf.

Anonymous:1877:IB

[9782] Anonymous. Industrial barbarians. Scientific American, 36(10):151, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-151.pdf.

Anonymous:1877:NAa

[9783] Anonymous. A new ambulance. Scientific American, 36(10):151, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-151a.pdf.

Anonymous:1877:DIY

[9784] Anonymous. A Day’s ice yachting on the Hudson. Scientific American, 36(10):152, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
152.pdf.

Anonymous:1877:NYAd

[9785] Anonymous. New York Academy of Sciences. Scientific American, 36(10):152, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
152a.pdf.

Anonymous:1877:RAFj

[9786] Anonymous. Recent American and foreign patents. Scientific American, 36(10):153, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
153.pdf.

Anonymous:1877:NTIa

[9787] Anonymous. New textile inventions. Scientific American, 36(10):153, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n10/pdf/scientificamerican03101877-153a.pdf.

Anonymous:1877:NAIg

[9788] Anonymous. New agricultural inventions. Scientific American, 36(10):153, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
153b.pdf.

Anonymous:1877:NMIi

[9789] Anonymous. New miscellaneous inventions. Scientific American, 36 (10):153–154, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
153c.pdf.

Anonymous:1877:NMEf

[9790] Anonymous. New mechanical and engineering inventions. Scientific American, 36(10):154, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
154.pdf.

Anonymous:1877:NWHf

[9791] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(10):154, March 10, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n10/pdf/scientificamerican03101877-154a.pdf.

Anonymous:1877:CRj

[9792] Anonymous. Communications received. Scientific American, 36 (10):156, March 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n10/pdf/scientificamerican03101877-
156.pdf.

Anonymous:1877:BPk

[9793] Anonymous. Business and personal. Scientific American, 36(11):171, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-171.pdf.

Anonymous:1877:NQk

[9794] Anonymous. Notes and queries. Scientific American, 36(11):171–172, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-171a.pdf.

Anonymous:1877:IIj

[9795] Anonymous. Index of inventions. Scientific American, 36(11):172, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-172a.pdf.

Anonymous:1877:PNPj

[9796] Anonymous. Patent notes: Patents. Scientific American, 36(11):174, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-174.pdf.

Anonymous:1877:UST

[9797] Anonymous. The United States torpedo steamer Alarm. Scientific American, 36(11):159, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
159.pdf.

Anonymous:1877:WIC

[9798] Anonymous. What it costs to feed insects. Scientific American, 36(11):160, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
160.pdf.

Anonymous:1877:PNd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-160a.pdf.

Anonymous:1877:ELR

[9800] Anonymous. Explosions on lightning rod points. Scientific American, 36(11):160, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
160b.pdf.

Anonymous:1877:IRN

[9801] Anonymous. The industries and resources of New South Wales. Scientific American, 36(11):160–161, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
160c.pdf.

Anonymous:1877:WPa

[9802] Anonymous. The World’s population. Scientific American, 36(11):161, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-161.pdf.

Anonymous:1877:PRPa

[9803] Anonymous. Patent rights and patent wrongs. Scientific American, 36(11):161, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
161a.pdf.

Anonymous:1877:RLG

[9804] Anonymous. Robaugh’s little giant pump. Scientific American, 36 (11):162–163, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
162.pdf.

Anonymous:1877:SMa

[9805] Anonymous. Spike machinery. Scientific American, 36(11):163, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-163.pdf.

Anonymous:1877:Ck

[9806] Anonymous. Communications. Scientific American, 36(11):164, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-164.pdf.

Rose:1877:PMf

[9807] Joshua Rose. Practical mechanism. Scientific American, 36(11):165, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-165.pdf.

Anonymous:1877:ISEa

[9808] Anonymous. Improved steam engine cut-off. Scientific American, 36(11):166, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
166.pdf.

Anonymous:1877:NDS

[9809] Anonymous. A new dumping scow. Scientific American, 36(11):166, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n11/pdf/scientificamerican03171877-166a.pdf.

Anonymous:1877:BLW

[9810] Anonymous. The bad lands of White River. Scientific American, 36(11):167, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
167.pdf.

Anonymous:1877:OMI

[9811] Anonymous. The oleo-margarin industry. Scientific American, 36 (11):168–169, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
168.pdf.

Anonymous:1877:NMEg

[9812] Anonymous. New mechanical and engineering inventions. Scientific American, 36(11):170, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
170.pdf.

Anonymous:1877:NMIj

[9813] Anonymous. New miscellaneous inventions. Scientific American, 36(11):170, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
170a.pdf.

Anonymous:1877:NAIh

[9814] Anonymous. New agricultural inventions. Scientific American, 36(11):170, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
170b.pdf.

Anonymous:1877:NHIf

[9815] Anonymous. New household inventions. Scientific American, 36 (11):170, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
170c.pdf.

Anonymous:1877:NWHg

[9816] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(11):170, March 17, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n11/pdf/scientificamerican03171877-170d.pdf.

Anonymous:1877:CRk

[9817] Anonymous. Communications received. Scientific American, 36 (11):172, March 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n11/pdf/scientificamerican03171877-
172.pdf.

Anonymous:1877:BPl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-186.pdf.

Anonymous:1877:NQl

[9819] Anonymous. Notes and queries. Scientific American, 36(12):186–188, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-186a.pdf.

Anonymous:1877:IIk

[9820] Anonymous. Index of inventions. Scientific American, 36(12):188, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-188a.pdf.

Anonymous:1877:PNPk

[9821] Anonymous. Patent notes: Patents. Scientific American, 36(12):190, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-190.pdf.

Prescott:1877:PT

[9822] George B. Prescott. Pneumatic transmission. Scientific American, 36(12):175, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
175.pdf.

Anonymous:1877:SAM

[9823] Anonymous. Selection applied to man. Scientific American, 36(12):176, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-176.pdf.

Anonymous:1877:EKL

[9824] Anonymous. Explosion of kerosene lamps. Scientific American, 36 (12):176–177, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
176a.pdf.

Anonymous:1877:LANb

[9825] Anonymous. A lesson in Arctic navigation. Scientific American, 36(12):177, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
177.pdf.

Anonymous:1877:LDS

[9826] Anonymous. Light and the distances of the stars. Scientific American, 36(12):177, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
177a.pdf.

Anonymous:1877:NEP

[9827] Anonymous. The new English patent bill. Scientific American, 36(12):177, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
177b.pdf.

Anonymous:1877:NEMa

[9828] Anonymous. New experiments on mechanical flight. Scientific American, 36(12):179, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
179.pdf.

Anonymous:1877:NKM

[9829] Anonymous. A new Keely motor deception. Scientific American, 36(12):179, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
179a.pdf.

Anonymous:1877:Cl

[9830] Anonymous. Communications. Scientific American, 36(12):180–181, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-180.pdf.

Anonymous:1877:SWMb

[9831] Anonymous. Straightening wrought metal plates. Scientific American, 36(12):181, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
181.pdf.

Anonymous:1877:NRP

[9832] Anonymous. New rotary pump. Scientific American, 36(12):182, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-182.pdf.

Anonymous:1877:NPR

[9833] Anonymous. A new Phylloxera remedy — decortication. Scientific American, 36(12):182, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
182a.pdf.

Anonymous:1877:NB

[9834] Anonymous. A new bathometer. Scientific American, 36(12):182, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-182b.pdf.

Anonymous:1877:PD

[9835] Anonymous. A Persian dwelling. Scientific American, 36(12):183–184, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-183.pdf.

Ricco:1877:NSD

[9836] A. Ricco. New spectroscope for direct vision. Scientific American, 36(12):184, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
184.pdf.

Anonymous:1877:PNe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-184a.pdf.

Anonymous:1877:RAFk

[9838] Anonymous. Recent American and foreign patents. Scientific American, 36(12):184, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
184b.pdf.

Anonymous:1877:NTIb

[9839] Anonymous. New textile invention. Scientific American, 36(12):184, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n12/pdf/scientificamerican03241877-184c.pdf.

Anonymous:1877:NMIk

[9840] Anonymous. New miscellaneous inventions. Scientific American, 36 (12):184–185, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
184d.pdf.

Anonymous:1877:NMEh

[9841] Anonymous. New mechanical and engineering inventions. Scientific American, 36(12):185, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
185.pdf.

Anonymous:1877:NAIi

[9842] Anonymous. New agricultural inventions. Scientific American, 36(12):185, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
185a.pdf.

Anonymous:1877:NHIg

[9843] Anonymous. New household inventions. Scientific American, 36 (12):185–186, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
185b.pdf.

Anonymous:1877:CRl

[9844] Anonymous. Communications received. Scientific American, 36 (12):188, March 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n12/pdf/scientificamerican03241877-
188.pdf.

Anonymous:1877:BNBg

[9845] Anonymous. Books: New books and publications. Scientific American, 36(13):201, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
201.pdf.

Anonymous:1877:BPm

[9846] Anonymous. Business and personal. Scientific American, 36(13):203, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-203.pdf.

Anonymous:1877:NQm

[9847] Anonymous. Notes and queries. Scientific American, 36(13):203, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-203a.pdf.

Anonymous:1877:IIl

[9848] Anonymous. Index of inventions. Scientific American, 36(13):204, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-204.pdf.

Anonymous:1877:Ta

[9849] Anonymous. The telephone. Scientific American, 36(13):191, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-191.pdf.

Anonymous:1877:PNMa

[9850] Anonymous. Publishers notice to mail subscribers. Scientific American, 36(13):192, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
192.pdf.

Anonymous:1877:ONT

[9851] Anonymous. Our national trial and its results. Scientific American, 36(13):192, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
192a.pdf.

Anonymous:1877:AV

[9852] Anonymous. The Ashtabula verdict. Scientific American, 36(13):192, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-192b.pdf.

Anonymous:1877:IFB

[9853] Anonymous. Iron front buildings. Scientific American, 36(13):192, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-192c.pdf.

Anonymous:1877:TLb

[9854] Anonymous. The theories of light. Scientific American, 36(13):192–193, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-192d.pdf.

Anonymous:1877:PEa

[9855] Anonymous. Photographs in the eye. Scientific American, 36(13):193, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-193.pdf.

Anonymous:1877:EPMa

[9856] Anonymous. An Editor’s perpetual motion. Scientific American, 36(13):193, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
193a.pdf.

Anonymous:1877:CPc

[9857] Anonymous. The Commissioner of Patents. Scientific American, 36(13):193, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
193b.pdf.

Anonymous:1877:RCCa

[9858] Anonymous. Rice culture in China. Scientific American, 36(13):194, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-194.pdf.

Anonymous:1877:M

[9859] Anonymous. Mushrooms. Scientific American, 36(13):194, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-194a.pdf.

Rose:1877:PMg

[9860] Joshua Rose. Practical mechanism. Scientific American, 36(13): 195–196, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
195.pdf.

Anonymous:1877:Cn

[9861] Anonymous. Communications. Scientific American, 36(13):196–197, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-196.pdf.

Anonymous:1877:SCP

[9862] Anonymous. The selden compound pump. Scientific American, 36(13):198, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
198.pdf.

Anonymous:1877:NTC

[9863] Anonymous. A Norwegian timber Church. Scientific American, 36(13):199, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
199.pdf.

Anonymous:1877:DCf

[9864] Anonymous. Decisions of the courts. Scientific American, 36(13):201, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n13/pdf/scientificamerican03311877-201a.pdf.

Anonymous:1877:RAFl

[9865] Anonymous. Recent American and foreign patents. Scientific American, 36(13):202, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
202.pdf.

Anonymous:1877:NMIl

[9866] Anonymous. New miscellaneous inventions. Scientific American, 36(13):202, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
202a.pdf.

Anonymous:1877:NWHh

[9867] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(13):202, March 31, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n13/pdf/scientificamerican03311877-202b.pdf.

Anonymous:1877:NAIj

[9868] Anonymous. New agricultural inventions. Scientific American, 36(13):202, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
202c.pdf.

Anonymous:1877:NMEi

[9869] Anonymous. New mechanical and engineering inventions. Scientific American, 36(13):202, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
202d.pdf.

Anonymous:1877:CRm

[9870] Anonymous. Communications received. Scientific American, 36 (13):203, March 31, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n13/pdf/scientificamerican03311877-
203b.pdf.

Anonymous:1877:BNBh

[9871] Anonymous. Books: New books and publications. Scientific American, 36(14):217, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
217a.pdf.

Anonymous:1877:BPn

[9872] Anonymous. Business and personal. Scientific American, 36(14):219, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-219.pdf.

Anonymous:1877:NQn

[9873] Anonymous. Notes and queries. Scientific American, 36(14):219–220, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-219a.pdf.

Anonymous:1877:IIm

[9874] Anonymous. Index of inventions. Scientific American, 36(14):220–221, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-220a.pdf.

Anonymous:1877:PNPl

[9875] Anonymous. Patent notes: Patents. Scientific American, 36(14):222, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-222.pdf.

Anonymous:1877:ISW

[9876] Anonymous. Improved steam wheel. Scientific American, 36(14):207, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-207.pdf.

Anonymous:1877:PNMb

[9877] Anonymous. Publishers’ notice to mail subscribers. Scientific American, 36(14):208, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
208.pdf.

Anonymous:1877:PL

[9878] Anonymous. The preservation of learning. Scientific American, 36(14):208, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
208a.pdf.

Anonymous:1877:BGE

[9879] Anonymous. The blue glass epidemic. Scientific American, 36(14):208, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-208b.pdf.

Anonymous:1877:CHO

[9880] Anonymous. A curious history of an old inventor. Scientific American, 36(14):208–209, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
208c.pdf.

Anonymous:1877:KI

[9881] Anonymous. On keeping an index. Scientific American, 36(14):209, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-209.pdf.

Anonymous:1877:AIP

[9882] Anonymous. American inventive progress. Scientific American, 36(14):209, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
209a.pdf.

Anonymous:1877:EMb

[9883] Anonymous. Exhibition of microscopes. Scientific American, 36 (14):209, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
209b.pdf.

Anonymous:1877:NFW

[9884] Anonymous. New French woodworking machines. Scientific American, 36(14):210, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
210.pdf.

Anonymous:1877:SMP

[9885] Anonymous. Straightening metal plates. Scientific American, 36 (14):211, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
211.pdf.

Anonymous:1877:Co

[9886] Anonymous. Communications. Scientific American, 36(14):212–213, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-212.pdf.

Anonymous:1877:SWI

[9887] Anonymous. Stone-working implements. Scientific American, 36 (14):213, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
213.pdf.

Anonymous:1877:NLT

[9888] Anonymous. New lubricant-testing machine. Scientific American, 36(14):214, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
214.pdf.

Anonymous:1877:IVS

[9889] Anonymous. Improved vehicle springs. Scientific American, 36(14):214, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-214a.pdf.

Anonymous:1877:SML

[9890] Anonymous. Some members of the lizard family. Scientific American, 36(14):215–216, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
215.pdf.

Anonymous:1877:ANd

[9891] Anonymous. Astronomical notes. Scientific American, 36(14):216, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n14/pdf/scientificamerican04071877-216.pdf.

Schuch:1877:SCN

[9892] Leonidas Schuch. The separation of cobalt from nickel by colorimetric test. Scientific American, 36(14):216–217, April 7, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n14/pdf/scientificamerican04071877-216a.pdf.

Anonymous:1877:NYAe

[9893] Anonymous. New York Academy of Sciences. Scientific American, 36(14):217, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
217.pdf.

Anonymous:1877:RAFm

[9894] Anonymous. Recent American and foreign patents. Scientific American, 36(14):218, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
218.pdf.

Anonymous:1877:NAIk

[9895] Anonymous. New agricultural inventions. Scientific American, 36 (14):218, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
218a.pdf.

Anonymous:1877:NHIh

[9896] Anonymous. New household inventions. Scientific American, 36 (14):218, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
218b.pdf.

Anonymous:1877:NWHi

[9897] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(14):218, April 7, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n14/pdf/scientificamerican04071877-218c.pdf.

Anonymous:1877:NMIm

[9898] Anonymous. New miscellaneous inventions. Scientific American, 36(14):218, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
218d.pdf.

Anonymous:1877:CRn

[9899] Anonymous. Communications received. Scientific American, 36 (14):220, April 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n14/pdf/scientificamerican04071877-
220.pdf.

Anonymous:1877:BNBi

[9900] Anonymous. Books: New books and publications. Scientific American, 36(15):234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
234.pdf.

Anonymous:1877:BPo

[9901] Anonymous. Business and personal. Scientific American, 36(15):235, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-235.pdf.

Anonymous:1877:NQo

[9902] Anonymous. Notes and queries. Scientific American, 36(15):235–236, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-235a.pdf.

Anonymous:1877:IIn

[9903] Anonymous. Index of inventions. Scientific American, 36(15):236, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-236a.pdf.

Anonymous:1877:PNPm

[9904] Anonymous. Patent notes: Patents. Scientific American, 36(15):238, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-238.pdf.

Anonymous:1877:IRGa

[9905] Anonymous. Improved railroad grader. Scientific American, 36 (15):223–224, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
223.pdf.

Anonymous:1877:PLC

[9906] Anonymous. The properties of light and color. Scientific American, 36(15):224, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
224.pdf.

Anonymous:1877:AECa

[9907] Anonymous. The American exhibit at the coming Paris Exposition. Scientific American, 36(15):224, April 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n15/pdf/scientificamerican04141877-224a.pdf.

Anonymous:1877:ACR

[9908] Anonymous. Adjusting connecting rod brasses. Scientific American, 36(15):225, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
225.pdf.

Anonymous:1877:EEN

[9909] Anonymous. The engineers’ examination of the Niagara Suspension Bridge. Scientific American, 36(15):225, April 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n15/pdf/scientificamerican04141877-225a.pdf.

Anonymous:1877:SEE

[9910] Anonymous. Singular explosion of an electrical machine. Scientific American, 36(15):225, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
225b.pdf.

Anonymous:1877:GSB

[9911] Anonymous. The great suspension bridge between New York and Brooklyn. Scientific American, 36(15):225, April 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n15/pdf/scientificamerican04141877-225c.pdf.

Anonymous:1877:SSP

[9912] Anonymous. A screw snow plow. Scientific American, 36(15):226, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-226.pdf.

Anonymous:1877:PRD

[9913] Anonymous. Photographer’s retouching desk. Scientific American, 36(15):226, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
226a.pdf.

Anonymous:1877:NSCc

[9914] Anonymous. A new soda-carbonating apparatus. Scientific American, 36(15):226, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
226b.pdf.

Anonymous:1877:PNEa

[9915] Anonymous. The propagation of nerve excitements. Scientific American, 36(15):227, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
227.pdf.

Anonymous:1877:IGSa

[9916] Anonymous. Improved grain sampler. Scientific American, 36 (15):227–228, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
227a.pdf.

Anonymous:1877:Cq

[9917] Anonymous. Communications. Scientific American, 36(15):228–229, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-228.pdf.

Anonymous:1877:NMSa

[9918] Anonymous. New mode of stopping leaks in boiler tubes. Scientific American, 36(15):230, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
230.pdf.

Anonymous:1877:NPB

[9919] Anonymous. A new perfume bottle. Scientific American, 36(15):230, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-230a.pdf.

Anonymous:1877:RCA

[9920] Anonymous. Robertson’s cradle attachment for bedsteads. Scientific American, 36(15):230, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
230b.pdf.

Anonymous:1877:AIT

[9921] Anonymous. Adee’s improved trap. Scientific American, 36(15):230, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-230c.pdf.

Anonymous:1877:CFI

[9922] Anonymous. The city of Florence, Italy. Scientific American, 36 (15):231–233, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
231.pdf.

Anonymous:1877:DCg

[9923] Anonymous. Decisions of the courts. Scientific American, 36 (15):233–234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
233.pdf.

Anonymous:1877:RAFn

[9924] Anonymous. Recent American and foreign patents. Scientific American, 36(15):234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
234a.pdf.

Anonymous:1877:NMIn

[9925] Anonymous. New miscellaneous inventions. Scientific American, 36(15):234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
234b.pdf.

Anonymous:1877:NWHj

[9926] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(15):234, April 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n15/pdf/scientificamerican04141877-234c.pdf.

Anonymous:1877:NTIc

[9927] Anonymous. New textile invention. Scientific American, 36(15):234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n15/pdf/scientificamerican04141877-234d.pdf.

Anonymous:1877:NAIl

[9928] Anonymous. New agricultural inventions. Scientific American, 36(15):234, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
234e.pdf.

Anonymous:1877:CRo

[9929] Anonymous. Communications received. Scientific American, 36 (15):236, April 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n15/pdf/scientificamerican04141877-
236.pdf.

Anonymous:1877:BPp

[9930] Anonymous. Business and personal. Scientific American, 36(16):250, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-250.pdf.

Anonymous:1877:NQp

[9931] Anonymous. Notes and queries. Scientific American, 36(16):250–252, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-250a.pdf.

Anonymous:1877:IIo

[9932] Anonymous. Index of inventions. Scientific American, 36(16):252, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-252a.pdf.

Anonymous:1877:PNPn

[9933] Anonymous. Patent notes: Patents. Scientific American, 36(16):254, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-254.pdf.

Anonymous:1877:TLW

[9934] Anonymous. Torpedo launch warfare. Scientific American, 36 (16):239–240, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
239.pdf.

Anonymous:1877:DEC

[9935] Anonymous. Darwin on the effects of cross and self fertilization in plants. Scientific American, 36(16):240, April 21, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n16/pdf/scientificamerican04211877-240.pdf.

Anonymous:1877:TCT

[9936] Anonymous. The transmission of correct time. Scientific American, 36(16):240–241, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
240a.pdf.

Anonymous:1877:UR

[9937] Anonymous. The utilization of rats. Scientific American, 36(16):241, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-241.pdf.

Anonymous:1877:FTD

[9938] Anonymous. A fifty thousand dollar botch that the people paid for. Scientific American, 36(16):241, April 21, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n16/pdf/scientificamerican04211877-241a.pdf.

Rose:1877:PMh

[9939] Joshua Rose. Practical mechanism. Scientific American, 36(16):242, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-242.pdf.

Anonymous:1877:OON

[9940] Anonymous. Organs, old and new. Scientific American, 36(16):242–244, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-242a.pdf.

Anonymous:1877:Cs

[9941] Anonymous. Communications. Scientific American, 36(16):244–247, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n16/pdf/scientificamerican04211877-244.pdf.

Anonymous:1877:CAH

[9942] Anonymous. Central African habitations. Scientific American, 36 (16):247–249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
247.pdf.

Anonymous:1877:RAFo

[9943] Anonymous. Recent American and foreign patents. Scientific American, 36(16):249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249.pdf.

Anonymous:1877:NMIo

[9944] Anonymous. New miscellaneous inventions. Scientific American, 36(16):249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249a.pdf.

Anonymous:1877:NMEj

[9945] Anonymous. New mechanical and engineering inventions. Scientific American, 36(16):249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249b.pdf.

Anonymous:1877:NAIm

[9946] Anonymous. New agricultural inventions. Scientific American, 36(16):249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249c.pdf.

Anonymous:1877:NHIi

[9947] Anonymous. New household inventions. Scientific American, 36 (16):249, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249d.pdf.

Anonymous:1877:NWHk

[9948] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(16):249, April 21, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n16/pdf/scientificamerican04211877-249e.pdf.

Anonymous:1877:NTId

[9949] Anonymous. New textile invention. Scientific American, 36(16): 249–250, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
249f.pdf.

Anonymous:1877:CRp

[9950] Anonymous. Communications received. Scientific American, 36 (16):252, April 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n16/pdf/scientificamerican04211877-
252.pdf.

Anonymous:1877:BNBj

[9951] Anonymous. Books: New books and publications. Scientific American, 36(17):266, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
266.pdf.

Anonymous:1877:BPq

[9952] Anonymous. Business and personal. Scientific American, 36(17):267, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-267.pdf.

Anonymous:1877:NQNa

[9953] Anonymous. Notes and queries: Notes & queries. Scientific American, 36(17):267–268, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
267a.pdf.

Anonymous:1877:IIp

[9954] Anonymous. Index of inventions. Scientific American, 36(17):268, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-268a.pdf.

Anonymous:1877:PNPo

[9955] Anonymous. Patent notes: Patents. Scientific American, 36(17):270, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-270.pdf.

Anonymous:1877:IGL

[9956] Anonymous. Improved gap lathe. Scientific American, 36(17):255, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-255.pdf.

Anonymous:1877:BTa

[9957] Anonymous. A bull on the track. Scientific American, 36(17):256, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-256.pdf.

Anonymous:1877:DLa

[9958] Anonymous. Dual life. Scientific American, 36(17):256, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-256a.pdf.

Anonymous:1877:CSa

[9959] Anonymous. The Centennial surplus. Scientific American, 36 (17):256–257, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
256b.pdf.

Anonymous:1877:AMP

[9960] Anonymous. Alleged magneto-photography. Scientific American, 36(17):257, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
257.pdf.

Anonymous:1877:RRE

[9961] Anonymous. Remarkable results of evaporation and rainfall. Scientific American, 36(17):257–258, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
257a.pdf.

Anonymous:1877:SSa

[9962] Anonymous. Straightening saws. Scientific American, 36(17):259–260, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-259.pdf.

Anonymous:1877:Ct

[9963] Anonymous. Communications. Scientific American, 36(17):260, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-260.pdf.

Riley:1877:IOR

[9964] C. V. Riley. Important observations on the Rocky Mountain locust, or “grasshopper” pest of the West. Scientific American, 36(17):260–261, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-260a.pdf.

Anonymous:1877:BIA

[9965] Anonymous. The British ironclad Alexandra. Scientific American, 36(17):261, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
261.pdf.

Anonymous:1877:NHE

[9966] Anonymous. A new hydraulic engine. Scientific American, 36(17):262, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-262.pdf.

Anonymous:1877:MGR

[9967] Anonymous. Martin’s gas regulator for steam boilers. Scientific American, 36(17):262, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
262a.pdf.

Anonymous:1877:IYS

[9968] Anonymous. Improved yard stick and bias measure. Scientific American, 36(17):262, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
262b.pdf.

Anonymous:1877:PGT

[9969] Anonymous. Professor Gray’s telephone. Scientific American, 36 (17):263, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
263.pdf.

Anonymous:1877:FPS

[9970] Anonymous. The friction of plain slide valves. Scientific American, 36(17):264–265, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
264.pdf.

Anonymous:1877:CTa

[9971] Anonymous. The cat teaser. Scientific American, 36(17):265–266, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n17/pdf/scientificamerican04281877-265.pdf.

Anonymous:1877:RAFp

[9972] Anonymous. Recent American and foreign patents. Scientific American, 36(17):266, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
266a.pdf.

Anonymous:1877:NMIp

[9973] Anonymous. New miscellaneous inventions. Scientific American, 36(17):266, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
266b.pdf.

Anonymous:1877:NMEk

[9974] Anonymous. New mechanical and engineering inventions. Scientific American, 36(17):266, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
266c.pdf.

Anonymous:1877:CRq

[9975] Anonymous. Communications received. Scientific American, 36 (17):268, April 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n17/pdf/scientificamerican04281877-
268.pdf.

Anonymous:1877:BPr

[9976] Anonymous. Business and personal. Scientific American, 36(18):283, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-283.pdf.

Anonymous:1877:NQq

[9977] Anonymous. Notes and queries. Scientific American, 36(18):283–284, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-283a.pdf.

Anonymous:1877:IIq

[9978] Anonymous. Index of inventions. Scientific American, 36(18):284, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-284a.pdf.

Anonymous:1877:PNPp

[9979] Anonymous. Patent notes: Patents. Scientific American, 36(18):286, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-286.pdf.

Anonymous:1877:IBMa

[9980] Anonymous. Improved brick machine. Scientific American, 36 (18):271–272, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
271.pdf.

Anonymous:1877:EWA

[9981] Anonymous. The European war as affecting American inventors and farmers. Scientific American, 36(18):272, May 5, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n18/pdf/scientificamerican05051877-272.pdf.

Anonymous:1877:HF

[9982] Anonymous. Hotel fires. Scientific American, 36(18):272–273, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-272a.pdf.

Anonymous:1877:OIS

[9983] Anonymous. Our iron shipbuilding industry. Scientific American, 36(18):273, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
273.pdf.

Anonymous:1877:AMU

[9984] Anonymous. The advancement in microscopy in the United States. Scientific American, 36(18):273, May 5, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n18/pdf/scientificamerican05051877-273a.pdf.

Anonymous:1877:NLP

[9985] Anonymous. A novel life preserver. Scientific American, 36(18):274, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-274.pdf.

Anonymous:1877:IWE

[9986] Anonymous. Improved water elevator. Scientific American, 36 (18):274–275, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
274a.pdf.

Anonymous:1877:DSS

[9987] Anonymous. The Dexter scroll saw, emery grinder, and polisher. Scientific American, 36(18):275, May 5, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n18/pdf/scientificamerican05051877-275.pdf.

Anonymous:1877:DHD

[9988] Anonymous. Device for holding door knob screws. Scientific American, 36(18):275, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
275a.pdf.

Anonymous:1877:Cu

[9989] Anonymous. Communications. Scientific American, 36(18):276, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-276.pdf.

Riley:1877:ELE

[9990] C. V. Riley. Experiments with locust eggs, and conclusions drawn therefrom. Scientific American, 36(18):276–277, May 5, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n18/pdf/scientificamerican05051877-276a.pdf.

Anonymous:1877:EAMa

[9991] Anonymous. The exportation of American meat to England. Scientific American, 36(18):278–280, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
278.pdf.

Anonymous:1877:ANe

[9992] Anonymous. Astronomical notes. Scientific American, 36(18):280, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n18/pdf/scientificamerican05051877-280.pdf.

Anonymous:1877:NYAf

[9993] Anonymous. New York Academy of Sciences. Scientific American, 36(18):281–282, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
281.pdf.

Anonymous:1877:RAFq

[9994] Anonymous. Recent American and foreign patents. Scientific American, 36(18):282, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
282.pdf.

Anonymous:1877:NMIq

[9995] Anonymous. New miscellaneous inventions. Scientific American, 36(18):282, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
282a.pdf.

Anonymous:1877:NWHl

[9996] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(18):282, May 5, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n18/pdf/scientificamerican05051877-282b.pdf.

Anonymous:1877:NMEl

[9997] Anonymous. New mechanical and engineering inventions. Scientific American, 36(18):282, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
282c.pdf.

Anonymous:1877:NAIn

[9998] Anonymous. New agricultural inventions. Scientific American, 36(18):282, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
282d.pdf.

Anonymous:1877:CRr

[9999] Anonymous. Communications received. Scientific American, 36 (18):284, May 5, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n18/pdf/scientificamerican05051877-
284.pdf.

Anonymous:1877:BPs

[10000] Anonymous. Business and personal. Scientific American, 36(19):299, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-299.pdf.

Anonymous:1877:NQr

[10001] Anonymous. Notes and queries. Scientific American, 36(19):299–300, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-299a.pdf.

Anonymous:1877:IIr

[10002] Anonymous. Index of inventions. Scientific American, 36(19):300, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-300a.pdf.

Anonymous:1877:PNPq

[10003] Anonymous. Patent notes: Patents. Scientific American, 36(19):302, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-302.pdf.

Anonymous:1877:LBO

[10004] Anonymous. Lands below the ocean level. Scientific American, 36(19):288, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
288.pdf.

Anonymous:1877:TFL

[10005] Anonymous. The Thorneycroft fast launches. Scientific American, 36(19):288, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
288a.pdf.

Anonymous:1877:PSW

[10006] Anonymous. The permanent supply water works of Baltimore. Scientific American, 36(19):288–289, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
288b.pdf.

Anonymous:1877:WTS

[10007] Anonymous. What is a temporary star? Scientific American, 36(19):289–290, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
289.pdf.

Anonymous:1877:MSb

[10008] Anonymous. The manufacture of silverware. Scientific American, 36(19):290–293, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
290.pdf.

Anonymous:1877:Cv

[10009] Anonymous. Communications. Scientific American, 36(19):292, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-292.pdf.

Rose:1877:PMi

[10010] Joshua Rose. Practical mechanism. Scientific American, 36(19): 293–294, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
293.pdf.

Anonymous:1877:IFEa

[10011] Anonymous. Improved fire escape. Scientific American, 36(19):294, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-294.pdf.

Anonymous:1877:ICMb

[10012] Anonymous. Improved calculating machine. Scientific American, 36(19):294–295, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
294a.pdf.

Anonymous:1877:BB

[10013] Anonymous. Baby brutes. Scientific American, 36(19):295–296, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n19/pdf/scientificamerican05121877-295.pdf.

Anonymous:1877:CPM

[10014] Anonymous. A curious piece of mechanical manipulation. — cutting right or left hand threads with right hand dies. Scientific American, 36(19):296–297, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
296.pdf.

Anonymous:1877:DCh

[10015] Anonymous. Decisions of the courts. Scientific American, 36 (19):297–298, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
297.pdf.

Anonymous:1877:RAFr

[10016] Anonymous. Recent American and foreign patents. Scientific American, 36(19):298, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
298.pdf.

Anonymous:1877:NHIj

[10017] Anonymous. New household inventions. Scientific American, 36 (19):298, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
298a.pdf.

Anonymous:1877:NWHm

[10018] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(19):298, May 12, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n19/pdf/scientificamerican05121877-298b.pdf.

Anonymous:1877:NMIr

[10019] Anonymous. New miscellaneous inventions. Scientific American, 36(19):298, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
298c.pdf.

Anonymous:1877:NMEm

[10020] Anonymous. New mechanical and engineering inventions. Scientific American, 36(19):298, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
298d.pdf.

Anonymous:1877:NAIo

[10021] Anonymous. New agricultural inventions. Scientific American, 36 (19):298–300, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
298e.pdf.

Anonymous:1877:CRs

[10022] Anonymous. Communications received. Scientific American, 36 (19):300, May 12, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n19/pdf/scientificamerican05121877-
300.pdf.

Anonymous:1877:BNBk

[10023] Anonymous. Books: New books and publications. Scientific American, 36(20):313, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
313a.pdf.

Anonymous:1877:BPt

[10024] Anonymous. Business and personal. Scientific American, 36(20):315, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-315.pdf.

Anonymous:1877:NQs

[10025] Anonymous. Notes and queries. Scientific American, 36(20):315–316, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-315a.pdf.

Anonymous:1877:IIs

[10026] Anonymous. Index of inventions. Scientific American, 36(20):316, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-316a.pdf.

Anonymous:1877:PNPr

[10027] Anonymous. Patent notes: Patents. Scientific American, 36(20):318, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-318.pdf.

Anonymous:1877:NLSa

[10028] Anonymous. New lithographic stone-dressing machine. Scientific American, 36(20):303, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
303.pdf.

Anonymous:1877:PSE

[10029] Anonymous. Portable steam engine. Scientific American, 36(20):303, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-303a.pdf.

Anonymous:1877:RIP

[10030] Anonymous. Recent improvements in photography. Scientific American, 36(20):304, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
304.pdf.

Anonymous:1877:NSBa

[10031] Anonymous. Neuralgic storm belts. Scientific American, 36(20):304, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-304a.pdf.

Anonymous:1877:NEC

[10032] Anonymous. A new explosive compound for large guns. Scientific American, 36(20):304–305, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
304b.pdf.

Anonymous:1877:WFW

[10033] Anonymous. Why fresh water fish cannot live in salt water. Scientific American, 36(20):305, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
305.pdf.

Anonymous:1877:LHA

[10034] Anonymous. “Lost his ambition”. Scientific American, 36(20):305, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-305a.pdf.

Anonymous:1877:WHC

[10035] Anonymous. Water held in a cage and boiled in a sieve. Scientific American, 36(20):306, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
306.pdf.

Anonymous:1877:NES

[10036] Anonymous. A new experiment for the synthesis of sunlight. Scientific American, 36(20):306, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
306a.pdf.

Anonymous:1877:IFSa

[10037] Anonymous. Improved fish scaler. Scientific American, 36(20):306, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-306b.pdf.

Anonymous:1877:PPR

[10038] Anonymous. The papyrus or paper Reed. Scientific American, 36(20):307, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
307.pdf.

Anonymous:1877:Cw

[10039] Anonymous. Communications. Scientific American, 36(20):308, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-308.pdf.

Anonymous:1877:HDI

[10040] Anonymous. How to do it, and how not to do it. Scientific American, 36(20):309, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
309.pdf.

Anonymous:1877:IHS

[10041] Anonymous. Improved high speed dry air compressor. Scientific American, 36(20):310, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
310.pdf.

Anonymous:1877:KAR

[10042] Anonymous. Kenyon’s adjustable rubber bucket for chain pumps. Scientific American, 36(20):310, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
310a.pdf.

Anonymous:1877:CPT

[10043] Anonymous. Combined pipe tongs, wrench, and screwdriver. Scientific American, 36(20):310, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
310b.pdf.

Anonymous:1877:RK

[10044] Anonymous. A remarkable kingfisher. Scientific American, 36 (20):311–313, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
311.pdf.

Anonymous:1877:DCi

[10045] Anonymous. Decisions of the courts. Scientific American, 36(20):313, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-313.pdf.

Anonymous:1877:RAFs

[10046] Anonymous. Recent American and foreign patents. Scientific American, 36(20):313, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
313b.pdf.

Anonymous:1877:NMIs

[10047] Anonymous. New miscellaneous inventions. Scientific American, 36(20):313–314, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
313c.pdf.

Anonymous:1877:NMEn

[10048] Anonymous. New mechanical and engineering inventions. Scientific American, 36(20):314, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
314.pdf.

Anonymous:1877:NWHn

[10049] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(20):314, May 19, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n20/pdf/scientificamerican05191877-314a.pdf.

Anonymous:1877:NAIp

[10050] Anonymous. New agricultural inventions. Scientific American, 36 (20):314, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
314b.pdf.

Anonymous:1877:NTIe

[10051] Anonymous. New textile invention. Scientific American, 36(20):314, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n20/pdf/scientificamerican05191877-314c.pdf.

Anonymous:1877:NHIk

[10052] Anonymous. New household inventions. Scientific American, 36 (20):314, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
314d.pdf.

Anonymous:1877:CRt

[10053] Anonymous. Communications received. Scientific American, 36 (20):316, May 19, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n20/pdf/scientificamerican05191877-
316.pdf.

Anonymous:1877:BNBl

[10054] Anonymous. Books: New books and publications. Scientific American, 36(21):329, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
329a.pdf.

Anonymous:1877:BPu

[10055] Anonymous. Business and personal. Scientific American, 36(21):331, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-331.pdf.

Anonymous:1877:NQt

[10056] Anonymous. Notes and queries. Scientific American, 36(21):331–332, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-331a.pdf.

Anonymous:1877:IIt

[10057] Anonymous. Index of inventions. Scientific American, 36(21):332, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-332a.pdf.

Anonymous:1877:PNPs

[10058] Anonymous. Patent notes: Patents. Scientific American, 36(21):334, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-334.pdf.

Anonymous:1877:DLT

[10059] Anonymous. The Delaware and Lackawanna Tunnel through Bergen Hill, N.J. Scientific American, 36(21):319, May 26, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n21/pdf/scientificamerican05261877-319.pdf.

Anonymous:1877:RSM

[10060] Anonymous. The rise of the sewing machine. Scientific American, 36(21):320–321, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
320.pdf.

Anonymous:1877:FLM

[10061] Anonymous. The fate of the last man. Scientific American, 36(21):321, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-321.pdf.

Anonymous:1877:AEP

[10062] Anonymous. American exhibitors at Paris. Scientific American, 36(21):321, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
321a.pdf.

Anonymous:1877:AC

[10063] Anonymous. The alhydric chain. Scientific American, 36(21):322, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-322.pdf.

Anonymous:1877:ISCa

[10064] Anonymous. Improved steam chest seat milling machine. Scientific American, 36(21):322, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
322a.pdf.

Anonymous:1877:PRPb

[10065] Anonymous. A Persian Royal Pavilion. Scientific American, 36(21):323, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-323.pdf.

Anonymous:1877:Cx

[10066] Anonymous. Communications. Scientific American, 36(21):324–325, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-324.pdf.

Rose:1877:PMj

[10067] Joshua Rose. Practical mechanism. Scientific American, 36(21):325, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-325.pdf.

Anonymous:1877:Pb

[10068] Anonymous. Pompeii. Scientific American, 36(21):326–329, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-326.pdf.

Anonymous:1877:NYAg

[10069] Anonymous. New York Academy of Sciences. Scientific American, 36(21):329, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
329.pdf.

Anonymous:1877:RAFt

[10070] Anonymous. Recent American and foreign patents. Scientific American, 36(21):329, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
329b.pdf.

Anonymous:1877:NMIt

[10071] Anonymous. New miscellaneous inventions. Scientific American, 36(21):329–330, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
329c.pdf.

Anonymous:1877:NMEo

[10072] Anonymous. New mechanical and engineering inventions. Scientific American, 36(21):330, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
330.pdf.

Anonymous:1877:NTIf

[10073] Anonymous. New textile invention. Scientific American, 36(21):330, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n21/pdf/scientificamerican05261877-330a.pdf.

Anonymous:1877:NHIl

[10074] Anonymous. New household inventions. Scientific American, 36 (21):330, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
330b.pdf.

Anonymous:1877:CRu

[10075] Anonymous. Communications received. Scientific American, 36 (21):332, May 26, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n21/pdf/scientificamerican05261877-
332.pdf.

Anonymous:1877:BNBm

[10076] Anonymous. Books: New books and publications. Scientific American, 36(22):345, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
345b.pdf.

Anonymous:1877:BPv

[10077] Anonymous. Business and personal. Scientific American, 36(22):347, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-347.pdf.

Anonymous:1877:NQu

[10078] Anonymous. Notes and queries. Scientific American, 36(22):347–348, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-347a.pdf.

Anonymous:1877:IIu

[10079] Anonymous. Index of inventions. Scientific American, 36(22):348, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-348a.pdf.

Anonymous:1877:PNPt

[10080] Anonymous. Patent notes: Patents. Scientific American, 36(22):350, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-350.pdf.

Anonymous:1877:NLSb

[10081] Anonymous. New life-saving inventions. Scientific American, 36 (22):335, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
335.pdf.

Anonymous:1877:HMF

[10082] Anonymous. The human machine and its fuel. Scientific American, 36(22):336, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
336.pdf.

Anonymous:1877:WTD

[10083] Anonymous. Wanted-torpedo defences. Scientific American, 36 (22):336–337, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
336a.pdf.

Gakuma:1877:T

[10084] G. Gakuma. Torpedoes. Scientific American, 36(22):337, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-337.pdf.

Anonymous:1877:IFH

[10085] Anonymous. Improved friction hoisting engines. Scientific American, 36(22):338, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
338.pdf.

Anonymous:1877:TNU

[10086] Anonymous. Two new utilizations of paper pulp. Scientific American, 36(22):339, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
339.pdf.

Anonymous:1877:NGB

[10087] Anonymous. The new Godefroy burner. Scientific American, 36 (22):339, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
339a.pdf.

Anonymous:1877:EC

[10088] Anonymous. The electric candle. Scientific American, 36(22):339–340, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-339b.pdf.

Anonymous:1877:Cy

[10089] Anonymous. Communications. Scientific American, 36(22):340–341, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-340.pdf.

Morton:1877:PIC

[10090] Henry Morton. Projection of interference colors from soap films. Scientific American, 36(22):341, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
341.pdf.

Anonymous:1877:ANf

[10091] Anonymous. Astronomical notes. Scientific American, 36(22):341, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-341a.pdf.

Anonymous:1877:GIH

[10092] Anonymous. Grant’s improved horse hay fork. Scientific American, 36(22):341, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
341b.pdf.

Anonymous:1877:WW

[10093] Anonymous. A wonderful watch. Scientific American, 36(22):342, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-342.pdf.

Anonymous:1877:PRH

[10094] Anonymous. Professor R. H. Thurston’s autographic testing machine. Scientific American, 36(22):342, June 2, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n22/pdf/scientificamerican06021877-342a.pdf.

Anonymous:1877:FAb

[10095] Anonymous. The fountains at Aranjuez. Scientific American, 36 (22):343, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
343.pdf.

Anonymous:1877:CSL

[10096] Anonymous. Californian sea lions. Scientific American, 36(22):343–344, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n22/pdf/scientificamerican06021877-343a.pdf.

Anonymous:1877:NYAh

[10097] Anonymous. New York Academy of Sciences. Scientific American, 36(22):345, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
345.pdf.

Anonymous:1877:IPO

[10098] Anonymous. International postal order system. Scientific American, 36(22):345, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
345a.pdf.

Anonymous:1877:RAFu 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
345c.pdf.

Anonymous:1877:NMEp

[10100] Anonymous. New mechanical and engineering inventions. Scientific American, 36(22):345–346, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
345d.pdf.

Anonymous:1877:NAIq

[10101] Anonymous. New agricultural inventions. Scientific American, 36(22):346, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
346.pdf.

Anonymous:1877:NHIm

[10102] Anonymous. New household inventions. Scientific American, 36 (22):346, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
346a.pdf.

Anonymous:1877:NMIu

[10103] Anonymous. New miscellaneous inventions. Scientific American, 36(22):346, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
346b.pdf.

Anonymous:1877:CRv

[10104] Anonymous. Communications received. Scientific American, 36 (22):348, June 2, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n22/pdf/scientificamerican06021877-
348.pdf.

Anonymous:1877:BNBn

[10105] Anonymous. Books: New books and publications. Scientific American, 36(23):360, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
360a.pdf.

Anonymous:1877:BPw

[10106] Anonymous. Business and personal. Scientific American, 36(23):362, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-362.pdf.

Anonymous:1877:NQv

[10107] Anonymous. Notes and queries. Scientific American, 36(23):362–364, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-362a.pdf.

Anonymous:1877:IIv

[10108] Anonymous. Index of inventions. Scientific American, 36(23):364, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-364a.pdf.

Anonymous:1877:PNPu

[10109] Anonymous. Patent notes: Patents. Scientific American, 36(23):366, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-366.pdf.

Anonymous:1877:STN

[10110] Anonymous. The speaking telephone in New York. Scientific American, 36(23):351, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
351.pdf.

Anonymous:1877:WIR

[10111] Anonymous. Wiggins’ improved relief and safety stop valve. Scientific American, 36(23):351, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
351a.pdf.

Anonymous:1877:SSB

[10112] Anonymous. The South Street Bridge, Philadelphia, PA. Scientific American, 36(23):351, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
351b.pdf.

Anonymous:1877:SIS

[10113] Anonymous. Sewage irrigation on a small scale. Scientific American, 36(23):352, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
352.pdf.

Anonymous:1877:PMb

[10114] Anonymous. Prevalent manias. Scientific American, 36(23):352, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-352a.pdf.

Anonymous:1877:SVT

[10115] Anonymous. Safety valve tests. Scientific American, 36(23):352–353, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-352b.pdf.

Anonymous:1877:IMSa

[10116] Anonymous. Improved machine for stamping leather goods. Scientific American, 36(23):355, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
355.pdf.

Anonymous:1877:NIS

[10117] Anonymous. A new injector for steam boilers. Scientific American, 36(23):355, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
355a.pdf.

Anonymous:1877:NAb

[10118] Anonymous. A new axe. Scientific American, 36(23):355, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-355b.pdf.

Anonymous:1877:Cz

[10119] Anonymous. Communications. Scientific American, 36(23):356, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-356.pdf.

Rose:1877:PMk

[10120] Joshua Rose. Practical mechanism. Scientific American, 36(23):357, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-357.pdf.

Anonymous:1877:IRJa

[10121] Anonymous. Improved railroad joint and nut lock. Scientific American, 36(23):358, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
358.pdf.

Anonymous:1877:MIR

[10122] Anonymous. Mershon’s improved rock drill. Scientific American, 36(23):358, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
358a.pdf.

Anonymous:1877:NF

[10123] Anonymous. A new fumigator. Scientific American, 36(23):358, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-358b.pdf.

Anonymous:1877:SEF

[10124] Anonymous. A snake-eating frog. Scientific American, 36(23):359, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-359.pdf.

Anonymous:1877:BF

[10125] Anonymous. A beautiful fern. Scientific American, 36(23):359–360, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n23/pdf/scientificamerican06091877-359a.pdf.

Anonymous:1877:NYAi

[10126] Anonymous. New York Academy of Sciences. Scientific American, 36(23):360, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
360.pdf.

Anonymous:1877:DCj

[10127] Anonymous. Decisions of the courts. Scientific American, 36 (23):360–361, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
360b.pdf.

Anonymous:1877:RAFv

[10128] Anonymous. Recent American and foreign patents. Scientific American, 36(23):361, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
361.pdf.

Anonymous:1877:NMEq

[10129] Anonymous. New mechanical and engineering inventions. Scientific American, 36(23):361, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
361a.pdf.

Anonymous:1877:NMIv

[10130] Anonymous. New miscellaneous inventions. Scientific American, 36(23):361, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
361b.pdf.

Anonymous:1877:NWHo

[10131] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(23):361, June 9, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n23/pdf/scientificamerican06091877-361c.pdf.

Anonymous:1877:CRw

[10132] Anonymous. Communications received. Scientific American, 36 (23):362, June 9, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n23/pdf/scientificamerican06091877-
364.pdf.

Anonymous:1877:BNBo

[10133] Anonymous. Books: New books and publications. Scientific American, 36(24):377, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
377.pdf.

Anonymous:1877:BPx

[10134] Anonymous. Business and personal. Scientific American, 36(24):379, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-379.pdf.

Anonymous:1877:NQw

[10135] Anonymous. Notes and queries. Scientific American, 36(24):379–380, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-379a.pdf.

Anonymous:1877:IIw

[10136] Anonymous. Index of inventions. Scientific American, 36(24):380, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-380.pdf.

Anonymous:1877:PNPv

[10137] Anonymous. Patent notes: Patents. Scientific American, 36(24):382, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-382.pdf.

Anonymous:1877:SUN

[10138] Anonymous. Sewage utilization near Paris. Scientific American, 36(24):367, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
367.pdf.

Anonymous:1877:PEI

[10139] Anonymous. Progress of electric illumination. Scientific American, 36(24):368, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
368.pdf.

Anonymous:1877:SAW

[10140] Anonymous. Sailing against the wind. Scientific American, 36(24):368, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-368a.pdf.

Anonymous:1877:PNG

[10141] Anonymous. Passage of the new German patent law. Scientific American, 36(24):368–369, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
368b.pdf.

Anonymous:1877:TPL

[10142] Anonymous. Taxing power of the legislature. Scientific American, 36(24):369, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
369.pdf.

Anonymous:1877:ZLW

[10143] Anonymous. Zinc-lined water coolers. Scientific American, 36(24):369, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-369a.pdf.

Anonymous:1877:QT

[10144] Anonymous. Quadruplex telegraphy. Scientific American, 36(24):369, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-369b.pdf.

Riley:1877:LP

[10145] C. V. Riley. Locust prospects. Scientific American, 36(24):369, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-369c.pdf.

Anonymous:1877:NTIg

[10146] Anonymous. New Turkish ironclad. Scientific American, 36(24):370, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-370.pdf.

Anonymous:1877:PT

[10147] Anonymous. Pipe tools. Scientific American, 36(24):371, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-371.pdf.

Anonymous:1877:Cba

[10148] Anonymous. Communications. Scientific American, 36(24):372, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-372.pdf.

Anonymous:1877:Gb

[10149] Anonymous. Gauges. Scientific American, 36(24):372–373, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-372a.pdf.

Anonymous:1877:NFP

[10150] Anonymous. New feathering paddle wheel. Scientific American, 36(24):374, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
374.pdf.

Anonymous:1877:WIL

[10151] Anonymous. Wilkinson’s improved liquid measure. Scientific American, 36(24):374, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
374a.pdf.

Anonymous:1877:IMSb

[10152] Anonymous. Improved machine for shaping springs. Scientific American, 36(24):374, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
374b.pdf.

Anonymous:1877:PAa

[10153] Anonymous. Peruvian antiquities. Scientific American, 36(24):375–377, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n24/pdf/scientificamerican06161877-375.pdf.

Anonymous:1877:DCk

[10154] Anonymous. Decisions of the courts. Scientific American, 36 (24):377–378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
377a.pdf.

Anonymous:1877:RAFw

[10155] Anonymous. Recent American and foreign patents. Scientific American, 36(24):378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
378.pdf.

Anonymous:1877:NMEr

[10156] Anonymous. New mechanical and engineering inventions. Scientific American, 36(24):378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
378a.pdf.

Anonymous:1877:NHIn

[10157] Anonymous. New household inventions. Scientific American, 36 (24):378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
378b.pdf.

Anonymous:1877:NMIw

[10158] Anonymous. New miscellaneous inventions. Scientific American, 36(24):378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
378c.pdf.

Anonymous:1877:NWHp

[10159] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(24):378, June 16, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n24/pdf/scientificamerican06161877-378d.pdf.

Anonymous:1877:NAIr

[10160] Anonymous. New agricultural inventions. Scientific American, 36 (24):378, June 16, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n24/pdf/scientificamerican06161877-
378e.pdf.

Anonymous:1877:BPy

[10161] Anonymous. Business and personal. Scientific American, 36(25):395, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-395.pdf.

Anonymous:1877:NQx

[10162] Anonymous. Notes and queries. Scientific American, 36(25):395–396, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-395a.pdf.

Anonymous:1877:IIx

[10163] Anonymous. Index of inventions. Scientific American, 36(25):396, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-396a.pdf.

Anonymous:1877:PNPw

[10164] Anonymous. Patent notes: Patents. Scientific American, 36(25):398, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-398.pdf.

Anonymous:1877:IFEb

[10165] Anonymous. Improved fire-extinguishing apparatus for vessels. Scientific American, 36(25):383–384, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
383.pdf.

Anonymous:1877:BGB

[10166] Anonymous. Blue glass blindness. Scientific American, 36(25):384, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-384.pdf.

Anonymous:1877:DD

[10167] Anonymous. Drunk or diseased? Scientific American, 36(25):384, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-384a.pdf.

Anonymous:1877:AGa

[10168] Anonymous. About gravestones. Scientific American, 36(25):384–385, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-384b.pdf.

Anonymous:1877:HI

[10169] Anonymous. Helping inventors. Scientific American, 36(25):385, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-385.pdf.

Anonymous:1877:LR

[10170] Anonymous. Lightning rods. Scientific American, 36(25):385, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-385a.pdf.

Anonymous:1877:IDPa

[10171] Anonymous. Improved diagonal planing and polishing machine. Scientific American, 36(25):386, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
386.pdf.

Anonymous:1877:ICMc

[10172] Anonymous. Improved calf muzzle. Scientific American, 36(25):386, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-386a.pdf.

Anonymous:1877:IHB

[10173] Anonymous. Improved horse boot. Scientific American, 36(25):386, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-386b.pdf.

Rose:1877:PMl

[10174] Joshua Rose. Practical mechanism. Scientific American, 36(25): 387–388, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
387.pdf.

Anonymous:1877:Cbb

[10175] Anonymous. Communications. Scientific American, 36(25):388–389, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-388.pdf.

Anonymous:1877:NGE

[10176] Anonymous. The new grain elevator in New York City. Scientific American, 36(25):390–391, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
390.pdf.

Riley:1877:H

[10177] C. V. Riley. The hellgrammite. Scientific American, 36(25):392–393, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-392.pdf.

Anonymous:1877:DCl

[10178] Anonymous. Decisions of the courts. Scientific American, 36 (25):393–394, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
393.pdf.

Anonymous:1877:RAFx

[10179] Anonymous. Recent American and foreign patents. Scientific American, 36(25):394, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
394.pdf.

Anonymous:1877:NMEs

[10180] Anonymous. New mechanical and engineering inventions. Scientific American, 36(25):394, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
394a.pdf.

Anonymous:1877:NMIx

[10181] Anonymous. New miscellaneous inventions. Scientific American, 36(25):394, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
394b.pdf.

Anonymous:1877:NTIh

[10182] Anonymous. New textile invention. Scientific American, 36(25):394, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n25/pdf/scientificamerican06231877-394c.pdf.

Anonymous:1877:NAIs (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
394d.pdf.

Anonymous:1877:NWHq

[10184] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(25):394, June 23, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n25/pdf/scientificamerican06231877-394e.pdf.

Anonymous:1877:CRx

[10185] Anonymous. Communications received. Scientific American, 36 (25):396, June 23, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n25/pdf/scientificamerican06231877-
396.pdf.

Anonymous:1877:BNBp

[10186] Anonymous. Books: New books and publications. Scientific American, 36(26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409a.pdf.

Anonymous:1877:BPz

[10187] Anonymous. Business and personal. Scientific American, 36(26):410, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-410.pdf.

Anonymous:1877:NQy

[10188] Anonymous. Notes and queries. Scientific American, 36(26):410, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-410a.pdf.

Anonymous:1877:IIy

[10189] Anonymous. Index of inventions. Scientific American, 36(26):410–412, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-410c.pdf.

Anonymous:1877:PNPx

[10190] Anonymous. Patent notes: Patents. Scientific American, 36(26):414, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-414.pdf.

Anonymous:1877:SN

[10191] Anonymous. A steam navvy. Scientific American, 36(26):399, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-399.pdf.

Anonymous:1877:LPV

[10192] Anonymous. The limits and powers of vision. Scientific American, 36(26):400, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
400.pdf.

Anonymous:1877:Tb

[10193] Anonymous. Trance. Scientific American, 36(26):400–401, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-400a.pdf.

Anonymous:1877:BTb

[10194] Anonymous. The banian tree. Scientific American, 36(26):401, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-401.pdf.

Anonymous:1877:VNT

[10195] Anonymous. Volpicelli’s new theory of electro-static induction. Scientific American, 36(26):401, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
401a.pdf.

Anonymous:1877:MHTb

[10196] Anonymous. Manufacture of the Harvey torpedo. Scientific American, 36(26):402, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
402.pdf.

Anonymous:1877:IPP

[10197] Anonymous. Improved paper-pulp engine. Scientific American, 36(26):403, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
403.pdf.

Anonymous:1877:PEM

[10198] Anonymous. Pneumatic elevation in mines. Scientific American, 36(26):403–404, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
403a.pdf.

Anonymous:1877:Cbc

[10199] Anonymous. Communications. Scientific American, 36(26):404, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-404.pdf.

Hallock:1877:NLE

[10200] E. J. Hallock. A new lecture experiment. Scientific American, 36(26):405, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
405.pdf.

Anonymous:1877:SMC

[10201] Anonymous. The secrets of the mysterious clocks. Scientific American, 36(26):405, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
405a.pdf.

Anonymous:1877:FVS

[10202] Anonymous. Frost’s variety self-centering chuck. Scientific American, 36(26):406, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
406.pdf.

Anonymous:1877:NMSb

[10203] Anonymous. A new method of secret writing. Scientific American, 36(26):406, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
406a.pdf.

Anonymous:1877:IBTa

[10204] Anonymous. Improved ballasting tube for vessels. Scientific American, 36(26):406, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
406b.pdf.

Anonymous:1877:IG

[10205] Anonymous. Improved gauge. Scientific American, 36(26):406, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-406c.pdf.

Anonymous:1877:IMa

[10206] Anonymous. An interesting marmoset. Scientific American, 36(26):407, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-407.pdf.

Anonymous:1877:AMa

[10207] Anonymous. An artificial mammoth. Scientific American, 36 (26):407–409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
407a.pdf.

Anonymous:1877:DCm

[10208] Anonymous. Decisions of the courts. Scientific American, 36(26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v36/n26/pdf/scientificamerican06301877-409.pdf.

Anonymous:1877:RAFy 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409b.pdf.

Anonymous:1877:NWHr

[10210] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 36(26):409, June 30, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v36/
n26/pdf/scientificamerican06301877-409c.pdf.

Anonymous:1877:NHIo

[10211] Anonymous. New household inventions. Scientific American, 36 (26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409d.pdf.

Anonymous:1877:NMIy

[10212] Anonymous. New miscellaneous inventions. Scientific American, 36(26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409e.pdf.

Anonymous:1877:NMEt

[10213] Anonymous. New mechanical and engineering inventions. Scientific American, 36(26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409f.pdf.

Anonymous:1877:NAIt

[10214] Anonymous. New agricultural inventions. Scientific American, 36 (26):409, June 30, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
409g.pdf.

Anonymous:1877:CRy (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v36/n26/pdf/scientificamerican06301877-
410b.pdf.

Anonymous:1877:BPba

[10216] Anonymous. Business and personal. Scientific American, 37(1):11, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-11.pdf.

Anonymous:1877:NQNb

[10217] Anonymous. Notes and queries: Notes & queries. Scientific American, 37(1):11–12, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
11a.pdf.

Anonymous:1877:IIz

[10218] Anonymous. Index of inventions. Scientific American, 37(1):12, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-12a.pdf.

Anonymous:1877:PNPy

[10219] Anonymous. Patent notes: Patents. Scientific American, 37(1):14, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-14.pdf.

Anonymous:1877:MMS

[10220] Anonymous. A mew method of steamboat propulsion. Scientific American, 37(1):1, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
1.pdf.

Anonymous:1877:NPa

[10221] Anonymous. The nature of pain. Scientific American, 37(1):2, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-2.pdf.

Anonymous:1877:SSSa

[10222] Anonymous. Scientific sight-seeing. Scientific American, 37(1):2, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-2a.pdf.

Anonymous:1877:CDM

[10223] Anonymous. Croup due to miasma. Scientific American, 37(1):2–3, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-2b.pdf.

Anonymous:1877:TDP

[10224] Anonymous. The torpedo defence problem. Scientific American, 37(1): 3, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-3.pdf.

Anonymous:1877:NSPa

[10225] Anonymous. A new system of petroleum storage. Scientific American, 37(1):4–5, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
4.pdf.

Anonymous:1877:NMBa

[10226] Anonymous. A new mechanical butter-worker. Scientific American, 37(1):5, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
5.pdf.

Anonymous:1877:ISBa

[10227] Anonymous. Improved shot bag and charger. Scientific American, 37(1):6, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
6.pdf.

Anonymous:1877:NHA

[10228] Anonymous. A new hot-air engine. Scientific American, 37(1):6, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-6a.pdf.

Anonymous:1877:IPGb

[10229] Anonymous. Improved pressure gauge. Scientific American, 37(1):6, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-6b.pdf.

Anonymous:1877:AGb

[10230] Anonymous. An Australian gull. Scientific American, 37(1):7, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-7.pdf.

Anonymous:1877:FAc

[10231] Anonymous. The fountains at Aranjuez. Scientific American, 37(1):7, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-7a.pdf.

Anonymous:1877:Pf

[10232] Anonymous. The penguins. Scientific American, 37(1):7, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-7b.pdf.

Rose:1877:PMm

[10233] Joshua Rose. Practical mechanism. Scientific American, 37(1):8, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-8.pdf.

Anonymous:1877:Cbd

[10234] Anonymous. Communications. Scientific American, 37(1):9, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-9.pdf.

Anonymous:1877:ANg

[10235] Anonymous. Astronomical notes. Scientific American, 37(1):9–10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-9a.pdf.

Anonymous:1877:DCn

[10236] Anonymous. Decisions of the courts. Scientific American, 37(1):10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-10.pdf.

Anonymous:1877:RAFz

[10237] Anonymous. Recent American and foreign patents. Scientific American, 37(1):10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
10a.pdf.

Anonymous:1877:NAIu

[10238] Anonymous. New agricultural inventions. Scientific American, 37(1): 10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-10b.pdf.

Anonymous:1877:NMEu

[10239] Anonymous. New mechanical and engineering inventions. Scientific American, 37(1):10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n1/pdf/scientificamerican07071877-
10c.pdf.

Anonymous:1877:NWHs

[10240] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(1):10, July 7, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n1/pdf/scientificamerican07071877-10d.pdf.

Anonymous:1877:NMIz

[10241] Anonymous. New miscellaneous inventions. Scientific American, 37(1): 10, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-10e.pdf.

Anonymous:1877:CRz

[10242] Anonymous. Communications received. Scientific American, 37(1):12, July 7, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n1/pdf/scientificamerican07071877-12.pdf.

Anonymous:1877:BNBq

[10243] Anonymous. Books: New books and publications. Scientific American, 37(2):25, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
25a.pdf.

Anonymous:1877:BPbb

[10244] Anonymous. Business and personal. Scientific American, 37(2):27, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-27.pdf.

Anonymous:1877:NQz

[10245] Anonymous. Notes and queries. Scientific American, 37(2):27–28, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-27a.pdf.

Anonymous:1877:IIba

[10246] Anonymous. Index of inventions. Scientific American, 37(2):28, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-28a.pdf.

Anonymous:1877:PNPz

[10247] Anonymous. Patent notes: Patents. Scientific American, 37(2):30, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-30.pdf.

Anonymous:1877:BPAa

[10248] Anonymous. Bower’s patent air compressor. Scientific American, 37(2):15, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
15.pdf.

Anonymous:1877:CPAa

[10249] Anonymous. Carrying peace into Africa. Scientific American, 37(2):16, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-16.pdf.

Anonymous:1877:ACK

[10250] Anonymous. Address of Clarence King on catastrophism. Scientific American, 37(2):16–17, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
16a.pdf.

Anonymous:1877:DEa

[10251] Anonymous. Duplex education. Scientific American, 37(2):17, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-17.pdf.

Anonymous:1877:CCa

[10252] Anonymous. The Carrigeen crop. Scientific American, 37(2):17, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-17a.pdf.

Anonymous:1877:CGb

[10253] Anonymous. Catastrophism in geology. Scientific American, 37(2):17, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-17b.pdf.

Anonymous:1877:NAS

[10254] Anonymous. A new apparatus for storing and utilizing solar heat. Scientific American, 37(2):18, July 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n2/pdf/scientificamerican07141877-18.pdf.

Anonymous:1877:UTSb

[10255] Anonymous. Utilization of tin scrap. Scientific American, 37(2):18, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-18a.pdf.

Anonymous:1877:NMBb

[10256] Anonymous. A new method of bookbinding. Scientific American, 37(2):19, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
19.pdf.

Anonymous:1877:ICF

[10257] Anonymous. Improved curtain fixture. Scientific American, 37(2):19, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-19a.pdf.

Anonymous:1877:BSM

[10258] Anonymous. Boot and shoe machinery. Scientific American, 37 (2):19–20, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
19b.pdf.

Rose:1877:SLS

[10259] Joshua Rose. Setting locomotive slide valves. Scientific American, 37(2):21, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
21.pdf.

Anonymous:1877:CIB

[10260] Anonymous. Collender’s improved billiard table. Scientific American, 37(2):22, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
22.pdf.

Anonymous:1877:IBB

[10261] Anonymous. Improved billiard ball holder. Scientific American, 37(2):22, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
22a.pdf.

Anonymous:1877:MCC

[10262] Anonymous. The monitor chalk cup. Scientific American, 37(2):22, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-22b.pdf.

Anonymous:1877:CCP

[10263] Anonymous. Curious carnivorous plants. Scientific American, 37 (2):23–25, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
23.pdf.

Anonymous:1877:DCo

[10264] Anonymous. Decisions of the courts. Scientific American, 37(2):25, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-25.pdf.

Anonymous:1877:RAFba

[10265] Anonymous. Recent American and foreign patents. Scientific American, 37(2):25, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
25b.pdf.

Anonymous:1877:NAIv

[10266] Anonymous. New agricultural inventions. Scientific American, 37(2):25–26, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
25c.pdf.

Anonymous:1877:NTIi

[10267] Anonymous. New textile inventions. Scientific American, 37(2):26, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-26.pdf.

Anonymous:1877:NWHt

[10268] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(2):26, July 14, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n2/pdf/scientificamerican07141877-26a.pdf.

Anonymous:1877:NHIp

[10269] Anonymous. New household inventions. Scientific American, 37(2):26, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-26b.pdf.

Anonymous:1877:NMEv

[10270] Anonymous. New mechanical and engineering inventions. Scientific American, 37(2):26, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
26c.pdf.

Anonymous:1877:NMIba

[10271] Anonymous. New miscellaneous inventions. Scientific American, 37(2):26, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n2/pdf/scientificamerican07141877-
26d.pdf.

Anonymous:1877:CRba

[10272] Anonymous. Communications received. Scientific American, 37(2):28, July 14, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n2/pdf/scientificamerican07141877-28.pdf.

Anonymous:1877:BPbc

[10273] Anonymous. Business and personal. Scientific American, 37(3):43, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-43.pdf.

Anonymous:1877:NQba

[10274] Anonymous. Notes and queries. Scientific American, 37(3):43–44, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-43a.pdf.

Anonymous:1877:IIbb

[10275] Anonymous. Index of inventions. Scientific American, 37(3):44, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-44a.pdf.

Anonymous:1877:PNPba

[10276] Anonymous. Patent notes: Patents. Scientific American, 37(3):46, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-46.pdf.

Anonymous:1877:MSL

[10277] Anonymous. The manufacture of sheet lead and lead pipe. Scientific American, 37(3):31, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
31.pdf.

Anonymous:1877:IS

[10278] Anonymous. An inexact science. Scientific American, 37(3):32, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-32.pdf.

Anonymous:1877:GPb

[10279] Anonymous. The great problem. Scientific American, 37(3):32–33, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-32a.pdf.

Anonymous:1877:PCAa

[10280] Anonymous. The physical characteristics of American children. Scientific American, 37(3):33, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
33.pdf.

Anonymous:1877:LLC

[10281] Anonymous. Large locomotive crane. Scientific American, 37(3):34–35, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-34.pdf.

Anonymous:1877:GEH

[10282] Anonymous. The great eruption at Hawaii. Scientific American, 37(3):35, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
35.pdf.

Anonymous:1877:Cbe

[10283] Anonymous. Communications. Scientific American, 37(3):36–37, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-36.pdf.

Rose:1877:PMn

[10284] Joshua Rose. Practical mechanism. Scientific American, 37(3):37, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-37.pdf.

Anonymous:1877:IHCb

[10285] Anonymous. Improved horse collar stuffing machine. Scientific American, 37(3):38, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
38.pdf.

Anonymous:1877:GBa

[10286] Anonymous. A gaspipe bedstead. Scientific American, 37(3):38, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-38a.pdf.

Anonymous:1877:NIT

[10287] Anonymous. New ice tongs. Scientific American, 37(3):38, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-38b.pdf.

Anonymous:1877:SGH

[10288] Anonymous. Steam glue-heating apparatus. Scientific American, 37(3):38, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
38c.pdf.

Anonymous:1877:TW

[10289] Anonymous. Torpedo warfare. Scientific American, 37(3):39–40, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-39.pdf.

Anonymous:1877:ADI

[10290] Anonymous. Are the aniline dyes injurious to health? Scientific American, 37(3):40–41, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
40.pdf.

Anonymous:1877:DCp

[10291] Anonymous. Decisions of the courts. Scientific American, 37(3):41, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-41.pdf.

Anonymous:1877:RAFbb

[10292] Anonymous. Recent American and foreign patents. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
42.pdf.

Anonymous:1877:NMEw

[10293] Anonymous. New mechanical and engineering inventions. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
42a.pdf.

Anonymous:1877:NTIj

[10294] Anonymous. New textile inventions. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-42b.pdf.

Anonymous:1877:NWHu

[10295] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(3):42, July 21, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n3/pdf/scientificamerican07211877-42c.pdf.

Anonymous:1877:NAIw

[10296] Anonymous. New agricultural inventions. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-42d.pdf.

Anonymous:1877:NHIq

[10297] Anonymous. New household inventions. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-42e.pdf.

Anonymous:1877:NMIbb

[10298] Anonymous. New miscellaneous inventions. Scientific American, 37(3):42, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n3/pdf/scientificamerican07211877-
42f.pdf.

Anonymous:1877:CRbb

[10299] Anonymous. Communications received. Scientific American, 37(3):44, July 21, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n3/pdf/scientificamerican07211877-44.pdf.

Anonymous:1877:BNBr

[10300] Anonymous. Books: New books and publications. Scientific American, 37(4):57, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
57.pdf.

Anonymous:1877:BPbd

[10301] Anonymous. Business and personal. Scientific American, 37(4):59, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-59.pdf.

Anonymous:1877:NQbb

[10302] Anonymous. Notes and queries. Scientific American, 37(4):59–60, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-59a.pdf.

Anonymous:1877:IIbc

[10303] Anonymous. Index of inventions. Scientific American, 37(4):60, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-60a.pdf.

Anonymous:1877:PNPbb

[10304] Anonymous. Patent notes: Patents. Scientific American, 37(4):62, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-62.pdf.

Anonymous:1877:MSc

[10305] Anonymous. The manufacture of shot. Scientific American, 37(4):47, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-47.pdf.

Anonymous:1877:LBN

[10306] Anonymous. The latest British Naval failure. Scientific American, 37(4):48, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
48.pdf.

Anonymous:1877:GAa

[10307] Anonymous. Glimpses of Atlantis. Scientific American, 37(4):48–49, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-48a.pdf.

Anonymous:1877:EMA

[10308] Anonymous. An English mechanic on American machines. Scientific American, 37(4):49, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
49.pdf.

Anonymous:1877:PMJ

[10309] Anonymous. Progress of the Mississippi jetty works. Scientific American, 37(4):49, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
49a.pdf.

Anonymous:1877:ARF

[10310] Anonymous. American representation at the French Exposition. Scientific American, 37(4):49, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
49b.pdf.

Anonymous:1877:FFR

[10311] Anonymous. Fast freight railway trains. Scientific American, 37(4):49, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-49c.pdf.

Anonymous:1877:SSTa

[10312] Anonymous. A Swiss steam tramway car. Scientific American, 37(4):50, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-50.pdf.

Anonymous:1877:IFa

[10313] Anonymous. Improved faucet. Scientific American, 37(4):50, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-50a.pdf.

Anonymous:1877:IMP

[10314] Anonymous. Improved metal punching machine. Scientific American, 37(4):50, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
50b.pdf.

Anonymous:1877:IRGb

[10315] Anonymous. Improved running gear for wagons. Scientific American, 37(4):51, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
51.pdf.

Anonymous:1877:AIB

[10316] Anonymous. Angle iron bending machine. Scientific American, 37(4):51, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
51a.pdf.

Anonymous:1877:SSc

[10317] Anonymous. Science and schools. Scientific American, 37(4):52, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-52.pdf.

Anonymous:1877:SWA

[10318] Anonymous. Setting work after casehardening. Scientific American, 37(4):53, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
53.pdf.

Anonymous:1877:GIO

[10319] Anonymous. Graves’ improved oil can. Scientific American, 37(4):54, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-54.pdf.

Anonymous:1877:HDb

[10320] Anonymous. Horse detacher. Scientific American, 37(4):54, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-54a.pdf.

Anonymous:1877:NLB

[10321] Anonymous. Novel life boat. Scientific American, 37(4):54, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-54b.pdf.

Anonymous:1877:Tc

[10322] Anonymous. The tunny. Scientific American, 37(4):55, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-55.pdf.

Anonymous:1877:LF

[10323] Anonymous. The lantern fly. Scientific American, 37(4):55, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-55a.pdf.

Anonymous:1877:AFB

[10324] Anonymous. An artificial flower that blooms. Scientific American, 37(4):56, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
56.pdf.

Anonymous:1877:Ba

[10325] Anonymous. Brushes. Scientific American, 37(4):56, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-56a.pdf.

Anonymous:1877:CCC

[10326] Anonymous. A cigar-cutting cigar box. Scientific American, 37 (4):56–57, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
56b.pdf.

Anonymous:1877:RAFbc

[10327] Anonymous. Recent American and foreign patents. Scientific American, 37(4):58, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
58.pdf.

Anonymous:1877:NAIx

[10328] Anonymous. New agricultural inventions. Scientific American, 37(4):58, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-58a.pdf.

Anonymous:1877:NMEx

[10329] Anonymous. New mechanical and engineering inventions. Scientific American, 37(4):58, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
58b.pdf.

Anonymous:1877:NWHv

[10330] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(4):58, July 28, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n4/pdf/scientificamerican07281877-58c.pdf.

Anonymous:1877:NMIbc

[10331] Anonymous. New miscellaneous inventions. Scientific American, 37(4):58, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n4/pdf/scientificamerican07281877-
58d.pdf.

Anonymous:1877:NHIr

[10332] Anonymous. New household inventions. Scientific American, 37(4):58, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-58e.pdf.

Anonymous:1877:CRbc

[10333] Anonymous. Communications received. Scientific American, 37(4):60, July 28, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n4/pdf/scientificamerican07281877-60.pdf.

Anonymous:1877:BPbe

[10334] Anonymous. Business and personal. Scientific American, 37(5):75, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-75.pdf.

Anonymous:1877:NQbc

[10335] Anonymous. Notes and queries. Scientific American, 37(5):75–76, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-75a.pdf.

Anonymous:1877:IIbd

[10336] Anonymous. Index of inventions. Scientific American, 37(5):76, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-76a.pdf.

Anonymous:1877:PNPbc

[10337] Anonymous. Patent notes: Patents. Scientific American, 37(5):78, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-78.pdf.

Anonymous:1877:CME

[10338] Anonymous. Cable making of the East River Bridge. Scientific American, 37(5):63, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
63.pdf.

Anonymous:1877:TDI

[10339] Anonymous. Torpedo defence — invention wanted. Scientific American, 37(5):64, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
64.pdf.

Anonymous:1877:NPOa

[10340] Anonymous. Notes of Patent Office decisions. Scientific American, 37(5):64, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
64a.pdf.

Anonymous:1877:RGI

[10341] Anonymous. Recent great improvements in electric illumination. Scientific American, 37(5):64–65, August 4, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n5/pdf/scientificamerican08041877-64b.pdf.

Anonymous:1877:NIC

[10342] Anonymous. New investigations on copper alloys. Scientific American, 37(5):65, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
65.pdf.

Anonymous:1877:BEW

[10343] Anonymous. Balancing emery wheels. Scientific American, 37(5):65, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-65a.pdf.

Anonymous:1877:GRS

[10344] Anonymous. The great railroad strikes. Scientific American, 37(5):65, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-65b.pdf.

Anonymous:1877:RMB

[10345] Anonymous. A rotary mould board plow. Scientific American, 37(5):66, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-66.pdf.

Anonymous:1877:APB

[10346] Anonymous. Apparatus for preparing butter for packing. Scientific American, 37(5):66, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
66a.pdf.

Anonymous:1877:IGSb

[10347] Anonymous. Improved grain scale. Scientific American, 37(5):66, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-66b.pdf.

Anonymous:1877:ND

[10348] Anonymous. A new dynamometer. Scientific American, 37(5):66, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-66c.pdf.

Anonymous:1877:AT

[10349] Anonymous. The American toad. Scientific American, 37(5):67, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-67.pdf.

Anonymous:1877:VD

[10350] Anonymous. The vapart disintegrator. Scientific American, 37 (5):67–68, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
67a.pdf.

Anonymous:1877:Cbf

[10351] Anonymous. Communications. Scientific American, 37(5):68–69, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-68.pdf.

Anonymous:1877:NPc

[10352] Anonymous. New pump. Scientific American, 37(5):70, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-70.pdf.

Anonymous:1877:ICCa

[10353] Anonymous. Improved chicken coop. Scientific American, 37(5):70, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-70a.pdf.

Anonymous:1877:SIR

[10354] Anonymous. Spelman’s improved rowlock. Scientific American, 37(5):70, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
70b.pdf.

Anonymous:1877:CLa

[10355] Anonymous. The Carolina lizard. Scientific American, 37(5):71, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-71.pdf.

Anonymous:1877:SMb

[10356] Anonymous. The sea monster. Scientific American, 37(5):71, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-71a.pdf.

Anonymous:1877:MCb

[10357] Anonymous. The mole Cricket. Scientific American, 37(5):71–72, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-71b.pdf.

Anonymous:1877:SCG

[10358] Anonymous. A self-cooling goblet. Scientific American, 37(5):72–74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-72.pdf.

Anonymous:1877:RAFbd

[10359] Anonymous. Recent American and foreign patents. Scientific American, 37(5):74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
74.pdf.

Anonymous:1877:NMEy

[10360] Anonymous. New mechanical and engineering inventions. Scientific American, 37(5):74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
74a.pdf.

Anonymous:1877:NMIbd

[10361] Anonymous. New miscellaneous inventions. Scientific American, 37(5):74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n5/pdf/scientificamerican08041877-
74b.pdf.

Anonymous:1877:NWHw

[10362] Anonymous. New woodworking house and carriage building inventions. Scientific American, 37(5):74, August 4, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n5/pdf/scientificamerican08041877-74c.pdf.

Anonymous:1877:NAIy

[10363] Anonymous. New agricultural inventions. Scientific American, 37(5):74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-74d.pdf.

Anonymous:1877:NHIs

[10364] Anonymous. New household inventions. Scientific American, 37(5):74, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-74e.pdf.

Anonymous:1877:CRbd

[10365] Anonymous. Communications received. Scientific American, 37(5):76, August 4, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n5/pdf/scientificamerican08041877-76.pdf.

Anonymous:1877:BPbf

[10366] Anonymous. Business and personal. Scientific American, 37(6):91, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-91.pdf.

Anonymous:1877:NQbd

[10367] Anonymous. Notes and queries. Scientific American, 37(6):91–92, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-91a.pdf.

Anonymous:1877:IIbe

[10368] Anonymous. Index of inventions. Scientific American, 37(6):92, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-92a.pdf.

Anonymous:1877:PNPbd

[10369] Anonymous. Patent notes: Patents. Scientific American, 37(6):94, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-94.pdf.

Anonymous:1877:DPE

[10370] Anonymous. Despaquls’ photo-engraving process. Scientific American, 37(6):79, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
79.pdf.

Anonymous:1877:CFE

[10371] Anonymous. Cable fastenings of the East River Bridge. Scientific American, 37(6):79, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
79a.pdf.

Anonymous:1877:DLb

[10372] Anonymous. The degradation of labor. Scientific American, 37 (6):80, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
80.pdf.

Anonymous:1877:ISR

[10373] Anonymous. Iron and steel in railway car construction. Scientific American, 37(6):80, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
80a.pdf.

Anonymous:1877:AS

[10374] Anonymous. About sparrows. Scientific American, 37(6):80–81, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-80b.pdf.

Anonymous:1877:SPA

[10375] Anonymous. Selling patents by auction. Scientific American, 37 (6):81, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
81.pdf.

Anonymous:1877:HPC

[10376] Anonymous. The Howgate polar colony. Scientific American, 37 (6):81, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
81a.pdf.

Anonymous:1877:Cbg

[10377] Anonymous. Chess. Scientific American, 37(6):81, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-81b.pdf.

Anonymous:1877:NT

[10378] Anonymous. A new telephone. Scientific American, 37(6):83, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-83.pdf.

Anonymous:1877:IDPb

[10379] Anonymous. Improved donkey pump. Scientific American, 37 (6):83, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
83a.pdf.

Anonymous:1877:Cbh

[10380] Anonymous. Communications. Scientific American, 37(6):84–85, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-84.pdf.

Anonymous:1877:ANh

[10381] Anonymous. Astronomical notes. Scientific American, 37(6):85, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-85.pdf.

Anonymous:1877:PSC

[10382] Anonymous. Postage stamp canceller. Scientific American, 37 (6):86, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
86.pdf.

Anonymous:1877:BBC

[10383] Anonymous. Browne’s book cover protector. Scientific American, 37(6):86, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
86a.pdf.

Anonymous:1877:ILTb

[10384] Anonymous. Improved log turner. Scientific American, 37(6):86, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-86b.pdf.

Anonymous:1877:MCd

[10385] Anonymous. Micrometer caliper. Scientific American, 37(6):86, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-86c.pdf.

Seiss:1877:BB

[10386] C. Few Seiss. The brown bat. Scientific American, 37(6):87, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-87.pdf.

Anonymous:1877:WEJ

[10387] Anonymous. Wood engraving in Japan. Scientific American, 37 (6):87, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
87a.pdf.

Anonymous:1877:K

[10388] Anonymous. Knives. Scientific American, 37(6):88, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-88.pdf.

Anonymous:1877:EMc

[10389] Anonymous. Exterminating moles. Scientific American, 37(6):88, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n6/pdf/scientificamerican08111877-88a.pdf.

Anonymous:1877:WIHa

[10390] Anonymous. The watch — its invention and history. Scientific American, 37(6):89, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
89.pdf.

Anonymous:1877:RAFbe

[10391] Anonymous. Recent American and foreign patents. Scientific American, 37(6):90, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
90.pdf.

Anonymous:1877:NMIbe

[10392] Anonymous. New miscellaneous inventions. Scientific American, 37(6):90, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
90a.pdf.

Anonymous:1877:NHIt

[10393] Anonymous. New household inventions. Scientific American, 37 (6):90, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
90b.pdf.

Anonymous:1877:NWHx

[10394] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(6):90, August 11, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n6/pdf/scientificamerican08111877-90c.pdf.

Anonymous:1877:NAIz

[10395] Anonymous. New agricultural inventions. Scientific American, 37(6):90, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
90d.pdf.

Anonymous:1877:NMEz

[10396] Anonymous. New mechanical and engineering inventions. Scientific American, 37(6):90, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
90e.pdf.

Anonymous:1877:CRbe

[10397] Anonymous. Communications received. Scientific American, 37 (6):92, August 11, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n6/pdf/scientificamerican08111877-
92.pdf.

Anonymous:1877:BNBs

[10398] Anonymous. Books: New books and publications. Scientific American, 37(7):105, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
105.pdf.

Anonymous:1877:BPbg

[10399] Anonymous. Business and personal. Scientific American, 37(7):107, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-107.pdf.

Anonymous:1877:NQbe

[10400] Anonymous. Notes and queries. Scientific American, 37(7):107– 108, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
107a.pdf.

Anonymous:1877:IIbf

[10401] Anonymous. Index of inventions. Scientific American, 37(7):108, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-108a.pdf.

Anonymous:1877:PNPbe

[10402] Anonymous. Patent notes: Patents. Scientific American, 37(7):110, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-110.pdf.

Anonymous:1877:IIMa

[10403] Anonymous. Improved ice machine. Scientific American, 37(7):95, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-95.pdf.

Anonymous:1877:PIR

[10404] Anonymous. Poison ivy and its remedies. Scientific American, 37(7):96, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
96.pdf.

Anonymous:1877:NTS

[10405] Anonymous. A new theory of the solar spectrum. Scientific American, 37(7):96, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
96a.pdf.

Anonymous:1877:DOS

[10406] Anonymous. Discovery of oxygen in the Sun by photography. Scientific American, 37(7):96, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
96b.pdf.

Anonymous:1877:EF

[10407] Anonymous. An electric fire. Scientific American, 37(7):96–97, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-96c.pdf.

Anonymous:1877:SLN

[10408] Anonymous. Sunshine in London and New York. Scientific American, 37(7):97, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
97.pdf.

Anonymous:1877:DEb

[10409] Anonymous. Dimness of the eyes. Scientific American, 37(7):97, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-97a.pdf.

Anonymous:1877:IW

[10410] Anonymous. Improved washboiler. Scientific American, 37(7):98, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-98.pdf.

Anonymous:1877:IDb

[10411] Anonymous. Improved dustpan. Scientific American, 37(7):98, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-98a.pdf.

Anonymous:1877:IFCa

[10412] Anonymous. Improved feed cutter. Scientific American, 37(7):98, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-98b.pdf.

Anonymous:1877:ICMd

[10413] Anonymous. Improved concrete mixer. Scientific American, 37 (7):99, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
99.pdf.

Anonymous:1877:Cbi

[10414] Anonymous. Communications. Scientific American, 37(7):100, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-100.pdf.

Anonymous:1877:WIHb

[10415] Anonymous. The watch — its invention and history. Scientific American, 37(7):101, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
101.pdf.

Anonymous:1877:ICL

[10416] Anonymous. Improved carpet loom. Scientific American, 37(7): 102–103, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
102.pdf.

Seiss:1877:WSH

[10417] C. Few Seiss. The water snake — how it eats. Scientific American, 37(7):103, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
103.pdf.

Anonymous:1877:CHA

[10418] Anonymous. The curly-haired antelope. Scientific American, 37 (7):103–104, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
103a.pdf.

Anonymous:1877:KG

[10419] Anonymous. Kid gloves. Scientific American, 37(7):104, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n7/pdf/scientificamerican08181877-104.pdf.

Anonymous:1877:FFB

[10420] Anonymous. A floating flower bed. Scientific American, 37(7): 104–105, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
104a.pdf.

Anonymous:1877:RAFbf

[10421] Anonymous. Recent American and foreign patents. Scientific American, 37(7):106, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
106.pdf.

Anonymous:1877:NMEba

[10422] Anonymous. New mechanical and engineering inventions. Scientific American, 37(7):106, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
106a.pdf.

Anonymous:1877:NMIbf

[10423] Anonymous. New miscellaneous inventions. Scientific American, 37(7):106, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
106b.pdf.

Anonymous:1877:NHIu

[10424] Anonymous. New household inventions. Scientific American, 37 (7):106, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
106c.pdf.

Anonymous:1877:NWHy

[10425] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(7):106, August 18, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n7/pdf/scientificamerican08181877-106d.pdf.

Anonymous:1877:NAIba

[10426] Anonymous. New agricultural inventions. Scientific American, 37(7):106, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
106e.pdf.

Anonymous:1877:CRbf

[10427] Anonymous. Communications received. Scientific American, 37 (7):108, August 18, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n7/pdf/scientificamerican08181877-
108.pdf.

Anonymous:1877:BPbh

[10428] Anonymous. Business and personal. Scientific American, 37(8):123, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-123.pdf.

Anonymous:1877:NQbf

[10429] Anonymous. Notes and queries. Scientific American, 37(8):123– 124, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
123a.pdf.

Anonymous:1877:IIbg

[10430] Anonymous. Index of inventions. Scientific American, 37(8):124, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-124a.pdf.

Anonymous:1877:PNPbf

[10431] Anonymous. Patent notes: Patents. Scientific American, 37(8):126, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-126.pdf.

Anonymous:1877:RDM

[10432] Anonymous. Radial drilling machine. Scientific American, 37 (8):111, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
111.pdf.

Anonymous:1877:FE

[10433] Anonymous. Fire engine. Scientific American, 37(8):111–112, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-111a.pdf.

Anonymous:1877:DWR

[10434] Anonymous. Dead weight in railway cars. Scientific American, 37(8):112, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
112.pdf.

Anonymous:1877:PIP

[10435] Anonymous. Pens, ink, and pencils. Scientific American, 37 (8):112–113, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
112a.pdf.

Anonymous:1877:SNI

[10436] Anonymous. Some new inventions necessary for future astronomical observations. Scientific American, 37(8):113, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n8/pdf/scientificamerican08251877-113.pdf.

Anonymous:1877:HFP

[10437] Anonymous. Hereditary as a factor in pauperism and crime. Scientific American, 37(8):113, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
113a.pdf.

Anonymous:1877:KMS

[10438] Anonymous. Keely motor stock at a discount. Scientific American, 37(8):113–114, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
113b.pdf.

Anonymous:1877:MFS

[10439] Anonymous. Machine for forging screw threads. Scientific American, 37(8):114–115, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
114.pdf.

Anonymous:1877:IFCb

[10440] Anonymous. Improved fodder cutter. Scientific American, 37 (8):115, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
115.pdf.

Anonymous:1877:Cbj

[10441] Anonymous. Communications. Scientific American, 37(8):116–117, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-116.pdf.

Rose:1877:PMo

[10442] Joshua Rose. Practical mechanism. Scientific American, 37(8):117, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-117.pdf.

Anonymous:1877:IVW

[10443] Anonymous. Improved vehicle wheel. Scientific American, 37 (8):118, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
118.pdf.

Anonymous:1877:IAFa

[10444] Anonymous. Improved automatic fog signal. Scientific American, 37(8):118, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
118a.pdf.

Anonymous:1877:IUS

[10445] Anonymous. Improved universal square. Scientific American, 37 (8):118, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
118b.pdf.

Anonymous:1877:HMR

[10446] Anonymous. How to make rustic work. Scientific American, 37 (8):119–120, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
119.pdf.

Anonymous:1877:GPc

[10447] Anonymous. A geometrical puzzle. Scientific American, 37(8):120, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-120.pdf.

Anonymous:1877:EG

[10448] Anonymous. Engraving on glass. Scientific American, 37(8):120, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n8/pdf/scientificamerican08251877-120a.pdf.

Seiss:1877:ACS

[10449] C. Few Seiss. The ash-colored Salamander. Scientific American, 37(8):120, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
120b.pdf.

Anonymous:1877:AGT

[10450] Anonymous. The aleurometer — gluten tester. Scientific American, 37(8):120, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
120c.pdf.

Anonymous:1877:WIHc

[10451] Anonymous. The watch — its invention and history. Scientific American, 37(8):121, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
121.pdf.

Anonymous:1877:RAFbg

[10452] Anonymous. Recent American and foreign patents. Scientific American, 37(8):121, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
121a.pdf.

Anonymous:1877:NMEbb

[10453] Anonymous. New mechanical and engineering inventions. Scientific American, 37(8):121–122, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
121b.pdf.

Anonymous:1877:NMIbg

[10454] Anonymous. New miscellaneous inventions. Scientific American, 37(8):122, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
122.pdf.

Anonymous:1877:NAIbb

[10455] Anonymous. New agricultural inventions. Scientific American, 37(8):122, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
122a.pdf.

Anonymous:1877:NWHz

[10456] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(8):122, August 25, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n8/pdf/scientificamerican08251877-122b.pdf.

Anonymous:1877:CRbg

[10457] Anonymous. Communications received. Scientific American, 37 (8):124, August 25, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n8/pdf/scientificamerican08251877-
124.pdf.

Anonymous:1877:BPbi

[10458] Anonymous. Business and personal. Scientific American, 37(9): 140, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
140.pdf.

Anonymous:1877:NQbg

[10459] Anonymous. Notes and queries. Scientific American, 37(9):140, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n9/pdf/scientificamerican09011877-140a.pdf.

Anonymous:1877:IIbh

[10460] Anonymous. Index of inventions. Scientific American, 37(9): 140, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
140c.pdf.

Anonymous:1877:PNPbg

[10461] Anonymous. Patent notes: Patents. Scientific American, 37 (9):142, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
142.pdf.

Anonymous:1877:EPMb

[10462] Anonymous. Electro-plating machinery. Scientific American, 37 (9):127, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
127.pdf.

Anonymous:1877:HSR

[10463] Anonymous. How shall railway workmen live? Scientific American, 37(9):128, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
128.pdf.

Anonymous:1877:ECM

[10464] Anonymous. Early Connecticut manufactures. Scientific American, 37(9):128–129, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
128a.pdf.

Anonymous:1877:NPOb

[10465] Anonymous. Notes of Patent Office decisions. Scientific American, 37(9):129, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
129.pdf.

Anonymous:1877:MCE

[10466] Anonymous. Muscular contraction and electricity. Scientific American, 37(9):129, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
129a.pdf.

Anonymous:1877:BTC

[10467] Anonymous. Boiler test challenge. Scientific American, 37(9): 129, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
129b.pdf.

Anonymous:1877:IHH

[10468] Anonymous. Improved hydraulic hauling machinery. Scientific American, 37(9):130, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
130.pdf.

Anonymous:1877:MN

[10469] Anonymous. The manufacture of needles. Scientific American, 37 (9):131–132, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
131.pdf.

Anonymous:1877:Cbk

[10470] Anonymous. Communications. Scientific American, 37(9):132– 133, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
132.pdf.

Anonymous:1877:HRB

[10471] Anonymous. How to re-bore the ends of steam cylinders. Scientific American, 37(9):133, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
133.pdf.

Anonymous:1877:ANi

[10472] Anonymous. Astronomical notes. Scientific American, 37(9): 133, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
133a.pdf.

Anonymous:1877:CTb

[10473] Anonymous. The coventry tricycle. Scientific American, 37(9): 134, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
134.pdf.

Anonymous:1877:IME

[10474] Anonymous. Improved mining engine. Scientific American, 37 (9):134, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
134a.pdf.

Anonymous:1877:AECb

[10475] Anonymous. Anderson’s Equilibrio couch. Scientific American, 37(9):134, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
134b.pdf.

Anonymous:1877:SLL

[10476] Anonymous. Snowy-lipped Lady’s slipper. Scientific American, 37(9):135, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
135.pdf.

Anonymous:1877:LNV

[10477] Anonymous. Low’s necklace vanda. Scientific American, 37(9): 135, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
135a.pdf.

Anonymous:1877:Gc

[10478] Anonymous. The gorilla. Scientific American, 37(9):135–136, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n9/pdf/scientificamerican09011877-135b.pdf.

Anonymous:1877:SDH

[10479] Anonymous. The study of the diseases of the human voice. Scientific American, 37(9):136, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
136.pdf.

Huger:1877:RHI

[10480] Colonel C. K. Huger. Rice — how it is prepared for market. Scientific American, 37(9):137, September 1, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n9/pdf/scientificamerican09011877-137.pdf.

Anonymous:1877:KMD

[10481] Anonymous. The Keely motor deception. Scientific American, 37(9):137, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
137a.pdf.

Anonymous:1877:E

[10482] Anonymous. Eburine. Scientific American, 37(9):137, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n9/pdf/scientificamerican09011877-137b.pdf.

Anonymous:1877:ES

[10483] Anonymous. Engine “spurts”. Scientific American, 37(9):137– 138, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
137c.pdf.

Anonymous:1877:RAFbh

[10484] Anonymous. Recent American and foreign patents. Scientific American, 37(9):138, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
138.pdf.

Anonymous:1877:NMEbc

[10485] Anonymous. New mechanical and engineering inventions. Scientific American, 37(9):138, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
138a.pdf.

Anonymous:1877:NMIbh

[10486] Anonymous. New miscellaneous inventions. Scientific American, 37 (9):138–139, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
138b.pdf.

Anonymous:1877:NAIbc

[10487] Anonymous. New agricultural inventions. Scientific American, 37(9):139, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
139.pdf.

Anonymous:1877:NTIk

[10488] Anonymous. New textile inventions. Scientific American, 37 (9):139, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
139a.pdf.

Anonymous:1877:NHIv

[10489] Anonymous. New household inventions. Scientific American, 37 (9):139, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
139b.pdf.

Anonymous:1877:CRbh

[10490] Anonymous. Communications received. Scientific American, 37 (9):140, September 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n9/pdf/scientificamerican09011877-
140b.pdf.

Anonymous:1877:BNBt

[10491] Anonymous. Books: New books and publications. Scientific American, 37(10):153, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
153.pdf.

Anonymous:1877:BPbj

[10492] Anonymous. Business and personal. Scientific American, 37 (10):155, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
155.pdf.

Anonymous:1877:NQbh

[10493] Anonymous. Notes and queries. Scientific American, 37(10): 155–156, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
155a.pdf.

Anonymous:1877:IIbi

[10494] Anonymous. Index of inventions. Scientific American, 37(10): 156, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
156a.pdf.

Anonymous:1877:PNPbh

[10495] Anonymous. Patent notes: Patents. Scientific American, 37 (10):158, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
158.pdf.

Anonymous:1877:IBCa

[10496] Anonymous. Improved board cutting and seasoning machines. Scientific American, 37(10):143–144, September 8, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n10/pdf/scientificamerican09081877-143.pdf.

Anonymous:1877:WSEa

[10497] Anonymous. The Woodruff Scientific Expedition. Scientific American, 37(10):144, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
144.pdf.

Anonymous:1877:DSM

[10498] Anonymous. Discovery of satellites of Mars. Scientific American, 37(10):144, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
144a.pdf.

Anonymous:1877:MHC

[10499] Anonymous. Machine honesty and its circumvention. Scientific American, 37(10):144–145, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
144b.pdf.

Anonymous:1877:F

[10500] Anonymous. Food. Scientific American, 37(10):145, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n10/pdf/scientificamerican09081877-145.pdf.

Anonymous:1877:IOS

[10501] Anonymous. Improved overhead sewing machine. Scientific American, 37(10):146–147, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
146.pdf.

Anonymous:1877:IAFb

[10502] Anonymous. Improved automatic fly fan. Scientific American, 37 (10):147, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
147.pdf.

Anonymous:1877:CPP

[10503] Anonymous. Combined push-pin and watch key. Scientific American, 37(10):147, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
147a.pdf.

Anonymous:1877:IFD

[10504] Anonymous. Improved fastening device for billiard cushion rails. Scientific American, 37(10):147, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
147b.pdf.

Anonymous:1877:Cbl

[10505] Anonymous. Communications. Scientific American, 37(10):148– 149, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
148.pdf.

Anonymous:1877:FEA

[10506] Anonymous. A fire escape accident. Scientific American, 37 (10):149, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
149.pdf.

Anonymous:1877:ISEb

[10507] Anonymous. Improved steam engine governor. Scientific American, 37(10):150, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
150.pdf.

Anonymous:1877:IRMa

[10508] Anonymous. Improved rivet-making machine. Scientific American, 37(10):150, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
150a.pdf.

Seiss:1877:CF

[10509] C. Few Seiss. The clamorous frog. Scientific American, 37 (10):151, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
151.pdf.

Anonymous:1877:IBJ

[10510] Anonymous. The Iwakuni Bridge, Japan. Scientific American, 37 (10):151–152, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
151a.pdf.

Anonymous:1877:IGA

[10511] Anonymous. Improved germinating apparatus. Scientific American, 37(10):152, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
152.pdf.

Anonymous:1877:NPOc

[10512] Anonymous. Notes of Patent Office decisions. Scientific American, 37(10):152, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
152a.pdf.

Anonymous:1877:GL

[10513] Anonymous. Giant Lily. Scientific American, 37(10):152–153, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n10/pdf/scientificamerican09081877-152b.pdf.

Anonymous:1877:RAFbi

[10514] Anonymous. Recent American and foreign patents. Scientific American, 37(10):153, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
153a.pdf.

Anonymous:1877:NHIw

[10515] Anonymous. New household inventions. Scientific American, 37 (10):153, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
153b.pdf.

Anonymous:1877:NMEbd

[10516] Anonymous. New mechanical and engineering inventions. Scientific American, 37(10):153–154, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
153c.pdf.

Anonymous:1877:NMIbi

[10517] Anonymous. New miscellaneous inventions. Scientific American, 37(10):154, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
154.pdf.

Anonymous:1877:NAIbd

[10518] Anonymous. New agricultural inventions. Scientific American, 37 (10):154, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
154a.pdf.

Anonymous:1877:CRbi

[10519] Anonymous. Communications received. Scientific American, 37 (10):156, September 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n10/pdf/scientificamerican09081877-
156.pdf.

Anonymous:1877:BPbk

[10520] Anonymous. Business and personal. Scientific American, 37 (11):172, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
172.pdf.

Anonymous:1877:NQbi

[10521] Anonymous. Notes and queries. Scientific American, 37(11): 172, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
172a.pdf.

Anonymous:1877:IIbj

[10522] Anonymous. Index of inventions. Scientific American, 37(11): 172, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
172c.pdf.

Anonymous:1877:PNPbi

[10523] Anonymous. Patent notes: Patents. Scientific American, 37 (11):174, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
174.pdf.

Anonymous:1877:IRJb

[10524] Anonymous. Improved railway journal boxes wanted. Scientific American, 37(11):159, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
159.pdf.

Anonymous:1877:IMC

[10525] Anonymous. Improved method of compressing cotton. Scientific American, 37(11):159, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
159a.pdf.

Anonymous:1877:AFF

[10526] Anonymous. American fruit in foreign markets. Scientific American, 37(11):160, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
160.pdf.

Anonymous:1877:IMb

[10527] Anonymous. Indigestible medicines. Scientific American, 37(11): 160, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
160a.pdf.

Orton:1877:SIP

[10528] James Orton. The sugar interest in Peru. Scientific American, 37 (11):160–161, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
160b.pdf.

Anonymous:1877:EAMb

[10529] Anonymous. The emigration of American mechanics to England. Scientific American, 37(11):161, September 15, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n11/pdf/scientificamerican09151877-161.pdf.

Anonymous:1877:PSA

[10530] Anonymous. Progress of the Scientific American. Scientific American, 37(11):161, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
161a.pdf.

Anonymous:1877:IFSb

[10531] Anonymous. Increasing the flow of springs. Scientific American, 37(11):161, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
161b.pdf.

Anonymous:1877:IFL

[10532] Anonymous. Improved farm locomotives and steam road rollers. Scientific American, 37(11):162, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
162.pdf.

Anonymous:1877:CSF

[10533] Anonymous. Combined skimmer and fork. Scientific American, 37 (11):163, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
163.pdf.

Anonymous:1877:IHCc

[10534] Anonymous. Improved hop cultivator. Scientific American, 37 (11):163, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
163a.pdf.

Anonymous:1877:AIF

[10535] Anonymous. Allen’s improved fire escape. Scientific American, 37 (11):163–164, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
163b.pdf.

Anonymous:1877:Cbm

[10536] Anonymous. Communications. Scientific American, 37(11):164, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n11/pdf/scientificamerican09151877-164.pdf.

Rose:1877:PMp

[10537] Joshua Rose. Practical mechanism. Scientific American, 37(11): 165, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
165.pdf.

Anonymous:1877:HIB

[10538] Anonymous. Hoffman’s improved beer cask. Scientific American, 37(11):166, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
166.pdf.

Anonymous:1877:EAI

[10539] Anonymous. Electrical apparatus to indicate overstrain or weakness in bridges or other structures. Scientific American, 37(11):166, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n11/pdf/scientificamerican09151877-166a.pdf.

Seiss:1877:BT

[10540] C. Few Seiss. The box tortoise. Scientific American, 37(11): 167, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
167.pdf.

Anonymous:1877:WHb

[10541] Anonymous. Water hogs. Scientific American, 37(11):167, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n11/pdf/scientificamerican09151877-167a.pdf.

Anonymous:1877:CCF

[10542] Anonymous. The cordonnier, or cobbler fish. Scientific American, 37(11):168, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
168.pdf.

Anonymous:1877:HCB

[10543] Anonymous. How to chuck and bore an engine crosshead. Scientific American, 37(11):168, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
168a.pdf.

Anonymous:1877:SWMc

[10544] Anonymous. Simple way to make ice. Scientific American, 37 (11):168, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
168b.pdf.

Anonymous:1877:IGP

[10545] Anonymous. Improved glue pot. Scientific American, 37(11): 168–170, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
168c.pdf.

Anonymous:1877:RAFbj

[10546] Anonymous. Recent American and foreign patents. Scientific American, 37(11):170, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
170.pdf.

Anonymous:1877:NHIx

[10547] Anonymous. New household inventions. Scientific American, 37 (11):170, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
170a.pdf.

Anonymous:1877:NWHba

[10548] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(11):170, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n11/pdf/scientificamerican09151877-170b.pdf.

Anonymous:1877:NMEbe

[10549] Anonymous. New mechanical and engineering inventions. Scientific American, 37(11):170–171, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
170c.pdf.

Anonymous:1877:NMIbj

[10550] Anonymous. New miscellaneous inventions. Scientific American, 37(11):171, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
171.pdf.

Anonymous:1877:NAIbe

[10551] Anonymous. New agricultural inventions. Scientific American, 37 (11):171, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
171a.pdf.

Anonymous:1877:NTIl

[10552] Anonymous. New textile inventions. Scientific American, 37 (11):171, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
171b.pdf.

Anonymous:1877:CRbj

[10553] Anonymous. Communications received. Scientific American, 37 (11):172, September 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n11/pdf/scientificamerican09151877-
172b.pdf.

Anonymous:1877:BNBu

[10554] Anonymous. Books: New books and publications. Scientific American, 37(12):185, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
185.pdf.

Anonymous:1877:BPbl

[10555] Anonymous. Business and personal. Scientific American, 37 (12):187, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
187.pdf.

Anonymous:1877:NQbj

[10556] Anonymous. Notes and queries. Scientific American, 37(12): 187–188, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
187a.pdf.

Anonymous:1877:IIbk

[10557] Anonymous. Index of inventions. Scientific American, 37(12): 188, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
188a.pdf.

Anonymous:1877:PNPbj

[10558] Anonymous. Patent notes: Patents. Scientific American, 37 (12):190, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
190.pdf.

Anonymous:1877:MCBa

[10559] Anonymous. The manufacture of covered buttons. Scientific American, 37(12):175–176, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
175.pdf.

Anonymous:1877:TCS

[10560] Anonymous. Thiers and contemporary science in France. Scientific American, 37(12):176, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
176.pdf.

Anonymous:1877:SCb

[10561] Anonymous. Sargent’s case. Scientific American, 37(12):176– 177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
176a.pdf.

Anonymous:1877:NPEa

[10562] Anonymous. New process for electro-plating. Scientific American, 37(12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177.pdf.

Anonymous:1877:ESL

[10563] Anonymous. The electro-silicic light. Scientific American, 37 (12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177a.pdf.

Anonymous:1877:PHG

[10564] Anonymous. Progress of hardened glass making. Scientific American, 37(12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177b.pdf.

Anonymous:1877:LMM

[10565] Anonymous. Is life a mode of motion? Scientific American, 37 (12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177c.pdf.

Anonymous:1877:TDQ

[10566] Anonymous. The torpedo defense question. Scientific American, 37(12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177d.pdf.

Anonymous:1877:DTE

[10567] Anonymous. Dr. Thomson on embryology and evolution. Scientific American, 37(12):177, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
177e.pdf.

Anonymous:1877:IVG

[10568] Anonymous. Improved vertical gas engine. Scientific American, 37 (12):178, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
178.pdf.

Anonymous:1877:NRT

[10569] Anonymous. New registering thermometer. Scientific American, 37 (12):179, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
179.pdf.

Anonymous:1877:NHC

[10570] Anonymous. New hoisting clamp for building stone. Scientific American, 37(12):179, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
179a.pdf.

Anonymous:1877:SBAa

[10571] Anonymous. Science at the British Association. Scientific American, 37(12):179, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
179b.pdf.

Anonymous:1877:Cbn

[10572] Anonymous. Communications. Scientific American, 37(12):180, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n12/pdf/scientificamerican09221877-180.pdf.

Anonymous:1877:ALA

[10573] Anonymous. Ancient life in America. Scientific American, 37(12): 180–181, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
180a.pdf.

Anonymous:1877:IBCb

[10574] Anonymous. Improved ball check valve for sewers. Scientific American, 37(12):182, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
182.pdf.

Anonymous:1877:SATa

[10575] Anonymous. Sherwood’s automatic tea and coffee pot. Scientific American, 37(12):182, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
182a.pdf.

Anonymous:1877:INC

[10576] Anonymous. Improved needle clamp for sewing machines. Scientific American, 37(12):182, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
182b.pdf.

Anonymous:1877:WSEb

[10577] Anonymous. The Woodruff Scientific Expedition steamer Ontario. Scientific American, 37(12):183, September 22, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n12/pdf/scientificamerican09221877-183.pdf.

Anonymous:1877:HPG

[10578] Anonymous. How to prevent Grasshopper raids. Scientific American, 37(12):183, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
183a.pdf.

Anonymous:1877:LME

[10579] Anonymous. The Lontin magneto-electric machine. Scientific American, 37(12):184, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
184.pdf.

Anonymous:1877:ABP

[10580] Anonymous. American borax production. Scientific American, 37 (12):184–185, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
184a.pdf.

Anonymous:1877:RAFbk

[10581] Anonymous. Recent American and foreign patents. Scientific American, 37(12):185, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
185a.pdf.

Anonymous:1877:NMEbf

[10582] Anonymous. New mechanical and engineering inventions. Scientific American, 37(12):185, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
185b.pdf.

Anonymous:1877:NHIy

[10583] Anonymous. New household inventions. Scientific American, 37 (12):185, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
185c.pdf.

Anonymous:1877:NMIbk

[10584] Anonymous. New miscellaneous inventions. Scientific American, 37 (12):185–186, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
185d.pdf.

Anonymous:1877:NAIbf

[10585] Anonymous. New agricultural inventions. Scientific American, 37 (12):186, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
186.pdf.

Anonymous:1877:NWHbb

[10586] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(12):186, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n12/pdf/scientificamerican09221877-186a.pdf.

Anonymous:1877:CRbk

[10587] Anonymous. Communications received. Scientific American, 37 (12):188, September 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n12/pdf/scientificamerican09221877-
188.pdf.

Anonymous:1877:BNBv

[10588] Anonymous. Books: New books and publications. Scientific American, 37(13):201, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
201b.pdf.

Anonymous:1877:BPbm

[10589] Anonymous. Business and personal. Scientific American, 37 (13):203, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
203.pdf.

Anonymous:1877:NQbk

[10590] Anonymous. Notes and queries. Scientific American, 37(13): 203–204, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
203a.pdf.

Anonymous:1877:IIbl

[10591] Anonymous. Index of inventions. Scientific American, 37(13): 204, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
204a.pdf.

Anonymous:1877:PNPbk

[10592] Anonymous. Patent notes: Patents. Scientific American, 37 (13):206, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
206.pdf.

Anonymous:1877:SAC

[10593] Anonymous. Self-acting car couplers, preparation of oxalic acid from parchment paper. Scientific American, 37(13):191, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-191.pdf.

Anonymous:1877:VCE

[10594] Anonymous. Vertical compound engine with Allison’s boiler. Scientific American, 37(13):191, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
191a.pdf.

Anonymous:1877:Da

[10595] Anonymous. Disgust. Scientific American, 37(13):192, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-192.pdf.

Anonymous:1877:ATB

[10596] Anonymous. Anomalies in the temperature of the boiling point. Scientific American, 37(13):192, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
192a.pdf.

Anonymous:1877:SSFb

[10597] Anonymous. Sun spots and famine. Scientific American, 37(13): 192–193, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
192b.pdf.

Anonymous:1877:AVW

[10598] Anonymous. The agricultural value of worms. Scientific American, 37(13):193, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
193.pdf.

Anonymous:1877:IRC

[10599] Anonymous. The international rifle contest. Scientific American, 37(13):193, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
193a.pdf.

Anonymous:1877:MH

[10600] Anonymous. Meteoric heat. Scientific American, 37(13):193, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-193b.pdf.

Anonymous:1877:ILE

[10601] Anonymous. Influence of light on the electric condition of metals in saline solutions. Scientific American, 37(13):193, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-193c.pdf.

Berdell:1877:T

[10602] Theodore Berdell. Tellurium. Scientific American, 37(13):194, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-194.pdf.

Anonymous:1877:AMC

[10603] Anonymous. Another mysterious clock. Scientific American, 37 (13):194, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
194a.pdf.

Anonymous:1877:IAE

[10604] Anonymous. The industrial applications of the electric light. Scientific American, 37(13):195–196, September 29, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n13/pdf/scientificamerican09291877-195.pdf.

Anonymous:1877:Cbo

[10605] Anonymous. Communications. Scientific American, 37(13):196, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-196.pdf.

Anonymous:1877:SBAb

[10606] Anonymous. Science at the British Association. Scientific American, 37(13):197, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
197.pdf.

Anonymous:1877:ICCb

[10607] Anonymous. Improved car coupling. Scientific American, 37 (13):198, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
198.pdf.

Anonymous:1877:DHC

[10608] Anonymous. Dillon’s halter chain, and trace coupling. Scientific American, 37(13):198, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
198a.pdf.

Riley:1877:CPB

[10609] C. V. Riley. The Colorado potato-beetle in Europe. — German thoroughness. Scientific American, 37(13):198, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n13/pdf/scientificamerican09291877-198b.pdf.

Anonymous:1877:ISBb

[10610] Anonymous. Improved sash balance. Scientific American, 37 (13):198, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
198c.pdf.

Anonymous:1877:BPbn

[10611] Anonymous. The birds of paradise. Scientific American, 37(13): 199–200, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
199.pdf.

Anonymous:1877:CSb

[10612] Anonymous. A chimneyless stove. Scientific American, 37(13): 200, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
200.pdf.

Anonymous:1877:ALS

[10613] Anonymous. The American life saver or surf car. Scientific American, 37(13):200, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
200a.pdf.

Anonymous:1877:TMA

[10614] Anonymous. Trial of military appliances. Scientific American, 37 (13):200, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
200b.pdf.

Anonymous:1877:SBAc

[10615] Anonymous. Science at the British Association. Scientific American, 37(13):201, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
201.pdf.

Goodrich:1877:IMB

[10616] John C. Goodrich, Jr. Improvement in manufacture of beton. Scientific American, 37(13):201, September 29, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n13/pdf/scientificamerican09291877-201a.pdf.

Anonymous:1877:RAFbl

[10617] Anonymous. Recent American and foreign patents. Scientific American, 37(13):201, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
201c.pdf.

Anonymous:1877:NMEbg

[10618] Anonymous. New mechanical and engineering inventions. Scientific American, 37(13):201–202, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
201d.pdf.

Anonymous:1877:NMIbl

[10619] Anonymous. New miscellaneous inventions. Scientific American, 37(13):202, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
202.pdf.

Anonymous:1877:CRbl

[10620] Anonymous. Communications received. Scientific American, 37 (13):204, September 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n13/pdf/scientificamerican09291877-
204.pdf.

Anonymous:1877:BPbo

[10621] Anonymous. Business and personal. Scientific American, 37(14):220, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-220.pdf.

Anonymous:1877:NQbl

[10622] Anonymous. Notes and queries. Scientific American, 37(14):220, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-220a.pdf.

Anonymous:1877:IIbm

[10623] Anonymous. Index of inventions. Scientific American, 37(14):220, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-220c.pdf.

Anonymous:1877:PNPbl

[10624] Anonymous. Patent notes: Patents. Scientific American, 37(14):222, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-222.pdf.

Anonymous:1877:NBT

[10625] Anonymous. The new Bell telephone. Scientific American, 37(14):207, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-207.pdf.

Anonymous:1877:SRS

[10626] Anonymous. Steam railways in the streets of New York. Scientific American, 37(14):208, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
208.pdf.

Anonymous:1877:ETC

[10627] Anonymous. The extension of tea culture. Scientific American, 37(14):208, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
208a.pdf.

Anonymous:1877:CRbm

[10628] Anonymous. The Congo River. Scientific American, 37(14):208–209, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-208b.pdf.

Anonymous:1877:DWS

[10629] Anonymous. Domestic water supply in the country. Scientific American, 37(14):209, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
209.pdf.

Anonymous:1877:NPOd

[10630] Anonymous. Notes of Patent Office decisions. Scientific American, 37(14):209, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
209a.pdf.

Anonymous:1877:LWS

[10631] Anonymous. Locomotive wheel sliding. Scientific American, 37(14):209, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-209b.pdf.

Anonymous:1877:TCM

[10632] Anonymous. Tax collection by machinery. Scientific American, 37(14):209, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
209c.pdf.

Anonymous:1877:Db

[10633] Anonymous. The diplograph. Scientific American, 37(14):210, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-210.pdf.

Anonymous:1877:INF

[10634] Anonymous. Improved nail forging machine. Scientific American, 37(14):210–211, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
210a.pdf.

Anonymous:1877:IMH

[10635] Anonymous. Improved mule head stocks. Scientific American, 37(14):211, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
211.pdf.

Anonymous:1877:DIB

[10636] Anonymous. Drew’s improved Baker. Scientific American, 37(14):211, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-211a.pdf.

Anonymous:1877:Cbp

[10637] Anonymous. Communication. Scientific American, 37(14):212–213, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-212.pdf.

Anonymous:1877:IMM

[10638] Anonymous. Improved miter machine. Scientific American, 37(14):214, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-214.pdf.

Anonymous:1877:NMG

[10639] Anonymous. New metal grinding machine. Scientific American, 37 (14):214–215, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
214a.pdf.

Anonymous:1877:WPb

[10640] Anonymous. The wine palm. Scientific American, 37(14):215, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-215.pdf.

Seiss:1877:ASI

[10641] C. Few Seiss. About some insects. Scientific American, 37(14):215–216, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-215a.pdf.

Anonymous:1877:SMH

[10642] Anonymous. The spontaneous movements of the hornwort. Scientific American, 37(14):216, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
216.pdf.

Anonymous:1877:BIH

[10643] Anonymous. Beaudet’s improved hinge. Scientific American, 37 (14):216, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
216a.pdf.

Anonymous:1877:ANj

[10644] Anonymous. Astronomical notes. Scientific American, 37(14):216–217, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n14/pdf/scientificamerican10061877-216b.pdf.

Anonymous:1877:RAFbm

[10645] Anonymous. Recent American and foreign patents. Scientific American, 37(14):217, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
217.pdf.

Anonymous:1877:NAIbg

[10646] Anonymous. New agricultural inventions. Scientific American, 37 (14):217–218, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
217a.pdf.

Anonymous:1877:NWHbc

[10647] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(14):218, October 6, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n14/pdf/scientificamerican10061877-218.pdf.

Anonymous:1877:NHIz

[10648] Anonymous. New household inventions. Scientific American, 37 (14):218–219, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
218a.pdf.

Anonymous:1877:NMEbh

[10649] Anonymous. New mechanical and engineering inventions. Scientific American, 37(14):219, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
219.pdf.

Anonymous:1877:NMIbm

[10650] Anonymous. New miscellaneous inventions. Scientific American, 37(14):219, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
219a.pdf.

Anonymous:1877:CRbn

[10651] Anonymous. Communications received. Scientific American, 37 (14):220, October 6, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n14/pdf/scientificamerican10061877-
220b.pdf.

Anonymous:1877:BPbp

[10652] Anonymous. Business and personal. Scientific American, 37 (15):236, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
236.pdf.

Anonymous:1877:NQbm

[10653] Anonymous. Notes and queries. Scientific American, 37(15):236, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n15/pdf/scientificamerican10131877-236a.pdf.

Anonymous:1877:IIbn

[10654] Anonymous. Index of inventions. Scientific American, 37(15):236, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n15/pdf/scientificamerican10131877-236c.pdf.

Anonymous:1877:PNPbm

[10655] Anonymous. Patent notes: Patents. Scientific American, 37 (15):238, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
238.pdf.

Anonymous:1877:ADa

[10656] Anonymous. Aniline dyes. Scientific American, 37(15):223, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n15/pdf/scientificamerican10131877-223.pdf.

Anonymous:1877:MWP

[10657] Anonymous. The manufacture of wall paper. Scientific American, 37(15):223, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
223a.pdf.

Anonymous:1877:FUP

[10658] Anonymous. The fire in the U.S. Patent Office. Scientific American, 37(15):224, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
224.pdf.

Anonymous:1877:LHW

[10659] Anonymous. Leverrier and his work. Scientific American, 37 (15):224, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
224a.pdf.

Anonymous:1877:MDP

[10660] Anonymous. Models destroyed by the Patent Office fire. Scientific American, 37(15):224–225, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
224b.pdf.

Anonymous:1877:NPOe

[10661] Anonymous. Notes of Patent Office decisions. Scientific American, 37(15):225, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
225.pdf.

Anonymous:1877:IAT

[10662] Anonymous. Improved apparatus for testing bank checks. Scientific American, 37(15):227, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
227.pdf.

Anonymous:1877:WBIa

[10663] Anonymous. Walker Brothers’ improved scroll saw. Scientific American, 37(15):227–228, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
227a.pdf.

Rose:1877:PMq

[10664] Joshua Rose. Practical mechanism. Scientific American, 37(15): 228–229, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
228.pdf.

Anonymous:1877:ISCb

[10665] Anonymous. Improved sleeping car. Scientific American, 37 (15):230, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
230.pdf.

Anonymous:1877:BIC

[10666] Anonymous. Brown’s improved cotton stalk puller. Scientific American, 37(15):230, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
230a.pdf.

Seiss:1877:SDF

[10667] C. Few Seiss. The skull of the domestic fowl. Scientific American, 37(15):231, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
231.pdf.

Anonymous:1877:LD

[10668] Anonymous. Lomaria Dalgairnsae. Scientific American, 37(15): 231–232, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
231a.pdf.

Anonymous:1877:SIF

[10669] Anonymous. Signor Ignazio’s “flying man”. Scientific American, 37(15):232, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
232.pdf.

Anonymous:1877:NSBb

[10670] Anonymous. New smoke-burning furnace for steam boilers. Scientific American, 37(15):232–233, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
232a.pdf.

Anonymous:1877:RAFbn

[10671] Anonymous. Recent American and foreign patents. Scientific American, 37(15):233, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
233.pdf.

Anonymous:1877:NMEbi

[10672] Anonymous. New mechanical and engineering inventions. Scientific American, 37(15):233–234, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
233a.pdf.

Anonymous:1877:NAIbh

[10673] Anonymous. New agricultural inventions. Scientific American, 37 (15):234–235, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
234.pdf.

Anonymous:1877:NWHbd

[10674] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(15):235, October 13, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n15/pdf/scientificamerican10131877-235.pdf.

Anonymous:1877:NMIbn

[10675] Anonymous. New miscellaneous inventions. Scientific American, 37(15):235, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
235a.pdf.

Anonymous:1877:CRbo

[10676] Anonymous. Communications received. Scientific American, 37 (15):236, October 13, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n15/pdf/scientificamerican10131877-
236b.pdf.

Anonymous:1877:BNBw

[10677] Anonymous. Books: New books and publications. Scientific American, 37(16):249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249.pdf.

Anonymous:1877:BPbq

[10678] Anonymous. Business and personal. Scientific American, 37 (16):251, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
251.pdf.

Anonymous:1877:NQbn

[10679] Anonymous. Notes and queries. Scientific American, 37(16): 251–252, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
251a.pdf.

Anonymous:1877:IIbo

[10680] Anonymous. Index of inventions. Scientific American, 37(16):252, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n16/pdf/scientificamerican10201877-252a.pdf.

Anonymous:1877:PNPbn

[10681] Anonymous. Patent notes: Patents. Scientific American, 37 (16):254, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
254.pdf.

Anonymous:1877:FLS

[10682] Anonymous. Fireless locomotives for street railways. Scientific American, 37(16):239, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
239.pdf.

Anonymous:1877:SNM

[10683] Anonymous. Star, nebula or meteorite. Scientific American, 37 (16):240, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
240.pdf.

Anonymous:1877:PEb

[10684] Anonymous. Practical evolution. Scientific American, 37(16):240, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n16/pdf/scientificamerican10201877-240a.pdf.

Anonymous:1877:AMb

[10685] Anonymous. The abolition of models. Scientific American, 37 (16):240–241, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
240b.pdf.

Anonymous:1877:PSI

[10686] Anonymous. Postage stamp inventions. Scientific American, 37 (16):241, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
241.pdf.

Anonymous:1877:BMN

[10687] Anonymous. The borax mines of Nevada. — a valuable discovery. Scientific American, 37(16):241, October 20, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n16/pdf/scientificamerican10201877-241a.pdf.

Anonymous:1877:PWM

[10688] Anonymous. Protected working men. Scientific American, 37 (16):241, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
241b.pdf.

Anonymous:1877:NPOf

[10689] Anonymous. Notes of Patent Office decisions. Scientific American, 37(16):241, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
241c.pdf.

Anonymous:1877:MPC

[10690] Anonymous. The manufacture of plumbago crucibles. Scientific American, 37(16):242, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
242.pdf.

Anonymous:1877:Bc

[10691] Anonymous. The baromotor. Scientific American, 37(16):242–243, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n16/pdf/scientificamerican10201877-242a.pdf.

Anonymous:1877:RBM

[10692] Anonymous. Roots’ blower for mine ventilation. Scientific American, 37(16):243–244, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
243.pdf.

Anonymous:1877:Cbq

[10693] Anonymous. Communications. Scientific American, 37(16):245, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n16/pdf/scientificamerican10201877-245.pdf.

Anonymous:1877:NSPb

[10694] Anonymous. New surface planing machine. Scientific American, 37(16):246, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
246.pdf.

Anonymous:1877:IAFc

[10695] Anonymous. Improved automatic fire extinguisher. Scientific American, 37(16):246–247, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
246a.pdf.

Anonymous:1877:SGL

[10696] Anonymous. Small geared locomotives. Scientific American, 37 (16):247, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
247.pdf.

Anonymous:1877:WCG

[10697] Anonymous. The White cheeked gibbon. Scientific American, 37 (16):247–248, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
247a.pdf.

Anonymous:1877:WB

[10698] Anonymous. The white Brugmansia. Scientific American, 37 (16):248, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
248.pdf.

Wagner:1877:MHS

[10699] Rudolf Wagner. The modern hyposulphite of soda. Scientific American, 37(16):248–249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
248a.pdf.

Anonymous:1877:RAFbo

[10700] Anonymous. Recent American and foreign patents. Scientific American, 37(16):249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249a.pdf.

Anonymous:1877:NWHbe

[10701] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(16):249, October 20, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n16/pdf/scientificamerican10201877-249b.pdf.

Anonymous:1877:NTIm

[10702] Anonymous. New textile inventions. Scientific American, 37 (16):249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249c.pdf.

Anonymous:1877:NAIbi

[10703] Anonymous. New agricultural inventions. Scientific American, 37(16):249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249d.pdf.

Anonymous:1877:NMEbj

[10704] Anonymous. New mechanical and engineering inventions. Scientific American, 37(16):249, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249e.pdf.

Anonymous:1877:NMIbo

[10705] Anonymous. New miscellaneous inventions. Scientific American, 37 (16):249–250, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
249f.pdf.

Anonymous:1877:CRbp

[10706] Anonymous. Communications received. Scientific American, 37 (16):252, October 20, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n16/pdf/scientificamerican10201877-
252.pdf.

Anonymous:1877:BNBx

[10707] Anonymous. Books: New books and publications. Scientific American, 37(17):265, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
265.pdf.

Anonymous:1877:BPbr

[10708] Anonymous. Business and personal. Scientific American, 37 (17):267, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
267.pdf.

Anonymous:1877:NQbo

[10709] Anonymous. Notes and queries. Scientific American, 37(17): 267–268, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
267a.pdf.

Anonymous:1877:IIbp

[10710] Anonymous. Index of inventions. Scientific American, 37(17):268, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-268a.pdf.

Anonymous:1877:PNPbo

[10711] Anonymous. Patent notes: Patents. Scientific American, 37 (17):270, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
270.pdf.

Anonymous:1877:IAFd

[10712] Anonymous. Improved automatic fire escape. Scientific American, 37(17):255, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
255.pdf.

Anonymous:1877:OS

[10713] Anonymous. An obelisk at sea. Scientific American, 37(17):255, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-255a.pdf.

Anonymous:1877:CAR

[10714] Anonymous. Curious animal relationships. Scientific American, 37(17):256, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
256.pdf.

Anonymous:1877:SPc

[10715] Anonymous. The science of play. Scientific American, 37(17):256, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-256a.pdf.

Anonymous:1877:DCPa

[10716] Anonymous. The duties of the Commissioner of Patents and the supervising powers of the Secretary of the Interior. Scientific American, 37(17):257, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
257.pdf.

Anonymous:1877:ISCc

[10717] Anonymous. Influence of the spinal cord on bodily temperature. Scientific American, 37(17):257, October 27, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n17/pdf/scientificamerican10271877-257a.pdf.

Anonymous:1877:AON

[10718] Anonymous. Amateur observations of the November meteors. Scientific American, 37(17):257, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
257b.pdf.

Anonymous:1877:ATE

[10719] Anonymous. Another traveling expedition. Scientific American, 37(17):257, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
257c.pdf.

Anonymous:1877:ITC

[10720] Anonymous. Imitation terra cotta. Scientific American, 37(17):257, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-257d.pdf.

Anonymous:1877:IDS

[10721] Anonymous. Improved double shedding harness Jacquard loom. Scientific American, 37(17):258, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
258.pdf.

Anonymous:1877:APG

[10722] Anonymous. Allan’s pressure gauge. Scientific American, 37 (17):258, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
258a.pdf.

Anonymous:1877:NDA

[10723] Anonymous. A new dredging apparatus. Scientific American, 37 (17):259, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
259.pdf.

Anonymous:1877:ISSa

[10724] Anonymous. Improved Ship’s speed indicator. Scientific American, 37(17):259–261, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
259a.pdf.

Anonymous:1877:NEMb

[10725] Anonymous. New electrical machines. Scientific American, 37 (17):261, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
261.pdf.

Anonymous:1877:IWF

[10726] Anonymous. Improved window fastener. Scientific American, 37 (17):262, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
262.pdf.

Anonymous:1877:Cbr

[10727] Anonymous. The cosmograph. Scientific American, 37(17):262, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-262a.pdf.

Anonymous:1877:ISA

[10728] Anonymous. Improved safety alarm and door Bell. Scientific American, 37(17):262–264, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
262b.pdf.

Anonymous:1877:SMd

[10729] Anonymous. The satellites of Mars. Scientific American, 37 (17):264, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
264.pdf.

Anonymous:1877:O

[10730] Anonymous. Orchids. Scientific American, 37(17):264–265, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n17/pdf/scientificamerican10271877-264a.pdf.

Anonymous:1877:RAFbp

[10731] Anonymous. Recent American and foreign patents. Scientific American, 37(17):265, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
265a.pdf.

Anonymous:1877:NMIbp

[10732] Anonymous. New miscellaneous inventions. Scientific American, 37 (17):265–266, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
265b.pdf.

Anonymous:1877:NHIba

[10733] Anonymous. New household inventions. Scientific American, 37 (17):266, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
266.pdf.

Anonymous:1877:NMEbk

[10734] Anonymous. New mechanical and engineering inventions. Scientific American, 37(17):266, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
266a.pdf.

Anonymous:1877:CRbq

[10735] Anonymous. Communications received. Scientific American, 37 (17):268, October 27, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n17/pdf/scientificamerican10271877-
268.pdf.

Anonymous:1877:BPbs

[10736] Anonymous. Business and personal. Scientific American, 37 (18):283, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
283.pdf.

Anonymous:1877:NQbp

[10737] Anonymous. Notes and queries. Scientific American, 37(18): 283–284, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
283a.pdf.

Anonymous:1877:IIbq (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
284a.pdf.

Anonymous:1877:PNPbp

[10739] Anonymous. Patent notes: Patents. Scientific American, 37 (18):286, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
286.pdf.

Anonymous:1877:NPEb

[10740] Anonymous. New process of embalming the dead. Scientific American, 37(18):271, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
271.pdf.

Anonymous:1877:IPF

[10741] Anonymous. Improved puddling furnace. Scientific American, 37 (18):271, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
271a.pdf.

Anonymous:1877:OI

[10742] Anonymous. An opportunity for inventors. —. Scientific American, 37(18):272, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
272.pdf.

Anonymous:1877:PECa

[10743] Anonymous. The Paris Exposition in Congress. Scientific American, 37(18):272, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
272a.pdf.

Anonymous:1877:IMW

[10744] Anonymous. Improved method of wintering cows. Scientific American, 37(18):272–273, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
272b.pdf.

Anonymous:1877:FO

[10745] Anonymous. The fortunes of the obelisks. Scientific American, 37(18):273, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
273.pdf.

Anonymous:1877:SATb

[10746] Anonymous. Speech automatically transmitted in short hand by the telegraph. Scientific American, 37(18):273, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n18/pdf/scientificamerican11031877-273a.pdf.

Anonymous:1877:APS

[10747] Anonymous. Alleged poison in sugars. Scientific American, 37 (18):273, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
273b.pdf.

Anonymous:1877:DCq

[10748] Anonymous. Decisions of the courts. Scientific American, 37 (18):273, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
273c.pdf.

Anonymous:1877:HOW

[10749] Anonymous. How our would looks from other worlds. Scientific American, 37(18):275–276, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
275.pdf.

Anonymous:1877:Cbs

[10750] Anonymous. Communication. Scientific American, 37(18):277– 278, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
277.pdf.

Anonymous:1877:AHW

[10751] Anonymous. An ancient hand warmer. Scientific American, 37 (18):278, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
278.pdf.

Anonymous:1877:ISF

[10752] Anonymous. Improves self-feeding drill. Scientific American, 37 (18):278, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
278a.pdf.

Anonymous:1877:NFO

[10753] Anonymous. A new floating oil burner. Scientific American, 37 (18):278–279, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
278b.pdf.

Anonymous:1877:RPU

[10754] Anonymous. The ramie plant and its utilization. Scientific American, 37(18):279–281, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
279.pdf.

Anonymous:1877:ANk

[10755] Anonymous. Astronomical notes. Scientific American, 37(18): 281, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
281.pdf.

Anonymous:1877:RAFbq

[10756] Anonymous. Recent American and foreign patents. Scientific American, 37(18):281, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
281a.pdf.

Anonymous:1877:NMIbq

[10757] Anonymous. New miscellaneous inventions. Scientific American, 37 (18):281–282, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
281b.pdf.

Anonymous:1877:NWHbf

[10758] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(18):282, November 3, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n18/pdf/scientificamerican11031877-282.pdf.

Anonymous:1877:NHIbb

[10759] Anonymous. New household inventions. Scientific American, 37 (18):282, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
282a.pdf.

Anonymous:1877:NMEbl

[10760] Anonymous. New mechanical and engineering inventions. Scientific American, 37(18):282, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
282b.pdf.

Anonymous:1877:CRbr

[10761] Anonymous. Communications received. Scientific American, 37 (18):284, November 3, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n18/pdf/scientificamerican11031877-
284.pdf.

Anonymous:1877:BPbt

[10762] Anonymous. Business and personal. Scientific American, 37 (19):299, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
299.pdf.

Anonymous:1877:NQbq

[10763] Anonymous. Notes and queries. Scientific American, 37(19): 299–300, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
299a.pdf.

Anonymous:1877:IIbr

[10764] Anonymous. Index of inventions. Scientific American, 37(19): 300, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
300a.pdf.

Anonymous:1877:PNPbq

[10765] Anonymous. Patent notes: Patents. Scientific American, 37 (19):302, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
302.pdf.

Anonymous:1877:IBMb

[10766] Anonymous. The imperial brick machine. Scientific American, 37 (19):287, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
287.pdf.

Anonymous:1877:ADb

[10767] Anonymous. Arrested development. Scientific American, 37(19): 288, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
288.pdf.

Anonymous:1877:PTP

[10768] Anonymous. Professor Tyndall on the phenomena of human life. Scientific American, 37(19):288–289, November 10, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n19/pdf/scientificamerican11101877-288a.pdf.

Anonymous:1877:ARS

[10769] Anonymous. The American railway system. — maintenance of way. Scientific American, 37(19):289, November 10, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n19/pdf/scientificamerican11101877-289.pdf.

Anonymous:1877:NPOg

[10770] Anonymous. Notes of Patent Office decisions. Scientific American, 37(19):289, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
289a.pdf.

Anonymous:1877:WPCb

[10771] Anonymous. The wooden pavements of Chicago. Scientific American, 37(19):289, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
289b.pdf.

Anonymous:1877:WGW

[10772] Anonymous. The Wren gas works. Scientific American, 37(19): 290, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
290.pdf.

Anonymous:1877:NSG

[10773] Anonymous. New standard grinding mill. Scientific American, 37 (19):291–292, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
291.pdf.

Anonymous:1877:MAS

[10774] Anonymous. The meeting of the Academy of Sciences. Scientific American, 37(19):292, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
292.pdf.

Anonymous:1877:SIH

[10775] Anonymous. Some inventions for husking corn. Scientific American, 37(19):293, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
293.pdf.

Anonymous:1877:SNB

[10776] Anonymous. Streit’s new bevel flange loose pulleys. Scientific American, 37(19):294, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
294.pdf.

Anonymous:1877:JIB

[10777] Anonymous. Johnson’s improved boiler plate shears. Scientific American, 37(19):294, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
294a.pdf.

Anonymous:1877:IBA

[10778] Anonymous. Improved burglar alarm window fastening. Scientific American, 37(19):294, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
294b.pdf.

Anonymous:1877:BIG

[10779] Anonymous. Borden’s improved gauge cock. Scientific American, 37 (19):294–295, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
294c.pdf.

Anonymous:1877:AF

[10780] Anonymous. Archer fishes. Scientific American, 37(19):295, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n19/pdf/scientificamerican11101877-295.pdf.

Seiss:1877:RTL

[10781] C. Few Seiss. The red and the two-lined Salamander. Scientific American, 37(19):295–297, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
295a.pdf.

Anonymous:1877:RAFbr

[10782] Anonymous. Recent American and foreign patents. Scientific American, 37(19):297, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
297.pdf.

Anonymous:1877:NMEbm

[10783] Anonymous. New mechanical and engineering inventions. Scientific American, 37(19):297, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
297a.pdf.

Anonymous:1877:NMIbr

[10784] Anonymous. New miscellaneous inventions. Scientific American, 37 (19):297–298, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
297b.pdf.

Anonymous:1877:NAIbj

[10785] Anonymous. New agricultural inventions. Scientific American, 37 (19):298, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
298.pdf.

Anonymous:1877:NWHbg

[10786] Anonymous. New woodworking and house and carriage building inventions. Scientific American, 37(19):298, November 10, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n19/pdf/scientificamerican11101877-298a.pdf.

Anonymous:1877:NHIbc

[10787] Anonymous. New household inventions. Scientific American, 37 (19):298, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
298b.pdf.

Anonymous:1877:CRbs

[10788] Anonymous. Communications received. Scientific American, 37 (19):300, November 10, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n19/pdf/scientificamerican11101877-
300.pdf.

Anonymous:1877:BNBy

[10789] Anonymous. Books: New books and publications. Scientific American, 37(20):313, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
313.pdf.

Anonymous:1877:BPbu

[10790] Anonymous. Business and personal. Scientific American, 37 (20):315, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
315.pdf.

Anonymous:1877:NQbr

[10791] Anonymous. Notes and queries. Scientific American, 37(20): 315–316, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
315a.pdf.

Anonymous:1877:IIbs

[10792] Anonymous. Index of inventions. Scientific American, 37(20): 316, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
316a.pdf.

Anonymous:1877:PNPbr

[10793] Anonymous. Patent notes: Patents. Scientific American, 37 (20):318, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
318.pdf.

Anonymous:1877:ETB

[10794] Anonymous. Experiments with the Turkish bath. Scientific American, 37(20):303, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
303.pdf.

Anonymous:1877:IPL

[10795] Anonymous. Improved portable lunch heater. Scientific American, 37(20):303, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
303a.pdf.

Anonymous:1877:IRMb

[10796] Anonymous. Ingram rotary machine for printing illustrated newspapers. Scientific American, 37(20):303, November 17, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n20/pdf/scientificamerican11171877-303b.pdf.

Anonymous:1877:WIS

[10797] Anonymous. A wonderful invention. — speech capable of indefinite repetition from automatic records. Scientific American, 37 (20):304, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
304.pdf.

Anonymous:1877:IHT

[10798] Anonymous. Influence of heat on the tenacity of metals. Scientific American, 37(20):304–305, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
304a.pdf.

Anonymous:1877:NRS

[10799] Anonymous. A new and remarkable strike. Scientific American, 37(20):305, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
305.pdf.

Anonymous:1877:NDCa

[10800] Anonymous. Notes of decisions of the courts. Scientific American, 37(20):305, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
305a.pdf.

Anonymous:1877:PJO

[10801] Anonymous. Professor James Orton. Scientific American, 37 (20):305, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
305b.pdf.

Anonymous:1877:IBTb

[10802] Anonymous. Improved boring and turning machine. Scientific American, 37(20):306, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
306.pdf.

Anonymous:1877:GPd

[10803] Anonymous. Graphic phonetics. Scientific American, 37(20): 307–308, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
307.pdf.

Anonymous:1877:Cbt

[10804] Anonymous. Communications. Scientific American, 37(20):308– 309, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
308.pdf.

Anonymous:1877:NSL

[10805] Anonymous. New standard light and medium grinding mills. Scientific American, 37(20):310, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
310.pdf.

Anonymous:1877:NIB

[10806] Anonymous. New improved band-sawing machine. Scientific American, 37(20):310, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
310a.pdf.

Anonymous:1877:SMR

[10807] Anonymous. The steamship “Massachusetts” on and off the rocks. Scientific American, 37(20):311, November 17, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n20/pdf/scientificamerican11171877-311.pdf.

Anonymous:1877:BBS

[10808] Anonymous. Base ball science. Scientific American, 37(20):312– 313, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
312.pdf.

Anonymous:1877:RAFbs

[10809] Anonymous. Recent American and foreign patents. Scientific American, 37(20):313, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
313a.pdf.

Anonymous:1877:NMEbn

[10810] Anonymous. New mechanical and engineering inventions. Scientific American, 37(20):313–314, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
313b.pdf.

Anonymous:1877:NMIbs

[10811] Anonymous. New miscellaneous inventions. Scientific American, 37(20):314, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
314.pdf.

Anonymous:1877:NAIbk

[10812] Anonymous. New agricultural inventions. Scientific American, 37 (20):314, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
314a.pdf.

Anonymous:1877:NHIbd

[10813] Anonymous. New household inventions. Scientific American, 37 (20):314, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
314b.pdf.

Anonymous:1877:CRbt

[10814] Anonymous. Communications received. Scientific American, 37 (20):316, November 17, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n20/pdf/scientificamerican11171877-
316.pdf.

Anonymous:1877:BNBz

[10815] Anonymous. Books: New books and publications. Scientific American, 37(21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330a.pdf.

Anonymous:1877:BPbv

[10816] Anonymous. Business and personal. Scientific American, 37 (21):331, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
331.pdf.

Anonymous:1877:NQNc

[10817] Anonymous. Notes and queries: Notes & queries. Scientific American, 37(21):331–332, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
331a.pdf.

Anonymous:1877:IIbt

[10818] Anonymous. Index of inventions. Scientific American, 37(21): 332, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
332a.pdf.

Anonymous:1877:PNPbs

[10819] Anonymous. Patent notes: Patents. Scientific American, 37 (21):334, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
334.pdf.

Anonymous:1877:SSTb

[10820] Anonymous. The 42d street sewage tunnel, New York City. Scientific American, 37(21):319–320, November 24, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n21/pdf/scientificamerican11241877-319.pdf.

Anonymous:1877:RCCb

[10821] Anonymous. $80,000 reward for a cure for cholera. Scientific American, 37(21):320, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
320.pdf.

Anonymous:1877:SSSb

[10822] Anonymous. Sun spots studied by solar photography. Scientific American, 37(21):320, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
320a.pdf.

Anonymous:1877:CTc

[10823] Anonymous. Cerebral thermometry. Scientific American, 37(21): 320, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
320b.pdf.

Anonymous:1877:LMD

[10824] Anonymous. Lessons in mechanical drawing. Scientific American, 37 (21):320–321, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
320c.pdf.

Anonymous:1877:GTD

[10825] Anonymous. The glandular theory of disease. Scientific American, 37(21):321, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
321.pdf.

Anonymous:1877:NDCb

[10826] Anonymous. Notes of decisions of the courts. Scientific American, 37(21):321, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
321a.pdf.

Anonymous:1877:ITM

[10827] Anonymous. Improved tile-making machine. Scientific American, 37(21):322, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
322.pdf.

Anonymous:1877:MTN

[10828] Anonymous. M. Trouve’s new moist battery. Scientific American, 37(21):323, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
323.pdf.

Anonymous:1877:HNP

[10829] Anonymous. Harrison’s new portable flouring mill. Scientific American, 37(21):323, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
323a.pdf.

Anonymous:1877:Cbu

[10830] Anonymous. Communications. Scientific American, 37(21):324, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n21/pdf/scientificamerican11241877-324.pdf.

Rose:1877:PMr

[10831] Joshua Rose. Practical mechanism. Scientific American, 37(21): 325–326, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
325.pdf.

Anonymous:1877:IBMc

[10832] Anonymous. Improved boring machine. Scientific American, 37 (21):326, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
326.pdf.

Anonymous:1877:ILH

[10833] Anonymous. Improved label holder. Scientific American, 37(21): 326–327, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
326a.pdf.

Anonymous:1877:NFE

[10834] Anonymous. The new French Exposition Palace. Scientific American, 37(21):327, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
327.pdf.

Anonymous:1877:CRDa

[10835] Anonymous. A curious result of denudation. Scientific American, 37(21):327, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
327a.pdf.

Anonymous:1877:RO

[10836] Anonymous. Raising obelisks. Scientific American, 37(21):328– 329, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
328.pdf.

Anonymous:1877:NAAa

[10837] Anonymous. New arrangement of the air reservoirs in pumps. Scientific American, 37(21):329–330, November 24, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n21/pdf/scientificamerican11241877-329.pdf.

Anonymous:1877:HAE

[10838] Anonymous. How to adjust engine guide bars. Scientific American, 37(21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330.pdf.

Anonymous:1877:RAFbt

[10839] Anonymous. Recent American and foreign patents. Scientific American, 37(21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330b.pdf.

Anonymous:1877:NAIbl

[10840] Anonymous. New agricultural inventions. Scientific American, 37 (21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330c.pdf.

Anonymous:1877:NMIbt

[10841] Anonymous. New miscellaneous inventions. Scientific American, 37(21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330d.pdf.

Anonymous:1877:NMEbo

[10842] Anonymous. New mechanical and engineering inventions. Scientific American, 37(21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330e.pdf.

Anonymous:1877:NHIbe

[10843] Anonymous. New household inventions. Scientific American, 37 (21):330, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
330f.pdf.

Anonymous:1877:WBIb

[10844] Anonymous. Wants and business inquiries. Scientific American, 37(21):332, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
332.pdf.

Anonymous:1877:SAa

[10845] Anonymous. The Scientific American. Scientific American, 37 (21):332, November 24, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n21/pdf/scientificamerican11241877-
332b.pdf.

Anonymous:1877:BPbw

[10846] Anonymous. Business and personal. Scientific American, 37 (22):347, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
347.pdf.

Anonymous:1877:NQbs

[10847] Anonymous. Notes and queries. Scientific American, 37(22):347, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n22/pdf/scientificamerican12011877-347a.pdf.

Anonymous:1877:IIbu

[10848] Anonymous. Index of inventions. Scientific American, 37(22):348, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n22/pdf/scientificamerican12011877-348.pdf.

Anonymous:1877:PNPbt

[10849] Anonymous. Patent notes: Patents. Scientific American, 37 (22):349, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
349.pdf.

Anonymous:1877:PIM

[10850] Anonymous. The Pictet ice machine. Scientific American, 37 (22):335, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
335.pdf.

Anonymous:1877:CDa

[10851] Anonymous. The cure of diphtheria. Scientific American, 37 (22):336, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
336.pdf.

Anonymous:1877:GRA

[10852] Anonymous. The geological relations of the atmosphere. Scientific American, 37(22):336, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
336a.pdf.

Anonymous:1877:VJ

[10853] Anonymous. Venerable journalists. Scientific American, 37(22): 336–337, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
336b.pdf.

Anonymous:1877:WSN

[10854] Anonymous. Water supply for New York City. Scientific American, 37(22):337, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
337.pdf.

Anonymous:1877:GCC

[10855] Anonymous. The Green corn case. Scientific American, 37(22):337, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n22/pdf/scientificamerican12011877-337a.pdf.

Anonymous:1877:HMH

[10856] Anonymous. How to make homes healthy. Scientific American, 37(22):337, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
337b.pdf.

Anonymous:1877:BSB

[10857] Anonymous. Breakage of a steamboat beam. Scientific American, 37(22):337, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
337c.pdf.

Anonymous:1877:CES

[10858] Anonymous. Combined engine and sugar cane grinding mill. Scientific American, 37(22):338–339, December 1, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n22/pdf/scientificamerican12011877-338.pdf.

Anonymous:1877:RBSa

[10859] Anonymous. The ross Burr stone grinding mill. Scientific American, 37(22):339, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
339.pdf.

Anonymous:1877:Cbv

[10860] Anonymous. Communications. Scientific American, 37(22):340– 341, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
340.pdf.

Anonymous:1877:ITP

[10861] Anonymous. Improved tool post for lathes. Scientific American, 37(22):342, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
342.pdf.

Anonymous:1877:HFB

[10862] Anonymous. Hill’s folding book case. Scientific American, 37 (22):342, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
342a.pdf.

Anonymous:1877:Cbw

[10863] Anonymous. The capercailzie. Scientific American, 37(22):343, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n22/pdf/scientificamerican12011877-343.pdf.

Michels:1877:PLS

[10864] J. Michels. The proboscis and lancets of the stable fly. Scientific American, 37(22):344, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
344.pdf.

Anonymous:1877:IT

[10865] Anonymous. Improved tricycle. Scientific American, 37(22):344– 345, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
344a.pdf.

Plelsey:1877:SPI

[10866] T. C. Plelsey. Scientific and practical information. Scientific American, 37(22):345, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
345.pdf.

Riley:1877:CLHa

[10867] C. V. Riley. The curious life-history of our blister beetles. Scientific American, 37(22):346, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
346.pdf.

Anonymous:1877:CRbu

[10868] Anonymous. Communications received. Scientific American, 37 (22):347, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
347b.pdf.

Anonymous:1877:WBIc

[10869] Anonymous. Wants and business inquiries. Scientific American, 37 (22):347–348, December 1, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n22/pdf/scientificamerican12011877-
347c.pdf.

Anonymous:1877:INI

[10870] Anonymous. Inventions: New inventions. Scientific American, 37 (23):360–362, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
360c.pdf.

Anonymous:1877:BPbx

[10871] Anonymous. Business and personal. Scientific American, 37 (23):363, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
363.pdf.

Anonymous:1877:NQbt

[10872] Anonymous. Notes and queries. Scientific American, 37(23): 363–364, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
363a.pdf.

Anonymous:1877:IIbv

[10873] Anonymous. Index of inventions. Scientific American, 37(23):364, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n23/pdf/scientificamerican12081877-364b.pdf.

Anonymous:1877:PNPbu

[10874] Anonymous. Patent notes: Patents. Scientific American, 37 (23):366, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
366.pdf.

Anonymous:1877:IFf

[10875] Anonymous. Improved faucet. Scientific American, 37(23):351, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n23/pdf/scientificamerican12081877-351.pdf.

Anonymous:1877:PMN

[10876] Anonymous. Professor Mayer’s new method of finding the coefficient of expansion of metals and alloys. Scientific American, 37 (23):351–352, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
351a.pdf.

Anonymous:1877:TLI

[10877] Anonymous. Is there a law of invention? Scientific American, 37(23):352, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
352.pdf.

Anonymous:1877:USC

[10878] Anonymous. United states coinage and the great pyramid. Scientific American, 37(23):352–353, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
352a.pdf.

Anonymous:1877:NPOh

[10879] Anonymous. Notes of Patent Office decisions. Scientific American, 37(23):353, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
353.pdf.

Anonymous:1877:SSd

[10880] Anonymous. Shooting stars. Scientific American, 37(23):353, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n23/pdf/scientificamerican12081877-353a.pdf.

Anonymous:1877:ASC

[10881] Anonymous. Adam Scott Cameron. Scientific American, 37(23): 353, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
353b.pdf.

Anonymous:1877:DPS

[10882] Anonymous. The distances of the planets from the Sun. Scientific American, 37(23):353, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
353c.pdf.

Anonymous:1877:ACS

[10883] Anonymous. The abandonment of the Cleopatra at sea. Scientific American, 37(23):354, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
354.pdf.

Anonymous:1877:HPMb

[10884] Anonymous. How to paint magic lanterns. Scientific American, 37 (23):355–356, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
355.pdf.

Rose:1877:PMs

[10885] Joshua Rose. Practical mechanism. Scientific American, 37(23): 356–357, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
356.pdf.

Anonymous:1877:SWF

[10886] Anonymous. Self-watering flower pots. Scientific American, 37 (23):358, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
358.pdf.

Anonymous:1877:ICP

[10887] Anonymous. Improved construction of portable frame buildings. Scientific American, 37(23):358, December 8, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n23/pdf/scientificamerican12081877-358a.pdf.

Seiss:1877:CM

[10888] C. Few Seiss. The Carolina mantis. Scientific American, 37(23):359, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n23/pdf/scientificamerican12081877-359.pdf.

Anonymous:1877:TPa

[10889] Anonymous. The truncated pichiciago. Scientific American, 37(23): 359, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print),

1946-7087 (electronic).  URL http://en.wikipedia.org/wiki/
Pichiciago;   http://www.nature.com/scientificamerican/journal/
v37/n23/pdf/scientificamerican12081877-359a.pdf.

Anonymous:1877:NTW

[10890] Anonymous. New trapless watercloset. Scientific American, 37 (23):360, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
360.pdf.

Anonymous:1877:ANl

[10891] Anonymous. Astronomical notes. Scientific American, 37(23):360, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n23/pdf/scientificamerican12081877-360a.pdf.

Anonymous:1877:DAC

[10892] Anonymous. Dorr’s adjustable hat and coat rack. Scientific American, 37(23):360, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
360b.pdf.

Anonymous:1877:CRbv

[10893] Anonymous. Communications received. Scientific American, 37 (23):364, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
364.pdf.

Anonymous:1877:WBId

[10894] Anonymous. Wants and business inquiries. Scientific American, 37(23):364, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
364a.pdf.

Anonymous:1877:SAb

[10895] Anonymous. The Scientific American. Scientific American, 37 (23):364, December 8, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n23/pdf/scientificamerican12081877-
364c.pdf.

Anonymous:1877:BPby

[10896] Anonymous. Business and personal. Scientific American, 37 (24):379, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
379.pdf.

Anonymous:1877:NQbu

[10897] Anonymous. Notes and queries. Scientific American, 37(24): 379, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
379a.pdf.

Anonymous:1877:IIbw

[10898] Anonymous. Index of inventions. Scientific American, 37(24): 380, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
380.pdf.

Anonymous:1877:PNPbv

[10899] Anonymous. Patent notes: Patents. Scientific American, 37 (24):382, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
382.pdf.

Anonymous:1877:APL

[10900] Anonymous. Amending (?) the patent laws. Scientific American, 37(24):368, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
368.pdf.

Anonymous:1877:CSc

[10901] Anonymous. Curiosities of suicide. Scientific American, 37(24): 368, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
368a.pdf.

Anonymous:1877:CAM

[10902] Anonymous. A common ailment-malarial poisoning. Scientific American, 37(24):368, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
368b.pdf.

Anonymous:1877:NPDa

[10903] Anonymous. Notes of patent decisions of the courts. Scientific American, 37(24):368–369, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
368c.pdf.

Anonymous:1877:PAP

[10904] Anonymous. The proposed amendments to the patent laws. Scientific American, 37(24):369, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
369.pdf.

Anonymous:1877:HTC

[10905] Anonymous. How to true up a Crank pin. Scientific American, 37(24):369, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
369a.pdf.

Anonymous:1877:MNS

[10906] Anonymous. More news from the Sun. Scientific American, 37 (24):369, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
369b.pdf.

Anonymous:1877:HCE

[10907] Anonymous. Horizontal compound engine. Scientific American, 37 (24):370–371, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
370.pdf.

Anonymous:1877:IPN

[10908] Anonymous. Improved protected non-recoil gun. Scientific American, 37(24):371, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
371.pdf.

Anonymous:1877:IPD

[10909] Anonymous. Improved pneumatic dredging tube. Scientific American, 37(24):371, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
371a.pdf.

Anonymous:1877:Cbx

[10910] Anonymous. Communications. Scientific American, 37(24):372, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n24/pdf/scientificamerican12151877-372.pdf.

Riley:1877:CLHb

[10911] C. V. Riley. The curious life-history of our blister beetles number II. Scientific American, 37(24):373–374, December 15, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n24/pdf/scientificamerican12151877-373.pdf.

Anonymous:1877:CIA

[10912] Anonymous. The corrugated iron air bridge and fuel economizer. Scientific American, 37(24):374, December 15, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n24/pdf/scientificamerican12151877-374.pdf.

Anonymous:1877:CLS

[10913] Anonymous. Combination lathe, scroll saw, etc. Scientific American, 37(24):374, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
374a.pdf.

Anonymous:1877:DBP

[10914] Anonymous. The duck-billed platypus. Scientific American, 37 (24):375, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
375.pdf.

Anonymous:1877:PBV

[10915] Anonymous. Proposed balloon voyage to the North Pole. Scientific American, 37(24):375, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
375a.pdf.

Anonymous:1877:MHW

[10916] Anonymous. Machines that hear and write. Scientific American, 37 (24):376–378, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
376.pdf.

Anonymous:1877:JWD

[10917] Anonymous. John William Draper. Scientific American, 37(24): 378, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
378.pdf.

Anonymous:1877:ASO

[10918] Anonymous. Automatic shaft oiler. Scientific American, 37(24): 378, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
378a.pdf.

Anonymous:1877:CRbw

[10919] Anonymous. Communications received. Scientific American, 37 (24):379, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
379b.pdf.

Anonymous:1877:WBIe

[10920] Anonymous. Wants and business inquiries. Scientific American, 37 (24):379–380, December 15, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n24/pdf/scientificamerican12151877-
379c.pdf.

Anonymous:1877:BNBba

[10921] Anonymous. Books: New books and publications. Scientific American, 37(25):394, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
394.pdf.

Anonymous:1877:BPbz

[10922] Anonymous. Business and personal. Scientific American, 37 (25):395, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
395.pdf.

Anonymous:1877:NQbv

[10923] Anonymous. Notes and queries. Scientific American, 37(25): 395–396, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
395a.pdf.

Anonymous:1877:PNPbw

[10924] Anonymous. Patent notes: Patents. Scientific American, 37 (25):398, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
398.pdf.

Anonymous:1877:TFK

[10925] Anonymous. The tuttle family knitter. Scientific American, 37 (25):383, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
383.pdf.

Anonymous:1877:FFA

[10926] Anonymous. The Forster–Firmin amalgamator and ore washer. Scientific American, 37(25):383–384, December 22, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n25/pdf/scientificamerican12221877-383a.pdf.

Anonymous:1877:CIH

[10927] Anonymous. Constructing ice houses. Scientific American, 37 (25):384, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
384.pdf.

Anonymous:1877:TPb

[10928] Anonymous. The talking phonograph. Scientific American, 37(25): 384–385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
384a.pdf.

Anonymous:1877:EVX

[10929] Anonymous. The end of volume XXXVII. Scientific American, 37(25):385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
385.pdf.

Anonymous:1877:WSW

[10930] Anonymous. Winter shop work — building boats. Scientific American, 37(25):385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
385a.pdf.

Anonymous:1877:BIB

[10931] Anonymous. Building ice boats. Scientific American, 37(25): 385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
385b.pdf.

Anonymous:1877:RYP

[10932] Anonymous. Receiving your paper. Scientific American, 37(25): 385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
385c.pdf.

Anonymous:1877:BS

[10933] Anonymous. Build a steamboat. Scientific American, 37(25): 385, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
385d.pdf.

Anonymous:1877:PECb

[10934] Anonymous. The photo-engraving company of New York. Scientific American, 37(25):386–387, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
386.pdf.

Anonymous:1877:Cby

[10935] Anonymous. Communications. Scientific American, 37(25):388– 389, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
388.pdf.

Anonymous:1877:SCM

[10936] Anonymous. Some curious methods of boring tapers. Scientific American, 37(25):389, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
389.pdf.

Anonymous:1877:ISSb

[10937] Anonymous. Improved scroll sawing machine. Scientific American, 37(25):390, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
390.pdf.

Anonymous:1877:DWM

[10938] Anonymous. The Denney washing machine. Scientific American, 37(25):390, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
390a.pdf.

Anonymous:1877:IOG

[10939] Anonymous. Improved oscillating grate bar. Scientific American, 37(25):390, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
390b.pdf.

Anonymous:1877:RBSb

[10940] Anonymous. Revolving boot and shoe heel protector. Scientific American, 37(25):390, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
390c.pdf.

Anonymous:1877:BCa

[10941] Anonymous. The banded chersydrus. Scientific American, 37 (25):391, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
391.pdf.

Anonymous:1877:SEY

[10942] Anonymous. Storks eating Young rabbits. Scientific American, 37 (25):391–392, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
391a.pdf.

Anonymous:1877:EMP

[10943] Anonymous. An easy method of producing bas reliefs. Scientific American, 37(25):392–393, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
392.pdf.

Anonymous:1877:LNK

[10944] Anonymous. The limits of natural knowledge. Scientific American, 37(25):393–394, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
393.pdf.

Anonymous:1877:CRbx

[10945] Anonymous. Communications received. Scientific American, 37 (25):396, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
396.pdf.

Anonymous:1877:WBIf

[10946] Anonymous. Wants and business inquiries. Scientific American, 37(25):396, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
396a.pdf.

Anonymous:1877:SAd

[10947] Anonymous. The Scientific American. Scientific American, 37 (25):396, December 22, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n25/pdf/scientificamerican12221877-
396b.pdf.

Anonymous:1877:BNBbb

[10948] Anonymous. Books: New books and publications. Scientific American, 37(26):409, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
409a.pdf.

Anonymous:1877:BPca

[10949] Anonymous. Business and personal. Scientific American, 37 (26):410, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
410.pdf.

Anonymous:1877:NQbw

[10950] Anonymous. Notes and queries. Scientific American, 37(26): 410, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
410a.pdf.

Anonymous:1877:IIbx

[10951] Anonymous. Index of inventions. Scientific American, 37(26): 410, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
410d.pdf.

Anonymous:1877:I

[10952] Anonymous. Index. Scientific American, 37(26):411–412, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n26/pdf/scientificamerican12291877-411.pdf.

Anonymous:1877:PNPbx

[10953] Anonymous. Patent notes: Patents. Scientific American, 37 (26):414, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
414.pdf.

Anonymous:1877:CPd

[10954] Anonymous. A curious pocketbook. Scientific American, 37(26): 399, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
399.pdf.

Anonymous:1877:BMM

[10955] Anonymous. Brick-making machine. Scientific American, 37(26): 399, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
399a.pdf.

Anonymous:1877:BCG

[10956] Anonymous. The best Christmas gift. Scientific American, 37 (26):400, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
400.pdf.

Anonymous:1877:STE

[10957] Anonymous. Scientific traveling expeditions. Scientific American, 37(26):400, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
400a.pdf.

Anonymous:1877:AMI

[10958] Anonymous. The American machinery and inventions, wanted abroad. Scientific American, 37(26):400–401, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n26/pdf/scientificamerican12291877-400b.pdf.

Anonymous:1877:CSS

[10959] Anonymous. Cuca as a strength sustainer. Scientific American, 37(26):401, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
401.pdf.

Anonymous:1877:TSH

[10960] Anonymous. 30-ton steam hammer. Scientific American, 37 (26):402, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
402.pdf.

Anonymous:1877:SSRb

[10961] Anonymous. Stevenson’s suspension railway. Scientific American, 37(26):402, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
402a.pdf.

Anonymous:1877:UBS

[10962] Anonymous. The uchatius bronze steel gun. Scientific American, 37 (26):403–404, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
403.pdf.

Anonymous:1877:Cbz

[10963] Anonymous. Communications. Scientific American, 37(26):404, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v37/n26/pdf/scientificamerican12291877-404.pdf.

Rose:1877:PCE

[10964] Joshua Rose. Producing cutting edges for tools and instruments. Scientific American, 37(26):404–405, December 29, 1877. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v37/
n26/pdf/scientificamerican12291877-404a.pdf.

Anonymous:1877:SIG

[10965] Anonymous. Stoy’s improved gasket. Scientific American, 37 (26):406, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
406.pdf.

Anonymous:1877:PHE

[10966] Anonymous. Portable hoisting engine. Scientific American, 37 (26):406, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
406a.pdf.

Anonymous:1877:ISSc

[10967] Anonymous. Improved stand for smoothing fabrics. Scientific American, 37(26):406, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
406b.pdf.

Anonymous:1877:FB

[10968] Anonymous. The fringed birthwort. Scientific American, 37(26): 407, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
407.pdf.

Anonymous:1877:CHb

[10969] Anonymous. A curious hybrid. Scientific American, 37(26):407– 408, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
407a.pdf.

Anonymous:1877:NPDb

[10970] Anonymous. Notes of the patent, decisions of the courts. Scientific American, 37(26):409, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
409.pdf.

Anonymous:1877:CRby

[10971] Anonymous. Communications received. Scientific American, 37 (26):410, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
410b.pdf.

Anonymous:1877:WBIg

[10972] Anonymous. Wants and business inquiries. Scientific American, 37(26):410, December 29, 1877. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v37/n26/pdf/scientificamerican12291877-
410c.pdf.

Anonymous:1878:BPa

[10973] Anonymous. Business and personal. Scientific American, 38(1):11, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-11.pdf.

Anonymous:1878:NQa

[10974] Anonymous. Notes and queries. Scientific American, 38(1):11–12, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-11a.pdf.

Anonymous:1878:IIa

[10975] Anonymous. Index of inventions. Scientific American, 38(1):12, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-12b.pdf.

Anonymous:1878:PNPa

[10976] Anonymous. Patent notes: Patents. Scientific American, 38(1):14, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-14.pdf.

Anonymous:1878:UHS

[10977] Anonymous. Utilization of the heat of the sun, a new cement, and more. Scientific American, 38(1):1, January 5, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n1/pdf/scientificamerican01051878-1.pdf.

Anonymous:1878:MCPa

[10978] Anonymous. The Meridian Circle in the Paris Observatory. Scientific American, 38(1):1, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
1a.pdf.

Anonymous:1878:CHE

[10979] Anonymous. Curious hydraulic experiments. Scientific American, 38(1):2, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
2.pdf.

Anonymous:1878:PEa

[10980] Anonymous. The Paris Exposition. Scientific American, 38(1):2, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-2a.pdf.

Anonymous:1878:RSA

[10981] Anonymous. Rubies and sapphires, artificially produced. Scientific American, 38(1):2, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
2b.pdf.

Anonymous:1878:ECE

[10982] Anonymous. The education of a civil engineer. Scientific American, 38(1):2–3, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
2c.pdf.

Anonymous:1878:NYAa

[10983] Anonymous. New York Academy of Sciences. Scientific American, 38(1):3, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
3.pdf.

Anonymous:1878:SG

[10984] Anonymous. Silvering glass. Scientific American, 38(1):3, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-3a.pdf.

Anonymous:1878:IMCa

[10985] Anonymous. An ingenious method of cutting spiral or rifle grooves with an ordinary planer. Scientific American, 38(1):4, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-4.pdf.

Anonymous:1878:NDT

[10986] Anonymous. New diamond tool millstone dressing machine. Scientific American, 38(1):5, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
5.pdf.

Anonymous:1878:RIB

[10987] Anonymous. Rice’s improved boiler feeder. Scientific American, 38(1):6, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-6.pdf.

Anonymous:1878:RIS

[10988] Anonymous. Roth’s improved saw file guide. Scientific American, 38(1):6, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
6a.pdf.

Anonymous:1878:KIF

[10989] Anonymous. Konz’s improved furniture Caster. Scientific American, 38(1):6, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
6b.pdf.

Anonymous:1878:DSJ

[10990] Anonymous. Draper’s stovepipe joint. Scientific American, 38(1):6, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-6c.pdf.

Anonymous:1878:SMa

[10991] Anonymous. The silky marmoset. Scientific American, 38(1):7, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-7.pdf.

Anonymous:1878:BO

[10992] Anonymous. A beautiful orchid. Scientific American, 38(1):7, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-7a.pdf.

Anonymous:1878:HCCa

[10993] Anonymous. Hymer’s coal cabinet. Scientific American, 38(1):8, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-8.pdf.

Anonymous:1878:AN

[10994] Anonymous. Astronomical notes. Scientific American, 38(1):8, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-8a.pdf.

Anonymous:1878:REa

[10995] Anonymous. Reeves’ excavator. Scientific American, 38(1):8, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-8b.pdf.

Anonymous:1878:Ca

[10996] Anonymous. Communications. Scientific American, 38(1):9–10, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-9.pdf.

Anonymous:1878:CRa

[10997] Anonymous. Communications received. Scientific American, 38(1):12, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-12.pdf.

Anonymous:1878:WBIa

[10998] Anonymous. Wants and business inquiries. Scientific American, 38(1):12, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n1/pdf/scientificamerican01051878-
12a.pdf.

Anonymous:1878:SAa

[10999] Anonymous. The Scientific American. Scientific American, 38(1):12, January 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n1/pdf/scientificamerican01051878-12c.pdf.

Anonymous:1878:BNBa

[11000] Anonymous. Books: New books and publications. Scientific American, 38(2):26, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
26.pdf.

Anonymous:1878:BPb

[11001] Anonymous. Business and personal. Scientific American, 38(2):27, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-27.pdf.

Anonymous:1878:NQb

[11002] Anonymous. Notes and queries. Scientific American, 38(2):27–28, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-27a.pdf.

Anonymous:1878:IIb

[11003] Anonymous. Index of inventions. Scientific American, 38(2):28, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-28b.pdf.

Anonymous:1878:PNPb

[11004] Anonymous. Patent notes: Patents. Scientific American, 38(2):30, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-30.pdf.

Anonymous:1878:RS

[11005] Anonymous. The road to success. Scientific American, 38(2):16, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-16.pdf.

Anonymous:1878:MTF

[11006] Anonymous. The mechanical theory of forgetfulness. Scientific American, 38(2):16–17, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
16a.pdf.

Anonymous:1878:PMP

[11007] Anonymous. The Patent Middlings Purifier Case. — another important decision by the Supreme Court. Scientific American, 38(2):17, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-17.pdf.

Anonymous:1878:Cb

[11008] Anonymous. Chloroform. Scientific American, 38(2):17, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-17a.pdf.

Anonymous:1878:TMTa

[11009] Anonymous. The trouve multiple telephone. Scientific American, 38(2):17, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
17b.pdf.

Anonymous:1878:NPD

[11010] Anonymous. Notes of patent decisions. Scientific American, 38 (2):17, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
17c.pdf.

Anonymous:1878:OSa

[11011] Anonymous. To our subscribers. Scientific American, 38(2):17–18, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-17d.pdf.

Anonymous:1878:NYE

[11012] Anonymous. The New York elevated railway. Scientific American, 38(2):18–20, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
18.pdf.

Anonymous:1878:MWa

[11013] Anonymous. Mineral wool. Scientific American, 38(2):20, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-20.pdf.

Anonymous:1878:Cc

[11014] Anonymous. Communications. Scientific American, 38(2):21, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-21.pdf.

Anonymous:1878:BPE

[11015] Anonymous. Buildings of the Paris Exposition of 1878. Scientific American, 38(2):22, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
22.pdf.

Anonymous:1878:SCG

[11016] Anonymous. Stolp’s complete gear instrument. Scientific American, 38(2):22–23, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
22a.pdf.

Anonymous:1878:MWS

[11017] Anonymous. Mery’s weighing scoop. Scientific American, 38 (2):23–24, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
23.pdf.

Michels:1878:CBD

[11018] John Michels. Curious blood disks. Scientific American, 38(2):24, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n2/pdf/scientificamerican01121878-24.pdf.

Anonymous:1878:SSP

[11019] Anonymous. Steam syphon pumps. Scientific American, 38(2): 25–26, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
25.pdf.

Anonymous:1878:CRb

[11020] Anonymous. Communications received. Scientific American, 38 (2):28, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
28.pdf.

Anonymous:1878:WBIb

[11021] Anonymous. Wants and business inquiries. Scientific American, 38(2):28, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
28a.pdf.

Anonymous:1878:SAb

[11022] Anonymous. The Scientific American. Scientific American, 38 (2):28, January 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n2/pdf/scientificamerican01121878-
28c.pdf.

Anonymous:1878:BPc

[11023] Anonymous. Business and personal. Scientific American, 38(3):43, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-43.pdf.

Anonymous:1878:NQc

[11024] Anonymous. Notes and queries. Scientific American, 38(3):43–44, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-43b.pdf.

Anonymous:1878:IIc

[11025] Anonymous. Index of inventions. Scientific American, 38(3):44, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-44b.pdf.

Anonymous:1878:PNPc

[11026] Anonymous. Patent notes: Patents. Scientific American, 38(3):46, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-46.pdf.

Anonymous:1878:MDPa

[11027] Anonymous. The manufacture of drain pipe. Scientific American, 38(3):31, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
31.pdf.

Anonymous:1878:NHC

[11028] Anonymous. Natural history collections as educators. Scientific American, 38(3):32, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
32.pdf.

Anonymous:1878:PCS

[11029] Anonymous. Preventing collisions at sea. Scientific American, 38(3):32, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
32a.pdf.

Anonymous:1878:CBE

[11030] Anonymous. Carbon burned in an electro-chemical battery. Scientific American, 38(3):32, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
32b.pdf.

Anonymous:1878:DB

[11031] Anonymous. Drawing on the blackboard. Scientific American, 38 (3):32–33, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
32c.pdf.

Anonymous:1878:CDFa

[11032] Anonymous. Coal dust fuel. Scientific American, 38(3):33, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-33.pdf.

Anonymous:1878:GSW

[11033] Anonymous. Government scientific work. Scientific American, 38(3):33, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
33a.pdf.

Anonymous:1878:OSb

[11034] Anonymous. To our subscribers. Scientific American, 38(3):33–34, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-33b.pdf.

Anonymous:1878:CMVa

[11035] Anonymous. The Colorado mud volcanoes. Scientific American, 38(3):34, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
34.pdf.

Anonymous:1878:REL

[11036] Anonymous. The regnier electric light. Scientific American, 38 (3):34, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
34a.pdf.

Anonymous:1878:LTH

[11037] Anonymous. Lovegrove’s two-horse power engine. Scientific American, 38(3):34, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
34b.pdf.

Anonymous:1878:GHP

[11038] Anonymous. The Godfrey and Howson puddling furnace. Scientific American, 38(3):35–36, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
35.pdf.

Anonymous:1878:Cd

[11039] Anonymous. Communications. Scientific American, 38(3):36, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-36.pdf.

Rose:1878:PMa

[11040] Joshua Rose. Practical mechanism. Scientific American, 38(3): 36–37, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
36a.pdf.

Anonymous:1878:IME

[11041] Anonymous. Improved marine engine governor. Scientific American, 38(3):38, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
38.pdf.

Anonymous:1878:IFM

[11042] Anonymous. Improved fulling mill. Scientific American, 38(3):38, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-38a.pdf.

Anonymous:1878:CCa

[11043] Anonymous. The cultivation of cinchona. Scientific American, 38 (3):39–40, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
39.pdf.

Anonymous:1878:WHB

[11044] Anonymous. Water heater for baths. Scientific American, 38(3):40, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-40.pdf.

Anonymous:1878:NSG

[11045] Anonymous. A novel steam generator. Scientific American, 38 (3):40–41, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
40a.pdf.

Anonymous:1878:NTM

[11046] Anonymous. A new tide mill. Scientific American, 38(3):41, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-41.pdf.

Anonymous:1878:DT

[11047] Anonymous. The Dreadnought tricycle. Scientific American, 38 (3):41–42, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
41a.pdf.

Anonymous:1878:AO

[11048] Anonymous. An astonishing offer. Scientific American, 38(3):43, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n3/pdf/scientificamerican01191878-43a.pdf.

Anonymous:1878:CRc

[11049] Anonymous. Communications received. Scientific American, 38 (3):44, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
44.pdf.

Anonymous:1878:WBIc

[11050] Anonymous. Wants and business inquiries. Scientific American, 38(3):44, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
44a.pdf.

Anonymous:1878:SAc

[11051] Anonymous. The Scientific American. Scientific American, 38 (3):44, January 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n3/pdf/scientificamerican01191878-
44c.pdf.

Anonymous:1878:BPd

[11052] Anonymous. Business and personal. Scientific American, 38(4):59, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-59.pdf.

Anonymous:1878:NQNa

[11053] Anonymous. Notes and queries: Notes & queries. Scientific American, 38(4):59–60, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
59a.pdf.

Anonymous:1878:IId

[11054] Anonymous. Index of inventions. Scientific American, 38(4):60, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-60.pdf.

Anonymous:1878:PNPd

[11055] Anonymous. Patent notes: Patents. Scientific American, 38(4):62, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-62.pdf.

Anonymous:1878:BPF

[11056] Anonymous. The Berryman patent feed water heater. Scientific American, 38(4):47, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
47.pdf.

Anonymous:1878:SPG

[11057] Anonymous. Successful progress of the great jetties. Scientific American, 38(4):48, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
48.pdf.

Anonymous:1878:US

[11058] Anonymous. Unscientific scientists. Scientific American, 38(4):48, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-48a.pdf.

Anonymous:1878:NSA

[11059] Anonymous. A new system of armor plating needed. Scientific American, 38(4):48–49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
48b.pdf.

Anonymous:1878:COR

[11060] Anonymous. Confidentially, with our readers. Scientific American, 38(4):49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
49.pdf.

Anonymous:1878:GTM

[11061] Anonymous. Government tests of magazine guns. Scientific American, 38(4):49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
49a.pdf.

Anonymous:1878:RDE

[11062] Anonymous. A remarkable and disastrous explosion. Scientific American, 38(4):49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
49b.pdf.

Anonymous:1878:AEP

[11063] Anonymous. The American exhibit at the Paris Exposition. Scientific American, 38(4):49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
49c.pdf.

Anonymous:1878:OSc

[11064] Anonymous. To our subscribers. Scientific American, 38(4):49, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-49d.pdf.

Anonymous:1878:UGQ

[11065] Anonymous. Utilizations of the goose quill. Scientific American, 38(4):50–51, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
50.pdf.

Anonymous:1878:RDMa

[11066] Anonymous. Radial drilling machine. Scientific American, 38 (4):51, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
51.pdf.

Anonymous:1878:Ce

[11067] Anonymous. Communications. Scientific American, 38(4):52–53, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-52.pdf.

Anonymous:1878:FTGa

[11068] Anonymous. Fisher’s torpedo guard. Scientific American, 38(4):54, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n4/pdf/scientificamerican01261878-54.pdf.

Anonymous:1878:IOC

[11069] Anonymous. Improved ore crusher and pulverizer. Scientific American, 38(4):54, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
54a.pdf.

Anonymous:1878:NBC

[11070] Anonymous. The no. 10 bolt cutter and nut tapper. Scientific American, 38(4):54–55, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
54b.pdf.

Anonymous:1878:CSHa

[11071] Anonymous. A celebrated short-horned cow. Scientific American, 38(4):55, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
55.pdf.

Anonymous:1878:CPC

[11072] Anonymous. Chapell’s pipe coupling. Scientific American, 38 (4):55–56, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
55a.pdf.

Anonymous:1878:BIK

[11073] Anonymous. Boland’s improved kneading machine. Scientific American, 38(4):56–57, January 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
56.pdf.

Anonymous:1878:RMT

[11074] Anonymous. The reproduction of musical tones by electro-magnetism. Scientific American, 38(4):57–59, January 26, 1878. CODEN SCA-MAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://

www.nature.com/scientificamerican/journal/v38/n4/pdf/scientificamerican01261878-
57.pdf.

Anonymous:1878:BPg

[11075] Anonymous. Business and personal. Scientific American, 38 (5):75, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
75.pdf.

Anonymous:1878:BNBb

[11076] Anonymous. Books: New books and publications. Scientific American, 38(5):75, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
75a.pdf.

Anonymous:1878:NQNb

[11077] Anonymous. Notes and queries: Notes & queries. Scientific American, 38(5):75–76, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
75b.pdf.

Anonymous:1878:IIe

[11078] Anonymous. Index of inventions. Scientific American, 38(5):76, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n5/pdf/scientificamerican02021878-76b.pdf.

Anonymous:1878:PNPe

[11079] Anonymous. Patent notes: Patents. Scientific American, 38 (5):78, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
78.pdf.

Anonymous:1878:GER

[11080] Anonymous. The Gilbert elevated railway. Scientific American, 38(5):63, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
63.pdf.

Anonymous:1878:TSQ

[11081] Anonymous. A three sided question. Scientific American, 38 (5):64, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
64.pdf.

Anonymous:1878:LAA

[11082] Anonymous. The liquefaction of air and all the permanent gases. Scientific American, 38(5):64–65, February 2, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n5/pdf/scientificamerican02021878-64a.pdf.

Anonymous:1878:MEa

[11083] Anonymous. Mysterious explosions. Scientific American, 38(5): 65, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
65.pdf.

Anonymous:1878:NPLa

[11084] Anonymous. Notes of patent law decisions. Scientific American, 38(5):65, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
65a.pdf.

Anonymous:1878:CBF

[11085] Anonymous. Corrugated boiler flues. Scientific American, 38 (5):67–68, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
67.pdf.

Wright:1878:ANa

[11086] Berlin H. Wright. Astronomical notes. Scientific American, 38 (5):68–69, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
68.pdf.

Anonymous:1878:QSD

[11087] Anonymous. Quick-speed drill. Scientific American, 38(5):69, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n5/pdf/scientificamerican02021878-69.pdf.

Anonymous:1878:NYAb

[11088] Anonymous. New York Academy of Sciences. Scientific American, 38(5):69, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
69a.pdf.

Anonymous:1878:PWT

[11089] Anonymous. Paper and wire testers. Scientific American, 38 (5):69–70, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
69b.pdf.

Anonymous:1878:FIL

[11090] Anonymous. Fawcett’s improved ladle. Scientific American, 38 (5):70, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
70.pdf.

Anonymous:1878:NBPa

[11091] Anonymous. The new Baxter portable engine. Scientific American, 38(5):70, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
70a.pdf.

Anonymous:1878:MAT

[11092] Anonymous. Metallic arches for tunnels. Scientific American, 38(5):70, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
70b.pdf.

Anonymous:1878:CIB

[11093] Anonymous. Carnes’ improved basin stopper. Scientific American, 38(5):70, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
70c.pdf.

Pictet:1878:LO

[11094] M. Raoul Pictet. The liquefaction of oxygen. Scientific American, 38(5):71, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
71.pdf.

Anonymous:1878:MPR

[11095] Anonymous. Machine for preparing ramie fiber. Scientific American, 38(5):72, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
72.pdf.

Anonymous:1878:HSCa

[11096] Anonymous. How submarine cables are destroyed. Scientific American, 38(5):73, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
73.pdf.

Anonymous:1878:NAC

[11097] Anonymous. New apparatus for the compression of hydrogen and oxygen. Scientific American, 38(5):73–75, February 2, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n5/pdf/scientificamerican02021878-73a.pdf.

Anonymous:1878:CRd

[11098] Anonymous. Communications received. Scientific American, 38 (5):76, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
76.pdf.

Anonymous:1878:WBId

[11099] Anonymous. Wants and business inquiries. Scientific American, 38(5):76, February 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n5/pdf/scientificamerican02021878-
76a.pdf.

Anonymous:1878:BPh

[11100] Anonymous. Business and personal. Scientific American, 38 (6):91, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
91.pdf.

Anonymous:1878:NQd

[11101] Anonymous. Notes and queries. Scientific American, 38(6):91– 92, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
91a.pdf.

Anonymous:1878:IIf

[11102] Anonymous. Index of inventions. Scientific American, 38(6):92, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n6/pdf/scientificamerican02091878-92a.pdf.

Anonymous:1878:PNPf

[11103] Anonymous. Patent notes: Patents. Scientific American, 38 (6):94, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
94.pdf.

Anonymous:1878:HPb

[11104] Anonymous. The horizontal pendulum. Scientific American, 38 (6):79–80, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
79.pdf.

Anonymous:1878:MP

[11105] Anonymous. Medical progress in 1877. Scientific American, 38 (6):80, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
80.pdf.

Anonymous:1878:GMM

[11106] Anonymous. Is gravity a mode of motion? Scientific American, 38(6):80, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
80a.pdf.

Anonymous:1878:SLa

[11107] Anonymous. Stick to the law. Scientific American, 38(6):80–81, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n6/pdf/scientificamerican02091878-80b.pdf.

Anonymous:1878:POE

[11108] Anonymous. Patent Office employees to be disgraced. Scientific American, 38(6):81, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
81.pdf.

Anonymous:1878:MM

[11109] Anonymous. The mail of the Metropolis. Scientific American, 38(6):81, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
81a.pdf.

Anonymous:1878:HDR

[11110] Anonymous. Heinrich Daniel Ruhmkorff. Scientific American, 38(6):81, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
81b.pdf.

Anonymous:1878:MTW

[11111] Anonymous. Modern torpedo warfare. Scientific American, 38 (6):82, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
82.pdf.

Anonymous:1878:FDF

[11112] Anonymous. Folding-door fireproof safe. Scientific American, 38 (6):82, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
82a.pdf.

Anonymous:1878:SIW

[11113] Anonymous. Smith’s improved windmill. Scientific American, 38 (6):83, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
83.pdf.

Anonymous:1878:NDM

[11114] Anonymous. New digging machine. Scientific American, 38(6): 83–84, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
83a.pdf.

Anonymous:1878:Cf

[11115] Anonymous. Communications. Scientific American, 38(6):84, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n6/pdf/scientificamerican02091878-84.pdf.

Anonymous:1878:MC

[11116] Anonymous. Making callipers. Scientific American, 38(6):85, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n6/pdf/scientificamerican02091878-85.pdf.

Anonymous:1878:IPC

[11117] Anonymous. Improved portable cider mill. Scientific American, 38(6):86, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
86.pdf.

Anonymous:1878:PICa

[11118] Anonymous. Powell’s improved cotton harvester. Scientific American, 38(6):86, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
86a.pdf.

Seiss:1878:LNC

[11119] C. Few Seiss. The long-necked chelodine. Scientific American, 38(6):87, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
87.pdf.

Anonymous:1878:HSK

[11120] Anonymous. The high school of Kolapore, India. Scientific American, 38(6):87, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
87a.pdf.

Anonymous:1878:WW

[11121] Anonymous. Whales and whalebone. Scientific American, 38 (6):88–89, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
88.pdf.

Anonymous:1878:BTC

[11122] Anonymous. Beard’s thill coupling. Scientific American, 38(6): 89, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
89.pdf.

Michels:1878:FM

[11123] John Michels. A fowl monstrosity. Scientific American, 38(6): 89–90, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
89a.pdf.

Wright:1878:ANb

[11124] Berlin H. Wright. Astronomical notes. Scientific American, 38 (6):90, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
90.pdf.

Anonymous:1878:WBIe

[11125] Anonymous. Wants and business inquiries. Scientific American, 38(6):92, February 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n6/pdf/scientificamerican02091878-
92.pdf.

Anonymous:1878:BPi

[11126] Anonymous. Business and personal. Scientific American, 38(7): 107, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
107.pdf.

Anonymous:1878:NQe

[11127] Anonymous. Notes and queries. Scientific American, 38(7):107– 108, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
107a.pdf.

Anonymous:1878:IIg

[11128] Anonymous. Index of inventions. Scientific American, 38(7):108, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n7/pdf/scientificamerican02161878-108a.pdf.

Anonymous:1878:PNPg

[11129] Anonymous. Patent notes: Patents. Scientific American, 38 (7):110, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
110.pdf.

Anonymous:1878:CRM

[11130] Anonymous. The cash recording machine. Scientific American, 38(7):95, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
95.pdf.

Anonymous:1878:POC

[11131] Anonymous. A Patent Office Court. Scientific American, 38 (7):96, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
96.pdf.

Anonymous:1878:ARC

[11132] Anonymous. The Annual Report of the Commissioner of Patents. Scientific American, 38(7):96, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
96a.pdf.

Anonymous:1878:TW

[11133] Anonymous. Telegraphing without wires. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
97.pdf.

Anonymous:1878:CBR

[11134] Anonymous. Color-blindness in its relations to the safety of the traveling public. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n7/pdf/scientificamerican02161878-97a.pdf.

Anonymous:1878:ELP

[11135] Anonymous. The electric light in photography. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
97b.pdf.

Anonymous:1878:NE

[11136] Anonymous. “That is near enough”. Scientific American, 38 (7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
97c.pdf.

Anonymous:1878:DSD

[11137] Anonymous. Dangerous smoke and drink. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
97d.pdf.

Anonymous:1878:SCa

[11138] Anonymous. The shop clock. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n7/pdf/scientificamerican02161878-97e.pdf.

Anonymous:1878:BPj

[11139] Anonymous. The bullion product of 1877. Scientific American, 38(7):97, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
97f.pdf.

Anonymous:1878:NASa

[11140] Anonymous. New adjustable swivel hanger. Scientific American, 38(7):98, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
98.pdf.

Anonymous:1878:SIE

[11141] Anonymous. Shaw’s improved exhaust nozzle. Scientific American, 38(7):98, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
98a.pdf.

Anonymous:1878:ASL

[11142] Anonymous. Appliances for supporting life in irrespirable media. Scientific American, 38(7):99–100, February 16, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n7/pdf/scientificamerican02161878-99.pdf.

Anonymous:1878:Cg

[11143] Anonymous. Communications. Scientific American, 38(7):100– 101, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
100.pdf.

Anonymous:1878:WIC

[11144] Anonymous. Wise’s ice cream beater. Scientific American, 38 (7):102, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
102.pdf.

Anonymous:1878:IHMa

[11145] Anonymous. An improved hoisting machine. Scientific American, 38(7):102, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
102a.pdf.

Anonymous:1878:TWSa

[11146] Anonymous. The Thompson wooden Spring for vehicles. Scientific American, 38(7):103, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
103.pdf.

Anonymous:1878:IRT

[11147] Anonymous. The India rubber tree. Scientific American, 38 (7):103, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
103a.pdf.

Anonymous:1878:CM

[11148] Anonymous. The codling moth. Scientific American, 38(7):103– 104, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
103b.pdf.

Anonymous:1878:MHG

[11149] Anonymous. M. Henri Giffard’s new apparatus for the production of hydrogen in large quantities. Scientific American, 38(7):104–105, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n7/pdf/scientificamerican02161878-104.pdf.

Anonymous:1878:PGE

[11150] Anonymous. Prevention of gaseous emanations from drains and sewers. Scientific American, 38(7):105–106, February 16, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n7/pdf/scientificamerican02161878-105.pdf.

Anonymous:1878:WBIf

[11151] Anonymous. Wants and business inquiries. Scientific American, 38(7):108, February 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n7/pdf/scientificamerican02161878-
108.pdf.

Anonymous:1878:BPk

[11152] Anonymous. Business and personal. Scientific American, 38(8): 123, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
123.pdf.

Anonymous:1878:BNBc

[11153] Anonymous. Books: New books and publications. Scientific American, 38(8):123, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
123a.pdf.

Anonymous:1878:NQf

[11154] Anonymous. Notes and queries. Scientific American, 38(8):123– 124, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
123b.pdf.

Anonymous:1878:IIh

[11155] Anonymous. Index of inventions. Scientific American, 38(8):124, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n8/pdf/scientificamerican02231878-124.pdf.

Anonymous:1878:PNPh

[11156] Anonymous. Patent notes: Patents. Scientific American, 38 (8):126, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
126.pdf.

Anonymous:1878:LGC

[11157] Anonymous. Liquefaction of gases — Cailletet’s experiments. Scientific American, 38(8):111–112, February 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n8/pdf/scientificamerican02231878-111.pdf.

Anonymous:1878:PSa

[11158] Anonymous. Paradoxes in steam. Scientific American, 38(8):112, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n8/pdf/scientificamerican02231878-112.pdf.

Anonymous:1878:AE

[11159] Anonymous. Arctic explorations. Scientific American, 38(8):112– 113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
112a.pdf.

Anonymous:1878:WSD

[11160] Anonymous. What steam has done for France. Scientific American, 38(8):113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
113.pdf.

Anonymous:1878:IRa

[11161] Anonymous. Improvement of rivers. Scientific American, 38(8): 113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
113a.pdf.

Anonymous:1878:PPa

[11162] Anonymous. Printing the patents. Scientific American, 38(8): 113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
113b.pdf.

Anonymous:1878:CAA

[11163] Anonymous. The carbolic acid or antiseptic treatment of wounds. Scientific American, 38(8):113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
113c.pdf.

Anonymous:1878:WNO

[11164] Anonymous. The Washington Naval Observatory. Scientific American, 38(8):113, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
113d.pdf.

Anonymous:1878:IDT

[11165] Anonymous. Improved Draughtsman’s table. Scientific American, 38(8):114, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
114.pdf.

Anonymous:1878:IMD

[11166] Anonymous. Improved method for drying lumber. Scientific American, 38(8):114, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
114a.pdf.

Anonymous:1878:BTG

[11167] Anonymous. The bon ton game. Scientific American, 38(8):114, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n8/pdf/scientificamerican02231878-114b.pdf.

Anonymous:1878:VIB

[11168] Anonymous. Vanier’s improved brick machine. Scientific American, 38(8):115, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
115.pdf.

Anonymous:1878:BPS

[11169] Anonymous. Baugh’s patent sectional mill. Scientific American, 38 (8):115–116, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
115a.pdf.

Anonymous:1878:Ch

[11170] Anonymous. Communications. Scientific American, 38(8):116– 117, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
116.pdf.

Anonymous:1878:PCN

[11171] Anonymous. Professor S. N. Carvalho’s new system of superheating steam. Scientific American, 38(8):118, February 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n8/pdf/scientificamerican02231878-118.pdf.

Anonymous:1878:VIG

[11172] Anonymous. Vitt’s improved grain toller. Scientific American, 38(8):118, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
118a.pdf.

Anonymous:1878:HT

[11173] Anonymous. The horned tragopan. Scientific American, 38(8): 119, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
119.pdf.

Anonymous:1878:TIa

[11174] Anonymous. A Turkish ironclad. Scientific American, 38(8):119, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n8/pdf/scientificamerican02231878-119a.pdf.

Anonymous:1878:SNM

[11175] Anonymous. Sadlier’s new method of raising and lowering screw propellers. Scientific American, 38(8):120, February 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n8/pdf/scientificamerican02231878-120.pdf.

Wright:1878:ANc

[11176] Berlin H. Wright. Astronomical notes. Scientific American, 38 (8):120, February 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n8/pdf/scientificamerican02231878-
120a.pdf.

Anonymous:1878:PIBa

[11177] Anonymous. The physiological influence of barometric pressure. Scientific American, 38(8):121–122, February 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n8/pdf/scientificamerican02231878-121.pdf.

Anonymous:1878:BPl

[11178] Anonymous. Business and personal. Scientific American, 38(9):139, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-139.pdf.

Anonymous:1878:BNBd

[11179] Anonymous. Books: New books and publications. Scientific American, 38(9):139, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
139a.pdf.

Anonymous:1878:NQNc

[11180] Anonymous. Notes and queries: Notes & queries. Scientific American, 38(9):139–140, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
139b.pdf.

Anonymous:1878:IIi

[11181] Anonymous. Index of inventions. Scientific American, 38(9):140, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-140.pdf.

Anonymous:1878:PNPi

[11182] Anonymous. Patent notes: Patents. Scientific American, 38(9):142, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-142.pdf.

Anonymous:1878:KHP

[11183] Anonymous. Kinney’s hydro-pneumatic accumulator. Scientific American, 38(9):127, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
127.pdf.

Anonymous:1878:OPN

[11184] Anonymous. Opposition to patents natural. Scientific American, 38(9):128, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
128.pdf.

Anonymous:1878:AMP

[11185] Anonymous. American manufactures and produce abroad. Scientific American, 38(9):128–129, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
128a.pdf.

Anonymous:1878:RPa

[11186] Anonymous. The rail puzzle. Scientific American, 38(9):129, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-129.pdf.

Anonymous:1878:POMa

[11187] Anonymous. Patent Office models. Scientific American, 38(9):129, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-129a.pdf.

Ricco:1878:TBO

[11188] A. Ricco. Two beautiful optical experiments. Scientific American, 38(9):129, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
129b.pdf.

Anonymous:1878:SJa

[11189] Anonymous. Scientific jugglery. Scientific American, 38(9):129, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-129c.pdf.

Anonymous:1878:BTA

[11190] Anonymous. Bailey’s testing apparatus. Scientific American, 38(9):130, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-130.pdf.

Anonymous:1878:RCP

[11191] Anonymous. Rider compression pumping engine. Scientific American, 38(9):131–132, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
131.pdf.

Anonymous:1878:Ci

[11192] Anonymous. Communications. Scientific American, 38(9):132, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-132.pdf.

Anonymous:1878:EPW

[11193] Anonymous. Economy in pattern work. Scientific American, 38(9):133, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-133.pdf.

Anonymous:1878:LDS

[11194] Anonymous. Law’s double shears and engine. Scientific American, 38(9):134, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
134.pdf.

Anonymous:1878:ISP

[11195] Anonymous. Imhof’s stove pipe expander. Scientific American, 38(9):134, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
134a.pdf.

Anonymous:1878:TCT

[11196] Anonymous. Teeth, claws, tongue, and whiskers. Scientific American, 38(9):135, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
135.pdf.

Anonymous:1878:NBa

[11197] Anonymous. A Nile bridge. Scientific American, 38(9):135–136, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-135a.pdf.

Anonymous:1878:ML

[11198] Anonymous. A mechanical larynx. Scientific American, 38(9):136–137, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-136.pdf.

Anonymous:1878:AGL

[11199] Anonymous. Automatic gas lighting. Scientific American, 38 (9):137–138, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n9/pdf/scientificamerican03021878-
137.pdf.

Wright:1878:ANd

[11200] Berlin H. Wright. Astronomical notes. Scientific American, 38(9):138, March 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n9/pdf/scientificamerican03021878-138.pdf.

Anonymous:1878:BPm

[11201] Anonymous. Business and personal. Scientific American, 38(10):155, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-155.pdf.

Anonymous:1878:NQg

[11202] Anonymous. Notes and queries. Scientific American, 38(10):155, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-155a.pdf.

Anonymous:1878:IIj

[11203] Anonymous. Index of inventions. Scientific American, 38(10):155, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-155b.pdf.

Anonymous:1878:PNPj

[11204] Anonymous. Patent notes: Patents. Scientific American, 38(10):158, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-158.pdf.

Anonymous:1878:CDa

[11205] Anonymous. The hat conformator and its diagrams. Scientific American, 38(10):143, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
143.pdf.

Anonymous:1878:IMCb

[11206] Anonymous. Invention as a means of culture. Scientific American, 38(10):144, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
144.pdf.

Anonymous:1878:GTa

[11207] Anonymous. The genesis of the telephone. Scientific American, 38(10):144, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
144a.pdf.

Anonymous:1878:TIb

[11208] Anonymous. The treatment of the insane. Scientific American, 38(10):144–145, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
144b.pdf.

Anonymous:1878:FTS

[11209] Anonymous. Fish teeth set on hinges. Scientific American, 38(10):145, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-145.pdf.

Anonymous:1878:BHO

[11210] Anonymous. A brief history of our patent system. Scientific American, 38(10):145, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
145a.pdf.

Anonymous:1878:AFC

[11211] Anonymous. The astounding famine in China. Scientific American, 38(10):145, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
145b.pdf.

Anonymous:1878:HWP

[11212] Anonymous. Hayford’s wood-preserving process. Scientific American, 38(10):145, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
145c.pdf.

Anonymous:1878:IFE

[11213] Anonymous. Improved fire escape. Scientific American, 38(10):147, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-147.pdf.

Anonymous:1878:LGM

[11214] Anonymous. The liquefaction of the gases — M. Pictet’s apparatus. Scientific American, 38(10):147, March 9, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n10/pdf/scientificamerican03091878-147a.pdf.

Anonymous:1878:Cj

[11215] Anonymous. Communications. Scientific American, 38(10):148, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-148.pdf.

Rose:1878:PMb

[11216] Joshua Rose and M. E. Practical mechanism. Scientific American, 38(10):149, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
149.pdf.

Anonymous:1878:THE

[11217] Anonymous. Talley’s hydraulic engine. Scientific American, 38(10):150, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-150.pdf.

Anonymous:1878:RSCa

[11218] Anonymous. The ransom siphon condenser. Scientific American, 38(10):150–151, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
150a.pdf.

Anonymous:1878:CCb

[11219] Anonymous. The Cereus Candicans. Scientific American, 38(10):151, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-151.pdf.

Anonymous:1878:BCP

[11220] Anonymous. Black cock and ptarmigan. Scientific American, 38 (10):151–152, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
151a.pdf.

Anonymous:1878:ADAa

[11221] Anonymous. The “Acme” double acting steam pump. Scientific American, 38(10):152–153, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
152.pdf.

Anonymous:1878:SI

[11222] Anonymous. Submarine illumination. Scientific American, 38 (10):153–154, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n10/pdf/scientificamerican03091878-
153.pdf.

Wright:1878:ANe

[11223] Berlin H. Wright. Astronomical notes. Scientific American, 38(10):154, March 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n10/pdf/scientificamerican03091878-154.pdf.

Anonymous:1878:BPn

[11224] Anonymous. Business and personal. Scientific American, 38(11):171, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-171.pdf.

Anonymous:1878:NQh

[11225] Anonymous. Notes and queries. Scientific American, 38(11):171–172, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-171a.pdf.

Anonymous:1878:IIk

[11226] Anonymous. Index of inventions. Scientific American, 38(11):172, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-172a.pdf.

Anonymous:1878:PNPk

[11227] Anonymous. Patent notes: Patents. Scientific American, 38(11):172, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-172b.pdf.

Anonymous:1878:HIM

[11228] Anonymous. The holden ice machine. Scientific American, 38(11):159, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-159.pdf.

Anonymous:1878:WLP

[11229] Anonymous. Wanted: A legal process for confiscating inventions. Scientific American, 38(11):160, March 16, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n11/pdf/scientificamerican03161878-160.pdf.

Anonymous:1878:GP

[11230] Anonymous. Geographical progress. Scientific American, 38(11):160, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-160a.pdf.

Anonymous:1878:CSEa

[11231] Anonymous. A Congressional scientific expedition. Scientific American, 38(11):160–161, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
160b.pdf.

Anonymous:1878:EBN

[11232] Anonymous. Educated brutes at the New York aquarium. Scientific American, 38(11):161, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
161.pdf.

Anonymous:1878:HLC

[11233] Anonymous. “Heat is life — cold is death”. Scientific American, 38(11):161, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
161a.pdf.

Anonymous:1878:LCF

[11234] Anonymous. The lesson of the Chinese famine. Scientific American, 38(11):161, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
161b.pdf.

Anonymous:1878:NSPa

[11235] Anonymous. The National steam pump. Scientific American, 38(11):162, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
162.pdf.

Hille:1878:H

[11236] L. Hille. The helioscope. Scientific American, 38(11):163, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-163.pdf.

Anonymous:1878:PSMa

[11237] Anonymous. Punching and shearing machine. Scientific American, 38(11):163, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
163a.pdf.

Anonymous:1878:Ck

[11238] Anonymous. Communication. Scientific American, 38(11):164–165, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-164.pdf.

Anonymous:1878:NBPb

[11239] Anonymous. A new boat propelling device. Scientific American, 38(11):166, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
166.pdf.

Anonymous:1878:PIBb

[11240] Anonymous. Phillips’ improved buckboard wagon. Scientific American, 38(11):166, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
166a.pdf.

Anonymous:1878:MHR

[11241] Anonymous. Millington’s hydraulic ram. Scientific American, 38(11):166, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
166b.pdf.

Seiss:1878:SD

[11242] C. Few Seiss. Serpents at dinner. Scientific American, 38(11):167, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-167.pdf.

Anonymous:1878:GTb

[11243] Anonymous. Gathering toddy. Scientific American, 38(11):167–168, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-167a.pdf.

Anonymous:1878:CDb

[11244] Anonymous. Colmer’s dosimeter. Scientific American, 38(11):168, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-168.pdf.

Anonymous:1878:LMP

[11245] Anonymous. The life of a million people. Scientific American, 38 (11):168–169, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
168a.pdf.

Anonymous:1878:MPNa

[11246] Anonymous. M. Plante’s new rheostatic machine. Scientific American, 38(11):169, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
169.pdf.

Wright:1878:ANf

[11247] Berlin H. Wright. Astronomical notes. Scientific American, 38(11):169, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-169a.pdf.

Anonymous:1878:Ea

[11248] Anonymous. Evolution. Scientific American, 38(11):170–171, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n11/pdf/scientificamerican03161878-170.pdf.

Anonymous:1878:CRe

[11249] Anonymous. Communications received. Scientific American, 38 (11):172, March 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n11/pdf/scientificamerican03161878-
172.pdf.

Anonymous:1878:BPo

[11250] Anonymous. Business and personal. Scientific American, 38(12):187, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-187.pdf.

Anonymous:1878:BNBe

[11251] Anonymous. Books: New books and publications. Scientific American, 38(12):187, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
187a.pdf.

Anonymous:1878:NQNd

[11252] Anonymous. Notes and queries: Notes & queries. Scientific American, 38(12):187–188, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
187b.pdf.

Anonymous:1878:IIl

[11253] Anonymous. Index of inventions. Scientific American, 38(12):188, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-188a.pdf.

Anonymous:1878:TCG

[11254] Anonymous. Turning a corner on the Gilbert elevated railway. Scientific American, 38(12):175, March 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n12/pdf/scientificamerican03231878-175.pdf.

Anonymous:1878:IWW

[11255] Anonymous. Improvements wanted in working gold and silver ores. Scientific American, 38(12):176, March 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n12/pdf/scientificamerican03231878-176.pdf.

Anonymous:1878:WOM

[11256] Anonymous. Will our Moon ever rise in the west? Scientific American, 38(12):176, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
176a.pdf.

Anonymous:1878:WAI

[11257] Anonymous. Work ahead for inventors. Scientific American, 38 (12):176–177, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
176b.pdf.

Anonymous:1878:NYH

[11258] Anonymous. The New York Herald’s best work. Scientific American, 38(12):177, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
177.pdf.

Anonymous:1878:HS

[11259] Anonymous. The harvest of the sea. Scientific American, 38(12):177, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-177a.pdf.

Anonymous:1878:POMb

[11260] Anonymous. Patent Office models. Scientific American, 38(12):177, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-177b.pdf.

Riley:1878:CM

[11261] C. V. Riley. Clothes moths. Scientific American, 38(12):177, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-177c.pdf.

Anonymous:1878:SBE

[11262] Anonymous. The Steam Boiler Explosion Commission. Scientific American, 38(12):178, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
178.pdf.

Wright:1878:ANg

[11263] Berlin H. Wright. Astronomical notes. Scientific American, 38(12):179, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-179.pdf.

Anonymous:1878:LSPa

[11264] Anonymous. Large steam pumping engines — a duplex pressure pump. Scientific American, 38(12):179, March 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n12/pdf/scientificamerican03231878-179a.pdf.

Anonymous:1878:BNS

[11265] Anonymous. Bozerian’s new shower bath. Scientific American, 38 (12):179–181, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
179b.pdf.

Anonymous:1878:NT

[11266] Anonymous. A new tool. Scientific American, 38(12):181, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-181.pdf.

Anonymous:1878:WNC

[11267] Anonymous. Why nuts come loose and bolts drop out. Scientific American, 38(12):181, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
181a.pdf.

Bottone:1878:S

[11268] Selimo Bottone. The saccharometer. Scientific American, 38(12):181, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-181b.pdf.

Anonymous:1878:IHC

[11269] Anonymous. Improved hose coupling. Scientific American, 38(12):182, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-182.pdf.

Anonymous:1878:CPD

[11270] Anonymous. Curtis’ propelling and dry dock attachment for vessels. Scientific American, 38(12):182, March 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n12/pdf/scientificamerican03231878-182a.pdf.

Anonymous:1878:EPD

[11271] Anonymous. Easel for pottery decoration. Scientific American, 38(12):183, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
183.pdf.

Anonymous:1878:RW

[11272] Anonymous. The red wolf. Scientific American, 38(12):183–184, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-184.pdf.

Anonymous:1878:NMD

[11273] Anonymous. A new mode of ditching. Scientific American, 38(12):185, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n12/pdf/scientificamerican03231878-185.pdf.

Anonymous:1878:NAM

[11274] Anonymous. New apparatus for measuring the vaporizing heat of liquids. Scientific American, 38(12):185, March 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n12/pdf/scientificamerican03231878-185a.pdf.

Anonymous:1878:SGC

[11275] Anonymous. The spontaneous generation controversy. Scientific American, 38(12):186, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
186.pdf.

Anonymous:1878:CRf

[11276] Anonymous. Communications received. Scientific American, 38 (12):188, March 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n12/pdf/scientificamerican03231878-
188.pdf.

Anonymous:1878:BPp

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-203.pdf.

Anonymous:1878:NQNe

[11278] Anonymous. Notes and queries: Notes & queries. Scientific American, 38(13):203–204, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
203a.pdf.

Anonymous:1878:IIm

[11279] Anonymous. Index of inventions. Scientific American, 38(13):204, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-204a.pdf.

Anonymous:1878:PNPl

[11280] Anonymous. Patent notes: Patents. Scientific American, 38(13):204, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-204b.pdf.

Anonymous:1878:PSE

[11281] Anonymous. The Peerless steam engine. Scientific American, 38(13):191, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
191.pdf.

Anonymous:1878:IBP

[11282] Anonymous. Improved bench plane. Scientific American, 38(13):191, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-191a.pdf.

Anonymous:1878:IPSa

[11283] Anonymous. Improvements in the postal service. Scientific American, 38(13):192, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
192.pdf.

Anonymous:1878:BWE

[11284] Anonymous. The best way to encourage invention. Scientific American, 38(13):192, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
192a.pdf.

Anonymous:1878:PPb

[11285] Anonymous. A popular project. Scientific American, 38(13):192–193, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-192b.pdf.

Anonymous:1878:Pa

[11286] Anonymous. The phonograph. Scientific American, 38(13):193, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-193.pdf.

Anonymous:1878:TWSb

[11287] Anonymous. Tree waste and its sequence. Scientific American, 38(13):193, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
193a.pdf.

Wright:1878:ANh

[11288] Berlin H. Wright. Astronomical notes. Scientific American, 38 (13):193–194, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
193b.pdf.

Anonymous:1878:LSM

[11289] Anonymous. A large sugar mill. Scientific American, 38(13):194, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-194.pdf.

Anonymous:1878:IJL

[11290] Anonymous. Independent jaw for lathe chucks. Scientific American, 38(13):194, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
194a.pdf.

Anonymous:1878:NOG

[11291] Anonymous. The new Otto gas engine. Scientific American, 38 (13):195–196, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
195.pdf.

Anonymous:1878:Cl

[11292] Anonymous. Communications. Scientific American, 38(13):196, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-196.pdf.

Anonymous:1878:PMa

[11293] Anonymous. Plant mind. Scientific American, 38(13):196–197, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-196a.pdf.

Anonymous:1878:WBE

[11294] Anonymous. Wild beast exterminators wanted. Scientific American, 38(13):197, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
197.pdf.

Anonymous:1878:PAPa

[11295] Anonymous. The progress of astronomical photography. Scientific American, 38(13):197, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
197a.pdf.

Anonymous:1878:NEC

[11296] Anonymous. The new Egyptian cotton. Scientific American, 38 (13):197–198, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
197b.pdf.

Anonymous:1878:ISMa

[11297] Anonymous. Improved saw mill dog. Scientific American, 38(13):198, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-198.pdf.

Anonymous:1878:IFR

[11298] Anonymous. Improved feed regulator for boilers. Scientific American, 38(13):198, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
198a.pdf.

Anonymous:1878:DIH

[11299] Anonymous. Dobbin’s improved harrow. Scientific American, 38(13):199, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
199.pdf.

Anonymous:1878:CN

[11300] Anonymous. Cleopatra’s needle. Scientific American, 38(13):199–201, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-199a.pdf.

Anonymous:1878:ITa

[11301] Anonymous. Improved tricycle. Scientific American, 38(13):201, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-201.pdf.

Anonymous:1878:SBa

[11302] Anonymous. Sea beans. Scientific American, 38(13):201, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n13/pdf/scientificamerican03301878-201a.pdf.

Anonymous:1878:NPOa

[11303] Anonymous. Notes of Patent Office decisions. Scientific American, 38(13):202, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
202.pdf.

Anonymous:1878:CRg

[11304] Anonymous. Communications received. Scientific American, 38 (13):204, March 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n13/pdf/scientificamerican03301878-
204.pdf.

Anonymous:1878:BPq

[11305] Anonymous. Business and personal. Scientific American, 38(14):219, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-219.pdf.

Anonymous:1878:BNBf

[11306] Anonymous. Books: New books and publications. Scientific American, 38(14):219, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
219a.pdf.

Anonymous:1878:NQi

[11307] Anonymous. Notes and queries. Scientific American, 38(14):219–220, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-219b.pdf.

Anonymous:1878:IIn

[11308] Anonymous. Index of inventions. Scientific American, 38(14):220, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-220a.pdf.

Anonymous:1878:PNPm

[11309] Anonymous. Patent notes: Patents. Scientific American, 38(14):220, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-220b.pdf.

Anonymous:1878:NYAc

[11310] Anonymous. The New York approach to the East River Bridge. Scientific American, 38(14):207–208, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
207.pdf.

Anonymous:1878:CPG

[11311] Anonymous. City parks as garden schools. Scientific American, 38(14):208, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
208.pdf.

Anonymous:1878:DIR

[11312] Anonymous. Disrespect of inventors’ rights. Scientific American, 38(14):208–209, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
208a.pdf.

Anonymous:1878:WM

[11313] Anonymous. A working man. Scientific American, 38(14):209, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-209.pdf.

Anonymous:1878:TSE

[11314] Anonymous. The telephone a sensitive electroscope. Scientific American, 38(14):209, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
209a.pdf.

Anonymous:1878:NPa

[11315] Anonymous. Nickel plating. Scientific American, 38(14):209, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-209b.pdf.

Anonymous:1878:WWH

[11316] Anonymous. The working Women’s hotel. Scientific American, 38(14):209, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
209c.pdf.

Anonymous:1878:ICCa

[11317] Anonymous. Improved centering chuck. Scientific American, 38 (14):210, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
210.pdf.

Anonymous:1878:CVC

[11318] Anonymous. Combined ventilator and chimney. Scientific American, 38(14):210–211, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
210a.pdf.

Anonymous:1878:BBM

[11319] Anonymous. Brownhill’s brick machine. Scientific American, 38 (14):211, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
211.pdf.

Anonymous:1878:IPSb

[11320] Anonymous. Improved paper shutter for Windows. Scientific American, 38(14):211–212, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
211a.pdf.

Anonymous:1878:Co

[11321] Anonymous. Communications. Scientific American, 38(14):212–213, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-212.pdf.

Anonymous:1878:SFa

[11322] Anonymous. The star feed. Scientific American, 38(14):213, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-213.pdf.

Anonymous:1878:ESE

[11323] Anonymous. The eclipse steam engine. Scientific American, 38(14):214, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-214.pdf.

Riley:1878:MB

[11324] C. V. Riley. Migratory butterflies. Scientific American, 38(14):215, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-215.pdf.

Anonymous:1878:BFP

[11325] Anonymous. A beton fountain, Prospect Park, Brooklyn. Scientific American, 38(14):215–216, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
215a.pdf.

Anonymous:1878:PMIa

[11326] Anonymous. Plant mind II. Scientific American, 38(14):216, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-216.pdf.

Anonymous:1878:FS

[11327] Anonymous. The first shad. Scientific American, 38(14):216, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-216a.pdf.

Anonymous:1878:NPOb

[11328] Anonymous. Notes of Patent Office decisions. Scientific American, 38(14):217, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
217.pdf.

Anonymous:1878:PTS

[11329] Anonymous. Platinum-tipped settings for diamonds. Scientific American, 38(14):217, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
217a.pdf.

Anonymous:1878:SSS

[11330] Anonymous. The Stafford scroll saw. Scientific American, 38 (14):217–218, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
217b.pdf.

Wright:1878:ANi

[11331] Berlin H. Wright. Astronomical notes. Scientific American, 38(14):218, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n14/pdf/scientificamerican04061878-218.pdf.

Anonymous:1878:CRh

[11332] Anonymous. Communications received. Scientific American, 38 (14):220, April 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n14/pdf/scientificamerican04061878-
220.pdf.

Anonymous:1878:BPr

[11333] Anonymous. Business and personal. Scientific American, 38(15):235, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-235.pdf.

Anonymous:1878:BNBg

[11334] Anonymous. Books: New books and publications. Scientific American, 38(15):235, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
235a.pdf.

Anonymous:1878:NQj

[11335] Anonymous. Notes and queries. Scientific American, 38(15):235–236, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-235b.pdf.

Anonymous:1878:IIo

[11336] Anonymous. Index of inventions. Scientific American, 38(15):236, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-236a.pdf.

Anonymous:1878:PNPn

[11337] Anonymous. Patent notes: Patents. Scientific American, 38(15):236, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-236b.pdf.

Anonymous:1878:IHMb

[11338] Anonymous. Improved hoisting machinery. Scientific American, 38(15):223, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
223.pdf.

Anonymous:1878:PEP

[11339] Anonymous. The proposed emasculation of the patent law. Scientific American, 38(15):224, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
224.pdf.

Anonymous:1878:KWW

[11340] Anonymous. “Keramics” and woman’s work. Scientific American, 38(15):224–225, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
224a.pdf.

Anonymous:1878:MD

[11341] Anonymous. The manufacture of daubs. Scientific American, 38 (15):225, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
225.pdf.

Anonymous:1878:SAP

[11342] Anonymous. The scientific applications of photography. Scientific American, 38(15):225, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
225a.pdf.

Anonymous:1878:ILC

[11343] Anonymous. Improved lathe chucks. Scientific American, 38(15):226, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-226.pdf.

Anonymous:1878:ECT

[11344] Anonymous. English cab trucks and axles. Scientific American, 38(15):226, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
226a.pdf.

Anonymous:1878:GGA

[11345] Anonymous. The Gatling gun aloft. Scientific American, 38(15):227, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-227.pdf.

Anonymous:1878:Cp

[11346] Anonymous. Communications. Scientific American, 38(15):228–229, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-228.pdf.

Anonymous:1878:IBT

[11347] Anonymous. Improved bale tie. Scientific American, 38(15):229, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-229.pdf.

Wright:1878:ANj

[11348] Berlin H. Wright. Astronomical notes. Scientific American, 38(15):230, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-230.pdf.

Anonymous:1878:ISMb

[11349] Anonymous. Improved scrubbing machine. Scientific American, 38(15):230, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
230a.pdf.

Anonymous:1878:IL

[11350] Anonymous. Improved lubricator. Scientific American, 38(15):230, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-230b.pdf.

Hopkins:1878:SE

[11351] M. Hopkins. Spider engineering. Scientific American, 38(15):231, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-231.pdf.

Anonymous:1878:APPa

[11352] Anonymous. The Algerian Palace at the Paris Exposition. Scientific American, 38(15):231–232, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
231a.pdf.

Anonymous:1878:SPM

[11353] Anonymous. Steam power moulding machinery. Scientific American, 38(15):232, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
232.pdf.

Anonymous:1878:PMIb

[11354] Anonymous. Plant mind III. Scientific American, 38(15):233–234, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n15/pdf/scientificamerican04131878-233.pdf.

Anonymous:1878:SSH

[11355] Anonymous. Some seasonable hints on plant cultivation. Scientific American, 38(15):234, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
234.pdf.

Anonymous:1878:CRi

[11356] Anonymous. Communications received. Scientific American, 38 (15):236, April 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n15/pdf/scientificamerican04131878-
236.pdf.

Anonymous:1878:BPt

[11357] Anonymous. Business and personal. Scientific American, 38(16):250, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-250.pdf.

Anonymous:1878:NQk

[11358] Anonymous. Notes and queries. Scientific American, 38(16):250–252, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-250a.pdf.

Anonymous:1878:Ia

[11359] Anonymous. Index. Scientific American, 38(16):252, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-252a.pdf.

Anonymous:1878:PNPo

[11360] Anonymous. Patent notes: Patents. Scientific American, 38(16):252, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-252b.pdf.

Anonymous:1878:RDC

[11361] Anonymous. Report of the Director of the Central Park Museum, 1878. Scientific American, 38(16):239, April 20, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n16/pdf/scientificamerican04201878-239.pdf.

Anonymous:1878:SRC

[11362] Anonymous. A sailing railway car. Scientific American, 38(16):239–240, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-239a.pdf.

Anonymous:1878:EMH

[11363] Anonymous. Every man his own Keely motor. Scientific American, 38(16):240–241, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
240.pdf.

Anonymous:1878:HFC

[11364] Anonymous. How fast can you think. Scientific American, 38(16):241, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-241.pdf.

Anonymous:1878:ITSa

[11365] Anonymous. The intensity of telephonic sounds. Scientific American, 38(16):241, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
241a.pdf.

Anonymous:1878:PPO

[11366] Anonymous. Progress at the Patent Office. Scientific American, 38(16):241, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
241b.pdf.

Anonymous:1878:CBa

[11367] Anonymous. Color blindness. Scientific American, 38(16):241, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-241c.pdf.

Anonymous:1878:SMb

[11368] Anonymous. Slotting machine. Scientific American, 38(16):242, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-242.pdf.

Anonymous:1878:SWB

[11369] Anonymous. Sea water as a beverage. Scientific American, 38(16):242, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-242a.pdf.

Anonymous:1878:ICGa

[11370] Anonymous. Improved circulating generator for steam boilers. Scientific American, 38(16):243, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
243.pdf.

Anonymous:1878:ISG

[11371] Anonymous. Improved sewer gas check valve for stationary wash basins, etc. Scientific American, 38(16):243–244, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n16/pdf/scientificamerican04201878-243a.pdf.

Anonymous:1878:INL

[11372] Anonymous. Improved nut lock. Scientific American, 38(16):244, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-244.pdf.

Anonymous:1878:Cq

[11373] Anonymous. Communications. Scientific American, 38(16):244, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-244a.pdf.

Anonymous:1878:SNW

[11374] Anonymous. Some new workshop appliances. Scientific American, 38(16):245, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
245.pdf.

Anonymous:1878:IDM

[11375] Anonymous. Improved dovetailing machine. Scientific American, 38(16):246, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
246.pdf.

Anonymous:1878:CPE

[11376] Anonymous. Centrifugal pumping engine. Scientific American, 38(16):246, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
246a.pdf.

Anonymous:1878:NIG

[11377] Anonymous. New investigations on gallium. Scientific American, 38(16):246, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
246b.pdf.

Anonymous:1878:VRa

[11378] Anonymous. The Victoria Regia. Scientific American, 38(16):247, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-247.pdf.

Anonymous:1878:RNR

[11379] Anonymous. Redier’s new registering thermometer. Scientific American, 38(16):248–249, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
248.pdf.

Anonymous:1878:PLM

[11380] Anonymous. The Philadelphia lawn mowers. Scientific American, 38(16):249, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
249.pdf.

Riley:1878:HB

[11381] C. V. Riley. The Horn bug. Scientific American, 38(16):249, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-249a.pdf.

Wright:1878:ANk

[11382] Berlin H. Wright. Astronomical notes. Scientific American, 38(16):249, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-249b.pdf.

Anonymous:1878:CRj

[11383] Anonymous. Communications received. Scientific American, 38 (16):252, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n16/pdf/scientificamerican04201878-
252.pdf.

Anonymous:1878:SAd

[11384] Anonymous. The Scientific American. Scientific American, 38(16):254, April 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n16/pdf/scientificamerican04201878-254.pdf.

Anonymous:1878:BPu

[11385] Anonymous. Business and personal. Scientific American, 38(17):267, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-267.pdf.

Anonymous:1878:BNBh

[11386] Anonymous. Books: New books and publications. Scientific American, 38(17):267, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
267a.pdf.

Anonymous:1878:NQl

[11387] Anonymous. Notes and queries. Scientific American, 38(17):267–268, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-267b.pdf.

Anonymous:1878:Ib

[11388] Anonymous. Index. Scientific American, 38(17):268, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-268a.pdf.

Anonymous:1878:PNPp

[11389] Anonymous. Patent notes: Patents. Scientific American, 38(17):268, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-268b.pdf.

Anonymous:1878:MEW

[11390] Anonymous. The manufacture of emery wheels. Scientific American, 38(17):255, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
255.pdf.

Anonymous:1878:SCP

[11391] Anonymous. Science and commerce as peacemakers. Scientific American, 38(17):256, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
256.pdf.

Anonymous:1878:UWM

[11392] Anonymous. The utilization of waste matters. Scientific American, 38(17):256–257, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
256a.pdf.

Anonymous:1878:CCc

[11393] Anonymous. “Convict competition”. Scientific American, 38(17):257, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-257.pdf.

Anonymous:1878:SLb

[11394] Anonymous. Sterilization by light. Scientific American, 38(17):257, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-257a.pdf.

Anonymous:1878:DIa

[11395] Anonymous. A dangerous item. Scientific American, 38(17):257, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-257b.pdf.

Anonymous:1878:CHP

[11396] Anonymous. The culture of house plants. Scientific American, 38(17):258, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
258.pdf.

Hopkins:1878:WLM

[11397] Geo. M. Hopkins. A water lens microscope. Scientific American, 38(17):258, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
258a.pdf.

Anonymous:1878:CME

[11398] Anonymous. Compound marine engines. Scientific American, 38 (17):259–260, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
259.pdf.

Anonymous:1878:Cs

[11399] Anonymous. Communications. Scientific American, 38(17):260, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-260.pdf.

Anonymous:1878:PMb

[11400] Anonymous. Plant mind. Scientific American, 38(17):260, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-260a.pdf.

Anonymous:1878:SBI

[11401] Anonymous. Steam boiler inspection. Scientific American, 38 (17):261–262, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
261.pdf.

Anonymous:1878:IVE

[11402] Anonymous. Improved variable exhaust. Scientific American, 38 (17):262, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
262.pdf.

Anonymous:1878:CDFb

[11403] Anonymous. Combined dinner fail and lantern. Scientific American, 38(17):262, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
262a.pdf.

Anonymous:1878:TTB

[11404] Anonymous. The three-tailed birthwort. Scientific American, 38 (17):263, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
263.pdf.

Anonymous:1878:BT

[11405] Anonymous. The baobab tree. Scientific American, 38(17):263, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-263a.pdf.

Wright:1878:ANl

[11406] Berlin H. Wright. Astronomical notes. Scientific American, 38(17):264, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-264.pdf.

Anonymous:1878:CIP

[11407] Anonymous. Coleman’s improved pipe wrench. Scientific American, 38(17):264–265, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
264a.pdf.

Anonymous:1878:NV

[11408] Anonymous. A new vehicle. Scientific American, 38(17):265–266, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-265.pdf.

Anonymous:1878:FBS

[11409] Anonymous. French band saw blades. Scientific American, 38(17):266, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-266.pdf.

Anonymous:1878:CRk

[11410] Anonymous. Communications received. Scientific American, 38 (17):268, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n17/pdf/scientificamerican04271878-
268.pdf.

Anonymous:1878:SAe

[11411] Anonymous. The Scientific American. Scientific American, 38(17):270, April 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n17/pdf/scientificamerican04271878-270.pdf.

Anonymous:1878:BPv

[11412] Anonymous. Business and personal. Scientific American, 38(18):283, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-283.pdf.

Anonymous:1878:BNBi

[11413] Anonymous. Books: New books and publications. Scientific American, 38(18):283, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
283a.pdf.

Anonymous:1878:NQm

[11414] Anonymous. Notes and queries. Scientific American, 38(18):283, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-283b.pdf.

Anonymous:1878:Ic

[11415] Anonymous. Index. Scientific American, 38(18):284, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-284.pdf.

Anonymous:1878:PNPq

[11416] Anonymous. Patent notes: Patents. Scientific American, 38(18):286, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-286.pdf.

Anonymous:1878:NMI

[11417] Anonymous. New mechanical inventions. Scientific American, 38 (18):271, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
271.pdf.

Anonymous:1878:IMTa

[11418] Anonymous. Improved machine tools. Scientific American, 38(18):271, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-271a.pdf.

Anonymous:1878:HPA

[11419] Anonymous. How patents affect prices. Scientific American, 38(18):272, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-272.pdf.

Anonymous:1878:WL

[11420] Anonymous. What is life? Scientific American, 38(18):272–273, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-272a.pdf.

Anonymous:1878:DIb

[11421] Anonymous. The decline of the ironclad. Scientific American, 38(18):273, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
273.pdf.

Anonymous:1878:CPO

[11422] Anonymous. Congress and the Patent Office funds. Scientific American, 38(18):273–274, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
273a.pdf.

Anonymous:1878:IMTb

[11423] Anonymous. Improved machine tools. Scientific American, 38(18):274, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-274.pdf.

Anonymous:1878:HMC

[11424] Anonymous. Hydraulic mining in California. Scientific American, 38(18):274–275, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
274a.pdf.

Anonymous:1878:ITSb

[11425] Anonymous. Improved tool stocks for dental engines. Scientific American, 38(18):275, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
275.pdf.

Anonymous:1878:Ct

[11426] Anonymous. Communications. Scientific American, 38(18):276, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-276.pdf.

Wright:1878:ANm

[11427] Berlin H. Wright. Astronomical notes. Scientific American, 38(18):277, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-277.pdf.

Anonymous:1878:NYAd

[11428] Anonymous. New York Academy of Sciences. Scientific American, 38(18):277–278, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
277a.pdf.

Anonymous:1878:LRM

[11429] Anonymous. Lining roofs with mineral wool. Scientific American, 38(18):278, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
278.pdf.

Anonymous:1878:IIP

[11430] Anonymous. Improved insect powder blower. Scientific American, 38(18):278, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
278a.pdf.

Anonymous:1878:BLD

[11431] Anonymous. The boiling lake of Dominica. Scientific American, 38(18):279, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
279.pdf.

Anonymous:1878:ICB

[11432] Anonymous. Improved clam Baker. Scientific American, 38(18): 279–280, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
279a.pdf.

Anonymous:1878:PPc

[11433] Anonymous. The Pyrenean pine. Scientific American, 38(18):280, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-280.pdf.

Anonymous:1878:OBS

[11434] Anonymous. The oil bearing sands of Pennsylvania. Scientific American, 38(18):280–281, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
280a.pdf.

Anonymous:1878:IPT

[11435] Anonymous. Improved pipe tongs. Scientific American, 38(18):281, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-281.pdf.

Anonymous:1878:IRS

[11436] Anonymous. Improved receiver and stench trap. Scientific American, 38(18):281, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
281a.pdf.

Anonymous:1878:ICM

[11437] Anonymous. Influence of cosmical matter. Scientific American, 38(18):281, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
281b.pdf.

Anonymous:1878:PII

[11438] Anonymous. Proposed injustice to inventors. Scientific American, 38(18):282, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
282.pdf.

Anonymous:1878:HA

[11439] Anonymous. Hydrobromic acid. Scientific American, 38(18):282, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-282a.pdf.

Anonymous:1878:CRl

[11440] Anonymous. Communications received. Scientific American, 38 (18):283, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n18/pdf/scientificamerican05041878-
283c.pdf.

Anonymous:1878:SAf

[11441] Anonymous. The Scientific American. Scientific American, 38(18):284, May 4, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n18/pdf/scientificamerican05041878-284a.pdf.

Anonymous:1878:BPw

[11442] Anonymous. Business and personal. Scientific American, 38(19):299, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-299.pdf.

Anonymous:1878:BNBj

[11443] Anonymous. Books: New books and publications. Scientific American, 38(19):299, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
299a.pdf.

Anonymous:1878:NQo

[11444] Anonymous. Notes and queries. Scientific American, 38(19):299–300, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-299b.pdf.

Anonymous:1878:PNPr

[11445] Anonymous. Patent notes: Patents. Scientific American, 38(19):302, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-302.pdf.

Anonymous:1878:AW

[11446] Anonymous. Aerial warfare. Scientific American, 38(19):287–288, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-287.pdf.

Anonymous:1878:AS

[11447] Anonymous. Arguments for section 11. Scientific American, 38 (19):288–299, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
288.pdf.

Anonymous:1878:NSPb

[11448] Anonymous. A national sanitary precaution. Scientific American, 38(19):289, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
289.pdf.

Anonymous:1878:TS

[11449] Anonymous. Teaching science. Scientific American, 38(19):289, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-289a.pdf.

Anonymous:1878:RME

[11450] Anonymous. A ready means of estimating the valuable constituents of cereals, etc. Scientific American, 38(19):289, May 11, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n19/pdf/scientificamerican05111878-289b.pdf.

Anonymous:1878:EI

[11451] Anonymous. The eclipse injector. Scientific American, 38(19):290, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-290.pdf.

Anonymous:1878:IWB

[11452] Anonymous. Improved wagon box adjuster and fastenings. Scientific American, 38(19):290–291, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
290a.pdf.

Anonymous:1878:ISMc

[11453] Anonymous. Improved saw mill head block. Scientific American, 38(19):291, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
291.pdf.

Anonymous:1878:GWB

[11454] Anonymous. Gilbert’s wash basin valve. Scientific American, 38 (19):291–292, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
291a.pdf.

Anonymous:1878:Cu

[11455] Anonymous. Communications. Scientific American, 38(19):292–293, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-292.pdf.

Wright:1878:ANn

[11456] Berlin H. Wright. Astronomical notes. Scientific American, 38(19):293, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-293.pdf.

Anonymous:1878:OFS

[11457] Anonymous. Ostrich farming in South Africa. Scientific American, 38(19):294–296, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
294.pdf.

Anonymous:1878:PTa

[11458] Anonymous. The premier tricycle. Scientific American, 38(19):296, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-296.pdf.

Anonymous:1878:PMc

[11459] Anonymous. Plant mind. Scientific American, 38(19):296, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-296a.pdf.

Anonymous:1878:IWP

[11460] Anonymous. Improved wood planer. Scientific American, 38(19):297, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-297.pdf.

Anonymous:1878:ILJ

[11461] Anonymous. Improved lifting jack. Scientific American, 38(19):297, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-297a.pdf.

Anonymous:1878:AASa

[11462] Anonymous. American Academy of Sciences. Scientific American, 38(19):297–298, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
297b.pdf.

Anonymous:1878:CBI

[11463] Anonymous. The cause of brick incrustations. Scientific American, 38(19):298, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
298.pdf.

Anonymous:1878:CRn

[11464] Anonymous. Communications received. Scientific American, 38 (19):300, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
300.pdf.

Anonymous:1878:LSPb

[11465] Anonymous. Lists and specifications of patents. Scientific American, 38(19):300, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n19/pdf/scientificamerican05111878-
300a.pdf.

Anonymous:1878:SAg

[11466] Anonymous. The Scientific American. Scientific American, 38(19):300, May 11, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n19/pdf/scientificamerican05111878-300b.pdf.

Anonymous:1878:BPx

[11467] Anonymous. Business and personal. Scientific American, 38(20):316, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-316.pdf.

Anonymous:1878:BNBk

[11468] Anonymous. Books: New books and publications. Scientific American, 38(20):316, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
316a.pdf.

Anonymous:1878:NQp

[11469] Anonymous. Notes and queries. Scientific American, 38(20):316, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-316b.pdf.

Anonymous:1878:PNPs

[11470] Anonymous. Patent notes: Patents. Scientific American, 38(20):318, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-318.pdf.

Anonymous:1878:CSEb

[11471] Anonymous. Combining the strands of the East River Bridge cables. Scientific American, 38(20):303, May 18, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n20/pdf/scientificamerican05181878-303.pdf.

Anonymous:1878:RIR

[11472] Anonymous. A raid on inventors’ rights. Scientific American, 38 (20):304, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
304.pdf.

Anonymous:1878:MBa

[11473] Anonymous. The mechanics of the bicycle. Scientific American, 38(20):304–305, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
304a.pdf.

Anonymous:1878:CEP

[11474] Anonymous. Captain Eads’ proposed improvements of the Mississippi and St. John’s Rivers. Scientific American, 38(20):305, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-305.pdf.

Anonymous:1878:CRLa

[11475] Anonymous. The confectioner regarded in the light of a colorman. Scientific American, 38(20):305, May 18, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n20/pdf/scientificamerican05181878-305a.pdf.

Anonymous:1878:CMP

[11476] Anonymous. Congress to be made a patent mill. Scientific American, 38(20):305, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
305b.pdf.

Anonymous:1878:AMM

[11477] Anonymous. Advertising — a mode of motion. Scientific American, 38(20):306, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
306.pdf.

Anonymous:1878:RSCb

[11478] Anonymous. A rudder and screw combined. Scientific American, 38(20):306, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
306a.pdf.

Anonymous:1878:MCB

[11479] Anonymous. M. Claude Bernard. Scientific American, 38(20):307, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-307.pdf.

Anonymous:1878:CTP

[11480] Anonymous. The Canadian trophy at the Paris Exhibition. Scientific American, 38(20):307–308, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
307a.pdf.

Anonymous:1878:DSV

[11481] Anonymous. Detachable spoke vehicle wheel. Scientific American, 38(20):308–309, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
308.pdf.

Anonymous:1878:SAS

[11482] Anonymous. Shute’s adjustable safety stilt. Scientific American, 38(20):309, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
309.pdf.

Anonymous:1878:TMN

[11483] Anonymous. Too many names. Scientific American, 38(20):309, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-309a.pdf.

Anonymous:1878:ICCb

[11484] Anonymous. Improved calender clock. Scientific American, 38(20):309, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-309b.pdf.

Anonymous:1878:BAS

[11485] Anonymous. The buckeye automatic steam engine. Scientific American, 38(20):310, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
310.pdf.

Anonymous:1878:BIE

[11486] Anonymous. The British Indian exhibit. Scientific American, 38 (20):311–312, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
311.pdf.

Anonymous:1878:PAMa

[11487] Anonymous. The production of artificial monsters. Scientific American, 38(20):312, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
312.pdf.

Anonymous:1878:MPNb

[11488] Anonymous. M. Plante’s new investigations on the effects of high tension currents. Scientific American, 38(20):313–315, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n20/pdf/scientificamerican05181878-313.pdf.

Anonymous:1878:CRo

[11489] Anonymous. Communications received. Scientific American, 38 (20):316, May 18, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n20/pdf/scientificamerican05181878-
316c.pdf.

Anonymous:1878:BPy

[11490] Anonymous. Business and personal. Scientific American, 38(21):331, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-331.pdf.

Anonymous:1878:BNBl

[11491] Anonymous. Books: New books and publications. Scientific American, 38(21):331, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
331a.pdf.

Anonymous:1878:NQq

[11492] Anonymous. Notes and queries. Scientific American, 38(21):331, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-331b.pdf.

Anonymous:1878:IIq

[11493] Anonymous. Index of inventions. Scientific American, 38(21):332, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-332.pdf.

Anonymous:1878:PNPt

[11494] Anonymous. Patent notes: Patents. Scientific American, 38(21):334, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-334.pdf.

Anonymous:1878:IFW

[11495] Anonymous. Improved feed water heater and purifier. Scientific American, 38(21):319, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
319.pdf.

Anonymous:1878:BSD

[11496] Anonymous. The black sile deception. Scientific American, 38(21):320, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-320.pdf.

Anonymous:1878:NOSa

[11497] Anonymous. A new opportunity to secure foreign trade. — Scientific American export edition. Scientific American, 38(21):321, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-321.pdf.

Anonymous:1878:SSM

[11498] Anonymous. Spirit strings and morbid minds. Scientific American, 38(21):321, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
321a.pdf.

Anonymous:1878:TMI

[11499] Anonymous. The transit of Mercury and the inter-Mercurial body. Scientific American, 38(21):322, May 25, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n21/pdf/scientificamerican05251878-322.pdf.

Anonymous:1878:ONN

[11500] Anonymous. Our Naval necessities. Scientific American, 38(21):322, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-322a.pdf.

Anonymous:1878:AMS

[11501] Anonymous. The accident to the machinery of the steamer Old Colony. Scientific American, 38(21):322–323, May 25, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n21/pdf/scientificamerican05251878-322b.pdf.

Anonymous:1878:WCA

[11502] Anonymous. What are the causes that affect the taste of drinking waters? Scientific American, 38(21):323–324, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n21/pdf/scientificamerican05251878-323.pdf.

Anonymous:1878:Cv

[11503] Anonymous. Communications. Scientific American, 38(21):324, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-324.pdf.

Anonymous:1878:OSV

[11504] Anonymous. Our simian visitors. Scientific American, 38(21):325, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-325.pdf.

Anonymous:1878:Id

[11505] Anonymous. The injector. Scientific American, 38(21):325, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-325a.pdf.

Anonymous:1878:WRO

[11506] Anonymous. Winterburn’s regulator for oil stove wicks. Scientific American, 38(21):326, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
326.pdf.

Anonymous:1878:IDP

[11507] Anonymous. Improved differential pressure regulator. Scientific American, 38(21):326, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
326a.pdf.

Anonymous:1878:IFG

[11508] Anonymous. Improves farm gate. Scientific American, 38(21):326, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-326b.pdf.

Seiss:1878:FS

[11509] C. Few Seiss. Fangs of serpents. Scientific American, 38(21):327, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-327.pdf.

Anonymous:1878:W

[11510] Anonymous. Waterspouts. Scientific American, 38(21):327–328, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-327a.pdf.

Anonymous:1878:ISBa

[11511] Anonymous. Improved Spring bed. Scientific American, 38(21):328, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-328.pdf.

Anonymous:1878:IGM

[11512] Anonymous. Improved grinding mills. Scientific American, 38(21):328, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-328a.pdf.

Anonymous:1878:MVM

[11513] Anonymous. Machinery vs. manual labor. Scientific American, 38(21):329, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
329.pdf.

Anonymous:1878:MVR

[11514] Anonymous. M. Victor Regnault. Scientific American, 38(21):329, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-329a.pdf.

Anonymous:1878:PBT

[11515] Anonymous. Proposed bridge over the Thames. Scientific American, 38(21):329, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
329b.pdf.

Anonymous:1878:BPRa

[11516] Anonymous. Bastard patent rights. Scientific American, 38(21): 329–330, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
329c.pdf.

Wright:1878:ANo

[11517] Berlin H. Wright. Astronomical notes. Scientific American, 38(21):330, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n21/pdf/scientificamerican05251878-330.pdf.

Anonymous:1878:CRp

[11518] Anonymous. Communications received. Scientific American, 38 (21):331–332, May 25, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n21/pdf/scientificamerican05251878-
331c.pdf.

Anonymous:1878:BPz

[11519] Anonymous. Business and personal. Scientific American, 38(22):348, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-348.pdf.

Anonymous:1878:NQr

[11520] Anonymous. Notes and queries. Scientific American, 38(22):348, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-348a.pdf.

Anonymous:1878:PNPu

[11521] Anonymous. Patent notes: Patents. Scientific American, 38(22):350, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-350.pdf.

Anonymous:1878:REb

[11522] Anonymous. The Roberts engine. Scientific American, 38(22):242–243, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-342a.pdf.

Hopkins:1878:EG

[11523] Geo. M. Hopkins. An electrical gyroscope. Scientific American, 38(22):335, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
335.pdf.

Anonymous:1878:DII

[11524] Anonymous. Do inventions injure the laborer? Scientific American, 38(22):336, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
336.pdf.

Anonymous:1878:PJH

[11525] Anonymous. Professor Joseph Henry. Scientific American, 38 (22):336–337, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
336a.pdf.

Anonymous:1878:MAW

[11526] Anonymous. Meddlers in arts in which they are not skilled. Scientific American, 38(22):337, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
337.pdf.

Anonymous:1878:CCd

[11527] Anonymous. Clearance and compression. Scientific American, 38 (22):337, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
337a.pdf.

Wright:1878:ANp

[11528] Berlin H. Wright. Astronomical notes. Scientific American, 38(22):338, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-338.pdf.

Anonymous:1878:RTB

[11529] Anonymous. Russian torpedo boats. Scientific American, 38(22):338, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-338a.pdf.

Anonymous:1878:EDA

[11530] Anonymous. The Empire dustless ash sifter. Scientific American, 38(22):338–339, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
338b.pdf.

Anonymous:1878:ADAb

[11531] Anonymous. Andrews’ draught apparatus for stoves. Scientific American, 38(22):339, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
339.pdf.

Anonymous:1878:IMN

[11532] Anonymous. Improved mosquito net frame. Scientific American, 38(22):339, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
339a.pdf.

Anonymous:1878:Cw

[11533] Anonymous. Communications. Scientific American, 38(22):340, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-340.pdf.

Duerden:1878:NT

[11534] L. L. Duerden. Notes on the telephone. Scientific American, 38(22):341, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-341.pdf.

Anonymous:1878:IVH

[11535] Anonymous. Improved vertical and horizontal vise. Scientific American, 38(22):342, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
342.pdf.

Anonymous:1878:NCBa

[11536] Anonymous. New collapsing boat. Scientific American, 38(22):343, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-343.pdf.

Anonymous:1878:AB

[11537] Anonymous. The ant bear. Scientific American, 38(22):343, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-343a.pdf.

Anonymous:1878:CMM

[11538] Anonymous. Cigarette making by machinery. Scientific American, 38(22):344, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
344.pdf.

Anonymous:1878:REM

[11539] Anonymous. Recent eruption of Mount Hecla. Scientific American, 38(22):345, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
345.pdf.

Anonymous:1878:ASF

[11540] Anonymous. American steam fire engines. Scientific American, 38(22):345, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
345a.pdf.

Anonymous:1878:SRF

[11541] Anonymous. Some recent facts about common poisons. Scientific American, 38(22):346, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n22/pdf/scientificamerican06011878-
346.pdf.

Anonymous:1878:PMd

[11542] Anonymous. Perpetual motion. Scientific American, 38(22):346–347, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-346a.pdf.

Wood:1878:OTM

[11543] J. Walter Wood and Alfred M. Mayer. Observations on the Transit of Mercury of May 6, 1878. Scientific American, 38(22):347, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-347.pdf.

Anonymous:1878:CRq

[11544] Anonymous. Communication received. Scientific American, 38(22):348, June 1, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n22/pdf/scientificamerican06011878-348b.pdf.

Anonymous:1878:BPba

[11545] Anonymous. Business and personal. Scientific American, 38(23):363, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-363.pdf.

Anonymous:1878:NQs

[11546] Anonymous. Notes and queries. Scientific American, 38(23):363, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-363a.pdf.

Anonymous:1878:IIr

[11547] Anonymous. Index of inventions. Scientific American, 38(23):363–364, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-363c.pdf.

Anonymous:1878:PNPv

[11548] Anonymous. Patent notes: Patents. Scientific American, 38(23):366, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-366.pdf.

Anonymous:1878:BMM

[11549] Anonymous. Brush making by machinery. Scientific American, 38(23):351–352, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
351.pdf.

Anonymous:1878:APA

[11550] Anonymous. American productions abroad. Scientific American, 38(23):352–353, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
352.pdf.

Anonymous:1878:CTa

[11551] Anonymous. The carbon telephone. Scientific American, 38(23): 353–354, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
353.pdf.

Anonymous:1878:BF

[11552] Anonymous. The Buffalo forge. Scientific American, 38(23):354, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-354.pdf.

Anonymous:1878:ISE

[11553] Anonymous. Improved signal egg boiler. Scientific American, 38 (23):354, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
354a.pdf.

Anonymous:1878:BM

[11554] Anonymous. The Brownian movement. Scientific American, 38(23):354, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-354b.pdf.

Anonymous:1878:RTG

[11555] Anonymous. A reliable test gold and silver coin. Scientific American, 38(23):355, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
355.pdf.

Anonymous:1878:IWJ

[11556] Anonymous. Improved wagon jack. Scientific American, 38(23):355, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-355a.pdf.

Anonymous:1878:IS

[11557] Anonymous. Improved sled. Scientific American, 38(23):355, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-355b.pdf.

Anonymous:1878:IAL

[11558] Anonymous. Improved Argand lamp burner. Scientific American, 38(23):355–356, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
355c.pdf.

Anonymous:1878:GM

[11559] Anonymous. Gas motors. Scientific American, 38(23):356, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-356.pdf.

Anonymous:1878:BC

[11560] Anonymous. Boiler corrosion. Scientific American, 38(23):356–357, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-356a.pdf.

Anonymous:1878:STA

[11561] Anonymous. Swiss testimony to the advantages of our patent system. Scientific American, 38(23):357, June 8, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v38/
n23/pdf/scientificamerican06081878-357.pdf.

Anonymous:1878:RPM

[11562] Anonymous. A remarkable picture of the Moon. Scientific American, 38(23):357, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
357a.pdf.

Anonymous:1878:IRF

[11563] Anonymous. Improved rotary force pump. Scientific American, 38(23):358, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
358.pdf.

Anonymous:1878:IGF

[11564] Anonymous. Improved geared friction shaper. Scientific American, 38(23):358, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
358a.pdf.

Anonymous:1878:PEB

[11565] Anonymous. The Paris Exhibition building at the Trocadero. Scientific American, 38(23):358, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
358b.pdf.

Anonymous:1878:ADb

[11566] Anonymous. African diamonds. Scientific American, 38(23):360, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-360.pdf.

Anonymous:1878:PET

[11567] Anonymous. Physiological effects of thirst. Scientific American, 38(23):360, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
360a.pdf.

Wright:1878:ANq

[11568] Berlin H. Wright. Astronomical notes. Scientific American, 38(23):361, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-361.pdf.

Anonymous:1878:PTB

[11569] Anonymous. Proposed Thames bridge. Scientific American, 38(23):361, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-361a.pdf.

Anonymous:1878:MBb

[11570] Anonymous. M. Becquerel. Scientific American, 38(23):361–362, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-361b.pdf.

Anonymous:1878:NPOc

[11571] Anonymous. Notes of Patent Office decisions. Scientific American, 38(23):362, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
362.pdf.

Anonymous:1878:IMCc

[11572] Anonymous. Is it a mere coincidence? Scientific American, 38(23):362, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n23/pdf/scientificamerican06081878-362a.pdf.

Anonymous:1878:CRr

[11573] Anonymous. Communications received. Scientific American, 38 (23):363, June 8, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n23/pdf/scientificamerican06081878-
363b.pdf.

Anonymous:1878:BPbb

[11574] Anonymous. Business and personal. Scientific American, 38(24):379, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-379.pdf.

Anonymous:1878:BNBm

[11575] Anonymous. Books: New books and publications. Scientific American, 38(24):379, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
379a.pdf.

Anonymous:1878:NQt

[11576] Anonymous. Notes and queries. Scientific American, 38(24):379, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-379b.pdf.

Anonymous:1878:IIs

[11577] Anonymous. Index of inventions. Scientific American, 38(24):379–380, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-379d.pdf.

Anonymous:1878:PNPw

[11578] Anonymous. Patent notes: Patents. Scientific American, 38(24):382, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-382.pdf.

Anonymous:1878:HLS

[11579] Anonymous. The high-level street railways of New York. Scientific American, 38(24):367, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
367.pdf.

Anonymous:1878:SMM

[11580] Anonymous. Silver mining in Massachusetts. Scientific American, 38(24):368, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
368.pdf.

Anonymous:1878:CCe

[11581] Anonymous. Counterfeit coin. Scientific American, 38(24):368–369, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-368a.pdf.

Anonymous:1878:AF

[11582] Anonymous. Artificial fuel. Scientific American, 38(24):369, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-369.pdf.

Anonymous:1878:SAh

[11583] Anonymous. Silver in art. Scientific American, 38(24):369, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-369a.pdf.

Anonymous:1878:FPR

[11584] Anonymous. Formation of planetary rings and satellites. Scientific American, 38(24):369–370, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
369b.pdf.

Anonymous:1878:VEB

[11585] Anonymous. Vertical engines and boilers. Scientific American, 38(24):371, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
371.pdf.

Anonymous:1878:SSa

[11586] Anonymous. School shops. Scientific American, 38(24):371, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-371a.pdf.

Anonymous:1878:IAK

[11587] Anonymous. Improved automatic knife grinder. Scientific American, 38(24):371–372, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
371b.pdf.

Wright:1878:ANr

[11588] Berlin H. Wright. Astronomical notes. Scientific American, 38(24):372, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-372.pdf.

Anonymous:1878:NPOd

[11589] Anonymous. Notes of Patent Office decisions. Scientific American, 38(24):372, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
372a.pdf.

Rose:1878:MCL

[11590] Joshua Rose. Method of correcting a leasing screw. Scientific American, 38(24):373, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
373.pdf.

Anonymous:1878:FM

[11591] Anonymous. Firework making. Scientific American, 38(24):373–374, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-373a.pdf.

Anonymous:1878:GBC

[11592] Anonymous. The glass ball Caster. Scientific American, 38(24):374, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-374.pdf.

Anonymous:1878:IWa

[11593] Anonymous. The industrial west. Scientific American, 38(24):374, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-374a.pdf.

Anonymous:1878:PDU

[11594] Anonymous. The prairie dog and its uninvited quests. Scientific American, 38(24):375, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
375.pdf.

Anonymous:1878:RGA

[11595] Anonymous. Remains of gigantic antediluvians. Scientific American, 38(24):375–376, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
375a.pdf.

Anonymous:1878:CSHb

[11596] Anonymous. Cooking by solar heat. Scientific American, 38(24):376, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-376.pdf.

Anonymous:1878:PAS

[11597] Anonymous. Producing accurate spheres. Scientific American, 38 (24):376, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
376a.pdf.

Anonymous:1878:OVA

[11598] Anonymous. Organography, or vegetable anatomy. Scientific American, 38(24):377, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
377.pdf.

Anonymous:1878:MF

[11599] Anonymous. Mexican flora. Scientific American, 38(24):377–378, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n24/pdf/scientificamerican06151878-377a.pdf.

Anonymous:1878:CRs

[11600] Anonymous. Communications received. Scientific American, 38 (24):379, June 15, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n24/pdf/scientificamerican06151878-
379c.pdf.

Anonymous:1878:BPbc

[11601] Anonymous. Business and personal. Scientific American, 38(25):396, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-396.pdf.

Anonymous:1878:BNBn

[11602] Anonymous. Books: New books and publications. Scientific American, 38(25):396, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
396a.pdf.

Anonymous:1878:NQu

[11603] Anonymous. Notes and queries. Scientific American, 38(25):396, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-396b.pdf.

Anonymous:1878:PNPx

[11604] Anonymous. Patent notes: Patents. Scientific American, 38(25):398, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-398.pdf.

Anonymous:1878:PEc

[11605] Anonymous. The Paris Exhibition. Scientific American, 38(25):383, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-383.pdf.

Anonymous:1878:PUS

[11606] Anonymous. The problem of unsinkable ships. Scientific American, 38(25):384, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
384.pdf.

Anonymous:1878:PWV

[11607] Anonymous. The phonograph wins a victory. Scientific American, 38(25):384, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
384a.pdf.

Anonymous:1878:RMP

[11608] Anonymous. Reapers and mowers at the Paris Exposition. Scientific American, 38(25):384, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
384b.pdf.

Anonymous:1878:EMT

[11609] Anonymous. Edison’s micro-tasimeter. Scientific American, 38(25):385, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-385.pdf.

Anonymous:1878:GWC

[11610] Anonymous. Good work from cheap patents. Scientific American, 38(25):385, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
385a.pdf.

Anonymous:1878:CVG

[11611] Anonymous. Collmann’s valve gear. Scientific American, 38 (25):386–387, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
386.pdf.

Anonymous:1878:EIM

[11612] Anonymous. An ether ice-making machine. Scientific American, 38(25):387, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
387.pdf.

Anonymous:1878:ISBb

[11613] Anonymous. Improved shoe brush. Scientific American, 38(25): 387–388, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
387a.pdf.

Anonymous:1878:Cx

[11614] Anonymous. Communications. Scientific American, 38(25):388, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-388.pdf.

Anonymous:1878:HMA

[11615] Anonymous. Hughes’ microphone an alleged piracy. Scientific American, 38(25):388, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
388a.pdf.

Michels:1878:TS

[11616] John Michels. The trichinze in shad. Scientific American, 38(25):389, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-389.pdf.

Anonymous:1878:Ma

[11617] Anonymous. Microscopy. Scientific American, 38(25):389–390, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-389a.pdf.

Anonymous:1878:CID

[11618] Anonymous. Carr’s improved dredger. Scientific American, 38(25):390, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-390.pdf.

Anonymous:1878:IHA

[11619] Anonymous. Improved hot air furnace. Scientific American, 38 (25):390–391, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
390a.pdf.

Anonymous:1878:BPbd

[11620] Anonymous. Birds of paradise. Scientific American, 38(25):391–392, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-391.pdf.

Anonymous:1878:MTN

[11621] Anonymous. M. Trouve’s new telephone improvements. Scientific American, 38(25):392, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
392.pdf.

Anonymous:1878:PMe

[11622] Anonymous. Plant mind. Scientific American, 38(25):393, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-393.pdf.

Anonymous:1878:NCPa

[11623] Anonymous. A novel chain pump. Scientific American, 38(25):393, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-393a.pdf.

Anonymous:1878:LPE

[11624] Anonymous. Letters from the Paris Exhibition. Scientific American, 38(25):394, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
394.pdf.

Anonymous:1878:CWa

[11625] Anonymous. Catalpa wood. Scientific American, 38(25):395, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-395.pdf.

Anonymous:1878:FE

[11626] Anonymous. Fence economy. Scientific American, 38(25):395, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n25/pdf/scientificamerican06221878-395a.pdf.

Anonymous:1878:CRt

[11627] Anonymous. Communications received. Scientific American, 38 (25):396, June 22, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n25/pdf/scientificamerican06221878-
396c.pdf.

Anonymous:1878:BPbe

[11628] Anonymous. Business and personal. Scientific American, 38(26):410, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-410.pdf.

Anonymous:1878:BNBo

[11629] Anonymous. Books: New books and publications. Scientific American, 38(26):410, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
410a.pdf.

Anonymous:1878:NQv

[11630] Anonymous. Notes and queries. Scientific American, 38(26):410, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-410b.pdf.

Anonymous:1878:Ie

[11631] Anonymous. Index. Scientific American, 38(26):411–412, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-411.pdf.

Anonymous:1878:PNPy

[11632] Anonymous. Patent notes: Patents. Scientific American, 38(26):414, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-414.pdf.

Anonymous:1878:FSSa

[11633] Anonymous. The Firmenich safety steam boiler. Scientific American, 38(26):399, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
399.pdf.

Anonymous:1878:MMM

[11634] Anonymous. Mind, muscle, and machinery. Scientific American, 38(26):400, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
400.pdf.

Anonymous:1878:NI

[11635] Anonymous. A neglected industry. Scientific American, 38(26):400–401, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-400a.pdf.

Anonymous:1878:KS

[11636] Anonymous. A Keely seance. Scientific American, 38(26):401, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-401.pdf.

Anonymous:1878:IP

[11637] Anonymous. An industrial paradox. Scientific American, 38(26):401, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-401a.pdf.

Wright:1878:ANs

[11638] Berlin H. Wright. Astronomical notes. Scientific American, 38(26):402, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-402.pdf.

Anonymous:1878:FTGb

[11639] Anonymous. Fire telegraphs in Germany. Scientific American, 38 (26):402–404, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
402a.pdf.

Anonymous:1878:OII

[11640] Anonymous. Our iron industry. Scientific American, 38(26):404, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-404.pdf.

Anonymous:1878:CNT

[11641] Anonymous. Combination needle and thread cutter. Scientific American, 38(26):404, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
404a.pdf.

Anonymous:1878:SFb

[11642] Anonymous. A simple furnace. Scientific American, 38(26):404, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-404b.pdf.

Anonymous:1878:Cy

[11643] Anonymous. Communications. Scientific American, 38(26):404–405, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-404c.pdf.

Anonymous:1878:NPE

[11644] Anonymous. Notes from the Paris Exhibition. Scientific American, 38(26):405–406, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
405.pdf.

Anonymous:1878:ISMd

[11645] Anonymous. Improved shears for marking stock. Scientific American, 38(26):406, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
406.pdf.

Anonymous:1878:FSSb

[11646] Anonymous. File sharpening by the sand blast. Scientific American, 38(26):406, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
406a.pdf.

Anonymous:1878:DP

[11647] Anonymous. The doum palm. Scientific American, 38(26):407, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-407.pdf.

Anonymous:1878:OP

[11648] Anonymous. Official prejudice. Scientific American, 38(26):407, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-407a.pdf.

Anonymous:1878:IY

[11649] Anonymous. Improved yielding hat. Scientific American, 38(26):407, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-407b.pdf.

Anonymous:1878:NPLb

[11650] Anonymous. Notes of patent law-decisions of the courts. Scientific American, 38(26):407–408, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
407c.pdf.

Anonymous:1878:NLS

[11651] Anonymous. The new Laurent saccharometer. Scientific American, 38(26):408, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
408.pdf.

Anonymous:1878:PGD

[11652] Anonymous. Patent gear-dressing machine. Scientific American, 38(26):409, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
409.pdf.

Anonymous:1878:IGP

[11653] Anonymous. An improved gang plow. Scientific American, 38(26):409, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v38/n26/pdf/scientificamerican06291878-409a.pdf.

Anonymous:1878:PLPa

[11654] Anonymous. Poisonous leguminous plants. Scientific American, 38(26):409, June 29, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v38/n26/pdf/scientificamerican06291878-
409b.pdf.

Anonymous:1878:BPbf

[11655] Anonymous. Business and personal. Scientific American, 39(1):12, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-12.pdf.

Anonymous:1878:NQNf

[11656] Anonymous. Notes and queries: Notes and quries. Scientific American, 39(1):12, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
12a.pdf.

Hopkins:1878:EISa

[11657] George M. Hopkins. Electrical indicator for showing the rotation of the Earth. Scientific American, 39(1):1–5, July 6, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n1/pdf/scientificamerican07061878-1.pdf.

Anonymous:1878:ERN

[11658] Anonymous. The elevated railroad nuisance. Scientific American, 39(1): 2, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-2.pdf.

Anonymous:1878:SBb

[11659] Anonymous. Steam boilers. Scientific American, 39(1):2, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-2a.pdf.

Anonymous:1878:POW

[11660] Anonymous. Progress of our Western industries. Scientific American, 39(1):2–3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
2b.pdf.

Anonymous:1878:DWI

[11661] Anonymous. The decline of the whaling industry. Scientific American, 39(1):3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
3.pdf.

Anonymous:1878:TPE

[11662] Anonymous. Transmitting power by electricity. Scientific American, 39(1):3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
3a.pdf.

Anonymous:1878:SAEa

[11663] Anonymous. Scientific American export edition for June. Scientific American, 39(1):3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
3b.pdf.

Anonymous:1878:EAE

[11664] Anonymous. The Eothen Arctic Expedition. Scientific American, 39(1): 3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-3c.pdf.

Anonymous:1878:PMCa

[11665] Anonymous. Patent matters in Congress. Scientific American, 39(1):3, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-3d.pdf.

Anonymous:1878:HCE

[11666] Anonymous. Horizontal condensing engine at the Paris Exhibition. Scientific American, 39(1):4, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
4.pdf.

Anonymous:1878:MTE

[11667] Anonymous. Mr. Thomas A. Edison. Scientific American, 39(1):5, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-5.pdf.

Anonymous:1878:PICb

[11668] Anonymous. Patteson’s improved car coupling. Scientific American, 39(1):5–6, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
5a.pdf.

Wright:1878:ANt

[11669] Berlin H. Wright. Astronomical notes. Scientific American, 39(1):7, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-7.pdf.

Anonymous:1878:IWc

[11670] Anonymous. Improved wrench. Scientific American, 39(1):8, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-8.pdf.

Anonymous:1878:RDMb

[11671] Anonymous. Radial drilling machine. Scientific American, 39(1):8, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-8a.pdf.

Anonymous:1878:ISO

[11672] Anonymous. Improved self-oiling car wheel. Scientific American, 39(1):8–9, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
8b.pdf.

Anonymous:1878:WHT

[11673] Anonymous. The White head torpedo. Scientific American, 39(1):9, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n1/pdf/scientificamerican07061878-9.pdf.

Anonymous:1878:ETR

[11674] Anonymous. Edison’s telephonic researches. Scientific American, 39(1):10–11, July 6, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n1/pdf/scientificamerican07061878-
10.pdf.

Anonymous:1878:BPbg

[11675] Anonymous. Business and personal. Scientific American, 39(2):27, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-27.pdf.

Anonymous:1878:BNBp

[11676] Anonymous. Books: New books and publications. Scientific American, 39(2):27, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
27a.pdf.

Anonymous:1878:NQw

[11677] Anonymous. Notes and queries. Scientific American, 39(2):27–28, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-27b.pdf.

Anonymous:1878:IIt

[11678] Anonymous. Index of inventions. Scientific American, 39(2):28, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-28.pdf.

Anonymous:1878:AGP

[11679] Anonymous. The Adams gas process. Scientific American, 39(2):15–16, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-15.pdf.

Anonymous:1878:IM

[11680] Anonymous. The invention of the microphone. Scientific American, 39(2):16, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
16.pdf.

Anonymous:1878:PIF

[11681] Anonymous. Preparation of iron fuels. Scientific American, 39(2):16, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-16a.pdf.

Anonymous:1878:Mg

[11682] Anonymous. Millstones. Scientific American, 39(2):16, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-16b.pdf.

Anonymous:1878:HE

[11683] Anonymous. An hour with edison. Scientific American, 39(2):17, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-17.pdf.

Anonymous:1878:USV

[11684] Anonymous. Unsinkable steam vessels. Scientific American, 39(2):18, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-18.pdf.

Anonymous:1878:LCVa

[11685] Anonymous. Lever and cam valve. Scientific American, 39(2):19, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-19.pdf.

Anonymous:1878:ITb

[11686] Anonymous. Ingenious toy. Scientific American, 39(2):19, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-19a.pdf.

Anonymous:1878:MSB

[11687] Anonymous. Milk as a substitute for blood transfusion. Scientific American, 39(2):19, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
19b.pdf.

Anonymous:1878:SFE

[11688] Anonymous. A simple fire escape. Scientific American, 39(2):19, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-19c.pdf.

Anonymous:1878:Cz

[11689] Anonymous. Communication. Scientific American, 39(2):20, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-20.pdf.

Bellerophon:1878:EAC

[11690] Bellerophon. The etiology of Asiatic cholera. — a new theory. Scientific American, 39(2):20, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
20a.pdf.

Blair:1878:PPF

[11691] John Blair. Proposed process for the fixation of atmospheric nitrogen. Scientific American, 39(2):21, July 13, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n2/pdf/scientificamerican07131878-21.pdf.

Anonymous:1878:IBa

[11692] Anonymous. Improved beehive. Scientific American, 39(2):22, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-22.pdf.

Anonymous:1878:CIG

[11693] Anonymous. Cross’ improved gas condenser. Scientific American, 39(2):22–23, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
22a.pdf.

Anonymous:1878:SHP

[11694] Anonymous. The Swiss House at the Paris Exposition. Scientific American, 39(2):23, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n2/pdf/scientificamerican07131878-
23.pdf.

Wright:1878:ANu

[11695] Berlin H. Wright. Astronomical notes. Scientific American, 39(2):23, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-23a.pdf.

Anonymous:1878:SBT

[11696] Anonymous. The St. Benoît twins. Scientific American, 39(2):24–25, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-24.pdf.

Anonymous:1878:CHF

[11697] Anonymous. Curious hedge figures. Scientific American, 39(2):25, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-25.pdf.

Anonymous:1878:LGS

[11698] Anonymous. The Leona goat sucker. Scientific American, 39(2):25–26, July 13, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n2/pdf/scientificamerican07131878-25a.pdf.

Anonymous:1878:BPbh

[11699] Anonymous. Business and personal. Scientific American, 39(3):43, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-43.pdf.

Anonymous:1878:NQx

[11700] Anonymous. Notes and queries. Scientific American, 39(3):43, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-43a.pdf.

Anonymous:1878:IIu

[11701] Anonymous. Index of inventions. Scientific American, 39(3):43–44, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-43b.pdf.

Anonymous:1878:IVA

[11702] Anonymous. Improved variable automatic cut-off. Scientific American, 39(3):31–32, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
31.pdf.

Anonymous:1878:UM

[11703] Anonymous. The uses of mechanism. Scientific American, 39(3):32, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-32.pdf.

Anonymous:1878:WGO

[11704] Anonymous. Working gold ores. Scientific American, 39(3):32–33, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-32a.pdf.

Langley:1878:Sa

[11705] S. P. Langley. The Sun. Scientific American, 39(3):33–34, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-33.pdf.

Anonymous:1878:PEN

[11706] Anonymous. Professor Edison’s new carbon rheostat. Scientific American, 39(3):35, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
35.pdf.

Anonymous:1878:CEG

[11707] Anonymous. The chase elemental governor. Scientific American, 39(3):35–37, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
35a.pdf.

Anonymous:1878:Cbb

[11708] Anonymous. Communications. Scientific American, 39(3):37, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-37.pdf.

Anonymous:1878:IPR

[11709] Anonymous. Improved piston rod stuffing box. Scientific American, 39(3):38, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
38.pdf.

Anonymous:1878:IAF

[11710] Anonymous. Improved automatic fan. Scientific American, 39(3):38, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-38a.pdf.

Anonymous:1878:ISBc

[11711] Anonymous. Improved step box. Scientific American, 39(3):38–39, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-38b.pdf.

Anonymous:1878:WCS

[11712] Anonymous. Wood carver of Simla. Scientific American, 39(3):39–40, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-39.pdf.

Wright:1878:ANv

[11713] Berlin H. Wright. Astronomical notes. Scientific American, 39(3):40, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-40.pdf.

Anonymous:1878:Pb

[11714] Anonymous. The phonograph. Scientific American, 39(3):40, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-40a.pdf.

Anonymous:1878:AAM

[11715] Anonymous. Apparatus for administering medicine to horses. Scientific American, 39(3):41, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
41.pdf.

Anonymous:1878:BOT

[11716] Anonymous. A boat older than the Ark. Scientific American, 39(3):41–42, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n3/pdf/scientificamerican07201878-
41a.pdf.

Anonymous:1878:OB

[11717] Anonymous. The Otto bicycle. Scientific American, 39(3):42, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-42.pdf.

Anonymous:1878:SGG

[11718] Anonymous. A simple gas generator. Scientific American, 39(3):42, July 20, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n3/pdf/scientificamerican07201878-42a.pdf.

Anonymous:1878:BPbi

[11719] Anonymous. Business and personal. Scientific American, 39(4):59, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-59.pdf.

Anonymous:1878:NQy

[11720] Anonymous. Notes and queries. Scientific American, 39(4):59–60, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-59a.pdf.

Anonymous:1878:IIv

[11721] Anonymous. Index of inventions. Scientific American, 39(4):60, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-60a.pdf.

Anonymous:1878:NTS

[11722] Anonymous. The new twin steamer “Calais-Douvées”. Scientific American, 39(4):47, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
47.pdf.

Anonymous:1878:VRb

[11723] Anonymous. Vulcanizing rubber. Scientific American, 39(4):48, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-48.pdf.

Anonymous:1878:SCb

[11724] Anonymous. Strawberries and constipation. Scientific American, 39(4):48, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
48a.pdf.

Anonymous:1878:PLPb

[11725] Anonymous. Professor Langley’s papers on the Sun. Scientific American, 39(4):48, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
48b.pdf.

Anonymous:1878:DVC

[11726] Anonymous. Destruction vs construction of ironclads. Scientific American, 39(4):48–49, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
48c.pdf.

Langley:1878:Sb

[11727] S. P. Langley. The Sun. Scientific American, 39(4):49–50, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-49.pdf.

Anonymous:1878:EP

[11728] Anonymous. Edison’s phonomoter. Scientific American, 39(4):51, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-51.pdf.

Anonymous:1878:ESN

[11729] Anonymous. Excavating scoop, new mechanical inventions, and more. Scientific American, 39(4):51–52, July 27, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n4/pdf/scientificamerican07271878-51a.pdf.

Anonymous:1878:Cbc

[11730] Anonymous. Communications. Scientific American, 39(4):52, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-52.pdf.

Rose:1878:HUF

[11731] Joshua Rose. How to use a file. Scientific American, 39(4):53, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-53.pdf.

Anonymous:1878:NSC

[11732] Anonymous. New screw cutting lathe. Scientific American, 39(4):53–54, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-53a.pdf.

Anonymous:1878:NCM

[11733] Anonymous. New cloth measuring apparatus. Scientific American, 39(4):54, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
54.pdf.

Anonymous:1878:GWS

[11734] Anonymous. Gampert’s wood-sole shoe. Scientific American, 39 (4):54–55, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
54a.pdf.

Anonymous:1878:LCN

[11735] Anonymous. Landing of Cleopatra’s needle. Scientific American, 39(4):55–56, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
55.pdf.

Anonymous:1878:TES

[11736] Anonymous. The total eclipse of the Sun. Scientific American, 39(4):56–57, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
56.pdf.

Anonymous:1878:NIF

[11737] Anonymous. New iron fence. Scientific American, 39(4):57, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-57.pdf.

Anonymous:1878:Ac

[11738] Anonymous. The Adjutant. Scientific American, 39(4):57–58, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-57a.pdf.

Wright:1878:ANw

[11739] Berlin H. Wright. Astronomical notes. Scientific American, 39(4):58, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-58.pdf.

Anonymous:1878:CTb

[11740] Anonymous. The curiosities of tobacco. Scientific American, 39(4):58, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-58a.pdf.

Anonymous:1878:PFH

[11741] Anonymous. Preserving fish by hydraulic pressure. Scientific American, 39(4):58, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n4/pdf/scientificamerican07271878-
58b.pdf.

Anonymous:1878:CRu

[11742] Anonymous. Communications received. Scientific American, 39(4):60, July 27, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n4/pdf/scientificamerican07271878-60.pdf.

Anonymous:1878:BPbj

[11743] Anonymous. Business and personal. Scientific American, 39(5):75, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-75.pdf.

Anonymous:1878:BNBq

[11744] Anonymous. Books: New books and publications. Scientific American, 39(5):75, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
75a.pdf.

Anonymous:1878:NQz

[11745] Anonymous. Notes and queries. Scientific American, 39(5):75, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-75b.pdf.

Anonymous:1878:IIw

[11746] Anonymous. Index of inventions. Scientific American, 39(5):76, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-76.pdf.

Anonymous:1878:PNPz

[11747] Anonymous. Patent notes: Patents. Scientific American, 39(5):78, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-78.pdf.

Anonymous:1878:SLG

[11748] Anonymous. “Snyder’s Little Giant” engine. Scientific American, 39(5):63, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
63.pdf.

Anonymous:1878:SSb

[11749] Anonymous. Smoke and sparks. Scientific American, 39(5):64, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-64.pdf.

Anonymous:1878:CSA

[11750] Anonymous. Costs of silver amalgamation. Scientific American, 39(5):64, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
64a.pdf.

Anonymous:1878:AFT

[11751] Anonymous. Advertising for foreign trade. Scientific American, 39(5):64, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
64b.pdf.

Anonymous:1878:FPa

[11752] Anonymous. French pavements. Scientific American, 39(5):65–66, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-65.pdf.

Anonymous:1878:EGM

[11753] Anonymous. Emery grinding machinery. Scientific American, 39(5):67, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-67.pdf.

Anonymous:1878:NPOe

[11754] Anonymous. Notes of Patent Office decisions. Scientific American, 39(5):67–68, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
67a.pdf.

Anonymous:1878:Cbd

[11755] Anonymous. Communication. Scientific American, 39(5):68–69, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-68.pdf.

Anonymous:1878:NPP

[11756] Anonymous. New power press. Scientific American, 39(5):69–70, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-69.pdf.

Anonymous:1878:NPM

[11757] Anonymous. New portable mill. Scientific American, 39(5):70, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-70.pdf.

Anonymous:1878:NSPc

[11758] Anonymous. The new Spanish process for silver and copper. Scientific American, 39(5):70, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
70a.pdf.

Anonymous:1878:AJ

[11759] Anonymous. The Australian jabiru. Scientific American, 39(5):70, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-70b.pdf.

Anonymous:1878:ZGFa

[11760] Anonymous. Zoological Garden at Fairmount Park Philadelphia. Scientific American, 39(5):71, August 3, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n5/pdf/scientificamerican08031878-71.pdf.

Anonymous:1878:SMW

[11761] Anonymous. Sheet metal working presses. Scientific American, 39(5):72, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
72.pdf.

Anonymous:1878:RAR

[11762] Anonymous. Rules for artificial respiration in the treatment of the drowned. Scientific American, 39(5):73, August 3, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n5/pdf/scientificamerican08031878-73.pdf.

Anonymous:1878:CIa

[11763] Anonymous. A camera improvement. Scientific American, 39(5):73, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-73a.pdf.

Anonymous:1878:SOA

[11764] Anonymous. Safety oxygen apparatus. Scientific American, 39(5):73, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-73b.pdf.

Wright:1878:ANx

[11765] Berlin H. Wright. Astronomical notes. Scientific American, 39(5):74, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n5/pdf/scientificamerican08031878-74.pdf.

Anonymous:1878:CRv

[11766] Anonymous. Communications received. Scientific American, 39 (5):75–76, August 3, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n5/pdf/scientificamerican08031878-
75c.pdf.

Anonymous:1878:BPbk

[11767] Anonymous. Business and personal. Scientific American, 39(6):91, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-91.pdf.

Anonymous:1878:BNBr

[11768] Anonymous. Books: New books and publications. Scientific American, 39(6):91, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
91a.pdf.

Anonymous:1878:NQba

[11769] Anonymous. Notes and queries. Scientific American, 39(6):91, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-91b.pdf.

Anonymous:1878:IIx

[11770] Anonymous. Index of inventions. Scientific American, 39(6):91–92, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-91d.pdf.

Anonymous:1878:PNPba

[11771] Anonymous. Patent notes: Patents. Scientific American, 39(6):94, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-94.pdf.

Anonymous:1878:PESa

[11772] Anonymous. The Paris Exhibition. — a sketch in the Park. Scientific American, 39(6):79, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
79.pdf.

Anonymous:1878:PIM

[11773] Anonymous. Progress of iron making. Scientific American, 39 (6):80, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
80.pdf.

Anonymous:1878:TPS

[11774] Anonymous. The telephone as a promoter of science. Scientific American, 39(6):80, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
80a.pdf.

Anonymous:1878:LPH

[11775] Anonymous. Letter from Professor Hughes. Scientific American, 39(6):80, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
80b.pdf.

Anonymous:1878:SAEb

[11776] Anonymous. The Scientific American export edition. Scientific American, 39(6):80, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
80c.pdf.

Anonymous:1878:FPb

[11777] Anonymous. Fountain pens. Scientific American, 39(6):80, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-80d.pdf.

Langley:1878:Sc

[11778] S. P. Langley. The Sun. Scientific American, 39(6):80–81, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-80e.pdf.

Anonymous:1878:MPS

[11779] Anonymous. Mechanical puddling in Sweden. Scientific American, 39(6):82–83, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
82.pdf.

Anonymous:1878:NDS

[11780] Anonymous. A new deep sea thermometer. Scientific American, 39(6):83, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
83.pdf.

Anonymous:1878:Cbg

[11781] Anonymous. Communication. Scientific American, 39(6):84, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-84.pdf.

Anonymous:1878:DC

[11782] Anonymous. The distillation of coal. Scientific American, 39 (6):85–86, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
85.pdf.

Anonymous:1878:NCBb

[11783] Anonymous. New cutting and boring attachment for lathes. Scientific American, 39(6):86, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
86.pdf.

Anonymous:1878:NSV

[11784] Anonymous. New steam valve. Scientific American, 39(6):86, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-86a.pdf.

Anonymous:1878:RH

[11785] Anonymous. The rhinoceros hornbill. Scientific American, 39 (6):87, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
87.pdf.

Anonymous:1878:JBP

[11786] Anonymous. The Japanese building at the Paris Exhibition. Scientific American, 39(6):87–88, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
87a.pdf.

Anonymous:1878:CJa

[11787] Anonymous. Crooked journalism. Scientific American, 39(6):88–89, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n6/pdf/scientificamerican08101878-88.pdf.

Anonymous:1878:VFE

[11788] Anonymous. A velocipede feat extraordinary. Scientific American, 39(6):89–90, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
89.pdf.

Wright:1878:ANy

[11789] Berlin H. Wright. Astronomical notes. Scientific American, 39 (6):90, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
90.pdf.

Anonymous:1878:CRw

[11790] Anonymous. Communications received. Scientific American, 39 (6):91, August 10, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n6/pdf/scientificamerican08101878-
91c.pdf.

Anonymous:1878:BPbl

[11791] Anonymous. Business and personal. Scientific American, 39(7):107, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-107.pdf.

Anonymous:1878:BNBs 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
107a.pdf.

Anonymous:1878:NQbb

[11793] Anonymous. Notes and queries. Scientific American, 39(7):107, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-107b.pdf.

Anonymous:1878:IIy

[11794] Anonymous. Index of inventions. Scientific American, 39(7):108, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-108.pdf.

Anonymous:1878:PNPbb

[11795] Anonymous. Patent notes: Patents. Scientific American, 39(7):110, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-110.pdf.

Anonymous:1878:HNP

[11796] Anonymous. Holly’s new pumping engine and automatic pressure regulator. Scientific American, 39(7):95–96, August 17, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n7/pdf/scientificamerican08171878-95.pdf.

Anonymous:1878:SCRa

[11797] Anonymous. Steam on common roads. Scientific American, 39 (7):96, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
96.pdf.

Anonymous:1878:ESJ

[11798] Anonymous. The eclipse of the sun, July 29, 1878. Scientific American, 39(7):96, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
96a.pdf.

Anonymous:1878:DIc

[11799] Anonymous. The discouragement of invention. Scientific American, 39(7):97, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
97.pdf.

Anonymous:1878:WOI

[11800] Anonymous. Where our inventors live. Scientific American, 39 (7):97, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
97a.pdf.

Anonymous:1878:MWb

[11801] Anonymous. Not so many out of work. Scientific American, 39 (7):97, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
97b.pdf.

Anonymous:1878:PTc

[11802] Anonymous. The partition of Turkey. Scientific American, 39 (7):97, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
97c.pdf.

Anonymous:1878:AI

[11803] Anonymous. Artificial indigo. Scientific American, 39(7):97, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-97d.pdf.

Anonymous:1878:PLS

[11804] Anonymous. Progress of labor saving machinery in the South. Scientific American, 39(7):97–98, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
97e.pdf.

Anonymous:1878:NCS

[11805] Anonymous. New cotton spooling machine. Scientific American, 39(7):98, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
98.pdf.

Anonymous:1878:NWG

[11806] Anonymous. A new working glove. Scientific American, 39(7):98, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-98a.pdf.

Anonymous:1878:EBA

[11807] Anonymous. Electromagnetic burglar alarm safe. Scientific American, 39(7):99, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
99.pdf.

Anonymous:1878:NCPb

[11808] Anonymous. New corn plow and marker. Scientific American, 39(7):99–100, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
99a.pdf.

Hopkins:1878:EISb

[11809] Geo. M. Hopkins. Electrical indicator for showing the rotation of the Earth. Scientific American, 39(7):100, August 17, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n7/pdf/scientificamerican08171878-100.pdf.

Anonymous:1878:TSC

[11810] Anonymous. Tin shafting cup. Scientific American, 39(7):100, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-100a.pdf.

Anonymous:1878:MMC

[11811] Anonymous. The musical mechanism of the Cincinnati Organ. Scientific American, 39(7):100, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
100b.pdf.

Anonymous:1878:FIEa

[11812] Anonymous. First impressions of the eclipse observations. Scientific American, 39(7):101, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
101.pdf.

Anonymous:1878:LMM

[11813] Anonymous. The Lechner mining machine. Scientific American, 39(7):102, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
102.pdf.

Anonymous:1878:LB

[11814] Anonymous. The lyre bird. Scientific American, 39(7):103, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-103.pdf.

Anonymous:1878:NBD

[11815] Anonymous. New bridge over the Douro River, Portugal. Scientific American, 39(7):103–104, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
103a.pdf.

Anonymous:1878:IICa

[11816] Anonymous. Ices and ice creaks. Scientific American, 39(7):104, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n7/pdf/scientificamerican08171878-104.pdf.

Anonymous:1878:MIR

[11817] Anonymous. The manufacture of India rubber. Scientific American, 39(7):105–106, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
105.pdf.

Wright:1878:ANz

[11818] Berlin H. Wright. Astronomical notes. Scientific American, 39 (7):106, August 17, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n7/pdf/scientificamerican08171878-
106.pdf.

Anonymous:1878:BPbm

[11819] Anonymous. Business and personal. Scientific American, 39(8):123, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-123.pdf.

Anonymous:1878:BNBt

[11820] Anonymous. Books: New books and publications. Scientific American, 39(8):123, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
123a.pdf.

Anonymous:1878:NQbc

[11821] Anonymous. Notes and queries. Scientific American, 39(8):123– 124, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
123b.pdf.

Anonymous:1878:IIz

[11822] Anonymous. Index of inventions. Scientific American, 39(8):124, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-124a.pdf.

Anonymous:1878:PNPbc

[11823] Anonymous. Patent notes: Patents. Scientific American, 39(8):126, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-126.pdf.

Anonymous:1878:CIc

[11824] Anonymous. A chance for inventors. Scientific American, 39(8):111, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-111.pdf.

Anonymous:1878:DSCa

[11825] Anonymous. The Delaware Ship Canal. Scientific American, 39 (8):111, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
111a.pdf.

Anonymous:1878:AIB

[11826] Anonymous. American inventions in Bavaria. Scientific American, 39(8):111, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
111b.pdf.

Anonymous:1878:EM

[11827] Anonymous. Edison’s megaphone. Scientific American, 39(8):111, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-111c.pdf.

Anonymous:1878:EUU

[11828] Anonymous. Edison and the unseen universe. Scientific American, 39(8):112, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
112.pdf.

Anonymous:1878:INS

[11829] Anonymous. Improvements needed in salt making. Scientific American, 39(8):112, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
112a.pdf.

Anonymous:1878:NPLc

[11830] Anonymous. The new patent law of Spain. Scientific American, 39(8):112, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
112b.pdf.

Anonymous:1878:LEM

[11831] Anonymous. Local encouragement of manufactures. Scientific American, 39(8):112–113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
112c.pdf.

Anonymous:1878:IWd

[11832] Anonymous. Independent workers. Scientific American, 39(8):113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-113.pdf.

Anonymous:1878:CIe

[11833] Anonymous. A curious insect. Scientific American, 39(8):113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-113a.pdf.

Anonymous:1878:POV

[11834] Anonymous. Photographic and other views of the eclipse. Scientific American, 39(8):113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
113b.pdf.

Anonymous:1878:ENP

[11835] Anonymous. The eclipse. — a note from Professor Mitchell. Scientific American, 39(8):113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
113c.pdf.

Anonymous:1878:SRL

[11836] Anonymous. The study of real life in schools. Scientific American, 39(8):113, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
113d.pdf.

Anonymous:1878:FEA

[11837] Anonymous. Further evidence of Atlantis. Scientific American, 39 (8):114–115, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
114.pdf.

Anonymous:1878:NTP

[11838] Anonymous. New trace and pad buckle. Scientific American, 39(8):115, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
115.pdf.

Anonymous:1878:QSH

[11839] Anonymous. Quick speed hand drill. Scientific American, 39 (8):115–116, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
115a.pdf.

Anonymous:1878:Cbh

[11840] Anonymous. Communications. Scientific American, 39(8):116, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-116.pdf.

Anonymous:1878:FR

[11841] Anonymous. Files and rasps. Scientific American, 39(8):117–118, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-117.pdf.

Anonymous:1878:SPb

[11842] Anonymous. A simple phonograph. Scientific American, 39(8):118, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-118.pdf.

Anonymous:1878:FSH

[11843] Anonymous. Folding shaving horse. Scientific American, 39(8): 118–119, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
118a.pdf.

Anonymous:1878:WLR

[11844] Anonymous. Wood’s lap ring. Scientific American, 39(8):119, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n8/pdf/scientificamerican08241878-119.pdf.

Anonymous:1878:PEI

[11845] Anonymous. Paris Exhibition. — the Italian facade. Scientific American, 39(8):119–120, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
119a.pdf.

Wright:1878:ANba

[11846] Berlin H. Wright. Astronomical notes. Scientific American, 39 (8):120–121, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
120.pdf.

Anonymous:1878:OSH

[11847] Anonymous. Oliver’s screw-headed key. Scientific American, 39 (8):121, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
121.pdf.

Anonymous:1878:NTR

[11848] Anonymous. New tin roving can. Scientific American, 39(8): 121–122, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
121a.pdf.

Anonymous:1878:CRx

[11849] Anonymous. Communications received. Scientific American, 39 (8):124, August 24, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n8/pdf/scientificamerican08241878-
124.pdf.

Anonymous:1878:BPbn

[11850] Anonymous. Business and personal. Scientific American, 39(9):139, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-139.pdf.

Anonymous:1878:BNBu

[11851] Anonymous. Books: New books and publications. Scientific American, 39(9):139, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
139a.pdf.

Anonymous:1878:NQbd

[11852] Anonymous. Notes and queries. Scientific American, 39(9):139– 140, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
139b.pdf.

Anonymous:1878:IIba

[11853] Anonymous. Index of inventions. Scientific American, 39(9):140, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-140a.pdf.

Anonymous:1878:PNPbd

[11854] Anonymous. Patent notes: Patents. Scientific American, 39(9):142, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-142.pdf.

Anonymous:1878:BD

[11855] Anonymous. The balance dynamometer. Scientific American, 39 (9):127, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
127.pdf.

Anonymous:1878:RI

[11856] Anonymous. The rights of investigators. Scientific American, 39(9):128, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
128.pdf.

Anonymous:1878:PV

[11857] Anonymous. The Planet Vulcan. Scientific American, 39(9):128, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-128a.pdf.

Anonymous:1878:AA

[11858] Anonymous. Ammonia in the air. Scientific American, 39(9):128, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-128b.pdf.

Anonymous:1878:SVH

[11859] Anonymous. Soft vs hard iron. Scientific American, 39(9):128–129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-128c.pdf.

Anonymous:1878:ND

[11860] Anonymous. The new divinity. Scientific American, 39(9):129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-129.pdf.

Anonymous:1878:WFM

[11861] Anonymous. The West as a field for manufactures. Scientific American, 39(9):129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
129a.pdf.

Anonymous:1878:PESb

[11862] Anonymous. Pneumatic engines for street cabs. Scientific American, 39(9):129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
129b.pdf.

Anonymous:1878:FA

[11863] Anonymous. A false alarm. Scientific American, 39(9):129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-129c.pdf.

Anonymous:1878:WT

[11864] Anonymous. The Wallingford tornado. Scientific American, 39 (9):129, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
129d.pdf.

Anonymous:1878:CBj

[11865] Anonymous. Captive balloon of 1878. Scientific American, 39 (9):130–132, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
130.pdf.

Wright:1878:ANbb

[11866] Berlin H. Wright. Astronomical notes. Scientific American, 39 (9):133, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
133.pdf.

Anonymous:1878:NOV

[11867] Anonymous. New oscillating valve. Scientific American, 39(9):134, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-134.pdf.

Anonymous:1878:NRH

[11868] Anonymous. New rein holder. Scientific American, 39(9):134, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-134a.pdf.

Anonymous:1878:NMC

[11869] Anonymous. The Needham musical cabinet. Scientific American, 39(9):134–135, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
134b.pdf.

Hopkins:1878:SF

[11870] Geo. M. Hopkins. The speaking flame. Scientific American, 39 (9):135–136, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
135.pdf.

Anonymous:1878:Mh

[11871] Anonymous. The mesquite. Scientific American, 39(9):137–138, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n9/pdf/scientificamerican08311878-137.pdf.

Anonymous:1878:CRy

[11872] Anonymous. Communications received. Scientific American, 39 (9):140, August 31, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n9/pdf/scientificamerican08311878-
140.pdf.

Anonymous:1878:Cbk

[11873] Anonymous. Correspondence. Scientific American, 39(10):148– 149, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
148.pdf.

Anonymous:1878:BPbo

[11874] Anonymous. Business and personal. Scientific American, 39 (10):155, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
155.pdf.

Anonymous:1878:BNBv

[11875] Anonymous. Books: New books and publications. Scientific American, 39(10):155, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
155a.pdf.

Anonymous:1878:NQbe

[11876] Anonymous. Notes and queries. Scientific American, 39(10): 155–156, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
155b.pdf.

Anonymous:1878:IIbb

[11877] Anonymous. Index of inventions. Scientific American, 39(10): 156, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
156a.pdf.

Anonymous:1878:PNPbe

[11878] Anonymous. Patent notes: Patents. Scientific American, 39 (10):158, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
158.pdf.

Anonymous:1878:CBP

[11879] Anonymous. The Chinese building at the Paris Exhibition. Scientific American, 39(10):143, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
143.pdf.

Anonymous:1878:IAA

[11880] Anonymous. Independence among artisans. Scientific American, 39(10):144, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
144.pdf.

Anonymous:1878:WAD

[11881] Anonymous. What Americans have done for Turkey. Scientific American, 39(10):144–145, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
144a.pdf.

Anonymous:1878:SEI

[11882] Anonymous. Some eggstraordinary inventions. Scientific American, 39(10):145, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
145.pdf.

Anonymous:1878:MO

[11883] Anonymous. The mining outlook. Scientific American, 39(10): 145, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
145a.pdf.

Anonymous:1878:NPT

[11884] Anonymous. New pulley turning machine. Scientific American, 39(10):146, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
146.pdf.

Anonymous:1878:SBAa

[11885] Anonymous. Sheaf-binding apparatus. Scientific American, 39(10): 146–147, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
146a.pdf.

Anonymous:1878:BSL

[11886] Anonymous. Bartholdi’s statue of Liberty. Scientific American, 39(10):147, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
147.pdf.

Anonymous:1878:WNA

[11887] Anonymous. Wenzel’s nickel anode. Scientific American, 39(10): 150, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
150.pdf.

Anonymous:1878:NAT

[11888] Anonymous. A novel animal trap. Scientific American, 39(10): 150, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
150a.pdf.

Anonymous:1878:NC

[11889] Anonymous. A new candlestick. Scientific American, 39(10): 150, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
150b.pdf.

Anonymous:1878:SCc

[11890] Anonymous. The spider crab. Scientific American, 39(10):151– 152, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
151.pdf.

Wright:1878:ANbc

[11891] Berlin H. Wright. Astronomical notes. Scientific American, 39 (10):152, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
152.pdf.

Anonymous:1878:NHB

[11892] Anonymous. The new harbor for Boulogne. Scientific American, 39(10):153, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
153.pdf.

Anonymous:1878:NTC

[11893] Anonymous. New telephone call signal. Scientific American, 39 (10):153, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
153a.pdf.

Anonymous:1878:FTI

[11894] Anonymous. Flowers. — their industrial and medicinal uses. Scientific American, 39(10):154, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
154.pdf.

Anonymous:1878:CCf

[11895] Anonymous. The chlorination of copper. Scientific American, 39 (10):154, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
154a.pdf.

Anonymous:1878:CRz

[11896] Anonymous. Communication received. Scientific American, 39 (10):156, September 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n10/pdf/scientificamerican09071878-
156.pdf.

Anonymous:1878:BPbp

[11897] Anonymous. Business and personal. Scientific American, 39 (11):171, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
171.pdf.

Anonymous:1878:BNBw

[11898] Anonymous. Books: New books and publications. Scientific American, 39(11):171, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
171a.pdf.

Anonymous:1878:NQbf

[11899] Anonymous. Notes and queries. Scientific American, 39(11): 171, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
171b.pdf.

Anonymous:1878:IIbc

[11900] Anonymous. Index of inventions. Scientific American, 39(11): 171–172, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
171c.pdf.

Anonymous:1878:WST

[11901] Anonymous. Waiting for something to turn up. Scientific American, 39(11):160, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
160.pdf.

Anonymous:1878:CCg

[11902] Anonymous. Charitable colonizing. Scientific American, 39(11): 160, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
160a.pdf.

Anonymous:1878:VOI

[11903] Anonymous. Value of observation in invention. Scientific American, 39(11):160, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
160b.pdf.

Anonymous:1878:SMc

[11904] Anonymous. Speculative mining. Scientific American, 39(11): 160–161, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
160c.pdf.

Anonymous:1878:AASb

[11905] Anonymous. The American Association at St. Louis. Scientific American, 39(11):161, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
161.pdf.

Anonymous:1878:FDS

[11906] Anonymous. Facade of the Danish section at the Paris Exhibition. Scientific American, 39(11):161, September 14, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n11/pdf/scientificamerican09141878-161a.pdf.

Anonymous:1878:APF

[11907] Anonymous. American plows in France. Scientific American, 39 (11):161, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
161b.pdf.

Anonymous:1878:PEd

[11908] Anonymous. Paris Exhibition of 1878. Scientific American, 39(11): 162–165, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
162.pdf.

Anonymous:1878:NSPd

[11909] Anonymous. A new sheep protector. Scientific American, 39 (11):166, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
166.pdf.

Anonymous:1878:EMS

[11910] Anonymous. Elwell’s middlings separator. Scientific American, 39 (11):166, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
166a.pdf.

Anonymous:1878:TB

[11911] Anonymous. The tailor bird. Scientific American, 39(11):167, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n11/pdf/scientificamerican09141878-167.pdf.

Anonymous:1878:AAEa

[11912] Anonymous. The American Arctic Expedition. Scientific American, 39(11):167–168, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
167a.pdf.

Anonymous:1878:NRC

[11913] Anonymous. A novel railway car window. Scientific American, 39 (11):168, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
168.pdf.

Anonymous:1878:NCPc

[11914] Anonymous. New coffee pot. Scientific American, 39(11):168, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n11/pdf/scientificamerican09141878-168a.pdf.

Anonymous:1878:LAF

[11915] Anonymous. Lescale’s automatic fire escape. Scientific American, 39 (11):169–170, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
169.pdf.

Hopking:1878:MT

[11916] Geo. M. Hopking. The micro-telephone. Scientific American, 39 (11):170, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
170.pdf.

Wright:1878:ANbd

[11917] Berlin H. Wright. Astronomical notes. Scientific American, 39 (11):170, September 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n11/pdf/scientificamerican09141878-
170a.pdf.

Anonymous:1878:BPbq

[11918] Anonymous. Business and personal. Scientific American, 39 (12):187, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
187.pdf.

Anonymous:1878:NQbg

[11919] Anonymous. Notes and queries. Scientific American, 39(12): 187–188, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
187a.pdf.

Anonymous:1878:IIbd

[11920] Anonymous. Index of inventions. Scientific American, 39(12): 188, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
188a.pdf.

Anonymous:1878:ELa

[11921] Anonymous. Electric lighting. Scientific American, 39(12):175– 178, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
175.pdf.

Anonymous:1878:OAM

[11922] Anonymous. The origin of American mechanical genius. Scientific American, 39(12):176, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
176.pdf.

Anonymous:1878:UW

[11923] Anonymous. The utilization of weeds. Scientific American, 39 (12):176, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
176a.pdf.

Anonymous:1878:MEb

[11924] Anonymous. Mental experimenting. Scientific American, 39(12): 176–177, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
176b.pdf.

Anonymous:1878:SBN

[11925] Anonymous. Senate Bill No. 300. Scientific American, 39(12): 177, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
177.pdf.

Anonymous:1878:BA

[11926] Anonymous. The British Association. Scientific American, 39 (12):177, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
177a.pdf.

Anonymous:1878:ICA

[11927] Anonymous. The imagination in the constructive arts. Scientific American, 39(12):177, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
177b.pdf.

Anonymous:1878:BWM

[11928] Anonymous. The Bag-Worm’s mother. Scientific American, 39 (12):177, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
177c.pdf.

Anonymous:1878:DGC

[11929] Anonymous. The dynamometer of the great captive balloon at the Paris Exhibition. Scientific American, 39(12):179, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n12/pdf/scientificamerican09211878-179.pdf.

Anonymous:1878:SBc

[11930] Anonymous. Sheaf binding. Scientific American, 39(12):179– 180, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
179a.pdf.

Anonymous:1878:GPT

[11931] Anonymous. Gang plow trials — Paris Exhibition. Scientific American, 39(12):180–181, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
180.pdf.

Anonymous:1878:NCL

[11932] Anonymous. A new Camera Lucida. Scientific American, 39 (12):181, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
181.pdf.

Anonymous:1878:NPLd

[11933] Anonymous. Notes of patent law. Scientific American, 39(12): 181, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
181a.pdf.

Anonymous:1878:NHP

[11934] Anonymous. A new head protector. Scientific American, 39 (12):182, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
182.pdf.

Anonymous:1878:Pc

[11935] Anonymous. The polyscope. Scientific American, 39(12):182, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n12/pdf/scientificamerican09211878-182a.pdf.

Anonymous:1878:PDa

[11936] Anonymous. Potato digger. Scientific American, 39(12):182– 183, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
182b.pdf.

Anonymous:1878:GSM

[11937] Anonymous. Gathering the sap of the maguey. Scientific American, 39(12):183, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
183.pdf.

Anonymous:1878:TC

[11938] Anonymous. The tufted coquette. Scientific American, 39(12): 183, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
183a.pdf.

Anonymous:1878:LLB

[11939] Anonymous. The lagetto or lace-bark tree. Scientific American, 39(12):184, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
184.pdf.

Anonymous:1878:PPDa

[11940] Anonymous. Professor Palmieri’s diagometer. Scientific American, 39 (12):185–186, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
185.pdf.

Wright:1878:ANbe

[11941] Berlin H. Wright. Astronomical notes. Scientific American, 39 (12):186, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
186.pdf.

Anonymous:1878:CRba

[11942] Anonymous. Communications received. Scientific American, 39 (12):188, September 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n12/pdf/scientificamerican09211878-
188.pdf.

Anonymous:1878:Cbl

[11943] Anonymous. Correspondence. Scientific American, 39(13):196– 198, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
196.pdf.

Anonymous:1878:BPbr

[11944] Anonymous. Business and personal. Scientific American, 39 (13):203, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
203.pdf.

Anonymous:1878:BNBx

[11945] Anonymous. Books: New books and publications. Scientific American, 39(13):203, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
203a.pdf.

Anonymous:1878:NQbh

[11946] Anonymous. Notes and queries. Scientific American, 39(13): 203–204, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
203b.pdf.

Anonymous:1878:IIbe

[11947] Anonymous. Index of inventions. Scientific American, 39(13): 204, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
204a.pdf.

Anonymous:1878:PEFa

[11948] Anonymous. Paris Exhibition — the facades of Annam, Persia, Siam, Tunis, Monaco, and San Marino, new explosive, and more. Scientific American, 39(13):191, September 28, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n13/pdf/scientificamerican09281878-191.pdf.

Anonymous:1878:PLB

[11949] Anonymous. Pocket life buoys. Scientific American, 39(13): 192, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
192.pdf.

Anonymous:1878:HPc

[11950] Anonymous. The heroes of the pestilence. Scientific American, 39 (13):192, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
192a.pdf.

Anonymous:1878:SMH

[11951] Anonymous. Silver mining here and abroad. Scientific American, 39 (13):192–193, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
192b.pdf.

Anonymous:1878:COD

[11952] Anonymous. Corundum. — its occurrence and distribution. Scientific American, 39(13):193, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
193.pdf.

Anonymous:1878:NPOf

[11953] Anonymous. Notes of Patent Office decisions. Scientific American, 39(13):193, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
193a.pdf.

Anonymous:1878:CBAa

[11954] Anonymous. The captive balloon-apparatus for testing the strength of materials. Scientific American, 39(13):194–195, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n13/pdf/scientificamerican09281878-194.pdf.

Anonymous:1878:ASY

[11955] Anonymous. Ancient stand for Yule log. Scientific American, 39 (13):195, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
195.pdf.

Anonymous:1878:Cbm

[11956] Anonymous. A cabinet. Scientific American, 39(13):195–196, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n13/pdf/scientificamerican09281878-195a.pdf.

Anonymous:1878:REc

[11957] Anonymous. The raising of the Eurydice. Scientific American, 39 (13):198–199, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
198.pdf.

Hopkins:1878:SEL

[11958] Geo. M. Hopkins. Simple electric light apparatus. Scientific American, 39(13):200, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
200.pdf.

Anonymous:1878:IEL

[11959] Anonymous. Interesting experiments with lycopodium seeds. Scientific American, 39(13):200, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
200a.pdf.

Anonymous:1878:MDPb

[11960] Anonymous. Method of determining the position of vessels. Scientific American, 39(13):200–201, September 28, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n13/pdf/scientificamerican09281878-200b.pdf.

Anonymous:1878:N

[11961] Anonymous. The neomorpha. Scientific American, 39(13):201– 202, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
201.pdf.

Wright:1878:ANbf

[11962] Berlin H. Wright. Astronomical notes. Scientific American, 39 (13):202, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
202.pdf.

Anonymous:1878:CRbb

[11963] Anonymous. Communications received. Scientific American, 39 (13):204, September 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n13/pdf/scientificamerican09281878-
204.pdf.

Anonymous:1878:BPbs

[11964] Anonymous. Business and personal. Scientific American, 39(14):219, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-219.pdf.

Anonymous:1878:NQbi

[11965] Anonymous. Notes and queries. Scientific American, 39(14):219–220, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-219a.pdf.

Anonymous:1878:IIbf

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-220.pdf.

Anonymous:1878:PCD

[11967] Anonymous. The pneumatic clock, depth to which roots penetrate, and more. Scientific American, 39(14):207, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-207.pdf.

Anonymous:1878:SCRb

[11968] Anonymous. Steam on the common roads. Scientific American, 39(14):208, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
208.pdf.

Anonymous:1878:PWCa

[11969] Anonymous. People who can’t be helped. Scientific American, 39(14):208, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
208a.pdf.

Anonymous:1878:WCO

[11970] Anonymous. Working of copper ores. Scientific American, 39(14):208, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-208b.pdf.

Anonymous:1878:CSW

[11971] Anonymous. Conference of Swiss wrists at Geneva. Scientific American, 39(14):209, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
209.pdf.

Anonymous:1878:TT

[11972] Anonymous. The “true thread”. Scientific American, 39(14):209, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-209a.pdf.

Anonymous:1878:PFI

[11973] Anonymous. A promising field for invention. Scientific American, 39(14):209, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
209b.pdf.

Anonymous:1878:ADc

[11974] Anonymous. American dexterity. Scientific American, 39(14):209, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-209c.pdf.

Anonymous:1878:CFE

[11975] Anonymous. Cuba as a field for enterprise. Scientific American, 39(14):209, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
209d.pdf.

Anonymous:1878:IDE

[11976] Anonymous. An Inventor’s difficulties in england, the secret of soap and water, and more. Scientific American, 39(14):209–210, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-209e.pdf.

Anonymous:1878:TMP

[11977] Anonymous. Testing machines at the Paris Exhibition. Scientific American, 39(14):210–212, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
210.pdf.

Anonymous:1878:RES

[11978] Anonymous. The recent eclipse of the Sun. Scientific American, 39(14):212, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
212.pdf.

Wright:1878:ANbg

[11979] Berlin H. Wright. Astronomical notes. Scientific American, 39 (14):212–213, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
212a.pdf.

Anonymous:1878:HTL

[11980] Anonymous. How to test a lathe. Scientific American, 39(14):213, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-213.pdf.

Anonymous:1878:SPA

[11981] Anonymous. The stylographic pen, an accident on the Mt Washington Railroad, and more. Scientific American, 39(14):213–214, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-213a.pdf.

Anonymous:1878:ZGFb

[11982] Anonymous. Zoological Garden at Fairmount Park, Philadelphia, how calicoes are made, and more. Scientific American, 39(14):214–216, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-214.pdf.

Anonymous:1878:PAE

[11983] Anonymous. The porcupine ant eaters recently discovered in New Guinea, hair eels, and more. Scientific American, 39(14):217–218, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n14/pdf/scientificamerican10051878-217.pdf.

Anonymous:1878:NCBc

[11984] Anonymous. The new carpet beetle, lac, and more. Scientific American, 39(14):218, October 5, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n14/pdf/scientificamerican10051878-
218.pdf.

Anonymous:1878:BPbt

[11985] Anonymous. Business and personal. Scientific American, 39 (15):234, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
234.pdf.

Anonymous:1878:NQbj

[11986] Anonymous. Notes and queries. Scientific American, 39(15): 234–235, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
234a.pdf.

Anonymous:1878:IIbg

[11987] Anonymous. Index of inventions. Scientific American, 39(15): 235–236, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
235a.pdf.

Hopkins:1878:NED

[11988] Geo. M. Hopkins. New electrical diapason. Scientific American, 39(15):223, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
223.pdf.

Anonymous:1878:RBG

[11989] Anonymous. Relation between the grate surface and the heating surface of boilers. Scientific American, 39(15):224, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-224.pdf.

Anonymous:1878:WCL

[11990] Anonymous. Wages and the cost of living. Scientific American, 39(15):224, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
224a.pdf.

Anonymous:1878:MLV

[11991] Anonymous. Magnesian lice vs. pure lime for mortar. Scientific American, 39(15):224–225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
224b.pdf.

Anonymous:1878:NAD

[11992] Anonymous. Novel applications of dynamite. Scientific American, 39(15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225.pdf.

Anonymous:1878:PWM

[11993] Anonymous. Progress of Western mining operations. Scientific American, 39(15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225a.pdf.

Anonymous:1878:BWS

[11994] Anonymous. The Boston Whittling Schools. Scientific American, 39(15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225b.pdf.

Anonymous:1878:RC

[11995] Anonymous. The restoration of Cuba. Scientific American, 39 (15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225c.pdf.

Anonymous:1878:OSE

[11996] Anonymous. Our September export edition. Scientific American, 39(15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225d.pdf.

Anonymous:1878:RWN

[11997] Anonymous. A remarkable wasps’ nest. Scientific American, 39 (15):225, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225e.pdf.

Anonymous:1878:APPb

[11998] Anonymous. The American prizes at Paris. Scientific American, 39 (15):225–226, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
225f.pdf.

Anonymous:1878:NED

[11999] Anonymous. New electrical diapason. Scientific American, 39 (15):226, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
226.pdf.

Anonymous:1878:L

[12000] Anonymous. Lampblack. Scientific American, 39(15):226, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-226a.pdf.

Anonymous:1878:HWE

[12001] Anonymous. Hastie’s water engine. Scientific American, 39(15): 227–228, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
227.pdf.

Anonymous:1878:RNE

[12002] Anonymous. Reynier’s new electric lamp. Scientific American, 39 (15):228–229, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
228.pdf.

Anonymous:1878:BSR

[12003] Anonymous. Bauer’s steel restoring and refining compound, a queer boat, and more. Scientific American, 39(15):229, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-229.pdf.

Anonymous:1878:NNL

[12004] Anonymous. A new nut lock. Scientific American, 39(15):230, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-230.pdf.

Anonymous:1878:ST

[12005] Anonymous. Simple telephones. Scientific American, 39(15):230, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-230a.pdf.

Anonymous:1878:NBL

[12006] Anonymous. New breech-loading rifle. Scientific American, 39 (15):230–231, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
230b.pdf.

Anonymous:1878:BS

[12007] Anonymous. The bearded saki. Scientific American, 39(15):231, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-231.pdf.

Anonymous:1878:RBP

[12008] Anonymous. Red bird of paradise. Scientific American, 39(15):231, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n15/pdf/scientificamerican10121878-231a.pdf.

Anonymous:1878:TBC

[12009] Anonymous. The torrey botanical club. Scientific American, 39 (15):231–232, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
231b.pdf.

Wright:1878:ANbh

[12010] Berlin H. Wright. Astronomical notes. Scientific American, 39 (15):232, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
232.pdf.

Anonymous:1878:NSM

[12011] Anonymous. New stethoscopic microphone. Scientific American, 39(15):233, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
233.pdf.

Anonymous:1878:PSC

[12012] Anonymous. A platform scale of constant equtlibrium. Scientific American, 39(15):233, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
233a.pdf.

Anonymous:1878:CRbc

[12013] Anonymous. Communications received. Scientific American, 39 (15):235, October 12, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n15/pdf/scientificamerican10121878-
235.pdf.

Anonymous:1878:Cbn

[12014] Anonymous. Correspondence. Scientific American, 39(16):244–245, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-244.pdf.

Anonymous:1878:BPbu

[12015] Anonymous. Business and personal. Scientific American, 39 (16):250, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
250a.pdf.

Anonymous:1878:NQbk

[12016] Anonymous. Notes and queries. Scientific American, 39(16): 250–251, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
250b.pdf.

Anonymous:1878:IIbh

[12017] Anonymous. Index of inventions. Scientific American, 39(16): 251–252, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
251a.pdf.

Anonymous:1878:ISMe

[12018] Anonymous. Improvements in sugar making wanted. Scientific American, 39(16):240, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
240.pdf.

Anonymous:1878:DE

[12019] Anonymous. Dust explosions. Scientific American, 39(16):240, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-240a.pdf.

Anonymous:1878:SMV

[12020] Anonymous. A second Matthew Vassar. Scientific American, 39 (16):240, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
240b.pdf.

Anonymous:1878:PHG

[12021] Anonymous. Poisonous hats, gloves, stockings, and clothing. Scientific American, 39(16):240–241, October 19, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n16/pdf/scientificamerican10191878-240c.pdf.

Anonymous:1878:PTd

[12022] Anonymous. The pauper tribe. Scientific American, 39(16):241, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-241.pdf.

Anonymous:1878:PSb

[12023] Anonymous. Public sanitation. Scientific American, 39(16):241, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-241a.pdf.

Anonymous:1878:HMS

[12024] Anonymous. A Horse’s motion scientifically determined. Scientific American, 39(16):241, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
241b.pdf.

Anonymous:1878:IEa

[12025] Anonymous. Improved education. Scientific American, 39(16):241, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-241c.pdf.

Anonymous:1878:SSWa

[12026] Anonymous. The spectroscope in solar work. Scientific American, 39(16):242–244, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
242.pdf.

Anonymous:1878:WF

[12027] Anonymous. A wrinkle in filing. Scientific American, 39(16): 245–246, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
245.pdf.

Anonymous:1878:ISF

[12028] Anonymous. Improved steam fire engine. Scientific American, 39(16):246, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
246.pdf.

Wright:1878:ANbi

[12029] Berlin H. Wright. Astronomical notes. Scientific American, 39 (16):246, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
246a.pdf.

Anonymous:1878:HPT

[12030] Anonymous. Hairy prehensile-tailed porcupine. Scientific American, 39(16):247, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
247.pdf.

Anonymous:1878:GBP

[12031] Anonymous. Golden bird of paradise. Scientific American, 39 (16):247, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
247a.pdf.

Anonymous:1878:MEBa

[12032] Anonymous. Mechanical exhibition at Boston. Scientific American, 39(16):248, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
248.pdf.

Anonymous:1878:AGB

[12033] Anonymous. American goods in Brazil. Scientific American, 39 (16):248–249, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
248a.pdf.

Anonymous:1878:If

[12034] Anonymous. To inventors. Scientific American, 39(16):250, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n16/pdf/scientificamerican10191878-250.pdf.

Anonymous:1878:CRbd

[12035] Anonymous. Communications received. Scientific American, 39 (16):251, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
251.pdf.

Anonymous:1878:VPH

[12036] Anonymous. Value of patents and how to obtain them. Scientific American, 39(16):253, October 19, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n16/pdf/scientificamerican10191878-
253.pdf.

Anonymous:1878:BPbv

[12037] Anonymous. Business and personal. Scientific American, 39 (17):267, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
267a.pdf.

Anonymous:1878:BNBy

[12038] Anonymous. Books: New books and publications. Scientific American, 39(17):267, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
267b.pdf.

Anonymous:1878:NQbl

[12039] Anonymous. Notes and queries. Scientific American, 39(17): 267–268, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
267c.pdf.

Anonymous:1878:IIbi

[12040] Anonymous. Index of inventions. Scientific American, 39(17):268, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n17/pdf/scientificamerican10261878-268.pdf.

Anonymous:1878:PNPbf

[12041] Anonymous. Patent notes: Patents. Scientific American, 39 (17):270, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
270.pdf.

Anonymous:1878:PTM

[12042] Anonymous. Patents and trade marks in England. Scientific American, 39(17):255, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
255.pdf.

Anonymous:1878:MCPb

[12043] Anonymous. The manufacture of the chickering piano. Scientific American, 39(17):255, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
255a.pdf.

Anonymous:1878:FWT

[12044] Anonymous. Flow of water through pipes. Scientific American, 39(17):256, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
256.pdf.

Anonymous:1878:CPB

[12045] Anonymous. A cockney plan to banish smoke. Scientific American, 39(17):256, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
256a.pdf.

Anonymous:1878:DBC

[12046] Anonymous. Decision of the to barrel courts relating machinery. Scientific American, 39(17):256–257, October 26, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n17/pdf/scientificamerican10261878-256b.pdf.

Anonymous:1878:MEc

[12047] Anonymous. The modulus of elasticity. Scientific American, 39 (17):257, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
257.pdf.

Anonymous:1878:NRTa

[12048] Anonymous. The noise of rapid transit — a chance for inventors. Scientific American, 39(17):257, October 26, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n17/pdf/scientificamerican10261878-257a.pdf.

Anonymous:1878:AAEb

[12049] Anonymous. American agricultural exhibits at Paris. Scientific American, 39(17):257, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
257b.pdf.

Anonymous:1878:PAMb

[12050] Anonymous. Pneumatic appliances for mines. Scientific American, 39(17):260–261, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
260.pdf.

Anonymous:1878:Ag

[12051] Anonymous. The Alhambra. Scientific American, 39(17):261, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n17/pdf/scientificamerican10261878-261.pdf.

Anonymous:1878:NPb

[12052] Anonymous. New protractor. Scientific American, 39(17):262, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n17/pdf/scientificamerican10261878-262.pdf.

Anonymous:1878:TTR

[12053] Anonymous. True time regulator. Scientific American, 39(17): 262–263, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
262a.pdf.

Anonymous:1878:MFC

[12054] Anonymous. Mr. Fowler crossing the Channel. Scientific American, 39(17):263, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
263.pdf.

Anonymous:1878:NBb

[12055] Anonymous. A new beehive. Scientific American, 39(17):263, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n17/pdf/scientificamerican10261878-263a.pdf.

Anonymous:1878:SCCa

[12056] Anonymous. Statue of Captain Cook. Scientific American, 39 (17):264–265, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
264.pdf.

Anonymous:1878:NGS

[12057] Anonymous. New guard for sheep shears. Scientific American, 39(17):265, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
265.pdf.

Anonymous:1878:NMS

[12058] Anonymous. New measuring stopper. Scientific American, 39 (17):265–266, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n17/pdf/scientificamerican10261878-
265a.pdf.

Anonymous:1878:Ig

[12059] Anonymous. To inventors. Scientific American, 39(17):267, October 26, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n17/pdf/scientificamerican10261878-267.pdf.

Anonymous:1878:Cbo

[12060] Anonymous. Correspondence. Scientific American, 39(18):276, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n18/pdf/scientificamerican11021878-276.pdf.

Anonymous:1878:BPbw

[12061] Anonymous. Business and personal. Scientific American, 39 (18):283, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
283a.pdf.

Anonymous:1878:NQbm

[12062] Anonymous. Notes and queries. Scientific American, 39(18): 283–284, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
283b.pdf.

Anonymous:1878:IIbj

[12063] Anonymous. Index of inventions. Scientific American, 39(18): 284, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
284a.pdf.

Anonymous:1878:PNPbg

[12064] Anonymous. Patent notes: Patents. Scientific American, 39 (18):284, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
284b.pdf.

Anonymous:1878:CLC

[12065] Anonymous. Chard’s lubricene and cups. Scientific American, 39 (18):271, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
271.pdf.

Anonymous:1878:FFA

[12066] Anonymous. The Forster–Firmin amalgamator. Scientific American, 39(18):271, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
271a.pdf.

Anonymous:1878:ELG

[12067] Anonymous. The electric light and the gas companies — remarkable effect of a new invention in the stock markets. Scientific American, 39(18):272, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
272.pdf.

Anonymous:1878:PEAa

[12068] Anonymous. Progress in England and America. Scientific American, 39(18):272, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
272a.pdf.

Anonymous:1878:ICP

[12069] Anonymous. The incoming Commissioner of Patents. Scientific American, 39(18):272–273, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
272b.pdf.

Anonymous:1878:SAEc

[12070] Anonymous. Success of American exhibitors at Paris. Scientific American, 39(18):273, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
273.pdf.

Anonymous:1878:SNE

[12071] Anonymous. Should the nation engage in manufactures. Scientific American, 39(18):273, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
273a.pdf.

Anonymous:1878:FG

[12072] Anonymous. Fuel gas. Scientific American, 39(18):273, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n18/pdf/scientificamerican11021878-273b.pdf.

Anonymous:1878:ICd

[12073] Anonymous. An improvement on the Chromos. Scientific American, 39(18):273, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
273c.pdf.

Anonymous:1878:SAO

[12074] Anonymous. A South Australian offer for an improvement. Scientific American, 39(18):273–274, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
273d.pdf.

Anonymous:1878:LT

[12075] Anonymous. Lymans’s trigonometer. Scientific American, 39 (18):274, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
274.pdf.

Anonymous:1878:PEe

[12076] Anonymous. The Paris Exhibition. Scientific American, 39(18): 274, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
274a.pdf.

Wright:1878:ANbj

[12077] Berlin H. Wright. Astronomical notes. Scientific American, 39 (18):277–278, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
277.pdf.

Anonymous:1878:NWO

[12078] Anonymous. The new Wilson oscillating Shuttle sewing machine. Scientific American, 39(18):278–280, November 2, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n18/pdf/scientificamerican11021878-278.pdf.

Anonymous:1878:SSWb

[12079] Anonymous. Serpulas, or sea worms. Scientific American, 39 (18):281, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
281.pdf.

Anonymous:1878:KT

[12080] Anonymous. King Tody. Scientific American, 39(18):281–283, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n18/pdf/scientificamerican11021878-281a.pdf.

Anonymous:1878:Ih

[12081] Anonymous. To inventors. Scientific American, 39(18):283, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n18/pdf/scientificamerican11021878-283.pdf.

Anonymous:1878:CRbe

[12082] Anonymous. Communications received. Scientific American, 39 (18):284, November 2, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n18/pdf/scientificamerican11021878-
284.pdf.

Anonymous:1878:BPbx

[12083] Anonymous. Business and personal. Scientific American, 39 (19):298, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
298a.pdf.

Anonymous:1878:BNBz 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
298b.pdf.

Anonymous:1878:NQbn

[12085] Anonymous. Notes and queries. Scientific American, 39(19): 298–299, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
298c.pdf.

Anonymous:1878:IIbk

[12086] Anonymous. Index of inventions. Scientific American, 39(19): 299–300, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
299a.pdf.

Anonymous:1878:PPE

[12087] Anonymous. Progress and prospects of the East River Bridge. Scientific American, 39(19):287, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
287.pdf.

Anonymous:1878:PMEa

[12088] Anonymous. Professor Morton on the electric light. Scientific American, 39(19):288–289, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
288.pdf.

Anonymous:1878:EDM

[12089] Anonymous. The Electrical Department in the Mechanics Fair, Boston, Mass. Scientific American, 39(19):289, November 9, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n19/pdf/scientificamerican11091878-289.pdf.

Anonymous:1878:FIEb

[12090] Anonymous. The French industrial exhibition of 1878. Scientific American, 39(19):289, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
289a.pdf.

Anonymous:1878:AHP

[12091] Anonymous. Awards and honors at Paris. Scientific American, 39(19):289, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
289b.pdf.

Anonymous:1878:WWI

[12092] Anonymous. Who will invent a satisfactory milking machine. Scientific American, 39(19):289, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
289c.pdf.

Anonymous:1878:HSCb

[12093] Anonymous. The hermetical sanitary closet. Scientific American, 39(19):290, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
290.pdf.

Anonymous:1878:TP

[12094] Anonymous. The trial of the “Pyx”. Scientific American, 39 (19):290, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
290a.pdf.

Anonymous:1878:NRB

[12095] Anonymous. New refrigerator basket. Scientific American, 39 (19):290, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
290b.pdf.

Anonymous:1878:NFS

[12096] Anonymous. A new fireproof shutter. Scientific American, 39 (19):290, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
290c.pdf.

Anonymous:1878:IRP

[12097] Anonymous. Ingram rotary press at the Paris Exhibition. Scientific American, 39(19):291, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
291.pdf.

Wright:1878:ANbk

[12098] Berlin H. Wright. Astronomical notes. Scientific American, 39 (19):292, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
292.pdf.

Anonymous:1878:CMVb

[12099] Anonymous. California mining vs farming. Scientific American, 39(19):292, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
292a.pdf.

Anonymous:1878:POF

[12100] Anonymous. Progress of our foreign trade. Scientific American, 39 (19):292–293, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
292b.pdf.

Anonymous:1878:NFPa

[12101] Anonymous. A new foot power. Scientific American, 39(19): 294, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
294.pdf.

Anonymous:1878:NWS

[12102] Anonymous. New wool scouring and rinsing machine. Scientific American, 39(19):294, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
294a.pdf.

Anonymous:1878:WMN

[12103] Anonymous. Washington memorials in Northamptonshire. Scientific American, 39(19):295, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
295.pdf.

Anonymous:1878:NMJ

[12104] Anonymous. A new measuring jacket. Scientific American, 39 (19):295–296, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
295a.pdf.

Anonymous:1878:NR

[12105] Anonymous. A new rheostat. Scientific American, 39(19):296, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n19/pdf/scientificamerican11091878-296.pdf.

Anonymous:1878:UB

[12106] Anonymous. The umbrella bird. Scientific American, 39(19): 297, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
297.pdf.

Anonymous:1878:Ij

[12107] Anonymous. To inventors. Scientific American, 39(19):298, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n19/pdf/scientificamerican11091878-298.pdf.

Anonymous:1878:CRbf

[12108] Anonymous. Communications received. Scientific American, 39 (19):299, November 9, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n19/pdf/scientificamerican11091878-
299.pdf.

Anonymous:1878:Cbq

[12109] Anonymous. Correspondence. Scientific American, 39(20):308– 309, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
308.pdf.

Anonymous:1878:BPby

[12110] Anonymous. Business and personal. Scientific American, 39 (20):315, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
315a.pdf.

Anonymous:1878:NQbo

[12111] Anonymous. Notes and queries. Scientific American, 39(20): 315, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
315b.pdf.

Anonymous:1878:IIbl

[12112] Anonymous. Index of inventions. Scientific American, 39(20): 315–316, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
315d.pdf.

Anonymous:1878:PNPbh

[12113] Anonymous. Patent notes: Patents. Scientific American, 39 (20):316, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
316.pdf.

Anonymous:1878:TVD

[12114] Anonymous. The torpedo vessel, destroyer. Scientific American, 39(20):303, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
303.pdf.

Anonymous:1878:SPc

[12115] Anonymous. Steam from petroleum. Scientific American, 39 (20):304, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
304.pdf.

Anonymous:1878:ANE

[12116] Anonymous. Another new electric light. Scientific American, 39 (20):304, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
304a.pdf.

Anonymous:1878:RBR

[12117] Anonymous. A remarkable bank robbery. — scientific guards neglected. Scientific American, 39(20):304–305, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n20/pdf/scientificamerican11161878-304b.pdf.

Anonymous:1878:GWF

[12118] Anonymous. A grand World’s fair in New York. Scientific American, 39(20):305, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
305.pdf.

Anonymous:1878:MWW

[12119] Anonymous. The motion of a wagon wheel. Scientific American, 39(20):305, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
305a.pdf.

Anonymous:1878:PIP

[12120] Anonymous. The Paris international patent Congress. Scientific American, 39(20):305, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
305b.pdf.

Anonymous:1878:PRW

[12121] Anonymous. Patent rights, and who oppose them. Scientific American, 39(20):305, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
305c.pdf.

Anonymous:1878:SJb

[12122] Anonymous. A steam juryman. Scientific American, 39(20): 305–306, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
305d.pdf.

Anonymous:1878:NGR

[12123] Anonymous. A new gas regulator. Scientific American, 39(20): 306, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
306.pdf.

Anonymous:1878:CTE

[12124] Anonymous. Combined traction engine and steam fire engine. Scientific American, 39(20):306–307, November 16, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n20/pdf/scientificamerican11161878-306a.pdf.

Anonymous:1878:VRC

[12125] Anonymous. Van Rennes’ caloric engine and pump. Scientific American, 39(20):307–308, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
307.pdf.

Anonymous:1878:WP

[12126] Anonymous. The Watson pump. Scientific American, 39(20): 309–310, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
309.pdf.

Anonymous:1878:PWCb

[12127] Anonymous. Pipe wrench and cutter. Scientific American, 39 (20):310, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
310.pdf.

Anonymous:1878:LME

[12128] Anonymous. Locomotive for the metropolitan elevated railway. Scientific American, 39(20):310–311, November 16, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n20/pdf/scientificamerican11161878-310a.pdf.

Anonymous:1878:DSH

[12129] Anonymous. Drilling square holes. Scientific American, 39(20): 311, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
311.pdf.

Anonymous:1878:RPb

[12130] Anonymous. Russian pottery. Scientific American, 39(20):311, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n20/pdf/scientificamerican11161878-311a.pdf.

Anonymous:1878:NMM

[12131] Anonymous. New mortising machine. Scientific American, 39(20): 311–312, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
311b.pdf.

Weight:1878:AN

[12132] Berlin H. Weight. Astronomical notes. Scientific American, 39 (20):312–313, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
312.pdf.

Hopkins:1878:SMM

[12133] Geo. M. Hopkins. Some modifications of the microphone and telephone. Scientific American, 39(20):313–314, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n20/pdf/scientificamerican11161878-313.pdf.

Anonymous:1878:RGQ

[12134] Anonymous. The Rockport granite quarries. Scientific American, 39(20):314, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
314.pdf.

Anonymous:1878:Ik

[12135] Anonymous. To inventors. Scientific American, 39(20):315, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n20/pdf/scientificamerican11161878-315.pdf.

Anonymous:1878:CRbg

[12136] Anonymous. Communications received. Scientific American, 39 (20):315, November 16, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n20/pdf/scientificamerican11161878-
315c.pdf.

Anonymous:1878:Cbs

[12137] Anonymous. Correspondence. Scientific American, 39(21):325, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n21/pdf/scientificamerican11231878-325.pdf.

Anonymous:1878:BPbz

[12138] Anonymous. Business and personal. Scientific American, 39 (21):331, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
331a.pdf.

Anonymous:1878:BNBba 8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
331b.pdf.

Anonymous:1878:NQbp

[12140] Anonymous. Notes and queries. Scientific American, 39(21): 331–332, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
331c.pdf.

Anonymous:1878:IIbm

[12141] Anonymous. Index of inventions. Scientific American, 39(21): 332, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
332a.pdf.

Anonymous:1878:PNPbi

[12142] Anonymous. Patent notes: Patents. Scientific American, 39 (21):332, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
332b.pdf.

Anonymous:1878:NSFa

[12143] Anonymous. A new steak fire engine. Scientific American, 39 (21):319, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
319.pdf.

Anonymous:1878:FLN

[12144] Anonymous. “Forney” locomotive for the New York Elevated Railroad. Scientific American, 39(21):319, November 23, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n21/pdf/scientificamerican11231878-319a.pdf.

Anonymous:1878:SVO

[12145] Anonymous. The steam value of oil fuels. Scientific American, 39(21):320, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
320.pdf.

Anonymous:1878:NBN

[12146] Anonymous. A new bank note paper wanted. Scientific American, 39(21):320, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
320a.pdf.

Anonymous:1878:TMTb

[12147] Anonymous. Trade mark treaty with Brazil. Scientific American, 39(21):320, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
320b.pdf.

Anonymous:1878:TH

[12148] Anonymous. The treatment of hydrophobia. Scientific American, 39(21):320, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
320c.pdf.

Anonymous:1878:MOP

[12149] Anonymous. The mechanical and other properties of iron and mild steel. Scientific American, 39(21):320–321, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n21/pdf/scientificamerican11231878-320d.pdf.

Anonymous:1878:PAPb

[12150] Anonymous. The proposed addition to the Patent Office. Scientific American, 39(21):321, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
321.pdf.

Anonymous:1878:TF

[12151] Anonymous. Table forks. Scientific American, 39(21):321, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n21/pdf/scientificamerican11231878-321a.pdf.

Anonymous:1878:YWP

[12152] Anonymous. A Year’s work in the Patent Office. Scientific American, 39(21):321, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
321b.pdf.

Anonymous:1878:NRTb

[12153] Anonymous. New rule in trade hark cases. Scientific American, 39(21):321, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
321c.pdf.

Anonymous:1878:FW

[12154] Anonymous. French wheelbarrows. Scientific American, 39(21): 322–324, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
322.pdf.

Anonymous:1878:ELb

[12155] Anonymous. The electric light. Scientific American, 39(21):326– 327, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
326.pdf.

Anonymous:1878:HMP

[12156] Anonymous. Hydraulic motors at the Paris Exhibition. Scientific American, 39(21):327, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
327.pdf.

Anonymous:1878:BEL

[12157] Anonymous. Burner for electric light. Scientific American, 39(21): 327–328, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
327a.pdf.

Anonymous:1878:NEO

[12158] Anonymous. A novel egg opener. Scientific American, 39(21): 328, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
328.pdf.

Anonymous:1878:ABL

[12159] Anonymous. Artificial ball lighting. Scientific American, 39(21): 328–330, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
328a.pdf.

Anonymous:1878:NASb

[12160] Anonymous. The National Academy of Sciences. Scientific American, 39(21):330, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
330.pdf.

Anonymous:1878:In

[12161] Anonymous. To inventors. Scientific American, 39(21):331, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n21/pdf/scientificamerican11231878-331.pdf.

Anonymous:1878:CRbh

[12162] Anonymous. Communications received. Scientific American, 39 (21):332, November 23, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n21/pdf/scientificamerican11231878-
332.pdf.

Anonymous:1878:BPca

[12163] Anonymous. Business and personal. Scientific American, 39 (22):347, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
347a.pdf.

Anonymous:1878:BNBbb

[12164] Anonymous. Books: New books and publications. Scientific American, 39(22):347, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
347b.pdf.

Anonymous:1878:NQNg

[12165] Anonymous. Notes and queries: Notes and querise. Scientific American, 39(22):347–348, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
347c.pdf.

Anonymous:1878:IIbn

[12166] Anonymous. Index of inventions. Scientific American, 39(22): 348, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
348a.pdf.

Anonymous:1878:PNPbj

[12167] Anonymous. Patent notes: Patents. Scientific American, 39 (22):348, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
348b.pdf.

Anonymous:1878:ETS

[12168] Anonymous. Electric time service for New York. Scientific American, 39(22):335, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
335.pdf.

Anonymous:1878:HM

[12169] Anonymous. The Hosmer motor. Scientific American, 39(22): 336, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
336.pdf.

Anonymous:1878:SFP

[12170] Anonymous. The Salisbury furnace for petroleum. Scientific American, 39(22):336, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
336a.pdf.

Anonymous:1878:PPDb

[12171] Anonymous. Patents for protecting the dead. Scientific American, 39(22):337, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
337.pdf.

Anonymous:1878:ORP

[12172] Anonymous. The official reports of the Paris Exhibition. Scientific American, 39(22):337, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
337a.pdf.

Anonymous:1878:SLN

[12173] Anonymous. Skilled labor in New York City. Scientific American, 39(22):337, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
337b.pdf.

Anonymous:1878:CIf

[12174] Anonymous. The cost of insecurity. Scientific American, 39 (22):337, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
337c.pdf.

Anonymous:1878:NPG

[12175] Anonymous. A new platen gauge. Scientific American, 39(22): 338, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
338.pdf.

Anonymous:1878:TRE

[12176] Anonymous. Tramway rail experiments. Scientific American, 39 (22):338, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
338a.pdf.

Anonymous:1878:NDP

[12177] Anonymous. A new draughty pencil. Scientific American, 39 (22):338, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
338b.pdf.

Anonymous:1878:SEG

[12178] Anonymous. Steam engine governor. Scientific American, 39 (22):339, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
339.pdf.

Anonymous:1878:NRA

[12179] Anonymous. New regulating attachment for clock pendulums. Scientific American, 39(22):339, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
339a.pdf.

Anonymous:1878:FSC

[12180] Anonymous. French subsoil and clearing plow. Scientific American, 39(22):340, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
340.pdf.

Anonymous:1878:BPCb

[12181] Anonymous. The blue process of copying tracings. Scientific American, 39(22):340, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
340a.pdf.

Anonymous:1878:MME

[12182] Anonymous. The De Meritens magneto-electric machine. Scientific American, 39(22):341, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
341.pdf.

Anonymous:1878:ICGb

[12183] Anonymous. Improved crusher and grinding mill. Scientific American, 39(22):341–342, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
341a.pdf.

Anonymous:1878:PIV

[12184] Anonymous. The poison ivy and Virginia creeper. Scientific American, 39(22):342–343, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
342.pdf.

Anonymous:1878:CHC

[12185] Anonymous. The crafty hermit crab. Scientific American, 39(22): 343–345, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
343.pdf.

Strong:1878:GSS

[12186] M. H. Strong. Gas a substitute for solid fuel. Scientific American, 39(22):346, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
346.pdf.

Anonymous:1878:Io

[12187] Anonymous. To inventors. Scientific American, 39(22):347, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n22/pdf/scientificamerican11301878-347.pdf.

Anonymous:1878:CRbi

[12188] Anonymous. Communications received. Scientific American, 39 (22):348, November 30, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n22/pdf/scientificamerican11301878-
348.pdf.

Anonymous:1878:BPcc

[12189] Anonymous. Business and personal. Scientific American, 39 (23):362, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
362a.pdf.

Anonymous:1878:BNBbc

[12190] Anonymous. Books: New books and publications. Scientific American, 39(23):362, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
362b.pdf.

Anonymous:1878:NQbq

[12191] Anonymous. Notes and queries. Scientific American, 39(23): 362–363, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
362c.pdf.

Anonymous:1878:IIbo

[12192] Anonymous. Index of inventions. Scientific American, 39(23):363, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-363a.pdf.

Anonymous:1878:PNPbk

[12193] Anonymous. Patent notes: Patents. Scientific American, 39 (23):363, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
363b.pdf.

Anonymous:1878:SME

[12194] Anonymous. The Sawyer–Man electric lamp. Scientific American, 39(23):351, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
351.pdf.

Anonymous:1878:UP

[12195] Anonymous. The utilization of petroleum. Scientific American, 39(23):352, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
352.pdf.

Anonymous:1878:TMD

[12196] Anonymous. A trade mark decision. Scientific American, 39 (23):352, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
352a.pdf.

Anonymous:1878:WCE

[12197] Anonymous. The wastage of carbon in electric lamps. Scientific American, 39(23):352, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
352b.pdf.

Anonymous:1878:VV

[12198] Anonymous. Ventilation of vessels. Scientific American, 39(23): 352–353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
352c.pdf.

Anonymous:1878:LCVb

[12199] Anonymous. The largest cash vault in the world. Scientific American, 39(23):353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
353.pdf.

Anonymous:1878:ITC

[12200] Anonymous. Increasing trade with Chili. Scientific American, 39 (23):353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
353a.pdf.

Anonymous:1878:ACM

[12201] Anonymous. American contributions to modern artillery. Scientific American, 39(23):353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
353b.pdf.

Anonymous:1878:FC

[12202] Anonymous. The fusing of carbon. Scientific American, 39(23):353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-353c.pdf.

Anonymous:1878:ABP

[12203] Anonymous. Alum in baking powders. Scientific American, 39 (23):353, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
353d.pdf.

Anonymous:1878:V

[12204] Anonymous. Valvoline. Scientific American, 39(23):356, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-356.pdf.

Robertson:1878:CP

[12205] D. D. Robertson. Car painting. Scientific American, 39(23):357, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-357.pdf.

Anonymous:1878:NTT

[12206] Anonymous. New trunnion and trimmer. Scientific American, 39(23):358, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
358.pdf.

Anonymous:1878:NTI

[12207] Anonymous. New time-indicating night lamp. Scientific American, 39(23):358, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
358a.pdf.

Anonymous:1878:IRb

[12208] Anonymous. An improved roofing. Scientific American, 39(23):358, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-358b.pdf.

Anonymous:1878:NLE

[12209] Anonymous. A new lamp extinguisher. Scientific American, 39 (23):358–359, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
358c.pdf.

Anonymous:1878:PEFb

[12210] Anonymous. Paris Exhibition — facade of Austria and Hungary. Scientific American, 39(23):359, December 7, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n23/pdf/scientificamerican12071878-359.pdf.

Anonymous:1878:CJb

[12211] Anonymous. The common jacana. Scientific American, 39(23): 359–360, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
359a.pdf.

Wright:1878:ANbl

[12212] Berlin H. Wright. Astronomical notes. Scientific American, 39 (23):361, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
361.pdf.

Anonymous:1878:AIF

[12213] Anonymous. American Institute Fair. Scientific American, 39 (23):361, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
361a.pdf.

Anonymous:1878:Iq

[12214] Anonymous. To inventors. Scientific American, 39(23):362, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n23/pdf/scientificamerican12071878-362.pdf.

Anonymous:1878:CRbj

[12215] Anonymous. Communications received. Scientific American, 39 (23):363, December 7, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n23/pdf/scientificamerican12071878-
363.pdf.

Anonymous:1878:Cbt

[12216] Anonymous. Correspondence. Scientific American, 39(24):376, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n24/pdf/scientificamerican12141878-376a.pdf.

Anonymous:1878:BPcd

[12217] Anonymous. Business and personal. Scientific American, 39 (24):378, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
378a.pdf.

Anonymous:1878:NQbr

[12218] Anonymous. Notes and queries. Scientific American, 39(24): 378–379, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
378b.pdf.

Anonymous:1878:IIbp

[12219] Anonymous. Index of inventions. Scientific American, 39(24): 379–380, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
379.pdf.

Anonymous:1878:PNPbl

[12220] Anonymous. Patent notes: Patents. Scientific American, 39 (24):380, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
380.pdf.

Anonymous:1878:BSC

[12221] Anonymous. The Belgian Ship Canal. Scientific American, 39 (24):367, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
367.pdf.

Anonymous:1878:PPf

[12222] Anonymous. Progress of petroleum. Scientific American, 39 (24):368, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
368.pdf.

Anonymous:1878:LW

[12223] Anonymous. The limit of work. Scientific American, 39(24):368, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n24/pdf/scientificamerican12141878-368a.pdf.

Anonymous:1878:IGa

[12224] Anonymous. Iridescent glass. Scientific American, 39(24):368, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n24/pdf/scientificamerican12141878-368b.pdf.

Anonymous:1878:RRA

[12225] Anonymous. Rails and railway accidents — New York Academy of Sciences. Scientific American, 39(24):368–369, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n24/pdf/scientificamerican12141878-368c.pdf.

Anonymous:1878:SSU

[12226] Anonymous. Sanitary science in the United States. Scientific American, 39(24):369–370, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
369.pdf.

Anonymous:1878:IV

[12227] Anonymous. An improved vise. Scientific American, 39(24): 370, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
370.pdf.

Anonymous:1878:NSFb

[12228] Anonymous. New shutter fastener. Scientific American, 39(24): 370, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
370a.pdf.

Anonymous:1878:IGS

[12229] Anonymous. An improved garden sprinkler. Scientific American, 39(24):370, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
370b.pdf.

Anonymous:1878:NFPb

[12230] Anonymous. A new foot power. Scientific American, 39(24): 370, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
370c.pdf.

Anonymous:1878:AMa

[12231] Anonymous. Amateur mechanics. Scientific American, 39(24): 371–373, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
371.pdf.

Anonymous:1878:WEL

[12232] Anonymous. The Werdermann electric light. Scientific American, 39(24):373, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
373.pdf.

Anonymous:1878:CSB

[12233] Anonymous. Cassons saw bench. Scientific American, 39(24): 374–375, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
374.pdf.

Anonymous:1878:APN

[12234] Anonymous. The argonaut, or paper nautilus. Scientific American, 39(24):375, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
375.pdf.

Anonymous:1878:TDS

[12235] Anonymous. The trap door spider of Jamaica. Scientific American, 39(24):375–376, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
375a.pdf.

Anonymous:1878:PSMb

[12236] Anonymous. The progress of science in Mexico. Scientific American, 39(24):376, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
376.pdf.

Wright:1878:ANbm

[12237] Berlin H. Wright. Astronomical notes. Scientific American, 39 (24):377, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
377.pdf.

Anonymous:1878:NWC

[12238] Anonymous. New wire clothing for burring cylinders. Scientific American, 39(24):377, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n24/pdf/scientificamerican12141878-
377a.pdf.

Anonymous:1878:Iu

[12239] Anonymous. To inventors. Scientific American, 39(24):378, December 14, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n24/pdf/scientificamerican12141878-378.pdf.

Anonymous:1878:Cbu

[12240] Anonymous. Correspondence. Scientific American, 39(25):388– 389, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
388c.pdf.

Anonymous:1878:BPce

[12241] Anonymous. Business and personal. Scientific American, 39 (25):394, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
394a.pdf.

Anonymous:1878:BNBbd

[12242] Anonymous. Books: New books and publications. Scientific American, 39(25):394, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
394b.pdf.

Anonymous:1878:NQbs

[12243] Anonymous. Notes and queries. Scientific American, 39(25): 394–395, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
394c.pdf.

Anonymous:1878:IIbq

[12244] Anonymous. Index of inventions. Scientific American, 39(25): 395, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
395a.pdf.

Anonymous:1878:NWE

[12245] Anonymous. A new water elevator. Scientific American, 39 (25):383, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
383.pdf.

Anonymous:1878:IDF

[12246] Anonymous. The international dairy fair. Scientific American, 39 (25):384, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
384.pdf.

Anonymous:1878:PDb

[12247] Anonymous. The Pomerania disaster. Scientific American, 39 (25):384, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
384a.pdf.

Anonymous:1878:SH

[12248] Anonymous. Short hand. Scientific American, 39(25):384, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n25/pdf/scientificamerican12211878-384b.pdf.

Anonymous:1878:ATF

[12249] Anonymous. Amalgam for tooth filling. Scientific American, 39 (25):385, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
385.pdf.

Anonymous:1878:AVM

[12250] Anonymous. An active volcano in the Moon. Scientific American, 39(25):385, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
385a.pdf.

Anonymous:1878:ES

[12251] Anonymous. Expiring subscriptions. Scientific American, 39(25): 385, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
385b.pdf.

Anonymous:1878:QSB

[12252] Anonymous. Quicklime as a substitute for blasting powder. Scientific American, 39(25):385, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
385c.pdf.

Anonymous:1878:NOSb

[12253] Anonymous. The new Otto silent gas engine. Scientific American, 39(25):386, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
386.pdf.

Anonymous:1878:DSCb

[12254] Anonymous. The diamond self-clamp paper cutter. Scientific American, 39(25):386, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
386a.pdf.

Anonymous:1878:AMD

[12255] Anonymous. Amateur mechanics. — drills and drilling. Scientific American, 39(25):387–388, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
387.pdf.

Anonymous:1878:FPc

[12256] Anonymous. The Fairbanks Prizes. Scientific American, 39(25): 388, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
388.pdf.

Anonymous:1878:NCN

[12257] Anonymous. New cutting nippers. Scientific American, 39(25): 388, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
388a.pdf.

Anonymous:1878:ATM

[12258] Anonymous. Another trade mark decision. Scientific American, 39(25):388, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
388b.pdf.

Mott:1878:PRS

[12259] Henry A. Mott, Jr. The purity of refined sugar. Scientific American, 39(25):389, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
389.pdf.

Anonymous:1878:CVH

[12260] Anonymous. Chapman’s valves and hydrants. Scientific American, 39(25):389, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
389a.pdf.

Anonymous:1878:NGE

[12261] Anonymous. A new gas engine. Scientific American, 39(25):390, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n25/pdf/scientificamerican12211878-390.pdf.

Anonymous:1878:NPV

[12262] Anonymous. A new parallel vise. Scientific American, 39(25): 390, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
390a.pdf.

Anonymous:1878:SRW

[12263] Anonymous. Some reasons why every manufacturer, mechanic, inventor, and artisan should become a patron of the Scientific American. Scientific American, 39(25):390, December 21, 1878. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v39/
n25/pdf/scientificamerican12211878-390b.pdf.

Anonymous:1878:HWT

[12264] Anonymous. A hairy water tortoise from China. Scientific American, 39(25):391, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
391.pdf.

Beard:1878:PFM

[12265] Daniel C. Beard. The paddle fish of the Mississippi. Scientific American, 39(25):391–392, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
391a.pdf.

Wright:1878:ANbn

[12266] Berlin H. Wright. Astronomical notes. Scientific American, 39 (25):392, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
392.pdf.

Anonymous:1878:PG

[12267] Anonymous. Pottery and glassware. Scientific American, 39 (25):393, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
393.pdf.

Anonymous:1878:Ix

[12268] Anonymous. To inventors. Scientific American, 39(25):394, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n25/pdf/scientificamerican12211878-394.pdf.

Anonymous:1878:CRbk

[12269] Anonymous. Communications received. Scientific American, 39 (25):395, December 21, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n25/pdf/scientificamerican12211878-
395.pdf.

Anonymous:1878:Cbv

[12270] Anonymous. Correspondence. Scientific American, 39(26):404, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-404.pdf.

Anonymous:1878:Iz

[12271] Anonymous. Index. Scientific American, 39(26):409–410, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-409.pdf.

Anonymous:1878:BPcf

[12272] Anonymous. Business and personal. Scientific American, 39 (26):411, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
411a.pdf.

Anonymous:1878:NQbt

[12273] Anonymous. Notes and queries. Scientific American, 39(26): 411, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
411b.pdf.

Anonymous:1878:IIbr

[12274] Anonymous. Index of inventions. Scientific American, 39(26): 412, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
412a.pdf.

Anonymous:1878:PNPbm

[12275] Anonymous. Patent notes: Patents. Scientific American, 39 (26):414, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
414.pdf.

Anonymous:1878:TEC

[12276] Anonymous. The trouble with English cottons. Scientific American, 39(26):399, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
399.pdf.

Anonymous:1878:CMT

[12277] Anonymous. A curious mode of taking turtle. Scientific American, 39(26):399, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
399a.pdf.

Anonymous:1878:Eb

[12278] Anonymous. The end of 1878. Scientific American, 39(26):400, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-400.pdf.

Anonymous:1878:PSd

[12279] Anonymous. Patent suits. Scientific American, 39(26):400, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-400a.pdf.

Anonymous:1878:SAEd

[12280] Anonymous. The Scientific American as an educator. Scientific American, 39(26):400–401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
400b.pdf.

Anonymous:1878:NO

[12281] Anonymous. The National Observatory. Scientific American, 39 (26):401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
401.pdf.

Anonymous:1878:ADI

[12282] Anonymous. An amendment to discourage invention. Scientific American, 39(26):401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
401a.pdf.

Anonymous:1878:TUB

[12283] Anonymous. The tunnel under the British channel. Scientific American, 39(26):401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
401b.pdf.

Anonymous:1878:PBa

[12284] Anonymous. Protection to banks. Scientific American, 39(26): 401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
401c.pdf.

Anonymous:1878:HCCb

[12285] Anonymous. The Hog Cholera Commission. Scientific American, 39(26):401, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
401d.pdf.

Anonymous:1878:EAa

[12286] Anonymous. The Eagle anvil. Scientific American, 39(26):402, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-402.pdf.

Anonymous:1878:FSW

[12287] Anonymous. Furnace for spelter and white of zinc. Scientific American, 39(26):402, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
402a.pdf.

Anonymous:1878:AMb

[12288] Anonymous. Animal mechanics. Scientific American, 39(26): 403–404, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
403.pdf.

Anonymous:1878:IBC

[12289] Anonymous. Improved bolt cutter. Scientific American, 39(26): 404, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
404a.pdf.

Wright:1878:ANbo

[12290] Berlin H. Wright. Astronomical notes. Scientific American, 39 (26):404–405, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
404b.pdf.

Anonymous:1878:NBS

[12291] Anonymous. A new bottle stopper. Scientific American, 39(26): 405–406, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
405.pdf.

Anonymous:1878:ICDa

[12292] Anonymous. Incline cutting, drawing, and stamping power press. Scientific American, 39(26):406, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
406.pdf.

Anonymous:1878:NIS

[12293] Anonymous. A new inserted saw tooth. Scientific American, 39 (26):406–407, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
406a.pdf.

Anonymous:1878:GHW

[12294] Anonymous. The great Hungarian wine cask. Scientific American, 39 (26):407–408, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
407.pdf.

Anonymous:1878:Ibb

[12295] Anonymous. To inventors. Scientific American, 39(26):411, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v39/n26/pdf/scientificamerican12281878-411.pdf.

Anonymous:1878:CRbl

[12296] Anonymous. Communications received. Scientific American, 39 (26):412, December 28, 1878. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v39/n26/pdf/scientificamerican12281878-
412.pdf.

Anonymous:1879:Ca

[12297] Anonymous. Correspondence. Scientific American, 40(1):7, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-7a.pdf.

Anonymous:1879:BPa

[12298] Anonymous. Business and personal. Scientific American, 40(1):11, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-11a.pdf.

Anonymous:1879:NQa

[12299] Anonymous. Notes and queries. Scientific American, 40(1):11–12, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-11b.pdf.

Anonymous:1879:IIa

[12300] Anonymous. Index of inventions. Scientific American, 40(1):12, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-12a.pdf.

Anonymous:1879:PNPa

[12301] Anonymous. Patent notes: Patents. Scientific American, 40(1):14, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-14.pdf.

Anonymous:1879:LBO

[12302] Anonymous. London, Berry & Orton’s band saw mill. Scientific American, 40(1):1, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
1.pdf.

Anonymous:1879:NCPa

[12303] Anonymous. A new Carpenter’s plane. Scientific American, 40(1):1, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-1a.pdf.

Anonymous:1879:CAP

[12304] Anonymous. Confiscation of American patents. Scientific American, 40(1):2, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
2.pdf.

Anonymous:1879:AHE

[12305] Anonymous. American history of the electric light. Scientific American, 40(1):2, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
2a.pdf.

Anonymous:1879:SCEa

[12306] Anonymous. The South as a competitor of England. Scientific American, 40(1):2, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
2b.pdf.

Anonymous:1879:WHR

[12307] Anonymous. William H. Rulofson. Scientific American, 40(1):2, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-2c.pdf.

Anonymous:1879:SIWa

[12308] Anonymous. A scientific invention wanted. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3.pdf.

Anonymous:1879:MN

[12309] Anonymous. Mining notes. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3a.pdf.

Anonymous:1879:Aa

[12310] Anonymous. Alloys. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3b.pdf.

Anonymous:1879:BEb

[12311] Anonymous. The breeding of eels. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3c.pdf.

Anonymous:1879:SUG

[12312] Anonymous. Sanitary uses of gunpowder. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3d.pdf.

Anonymous:1879:HMT

[12313] Anonymous. Hair on the musk turtle. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-3e.pdf.

Anonymous:1879:CMGa

[12314] Anonymous. The constitution of matter in the gaseous state. Scientific American, 40(1):3, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
3f.pdf.

Anonymous:1879:RELa

[12315] Anonymous. The Rapieff electric light. Scientific American, 40(1):4, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-4.pdf.

Anonymous:1879:NFL

[12316] Anonymous. A new foot lathe. Scientific American, 40(1):5, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-5.pdf.

Anonymous:1879:NBP

[12317] Anonymous. A new bench plane. Scientific American, 40(1):5–6, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-5a.pdf.

Anonymous:1879:IMC

[12318] Anonymous. Improved musical condenser. Scientific American, 40(1):6–7, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n1/pdf/scientificamerican01041879-
6.pdf.

Anonymous:1879:DSPa

[12319] Anonymous. Double screw parallel vise. Scientific American, 40(1):7, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-7.pdf.

Anonymous:1879:PEa

[12320] Anonymous. Paris Exhibition. Scientific American, 40(1):8–9, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-8.pdf.

Anonymous:1879:VBW

[12321] Anonymous. Vienna bronze work. Scientific American, 40(1):10, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-10.pdf.

Anonymous:1879:Ia

[12322] Anonymous. To inventors. Scientific American, 40(1):11, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-11.pdf.

Anonymous:1879:CRa

[12323] Anonymous. Communications received. Scientific American, 40(1):12, January 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n1/pdf/scientificamerican01041879-12.pdf.

Anonymous:1879:Cb

[12324] Anonymous. Correspondence. Scientific American, 40(2):21, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-21.pdf.

Anonymous:1879:BPb

[12325] Anonymous. Business and personal. Scientific American, 40(2):27, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-27a.pdf.

Anonymous:1879:NQb

[12326] Anonymous. Notes and queries. Scientific American, 40(2):27, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-27b.pdf.

Anonymous:1879:IIb

[12327] Anonymous. Index of inventions. Scientific American, 40(2):27–28, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-27d.pdf.

Anonymous:1879:PNPb

[12328] Anonymous. Patent notes: Patents. Scientific American, 40(2):30, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-30.pdf.

Anonymous:1879:RWT

[12329] Anonymous. A Remedy worse than the disease. Scientific American, 40(2):16, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
16.pdf.

Anonymous:1879:PCO

[12330] Anonymous. A plan of co-operative charity. Scientific American, 40(2):16, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
16a.pdf.

Anonymous:1879:MHMa

[12331] Anonymous. Miss Hosmer’s motor. Scientific American, 40(2):16, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-16b.pdf.

Anonymous:1879:FELa

[12332] Anonymous. The first electric lamps. Scientific American, 40 (2):17, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
17.pdf.

Wicks:1879:AINa

[12333] Hamilton S. Wicks. American industries — no. 1. Scientific American, 40(2):17, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
17a.pdf.

Anonymous:1879:POP

[12334] Anonymous. Patent Office practice. Scientific American, 40(2):17, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-17b.pdf.

Anonymous:1879:CWW

[12335] Anonymous. Chichester Water Works. Scientific American, 40 (2):18, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
18.pdf.

Anonymous:1879:NRI

[12336] Anonymous. A new revolving index. Scientific American, 40(2):18, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-18a.pdf.

Anonymous:1879:SHL

[12337] Anonymous. Sack holders and lifters at the exhibition. Scientific American, 40(2):19–20, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
19.pdf.

Anonymous:1879:NCa

[12338] Anonymous. A novel cooler. Scientific American, 40(2):22, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-22.pdf.

Anonymous:1879:CDF

[12339] Anonymous. The Casson–Dormoy furnace. Scientific American, 40(2):22, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
22a.pdf.

Anonymous:1879:TLW

[12340] Anonymous. A three legged woodcock. Scientific American, 40 (2):23, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
23.pdf.

Anonymous:1879:SBA

[12341] Anonymous. Studies from the Brighton Aquarium. Scientific American, 40(2):23–24, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
23a.pdf.

Anonymous:1879:BLF

[12342] Anonymous. Bronze lattice or fire guard. Scientific American, 40(2):24, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
24.pdf.

Trowrridge:1879:EI

[12343] John Trowrridge. Electric illumination. Scientific American, 40 (2):25, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
25.pdf.

Anonymous:1879:MSFa

[12344] Anonymous. The manufacture of soft felt hats. Scientific American, 40(2):26, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
26.pdf.

Anonymous:1879:Ib

[12345] Anonymous. To inventors. Scientific American, 40(2):27, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n2/pdf/scientificamerican01111879-27.pdf.

Anonymous:1879:CRb

[12346] Anonymous. Communications received. Scientific American, 40 (2):27, January 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n2/pdf/scientificamerican01111879-
27c.pdf.

Anonymous:1879:IRI

[12347] Anonymous. Inventions: Recent inventions. Scientific American, 40(3):38, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
38.pdf.

Anonymous:1879:Cc

[12348] Anonymous. Correspondence. Scientific American, 40(3):38, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-38a.pdf.

Anonymous:1879:BPc

[12349] Anonymous. Business and personal. Scientific American, 40(3):43, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-43a.pdf.

Anonymous:1879:BNBa

[12350] Anonymous. Books: New books and publications. Scientific American, 40(3):43, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
43b.pdf.

Anonymous:1879:NQc

[12351] Anonymous. Notes and queries. Scientific American, 40(3):43–44, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-43c.pdf.

Anonymous:1879:IIc

[12352] Anonymous. Index of inventions. Scientific American, 40(3):44, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-44.pdf.

Anonymous:1879:PNPc

[12353] Anonymous. Patent notes: Patents. Scientific American, 40(3):46, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-46.pdf.

Anonymous:1879:ASS

[12354] Anonymous. The adulteration of soft soap. Scientific American, 40(3):31, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
31.pdf.

Wicks:1879:AINb

[12355] Hamilton S. Wicks. American industries — no. 2. Scientific American, 40(3):31, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
31a.pdf.

Anonymous:1879:DPC

[12356] Anonymous. Over-density of population in cities. Scientific American, 40(3):32, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
32.pdf.

Anonymous:1879:PNS

[12357] Anonymous. Patents in New South Wales. Scientific American, 40(3):32, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
32a.pdf.

Anonymous:1879:SSCa

[12358] Anonymous. Sun spots and commercial crises. Scientific American, 40(3):32–33, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
32b.pdf.

Anonymous:1879:PSAa

[12359] Anonymous. Progress of the Swedish Arctic Expedition. Scientific American, 40(3):33, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
33.pdf.

Anonymous:1879:TQ

[12360] Anonymous. The trademarks quandary. Scientific American, 40 (3):33, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
33a.pdf.

Anonymous:1879:HUO

[12361] Anonymous. How to utilize old fruit cans. Scientific American, 40(3):33, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
33b.pdf.

Anonymous:1879:SBS

[12362] Anonymous. Senate Bill 300. — section 2. Scientific American, 40(3):33, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
33c.pdf.

Anonymous:1879:BFL

[12363] Anonymous. Boiler feeders for locomotives, etc. Scientific American, 40(3):33–34, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
33d.pdf.

Anonymous:1879:SSG

[12364] Anonymous. Straub’s scientific grain mill. Scientific American, 40(3):34, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
34.pdf.

Anonymous:1879:CBa

[12365] Anonymous. The Columbia bicycle. Scientific American, 40(3):35, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-35.pdf.

Hopkins:1879:EMP

[12366] M. Hopkins. Easily made physical apparatus. Scientific American, 40(3):35–36, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
35a.pdf.

Hovey:1879:LCa

[12367] H. C. Hovey. The Luray Cavern. Scientific American, 40(3):36–37, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-36.pdf.

Anonymous:1879:DSPb

[12368] Anonymous. Duplex steam pumping engines. Scientific American, 40(3):37, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
37.pdf.

Anonymous:1879:CIa

[12369] Anonymous. The Clipper injector. Scientific American, 40(3):37, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-37a.pdf.

Anonymous:1879:Cd

[12370] Anonymous. The chætodon. Scientific American, 40(3):39–40, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-39.pdf.

Anonymous:1879:EES

[12371] Anonymous. Early English style of piano at the Paris Exhibition. Scientific American, 40(3):40–41, January 18, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n3/pdf/scientificamerican01181879-40.pdf.

Anonymous:1879:SDK

[12372] Anonymous. Spreading diphtheria by kisses. Scientific American, 40(3):41, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
41.pdf.

Anonymous:1879:NMS

[12373] Anonymous. A new method of setting boiler tubes. Scientific American, 40(3):41, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n3/pdf/scientificamerican01181879-
41a.pdf.

Anonymous:1879:GGG

[12374] Anonymous. The Gruey gyroscope and the gyroscopic pendulum. Scientific American, 40(3):41–42, January 18, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n3/pdf/scientificamerican01181879-41b.pdf.

Anonymous:1879:Ic

[12375] Anonymous. To inventors. Scientific American, 40(3):43, January 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n3/pdf/scientificamerican01181879-43.pdf.

Anonymous:1879:Ce

[12376] Anonymous. Correspondence. Scientific American, 40(4):53, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-53b.pdf.

Anonymous:1879:BPd

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-59a.pdf.

Anonymous:1879:BNBb

[12378] Anonymous. Books: New books and publications. Scientific American, 40(4):59, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
59b.pdf.

Anonymous:1879:NQd

[12379] Anonymous. Notes and queries. Scientific American, 40(4):59–60, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-59c.pdf.

Anonymous:1879:IId

[12380] Anonymous. Index of inventions. Scientific American, 40(4):60, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-60a.pdf.

Anonymous:1879:PNPd

[12381] Anonymous. Patent notes: Patents. Scientific American, 40(4):62, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-62.pdf.

Anonymous:1879:TE

[12382] Anonymous. The transmutation of elements. Scientific American, 40(4):48, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
48.pdf.

Anonymous:1879:FSM

[12383] Anonymous. The fourth state of matter. Scientific American, 40(4):48, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
48a.pdf.

Wicks:1879:AINc

[12384] Hamilton S. Wicks. American industries. — no. 3. Scientific American, 40(4):48–49, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
48b.pdf.

Anonymous:1879:BDI

[12385] Anonymous. The bill to discourage inventions. Scientific American, 40(4):49, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
49.pdf.

Anonymous:1879:IT

[12386] Anonymous. An improved table. Scientific American, 40(4):50–51, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-50.pdf.

Anonymous:1879:SRMa

[12387] Anonymous. Saturator and regulator of Mekarski’s compressed air motor. Scientific American, 40(4):51, January 25, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n4/pdf/scientificamerican01251879-51.pdf.

Anonymous:1879:RLP

[12388] Anonymous. Rotary lift pump. Scientific American, 40(4):51–52, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-51a.pdf.

Anonymous:1879:AMa

[12389] Anonymous. Amateur mechanics. Scientific American, 40(4): 52–53, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
52.pdf.

Anonymous:1879:NBBa

[12390] Anonymous. A novel bridle bit. Scientific American, 40(4):53, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-53.pdf.

Anonymous:1879:NCC

[12391] Anonymous. A new continuous calk horseshoe. Scientific American, 40(4):53, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
53a.pdf.

Anonymous:1879:WFE

[12392] Anonymous. The Wallace–Farmer electric light. Scientific American, 40(4):54, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
54.pdf.

Anonymous:1879:EGT

[12393] Anonymous. Excavator on the Ghent and Terneu-Zen Ship Canal, Belgium. Scientific American, 40(4):54–55, January 25, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n4/pdf/scientificamerican01251879-54a.pdf.

Anonymous:1879:PEH

[12394] Anonymous. The penciled ear hog. Scientific American, 40(4): 55–56, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
55.pdf.

Anonymous:1879:NF

[12395] Anonymous. A new firearm. Scientific American, 40(4):56, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-56.pdf.

Cooley:1879:NFS

[12396] R. C. Cooley. A new form of sulphureted hydrogen apparatus. Scientific American, 40(4):56, January 25, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n4/pdf/scientificamerican01251879-56a.pdf.

Anonymous:1879:EF

[12397] Anonymous. Elizabethan fireplace. Scientific American, 40(4): 57–58, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
57.pdf.

Anonymous:1879:Id

[12398] Anonymous. To inventors. Scientific American, 40(4):59, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n4/pdf/scientificamerican01251879-59.pdf.

Anonymous:1879:CRc

[12399] Anonymous. Communications received. Scientific American, 40 (4):60, January 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n4/pdf/scientificamerican01251879-
60.pdf.

Anonymous:1879:Cf

[12400] Anonymous. Correspondence. Scientific American, 40(5):68, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-68.pdf.

Anonymous:1879:BPe

[12401] Anonymous. Business and personal. Scientific American, 40 (5):75, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
75a.pdf.

Anonymous:1879:NQe

[12402] Anonymous. Notes and queries. Scientific American, 40(5):75, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-75b.pdf.

Anonymous:1879:IIe

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-75d.pdf.

Anonymous:1879:PNPe

[12404] Anonymous. Patent notes: Patents. Scientific American, 40 (5):77, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
78.pdf.

Anonymous:1879:VP

[12405] Anonymous. The Voltaic pencil. Scientific American, 40(5):63, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-63.pdf.

Anonymous:1879:ETV

[12406] Anonymous. Engine of the torpedo vessel destroyer. Scientific American, 40(5):63, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
63a.pdf.

Anonymous:1879:NSPa

[12407] Anonymous. A new switch pedal. Scientific American, 40(5):63, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-63b.pdf.

Anonymous:1879:SPS

[12408] Anonymous. Some patents that section 11 would have killed. Scientific American, 40(5):64, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
64.pdf.

Anonymous:1879:GH

[12409] Anonymous. Glucose honey. Scientific American, 40(5):64, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-64a.pdf.

Anonymous:1879:VGC

[12410] Anonymous. A vegetable Green for confectioners. Scientific American, 40(5):64, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
64b.pdf.

Anonymous:1879:NI

[12411] Anonymous. A new industry. Scientific American, 40(5):64–65, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-64c.pdf.

Anonymous:1879:LNY

[12412] Anonymous. The laundries of New York. Scientific American, 40(5):65, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
65.pdf.

Anonymous:1879:HE

[12413] Anonymous. Haeckel on evolution. Scientific American, 40(5): 65, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
65a.pdf.

Anonymous:1879:GMO

[12414] Anonymous. A great market for our cattle and other produce. Scientific American, 40(5):65, February 1, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n5/pdf/scientificamerican02011879-65b.pdf.

Anonymous:1879:HIa

[12415] Anonymous. A hint for an invention. Scientific American, 40 (5):65, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
65c.pdf.

Anonymous:1879:DEB

[12416] Anonymous. The demenge engine and boiler. Scientific American, 40(5):66, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
66.pdf.

Anonymous:1879:NPDa

[12417] Anonymous. A new potato digger. Scientific American, 40(5):67, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-67.pdf.

Anonymous:1879:ML

[12418] Anonymous. The manufacture of lard. Scientific American, 40 (5):67–68, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
67a.pdf.

Stillman:1879:AMG

[12419] W. J. Stillman. American mechanical genius. Scientific American, 40(5):69, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
69.pdf.

Anonymous:1879:AAV

[12420] Anonymous. Atmospheric air and the vacuum. Scientific American, 40(5):70, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
70.pdf.

Anonymous:1879:Pa

[12421] Anonymous. The porcupine. Scientific American, 40(5):71–72, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-71.pdf.

Anonymous:1879:IWA

[12422] Anonymous. Improved window awning. Scientific American, 40 (5):72–73, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
72.pdf.

Anonymous:1879:DV

[12423] Anonymous. Designs for vases. Scientific American, 40(5):73–74, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-73.pdf.

Anonymous:1879:Ie

[12424] Anonymous. To inventors. Scientific American, 40(5):75, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n5/pdf/scientificamerican02011879-75.pdf.

Anonymous:1879:CRd

[12425] Anonymous. Communications received. Scientific American, 40 (5):75, February 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n5/pdf/scientificamerican02011879-
75c.pdf.

Anonymous:1879:Cg

[12426] Anonymous. Correspondence. Scientific American, 40(6):83–85, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-83a.pdf.

Anonymous:1879:BPf

[12427] Anonymous. Business and personal. Scientific American, 40 (6):91, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
91a.pdf.

Anonymous:1879:BNBc

[12428] Anonymous. Books: New books and publications. Scientific American, 40(6):91, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
91b.pdf.

Anonymous:1879:NQf

[12429] Anonymous. Notes and queries. Scientific American, 40(6):91– 92, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
91c.pdf.

Anonymous:1879:IIf

[12430] Anonymous. Index of inventions. Scientific American, 40(6):92, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-92a.pdf.

Anonymous:1879:PNPf

[12431] Anonymous. Patent notes: Patents. Scientific American, 40 (6):94, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
94.pdf.

Wicks:1879:AINd

[12432] Hamilton S. Wicks. American industries. — no. 4. Scientific American, 40(6):79–80, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
79.pdf.

Anonymous:1879:KP

[12433] Anonymous. Keep to the point. Scientific American, 40(6):80, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-80.pdf.

Anonymous:1879:EPT

[12434] Anonymous. The establishment of public town libraries. Scientific American, 40(6):80–81, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
80a.pdf.

Anonymous:1879:FCP

[12435] Anonymous. Forest culture pays. Scientific American, 40(6):81, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-81.pdf.

Anonymous:1879:HRP

[12436] Anonymous. Harbor of refuge on the Pacific coast. Scientific American, 40(6):81, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
81a.pdf.

Anonymous:1879:SSCb

[12437] Anonymous. The social science convention. Scientific American, 40(6):81, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
81b.pdf.

Anonymous:1879:SDG

[12438] Anonymous. Settlement of a doubtful geological point. Scientific American, 40(6):81–82, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
81c.pdf.

Anonymous:1879:NSSa

[12439] Anonymous. A new squaring shear. Scientific American, 40 (6):82, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
82.pdf.

Anonymous:1879:CPT

[12440] Anonymous. Clay pipes and their manufacture. Scientific American, 40(6):82–83, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
82a.pdf.

Anonymous:1879:NWS

[12441] Anonymous. The new Woodruff Scientific Expedition. Scientific American, 40(6):83, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
83.pdf.

Anonymous:1879:NO

[12442] Anonymous. A novel oiler. Scientific American, 40(6):85, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-85.pdf.

Anonymous:1879:NBS

[12443] Anonymous. A novel boot and shoe heel. Scientific American, 40(6):86, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
86.pdf.

Anonymous:1879:ILS

[12444] Anonymous. An improved locomotive smoke stack. Scientific American, 40(6):86, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
86a.pdf.

Anonymous:1879:CIb

[12445] Anonymous. A curious invention. Scientific American, 40(6):86, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-86b.pdf.

Anonymous:1879:NCSa

[12446] Anonymous. A new car stove. Scientific American, 40(6):86, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-86c.pdf.

Anonymous:1879:GSL

[12447] Anonymous. The great shielded locust of papua. Scientific American, 40(6):87–88, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
87.pdf.

Anonymous:1879:IPF

[12448] Anonymous. Improved power feed sandpapering machine. Scientific American, 40(6):89, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
89.pdf.

Anonymous:1879:HRI

[12449] Anonymous. Hard rolled iron and steel Rim pulleys. Scientific American, 40(6):89–90, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
89a.pdf.

Anonymous:1879:If

[12450] Anonymous. To inventors. Scientific American, 40(6):91, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n6/pdf/scientificamerican02081879-91.pdf.

Anonymous:1879:CRe

[12451] Anonymous. Communications received. Scientific American, 40 (6):92, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
92.pdf.

Anonymous:1879:COP

[12452] Anonymous. Can I obtain a patent? Scientific American, 40 (6):92, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
92b.pdf.

Anonymous:1879:PSAb

[12453] Anonymous. Prospectus of the Scientific American for 1879. Scientific American, 40(6):92, February 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n6/pdf/scientificamerican02081879-
92c.pdf.

Anonymous:1879:Ch

[12454] Anonymous. Correspondence. Scientific American, 40(7):101– 102, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
101.pdf.

Anonymous:1879:BPg

[12455] Anonymous. Business and personal. Scientific American, 40(7): 107, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
107a.pdf.

Anonymous:1879:NQg

[12456] Anonymous. Notes and queries. Scientific American, 40(7):107, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-107b.pdf.

Anonymous:1879:IIg

[12457] Anonymous. Index of inventions. Scientific American, 40(7):108, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-108a.pdf.

Anonymous:1879:NPBa

[12458] Anonymous. The new Patent Bill as passed by the Senate. Scientific American, 40(7):96, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
96.pdf.

Anonymous:1879:WBT

[12459] Anonymous. Will blood tell? Scientific American, 40(7):96, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-96a.pdf.

Anonymous:1879:ATa

[12460] Anonymous. American tea. Scientific American, 40(7):96, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-96b.pdf.

Anonymous:1879:LFM

[12461] Anonymous. The license fee as a measure of damages. Scientific American, 40(7):97, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
97.pdf.

Anonymous:1879:AINa

[12462] Anonymous. American industries. — no. 5. Scientific American, 40(7):97, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
97a.pdf.

Anonymous:1879:FELb

[12463] Anonymous. A floating electric light. Scientific American, 40 (7):97–98, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
97b.pdf.

Anonymous:1879:NRTa

[12464] Anonymous. A novel railway trace. Scientific American, 40(7): 98, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
98.pdf.

Anonymous:1879:ITW

[12465] Anonymous. An improved type writes. Scientific American, 40 (7):98, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
98a.pdf.

Anonymous:1879:HIb

[12466] Anonymous. The Hancock inspirator. Scientific American, 40 (7):99, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
99.pdf.

Anonymous:1879:NCRa

[12467] Anonymous. A new check row corn planter and drill. Scientific American, 40(7):99–100, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
99a.pdf.

Anonymous:1879:AMb

[12468] Anonymous. Amateur mechanics. Scientific American, 40(7): 100, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
100.pdf.

Anonymous:1879:MSFb

[12469] Anonymous. The manufacture of sewing and floss silk. Scientific American, 40(7):100–101, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
100a.pdf.

Anonymous:1879:APR

[12470] Anonymous. Aveling & Porter’s road and farm locomotive. Scientific American, 40(7):102, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
102.pdf.

Anonymous:1879:FWI

[12471] Anonymous. Fire and water indicator. Scientific American, 40 (7):102, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
102a.pdf.

Anonymous:1879:OBA

[12472] Anonymous. The octopus at the Berlin aquarium. Scientific American, 40(7):103, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
103.pdf.

Anonymous:1879:BSB

[12473] Anonymous. Barrel for shipping bottled liquors. Scientific American, 40(7):103–105, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
103a.pdf.

Anonymous:1879:CEa

[12474] Anonymous. Chest of ebony. Scientific American, 40(7):105–106, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-105.pdf.

Anonymous:1879:Ig

[12475] Anonymous. To inventors. Scientific American, 40(7):107, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n7/pdf/scientificamerican02151879-107.pdf.

Anonymous:1879:CRf

[12476] Anonymous. Communications received. Scientific American, 40 (7):108, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
108.pdf.

Anonymous:1879:VPH

[12477] Anonymous. Value of patents and how to obtain them. Scientific American, 40(7):108, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
108b.pdf.

Anonymous:1879:PSAc

[12478] Anonymous. Prospective of the Scientific American for 1879. Scientific American, 40(7):110, February 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n7/pdf/scientificamerican02151879-
110.pdf.

Anonymous:1879:BPh

[12479] Anonymous. Business and personal. Scientific American, 40(8): 123, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
123a.pdf.

Anonymous:1879:NQh

[12480] Anonymous. Notes and queries. Scientific American, 40(8):123– 124, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
123b.pdf.

Anonymous:1879:IIh

[12481] Anonymous. Index of inventions. Scientific American, 40(8):124, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n8/pdf/scientificamerican02221879-124a.pdf.

Anonymous:1879:PNPg

[12482] Anonymous. Patent notes: Patents. Scientific American, 40 (8):126, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
126.pdf.

Anonymous:1879:AINb

[12483] Anonymous. American industries. — no. 6. Scientific American, 40(8):111, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
111.pdf.

Anonymous:1879:NPBb

[12484] Anonymous. The new patent bill. — shall it pass the house. Scientific American, 40(8):112, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
112.pdf.

Anonymous:1879:CWP

[12485] Anonymous. Creosoted wood as a protection against teredoes. Scientific American, 40(8):112, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
112a.pdf.

Anonymous:1879:SAI

[12486] Anonymous. Seating on artificial ice. Scientific American, 40 (8):112, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
112b.pdf.

Anonymous:1879:RELb

[12487] Anonymous. Recent experiments with “laughing gas”. Scientific American, 40(8):113, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
113.pdf.

Anonymous:1879:POD

[12488] Anonymous. A paper observatory dome. Scientific American, 40(8):113, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
113a.pdf.

Anonymous:1879:WPM

[12489] Anonymous. The Woodbury planing machine case. Scientific American, 40(8):113, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
113b.pdf.

Anonymous:1879:NIM

[12490] Anonymous. A new and important mineral. Scientific American, 40(8):113, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
113c.pdf.

Anonymous:1879:RAPa

[12491] Anonymous. Recent American patents. Scientific American, 40 (8):115, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
115.pdf.

Anonymous:1879:NIC

[12492] Anonymous. A new ice creeper. Scientific American, 40(8):115, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n8/pdf/scientificamerican02221879-115a.pdf.

Anonymous:1879:IFC

[12493] Anonymous. An improved fruit can. Scientific American, 40 (8):115, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
115b.pdf.

Anonymous:1879:INY

[12494] Anonymous. An improvement in neck yokes. Scientific American, 40(8):115–116, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
115c.pdf.

Anonymous:1879:NSW

[12495] Anonymous. A new steam wagon. Scientific American, 40(8): 116–117, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
116.pdf.

Anonymous:1879:SMH

[12496] Anonymous. Stop motion hand piece for dental engines. Scientific American, 40(8):117, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
117.pdf.

Anonymous:1879:IPRa

[12497] Anonymous. An improved parallel ruler. Scientific American, 40(8):118, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
118.pdf.

Anonymous:1879:CCC

[12498] Anonymous. A combined chopper and cultivator. Scientific American, 40(8):118, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
118a.pdf.

Anonymous:1879:NWH

[12499] Anonymous. A new whiffletree hook. Scientific American, 40 (8):118, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
118b.pdf.

Anonymous:1879:MIa

[12500] Anonymous. Miscellaneous inventions. Scientific American, 40 (8):118, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
118c.pdf.

Beard:1879:GCA

[12501] Daniel C. Beard. Green Carolina anolis; or, the American chameleon. Scientific American, 40(8):119, February 22, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n8/pdf/scientificamerican02221879-119.pdf.

Anonymous:1879:NTB

[12502] Anonymous. A new torpedo boat. Scientific American, 40(8): 119–120, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
119a.pdf.

Anonymous:1879:WTa

[12503] Anonymous. A wonderful tree. Scientific American, 40(8):121, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n8/pdf/scientificamerican02221879-121.pdf.

Anonymous:1879:SEP

[12504] Anonymous. A simple electric pen. Scientific American, 40(8): 121–122, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
121a.pdf.

Anonymous:1879:Ih

[12505] Anonymous. To inventors. Scientific American, 40(8):123, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n8/pdf/scientificamerican02221879-123.pdf.

Anonymous:1879:CRg

[12506] Anonymous. Communications received. Scientific American, 40 (8):124, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
124.pdf.

Anonymous:1879:PSAd

[12507] Anonymous. Prospectus of the Scientific American for 1879. Scientific American, 40(8):124, February 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n8/pdf/scientificamerican02221879-
124b.pdf.

Anonymous:1879:Cj

[12508] Anonymous. Correspondence. Scientific American, 40(9):136–137, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-136.pdf.

Anonymous:1879:BPi

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-139a.pdf.

Anonymous:1879:BNBd

[12510] Anonymous. Books: New books and publications. Scientific American, 40(9):139, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
139b.pdf.

Anonymous:1879:NQi

[12511] Anonymous. Notes and queries. Scientific American, 40(9):139–140, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-139c.pdf.

Anonymous:1879:IIi

[12512] Anonymous. Index of inventions. Scientific American, 40(9):140, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-140a.pdf.

Anonymous:1879:PNPh

[12513] Anonymous. Patent notes: Patents. Scientific American, 40(9):142, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-142.pdf.

Anonymous:1879:AINc

[12514] Anonymous. American industries. — no. 7. Scientific American, 40(9):127, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
127.pdf.

Anonymous:1879:SLL

[12515] Anonymous. Strikes in Liverpool and London. Scientific American, 40(9):128, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
128.pdf.

Anonymous:1879:HPCa

[12516] Anonymous. Honor to Peter Cooper. Scientific American, 40(9):128, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-128a.pdf.

Anonymous:1879:SB

[12517] Anonymous. Samuel Batchelder. Scientific American, 40(9):128, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-128b.pdf.

Anonymous:1879:PR

[12518] Anonymous. The plague in Russia. Scientific American, 40(9):128–129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-128c.pdf.

Anonymous:1879:FAa

[12519] Anonymous. Food adulteration. Scientific American, 40(9):129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-129.pdf.

Anonymous:1879:FBF

[12520] Anonymous. The farmer’s best friend. Scientific American, 40(9):129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-129a.pdf.

Anonymous:1879:CJa

[12521] Anonymous. A cable jubilee. Scientific American, 40(9):129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-129b.pdf.

Anonymous:1879:TAE

[12522] Anonymous. Trade arbitration in England. Scientific American, 40(9):129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
129c.pdf.

Anonymous:1879:ICM

[12523] Anonymous. An improved cane mill wanted. Scientific American, 40(9):129, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
129d.pdf.

Anonymous:1879:PRC

[12524] Anonymous. Pneumatic regulator for clocks. Scientific American, 40(9):130–131, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
130.pdf.

Anonymous:1879:NGT

[12525] Anonymous. A new gun tool. Scientific American, 40(9):131, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-131.pdf.

Anonymous:1879:NDC

[12526] Anonymous. A novel door check. Scientific American, 40(9):131, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-131a.pdf.

Anonymous:1879:NPBc

[12527] Anonymous. A new protector for book covers. Scientific American, 40(9):131, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
131b.pdf.

Anonymous:1879:IFG

[12528] Anonymous. An improved file guard. Scientific American, 40(9):131, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-131c.pdf.

Anonymous:1879:REP

[12529] Anonymous. Recent engineering patents. Scientific American, 40 (9):131–132, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
131d.pdf.

Anonymous:1879:NRD

[12530] Anonymous. A new ratchet drill. Scientific American, 40(9):132–133, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-132.pdf.

Anonymous:1879:PPH

[12531] Anonymous. Patents pertaining to the household. Scientific American, 40(9):133, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
133.pdf.

Anonymous:1879:NBMa

[12532] Anonymous. A new beetling machine. Scientific American, 40(9):134, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-134.pdf.

Anonymous:1879:NCRb

[12533] Anonymous. A new cream-raising apparatus. Scientific American, 40(9):134, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
134a.pdf.

Anonymous:1879:MIb

[12534] Anonymous. Miscellaneous intentions. Scientific American, 40(9):134, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-134b.pdf.

Anonymous:1879:OEO

[12535] Anonymous. The oriental eagle owl. Scientific American, 40(9):135, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-135.pdf.

Anonymous:1879:HW

[12536] Anonymous. The humpback whale. Scientific American, 40(9):135–136, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-135a.pdf.

Anonymous:1879:AV

[12537] Anonymous. The Alhambra vase. Scientific American, 40(9):137, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-137.pdf.

Anonymous:1879:EL

[12538] Anonymous. Electric lighting. Scientific American, 40(9):138, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-138.pdf.

Anonymous:1879:AWP

[12539] Anonymous. Ancient works of Peru. Scientific American, 40(9):138, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-138a.pdf.

Anonymous:1879:IGW

[12540] Anonymous. To inventors. — the great want. Scientific American, 40(9):138, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n9/pdf/scientificamerican03011879-
138b.pdf.

Anonymous:1879:Ij

[12541] Anonymous. To inventors. Scientific American, 40(9):139, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-139.pdf.

Anonymous:1879:CRh

[12542] Anonymous. Communications received. Scientific American, 40(9):140, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-140.pdf.

Anonymous:1879:SAa

[12543] Anonymous. The Scientific American. Scientific American, 40(9):140, March 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n9/pdf/scientificamerican03011879-140b.pdf.

Anonymous:1879:RPI

[12544] Anonymous. Recently patented inventions. Scientific American, 40(10):150, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
150.pdf.

Anonymous:1879:BPj

[12545] Anonymous. Business and personal. Scientific American, 40(10):155, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-155a.pdf.

Anonymous:1879:BNBe

[12546] Anonymous. Books: New books and publications. Scientific American, 40(10):155, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
155b.pdf.

Anonymous:1879:NQj

[12547] Anonymous. Notes and queries. Scientific American, 40(10):155–156, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-155c.pdf.

Anonymous:1879:IIj

[12548] Anonymous. Index of inventions. Scientific American, 40(10):156, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-156a.pdf.

Anonymous:1879:PNPi

[12549] Anonymous. Patent notes: Patents. Scientific American, 40(10):158, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-158.pdf.

Anonymous:1879:SSW

[12550] Anonymous. The smallest steamboat in the world. Scientific American, 40(10):143, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
143.pdf.

Anonymous:1879:GMa

[12551] Anonymous. Gary’s motor. Scientific American, 40(10):144, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-144.pdf.

Anonymous:1879:SPBa

[12552] Anonymous. Senate Patent Bill No. 300. — shall it pass the House of Representatives? Scientific American, 40(10):144–145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-144a.pdf.

Anonymous:1879:SME

[12553] Anonymous. The Sawyer–Man electric light. Scientific American, 40(10):145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
145.pdf.

Anonymous:1879:NFCa

[12554] Anonymous. A new form of carbon. Scientific American, 40(10):145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-145a.pdf.

Anonymous:1879:GMMa

[12555] Anonymous. The gart magnetic-motor deception. Scientific American, 40(10):145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
145b.pdf.

Anonymous:1879:PFa

[12556] Anonymous. The preservation of forests. Scientific American, 40(10):145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
145c.pdf.

Anonymous:1879:TWB

[12557] Anonymous. The trouble with wire binders. Scientific American, 40(10):145, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
145d.pdf.

Anonymous:1879:WF

[12558] Anonymous. The waste of fire. Scientific American, 40(10):145–146, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-145e.pdf.

Anonymous:1879:SPM

[12559] Anonymous. The sweepstakes planer and matcher. Scientific American, 40(10):147, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
147.pdf.

Anonymous:1879:NDF

[12560] Anonymous. Nicolaieff depositing floating dock. Scientific American, 40(10):147–148, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
147a.pdf.

Anonymous:1879:WWF

[12561] Anonymous. What we find in honey. Scientific American, 40(10):149, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-149.pdf.

Anonymous:1879:MSSa

[12562] Anonymous. Machine for softening silk. Scientific American, 40 (10):150, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
150a.pdf.

Anonymous:1879:GMMb

[12563] Anonymous. Gary’s magnetic motor. Scientific American, 40 (10):151–153, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
151.pdf.

Hovey:1879:LCb

[12564] H. C. Hovey. The Luray Cavern. Scientific American, 40(10):153–154, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-153.pdf.

Anonymous:1879:SRIa

[12565] Anonymous. Some recent inventions. Scientific American, 40(10):154, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-154.pdf.

Anonymous:1879:Ik

[12566] Anonymous. To inventors. Scientific American, 40(10):155, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n10/pdf/scientificamerican03081879-155.pdf.

Anonymous:1879:CRi

[12567] Anonymous. Communications received. Scientific American, 40 (10):156, March 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n10/pdf/scientificamerican03081879-
156.pdf.

Anonymous:1879:Ck

[12568] Anonymous. Correspondence. Scientific American, 40(11):165, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-165.pdf.

Anonymous:1879:BPk

[12569] Anonymous. Business and personal. Scientific American, 40(11):171, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-171a.pdf.

Anonymous:1879:BNBf

[12570] Anonymous. Books: New books and publications. Scientific American, 40(11):171, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
171b.pdf.

Anonymous:1879:NQNa

[12571] Anonymous. Notes and queries: Notes and queris. Scientific American, 40(11):171, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
171c.pdf.

Anonymous:1879:IIk

[12572] Anonymous. Index of inventions. Scientific American, 40(11):171–172, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-171e.pdf.

Anonymous:1879:PNPj

[12573] Anonymous. Patent notes: Patents. Scientific American, 40(11):172, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-172.pdf.

Anonymous:1879:AINd

[12574] Anonymous. American industries — no. 8. Scientific American, 40(11):159, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
159.pdf.

Anonymous:1879:SSI

[12575] Anonymous. Science to solve industrial problems. Scientific American, 40(11):160, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
160.pdf.

Anonymous:1879:EGB

[12576] Anonymous. Economy in gas burning. Scientific American, 40(11):160, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-160a.pdf.

Anonymous:1879:EMa

[12577] Anonymous. Emigration and migration. Scientific American, 40 (11):160, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
160b.pdf.

Anonymous:1879:CFa

[12578] Anonymous. The canary fisheries. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-161.pdf.

Anonymous:1879:IOP

[12579] Anonymous. The iron ocean pier at long branch. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
161a.pdf.

Anonymous:1879:HSa

[12580] Anonymous. Honey sugar. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-161b.pdf.

Anonymous:1879:UCW

[12581] Anonymous. The use of cold water in cold weather. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
161c.pdf.

Anonymous:1879:GIa

[12582] Anonymous. Guatemala indigo. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-161d.pdf.

Anonymous:1879:LL

[12583] Anonymous. Light and life. Scientific American, 40(11):161, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-161e.pdf.

Anonymous:1879:NID

[12584] Anonymous. New instrument to determine the presence of metals in ores. Scientific American, 40(11):161, March 15, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n11/pdf/scientificamerican03151879-161f.pdf.

Anonymous:1879:NAIa

[12585] Anonymous. New agricultural inventions. Scientific American, 40(11):163, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
163.pdf.

Anonymous:1879:IWGa

[12586] Anonymous. An improved washer gutter. Scientific American, 40(11):163, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
163a.pdf.

Anonymous:1879:NHF

[12587] Anonymous. A new hurdle fence. Scientific American, 40(11):163, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-163b.pdf.

Anonymous:1879:AMc

[12588] Anonymous. Amateur mechanics. Scientific American, 40(11):164–165, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-164.pdf.

Anonymous:1879:SNP

[12589] Anonymous. Some new points in the diagnosis and prognosis of typhoid fever. Scientific American, 40(11):165, March 15, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n11/pdf/scientificamerican03151879-165a.pdf.

Anonymous:1879:NFMa

[12590] Anonymous. A new forging mandrel. Scientific American, 40(11):166, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-166.pdf.

Anonymous:1879:IIM

[12591] Anonymous. Improved ice-making apparatus. Scientific American, 40(11):166, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
166a.pdf.

Anonymous:1879:VT

[12592] Anonymous. The Victor turbine. Scientific American, 40(11):166, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-166b.pdf.

Anonymous:1879:RMIa

[12593] Anonymous. Recent mechanical inventions. Scientific American, 40(11):166, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
166c.pdf.

Anonymous:1879:SFA

[12594] Anonymous. Sponge fishing in the Adriatic — artificial sponge raising. Scientific American, 40(11):167–169, March 15, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n11/pdf/scientificamerican03151879-167.pdf.

Anonymous:1879:BGT

[12595] Anonymous. The Bawlkin Green thunderbolt. Scientific American, 40(11):169, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
169.pdf.

Anonymous:1879:SNI

[12596] Anonymous. The straw-necked ibis. Scientific American, 40(11): 169–170, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
169a.pdf.

Anonymous:1879:CCa

[12597] Anonymous. Catarrh and its cure. Scientific American, 40(11):171, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n11/pdf/scientificamerican03151879-171.pdf.

Anonymous:1879:CRj

[12598] Anonymous. Communications received. Scientific American, 40 (11):171, March 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n11/pdf/scientificamerican03151879-
171d.pdf.

Anonymous:1879:BPl

[12599] Anonymous. Business and personal. Scientific American, 40(12):187, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-187a.pdf.

Anonymous:1879:NQk

[12600] Anonymous. Notes and queries. Scientific American, 40(12):187–188, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-187b.pdf.

Anonymous:1879:IIl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-188a.pdf.

Anonymous:1879:PNPk

[12602] Anonymous. Patent notes: Patents. Scientific American, 40(12):188, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-188b.pdf.

Anonymous:1879:PBD

[12603] Anonymous. The patent bill defeated. Scientific American, 40(12):176, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-176.pdf.

Anonymous:1879:MT

[12604] Anonymous. Model tenements. Scientific American, 40(12):176, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-176a.pdf.

Anonymous:1879:PT

[12605] Anonymous. Progress of the telephone. Scientific American, 40(12):176, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-176b.pdf.

Anonymous:1879:ME

[12606] Anonymous. The menace to Europe. Scientific American, 40 (12):176–177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
176c.pdf.

Anonymous:1879:HPB

[12607] Anonymous. How the patent bill died. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-177.pdf.

Anonymous:1879:MRG

[12608] Anonymous. Mr. Robert G. Hatfield. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-177a.pdf.

Anonymous:1879:AINe

[12609] Anonymous. American industries — no. 9. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
177b.pdf.

Anonymous:1879:GMR

[12610] Anonymous. A glass mountain and road. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
177c.pdf.

Anonymous:1879:SMa

[12611] Anonymous. A surprise to millers. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-177d.pdf.

Anonymous:1879:SSa

[12612] Anonymous. Safety at sea. Scientific American, 40(12):177, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-177e.pdf.

Anonymous:1879:NVF

[12613] Anonymous. New vertical French Burr mills. Scientific American, 40(12):178, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
178.pdf.

Anonymous:1879:RAPb

[12614] Anonymous. Recent American patents. Scientific American, 40(12):178, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-178a.pdf.

Anonymous:1879:NTMa

[12615] Anonymous. New testing machine. Scientific American, 40(12):179, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-179.pdf.

Anonymous:1879:NHE

[12616] Anonymous. A new honey extractor. Scientific American, 40(12):179, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-179a.pdf.

Anonymous:1879:RAIa

[12617] Anonymous. Recent American inventions. Scientific American, 40(12):179, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
179b.pdf.

Anonymous:1879:SAP

[12618] Anonymous. Some anatomical peculiarities in the male chimpanzee. Scientific American, 40(12):180–181, March 22, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n12/pdf/scientificamerican03221879-180.pdf.

Anonymous:1879:Pb

[12619] Anonymous. The praxinoscope. Scientific American, 40(12):181, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-181.pdf.

Anonymous:1879:SRIb

[12620] Anonymous. Some recent inventions. Scientific American, 40(12):182, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-182.pdf.

Anonymous:1879:NPBd

[12621] Anonymous. A new packing box. Scientific American, 40(12):182, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-182a.pdf.

Anonymous:1879:EIa

[12622] Anonymous. Engineering inventions. Scientific American, 40(12):182, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-182b.pdf.

Anonymous:1879:PAT

[12623] Anonymous. The parasol ants of Texas. Scientific American, 40(12):183, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-183.pdf.

Anonymous:1879:GST

[12624] Anonymous. The giant snapping turtle. Scientific American, 40(12):183, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
183a.pdf.

Anonymous:1879:SEC

[12625] Anonymous. Siemens’ electric current governor. Scientific American, 40(12):184, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
184.pdf.

Anonymous:1879:NELa

[12626] Anonymous. New electric lamp. Scientific American, 40(12):184, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-184a.pdf.

Anonymous:1879:EELa

[12627] Anonymous. Edison’s electric light. Scientific American, 40(12):184, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-184b.pdf.

Anonymous:1879:EELb

[12628] Anonymous. The Edison electric light. Scientific American, 40 (12):185–186, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
185.pdf.

Anonymous:1879:RT

[12629] Anonymous. Righi’s telephone. Scientific American, 40(12):186, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-186.pdf.

Anonymous:1879:MIE

[12630] Anonymous. The Mexican Industrial Exhibition. Scientific American, 40(12):186, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
186a.pdf.

Anonymous:1879:AEL

[12631] Anonymous. The agricultural exhibition in London. Scientific American, 40(12):186, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
186b.pdf.

Anonymous:1879:Il

[12632] Anonymous. To inventors. Scientific American, 40(12):187, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n12/pdf/scientificamerican03221879-187.pdf.

Anonymous:1879:CRk

[12633] Anonymous. Communications received. Scientific American, 40 (12):188, March 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n12/pdf/scientificamerican03221879-
188.pdf.

Anonymous:1879:Cl

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-196.pdf.

Anonymous:1879:BPm

[12635] Anonymous. Business and personal. Scientific American, 40(13):203, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-203a.pdf.

Anonymous:1879:NQl

[12636] Anonymous. Notes and queries. Scientific American, 40(13):203, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-203b.pdf.

Anonymous:1879:IIn

[12637] Anonymous. Index of inventions. Scientific American, 40(13):203–204, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-203d.pdf.

Anonymous:1879:PNPl

[12638] Anonymous. Patent notes: Patents. Scientific American, 40(13):204, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-204.pdf.

Anonymous:1879:NSTa

[12639] Anonymous. A new steam tiller. Scientific American, 40(13):191–192, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-191.pdf.

Anonymous:1879:HOP

[12640] Anonymous. How our patent laws promote and improve American industries. Scientific American, 40(13):192, March 29, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n13/pdf/scientificamerican03291879-192.pdf.

Anonymous:1879:PPb

[12641] Anonymous. Pleuro-pneumonia. Scientific American, 40(13):192–193, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-192a.pdf.

Anonymous:1879:SFM

[12642] Anonymous. Spain a field for machinery and patents. Scientific American, 40(13):193, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
193.pdf.

Anonymous:1879:PCI

[12643] Anonymous. Petek coofer as an inventor. Scientific American, 40(13):193, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
193a.pdf.

Anonymous:1879:CFM

[12644] Anonymous. Curious facts in magnetism. Scientific American, 40(13):194, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
194.pdf.

Anonymous:1879:NOC

[12645] Anonymous. A new ore crusher. Scientific American, 40(13):194, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-194a.pdf.

Anonymous:1879:RAPc

[12646] Anonymous. Recent American patents. Scientific American, 40(13):194, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-194b.pdf.

Anonymous:1879:NIT

[12647] Anonymous. A new ironing table. Scientific American, 40(13):194, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-194c.pdf.

Anonymous:1879:NER

[12648] Anonymous. A novel engine regulator. Scientific American, 40(13):195, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-195.pdf.

Anonymous:1879:RNC

[12649] Anonymous. Remedy for the new carpet beetle. Scientific American, 40(13):195, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
195a.pdf.

Anonymous:1879:NSPb

[12650] Anonymous. New stencil pen. Scientific American, 40(13):195–196, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-195b.pdf.

Anonymous:1879:WTb

[12651] Anonymous. The writing telegraph. Scientific American, 40(13):196, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-196a.pdf.

Anonymous:1879:CWT

[12652] Anonymous. Cowper’s writing telegraph. Scientific American, 40(13):197, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
197.pdf.

Anonymous:1879:Ab

[12653] Anonymous. Aluminum. Scientific American, 40(13):197, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-197a.pdf.

Anonymous:1879:IDB

[12654] Anonymous. An improved door bolt. Scientific American, 40(13):198, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-198.pdf.

Anonymous:1879:IFB

[12655] Anonymous. Improved furnace for burning garbage. Scientific American, 40(13):198, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
198a.pdf.

Anonymous:1879:AGV

[12656] Anonymous. An ancient Greek vase. Scientific American, 40(13):199, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-199.pdf.

Anonymous:1879:NAIb

[12657] Anonymous. New agricultural inventions. Scientific American, 40 (13):199–200, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
199a.pdf.

Anonymous:1879:GS

[12658] Anonymous. Glass sponges. Scientific American, 40(13):200–202, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-200.pdf.

Anonymous:1879:Im

[12659] Anonymous. To inventors. Scientific American, 40(13):203, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n13/pdf/scientificamerican03291879-203.pdf.

Anonymous:1879:CRl

[12660] Anonymous. Communications received. Scientific American, 40 (13):203, March 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n13/pdf/scientificamerican03291879-
203c.pdf.

Anonymous:1879:BPn

[12661] Anonymous. Business and personal. Scientific American, 40(14):218, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-218a.pdf.

Anonymous:1879:NQm

[12662] Anonymous. Notes and queries. Scientific American, 40(14):218–220, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-218b.pdf.

Anonymous:1879:IIo

[12663] Anonymous. Index of inventions. Scientific American, 40(14):220, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-220a.pdf.

Anonymous:1879:PNPm

[12664] Anonymous. Patent notes: Patents. Scientific American, 40(14):220, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-220b.pdf.

Hopkins:1879:NAA

[12665] Geo. M. Hopkins. New acoustic apparatus. Scientific American, 40(14):207, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
207.pdf.

Anonymous:1879:LCL

[12666] Anonymous. Labor and the cost of living. Scientific American, 40(14):208, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
208.pdf.

Anonymous:1879:DLP

[12667] Anonymous. Dangers of lead poisoning from the use of tinned ware. Scientific American, 40(14):208, April 5, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n14/pdf/scientificamerican04051879-208a.pdf.

Anonymous:1879:GAN

[12668] Anonymous. Gary’s alleges neutral line. Scientific American, 40 (14):208, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
208b.pdf.

Anonymous:1879:CBAa

[12669] Anonymous. A contrast between American and English artisans. Scientific American, 40(14):209, April 5, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n14/pdf/scientificamerican04051879-209.pdf.

Anonymous:1879:SER

[12670] Anonymous. The security of the East River Bridge. Scientific American, 40(14):209, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
209a.pdf.

Anonymous:1879:NLN

[12671] Anonymous. The neutral line. — a note from Mr. Gary. Scientific American, 40(14):209, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
209b.pdf.

Anonymous:1879:MHMb

[12672] Anonymous. Miss Hosmer’s motor. Scientific American, 40(14):209, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-209c.pdf.

Anonymous:1879:HG

[12673] Anonymous. Heinrich Geissler. Scientific American, 40(14):209, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-209d.pdf.

Anonymous:1879:HTL

[12674] Anonymous. The Herreshoff torpedo launch. Scientific American, 40(14):210–211, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
210.pdf.

Anonymous:1879:ILM

[12675] Anonymous. An improved lawn mower. Scientific American, 40(14):211, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-211.pdf.

Anonymous:1879:NSTb

[12676] Anonymous. A new shipping tag. Scientific American, 40(14):211–212, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-211a.pdf.

Anonymous:1879:SRAa

[12677] Anonymous. Some recent American patents. Scientific American, 40(14):212–213, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
212.pdf.

Anonymous:1879:WIC

[12678] Anonymous. Western Iowa coal fields. Scientific American, 40 (14):213–214, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
213.pdf.

Anonymous:1879:IP

[12679] Anonymous. An improved press. Scientific American, 40(14):214, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-214.pdf.

Anonymous:1879:RAPd

[12680] Anonymous. Recent American patents. Scientific American, 40 (14):214–215, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
214a.pdf.

Anonymous:1879:B

[12681] Anonymous. The bobac. Scientific American, 40(14):215–216, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-215.pdf.

Anonymous:1879:MIc

[12682] Anonymous. Miscellaneous inventions. Scientific American, 40(14):216, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-216.pdf.

Anonymous:1879:NASa

[12683] Anonymous. New automatic siphon. Scientific American, 40(14):216, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-216a.pdf.

Anonymous:1879:SCD

[12684] Anonymous. Supreme Court decision. Scientific American, 40 (14):216–217, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
216b.pdf.

Anonymous:1879:IWM

[12685] Anonymous. An improved washing machine. Scientific American, 40(14):217, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
217.pdf.

Seiss:1879:KS

[12686] C. Few Seiss. De Kay’s snake. Scientific American, 40(14):217, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-217a.pdf.

Anonymous:1879:In

[12687] Anonymous. To inventors. Scientific American, 40(14):218, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n14/pdf/scientificamerican04051879-218.pdf.

Anonymous:1879:CRm

[12688] Anonymous. Communications received. Scientific American, 40 (14):220, April 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n14/pdf/scientificamerican04051879-
220.pdf.

Anonymous:1879:Cm

[12689] Anonymous. Correspondence. Scientific American, 40(15):228, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-228a.pdf.

Anonymous:1879:BPo

[12690] Anonymous. Business and personal. Scientific American, 40(15):235, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-235a.pdf.

Anonymous:1879:BNBg

[12691] Anonymous. Books: New books and publications. Scientific American, 40(15):235, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
235b.pdf.

Anonymous:1879:NQo

[12692] Anonymous. Notes and queries. Scientific American, 40(15):235–236, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-235c.pdf.

Anonymous:1879:IIp

[12693] Anonymous. Index of inventions. Scientific American, 40(15):236, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-236a.pdf.

Anonymous:1879:PNPn

[12694] Anonymous. Patent notes: Patents. Scientific American, 40(15):238, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-238.pdf.

Anonymous:1879:AMd

[12695] Anonymous. Amateur mechanics. Scientific American, 40(15):223–224, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-223.pdf.

Anonymous:1879:HM

[12696] Anonymous. The home of the meteors. Scientific American, 40(15):224, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-224.pdf.

Anonymous:1879:AAM

[12697] Anonymous. American affluence of mechanical invention. Scientific American, 40(15):224, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
224a.pdf.

Anonymous:1879:JH

[12698] Anonymous. Jaded heads. Scientific American, 40(15):224–225, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-224b.pdf.

Anonymous:1879:WPP

[12699] Anonymous. Woman the primitive pottery-maker and decorator. Scientific American, 40(15):225, April 12, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n15/pdf/scientificamerican04121879-225.pdf.

Anonymous:1879:WW

[12700] Anonymous. Wire in wheat. Scientific American, 40(15):225, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-225a.pdf.

Anonymous:1879:ACa

[12701] Anonymous. Applications of celluloid. Scientific American, 40(15):225, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-225b.pdf.

Anonymous:1879:NFMb

[12702] Anonymous. Notes on faience and its manufacture. Scientific American, 40(15):226, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
226.pdf.

Anonymous:1879:RMIb

[12703] Anonymous. Recent mechanical inventions. Scientific American, 40(15):226, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
226a.pdf.

Anonymous:1879:MSSb

[12704] Anonymous. Machine for stretching and softening cloth. Scientific American, 40(15):227, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
227.pdf.

Anonymous:1879:NAB

[12705] Anonymous. A new attachment for boilers. Scientific American, 40(15):227, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
227a.pdf.

Anonymous:1879:FWT

[12706] Anonymous. Fountain watering trough for cattle. Scientific American, 40(15):227–228, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
227b.pdf.

Anonymous:1879:MId

[12707] Anonymous. Miscellaneous inventions. Scientific American, 40(15):228, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-228.pdf.

Anonymous:1879:AIC

[12708] Anonymous. Australian insects in California. Scientific American, 40(15):228, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
228b.pdf.

Anonymous:1879:HSE

[12709] Anonymous. The Heslop steam engine. Scientific American, 40(15):229, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-229.pdf.

Anonymous:1879:AIa

[12710] Anonymous. Agricultural inventions. Scientific American, 40 (15):229–230, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
229a.pdf.

Anonymous:1879:RI

[12711] Anonymous. Recent inventions. Scientific American, 40(15):230, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-230.pdf.

Anonymous:1879:NLA

[12712] Anonymous. New loading apparatus for ordnance. Scientific American, 40(15):230, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
230a.pdf.

Anonymous:1879:NFV

[12713] Anonymous. A new fender for vessels. Scientific American, 40(15):230, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-230b.pdf.

Anonymous:1879:FC

[12714] Anonymous. The forceps crab. Scientific American, 40(15):231, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-231.pdf.

Anonymous:1879:CBb

[12715] Anonymous. The Cape buffalo. Scientific American, 40(15):231, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-231a.pdf.

Anonymous:1879:BTD

[12716] Anonymous. Better than dime novels. Scientific American, 40 (15):232–233, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
232.pdf.

Anonymous:1879:FCI

[12717] Anonymous. Fine cast iron work. Scientific American, 40(15):233, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-233.pdf.

Anonymous:1879:WSL

[12718] Anonymous. What should be the legal standard of kerosene. Scientific American, 40(15):234, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
234.pdf.

Anonymous:1879:ATb

[12719] Anonymous. The allanthus as timber. Scientific American, 40(15):234, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-234a.pdf.

Anonymous:1879:Io

[12720] Anonymous. To inventors. Scientific American, 40(15):235, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n15/pdf/scientificamerican04121879-235.pdf.

Anonymous:1879:CRn

[12721] Anonymous. Communications received. Scientific American, 40 (15):236, April 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n15/pdf/scientificamerican04121879-
236.pdf.

Anonymous:1879:Cn

[12722] Anonymous. Correspondence. Scientific American, 40(16):242, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-242a.pdf.

Anonymous:1879:BPp

[12723] Anonymous. Business and personal. Scientific American, 40(16):251, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-251a.pdf.

Anonymous:1879:NQp

[12724] Anonymous. Notes and queries. Scientific American, 40(16):251–252, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-251b.pdf.

Anonymous:1879:IIq

[12725] Anonymous. Index of inventions. Scientific American, 40(16):252, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-252a.pdf.

Anonymous:1879:PNPo

[12726] Anonymous. Patent notes: Patents. Scientific American, 40(16):254, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-254.pdf.

Anonymous:1879:TN

[12727] Anonymous. Toxicological notes. Scientific American, 40(16):239, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-239.pdf.

Anonymous:1879:EYIa

[12728] Anonymous. The engines of the yacht Isa. Scientific American, 40(16):239–240, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
239a.pdf.

Anonymous:1879:MBA

[12729] Anonymous. Mathews’ boiler attachment. Scientific American, 40(16):240, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
240.pdf.

Anonymous:1879:WFN

[12730] Anonymous. The World’s fair of 1883 at New York. Scientific American, 40(16):240, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
240a.pdf.

Anonymous:1879:WFI

[12731] Anonymous. The Worcester Free Institute of Industrial Science. Scientific American, 40(16):240–241, April 19, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n16/pdf/scientificamerican04191879-240b.pdf.

Anonymous:1879:BWW

[12732] Anonymous. The Baltimore Water Works. Scientific American, 40(16):241, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
241.pdf.

Anonymous:1879:EYIb

[12733] Anonymous. An example for Young inventors. Scientific American, 40(16):241, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
241a.pdf.

Anonymous:1879:GTA

[12734] Anonymous. Good times for American farmers. Scientific American, 40(16):241, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
241b.pdf.

Anonymous:1879:PIS

[12735] Anonymous. A possible improvement in sugar making. Scientific American, 40(16):241, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
241c.pdf.

Anonymous:1879:AAI

[12736] Anonymous. Art as an aid to industry. Scientific American, 40(16):241, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-241d.pdf.

Anonymous:1879:NYAa

[12737] Anonymous. New York Academy of Sciences. Scientific American, 40(16):242, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
242.pdf.

Anonymous:1879:DBS

[12738] Anonymous. Dean Brothers’ steam pump. Scientific American, 40(16):243, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
243.pdf.

Anonymous:1879:RMIc

[12739] Anonymous. Recent mechanical inventions. Scientific American, 40(16):243, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
243a.pdf.

Anonymous:1879:BMM

[12740] Anonymous. Brick making by machinery. Scientific American, 40(16):243, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
243b.pdf.

Anonymous:1879:PHI

[12741] Anonymous. Professor Hughes’ induction balance. Scientific American, 40(16):244–245, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
244.pdf.

Anonymous:1879:PEI

[12742] Anonymous. Prof. Edison’s induction balance for telephone lines. Scientific American, 40(16):245, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
245.pdf.

Anonymous:1879:NPf

[12743] Anonymous. A new pedometer. Scientific American, 40(16):246, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-246.pdf.

Anonymous:1879:NBF

[12744] Anonymous. Novel baggage fastener and shawl carrier. Scientific American, 40(16):246, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
246a.pdf.

Anonymous:1879:NDK

[12745] Anonymous. A new dough kneader. Scientific American, 40(16):246, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-246b.pdf.

Anonymous:1879:RAPe

[12746] Anonymous. Recent American patents. Scientific American, 40(16):246, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-246c.pdf.

Anonymous:1879:FFa

[12747] Anonymous. The flying frog. Scientific American, 40(16):247, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-247.pdf.

Anonymous:1879:FSF

[12748] Anonymous. The fennec, or Sahara fox. Scientific American, 40(16):247, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-247a.pdf.

Anonymous:1879:NHNa

[12749] Anonymous. Natural history notes. Scientific American, 40(16): 247–249, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
247b.pdf.

Anonymous:1879:SWB

[12750] Anonymous. Silver water bottle. Scientific American, 40(16):249–250, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-249.pdf.

Anonymous:1879:Iq

[12751] Anonymous. To inventors. Scientific American, 40(16):251, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n16/pdf/scientificamerican04191879-251.pdf.

Anonymous:1879:CRo

[12752] Anonymous. Communications received. Scientific American, 40 (16):252, April 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n16/pdf/scientificamerican04191879-
252.pdf.

Anonymous:1879:Co

[12753] Anonymous. Correspondence. Scientific American, 40(17):261, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-261.pdf.

Anonymous:1879:BPq

[12754] Anonymous. Business and personal. Scientific American, 40(17):267, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-267a.pdf.

Anonymous:1879:NQq

[12755] Anonymous. Notes and queries. Scientific American, 40(17):267–268, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-267b.pdf.

Anonymous:1879:IIr

[12756] Anonymous. Index of inventions. Scientific American, 40(17):268, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-268a.pdf.

Anonymous:1879:PNPp

[12757] Anonymous. Patent notes: Patents. Scientific American, 40(17):270, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-270.pdf.

Anonymous:1879:NFMc

[12758] Anonymous. Notes on faience and its manufacture. Scientific American, 40(17):255, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
255.pdf.

Anonymous:1879:SGN

[12759] Anonymous. The site of the great New York exhibition. Scientific American, 40(17):256, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
256.pdf.

Anonymous:1879:FYF

[12760] Anonymous. Frost and yellow fever. Scientific American, 40(17):256, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-256a.pdf.

Anonymous:1879:AEa

[12761] Anonymous. The Australian exhibitions. Scientific American, 40 (17):256–257, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
256b.pdf.

Anonymous:1879:BPR

[12762] Anonymous. The bicycle as a practical road vehicle. Scientific American, 40(17):257, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
257.pdf.

Anonymous:1879:MIf

[12763] Anonymous. Is the Moon inhabited! Scientific American, 40(17):257, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-257a.pdf.

Anonymous:1879:IBN

[12764] Anonymous. Interest-bearing notes. Scientific American, 40(17):257, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-257b.pdf.

Anonymous:1879:ACb

[12765] Anonymous. The aniline colors. Scientific American, 40(17):257, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-257c.pdf.

Anonymous:1879:BCA

[12766] Anonymous. Blasting by compressed air. Scientific American, 40(17):258, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
258.pdf.

Anonymous:1879:RAPf

[12767] Anonymous. Recent American patents. Scientific American, 40 (17):258–259, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
258a.pdf.

Anonymous:1879:NREa

[12768] Anonymous. A novel rotary engine. Scientific American, 40(17):259, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-259.pdf.

Anonymous:1879:EEC

[12769] Anonymous. Edison’s electro-chemical telephone. Scientific American, 40(17):260, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
260.pdf.

Anonymous:1879:EIb

[12770] Anonymous. Engineering inventions. Scientific American, 40(17):261, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-261a.pdf.

Anonymous:1879:PBa

[12771] Anonymous. Premiums for Boys. Scientific American, 40(17):261, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-261b.pdf.

Anonymous:1879:CBAb

[12772] Anonymous. A convenient box for artists. Scientific American, 40(17):262, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
262.pdf.

Anonymous:1879:NCb

[12773] Anonymous. A new carbureter. Scientific American, 40(17):262–263, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-262a.pdf.

Anonymous:1879:IBB

[12774] Anonymous. Improved barrel for breech-loading shot guns. Scientific American, 40(17):263, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
263.pdf.

Anonymous:1879:GCS

[12775] Anonymous. The great crab-spider. Scientific American, 40(17): 263–264, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
263a.pdf.

Anonymous:1879:GP

[12776] Anonymous. Gray pottery. Scientific American, 40(17):264, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-264.pdf.

Anonymous:1879:NAIc

[12777] Anonymous. New agricultural inventions. Scientific American, 40(17):264, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
264a.pdf.

Anonymous:1879:OHM

[12778] Anonymous. Old homes made new. Scientific American, 40 (17):265–266, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
265.pdf.

Anonymous:1879:Ir

[12779] Anonymous. To inventors. Scientific American, 40(17):267, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n17/pdf/scientificamerican04261879-267.pdf.

Anonymous:1879:CRp

[12780] Anonymous. Communications received. Scientific American, 40 (17):268, April 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n17/pdf/scientificamerican04261879-
268.pdf.

Anonymous:1879:Cp

[12781] Anonymous. Correspondence. Scientific American, 40(18):282, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-282.pdf.

Anonymous:1879:BPs

[12782] Anonymous. Business and personal. Scientific American, 40(18):283, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-283a.pdf.

Anonymous:1879:BNBh

[12783] Anonymous. Books: New books and publications. Scientific American, 40(18):283, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
283b.pdf.

Anonymous:1879:NQr

[12784] Anonymous. Notes and queries. Scientific American, 40(18):283–284, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-283c.pdf.

Anonymous:1879:IIs

[12785] Anonymous. Index of inventions. Scientific American, 40(18):284, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-284a.pdf.

Anonymous:1879:PN

[12786] Anonymous. Patent notes. Scientific American, 40(18):286, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-286.pdf.

Anonymous:1879:AINf

[12787] Anonymous. American industries — no. 10. Scientific American, 40(18):271, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
271.pdf.

Anonymous:1879:PEL

[12788] Anonymous. Progress of electric lighting. Scientific American, 40(18):272, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
272.pdf.

Anonymous:1879:SFb

[12789] Anonymous. Scarlet fever. Scientific American, 40(18):272, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-272a.pdf.

Anonymous:1879:BLT

[12790] Anonymous. Better late than never. Scientific American, 40(18):272, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-272b.pdf.

Anonymous:1879:MEM

[12791] Anonymous. Museums of Europe — Metropolitan Museum of Art. Scientific American, 40(18):273, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
273.pdf.

Anonymous:1879:BNP

[12792] Anonymous. Barff’s new process for preserving iron. Scientific American, 40(18):273, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
273a.pdf.

Anonymous:1879:VM

[12793] Anonymous. Vermont marble. Scientific American, 40(18):273, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-273b.pdf.

Anonymous:1879:RMId

[12794] Anonymous. Recent mechanical inventions. Scientific American, 40(18):275, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
275.pdf.

Anonymous:1879:IAa

[12795] Anonymous. An improved amalgamator. Scientific American, 40 (18):275, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
275a.pdf.

Anonymous:1879:RNa

[12796] Anonymous. Railway notes. Scientific American, 40(18):276, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-276.pdf.

Anonymous:1879:HIc

[12797] Anonymous. Household inventions. Scientific American, 40(18):277, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-277.pdf.

Anonymous:1879:NCPb

[12798] Anonymous. A new clutch pulley. Scientific American, 40(18):277, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-277a.pdf.

Anonymous:1879:EBP

[12799] Anonymous. Emery band polishing machine. Scientific American, 40(18):278, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
278.pdf.

Anonymous:1879:NAFa

[12800] Anonymous. A new anti-friction journal bearing. Scientific American, 40(18):278, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
278a.pdf.

Anonymous:1879:MIg

[12801] Anonymous. Miscellaneous inventions. Scientific American, 40(18):278, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-278b.pdf.

Anonymous:1879:IWE

[12802] Anonymous. An improved water elevator. Scientific American, 40(18):278, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
278c.pdf.

Anonymous:1879:RAR

[12803] Anonymous. Remarkable accident to a red deer. Scientific American, 40(18):279–280, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
279.pdf.

Anonymous:1879:NAId

[12804] Anonymous. New agricultural inventions. Scientific American, 40(18):280, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
280.pdf.

Anonymous:1879:JM

[12805] Anonymous. Japanese mirrors. Scientific American, 40(18):280, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-280a.pdf.

Anonymous:1879:CSa

[12806] Anonymous. Calcareous sponges. Scientific American, 40(18):281–282, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-281.pdf.

Anonymous:1879:Is

[12807] Anonymous. To inventors. Scientific American, 40(18):283, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n18/pdf/scientificamerican05031879-283.pdf.

Anonymous:1879:CRq

[12808] Anonymous. Communications received. Scientific American, 40 (18):284, May 3, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n18/pdf/scientificamerican05031879-
284.pdf.

Anonymous:1879:Cq

[12809] Anonymous. Correspondence. Scientific American, 40(19):296, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-296a.pdf.

Anonymous:1879:BPt

[12810] Anonymous. Business and personal. Scientific American, 40(19):299, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-299a.pdf.

Anonymous:1879:NQs

[12811] Anonymous. Notes and queries. Scientific American, 40(19):299, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-299b.pdf.

Anonymous:1879:IIt

[12812] Anonymous. Index of inventions. Scientific American, 40(19):299–300, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-299d.pdf.

Anonymous:1879:PNPq

[12813] Anonymous. Patent notes: Patents. Scientific American, 40(19):300, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-300.pdf.

Anonymous:1879:SIa

[12814] Anonymous. Study to have ideas. Scientific American, 40(19):288, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-288.pdf.

Anonymous:1879:PRN

[12815] Anonymous. “The patent right nuisance”. Scientific American, 40(19):288, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
288a.pdf.

Anonymous:1879:EVA

[12816] Anonymous. English vs. American rails. Scientific American, 40 (19):288–289, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
288b.pdf.

Anonymous:1879:RCR

[12817] Anonymous. A remarkable conflagration — the river between New York and Brooklyn set on fire. Scientific American, 40(19):289, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-289.pdf.

Anonymous:1879:AINg

[12818] Anonymous. American industries — no. 11. Scientific American, 40(19):289–290, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
289a.pdf.

Anonymous:1879:TMa

[12819] Anonymous. Therapeutic machinery. Scientific American, 40(19):290, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-290.pdf.

Anonymous:1879:MIh

[12820] Anonymous. Miscellaneous inventions. Scientific American, 40 (19):290–291, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
290a.pdf.

Anonymous:1879:RCa

[12821] Anonymous. Reflector for candles. Scientific American, 40(19):291, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-291.pdf.

Anonymous:1879:NHG

[12822] Anonymous. New hydraulic grid. Scientific American, 40(19):291–292, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-291a.pdf.

Anonymous:1879:NL

[12823] Anonymous. A new locomotive. Scientific American, 40(19):292, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-292.pdf.

Anonymous:1879:MCa

[12824] Anonymous. Milk cooler. Scientific American, 40(19):292, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-292a.pdf.

Anonymous:1879:EIc

[12825] Anonymous. Engineering inventions. Scientific American, 40(19):293, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-293.pdf.

Anonymous:1879:RIL

[12826] Anonymous. Rowell’s improved lens for spectacles. Scientific American, 40(19):293, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
293a.pdf.

Anonymous:1879:ET

[12827] Anonymous. The explosion on the thunderer. Scientific American, 40(19):293, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
293b.pdf.

Anonymous:1879:IMa

[12828] Anonymous. An improved mangle. Scientific American, 40(19):294, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-294.pdf.

Anonymous:1879:IEW

[12829] Anonymous. An improved emery wheel stand. Scientific American, 40(19):294–295, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
294a.pdf.

Anonymous:1879:WG

[12830] Anonymous. The window garden. Scientific American, 40(19):295, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-295.pdf.

Anonymous:1879:Y

[12831] Anonymous. The yak. Scientific American, 40(19):295–296, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-295a.pdf.

Anonymous:1879:ALP

[12832] Anonymous. Artificial lighting for photography. Scientific American, 40(19):296, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
296.pdf.

Anonymous:1879:CIc

[12833] Anonymous. A chance for inventors. Scientific American, 40(19):296, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-296b.pdf.

Anonymous:1879:EC

[12834] Anonymous. An elegant cabinet. Scientific American, 40(19):297, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-297.pdf.

Anonymous:1879:CTT

[12835] Anonymous. The catalpa as a timber tree. Scientific American, 40(19):297, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
297a.pdf.

Anonymous:1879:IEI

[12836] Anonymous. An international exhibition of industrial science. Scientific American, 40(19):298, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
298.pdf.

Anonymous:1879:RAPg

[12837] Anonymous. Recent American patents. Scientific American, 40(19):298, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-298a.pdf.

Anonymous:1879:Iu

[12838] Anonymous. To inventors. Scientific American, 40(19):299, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n19/pdf/scientificamerican05101879-299.pdf.

Anonymous:1879:CRr

[12839] Anonymous. Communications received. Scientific American, 40 (19):299, May 10, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n19/pdf/scientificamerican05101879-
299c.pdf.

Anonymous:1879:BPu

[12840] Anonymous. Business and personal. Scientific American, 40(20):315, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-315a.pdf.

Anonymous:1879:BNBi

[12841] Anonymous. Books: New books and publications. Scientific American, 40(20):315, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
315b.pdf.

Anonymous:1879:NQt

[12842] Anonymous. Notes and queries. Scientific American, 40(20):315–316, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-315c.pdf.

Anonymous:1879:IIu

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-316a.pdf.

Anonymous:1879:PNPr

[12844] Anonymous. Patent notes: Patents. Scientific American, 40(20):318, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-318.pdf.

Anonymous:1879:NEW

[12845] Anonymous. A new estimate of the World’s age. Scientific American, 40(20):304, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
304.pdf.

Anonymous:1879:LTW

[12846] Anonymous. The longest tunnel in the world. Scientific American, 40(20):304, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
304a.pdf.

Anonymous:1879:PVH

[12847] Anonymous. The power of vibration harnessed. Scientific American, 40(20):304–305, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
304b.pdf.

Anonymous:1879:WLD

[12848] Anonymous. When are laws discovered! Scientific American, 40(20):304–305, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
304c.pdf.

Anonymous:1879:MCNa

[12849] Anonymous. Molecular chemistry — no. 1. Scientific American, 40(20):305, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
305.pdf.

Anonymous:1879:EEI

[12850] Anonymous. Edison’s electric illuminator and Dr. Draper’s experiments thirty years ago. Scientific American, 40(20):305, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-305a.pdf.

Anonymous:1879:AINh

[12851] Anonymous. American industries — no. 12. Scientific American, 40(20):305, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
305b.pdf.

Anonymous:1879:TL

[12852] Anonymous. Testing lubricants. Scientific American, 40(20):306–307, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-306.pdf.

Anonymous:1879:FN

[12853] Anonymous. A few novelties. Scientific American, 40(20):307–308, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-307.pdf.

Anonymous:1879:NAIe

[12854] Anonymous. New agricultural inventions. Scientific American, 40 (20):309, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
309.pdf.

Anonymous:1879:DRT

[12855] Anonymous. The Detroit River Tunnel and Bridge. Scientific American, 40(20):309, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
309a.pdf.

Anonymous:1879:PPA

[12856] Anonymous. The patentees’ protective association. Scientific American, 40(20):309, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
309b.pdf.

Anonymous:1879:Da

[12857] Anonymous. Disinfection. Scientific American, 40(20):309, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-309c.pdf.

Anonymous:1879:IPC

[12858] Anonymous. Improved paper cutting and winding machine. Scientific American, 40(20):310, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
310.pdf.

Anonymous:1879:NCSb

[12859] Anonymous. New cut-off for steam engines. Scientific American, 40(20):310–311, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
310a.pdf.

Beard:1879:KPM

[12860] Daniel C. Beard. The kanchil, or pygmy musk. Scientific American, 40(20):311–312, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
311.pdf.

Anonymous:1879:RMIe

[12861] Anonymous. Recent mechanical inventions. Scientific American, 40(20):312, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
312.pdf.

Anonymous:1879:NHNb

[12862] Anonymous. Natural history notes. Scientific American, 40(20): 312–313, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
312a.pdf.

Anonymous:1879:Ct

[12863] Anonymous. The calamar. Scientific American, 40(20):313–314, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-313.pdf.

Anonymous:1879:Iv

[12864] Anonymous. To inventors. Scientific American, 40(20):315, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n20/pdf/scientificamerican05171879-315.pdf.

Anonymous:1879:CRs

[12865] Anonymous. Communications received. Scientific American, 40 (20):316, May 17, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n20/pdf/scientificamerican05171879-
316.pdf.

Anonymous:1879:Cu

[12866] Anonymous. Correspondence. Scientific American, 40(21):325, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-325.pdf.

Anonymous:1879:BPv

[12867] Anonymous. Business and personal. Scientific American, 40(21):331, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-331a.pdf.

Anonymous:1879:BNBj

[12868] Anonymous. Books: New books and publications. Scientific American, 40(21):331, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
331b.pdf.

Anonymous:1879:NQNb

[12869] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(21):331–332, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
331c.pdf.

Anonymous:1879:IIv

[12870] Anonymous. Index of inventions. Scientific American, 40(21):332, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-332a.pdf.

Anonymous:1879:Pc

[12871] Anonymous. Patents. Scientific American, 40(21):334, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-334.pdf.

Anonymous:1879:PRB

[12872] Anonymous. Purifying rancid butter. Scientific American, 40(21):319, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-319.pdf.

Anonymous:1879:AINi

[12873] Anonymous. American industries, no. 13. Scientific American, 40(21):319, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
319a.pdf.

Anonymous:1879:MRM

[12874] Anonymous. The mental requirements of modern enterprises. Scientific American, 40(21):320, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
320.pdf.

Anonymous:1879:TSE

[12875] Anonymous. Three successful efforts. Scientific American, 40(21):320, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-320a.pdf.

Anonymous:1879:ELPa

[12876] Anonymous. The electric light in Paris. Scientific American, 40(21):320, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-320b.pdf.

Anonymous:1879:NPP

[12877] Anonymous. New process for preserving and ornamenting iron. Scientific American, 40(21):320–321, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
320c.pdf.

Anonymous:1879:WOA

[12878] Anonymous. Who originated the Atlantic cable? Scientific American, 40(21):321, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
321.pdf.

Anonymous:1879:WCT

[12879] Anonymous. The world-circuit and time puzzle. Scientific American, 40(21):321, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
321a.pdf.

Anonymous:1879:PBN

[12880] Anonymous. Prof. Bert’s new anæsthetic. Scientific American, 40(21):321, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
321b.pdf.

Anonymous:1879:NSDa

[12881] Anonymous. A new satchel desk. Scientific American, 40(21):323, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-323.pdf.

Anonymous:1879:MCP

[12882] Anonymous. Moulding, carving, paneling and dovetailing machine. Scientific American, 40(21):323, May 24, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v40/
n21/pdf/scientificamerican05241879-323a.pdf.

Anonymous:1879:EId

[12883] Anonymous. Engineering inventions. Scientific American, 40(21):324, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-324.pdf.

Anonymous:1879:IMb

[12884] Anonymous. Improved microphones. Scientific American, 40(21):324, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-324a.pdf.

Balch:1879:SE

[12885] S. W. Balch. A simple ellipsograph. Scientific American, 40(21):324, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-324b.pdf.

Anonymous:1879:ISDa

[12886] Anonymous. Improved stock and dies for pipes and bolts. Scientific American, 40(21):325, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
325a.pdf.

Anonymous:1879:RAPh

[12887] Anonymous. Recent American patents. Scientific American, 40 (21):325–326, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
325b.pdf.

Anonymous:1879:NFG

[12888] Anonymous. A new flour, grain, and bolting cloth inspector. Scientific American, 40(21):326, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
326.pdf.

Anonymous:1879:NREb

[12889] Anonymous. A new rotary engine. Scientific American, 40(21):326, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-326a.pdf.

Anonymous:1879:MIi

[12890] Anonymous. Miscellaneous inventions. Scientific American, 40(21):326, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-326b.pdf.

Anonymous:1879:OF

[12891] Anonymous. The Orchis family. Scientific American, 40(21):327–328, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-327.pdf.

Anonymous:1879:NYAb

[12892] Anonymous. The New York Academy of Sciences. Scientific American, 40(21):328, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
328.pdf.

Seiss:1879:BD

[12893] C. Few Seiss. The brown desmognath. Scientific American, 40(21):329, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-329.pdf.

Anonymous:1879:RCb

[12894] Anonymous. A rich chair. Scientific American, 40(21):329, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-329a.pdf.

Anonymous:1879:SFc

[12895] Anonymous. A sensible fashion. Scientific American, 40(21):330, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-330.pdf.

Anonymous:1879:Iw

[12896] Anonymous. To inventors. Scientific American, 40(21):331, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n21/pdf/scientificamerican05241879-331.pdf.

Anonymous:1879:CRt

[12897] Anonymous. Communications received. Scientific American, 40 (21):332, May 24, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n21/pdf/scientificamerican05241879-
332.pdf.

Anonymous:1879:BPw

[12898] Anonymous. Business and personal. Scientific American, 40(22):347, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-347a.pdf.

Anonymous:1879:BNBk

[12899] Anonymous. Books: New books and publications. Scientific American, 40(22):347, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
347b.pdf.

Anonymous:1879:NQNc

[12900] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(22):347–348, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
347c.pdf.

Anonymous:1879:IIw

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-348a.pdf.

Anonymous:1879:Pd

[12902] Anonymous. Patents. Scientific American, 40(22):350, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-350.pdf.

Anonymous:1879:DPa

[12903] Anonymous. Design patents. Scientific American, 40(22):336, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-336.pdf.

Anonymous:1879:IY

[12904] Anonymous. The impatience of youth. Scientific American, 40(22):336, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-336a.pdf.

Anonymous:1879:CMH

[12905] Anonymous. Common mistakes in house building. Scientific American, 40(22):337, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
337.pdf.

Anonymous:1879:CMGb

[12906] Anonymous. Chicle, or Mexican gum. Scientific American, 40(22):337, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-337a.pdf.

Anonymous:1879:CCb

[12907] Anonymous. The cause of consumption. Scientific American, 40 (22):337, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
337b.pdf.

Anonymous:1879:AINj

[12908] Anonymous. American industries — no. 14. Scientific American, 40(22):337–338, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
337c.pdf.

Anonymous:1879:NM

[12909] Anonymous. A novel motor. Scientific American, 40(22):338, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-338.pdf.

Anonymous:1879:TSS

[12910] Anonymous. Thrasher, straw scale, and sheaf binder. Scientific American, 40(22):339, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
339.pdf.

Anonymous:1879:MIj

[12911] Anonymous. Miscellaneous inventions. Scientific American, 40(22):339, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-339a.pdf.

Anonymous:1879:AMe

[12912] Anonymous. Amateur mechanics. Scientific American, 40(22):340, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-340.pdf.

Anonymous:1879:RNb

[12913] Anonymous. Railway notes. Scientific American, 40(22):341, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-341.pdf.

Anonymous:1879:NSBa

[12914] Anonymous. New steam boiler. Scientific American, 40(22):342, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-342.pdf.

Anonymous:1879:EIe

[12915] Anonymous. Engineering inventions. Scientific American, 40(22): 342–343, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
342a.pdf.

Anonymous:1879:EA

[12916] Anonymous. The equine antelope. Scientific American, 40(22):343–344, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-343.pdf.

Anonymous:1879:GDC

[12917] Anonymous. Greek drinking cup. Scientific American, 40(22):344, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-344.pdf.

Anonymous:1879:NPSa

[12918] Anonymous. New provision safe. Scientific American, 40(22):344–345, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-344a.pdf.

Anonymous:1879:EMb

[12919] Anonymous. The edible mussel. Scientific American, 40(22):345, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-345.pdf.

Anonymous:1879:NAIf

[12920] Anonymous. New agricultural inventions. Scientific American, 40 (22):345–346, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
345a.pdf.

Anonymous:1879:GMC

[12921] Anonymous. Gerard Mercator, the cosmographer. Scientific American, 40(22):346, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
346.pdf.

Anonymous:1879:DPb

[12922] Anonymous. David Page. Scientific American, 40(22):346, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-346a.pdf.

Anonymous:1879:Ix

[12923] Anonymous. To inventors. Scientific American, 40(22):347, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n22/pdf/scientificamerican05311879-347.pdf.

Anonymous:1879:CRu

[12924] Anonymous. Communications received. Scientific American, 40 (22):348, May 31, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n22/pdf/scientificamerican05311879-
348.pdf.

Anonymous:1879:BPx

[12925] Anonymous. Business and personal. Scientific American, 40(23):363, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-363.pdf.

Anonymous:1879:NQNd

[12926] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(23):363, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
363a.pdf.

Anonymous:1879:IIx

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-363c.pdf.

Anonymous:1879:PNPs

[12928] Anonymous. Patent notes: Patents. Scientific American, 40(23):366, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-366.pdf.

Anonymous:1879:Ad

[12929] Anonymous. The autophone. Scientific American, 40(23):351, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-351.pdf.

Anonymous:1879:TMb

[12930] Anonymous. The trajectory of molecules. Scientific American, 40(23):352, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
352.pdf.

Anonymous:1879:ICC

[12931] Anonymous. The International Canal Congress. Scientific American, 40(23):352, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
352a.pdf.

Anonymous:1879:SD

[12932] Anonymous. Science as a detective. Scientific American, 40(23):352, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-352b.pdf.

Anonymous:1879:NRL

[12933] Anonymous. A new refrigerating liquid from beets. Scientific American, 40(23):352, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
352c.pdf.

Anonymous:1879:MMa

[12934] Anonymous. Magnetic motors. Scientific American, 40(23):352–353, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-352d.pdf.

Anonymous:1879:UTW

[12935] Anonymous. Underground telegraph wires. Scientific American, 40(23):353, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
353.pdf.

Anonymous:1879:LTC

[12936] Anonymous. Localizing telephone calls. Scientific American, 40 (23):353, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
353a.pdf.

Anonymous:1879:MCNb

[12937] Anonymous. Molecular chemistry — no. II. Scientific American, 40(23):353, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
353b.pdf.

Anonymous:1879:DRP

[12938] Anonymous. The Donnadieu reciprocating pump. Scientific American, 40(23):355–356, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
355.pdf.

Anonymous:1879:SPa

[12939] Anonymous. A sixpenny phonograph. Scientific American, 40 (23):356–357, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
356.pdf.

Anonymous:1879:NYAc

[12940] Anonymous. The New York Academy of Sciences. Scientific American, 40(23):357, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
357.pdf.

Anonymous:1879:RMIf

[12941] Anonymous. Recent mechanical inventions. Scientific American, 40(23):357–358, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
357a.pdf.

Anonymous:1879:NSHa

[12942] Anonymous. A new steam hammer. Scientific American, 40(23):358, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-358.pdf.

Anonymous:1879:ABF

[12943] Anonymous. Attachment for boiler feed pumps. Scientific American, 40(23):358, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
358a.pdf.

Anonymous:1879:RAPi

[12944] Anonymous. Recent American patents. Scientific American, 40(23):358, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-358b.pdf.

Anonymous:1879:TGa

[12945] Anonymous. Terrestrial gasteropods. Scientific American, 40(23):359, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-359.pdf.

Anonymous:1879:Db

[12946] Anonymous. Drainage. Scientific American, 40(23):360–361, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-360.pdf.

Anonymous:1879:SIWb

[12947] Anonymous. Stool of inlaid wood and embroidered cloth. Scientific American, 40(23):361, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
361.pdf.

Anonymous:1879:STPa

[12948] Anonymous. Specimens of Turkish pottery. Scientific American, 40(23):361–362, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
361a.pdf.

Anonymous:1879:APa

[12949] Anonymous. Asa packer. Scientific American, 40(23):362, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n23/pdf/scientificamerican06071879-362.pdf.

Anonymous:1879:CRv

[12950] Anonymous. Communications received. Scientific American, 40 (23):363, June 7, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n23/pdf/scientificamerican06071879-
363b.pdf.

Anonymous:1879:BPy

[12951] Anonymous. Business and personal. Scientific American, 40(24):379, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-379.pdf.

Anonymous:1879:BNBl

[12952] Anonymous. Books: New books and publications. Scientific American, 40(24):379, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
379a.pdf.

Anonymous:1879:NQNe

[12953] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(24):379–380, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
379b.pdf.

Anonymous:1879:IIy

[12954] Anonymous. Index of inventions. Scientific American, 40(24):380, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-380a.pdf.

Anonymous:1879:ONP

[12955] Anonymous. Opening of the Northeast Passage. Scientific American, 40(24):368, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
368.pdf.

Anonymous:1879:AOL

[12956] Anonymous. Another objection to the license system. Scientific American, 40(24):368, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
368a.pdf.

Anonymous:1879:HCC

[12957] Anonymous. How coffee is cleaned. Scientific American, 40(24):368, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-368b.pdf.

Anonymous:1879:RCc

[12958] Anonymous. Robert Crawshay. Scientific American, 40(24):368, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-368c.pdf.

Anonymous:1879:TME

[12959] Anonymous. Triumphs of modern engineering. Scientific American, 40(24):368–369, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
368d.pdf.

Anonymous:1879:MCNc

[12960] Anonymous. Molecular chemistry — no. III. Scientific American, 40(24):369, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
369.pdf.

Anonymous:1879:GG

[12961] Anonymous. The Gatling gun. Scientific American, 40(24):370, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-370.pdf.

Anonymous:1879:NAIg

[12962] Anonymous. New agricultural inventions. Scientific American, 40 (24):370, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
370a.pdf.

Anonymous:1879:NHP

[12963] Anonymous. New hay press. Scientific American, 40(24):371–372, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-371.pdf.

Anonymous:1879:ISF

[12964] Anonymous. Improved snow flanger. Scientific American, 40(24):372, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-372.pdf.

Anonymous:1879:NSCa

[12965] Anonymous. New sample case. Scientific American, 40(24):372, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-372a.pdf.

Anonymous:1879:RAPj

[12966] Anonymous. Recent American patents. Scientific American, 40 (24):372–373, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
372b.pdf.

Anonymous:1879:NPC

[12967] Anonymous. New pipe cutting and threading machine. Scientific American, 40(24):374, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
374.pdf.

Anonymous:1879:NCd

[12968] Anonymous. A new condenser. Scientific American, 40(24):374, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-374a.pdf.

Anonymous:1879:Cv

[12969] Anonymous. The chameleon. Scientific American, 40(24):375–376, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-375.pdf.

Anonymous:1879:RMIg

[12970] Anonymous. Recent mechanical inventions. Scientific American, 40(24):376, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
376.pdf.

Anonymous:1879:NWP

[12971] Anonymous. A new winged projectile. Scientific American, 40(24):376, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-376a.pdf.

Anonymous:1879:WTc

[12972] Anonymous. The writing telegraph. Scientific American, 40(24):376, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n24/pdf/scientificamerican06141879-376b.pdf.

Anonymous:1879:WHC

[12973] Anonymous. Wilson’s “horograph,” or clockwork pen. Scientific American, 40(24):377, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
377.pdf.

Anonymous:1879:RDA

[12974] Anonymous. Recent discoveries in Assyria. Scientific American, 40(24):377–378, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
377a.pdf.

Anonymous:1879:CRw

[12975] Anonymous. Communications received. Scientific American, 40 (24):380, June 14, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n24/pdf/scientificamerican06141879-
380.pdf.

Anonymous:1879:BPz

[12976] Anonymous. Business and personal. Scientific American, 40(25):395, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-395.pdf.

Anonymous:1879:NQNf

[12977] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(25):395–396, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
395a.pdf.

Anonymous:1879:IIz

[12978] Anonymous. Index of inventions. Scientific American, 40(25):396, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-396a.pdf.

Anonymous:1879:PDE

[12979] Anonymous. Practical divisibility of the electric light. Scientific American, 40(25):383, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
383.pdf.

Anonymous:1879:PBE

[12980] Anonymous. A prairie Boy’s experience. Scientific American, 40 (25):384, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
384.pdf.

Anonymous:1879:UNY

[12981] Anonymous. Unknown New York. Scientific American, 40(25):384, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-384a.pdf.

Anonymous:1879:SPP

[12982] Anonymous. Strike of Pittsburg puddlers. Scientific American, 40(25):384, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
384b.pdf.

Anonymous:1879:UB

[12983] Anonymous. The utility of bees. Scientific American, 40(25):384–385, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-384c.pdf.

Anonymous:1879:PHR

[12984] Anonymous. Philander Higley Roots. Scientific American, 40(25):385, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-385.pdf.

Anonymous:1879:ICa

[12985] Anonymous. The Isthmus Canal. Scientific American, 40(25):385, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-385a.pdf.

Anonymous:1879:IFP

[12986] Anonymous. Improved finishing press. Scientific American, 40(25):386, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-386.pdf.

Anonymous:1879:MMb

[12987] Anonymous. Milling machine. Scientific American, 40(25):386–387, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-386a.pdf.

Anonymous:1879:IWC

[12988] Anonymous. Improved window cornice. Scientific American, 40 (25):387, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
387.pdf.

Anonymous:1879:NCT

[12989] Anonymous. New combination tool for merchants. Scientific American, 40(25):387, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
387a.pdf.

Anonymous:1879:NSHb

[12990] Anonymous. A new scuffle hoe. Scientific American, 40(25):387, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-387b.pdf.

Anonymous:1879:EIf

[12991] Anonymous. Engineering inventions. Scientific American, 40(25): 387–388, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
387c.pdf.

Anonymous:1879:RMIh

[12992] Anonymous. Recent mechanical inventions. Scientific American, 40(25):388, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
388.pdf.

Anonymous:1879:PSB

[12993] Anonymous. Preventive for slipping belts. Scientific American, 40(25):388–389, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
388a.pdf.

Anonymous:1879:IRP

[12994] Anonymous. Improved road plane. Scientific American, 40(25):390, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-390.pdf.

Anonymous:1879:IHa

[12995] Anonymous. Improved hydrant. Scientific American, 40(25):390, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-390a.pdf.

Anonymous:1879:RAPk

[12996] Anonymous. Recent American patents. Scientific American, 40(25):390, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-390b.pdf.

Anonymous:1879:SIE

[12997] Anonymous. Some interesting echinodermata. Scientific American, 40(25):391–392, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
391.pdf.

Anonymous:1879:NHNc

[12998] Anonymous. Natural history notes. Scientific American, 40(25): 392–393, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
392.pdf.

Anonymous:1879:FSa

[12999] Anonymous. Flower stand. Scientific American, 40(25):393–394, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n25/pdf/scientificamerican06211879-393.pdf.

Anonymous:1879:CRx

[13000] Anonymous. Communications received. Scientific American, 40 (25):396, June 21, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n25/pdf/scientificamerican06211879-
396.pdf.

Anonymous:1879:BPba

[13001] Anonymous. Business and personal. Scientific American, 40(26):409, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-409a.pdf.

Anonymous:1879:BNBm

[13002] Anonymous. Books: New books and publications. Scientific American, 40(26):409, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
409b.pdf.

Anonymous:1879:NQNg

[13003] Anonymous. Notes and queries: Notes & queries. Scientific American, 40(26):409–410, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
409c.pdf.

Anonymous:1879:IIba

[13004] Anonymous. Index of inventions. Scientific American, 40(26):410, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-410a.pdf.

Anonymous:1879:PNPt

[13005] Anonymous. Patent notes: Patents. Scientific American, 40(26):414, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-414.pdf.

Anonymous:1879:NSCb

[13006] Anonymous. A new switch controller. Scientific American, 40(26):399, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-399.pdf.

Anonymous:1879:SMS

[13007] Anonymous. Six months of scientific progress. Scientific American, 40(26):400, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
400.pdf.

Anonymous:1879:CALa

[13008] Anonymous. Citric acid-from the lime and lemon. Scientific American, 40(26):400, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
400a.pdf.

Anonymous:1879:TWO

[13009] Anonymous. The tide water oil pipe line completed. Scientific American, 40(26):400, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
400b.pdf.

Anonymous:1879:ISM

[13010] Anonymous. Improvement in sugar manufacture. Scientific American, 40(26):400–401, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
400c.pdf.

Anonymous:1879:BC

[13011] Anonymous. The Brewers’ Convention. Scientific American, 40(26):401, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-401.pdf.

Anonymous:1879:AEb

[13012] Anonymous. The Australian exhibitions. Scientific American, 40 (26):401, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
401a.pdf.

Anonymous:1879:NSA

[13013] Anonymous. The new steamship Arizona. Scientific American, 40(26):401, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
401b.pdf.

Anonymous:1879:IMd

[13014] Anonymous. Insects as medicine. Scientific American, 40(26):401, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-401c.pdf.

Anonymous:1879:KEL

[13015] Anonymous. The Krupp electric lamp. Scientific American, 40(26):402, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-402.pdf.

Anonymous:1879:NSSb

[13016] Anonymous. New shop saw. Scientific American, 40(26):403, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-403.pdf.

Anonymous:1879:NIBa

[13017] Anonymous. A new invalid bed. Scientific American, 40(26):403, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-403a.pdf.

Anonymous:1879:AMf

[13018] Anonymous. Amateur mechanics. Scientific American, 40(26):404, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-404.pdf.

Anonymous:1879:MIk

[13019] Anonymous. Miscellaneous inventions. Scientific American, 40(26):404, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-404a.pdf.

Anonymous:1879:RMIi

[13020] Anonymous. Recent mechanical inventions. Scientific American, 40(26):405–406, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
405.pdf.

Anonymous:1879:NREc

[13021] Anonymous. A new rotary engine. Scientific American, 40(26):406, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-406.pdf.

Anonymous:1879:Cw

[13022] Anonymous. The cyclodes. Scientific American, 40(26):407, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-407.pdf.

Anonymous:1879:FRD

[13023] Anonymous. Fastening and releasing device for cattle stalls. Scientific American, 40(26):407–408, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
407a.pdf.

Anonymous:1879:EP

[13024] Anonymous. The electric pen. Scientific American, 40(26):408, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-408.pdf.

Anonymous:1879:NSe

[13025] Anonymous. A new sounder. Scientific American, 40(26):408, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v40/n26/pdf/scientificamerican06281879-408a.pdf.

Anonymous:1879:NAIh

[13026] Anonymous. New agricultural inventions. Scientific American, 40 (26):408, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
408b.pdf.

Anonymous:1879:OCRa

[13027] Anonymous. An old concern re-established. Scientific American, 40(26):409, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
409.pdf.

Anonymous:1879:CRy

[13028] Anonymous. Communications received. Scientific American, 40 (26):410, June 28, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v40/n26/pdf/scientificamerican06281879-
410.pdf.

Anonymous:1879:Cx

[13029] Anonymous. Correspondence. Scientific American, 41(1):10, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-10.pdf.

Anonymous:1879:BPbb

[13030] Anonymous. Business and personal. Scientific American, 41(1):11, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-11a.pdf.

Anonymous:1879:BNBn

[13031] Anonymous. Books: New books and publications. Scientific American, 41(1):11, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
11b.pdf.

Anonymous:1879:NQu

[13032] Anonymous. Notes and queries. Scientific American, 41(1):11–12, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-11c.pdf.

Anonymous:1879:IIbb

[13033] Anonymous. Index of inventions. Scientific American, 41(1):12, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-12a.pdf.

Anonymous:1879:PNPu

[13034] Anonymous. Patent notes: Patents. Scientific American, 41(1):14, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-14.pdf.

Anonymous:1879:KW

[13035] Anonymous. The Kansas whirlwinds. Scientific American, 41(1):1–2, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-1.pdf.

Anonymous:1879:PRD

[13036] Anonymous. A patent right discussion in England. Scientific American, 41(1):2, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
2.pdf.

Anonymous:1879:MMI

[13037] Anonymous. Magnetizing molten iron. Scientific American, 41(1):2, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-2a.pdf.

Anonymous:1879:NTEa

[13038] Anonymous. A new theory of the Earth’s magnetic poles. Scientific American, 41(1):2–3, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
2b.pdf.

Anonymous:1879:OSP

[13039] Anonymous. Objections to self-propelling fire engines. Scientific American, 41(1):3, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
3.pdf.

Anonymous:1879:PTC

[13040] Anonymous. The prospects of tea culture. Scientific American, 41(1): 3, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-3a.pdf.

Anonymous:1879:DBT

[13041] Anonymous. The diseases of building timber. Scientific American, 41(1):3, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
3b.pdf.

Anonymous:1879:SC

[13042] Anonymous. “scientific credulity”. Scientific American, 41(1):3, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-3c.pdf.

Anonymous:1879:RAPl

[13043] Anonymous. Recent American patents. Scientific American, 41(1):4, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-4.pdf.

Anonymous:1879:NELb

[13044] Anonymous. New electric lamp. Scientific American, 41(1):4, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-4a.pdf.

Anonymous:1879:LUT

[13045] Anonymous. Luedtge’s universal telephone. Scientific American, 41(1): 4, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-4b.pdf.

Anonymous:1879:HA

[13046] Anonymous. Hydriodic acid. Scientific American, 41(1):5, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-5.pdf.

Anonymous:1879:PCa

[13047] Anonymous. Pneumatic clock. Scientific American, 41(1):5–6, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-5a.pdf.

Anonymous:1879:RMIj

[13048] Anonymous. Recent mechanical inventions. Scientific American, 41(1): 6, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-6.pdf.

Anonymous:1879:NRPa

[13049] Anonymous. A novel rotary pump. Scientific American, 41(1):6–7, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-6a.pdf.

Anonymous:1879:ES

[13050] Anonymous. The elephant shrew. Scientific American, 41(1):7, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-7.pdf.

Riley:1879:TTW

[13051] C. V. Riley. The thick-thighed walking stick. Scientific American, 41(1):7–9, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
7a.pdf.

Anonymous:1879:NDSa

[13052] Anonymous. New device for separating cream from milk. Scientific American, 41(1):9, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n1/pdf/scientificamerican07051879-
9.pdf.

Anonymous:1879:IVa

[13053] Anonymous. An inventor victorious. Scientific American, 41(1):9, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-9a.pdf.

Anonymous:1879:IAb

[13054] Anonymous. Improved anchor. Scientific American, 41(1):9, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-9b.pdf.

Anonymous:1879:EIg

[13055] Anonymous. Engineering inventions. Scientific American, 41(1):10, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-10a.pdf.

Anonymous:1879:OCRb

[13056] Anonymous. An old concern re-established. Scientific American, 41(1): 11, July 5, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-11.pdf.

Anonymous:1879:CRz

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n1/pdf/scientificamerican07051879-12.pdf.

Anonymous:1879:Cy

[13058] Anonymous. Correspondence. Scientific American, 41(2):18, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-18a.pdf.

Anonymous:1879:BPbc

[13059] Anonymous. Business and personal. Scientific American, 41(2):27, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-27.pdf.

Anonymous:1879:BNBo

[13060] Anonymous. Books: New books and publications. Scientific American, 41(2):27, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
27a.pdf.

Anonymous:1879:NQv

[13061] Anonymous. Notes and queries. Scientific American, 41(2):27–28, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-27b.pdf.

Anonymous:1879:IIbc

[13062] Anonymous. Index of inventions. Scientific American, 41(2):28, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-28a.pdf.

Anonymous:1879:PNPv

[13063] Anonymous. Patent notes: Patents. Scientific American, 41(2):30, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-30.pdf.

Anonymous:1879:IHC

[13064] Anonymous. Improved hose coupling. Scientific American, 41(2):15, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-15.pdf.

Anonymous:1879:NGE

[13065] Anonymous. A new gas exhauster governor. Scientific American, 41(2):15, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
15a.pdf.

Anonymous:1879:NN

[13066] Anonymous. The North-Northwest. Scientific American, 41(2):16, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-16.pdf.

Anonymous:1879:PES

[13067] Anonymous. Pluck and energy sufficient capital for pioneers. Scientific American, 41(2):16, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
16a.pdf.

Anonymous:1879:CDW

[13068] Anonymous. Contamination of drinking water. Scientific American, 41(2):16, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
16b.pdf.

Anonymous:1879:PGS

[13069] Anonymous. Pah-Gosa Springs, Colorado. Scientific American, 41(2):16–17, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
16c.pdf.

Anonymous:1879:DCP

[13070] Anonymous. Defeat of the cochrane patents. Scientific American, 41(2):17, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
17.pdf.

Anonymous:1879:RAPm

[13071] Anonymous. Recent American patents. Scientific American, 41(2):18, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-18.pdf.

Anonymous:1879:MW

[13072] Anonymous. A magnetized watch. Scientific American, 41(2):19, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-19.pdf.

Anonymous:1879:NSTc

[13073] Anonymous. A new sewer trap. Scientific American, 41(2):19, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-19a.pdf.

Anonymous:1879:RMIk

[13074] Anonymous. Recent mechanical inventions. Scientific American, 41(2):19, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
19b.pdf.

Anonymous:1879:NPE

[13075] Anonymous. A new portable engine. Scientific American, 41(2):19, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-19c.pdf.

Anonymous:1879:NCH

[13076] Anonymous. New cuff holder. Scientific American, 41(2):19–20, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-19d.pdf.

Anonymous:1879:AMg

[13077] Anonymous. Amateur mechanics. Scientific American, 41(2):20–21, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-20.pdf.

Anonymous:1879:NDSb

[13078] Anonymous. New drag sawing machine. Scientific American, 41(2):21, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-21.pdf.

Anonymous:1879:MIl

[13079] Anonymous. Miscellaneous inventions. Scientific American, 41(2):22, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-22.pdf.

Anonymous:1879:MMS

[13080] Anonymous. Machinery for making shingles. Scientific American, 41(2):22, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
22a.pdf.

Anonymous:1879:JNP

[13081] Anonymous. Jones’ new press machine and process. Scientific American, 41(2):22–23, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n2/pdf/scientificamerican07121879-
22b.pdf.

Anonymous:1879:M

[13082] Anonymous. The mara. Scientific American, 41(2):23, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-23.pdf.

Beard:1879:SQG

[13083] Daniel C. Beard. Scarus Quacamaia, or great American parrot fish. Scientific American, 41(2):23–25, July 12, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n2/pdf/scientificamerican07121879-23a.pdf.

Anonymous:1879:AK

[13084] Anonymous. Ancient keys. Scientific American, 41(2):25–26, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-25.pdf.

Anonymous:1879:CRba

[13085] Anonymous. Communications received. Scientific American, 41(2):28, July 12, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n2/pdf/scientificamerican07121879-28.pdf.

Anonymous:1879:BPbd

[13086] Anonymous. Business and personal. Scientific American, 41(3):43, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-43.pdf.

Anonymous:1879:BNBp

[13087] Anonymous. Books: New books and publications. Scientific American, 41(3):43, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
43a.pdf.

Anonymous:1879:NQw

[13088] Anonymous. Notes and queries. Scientific American, 41(3):43–44, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-43b.pdf.

Anonymous:1879:Iz

[13089] Anonymous. Index. Scientific American, 41(3):44, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n3/pdf/scientificamerican07191879-44a.pdf.

Anonymous:1879:PNPw

[13090] Anonymous. Patent notes: Patents. Scientific American, 41(3):46, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-46.pdf.

Anonymous:1879:USD

[13091] Anonymous. The United States drilling scow, East River. Scientific American, 41(3):31, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
31.pdf.

Anonymous:1879:ABP

[13092] Anonymous. Alum in baking powders. Scientific American, 41(3):32, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-32.pdf.

Anonymous:1879:CBC

[13093] Anonymous. The captive balloon at coney Island. Scientific American, 41(3):32, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
32a.pdf.

Anonymous:1879:TLI

[13094] Anonymous. The telephone as a lightning indicator. Scientific American, 41(3):32–33, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
32b.pdf.

Anonymous:1879:PSW

[13095] Anonymous. Peculiar steam whistling. Scientific American, 41(3):33, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-33.pdf.

Anonymous:1879:MCNd

[13096] Anonymous. Molecular chemistry. — no. 4. Scientific American, 41(3):33, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
33a.pdf.

Anonymous:1879:ITC

[13097] Anonymous. An improved tug coupling. Scientific American, 41(3):34, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-34.pdf.

Anonymous:1879:NWMa

[13098] Anonymous. A new water meter. Scientific American, 41(3):34–35, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-34a.pdf.

Anonymous:1879:IWW

[13099] Anonymous. Improved water wheel. Scientific American, 41(3):35, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-35.pdf.

Anonymous:1879:IWV

[13100] Anonymous. Improved washbasin valves. Scientific American, 41(3):35, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-35a.pdf.

Anonymous:1879:RAPn

[13101] Anonymous. Recent American patents. Scientific American, 41(3):35, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-35b.pdf.

Anonymous:1879:RMIl

[13102] Anonymous. Recent mechanical inventions. Scientific American, 41(3):36, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
36.pdf.

Anonymous:1879:RNd

[13103] Anonymous. Railway notes. Scientific American, 41(3):36, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-36a.pdf.

Anonymous:1879:GA

[13104] Anonymous. The geysers of the Azores. Scientific American, 41(3):37, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-37.pdf.

Anonymous:1879:NPA

[13105] Anonymous. New propelling apparatus. Scientific American, 41(3):38, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-38.pdf.

Anonymous:1879:NMA

[13106] Anonymous. New milling attachment for lathes. Scientific American, 41(3):38–39, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
38a.pdf.

Anonymous:1879:MIm

[13107] Anonymous. Miscellaneous inventions. Scientific American, 41 (3):39–40, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
39.pdf.

Anonymous:1879:NSGa

[13108] Anonymous. New sewer gas stopper. Scientific American, 41(3):40, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-40.pdf.

Anonymous:1879:PEb

[13109] Anonymous. Plowing by electricity. Scientific American, 41(3):41, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-41.pdf.

Anonymous:1879:NAIi

[13110] Anonymous. New agricultural inventions. Scientific American, 41(3):41–42, July 19, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n3/pdf/scientificamerican07191879-
41a.pdf.

Anonymous:1879:CRbb

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n3/pdf/scientificamerican07191879-44.pdf.

Anonymous:1879:Cz

[13112] Anonymous. Correspondence. Scientific American, 41(4):56–57, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-56.pdf.

Anonymous:1879:BPbe

[13113] Anonymous. Business and personal. Scientific American, 41(4):59, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-59.pdf.

Anonymous:1879:BNBq

[13114] Anonymous. Books: New books and publications. Scientific American, 41(4):59, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
59a.pdf.

Anonymous:1879:NQx

[13115] Anonymous. Notes and queries. Scientific American, 41(4):59–60, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-59b.pdf.

Anonymous:1879:IIbd

[13116] Anonymous. Index of inventions. Scientific American, 41(4):60, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-60a.pdf.

Anonymous:1879:PNPx

[13117] Anonymous. Patent notes: Patents. Scientific American, 41(4):62, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-62.pdf.

Anonymous:1879:ELPb

[13118] Anonymous. The electric light in photography. Scientific American, 41(4):47, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
47.pdf.

Anonymous:1879:APE

[13119] Anonymous. The American polar expedition. Scientific American, 41(4):48, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
48.pdf.

Anonymous:1879:PEW

[13120] Anonymous. Proposed exploration of Western Asia. Scientific American, 41(4):48, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
48a.pdf.

Anonymous:1879:FIA

[13121] Anonymous. Flint implements of the Aborigines. Scientific American, 41(4):48–49, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
48b.pdf.

Anonymous:1879:TF

[13122] Anonymous. A tropical fruit. Scientific American, 41(4):49, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-49.pdf.

Anonymous:1879:FFb

[13123] Anonymous. Fire-flies. Scientific American, 41(4):49, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-49a.pdf.

Anonymous:1879:NSCc

[13124] Anonymous. A novel steam car. Scientific American, 41(4):50, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-50.pdf.

Anonymous:1879:NBMb

[13125] Anonymous. A new brick machine. Scientific American, 41(4):50, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-50a.pdf.

Anonymous:1879:RAPo

[13126] Anonymous. Recent American patents. Scientific American, 41 (4):50–51, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
50b.pdf.

Anonymous:1879:ESE

[13127] Anonymous. The exeter steam engine. Scientific American, 41(4):51–52, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-51.pdf.

Anonymous:1879:PMP

[13128] Anonymous. Progress at Menlo Park. Scientific American, 41(4):52, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-52.pdf.

Anonymous:1879:SRE

[13129] Anonymous. The Sun’s radiant energy. Scientific American, 41 (4):53–54, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n4/pdf/scientificamerican07261879-
53.pdf.

Anonymous:1879:AMh

[13130] Anonymous. Amateur mechanics. Scientific American, 41(4):54, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-54.pdf.

Anonymous:1879:IMe

[13131] Anonymous. An improved mill. Scientific American, 41(4):54–55, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-54a.pdf.

Anonymous:1879:Af

[13132] Anonymous. The Aswail. Scientific American, 41(4):55, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-55.pdf.

Anonymous:1879:NHNd

[13133] Anonymous. Natural history notes. Scientific American, 41(4):55–56, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-55a.pdf.

Anonymous:1879:CDR

[13134] Anonymous. Curious discoveries in regard to the manner of making flint implements by the Aborigines and prehistoric inhabitants of America. Scientific American, 41(4):57–58, July 26, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n4/pdf/scientificamerican07261879-57.pdf.

Anonymous:1879:EIh

[13135] Anonymous. Engineering inventions. Scientific American, 41(4):58, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-58.pdf.

Anonymous:1879:NIBb

[13136] Anonymous. New ice breaker. Scientific American, 41(4):58, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-58a.pdf.

Anonymous:1879:CRbc

[13137] Anonymous. Communications received. Scientific American, 41(4):60, July 26, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n4/pdf/scientificamerican07261879-60.pdf.

Anonymous:1879:BPbf

[13138] Anonymous. Business and personal. Scientific American, 41(5):75, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-75.pdf.

Anonymous:1879:BNBr

[13139] Anonymous. Books: New books and publications. Scientific American, 41(5):75, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
75a.pdf.

Anonymous:1879:NQy

[13140] Anonymous. Notes and queries. Scientific American, 41(5):75, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-75b.pdf.

Anonymous:1879:IIbe

[13141] Anonymous. Index of inventions. Scientific American, 41(5):75–76, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-75c.pdf.

Anonymous:1879:PNPy

[13142] Anonymous. Patent notes: Patents. Scientific American, 41(5):78, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-78.pdf.

Anonymous:1879:GPD

[13143] Anonymous. The grand prize diploma of the French exhibition. Scientific American, 41(5):63, August 2, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n5/pdf/scientificamerican08021879-63.pdf.

Anonymous:1879:SWC

[13144] Anonymous. Shall we have a canal or ship railway? Scientific American, 41(5):64, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
64.pdf.

Anonymous:1879:CIE

[13145] Anonymous. The Cincinnati Industrial Exhibition. Scientific American, 41(5):64, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
64a.pdf.

Anonymous:1879:OM

[13146] Anonymous. The opening of the Mississippi. Scientific American, 41(5):64, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
64b.pdf.

Anonymous:1879:RML

[13147] Anonymous. The reflecting magic lantern in court. Scientific American, 41(5):64–65, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
64c.pdf.

Anonymous:1879:PS

[13148] Anonymous. Progress in spectroscopy. Scientific American, 41(5):65, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-65.pdf.

Anonymous:1879:WSC

[13149] Anonymous. Where to study chemistry in Germany. Scientific American, 41(5):65, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
65a.pdf.

Anonymous:1879:PPc

[13150] Anonymous. The polar pantograph. Scientific American, 41(5):66, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-66.pdf.

Anonymous:1879:CFS

[13151] Anonymous. Combined fork and spoon. Scientific American, 41(5):66, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-66a.pdf.

Anonymous:1879:IPD

[13152] Anonymous. An improved potato digger. Scientific American, 41(5):66, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-66b.pdf.

Anonymous:1879:RMIm

[13153] Anonymous. Recent mechanical inventions. Scientific American, 41(5):66, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
66c.pdf.

Langley:1879:SSP

[13154] S. P. Langley. The Sun a source of power. Scientific American, 41(5):67–68, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
67.pdf.

Anonymous:1879:SRP

[13155] Anonymous. A ship railway at Panama. Scientific American, 41(5):68, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-68.pdf.

Anonymous:1879:STV

[13156] Anonymous. Spar torpedo vessel Acheron. Scientific American, 41(5):69, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
69.pdf.

Anonymous:1879:PFM

[13157] Anonymous. Portable flour mill. Scientific American, 41(5):70, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-70.pdf.

Anonymous:1879:NTMb

[13158] Anonymous. New thrashing machine. Scientific American, 41(5):70–71, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-70a.pdf.

Anonymous:1879:Ta

[13159] Anonymous. Tornelia. Scientific American, 41(5):71, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-71.pdf.

Anonymous:1879:NHNe

[13160] Anonymous. Natural history notes. Scientific American, 41(5):71, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-71a.pdf.

Anonymous:1879:PFb

[13161] Anonymous. The phylloxera in France. Scientific American, 41 (5):72–73, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
72.pdf.

Anonymous:1879:NBBb

[13162] Anonymous. New bubble blower. Scientific American, 41(5):73, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n5/pdf/scientificamerican08021879-73.pdf.

Anonymous:1879:FAE

[13163] Anonymous. Future of American engineering. Scientific American, 41(5):74, August 2, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n5/pdf/scientificamerican08021879-
74.pdf.

Anonymous:1879:Cbd

[13164] Anonymous. Correspondence. Scientific American, 41(6):84, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-84.pdf.

Anonymous:1879:BPbg

[13165] Anonymous. Business and personal. Scientific American, 41(6):91, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-91.pdf.

Anonymous:1879:BNBs

[13166] Anonymous. Books: New books and publications. Scientific American, 41(6):91, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
91a.pdf.

Anonymous:1879:NQz

[13167] Anonymous. Notes and queries. Scientific American, 41(6):91–92, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-91b.pdf.

Anonymous:1879:IIbf

[13168] Anonymous. Index of inventions. Scientific American, 41(6):92, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-92a.pdf.

Anonymous:1879:PNPz

[13169] Anonymous. Patent notes: Patents. Scientific American, 41(6):94, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-94.pdf.

Anonymous:1879:MV

[13170] Anonymous. The manufacture of valves. Scientific American, 41(6):79–80, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
79.pdf.

Anonymous:1879:WDN

[13171] Anonymous. What to do next winter. Scientific American, 41(6):80–81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-80.pdf.

Anonymous:1879:PON

[13172] Anonymous. Probable opportunities for new discoveries in acoustics. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
81.pdf.

Anonymous:1879:NBN

[13173] Anonymous. The new Bureau of National Survey. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
81a.pdf.

Anonymous:1879:HPCb

[13174] Anonymous. How to preserve cider. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-81b.pdf.

Anonymous:1879:CPa

[13175] Anonymous. Changes in photography. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-81c.pdf.

Anonymous:1879:SPc

[13176] Anonymous. Spirit photographs. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-81d.pdf.

Anonymous:1879:EAC

[13177] Anonymous. The electrical alarm compass. Scientific American, 41(6):81, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
81e.pdf.

Anonymous:1879:IPM

[13178] Anonymous. Improved pressing machine. Scientific American, 41(6):82, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-82.pdf.

Anonymous:1879:IGM

[13179] Anonymous. Improved grinding mill. Scientific American, 41(6):83, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-83.pdf.

Anonymous:1879:NRTb

[13180] Anonymous. A new railroad tie. Scientific American, 41(6):83, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-83a.pdf.

Anonymous:1879:EIi

[13181] Anonymous. Engineering inventions. Scientific American, 41(6):83–84, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-83b.pdf.

Anonymous:1879:NAIj

[13182] Anonymous. New agricultural inventions. Scientific American, 41(6):84–85, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
84a.pdf.

Anonymous:1879:NOD

[13183] Anonymous. A new optical delusion. Scientific American, 41(6):85, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-85.pdf.

Anonymous:1879:ISG

[13184] Anonymous. Improvement in saw gumming. Scientific American, 41(6):86, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
86.pdf.

Anonymous:1879:INL

[13185] Anonymous. Improved nut lock. Scientific American, 41(6):86, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-86a.pdf.

Anonymous:1879:N

[13186] Anonymous. The narica. Scientific American, 41(6):87–88, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-87.pdf.

Anonymous:1879:MIn

[13187] Anonymous. Miscellaneous inventions. Scientific American, 41 (6):88–90, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
88.pdf.

Anonymous:1879:RMIn

[13188] Anonymous. Recent mechanical inventions. Scientific American, 41(6):90, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n6/pdf/scientificamerican08091879-
90.pdf.

Anonymous:1879:CRbd

[13189] Anonymous. Communications received. Scientific American, 41(6):92, August 9, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n6/pdf/scientificamerican08091879-92.pdf.

Anonymous:1879:Cbe

[13190] Anonymous. Correspondence. Scientific American, 41(7):100, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-100.pdf.

Anonymous:1879:BPbh

[13191] Anonymous. Business and personal. Scientific American, 41(7):107, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-107.pdf.

Anonymous:1879:BNBt

[13192] Anonymous. Books: New books and publications. Scientific American, 41(7):107, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
107a.pdf.

Anonymous:1879:NQba

[13193] Anonymous. Notes and queries. Scientific American, 41(7):107– 108, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
107b.pdf.

Anonymous:1879:IIbg

[13194] Anonymous. Index of inventions. Scientific American, 41(7):108, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-108a.pdf.

Anonymous:1879:PNPba

[13195] Anonymous. Patent notes: Patents. Scientific American, 41(7):110, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-110.pdf.

Anonymous:1879:AMi

[13196] Anonymous. Amateur mechanics. Scientific American, 41(7): 95–96, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
95.pdf.

Anonymous:1879:AAA

[13197] Anonymous. The American Association for the Advancement of Science. Scientific American, 41(7):96–97, August 16, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n7/pdf/scientificamerican08161879-96.pdf.

Anonymous:1879:SW

[13198] Anonymous. Sandal wood. Scientific American, 41(7):97, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-97.pdf.

Anonymous:1879:SSb

[13199] Anonymous. The Système Sèbillot. Scientific American, 41(7):97, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-97a.pdf.

Anonymous:1879:SCEb

[13200] Anonymous. A sanitary Captain Eads wanted. Scientific American, 41(7):97–98, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
97b.pdf.

Anonymous:1879:NPR

[13201] Anonymous. A new piston rod packing. Scientific American, 41 (7):98, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
98.pdf.

Anonymous:1879:ICb

[13202] Anonymous. An improved churn. Scientific American, 41(7):98, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-98a.pdf.

Anonymous:1879:EIj

[13203] Anonymous. Engineering inventions. Scientific American, 41(7):98, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-98b.pdf.

Anonymous:1879:RAIb

[13204] Anonymous. Recent agricultural inventions. Scientific American, 41(7):99, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
99.pdf.

Hopkins:1879:NM

[13205] Geo. M. Hopkins. New magnetometer. Scientific American, 41 (7):99–100, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
99a.pdf.

Anonymous:1879:NEC

[13206] Anonymous. New electric current regulator. Scientific American, 41(7):101, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
101.pdf.

Anonymous:1879:RMIo

[13207] Anonymous. Recent mechanical inventions. Scientific American, 41(7):101, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
101a.pdf.

Anonymous:1879:NCe

[13208] Anonymous. New cut-off. Scientific American, 41(7):102, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-102.pdf.

Anonymous:1879:CSI

[13209] Anonymous. Combined sad-iron and fluting roller. Scientific American, 41(7):102, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
102a.pdf.

Anonymous:1879:MIo

[13210] Anonymous. Miscellaneous inventions. Scientific American, 41 (7):102, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
102b.pdf.

Anonymous:1879:F

[13211] Anonymous. Faience. Scientific American, 41(7):103–106, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n7/pdf/scientificamerican08161879-103.pdf.

Anonymous:1879:CRbe

[13212] Anonymous. Communications received. Scientific American, 41 (7):108, August 16, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n7/pdf/scientificamerican08161879-
108.pdf.

Anonymous:1879:Cbf

[13213] Anonymous. Correspondence. Scientific American, 41(8):116, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-116.pdf.

Anonymous:1879:BPbi

[13214] Anonymous. Business and personal. Scientific American, 41(8):123, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-123.pdf.

Anonymous:1879:BNBu

[13215] Anonymous. Books: New books and publications. Scientific American, 41(8):123, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
123a.pdf.

Anonymous:1879:NQbb

[13216] Anonymous. Notes and queries. Scientific American, 41(8):123– 124, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
123b.pdf.

Anonymous:1879:IIbh

[13217] Anonymous. Index of inventions. Scientific American, 41(8):124, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-124a.pdf.

Anonymous:1879:PNPbb

[13218] Anonymous. Patent notes: Patents. Scientific American, 41(8):126, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-126.pdf.

Anonymous:1879:NMCa

[13219] Anonymous. New meridian circle. Scientific American, 41(8):111, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-111.pdf.

Anonymous:1879:SIb

[13220] Anonymous. A subject for investigation. Scientific American, 41(8):112, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
112.pdf.

Anonymous:1879:SYF

[13221] Anonymous. Stillè on yellow fever. Scientific American, 41(8):112, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-112a.pdf.

Anonymous:1879:NML

[13222] Anonymous. A new method of locating lightning rods. Scientific American, 41(8):112–113, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
112b.pdf.

Anonymous:1879:LHA

[13223] Anonymous. Labor at home and abroad. Scientific American, 41(8):113, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
113.pdf.

Anonymous:1879:DF

[13224] Anonymous. Dragon flies. Scientific American, 41(8):113, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-113a.pdf.

Anonymous:1879:AOS

[13225] Anonymous. The American onion-smut in France. Scientific American, 41(8):113, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
113b.pdf.

Anonymous:1879:NEa

[13226] Anonymous. Nordenskjold’s expedition. Scientific American, 41 (8):113, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
113c.pdf.

Anonymous:1879:HSS

[13227] Anonymous. The Holly system of steam heating. Scientific American, 41(8):114–116, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
114.pdf.

Anonymous:1879:MIp

[13228] Anonymous. Miscellaneous inventions. Scientific American, 41 (8):116–117, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
116a.pdf.

Anonymous:1879:ISW

[13229] Anonymous. Improved steam whistle. Scientific American, 41 (8):117, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
117.pdf.

Anonymous:1879:LLP

[13230] Anonymous. Lock letter press. Scientific American, 41(8):117, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n8/pdf/scientificamerican08231879-117a.pdf.

Anonymous:1879:NWT

[13231] Anonymous. Novel work table implement. Scientific American, 41(8):118, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
118.pdf.

Anonymous:1879:IPRb

[13232] Anonymous. Improved portable railroad lock. Scientific American, 41(8):118, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
118a.pdf.

Anonymous:1879:ISCa

[13233] Anonymous. Improved straw cutter. Scientific American, 41 (8):118, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
118b.pdf.

Anonymous:1879:RMIp

[13234] Anonymous. Recent mechanical inventions. Scientific American, 41(8):118, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
118c.pdf.

Anonymous:1879:FD

[13235] Anonymous. The flying dragon. Scientific American, 41(8):119– 120, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
119.pdf.

Anonymous:1879:MCNe

[13236] Anonymous. Molecular chemistry — no. 5. Scientific American, 41(8):120–121, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
120.pdf.

Anonymous:1879:EIk

[13237] Anonymous. Engineering inventions. Scientific American, 41(8): 121, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
121.pdf.

Anonymous:1879:RDRa

[13238] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(8):122, August 23, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n8/pdf/scientificamerican08231879-122.pdf.

Anonymous:1879:CRbf

[13239] Anonymous. Communications received. Scientific American, 41 (8):124, August 23, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n8/pdf/scientificamerican08231879-
124.pdf.

Anonymous:1879:Cbg

[13240] Anonymous. Correspondence. Scientific American, 41(9):133, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-133.pdf.

Anonymous:1879:BPbj

[13241] Anonymous. Business and personal. Scientific American, 41(9):139, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-139.pdf.

Anonymous:1879:BNBv

[13242] Anonymous. Books: New books and publications. Scientific American, 41(9):139, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
139a.pdf.

Anonymous:1879:NQbc

[13243] Anonymous. Notes and queries. Scientific American, 41(9):139– 140, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
139b.pdf.

Anonymous:1879:IIbi

[13244] Anonymous. Index of inventions. Scientific American, 41(9):140, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-140a.pdf.

Anonymous:1879:PNPbc

[13245] Anonymous. Patent notes: Patents. Scientific American, 41(9):142, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-142.pdf.

Anonymous:1879:BS

[13246] Anonymous. Bessemer steel. Scientific American, 41(9):127, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-127.pdf.

Anonymous:1879:BTM

[13247] Anonymous. Balance of trade in men. Scientific American, 41 (9):128, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
128.pdf.

Anonymous:1879:LP

[13248] Anonymous. A little paradox. Scientific American, 41(9):128, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-128a.pdf.

Anonymous:1879:FM

[13249] Anonymous. Fungi in man. Scientific American, 41(9):128, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-128b.pdf.

Riley:1879:FTR

[13250] C. V. Riley. Failure of tea roses. Scientific American, 41(9):129, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-129.pdf.

Anonymous:1879:NPSb

[13251] Anonymous. New propeller screw. Scientific American, 41(9):130, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-130.pdf.

Anonymous:1879:TGb

[13252] Anonymous. The Thunderer’s gun. Scientific American, 41(9):130, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-130a.pdf.

Anonymous:1879:Cbh

[13253] Anonymous. The cycloscope. Scientific American, 41(9):131, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-131.pdf.

Anonymous:1879:AWC

[13254] Anonymous. Arsenic in water colors. Scientific American, 41 (9):131–132, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
131a.pdf.

Anonymous:1879:MIq

[13255] Anonymous. Mechanical inventions. Scientific American, 41(9): 132–133, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
132.pdf.

Anonymous:1879:NMCb

[13256] Anonymous. A new miter cutter. Scientific American, 41(9):134, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-134.pdf.

Anonymous:1879:SSd

[13257] Anonymous. Screw scaffolding. Scientific American, 41(9):134, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-134a.pdf.

Anonymous:1879:MIr

[13258] Anonymous. Miscellaneous inventions. Scientific American, 41 (9):134, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
134b.pdf.

Anonymous:1879:JGA

[13259] Anonymous. The Japanese goat antelope. Scientific American, 41(9):135, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
135.pdf.

Anonymous:1879:GTb

[13260] Anonymous. Giant tree. Scientific American, 41(9):135, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n9/pdf/scientificamerican08301879-135a.pdf.

Anonymous:1879:NHNf

[13261] Anonymous. Natural history notes. Scientific American, 41(9): 135–136, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
135b.pdf.

Anonymous:1879:EIl

[13262] Anonymous. Engineering inventions. Scientific American, 41(9): 137–138, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
137.pdf.

Anonymous:1879:RDRb

[13263] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(9):138, August 30, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n9/pdf/scientificamerican08301879-138.pdf.

Anonymous:1879:CRbg

[13264] Anonymous. Communications received. Scientific American, 41 (9):140, August 30, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n9/pdf/scientificamerican08301879-
140.pdf.

Anonymous:1879:Cbi

[13265] Anonymous. Correspondence. Scientific American, 41(10):148, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n10/pdf/scientificamerican09061879-148.pdf.

Anonymous:1879:BPbk

[13266] Anonymous. Business and personal. Scientific American, 41 (10):155, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
155.pdf.

Anonymous:1879:BNBw

[13267] Anonymous. Books: New books and publications. Scientific American, 41(10):155, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
155a.pdf.

Anonymous:1879:NQbd

[13268] Anonymous. Notes and queries. Scientific American, 41(10): 155, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
155b.pdf.

Anonymous:1879:IIbj

[13269] Anonymous. Index of inventions. Scientific American, 41(10): 155–156, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
155d.pdf.

Anonymous:1879:PNPbd

[13270] Anonymous. Patent notes: Patents. Scientific American, 41 (10):158, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
158.pdf.

Anonymous:1879:BSCa

[13271] Anonymous. The Belgian system of canal towing. Scientific American, 41(10):143, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
143.pdf.

Anonymous:1879:IWN

[13272] Anonymous. Intelligent workmen needed. Scientific American, 41 (10):144, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
144.pdf.

Anonymous:1879:HST

[13273] Anonymous. Hydrophobia successfully treated with curare. Scientific American, 41(10):144, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
144a.pdf.

Anonymous:1879:EAS

[13274] Anonymous. Early advocates of ship railways. Scientific American, 41(10):144, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
144b.pdf.

Anonymous:1879:AVFa

[13275] Anonymous. American vines in France. Scientific American, 41 (10):144–145, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
144c.pdf.

Anonymous:1879:RSH

[13276] Anonymous. Reading at seven and a half miles distance from the candle. Scientific American, 41(10):145, September 6, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n10/pdf/scientificamerican09061879-145.pdf.

Anonymous:1879:CWC

[13277] Anonymous. A covering wanted for cotton bales. Scientific American, 41(10):145, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
145a.pdf.

Anonymous:1879:SMb

[13278] Anonymous. A singular memory. Scientific American, 41(10): 145, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
145b.pdf.

Anonymous:1879:FHD

[13279] Anonymous. Four hours in the dark. Scientific American, 41 (10):145, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
145c.pdf.

Anonymous:1879:MIs

[13280] Anonymous. Miscellaneous inventions. Scientific American, 41 (10):146, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
146.pdf.

Anonymous:1879:TP

[13281] Anonymous. Tubular piles. Scientific American, 41(10):147, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n10/pdf/scientificamerican09061879-147.pdf.

Anonymous:1879:IRa

[13282] Anonymous. Improved rheostat. Scientific American, 41(10): 147, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
147a.pdf.

Anonymous:1879:IIRa

[13283] Anonymous. Improved instrument for ringing swine. Scientific American, 41(10):147, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
147b.pdf.

Anonymous:1879:CPC

[13284] Anonymous. Cutting packing company. — one of the largest establishments of the kind in the country. Scientific American, 41 (10):148–149, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
148a.pdf.

Anonymous:1879:NIF

[13285] Anonymous. New iron fence post. Scientific American, 41(10): 149, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
149.pdf.

Anonymous:1879:IWGb

[13286] Anonymous. Iron and wood gear-dressing machine. Scientific American, 41(10):150, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
150.pdf.

Anonymous:1879:IBF

[13287] Anonymous. Improved boiler feed pump. Scientific American, 41 (10):150–151, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
150a.pdf.

Anonymous:1879:STPb

[13288] Anonymous. The swelled trunk palm. Scientific American, 41 (10):151, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
151.pdf.

Anonymous:1879:MIt

[13289] Anonymous. Mechanical inventions. Scientific American, 41(10): 151–153, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
151a.pdf.

Anonymous:1879:EIm

[13290] Anonymous. Engineering inventions. Scientific American, 41(10): 153–154, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
153.pdf.

Anonymous:1879:CRbh

[13291] Anonymous. Communications received. Scientific American, 41 (10):155, September 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n10/pdf/scientificamerican09061879-
155c.pdf.

Anonymous:1879:Cbj

[13292] Anonymous. Correspondence. Scientific American, 41(11):165– 166, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
165.pdf.

Anonymous:1879:BPbl

[13293] Anonymous. Business and personal. Scientific American, 41 (11):171, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
171.pdf.

Anonymous:1879:BNBx

[13294] Anonymous. Books: New books and publications. Scientific American, 41(11):171, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
171a.pdf.

Anonymous:1879:NQbe

[13295] Anonymous. Notes and queries. Scientific American, 41(11): 171–172, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
171b.pdf.

Anonymous:1879:IIbk

[13296] Anonymous. Index of inventions. Scientific American, 41(11): 172, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
172a.pdf.

Anonymous:1879:PNPbe

[13297] Anonymous. Patent notes: Patents. Scientific American, 41 (11):174, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
174.pdf.

Anonymous:1879:L

[13298] Anonymous. Lenses. Scientific American, 41(11):159, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n11/pdf/scientificamerican09131879-159.pdf.

Anonymous:1879:AW

[13299] Anonymous. The American way. Scientific American, 41(11): 160, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
160.pdf.

Anonymous:1879:PPS

[13300] Anonymous. The proposed Panama ship-railway. Scientific American, 41(11):160, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
160a.pdf.

Anonymous:1879:RIP

[13301] Anonymous. Recent industrial progress. Scientific American, 41 (11):160–161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
160b.pdf.

Anonymous:1879:CPD

[13302] Anonymous. The Commissioner of Patents in a dilemma. Scientific American, 41(11):161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161.pdf.

Anonymous:1879:RSR

[13303] Anonymous. The red spider on roses. Scientific American, 41 (11):161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161a.pdf.

Anonymous:1879:RP

[13304] Anonymous. Rapid photographing. Scientific American, 41(11): 161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161b.pdf.

Anonymous:1879:CBIa

[13305] Anonymous. Curing beef by injecting brine. Scientific American, 41(11):161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161c.pdf.

Anonymous:1879:LLL

[13306] Anonymous. Lift locks and locking-time. Scientific American, 41 (11):161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161d.pdf.

Anonymous:1879:ASA

[13307] Anonymous. The American science association. Scientific American, 41(11):161, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
161e.pdf.

Anonymous:1879:DM

[13308] Anonymous. Drying machines. Scientific American, 41(11):162– 163, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
162.pdf.

Anonymous:1879:EIn

[13309] Anonymous. Engineering inventions. Scientific American, 41(11): 163, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
163.pdf.

Anonymous:1879:SLT

[13310] Anonymous. Self-laying tramway. Scientific American, 41(11): 163, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
163a.pdf.

Anonymous:1879:NMT

[13311] Anonymous. A new metallic thermometer. Scientific American, 41 (11):163–164, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
163b.pdf.

Anonymous:1879:MIu

[13312] Anonymous. Mechanical inventions. Scientific American, 41(11): 164, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
164.pdf.

Anonymous:1879:RAIc

[13313] Anonymous. Recent agricultural inventions. Scientific American, 41(11):164, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
164a.pdf.

Anonymous:1879:NPg

[13314] Anonymous. A new press. Scientific American, 41(11):166, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n11/pdf/scientificamerican09131879-166.pdf.

Anonymous:1879:NFEa

[13315] Anonymous. New form of electric lamp. Scientific American, 41 (11):166, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
166a.pdf.

Anonymous:1879:NDT

[13316] Anonymous. Novel device for transmitting motion. Scientific American, 41(11):166–167, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
166b.pdf.

Anonymous:1879:GWG

[13317] Anonymous. The Greenland whale and the Grampus. Scientific American, 41(11):167–168, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
167.pdf.

Anonymous:1879:MIv

[13318] Anonymous. Miscellaneous inventions. Scientific American, 41(11): 168–169, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
168.pdf.

Anonymous:1879:AVA

[13319] Anonymous. An ascent of the volcano of Agua, Central America. Scientific American, 41(11):170, September 13, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n11/pdf/scientificamerican09131879-170.pdf.

Anonymous:1879:CRbi

[13320] Anonymous. Communications received. Scientific American, 41 (11):172, September 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n11/pdf/scientificamerican09131879-
172.pdf.

Anonymous:1879:BPbm

[13321] Anonymous. Business and personal. Scientific American, 41 (12):187, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
187.pdf.

Anonymous:1879:BNBy

[13322] Anonymous. Books: New books and publications. Scientific American, 41(12):187, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
187a.pdf.

Anonymous:1879:NQbf

[13323] Anonymous. Notes and queries. Scientific American, 41(12): 187–188, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
187b.pdf.

Anonymous:1879:IIbl

[13324] Anonymous. Index of inventions. Scientific American, 41(12): 188, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
188a.pdf.

Anonymous:1879:PNPbf

[13325] Anonymous. Patent notes: Patents. Scientific American, 41 (12):190, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
190.pdf.

Anonymous:1879:MHC

[13326] Anonymous. The manufacture of the Harris–Corliss engine. Scientific American, 41(12):175–176, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.

nature.com/scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
175.pdf.

Anonymous:1879:AAb

[13327] Anonymous. The American Association. Scientific American, 41 (12):176, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
176.pdf.

Anonymous:1879:PYI

[13328] Anonymous. Patent your inventions abroad. Scientific American, 41 (12):176–177, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
176a.pdf.

Anonymous:1879:GN

[13329] Anonymous. Gelatine negatives. Scientific American, 41(12): 177, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
177.pdf.

Anonymous:1879:DC

[13330] Anonymous. The dissociation of chlorine. Scientific American, 41 (12):177, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
177a.pdf.

Anonymous:1879:NPSc

[13331] Anonymous. The Northwest Passage successfully made. Scientific American, 41(12):177, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
177b.pdf.

Anonymous:1879:PUS

[13332] Anonymous. Platinum in the United States. Scientific American, 41 (12):177–178, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
177c.pdf.

Anonymous:1879:SMC

[13333] Anonymous. Skimming milk by centrifugal action. Scientific American, 41(12):178–179, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
178.pdf.

Anonymous:1879:NPSd

[13334] Anonymous. A new punching and shearing press. Scientific American, 41(12):179, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
179.pdf.

Anonymous:1879:NFCb

[13335] Anonymous. New flange coupling. Scientific American, 41(12): 179, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
179a.pdf.

Anonymous:1879:FWY

[13336] Anonymous. A few words to Young steam fitters. Scientific American, 41(12):179–180, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
179b.pdf.

Anonymous:1879:PAA

[13337] Anonymous. Proceedings of the American Association. Scientific American, 41(12):180, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
180.pdf.

Anonymous:1879:AN

[13338] Anonymous. Astronomical notes. Scientific American, 41(12): 180, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
180a.pdf.

Anonymous:1879:MIw

[13339] Anonymous. Miscellaneous inventions. Scientific American, 41 (12):181, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
181.pdf.

Anonymous:1879:NEb

[13340] Anonymous. A novel exhibition. Scientific American, 41(12): 181–182, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
181a.pdf.

Anonymous:1879:NFT

[13341] Anonymous. New form of transit instrument. Scientific American, 41(12):182, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
182.pdf.

Anonymous:1879:PNB

[13342] Anonymous. Proposed new British polar expedition. Scientific American, 41(12):183, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
183.pdf.

Anonymous:1879:ISQ

[13343] Anonymous. Iodine as a substitute for quinine. Scientific American, 41(12):183–184, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
183a.pdf.

Anonymous:1879:TRA

[13344] Anonymous. Three rules for abbreviating multiplication. Scientific American, 41(12):184, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
184.pdf.

Anonymous:1879:EIo

[13345] Anonymous. Engineering inventions. Scientific American, 41(12): 184–185, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
184a.pdf.

Anonymous:1879:GOP

[13346] Anonymous. Gothic oak press. Scientific American, 41(12): 185, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
185.pdf.

Anonymous:1879:RAId

[13347] Anonymous. Recent agricultural inventions. Scientific American, 41 (12):185–186, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
185a.pdf.

Anonymous:1879:CRbj

[13348] Anonymous. Communications received. Scientific American, 41 (12):188, September 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n12/pdf/scientificamerican09201879-
188.pdf.

Anonymous:1879:BPbn

[13349] Anonymous. Business and personal. Scientific American, 41 (13):203, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
203.pdf.

Anonymous:1879:BNBz

[13350] Anonymous. Books: New books and publications. Scientific American, 41(13):203, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
203a.pdf.

Anonymous:1879:NQbg

[13351] Anonymous. Notes and queries. Scientific American, 41(13): 203–204, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
203b.pdf.

Anonymous:1879:IIbm

[13352] Anonymous. Index of inventions. Scientific American, 41(13): 204, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
204a.pdf.

Anonymous:1879:PNPbg

[13353] Anonymous. Patent notes: Patents. Scientific American, 41 (13):206, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
206.pdf.

Anonymous:1879:WC

[13354] Anonymous. A wonderful clock. Scientific American, 41(13): 191, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
191.pdf.

Anonymous:1879:TSN

[13355] Anonymous. Technical science in New Zealand. Scientific American, 41(13):192, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
192.pdf.

Anonymous:1879:BSI

[13356] Anonymous. The Bessemer steel interest. Scientific American, 41 (13):192, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
192a.pdf.

Anonymous:1879:ICd

[13357] Anonymous. Inventions in China. Scientific American, 41(13): 192–193, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
192b.pdf.

Anonymous:1879:CFb

[13358] Anonymous. Copyrighting fruit. Scientific American, 41(13): 193, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
193.pdf.

Anonymous:1879:SSAa

[13359] Anonymous. The social science association. Scientific American, 41(13):193, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
193a.pdf.

Anonymous:1879:RIA

[13360] Anonymous. Recent investigation into the action of anæsthetics. Scientific American, 41(13):193–194, September 27, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n13/pdf/scientificamerican09271879-193b.pdf.

Anonymous:1879:NFR

[13361] Anonymous. New form of Reynier’s electric lamp. Scientific American, 41(13):194, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
194.pdf.

Anonymous:1879:NCPc

[13362] Anonymous. A new combined punch and shear. Scientific American, 41(13):194, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
194a.pdf.

Anonymous:1879:MIx

[13363] Anonymous. Mechanical inventions. Scientific American, 41(13): 194, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
194b.pdf.

Mayer:1879:PEMa

[13364] Alfred M. Mayer. Practical experiments in magnetism, with special reference to the demagnetization of watches. — no. 1. Scientific American, 41(13):195–197, September 27, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n13/pdf/scientificamerican09271879-195.pdf.

Anonymous:1879:NSBb

[13365] Anonymous. A new surf boat. Scientific American, 41(13):198, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n13/pdf/scientificamerican09271879-198.pdf.

Anonymous:1879:ELT

[13366] Anonymous. Edison’s latest telephone. Scientific American, 41 (13):198, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
198a.pdf.

Anonymous:1879:MIy

[13367] Anonymous. Miscellaneous inventions. Scientific American, 41(13): 198–199, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
198b.pdf.

Anonymous:1879:GIb

[13368] Anonymous. Giant insects. Scientific American, 41(13):199– 200, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
199.pdf.

Anonymous:1879:RDRc

[13369] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(13):200–201, September 27, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n13/pdf/scientificamerican09271879-200.pdf.

Anonymous:1879:FSb

[13370] Anonymous. Fire screen. Scientific American, 41(13):201–202, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n13/pdf/scientificamerican09271879-201.pdf.

Anonymous:1879:CRbk

[13371] Anonymous. Communications received. Scientific American, 41 (13):204, September 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n13/pdf/scientificamerican09271879-
204.pdf.

Anonymous:1879:Cbk

[13372] Anonymous. Correspondence. Scientific American, 41(14):216, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-216.pdf.

Anonymous:1879:BPbo

[13373] Anonymous. Business and personal. Scientific American, 41(14):219, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-219.pdf.

Anonymous:1879:BNBba

[13374] Anonymous. Books: New books and publications. Scientific American, 41(14):219, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
219a.pdf.

Anonymous:1879:NQbh

[13375] Anonymous. Notes and queries. Scientific American, 41(14):219, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-219b.pdf.

Anonymous:1879:IIbn

[13376] Anonymous. Index of inventions. Scientific American, 41(14):219–220, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-219d.pdf.

Anonymous:1879:PNPbh

[13377] Anonymous. Patent notes: Patents. Scientific American, 41(14):222, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-222.pdf.

Anonymous:1879:GTN

[13378] Anonymous. The grain trade of New York. Scientific American, 41(14):208, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
208.pdf.

Anonymous:1879:FEN

[13379] Anonymous. Former extension northward of South America. Scientific American, 41(14):208–209, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
208a.pdf.

Anonymous:1879:FOC

[13380] Anonymous. The future of organic chemistry. Scientific American, 41(14):209, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
209.pdf.

Anonymous:1879:UJE

[13381] Anonymous. The use of the Jew’s ear fungus in China. Scientific American, 41(14):209, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
209a.pdf.

Anonymous:1879:CRCa

[13382] Anonymous. The cost of railway cars. Scientific American, 41(14):209, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-209b.pdf.

Anonymous:1879:VCW

[13383] Anonymous. Vertical car wheel borer. Scientific American, 41(14):210, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-210.pdf.

Anonymous:1879:NHC

[13384] Anonymous. New horse clipping machine. Scientific American, 41(14):210, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
210a.pdf.

Anonymous:1879:MIz

[13385] Anonymous. Miscellaneous inventions. Scientific American, 41(14):210, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-210b.pdf.

Mayer:1879:PEMb

[13386] Alfred M. Mayer. Practical experiments in magnetism, with special reference to the demagnetization of watches. — no. 2. Scientific American, 41(14):211–214, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
211.pdf.

Anonymous:1879:IET

[13387] Anonymous. Improved Engineer’s transit. Scientific American, 41(14):214, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
214.pdf.

Anonymous:1879:MIba

[13388] Anonymous. Mechanical inventions. Scientific American, 41(14):214, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-214a.pdf.

Anonymous:1879:HSb

[13389] Anonymous. The Holloway Sanatorium. Scientific American, 41 (14):215, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
215.pdf.

Anonymous:1879:Tb

[13390] Anonymous. The tamarind. Scientific American, 41(14):215–216, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n14/pdf/scientificamerican10041879-215a.pdf.

Anonymous:1879:JBV

[13391] Anonymous. Japanese bronze vases. Scientific American, 41 (14):217–218, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
217.pdf.

Anonymous:1879:RDRd

[13392] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(14):218, October 4, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n14/pdf/scientificamerican10041879-218.pdf.

Anonymous:1879:CRbl

[13393] Anonymous. Communications received. Scientific American, 41 (14):219, October 4, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n14/pdf/scientificamerican10041879-
219c.pdf.

Anonymous:1879:BPbp

[13394] Anonymous. Business and personal. Scientific American, 41 (15):235, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
235.pdf.

Anonymous:1879:BNBbb

[13395] Anonymous. Books: New books and publications. Scientific American, 41(15):235, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
235a.pdf.

Anonymous:1879:NQbi

[13396] Anonymous. Notes and queries. Scientific American, 41(15): 235–236, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
235b.pdf.

Anonymous:1879:IIbo

[13397] Anonymous. Index of inventions. Scientific American, 41(15):236, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n15/pdf/scientificamerican10111879-236a.pdf.

Anonymous:1879:PNPbi

[13398] Anonymous. Patent notes: Patents. Scientific American, 41 (15):238, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
238.pdf.

Anonymous:1879:H

[13399] Anonymous. The hydromotor. Scientific American, 41(15):223–224, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n15/pdf/scientificamerican10111879-223.pdf.

Anonymous:1879:TWS

[13400] Anonymous. The temperature of water supply. Scientific American, 41(15):224, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
224.pdf.

Anonymous:1879:SE

[13401] Anonymous. The Sydney exhibition. Scientific American, 41 (15):224, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
224a.pdf.

Anonymous:1879:WIU

[13402] Anonymous. Water for industrial uses in New York. Scientific American, 41(15):224, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
224b.pdf.

Anonymous:1879:CNP

[13403] Anonymous. Conspiracies to nullify the patent laws. Scientific American, 41(15):224–225, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
224c.pdf.

Anonymous:1879:EOM

[13404] Anonymous. Extraction of oils by means of solvents. Scientific American, 41(15):225, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
225.pdf.

Anonymous:1879:PPL

[13405] Anonymous. The policy of patent laws. Scientific American, 41 (15):225, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
225a.pdf.

Anonymous:1879:MIbb

[13406] Anonymous. Mechanical inventions. Scientific American, 41(15): 226, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
226.pdf.

Anonymous:1879:RCd

[13407] Anonymous. Rosset’s clock. Scientific American, 41(15):226, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n15/pdf/scientificamerican10111879-226a.pdf.

Mayer:1879:PEMc

[13408] Alfred M. Mayer. Practical experiments in magnetism, with special reference to the demagnetization of watches. — no. 3. Scientific American, 41(15):227–229, October 11, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n15/pdf/scientificamerican10111879-227.pdf.

Anonymous:1879:NDSc

[13409] Anonymous. A new drag sawing machine. Scientific American, 41(15):230, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
230.pdf.

Anonymous:1879:NFEb

[13410] Anonymous. A new fire escape. Scientific American, 41(15):230, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n15/pdf/scientificamerican10111879-230a.pdf.

Anonymous:1879:MIbc

[13411] Anonymous. Miscellaneous inventions. Scientific American, 41 (15):230–231, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
230b.pdf.

Anonymous:1879:NCAa

[13412] Anonymous. New calculating attachment for weighing scales. Scientific American, 41(15):231, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
231.pdf.

Anonymous:1879:Oa

[13413] Anonymous. The otocyon. Scientific American, 41(15):231, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-

7087 (electronic). URL http://en.wikipedia.org/wiki/Otocyon;

 http://www.nature.com/scientificamerican/journal/v41/n15/
pdf/scientificamerican10111879-231a.pdf.

Anonymous:1879:RAIe

[13414] Anonymous. Recent agricultural inventions. Scientific American, 41(15):231, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
231b.pdf.

Anonymous:1879:NHNg

[13415] Anonymous. Natural history notes. Scientific American, 41(15): 231–233, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
231c.pdf.

Anonymous:1879:FFc

[13416] Anonymous. French faience. Scientific American, 41(15):233, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n15/pdf/scientificamerican10111879-233.pdf.

Anonymous:1879:BEA

[13417] Anonymous. The baneful effects of absinthe. Scientific American, 41(15):234, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
234.pdf.

Anonymous:1879:CRbm

[13418] Anonymous. Communications received. Scientific American, 41 (15):236, October 11, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n15/pdf/scientificamerican10111879-
236.pdf.

Anonymous:1879:Cbl

[13419] Anonymous. Correspondence. Scientific American, 41(16):245, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-245a.pdf.

Anonymous:1879:BPbq

[13420] Anonymous. Business and personal. Scientific American, 41 (16):251, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
251.pdf.

Anonymous:1879:BNBbc

[13421] Anonymous. Books: New books and publications. Scientific American, 41(16):251, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
251a.pdf.

Anonymous:1879:NQbj

[13422] Anonymous. Notes and queries. Scientific American, 41(16): 251–252, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
251b.pdf.

Anonymous:1879:IIbp

[13423] Anonymous. Index of inventions. Scientific American, 41(16):252, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-252a.pdf.

Anonymous:1879:PNPbj

[13424] Anonymous. Patent notes: Patents. Scientific American, 41 (16):254, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
254.pdf.

Anonymous:1879:CWS

[13425] Anonymous. Concrete work at the sea ends of Mississippi jetties. Scientific American, 41(16):240, October 18, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n16/pdf/scientificamerican10181879-240.pdf.

Anonymous:1879:UPW

[13426] Anonymous. Use of photography in wood engraving. Scientific American, 41(16):240, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
240a.pdf.

Anonymous:1879:HRT

[13427] Anonymous. The Hudson River Tunnel. Scientific American, 41 (16):240, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
240b.pdf.

Anonymous:1879:PBU

[13428] Anonymous. The production of bromine in the United States. Scientific American, 41(16):240–241, October 18, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n16/pdf/scientificamerican10181879-240c.pdf.

Anonymous:1879:BWB

[13429] Anonymous. Bad water in Baltimore. Scientific American, 41 (16):241, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
241.pdf.

Anonymous:1879:ESM

[13430] Anonymous. England’s sources of motive power. Scientific American, 41(16):241, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
241a.pdf.

Anonymous:1879:CCT

[13431] Anonymous. Cold climates in the treatment of consumption. Scientific American, 41(16):241, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
241b.pdf.

Anonymous:1879:EEGa

[13432] Anonymous. Edison’s electrical generator. Scientific American, 41(16):242, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
242.pdf.

Anonymous:1879:MIbd

[13433] Anonymous. Mechanical inventions. Scientific American, 41(16): 242, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
242a.pdf.

Anonymous:1879:FHF

[13434] Anonymous. French horizontal flour mills. Scientific American, 41(16):243, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
243.pdf.

Anonymous:1879:CALb

[13435] Anonymous. Carving attachment for lathes. Scientific American, 41(16):243, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
243a.pdf.

Anonymous:1879:MIbe

[13436] Anonymous. Miscellaneous inventions. Scientific American, 41 (16):243–244, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
243b.pdf.

Anonymous:1879:TS

[13437] Anonymous. Tortoise shell. Scientific American, 41(16):244, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-244.pdf.

Anonymous:1879:LAB

[13438] Anonymous. The localization of arsenic in the brain. Scientific American, 41(16):244, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
244a.pdf.

Anonymous:1879:Cbm

[13439] Anonymous. Chlorophyl. Scientific American, 41(16):244, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-244b.pdf.

Anonymous:1879:RDRe

[13440] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(16):245, October 18, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n16/pdf/scientificamerican10181879-245.pdf.

Anonymous:1879:SAb

[13441] Anonymous. A suggestion to advertisers. Scientific American, 41(16):245, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
245b.pdf.

Anonymous:1879:EIp

[13442] Anonymous. Engineering inventions. Scientific American, 41(16): 246, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
246.pdf.

Anonymous:1879:NRPb

[13443] Anonymous. New recording pocketbook. Scientific American, 41 (16):246, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
246a.pdf.

Anonymous:1879:TC

[13444] Anonymous. The Tyler cartridge. Scientific American, 41(16):246, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-246b.pdf.

Anonymous:1879:NBC

[13445] Anonymous. New Barber’s chair. Scientific American, 41(16):246, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-246c.pdf.

Anonymous:1879:Td

[13446] Anonymous. The tarsier. Scientific American, 41(16):247, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-247.pdf.

Anonymous:1879:DWF

[13447] Anonymous. Decoys for wild fowl. Scientific American, 41(16):247, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n16/pdf/scientificamerican10181879-247a.pdf.

Anonymous:1879:SRAb

[13448] Anonymous. Some recent American patents. Scientific American, 41(16):247–249, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
247b.pdf.

Anonymous:1879:AIb

[13449] Anonymous. Agricultural inventions. Scientific American, 41(16): 249–251, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
249.pdf.

Anonymous:1879:CRbn

[13450] Anonymous. Communications received. Scientific American, 41 (16):252, October 18, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n16/pdf/scientificamerican10181879-
252.pdf.

Anonymous:1879:BPbr

[13451] Anonymous. Business and personal. Scientific American, 41 (17):267, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
267.pdf.

Anonymous:1879:NQbk

[13452] Anonymous. Notes and queries. Scientific American, 41(17):267, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-267a.pdf.

Anonymous:1879:IIbq

[13453] Anonymous. Index of inventions. Scientific American, 41(17):268, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-268.pdf.

Anonymous:1879:PNPbk

[13454] Anonymous. Patent notes: Patents. Scientific American, 41 (17):270, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
270.pdf.

Anonymous:1879:PER

[13455] Anonymous. The progress of elevated railways. Scientific American, 41(17):255, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
255.pdf.

Anonymous:1879:AIF

[13456] Anonymous. The American Institute Fair. Scientific American, 41(17):256, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
256.pdf.

Anonymous:1879:ZGN

[13457] Anonymous. A Zoological Garden for New York. Scientific American, 41(17):256, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
256a.pdf.

Anonymous:1879:OMM

[13458] Anonymous. The origin of machine-made pens. Scientific American, 41(17):256, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
256b.pdf.

Anonymous:1879:NTEb

[13459] Anonymous. New torpedo experiments by Capt. Ericsson. Scientific American, 41(17):257, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
257.pdf.

Anonymous:1879:MF

[13460] Anonymous. Metallic fences. Scientific American, 41(17):257, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-257a.pdf.

Anonymous:1879:MIbf

[13461] Anonymous. Mechanical inventions. Scientific American, 41(17): 259, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
259.pdf.

Anonymous:1879:SRMb

[13462] Anonymous. Shearing and riveting machinery. Scientific American, 41(17):259–260, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
259a.pdf.

Anonymous:1879:AId

[13463] Anonymous. Agricultural inventions. Scientific American, 41(17): 260–262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
260.pdf.

Anonymous:1879:ISa

[13464] Anonymous. An improvement in stovepipes. Scientific American, 41(17):262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
262.pdf.

Anonymous:1879:NHR

[13465] Anonymous. New hydraulic ram. Scientific American, 41(17):262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-262a.pdf.

Anonymous:1879:NAa

[13466] Anonymous. New amalgamator. Scientific American, 41(17):262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-262b.pdf.

Anonymous:1879:NCf

[13467] Anonymous. A novel candle. Scientific American, 41(17):262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-262c.pdf.

Anonymous:1879:EIq

[13468] Anonymous. Engineering inventions. Scientific American, 41(17): 262, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
262d.pdf.

Anonymous:1879:EJ

[13469] Anonymous. Electric jewelry. Scientific American, 41(17):263, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-263.pdf.

Anonymous:1879:PPd

[13470] Anonymous. Palissy plate. Scientific American, 41(17):263, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-263a.pdf.

Anonymous:1879:MIbg

[13471] Anonymous. Miscellaneous inventions. Scientific American, 41 (17):263–265, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
263b.pdf.

Anonymous:1879:BF

[13472] Anonymous. Baptismal font. Scientific American, 41(17):265–267, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n17/pdf/scientificamerican10251879-265.pdf.

Anonymous:1879:CRbo

[13473] Anonymous. Communications received. Scientific American, 41 (17):267, October 25, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n17/pdf/scientificamerican10251879-
267b.pdf.

Anonymous:1879:Cbn

[13474] Anonymous. Correspondence. Scientific American, 41(18):276, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n18/pdf/scientificamerican11011879-276.pdf.

Anonymous:1879:BPbs

[13475] Anonymous. Business and personal. Scientific American, 41 (18):283, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
283.pdf.

Anonymous:1879:BNBbd

[13476] Anonymous. Books: New books and publications. Scientific American, 41(18):283, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
283a.pdf.

Anonymous:1879:NQbl

[13477] Anonymous. Notes and queries. Scientific American, 41(18): 283–284, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
283b.pdf.

Anonymous:1879:IIbr

[13478] Anonymous. Index of inventions. Scientific American, 41(18): 284, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
284a.pdf.

Anonymous:1879:PNPbl

[13479] Anonymous. Patent notes: Patents. Scientific American, 41 (18):286, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
286.pdf.

Anonymous:1879:PL

[13480] Anonymous. Protection from lightning. Scientific American, 41 (18):272, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
272.pdf.

Anonymous:1879:FFd

[13481] Anonymous. Freezing in fire. Scientific American, 41(18):272, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n18/pdf/scientificamerican11011879-272a.pdf.

Anonymous:1879:EHC

[13482] Anonymous. Effects of heat in the Comstock Mines. Scientific American, 41(18):272–273, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
272b.pdf.

Anonymous:1879:PIB

[13483] Anonymous. Pushing an iron bridge across a river. Scientific American, 41(18):273, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
273.pdf.

Anonymous:1879:ERE

[13484] Anonymous. The elevated railway extension. — details of construction. Scientific American, 41(18):273, November 1, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n18/pdf/scientificamerican11011879-273a.pdf.

Anonymous:1879:AINk

[13485] Anonymous. American industries. — no. 21. Scientific American, 41 (18):273–274, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
273b.pdf.

Anonymous:1879:BT

[13486] Anonymous. The Blake transmitter. Scientific American, 41 (18):274, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
274.pdf.

Hopkins:1879:MRW

[13487] Geo. M. Hopkins. Modification of the Reynier and Werdermann electric lamp. Scientific American, 41(18):274, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n18/pdf/scientificamerican11011879-274a.pdf.

Anonymous:1879:MIbh

[13488] Anonymous. Mechanical inventions. Scientific American, 41(18): 274, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
274b.pdf.

Anonymous:1879:NFW

[13489] Anonymous. A new fish way. Scientific American, 41(18):275, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n18/pdf/scientificamerican11011879-275.pdf.

Anonymous:1879:MIbi

[13490] Anonymous. Miscellaneous inventions. Scientific American, 41 (18):275–276, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
275a.pdf.

Wheeler:1879:RDR

[13491] Judge Wheeler. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(18):277, November 1, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n18/pdf/scientificamerican11011879-277.pdf.

Anonymous:1879:NMM

[13492] Anonymous. New motor and meter. Scientific American, 41 (18):278, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
278.pdf.

Anonymous:1879:EIr

[13493] Anonymous. Engineering inventions. Scientific American, 41(18): 278, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
278a.pdf.

Anonymous:1879:NW

[13494] Anonymous. A new wrench. Scientific American, 41(18):278, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n18/pdf/scientificamerican11011879-278b.pdf.

Beard:1879:GSC

[13495] Daniel C. Beard. The great sea cow of Florida. Scientific American, 41(18):279–280, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
279.pdf.

Anonymous:1879:APb

[13496] Anonymous. Arrow poisons. Scientific American, 41(18):280, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n18/pdf/scientificamerican11011879-280.pdf.

Anonymous:1879:ACc

[13497] Anonymous. Arabesque cabinet. Scientific American, 41(18): 281–282, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
281.pdf.

Anonymous:1879:CRbp

[13498] Anonymous. Communications received. Scientific American, 41 (18):284, November 1, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n18/pdf/scientificamerican11011879-
284.pdf.

Anonymous:1879:Cbo

[13499] Anonymous. Correspondence. Scientific American, 41(19):292– 293, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
292.pdf.

Anonymous:1879:BPbt

[13500] Anonymous. Business and personal. Scientific American, 41 (19):299, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
299.pdf.

Anonymous:1879:BNBbe

[13501] Anonymous. Books: New books and publications. Scientific American, 41(19):299, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
299a.pdf.

Anonymous:1879:NQbm

[13502] Anonymous. Notes and queries. Scientific American, 41(19): 299–300, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
299b.pdf.

Anonymous:1879:IIbs

[13503] Anonymous. Index of inventions. Scientific American, 41(19): 300, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
300a.pdf.

Anonymous:1879:PNPbm

[13504] Anonymous. Patent notes: Patents. Scientific American, 41 (19):302, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
302.pdf.

Anonymous:1879:SIY

[13505] Anonymous. The speed of ice yachts. Scientific American, 41 (19):288, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
288.pdf.

Anonymous:1879:SRAc

[13506] Anonymous. Some reasons for American success. Scientific American, 41(19):288, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
288a.pdf.

Hovet:1879:FUS

[13507] H. C. Hovet. The fisheries of the United States. Scientific American, 41(19):289, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
289.pdf.

Anonymous:1879:CTM

[13508] Anonymous. Coal tipping machines and screens. Scientific American, 41(19):290, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
290.pdf.

Anonymous:1879:EIs

[13509] Anonymous. Engineering inventions. Scientific American, 41(19): 291, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
291.pdf.

Anonymous:1879:NBH

[13510] Anonymous. Novel barrel hoop. Scientific American, 41(19): 291, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
291a.pdf.

Anonymous:1879:NWMb

[13511] Anonymous. New wind mill. Scientific American, 41(19):291, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n19/pdf/scientificamerican11081879-291b.pdf.

Anonymous:1879:CCBa

[13512] Anonymous. Combined corkscrew and bottle faucet. Scientific American, 41(19):291, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
291c.pdf.

Anonymous:1879:ELE

[13513] Anonymous. An electric lamp for an English shilling. Scientific American, 41(19):293, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
293.pdf.

Edison:1879:RTF

[13514] Th. A. Edison. On a resonant tuning fork. Scientific American, 41(19):293, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
293a.pdf.

Anonymous:1879:SPR

[13515] Anonymous. Steak pavement rammer. Scientific American, 41 (19):294, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
294.pdf.

Anonymous:1879:MP

[13516] Anonymous. A mammoth plow. Scientific American, 41(19): 294, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
294a.pdf.

Anonymous:1879:MIbj

[13517] Anonymous. Miscellaneous inventions. Scientific American, 41 (19):294, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
294b.pdf.

Anonymous:1879:FFe

[13518] Anonymous. The flying Fox. Scientific American, 41(19):295, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n19/pdf/scientificamerican11081879-295.pdf.

Anonymous:1879:AINl

[13519] Anonymous. American industries. — no. 22. Scientific American, 41(19):296, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
296.pdf.

Anonymous:1879:AIe

[13520] Anonymous. Agricultural inventions. Scientific American, 41(19): 296–297, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
296a.pdf.

Anonymous:1879:MIbk

[13521] Anonymous. Mechanical inventions. Scientific American, 41(19): 297, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
297.pdf.

Anonymous:1879:MCb

[13522] Anonymous. Mysterious clock. Scientific American, 41(19):297– 298, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
297a.pdf.

Anonymous:1879:CRbq

[13523] Anonymous. Communications received. Scientific American, 41 (19):300, November 8, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n19/pdf/scientificamerican11081879-
300.pdf.

Anonymous:1879:Cbp

[13524] Anonymous. Correspondence. Scientific American, 41(20):308, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n20/pdf/scientificamerican11151879-308.pdf.

Anonymous:1879:BPbu

[13525] Anonymous. Business and personal. Scientific American, 41 (20):315, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
315.pdf.

Anonymous:1879:NQbn

[13526] Anonymous. Notes and queries. Scientific American, 41(20): 315, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
315a.pdf.

Anonymous:1879:IIbt

[13527] Anonymous. Index of inventions. Scientific American, 41(20): 315–316, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
315c.pdf.

Anonymous:1879:PNPbn

[13528] Anonymous. Patent notes: Patents. Scientific American, 41 (20):316, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
316.pdf.

Anonymous:1879:AINm

[13529] Anonymous. American industries. — no. 23. Scientific American, 41(20):303, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
303.pdf.

Anonymous:1879:FTD

[13530] Anonymous. Five thousand dollars reward for the invention of a stock car. Scientific American, 41(20):304, November 15, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n20/pdf/scientificamerican11151879-304.pdf.

Anonymous:1879:AVFb

[13531] Anonymous. Anti-vaccination folly. Scientific American, 41(20): 304, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
304a.pdf.

Anonymous:1879:FWS

[13532] Anonymous. The future water supply of Philadelphia. Scientific American, 41(20):304–305, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
304b.pdf.

Seeley:1879:EEG

[13533] Charles A. Seeley. Edison’s electrical generator. Scientific American, 41(20):305, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
305.pdf.

Anonymous:1879:BSCb

[13534] Anonymous. Blake’s sectional cushioned crusher. Scientific American, 41(20):306, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
306.pdf.

Anonymous:1879:MIbl

[13535] Anonymous. Miscellaneous inventions. Scientific American, 41 (20):306, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
306a.pdf.

Anonymous:1879:ISCb

[13536] Anonymous. Improved sleeping-car berth. Scientific American, 41 (20):307, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
307.pdf.

Anonymous:1879:NPDb

[13537] Anonymous. New perspective drawing apparatus. Scientific American, 41(20):307, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
307a.pdf.

Anonymous:1879:MIbm

[13538] Anonymous. Mechanical inventions. Scientific American, 41(20): 307–308, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
307b.pdf.

Anonymous:1879:ALS

[13539] Anonymous. The abuse of live stock on the way to market. Scientific American, 41(20):309, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
309.pdf.

Anonymous:1879:FRI

[13540] Anonymous. A few recent inventions. Scientific American, 41 (20):310, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
310.pdf.

Anonymous:1879:CTN

[13541] Anonymous. Cost and traffic of the New York elevated railways. Scientific American, 41(20):310, November 15, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n20/pdf/scientificamerican11151879-310a.pdf.

Anonymous:1879:RG

[13542] Anonymous. Roman glassware. Scientific American, 41(20):311– 312, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
311.pdf.

Anonymous:1879:Ag

[13543] Anonymous. Attraction. Scientific American, 41(20):312, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n20/pdf/scientificamerican11151879-312.pdf.

Anonymous:1879:NASb

[13544] Anonymous. The National Academy of Sciences. Scientific American, 41(20):313, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
313.pdf.

Anonymous:1879:RSM

[13545] Anonymous. A recent study of Mexican earthquakes. Scientific American, 41(20):314, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
314.pdf.

Anonymous:1879:CRbr

[13546] Anonymous. Communication received. Scientific American, 41 (20):315, November 15, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n20/pdf/scientificamerican11151879-
315b.pdf.

Anonymous:1879:Cbq

[13547] Anonymous. Correspondence. Scientific American, 41(21):324– 325, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
324.pdf.

Anonymous:1879:BPbv

[13548] Anonymous. Business and personal. Scientific American, 41 (21):331, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
331.pdf.

Anonymous:1879:BNBbf

[13549] Anonymous. Books: New books and publications. Scientific American, 41(21):331, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
331a.pdf.

Anonymous:1879:NQbo

[13550] Anonymous. Notes and queries. Scientific American, 41(21): 331–332, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
331b.pdf.

Anonymous:1879:IIbu

[13551] Anonymous. Index of inventions. Scientific American, 41(21): 332, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
332.pdf.

Anonymous:1879:PNPbo

[13552] Anonymous. Patent notes: Patents. Scientific American, 41 (21):334, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
334.pdf.

Anonymous:1879:FE

[13553] Anonymous. Friedmann’s ejector. Scientific American, 41(21): 319, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
319.pdf.

Anonymous:1879:ISb

[13554] Anonymous. Incendiary silk. Scientific American, 41(21):320, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n21/pdf/scientificamerican11221879-320.pdf.

Anonymous:1879:OSM

[13555] Anonymous. Oriental sand and mud baths. Scientific American, 41(21):320, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
320a.pdf.

Anonymous:1879:NAc

[13556] Anonymous. The National Academy. Scientific American, 41(21): 320–321, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
320b.pdf.

Anonymous:1879:BD

[13557] Anonymous. Boadmasters’ difficulties. Scientific American, 41 (21):321, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
321.pdf.

Anonymous:1879:CPb

[13558] Anonymous. The “conch peael”. Scientific American, 41(21): 321, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
321a.pdf.

Anonymous:1879:BEc

[13559] Anonymous. Blast engine. Scientific American, 41(21):322–323, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n21/pdf/scientificamerican11221879-322.pdf.

Anonymous:1879:NDP

[13560] Anonymous. A new device for preventing journals from heating. Scientific American, 41(21):323, November 22, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n21/pdf/scientificamerican11221879-323.pdf.

Anonymous:1879:NSGb

[13561] Anonymous. Novel steak generator. Scientific American, 41(21): 323, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
323a.pdf.

Anonymous:1879:AIg

[13562] Anonymous. Agricultural inventions. Scientific American, 41 (21):323, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
323b.pdf.

Anonymous:1879:MIbn

[13563] Anonymous. Mechanical inventions. Scientific American, 41(21): 323–324, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
323c.pdf.

Anonymous:1879:HWN

[13564] Anonymous. How to work the new copying process. Scientific American, 41(21):325, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
325.pdf.

Anonymous:1879:ICS

[13565] Anonymous. Improved cabinet seat. Scientific American, 41 (21):325, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
325a.pdf.

Anonymous:1879:MIbo

[13566] Anonymous. Miscellaneous intentions. Scientific American, 41(21): 325–326, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
325b.pdf.

Anonymous:1879:CTa

[13567] Anonymous. The cake of tools. Scientific American, 41(21):326, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n21/pdf/scientificamerican11221879-326.pdf.

Anonymous:1879:ISe

[13568] Anonymous. An improved spindle. Scientific American, 41(21): 327, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
327.pdf.

Anonymous:1879:Ob

[13569] Anonymous. The otter. Scientific American, 41(21):327, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n21/pdf/scientificamerican11221879-327a.pdf.

Anonymous:1879:NBL

[13570] Anonymous. New barrel lifter. Scientific American, 41(21):327, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n21/pdf/scientificamerican11221879-327b.pdf.

Anonymous:1879:NHNh

[13571] Anonymous. Natural history notes. Scientific American, 41(21): 328–330, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
328.pdf.

Anonymous:1879:EIt

[13572] Anonymous. Engineering inventions. Scientific American, 41(21): 330, November 22, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n21/pdf/scientificamerican11221879-
330.pdf.

Anonymous:1879:Cbr (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
340b.pdf.

Anonymous:1879:BPbw

[13574] Anonymous. Business and personal. Scientific American, 41 (22):347, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
347.pdf.

Anonymous:1879:BNBbg

[13575] Anonymous. Books: New books and publications. Scientific American, 41(22):347, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
347a.pdf.

Anonymous:1879:NQbp

[13576] Anonymous. Notes and queries. Scientific American, 41(22): 347, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
347b.pdf.

Anonymous:1879:IIbv

[13577] Anonymous. Index of inventions. Scientific American, 41(22): 347, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
347c.pdf.

Anonymous:1879:PNPbp

[13578] Anonymous. Patent notes: Patents. Scientific American, 41 (22):350, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
350.pdf.

Anonymous:1879:HTT

[13579] Anonymous. The Holyoke turbine tests. Scientific American, 41 (22):336, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
336.pdf.

Anonymous:1879:CSb

[13580] Anonymous. Collisions at sea. Scientific American, 41(22):336, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n22/pdf/scientificamerican11291879-336a.pdf.

Anonymous:1879:APH

[13581] Anonymous. The American public health association. Scientific American, 41(22):336–337, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
336b.pdf.

Anonymous:1879:USB

[13582] Anonymous. The use of steel for bridges. Scientific American, 41(22):337, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
337.pdf.

Anonymous:1879:EEGb

[13583] Anonymous. Edison’s electric generator. Scientific American, 41 (22):337, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
337a.pdf.

Anonymous:1879:HFC

[13584] Anonymous. How far can we hear with the telephone? Scientific American, 41(22):337, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
337b.pdf.

Anonymous:1879:NV

[13585] Anonymous. A new velocipede. Scientific American, 41(22): 338, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
338.pdf.

Anonymous:1879:NFP

[13586] Anonymous. New foundation plate for elevated railways. Scientific American, 41(22):338, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
338a.pdf.

Anonymous:1879:MIbp

[13587] Anonymous. Miscellaneous inventions. Scientific American, 41 (22):338, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
338b.pdf.

Anonymous:1879:NSf

[13588] Anonymous. A new sweeper. Scientific American, 41(22):339, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n22/pdf/scientificamerican11291879-339.pdf.

Anonymous:1879:NAFb

[13589] Anonymous. New automatic fire lighter. Scientific American, 41 (22):339, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
339a.pdf.

Anonymous:1879:AINn

[13590] Anonymous. American industries — no. 24. Scientific American, 41(22):340, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
340.pdf.

Anonymous:1879:SRIc

[13591] Anonymous. Some recent inventions. Scientific American, 41 (22):340, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
340a.pdf.

Anonymous:1879:NMI

[13592] Anonymous. New musical instrument. Scientific American, 41 (22):342, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
342.pdf.

Anonymous:1879:NWC

[13593] Anonymous. New window cleaning chair. Scientific American, 41 (22):342–343, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
342a.pdf.

Anonymous:1879:Th

[13594] Anonymous. Tadpoles. Scientific American, 41(22):343, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n22/pdf/scientificamerican11291879-343.pdf.

Anonymous:1879:HBM

[13595] Anonymous. House-builder moth. Scientific American, 41(22): 343–345, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
343a.pdf.

Anonymous:1879:RDRf

[13596] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(22):346, November 29, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n22/pdf/scientificamerican11291879-346.pdf.

Anonymous:1879:MIbq

[13597] Anonymous. Mechanical inventions. Scientific American, 41(22): 346, November 29, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n22/pdf/scientificamerican11291879-
346a.pdf.

Anonymous:1879:Cbs

[13598] Anonymous. Correspondence. Scientific American, 41(23):360, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-360.pdf.

Anonymous:1879:BPbx

[13599] Anonymous. Business and personal. Scientific American, 41 (23):363, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
363.pdf.

Anonymous:1879:BNBbh

[13600] Anonymous. Books: New books and publications. Scientific American, 41(23):363, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
363a.pdf.

Anonymous:1879:NQbq

[13601] Anonymous. Notes and queries. Scientific American, 41(23): 363–364, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
363b.pdf.

Anonymous:1879:IIbw

[13602] Anonymous. Index of inventions. Scientific American, 41(23):364, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-364a.pdf.

Anonymous:1879:PNPbq

[13603] Anonymous. Patent notes: Patents. Scientific American, 41 (23):364, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
364b.pdf.

Anonymous:1879:TMD

[13604] Anonymous. The trade mark decision. Scientific American, 41 (23):352, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
352.pdf.

Anonymous:1879:WTBa

[13605] Anonymous. A Western tow-boat. Scientific American, 41(23):352, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-352a.pdf.

Anonymous:1879:GGW

[13606] Anonymous. The great gas well and carbon factory at Murraysville, PA. Scientific American, 41(23):352–353, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-352b.pdf.

Anonymous:1879:NYAd

[13607] Anonymous. New York Academy of Sciences. Scientific American, 41(23):353, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
353.pdf.

Anonymous:1879:NPHa

[13608] Anonymous. The National Public Health Association. Scientific American, 41(23):353, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
353a.pdf.

Anonymous:1879:NTMc

[13609] Anonymous. New tramway motor. Scientific American, 41(23):353, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-354.pdf.

Baldwin:1879:HYS

[13610] William J. Baldwin. Hints to the Young steam fitter. Scientific American, 41(23):355–356, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
355.pdf.

Anonymous:1879:AINo

[13611] Anonymous. American industries — no. 25. Scientific American, 41(23):357, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
357.pdf.

Anonymous:1879:MIbr

[13612] Anonymous. Miscellaneous inventions. Scientific American, 41 (23):357–358, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
357a.pdf.

Anonymous:1879:AH

[13613] Anonymous. Artificial haymaking. Scientific American, 41(23):358, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-358.pdf.

Anonymous:1879:ISDb

[13614] Anonymous. An improved screw driver. Scientific American, 41 (23):358, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
358a.pdf.

Anonymous:1879:MIbs

[13615] Anonymous. Mechanical inventions. Scientific American, 41(23): 358, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
358b.pdf.

Anonymous:1879:Iba

[13616] Anonymous. The ibex. Scientific American, 41(23):359, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-359.pdf.

Anonymous:1879:NHNi

[13617] Anonymous. Natural history notes. Scientific American, 41(23): 359–360, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
359a.pdf.

Anonymous:1879:CBAc

[13618] Anonymous. Cross-breeding among fish. Scientific American, 41 (23):360, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
360a.pdf.

Anonymous:1879:IDF

[13619] Anonymous. The international dairy fair. Scientific American, 41(23):360, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
360b.pdf.

Anonymous:1879:STE

[13620] Anonymous. Steam towing on the Erie Canal. Scientific American, 41(23):360, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
360c.pdf.

Anonymous:1879:GTc

[13621] Anonymous. Glass tubing. Scientific American, 41(23):361, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n23/pdf/scientificamerican12061879-361.pdf.

Anonymous:1879:AIh

[13622] Anonymous. Agricultural inventions. Scientific American, 41(23): 361–362, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
361a.pdf.

Anonymous:1879:CRbs

[13623] Anonymous. Communications received. Scientific American, 41 (23):364, December 6, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n23/pdf/scientificamerican12061879-
364.pdf.

Anonymous:1879:Cbt

[13624] Anonymous. Correspondence. Scientific American, 41(24):380– 381, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
380a.pdf.

Anonymous:1879:BPby

[13625] Anonymous. Business and personal. Scientific American, 41 (24):387, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
387.pdf.

Anonymous:1879:BNBbi

[13626] Anonymous. Books: New books and publications. Scientific American, 41(24):387, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
387a.pdf.

Anonymous:1879:NQbr

[13627] Anonymous. Notes and queries. Scientific American, 41(24): 387–388, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
387b.pdf.

Anonymous:1879:IIbx

[13628] Anonymous. Index of inventions. Scientific American, 41(24): 388, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
388a.pdf.

Anonymous:1879:PNPbr

[13629] Anonymous. Patent notes: Patents. Scientific American, 41 (24):388, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
388b.pdf.

Anonymous:1879:SAE

[13630] Anonymous. The Scientific American as an educator of the young. Scientific American, 41(24):376, December 13, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n24/pdf/scientificamerican12131879-376.pdf.

Anonymous:1879:RTM

[13631] Anonymous. The registration of trade marks. Scientific American, 41(24):376, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
376a.pdf.

Anonymous:1879:HJL

[13632] Anonymous. How to judge of leather in belts. Scientific American, 41(24):376–377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
376b.pdf.

Anonymous:1879:YPG

[13633] Anonymous. The year product of gold and silver. Scientific American, 41(24):377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
377.pdf.

Anonymous:1879:SPd

[13634] Anonymous. Steam pilotage. Scientific American, 41(24):377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n24/pdf/scientificamerican12131879-377a.pdf.

Anonymous:1879:Cbu

[13635] Anonymous. Chicle. Scientific American, 41(24):377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n24/pdf/scientificamerican12131879-377b.pdf.

Anonymous:1879:IWU

[13636] Anonymous. Inventions wanted to utilize sawdust. Scientific American, 41(24):377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
377c.pdf.

Anonymous:1879:NACa

[13637] Anonymous. The new Atlantic cable. Scientific American, 41 (24):377, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
377d.pdf.

Anonymous:1879:MIbt

[13638] Anonymous. Miscellaneous inventions. Scientific American, 41 (24):378, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
378.pdf.

Anonymous:1879:NCg

[13639] Anonymous. New candlestick. Scientific American, 41(24):378, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n24/pdf/scientificamerican12131879-378a.pdf.

Anonymous:1879:NHD

[13640] Anonymous. New horse detacher. Scientific American, 41(24): 378–379, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
378b.pdf.

Anonymous:1879:SLB

[13641] Anonymous. The St. Louis Bridge. Scientific American, 41 (24):379, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
379.pdf.

Anonymous:1879:ICP

[13642] Anonymous. Improvement in chain pumps. Scientific American, 41 (24):379–380, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
379a.pdf.

Anonymous:1879:AINp

[13643] Anonymous. American industries — no. 25. Scientific American, 41(24):380, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
380.pdf.

Anonymous:1879:MIbu

[13644] Anonymous. Mechanical inventions. Scientific American, 41(24): 381, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
381.pdf.

Anonymous:1879:NTEc

[13645] Anonymous. Novel traction engine. Scientific American, 41(24): 382, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
382.pdf.

Anonymous:1879:IOC

[13646] Anonymous. Improved oil cabinet. Scientific American, 41(24): 382, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
382a.pdf.

Anonymous:1879:SHW

[13647] Anonymous. A singular habit of the woodcock. Scientific American, 41(24):383, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
383.pdf.

Anonymous:1879:NHNj

[13648] Anonymous. Natural history notes. Scientific American, 41(24): 383–384, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
383a.pdf.

Anonymous:1879:RDRg

[13649] Anonymous. Recent decisions relating to patents, trade marks, etc. Scientific American, 41(24):384, December 13, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n24/pdf/scientificamerican12131879-384.pdf.

Anonymous:1879:PTA

[13650] Anonymous. Platt’s tracing apparatus. Scientific American, 41 (24):385, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
385.pdf.

Anonymous:1879:AGa

[13651] Anonymous. Ancient glassware. Scientific American, 41(24): 385, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
385a.pdf.

Anonymous:1879:MR

[13652] Anonymous. Measuring resistances. Scientific American, 41(24): 385–386, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
385b.pdf.

Anonymous:1879:CRbt

[13653] Anonymous. Communications received. Scientific American, 41 (24):388, December 13, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n24/pdf/scientificamerican12131879-
388.pdf.

Anonymous:1879:BPbz

[13654] Anonymous. Business and personal. Scientific American, 41 (25):403, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
403.pdf.

Anonymous:1879:BNBbj

[13655] Anonymous. Books: New books and publications. Scientific American, 41(25):403, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
403a.pdf.

Anonymous:1879:NQbs

[13656] Anonymous. Notes and queries. Scientific American, 41(25): 403, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
403b.pdf.

Anonymous:1879:IIby

[13657] Anonymous. Index of inventions. Scientific American, 41(25): 404, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
404.pdf.

Anonymous:1879:PNPbs

[13658] Anonymous. Patent notes: Patents. Scientific American, 41 (25):404, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
404a.pdf.

Anonymous:1879:SAT

[13659] Anonymous. Science as applied to tanning. Scientific American, 41(25):392, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
392.pdf.

Anonymous:1879:GCD

[13660] Anonymous. The great Chanoine Dam at Pittsburg. Scientific American, 41(25):392–393, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
392a.pdf.

Anonymous:1879:ASP

[13661] Anonymous. The Angamar steam passenger car. Scientific American, 41(25):393, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
393.pdf.

Anonymous:1879:NPHb

[13662] Anonymous. A new plan for heating horse cars. Scientific American, 41(25):393, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
393a.pdf.

Anonymous:1879:HWR

[13663] Anonymous. A hot water river. Scientific American, 41(25):393, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n25/pdf/scientificamerican12201879-393b.pdf.

Anonymous:1879:NYAe

[13664] Anonymous. New York Academy of Sciences. Scientific American, 41 (25):393–394, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
393c.pdf.

Anonymous:1879:SPBb

[13665] Anonymous. The steam pilot boat Hercules. Scientific American, 41(25):394, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
394.pdf.

Anonymous:1879:NTMd

[13666] Anonymous. Novel treadle motion. Scientific American, 41(25): 394, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
394a.pdf.

Anonymous:1879:MIbv

[13667] Anonymous. Miscellaneous inventions. Scientific American, 41 (25):394, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
394b.pdf.

Anonymous:1879:NPi

[13668] Anonymous. New projectile. Scientific American, 41(25):395, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n25/pdf/scientificamerican12201879-395.pdf.

Anonymous:1879:ESI

[13669] Anonymous. Electrical signaling instrument. Scientific American, 41(25):395, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
395a.pdf.

Anonymous:1879:IHN

[13670] Anonymous. Improved hose nozzle. Scientific American, 41(25): 395, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
395b.pdf.

Anonymous:1879:AINq

[13671] Anonymous. American industries — no. 26. Scientific American, 41(25):396, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
396.pdf.

Anonymous:1879:AIi

[13672] Anonymous. Agricultural inventions. Scientific American, 41(25): 396–397, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
396a.pdf.

Mott:1879:FGC

[13673] Henry A. Mott, Jr. A fungoid geowth — the cause of whooping cough. Scientific American, 41(25):397, December 20, 1879. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/v41/
n25/pdf/scientificamerican12201879-397.pdf.

Anonymous:1879:PGI

[13674] Anonymous. Pruning and grafting implement. Scientific American, 41(25):398, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
398.pdf.

Anonymous:1879:IHT

[13675] Anonymous. Improved hardening tongs. Scientific American, 41 (25):398, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
398a.pdf.

Anonymous:1879:TMc

[13676] Anonymous. Tying-in machine. Scientific American, 41(25): 398, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
398b.pdf.

Anonymous:1879:MIbw

[13677] Anonymous. Mechanical inventions. Scientific American, 41(25): 398–399, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
398c.pdf.

Anonymous:1879:NEH

[13678] Anonymous. New egg holder. Scientific American, 41(25):399, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n25/pdf/scientificamerican12201879-399.pdf.

Anonymous:1879:GMb

[13679] Anonymous. The gila monster. Scientific American, 41(25):399– 401, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
399a.pdf.

Anonymous:1879:EIu

[13680] Anonymous. Engineering inventions. Scientific American, 41(25): 401–402, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
401.pdf.

Anonymous:1879:CRbu

[13681] Anonymous. Communications received. Scientific American, 41 (25):403, December 20, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n25/pdf/scientificamerican12201879-
403c.pdf.

Anonymous:1879:BPca

[13682] Anonymous. Business and personal. Scientific American, 41 (26):417, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
417.pdf.

Anonymous:1879:NQbt

[13683] Anonymous. Notes and queries. Scientific American, 41(26): 417–418, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
417a.pdf.

Anonymous:1879:IIbz

[13684] Anonymous. Index of inventions. Scientific American, 41(26): 418, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
418.pdf.

Anonymous:1879:PNPbt

[13685] Anonymous. Patent notes: Patents. Scientific American, 41 (26):418, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
418a.pdf.

Anonymous:1879:YP

[13686] Anonymous. A Year’s progress. Scientific American, 41(26): 408, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
408.pdf.

Anonymous:1879:SWF

[13687] Anonymous. Strength, weight, and fineness of woven fabrics. Scientific American, 41(26):408, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
408a.pdf.

Anonymous:1879:EBB

[13688] Anonymous. Erastus Brigham Bigelow. Scientific American, 41 (26):409, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
409.pdf.

Anonymous:1879:RFW

[13689] Anonymous. Remarkable fly wheel explosion. Scientific American, 41(26):409, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
409a.pdf.

Anonymous:1879:SAc

[13690] Anonymous. The Scientific American for 1880. Scientific American, 41(26):409, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
409b.pdf.

Anonymous:1879:FAb

[13691] Anonymous. The future of Africa. Scientific American, 41(26): 409, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
409c.pdf.

Anonymous:1879:STM

[13692] Anonymous. Stamps for trade marks. Scientific American, 41(26): 409–410, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
409d.pdf.

Anonymous:1879:DAC

[13693] Anonymous. The duplex air compressor. Scientific American, 41 (26):410, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
410.pdf.

Anonymous:1879:NT

[13694] Anonymous. A new telephone. Scientific American, 41(26):410, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n26/pdf/scientificamerican12271879-410a.pdf.

Anonymous:1879:MIbx

[13695] Anonymous. Miscellaneous inventions. Scientific American, 41(26): 410–411, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
410b.pdf.

Anonymous:1879:CSH

[13696] Anonymous. The climax safety hammerless gun. Scientific American, 41(26):411, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
411.pdf.

Anonymous:1879:AIj

[13697] Anonymous. Agricultural inventions. Scientific American, 41 (26):412, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
412.pdf.

Anonymous:1879:BLD

[13698] Anonymous. Boat-lowering and detaching apparatus. Scientific American, 41(26):412, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
412a.pdf.

Anonymous:1879:AINr

[13699] Anonymous. American industries — no. 27. Scientific American, 41(26):413, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
413.pdf.

Anonymous:1879:MIby

[13700] Anonymous. Mechanical inventions. Scientific American, 41(26): 413–414, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
413a.pdf.

Anonymous:1879:NCSc

[13701] Anonymous. New car step. Scientific American, 41(26):414, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v41/n26/pdf/scientificamerican12271879-414.pdf.

Anonymous:1879:NSDb

[13702] Anonymous. New stone-dressing tool. Scientific American, 41(26): 414–415, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
414a.pdf.

Beard:1879:SSA

[13703] Daniel C. Beard. The Sea Serpent accounted for. Scientific American, 41(26):415–416, December 27, 1879. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v41/n26/pdf/scientificamerican12271879-
415.pdf.